A Complete Bibliography of Scientific American: 1900–1909

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.06

Title word cross-reference

1/2 [9281]. 1/4 [9330, 9486]. $25 [11277]. $250 [599]. $3 [4079]. $300 [11250]. _{1} [7293].

-Knot [9281, 9330]. -Rays [7293].

000-Bushel [10024]. 000-Horse-Power [9248]. 000-K [2709]. 000-Mile [7630]. 000-Ton [5379, 5837, 9501, 147, 93, 922, 941]. 000-Volt [4690, 8963, 3766].

10-Inch [3349]. 100-Horse-Power [8685]. 1031/2Miles [7307]. 12-Inch [2312, 5470, 5157, 8617, 8858, 11144]. 14-Inch [10340]. 15-Inch [3892]. 150-Horsepower [6607]. 150-Ton [2157]. 16-Inch [2105, 2036, 2955, 4707, 5457]. 18-Inch [2807]. 180-Ton [8082]. 1900 [1680]. 1901 [1793, 2594, 3099]. 1902 [3514, 3773, 4408, 4326]. 1903 [2888, 6063].

1905 [7698, 7932, 8378]. 1906 [9047, 9141, 8626]. 1907 [10521, 11060, 10297, 10007, 10672, 9485]. 1908 [10852, 11430, 11633, 11580, 10654, 11086, 11702]. 1909 [313].

20-000-Ton [10597]. 20-Inch [2105]. 20-Ton [4107]. 200-Mile [435]. 24-Hour [10771]. 24-Inch [1990]. 25-Knot [7910, 10736, 10729, 11816]. 250-Ton [9842]. 26-Foot [4169]. 2Miles [7307].

3 [3806, 4028]. 30-Horse-Power [8799]. 35-Horse-Power [9945]. 38-Knot [11415].

4-Cylinder [8801]. 40-Foot [2784, 7490]. 40-Knot [2144]. 400-Horse-Power [8231, 11179]. 4500 [9801].

5-Inch [3160]. 50-Caliber [3270]. 50-Horse-Power [9926, 9949].

6 [3221, 3250]. 6-Cylinder [10456]. 6-Inch [6396]. 60-Horse-Power [9925]. 612-Foot [10301]. 65-Ton [2848].

7 [4262, 8897]. 70-Foot [1700]. 700-Foot [1492, 11353]. 750-Ton [7737].

8-Cylinder [10457].

9 [4711]. 9.2-Inch [6267]. 90-Degree [10457]. 90-Foot [2668, 6162].

Abandonment [1975]. Aberfoyle [1608]. Abolish [5227]. Abolition [2793]. Aborigines [2037, 6692]. Above [1020, 4578, 1094, 3818]. Abram [5455]. Abroad [1646, 2906, 5203, 1041, 3911]. Abruzzi [1981, 6268, 6295]. Absorbers [8811, 9953, 10465]. Absorbing [5834]. Absorption [2886]. Abstracts [5181]. Absurdum [11817]. Abu [8382]. Abuse [5304]. Abydos [3242, 11432]. Abyssinia [3837]. Academies [1846, 6497]. Academy [1506, 2703, 3421, 4637, 3884, 5181, 7557, 8740, 11036, 2236, 9462, 1529, 10860, 7831, 3879, 10591, 11437, 1729]. Accelerating [9874]. Acceleration [5178, 6993, 983, 2084]. Accession [5105]. Accident [2396, 2181, 3349, 2937, 2599, 4158, 5101, 6297, 6805, 8531, 10196, 10224, 892, 452, 9214]. Accidents [3354, 5203, 6689, 8469, 8493, 9427, 8894, 9834, 9719, 10077, 10935, 11058, 11127, 11316, 11317]. Accompanying [1634]. Accomplished [835]. Account [5910, 8375]. Accumulator [6850]. Accumulators [1958, 3987, 6334, 7806]. Accuracy [11650, 88]. Acetylene [11429, 1983, 1724, 1731, 1850, 2306, 1512, 4603, 3893, 3763, 9749, 9755, 9782, 10895, 10944, 1320]. Acetylene-Generator [1983]. Acheson [9897]. Achievement [6484]. Achievements [1694]. Acid [3110, 4946, 11748, 7276, 3762]. Acids [10049]. Acoustic [7786, 436, 5835, 1249]. Acoustics [9734]. Acre [5377]. Acres [7588].

Across [4809, 5769, 6603, 3843, 8591, 1606, 3317, 3593, 4502, 5507, 6626, 6479, 11282, 851, 33, 171, 4508, 1990]. Act [6020, 7214, 7158, 11426, 652]. Acting [9906, 1021]. Actinic [2539, 5367]. Actino [2213, 6555]. Actino-Electric [2213, 6555]. Action [2045, 3037, 4959, 6334, 6392, 6294, 6220, 8375, 9246, 10235, 11659, 11411, 4103, 4940]. Active [1345, 7134, 4050]. Activity [7317, 7161, 9136, 10609, 6438]. Actual [6080]. ad [11817]. Adam [6201]. Adaptation [7194]. Adaptations [10417]. Added [11278]. Adding [8925]. Addition [10351]. Additions [7263]. Adhesion [8086]. Adjustable [1703, 3204, 2911, 3365, 3916, 4373, 7093, 7233, 8993, 8837, 9546, 8802, 9821, 9845, 10836, 11160, 120, 685, 639, 5702]. Adjuster [8240, 1304]. Adjustment [7627]. Admirable [3384]. Admiral [2122, 5974, 3107, 3978, 7327, 8294, 7705, 7565, 11166]. Admiralty [1384, 5667]. Adopt [6008]. Adopted [10496]. Adriatic [10005]. Adulterated [3657]. Adulteration [9163]. Advance [4187, 9544, 4004]. Advancement [1935, 2026, 6602, 9368, 4751, 5413, 5207, 3175]. Advances [8102]. Advancing [2635]. Advantage [10872]. Advantages [3681, 5039, 6353, 8764, 8588, 9207, 9837]. Advertising [10324, 11243]. Advice [8077]. Aerial [1510, 2613, 4622, 5500, 8701, 10217, 11259, 143, 861, 263, 1049, 590, 8500, 7702, 1126, 8029, 1609, 7189, 11757, 11134]. Aero [9487, 9756, 10962, 11198, 9011, 10474]. Aerodrome [5741, 6141]. Aerolite [11605, 557]. Aeronautic [9442, 9833, 11188, 10941, 11522, 944, 11660, 11401, 434, 1275]. Aeronautical [4266, 8321, 9011, 9327, 10320, 11727, 11148, 11815, 509, 60, 173, 531, 963, 3108]. Aeronautics [1892, 9398, 11817, 11583, 11771, 10916, 11363, 108, 253, 339, 361, 552, 607, 654, 771, 794, 847, 1015, 1091, 1162, 1248, 581, 650]. Aeronauts [10649]. Aeroplane [3117, 3858, 3905, 5671, 5723, 6013, 7727, 7959, 7399, 7631, 7564, 9443, 9479, 9510, 9325, 8664, 9899, 9820, 10452, 10286, 10613, 10313, 11345, 11776, 11425, 11481, 10991, 11757, 11194, 11253, 11802, 11134, 11434, 10876, 11641, 10943, 11400, 11472, 11452, 11490, 11675, 11105, 11300, 11046, 11756, 11083, 11135, 11937, 11945, 11830, 434, 530, 730, 167, 549, 1172, 388, 1094, 62, 741, 380, 1291, 1242, 7281, 10524, 11684, 7754, 8381, 11167]. Aeroplanes [5651, 7166, 9624, 10525, 9771, 10320, 10830, 11685, 11522, 11236, 11922, 509, 511, 591, 464, 389]. Aerostatic [2159]. Aerostatics [3756]. Affixing [10835]. Affront [1840]. Afloat [9252, 10769, 11752, 10810, 11275, 11385, 6379]. Africa [1978, 2415, 1931, 3291, 2994, 3305, 4076, 4483, 5408, 7163, 8086, 9550, 8596, 8434, 1803, 1388]. African [1765]. After [8339, 10313, 7405, 9005, 335]. Aftermath [3283, 4499, 8126, 9616]. Afterthoughts [5474]. Again [2168, 1767, 1359, 3928, 6089, 4968, 4928, 5593, 6306, 6905, 9348, 557]. Against [359, 2567, 3976, 6707, 7400, 8450, 11317, 11596]. Age [6976, 11888, 2735]. Agent [2749, 4689]. Agitating [9993]. Ago [8691]. Agree [2275]. Agreement [5620]. Agricultural [1788, 4958, 6822, 6654, 8826, 10814, 5282]. Agriculture [2412, 1361, 4340, 5165, 7319, 9224, 9262, 1065, 5628, 3814]. Aid [1412, 10916, 769, 92, 5628, 9357]. Air [2197, 1943, 1907, 2508, 2487, 1337, 1515, 2264, 1359, 1743, 1421, 1626, 3034, 3576, 2859, 2749, 4693, 3940, 5080, 5232, 5323, 6156, 7232, 6650, 7208, 6761, 6648, 9249, 8623, 9664, 8812, 8791, 8792, 8920, 9530, 8793, 8795, 8717, 10244, 10299, 10226, 9729, 9751, 10455, 10321, 10357, 10800, 11598, 11573, 11257, 11153, 11522, 11283, 1298, 539, 163, 426, 804, 832, 1020, 1016, 11111, 10758, 11088, 4689, 635, 1938, 1362, 802, 9840, 8129, 10964]. Air-Brake [7232]. Air-Brakes [2197]. Air-Compressing [9729]. Air-Cooled [6650, 6648, 8812, 8791, 8792, 8793, 8795, 10455]. Air-Directing [1943]. Air-Propelled [9249]. Air-Pump [1626]. Air-Resistance [10244]. Air-Ship [6156]. Airship [2443, 1628, 2336, 2118, 3265, 2981, 3631, 3858, 4606, 4506, 3905, 3908, 4054, 3793, 4442, 4262, 4776, 4458, 4003, 5935, 5208, 6164, 5813, 5161, 6997, 6340, 6607, 6668, 7941, 7986, 8002, 9128, 8962, 9184, 9510, 9148, 10335, 10605, 10124, 10286, 10086, 11164, 11178, 11897, 997, 56, 465, 184, 158, 776, 7347, 7902, 9352, 5811, 10717, 10378, 5602, 4492, 6984]. Airships [2536, 4190, 5832, 6265, 7043, 7295, 9631, 9461, 9624, 10709, 9910, 10556, 11685, 11522, 11236, 11828, 11923]. Akouphone [6047]. Akropolis [1370]. Alabama [1966, 3479, 10580]. Aladdin [5089]. Alarm [1389, 2532, 4192, 5599, 5986, 6588, 6442, 7135, 7184, 7839, 8839, 9867, 409, 1235, 6962]. Alarms [7515]. Alaska [2059, 4798, 11278, 4265, 7329, 195, 6023]. Alaskan [1481, 1129]. Albany [1392, 1347, 1762, 3562]. Albatross [5958]. Albino [6180]. Alcohol [2371, 2302, 3529, 2847, 3399, 4144, 3951, 4574, 5642, 7035, 6570, 7498, 7931, 9209, 8653, 8889, 8870, 9543, 9546, 9749, 10193, 9792, 966, 8935]. Alcohol-Acetylene [9749]. Alcohol-Consuming [3399]. Alcoline [618]. Alderman [5559]. Aldermen [6042]. Aldershot [3174]. Aleuts [2059]. Alexander [5569, 5164]. Alexandra [3268]. Alexandre [1607, 1344]. Alfalfa [6147]. Algerian [10267]. Algiers [3362, 7679, 4831, 3721]. Algol [9367]. Aliner [119]. Alkali [5306]. Alkaline [6850]. All-Day [8313]. Alleged [9142, 10828, 10915]. Allen [6416]. Alley [6125]. Alligators [8104]. Allowed [4870]. Alloy [9668, 475]. Aloft [11659]. along [11176]. Alps [9096]. Altamira [5784]. Alternating [4690, 5363, 4128, 10776, 1202, 866]. Alternating-Current [10776]. Alternative [7669, 8979]. Alternator [2709]. Alternators [3461]. Altitude [1981, 5880]. Altitudes [11373]. Alumina [2303]. Aluminium [3055, 5766, 9668, 10377, 11188, 8591, 571, 6438, 6360]. Alva [4736]. Alvares [7727]. always [258]. Amalfi [1317]. Amalgamation [1222]. Amalgamator [7892]. Amalgams [2049]. Amateur [1367, 5906, 2829]. Amateurs [5786, 5525]. Ambitious [10393]. Ambrose [8517, 10748]. Ambulance [1341, 4620, 4733, 7884, 9165]. Amendment [4936, 4957, 6869, 8738]. America [2323, 3283, 3271, 3667, 3252, 5472, 10374, 10708, 2117, 1577, 1338, 1965, 2299, 4362, 4062, 3805, 4571, 4119, 6028, 5106, 5956, 6807, 8463, 7650, 8448, 8137, 9011, 9540, 9487, 9480, 9644, 8832, 10334, 10962, 11216, 11376, 11363, 82, 911, 513, 535, 23, 2864, 4102, 4409, 4469, 10268, 2773, 4162, 6482, 9859, 10466, 4830, 4372, 5510, 7059, 7004, 7519, 5365]. American [3882, 5206, 5496, 9896, 1039, 9368, 5207, 3154, 1935, 1999, 1507, 1952, 2286, 2366, 2436, 2169, 1383, 1419, 1764, 1549, 1861, 1510, 1646, 1480, 2289, 2005, 1443, 2423, 1793, 2434, 1527, 1537, 3657, 2590, 2906, 3131, 3591, 3563, 3400, 3042, 3544, 3385, 3388, 2729, 2839, 2870, 3932, 3835, 4677, 3856, 4055, 4727, 4878, 4481, 4310, 3860, 5454, 5576, 5280, 5247, 5250, 5159, 5031, 5177, 5795, 5414, 5322, 5498, 5108, 5113, 6116, 6602, 7157, 6532, 6527, 7764, 8225, 8525, 7952, 7814, 7658, 7804, 7937, 7853, 7597, 7790, 7791]. American [8335, 9048, 8848, 9562, 9202, 8800, 8644, 10104, 10188, 10197, 10713, 9920, 11002, 10935, 10856, 10876, 11386, 11042, 11010, 11233, 22, 1075, 756, 761, 1298, 415, 636, 726, 79, 389, 3318, 2600, 3988, 4751, 5413, 10524, 10946, 2179, 2245, 2182, 9489, 2009, 2180, 2968, 3552, 4638, 3154, 4312, 1422, 3175, 6535, 8229, 9704, 10860, 2401, 829, 5020, 10992, 10026, 6499, 1171, 10003, 4662, 4686, 4615, 4871, 1256, 342, 10299, 10794, 10423, 9857, 10321, 10438, 11641, 11400, 11830, 82, 1172, 1198, 104, 113, 3347, 4933, 5393, 5865, 6957, 7336, 8228, 10320, 9833, 9789]. American [11150, 10974, 11815, 11011, 855]. American-Built [5113]. Americanists [5972]. Americans [5072, 4830]. America’s [9161]. Amerika [8190]. Ammunition [3881, 10749]. Amœba [5838]. Among [3493, 10641]. Amount [4849, 9029]. Amphibicycle [237]. Amphitheater [7522]. Amputation [9858]. Amur [4008]. Amusement [5188, 5335, 6127, 9374]. Amusement-Way [5188]. ana [8195]. Anachronism [11875]. Anæsthetic [7421]. Anæsthetics [8587]. Analogue [7418]. Analysis [4575, 4097, 7722, 5790, 7912]. Ancestors [9449, 799, 800, 830, 859, 921, 1146, 1299]. Anchor [1343, 5387, 6741, 8151, 9631, 619, 4101]. Anchoring [9710, 1289]. Ancient [1753, 4528, 3914, 8054, 7429, 9708, 1098, 9962, 3318, 4919, 4065, 8495, 10403, 7086, 7116, 10734, 2910, 8210, 8191, 8186, 8825, 8452, 6445, 1681, 4785]. Androids [5116]. Angeles [4626, 435]. Angled [194]. Anglesey [5928]. Aniline [9353]. Animal [2571, 3295, 6966, 7772, 9171, 10918, 11843]. Animals [2415, 3306, 5784, 6354, 6314, 7282, 6223, 7400, 9369, 5238, 5213, 8304, 8237, 8359, 6579, 2328, 10858, 7680, 10090, 192, 2820, 3944, 9055, 9147, 10417, 597, 4493, 5237, 4435]. Animation [8960]. Annapolis [7831, 11417]. Anne [1623]. Anniversary [4484, 5301, 9746, 5310]. Announcement [11782]. Announcers [357]. Annoyances [4843]. Annual [3327, 5104, 5460, 6638, 7605, 7785, 8170, 9487, 9480, 10713, 9977, 11841, 525, 1096, 1529, 2773, 3523, 2233, 1335, 2993, 3058, 7834, 4162]. Annuals [1187]. Annular [11086]. Annunciator [9056]. Anode [7912]. Anomalies [831]. Anomaly [11165]. Answer [8705, 10695]. Ant [2715]. Antarctic [5043, 6457, 7906, 10037, 135, 11936]. Antelopes [7089]. Antennæ [7860]. Anthracite [3134, 10643, 11317, 2263]. Anthropometry [10113]. Anti [2371, 4813, 11705, 7301]. Anti-Alcohol [2371]. Anti-Bubonic [7301]. Anti-Plague [11705]. Anti-Telescoping [4813]. Antimonite [10038]. Antinoë [8171]. Antique [4064, 1213, 7915, 11417]. Antitoxin [7470]. Ants [8662, 10443, 8192]. any [4579]. Apache [2933]. Apartment [6215].

Appalling [5278]. Apparatus [1463, 1487, 2212, 1896, 1823, 2156, 3157, 2801, 2823, 3893, 4874, 4345, 5232, 5117, 5255, 6103, 5599, 6127, 5311, 5021, 5572, 7166, 7186, 6787, 6556, 6605, 6624, 6858, 8234, 8504, 8107, 7637, 8508, 7786, 7700, 8057, 7722, 8060, 9011, 9291, 8624, 8876, 8874, 9629, 9424, 8658, 8872, 9993, 10471, 10505, 10658, 10300, 9990, 10357, 10480, 9705, 11266, 11063, 10886, 10923, 443, 720, 1179, 121, 266, 12, 595, 11066, 413, 11192, 3185, 9589, 9529, 5574, 6465, 1028, 5388, 7209, 7520, 8683, 9674, 9470, 8919, 10114, 9484, 6420, 6892, 6341]. Apparatus [7958, 8090, 8030, 10995, 225, 10279, 7661, 1397, 6443]. Apparent [9429]. Appeals [2593]. Appear [2640]. Apple [7830]. Appliance [9912]. Appliances [5036, 6762, 10691]. Application [2679, 3328, 4098, 8654, 1280, 4684, 2076]. Applications [4369, 5751, 8892]. Applied [1458, 2098, 2097, 550]. Applying [5861]. Appreciation [3521]. Approach [6029]. Approaches [1561, 3404, 11302]. Appropriation [3931, 6754]. Appropriations [6042]. Approved [579]. Approves [11747]. Appulse [11702]. April [7932, 1398, 2626, 3773, 5016, 6279, 7431, 8626, 9759, 10889, 130]. Aquarium [4758, 7659]. Aqueduct [10622, 6373, 10750]. Arabic [6030]. Arable [11114]. Arago [3849]. Ararat [1980]. Arc [2693, 3111, 8547, 9570, 9071, 2867, 3294]. Arcade [6506]. Arch [10041, 3425, 2565]. Archaeological [4873, 962, 8171, 10606, 10631]. Archdeacon [11253, 7166, 9249]. Arched [9078, 9313]. Archimède [10471]. Archimedean [7194]. Architects [2233]. Architectural [6680, 6809, 9742]. Architecture [4596, 8745]. Arco [3525]. Arcs [6394]. Arctic [8566, 3218, 6917, 8385, 116]. Ardennes [4905]. Area [11371]. Areas [1911]. Arena [11376]. Argon [6375]. Argyll [1532]. Arid [1071, 11919]. Arizona [8333, 2445]. Armada [6560]. Armament [1778, 1974]. Armchair [4977]. Armies [6446]. Armor [2231, 1693, 2116, 1605, 1679, 2199, 2485, 3486, 3484, 3784, 4844, 5950, 8705, 9365, 8642, 9231, 10594, 11482, 5947, 6916, 6856]. Armor-Piercing [8642, 9231, 11482]. Armor-Plate [2116, 9365]. Armorclad [1394]. Armorclads [6766]. Armored [1572, 1726, 2266, 1439, 3487, 2537, 3488, 4490, 4738, 4786, 3986, 6803, 7155, 6625, 7042, 8593, 8590, 9824, 9812, 11002, 11033, 2339, 4582]. Armored-Concrete [9824]. Armories [3792]. Armory [5260]. Arms [5957, 6446]. Armstrong [2257, 2776, 4867]. Army [2616, 10503, 1866, 2036, 2711, 3160, 2793, 4730, 3745, 5560, 5804, 6396, 6485, 6216, 6462, 8519, 9461, 9004, 8658, 9160, 9696, 11345, 1116, 530, 730, 981, 9165, 10265, 10674, 9211, 342]. Arnold [4597]. Arrangement [468]. Arrested [9415]. Arrester [10048]. Arresting [5621]. Arrow [4383, 4392]. Arrowheads [8575]. Arsenic [5924, 6930]. Arsenious [7276]. Art [4214, 6821, 7340, 8834, 8853, 9007, 10795, 9607, 10088, 8939]. Artemiev [5239]. Artesian [4881, 3383]. Arthur [7258, 7822, 6307, 6476, 7078, 6362, 8126, 7405, 7371, 7385, 7609]. Articles [8077, 11630, 11949, 11715]. Artificial [9253, 2260, 4945, 7625, 8937, 10628, 10437, 11779, 11884, 11599, 11347, 912, 417, 6895, 9727, 3359, 7728, 5882, 8007, 10443, 8409, 9127, 9733, 10091, 8661, 8087, 8336]. Artificially [11698]. Artillery [2417, 2457, 5624, 11085, 997, 143, 722]. Artisan [7283]. Artist [9467]. Artistic [5246, 8774, 8641]. Arts [8022, 7514, 7538, 7558]. Asaph [4313]. Asbestos [11348, 6691, 1442]. Ascended [1980]. Ascension [2443]. Ascensions [8680, 10737]. Ascent [3183, 7579, 860]. Ascents [4458]. Ash [3939, 7989, 9492]. Ash-Pan [9492]. Ash-Sifter [3939]. Asia [1696]. Asiatic [1364]. Asie [3845]. Asking [1250]. Askold [6982]. Asparagus [4624]. Aspect [510]. Aspects [6337]. Aspen [4218]. Asphalt [2576, 5542, 7782]. Asphyxiated [4782]. Assay [3743, 4243, 1598]. Assaying [1653]. Assets [7781]. Assignments [3543]. Assistant [1432, 1495, 2728]. Association [1935, 2026, 1846, 6602, 10713, 9368, 4751, 5413, 5207, 3175, 2044, 5666, 5354, 6043, 11757, 11134]. Assouan [7554, 8282, 8932, 10062, 9855, 11274]. Assumed [1888]. Assured [2883]. Asteroid [7088, 9646]. Astor [3364, 4561, 10797, 346, 864]. Astrolabe [8475]. Astrometry [11597]. Astronomical [7320, 7627, 7401, 10405, 7322]. Astronomy [9026]. Astrophysical [7771]. Aswan [11390]. Asyut [5545]. Atbara [3740]. Athletic [11376]. Atkinson [8352]. Atlanta [957]. Atlantic [2167, 2002, 2486, 2170, 2657, 3177, 3462, 2793, 4406, 4809, 5365, 6028, 5761, 5168, 7535, 7474, 9344, 9670, 613]. Atmosphere [9253, 716]. Atmospheric [1668, 1676, 4487, 7953, 8937, 9243, 3960, 10399, 9029, 10114, 8411]. Atom [8828]. Atoms [2792, 8634, 6824]. Attaching [8301]. Attachment [1462, 1677, 2308, 2154, 1728, 2799, 3040, 3332, 4010, 4002, 4814, 4375, 5264, 5263, 7008, 6323, 6590, 6540, 6511, 6695, 6742, 6826, 6322, 8215, 7612, 7802, 7838, 9654, 9015, 9867, 9991, 9911, 10694, 10447, 9757, 10885, 11467, 11643, 11021, 11019, 11331, 11359, 10814, 11572, 902, 11612, 10969, 1101]. Attack [2379, 7670, 11231, 11261, 11217, 11364, 997]. Attained [11848]. Attempt [3171, 6938, 1193]. Attempts [3230]. Attitude [924]. Attorneys [2232]. Attraction [5091]. Attractions [1702]. Attractive [1684]. Attuned [3930]. Audible [6122]. Auditorium [9901]. Audubon [11366]. Augers [400]. August [11430, 10297, 3555, 2393, 4791, 6063, 7242, 8415, 9554, 10676, 11849, 1182]. Aurora [5134, 5923, 593, 594]. Auroras [391]. Austin [1518, 1578]. Australia [4314, 7450]. Australian [7214, 9862]. Austrian [2011, 2010, 10612]. Authentic [3926]. Authorities [7016]. Auto [1873, 3041, 3141, 3860, 6075, 6896, 6532, 7793, 9079, 10471, 9826]. Auto-Boat [6532]. Auto-Bolide [6896]. Auto-Carte [9079]. Auto-Commutator [3041]. Auto-Quadricycle [1873]. Auto-Slip [10471]. Auto-Stage [7793]. Auto-Tricycle [3141]. Autocar [6657, 9967]. Autogenous [9755, 11188, 10261]. Automatic [1983, 2197, 1848, 1582, 1816, 2007, 1465, 2418, 1626, 2713, 2756, 3112, 2603, 3289, 3202, 4374, 3964, 3953, 3918, 4586, 4555, 4466, 4390, 4850, 5188, 6125, 5858, 6161, 6032, 5052, 4927, 5501, 5283, 7232, 6557, 6588, 6442, 6313, 7083, 6694, 6888, 7092, 8234, 8261, 7839, 8110, 8262, 8502, 9626, 9552, 9250, 9651, 8549, 8706, 8855, 9169, 8839, 10545, 10470, 10064, 10716, 11503, 10902, 11242, 11137, 534, 1175, 720, 444, 775, 834, 7627, 1201, 1560, 4806, 10659, 8277, 9674, 9484, 2555, 4783, 5573, 8741]. Automatically [6890, 8107, 9194, 8090]. Automaton [8816]. Automobile [1507, 1555, 2028, 1490, 1791, 1341, 1340, 1447, 1904, 2264, 1752, 2469, 2287, 2063, 3267, 2570, 3683, 3353, 2865, 2931, 3502, 2550, 3550, 3184, 3444, 3016, 3378, 3354, 3409, 3428, 4620, 3796, 3794, 3985, 3800, 4279, 4302, 4097, 4301, 4276, 4292, 4300, 4087, 3810, 4574, 4290, 3797, 4488, 4507, 4257, 4283, 4576, 4808, 3933, 3987, 4833, 5454, 5080, 6060, 5084, 5092, 5737, 6053, 5751, 5795, 5108, 5736, 5050, 5411, 5366, 6095, 5474, 6090, 5045, 5460, 5791, 5087, 6382, 6825, 6910, 6933, 6954, 7255, 7298, 6581, 6221, 6638]. Automobile [6963, 6781, 6571, 6653, 6640, 6639, 6636, 7307, 6888, 6447, 6652, 6648, 7064, 6637, 6717, 6851, 7884, 7834, 8258, 7605, 8039, 7785, 7804, 7863, 7586, 7799, 7564, 8170, 8383, 8384, 7808, 7823, 8783, 8803, 8936, 9540, 8854, 8811, 8787, 8659, 9053, 9006, 9035, 9511, 9564, 8784, 8896, 8894, 9480, 9898, 9924, 10084, 10451, 10477, 10469, 10193, 9782, 10517, 10647, 10495, 10771, 9928, 10266, 10281, 9954, 10713, 10625, 10200, 9929, 9919, 9961, 9977, 10061, 11727, 11562, 11132, 10816, 11107, 11841, 11579, 11603, 11724, 10980, 10963, 10939, 10944, 11934]. Automobile [11372, 11502, 493, 475, 1301, 410, 472, 185, 478, 474, 105, 582, 86, 525, 442, 490, 489, 473, 10670, 11706, 5330, 9588, 5078, 418, 10467, 482, 5083, 2062, 9099, 5648, 5439, 9859, 483, 10241, 6931, 7059, 4413, 4278, 2222, 8832, 10474]. Automobiles [1853, 2131, 2130, 2350, 1411, 3529, 2637, 3092, 3201, 2992, 4144, 4603, 4285, 4288, 4739, 5088, 5072, 5414, 5077, 6660, 6423, 6359, 7002, 6574, 6889, 7764, 7540, 7674, 7792, 7612, 8091, 7800, 7758, 7732, 8476, 7789, 7807, 8149, 8759, 8590, 9324, 9082, 10471, 10178, 9920, 10895, 491, 10465, 3358]. Automobiling [8879]. Automobilism [1430]. Automobilists [8804]. Autoplastic [6577]. Autoplate [5851]. Autos [11153]. Autumn [3356]. Auxetophone [7582]. Auxiliary [3336, 4682, 4703, 5070, 2989]. Available [6436]. Avalanches [7914]. Avenue [2408, 3161, 3791, 3735, 3815, 3737]. Aversion [8654]. Aviation [7445, 10962, 82, 851, 670, 707, 680]. Avoiding [9684]. Award [8482, 11345, 11150, 11635, 427]. Awarded [10493]. Awards [1999, 2515, 2052]. Ax [10885]. Axis [10971]. Axle [3201, 2991, 4301, 6060, 9653, 8809, 9948, 2989]. Axles [3272, 5265, 10353, 9207].

B [9621, 11168]. B. [11519, 218]. B.C. [9801]. Babbittin [1262]. Babbitting [4915]. Baby [3591, 902, 1129]. Babylonian [9801]. Babylonians [8452]. Bach [4872]. Bacillus [11492]. Back [4685, 267]. Backbone [11371]. Background [6556]. Backing [7407]. Backward [2656]. Bacteria [5249, 6822]. Bacterial [7396, 8317]. Bad [6480, 7355]. Bag [5022, 5344, 6404, 811]. Bagging [10174]. Bahama [5325]. Bahamas [3912]. Baikal [2752, 7695]. Bail [438, 460]. Bait [2556]. Baker [4087, 10059].

Bakery [6932]. Balance [1830, 8240, 2539]. Balanced [7560]. Balancing [6319, 9194]. Baldwin [1731, 2727, 7129, 7347]. Bale [11263, 1894]. Ball [1446, 1805, 7501, 8802, 919, 10063]. Ball-Bearing [1446, 919]. Ballons [7650, 635]. Ballons-Sondes [7650]. Balloon [3353, 3317, 3221, 3620, 3163, 4502, 6668, 8252, 9308, 8680, 9264, 10335, 10441, 11555, 10907, 11189, 11110, 724, 57, 6159]. Ballooning [11900]. Balloonists [172]. Balloons [2666, 7235, 8519, 10505, 10907, 11938, 10949, 998, 6440, 10503]. Ballroom [9901]. Baltic [6867, 7211, 7005, 9597, 5977]. Baltimore [6736, 2577, 7318, 9345, 5793, 6224, 11457]. Bananas [7813]. Bancroft [5974]. Band [7836]. Bandelier [6608]. Bands [6574]. Bank [5143]. Banker [8705]. Banks [2073]. Banyan [1854]. Bar [2262, 3916, 5638, 5797, 6541, 9451, 10567, 4167]. Bar-Men-Vohwinkel-El-Berfeld [4167]. Barbara [4319]. Barber [4914]. Barbette [7832]. Barge [8070, 11496, 11294, 1159, 8121]. Barges [4052]. Bark [7539, 3845]. Barnard [6118]. Barney [5719]. Barometer [10027, 4551, 3018, 7933, 3621]. Barometric [8934]. Baron [3658]. Barossa [7450]. Barrage [6333, 33]. Barrages [9815]. Barrel [10784, 870, 7766]. Barreled [5502]. Barrels [7607, 9014, 8751]. Barretter [9284]. Barrington [4312]. Bars [6589]. Barton [6607, 8002]. Bas [9957, 11260]. Bas-Relief [9957]. Bas-Reliefs [11260]. Basaltic [3011]. Base [949]. Baseball [1192]. Bases [9666]. Basin [10592, 2844, 9348, 8690]. Basins [9293]. Basis [1065]. Basket [4323, 3741, 6814, 11308]. Basket-Making [3741]. Basketry [2401]. Baskets [4037]. Bataclan [3745]. Bates [9030]. Bath [6891, 2913]. Baths [9127]. Battenberg [8294]. Batteries [5639, 6246, 9071, 1235, 1104, 1260]. Battery [2510, 1513, 1650, 3386, 2837, 2895, 2803, 3599, 6000, 5015, 6093, 5475, 7190, 7306, 6715, 6391, 6665, 8279, 9545, 10416, 10565, 488, 11688, 686]. Battle [3094, 3054, 8655, 11144, 5392, 11774, 7704]. Battle-Ships [3054]. Battleship [1718, 1966, 2288, 2518, 1393, 1517, 1682, 3479, 3610, 3480, 3405, 2868, 3481, 3622, 2640, 2908, 3314, 3482, 3958, 3935, 4350, 4337, 5033, 5739, 5523, 5183, 5517, 6297, 6294, 6352, 6390, 6374, 7170, 7303, 7648, 8051, 7422, 8355, 7832, 9649, 9281, 8569, 9501, 9585, 10682, 10574, 10580, 10581, 10582, 10583, 9685, 11752, 11896, 11844, 10921, 604, 365, 191, 1012, 10201, 10221, 1601, 7237, 5333, 6552, 11399, 10343]. Battleships [1974, 1407, 2219, 2367, 3478, 3443, 3549, 4880, 4869, 5129, 5639, 5745, 5379, 6068, 7169, 6219, 8313, 8459, 9309, 9586, 8700, 10578, 10579, 10294, 10595, 10596, 10192, 11747, 11173, 10849, 10973, 335, 933, 53, 459, 147, 9837, 9548]. Bavarian [9751, 9703]. Bay [1600, 1884, 1396, 7449]. Bayonet [2822, 7656]. Be [6004, 10433, 3377, 3193, 3043, 6414, 7822, 7749, 8857, 9283, 10243, 10255, 11293, 11797, 11108, 232, 1194, 487, 11014]. Beach [4202, 10771]. Beaches [10869]. Beachy [4641, 3360]. Beads [3545]. Beam [4370, 8058, 761]. Bean [2763]. Beans [5984, 5743]. Bear [1129]. Bearing [1446, 4489, 7767, 9564, 8802, 10136, 452, 4940, 4409, 919]. Bearings [2780, 4301]. Bears [7473]. Beast [3817]. Beating [10401]. Beating-Wing [10401]. Beats [5089]. Beaumont [3019]. Beautifying [9448]. Beauty [3817]. Beaver [6143]. Becquerel [7114]. Bed [8315, 10783, 1058]. Beds [558]. Bee [7283, 6707, 7587, 3338, 4085]. Bee-Culture [4085]. Bee-Keeping [3338]. Beedle [6340]. Beehive [7738]. Beer [2984]. Beer-Cooling [2984]. Bees [935, 6226]. Beet [1339, 4585, 791]. Beet-Topping [4585]. Beetle [1659]. before [10515]. Beginnings [5357]. Behavior [289]. Behind [6728]. Behr [2701, 9389]. Behring [6089, 8651]. Beinn [10382]. Belgian [1918, 6499]. Belgium [11194]. Belin [10627, 413]. Bell [4422, 5233, 8704, 10887, 1280, 9189, 62, 5164]. Bell-Buoy [5233]. Belleisle [2379, 2367, 4893, 4846, 5667]. Belleville [3245, 3420, 5899]. Bells [9800]. Below [8398, 5251, 3818]. Belt [2577, 5438, 8398, 7525, 7602]. Belt-Stud [7525]. Belting [665]. Ben [2494]. Benbow [6997]. Bench [3062, 10721, 11279, 201, 563]. Bend [4467]. Bending [5524, 6274, 10356]. Beneath [11251]. Benjamin [10059]. Bennett [1466, 5576, 6053, 7023, 8383]. Bennington [8480, 8446]. Berfeld [4167]. Berger [7257]. Berkeley [7522]. Berley [1804]. Berlin [3378, 3575, 2051, 3398, 3039, 4690, 5910, 5110, 11632, 11106, 851, 827, 776, 8305, 6006, 11016, 9568, 8004]. Berliner [3377, 2567, 5671, 5456]. Bermuda [10083, 10042, 11193, 4831, 3816, 4580]. Berry [3114]. Berthelot [9769]. Berthier [6555]. Berthoud [4081]. Bertillon [10113]. Bessemer [5673]. Best [3176, 10192]. Bethlehem [1605, 1679, 2199, 2456]. Beton [10940]. Better [5495, 10707, 10295, 11075]. Betts [8275]. Between [3395, 4964, 9345, 5392, 1930, 4728, 5620, 8313, 8719, 5259, 1939]. Bevel [9933]. Bhreagh [10382]. Bible [9395]. Bicycle [1873, 1731, 1914, 1852, 2426, 2262, 2100, 1774, 2993, 3315, 2824, 2958, 4293, 4763, 3822, 4294, 4090, 6589, 8261, 8039, 7801, 7354, 9249, 8924, 10197, 10478, 641]. Bicycle-Driven [4293]. Bicycle-Lamp [1731]. Bicycles [5582, 5076, 6540, 6738, 7802, 8788, 567, 7908]. Bids [2485, 8491, 9879]. Big [1336, 2211, 3632, 3326, 5409, 6928, 10192, 4319, 5913, 4462, 4467]. Biggest [2930, 933]. Bilge [5117]. Bill [2275, 1841, 1840, 1548, 3004, 2723, 2772, 4048, 3931, 3786, 4118, 4188, 8635, 8585, 9090, 10847]. Billed [10950]. Billiard [1805, 10653, 1142]. Billiards [8436]. Billion [1363]. Bills [1622]. Bin [9909, 1003]. Binders [9867]. Binocular [8150]. Bins [440]. Bioscope [7167]. Biplane [1275]. Bird [6323, 6415, 7959, 1221, 4940]. Bird’s [2289, 1796, 11042, 9471, 925, 11848, 9420, 5718, 1519, 8505]. Bird’s-Eye [2289]. Birkeland [9253]. Birth [1075]. Birthday [6498]. Bit [9783, 8643]. Bite [10153]. Bits [8324]. Bitumens [927]. Bituminous [3286]. Black [10795, 10125, 1445]. Black-Wall [1445]. Blackberry [5190]. Blackboard [2694]. Blackmail [3540]. Blacksmith [11547]. Blackwell [5650, 7451, 10738, 10867, 10906, 11933, 11103]. Blanc [4316, 8471]. Blas [4254]. Blast [1600, 7408, 11448, 7909, 3577]. Blasting [2749, 4349, 9425, 11319]. Blasts [1122]. Blaze [6342]. Blending [1197]. Blériot [9899, 11829, 1210, 1254]. Blind [10179, 11087, 11658, 754]. Blindness [4572, 6287, 11899]. Blinds [1304]. Block [4688, 3768, 4927, 5308, 8769, 9473, 9577, 10696, 10688, 11316, 534, 989]. Blocking [6307, 9082]. Blocks [10970, 8456]. Blondlot [6140, 7153]. Blood [5030, 8471, 10404, 10259]. Bloom [2328, 5118]. Bloomington [7827]. Bloomless [7830]. Blow [7091, 11903]. Blow-Off [7091]. Blowing [5694]. Blown [3574, 4192, 9687]. Blowpipe [3572, 6624, 9138, 9755, 68]. Blue [5931, 6720, 8587, 327, 7421, 6027, 1031, 7773, 6139]. Blue-Print [5931, 327, 6027]. Bluefish [8967]. Blunder [10703, 11674]. Board [1479, 3384, 6042, 10494, 1384, 2704, 6216]. Boat [9507, 2021, 2396, 1493, 1985, 1769, 2244, 2402, 3203, 2912, 3492, 2734, 4212, 4636, 4074, 4392, 4809, 4233, 4525, 5901, 5260, 6011, 6100, 4965, 4976, 5794, 6201, 6382, 6390, 6885, 6221, 6532, 6195, 6812, 6225, 6216, 7064, 6580, 6444, 8463, 7679, 7490, 8050, 8081, 7942, 7940, 7720, 8231, 9203, 8685, 9265, 9123, 9118, 10439, 10414, 10243, 10236, 10626, 10576, 10107, 10066, 10237, 10083, 10069, 10589, 10042, 10234, 11179, 11924, 11906, 11193, 10855, 11761, 11040, 11895, 1054, 170, 583, 1276, 10670, 1274, 4625, 9124, 10778, 10263, 661, 10025, 3867]. Boat-Propelling [11761]. Boating [11374]. Boats [3142, 3490, 3491, 4640, 6464, 8020, 8469, 9504, 9119, 9116, 9115, 8913, 10239, 10240, 10238, 10379, 11606, 11413, 262, 8109, 5259, 9114, 8901]. Bob [11666]. Bodies [2508, 2792, 2886, 10158, 8233]. Bodleian [4484]. Body [6930, 6756, 7565, 9076, 8934, 8894, 10364]. Boer [1765]. Bogatyr [6507]. Bogue [3687]. Bogus [2232]. Boiler [2094, 2713, 2663, 3056, 3396, 3061, 4553, 4019, 4857, 3790, 4336, 4410, 3824, 4256, 4321, 4455, 3844, 5279, 7295, 7865, 7604, 7537, 8920, 8554, 8941, 9065, 265, 29, 918, 811, 7825, 6714, 324, 518, 2783, 3420, 5899, 6941]. Boiler-Feeder [2713]. Boiler-Tube [265]. Boilers [809, 1389, 2512, 3245, 2633, 3635, 3836, 3799, 5292, 7205, 10170, 564]. Bolas [5606]. Bolas-Throwing [5606]. Bolide [6896]. Bolometer [2539]. Bolometric [9582]. Bolt [3231, 7008, 948]. Bolt-Heading [7008]. Bolton [1908]. Bombs [8494]. Bond [9153]. Bones [11888, 7811]. Bonner [1584]. Book [1659, 3204, 2558, 4589, 4766, 4108, 5861, 6022, 5857, 6957, 8752, 11642, 780, 11823, 6911]. Book-Covering [5861]. Book-Typewriter [4108]. Bookbinding [5862]. Bookcase [6561]. Books [1311, 1326, 2067, 2088, 2113, 2140, 2164, 2185, 2227, 2250, 2272, 2293, 2319, 2343, 2361, 2384, 2429, 2450, 2480, 2503, 1354, 1378, 1401, 1473, 1500, 1524, 1567, 1589, 1617, 1639, 1665, 1688, 1713, 1739, 1756, 1782, 1809, 1835, 1856, 1878, 1899, 1946, 1996, 2039, 1863, 2561, 2586, 3415, 3453, 3498, 3518, 3536, 3557, 3581, 3604, 3628, 3649, 2607, 2696, 2745, 2769, 2788, 2855, 2902, 2925, 2949, 2975, 3000, 3026, 3050, 3074, 3102, 3125, 3146, 3167, 3236, 3256, 3279, 3298, 3346, 3370, 3731, 3751]. Books [4675, 4770, 4795, 4820, 4840, 4862, 4888, 3778, 3876, 3899, 3993, 4015, 4042, 4071, 4095, 4181, 4207, 4226, 4250, 4271, 4332, 4359, 4402, 4429, 4449, 4498, 4521, 4545, 4595, 4633, 4657, 5861, 4988, 5011, 5874, 5896, 5917, 5942, 5965, 5993, 6039, 6065, 6086, 6111, 6132, 5060, 5124, 5153, 5173, 5200, 5220, 5298, 5319, 5401, 5424, 5450, 5466, 5491, 5536, 5588, 5613, 5660, 5687, 5710, 5729, 5756, 5779, 5820, 5844, 4949, 4981, 5244, 5271, 5394, 6173, 6192, 7052, 7075, 7125, 7149, 7177, 7201, 7246]. Books [7289, 7313, 6232, 6254, 6284, 6303, 6349, 6387, 6452, 6472, 6492, 6524, 6596, 6632, 6702, 6725, 6771, 6794, 6818, 6840, 6864, 6902, 6924, 6952, 6990, 7032, 7368, 7390, 8267, 8330, 8346, 8369, 8419, 8443, 8514, 7413, 7435, 7464, 7485, 7573, 7593, 7618, 7642, 7689, 7714, 7743, 7779, 7819, 7845, 7872, 7926, 7949, 7972, 8013, 8036, 8067, 8156, 8199, 8222, 8009, 8537, 8560, 9498, 9518, 9536, 9612, 9661, 8579, 8606, 8629, 8713, 8730, 8756, 8778, 8844, 8864, 8884, 8929, 8974, 8999, 9021, 9061, 9086]. Books [9199, 9278, 9300, 9318, 9362, 9385, 9410, 9437, 8636, 9716, 9739, 10618, 10700, 10726, 10744, 10764, 10786, 9809, 9830, 9850, 9890, 9971, 9997, 10015, 10032, 10053, 10074, 10099, 10118, 10144, 10166, 10185, 10289, 10309, 10347, 10370, 10410, 10487, 10512, 10571, 10600, 10804, 10841, 11672, 11695, 11744, 11766, 11792, 11853, 11872, 11892, 11910, 11929, 10864, 10892, 10954, 11007, 11026, 11050, 11071, 11099, 11124, 11207, 11228, 11247, 11290, 11338, 11382, 11422, 11444, 11488, 11508, 11531, 11553, 11594, 11622, 11648, 4, 42, 1036, 1061, 1109, 1133]. Books [1156, 1207, 1265, 1286, 75, 101, 208, 247, 272, 300, 377, 404, 449, 545, 576, 625, 647, 667, 695, 786, 817, 840, 877, 908, 929, 953, 979, 1008, 204, 17, 11155, 1170]. Boom [97]. Boomerang [8873]. Boonton [2574]. Boot [9049]. Boots [11369]. Borax [2099, 1629, 7207]. Bordeaux [2931, 9746]. Bore [2105, 9322]. Borealis [5923, 593]. Borelly [5568, 6071, 6145, 6144]. Boring [5879, 7468, 8186, 9821, 10567, 439, 281, 256, 1027, 4376, 8640]. Borings [6260]. Boston [1919, 4488, 4507, 7774, 8679, 10297, 2323, 1942, 1705, 6581, 7542, 8532, 7961, 218, 555, 2601, 3012, 6959]. Botanical [2661, 2051]. Botany [9580]. Bottle [1440, 4836, 6629, 8285, 9628, 9430, 10203, 11868, 537, 517]. Bottle-Washing [4836]. Bottles [10020, 1178]. Bottoms [1869, 11607]. Botts [6835]. Boudin [6201]. Bound [1102]. Bow [1295, 10529]. Bowlders [10247, 5630]. Bowles [1920]. Bowling [6125]. Box [3613, 2936, 4057, 4621, 5193, 5565, 6587, 6276, 6558, 8363, 8623, 9675, 10048, 9866, 11524]. Box-Cap [3613]. Boxes [9213, 11467, 11334]. Boy [8228, 10338]. Boys [7930, 7939]. Brace [6467, 8324, 9783, 397]. Bracing [5921]. Bracket [11201, 11611, 11834, 1228]. Brackets [8300]. Bradsky [4606, 4506]. Brain [6731]. Brains [9643]. Brake [2285, 2952, 2644, 2750, 3569, 3822, 5653, 7232, 6734, 7732, 9376, 10545, 9865, 11666, 11762, 486, 10964]. Brakes [2197, 4693, 9939, 10800, 496]. Branch [9818, 10496]. Branger [10239]. Branly [4622, 10658]. Brass [5331, 439, 71]. Brasses [2780]. Braun [9031]. Braunschweig [11399]. Brazilian [1394, 2290, 11752]. Bread [3700, 4027, 7672, 8666, 10898, 10530]. Bread-Making [3700, 4027]. Breads [536]. Break [801, 672, 896, 993, 11353]. Breaker [3285, 6395, 8111]. Breakfast [9163, 9393, 9285]. Breaking [1408, 3038, 3892, 6662, 8110, 8128, 11372, 898, 964, 8090, 10343]. Breaks [1979, 1767]. Breakwater [5047]. Breathing [10176]. Bredsdorff [10523]. Breech [3485, 6163, 8858]. Breech-Loading [8858]. Breeches [10809]. Breed [4779]. Breeding [1481, 8027]. Brennan [10023]. Breton [4832]. Brewing [6317, 6559]. Briarcliff [11013]. Brick [2897, 5719, 4966, 185, 9672, 4464, 642, 8106]. Brick-Carrying [2897]. Brick-making [4966]. Bricks [4639, 10829]. Bridge [1607, 1390, 1550, 1705, 2106, 1890, 3682, 2723, 3302, 3463, 3740, 4059, 3806, 4612, 4080, 4623, 3834, 4263, 4570, 4028, 5246, 4999, 4929, 5952, 5670, 5224, 5409, 6479, 6617, 8531, 8168, 8611, 10267, 10149, 10041, 10553, 10528, 10312, 10292, 10318, 10844, 11344, 11797, 11933, 11475, 10988, 1019, 250, 1344, 6940, 11940, 10558, 11519, 1800, 2145, 1606, 1561, 3404, 2688, 3106, 3284, 3600, 3585, 2893, 3044, 3608, 3181, 3161, 4635, 4232, 4214, 4866, 4784, 3781, 4197, 3755, 6029, 5650, 4972, 6045, 5982, 5519, 5376, 5878, 7189, 6484, 6178]. Bridge [6626, 6798, 7210, 6599, 8491, 8073, 8410, 8003, 7555, 8185, 7451, 9043, 9227, 8744, 9159, 10147, 10738, 10331, 10415, 10400, 11864, 11282, 10936, 11211, 10947, 10867, 10906, 11732, 11479, 11302, 11733, 11447, 10958, 11103, 11731, 363, 453, 7402, 4508, 3643, 11457]. Bridges [1369, 2014, 3522, 3326, 3593, 4314, 5664, 6365, 8160, 7374, 7876, 10332, 11730, 11146, 3425, 3228, 10798, 10026, 706, 7998, 171]. Bridging [9459, 9700]. Bridle [9986]. Bridles [2056]. Brief [3812]. Brierley [8874]. Bright [10707]. Brighton [10771]. Brilliant [9304, 11213]. Briquettes [2286]. Britain [2232, 2147, 3833, 4411, 5558, 6732, 7231, 8031, 8019, 11170, 3954]. British [1952, 3376, 2026, 3244, 3153, 3586, 2702, 2633, 3932, 4727, 4481, 4453, 3738, 4459, 4084, 4490, 4577, 3908, 4703, 5620, 5033, 5763, 5250, 6008, 5257, 5043, 5210, 5804, 5667, 6238, 6533, 7158, 6651, 7419, 7422, 7910, 7729, 8082, 8255, 9649, 9641, 8916, 8569, 8677, 10790, 10935, 11179, 11415, 10990, 10832, 11426, 10938, 11923, 22, 530, 730, 844, 982, 4000, 1765, 10968, 11826, 2689, 6465, 7234, 5139, 6417, 11133, 11774, 11585, 5020, 10343, 10378, 8407]. Broaching [2780]. Broadside [3840]. Broadway [2085, 4613, 6114, 8969]. Brockton [8679]. Broken [3096, 7607, 9834, 9982, 9974, 408, 495, 880, 734, 868, 1203]. Bromelia [8986]. Bromide [1386]. Brontosaur [7772]. Bronze [4064, 6939, 7394, 7504, 11417]. Bronzes [7480]. Brooklyn [1562, 1395, 4708, 1749, 2070, 1492, 1921, 2655, 3284, 3600, 3585, 3302, 2883, 3608, 4232, 4866, 3781, 3755, 5544, 5224, 4994, 6599, 7372, 9389, 10149, 9722, 10331, 10415, 11344, 11302, 10986, 10958, 7402, 2805]. Brooks [2520, 4082]. Brooms [11949]. Brothers [11945, 306]. Brown [3349, 6661]. Brunel [5997, 10079]. Brunt [2284]. Brush [4836, 7795, 11200, 1178, 373, 11111]. Brushes [11948]. Brussels [887, 8597]. Bubbles [242]. Bubonic [7301]. Buck [1212]. Buckboard [6640]. Bucket [3202, 10247, 11435, 11942, 8278]. Buckets [5546]. Buckeye [6829]. Buckle [6584]. Buckled [572]. Budapest [7998, 8003]. Buddhist [6535]. Buds [1895]. Buffalo [3220, 5047, 671, 3637]. Buffer [8476, 10225]. Bug [11400, 11830]. Buggy [478, 86]. Buggy-Type [86]. Build [3084, 8248, 9696, 10639, 10490, 10078, 291]. Builder [2614, 3507, 6143, 5739, 6156]. Builders [380]. Building [1922, 1912, 2256, 2102, 1563, 1890, 3404, 2781, 4055, 4727, 4878, 6160, 5952, 5404, 5566, 6808, 8070, 8517, 7765, 8772, 9167, 10696, 10686, 10154, 10542, 10223, 11273, 11413, 10829, 11480, 4706, 596, 2076, 8456, 3360, 7773, 8503, 10008, 5996, 8772, 10301]. Building-Bricks [10829]. Buildings [1349, 4990, 6468, 7876, 9723, 11845, 11734, 8836]. Built [2470, 3796, 5113, 6807, 8051, 9283, 10243, 10108, 312, 10382, 9565, 11833, 10304, 11807, 4753]. Bulbs [6105]. Bulk [2566, 1072]. Bulkhead [8049]. Bulkheads [3953, 6374]. Bull [4997]. Bullet [9908]. Bunk [2581]. Buoy [5233, 10523]. Buoyancy [8020]. Buoys [3089, 10809, 10337]. Burbank [10948, 8027]. Burchell [9747]. Bureau [2326, 4074, 4634, 1115, 8299, 11301, 10999, 3327, 3809]. Bureaus [630]. Burglar [7135, 7184, 8705, 1161]. Burglar-Proof [1161]. Burglary [9007]. Burial [3291, 9396]. Burke [8375, 8021]. Burner [1438, 2647, 3770, 4537, 4165, 5345, 5504, 7497]. Burners [2050, 4509, 5264]. Burning [2198, 4883, 3790, 8767, 11176, 8571, 3639, 4753]. Burns [3975]. Burrowers [2339]. Burry [2597, 10875]. Bursting [387]. Burton [2987]. Bus [11932, 11372]. Bushel [10024]. Business [1312, 1327, 1924, 1947, 1970, 1994, 2017, 2040, 2065, 2089, 2111, 2138, 2162, 2186, 2205, 2225, 2251, 2270, 2294, 2317, 2344, 2362, 2385, 2404, 2430, 2451, 2478, 2504, 1355, 1376, 1402, 1425, 1450, 1474, 1498, 1522, 1544, 1568, 1590, 1615, 1637, 1663, 1689, 1714, 1737, 1757, 1783, 1810, 1836, 1857, 1879, 1900, 1987, 1599, 2559, 2582, 3276, 3296, 3319, 3341, 3368, 3390, 3412, 3430, 3448, 3470, 3494, 3513, 3533, 3554, 3578, 3601, 3623, 3647, 3670, 3690, 2604, 2625, 2649, 2670, 2674, 2717, 2740, 2764]. Business [2809, 2831, 2851, 2874, 2899, 2919, 2944, 3071, 3098, 3122, 3144, 3164, 3186, 3210, 3233, 3254, 2526, 2767, 2947, 2973, 2998, 3027, 3086, 3705, 3710, 4325, 4354, 4377, 4398, 4423, 4445, 4473, 4494, 4514, 4539, 4563, 4588, 4607, 4627, 4650, 4672, 4694, 4718, 4740, 4765, 4790, 4815, 4837, 4858, 4884, 4907, 3746, 3772, 3776, 3826, 3850, 3872, 3895, 3920, 3945, 3966, 3989, 4012, 4038, 4066, 4091, 4110, 4132, 4152, 4176, 4203, 4224, 4247, 4269, 4303, 4922, 4924, 5584, 5611, 5632, 5655, 5683, 5707]. Business [5725, 5752, 5775, 5799, 5816, 5839, 5871, 5891, 5914, 5940, 5961, 5989, 6015, 6035, 6062, 6082, 6107, 6129, 6149, 6168, 4980, 5002, 5007, 5028, 5059, 5119, 5147, 5169, 5195, 5198, 5241, 5268, 5293, 5315, 5346, 5350, 5371, 5418, 5445, 5462, 5486, 5509, 5530, 5553, 6187, 6209, 6788, 6792, 6836, 6861, 6880, 6898, 6920, 6947, 6967, 6986, 7010, 7027, 7046, 7070, 7095, 7119, 7146, 7172, 7196, 7222, 7241, 7264, 7285, 7309, 7330, 6212, 7348, 6230, 6278, 6299, 6325, 6343, 6347, 6383, 6407, 6426]. Business [6429, 6469, 6487, 6519, 6544, 6562, 6591, 6612, 6631, 6672, 6698, 6721, 6746, 6767, 8195, 7364, 7386, 7992, 8008, 8032, 8063, 8092, 8112, 8131, 8152, 8177, 8218, 8242, 8263, 8288, 8306, 8326, 8341, 8365, 8392, 8414, 8439, 8464, 8483, 8510, 8533, 7409, 7430, 7460, 7481, 7505, 7528, 7548, 7569, 7589, 7614, 7638, 7662, 7686, 7710, 7739, 7757, 7775, 7815, 7841, 7867, 7894, 7921, 7944, 7967, 8555, 8558, 9154, 9174, 9195, 9218, 9234, 9254, 9273, 9295, 9315, 9339, 9358, 9381, 9406, 9433, 9453]. Business [9472, 9494, 9514, 9532, 9553, 9572, 9590, 9608, 9633, 9656, 9677, 8601, 8604, 8645, 8667, 8692, 8709, 8726, 8753, 8775, 8817, 8840, 8859, 8880, 8906, 8926, 8949, 8970, 8994, 9018, 9036, 9057, 9083, 9103, 9131, 10181, 9357]. Butt [8905]. Butte [4388]. Butter [7019]. Butterflies [6985, 10336]. Butterfly [5286, 5959]. Buttes [3984]. Button [10508]. Buy [11898].

C [2935, 6048, 9252]. Cab [1657, 6666, 8099, 10688, 10767]. Cabin [6630]. Cabinet [5341, 7305, 8666, 10226, 10884]. Cable [2078, 1951, 3337, 2838, 2572, 2821, 4895, 5340, 6167, 6028, 5459, 5211, 8175, 9066, 9096, 9709, 11211, 11315, 1725, 8963, 3424, 7278, 7681, 1861]. Cable-Cutting [9709]. Cable-Grip [2078]. Cables [1800, 3106, 4784, 4219, 6045, 5638, 4929, 5519, 6515, 8073, 7444]. Caboose [5384]. Cabs [5695, 10466]. Cactus [2051, 10942]. Cadillac [5079, 6659, 8796]. Cadmium [7426]. Cælostat [7771]. Cage [5340]. Cages [6323]. Cahokia [6250]. Caine [4780]. Cairo [3441, 9397, 9391]. Caisson [9483]. Caissons [4028, 514, 10304]. Cal [3225, 4626, 7522, 4142]. Calabash [687]. Calais [5831]. Calamity [2190]. Calcium [2097, 2303, 6530]. Calculator [2398]. Calendar [11879, 1243, 7627]. Caliber [3270, 9241, 10716, 10192, 872, 9322]. California [1918, 1336, 2211, 1602, 2266, 1940, 2530, 3570, 3565, 5812, 5984, 5360, 7057, 8295, 10080, 11349, 7751, 3333, 5529, 8460, 7115, 6505, 8084, 11397, 5444, 2201, 2624, 2552, 4318, 4199, 7321, 7168, 5235]. Caliper [1511]. Caliper-Scale [1511]. Calipers [6321, 11359]. Calking [8684]. Call [4437, 6413, 6468]. Called [5102, 137, 232, 10398]. Callitypy [5035, 5677]. Calumet [3508]. Cambodia [2123]. Cameos [6573]. Camera [1367, 2425, 1540, 3113, 9493, 9735, 283, 7088, 9646, 6911, 7701]. Cameras [7475]. Camp [11020]. Campagna [1720]. Campaign [2493]. Campanile [4434, 9508, 9773, 11015]. Campers [10383]. Camphor [5882]. Camping [6693]. Can [3615, 6414, 7822, 7749, 8479, 7840, 9432, 10243, 11356, 371, 127, 1232, 3043]. Can-Filling [371]. Can-Opener [3615]. Canada [2818, 2751, 9645, 382, 3045, 1442, 10283]. Canadian [7833, 2858, 5571, 8424, 9366, 4667, 9288]. Canaigre [3196]. Canal [1406, 1840, 1642, 1548, 2298, 3890, 4137, 4138, 3902, 4118, 5785, 5012, 5562, 5469, 6308, 6797, 8203, 8138, 7416, 9030, 9024, 9110, 8957, 8887, 8573, 628, 334, 671, 112, 9676, 5721, 560, 10761, 5310, 1865, 2277, 1814, 2484, 1763, 1429, 1830, 2331, 1843, 2748, 2678, 2774, 2632, 2636, 4148, 4254, 4546, 4136, 3886, 4600, 4309, 4045, 4529, 3787, 4052, 4747, 4452, 4569, 4787, 3995, 5355, 6155, 5595, 5571, 5324, 6094, 5406, 6421, 7152, 6479, 7228, 7669, 8070, 7677, 8248, 8101, 8016, 8249, 8226, 8979, 8565, 9283, 8822]. Canal [8984, 9149, 9522, 8679, 9460, 9271, 8572, 9696, 9917, 10190, 10433, 9721, 10373, 11409, 11077, 11053, 11496, 11250, 11152, 11294, 11755, 11846, 11272, 452, 651, 579, 1159, 1011, 4756, 1630, 1888, 4100, 1148, 4198, 9547, 11237]. Canals [2751, 8461, 9596, 8827, 3045, 28, 5385, 8702, 11493, 8454]. Canard [7181]. Canceling [9191]. Cancer [8891]. Candle [7093, 8942]. Candles [11713]. Candlestick [2995]. Cane [2641, 5575, 5386, 9629]. Canet [2373, 4904, 2257, 4345]. Canned [6787]. Cannibal [6124, 11810]. Cannon [2707, 5502, 163, 11417]. Cannot [487]. Canoes [8213]. Cañon [6578]. Canovetti [10244]. Cantilever [1550, 2106, 3682, 11733, 708, 1851]. Cantilevers [10867]. Canvases [7492]. Cap [3613, 4618, 5746, 11093, 11667].

Capacity [1382, 2692, 4659, 11145, 11112]. Cape [4832, 1552, 1625, 1610, 1368, 3441, 6537, 9391, 9397, 9917, 651, 1819, 4268]. Capillary [7735]. Capitaine [7936, 8544]. Capital [9345]. Capitol [4371, 7398]. Capitulation [7405]. Capped [1594]. Caprice [5444]. Capsized [9483]. Capsizing [3845]. Capsuling [3121]. Capsuling-Machines [3121]. Capt [730]. Capt. [6486, 9820]. Captive [8260, 10503]. Capture [9137, 7609]. Captured [9569]. Car [2238, 1655, 1627, 2891, 2644, 3313, 2915, 2737, 2991, 3138, 3964, 3719, 4295, 3986, 4158, 5141, 5382, 5713, 6106, 5265, 5074, 5692, 5971, 5025, 6240, 6642, 6186, 6907, 6664, 7341, 6918, 6643, 6645, 7683, 8318, 7798, 7803, 7918, 7790, 8335, 8449, 9653, 9627, 8796, 8798, 8793, 9312, 8801, 8799, 8794, 9240, 8599, 8897, 9566, 10353, 10325, 9797, 10458, 9922, 9925, 9959, 9964, 9949, 10363, 10461, 9966, 9947, 11021, 11483, 11495, 11459, 11513, 11013, 11243, 11535, 476, 477, 10425, 2968, 5384, 10739, 10220, 5439, 9623, 8503]. Car-Fender [2891]. Car-Platform [2238]. Carabao [1866]. Carable [7836]. Caramels [3207]. Carbide [2097, 6530]. Carbides [11254]. Carbon [2596, 6776, 9244, 11598, 485, 10903]. Carbonated [5598]. Carbonic [3762]. Carborundum [1706, 9732]. Carboy [7060]. Carbureter [4368, 5501, 7092, 7966, 8706, 9546, 10479, 494]. Cards [11243, 9484]. Carl [9355]. Carlo [6382]. Carmania [8358]. Carnegie [5520, 6684, 9323, 11597, 615, 9819]. Carnival [8896, 10963]. Carnivorous [1097, 10614]. Carolina [9817, 9309, 9586, 10596, 11135, 10808]. Caronia [7383]. Carpet [5680, 9780, 11617]. Carquinez [3337]. Carrara [7004, 10527]. Carriage [2496, 1802, 1806, 3418, 3388, 3575, 4296, 4286, 4282, 8057]. Carriage-Washing [8057]. Carriages [2302, 2847, 4280, 8805, 902]. Carried [1072]. Carrier [5252, 6196, 9014, 8961, 7]. Carriers [5249]. Carries [9650]. Carroll [6685]. Carry [55]. Carrying [2961, 2897, 5929, 6079, 490, 468, 8124]. Cars [2181, 2078, 2285, 2369, 2581, 2942, 3273, 4018, 4193, 4774, 3906, 5576, 5900, 5705, 5578, 6413, 6647, 6650, 6689, 6644, 8525, 8482, 7658, 7929, 7976, 7791, 8401, 9664, 9560, 8812, 8800, 8797, 9205, 8912, 8832, 9927, 10496, 9935, 10021, 9697, 10603, 10159, 9939, 9730, 10826, 11510, 11867, 11474, 11625, 157, 266, 144, 994, 10403, 6623, 4060]. Carson [8025]. Cart [1959]. Carte [9079]. Carter [10458]. Carthage [1829, 7250, 8054]. Cartridge [8277]. Cartridge-Counter [8277]. Carts [1101]. Carving [943]. Carvings [5001]. Cascades [7212]. Case [2467, 3502, 4967, 5528, 5937, 6480, 7355, 9142, 15, 490, 11395, 743]. Casements [9491]. Cases [11453]. Casings [4536]. Cast [3892, 6310, 9800]. Cast-Iron [3892, 6310]. Casting [1893, 5331, 5981, 7069, 6939, 6484, 10928, 7504, 9189, 219]. Castings [572, 9684]. Casts [2690, 11195]. Casualties [7438]. Casualty [10188]. Catacombs [6354]. Catalina [5185]. Catalogue [5393]. Catalpa [6458]. Catalytic [4145]. Catamaran [4293]. Catastrophe [4054, 8722]. Catch [5192, 9293]. Catcher [6512, 6404, 8682]. Catechism [9523]. Caterpillar [11084]. Cathedral [7039, 3655, 11233]. Catherine [4468]. Cathode [1964, 6601, 2172]. Cats [1681]. Catskill [10622, 10111, 10293, 10750, 11029, 11729, 10831, 10827, 980]. Cattle [4396, 7061, 9825, 6147]. Cause [4434, 5641, 4968, 8469, 10732, 9982, 10844, 880, 8076]. Caused [8282, 8690]. Causes [3199, 11767, 3616]. Caution [7316]. Caux [117]. Cavalry [3142, 7446]. Cave [3458, 3791, 3815, 5506, 11088, 5784, 6159, 340]. Cave-in [3791, 3815]. Caves [7723, 6355]. Caviar [2465]. Cavité [8550]. Cecilie [10769]. Cecille [2155]. Cedar [11781]. Ceiling [2580]. Ceiling-Switch [2580]. Celebration [10375, 820, 213]. Celestial [2569]. Cell [9012, 62, 871, 70, 940]. Cellar [869]. ’Cello [832]. Cells [5700, 10382, 9733]. Cellulith [1509]. Celluloid [3545]. Cellulose [2462, 2725]. Celtic [2731, 3642]. Cement [1226, 7856, 8774, 9666, 8762, 8767, 8890, 10494, 11475, 292]. Censor [7726]. Census [3866]. Cent [2055]. Centenary [11515]. Centennial [10643]. Center [157, 144, 1153]. Centering [9823, 4599]. Centers [6731, 11788]. Central [2255, 1585, 3352, 3609, 4139, 4170, 4130, 4191, 5412, 7254, 7885, 8028, 7769, 9561, 9506, 9272, 10474, 10626, 10276, 11739, 11177, 6006, 1671, 5408, 6236, 7041, 8492, 9877, 9699, 10132, 10264, 7192]. Centrifugal [2168, 2012, 2230, 5188, 5938, 5267, 7722, 8497, 8919]. Centuries [9005]. Century [2390, 1852, 1573, 2146, 4685, 9773, 2348, 5574, 10089, 5118, 2193, 2121, 2979, 11875]. Cereal [7021, 9163, 9285, 790]. Cereals [11296]. Ceremonial [5774]. Certain [4616, 5715]. Ceylon [221, 7066]. Chain [1775, 1366, 3571, 5387, 5955, 6045, 7221, 8149, 9066, 8611, 9936]. Chain-Cable [9066]. Chain-Stopper [3571]. Chainless [1774, 3822]. Chainless-Bicycle [3822]. Chains [8073, 4101]. Chair [5334, 5738, 5024, 6513, 7802, 9102, 8987, 9797, 10884, 11280, 864]. Chairs [6075, 9912]. Chaldean [8210]. Chalk [3178]. Chalk-Line [3178]. Challenge [6885, 10392, 10374]. Challenger [1965, 3667, 9161, 11179]. Challenging [6441]. Chamber [4946, 7110]. Chamberlin [11]. Chambers [10121, 2085]. Chameleon [9290]. Champions [3541]. Chances [8655]. Change [8563, 11053]. Changed [10750]. Changer [3822]. Changes [1432, 4991, 6639, 9012, 3616]. Changing [492, 414]. Channel [1705, 8698, 10169, 10554, 1210, 1193, 1298, 1218, 5362, 7137, 6603, 8517, 10748, 1254]. Channels [2784, 5877, 10880]. Chapel [7831]. Character [3060]. Characteristic [2600]. Charcoal [10918]. Charcot [11132, 7906]. Charges [4346]. Charging [9071, 1260]. Chariot [5908]. Charlemagne [3405]. Charles [4508, 6048, 6498, 10006, 3679, 555]. Charleston [10585]. Charlottenburg [8004]. Charts [7543]. Chassis [9941, 9945, 10463, 10464, 9944, 10459]. Chattanooga [10586]. Cheap [1416, 1413, 7534, 8653, 8851, 11041, 736, 4242]. Cheaply [65]. Cheaply-Constructed [65]. Check [2373, 9654, 9404, 8553, 9965, 11567, 11019]. Cheese [11119]. Chemical [2192, 1694, 2682, 3151, 4374, 5563, 6846, 6973, 11234, 888, 4077, 970, 1350, 9423, 8938, 10650, 684, 198]. Chemicals [11688]. Chemist [10334]. Chemistry [2098, 2489, 1621, 1361, 4340, 6044, 6136, 9245, 1244, 7492, 2097]. Cherry [1040]. Chess [11281]. Chester [1318]. Chestnut [9489]. Chevallier [4124]. Chicago [1080, 2028, 1865, 2394, 2748, 4194, 4685, 4277, 4120, 5355, 5134, 7384, 11259, 1078, 614, 1114, 2939, 2062, 10341, 4127, 11237, 10920]. Chicken [7770]. Chicken-Feeding [7770]. Chief [2278, 715, 2326, 4634, 11077]. Chiefs [4074]. Children [1412, 5750, 9511, 1562]. Chile [4617]. Chilian [5523]. Chilling [1245]. Chime [2246]. Chimney [4618, 5746, 5581, 8380, 2800, 4464, 8298, 5230]. Chimneys [1699, 2781, 3574]. China [2376, 2513, 2411, 1599, 9329, 7118, 4372]. Chinese [2457, 2376, 1371, 2024, 1433, 2894, 3422, 5933, 7480, 2493, 2127, 3060, 4007, 4469, 11136]. Chiselhurst [6355]. Chlorides [2596]. Chloroforming [6420]. Chocks [9654]. Choir [7039]. Chokes [704]. Chord [11933]. Christian [1388]. Christie [8896]. Christmas [11718, 11714, 11712, 11711, 11709, 11713]. Chromatic [10632]. Chrono [8388]. Chrono-Photography [8388]. Chronograph [7836]. Chronometer [6359, 7675]. Chronoscope [2416]. Chronosphere [1224]. Chrowder [11]. Chuck [10180]. Church [5703, 9687, 7627, 9800, 5494]. Cigar [6046, 9375, 9213, 9338, 9150]. Cigar-Vending [9375]. Cigarette [6587]. Cigars [9371]. Cinematograph [10873]. Circuit [8110, 8111, 8090, 4905]. Circuit-Breaker [8111]. Circular [11300]. Circulating [807]. Cire [6939]. Cire-Perdue [6939]. Cirque [3530]. Cisterns [972]. Cities [8191, 4238, 10354]. City [1787, 3386, 3792, 3914, 5750, 5823, 5714, 5666, 5284, 5103, 6373, 6455, 7954, 8399, 8721, 8660, 9765, 9903, 11730, 11104, 11697, 1158, 164, 504, 148, 8452, 9565, 9373, 5676, 2487, 1644, 1530, 1929, 1772, 3205, 3039, 3118, 3677, 4826, 4086, 4700, 5355, 5690, 5167, 5540, 5301, 5277, 6506, 6606, 6753, 8118, 7901, 7583, 8377, 9025, 9344, 8618, 8856, 9742, 10103, 10339, 10351, 851, 7401, 4873, 8836, 9166, 9669]. Civil [1798, 7438, 1243, 6044, 8048]. Civilian [11770]. Civilization [6608]. Clad [11079]. Claims [1362]. Clair [9067]. Clam [11942, 8084]. Clam-Shell [11942, 8084]. Clamp [4353, 7221, 6761, 9173, 1261]. Clams [4376]. Clarence [3955]. Clark [8159, 7563, 1648, 5569]. Clarke [2935]. Clasp [1581]. Class [1966, 2288, 2518, 2338, 1883, 1682, 3479, 3176, 3405, 2840, 3481, 3488, 3482, 10580, 10581, 10582, 10583, 10595, 10596, 10585, 10587, 10586, 238, 5906, 3479]. Class-Room [5906]. Classic [9251, 9267]. Claus [1102]. Claw [3916]. Claw-Bar [3916]. Clay [5132, 9667, 533, 6579]. Cleaner [1707, 5018, 5581, 4947, 9337, 265, 11618, 293]. Cleaning [1533, 1869, 2694, 3936, 8174, 11607, 286, 373, 5770]. Cleansing [9293, 1178]. Clearing [4194, 7660]. Clergy [965]. Clermont [756, 758]. Cleveland [11494]. Clever [2425, 3309, 3382, 8816, 11224]. Clifden [10541]. Cliff [7044, 1777, 2762]. Cliff-Dwelling [2762]. Climate [9642, 10127, 1991]. Climb [10963]. Climbing [8435, 10323, 11373, 4943]. Climbs [4579].

Clinical [8123]. Clinton [2464]. Clipper [3097, 5782]. Clock [2268, 2511, 2730, 2534, 6442, 6483, 7280, 6401, 10205, 11116, 1224, 341, 7627, 7401, 9863, 10405, 10997, 6399, 11783, 4939]. Clocks [7251]. Clockwork [2985]. Close [2234, 2146, 7128, 106, 973]. Close-Corner [973]. Closed [8805]. Closes [11567]. Closet [1233]. Closing [2938, 3953, 4056, 7848, 9626, 11353]. Closure [8285, 9886]. Cloth [1907, 3464, 1100, 9839]. Clothes [2126, 8992, 10092, 1233]. Clothes-Line [8992]. Clothes-Washing [2126]. Clothing [2910, 6516, 6463, 11877]. Clouds [2301, 1721]. Club [4011, 9011, 9487, 9480, 8832, 9756, 10474, 10962, 9859, 5065, 5064, 10061]. Clubs [4879]. Clusters [3457]. Clutch [4036, 3888, 6167, 8837, 8809, 9943, 9935, 9938, 9937, 9942, 11332, 903, 328, 48]. Clutches [9934, 9932, 1871]. Clyde [6094, 2933]. Co [8164]. Coach [9584]. Coaches [5257, 10686, 47, 7342]. Coal [3134, 2286, 2005, 2978, 4499, 4241, 4385, 4511, 5407, 7156, 6845, 7647, 8520, 10643, 10174, 9909, 10171, 11063, 11030, 383, 8571, 2263, 4491, 2310, 814, 3286, 5631, 5258, 8500, 1727, 7140, 5573]. Coal-Bagging [10174]. Coal-Conveying [7140]. Coal-Loading [4241]. Coal-Mining [1727]. Coal-Porcelain [5631]. Coal-Testing [8520]. Coaling [7700, 9424, 10174, 5389, 8983]. Coast [2379, 1978, 2794, 4234, 6216, 7622, 7518, 9502, 11085, 9983, 10811, 10340, 2256, 3506, 10215, 614, 9095, 10710, 3140, 4376, 3355, 10692]. Coast-Defense [2379, 7518]. Coaster [3822]. Coaster-Brake [3822]. Coasting [641, 294]. Coasts [11707]. Coat [5239]. Coating [4705, 812]. Coats [11158]. Cobalt [3133]. Cobra [7654, 3260]. Coburg [4267]. Cock [1389, 9781]. Cocoa [2869, 5743]. Cod [9917, 651, 4268]. Cody [730]. Coffee [9758, 958, 6854, 1229]. Coherer [2679]. Coherers [5890, 6197, 2258]. Cohesive [9745]. Coil [9125, 141, 562, 2535, 420]. Coiled [11545]. Coils [6961, 10480, 5261]. Coin [6696, 11057, 1055, 9146]. Coin-Operated [1055]. Coke [3134, 3617, 5518, 11824, 10406]. Col. [3364, 4561, 11213, 346, 864]. Cold [10305, 732]. Colds [8253]. Coldwell [6655]. Collapse [4434, 5128, 6310]. Collapsible [2101, 9492, 11665]. Collar [2579]. Collecting [9452, 11035, 3960]. Collection [1770, 4267, 5189, 8505, 11640, 6856, 7915, 6916]. Collectors [6181]. College [5442, 10753, 10673]. Colleges [4958]. Colliers [4459]. Collins [11458, 4755]. Collision [4626, 4130, 7187, 9374, 6958]. Collisions [6455, 11579]. Colloids [8227]. Colman [4914]. Cologne [7987]. Colombia [4339, 6155]. Colon [7417]. Colonial [5978]. Colonies [10443, 11170]. Colonists [5432]. Colonnade [7212, 6809]. Colony [11222]. Color [2013, 1645, 1536, 1435, 5864, 7063, 8876, 9600, 9369, 10667, 11320, 311, 5626, 2179, 970, 11899]. Color-Blindness [11899]. Color-Photography [1645, 311]. Colorado [10730, 9215, 3456, 7042, 9509, 2441, 8690, 8707, 802, 2762, 9886]. Colors [1367, 6806, 7344, 796, 5626, 4616]. Colossus [4463]. Colt [1465, 4783]. Columbia [2523, 3219, 3172, 3654, 5075, 5225, 6664, 11305]. Column [3086, 7499, 11882, 8283]. Columns [2782, 4053, 5698, 6310, 7039, 3011, 8640]. Combating [769]. Combination [2303, 2738, 3062, 4166, 5860, 6271, 6660, 9102, 11241, 619, 10260]. Combinations [1104]. Combine [3903, 4047, 5620, 4453]. Combined [3159, 1490, 1342, 3613, 2995, 3528, 4422, 4897, 5680, 6740, 6406, 6197, 8086, 8675, 8659, 9967, 10440, 10048, 10789, 9797, 10566, 9943, 9931, 9937, 10286, 10884, 10994, 11662, 10886, 15, 903, 438, 1001]. Combining [3905]. Combs [6046]. Combustible [7272]. Combustion [5580, 9749, 1001, 991]. Come [11216]. Comedy [10147]. Comet [7117, 798, 2520, 5568, 6071, 10359, 11456, 593, 828, 742, 4082, 6145, 289, 2887, 6144, 1252]. Comets [11885]. Coming [1489, 3810, 6607, 6482]. Command [8202]. Commander [728]. Commanders [7385]. Commemorating [9746]. Commemoration [6389]. Commencement [5310]. Commerce [1845, 1455, 4021, 6052, 5561, 5427, 9874, 10537, 9641]. Commerce-Destroyer [6052]. Commercial [1745, 3199, 4798, 4076, 4487, 4523, 6530, 9445, 9603, 9243, 8589, 9875, 10792, 11934, 167, 11609, 479]. Commission [3882, 6821, 9366, 8547, 1642, 2275, 2300, 2298, 5970, 6308, 9042, 9878, 11483, 11819, 1598, 5771]. Commissioner [1335, 1495, 3058, 2728, 2660, 2615, 2595, 9978, 11571]. Commissioners [6373]. Commissionership [1432]. Committee [8887]. Common [3386, 4138, 5870, 5884, 8571]. Common-Battery [3386]. Commune [5357]. Communication [5675, 998, 5259]. Commutator [3041]. Compact [7857]. Companies [1697, 10339]. Company [2331, 3240, 11210, 2217, 3508, 2665, 3688, 7833, 7862, 4667, 9215, 9288]. Comparative [1791, 2257, 3036, 2860, 7633, 9720, 10193, 10896, 610, 305]. Compared [1333, 2325, 2105, 3375, 3654, 6618, 7597, 11493, 11365, 857, 613, 276, 4103]. Comparison [1419, 4148, 3836, 5869, 6705, 8100, 8676, 9351, 8959, 11844, 358, 591, 803, 5333]. Comparisons [458]. Compass [1511, 11196, 9802]. Compass-Testing [11196]. Compasses [4535]. Compensating [6514, 7216]. Compensation [2200]. Compensator [8833]. Compete [5576]. Competing [10320]. Competition [2366, 2436, 2170, 5329, 5322, 6963, 7534, 10794, 9857, 10870, 211, 1952, 1168]. Competitions [10335]. Competitive [3442]. Competitor [2210, 11400]. Competitors [10042]. Complement [333]. Complete [7622, 7491]. Completed [2640, 4823, 9204, 9585, 10748, 11251, 11313]. Completely [2693]. Completing [6876, 6536, 7581]. Completion [4309, 4972, 6235, 6830, 6679, 7774, 8410, 9282, 9002, 10299, 9722, 1255, 7404]. Complications [2748]. Composing [5494, 8988]. Composite [6827, 8304]. Composition [5853, 8472, 616, 716]. Compound [2192, 1872, 3418, 3261, 3331, 4033, 6241, 7190, 7254, 7377, 8871, 333, 6575, 11549, 8621]. Compounds [2682, 2753, 9793, 10333]. Comprehensive [10604]. Compressed [2487, 3940, 6761, 9530, 9751, 539, 832, 1016, 4689, 9840].

Compressing [9729]. Compression [1515, 2859, 10332, 10058, 11146, 136, 8900, 364, 8043]. Compressor [5323, 7208]. Compressors [1175]. Compromise [1928]. Computing [11119, 3761]. Comstock [3573]. Concentrated [7502]. Concentrates [11869]. Concentrator [11465]. Conception [8828, 10974]. Conceptions [9026]. Concern [1987]. Concerning [2512, 10177]. Conclusions [10908]. Concourses [7166]. Concrete [2103, 6160, 5803, 6263, 6365, 8380, 8764, 8769, 9639, 8768, 8892, 8761, 8772, 9473, 8771, 8765, 9824, 10696, 10747, 9812, 10522, 10396, 9777, 10215, 10254, 10223, 11265, 11413, 10900, 11475, 10871, 514, 1296, 9095, 2800, 11716, 3425, 8283, 8773, 11704, 8405, 989, 1148, 917, 8456, 7628, 7831, 11806, 11326, 7522]. Concreting [8766]. Condenser [1140]. Condensers [3505, 11237]. Condensing [3664, 3418]. Condiment [5291]. Condition [3906, 5166, 8932, 9521, 9263, 11152, 11386]. Conditions [2683, 3305, 5663, 7230, 7953, 8847, 9414, 8868, 9447, 10404, 9876, 10791, 9813, 10957, 335, 1011, 5477, 560, 6023]. Condor [11397]. Conducted [5165, 6317, 6559, 7045]. Conductivity [10038]. Conductor [9768]. Conductors [8151, 10666]. Conduit [917, 435]. Cones [9667]. Coney [2012, 11904]. Conference [9326, 11578, 11747]. Congested [2862]. Congestion [2255, 5205, 9413, 10087, 10826]. Congo [2370, 8462]. Congress [2279, 4388, 6497, 10297, 2098, 2097, 2277, 1576, 1644, 2148, 2463, 2027, 2322, 1955, 3108, 5972, 6117, 7297, 6871, 7576, 7623, 9853, 11186, 10850, 581, 1067]. Conjunctions [1281]. Connect [808]. Connected [1579]. Connecticut [4880, 6938, 6893, 6906, 7536, 8617, 10583, 857]. Connecting [3526, 9918, 1104]. Connection [809, 5387, 5412, 5266, 7455, 10556, 11358, 1158]. Connections [3130]. Connellsville [3617]. Conqueror [4785]. Conquest [3576]. Cons [11840]. Conservation [11387, 11102]. Conservatory [1684]. Consigned [7439]. Consolidated [3240]. Consolidation [630]. Constant [7183]. Constitutes [10573]. Constitution [2759, 2806, 2795]. Construct [10907, 3185, 8746]. Constructed [8380, 9584, 10759, 10313, 558, 7912, 65]. Constructing [2533, 8147, 10622, 1148, 4151, 323]. Construction [1550, 2157, 1360, 2276, 1683, 2299, 2444, 2688, 2592, 2759, 3282, 3056, 3335, 2915, 3325, 3361, 3475, 4218, 3908, 4336, 5033, 5545, 5769, 5866, 4931, 5890, 5183, 5537, 5803, 4992, 6732, 6639, 6435, 6884, 7717, 8632, 8716, 9446, 9125, 9512, 9008, 9111, 8765, 10696, 10494, 9745, 10396, 9777, 10215, 10254, 10507, 10057, 11188, 10870, 11481, 11864, 10960, 11479, 11259, 11056, 1196, 774, 465, 1003, 79, 11690, 9981, 11440, 5699, 11704, 9798, 7025, 9805, 8900, 9983, 10739, 2076, 70, 4060, 10283, 11806, 1479]. Constructor [2278, 1920]. Consular [1331]. Consumed [10422]. Consumer [4311, 6183]. Consuming [1415, 3399]. Consumption [2978, 9595, 11388, 11030]. Contact [5565, 6406]. Contaminated [7397]. Contest [3271, 6885, 6532, 6984, 7536, 7399, 8435, 8832, 106, 184, 2117, 5472, 6164, 6812, 10708]. Contests [7621, 3252, 670]. Continent [8787]. Continental [2955]. Continuous [7000, 9799, 4128]. Continuous-Alternating [4128]. Continuous-Feed [7000]. Contract [8248, 9283, 10493, 10190, 1255, 4079, 1454]. Contractor [5156]. Contracts [11345]. Contrivance [3960]. Control [6557, 6202, 8049, 8120, 9552, 10658, 10064, 11586, 11085, 11113]. Controlled [7454, 9796, 775, 63]. Controller [4555, 11905]. Controlling [1487, 7840, 9509, 9752, 1019]. Controls [9946]. Controversy [2231, 3622, 6114]. Convalescents [8163]. Convenient [2079, 2421, 1540, 5191, 5382, 6323, 8217, 8749, 780, 468, 10442, 9051, 1233, 11689]. Convention [2982, 5999, 5105]. Conversion [3457]. Converter [5480, 7541]. Convertible [10996]. Converting [10106]. Conveyer [4025, 5719, 1220, 8741]. Conveyers [2311]. Conveying [2801, 678, 599, 7140]. Cook [729, 709, 696]. Cooker [8006, 65, 67]. Cooking [3447, 4511, 6693, 10383, 9422, 9956]. Cooled [6647, 6650, 6648, 8812, 8791, 8792, 8793, 8795, 10455]. Cooler [4397, 10566]. Cooling [2984, 5232, 11427, 8985]. Coop [7965]. Cooper [2716, 5538, 5480, 7855, 7541]. Copenhagen [363]. Copper [2045, 1386, 3742, 5673, 10685, 10733, 11079, 11860, 344, 11715, 203, 2864]. Copper-Clad [11079]. Copper-Plating [11715]. Coppering [326]. Copy [2558]. Copying [9735]. Copyright [652]. Copyrighted [8074]. Coral [6260, 8458]. Cordage [3287]. Cords [1371]. Core [6025, 27]. Cored [11012]. Coreless [7830]. Corinth [1932, 4389]. Cork [310, 1234]. Cork-Lined [310]. Corn [1913, 3115, 9016, 1017]. Corner [973, 8428, 11233]. Corners [2150]. Cornstalks [11558]. Cornu [10172, 10972]. Corona [8473]. Coronation [4899]. Corporation [4500]. Corps [4730, 11192, 7299]. Corpuscles [8471]. Corpuscular [2792]. Correct [6464, 6706]. Correcting [494, 436]. Correction [6047]. Correctly [558]. Correspondence [1309, 1324, 2136, 2203, 2248, 2359, 2448, 2501, 1352, 1399, 1423, 1496, 1542, 1565, 1613, 1661, 1686, 1711, 1735, 1754, 1807, 1833, 1876, 1944, 1968, 1992, 2086, 2109, 2584, 2627, 2672, 2742, 2765, 2852, 2876, 2923, 2971, 3024, 3048, 3100, 3123, 3165, 3188, 3213, 3277, 3321, 3449, 3495, 3534, 3579, 3602, 3625, 3691, 3707, 4517, 4564, 4608, 4652, 4696, 4721, 4741, 4767, 4817, 3747, 3774, 3827, 3873, 3921, 3946, 3990, 4039, 4067, 4092, 4112, 4153, 4177, 4204, 4304, 4355, 4378, 4399, 4426, 4475, 4908, 4951, 5840, 5872, 5892, 5962, 5990, 6108]. Correspondence [6150, 4983, 5004, 5094, 5121, 5150, 5170, 5196, 5217, 5273, 5349, 5397, 5420, 5447, 5464, 5487, 5512, 5554, 5656, 5726, 5753, 5817, 6169, 6188, 6815, 6899, 6921, 6949, 6968, 6987, 7012, 7029, 7048, 7072, 7096, 7121, 7174, 7198, 7243, 7266, 7286, 7310, 7331, 6210, 6227, 6251, 6280, 6300, 6326, 6368, 6408, 6427, 6448, 6489, 6520, 6545, 6563, 6592, 6673, 6722, 6747, 6768, 6790, 7365, 7387, 8064, 8094, 8113, 8132, 8153, 8179, 8196, 8219, 8243, 8264, 8289, 8327, 8343, 8366, 8393, 8416, 8440, 8465, 8484, 8511, 7410, 7432, 7461, 7482, 7506]. Correspondence [7549, 7570, 7590, 7615, 7639, 7664, 7711, 7740, 7816, 7842, 7869, 7923, 7946, 7969, 7993, 8010, 8033, 8534, 8556, 9255, 9275, 9297, 9359, 9382, 9407, 9454, 9474, 9495, 9515, 9533, 9555, 9573, 9609, 9634, 9658, 8602, 8627, 8646, 8668, 8693, 8727, 8841, 8861, 8881, 8907, 8951, 8971, 8996, 9037, 9058, 9084, 9105, 9132, 9155, 9175, 9196, 9220, 9235, 10115, 9691, 9713, 10367, 10387, 10407, 10429, 10484, 10509, 10531, 10549, 10598, 10615, 10660, 10677, 10723, 10742, 10762, 9736, 9760, 9806, 9827, 9869, 9887, 9915, 9968, 9994, 10012, 10029, 10050, 10071, 10141]. Correspondence [10163, 10182, 10249, 10269, 10306, 10326, 10344, 10822, 10838, 11485, 11505, 11528, 11550, 11574, 11591, 11645, 11683, 11691, 11741, 11789, 11811, 11835, 11850, 11889, 11907, 11926, 10861, 10910, 10931, 10951, 10983, 11004, 11023, 11047, 11068, 11096, 11121, 11140, 11204, 11225, 11244, 11268, 11287, 11335, 11379, 11403, 11441, 39, 72, 710, 737, 783, 815, 837, 874, 906, 950, 974, 1005, 1032, 1060, 1081, 1106, 1130, 1181, 1204, 1263, 1283, 98, 129, 151, 179, 205, 227, 244, 269, 297, 329, 351, 374, 401, 446, 498, 519, 542, 573, 600, 621, 690, 4327]. Correspondences [10115]. Corrosion [6680, 7205]. Corrugated [774]. Corsica [3419]. Cortlandt [10154]. Cost [1791, 2299, 3036, 2754, 3881, 3714, 5303, 7229, 6972, 11183, 11250, 476, 3185, 11064]. Costly [2882, 10703, 10103, 8981]. Cotton [1484, 2055, 1633, 2683, 2754, 2648, 4346, 7253, 7625, 657, 8748, 1654, 1894]. Cotton-Picker [1633]. Cotton-Picking [7253]. Could [1194]. Count [2443, 1628, 3317, 9128, 883, 968]. Counter [11057, 8277]. Counterpart [612]. Counterweighted [7278]. Counting [6760, 9146, 703]. Countries [2325, 3716, 790]. Country [1363, 6309, 11346, 1193, 1275, 1041, 10, 1276]. Coupé [3796, 7919]. Couple [3436]. Coupling [2197, 2779, 7232, 7218, 7233, 6184, 8217, 7456, 10545, 9905, 10977]. Courage [8733]. Course [6739]. Court [1841, 2593]. Courtesy [11935]. Cousins [800]. Cover [4560, 5889, 126]. Covering [3275, 3401, 5861, 10816, 7588]. Covers [5861]. Cow [5549, 7580, 10444]. Cow-Milking [7580]. Cows [5481]. Cracked [10353, 517]. Cracker [10898]. Craddock [3701]. Cradle [4788]. Craft [10712, 11179, 11651, 10088, 9114, 4372]. Crafts [4168]. Cranberry [3765]. Crane [3248, 4107, 5909, 8940, 1141, 10139, 4804, 7324, 9567]. Cranes [5133, 1851, 5697]. Crank [9207]. Crate [1231]. Crater [7579]. Craze [2956, 137, 4662]. Cream [4020, 6009, 441]. Creation [9332, 7728, 8087]. Creatures [9033, 7363, 7916, 10065]. Creek [5577]. Creeping [1343, 1390]. Crescent [11043, 11631]. Crib [9378]. Criminals [8088]. Crisis [2457, 3676, 7182, 10148]. Critical [5868]. Criticism [2955, 4481, 4635, 11029, 11721, 11343]. Criticisms [11295, 334]. Crocco [10439]. Crocker [3427]. Crookedness [10848]. Crookes [6098, 5131]. Crop [1940, 10024, 1077]. Crops [1821, 1095]. Cross [3419, 5703, 6914, 1193, 1298, 1275, 1218, 1041, 949, 1276, 5362]. Cross-Channel [1193, 1218, 5362]. Cross-Country [1193, 1275, 1041, 1276]. Crossing [8787, 10704]. Crossings [699, 6122]. Crossley [5236]. Croton [5284, 1772, 3205, 3194, 3455, 4991, 6876, 7848, 8017, 7496, 8959, 9002, 9647, 11112, 239]. Crowd [7517]. Crowding [7511]. Crown [123]. Crude [2498, 10406]. Cruise [10294, 335, 620, 604, 10577]. Cruiser [1347, 1762, 2537, 4865, 4490, 4786, 6056, 5544, 6764, 6982, 6803, 7155, 7042, 6507, 8593, 9835, 10757, 10585, 10587, 10586, 11002, 11896, 11033, 11637, 46, 1016, 11585]. Cruiser-Battleship [11896]. Cruisers [1974, 1407, 2155, 2266, 2338, 1750, 3487, 3489, 2934, 3488, 4738, 6625, 7852, 10584, 10896, 11451, 305, 53, 1136, 11133, 7858]. Cruising [3009, 5069, 10240]. Crusade [11410]. Crush [10415]. Crushed [6070, 9639, 1295]. Crushed-in [1295]. Crusher [4131, 11604]. Crushing [1930]. Crust [9579]. Crustaceans [3988]. Cruz [3702]. Crystalline [1953]. Crystallization [7276]. Crystals [9429, 6335, 7603]. Cuba [2280, 580, 5443, 596, 4031, 10284, 9331]. Cuban [4775]. Culebra [11887]. Culled [4847]. Cullen [1446]. Cullinan [11543, 985]. Culm [2073]. Cultivating [9193]. Cultivation [2869, 3359, 2798, 7361, 7066]. Cultivator [588]. Culture [2530, 5360, 7587, 4102, 6854, 7813, 4085, 5328, 10427]. Cultures [8589]. Cunard [3734, 4524, 5620, 5970, 6831, 8408, 7383, 9330]. Cunarder [7296]. Cunarders [4930, 5248, 5802, 6331, 8300, 8639]. Cup [2117, 1466, 3283, 2611, 3667, 3252, 3541, 2914, 4362, 5472, 5576, 6053, 4956, 6885, 6781, 7023, 7605, 7658, 7478, 8075, 8108, 8139, 9287, 10374, 10239, 10708, 11603, 11895, 11580, 957, 1201, 1059, 2323, 1965, 3271, 3172, 2634, 5365, 5285, 6100, 5063, 6963, 7467, 8170, 8448, 9230, 9161, 9622, 9270, 10797, 10392, 911]. Cup-Yacht [2611, 2634]. Cup-Yachts [2914]. Cupidity [10747]. Cuprene [1482]. Curb [8193]. Cure [5641, 9225, 10732, 9982, 10126, 5911]. Cures [10775]. Curie [3938, 5042, 7273, 6572, 6996]. Curing [2498]. Curiosities [59, 1052, 923, 1875, 11396, 8745]. Curiosity [2568, 3227, 4758]. Curious [2192, 2467, 1976, 2305, 1805, 1608, 2937, 3293, 3226, 3290, 3367, 4059, 4240, 4605, 4967, 5528, 4999, 4920, 5701, 6908, 7069, 8531, 7380, 9078, 10905, 11587, 11777, 888, 515, 920, 592, 705, 5001, 7943, 8236, 8919, 4464, 2356, 3208, 4552, 8229, 9423, 10009, 9214, 10530, 3789, 11221, 1721, 5715, 2783, 1709, 7024]. Curiously [1631, 7503]. Current [2591, 4460, 4690, 4971, 5435, 5541, 5567, 5765, 5787, 5830, 5976, 6003, 6120, 6157, 5619, 6357, 6393, 6711, 6849, 6236, 7443, 7696, 7724, 7787, 7879, 7904, 7980, 8023, 8045, 8078, 8166, 8376, 8241, 9328, 9349, 9394, 9482, 8781, 8542, 9725, 9748, 9794, 10128, 10194, 10419, 10519, 10645, 10711, 10755, 10772, 10853, 10874, 11918, 10919, 10966, 11061, 11081, 11109, 11149, 11171, 11190, 11235, 11258, 11299, 11322, 11350, 11393, 11431, 11455, 11478, 11634, 11822, 11881, 11901, 11568, 252, 721, 4128, 6122, 10776, 1202, 866, 3942]. Currents [5363, 10685, 9768, 11946, 1278, 10780, 6206, 4644]. Curtain [1603, 7081]. Curtain-Hanger [1603]. Curtains [6691]. Curtiss [434, 1172]. Curves [3780, 8041, 8449, 9640, 366, 5790, 6121]. Curzon [11479]. Cushing [1486]. Cushion [1774, 8990]. Cushion-Frame [1774]. Cushioned [10325]. Cushions [5080]. Customs [11433]. Cut [1631, 6381, 11887]. Cutler [3046]. Cutter [3139, 4173, 6541, 8472, 9233, 10046, 10718, 9844, 11119, 439, 781]. Cutter-Head [439]. Cutting [1414, 2178, 5542, 7343, 9709, 11334, 11525, 322, 11038]. Cycle [3232, 3016, 4291, 4882, 5637, 6033, 6264, 8797, 9120, 10483, 11658, 11808, 2468, 8259, 483, 3510]. Cycle-Whirls [6033]. Cycles [8528]. Cycleway [2201]. Cycling [827]. Cyclograph [6481]. Cyclone [4461, 11865, 7260]. Cyclones [11682]. Cylinder [1872, 3387, 6647, 6264, 6220, 6645, 7801, 8231, 7790, 8796, 8791, 8792, 8801, 8795, 8794, 9123, 10457, 10456, 9949, 10213]. Cylinders [9513, 8985]. Cylindrical [3245, 6339]. Cynips [4030]. Czarevitch [2640].

D. [4643]. D.C. [7628]. Dabbing [11200]. Dado [2053]. Dado-Head [2053]. Dairy [3268, 7706]. Dakota [6803, 10597, 11703, 10973, 93, 191, 1012, 967]. Dakota’s [345]. Dalmatia [10692]. Dam [9045, 1518, 6143, 6020, 8632, 10440, 10730, 10036, 5627, 10808, 1772, 3205, 3194, 3455, 5545, 4991, 5577, 5284, 6876, 7848, 8282, 8333, 7554, 8959, 8565, 8932, 10062, 9855, 10681, 11274, 11795, 11390, 11074, 11416, 524, 7450, 11398, 11590]. Dam-Builder [6143]. Damage [2031, 3502, 7380]. Damascening [1197]. Damp [8450, 11063]. Dams [9239, 10189, 11409, 6339]. Dance [5958]. Danger [3655, 8804]. Dangerous [7955]. Dangers [4020, 4189, 5102, 5733, 7468, 7376]. Daniel [9621, 10359]. Danilewsky [1780]. Danube [1984, 4508]. Danysz [6619]. Darien [4254]. Daring [4075, 5335, 7316]. Darkness [6354]. Darlington [6310, 6215]. Darwin [8737, 260]. Darwinian [11820]. Dashboard [8807]. Dastardly [6938]. Date [9580]. Dating [9801]. David [8498]. Davidson [6685]. Davis [11419]. Day [4488, 6115, 5737, 8313, 8319, 8048, 10414, 9791, 10832, 11460, 11886, 11565, 288, 10970]. Daylight [5193, 9881, 11858, 1134]. Days [10664]. Dayton [10481, 9944]. Dead [3658, 7676]. Deadlock [9975]. Deadly [5392]. Deaf [5332, 10044]. Deaf-Mutes [10044]. Dealing [2862]. Death [1532, 1723, 1486, 1648, 2174, 1387, 1795, 1822, 2514, 1437, 1460, 2464, 2935, 2776, 3014, 2705, 3380, 2706, 4661, 4366, 6048, 4934, 6050, 5973, 5867, 5805, 5974, 7130, 6460, 6416, 8352, 9621, 9769, 10689, 10641, 11168, 10875, 11169, 882, 1602]. Debris [7057]. Debt [580]. Decade [4255, 8914]. Decadence [3059, 3807]. Decadent [11054, 503]. December [11060, 8307, 11702, 7107, 2247, 3438, 4628, 5925, 9434, 10548, 11692, 1033]. Decision [1719, 1862, 1790, 2567, 2545, 5456, 1039]. Decisions [2593]. Deck [1444, 5713, 6240, 7477, 8684, 10087, 10339, 10959]. Deck-Calking [8684]. Decked [5127]. Decking [4232]. Decks [7256]. Decolorizing [10918]. Decorating [3864]. Decoration [11111]. Decrease [11510, 11127]. Dedicated [7880, 1123]. Dedicates [4561]. Dedication [1469]. Deductions [7318]. Deep [3632, 3241, 5879, 5507, 1073, 7943, 6424]. Deep-Level [5879]. Deep-Sea [6424]. Deepening [7137]. Deer [6180, 8212]. Defeat [7346]. Defective [3543]. Defects [8867, 10915, 146]. Defender [3172, 11895, 5285]. Defense [2379, 2794, 6216, 7518, 10588, 11805, 143, 359, 1274, 3208, 10811, 6909, 10340]. Defenses [4363, 7622, 9502]. Deferred [3059]. Definition [4022]. Deflector [9294]. Deflocculated [9897]. Deforestation [995]. Degradation [10733]. Degree [10457]. Degrees [8432]. Del [3140]. Delagrange [11216, 10991, 11046]. Delaware [10597, 10973, 789]. Delay [5866, 4931, 8517, 9043]. Delayed [3282, 11411, 303]. Delayed-Action [11411]. Delays [11276]. Delicate [5835]. Deliverer [10817]. Delivery [3332, 6276, 6558, 7796, 7840, 8363, 266, 9588]. Delos [1279]. Delphi [316]. Delta [3456]. Demagnetizer [7566]. Demand [9743, 8785]. Démocratie [7237]. Demolition [218]. Demon [2337]. Demonstrating [10114]. Denatured [8870, 9792]. Denaturized [8889]. Dene [6355]. Dene-Holes [6355]. Density [443]. Denver [5577]. Deodoro [1394]. Departement [7090]. Department [1455, 1828, 1361, 3347, 3445, 3467, 3531, 3794, 3985, 3767, 3821, 3870, 3915, 3963, 4009, 4088, 4417, 4646, 4713, 4760, 4811, 4854, 5165, 5051, 5262, 5337, 5416, 5482, 5548, 5579, 5887, 5936, 5985, 6057, 6102, 6270, 6320, 6400, 6466, 6510, 6582, 6610, 6942, 7006, 7142, 7345, 7452, 7524, 7610, 7634, 7682, 7864, 7890, 8056, 8173, 8214, 8238, 8284, 8389, 8506, 8552, 8903, 8943, 9334, 9450, 9490, 9625, 9985, 10134, 10322, 10504, 10544, 10564, 10693, 10834, 10926, 10975, 11091, 11306, 11330, 11464, 11500, 11566, 11661, 11759, 11832, 11866, 11947]. Department [118, 174, 264, 370, 437, 778, 901, 946, 1000, 1173, 1302, 3666, 4998, 2131, 5277, 11769, 1086, 1420]. Departments [11697]. Departure [4938, 6308]. Deposit [2645, 4789]. Depositing [5837]. Deposition [10152]. Deposits [1610, 3941, 6670]. Depot [8983]. Depressed [8141]. Depression [3199, 9552]. Depth [8324]. Derailment [8041]. Derelict [9893]. Derivatives [7037]. Derrick [3507, 2691, 5458, 8940]. Desaix [9701]. Descent [9620]. Describing [6121]. Description [5628]. Desert [431]. Design [3610, 3152, 3622, 3443, 4560, 3999, 4033, 3887, 5246, 6706, 7852, 9544, 11526, 10968, 919]. Designed [4299, 8987]. Designer [7130, 10708, 617]. Designers [11344]. Designing [2611, 5618, 8782]. Designs [1407, 10070, 11747, 9750]. Destroy [2953]. Destroyed [1934]. Destroyer [1493, 3492, 3783, 6052, 5794, 11415, 9641]. Destroyers [2023, 3491, 4212, 6222, 10066, 10589, 894]. Destroying [5563, 7632]. Destroys [11492]. Destruction [1336, 4549, 5823, 721, 5209, 11114]. Destructive [3231, 7187, 11781]. Destructor [8678]. Detachable [5085, 6658, 7447, 8798, 8574, 9566, 10137]. Details [5475, 293]. Detect [4935]. Detecting [2638, 165, 5835]. Detection [2679, 11598, 11748, 6714]. Detector [4471, 3929, 9798, 10833, 321]. Detectors [4439, 9732]. Deterioration [3284]. Determination [8547]. Determine [6464]. Determining [2130, 3587, 5311, 6605, 7767, 443, 9029, 5083, 7606, 8524]. Detonation [9097]. Detonator [1632]. Detroit [10630]. Deutsch [3408, 11253]. Deutschland [2368, 1979, 2423, 3568]. Develop [11843]. Developed [10277]. Developer [1674]. Developing [5193, 9881, 540]. Development [2169, 1645, 1852, 2668, 3476, 3197, 4682, 4454, 4255, 4748, 6162, 5969, 5179, 6530, 6221, 6637, 6391, 6843, 7753, 8701, 8544, 10593, 10234, 10916, 755, 433, 10968, 5443, 4681, 3286, 4100, 5135, 4467, 10241, 659, 7059]. Developments [2096, 1956, 3586, 8146, 8468, 8481, 8616, 8206, 5947, 2664]. Device [1943, 1848, 3660, 3567, 2961, 2897, 3505, 3137, 4585, 4171, 4320, 4587, 3864, 4146, 4812, 6125, 4915, 6167, 5088, 5549, 5704, 5681, 5077, 5417, 5986, 6274, 6661, 6693, 6514, 6517, 6945, 6977, 6272, 6889, 8390, 8301, 8091, 8362, 8363, 8239, 9626, 9250, 8944, 9377, 9232, 9116, 9452, 9046, 8751, 9335, 8810, 9082, 9710, 10362, 9709, 9960, 10760, 10324, 9976, 9823, 11264, 10813, 11018, 11761, 11616, 444, 1055, 35, 538, 540, 175, 441, 1029, 873, 6999, 4599, 867, 1203, 11788, 1030, 2061]. Devices [3367, 3399, 3446, 4750, 5550, 8058, 9466, 10156, 10039, 9789, 10157, 11321, 11185, 10961, 11819, 11476, 8302, 10003, 2267, 10965]. Devil [9290]. Devils [5859]. Dial [11156]. Dials [9077, 11174]. Diamond [1804, 4906, 6537, 7479, 8472, 8619, 11543]. Diamonds [6025, 11347, 8007]. Diaphanoscope [926]. Diaphragm [5054, 6859, 11014]. Diary [5433]. Diction [88]. Did [55]. Die-Sinker [2439]. Dietrich [3092]. Differences [1939]. Different [2886, 6930, 9291]. Differential [9931, 891]. Difficulties [7468, 10622]. Diffraction [8876]. Digestion [1149]. Digger [7526]. Digging [1539, 4106]. Dime [5479]. Dimension [427, 106, 159, 656, 942, 1168, 10879, 1143]. Dimensions [11281]. Dining [10741, 7376]. Dinner [9756]. Dino [6916, 6856]. Dinosaur [3644, 10614, 10950, 7853, 3697]. Dinosaurs [2600]. Dioxide [9244]. Dip [6080]. Diplodocus [924, 10065]. Dipper [3088]. Direct [2303, 1579, 6157, 5619, 6236, 6806, 7344, 7903, 8781, 8542, 252, 598, 3788, 1202]. Direct-Connected [1579]. Direct-Current [6236, 8781, 8542, 252, 1202]. Direct-Reading [598]. Directed [8849]. Directing [1943, 10505, 9031]. Directive [935]. Dirigible [2336, 2536, 5208, 9128, 11522, 11938, 9352]. Dirigibles [359]. Disabled [2848, 6764]. Disappearing [2496, 2761, 7518, 9615, 2337]. Disaster [2932, 4606, 4005, 6352, 6215, 6476, 8480, 8446, 8733, 8718, 10682, 9744, 10233, 10312, 11178, 11472, 11897, 1040, 275, 548, 2022, 8770, 10553, 10292, 10318, 10988]. Disasters [2863, 6501, 7720, 10607]. Disastrous [3763]. Disbursements [4215]. Discharge [2084, 6097]. Discharges [5930, 9010]. Discharging [5117]. Disclaimer [4825]. Discovered [3295, 4527, 6140, 7141, 7772, 751, 8594, 5676]. Discoverers [259]. Discoveries [2074, 1932, 3218, 6069, 5180, 8171, 9618, 9069, 10631, 6608, 5983, 4873, 340, 11754, 7734]. Discovery [5049, 6619, 6973, 7397, 8825, 10733, 728, 709, 6081, 1777, 9353, 9027]. Discredited [8183]. Disease [7400, 8760, 9489]. Diseases [5473, 10126]. Disengaging [8145]. Disfigurement [6992]. Dish [2007, 1889, 9776]. Dish-Washing [1889]. Disinfecting [7257]. Disinfection [1170]. Disintegrator [346]. Disk [6744, 7540, 9943]. Dismal [1494]. Dismantled [6004]. Dispatch [1644]. Dispelling [6810]. Dispersion [6687, 6289, 7859]. Displace [7133]. Display [2282, 5937]. Disposal [1929, 5772, 5877, 5793]. Disposes [9144]. Disposing [4734]. Disposition [6955]. Dissect [9643]. Dissociation [1337]. Distance [1842, 2980, 3713, 4186, 5795, 7231, 8463, 7718, 8400, 7793, 9504, 10080, 10441, 10236, 11458, 983, 8169, 1658, 5235]. Distances [5388]. Distant [2004, 10658]. Distillation [6461, 9112, 7133]. Distorted [7503, 250]. Distributer [3224, 1001, 440]. Distributing [6517].

Distribution [3159, 4386, 6005, 5640, 11740]. District [6119, 1819]. Districts [6309, 5720, 11919]. Disturbances [4045, 956]. Disulphide [10903]. Ditch [1539]. Ditch-Digging [1539]. Ditcher [6829]. Ditching [7684, 10547]. Divergent [3610]. Diverging [4928]. Divers [6533]. Diversion [8582, 5037]. Diversions [1045]. Divided [5571]. Dividend [6001]. Dividing [7836]. Divine [5698, 7039, 11233]. Diving [6935, 9465, 1280, 5574, 1733]. Division [1431]. Divisions [6994]. Divison [1361]. Do [8895, 801, 896, 7680, 10090, 9420, 672, 11925]. Dock [2276, 1492, 1395, 2103, 2244, 3596, 4973, 5837, 6179, 11736, 10945, 10877, 5697, 3816, 4580]. Docking [2937, 10710]. Docks [2198, 1530, 2190, 5277, 4831]. Dockyard [8082]. Doctoring [11377]. Doctors [5223]. Doctrine [260]. Doenvig [6486, 6848]. Does [5129, 9025, 10413, 753, 1221, 1046, 8296, 11843]. Dog [2101, 8322, 9142]. Dog-Shearing [8322]. Dollar [1363, 3240, 7600]. Dollars [4190, 7854]. Dolls [4668]. Dome [7579]. Domestic [4385, 6891, 10022, 11540]. Done [3193]. Door [3119, 4320, 4855, 5692, 7839, 7560, 9626, 9627, 9404, 8553, 9688, 9697, 11567, 10826, 10914, 157, 629, 7342]. Door-Knobs [4320]. Doors [8049, 10686, 11625, 144]. Double [1479, 1456, 2843, 3387, 4232, 3769, 4084, 5713, 5086, 5502, 5127, 6240, 6647, 6576, 6220, 7477, 9282, 10087, 10339, 10056, 9934, 10826, 10959, 11184, 629, 258, 8237, 2983]. Double-Action [6220]. Double-Barreled [5502]. Double-Cylinder [3387, 6647]. Double-Deck [5713, 6240, 7477, 10087, 10339, 10959]. Double-Decked [5127]. Double-Decking [4232]. Double-Door [10826]. Double-Flow [11184]. Double-Moldboard [3769]. Double-Side-Door [629]. Double-Track [2843]. Double-Turret [1456]. Double-Turreted [4084]. Doubtful [11453]. Dover [5831, 912]. Dovetailed [11524]. Dovetailing [5234]. Down [1289, 1167, 11925, 10342, 3093]. Down-Town [1167]. Dozen [5429]. Dr. [1723, 3422, 4934, 5494, 6049, 6619, 10658, 11132, 709, 696, 9900, 5811, 6482, 1658]. Draft [3348, 6742, 9770, 11333, 11139, 5230]. Drafting [8749]. Drafts [11783]. Drag [1053]. Dragon [9941]. Drain [1175]. Drainage [1865, 2748, 5355, 671, 3465, 5019]. Drained [8452]. Drainer [9776]. Draught [2978]. Drawbacks [6353]. Drawbridge [1942, 9388, 8969, 9671]. Drawer [14, 178]. Drawing [3070, 4535, 6292, 7454, 468, 1029, 10651]. Drawings [1463, 1470, 3458, 5784]. Draws [10070]. Dreadnought [9649, 9309, 9091, 9416, 9441, 8569, 11657, 137, 990, 211, 658, 967, 11133]. Dreadnoughts [10493, 844, 982, 231]. Dredge [2083, 3469, 3088, 6590, 7982, 8052, 10008, 9206]. Dredger [6554, 7608, 11435, 922, 4706, 8084, 7321]. Dredges [2784, 8984, 10247, 1937]. Dredging [4559]. Dress [9465]. Dressing [9568]. Dried [8918, 11920, 10424]. Driers [1675]. Drill [11758, 4962, 6511, 6604, 1257, 115, 3666, 642]. Drilling [6025, 11613, 11786, 4142]. Drills [10180, 11785]. Drinking [1749]. Drive [2377, 3918, 9927, 9928, 9933, 10464, 9936, 9929, 11753]. Drive-Gate [3918].

Driven [4293, 5651, 4965, 5607, 6500, 6569, 7821, 7797, 8600, 10022, 9987, 11844, 11212, 10939, 11449, 53, 10265, 7360, 1105]. Driver [2667, 7838, 9823, 973, 11785]. Drivers [4438]. Driving [1366, 3201, 2755, 6060, 6358, 10760, 398, 2989, 7140]. Driving-Gear [1366]. Driving-Wheel [10760]. Drlven [6013]. Drop [8739, 1238]. Drove [11341]. Droz [5116]. Drugs [9286]. Drunken [258]. Dry [1530, 2244, 3599, 4022, 4973, 8286, 11448, 871, 11919, 686, 1395, 2103]. Dry-Blast [11448]. Dry-Grinding [8286]. Drydock [2896, 3362, 5380, 10774, 1145, 11440, 11224, 3721, 7542, 8532, 8550]. Drydocking [11607]. Drydocks [933]. Drying [5743, 8508, 4391]. Dubuque [6827]. Duck [10950]. Duck-Billed [10950]. Ducretet [2823, 3717]. Ductile [8923]. Due [7511, 7517, 714]. Duel [7045]. Dueling [9350]. Duell [2595]. Dufaux [8143]. Duke [1981, 1532, 4267]. Duluth [7189, 9671]. Dummy [115]. Dumont [3221, 3419, 3631, 3620, 3250, 3408, 3163, 4262, 4711, 5813, 5832, 7129, 7016, 8127, 9419, 9325, 10286, 10613, 10426, 11756, 1938, 9050]. Dumping [5382, 6186, 7683, 4060]. Duodecimal [4701]. Duplex [2546, 2554]. Duplicating [2680]. Dupligraph [5863]. Dupont [2396]. Durability [475]. Durable [5339]. Duration [8613]. Durban [6179]. During [5471, 5921, 9956, 1892, 4255, 6044, 10352]. Duryea [4296, 6646]. Dust [4575, 5473, 5186, 5621, 5044, 3960, 9029]. Dust-Arresting [5621]. Duster [6943]. Dweller [336]. Dwellers [7044]. Dwelling [1684, 2762]. Dwellings [6878, 1777]. Dye [9353]. Dynamite [4865, 4220, 8599, 107, 1259, 7909, 4462, 6958]. Dynamiting [257]. Dynamo [3760, 4166, 4583, 9527]. Dynamometer [9080, 10652]. Dynasty [7427].

Eads [1390]. Ear [165]. Ear-Trumpet [165]. Early [1831, 4084, 4019, 3849, 5749, 5071, 5908, 5998, 8079, 9561, 9920, 11482, 11722, 1048, 724, 1054, 727, 6766]. Ears [10177]. Earth [2268, 8303, 1768, 3717, 5600, 8504, 8720, 9617, 9579, 9263, 9958, 11285, 10971, 414]. Earthly [10065]. Earthquake [4062, 8922, 8716, 8898, 9143, 9644, 10151, 10057, 275, 58, 8763, 8836, 8835, 8887, 1586]. Earthquake-Proof [8716]. Earthquakes [8632, 4670, 510]. Earths [10122]. Easily [10611, 7912]. Easily-Constructed [7912]. East [4994, 9418, 10845, 9983, 1800, 1978, 1561, 3404, 2688, 3106, 3640, 3078, 2893, 3181, 3593, 4635, 3806, 4214, 4612, 4784, 4623, 4028, 6029, 4972, 6045, 5982, 5519, 5376, 6178, 6626, 6870, 7774, 8000, 8867, 9587, 9364, 11282, 10936, 11232, 171, 4008, 6198, 6757, 11462]. Easter [8689]. Eastern [5178, 2143, 5997]. Eat [10777, 2328, 9420]. Eaten [8636, 11262]. Eats [6425]. Eccentricities [6097]. Eclipse [1489, 1704, 1669, 3427, 5700, 8026, 9075, 705, 7755, 10007, 10672, 10654, 11086, 11702, 11842, 2658, 7326, 9485, 1680, 7827]. Eclipses [9604, 11436, 313, 992]. Economic [1744, 11058, 698, 212]. Economical [7857, 11112, 10284]. Economics [7577, 8678]. Economies [2455, 7977]. Economy [3032, 5759, 5637, 6456, 8258, 8673, 8584, 9720, 10937]. Ecuador [8547]. Ed [9774]. Edem [6074]. Edison [3764, 4347, 1443, 2334, 2837, 2895, 4736, 6000, 5015, 6093, 5475, 6568, 6391, 6665, 9012, 8879, 10522, 1296, 636]. Educating [7930]. Education [3653, 4098, 11521, 7939, 7911]. Educational [7271, 820]. Edward [7422, 3551, 3619, 8352, 10644]. Eels [8654]. Effect [2978, 4929, 6354, 6982, 8282, 7717, 9642, 387, 1278, 4464, 7825, 8043, 9897]. Effective [1539]. Effects [2494, 3199, 7380, 8674, 10658, 9957, 8919, 8836, 11014, 10882, 10472, 5715, 1701, 2783, 9428]. Efficiency [2257, 2860, 3037, 2771, 3035, 4485, 5302, 8848, 8615, 8759, 8832, 11355]. Efficient [1416, 1913, 1707, 2891, 2580, 3097, 5092, 6128, 7540, 7963, 3185, 4324]. Effigies [4642]. Egg [11412, 7835, 11843]. Eggs [3618, 5924, 5390, 8689, 9598, 538, 8981]. Egleston [1795]. Egypt [3263, 3932, 8191, 8825, 10631, 9815, 11390, 11542, 9774, 1681]. Egyptian [3589, 6897, 588, 4712, 10761]. Egyptianized [5132]. Ehrlich [8891]. Eiffel [6004, 6708, 883, 6245]. Eight [2990, 7183]. Eight-Mile [2990]. Eighteenpenny [5763]. Eighteenth [5574, 10089]. Eighth [6871, 10517]. Eighty [3248]. Eighty-Ton [3248]. Ejector [8600, 11687]. Ekenberg [11920]. Elastic [10475]. Elbert [8124]. Electric [1906, 2488, 5643, 61, 1943, 1977, 2303, 1341, 1389, 2268, 1657, 2131, 2311, 1719, 1574, 2153, 2304, 2029, 1557, 2378, 2469, 2333, 1626, 1368, 2444, 2215, 3204, 3398, 2865, 3337, 3312, 2941, 2610, 2644, 3328, 3388, 3111, 2796, 3288, 3436, 3684, 2596, 3240, 3569, 3688, 4341, 4087, 4124, 4437, 4471, 4501, 4159, 4338, 4762, 4081, 3862, 4868, 4297, 4906, 3929, 4433, 4193, 4774, 4189, 4107, 5616, 6157, 5931, 5759, 5742, 6026, 4941, 5255, 5565, 5900, 5785, 5773, 5969, 6051, 5795, 4962, 5619, 5930, 5302, 6105, 6588, 7229, 6972]. Electric [6800, 7041, 7154, 6394, 6641, 6433, 7104, 6734, 6359, 6401, 7111, 6667, 6998, 7249, 6604, 6665, 7633, 7786, 7676, 8492, 7674, 8295, 8146, 8468, 8481, 7624, 7440, 7469, 7523, 8457, 7750, 7377, 8205, 7756, 7919, 8164, 9664, 9366, 8676, 9541, 8699, 9068, 8805, 8742, 9121, 9561, 8616, 9344, 8784, 9046, 8852, 8652, 9466, 9345, 8833, 9126, 10377, 10022, 10741, 9782, 10473, 10302, 10491, 10557, 9767, 9877, 10685, 10714, 9894, 9768, 10228, 10150, 10132, 10684, 10257, 10561, 9836, 10276, 10450, 9895, 10264, 11607, 11796, 10937, 11232, 11089, 11946]. Electric [11241, 11329, 720, 263, 932, 605, 1177, 834, 37, 1092, 721, 1088, 187, 1141, 6852, 2213, 9289, 689, 10715, 10245, 10342, 8206, 4260, 5038, 9602, 8663, 9034, 6555, 8090, 1578, 3573, 1235, 2246, 2291, 2314, 4239, 6027, 6623, 7709, 5251, 1229, 2216, 3911, 5385, 4917, 6499, 7324, 198, 3294, 9567, 9623, 11221, 1219, 7476, 1701]. Electric-Acoustic [7786]. Electric-Car [2644]. Electrical [1697, 3159, 5014, 2083, 2408, 2080, 2994, 3249, 3587, 2659, 3223, 3446, 4597, 3726, 4465, 4364, 6008, 5180, 5240, 7186, 7418, 8049, 7885, 8351, 8251, 7544, 8816, 8721, 8637, 10157, 9705, 10122, 11650, 11914, 11234, 752, 1069, 413, 11192, 4877, 4900, 5115, 2440, 2687, 887, 5239, 5628, 7883, 3154, 4312, 1727, 7702, 9314, 11297, 3616, 2397, 940, 67, 1955, 10397]. Electrically [2639, 6483, 6500, 11266, 8386, 10265, 1140]. Electrically-Driven [10265]. Electrically-Operated [6483, 6500, 8386]. Electricity [1596, 3974, 5306, 5595, 6687, 7319, 6247, 6528, 6755, 6421, 6246, 7672, 7693, 7859, 8587, 8827, 8724, 10106, 11517, 11537, 11560, 11582, 11601, 11653, 11679, 11700, 11726, 11750, 11772, 11799, 10872, 25, 385, 411, 429, 455, 506, 528, 585, 608, 632, 675, 701, 718, 745, 825, 848, 885, 915, 937, 960, 987, 1013, 1043, 1119, 50, 1138, 1163, 1190, 1215, 1247, 1272, 1293, 84, 109, 139, 161, 189, 216, 235, 254, 278, 308, 338, 92, 890, 3788, 6810, 7403, 3251, 8455, 5163, 5628, 8814, 9400, 9619, 4083, 5019, 7140, 10427]. Electrification [5160, 10393, 10105, 11628, 1114]. Electrified [4896, 8256]. Electrifying [9067]. Electro [1919, 1823, 2285, 2332, 2156, 4316, 4141, 4145, 5623, 7836, 7913, 8980, 11630, 10152, 9146, 9589, 1779, 4735, 7912]. Electro-Analysis [7912]. Electro-Catalytic [4145]. Electro-Chronograph [7836]. Electro-Deposition [10152]. Electro-Engraving [2156]. Electro-Magnet [7913]. Electro-Magnetic [1823, 2285, 4316, 4735, 1779]. Electro-Mechanical [9146]. Electro-Plating [11630, 9589]. Electro-Pneumatic [1919, 2332]. Electro-Technical [8980]. Electro-Telethermometer [5623]. Electrochemical [6069, 6503, 1222]. Electroengraving [2196]. Electrograph [2892, 4604]. Electrolysis [9812]. Electrolytic [3007, 3742, 5572, 6315, 10731, 8968, 8477, 1260]. Electromagnet [4960]. Electromagnetic [5923, 1257]. Electrometer [4061]. Electrons [8634]. Electrophone [5926]. Electroscope [9601]. Electrostatic [2491, 7259]. Electrotyping [5852]. Element [2509, 5246, 10811]. Elementary [482]. Elements [4487]. Elephant [9449, 8413, 5630]. Elephants [10360]. Elevated [2313, 4159, 4596, 3959, 3780, 5664, 5900, 5540, 5302, 8047, 7372, 8041, 11408, 10900, 233, 3012, 1942, 2220]. Elevated-Underground [3012]. Elevating [1444, 5290, 1144]. Elevation [9640]. Elevator [4486, 4996, 5458, 6927, 10546, 11869, 708, 1021, 9214, 11221]. Elevators [1487, 1366, 3663, 5192, 6202, 11332, 5886]. Eleventh [3667]. Elgin [5134]. Elimination [8075, 8108, 699, 9230]. Elisha [3014]. Elizabeth [8003]. Elk [4552]. Elliott [4108]. Ellipse [1196]. Ellipses [3070, 10651]. Elliptical [4999]. Emanation [11392]. Emanations [6218, 7273, 6572, 7317]. Embankments [4185]. Embossing [6912, 9356]. Emden [3423]. Emerald [4339, 5067]. Emergency [1059]. Emigration [1625]. Eminent [3565]. Eminently [11391]. Emission [6756]. Emitted [6731]. Emperor [4064, 7478, 7467, 4287]. Empire [1224, 6445]. Employed [1598]. Employer [10621]. Employing [1866]. Employment [10780]. Emptying [873]. Enabling [5332]. Enactment [4957]. Encke [2887]. Encouragement [3083]. Encouraging [11510]. Encyclopedia [1213]. End [2390, 3400, 4506, 4130, 5459, 144]. End-of [2390]. Endowment [7320]. Endurance [3220, 4097, 3810, 4574, 3933, 5737, 5791, 6844, 10625, 10061, 11627]. Enemies [10275, 769, 11155]. Enemy [9181]. Energy [2193, 3159, 9117, 29, 559]. Engine [1726, 1535, 2439, 1579, 1988, 1439, 2152, 2309, 2209, 3351, 2622, 3662, 3336, 4528, 3888, 3833, 3952, 3754, 3965, 4512, 4761, 4343, 4033, 4410, 5582, 5361, 5479, 5307, 5162, 5256, 5087, 6808, 7084, 7164, 6264, 6843, 7857, 7936, 8425, 8406, 8404, 9544, 9351, 8544, 8615, 8687, 9513, 9122, 9123, 9967, 10469, 10769, 10458, 9922, 9925, 9923, 9926, 10456, 10481, 9958, 10927, 11627, 11840, 11341, 761, 119, 1048, 768, 333, 187, 689, 5115, 2221, 8985, 1017, 8043, 1051]. Engineer [5559, 11077, 11297]. Engineering [1952, 2366, 2436, 2169, 1984, 2590, 4137, 4123, 4823, 5566, 7599, 7752, 7956, 10669, 10057, 11516, 11536, 11559, 11581, 11600, 11652, 11678, 11699, 11725, 11749, 11798, 24, 384, 428, 454, 505, 527, 551, 584, 631, 653, 674, 700, 717, 744, 49, 770, 793, 824, 884, 914, 936, 959, 986, 1042, 1068, 1089, 1118, 1137, 1189, 1214, 1246, 1271, 1292, 83, 138, 160, 188, 215, 234, 277, 307, 337, 360, 6718, 7115, 8236, 4902]. Engineers [1798, 2994, 8099, 10767, 9977, 11913, 580, 628, 4312, 4198, 2233, 9696, 9896, 3154]. Engines [3664, 1848, 1419, 1582, 1493, 1391, 1395, 3480, 2757, 3505, 3619, 3764, 3904, 3937, 5410, 6138, 5441, 5501, 5142, 5580, 6945, 6331, 8111, 8100, 7942, 8059, 8509, 7471, 9203, 9749, 10789, 9921, 10454, 10094, 9895, 1019, 1001, 991, 1534, 10338, 3645, 6783, 5347, 7878, 813, 2075, 10242, 2909]. England [3131, 3059, 3544, 4704, 3907, 5034, 6241, 6219, 9615, 9144, 10607, 724, 5911]. Englehardt [7957]. English [1480, 2002, 3262, 3835, 4868, 4683, 5624, 5159, 6000, 5361, 5162, 5806, 9443, 9560, 8698, 10169, 9984, 10554, 11775, 10673, 6198, 6757, 6360, 6962, 1140]. Engraving [1963, 2156, 4342, 10859, 8831]. Enlargement [2957, 10560, 11274, 4371, 9547]. Enlargements [10004]. Enlarging [7408, 9735]. Enormous [5226, 10641, 10422, 10350, 11512, 1095, 141]. Entablatures [6484]. Entanglements [7731]. Entering [4729]. Enterprise [3837]. Entertainment [3264, 11718]. Entirely [4058]. Entomology [1431]. Entrance [2460, 1485, 6307, 7359]. Entrances [10603]. Entries [7478]. Envelop [2079, 5023]. Envelope [4513]. Envelope-Sealing [4513]. Ephemeræformis [2830]. Epitome [3563]. Epoch [3458]. Equable [6681]. Equatorial [3744, 9229]. Equipment [1914, 2408, 2313, 2882, 4501, 7885, 8699, 8721, 9121, 8788, 10450, 3525, 9883, 1166, 7912, 9623]. Equivalent [8501]. Era [1787, 3754, 11675, 2009]. Eraser [7990]. Erasers [2694]. Erecting [7039, 11864, 11012]. Erection [3834, 5921, 9227, 10738, 10301, 11282, 797, 3643]. Erie [9045, 5595, 6421, 10450, 10220, 1814, 2484, 1429, 1830, 2678, 2632, 4546, 3886, 5406, 8070, 7393, 9721, 11493, 9547]. Ermack [6395]. Eros [2569, 2797]. Erosion [3082, 7355, 8282, 8250, 9182, 9225, 10121, 9814, 10296, 11778, 3140, 11114]. Errors [6994]. Eruption [4046, 4853]. Eruptions [3995, 606, 6078]. Escape [7963, 7837, 3720, 6366]. Eskimo [6983]. Especially [6756]. Essay [1143, 1168]. Essays [7856, 159]. Essen [3868]. Establishing [1841]. Establishment [9568]. Estates [7157]. Estimated [10056]. Estimates [3303]. Estimation [3110]. Etc [1311, 1326, 2067, 2088, 2113, 2140, 2164, 2185, 2227, 2250, 2272, 2293, 2319, 2343, 2361, 2384, 2429, 2450, 2480, 2503, 1354, 1378, 1401, 1473, 1500, 1524, 1567, 1589, 1617, 1639, 1665, 1688, 1713, 1739, 1756, 1782, 1809, 1835, 1856, 1878, 1899, 1946, 1996, 2039, 2561, 2586, 3415, 3453, 3498, 3536, 3557, 3581, 3604, 3628, 3649, 2607, 2696, 2745, 2769, 2788, 2855, 2902, 2925, 2949, 2975, 3000, 3026, 3050, 3074, 3102, 3125, 3146, 3167, 3236, 3256, 3279, 3298, 3346, 3370, 3731, 3751, 4675, 4770]. Etc [4795, 4820, 4840, 4862, 4888, 3778, 3876, 3899, 3993, 4015, 4042, 4181, 4207, 4226, 4250, 4271, 4332, 4359, 4402, 4429, 4449, 4498, 4521, 4545, 4595, 4633, 4657, 4988, 5011, 5874, 5896, 5917, 5942, 5965, 5993, 6039, 6065, 6086, 6111, 6132, 5060, 5124, 5153, 5173, 5200, 5220, 5298, 5319, 5401, 5424, 5450, 5466, 5491, 5536, 5588, 5613, 5660, 5687, 5710, 5729, 5756, 5779, 5820, 5844, 4949, 4981, 5244, 5784, 6173, 6192, 7052, 7075, 7125, 7149, 7177, 7201, 7246, 7289, 7313, 6232, 6254, 6284, 6303]. Etc [6349, 6387, 6452, 6472, 6492, 6524, 6596, 6632, 6702, 6725, 6771, 6794, 6818, 6840, 6864, 6902, 6924, 6952, 6990, 7032, 6354, 7368, 7390, 8267, 8330, 8346, 8369, 8419, 8443, 8514, 7413, 7435, 7464, 7485, 7573, 7593, 7618, 7642, 7689, 7714, 7743, 7779, 7819, 7845, 7872, 7926, 7949, 7972, 8013, 8036, 8067, 8156, 8199, 8222, 8009, 8537, 8560, 9498, 9518, 9536, 9612, 9661, 8579, 8606, 8629, 8713, 8730, 8756, 8778, 8844, 8864, 8884, 8929, 8974, 8999, 9021, 9061, 9086, 9199, 9278, 9300, 9318, 9362, 9385]. Etc [9410, 9437, 9716, 9739, 10618, 10700, 10726, 10744, 10764, 10786, 9809, 9830, 9850, 9890, 9971, 9997, 10015, 10032, 10053, 10074, 10099, 10118, 10144, 10166, 10185, 10289, 10309, 10347, 10370, 10410, 10487, 10512, 10571, 10600, 10804, 10841, 11672, 11695, 11744, 11766, 11792, 11853, 11872, 11892, 11910, 11929, 10864, 10892, 10954, 11007, 11026, 11050, 11071, 11099, 11124, 11207, 11228, 11247, 11290, 11338, 11382, 11422, 11444, 11488, 11508, 11531, 11553, 11594, 11622, 11648, 4, 42, 1036, 1061, 1109, 1133, 1156, 1207, 1265, 1286, 75, 101, 208]. Etc [247, 272, 300, 377, 404, 449, 545, 576, 625, 647, 667, 695, 786, 817, 840, 877, 908, 929, 953, 979, 1008, 204, 496]. Etching [4261]. Ether [3703]. Ether-Vapor [3703]. Ethics [11721]. Etruscan [1710]. Eucalypts [5668]. Eugene [6050]. Europe [2099, 2516, 2175, 2299, 2124, 2796, 4824, 4933, 7337, 8701, 9510]. European [2025, 4732, 5410, 9541, 8922, 11685, 11236, 758, 1117, 11325, 5697]. Evans [11166]. Evaporation [9793]. Evaporative [1140]. Evening [8592, 10199, 11327, 30]. Ever [2470, 6807, 8380, 10759, 9843, 10108, 312, 10908]. Ever-Shifting [10908]. Every [8428]. Evidence [4602]. Evil [995]. Evolution [8430, 9092, 10314, 821, 5336, 11524, 11903, 8084, 9093, 8121, 3510, 260]. Exact [5618]. Examination [1749, 3618, 6218, 9541, 10471, 3960]. Examining [9470]. Example [3293, 4485]. Examples [6879]. Excavated [3129]. Excavation [8875, 8984, 11887, 11251, 11365]. Excavations [1815, 1370, 1829, 3894, 6296, 7250, 8210, 11432, 1279, 4389, 9774, 3116, 316, 393]. Excavator [3890, 8325]. Excellent [4404, 8470, 10001]. Exchange [4466]. Excursion [3313]. Exercise [6249]. Exerciser [2029, 5054]. Exercises [2822]. Exercising [7143]. Exhaust [6945, 8665, 1087]. Exhauster [3823]. Exhaustion [8614]. Exhibit [7764, 9011, 1962, 1434, 7139, 7345]. Exhibited [1334, 2350, 11108]. Exhibition [2394, 2594, 4750, 9487, 10965, 944, 1096, 1529, 10671, 2153, 3529, 3399, 4144, 6682, 11660, 11632, 2062]. Exhibitions [6478]. Exhibitor [6104]. Exhibitors [1999]. Exhibits [2159, 2085, 2057, 3385, 3409, 3428, 6941, 7292]. Existence [2792, 7358, 1076]. Existing [10036]. Exit [144]. Exits [10603]. Expander [2987]. Expanding [1178]. Expansion [5643, 1745, 4021, 4483, 8404, 7607, 439, 3838]. Expected [2887]. Expediting [9008]. Expedition [2727, 3427, 5325, 5880, 7827, 4643, 10612, 1981, 1905, 4235, 4702, 4969, 5043, 5735, 6268, 6295, 6457, 8252, 7906, 8490, 8566, 9148, 10037, 10086, 11132, 11597, 1053, 135, 5189, 7755, 4943, 11936, 2658]. Experiences [10334]. Experiment [1383, 1720, 2367, 4666, 6013, 11757, 11134, 11346, 196, 9423, 9734, 95, 10427, 5048, 11305]. Experimental [1722, 1676, 2844, 4201, 10300, 11409, 11541, 852, 431, 6810, 9093, 8746, 970, 4583, 8129, 4671, 9527]. Experimentally [5083]. Experiments [2049, 1964, 1954, 2491, 1421, 2446, 2571, 2849, 2536, 4316, 3938, 3717, 3696, 4558, 4893, 4266, 4149, 4846, 4803, 5165, 5651, 5042, 6141, 5667, 6354, 6394, 6464, 6932, 7273, 7156, 6217, 7373, 9510, 9442, 8869, 10244, 9820, 11107, 10830, 11802, 11105, 11083, 416, 7281, 10609, 7259, 6929, 4468, 6245, 2913, 5367, 10420, 11638, 1202, 10563, 10903, 1249, 1165, 10908, 219, 866, 4690]. Expiration [1443]. Explained [7371, 656, 942, 10879]. Explanation [4005, 7723, 10869, 95, 10592]. Exploits [7986]. Exploration [3589, 747, 635]. Explorations [2171, 5408, 10998]. Exploring [7960]. Explosion [1724, 1957, 3226, 2663, 2757, 4220, 4553, 4211, 4410, 5470, 5142, 7905, 8249, 8941, 10256, 10094, 193, 918]. Explosion-Engines [2757, 5142]. Explosions [3763, 2783]. Explosive [3486, 3442, 3435, 6288, 8111, 29, 1320]. Explosives [3808, 6713, 7645, 10656, 11639]. Export [4231]. Exported [8077]. Exports [2005, 2325, 1436, 1695, 3680, 5226, 7824, 8826, 9390, 10751]. Exposition [1793, 2052, 1962, 2888, 10156, 11321, 9746, 1383, 2515, 1510, 2085, 2289, 2234, 1485, 2080, 2128, 2054, 2011, 2033, 2522, 1505, 2010, 2102, 1955, 2521, 1988, 1984, 1461, 1508, 1531, 1554, 1649, 1698, 1936, 2194, 2372, 2490, 1656, 2442, 2473, 1434, 2183, 2350, 2057, 1751, 2309, 1702, 2125, 3591, 3388, 3046, 3090, 2729, 2839, 2870, 3114, 6055, 7212, 7044, 7271, 6809, 6259, 7238, 7252, 6855, 6878, 7016, 7001, 8225, 8340, 9984, 9854, 11878, 2179, 2245, 2182, 9861, 2424, 195, 4709, 10025, 7563, 6398]. Express [2575, 4891, 4411, 5361, 5114, 5806, 6468, 6808, 8031, 10108, 10959, 11450, 983, 6575, 2989, 7254, 10174, 11839]. Extemporized [10846]. Extended [11293]. Extension [2070, 4942, 5540, 6258, 6869, 8635, 8660, 10446, 614, 1200, 2655]. Extensions [8377]. Extensive [3263, 6732, 11628]. Exterminating [6619]. Extermination [5999, 5130, 7018, 8451, 8703, 11349, 10942, 3155]. Exterminator [6158]. Extinct [4050, 4151]. Extinguisher [4374, 7636]. Extracting [11115]. Extraction [7206, 10752]. Extraordinary [1862, 2931, 4932, 8282, 9414, 10004, 4106, 5539]. Eyde [9253]. Eye [2289, 5638, 6556, 7220, 7749, 8271, 863, 392, 9531, 7165]. Eye-Bar [5638]. Eye-Shade [7220]. Eye-Spotting [9531]. Eye-Strain [7749]. Eyeball [10908]. Eyeglasses [1127]. Eyes [2755, 11539, 192]. Ezekiel [3265].

F.R.S. [6098, 6010]. Fabled [10354]. Fabric [310]. Fabrics [3657, 7556, 11234, 11111, 6691, 9602]. Fac [4604, 10219]. Fac-Simile [4604, 10219]. Facade [218]. Facilitating [9293, 5790]. Facilities [4826, 4942, 10710]. Facing [8397]. Facsimile [2892, 5570, 6482]. Fact [650]. Factor [2841]. Factory [1867, 7082, 6713, 9884, 11684, 150, 7878, 6814]. Facts [1976, 6391, 8890, 10905, 11587, 11777, 11795, 592, 10125, 5391, 10133, 1362, 11454, 10964]. Faculties [11439]. Fades [797]. Failed [10400]. Fails [3221, 3820, 7984]. Failure [2282, 1518, 3600, 5157, 5741, 5359, 8489, 10396, 10844, 7902]. Failures [6020]. Fair [2063, 6433, 6459, 6894, 1563, 6075, 6138, 6001, 7298, 7128, 6983, 7292, 6912, 7087, 7277, 7000, 7325, 7139, 7336, 7261, 7282, 7345, 8159, 7342, 5693]. Fakes [11454]. Fall [11519]. Fallacy [1743, 11599]. Faller [4466]. Falling [10158]. Falls [4454, 4659, 5044, 8410, 11210, 239, 4901, 4667, 5627, 4150, 2664, 6786, 8521, 1706, 8146, 8468, 8481, 7544, 7862, 7488, 8616, 8697, 10110, 9248, 31]. Famine [9025, 11596]. Famous [1965, 5066, 5617, 7130, 7915, 3011, 4579, 11286]. Fan [2082, 6647]. Fan-Cooled [6647]. Fancy [650]. Fans [1943]. Far [9594, 1290, 8000, 6198, 6757, 7304]. Far-Sighted [1290]. Fares [7534]. Farm [5812, 5750, 5417, 11902]. Farman [11194, 11253, 11802, 11300, 11876, 968, 276]. Farmer [11220]. Farmers [1413]. Farming [6569, 7118, 11919, 3838, 2988]. Farmobile [1171]. Farms [8541, 5286, 5628]. Farquhar [2122]. Farragut [1801]. Farthest [9392]. Fashions [11796]. Fast [4679, 7718, 7955, 11807]. Fastener [1365, 8151, 7568, 9627, 9866, 782]. Fastening [4587, 5681, 5343, 5215, 5145, 5550, 6272, 6586]. Fastest [3568, 2908, 3926, 3713, 4337, 4392, 7490, 8400, 10769, 10197, 10856, 10810, 11275, 11385]. Fastnet [10361]. Fatal [5101]. Fatalities [6927, 6475, 7055]. Fated [10343]. Father [1909, 10792]. Faucet [2938]. Faucets [807]. Faulty [4185, 3782, 436]. Favor [628]. Favorable [2683]. Fayum [11542]. Feat [2425, 1351, 3382, 3195, 4823, 4578, 5335, 6319, 6718, 8236]. Feathered [8207]. Feathering [4195]. Feathers [2573, 4847]. Feats [4902]. Features [2826, 4277, 9935, 2715, 8763, 3868]. February [10852, 2996, 4178, 5446, 6674, 8995, 10162, 11267, 520, 7776]. Federal [1796, 8120]. Feed [5987, 7000, 9130, 9546, 8839]. Feed-Water [5987]. Feeder [2713]. Feeders [6161]. Feeding [3306, 4169, 4812, 7770, 342]. Feeling [8720, 11396]. Feet [4578, 5880, 9167, 11015, 11371, 7183, 7068]. Felling [1699, 7186, 9884, 8283, 462]. Felt [8085]. Fender [1655, 2891, 5283, 7918, 10363, 11483, 11459, 11513]. Ferber [9820]. Ferment [11412]. Ferries [10554, 2968]. Ferris [6832]. Ferrocyanide [1386]. Ferry [61, 1444, 1144, 5259, 10220, 7702]. Ferryboat [2932, 4843, 7935]. Fertilizer [3224]. Fertilizer-Distributer [3224]. Fery [6960]. Fessenden [3929, 4870, 4916]. Festina [8293]. Festival [6809, 767, 8337]. Fever [2612, 5641, 5176, 8451]. Few [3383, 3406, 11476, 11454]. Fewer [5223, 11316, 408]. Fiala [8490]. Fiasco [1693, 4865, 4233, 5878, 6798, 11556]. Fiber [1533, 8986]. Fiber-Cleaning [1533]. Field [1596, 2491, 1345, 3862, 5804, 90, 10109, 3525, 6311, 3814]. Fields [1552, 5407, 8500, 8338]. Fiestas [10629]. Fife [10708]. Fifteen [2668]. Fifth [2739, 6498, 6781, 7785, 11841, 957]. Fifth-Wheel [2739]. Fiftieth [5301, 5310]. Fifty [3252, 6998, 9567, 4101]. Fifty-Ton [4101]. Fight [10315]. Fighting [1348, 3035, 4387, 6762, 9550, 9311, 11659, 11752, 11705, 2261]. Figured [6628]. Figurehead [1225]. Figures [2754, 4932, 895]. Fiji [4126]. Fijians [4126]. Filament [10150, 868]. File [5436, 568]. Filing [1261]. Filipinos [2960]. Fill [11563]. Filling [9348, 10475, 371]. Filter [3274, 5987, 10566, 9758, 11117, 972]. Filtered [660]. Filtering [558]. Filtration [2324, 1392, 3562, 5176, 8017, 8118, 164, 285, 8433, 4961]. Final [5031, 6001, 9002, 9053, 11152]. Finality [1503]. Finally [2231]. Finback [4978]. Finder [2569, 5210, 7294]. Finders [11085]. Finding [7565, 5163, 1153, 11788]. Fine [6552]. Finest [11376]. Finger [540, 751]. Finger-Developing [540]. Finish [7703]. Finished [5869, 10845, 7651]. Finishing [9647]. Finsen [5911]. Fire [2131, 2190, 1828, 1747, 3270, 2532, 3635, 4374, 4528, 4192, 3785, 3805, 4935, 4929, 5410, 6137, 5599, 5836, 6588, 6414, 6982, 6306, 6730, 7084, 6237, 6762, 6753, 6257, 7081, 7963, 7636, 7837, 7515, 8450, 9192, 8912, 9311, 10350, 11063, 11586, 10897, 11085, 11131, 11636, 570, 582, 913, 409, 889, 187, 7192, 8412, 5115, 6691, 3720, 3666, 3645, 6783, 6366, 2261, 8836, 9422, 6620, 6962, 11655, 11656, 4612, 4623, 3735, 7318, 6736, 9616, 8958, 11273, 9306, 6224]. Fire-Alarm [2532, 409]. Fire-Control [11586]. Fire-Engines [6783]. Fire-Fighting [9311]. Fire-Ladder [1747]. Fire-Plug [570]. Fire-Proof [2190]. Fire-Tube [3635]. Firearms [7567, 7429]. Fireballs [752]. Firebox [3447, 5702]. Fireless [8006, 11017, 65, 67]. Fireman [7234]. Firemen [1057, 8030]. Fireproof [2276, 2374, 3275, 4018, 4441, 4774, 5900, 6782, 8426, 8761, 9745, 6691]. Fireproofing [2389, 8899]. Fires [5036, 7020, 10002, 11512, 11746, 5212, 11062]. Firing [2777, 10705, 1122]. Firms [6636]. First [1340, 1966, 2288, 2518, 3405, 3551, 3481, 2754, 3482, 2953, 4829, 4488, 4691, 6028, 6008, 5737, 5229, 5285, 5761, 5494, 5211, 6532, 6963, 6847, 6679, 6207, 8463, 8423, 7650, 8050, 8162, 8188, 7474, 9562, 9308, 9445, 9479, 9264, 10580, 10581, 10582, 10583, 10595, 10596, 10687, 11660, 11012, 11211, 11134, 11434, 11816, 11829, 10876, 11641, 10943, 11657, 11346, 434, 1193, 1298, 92, 1275, 1198, 850, 828, 1218, 673, 658, 158, 10382, 1725, 9774, 5362, 4409, 3954, 4753, 9353, 10378, 3479, 509, 944]. First-Class [1966, 2288, 3405, 3481, 3482, 10580, 10581, 10582, 10583, 10595, 10596, 3479]. Fish [1574, 7195, 11642, 11357, 11377, 3512, 6101, 9290, 10268]. Fish-Line [11357]. Fisheries [8323, 5771]. Fishery [221]. Fishes [7943, 4810, 1634, 4598, 7547]. Fishing [4802, 10851, 11040, 569]. Fishskin [4008]. Fiske [3380]. Fitch [757, 789, 856]. Fits [4122]. Fitting [8940]. Fittings [11687]. Five [3182, 2889, 4524, 5732, 8432, 7586, 8164, 10582, 10936, 8691]. Five-Horned [3182]. Five-Wheeled [7586]. Fixation [10399]. Flag [9562, 9778, 11223]. Flags [8964]. Flagship [1801, 7327, 8040, 10575, 7705]. Flame [8767, 8567, 10279]. Flaming [9570]. Flamingo [7239]. Flange [9690]. Flap [11461]. Flapping [1128]. Flarebacks [11557, 11055]. Flash [3466, 10978]. Flashes [8613]. Flashing [7637]. Flashless [9289]. Flasks [198]. Flax [8319, 10752]. Fleet [3738, 8255, 3153, 6867, 7005, 7258, 7822, 7385, 9597, 8677, 9670, 10593, 10709, 9860, 10575, 616]. Fleets [3880, 5620, 9346, 613, 6198, 6757]. Fleming [2579]. Flexibility [3795]. Flexible [2056, 2779]. Flies [5249]. Flight [6997, 8188, 9578, 9419, 9325, 10313, 10991, 11434, 11641, 11300, 11046, 11830, 851, 1210, 1254, 883, 1193, 184, 1172, 1275, 944, 1128, 1218, 11848, 4940, 5718, 6376, 8336]. Flights [9979, 11194, 11134, 11829, 1198, 850, 1217, 968, 158, 823, 214]. Flip [11461]. Flip-Flap [11461]. Float [9546, 11357, 494]. Float-Feed [9546]. Floating [1530, 1492, 3248, 3596, 2896, 5645, 5909, 5837, 6179, 6682, 8225, 8675, 8984, 142, 4831, 6419, 3721, 5697, 3816, 8983]. Floatless [10479]. Floes [9816]. Flood [2418, 4644, 3553]. Flood-Gate [2418]. Flooded [5768]. Floods [5760]. Floor [2893, 4972, 1289]. Flooring [7782]. Floral [6259]. Florida [6717, 7834, 7863, 7823, 9035, 11217, 1295, 9983]. Flotilla [8677, 11632]. Flour [6763]. Flow [11184, 681]. Flower [11716]. Flowering [5766]. Flowers [6148, 9727, 11715, 326]. Flowing [539]. Flue [5581]. Fluid [8893, 10058, 364]. Fluids [4201]. Fluorides [1722]. Fluorine [5572]. Flush [122]. Fly [1556, 6512, 232, 559]. Fly-Catcher [6512]. Fly-Screen [1556]. Flying [1780, 2736, 2866, 3015, 3598, 3724, 7727, 8143, 8620, 10335, 10299, 10794, 9984, 9857, 10321, 11198, 1041, 104, 113, 276, 6835, 11370, 10401, 7025, 9753, 8207, 10420, 9050, 11367, 349]. Flying-Machine [3598, 7025]. Flywheel [2157, 11076]. Foaming [10161, 7825]. Fog [3229, 3570, 6074, 7859, 8874, 720, 4468, 5654, 6810]. Fog-Signaling [8874, 720]. Foggy [11555]. Fogs [6687]. Folding [2100, 2581, 3315, 4057, 5860, 5705, 5289, 5210, 7965, 9378, 10445, 10982, 11095, 11092, 8213, 11710, 1228]. Follies [11164]. Fomentations [3020]. Food [5165, 5716, 7207, 7375, 11599, 284, 8915, 9956, 10424]. Foods [6930, 6908, 7021, 9163, 9393, 9285, 1149, 2960]. Foodstuffs [11748]. Fool [3247]. Foolhardy [3230]. Foolishness [1761]. Foot [1492, 1799, 1700, 2784, 2668, 4169, 6162, 7490, 10079, 10301, 11353, 9676, 11549, 4245, 11940]. Foot-Power [1799, 11549]. Footways [3181, 11864]. Footwear [5053]. Forage [3267]. Forbidden [10612]. Force [2230, 3387, 7271, 11579, 8919]. Force-Pump [3387]. Forced [2978, 3348, 11139, 4122]. Forces [8145, 10274]. Forcing [6730]. Ford [6662]. Forecast [1596, 9260]. Forecaster [2120]. Foreign [1845, 1750, 2885, 4021, 5072, 6138, 5561, 7292, 7764, 7792, 9562, 8620, 9270, 11170, 11922, 11215, 11943, 10647]. Forerunners [8495]. Foreshore [9066]. Forest [1730, 5378, 7020, 7230, 7850, 10640, 11512, 11746, 10989, 7588, 11114, 11656, 4851, 6603, 432, 2445]. Forestry [2818, 11037, 900, 1256, 10999]. Forests [11521, 4199]. Foretold [10173].

Forge [2574]. Forger [4778]. Forging [97]. Forgotten [10317, 5885]. Fork [11118]. Forks [1023]. Forlorn [6867]. Form [1655, 2469, 1558, 2239, 2152, 2218, 3069, 4061, 4801, 4109, 6012, 7889, 7887, 8518, 8991, 8586, 9512, 9374, 9525, 9908, 9846, 9931, 904, 11681, 2913, 4757, 2421, 2462, 1533, 1391, 1488, 1732, 3112, 7111, 6502, 8706, 10462, 3621]. Formation [2305, 1722, 4906, 6334, 3984, 4789, 7024]. Formed [4977, 7377]. Former [11077]. Forming [2195]. Forms [1889, 7340, 6811, 86, 3264]. Formula [10395]. Formulæ [194]. Forns [5623]. Forres [2243]. Fort [1705, 11434, 1255, 1217]. Forth [6607, 6094]. Forth-Coming [6607]. Forthcoming [4257]. Fortifications [611]. Fortis [2029]. Fortress [9805]. Forty [9584, 10664, 10105, 2967, 3064]. Forty-Four-Mile [10105]. Forty-Fourth [2967]. Forum [3894, 6296]. Fossil [1575, 3136, 6081, 9882, 11542, 4151]. Fossils [7960]. Foster [3229]. Foucault [4558, 11305]. Found [6930, 8210]. Foundations [4974, 10001, 10282, 10189, 8900, 7628]. Founder [4785]. Founding [5850]. Foundry [8643]. Fountain [2080, 2147, 2827, 3668, 6103, 7454, 8409]. Fountains [2639]. Four [1872, 4286, 7292, 6264, 7801, 7797, 8101, 7790, 8791, 8789, 8795, 8797, 10414, 9791, 10105, 9949, 11514, 748, 7600]. Four-Cycle [6264, 8797]. Four-Cylinder [1872, 7801, 7790, 8791, 8795, 9949]. Four-Day [10414, 9791]. Four-Passenger [4286, 8789]. Four-Thousand-Dollar [7600]. Four-Track [11514]. Fourteen [981]. Fourteen-Inch [981]. Fourth [3064, 10713, 11757, 427, 106, 159, 656, 942, 1143, 1168, 10879, 2967, 6638, 11603]. Fourth-Dimension [427]. Fowls [4779]. Fox [1481]. Foxton [6797]. Frailty [4212]. Frame [1774, 3795, 4648, 6060, 10137]. Frames [4090, 8300]. Framework [4990]. Framing [2634, 8215]. France [2371, 1872, 7166, 9540, 9244, 9398, 9624, 9442, 9186, 11107, 11945, 1254, 549, 5286, 9464, 9902, 9746]. Francis [1920, 11168]. Francisco [1600, 1884, 1396, 9616, 8958, 8887, 8770, 7449, 9306, 4973, 8899, 8722, 10057, 8836]. Frank [1486]. Franklin [6649, 8812, 8725, 9005]. Franz [2727]. Frauds [10828, 10898, 532]. Frayer [8792]. Freak [6033]. Frederic [8275]. Free [7035, 6276, 7931, 9588]. Free-Alcohol [7931]. Freeboard [10921]. Freed [8870]. Freezing [4664, 4078, 5646, 9364, 11463]. Freight [1419, 2311, 3661, 4086, 6106, 7384, 7645, 11740, 11514, 1080]. Freighters [4678]. French [2155, 2370, 2077, 1457, 1447, 1412, 1393, 2537, 3405, 3315, 3198, 3745, 4574, 3837, 4739, 6052, 5361, 4976, 7399, 7597, 9461, 9347, 9603, 10709, 10525, 10605, 9771, 10452, 10633, 10413, 10124, 10238, 10830, 11557, 11460, 11828, 1193, 388, 464, 261, 8235, 8641, 8942, 8547, 369, 7237, 8621, 11090, 3845]. Frequency [10780, 5261]. Fresh [10226, 11527, 6379, 10877]. Fresh-Air [10226]. Fresh-Water [6379, 10877]. Friction [4036, 3888, 9927, 10464, 9929, 903, 48, 8411]. Friction-Drive [9927, 10464, 9929]. Friedrich [4661]. Friend [8551]. Frog [8211]. Frogs [3511]. Front [6663, 8563]. Frost [2756, 6686]. Frozen [6246]. Fruit [3457, 3468, 4549, 5716, 6944, 7375, 10365, 11308, 11349, 605, 10315, 869].

Fruit-Trees [3468]. Fruits [10424, 10671]. Frying [10982, 175]. Fuel [2073, 2437, 2302, 2847, 4465, 4432, 4547, 3997, 4892, 5307, 5158, 6709, 6550, 6823, 7497, 8759, 9595, 8832, 10770, 11387, 10938, 5476, 5668, 9215, 8966]. Fuels [10193, 11859]. Fulfillment [9260]. Full [9877, 11933, 136]. Full-Sized [11933, 136]. Fulton [3503, 10375, 820, 1211, 213, 767, 2912, 10792, 184, 749, 155]. Fumes [6517]. Fun [8274]. Fund [3589]. Fungicide [344]. Fungus [11781]. Funnels [46]. Furnace [1342, 2303, 1719, 3249, 4906, 6978, 7377, 10515, 10228, 71, 5138, 8571, 11783]. Furnaces [1415, 1891, 2596, 3723]. Furnishing [11787, 66, 640]. Furs [10305]. Further [3420, 4454, 4846, 5042, 293, 11631]. Fuse [5144, 10048, 11667]. Fusible [27]. Fusion [3998, 9068]. Futile [334]. Future [2167, 1718, 1484, 1382, 3564, 3585, 2700, 4891, 3849, 4826, 6022, 5948, 5690, 5134, 6928, 6258, 7647, 8118, 8425, 7901, 7748, 8781, 10574, 10639, 9835, 11170, 11451, 11847, 56, 932, 800, 1097, 8109, 8460, 10201, 10221].

Gabet [775]. Gage [2380, 5265, 6694, 6456, 7217, 6583, 8324, 7457, 9491, 8905, 10079, 10819, 804, 1152, 400]. Gages [2308]. Galapagos [3689]. Gale [1268]. Galleries [2815, 4613, 5691, 8977]. Gallery [11541, 7398]. Galleys [7141, 9372]. Gallon [8832]. Galvanic [2045]. Galvanizing [9668]. Galveston [2022, 2120, 9565, 8530]. Game [2434, 4728, 5191, 5031, 8436, 5913, 5206, 5496]. Games [3264]. Gammeter [10401]. Ganges [11479]. Ganz [2941, 2610, 5785]. Garbage [1929, 5877, 10021, 5793]. Garden [1853, 2222, 2051, 2661, 8351, 664, 11712, 2993, 10517, 10397, 10647, 9919, 525, 10263]. Gardens [2124, 3307, 7157, 7952, 1129]. Garment [7611, 38]. Garments [11644, 11158, 4008, 5774]. Gary [1076, 8124]. Gas [2073, 1697, 4024, 11429, 1724, 2373, 2128, 2058, 2510, 3351, 3437, 2967, 3064, 2803, 3812, 4192, 3888, 4416, 3763, 4509, 6138, 5264, 7102, 6890, 7000, 6843, 7936, 8425, 8050, 7821, 8390, 7868, 8194, 9250, 8588, 8544, 8615, 8685, 9687, 9720, 9782, 10262, 11503, 11844, 11212, 10927, 10907, 10872, 11139, 11809, 11840, 10939, 11449, 11859, 11358, 11329, 358, 53, 1236, 9503, 8235, 11588, 8356, 7809, 1235, 814, 5347, 8985, 10406, 7878, 813, 1990, 8043, 11039, 10242, 2909]. Gas-Driven [7821, 11844, 11449, 53]. Gas-Electric [11329]. Gas-Engine [3888, 8985]. Gas-Generating [2128]. Gas-Lighter [9782]. Gas-Producer [8425]. Gas-Recoil-Check [2373]. Gas-Saving [8390]. Gas-Tube [1236]. Gaseous [9619]. Gases [2173, 3157, 6054]. Gasoline [2550, 2992, 4296, 4285, 4288, 4619, 4140, 4295, 3887, 4369, 4165, 5079, 5075, 5086, 5069, 5441, 5501, 5751, 5074, 5695, 5077, 6666, 7084, 6464, 6265, 6264, 6646, 6643, 6220, 8525, 7794, 7797, 8318, 7805, 8231, 7790, 9203, 8595, 8830, 9402, 9546, 9564, 8852, 10193, 10496, 9770, 10476, 11944, 10930, 497, 5347, 1017]. Gasoline-Motor-Propelled [7084]. Gasoline-Propelled [5695, 9402, 11944]. Gate [2418, 3918, 9172]. Gates [3266]. Gathering [2498, 11118, 1123]. Gathmann [3928, 2807]. Gatling [4934]. Gatun [8565, 10639, 10001, 10056, 10681, 11795, 11074, 1289, 524]. Gear [1535, 1870, 1366, 3822, 4259, 6060, 5240, 5021, 5607, 6659, 6813, 9008, 9940, 9952, 9943, 9934, 10462, 492, 1305]. Gear-Changer [3822]. Gearless [7354, 10460]. Gelatin [8375, 11195]. Gelatine [4959]. Gem [5605, 9162]. Generated [11139]. Generating [2128, 8851, 2424, 10995]. Generation [9705]. Generations [11847]. Generator [1983, 1579, 1850, 3703, 9228, 9051]. Generators [6569, 2075]. Genesis [9479, 956]. Geodetic [8547]. Geographic [7297, 6871]. Geological [6117, 2773, 4162, 1577]. Geology [4789]. Geometrical [10070]. Geometrician [3869]. Geometry [10881]. George [1460, 8737, 4894, 7441]. Georgia [1682, 3482, 2884, 5044, 6814, 10682, 10582, 1941]. Geraes [11752]. German [2198, 1417, 9507, 1430, 2002, 2102, 3423, 3142, 2934, 3363, 4733, 5280, 5252, 5031, 5206, 5496, 5383, 5764, 6378, 7065, 6457, 7478, 7467, 10334, 11348, 11628, 11196, 11886, 11828, 11438, 1039, 850, 123, 682, 982, 314, 231, 8144, 9054, 10503, 10265, 9753, 10998, 7428, 5333, 10798, 2989, 4221, 7324, 11399]. Germania [11438]. Germany [1549, 1409, 2466, 3199, 2885, 3676, 4231, 4728, 4098, 5105, 11017, 11033, 158, 23, 6620, 6339, 900]. Get [773]. Getting [495, 488]. Geyser [8409, 7193, 7240]. Gherardi [5974]. Giant [3508, 5929, 7772, 7982, 9033, 11730]. Giantess [7500]. Giants [8148, 8523]. Gibier [1723]. Gifted [11391]. Gigantic [11178, 9863, 5364]. Gila [10153, 9188]. Gillespie [7754]. Gilman [1648]. Gimignano [2551]. Gin [2648]. Giraffe [3182]. Given [6218]. Glacial [2735]. Glacier [9880, 1934]. Glaciers [8965]. Gladiator [11585]. Gladiatorial [5957]. Glasgow [2594, 9724]. Glass [3249, 3574, 2804, 2681, 2538, 4338, 3864, 4367, 4959, 5309, 5291, 5905, 5694, 6663, 7340, 6381, 8215, 9068, 9445, 11080, 11234, 1161, 220, 566, 11609, 1228, 3817]. Glass-Framing [8215]. Glass-Front [6663]. Glass-Furnace [3249]. Glass-Lamp [3574]. Glass-Sponges [11609]. Glassware [3940]. Glehn [6241]. Gleniffer [6031]. Glidden [10713, 11841, 8482, 9599]. Glider [196]. Gliding [6834, 9265, 10712, 10069, 11606, 8109]. Glimpse [3094, 10612]. Globe [2006, 6486, 7308, 6848, 8904, 4579, 1056]. Globes [198]. Globular [6201, 8745]. Glow [6776]. Glucinium [4236]. Glue [11615]. Gluten [10898]. Gnu [6720]. Go [4482, 1101]. Go-Carts [1101]. Goat [7493]. Gobelin [7585, 11303]. Goblets [317]. God [6020, 3193]. Gods [2953]. Goes [6092]. Going [10379]. Gokteik [3643]. Gold [1552, 1610, 1885, 4704, 6119, 7206, 9112, 9789, 4559, 1941, 3251, 9549, 1653, 7938]. Golf [4879]. Good [1744, 9594, 9181, 10706, 10413, 11857, 1186, 284, 665, 5699, 4490, 957]. Goods [1764, 2471, 6787, 10380, 11108, 10246]. Goodspeed [5305]. Goose [5577]. Gordon [1466, 5576, 6053, 7023]. Gorgas [11213]. Gorge [8591]. Gorges [5507]. Gorter [4963]. Gotha [4267]. Gourds [687]. Governing [9857]. Government [11183, 1360, 1868, 2825, 3475, 3784, 4453, 4121, 5620, 5901, 4992, 6832, 7238, 6884, 8051, 8520, 9365, 8762, 8595, 9239, 10770, 9719, 10211, 9698, 10039, 10085, 10640, 10916, 11626, 11639, 11185, 11365, 11923, 1255, 4079, 2658].

Government-Built [8051]. Governmental [114]. Governor [2209, 10470, 10560]. Grab [3202]. Grade [10704, 11859, 699, 6122]. Grades [3661, 4556, 7978, 9291]. Gradient [2353]. Graduation [6994]. Græco [5908]. Graham [5164]. Grain [3663, 4855, 4996, 5458, 4947, 7219, 6243, 11760, 440, 1077, 708, 1003, 599, 1468, 9214]. Grain-Washing [11760]. Gramme [3085]. Gramophone [7582, 2708]. Granaries [5724]. Grand [10626, 11847, 2255, 5412, 8028, 7769, 9506, 9566, 10474, 10495, 10276, 11724, 11739, 11177, 6578]. Grandeur [1211]. Grandfather [7933]. Grangesberg [5929]. Granite [2782, 4053, 11589]. Grant [5520]. Grants [6684]. Grape [3457]. Grapevine [4319]. Graphic [5910, 7675]. Graphical [5869]. Graphite [11905, 9897]. Graphite-Resistance [11905]. Graphotype [5854]. Grass [10046, 4581, 4063]. Grasses [1647]. Grasshopper [5458]. Grate [3447, 5797, 7893]. Grate-Bar [5797]. Gratifying [11127]. Grave [6868]. Gravestone [2243]. Gravity [7920, 9096, 1306]. Gray [3014]. Grazing [2684]. Grease [3180]. Grease-Pump [3180]. Great [1484, 2190, 1600, 1541, 3195, 3266, 3593, 4341, 4501, 4823, 3903, 4678, 5616, 5750, 5637, 5594, 7180, 7338, 7252, 7630, 7479, 8138, 7621, 8699, 9795, 9903, 10937, 11033, 11252, 11515, 912, 504, 6718, 7983, 8763, 8773, 5388, 4312, 6622, 11067, 4781, 11543, 2232, 1980, 1887, 2147, 2858, 3530, 2665, 4194, 3833, 4411, 5178, 5997, 5128, 5577, 5284, 5558, 6732, 6736, 6241, 7231, 8031, 8019, 7554, 10537, 10622, 11170, 1056, 8337, 3843, 4681, 4831, 3697, 8530, 3954, 243, 7162, 7240, 11398]. Greater [10296, 9180]. Greatest [1987, 2327, 2881, 4869, 3907, 9162, 10065, 10740, 6363, 8124]. Greatly [6688]. Green [1818]. Grenade [280]. Grenades [7985]. Grenadiers [7985]. Greth [5811]. Grinder [3097, 6034, 11549]. Grinders [319]. Grinding [1580, 4002, 7017, 8286, 11019]. Grindstone [292]. Grip [2078, 6167, 8149, 11375]. Grisson [4128]. Gromoboi [4786]. Grooming [6165]. Grooving [2053]. Grooving-Saw [2053]. Gros [7602]. Gross [3766]. Gross-Lichterfelde [3766]. Ground [1408, 10385, 1094, 8900, 9396, 2352]. Group [2966, 5066, 4852, 8857]. Grouping [9390, 11223]. Groups [11042, 6250]. Grout [7799, 9945, 4284]. Grow [11925, 5000]. Growing [1911, 3196, 7850, 9262, 11323]. Grown [791, 4348, 687]. Growth [2817, 3437, 2861, 3330, 3118, 4240, 3910, 4977, 4920, 5561, 5825, 6527, 6397, 6332, 7693, 8314, 9202, 9665, 9190, 9415, 9874, 10477, 10751, 1075, 888, 474, 1350]. Growths [3597, 5701]. Groynes [9066]. Grubb [7678]. Gruhn [8208]. Guadalajara [917]. Guadeloupe [4964]. Guard [1773, 1556, 1418, 4813, 5283, 6058, 7061, 8312, 9571, 10179, 11459, 2417]. Guardian [2422]. Guarding [3976]. Guards [496]. Guarini [5599, 3722]. Guatemala [4670, 5676, 6078]. Guayule [939]. Guichen [6052]. Guide [11613, 399]. Guides [10836]. Guiding [11707]. Guilleminot [5261]. Gulf [11237, 1278, 1073]. Gum [7773]. Gummed [6543]. Gumming [3010]. Gun [1465, 1446, 2168, 2373, 1707, 2036, 2312, 3349, 2955, 3485, 3160, 2777, 2794, 3150, 2848, 3082, 3928, 4660, 4346, 4904, 4707, 5470, 5101, 5157, 5950, 5804, 5747, 5457, 6396, 6728, 6267, 7670, 7355, 7678, 8250, 8502, 9241, 9416, 9015, 8858, 9225, 9160, 10574, 10121, 9814, 10716, 10192, 11482, 11778, 78, 822, 612, 115, 9289, 6785, 2807, 872, 10340, 4783]. Gun-Cotton [4346]. Gun-Mounts [3485]. Gun-Sight [7678]. Gunboat [6827]. Gunboats [11944]. Gunfire [11364]. Gunmaker [9182]. Gunnery [4846, 5021, 10665, 10790]. Gunpowders [9744]. Guns [2257, 1605, 1679, 2199, 2860, 3484, 3270, 3892, 6163, 5021, 5453, 6928, 6485, 6784, 6462, 7518, 8048, 8617, 10594, 10296, 11055, 981, 1160, 1765, 6477, 5947, 10112]. Gurgless [4419]. Gurnards [349]. Gusher [4883]. Gutter [438]. Gymnasium [9080]. Gyrating [8406]. Gyratory [11604]. Gyroscope [9009, 10023, 10556, 9802, 8386]. Gyroscopes [2687]. Gyroscopic [10235, 9976, 1018]. Gyrostat [10067].

H [1486, 4692, 5805, 6498, 8275, 8737, 8124]. H.R.R [8164]. Habitations [336]. Habits [11059]. Hack [5907]. Hagenbeck [9355]. Hail [8494, 10949]. Hailstones [5390]. Half [1351, 1852, 1963, 10748, 985]. Half-Tone [1351, 1963]. Halford [2353]. Halifax [4824]. Hall [1908, 4704, 4053, 5461, 7401, 6809, 7853, 4313]. Hall-Marking [4704]. Halley [828, 742, 1252]. Halls [436]. Halske [6604]. Halyards [7561]. Hamburg [2423, 3042]. Hammer [1799, 2911, 6026, 7891]. Hammers [10697, 120]. Hammock [569]. Hand [2024, 1333, 3273, 6604, 9493, 8838, 9511, 9689, 10364, 10245, 8499, 10657, 1228]. Hand-Cars [3273]. Hand-Drill [6604]. Hand-Labor [2024]. Hand-Operated [8838, 9689]. Hand-Pressure [8499]. Hand-Propelled [9511]. Handle [2262, 6589, 10982, 11157, 323]. Handle-Bar [2262]. Handless [7280]. Handling [3129, 4086, 5807, 10049, 11824, 1077, 9981, 2263, 3723, 4006, 2310, 5364, 5573]. Handy [6696, 11544, 11610, 11677, 11686, 64, 94, 197, 223, 290, 318, 394, 419, 467, 516, 561, 638, 683, 731, 1099, 1150, 1199, 1227, 1258, 296, 149, 969, 10139, 150, 484, 201, 662, 872, 688, 805, 1022, 865]. Handyman [11708]. Hanger [1603, 1488, 4035, 8175, 7611, 10092, 10976, 38, 1180, 685, 1233]. Hanging [8437]. Hansom [6666]. Hanson [3871]. Harbor [3248, 3225, 5877, 5047, 7608, 10174, 10588, 11736, 10880, 912, 7773, 2784, 9425, 11856, 7359, 2601]. Harbor-Defense [10588]. Hardening [203]. Hardness [11941]. Hardships [11373]. Hardware [3957]. Hare [1918]. Harlem [3161, 5792, 6780, 1990]. Harmless [6609]. Harmonic [2472]. Harmsworth [11906, 6100, 11179]. Harness [1418, 2382, 7407]. Harnessing [7544]. Harrow [4143, 6272]. Harrow-Tooth [4143]. Harrows [2154, 2032]. Harry [5408]. Hartford [1801, 8742, 3823]. Harvard [5520, 7734]. Harvest [9838, 9268]. Harvester [4322, 4897, 9016]. Harvesting [1075]. Haschke [7190]. Haskin [2464]. Haswell [6498, 10006]. Hatching [11831]. Hathamite [4554]. Hats [6504]. Hatteras [6537, 1368]. Haulage [3528, 5785, 5385]. Havana [3996]. Haven [4158, 8742, 10714, 10132, 10684, 9918]. Hawaii [2461, 2641]. Hawaiian [2778]. Hawk [11167]. Hay [3507, 5641, 7917, 7829]. Haynes [10006, 9933]. Haystacker [2691]. Head [2053, 4064, 4626, 7187, 9233, 9172, 120, 439, 781, 947, 7183, 8150, 1701, 4641, 3360, 2352]. Head-On [4626, 7187]. Headers [7219]. Heading [7008]. Headlight [4370, 4260]. Heads [5279, 8175, 10732, 11697, 485]. Health [5714, 9876, 10791]. Healthy [3807, 10957]. Heap [4405, 7439]. Hear [5332]. Heard [5998]. Hearing [1642, 11396, 10915]. Hearse [9898]. Heart [5750]. Heat [3664, 3336, 8432, 7725, 8501, 9092, 11076, 723, 750, 3788, 8076, 200, 812, 3509]. Heated [2492]. Heater [3223, 3686, 3569, 4857, 3982, 5742, 5264, 8361, 10566, 10364, 1303]. Heating [1342, 3633, 2623, 3726, 4511, 7229, 8743, 9667, 10685, 10886, 8356, 9602, 802]. Heaton [7941, 7902]. Heavens [2842, 2921, 8307, 9145, 3555, 7107, 1398, 2393, 2247, 1612, 1734, 1520, 2135, 2015, 2500, 2626, 3438, 2996, 3514, 2718, 3099, 3320, 3212, 3671, 3773, 4791, 4628, 4178, 4023, 4111, 3706, 3896, 4515, 4408, 4326, 5016, 6063, 5925, 5446, 4909, 5272, 5381, 5148, 5531, 5800, 5684, 5585, 6279, 7242, 6674, 6170, 6488, 6777, 6189, 6384, 6975, 6881, 6789, 7431, 8415, 7663, 7945, 7529, 7922, 8178, 8093, 7968, 8626, 9554, 9434, 8995, 8860, 8950, 8710, 9104, 9296, 9219, 9657, 9759, 10676, 10548, 10162, 9678, 10011, 10096, 9868, 10248, 10428, 10366, 10793, 10889, 11849, 11692, 11267, 10801, 11120, 11203, 10821]. Heavens [11003, 11575, 11484, 11402, 130, 1182, 1033, 520, 1, 352, 644, 228, 622, 875, 763, 1284, 7776, 679]. Heavier [10299, 10321]. Heavier-Than-Air [10299]. Heaviest [4754]. Heavy [3661, 3846, 4053, 5773, 6163, 5453, 8676, 9894, 10825, 11934, 11389, 6852, 9684]. Hebrew [8825]. Hecla [3508]. Heel [5053]. Height [3132, 11845, 11676, 854, 1094, 6159]. Held [3377]. Helicopter [10172, 10972, 11638]. Helicoptère [9050]. Helicopters [11522]. Heliotropism [8083]. Helium [526]. Helmet [1057]. Helmets [10176]. Hemisphere [4643]. Hemp [4638]. Hendrick [8744]. Henri [7114]. Henry [2705, 9524, 6460, 10947, 11876, 968, 10652]. Her [1393, 4231, 3814, 5510, 1767]. Herbert [5973]. Here [7337, 5203]. Heritage [7514, 7538, 7558]. Heroic [10960]. Herreshoff [1700, 2634]. Heskett [7903]. Hewitt [2716, 5455, 5538, 5480, 7855, 10457, 7541]. Hichborn [3107]. Hiding [6226]. Hieroglyphs [679]. High [1841, 2484, 2435, 2093, 1750, 3486, 2701, 3398, 2860, 2840, 3081, 3442, 3350, 3143, 3435, 2750, 2557, 3745, 3974, 4289, 5303, 5177, 5323, 5836, 5783, 6221, 6378, 6288, 7164, 6195, 6290, 6396, 6496, 7645, 8081, 8250, 7629, 8046, 9110, 8800, 8543, 9755, 10243, 10656, 9775, 11015, 10897, 11946, 11636, 1041, 582, 894, 913, 10780, 9010, 8388, 6122, 5261, 11325, 9703, 9114, 2075, 4221, 7661, 6054, 11655, 7707, 480, 11373, 5235, 6237]. High-Class [2840]. High-Explosive [3486, 3442, 3435, 6288]. High-Frequency [10780, 5261]. High-Jumping [3745]. High-Level [9110]. High-Lift [2484]. High-Potential [9010]. High-Power [2860, 11325]. High-Powered [7164, 8800]. High-Pressure [2557, 5836, 9755, 10897, 11636, 582, 913, 11655, 6237]. High-Speed [2093, 1750, 2701, 3081, 3974, 5303, 5323, 5783, 6221, 6378, 6195, 6290, 8081, 7629, 10243, 9775, 894, 8388, 9703, 9114, 4221, 7707]. High-Tension [11946, 5261, 5235]. High-Velocity [6396]. Higher [3653, 4773, 4579].

Highest [1981, 5358, 860]. Highspeed [3937]. Highway [105]. Highways [3209]. Hildenbrand [8185]. Hill [8435, 10963, 4211, 10148, 11128]. Hill-Climbing [8435]. Hiller [1913]. Hilling [3769]. Hilling-Up [3769]. Himalaya [4943]. Himalaya-Climbing [4943]. Himalayas [5880, 11373]. Hindu [11499]. Hinge [3068, 9404, 16]. Hints [4830]. Hippodrome [9841, 7424]. Hippopotami [8260]. Historic [7020, 6856, 1709]. Historical [4136, 5945, 7653, 8989]. Histories [515]. History [2918, 3634, 2547, 3812, 6573, 7677, 7783, 8458, 9169, 11452, 748, 9962, 8597, 4686, 6894, 7937, 7853, 3988]. Hoarfrost [7887]. Hoboken [2198]. Hoe [11706]. Hofman [2736]. Hoist [10434, 5546, 10139]. Hoisting [1703, 2961, 10469, 678, 5886]. Hoists [6500, 3573]. Holder [2238, 3613, 3178, 2558, 3139, 6059, 5291, 7093, 6696, 8905, 10140, 9778, 11308, 11949, 780, 537, 319, 11689]. Holders [6046]. Holding [10203]. Holds [11896]. Holes [7408, 11613, 890, 256, 11786, 6355]. Holland [2023, 3244, 4662]. Hollow [5797, 9473, 11371, 8456]. Holsman [9928, 10461]. Home [11546, 2212, 3864, 5091, 5786, 5065, 6380, 10243, 11714, 495, 488, 791, 10027, 11862, 7188, 639, 71, 1027, 96, 10758, 11618, 293, 7448, 11833, 9229, 7655, 2221, 4239, 1023, 11717, 10776, 10402, 684, 10717, 1154, 9859, 687, 68, 663, 3156]. Home-Grown [791, 687]. Home-Made [11546, 2212, 3864, 11714, 10027, 11862, 639, 71, 1027, 96, 10758, 11618, 293, 7448, 11833, 9229, 2221, 4239, 11717, 10776, 10402, 684, 10717, 1154, 68, 663, 3156]. Homemade [1583]. Homes [4605, 7952]. Homing [9979]. Hon. [4692]. Honduras [5109]. Hone [2418]. Honey [10437]. Honeycomb [9963]. Honolulu [1364, 5459, 11256]. Honor [714]. Honorable [1143, 1168]. Hook [2382, 3384, 3442, 3435, 2816, 4707, 5470, 6760, 11644, 2825, 9502]. Hooks [11642]. Hooliganism [5593]. Hoop [5342]. Hope [6867, 6318, 4490]. Hopkins [4894, 7836]. Horizon [4945, 6080, 10628]. Horizontal [7771, 11613]. Horn [4422, 4240, 10980, 4552]. Horned [3182]. Horns [8212]. Horror [7620]. Horrors [5278]. Horse [1791, 1585, 1418, 3745, 4636, 3715, 4989, 6935, 6185, 6874, 6244, 8430, 8231, 8685, 8799, 9123, 9925, 9945, 9926, 9949, 11179, 5603, 5336, 9248, 5083, 5648, 11220, 479, 7024]. Horse-Power [6244, 9123]. Horseflesh [1338]. Horsepower [1419, 3032, 6607, 6642, 5693]. Horses [2151, 4587, 4799, 4396, 6165, 11562, 7446, 7379]. Horseshoer [4621]. Horseshoes [7499, 8301, 10815]. Horton [3741]. Hose [4058, 7232, 7218, 7456, 9192, 11201, 11093, 664]. Hospital [2714, 10858, 1388]. Hospitals [8814]. Hot [809, 3726, 8920, 10886, 807, 9187]. Hot-Air [8920]. Hot-Water [809, 3726, 10886, 807, 9187]. Hotel [3735, 11326, 6310]. Hotels [3785]. Hour [1769, 4231, 5732, 7307, 6998, 10771, 9113]. Hours [9584]. House [2306, 2990, 2957, 3203, 2936, 6160, 5814, 5960, 6629, 6215, 1296, 232, 222, 559, 11088, 1469, 5770, 1548, 4048, 6941, 6690, 7826, 9965, 6786]. House-Fly [559]. House-Lamp [2306]. House-to-House [222]. Houseboat [4298, 4299]. Household [6247, 10157]. Housekeeping [7977]. Houses [3307, 4896, 10522, 4581, 1469]. Howell [2496]. Hower [10713]. Hub [4714, 9954]. Hudson [4253, 6915, 8744, 10947, 3081, 6142, 6358, 6269, 6679, 7261, 9459, 9665, 9204, 9700, 10339, 10375, 10848, 10960, 11293, 10825, 11232, 11351, 259, 755, 820, 748, 1211, 213, 747, 155, 1167, 767]. Huge [4527, 7913, 9313, 9239, 10247, 11435, 660, 1077, 10614, 945]. Hughes [1822]. Huichol [4645, 11239]. Hull [5544, 287, 512]. Human [6756, 9076, 8934, 8894, 860, 1098, 11355, 10246]. Humane [11635]. Humanitarian [10435]. Humidity [3633, 4412, 8937]. Humors [5430]. Hundred [4190, 5909, 5732, 5301, 6998, 7836, 8128, 8940, 8652, 10061, 768, 762, 9567]. Hundred-Ton [5909]. Hundredth [4484]. Hungarian [1984]. Hungary [2473]. Hunt [2665]. Hunter [4973]. Hunting [9816]. Hur [2494]. Hurricane [9428, 2120]. Husker [9016]. Husking [1017]. Hyacinth [5563]. Hybridization [8545]. Hybrids [5815]. Hydrant [4664]. Hydraulic [1683, 1515, 2824, 2644, 2859, 4341, 4346, 5616, 7079, 6403, 6912, 6202, 9629, 9729, 11753, 11241, 148, 8497, 5721, 518, 5886, 8052]. Hydraulic-Electric [5616]. Hydride [4548]. Hydro [61]. Hydro-Electric [61]. Hydrobion [6177]. Hydrocarbon [1482, 3137, 3770]. Hydrogen [8519, 8851, 11041, 11254, 9051, 8477, 10995, 8043]. Hydrographic [3206, 7543]. Hydrometer [2177, 4242]. Hydroplane [10439, 10107, 10238, 10426, 10069, 81, 9114]. Hydroplanes [10712, 416]. Hydroscope [6712]. Hydrostatic [1028, 2076]. Hygiene [10156, 9648, 10003]. Hygienic [4027]. Hygroscope [2963]. Hyperbolas [10651]. Hypnosis [3511]. Hypnotic [10398]. Hypo [2638]. Hyposcope [5674]. Hypothesis [2792].

Ibis [1827]. Ibuki [11002]. Icarus [777]. Ice [2827, 2801, 4020, 4057, 4965, 6070, 7343, 6395, 7380, 7723, 7940, 7607, 9795, 9880, 11331, 10855, 11540, 317, 1053, 11371, 9816, 5477, 8455, 257, 9099, 243, 116]. Ice-Boat [4965]. Ice-Box [4057]. Ice-Breaker [6395]. Ice-Conveying [2801]. Ice-Cutting [7343]. Ice-Making [11540, 8455]. Iceberg [5190]. Iceless [7709, 1151]. Ichthyology [1875]. Idaho [3240, 7169, 10595, 1961]. Ideal [1904, 6034, 7773]. Identification [10674]. Identifying [8088]. Idle [11510]. If [10360]. Igniter [2757, 10927]. Ignition [3760, 7806, 8807, 9545, 9923, 10171, 10480, 1260, 480]. Iguana [7328]. II [3667, 10832, 2711, 2758, 3064, 3428, 3654, 4878, 3981, 4530, 5168, 4993, 5884, 6295, 6812, 6559, 8146, 7862, 8853, 9030, 11915, 415, 535, 458, 320, 7916, 9354, 10221, 4900, 7538, 5526, 7116, 9267, 6877, 2246, 8707, 6293, 6916, 9166]. III [1607, 1344, 4287, 5071, 5380, 5137, 8616, 11886, 11428, 673, 776, 640, 7558, 5547, 2291]. Ill [10343]. Ill-Fated [10343]. Illinois [2614, 6297]. Illuminant [3089, 5642]. Illuminated [2639, 4011, 5703]. Illuminating [2058, 10421]. Illumination [6178, 10880, 6459]. Illuminations [7259]. Illusion [8060, 8387, 169, 8919, 10009, 10632, 9772, 8141]. Illusions [11418]. Illusory [6495]. Illustrate [659]. Illustrating [1984, 1028, 11803]. Image [4022].

Immigration [4932]. Immortality [11861]. Impending [8614]. Imperfect [4022]. Imperial [10463, 6682]. Impinging [8767]. Implement [9193, 8575]. Implements [10447, 10814]. Importance [3006, 6822]. Important [1719, 2409, 3328, 3442, 2545, 2593, 4433, 4747, 5969, 7397, 8042, 8583, 10351, 11053, 90]. Importer [7792]. Imports [2325, 9390]. Impossible [1622]. Impression [5478]. Impressions [415, 636]. Impressive [11368]. Improve [3737]. Improved [2284, 1896, 1775, 2126, 1823, 1535, 1332, 2380, 2420, 1960, 1708, 1748, 1916, 2239, 3572, 3178, 3387, 3224, 2550, 3206, 3507, 2712, 2680, 2646, 3366, 4535, 4604, 4855, 4395, 4174, 4874, 4470, 4264, 4489, 3841, 4323, 4614, 6163, 5890, 6009, 5987, 5345, 5436, 5888, 5679, 6051, 5308, 6073, 5441, 6166, 6584, 6640, 6321, 7061, 6186, 7143, 7217, 6243, 6979, 6518, 6513, 6538, 6589, 7026, 6813, 6275, 7989, 7892, 8285, 7407, 7525, 7966, 7683, 7684, 8325, 7965, 7636, 7837, 7611, 7917, 7868, 8438, 7526, 7865, 7635, 8241, 8507, 8837]. Improved [9171, 9056, 9014, 9233, 8990, 8687, 8706, 8991, 9493, 8838, 9545, 9405, 8946, 9675, 9379, 9153, 9081, 8945, 9336, 9451, 9377, 9431, 9152, 9632, 8808, 9551, 9652, 9232, 8598, 8554, 9986, 9783, 10567, 10508, 10721, 9709, 10092, 9777, 10547, 10365, 10046, 10783, 10718, 10386, 10161, 9910, 10323, 9864, 9989, 10138, 9940, 10136, 9844, 9948, 10384, 9952, 9904, 10719, 11573, 11263, 11200, 11333, 11357, 10930, 11867, 10819, 11265, 10836, 11466, 11448, 10883, 11569, 11763, 11157, 10929, 11094, 11570, 11664, 10837, 781, 86, 38, 16, 947, 124, 12, 176]. Improved [1002, 34, 413, 9228, 10715, 8208, 8388, 4391, 6443]. Improvement [1675, 2191, 1415, 1429, 3225, 3121, 3469, 2669, 3886, 4244, 7180, 7567, 9742, 10697, 10781, 10665, 11187, 11428, 11915, 1290, 8454, 6123, 10992, 555]. Improvements [2411, 2774, 3677, 4835, 4503, 5412, 5879, 5056, 6639, 8028, 8803, 9012, 9006, 9125, 9272, 10103, 10455, 9959, 9963, 10478, 11736, 11378, 11276, 913, 11440, 6076, 432, 1658]. Improvised [10720, 1230]. Impulses [9140]. in-Turning [5039]. Inauguration [6971, 10714, 10111, 11232]. Inca [6445, 6608]. Incandescent [2284, 1438, 4509, 6890, 7317, 7868, 11139, 11214, 11809, 11588]. Incandescent-Lamp [2284]. Incas [4785]. Incased [948]. Inch [2105, 2036, 2312, 3349, 2955, 3160, 3892, 4707, 5470, 5157, 5457, 6396, 6267, 8617, 8858, 11144, 981, 2807, 1990, 10340]. Incident [7681]. Incised [6381]. Inclinator [7060]. Incline [3684]. Inclined [4600, 28]. Inclosed [2693]. Income [114]. Incomes [3757]. Inconclusive [1136]. Inconstancy [1888]. Increase [2392, 3394, 5204, 11252]. Increased [2692, 8934]. Increases [11186]. Increasing [3209, 2566, 6475, 7055, 7747, 9530, 8785]. Incubator [7738]. Incubators [3591, 10473]. Independence [2620]. Independent [9240]. Index [1314, 1329, 1902, 1926, 1949, 1972, 1997, 2019, 2042, 2068, 2091, 2114, 2141, 2165, 2188, 2207, 2228, 2253, 2273, 2296, 2320, 2346, 2364, 2387, 2406, 2432, 2453, 2481, 1357, 2506, 1379, 1404, 1427, 1452, 1476, 1501, 1525, 1546, 1570, 1592, 1618, 1640, 1666, 1691, 1716, 1740, 1759, 1785, 1812, 1838, 1859, 1881, 1741, 2528, 2542, 3127, 3148, 3169, 3191, 3216, 3238, 3258, 3281, 3300, 3324, 3344, 3372, 3392, 3414, 3433, 3451, 3472, 3497, 3516, 3538, 3559, 3583, 3606, 3629, 3651, 3674, 2563, 2588, 2608, 2630]. Index [2653, 2676, 2698, 2721, 2746, 2770, 2790, 2813, 2835, 2856, 2879, 2904, 2927, 2950, 2976, 3001, 3029, 3052, 3076, 3104, 2929, 3517, 3693, 3711, 4306, 4330, 4357, 4380, 4424, 4428, 4448, 4477, 4497, 4520, 4544, 4566, 4593, 4611, 4631, 4655, 4676, 4699, 4724, 4744, 4769, 4794, 4819, 4839, 4861, 4887, 3730, 3750, 3777, 3803, 3829, 3853, 3875, 3898, 3923, 3948, 3969, 3992, 4014, 4041, 4070, 4094, 4114, 4134, 4156, 4180, 4208, 4227, 4249, 4270, 4115, 4725, 5375, 4911, 4926, 5557, 5590, 5615, 5635, 5661, 5688, 5711]. Index [5730, 5757, 5780, 5801, 5821, 5845, 5875, 5897, 5918, 5943, 5966, 5994, 6018, 6040, 6066, 6087, 6112, 6133, 6153, 4954, 4986, 5009, 5029, 5062, 5097, 5125, 5154, 5174, 5201, 5221, 5270, 5275, 5299, 5320, 5352, 5373, 5402, 5422, 5449, 5468, 5492, 5515, 5535, 6174, 6193, 6772, 6795, 6819, 6841, 6865, 6883, 6903, 6925, 6953, 6969, 6991, 7014, 7033, 7053, 7076, 7099, 7126, 7150, 7178, 7202, 7226, 7247, 7269, 7290, 7314, 7334, 6213, 6233, 6255, 6285, 6304, 6329, 6350, 6371, 6388, 6411, 6431, 6453, 6473, 6493, 6525]. Index [6548, 6566, 6597, 6615, 6633, 6677, 6703, 6726, 6751, 7745, 7351, 7369, 7973, 7996, 8014, 8037, 8068, 8097, 8116, 8135, 8157, 8181, 8200, 8223, 8246, 8268, 8291, 8310, 8331, 8347, 8370, 8395, 8420, 8444, 8466, 8487, 8515, 7391, 7414, 7436, 7465, 7486, 7509, 7532, 7552, 7574, 7594, 7619, 7643, 7667, 7690, 7715, 7744, 7761, 7780, 7820, 7846, 7873, 7897, 7927, 7950, 8955, 8538, 8561, 9134, 9157, 9178, 9200, 9222, 9237, 9258, 9279, 9301, 9319, 9342, 9386, 9411, 9438, 9457, 9477, 9499, 9519, 9537, 9558, 9575, 9592]. Index [9613, 9637, 9662, 8580, 8607, 8630, 8649, 8671, 8695, 8714, 8731, 8757, 8779, 8820, 8845, 8865, 8885, 8910, 8930, 8954, 8975, 9000, 9022, 9040, 9062, 9088, 9108, 9681, 9694, 10272, 10290, 10310, 10328, 10348, 10371, 10390, 10411, 10431, 10448, 10488, 10513, 10533, 10551, 10570, 10601, 10619, 10636, 10662, 10679, 10701, 10727, 10745, 10765, 10787, 9717, 9740, 9763, 9787, 9810, 9831, 9851, 9872, 9891, 9916, 9972, 9998, 10016, 10033, 10054, 10075, 10100, 10119, 10145, 10167, 10186, 10209, 10231, 10252, 10101, 10637, 11794, 11246, 10805, 10823, 11383, 11406]. Index [11423, 11445, 11471, 11489, 11509, 11532, 11554, 11577, 11595, 11623, 11649, 11673, 11696, 11720, 11745, 11768, 11793, 11813, 11837, 11854, 11873, 11893, 11911, 11930, 10842, 10865, 10893, 10912, 10933, 10955, 10984, 11008, 11027, 11051, 11072, 11100, 11125, 11142, 11162, 11181, 11208, 11229, 11248, 11270, 11291, 11311, 11339, 11361, 5, 20, 669, 692, 713, 740, 766, 788, 819, 842, 879, 910, 931, 955, 978, 1010, 1038, 1063, 1084, 1111, 1157, 1185, 1209, 1241, 1266, 1288, 44, 77, 103, 134, 154, 183, 210, 230, 249, 274, 302, 332]. Index [355, 378, 405, 451, 471, 501, 523, 546, 578, 602, 626, 649, 424, 1112]. Indexing [4077]. India [2869, 2169, 2037, 2818, 5734, 11705, 11416, 8702]. Indian [2352, 2617, 4011, 7036, 5724, 5440, 2401, 4852]. Indians [3640, 4645, 11239, 4581, 2474]. Indican [2413]. Indicating [5088, 11063, 7322, 7958, 3847]. Indication [11888]. Indications [6551]. Indicator [1986, 3638, 3662, 4805, 3825, 7002, 7382, 8687, 891, 677, 991, 3461].

Indicators [10019]. Indies [4005, 9399]. Indigo [2123]. Individual [9938, 9932]. Indoor [10969, 1058]. Induced [5305, 7317, 8083, 6438]. Induction [6157, 6961, 7259, 141, 2535]. Industrial [3676, 2982, 6044, 7035, 6570, 9245, 9543, 10156, 3359, 8935, 554, 10003, 7059]. Industries [4386, 5280, 8980, 10188, 10422, 11234]. Industry [1339, 1455, 1483, 2861, 2683, 2681, 2576, 3765, 4339, 4777, 4573, 3972, 4748, 5250, 5948, 5518, 8319, 9414, 9473, 10262, 9706, 11804, 11521, 1075, 657, 474, 220, 513, 535, 7118, 9464, 2846, 2802, 4638, 7329, 3912, 2624, 3286, 939, 5258, 8005, 7085, 4127, 11523, 6846, 682]. Inefficient [8119]. Inexpensive [1028, 2340]. Infantry [1651]. Infection [1978]. Infested [7660]. Inflammation [8719]. Inflatable [4472]. Inflating [6888, 10907, 6440]. Inflexible [857]. Inflexibles [844]. Influence [4556, 7275, 6601, 8934, 10038, 11946, 10398, 8125, 10657, 10241]. Information [5621]. Infringement [11898, 1291]. Ingalls [8772]. Ingenious [2398, 1580, 1464, 3339, 5722, 7233, 6688, 9960, 10633, 11943, 3065]. Ingeniously [2082]. Ingenuity [2547, 8495]. Ingomar [6441]. Ingot [219]. Ingots [460, 364]. Inhabitant [7308]. Inhabitants [11879]. Inhalatorium [9100]. Inhaling [6341]. Inherit [7228]. Injection [1792, 11475, 1179]. Injured [5544, 92]. Injuries [6671]. Injurious [769]. Injury [348]. Ink [2692, 7454]. Ink-Controlled [7454]. Inkwell [3038]. Inlaid [9801]. Inland [9878, 142, 1874, 9817]. Inlaying [1197]. Innovation [8042]. Inoculating [7400]. Inoculation [7396]. Inorganic [8233]. Inquest [4191]. Inquiry [2512, 10857]. Insect [6937, 2918, 8148, 10004, 6158, 11155, 6879, 7167, 2830, 3869, 10779]. Insects [3891, 7003, 6729, 10360, 769, 4103, 10562, 11262, 2624, 2133]. Inspection [4024, 1867, 3785, 4866, 4256, 11210, 11044]. Install [661]. Installations [5019]. Instance [3789]. Instantaneous [1397]. Instinct [3820]. Instincts [4103]. Institute [2329, 3154, 4312, 2805, 282, 6959, 1562, 6619, 10526]. Institution [9323, 9100, 11256, 615, 1844, 5520, 6684, 7771]. Instruction [7341]. Instructive [17]. Instrument [3070, 3143, 2643, 6292, 8062, 8749, 8991, 9953, 7067, 7167, 7322, 10651, 5835, 7584, 6121, 662, 10472, 6786]. Instruments [1604, 3206, 5332, 947, 11862, 3406, 7606]. Insulating [5552, 6787]. Insulator [3069, 8947, 10048]. Intact [11045]. Integral [194]. Intensification [10202]. Interest [7850]. Interesting [2511, 3367, 3502, 2663, 3385, 2966, 2536, 3532, 4223, 3727, 3771, 4175, 7340, 6653, 6319, 6464, 7113, 6391, 6223, 7608, 9008, 10444, 10502, 226, 971, 6361, 7259, 3868, 8144, 9753, 5391, 8055, 2133, 5647, 11808, 2830, 95, 9684, 10427]. Interior [2521, 7579, 9446, 9263, 7165]. Interlocked [8212]. Interlocking [4688]. Intermediate [5639]. Internal [5580, 6605, 9749, 1001, 991]. Internal-Combustion [9749, 1001, 991]. International [2982, 3504, 2620, 6117, 5597, 6963, 6781, 6571, 6871, 7658, 8383, 7854, 9308, 8697, 9264, 9270, 10495, 11914, 11257, 10941, 11895, 11723, 9326, 10737, 10297, 1846, 2279, 2394, 2463, 2027, 4388, 5105, 6812, 7576, 7623, 9860, 10850, 11906, 9746, 11301]. Internationally [1194]. Interoceanic [4309].

Interrupter [5538, 6961, 8431]. Introducing [3544]. Introduction [4933, 3983]. Introductory [4275]. Inundation [8690]. Invading [724]. Invalid [3377, 10783]. Invented [11280]. Inventing [3264, 8939]. Invention [2021, 1503, 2121, 5855, 4979, 5609, 5682, 5798, 6627, 6860, 6946, 7009, 7094, 7458, 7613, 8948, 9708, 9712, 10722, 11348, 11319, 10820, 11504, 11669, 1052, 905, 1004, 923, 10753, 2108, 3789]. Inventions [1314, 1329, 1902, 1926, 1949, 1972, 1997, 2019, 2042, 2068, 2091, 2114, 2141, 2165, 2188, 2207, 2228, 2253, 2273, 2296, 2320, 2346, 2364, 2387, 2406, 2432, 2453, 2481, 1357, 2506, 1379, 1404, 1427, 1452, 1476, 1501, 1525, 1546, 1570, 1592, 1618, 1640, 1666, 1691, 1716, 1740, 1759, 1785, 1812, 1838, 1859, 1881, 1334, 1310, 1325, 1923, 1945, 1969, 1993, 2016, 2038, 2064, 2087, 2110, 2137, 2161, 2184, 2204, 2224, 2249, 2269, 2292, 2316, 2342, 2360, 2383, 2403, 2428, 2449, 2477, 2502, 1353, 1375, 1400]. Inventions [1424, 1449, 1472, 1497, 1521, 1543, 1566, 1588, 1614, 1636, 1662, 1687, 1712, 1736, 1755, 1781, 1808, 1834, 1855, 1877, 1898, 2283, 2528, 2542, 3127, 3148, 3169, 3191, 3216, 3238, 3258, 3281, 3300, 3324, 3344, 3372, 3392, 3414, 3433, 3451, 3472, 3497, 3516, 3538, 3559, 3583, 3606, 3629, 3651, 3674, 2563, 2588, 2608, 2630, 2653, 2676, 2698, 2721, 2746, 2770, 2790, 2813, 2835, 2856, 2879, 2904, 2927, 2950, 2976, 3001, 3029, 3052, 3076, 3104, 3021, 3083, 2525, 2540, 3124, 3145, 3166, 3189, 3214, 3235, 3255]. Inventions [3278, 3297, 3322, 3343, 3369, 3391, 3413, 3432, 3450, 3471, 3496, 3515, 3535, 3556, 3580, 3603, 3626, 3648, 3672, 2560, 2585, 2605, 2628, 2651, 2673, 2695, 2719, 2743, 2766, 2787, 2811, 2833, 2853, 2877, 2901, 2924, 2946, 2972, 2997, 3025, 3049, 3073, 3101, 3532, 3693, 3711, 4306, 4330, 4357, 4380, 4424, 4428, 4448, 4477, 4497, 4520, 4544, 4566, 4593, 4611, 4631, 4655, 4676, 4699, 4724, 4744, 4769, 4794, 4819, 4839, 4861, 4887, 3730, 3750, 3777, 3803, 3829, 3853, 3875, 3898, 3923, 3948, 3969, 3992]. Inventions [4014, 4041, 4070, 4094, 4114, 4134, 4156, 4180, 4208, 4227, 4249, 4270, 4223, 3917, 4352, 4562, 4647, 4222, 3692, 3709, 4305, 4329, 4356, 4379, 4401, 4427, 4447, 4476, 4495, 4518, 4542, 4565, 4591, 4609, 4629, 4653, 4674, 4697, 4722, 4743, 4768, 4793, 4818, 4838, 4860, 4886, 3729, 3749, 3775, 3802, 3828, 3852, 3874, 3897, 3922, 3947, 3968, 3991, 4013, 4040, 4069, 4093, 4113, 4133, 4155, 4179, 4206, 4225, 4248, 4274, 3727, 3771, 4175, 5494, 4911, 4926, 5557, 5590, 5615, 5635, 5661, 5688, 5711, 5730]. Inventions [5757, 5780, 5801, 5821, 5845, 5875, 5897, 5918, 5943, 5966, 5994, 6018, 6040, 6066, 6087, 6112, 6133, 6153, 4954, 4986, 5009, 5029, 5062, 5097, 5125, 5154, 5174, 5201, 5221, 5270, 5275, 5299, 5320, 5352, 5373, 5402, 5422, 5449, 5468, 5492, 5515, 5535, 5194, 5485, 5583, 5939, 5988, 6061, 4910, 4923, 5555, 5587, 5612, 5633, 5658, 5685, 5708, 5727, 5754, 5777, 5818, 5841, 5873, 5894, 5915, 5963, 5991, 6016, 6037, 6064, 6084, 6109, 6130, 6151, 4952, 4984, 5006, 5026, 5061, 5095, 5122, 5151, 5171, 5197]. Inventions [5218, 5242, 5274, 5296, 5317, 5351, 5370, 5399, 5421, 5448, 5465, 5489, 5513, 5533, 6174, 6193, 6772, 6795, 6819, 6841, 6865, 6883, 6903, 6925, 6953, 6969, 6991, 7014, 7033, 7053, 7076, 7099, 7126, 7150, 7178, 7202, 7226, 7247, 7269, 7290, 7314, 7334, 6213, 6233, 6255, 6285, 6304, 6329, 6350, 6371, 6388, 6411, 6431, 6453, 6473, 6493, 6525, 6548, 6566, 6597, 6615, 6633, 6677, 6703, 6726, 6751, 6324, 6611, 6745, 6171, 6190, 6769, 6791, 6816, 6838, 6862, 6882, 6900, 6922, 6950, 6988, 7013, 7030, 7050, 7073]. Inventions [7097, 7123, 7147, 7175, 7199, 7224, 7244, 7267, 7287, 7311, 7332, 6211, 6229, 6252, 6282, 6301, 6327, 6346, 6369, 6385, 6409, 6428, 6450, 6470, 6490, 6522, 6546, 6564, 6594, 6613, 6634, 6675, 6700, 6723, 6749, 7351, 7369, 7973, 7996, 8014, 8037, 8068, 8097, 8116, 8135, 8157, 8181, 8200, 8223, 8246, 8268, 8291, 8310, 8331, 8347, 8370, 8395, 8420, 8444, 8466, 8487, 8515, 7391, 7414, 7436, 7465, 7486, 7509, 7532, 7552, 7574, 7594, 7619, 7643, 7667, 7690, 7715, 7744, 7761, 7780, 7820, 7846, 7873, 7897, 7927]. Inventions [7950, 7527, 7866, 8176, 8391, 7349, 7366, 7970, 7994, 8011, 8034, 8065, 8095, 8114, 8133, 8154, 8180, 8197, 8220, 8244, 8265, 8290, 8308, 8328, 8344, 8367, 8394, 8417, 8441, 8485, 8512, 7388, 7411, 7433, 7462, 7483, 7507, 7530, 7550, 7571, 7591, 7616, 7640, 7665, 7687, 7712, 7741, 7759, 7777, 7817, 7843, 7870, 7895, 7924, 7947, 7855, 8538, 8561, 9134, 9157, 9178, 9200, 9222, 9237, 9258, 9279, 9301, 9319, 9342, 9386, 9411, 9438, 9457, 9477, 9499, 9519, 9537, 9558, 9575, 9592, 9613, 9637, 9662]. Inventions [8580, 8607, 8630, 8649, 8671, 8695, 8714, 8731, 8757, 8779, 8820, 8845, 8865, 8885, 8910, 8930, 8954, 8975, 9000, 9022, 9040, 9062, 9088, 9108, 8625, 9017, 9380, 8535, 8557, 9133, 9156, 9176, 9197, 9221, 9236, 9256, 9276, 9298, 9316, 9340, 9360, 9383, 9408, 9435, 9455, 9475, 9496, 9516, 9534, 9556, 9574, 9591, 9610, 9635, 9659, 8577, 8603, 8628, 8647, 8669, 8694, 8711, 8728, 8754, 8776, 8818, 8842, 8862, 8882, 8908, 8927, 8952, 8972, 8997, 9019, 9038, 9059, 9085, 9106, 9311, 9681, 9694, 10272, 10290]. Inventions [10310, 10328, 10348, 10371, 10390, 10411, 10431, 10448, 10488, 10513, 10533, 10551, 10570, 10601, 10619, 10636, 10662, 10679, 10701, 10727, 10745, 10765, 10787, 9717, 9740, 9763, 9787, 9810, 9831, 9851, 9872, 9891, 9916, 9972, 9998, 10016, 10033, 10054, 10075, 10100, 10119, 10145, 10167, 10186, 10209, 10231, 10252, 9784, 9679, 9692, 10270, 10287, 10307, 10327, 10345, 10368, 10388, 10408, 10430, 10485, 10510, 10532, 10550, 10568, 10599, 10616, 10634, 10661, 10678, 10698, 10724, 10743, 10763, 10785, 9714, 9737, 9761, 9785, 9807, 9828, 9848, 9870, 9888, 9913]. Inventions [9969, 9995, 10013, 10030, 10051, 10072, 10097, 10116, 10142, 10164, 10183, 10207, 10229, 10250, 11246, 10805, 10823, 11383, 11406, 11423, 11445, 11471, 11489, 11509, 11532, 11554, 11577, 11595, 11623, 11649, 11673, 11696, 11720, 11745, 11768, 11793, 11813, 11837, 11854, 11873, 11893, 11911, 11930, 10842, 10865, 10893, 10912, 10933, 10955, 10984, 11008, 11027, 11051, 11072, 11100, 11125, 11142, 11162, 11181, 11208, 11229, 11248, 11270, 11291, 11311, 11339, 11361, 10888, 11022, 10802, 10839, 11420, 11442, 11469, 11486, 11506, 11529, 11551, 11576, 11592, 11620, 11646, 11670, 11693, 11742]. Inventions [11764, 11790, 11812, 11836, 11851, 11870, 11890, 11908, 11927, 10862, 10890, 10911, 10932, 10952, 11005, 11024, 11048, 11069, 11097, 11122, 11141, 11161, 11180, 11205, 11226, 11245, 11269, 11288, 11310, 11336, 11360, 11380, 11404, 11943, 5, 20, 669, 692, 713, 740, 766, 788, 819, 842, 879, 910, 931, 955, 978, 1010, 1038, 1063, 1084, 1111, 1157, 1185, 1209, 1241, 1266, 1288, 44, 77, 103, 134, 154, 183, 210, 230, 249, 274, 302, 332, 355, 378, 405, 451, 471, 501, 523, 546, 578, 602, 626, 649, 128]. Inventions [268, 445, 2, 18, 666, 693, 711, 738, 764, 784, 816, 838, 876, 907, 928, 951, 975, 1006, 1034, 1062, 1082, 1107, 1131, 1155, 1183, 1205, 1239, 1264, 1285, 40, 73, 99, 131, 152, 180, 206, 245, 270, 298, 330, 353, 375, 402, 421, 447, 469, 499, 521, 543, 574, 601, 623, 645, 9962, 8495]. Inventive [10109]. Inventor [1387, 1771, 4506, 4534, 11515, 11770, 123]. Inventors [1457, 1323, 1374, 2499, 2524, 1448, 1587, 1611, 1635, 1660, 1685, 1832, 1967, 2134, 2160, 2202, 2223, 2315, 2341, 2358, 2427, 2447, 2476, 2650, 2920, 3187, 3211, 3234, 3624, 3859, 5055, 415, 636, 3757, 11476, 11401]. Investigated [11797]. Investigating [9208]. Investigation [4075, 9617, 9719, 11257, 90]. Investigations [8917, 11074]. Investment [7609]. Invoice [6482]. Involved [526, 11579]. Iodide [1953]. Iona [9333, 9354]. Ionic [10675]. Ions [8634]. Iowa’s [5157]. Ira [5207]. Ireland [3395, 6100, 9269]. Irish [6603]. Iron [1893, 2327, 1436, 3506, 2623, 2907, 3463, 3892, 5229, 6091, 5951, 5946, 6310, 7154, 6436, 6801, 7903, 9366, 9012, 8614, 9329, 1079, 1303, 1072, 10382, 3954, 9215, 1103, 2760, 7476, 2329]. Iron-Ore [6436]. Ironing [10445]. Iroquois [3685, 4065]. Irrigating [2461, 3263, 7983]. Irrigation [2143, 1911, 3266, 3456, 2778, 2733, 4849, 4121, 4384, 4393, 5545, 5733, 8333, 8959, 9239, 11390, 8025, 1961, 10761, 9288, 2988, 7168, 11398, 11863, 11590]. Isham [8489]. Island [3526, 4363, 5650, 7451, 8721, 8660, 8902, 10738, 10867, 10906, 11933, 11103, 1300, 9471, 8256, 9623, 7493, 2012, 2158, 3640, 3130, 3980, 5130, 8355, 8660, 10560, 10774, 10704, 11514, 11904]. Islands [1481, 2869, 2778, 5325]. Isna [33]. Isolated [11429]. Issue [296]. Isthmian [2277, 1406, 1763, 2298, 4136, 4045, 4787, 3995, 4198]. Italia [8962]. Italian [7605, 9805, 6367, 6575]. Italy [7206, 7235, 9444, 58, 214, 9032]. Itinerary [620]. itself [2399, 7783]. IV [5049, 11460, 11828, 66, 4944, 2314]. Ivel [6654, 5282]. Ives [1332, 1536]. Ivory [5738, 9801]. Iwami [10921].

J [2514, 1370, 3364, 4561, 4934, 346]. Jab [69]. Jacinto [1586]. Jack [4193, 6697, 9904]. Jacket [7234]. Jacketing [8917]. Jacob [864]. Jade [5001]. Jail [5506]. Jam [9795, 257, 243]. Jamaican [10151]. James [5106, 9206]. Jamestown [10770, 9860, 9854, 9861, 9885, 10025]. January [9047, 10654, 9485, 3514, 4909, 6170, 9678, 10801, 1]. Japan [1517, 4748, 6528, 7087, 6618, 7704, 7415, 8989, 9441, 11900, 6692, 6198, 6757]. Japanese [1647, 2034, 3935, 7182, 6844, 6709, 6485, 6222, 7316, 6476, 6716, 6501, 6784, 6625, 6446, 8339, 7596, 7731, 8459, 7999, 9143, 8888, 9226, 11002, 10921, 11878, 11565, 4779, 9648, 6249, 6293, 9885, 9548]. Jaquet [5116]. Jar [11548]. Jars [396]. Jaubert [4416]. Java [3890, 5720]. Jaw [9932, 903]. Jawed [3120]. Jaws [11160]. Jefferson [2397]. Jellyfishes [2475]. Jena [9744]. Jentsch [5041]. Jeremiah [6416]. Jericho [855, 393]. Jerome [3129, 2845, 4185, 5354, 7581, 8766]. Jersey [8478, 8320, 11187, 11428, 11915, 10339]. Jesuit [5442]. Jesup [4702, 5189].

Jet [7079, 10846, 11237]. Jetties [11313]. Jeypore [8429]. Job [5143]. Jobbing [5858]. John [3380, 5698, 7039, 11233, 864, 1820, 10875, 757, 760, 789]. Johnston [5408]. Joining [10867, 10906]. Joint [2824, 5679, 7406, 7455, 9081, 8808, 11410, 11526, 495, 201, 1103, 563]. Jointed [8213]. Joints [2410, 3031]. Jolo [6628]. Jonas [1648]. Jones [7565]. Josef [2727]. Joseph [6010, 6482]. Joys [11904]. Juan [5623]. Judgement [45]. Jug [4419]. July [2921, 10007, 10672, 1734, 4111, 5381, 6777, 7945, 8950, 10096, 11203, 644]. Jumping [3745]. June [8378, 11633, 2842, 11400, 11830, 11086, 313, 1612, 4023, 5272, 6488, 7663, 8860, 10011, 11120, 352]. Jungfrau [5883]. Jungle [9347]. Jungner [6850]. Junk [11833]. Jupiter [3269, 6508, 6206]. Jurassic [3697].

Kaemper [340]. Kaiser [1767, 5168, 11632]. Kaiserin [1417]. Kamm [5744]. Kanave [11416]. Kansas [3381]. Kasaga [6625]. Katori [9548]. Kea [10799]. Kearsarge [1467, 2848, 6088, 8718, 10579]. Keeler [2514]. Keep [356]. Keeping [5714, 7838, 1245, 3338]. Kelvin [1557, 3907, 8828, 10655]. Kentucky [2518, 10579, 6552]. Kept [1194]. Kerosene [3504, 4809, 5256, 7857, 9542, 10193]. Kerosene-Oil [9542]. Kettle [36]. Key [8565, 10203, 7877, 11314]. Keyboard [4129, 7171, 7067]. Keyways [2176]. Killed [5481]. Killer [3247]. Killing [10799]. Kiln [6979]. Kim [1804]. Kim-Berley [1804]. Kind [3594, 6731, 8426]. King [8498, 3955, 3619, 7422, 9622, 10797, 4508]. Kinzua [2129]. Kirkwood [7827]. Kishima [9548]. Kitchen [809, 126, 564]. Kite [5329, 8704, 10737]. Kites [2056, 5327, 9702, 5933]. Kitts [1854]. Kitty [11167]. Klondike [2570, 4147]. Knabenshue [7986]. Kniaz [4350]. Knife [4395, 9467]. Knippels [363]. Knob [1027]. Knobs [4320]. Knot [2021, 2305, 2144, 2889, 4524, 5922, 7910, 9281, 9330, 10736, 10729, 11415, 11816]. Knots [1769, 9486, 10540]. Knotter [4914]. Know [7888, 11285]. Knowledge [6248]. Known [7308]. Knox [4286, 6647]. Knudsen [11154, 11584]. Königsberg [706]. Korea [2662, 6528]. Korean [5140]. Korn [10219]. Koryaks [8337]. Krebs [5501]. Kress [2981, 3117]. Krieger [2865, 10266]. Kromarograph [9674]. Kromskop [1332]. Kronprinz [3177]. Krupp [2257, 2860, 4661, 288, 3868]. Kunzite [5605]. Kwakiutl [6124].

L [4099, 5848]. Labor [2024, 9466, 10535, 526, 8302]. Labor-Saving [9466, 8302]. Laboratories [999, 3588]. Laboratory [3157, 6354, 6624, 7733, 887, 8004, 2397]. Lace [7539]. Lace-Bark [7539]. Lack [9159]. Lackawanna [6182]. Ladder [1747, 10385, 11665, 11095, 1200]. Ladders [7083]. Ladle [11615]. Lafayette [2381]. Lake [9045, 2657, 2714, 5128, 5901, 6011, 6225, 6216, 9030, 9391, 9024, 9372, 9348, 11259, 336, 10220, 2171, 1595, 2576, 2752, 5769, 7141, 7393, 7695, 10056, 579, 1072, 3843, 1888]. Lake-Dweller [336]. Lakes [5427, 2858, 8203, 10537, 1073, 4681].

Lakes-to-the-Gulf [1073]. Lambert [10464, 883, 968]. Lambs [10068]. Lamp [1731, 2284, 2522, 2006, 2306, 3204, 2716, 2693, 3574, 3389, 4762, 4035, 3962, 5763, 5480, 5089, 7257, 7868, 7426, 7356, 9542, 8904, 10150, 9707, 11139, 11214, 10978, 1307, 9526, 10342, 4239, 11717, 1238, 736, 7300, 565]. Lamp-Hanger [4035]. Lamp-Posts [3389]. Lamprey [1471]. Lamps [4216, 11429, 1770, 3040, 5264, 6105, 6776, 9782, 10895, 639, 1989, 1219]. Lance [3142]. Lance-Boats [3142]. Lancehead [11035, 11115]. Land [5260, 7316, 6741, 7136, 2727, 10612]. Landing [5459]. Lands [4234, 5432, 9539, 1071, 536, 11114]. Landslide [1317]. Lane [4282]. Langen [2054, 1538, 4167]. Langley [5741, 6141, 6204, 8568, 10313, 11534]. Langston [2061]. Lansing [4557]. Lantern [1386, 6974, 7736, 6158, 7651]. Lanterns [5672]. Laocoön [8857]. Lap [1894]. Lapge [7758]. Lapland [7112]. Larchmont [10233]. Large [1369, 2395, 1988, 2782, 3311, 3644, 3952, 4080, 4875, 5571, 5437, 5905, 6468, 6463, 8588, 9905, 10360, 10121, 10774, 10296, 459, 1534, 2690, 2235, 4879]. Larger [5390, 9441]. Largest [3113, 3596, 3833, 4351, 4526, 5789, 5604, 5040, 6203, 6807, 7211, 7473, 7362, 7358, 8772, 8921, 9252, 10041, 10759, 9843, 10542, 11116, 1076, 4978, 5214, 6919, 5721, 7193, 6379, 10877, 11151, 6399, 7935]. Last [1407, 1965, 2838, 5617, 5692, 6318, 8423, 8705, 9182, 10729, 10906, 11315]. Latch [3119, 10508]. Late [4299, 1169]. Latest [1778, 1674, 2096, 1743, 1369, 1956, 1347, 3285, 3443, 3054, 3549, 4406, 4490, 4679, 3832, 5751, 5168, 5745, 5015, 6012, 5076, 8528, 8349, 8921, 9321, 9586, 9419, 10525, 10683, 10240, 10613, 11685, 10941, 11032, 11756, 862, 10968, 4004, 5699, 9750, 4943, 9114, 9050]. Lath [10718, 10160]. Latham [851]. Lathe [2101, 2782, 4010, 5359, 3156]. Lathe-Dog [2101]. Lathes [11612]. Latin [726, 7059]. Latin-American [726]. Latitude [729]. Latrodectus [350]. Latter [5083]. Latticed [11146]. Launch [1957, 2081, 2868, 2806, 2662, 2758, 2912, 2731, 3958, 4829, 5069, 5071, 5136, 6221, 6803, 7155, 6847, 7660, 9121, 10130, 11703, 11657, 10304, 5977, 10529, 10241]. Launched [11494]. Launches [9441]. Launching [1747, 4345, 6893, 6906, 11921, 87]. Launchings [6964, 3840]. Law [3999, 7459, 7932, 7826, 8738, 10077, 8915, 7598, 4775]. Lawn [6034, 6655, 5330]. Lawrence [2331, 2217, 8168]. Laws [2886, 10158, 125]. Laying [4895, 4639, 10384, 8963, 3424, 1990]. Laysan [5983, 5958]. Laziness [7470]. Lead [4364, 3758, 6146, 8586, 11363, 10125, 1103]. Leader [1232]. Leaders [5263]. Leading [2459, 3375, 4739, 5414, 7292, 7789]. Leadless [7520]. Leaf [9150, 7133]. League [9049, 11369, 10774]. Leaky [494]. Leamington [10644]. Leap [8211]. Leap-Frog [8211]. Leaping [7547]. Leather [665, 328, 10246, 3340]. Leaves [3356]. Lebaudy [5935, 5208, 6877]. Lecture [8873, 9804]. Lecture-Room [8873]. Lectures [8089]. Leduc [895, 9733]. Lee [3871]. Leffert [1212]. Lefferts [1212]. Leg [1200]. Legal [3342, 3431, 3708, 3728, 3748, 3801, 3851, 3967, 4068, 4154, 4205, 4328, 4400, 4541, 4717, 4742, 4792, 5005, 5146, 5295, 5369, 5398, 5488, 5532, 5586, 5657, 5776, 5893, 6036, 6083, 6228, 6281, 6345, 6449, 6521, 6593, 6699, 6748, 6837, 7049, 7122]. Leggings [6695]. Legion [481]. Legislation [1594, 8610, 9577, 11746]. Legislative [1761]. Legitimate [3083]. Legs [8082]. Legumes [8589]. Lemoine [11347, 11395]. Lemon [996]. Length [6289, 8570, 11284]. Lengthening [3685]. Lens [1673, 3040, 4601, 10922, 8150]. Lenses [8893, 5048]. Lensless [7213]. Lente [8293]. Leonids [2282]. Leper [11222]. Lepine [1096]. Lesson [1409, 2932, 4612, 3846, 8446, 8159, 10921, 1765, 9568, 6224]. Lessons [2230, 4211, 5031, 7596, 8564, 8770, 11273, 11472, 670, 604, 548]. Let [11783]. Letter [6403, 8656, 11467, 8691, 8141]. Letters [2281, 4560, 9395, 1055]. Levant [4802]. Levassor [2962]. Levavasseur [9265]. Levee [4706]. Levees [11353]. Level [1843, 3725, 3799, 4578, 5879, 7416, 9110, 112, 9676, 1888, 2076]. Leveling [1627, 10386]. Lever [4146]. Lever-Device [4146]. Leviathan [922]. Lewis [8159, 7563]. Leyden [11548, 396]. Liability [10621]. Libertad [5523]. Library [2003, 4484]. Lichterfelde [3766]. Lick [3427, 6145]. Lieut. [4407, 135]. Life [2918, 2267, 4405, 4525, 5663, 5036, 6486, 6832, 7340, 6415, 8458, 8161, 8360, 9102, 9702, 10850, 438, 515, 749, 163, 595, 32, 2715, 7728, 6785, 7602, 1093, 9529, 8192, 6465, 9429, 172, 7167, 8087, 6956, 7766]. Life-Belt [7602]. Life-Line [163]. Life-Saving [2267, 4525, 6486, 6832, 9702, 10850, 595, 32, 6785, 9529, 6465, 6956]. Lifebelts [10909]. Lifeboat [7882, 7957, 9803]. Lifeboats [6913, 8595]. Lift [2484, 1705, 6797, 9149, 3228]. Lift-Lock [9149]. Lifter [7219, 9432]. Lifting [3312, 4193, 6697, 8321, 10471, 11199, 554, 6499, 11589]. Light [2367, 1828, 2497, 3466, 3008, 2797, 3087, 4892, 4739, 3943, 6164, 5902, 5314, 6334, 6667, 7339, 7805, 8587, 8796, 8805, 9545, 8789, 9119, 8790, 8673, 8654, 9126, 10022, 10038, 10259, 10553, 11240, 10944, 11722, 587, 9570, 10796, 8076, 5911, 7421, 1238, 2035, 5367, 6335, 3983, 3739, 198, 1701]. Light-Cure [5911]. Light-Telephony [5314]. Light-Weight [4739, 6667, 7805, 9545, 11722]. Lighter [6442, 9338, 10174, 9782, 11503, 426]. Lighters [1235]. Lighthouse [4641, 6537, 8619, 10361, 8683, 9054, 8405, 9483, 3360, 4615]. Lighthouses [5672]. Lighting [1906, 3530, 2991]. Lightning [2260, 3231, 4759, 5287, 8151, 8613, 10048, 1297, 858, 725, 750, 5499, 4464, 9310, 11754, 3818, 5979, 10063]. Lightning-Shattered [4759]. Lights [10809, 142, 11707]. Lightship [1368, 4806]. Lightwood [7133]. Like [11359, 10068]. Lily [6965]. Limbless [10068]. Lime [6979, 10940]. Limit [11845, 984]. Limitations [8764, 81]. Limited [4685]. Limits [2488, 5619, 9894]. Limousine [8798, 9947]. Line [1887, 2466, 3055, 2941, 3178, 2577, 2884, 5335, 5836, 8992, 8898, 8902, 10496, 10092, 11357, 163, 3843, 8591, 11064, 4671, 2933, 11293]. Lined [310]. Lineman [7453]. Linen [11147]. Liner [2423, 1417, 2119, 2662, 2731, 3177, 3198, 2940, 3642, 4680, 5761, 5168, 5617, 5113, 6030, 6847, 7211, 7535, 7383, 8190, 7474, 9330, 9486, 10769, 9791, 10736, 10729, 10319, 10005, 11816, 11494, 5977]. Liners [2002, 4679, 3734, 5665, 8408]. Lines [3842, 6464, 6870, 397, 6853, 11807].

Lining [5138]. Linoleums [10501]. Linotype [6096]. Lipe [6661]. Lippmann [4061, 10922]. Liquefying [3157, 526]. Liquid [2264, 1359, 1743, 1421, 2749, 4547, 3997, 4892, 7111, 7497, 9284, 10140, 10357, 11475, 1362]. Liquid-Air [10357]. Liquids [1896, 2603, 2898, 7750, 9993, 10161]. List [10188]. Lister [1462]. Literature [8163]. Lithium [10733]. Litigation [1719, 2775, 9564]. Little [8551]. Littlefield [8087]. Live [2556, 9137, 9948]. Lives [5287]. Living [3907, 8655, 1989]. Lizards [4318]. Lloyd [2198, 1417, 4975]. Load [7511, 7517]. Loader [6313]. Loading [2311, 4241, 8858, 11118]. Loads [7082]. Lobster [7362, 829]. Lobsters [6895]. Local [2120, 2683, 956]. Lock [1915, 3641, 4174, 4716, 5382, 6271, 7110, 9030, 9110, 8957, 9149, 9688, 10203, 10047, 11867, 10929, 10818, 11495, 11307, 628, 334, 124, 1177, 442, 904, 112, 28, 579]. Lock-and-Lake [579]. Lock-Chamber [7110]. Locke [4962]. Locking [6811, 9232, 9960, 11018]. Locks [2484, 1830, 9024, 9578, 9460, 10433, 10189, 11053, 5721, 1148, 10639, 10001, 1289]. Lockyer [3646]. Locomobile [2400, 1986, 5074, 1802]. Locomotion [590, 8411]. Locomotive [2153, 2516, 2470, 3226, 3661, 3227, 2700, 2575, 3462, 3396, 2962, 3614, 3662, 3261, 3331, 3330, 3061, 4754, 4553, 4019, 3957, 3833, 4411, 4311, 3790, 3909, 4685, 3824, 4690, 3844, 5789, 5749, 5773, 5114, 5081, 6378, 7277, 7254, 6422, 6980, 6713, 7503, 8086, 7653, 7629, 8031, 8279, 8676, 8848, 8941, 10688, 10760, 10759, 10108, 10257, 10213, 9842, 10078, 932, 1088, 8434, 10968, 11826, 5111, 11325, 9703, 11044, 2989, 8621, 4686, 6894, 7109, 7707, 11352, 7045, 6575]. Locomotive-Testing [7277]. Locomotives [1419, 1549, 1582, 1872, 2906, 3131, 3932, 4055, 4727, 4878, 4941, 5807, 5806, 5158, 6557, 7041, 6855, 6402, 6273, 7633, 8492, 7821, 8447, 7597, 8164, 7537, 8742, 8665, 8830, 8542, 8644, 10302, 10019, 11017, 11450, 312, 1092, 9207, 2216]. Locust [90]. Lodge [6010, 5100, 9216, 5112]. Loftiest [9167]. Lofty [7472]. Log [6630]. Logging [10159, 10538, 7546, 9840, 2354]. Logs [4938, 10885]. Lombroso [882]. London [1671, 7162, 1977, 1445, 2304, 2378, 2444, 2817, 2610, 2862, 3397, 3044, 3304, 3941, 5247, 5160, 5879, 5695, 6782, 7395, 660, 1135, 11640]. Long [3059, 2980, 3526, 3713, 4186, 4405, 5795, 7204, 7231, 6185, 8463, 7718, 8400, 8160, 7793, 8659, 9504, 10080, 10441, 10236, 11458, 983, 335, 303, 4779, 7068, 7133, 1658, 5235, 4363, 3980, 5130, 8721, 8660, 8902, 10704, 1300, 8256, 9623]. Long-Deferred [3059]. Long-Delayed [303]. Long-Distance [5795, 7231, 8463, 7718, 8400, 7793, 9504, 10080, 10236, 11458, 983, 1658, 5235]. Long-Leaf [7133]. Long-Maned [6185]. Long-Span [8160]. Long-Tailed [4779]. Longdistance [5383]. Longer [5663, 1134, 425]. Longest [2572]. Longevity [11078]. Longitudinal [2634, 6374]. Loomis [4368]. Looms [8058]. Loop [6012, 5184, 6033, 6576, 7889, 7372, 10149]. Loop-the-Loops [6033]. Looping [6012, 6576, 7889, 5184]. Loops [6033, 1849]. Loose [6585, 202]. Lorain [5565]. Lord [1557, 2776, 3907, 5928, 8828, 10655]. Lorillard [4299]. Lorimer [8855]. Lorraine [2119]. Loschwitz [3179]. Loss [3260, 7258, 8040, 7725, 9117, 10605, 9843, 11512, 698, 212]. Losses [3805, 6207, 7415, 7999, 10350, 9956]. Lost [8022, 487]. Lotus [3109]. Loudoun [3203]. Louis [2888, 6075, 6164, 7044, 7128, 6259, 7252, 6912, 7087, 7000, 6855, 6878, 7325, 7139, 7336, 7001, 7345, 8294, 10341, 4709, 1390, 2338, 3488, 4776, 5355, 5426, 6941, 7065, 7043, 6984, 10335, 10441, 6894]. Louisiana [4880, 4883, 6055, 5183, 6389, 6809, 7536, 10028, 9583, 6398]. Love [1551]. Low [6128, 5706, 5543, 8833, 9923, 11859]. Low-Grade [11859]. Low-Priced [6128, 5706]. Low-Tension [9923]. Lower [5238, 5213, 1984]. Lowering [2486, 11129, 10920]. Lozier [9936]. Lubricating [1201]. Lubrication [482]. Lubricator [1582, 3686, 9513, 9822]. Lubricators [2669, 9950]. Ludlow [8188]. Luise [3042, 3009]. Lumber [2523, 5507, 7978, 4807, 5748]. Lumbering [4125, 8741]. Lumbermen [6058]. Lumière [11320, 10667]. Luminous [3668]. Lunar [6600, 367, 11702, 313, 992]. Lusitania [9595, 10736, 10319, 10768, 11030]. Lusol [9526]. Luther [10948, 8027]. Luxe [9922]. Lysistrata [2890].

M [4894, 6460, 6416]. M. [1954, 1370, 1435, 3221, 3419, 3938, 3717, 6140, 7166, 7273, 6572, 1938]. M.Curie [2849]. M.D. [4099]. Macadamized [11556]. Macaroni [7357]. Machine [11546, 2007, 1480, 1780, 1539, 1831, 1333, 2126, 1580, 1708, 1533, 1558, 2195, 2178, 1889, 2603, 2738, 2736, 3339, 2694, 2866, 3015, 3598, 3548, 2850, 2553, 3010, 3253, 4513, 4639, 4734, 4241, 4705, 4876, 3724, 5861, 5860, 5938, 6056, 5234, 5494, 6051, 5722, 5853, 5743, 6165, 6096, 5254, 5747, 7343, 7143, 6688, 6627, 6890, 7017, 7145, 6543, 6834, 6277, 7256, 7062, 7215, 7727, 8143, 7684, 8287, 7580, 8364, 8502, 9445, 9405, 9530, 9375, 8988, 8925, 9191, 10335, 10794, 10395, 10384, 10093, 10064, 9857, 10356, 11198, 11242, 10994]. Machine [11092, 11760, 10837, 11824, 11080, 11883, 10981, 11540, 11283, 10835, 371, 678, 372, 104, 113, 281, 941, 11355, 6835, 11176, 10970, 1027, 9371, 10859, 9146, 8831, 9356, 7259, 8125, 8748, 10401, 5625, 7025, 8499, 9753, 5790, 466, 10198, 7770, 11825, 6360, 11090, 1122, 347, 9050, 703, 10131]. Machine-Gun [8502]. Machinery [1816, 1893, 1773, 2885, 2618, 3700, 3741, 6161, 5931, 6941, 6844, 7325, 8080, 8768, 10413, 1075, 3961, 5364, 7140, 5573, 9052]. Machines [3121, 5607, 7008, 7253, 8769, 9230, 8620, 9270, 9911, 10694, 10299, 9984, 10321, 276, 11370, 2084, 8207, 1727, 6959]. Mad [2953]. Madagascar [2517]. Madame [6996]. Maddox [4099]. Made [11546, 1907, 2212, 1776, 3864, 4058, 6046, 6354, 6380, 7543, 8689, 9543, 9213, 10205, 10501, 10423, 11924, 11714, 10611, 10027, 11862, 996, 3162, 639, 71, 1027, 96, 10758, 11618, 293, 1127, 7448, 10338, 11174, 10246, 11833, 1231, 9229, 2221, 4239, 11687, 11717, 1238, 8661, 10776, 10402, 684, 10717, 1154, 68, 5017, 663, 8622, 3156, 3043, 5442, 729, 4879]. Madison [2222, 2993, 10517, 10397, 10647, 9919, 525, 10263, 1853]. Madres [3639]. Madrid [5366]. Magazine [1381, 5857, 5313, 5386, 7952, 7567, 7891, 10256, 11087, 8277, 9333, 9354]. Magazines [5861]. Magic [9605, 1074, 897, 11499, 8919]. Magical [10775].

Magnet [7913, 554]. Magnetic [1823, 2285, 1345, 4316, 6059, 6734, 7382, 8072, 9323, 615, 956, 4735, 1871, 9798, 11328, 1779]. Magnetically [4420]. Magnetically-Supported [4420]. Magnetism [1768, 594, 964]. Magneto [4960, 9923, 480]. Magneto-Optic [4960]. Magnets [3312, 6499]. Magnified [10779]. Magnitude [5427, 10262, 10110, 11804, 1159]. Mail [1883, 2662, 2936, 4018, 6276, 6558, 6404, 7674, 8363, 10817, 10847, 266, 1021, 4238, 5239, 8682]. Mail-Bag [6404]. Mail-Box [2936, 8363]. Mails [2659]. Main [898, 801, 1990]. Maine [3481, 10581, 5333, 4504]. Mains [6257]. Mainsprings [672]. Make [3334, 2953, 5575, 7907, 850, 689, 11716, 870, 1103, 2760, 4242, 2708, 7476, 686]. Maker [10317, 457, 811, 8412]. Makers [2116, 10191]. Makes [1769, 900]. Making [1463, 1963, 2265, 3617, 3506, 3022, 3700, 4027, 4784, 3741, 4601, 4844, 5955, 5869, 6146, 5309, 5498, 5954, 6890, 6804, 6974, 8110, 8085, 7585, 8080, 8834, 8853, 8666, 9074, 11334, 10907, 11452, 11080, 11698, 11540, 10901, 11563, 966, 283, 9, 8964, 732, 8942, 9727, 7449, 392, 2690, 6691, 8455, 11545, 3961, 9672, 8090, 4415, 989, 970, 6027, 11195, 8106, 7085, 11297, 565, 1219, 4101, 9331, 4966]. Malaria [2808]. Malarial [1720, 1978, 1939]. Malcotti [6853]. Malleable [4367]. Mallet [9842]. Malting [6317, 6559]. Mammoth [4897, 5912, 8408, 11604, 7321, 6622, 6159, 340]. Man [3090, 3066, 5495, 5287, 5210, 6728, 9620, 10515, 11544, 11610, 11677, 11686, 64, 94, 197, 223, 290, 318, 394, 419, 467, 516, 561, 638, 683, 731, 1099, 1150, 1199, 1227, 1258, 296, 149, 9774, 969, 150, 484, 201, 662, 872, 688, 11374, 805, 1022, 1051, 865, 4756]. Man-Engine [1051]. Man-of-War [4756]. Manchester [6479]. Mandatory [6139]. Maned [6185]. Maneuvering [7975]. Maneuvers [7521, 2122]. Manhattan [2313, 3130, 3959, 5205, 6626, 6798, 7210, 8491, 7998, 8073, 7555, 8185, 7372, 9043, 9159, 10147, 11864, 11282, 11211, 11514, 11732, 11797, 11251, 171]. Manholes [7604]. Manila [3881]. Manipulation [2441]. Manning [7561]. Manometric [10279]. Mantel [9213]. Mantis [3067]. Mantle [11809]. Mantles [1438, 6890, 11588]. Manufacture [1706, 2334, 1605, 1679, 2199, 2058, 2471, 1678, 3574, 3545, 3207, 3287, 3158, 3464, 2573, 2804, 2967, 3064, 3091, 4465, 4441, 4219, 3889, 3981, 4668, 4440, 4530, 3940, 5306, 5952, 5598, 5950, 5849, 5953, 5949, 6423, 6381, 6463, 7903, 9329, 8723, 8850, 10656, 10380, 11705, 11860, 9902, 11588, 10924, 369, 9686, 6873, 6504, 7357, 9839, 8456, 8302]. Manufacturer [9365, 8762, 11898]. Manufacturers [8077, 8611]. Manufactures [4482, 10751, 8402]. Manufacturing [2621, 3688, 4346, 6356, 7624, 11920, 11558, 11426, 1624]. Manure [5288]. Manxman [7360]. Many [7164, 1122]. Map [1776, 5356, 6259, 920, 4409, 9728]. Maps [6804, 9470]. Mar [3140]. Marage [9900]. Marble [2917, 5524, 10527]. Marbles [11786, 8305]. Marceau [5959]. March [3099, 3706, 5531, 6189, 7922, 9104, 10248, 10821, 622]. Marconi [1956, 3153, 3350, 3550, 4825, 3696, 4832, 5478, 5564, 5610, 6099, 5107, 10541, 4200]. Mareorama [2033]. Marfels [7915]. Mari [11416]. Mari-Kanave [11416]. Maria [1417, 1897]. Marie [2774, 4529, 4100]. Marine [9507, 1582, 2233, 1527, 1537, 2209, 3364, 4677, 4187, 4561, 4432, 3716, 3952, 4290, 4163, 5986, 6527, 7936, 8100, 8349, 7471, 8544, 9321, 9122, 9009, 9120, 10235, 10683, 10170, 11627, 11165, 11054, 11840, 358, 1048, 677, 503, 768, 3333, 6311, 7299, 10992, 4752, 10242, 2102, 2616, 9004, 9641, 1420]. Mariners [10109]. Mariposa [5476]. Maritime [9746]. Mark [2034, 2275, 3540, 4775, 7459, 7826, 8738, 7598, 7932]. Market [792]. Marking [2380, 2176, 4704, 8624, 9491, 10819]. Marking-Gage [2380]. Marks [2280, 3006, 5231, 6775, 8074, 11108, 751, 2300]. Marksmanship [9766]. Marmon [8793, 10455]. Marquess [6554]. Mars [2954, 6077, 8461, 10352, 10500, 368, 251, 461, 846, 1147, 9426, 589, 10878]. Marsh [9539, 2468]. Marshal [1394]. Martin [6048]. Martinique [4964]. Marvelous [6091, 7307, 1251]. Marvels [9168]. Maryland [2266]. Mask [7234]. Masks [10176, 9251, 9267, 4852]. Masonry [1518, 7542, 9167]. Mass [555]. Massachusetts [5643]. Massage [10093]. Massena [2217]. Masses [9194]. Masted [3292, 4829, 5070]. Masterpiece [6809, 10317]. Masterpieces [682, 10089]. Mastery [2048, 8760, 10623]. Masthead [97]. Mastodon [2805]. Mastodons [10740]. Masts [11586]. Match [3353, 5764, 6442, 6587, 8140]. Matches [3158, 9902]. Material [3129, 4441, 5552, 10035, 11189, 11413, 4006, 4317]. Materials [2212, 8287, 7403, 8004]. Mathematical [11277, 6121]. Mathematics [659]. Mathushek [9217]. Matter [1345, 1883, 10314, 7651, 9619]. Matterhorn [3183]. Mauretania [9330, 10543, 390, 1195]. Max [2174]. Maxim [6267, 8858, 10716]. Maximite [2825]. Maximum [4659, 7511, 7517, 11112, 8259]. May [7698, 10077, 11898, 11014, 9141, 1520, 2718, 3896, 5148, 6384, 7529, 8710, 9868, 11003, 228, 1680]. Mayor [8656]. M’Call [61]. Me [10045]. Mean [7258, 7322]. Means [2596, 3940, 5700, 8145, 9097, 9530, 9194, 9293, 148, 7728, 7183, 3208, 11374, 1153, 943, 1553, 7840, 635]. Measles [11810]. Measure [8633, 10161]. Measured [11240]. Measurement [3143, 5064, 2441]. Measurements [9582, 6159]. Measures [4701, 10960, 11284, 11301]. Measuring [2898, 4577, 4365, 6080, 8504, 8501, 890, 395, 1094, 681, 5625, 7730, 8169, 7584, 3509]. Meat [8838, 8869, 513, 535, 2328, 9568, 9208]. Meats [10777]. Mechanical [1633, 2471, 1891, 2001, 2104, 4027, 4173, 3698, 4966, 5859, 5018, 5833, 4948, 5694, 8501, 8447, 8816, 9151, 10658, 10380, 10070, 9896, 11904, 10924, 11780, 3866, 4940, 9146, 9589, 7514, 7538, 7558, 1319, 10089, 10881, 8900, 10417]. Mechanically [11848, 5718]. Mechanics [5786, 11153, 168, 10881]. Mechanism [11758, 6163, 6166, 7091, 6760, 6273, 7694, 9376, 9907, 10094, 10800, 11762, 11159, 10979, 3913]. Mechanisms [3485]. Medal [9789, 10438, 11150]. Medalist [10088]. Media [8375]. Medical [7306, 9100, 11213, 9314, 282, 420]. Medicinal [7587, 11115]. Medicine [10873, 11320]. Medievalism [1244]. Mediterranean [3317]. Meerschaum [4777, 8723]. Meet [2394, 7834, 851, 2062, 6717, 9053]. Meeting [2233, 6846, 9896, 9977, 680, 9368, 3884, 3154, 10297, 2773, 4162, 4388, 707, 2236, 4312]. Megaphone [8280]. Melted [9306]. Melting [3743]. Melville [5569].

Members [10332, 11146, 136]. Memorial [1607, 2145, 1606, 2549, 2565, 8744, 10947, 3503]. Men [3859, 6314, 8271, 9743, 258, 4167]. Menace [6907, 78, 11656]. Mend [569, 664]. Mending [517, 868]. Mens [6463]. Mental [10404]. Mention [1143, 1168]. Mercantile [3716, 5620]. Merchant [1527, 1537, 4677, 4432, 4163, 6527, 9641, 11054, 503, 2235, 10992, 2044, 5666, 5354, 6043, 9004]. Mercurial [3018]. Mercury [1953, 7907, 7541, 9071, 7038, 2913, 7300, 5810, 10060, 10559, 111]. Meredith [9375]. Mergenthaler [3741]. Meridian [11597, 8547, 10381, 7606]. Mersey [922]. Mescal [5649]. Message [2954, 6126, 9522]. Messages [1956, 5434, 8849, 11110, 9031]. Messina [79]. Metal [2738, 2872, 4058, 6914, 7317, 7144, 7737, 949, 11717, 11615, 11608, 565]. Metallic [1318, 5900, 6184, 9988, 10227, 9768, 11630, 2355, 868]. Metallographs [11608]. Metallurgical [1962, 2057]. Metals [1847, 4550, 6392, 7513, 9793, 10216, 1197, 11038, 8923, 9306, 10261, 6054]. Metchnikoff [11078]. Meteor [11039, 4287]. Meteorite [4527, 7937, 7188]. Meteorites [6181]. Meteorological [2077, 7132, 8378, 9047, 10521, 10852, 11060, 11430, 11633, 7112, 3406, 7698]. Meter [2307, 5054, 5338, 6744, 7279, 9826, 9757]. Meters [4024, 1177]. Method [2917, 1414, 2130, 2351, 3587, 2533, 2638, 3530, 2942, 2613, 2621, 2780, 4416, 3699, 4844, 3834, 4666, 4261, 5309, 5542, 5437, 7344, 6316, 7206, 6804, 6995, 7930, 8147, 8322, 8437, 7722, 8743, 9292, 9261, 9364, 9111, 10049, 10922, 11173, 11563, 436, 395, 1234, 1094, 5626, 852, 10152, 9031, 8571, 11545, 6440, 7939, 11260, 7651, 5083, 8985, 643, 324, 323, 7546, 1104, 8088, 565, 1894]. Methods [2169, 1629, 1610, 2444, 3563, 3544, 2862, 3361, 3835, 4335, 5846, 5498, 5428, 8162, 7706, 10300, 11241, 1598, 431, 8388, 9029, 10277, 11286, 10592, 4196, 7938]. Metric [2322, 4573, 3977, 5228, 8978, 10078]. Metropolitan [2495, 4436, 6690, 458]. Mexican [917, 4443, 10629]. Mexico [5001, 11239, 2798, 4873, 5649, 7086, 7116, 2474]. Miami [5595, 6421]. Mice [8891]. Michaud [6443]. Michel [5624]. Michigan [9586, 10596, 658, 11175]. Micro [11861, 9601]. Micro-Electroscope [9601]. Micro-Organisms [11861]. Microbe [11902]. Micrography [1397]. Micrometer [11359, 10611]. Microphone [2567]. Microphotography [8232]. Microphotoscope [9470]. Microscope [10690]. Microscopic [7340, 3960]. Mid [4845]. Middle [1078]. Miethe [7063]. Mikasa [3935, 7327, 6294, 8040, 10256]. Mikkelsen [8566]. Mile [2990, 7630, 8652, 10105, 10061, 4492, 435]. Miles [5732, 6998, 1080, 1270, 9113]. Military [1752, 1806, 2100, 3315, 3550, 2538, 4848, 6337, 6550, 6651, 7438, 7930, 8701, 9216, 10709, 11776, 10996, 11900, 11938, 581, 158, 9883, 7521, 7939, 1729, 10860, 10378, 9648]. Militia [4783]. Milk [6046, 7275, 7722, 7708, 8918, 10123, 10020, 10902, 11920, 11356, 1245, 537, 127, 441, 4757, 225]. Milk-Bottle [537]. Milker [10444]. Milking [7275, 7145, 7580, 9405]. Mill [2395, 2665, 8286, 9655]. Miller [8792]. Milligrade [1372]. Milling [6763, 10993]. Million [7836, 9074, 10950]. Mills [2754, 9629, 4643, 4504, 4969]. Milwaukee [614, 6893, 918]. Mimicry [6879]. Minas [11752]. Mind [5357, 11916]. Mine [1487, 2327, 3964, 6500, 6364, 7632, 8526, 9162, 11639, 11319, 1040, 508, 10734]. Mined [1072]. Mineola [1700]. Miner [2995]. Mineral [3587, 5598, 6572, 9541]. Minerals [7134, 11698, 4051]. Miners [10641, 11541, 1059]. Mines [1907, 1804, 2024, 1491, 1632, 5309, 6755, 6496, 7731, 9552, 9097, 9645, 10039, 10028, 11063, 1047, 1072, 1442, 9583]. Ming [7427]. Miniature [3388, 10239, 10, 7475, 11064]. Mining [2078, 1610, 2463, 1962, 3508, 3305, 3240, 4388, 3758, 7057, 8295, 11317, 11241, 1819, 8497, 1941, 3251, 2864, 2310, 1727, 2216, 5019, 10284, 7938, 10043]. Mining-Cars [2078]. Minnequa [9215]. Minnesota [5113, 6807]. Minnie [4780]. Minot [3679]. Mint [7001]. Minute [7381, 9470]. Mirages [315]. Miring [3381]. Mirograph [2965]. Mirror [1514, 222, 10563]. Mirrors [368]. Miscellaneous [4979, 5027]. Mishap [1805, 3631]. Misrepresentations [524]. Miss [11169]. Mission [10482]. Missions [4168]. Mississippi [2460, 1369, 2685, 7169, 10595, 11313, 365, 11353, 1937, 8785, 8052]. Missouri [3958, 6352]. Missouri’s [5739]. Mitchell [6650]. Mivart [1460]. Mixed [10443]. Mixer [7737, 11265]. Mixing [4104, 5142, 9291, 8768]. Mixture [9749]. Mixtures [2984]. M’Master [4872]. Mobile [1873, 9428]. Model [1563, 2844, 6195, 7000, 6915, 6918, 7959, 8404, 9984, 10313, 341, 7733, 5628, 9568, 10592]. Modeling [4202, 6579]. Models [2011, 1984, 4288, 5066, 5309, 5073, 7340, 7138, 226, 11803, 5647, 4709, 11640, 10779, 9606]. Moderate [459]. Moderate-Sized [459]. Modern [1458, 1381, 1410, 1621, 3663, 3656, 2566, 3987, 4833, 5855, 5663, 5846, 5036, 7170, 6763, 6855, 6205, 6444, 7629, 8273, 7706, 7495, 8655, 8657, 10300, 9708, 11369, 11055, 1076, 1244, 9962, 9607, 7115, 11440, 8495, 2689, 4438, 8433, 10403, 7171, 9251, 9267, 9189, 2310, 8527, 9357, 4083, 5019, 5135, 6135, 11220, 7661, 6833, 777, 4998, 7938]. Modification [9048]. Modifications [4898]. Moines [2843]. Moissan [1954, 5572]. Moistener [2079]. Moistening [6543]. Moldboard [3769]. Molding [3309, 8287, 11825, 11090]. Molds [11195]. Molecules [9073]. Mollusk [7753]. Molten [3463]. Moment [9901]. Mommsen [5867]. Monaco [9468, 262]. Money [1789, 3162]. Monitor [2081, 4669, 11217]. Monitors [3483, 4084, 10588]. Mono [2054, 2093, 1564, 263, 1018, 4671, 9389]. Mono-Rail [2054, 2093, 1564, 263, 4671, 9389]. Mono-Railroad [1018]. Monocycle [6739]. Monolith [5359, 11012]. Monoplane [11829, 850]. Monoplanes [261]. Monorail [2701, 10023, 11324]. Monoxide [11598]. Monreale [3684]. Monster [10153, 9569, 5838, 3988, 4151, 8413]. Monsters [9188]. Mont [4316, 5932, 7579, 4575, 4075, 5624, 8471, 4853]. Montagu [10343]. Montauk [10303]. Montcalm [2537]. Monte [6382]. Montgomery [7631]. Months [7900, 10416, 10276, 9792]. Montreal [5442, 4996, 1145]. Monument [856]. Monumental [5968, 7765, 11674]. Monuments [3291, 11913, 6719]. Moon [5700, 6248, 8360, 9964, 1223, 10672, 9268]. Moore [7903]. mooring [2061]. Mop [4421]. Mop-Wringers [4421]. Moral [2190]. Morehouse [11456, 593, 289]. Morgan [5620, 1370]. Morning [5848, 8592, 10199, 11327, 30]. Moro [8412]. Morocco [3796]. Morris [4600]. morrow [10173]. Mors [5928]. Morse [5433]. Mortar [11475]. Mortising [2850]. Mosaic [1652]. Mosely [5322]. Mosquito [1720, 5999, 5130, 7018, 8451, 8703, 212, 3155, 6929]. Mosquitoes [1976, 10942, 1939]. Most [2470, 2575, 4411, 5114, 8031, 10197, 10108, 11752, 1076, 1098, 312, 822, 3552, 11065]. Moth [2107, 5022, 8567]. Mother [6478]. Motion [2618, 2730, 2557, 4201, 4025, 4969]. Motive [6237, 9486, 9034]. Motogodille [9116]. Motor [10812, 2302, 1914, 1683, 1806, 1322, 2335, 2333, 2492, 1650, 2847, 3504, 3232, 3288, 3575, 3951, 4636, 4882, 3798, 3719, 4140, 3887, 4369, 4034, 5233, 6157, 5086, 5582, 5651, 5637, 5751, 6164, 5076, 6100, 5257, 4965, 5928, 5334, 5971, 6013, 5607, 5608, 6382, 6885, 7084, 6195, 6654, 6464, 6812, 6265, 7064, 6580, 7007, 6738, 6220, 8463, 8525, 7814, 7679, 7795, 7469, 7801, 7490, 8241, 7942, 7987, 8318, 7940, 7802, 8216, 7805, 7790, 9249, 8792, 9504, 9312, 8788, 9116, 9115, 8686, 9049, 8795, 8794, 9123, 9240, 8638, 9118]. Motor [9120, 8897, 8924, 9924, 10239, 10197, 10496, 9770, 10457, 10243, 10455, 10236, 9987, 10576, 10483, 10083, 10478, 10042, 10061, 10234, 11179, 10996, 11906, 11932, 11193, 11280, 11495, 10855, 11895, 11568, 146, 37, 583, 1276, 994, 1306, 262, 1242, 2440, 10425, 9165, 9124, 9750, 8746, 5282, 7735, 463, 200, 3595, 2291, 2314, 2340, 9310, 11808, 2468, 1202, 4583, 10778, 3702, 10466, 11374, 661, 9803, 483, 10025, 3510, 10626, 10263]. Motor-Boat [6382, 6885, 6195, 6812, 7064, 6580, 8463, 7679, 9123, 10236, 10234, 11179, 11193, 1276]. Motor-Car [11495]. Motor-Driven [5651, 4965, 5607, 9987]. Motor-Drlven [6013]. Motor-Paced [4882]. Motor-Propelled [7469]. Motor-Sleighs [146]. Motor-Wheel [1322]. Motoring [637]. Motorman [4361]. Motormen [2642]. Motors [2591, 3137, 4619, 3756, 5167, 6647, 7082, 6734, 7111, 8653, 9822, 10482, 11082, 11722, 509, 2578]. Motsinger [3760]. Mound [6250]. Mount [8053, 3570, 8435, 854, 7302, 5251]. Mountain [11839, 833, 11197, 11564, 7914]. Mountaineering [11373]. Mountaineers [9032]. Mountains [1917]. Mounting [5236, 6985, 8617, 10756, 9229, 11709, 8413]. Mounts [3485]. Movable [11093, 28, 6339]. Movement [9181, 11366]. Movements [5600, 10881]. Mover [2737]. Movers [3036]. Moving [2508, 2990, 2684, 2775, 4080, 4875, 3781, 3699, 5698, 5960, 5434, 7577, 7583, 7501, 8944, 10303, 10331, 833, 8, 283, 796, 457, 587, 5259, 9054, 4196]. Mower [3683, 6034, 6655, 5330]. Mr. [5498, 8185, 10148, 11128]. Much [5495, 8296]. Mud [1226, 5357, 496, 5720]. Mud-Puddle [5357]. Mud-Volcanoes [5720]. Muddle [2678]. Muffler [4368, 813]. Muffling [6945]. Muir [1934]. Muirhead [5100, 9216, 5112]. Mukden [7427]. Müller [2174]. Multi [7429, 287, 512]. Multi-Firearms [7429]. Multi-Hull [287, 512]. Multiface [6096]. Multiple [9943, 10480, 11522, 11658, 8359, 6852, 11825]. Multiple-Disk [9943]. Multiple-Unit [10480, 6852]. Multiplex [4510]. Mummies [1422]. Mummification [1681]. Municipal [3677, 4596, 6932, 10021, 753, 9568, 11902, 9724]. Municipalities [10946]. Munn [9701]. Murray [2732, 4211, 9169]. Muscles [6756]. Museum [1562, 7937, 7853, 9468, 682, 3988, 8305, 8597, 10003, 659, 6956, 2149, 2661, 9606, 11016, 2805]. Mushroom [4920, 6636, 9908]. Music [8637, 10106, 9705, 11823, 9674]. Music-Recording [9674]. Musical [7003, 2426, 3111, 947, 10472]. Musk [1129]. Must [7204]. Mutes [10044]. Mutoscope [2618, 10994]. Mutual [6601, 10881, 1412]. Mutually [1643]. My [7911, 4372]. Myer [11434, 1255, 1217]. Mysteries [3151, 8315]. Mysterious [10607]. Mystery [7371, 9841, 11885].

N [6140, 7153, 6289, 6756, 6995, 7293, 6217, 6495, 6437]. N-Ray [6437]. N-Rays [6140, 7153, 6289, 6756, 6995, 6217, 6495]. N.H [8164]. N.J. [2198]. N.Y [7698, 8378, 9047, 10521, 10852, 11060, 11430, 11633]. N.Y. [2217, 8164, 7089, 3067]. Naga [9774]. Naga-Ed-Der [9774]. Nail [2195, 3698, 5954, 9846, 398]. Nail-Puller [3698]. Nailless [8301]. Nails [6977]. Name [8609, 4409]. Naming [5903]. Naosaurus [9882]. Naphtha [1957]. Napier [6812]. Napkin [11547]. Narragansett [5040]. Narrow [6456]. Narrow-Gage [6456]. Narrowing [11479]. Narwhal [6919]. Nasmyth [11515]. Natal [5407, 2133]. Nation [1840, 4037]. National [1597, 2149, 7065, 10604, 10517, 10555, 10536, 9919, 10811, 11940, 8299, 2417, 2145, 1606, 7297, 7803, 9345, 10456, 1506, 2703, 3421, 4637, 3884, 5181, 7557, 8740, 11036, 6578, 2236, 9462, 3588]. Nations [8342, 2182]. Native [6878]. Natives [2133, 1803]. Natural [1367, 3348, 3437, 4642, 4977, 5306, 6806, 7937, 7853, 10267, 9687, 11128, 11102, 769, 796, 407, 3988, 10443, 6719, 7981, 3577, 10026, 8597, 8336]. Natural-Size [7981]. Nature [6985, 9531, 10045, 11370, 1251, 5048]. Nautical [5956]. Nautilus [836]. Navahos [1469]. Naval [2233, 1409, 1975, 2122, 3303, 3738, 4899, 4048, 4255, 6238, 6533, 7419, 7999, 8082, 8916, 9766, 10590, 10990, 11130, 1211, 1187, 381, 1066, 10591, 7886, 9333, 9354, 7088, 7755, 10673, 4903, 2352, 7831, 1479, 1920, 2312, 1385, 2072, 3521, 3282, 3094, 3004, 3542, 3476, 3493, 2565, 3477, 2889, 3362, 3635, 4660, 4547, 3931, 4074, 4904, 4728, 3736, 5866, 5909, 5021, 6501, 6784, 6705, 6754, 6683, 7717, 8048, 8209, 10274, 10665, 10573, 11490, 11343, 630, 822, 502, 1115, 9646, 8745, 6477, 11523]. Navies [3375, 6618, 7648, 610]. Navigation [3927, 8701, 10217, 10792, 8109, 9746]. Navipendulum [4666]. Navy [1921, 1395, 2103, 5252, 7542, 8532, 10490, 10304, 3426, 1384, 2278, 1360, 2392, 2840, 3244, 3270, 2896, 3527, 3417, 3394, 3474, 3586, 2702, 3475, 2633, 4459, 4703, 4480, 4160, 5101, 5745, 5405, 5204, 6709, 6397, 6716, 6737, 6964, 8339, 7729, 9304, 9351, 9190, 8548, 8824, 8888, 9226, 10624, 11557, 10856, 10832, 11460, 11886, 11343, 10938, 11769, 11231, 11261, 11295, 10915, 11010, 11318, 11565, 11056, 894, 862, 238, 1086, 981, 156, 617, 8379, 115, 11191, 7858, 10503, 11774, 5903, 10112, 10358]. Near [6679]. Nearly [5700]. Nebraska [1583, 9585, 11865]. Nebula [3459]. Necessary [10681]. Neck [4814, 8128]. Neck-Yoke [4814]. Need [1597, 4231, 4774, 4311, 4480, 5223, 7253, 8120, 10827, 10989, 8915, 9286]. Needed [5900, 5497, 6930, 10000]. Needle [11617]. Needles [2846, 9463]. Needless [4843]. Needs [1621, 11166]. Nefs [4267]. Negative [10204]. Negatives [3681]. Negras [5676]. Nemethy [3598, 7025]. Nemi [9372, 7141]. Nernst [4216, 2522, 3962]. Nerve [6731, 9140]. Nerves [6756]. Nervous [7730]. Nest [7239]. Nests [10562, 4582]. Net [569]. Nets [15]. Netting [9910]. Never [11898]. Newark [6182]. Newcastle [4263]. Newcastle-On-Tyne [4263]. Newcomb [1169]. Newfoundland [10043]. Newly [3295, 10825, 5676]. Newly-Discovered [3295]. Newly-Opened [10825]. Newmastic [10475]. Newport [2122, 1346, 3184, 11578, 11747, 7886]. News [1555, 6093, 5043, 5735, 7156, 1186]. Newspaper [5847, 5856, 10082, 5185]. Newspapers [11924, 9484]. Next [4255, 5972, 5050, 9161, 296]. Niagara [4667, 7493, 8521, 7984, 1706, 2014, 2106, 2402, 4454, 4659, 5324, 8146, 8468, 8481, 7833, 7862, 7488, 8582, 8616, 8697, 10211, 11626, 11210, 3230, 3425, 5477, 5499, 5037, 8236, 8283, 8591, 9248, 10421, 257, 31, 243, 2664, 6786]. Niasin [6625]. Nicaragua [1840, 1548, 1843, 1595, 3501, 4148, 1888]. Nice [3419, 4032, 6382]. Nichols [10644]. Nickel [9012, 9645, 5827]. Nickel-Iron [9012]. Nickel-Steel [5827]. Night [3313, 3842, 1192, 4239]. Nile [1574, 3266, 2733, 4384, 4393, 5545, 9815, 33, 704]. Nilgiris [2037]. Nine [3088]. Nine-Ton [3088]. Nineteenth [2121]. Ninety [6498]. Ninety-Fifth [6498]. Ninth [6117, 525, 1096]. Nippur [2003, 3116, 9731]. Nitrates [9253]. Nitro [2725]. Nitrogen [9243, 10399]. Nitroglycerine [2725]. Nivernais [4052]. No [1395, 2103, 3221, 3250, 3806, 4028, 4262, 4711, 8127, 8897, 9879, 11829, 11635, 11905]. No-Segment [11905]. Nobel [2074, 6005]. Noble [5304]. Nodes [1849]. Noise [11410, 233, 5979]. Noiseless [78]. Noises [2410, 3031]. Nome [1552, 1625, 1610, 1819]. Nomenclature [3811]. Non [1440, 4664, 5646, 7231, 6574, 8472, 7807, 8946, 8574, 9628, 8802, 9430, 11630, 11868, 958, 1939]. Non-Adjustable [8802]. Non-Freezing [4664, 5646]. Non-Malarial [1939]. Non-Metallic [11630]. Non-Poisonous [8472]. Non-Refillable [1440, 9628, 9430, 11868]. Non-Skid [7807, 8574]. Non-Skidding [8946]. Non-Stop [7231]. Non-Toxic [958]. Noon [7322]. Nordenfeldt [2734]. Nordenskiold [6070]. Nordenskjöld [3658]. Norfolk [10304]. Normal [4898]. Norman [3646]. North [9204, 5189, 2198, 1417, 3291, 2723, 5068, 6235, 7163, 8252, 8071, 8072, 8373, 9392, 9148, 10597, 11703, 11818, 11042, 10973, 11135, 728, 709, 696, 345, 1269, 93, 191, 1012, 967, 831, 9816, 7304, 11065, 10808, 9817]. Northeastern [4788]. Northern [2665, 1872, 3331, 7798, 9926, 9949, 5529]. Northwest [8350, 835, 10538, 4376, 2354, 8323, 10283, 10175]. Norwegian [4525]. Notable [1750, 6484, 7043, 11042, 10798]. Note [9871, 11487]. Noted [1387, 5600, 9006]. Notes [1313, 1328, 1925, 1948, 1971, 1995, 2018, 2041, 2066, 2090, 2112, 2139, 2163, 2187, 2206, 2226, 2252, 2271, 2295, 2318, 2345, 2363, 2386, 2405, 2431, 2452, 2479, 2505, 1356, 1377, 1403, 1426, 1451, 1475, 1499, 1523, 1545, 1569, 1591, 1616, 1638, 1664, 1690, 1715, 1738, 1758, 1784, 1811, 1837, 1858, 1880, 1901, 1461, 1508, 1531, 1554, 1649, 1698, 1936, 2194, 2372, 2490, 1315, 1330, 1927, 1950, 1973, 1998, 2020, 2043, 2069, 2092, 2115, 2142, 2166, 2189, 2208, 2229, 2254, 2274, 2297, 2321, 2347, 2365]. Notes [2388, 2407, 2433, 2454, 2482, 2507, 1358, 1380, 1405, 1428, 1453, 1477, 1502, 1526, 1547, 1571, 1593, 1619, 1641, 1667, 1692, 1717, 1742, 1760, 1786, 1813, 1839, 1860, 1882, 1903, 1982, 2214, 2259, 3342, 3431, 2527, 2541, 3126, 3147, 3168, 3190, 3215, 3237, 3257, 3280, 3299, 3323, 3345, 3371, 3393, 3416, 3434, 3454, 3473, 3499, 3519, 3537, 3558, 3582, 3605, 3627, 3650, 3673, 2562, 2587, 2606, 2629, 2652, 2675, 2697, 2720, 2744, 2768, 2789, 2812, 2834, 2854, 2878, 2903, 2926, 2948, 2974, 2999, 3028, 3051]. Notes [3075, 3103, 2529, 2543, 3149, 3170, 3192, 3217, 3239, 3259, 3301, 3373, 3452, 3539, 3560, 3584, 3607, 3630, 3652, 3675, 2564, 2589, 2609, 2631, 2654, 2677, 2699, 2722, 2747, 2791, 2814, 2836, 2857, 2880, 2905, 2928, 2951, 2977, 3002, 3030, 3053, 3077, 3105, 3128, 2583, 2671, 3566, 3590, 3612, 3636, 3659, 2741, 2786, 2810, 2832, 2875, 2900, 2922, 2945, 2970, 3023, 3047, 3072, 3135, 3200, 3222, 3246, 3308, 3329, 3357, 3379, 3402, 3439, 3460, 3546, 4279, 4123, 3708, 3728, 3748, 3801, 3851, 3967, 4068, 4154]. Notes [4205, 4328, 4400, 4541, 4717, 4742, 4792, 3694, 3712, 4308, 4333, 4358, 4381, 4403, 4431, 4450, 4478, 4496, 4519, 4543, 4568, 4592, 4610, 4630, 4654, 4673, 4698, 4723, 4746, 4772, 4796, 4822, 4841, 4863, 4890, 3733, 3752, 3779, 3804, 3831, 3855, 3878, 3900, 3925, 3950, 3971, 3994, 4017, 4044, 4073, 4096, 4116, 4135, 4157, 4183, 4209, 4229, 4252, 4272, 3732, 3753, 4594, 4632, 4656, 4726, 4745, 4771, 4797, 4821, 4842, 4864, 4889, 3830, 3854, 3877, 3901, 3924, 3949, 3970, 4016, 4043, 4072, 4117, 4182, 4210]. Notes [4228, 4251, 4273, 4307, 4331, 4360, 4382, 4430, 4451, 4479, 4567, 3718, 3759, 3813, 3839, 3861, 3885, 3934, 3956, 4001, 4217, 4237, 4258, 4425, 4446, 4474, 4516, 4540, 4590, 4651, 4695, 4720, 4816, 4859, 4885, 5005, 5146, 5295, 5369, 5398, 5488, 5532, 5586, 5657, 5776, 5893, 6036, 6083, 5027, 4925, 4953, 5659, 5686, 5709, 5728, 5755, 5778, 5819, 5842, 5895, 5916, 5941, 5964, 5992, 6017, 6038, 6085, 6110, 6131, 6152, 4985, 5008, 5058, 5096, 5123, 5152, 5172, 5199, 5219, 5243, 5297, 5318, 5372, 5400, 5425, 5451]. Notes [5467, 5490, 5514, 5534, 5556, 5589, 5614, 5634, 4912, 4955, 5712, 5731, 5758, 5781, 5822, 5843, 5876, 5898, 5919, 5944, 5967, 5995, 6019, 6041, 6067, 6113, 6134, 6154, 4987, 5010, 5057, 5098, 5126, 5155, 5175, 5202, 5222, 5245, 5276, 5300, 5321, 5353, 5374, 5403, 5423, 5452, 5493, 5516, 5591, 5636, 5662, 5689, 4937, 4950, 5829, 5881, 5904, 5927, 5975, 6024, 4982, 5003, 5093, 5120, 5149, 5182, 5216, 5253, 5281, 5294, 5316, 5348, 5368, 5396, 5419, 5463, 5511, 5596, 5622, 5644, 5669, 5717, 5740, 5788, 6825]. Notes [6910, 6933, 6639, 6228, 6281, 6345, 6449, 6521, 6593, 6699, 6748, 6837, 7049, 7122, 6172, 6191, 6793, 6817, 6839, 6863, 6901, 6923, 6951, 6989, 7031, 7051, 7074, 7098, 7124, 7148, 7176, 7200, 7225, 7245, 7268, 7288, 7312, 7333, 6231, 6253, 6283, 6302, 6328, 6348, 6370, 6386, 6410, 6430, 6451, 6471, 6491, 6523, 6547, 6565, 6595, 6614, 6635, 6676, 6701, 6724, 6750, 6770, 6175, 6194, 6842, 6866, 6904, 6926, 6970, 7015, 7034, 7054, 7077, 7100, 7127, 7151, 7179, 7203, 7227, 7248, 7270, 7291, 7315, 7335, 6214]. Notes [6234, 6256, 6286, 6305, 6330, 6351, 6372, 6412, 6432, 6454, 6474, 6494, 6526, 6549, 6567, 6598, 6616, 6678, 6704, 6727, 6752, 6773, 6820, 6199, 6239, 7120, 7160, 7173, 7197, 7223, 7265, 6262, 6291, 6312, 6336, 6344, 6377, 6418, 6439, 6531, 6553, 6621, 6733, 6758, 6779, 6802, 6872, 6887, 6948, 7011, 7028, 7047, 7071, 7108, 7599, 7752, 7956, 7350, 7367, 7971, 7995, 8012, 8035, 8066, 8096, 8115, 8134, 8155, 8198, 8221, 8245, 8266, 8309, 8329, 8345, 8368, 8418, 8442, 8486, 8513, 7389, 7412, 7434, 7463, 7484]. Notes [7508, 7531, 7551, 7572, 7592, 7617, 7641, 7666, 7688, 7713, 7742, 7760, 7778, 7818, 7844, 7871, 7896, 7925, 7948, 7352, 7370, 8069, 8098, 8117, 8136, 8158, 8182, 8201, 8224, 8247, 8269, 8292, 8311, 8348, 8371, 8396, 8421, 8445, 8467, 8488, 8516, 7392, 7437, 7466, 7487, 7510, 7533, 7553, 7575, 7595, 7644, 7668, 7691, 7716, 7746, 7762, 7847, 7874, 7898, 7928, 7951, 7974, 7997, 8015, 8038, 7420, 7442, 8354, 8403, 8427, 8453, 8474, 8496, 8522, 7494, 7516, 7559, 7626, 7652, 7673, 7699, 7768, 7788, 7828]. Notes [7934, 7979, 8024, 8103, 8122, 8142, 8165, 8187, 8230, 8254, 8276, 8297, 8316, 8936, 9327, 8536, 8559, 9257, 9277, 9299, 9317, 9341, 9361, 9384, 9409, 9436, 9456, 9476, 9497, 9517, 9535, 9557, 9611, 9636, 9660, 8578, 8605, 8648, 8670, 8712, 8729, 8755, 8777, 8819, 8843, 8863, 8883, 8909, 8928, 8953, 8973, 8998, 9020, 9039, 9060, 9087, 9107, 9177, 9198, 8539, 8562, 9158, 9179, 9201, 9223, 9238, 9259, 9280, 9302, 9320, 9343, 9363, 9387, 9412, 9439, 9458, 9478, 9500, 9520, 9538, 9559, 9576, 9614, 9638, 9663]. Notes [8581, 8608, 8631, 8650, 8672, 8696, 8732, 8758, 8780, 8821, 8846, 8866, 8886, 8911, 8931, 8956, 8976, 9001, 9023, 9041, 9063, 9089, 9109, 9135, 8546, 8576, 8681, 8786, 8829, 8982, 9028, 9070, 9094, 9266, 9274, 9370, 9421, 9505, 10669, 9871, 9680, 9693, 10271, 10288, 10308, 10346, 10369, 10389, 10409, 10486, 10511, 10569, 10617, 10635, 10699, 10725, 9715, 9738, 9762, 9786, 9808, 9829, 9849, 9889, 9914, 9970, 9996, 10014, 10031, 10052, 10073, 10098, 10117, 10143, 10165, 10184, 10208, 10230, 10251, 10237, 9682, 9695, 10291, 10311, 10329, 10349]. Notes [10372, 10391, 10412, 10432, 10449, 10489, 10514, 10534, 10552, 10572, 10602, 10620, 10638, 10663, 10680, 10702, 10728, 10746, 10766, 10788, 9718, 9741, 9764, 9788, 9811, 9832, 9852, 9873, 9892, 9973, 9999, 10017, 10034, 10055, 10076, 10102, 10120, 10146, 10168, 10187, 10210, 10232, 10253, 10273, 9726, 9980, 10040, 10218, 10278, 10298, 10316, 10355, 10497, 10520, 10539, 10608, 10646, 10735, 10754, 10773, 11148, 11246, 11487, 10803, 10840, 11671, 11694, 11743, 11765, 11791, 11852, 11871, 11891, 11909, 11928, 10863, 10891, 10953, 11006, 11025, 11049, 11070, 11098, 11123]. Notes [11206, 11227, 11289, 11337, 11381, 11405, 11421, 11443, 11470, 11507, 11530, 11552, 11593, 11621, 11647, 10806, 10824, 11384, 11407, 11424, 11446, 11624, 11767, 11814, 11838, 11855, 11874, 11894, 11912, 11931, 10843, 10866, 10894, 10913, 10934, 10956, 10985, 11009, 11028, 11052, 11073, 11101, 11126, 11143, 11163, 11182, 11209, 11230, 11249, 11271, 11292, 11312, 11340, 11362, 10967, 11172, 11298, 11394, 11477, 11773, 11821, 11880, 11917, 11939, 493, 60, 173, 531, 3, 41, 1132, 1206, 74, 100, 132, 181, 207, 246, 271, 299, 376, 403, 422, 448, 544].

Notes [575, 624, 646, 694, 785, 839, 952, 976, 1007, 1035, 1108, 19, 43, 668, 691, 712, 739, 765, 787, 818, 841, 878, 909, 930, 954, 977, 1009, 1037, 1064, 1083, 1110, 1184, 1208, 1240, 1267, 1287, 76, 102, 133, 153, 182, 209, 229, 248, 273, 301, 331, 354, 379, 406, 423, 450, 470, 500, 522, 547, 577, 603, 627, 648, 9503, 11631, 4686, 11882]. Notice [9901]. Notices [4589, 4766]. Nots [10045]. Nouveau [3530]. Nova [62, 3459]. Novel [1683, 2285, 2439, 1391, 2053, 2195, 1870, 1556, 2335, 2152, 2218, 1732, 2032, 3112, 3615, 3571, 3274, 2739, 3068, 2691, 2936, 3021, 3686, 2602, 2623, 3139, 4143, 3834, 4512, 4166, 3783, 4002, 3797, 4033, 3887, 6127, 5500, 6104, 5670, 5081, 5077, 7194, 7208, 6762, 6888, 6797, 6512, 6834, 6961, 6404, 6804, 7002, 6889, 6743, 7279, 6292, 6437, 7612, 7560, 8057, 8091, 7586, 7920, 8406, 8322, 7940, 7708, 7891, 8479, 8145, 7660, 7736, 7802, 8362, 7964, 8216, 9338, 8801, 9603, 9404, 9015, 8688, 8904, 8750, 9630, 9466, 8724]. Novel [9525, 10324, 9943, 9935, 9688, 10479, 9778, 9846, 10476, 9938, 10093, 9955, 9990, 10047, 10095, 9711, 10782, 10814, 11761, 11309, 11238, 11668, 10944, 11948, 864, 804, 1049, 490, 708, 283, 37, 388, 175, 285, 497, 595, 492, 1220, 11823, 3403, 5330, 5625, 5388, 8105, 7421, 7829, 7601, 9029, 10114, 1282, 6367, 6892, 9129, 6940, 200, 7546, 9314, 9355, 3942, 11806]. Novelties [2357, 2785, 2943, 3095, 3800, 4302, 5084, 7515, 8340, 8803, 9951, 10468, 489]. Novelty [9077, 310, 4505, 4504]. November [2135, 3320, 4515, 5800, 6975, 8178, 9296, 10428, 11575, 875]. Novik [6764]. Noxious [6517]. Nozzle [6073, 11093]. Nuisance [3609]. Number [3521, 4257, 5050, 9395, 10936, 799, 800, 830, 859, 921, 1146, 1299]. Numbers [10905, 11587, 11777, 59, 592]. Nurseries [10562]. Nut [1915, 3641, 4174, 904, 487]. Nut-Lock [1915, 3641, 4174]. Nutritive [9208]. Nuts [5716]. Nutshell [8585].

O [11519, 4643]. Oak [6765]. Oakland [3225]. Oar [124]. Oars [7561, 8213]. Obelisk [5932]. Object [2230, 3846]. Objections [3786]. Observation [5327, 8026, 11309]. Observations [5186, 7650, 10352, 729, 6145, 5810, 8835, 6118, 1680]. Observatories [7132]. Observatory [3427, 7827, 7771, 8429, 10904, 7755, 8053, 4903, 2077, 3376, 10381, 6145, 7088, 9646, 2913, 3523, 4245, 7734]. Observe [1050]. Observed [5700, 11456]. Observing [6995, 8107]. Obstacle [8411]. Obstacles [1429, 4918]. Obstructions [1396]. Obtain [11527]. Obtaining [194, 7183]. Occur [9604]. Ocean [2048, 3132, 4845, 5665, 5824, 6353, 7103, 6434, 7676, 7478, 7703, 8315, 7467, 7657, 7621, 9504, 8913, 9417, 10623, 10066, 10869, 425, 287, 1020, 9471, 5365, 8072, 10847, 11328]. Oceanic [5041, 1419]. Oceanographic [11016, 9468]. Oceanography [9069]. October [10521, 2015, 3212, 4408, 5684, 6881, 8093, 9219, 10366, 11484, 763]. Octopus [10490]. Odd [5981]. Oddities [3917, 4352, 4562, 4647, 5194, 5485, 5583, 5609, 5682, 5798, 5939, 5988, 6061, 6324, 6611, 6745, 6860, 6946, 7009, 7094, 7458, 7527, 7613, 7866, 8176, 8391, 8625, 8948, 9017, 9380, 9712, 9784, 10722, 10820, 10888, 11022, 11504, 11669, 128, 268, 445, 905, 1004, 6909]. Odorless [8194]. Odors [10216]. Off [7091, 6218, 9250]. Offer [8248]. Offered [5828]. Office [1434, 1563, 5996, 6468, 8772, 9167, 10542, 9606, 1868, 3634, 2615, 3908, 3743, 4243, 6568, 6651, 8353, 9413, 9481, 10828, 11186, 11255, 4077, 5139]. Officer [5280, 10860]. Officers [4163]. Offices [10928]. Official [1762, 4640, 6225, 8355, 8378, 9047, 10521, 10852, 11060, 11430, 11633]. Oft [7646]. Oft-Repeated [7646]. Ohio [11457, 2868, 4059, 7303, 8454, 6123, 5503]. Oil [1579, 3019, 3504, 3209, 3464, 2647, 4883, 4432, 4146, 4537, 3790, 4456, 5345, 5504, 5307, 5158, 6800, 6550, 6551, 8479, 7840, 8364, 8130, 9542, 10938, 539, 996, 5476, 5347, 1238, 4142, 10406, 4753, 5212, 6670, 8338, 8966]. Oil-Burning [3790, 4753]. Oil-Cloth [3464]. Oil-Rods [4146]. Oilcloths [10501]. Oiled [3316]. Oiler [6585, 9690]. Oiling [10136]. Oils [7637]. Okapi [3295, 9137]. Old [1629, 1395, 2942, 2617, 2724, 4800, 5433, 5502, 7875, 8518, 9144, 9976, 10413, 11218, 7480, 10950, 7284, 218, 11615, 2132, 7707, 2574, 3046, 3114, 4253, 4168, 7810, 10112, 11640]. Oldest [5356, 7308, 6422]. Oldsmobile [7796, 8797]. Oliver [6010]. Olives [5000]. Omnibus [1490, 7795]. Omnibuses [4917]. Omnigraph [2643]. One [5210, 5732, 6998, 7836, 8184, 7385, 8940, 10748, 10320, 10214, 10192, 768, 5214, 10811, 9567]. One-Caliber-Gun [10192]. One-Half [10748]. One-Hundred-and-Fifty-Ton [9567]. One-Hundred-Ton [8940]. One-Man [5210]. Onias [8825]. Onion [5812, 1231]. Onion-Seed [5812]. Ontario [7862]. Opacity [11888]. Open [8656, 11283, 1025, 802, 11783]. Open-Air [11283, 802]. Opened [10825]. Opener [3615]. Opening [1505, 2729, 2839, 3959, 5211, 5982, 6981, 7853, 8377, 7496, 7833, 8532, 9626, 9561, 9506, 10132, 10774, 10684, 11351, 1167, 9861]. Opera [5291, 6690, 11847]. Opera-Glass [5291]. Operated [5233, 5565, 6442, 6483, 6500, 7251, 8838, 10447, 9689, 1055, 8386, 6122, 1140]. Operating [2151, 9907, 35]. Operation [2050, 2438, 2754, 8184, 10276, 3403]. Operations [4349, 4125, 7385, 9702, 10923, 10284]. Operative [5479]. Ophthalmo [926]. Ophthalmo-Diaphanoscope [926]. Opinion [4310, 8734, 617]. Opinions [3610]. Opium [9706]. Opossum [3429]. Opportunity [79]. Opposition [10352, 9426, 589]. Oppressive [9856]. Optic [4532, 4960]. Optical [4141, 5980, 7913, 899, 54, 8919, 9772, 11418, 10202]. Oral [11138]. Orange [3043, 5360, 5529, 1165]. Orbits [2983]. Orchard [802]. Orchestra [3090]. Orchestra-Man [3090]. Ordinary [1337, 466]. Ordnance [3384, 3846, 1116, 715, 2689, 2326, 3327, 2704, 4634, 1115, 9013]. Ore [2327, 4025, 5255, 5929, 6079, 6313, 7154, 6436, 6277, 8286, 9645, 9655, 11869, 11465, 1072, 1319, 3723, 5163, 2555, 5364, 7040, 8124, 8278, 9052]. Ore-Carrying [5929, 6079, 8124]. Ore-Conveyer [4025]. Ore-Handling [5364]. Ore-Reducing [5255]. Ore-Sampling [6277]. Ore-Unloading [8278, 9052]. Oregon [2031, 2519, 8540, 2846]. Ores [1672, 3110, 4471, 6801, 1653, 9366]. Organ [7067]. Organic [8233]. Organisms [9246, 11861]. Organization [8677, 963, 4577]. Organs [6605, 10924, 6363]. Orient [10998]. Orientation [8961]. Origin [7887, 9164, 1117, 1223, 11043]. Original [9564, 11563, 790, 10277]. Originated [4342]. Orinoco [8281]. Orleans [3465]. Orling [4867]. Ormond [7834, 9053]. Ornamental [11408, 11716]. Ornamentation [336]. Ornithological [5983]. Orograph [1560]. Oroya [4579]. Orrery [10317]. Orson [9701]. Orthodiagraph [6605]. Orthopedic [10691]. Orthopter [10401]. Orvieto [1710]. Orville [1255, 1217, 823]. Osborne [10673]. Oscillating [4033]. Oscillations [9836, 8683]. Oscillograph [5363]. Osmon [7272]. Ostergren [5307]. Ostia [1753]. Ostrich [2573, 4847]. Ostriches [6298]. Other [2860, 8161, 11158, 985, 5238, 5630, 11715, 7908, 536, 4579, 8493, 1254, 706]. Otherwise [4995]. Ought [8857]. Our [1863, 2367, 1436, 1347, 1695, 2146, 1911, 2391, 2635, 3086, 2908, 3443, 3476, 3417, 3493, 3054, 4257, 4337, 4640, 3972, 4480, 4234, 4482, 4958, 5226, 6002, 5378, 5884, 5951, 6436, 6248, 6705, 7781, 7824, 8048, 8018, 8138, 9190, 9586, 9303, 8541, 10294, 10593, 10706, 10024, 10350, 10274, 10536, 10648, 10394, 11387, 10896, 11657, 11128, 11054, 11556, 11882, 11474, 11130, 11102, 658, 799, 800, 921, 503, 147, 304, 611, 881, 11707, 963, 7514, 7538, 7558, 481, 4435, 11167, 1778, 1974, 2324, 2392, 2840, 3561, 4021, 4731, 5432, 5948, 5561, 5825]. our [5204, 5870, 5946, 5978, 6068, 6397, 6332, 7159, 8314, 9224, 9501, 9874, 10422, 10751, 10191, 11166, 11231, 11261, 11295, 10915, 11056, 580, 933, 1077, 830, 859, 1146, 1299, 407, 8964, 10811, 1991, 703, 11284, 10453]. Ourdan [10859]. Ours [3193, 5238, 5213]. Outcome [11395]. Outcrop [10280]. Outer [9640]. Outfit [5809, 8807, 468, 663, 4278]. Outfits [7881]. Outgo [114]. Outing [4257, 7561]. Outlook [1788, 10707, 10382]. Outlooks [722]. Output [4500]. Oval [2007, 10653]. Oven [4391]. Overalls [5484]. Overcrowding [698]. Overflow [4586, 8707]. Overhead [2961, 6515, 10666, 9008]. Overman [4283]. Overshoes [6058]. Overworked [10109]. Ovipositor [11758]. Ovoids [1029]. Own [3863, 8375, 10434, 10421, 9650]. Ownership [753, 9724]. Ox [9403, 1129]. Ox-Shoeing [9403]. Oxalates [2047]. Oxford [3856, 3744, 4484]. Oxide [2303]. Oxy [9755]. Oxy-Acetylene [9755]. Oxygen [4416, 9503, 11038, 8477, 6341, 7655, 1059]. Oxygen-Inhaling [6341]. Oyster [5701, 8005, 2422]. Oysters [2175]. Ozone [6356, 1179, 9228, 5522].

P [2706, 218, 8141]. Pa [9819]. Paced [4882]. Pacific [1861, 3331, 3819, 8072, 8897, 10496, 5189, 7192, 9288, 7472, 2256, 3506, 2662, 2821, 4895, 4678, 5459, 5211, 8202, 10294, 10274, 10215, 10575, 11256, 614, 3843, 9095, 10710, 2475, 1634, 3140, 10577, 11328]. Pack [6070, 6740]. Package [11468]. Packard [4295]. Packing [2420, 5888, 7604, 9864, 9988, 10227, 4127]. Paddle [4195, 9115, 11606]. Paddle-Wheel [4195]. Paderborn [5522]. Padlock [948]. Paducah [6814]. Page [2597, 2732, 4510, 9169]. Page-Printing [2597, 2732, 4510, 9169]. Paige [2553]. Pail [2556, 7708]. Pain [8719]. Paint [2239, 4712].

Paint-Striper [2239]. Painter [11611]. Painting [3326, 6103, 5377, 7671, 10499, 7492]. Paintings [1251, 10202]. Pair [1608, 10068]. Pairs [8212]. Palace [2080, 3138, 315, 2010, 2473, 3817, 6941, 10474, 10626]. Palaces [1370]. Palæolithic [3458]. Palais [2521]. Palenke [4919]. Palermo [3684]. Palestine [8186, 790, 393]. Palettes [4712]. Palisade [1766]. Pallado [2047]. Pallado-Oxalates [2047]. Pampas [4063]. Pamphlets [5861]. Pan [9492, 10982, 1510, 2289, 1793, 3591, 3400, 3385, 3388, 2729, 2839, 2870, 11386, 2179, 2245, 2182, 3154]. Pan-American [1510, 2289, 1793, 3591, 3400, 3385, 3388, 2729, 2839, 2870, 11386, 2179, 2245, 2182, 3154]. Panama [4118, 3501, 4148, 4309, 3902, 4184, 4747, 4452, 4569, 6155, 5012, 5469, 6308, 7152, 7228, 7669, 7677, 7781, 8248, 8101, 8016, 7416, 8226, 8422, 7417, 8979, 9030, 8632, 8957, 8734, 9283, 8868, 9447, 8822, 8984, 9522, 9440, 8564, 8573, 9460, 8572, 9696, 9876, 10535, 10413, 10190, 10433, 10373, 10189, 11213, 11409, 11857, 10957, 11053, 11250, 11152, 11755, 11493, 11272, 452, 1186, 1011, 9676, 5885, 1148, 560]. Pango [1922]. Panhard [2962]. Panic [1594, 4230, 10987]. Panning [1176]. Panoramic [1540]. Pantheon [4558]. Pantry [2107]. Paper [3681, 3176, 2665, 6161, 5861, 6022, 5869, 5575, 5436, 5849, 6806, 6543, 7556, 10020, 11334, 11525, 11558, 11521, 1180, 685, 11111, 6873, 870, 3611]. Paper-Feeders [6161]. Paper-Making [5869]. Papers [5181]. Parade [6447, 1211]. Paradox [723, 971, 6361, 1028]. Parallax [7812]. Parallel [9379, 4662]. Parasite [2918, 6765]. Parasites [4549, 10315]. Parcel [11215]. Paris [509, 1999, 2159, 1383, 1607, 2515, 2085, 2098, 2234, 2148, 1657, 2131, 2153, 1485, 2080, 2128, 2279, 2329, 2054, 2011, 2033, 1505, 2010, 2052, 2102, 1955, 2521, 1988, 1984, 1461, 1508, 1531, 1554, 1649, 1698, 1936, 2194, 2372, 2490, 1962, 1656, 2442, 1334, 2495, 1434, 2350, 2381, 2057, 1751, 2309, 1702, 1797, 2125, 3529, 2931, 3530, 2819, 3378, 3575, 3399, 4144, 4750, 4436, 4576, 4808, 5411, 5366, 4921, 6639, 6652, 7251, 7565, 9209, 9350, 11219, 10381, 10946, 1096, 1344, 2913, 3523, 2424, 2075]. Parisian [8340]. Park [2408, 3129, 2845, 3791, 4185, 3815, 3737, 5354, 7581, 8766, 2762, 9355, 1321, 3735, 6966, 7263, 6223, 7089, 6578, 11940]. Parker [11273]. Parseval [9352]. Parsons [2709, 9048]. Parthenon [4663]. Partial [7579]. Parts [1773, 7836]. Pass [3682, 11153, 11313]. Passage [1548, 7826, 925, 8350, 835]. Passenger [3551, 3081, 3462, 4286, 4343, 5692, 5114, 6413, 6980, 7469, 7793, 8789, 9205, 10642, 10943, 11474, 145, 47, 4684, 1126, 7628]. Passengers [3699]. Passing [5461, 8540, 11659, 6579, 3230, 1373, 9704, 4871]. Past [5134, 8564, 8708, 10127, 11659]. Paste [10942]. Pasteur [6619]. Patch [564]. Patching [324]. Patent [2232, 1443, 1719, 1862, 2409, 1315, 1330, 1927, 1950, 1973, 1998, 2020, 2043, 2069, 2092, 2115, 2142, 2166, 2189, 2208, 2229, 2254, 2274, 2297, 2321, 2347, 2365, 2388, 2407, 2433, 2454, 2482, 2507, 1358, 1380, 1405, 1428, 1453, 1477, 1502, 1526, 1547, 1571, 1593, 1619, 1641, 1667, 1692, 1717, 1742, 1760, 1786, 1813, 1839, 1860, 1882, 1903, 2242, 3377, 2567, 2545, 3445, 3467, 3531, 2529, 2543, 3149, 3170, 3192, 3217, 3239, 3259, 3301, 3373, 3452, 3539, 3560, 3584, 3607, 3630, 3652, 3675, 2564]. Patent [2589, 2609, 2631, 2654, 2677, 2699, 2722, 2747, 2791, 2814, 2836, 2857, 2880, 2905, 2928, 2951, 2977, 3002, 3030, 3053, 3077, 3105, 3128, 2547, 2593, 3363, 4775, 3999, 3767, 3821, 3870, 3915, 3963, 4009, 4088, 4417, 4646, 4713, 4760, 4811, 4854, 3732, 3753, 4594, 4632, 4656, 4726, 4745, 4771, 4797, 4821, 4842, 4864, 4889, 3830, 3854, 3877, 3901, 3924, 3949, 3970, 4016, 4043, 4072, 4117, 4182, 4210, 4228, 4251, 4273, 4307, 4331, 4360, 4382, 4430, 4451, 4479, 4567, 4936, 4957, 5015, 5051, 5262, 5337]. Patent [5416, 5482, 5548, 5579, 5887, 5936, 5985, 6057, 6102, 4912, 4955, 5712, 5731, 5758, 5781, 5822, 5843, 5876, 5898, 5919, 5944, 5967, 5995, 6019, 6041, 6067, 6113, 6134, 6154, 4987, 5010, 5057, 5098, 5126, 5155, 5175, 5202, 5222, 5245, 5276, 5300, 5321, 5353, 5374, 5403, 5423, 5452, 5493, 5516, 5591, 5636, 5662, 5689, 5456, 5045, 7214, 7158, 6270, 6320, 6400, 6466, 6510, 6582, 6610, 6942, 7006, 7090, 7142, 6175, 6194, 6842, 6866, 6904, 6926, 6970, 7015, 7034, 7054, 7077, 7100, 7127, 7151, 7179, 7203]. Patent [7227, 7248, 7270, 7291, 7315, 7335, 6214, 6234, 6256, 6286, 6305, 6330, 6351, 6372, 6412, 6432, 6454, 6474, 6494, 6526, 6549, 6567, 6598, 6616, 6678, 6704, 6727, 6752, 6773, 6820, 7880, 7452, 7524, 7610, 7634, 7682, 7864, 7890, 8056, 8173, 8214, 8238, 8284, 8389, 8506, 7352, 7370, 8069, 8098, 8117, 8136, 8158, 8182, 8201, 8224, 8247, 8269, 8292, 8311, 8348, 8371, 8396, 8421, 8445, 8467, 8488, 8516, 7392, 7437, 7466, 7487, 7510, 7533, 7553, 7575, 7595, 7644, 7668, 7691, 7716, 7746, 7762, 7847, 7874]. Patent [7898, 7928, 7951, 7974, 7997, 8015, 8038, 9564, 8552, 8903, 8943, 9334, 9450, 9490, 9625, 8539, 8562, 9158, 9179, 9201, 9223, 9238, 9259, 9280, 9302, 9320, 9343, 9363, 9387, 9412, 9439, 9458, 9478, 9500, 9520, 9538, 9559, 9576, 9614, 9638, 9663, 8581, 8608, 8631, 8650, 8672, 8696, 8732, 8758, 8780, 8821, 8846, 8866, 8886, 8911, 8931, 8956, 8976, 9001, 9023, 9041, 9063, 9089, 9109, 9135, 9985, 10134, 10322, 10504, 10544, 10564, 10693, 9682, 9695, 10291, 10311, 10329, 10349, 10372, 10391, 10412, 10432, 10449]. Patent [10489, 10514, 10534, 10552, 10572, 10602, 10620, 10638, 10663, 10680, 10702, 10728, 10746, 10766, 10788, 9718, 9741, 9764, 9788, 9811, 9832, 9852, 9873, 9892, 9973, 9999, 10017, 10034, 10055, 10076, 10102, 10120, 10146, 10168, 10187, 10210, 10232, 10253, 10273, 11898, 11426, 10834, 10926, 10975, 11091, 11306, 11330, 11464, 11500, 11566, 11661, 11759, 11832, 11866, 11947, 10806, 10824, 11384, 11407, 11424, 11446, 11624, 11767, 11814, 11838, 11855, 11874, 11894, 11912, 11931, 10843, 10866, 10894, 10913, 10934, 10956, 10985, 11009, 11028, 11052, 11073, 11101, 11126]. Patent [11143, 11163, 11182, 11209, 11230, 11249, 11271, 11292, 11312, 11340, 11362, 1039, 118, 174, 264, 370, 437, 778, 901, 946, 1000, 1173, 1302, 1123, 19, 43, 668, 691, 712, 739, 765, 787, 818, 841, 878, 909, 930, 954, 977, 1009, 1037, 1064, 1083, 1110, 1184, 1208, 1240, 1267, 1287, 76, 102, 133, 153, 182, 209, 229, 248, 273, 301, 331, 354, 379, 406, 423, 450, 470, 500, 522, 547, 577, 603, 627, 648, 125, 743, 9606, 91, 1434, 2275, 3634, 2615, 4560, 5996, 6568]. Patent [8353, 9413, 9481, 10828, 11186, 11255, 4077]. Patented [1310, 1325, 1923, 1945, 1969, 1993, 2016, 2038, 2064, 2087, 2110, 2137, 2161, 2184, 2204, 2224, 2249, 2269, 2292, 2316, 2342, 2360, 2383, 2403, 2428, 2449, 2477, 2502, 1353, 1375, 1400, 1424, 1449, 1472, 1497, 1521, 1543, 1566, 1588, 1614, 1636, 1662, 1687, 1712, 1736, 1755, 1781, 1808, 1834, 1855, 1877, 1898, 2357, 3021, 2525, 2540, 3124, 3145, 3166, 3189, 3214, 3235, 3255, 3278, 3297, 3322, 3343, 3369, 3391, 3413, 3432, 3450, 3471, 3496, 3515, 3535, 3556, 3580, 3603, 3626, 3648, 3672]. Patented [2560, 2585, 2605, 2628, 2651, 2673, 2695, 2719, 2743, 2766, 2787, 2811, 2833, 2853, 2877, 2901, 2924, 2946, 2972, 2997, 3025, 3049, 3073, 3101, 2785, 2943, 3095, 4222, 3692, 3709, 4305, 4329, 4356, 4379, 4401, 4427, 4447, 4476, 4495, 4518, 4542, 4565, 4591, 4609, 4629, 4653, 4674, 4697, 4722, 4743, 4768, 4793, 4818, 4838, 4860, 4886, 3729, 3749, 3775, 3802, 3828, 3852, 3874, 3897, 3922, 3947, 3968, 3991, 4013, 4040, 4069, 4093, 4113, 4133, 4155, 4179, 4206, 4225, 4248, 4274, 5494, 4910, 4923]. Patented [5555, 5587, 5612, 5633, 5658, 5685, 5708, 5727, 5754, 5777, 5818, 5841, 5873, 5894, 5915, 5963, 5991, 6016, 6037, 6064, 6084, 6109, 6130, 6151, 4952, 4984, 5006, 5026, 5061, 5095, 5122, 5151, 5171, 5197, 5218, 5242, 5274, 5296, 5317, 5351, 5370, 5399, 5421, 5448, 5465, 5489, 5513, 5533, 6171, 6190, 6769, 6791, 6816, 6838, 6862, 6882, 6900, 6922, 6950, 6988, 7013, 7030, 7050, 7073, 7097, 7123, 7147, 7175, 7199, 7224, 7244, 7267, 7287, 7311, 7332, 6211, 6229, 6252, 6282, 6301, 6327, 6346]. Patented [6369, 6385, 6409, 6428, 6450, 6470, 6490, 6522, 6546, 6564, 6594, 6613, 6634, 6675, 6700, 6723, 6749, 7349, 7366, 7970, 7994, 8011, 8034, 8065, 8095, 8114, 8133, 8154, 8180, 8197, 8220, 8244, 8265, 8290, 8308, 8328, 8344, 8367, 8394, 8417, 8441, 8485, 8512, 7388, 7411, 7433, 7462, 7483, 7507, 7530, 7550, 7571, 7591, 7616, 7640, 7665, 7687, 7712, 7741, 7759, 7777, 7817, 7843, 7870, 7895, 7924, 7947, 8535, 8557, 9133, 9156, 9176, 9197, 9221, 9236, 9256, 9276, 9298, 9316, 9340, 9360, 9383]. Patented [9408, 9435, 9455, 9475, 9496, 9516, 9534, 9556, 9574, 9591, 9610, 9635, 9659, 8577, 8603, 8628, 8647, 8669, 8694, 8711, 8728, 8754, 8776, 8818, 8842, 8862, 8882, 8908, 8927, 8952, 8972, 8997, 9019, 9038, 9059, 9085, 9106, 9679, 9692, 10270, 10287, 10307, 10327, 10345, 10368, 10388, 10408, 10430, 10485, 10510, 10532, 10550, 10568, 10599, 10616, 10634, 10661, 10678, 10698, 10724, 10743, 10763, 10785, 9714, 9737, 9761, 9785, 9807, 9828, 9848, 9870, 9888, 9913, 9969, 9995, 10013, 10030, 10051, 10072, 10097, 10116, 10142]. Patented [10164, 10183, 10207, 10229, 10250, 10802, 10839, 11420, 11442, 11469, 11486, 11506, 11529, 11551, 11576, 11592, 11620, 11646, 11670, 11693, 11742, 11764, 11790, 11812, 11836, 11851, 11870, 11890, 11908, 11927, 10862, 10890, 10911, 10932, 10952, 11005, 11024, 11048, 11069, 11097, 11122, 11141, 11161, 11180, 11205, 11226, 11245, 11269, 11288, 11310, 11336, 11360, 11380, 11404, 2, 18, 666, 693, 711, 738, 764, 784, 816, 838, 876, 907, 928, 951, 975, 1006, 1034, 1062, 1082, 1107, 1131, 1155, 1183, 1205, 1239, 1264, 1285, 40]. Patented [73, 99, 131, 152, 180, 206, 245, 270, 298, 330, 353, 375, 402, 421, 447, 469, 499, 521, 543, 574, 601, 623, 645]. Patentees [5395]. Patents [1315, 1330, 1927, 1950, 1973, 1998, 2020, 2043, 2069, 2092, 2115, 2142, 2166, 2189, 2208, 2229, 2254, 2274, 2297, 2321, 2347, 2365, 2388, 2407, 2433, 2454, 2482, 2507, 1358, 1380, 1405, 1428, 1453, 1477, 1502, 1526, 1547, 1571, 1593, 1619, 1641, 1667, 1692, 1717, 1742, 1760, 1786, 1813, 1839, 1860, 1882, 1903, 2300, 2280, 3543, 2529, 2543, 3149, 3170, 3192, 3217, 3239, 3259, 3301, 3373, 3452, 3539, 3560, 3584, 3607, 3630, 3652, 3675, 2564, 2589, 2609, 2631, 2654, 2677, 2699, 2722, 2747]. Patents [2791, 2814, 2836, 2857, 2880, 2905, 2928, 2951, 2977, 3002, 3030, 3053, 3077, 3105, 3128, 4561, 4732, 4870, 3732, 3753, 4594, 4632, 4656, 4726, 4745, 4771, 4797, 4821, 4842, 4864, 4889, 3830, 3854, 3877, 3901, 3924, 3949, 3970, 4016, 4043, 4072, 4117, 4182, 4210, 4228, 4251, 4273, 4307, 4331, 4360, 4382, 4430, 4451, 4479, 4567, 4719, 4912, 4955, 5712, 5731, 5758, 5781, 5822, 5843, 5876, 5898, 5919, 5944, 5967, 5995, 6019, 6041, 6067, 6113, 6134, 6154, 4987, 5010, 5098, 5126, 5155, 5175, 5202]. Patents [5222, 5276, 5300, 5321, 5353, 5374, 5403, 5423, 5452, 5493, 5516, 5591, 5636, 5662, 5689, 6796, 6869, 6175, 6194, 6842, 6866, 6904, 6926, 6970, 7015, 7034, 7054, 7077, 7100, 7127, 7151, 7179, 7203, 7227, 7248, 7270, 7291, 7315, 7335, 6214, 6234, 6256, 6286, 6305, 6330, 6351, 6372, 6412, 6432, 6454, 6474, 6494, 6526, 6549, 6567, 6598, 6616, 6678, 6704, 6727, 6752, 6773, 6820, 8077, 7352, 7370, 8069, 8098, 8117, 8136, 8158, 8182, 8201, 8224, 8247, 8269, 8292, 8311, 8348, 8371, 8396, 8421, 8445]. Patents [8467, 8488, 8516, 7392, 7437, 7466, 7487, 7510, 7533, 7553, 7575, 7595, 7644, 7668, 7691, 7716, 7746, 7762, 7847, 7874, 7898, 7928, 7951, 7974, 7997, 8015, 8038, 8342, 8539, 8562, 9158, 9179, 9201, 9223, 9238, 9259, 9280, 9302, 9320, 9343, 9363, 9387, 9412, 9439, 9458, 9478, 9500, 9520, 9538, 9559, 9576, 9614, 9638, 9663, 8581, 8608, 8631, 8650, 8672, 8696, 8732, 8758, 8780, 8821, 8846, 8866, 8886, 8911, 8931, 8956, 8976, 9001, 9023, 9041, 9063, 9089, 9109, 9135, 8715, 9593, 9682, 9695, 10291]. Patents [10311, 10329, 10349, 10372, 10391, 10412, 10432, 10449, 10489, 10514, 10552, 10572, 10602, 10620, 10638, 10663, 10680, 10702, 10728, 10746, 10766, 10788, 9718, 9741, 9764, 9788, 9811, 9832, 9852, 9873, 9892, 9973, 9999, 10017, 10034, 10055, 10076, 10102, 10120, 10146, 10168, 10187, 10210, 10232, 10253, 10273, 10806, 10824, 11384, 11407, 11424, 11446, 11624, 11814, 11838, 11855, 11874, 11894, 11912, 11931, 10843, 10866, 10894, 10913, 10934, 10956, 10985, 11009, 11028, 11052, 11073, 11101, 11126, 11143, 11163, 11182, 11209, 11230, 11249, 11271, 11292, 11312, 11340]. Patents [11362, 11108, 19, 43, 668, 691, 712, 739, 765, 787, 818, 841, 878, 909, 930, 954, 977, 1009, 1037, 1064, 1083, 1110, 1184, 1208, 1240, 1267, 1287, 76, 102, 133, 153, 182, 209, 229, 248, 273, 301, 331, 354, 406, 423, 450, 470, 500, 522, 547, 577, 603, 627, 648, 1335, 1841, 1495, 3058, 2728, 2660, 2615, 8635, 9978, 11571]. Patrie [10605, 10124]. Patrol [1340, 7794]. Pattern [5484]. Patuxent [11921]. Paucity [11913]. Paul [1723, 3655, 7565, 614]. Pavement [5542, 7811]. Pavements [927]. Pavilion [7065]. Paving [4921, 10970]. Pay [5129, 753, 900]. Peace [3474, 8000, 7861, 10018, 1194, 356, 9853]. Peanut [1483]. Pearl [10851, 221, 8001, 8785]. Pearls [10258, 8001]. Peary [2281, 1905, 3218, 4407, 6917, 8385, 8373, 9392, 11818, 728, 1269, 7188]. Peas [4624]. Peasant [1816]. Peat [4465, 7272, 6823, 7624, 9444]. Peat-Fuel [4465]. Pecos [7472]. Peculiar [2599, 4759, 5342, 6342, 54, 9212, 289, 6965].

Peddled [4020]. Pedrail [5111]. Peel [3043]. Peerless [8799, 10482, 9966]. Pelé [5932]. Pelée [4575, 4075, 4046, 7579, 4853]. Pelicans [3848, 5391]. Pen [2684, 2692, 6857, 7454]. Pencil [6146, 325]. Pencils [10125]. Pendulum [3524, 6828, 11346, 11305, 4558]. Penetrating [7293]. Penetration [7038]. Penetrative [2035]. Peninsula [2059]. Pennsylvania [10528, 3462, 3526, 3500, 3786, 4197, 4658, 6915, 6606, 9282, 8875, 9418, 9185, 8660, 8856, 10302, 9918, 11276, 11342, 11738, 11251, 10845, 1300, 1092, 5258, 7810]. Pens [9686]. People [7517, 8636, 5885, 703]. Peoples [1117]. Pepper [1437]. Perdue [6939]. Perfect [6918]. Perfected [1438, 904]. Perfecting [4140]. Perforator [2419]. Performance [7597, 8676, 6552]. Performances [8664, 10124, 10238, 11685, 464]. Performed [11346, 4142]. Perfume [684]. Pergamon [8305]. Perhydrase [10123]. Peril [6137, 6306, 9982, 9893, 10257, 10002, 10914]. Perils [3927, 6799, 10704, 10705]. Periodic [5790]. Periods [748]. Perishable [11715]. Perkin [10689]. Permanence [4990]. Permanency [1888]. Permanent [2548, 10730, 11704, 11147]. Permitting [6869]. Perpetual [2730]. Perplexing [4584]. Perry [3492, 5794]. Persei [3459]. Perseus [2544, 3438]. Persia [1491, 4161]. Personal [1312, 1327, 1924, 1947, 1970, 1994, 2017, 2040, 2065, 2089, 2111, 2138, 2162, 2186, 2205, 2225, 2251, 2270, 2294, 2317, 2344, 2362, 2385, 2404, 2430, 2451, 2478, 2504, 1355, 1376, 1402, 1425, 1450, 1474, 1498, 1522, 1544, 1568, 1590, 1615, 1637, 1663, 1689, 1714, 1737, 1757, 1783, 1810, 1836, 1857, 1879, 1900, 2559, 2582, 3276, 3296, 3319, 3341, 3368, 3390, 3412, 3430, 3448, 3470, 3494, 3513, 3533, 3554, 3578, 3601, 3623, 3647, 3670, 3690, 2604, 2625, 2649, 2670, 2674, 2717, 2740, 2809, 2831, 2874]. Personal [2899, 2919, 2944, 3071, 3098, 3122, 3144, 3164, 3186, 3210, 3233, 3254, 2526, 2767, 2947, 2973, 2998, 3027, 3086, 3705, 3710, 4325, 4354, 4377, 4398, 4423, 4445, 4473, 4494, 4514, 4539, 4563, 4588, 4607, 4627, 4650, 4672, 4694, 4718, 4740, 4765, 4790, 4815, 4837, 4858, 4884, 4907, 3746, 3772, 3776, 3826, 3850, 3872, 3895, 3920, 3945, 3966, 3989, 4012, 4038, 4066, 4091, 4110, 4132, 4152, 4176, 4203, 4224, 4247, 4269, 4303, 4922, 4924, 5584, 5611, 5632, 5655, 5683, 5707, 5725, 5752, 5775, 5799]. Personal [5816, 5839, 5871, 5891, 5914, 5940, 5961, 5989, 6015, 6035, 6062, 6082, 6107, 6129, 6149, 6168, 4980, 5002, 5007, 5028, 5059, 5119, 5147, 5169, 5195, 5198, 5241, 5268, 5293, 5315, 5346, 5350, 5371, 5418, 5445, 5462, 5486, 5509, 5530, 5553, 6187, 6209, 6788, 6792, 6836, 6861, 6880, 6898, 6920, 6947, 6967, 6986, 7010, 7027, 7046, 7070, 7095, 7119, 7146, 7172, 7196, 7222, 7241, 7264, 7285, 7309, 7330, 6212, 7348, 6230, 6278, 6299, 6325, 6343, 6347, 6383, 6407, 6426, 6429, 6469, 6487, 6519]. Personal [6544, 6562, 6591, 6612, 6631, 6672, 6698, 6721, 6746, 6767, 8195, 7364, 7386, 7992, 8008, 8032, 8063, 8092, 8112, 8131, 8152, 8177, 8218, 8242, 8263, 8288, 8306, 8326, 8341, 8365, 8392, 8414, 8439, 8464, 8483, 8510, 8533, 7409, 7430, 7460, 7481, 7505, 7528, 7548, 7569, 7589, 7614, 7638, 7662, 7686, 7710, 7739, 7757, 7775, 7815, 7841, 7867, 7894, 7921, 7944, 7967, 8555, 8558, 9154, 9174, 9195, 9218, 9234, 9254, 9273, 9295, 9315, 9339, 9358, 9381, 9406, 9433, 9453, 9472, 9494, 9514, 9532].

Personal [9553, 9572, 9590, 9608, 9633, 9656, 9677, 8601, 8604, 8645, 8667, 8692, 8709, 8726, 8753, 8775, 8817, 8840, 8859, 8880, 8906, 8926, 8949, 8970, 8994, 9018, 9036, 9057, 9083, 9103, 9131, 10181, 2764, 2851]. Personalities [6996]. Perspectartigraph [6292]. Perspective [1463]. Perspectives [1751]. Peru [6608, 4579, 4785]. Peruvian [4969]. Perverted [11795]. Pests [11349, 10315, 10779]. Petard [10434]. Peterborough [9149]. Peters [10561]. Petit [2521]. Petrifaction [3293]. Petrified [7588, 2445]. Petroleum [2437, 3089, 2962, 956, 5720, 6505]. Petropavlovsk [1601]. Pewter [6014]. Phaetons [6663]. Pharmaceutical [9464]. Phase [4119, 4691, 6394, 8137, 9345, 8542, 10450, 252, 2075, 3766]. Phases [10672, 11001, 1125]. Phenomena [4141, 6437, 8227, 2213, 7112, 5790]. Phenomenal [1695, 11032]. Phenomenon [5326, 6315, 899, 54]. Philadelphia [5836, 789, 4961, 8357, 10558]. Philæ [4974]. Philippine [7238, 11222]. Philippines [1866, 2711, 2598, 5231, 11045, 4079, 3424, 4372, 10222]. Philological [4778]. Philosopher [8657]. Phonocard [8044]. Phonograph [8274, 2334, 2680, 2548, 7305, 6503, 9796, 6628]. Phosphate [2030]. Phosphorescence [11039]. Phosphorescent [9055, 7276]. Phosphorus [3007, 2596, 5764]. Photo [1673, 1397]. Photo-Micrography [1397]. Photobiography [8298]. Photochemical [845]. Photoelectric [5700]. Photograph [1704, 3863, 6203, 8989, 9185, 828, 4898, 6875]. Photographed [5767, 8461, 7909, 10518]. Photographers [7]. Photographic [2130, 2638, 4022, 4601, 5902, 6316, 7356, 9881, 9775, 751, 395, 10381, 7733, 6577, 11260, 10219, 1045, 3739, 7300, 10277, 8524, 5048]. Photographing [1367, 6556, 8473, 4260, 5388, 7547, 9290, 1124, 7962, 597, 3294]. Photographophone [3547]. Photographs [7344, 8876, 9957, 6999, 4602, 10882, 11619, 11608]. Photography [1645, 1909, 1536, 1435, 3849, 4848, 4749, 5930, 6806, 6508, 10667, 11320, 311, 314, 5626, 7213, 11498, 8388, 6735, 4737, 2035, 7981, 7165, 943, 1609, 1576]. Photometer [11827, 598]. Photomicrographs [10690]. Photomicrography [2829]. Photorama [3979]. Physical [2027, 559, 2397]. Physician [10944]. Physico [3588]. Physico-Technical [3588]. Physics [1694, 852]. Physiognomies [5440]. Physiology [4616]. Piano [4470, 4390, 8834, 8853, 9446, 9217, 10887, 10979, 120]. Piano-Player [4390]. Pianos [9294, 9845]. Picker [1633, 10365]. Pickers [11308]. Picking [7253, 372, 8748]. Picks [11414]. Pictet [1421]. Picture [2775, 3022, 4874, 6104, 457]. Pictures [2265, 4261, 6897, 11154, 8, 283, 796, 413, 587]. Picturesque [5933]. Pie [8080]. Pie-Making [8080]. Piece [3309, 2985, 7069]. Piedras [5676]. Pielock [6509]. Pier [11475]. Pierce [1774]. Piercing [8642, 9231, 11482]. Pierhead [11293]. Pierpont [6204, 8568]. Pierre [4105, 4299, 9769]. Piers [425]. Pig [1893]. Pigeon [2060, 4246, 11136]. Pigeons [5252, 6196, 8961, 9979, 7]. Pilatus [7302]. Pile [2667, 6358, 10282, 4438]. Piling [6263, 11883]. Pilot [1985]. Pin [8947]. Pine [1046, 2846, 7133].

Pinhole [11014]. Pinion [343]. Pins [6125]. Pint [4956, 8633]. Pioneers [9211]. Pipe [1887, 1775, 3062, 3275, 2646, 2533, 2621, 2602, 2815, 3120, 3782, 4613, 4764, 5953, 5691, 5836, 5889, 7218, 6184, 6541, 8217, 8438, 7964, 9630, 8977, 9989, 9847, 10356, 11466, 11199, 1026, 11614, 807, 11687, 6363, 1103, 733]. Pipe-Bending [10356]. Pipe-Wrench [1775, 2646, 3120]. Pipes [1907, 2756, 3332, 5681, 6246, 7881, 8723, 10545, 9905, 460, 806, 808]. Piston [5086, 5645, 5497, 8944, 9988, 10227, 485]. Pistons [9864]. Pitch [2576]. Pittsburgh [5737, 5791, 9819, 8203]. Pivoted [4714]. Pizarro [4785]. Place [6226]. Placed [7937]. Placer [11241, 5886]. Plague [1364, 11705]. Plagues [4871]. Plan [7491, 11152, 9030]. Plane [4600, 10836]. Planet [10854, 9210, 9507, 6077, 5810]. Planetarium [945]. Planetary [6659]. Planets [11001, 1281, 10518]. Planing [7256]. Plank [3613]. Plankton [1278]. Planning [8016]. Plans [3419, 7210, 7669, 9501, 4781]. Plant [61, 1392, 2128, 2217, 3240, 4341, 3715, 3832, 5616, 5594, 4989, 7277, 7693, 8520, 7833, 7862, 9216, 10440, 10022, 9729, 10461, 285, 9168, 4807, 8027, 431, 5693, 4667, 9248, 8433, 5118, 5627, 2424, 5251, 11304, 1140, 10427, 2397, 2665, 3562, 4901, 5499, 4150, 2664]. Plantations [2461]. Planter [4089]. Planters [1462]. Plants [3351, 5766, 5176, 5834, 6972, 6730, 8234, 8588, 10491, 11539, 11946, 358, 1097, 4103, 2328, 8545, 9733, 2552, 11439]. Plaster [2819, 2690]. Plastic [2987, 8287]. Plate [1693, 2116, 2200, 2485, 3486, 2804, 3784, 3823, 4489, 4440, 4530, 4844, 5950, 8705, 9365, 10667, 10815, 9356]. Plates [2268, 4022, 9881, 8831]. Platform [2238, 7577, 7583, 11611]. Platforms [10331]. Plating [11630, 11715, 9589, 1023]. Platinum [4805, 792]. Platt [11940]. Play [3060, 10917]. Player [4390, 9446, 11137, 10659]. Players [10979]. Playing [832, 889]. Plea [7320]. Pleasure [6637]. Plecher [2332]. Pledge [3737]. Plow [1677, 2712, 3769, 10547, 11238, 11668, 11663, 10155, 463, 5648]. Plowing [9193]. Plows [2516]. Plug [4619, 4145, 5144, 5090, 8174, 10695, 10506, 9847, 11094, 570, 688]. Plumb [3725]. Plumbing [6891]. Pluvy [9465]. Pneumatic [1644, 1919, 2332, 1830, 2871, 4705, 4834, 5809, 6007, 6977, 6943, 7185, 7990, 8684, 10093, 10475, 9954, 11934, 10979, 127, 3720, 4315, 4998]. Poas [8257]. Pocket [5344, 10442, 9467]. Pocket-Knife [9467]. Pockets [9244]. Pod [3115]. Poetry [10857]. Point [1705, 3056, 3243, 4336, 3879, 4973, 7489, 4335, 10860]. Pointers [5395]. Points [7637, 9463]. Poison [3109, 10203]. Poisoning [4364, 6787, 6707]. Poisonous [8472, 9598, 2624, 350]. Polar [6268, 6295, 9148, 10086, 4235, 6529, 1278]. Pole [5090, 8252, 10323, 8373, 11818, 728, 709, 696, 1269, 831, 10425]. Pole-Climbing [10323]. Poles [2799, 9666, 11021, 397]. Police [7794, 11437]. Policy [1290]. Polished [733]. Polishing [11543]. Pollak [2267, 1559]. Pollen [9452, 1123]. Pollen-Collecting [9452]. Pollen-Gathering [1123]. Pollok [2549]. Pollution [11856]. Polonium [2446]. Polycera [4030].

Polyglot [2513]. Polyphase [8746]. Polytechnic [8004]. Pompeii [4029]. Pompeiian [5957, 7545]. Poncho [11662]. Ponderous [11474]. Pontoon [5837]. Pools [438]. Poor [8372]. Pope [6642, 9937, 798]. Popoff [2823]. Popp [4622]. Popular [5269, 887]. Population [7101, 7440]. Porcelain [5631]. Port [1753, 3423, 1975, 6307, 6476, 7078, 6362, 7258, 8126, 7822, 7405, 7371, 7609]. Portable [1651, 1828, 2178, 2306, 2850, 4241, 4410, 4531, 5363, 5417, 5809, 6518, 6943, 7963, 8361, 9216, 9172, 10471, 9909, 9907, 11265, 1307, 1174, 3525, 10139, 8815, 722, 6341]. Portage [7810]. Portal [1300]. Portentous [10312]. Porter [2622]. Portland [8890]. Porto [3680, 7274]. Portraits [6577]. Portraiture [3466]. Ports [2025, 5640]. Position [7294, 11820, 11085, 6198, 6757]. Possessions [5978, 1991]. Possibilities [3564, 3951, 4230, 6823, 8774, 11522, 167, 27, 11367]. Possible [1796]. Post [15, 5139, 9484]. Postage [11170, 10981, 369, 9484]. Postal [1430, 2060, 6641, 9856, 11107, 11106, 11511, 11215, 10739, 1623]. Posts [3389, 9710]. Pot [8549, 8633, 1229, 11615]. Potato [1913, 6944, 9183]. Potemkin [4350]. Potential [9010]. Potentiality [605]. Potomac [1606]. Pots [9758, 11716]. Potter [3114]. Pottery [2180]. Poulsen [2000, 2008, 9488]. Poultry [4812, 5312, 372]. Poultry-Feeding [4812]. Poultry-Picking [372]. Pound [3032]. Pound-Per-Horsepower [3032]. Pouring [5331]. Powder [2509, 3082, 10121, 10749, 612]. Powdered [4757]. Powderless [9289]. Powders [2725]. Power [5014, 8521, 61, 1799, 1585, 2331, 2217, 1886, 2911, 2860, 2881, 3328, 4341, 3764, 4636, 3715, 4347, 3959, 4881, 4456, 4529, 5616, 5594, 4989, 5324, 5417, 6972, 6800, 6237, 7103, 7293, 6828, 6244, 6886, 8146, 8468, 8481, 8251, 8081, 7544, 8321, 7975, 7660, 7833, 7862, 7732, 8231, 8588, 8675, 8721, 8616, 8685, 9530, 9119, 9486, 8799, 8660, 8902, 9123, 8543, 10440, 9925, 10422, 10491, 9945, 9926, 10240, 9949, 10237, 10461, 11179, 10918, 11626, 11210, 11388, 11159, 991, 23, 3788, 4901, 7183, 5499, 9248, 10421, 1630, 9034, 5083, 11325, 1105]. Power [5627, 2424, 5251, 7878, 11549, 2075, 3983, 4150, 11304, 4467, 7162, 2664, 6786, 10808, 10692, 5235, 7493, 4454, 10211, 11130]. Power-Boat [8081]. Power-Driven [1105]. Power-Hammer [2911]. Power-Means [7183]. Powered [7164, 8800]. Powerful [2470, 2575, 3088, 3474, 4411, 3909, 5114, 5162, 5133, 5806, 6761, 7608, 8031, 7805, 10197, 10108, 11752, 11450, 312, 822, 678, 1578, 7428]. Powers [2004, 3493, 6683, 8209, 10590, 1066]. Practicable [11425]. Practical [4216, 1892, 4098, 5395, 6715, 8077, 8654, 10712, 10067, 11037, 17, 590, 2076, 940, 10999, 4278]. Practice [2417, 1346, 3174, 3363, 3792, 4160, 5879, 10575, 9790, 11774, 10811]. Precaution [11596]. Precautions [6707, 11317]. Precedent [8470]. Precious [1847, 9563, 9801]. Predecessor [1393]. Predicting [3913]. Predicts [10131]. Prehistoric [5784, 7772, 9033, 9774, 10950, 4443, 9396, 5676]. Preliminary [1642, 4079, 10420]. Premier [7479]. Preparation [3007, 2969, 7021, 6375, 7671]. Preparations [6984]. Preparatory [5956]. Preparing [3020, 7301]. Prepayment [5338]. Present [2167, 2457, 1382, 5134, 6116, 7491, 7677, 8119, 7789, 8932, 9521, 8868, 9447, 8708, 10127, 10516, 11152, 11386, 11820, 1011, 10710, 560, 8259, 2686, 6023]. Presented [5181]. Presents [11714]. Preservation [1730, 1766, 2415, 7850, 8273, 9581, 8869, 9723, 10989, 7492]. Preservatives [5165, 7578]. Preserved [7851]. Preserver [9102, 172]. Preserves [10640]. Preserving [11847]. President [5207, 1642, 3171, 6126, 9522, 9440, 9670, 10621, 9878, 11102, 11769]. Press [3020, 5858, 5855, 5856, 6403, 6912, 7726, 8838, 11833, 4469]. Pressed [2369, 2915]. Pressed-Steel [2915]. Presses [2419, 5857]. Pressing [4480]. Pressure [1337, 2987, 2557, 5836, 7374, 7876, 8934, 8543, 9755, 10404, 10800, 10897, 11636, 267, 582, 913, 8499, 10114, 6245, 11655, 6237]. Pressures [9065]. Preussen [4526]. Prevail [7649]. Prevent [1343, 6787, 10077, 10396, 11746, 814, 11711]. Prevented [6414, 8451]. Preventer [892]. Preventing [4171, 5549, 5704, 6686, 8091, 7209]. Prevention [2808, 2863, 6927, 8253, 8987, 8760, 11055, 10949, 6620]. Preventive [5760]. Prey [6148, 166]. Price [11497, 284]. Priced [6128, 5706]. Priestley [6482]. Primary [6715, 8110]. Prime [3036]. Primitive [4788, 3552, 8429]. Prince [8294, 123]. Princess [1388]. Principal [10590, 1066]. Principia [8374]. Principle [9594, 7025]. Principles [6317, 6559, 482, 919]. Print [5931, 327, 6027, 6139]. Printed [7651, 11147]. Printer [5859]. Printing [1868, 2419, 2597, 2732, 4648, 4510, 6161, 5869, 5846, 5808, 6688, 9169, 11643, 10928, 9356, 11833, 6006, 4083]. Printing-Telegraph [6006]. Prints [1386, 5902, 10277, 1030]. Prinzessin [3042, 3009]. Prior [756, 758]. Prism [6080]. Private [3282, 3307, 4992, 11365, 397]. Prix [10495, 9566]. Prize [2267, 2984, 5621, 5828, 6651, 8472, 9046, 10752, 11724, 10941, 11635, 11277, 11300, 427, 599, 2549, 3408, 11253]. Prize-Winning [6651, 11300]. Prizefighter [9151]. Prizes [1457, 3401, 3083, 4776, 4190, 7856, 9443, 10217, 2074, 6005]. Probable [11724]. Problem [1814, 1884, 1441, 1541, 3542, 3396, 4170, 4570, 4700, 4455, 6156, 5284, 7057, 6601, 8185, 8518, 9182, 8565, 9242, 10036, 10492, 9974, 11449, 11342, 168, 343, 4940, 7402, 1282, 9528]. Problems [4747, 7152, 2493]. Process [9253, 1672, 2680, 2819, 4346, 5673, 5563, 5331, 5694, 5864, 6939, 6688, 6356, 6974, 7624, 7903, 8145, 8918, 9445, 9138, 8851, 9243, 9581, 10667, 11920, 11154, 10901, 10993, 11860, 11463, 311, 1197, 27, 732, 10627, 6577, 8105, 7421, 7730, 1824, 9684, 11589]. Processes [8356]. Procession [8523]. Produced [5847, 2213, 8919, 11014]. Producer [7000, 7936, 8425, 8050, 8588, 8544, 9720, 10939, 11859, 358, 8356]. Producer-Gas [7936, 8050, 358]. Producer-Gas-Driven [10939]. Produces [10472]. Producing [2517, 4416, 8046, 10094, 6577, 11260, 10279]. Production [9253, 1629, 1333, 1885, 1847, 2596, 6091, 5572, 5543, 5951, 7344, 6845, 6801, 7101, 7758, 9445, 9563, 11296, 8007, 8125, 2180, 8477]. Products [2681, 5869, 8826, 533]. Prof [1169]. Prof. [1822, 2514, 1437, 1421, 3014, 2705, 3380, 4366, 5867, 5805, 5305, 6098, 7153, 8737, 9617, 882, 6835, 5207, 10881, 6118]. Profession [10657, 5652]. Professor [1795, 6996]. Profile [8688, 1560]. Profit [4222]. Profitable [1643, 11932]. Program [3738]. Programme [6732, 381, 1409, 1385, 5033, 6238, 7419, 8916, 10990, 502, 6417]. Progress [2121, 1620, 2483, 1892, 1504, 2241, 2072, 2590, 3335, 3033, 4404, 6055, 5826, 7018, 6710, 7445, 8320, 8019, 9159, 9169, 9398, 8784, 9473, 9272, 9418, 10293, 9721, 10373, 10831, 11887, 11449, 11314, 11294, 11755, 11493, 11105, 11272, 93, 2245, 2348, 5699, 10129]. Progressive [4025]. Progressive-Motion [4025]. Progresso [5209]. Project [3263, 9288, 8025]. Projectile [11419]. Projectiles [1594, 6288, 9117]. Projecting [1773, 4874, 587]. Projection [8089]. Projections [6077]. Projector [7063, 52]. Projects [8138]. Prominent [5072]. Promising [3150]. Promoter [3883]. Promotion [1359]. Proof [2190, 4619, 4472, 7685, 8716, 1161, 751, 1100, 6810]. Propagating [11946]. Propagation [9140, 6895]. Propelled [1410, 4683, 4636, 5695, 7084, 6847, 7469, 9249, 9402, 8685, 9511, 9673, 9141, 11944, 280]. Propeller [1825, 2218, 3288, 3524, 4665, 6706, 8321, 11522, 1121]. Propellers [3152, 11805, 677]. Propelling [3202, 9116, 9292, 11761, 3645]. Properties [5834, 6782, 436, 5539]. Property [2982]. Prophecy [3849, 5433]. Proportions [9239]. Proposals [8583, 11427]. Proposed [9045, 1429, 1975, 1455, 1606, 1830, 2862, 2793, 3634, 3394, 2774, 2565, 3302, 3194, 3044, 4701, 4170, 5224, 6094, 5204, 5996, 5519, 5540, 6869, 6537, 6754, 6599, 7372, 7877, 8738, 8635, 8984, 8679, 9571, 10640, 10375, 10624, 11933, 1073, 1145, 382, 7598]. Propulsion [1441, 3033, 7079, 10846, 9837, 4684]. Pros [11840]. Prospecting [7938]. Prospective [4309, 10710]. Prospects [1905, 1504, 2632, 4076, 10415]. Prosperity [9262, 1065]. Protect [5099]. Protected [1347, 2338, 3489, 5900, 9936, 10585]. Protecting [11462, 4435, 4644]. Protection [2434, 1796, 1334, 2659, 2982, 2907, 6753, 6257, 7081, 8450, 9388, 9066, 8697, 10897, 11108, 7192, 11713]. Protective [6574, 10062]. Protector [6663, 6743, 8574, 6011, 6225, 4625]. Protest [1223]. Proved [9599]. Provided [6760, 10836]. Providence [2236]. Proving [2352]. Provisioning [2940, 4680]. Prune [7361]. Pruner [6765]. Pruning [6242]. Prussian [4310, 3882]. Pseudo [606]. Pseudo-Volcanic [606]. Psychology [10300]. Pt [7385]. Ptomaine [6787]. Public [4561, 5432, 6042, 5537, 5156, 6821, 7180, 7880, 10206, 10904, 11459, 436, 1123, 11483, 11819]. Puck [10974]. Puddle [5357]. Pueblos [3914]. Puerto [1788]. Puff [9675]. Puget [3426]. Pull [2580]. Puller [4764, 3698, 9987]. Pulley [1703, 3334, 3311, 4036, 6585]. Pulleys [9654]. Pulling [6977, 6314, 8364, 11264, 1234]. Pullman [5928, 5578, 6907, 6689]. Pulp [5849]. Pulque [5649]. Pulse [8720]. Pump [1626, 2497, 3180, 3387, 2898, 4104, 6105, 6888, 6859, 7185, 8261, 8944, 8600, 11573, 804, 1578]. Pumping [1395, 5608, 7325]. Pumps [2557, 4056, 4259, 8497, 8129]. Punch [4599]. Punching [5344]. Puncture [4472, 7685]. Puncture-Proof [4472, 7685].

Pupils [11391]. Pupin [2980, 4732, 5383, 5041, 1658]. Puppet [2183]. Purchase [6055, 6809, 6778, 6398, 6389]. Pure [6375, 11254, 4995, 8915, 9286]. Pure-Food [8915]. Purification [2173, 7397, 8317]. Purifying [1179]. Purity [980]. Purpose [9854, 6398]. Purposes [1866, 3550, 4547, 6693, 7035, 6936, 6570, 7587, 8701, 9216, 10122, 3359, 8477, 8129]. Purse [8752]. Pushing [9594]. Put [4956, 3383]. Putting [328, 634, 7651]. Putting-On [634]. Puzzle [4584, 5920, 10695, 10506, 10650]. Puzzles [706]. Puzzling [8643]. Pygmies [8462]. Pyramids [8303, 3140]. Pyrites [7206]. Pyrometer [5980, 8833]. Pyrometric [6960]. Pyrotechnic [2532]. Python [4169, 6425]. Pythons [166].

Quadricycle [1873]. Quadruple [10543]. Quaint [5574]. Qualified [9743]. Qualities [8761]. Quality [7275, 2035]. Quantities [4879]. Quantity [7275]. Quarantine [8120]. Quarries [4349, 5503, 10527]. Quarry [3136]. Quarrying [9880, 11589]. Quart [4956, 8633]. Quartz [1933, 3998, 8286, 5017]. Quebec [8168, 9227, 10553, 10292, 10318, 10400, 10844, 10988, 11447]. Queen [4284, 6707, 1623, 2416, 3268, 5738]. Queensboro [453, 171]. Queensborough [11733]. Queer [7239, 5444, 4007, 6226, 4372]. Queerest [7363, 7916]. Queries [1313, 1328, 1925, 1948, 1971, 1995, 2018, 2041, 2066, 2090, 2112, 2139, 2163, 2187, 2206, 2226, 2252, 2271, 2295, 2318, 2345, 2363, 2386, 2405, 2431, 2452, 2479, 2505, 1356, 1377, 1403, 1426, 1451, 1475, 1499, 1523, 1545, 1569, 1591, 1616, 1638, 1664, 1690, 1715, 1738, 1758, 1784, 1811, 1837, 1858, 1880, 1901, 2527, 2541, 3126, 3147, 3168, 3190, 3215, 3237, 3257, 3280, 3299, 3323, 3345, 3371, 3393, 3416, 3434, 3454, 3473, 3499, 3519, 3537, 3558, 3582, 3605, 3627, 3650, 3673, 2562, 2587, 2606, 2629, 2652]. Queries [2675, 2697, 2720, 2744, 2768, 2789, 2812, 2834, 2854, 2878, 2903, 2926, 2948, 2974, 2999, 3028, 3051, 3075, 3103, 3694, 3712, 4308, 4333, 4358, 4381, 4403, 4431, 4450, 4478, 4496, 4519, 4543, 4568, 4592, 4610, 4630, 4654, 4673, 4698, 4723, 4746, 4772, 4796, 4822, 4841, 4863, 4890, 3733, 3752, 3779, 3804, 3831, 3855, 3878, 3900, 3925, 3950, 3971, 3994, 4017, 4044, 4073, 4096, 4116, 4135, 4157, 4183, 4209, 4229, 4252, 4272, 4925, 4953, 5659, 5686, 5709, 5728, 5755, 5778, 5819, 5842, 5895, 5916, 5941, 5964, 5992]. Queries [6017, 6038, 6085, 6110, 6131, 6152, 4985, 5008, 5058, 5096, 5123, 5152, 5172, 5199, 5219, 5243, 5297, 5318, 5372, 5400, 5425, 5451, 5467, 5490, 5514, 5534, 5556, 5589, 5614, 5634, 6172, 6191, 6793, 6817, 6839, 6863, 6901, 6923, 6951, 6989, 7031, 7051, 7074, 7098, 7124, 7148, 7176, 7200, 7225, 7245, 7268, 7288, 7312, 7333, 6231, 6253, 6283, 6302, 6328, 6348, 6370, 6386, 6410, 6430, 6451, 6471, 6491, 6523, 6547, 6565, 6595, 6614, 6635, 6676, 6701, 6724, 6750, 6770, 7350, 7367, 7971, 7995, 8012, 8035, 8066, 8096]. Queries [8115, 8134, 8155, 8198, 8221, 8245, 8266, 8309, 8329, 8345, 8368, 8418, 8442, 8486, 8513, 7389, 7412, 7434, 7463, 7484, 7508, 7531, 7551, 7572, 7592, 7617, 7641, 7666, 7688, 7713, 7742, 7760, 7778, 7818, 7844, 7871, 7896, 7925, 7948, 8536, 8559, 9257, 9277, 9299, 9317, 9341, 9361, 9384, 9409, 9436, 9456, 9476, 9497, 9517, 9535, 9557, 9611, 9636, 9660, 8578, 8605, 8648, 8670, 8712, 8729, 8755, 8777, 8819, 8843, 8863, 8883, 8909, 8928, 8953, 8973, 8998, 9020, 9039, 9060, 9087, 9107, 9177, 9198, 9871, 9680, 9693]. Queries [10271, 10288, 10308, 10346, 10369, 10389, 10409, 10486, 10511, 10569, 10617, 10635, 10699, 10725, 9715, 9738, 9762, 9786, 9808, 9829, 9849, 9889, 9914, 9970, 9996, 10014, 10031, 10052, 10073, 10098, 10117, 10143, 10165, 10184, 10208, 10230, 10251, 11246, 11487, 10803, 10840, 11671, 11694, 11743, 11765, 11791, 11852, 11871, 11891, 11909, 11928, 10863, 10891, 10953, 11006, 11025, 11049, 11070, 11098, 11123, 11206, 11227, 11289, 11337, 11381, 11405, 11421, 11443, 11470, 11507, 11530, 11552, 11593, 11621, 11647, 3, 41, 1132, 1206, 74, 100, 132, 181, 207, 246, 271]. Queries [299, 376, 403, 422, 448, 544, 575, 624, 646, 694, 785, 839, 952, 976, 1007, 1035, 1108, 11882]. Quest [5763, 8350, 11818]. Question [1883, 4546, 4892, 7931, 7646, 10767, 11629, 1888]. Quick [2777, 9906, 10170, 1153]. Quick-Acting [9906]. Quick-Firing [2777]. Quick-Steaming [10170]. Quito [8547]. Quoin [6811].

R [5805]. R. [9252]. R.I. [1346]. Race [1507, 2394, 4882, 5285, 6781, 6571, 6717, 6580, 7834, 8313, 8463, 7658, 7676, 7703, 7630, 7467, 8170, 8383, 7657, 7854, 9308, 9504, 9230, 9264, 9622, 9270, 10771, 10281, 10066, 10083, 10042, 11906, 11193, 11013, 2062, 1466, 2931, 3220, 3378, 3575, 4032, 4576, 4808, 5576, 6053, 5366, 6581, 7023, 7605, 7679, 7478, 8108, 8139, 9287, 10239, 10495, 11603, 11724, 11580, 957]. Race-Meet [2394]. Racehorse [10756]. Racer [4444, 6662, 8896, 10196]. Races [5030, 6100, 5601, 5597, 6382, 8081, 7808, 9053, 8913, 10084, 10687, 10441, 10236, 10797, 7086, 7116, 2063, 3283, 3184, 4362, 7863]. Racial [7139]. Racing [2668, 2620, 3005, 4087, 6056, 6162, 5414, 6954, 6637, 7658, 8231, 9540, 11179, 11438, 11723, 11491, 1268]. Racing-Yacht [2620]. Rack [2158, 6516, 8086, 7917, 11158, 343]. Racket [6539]. Radiation [4801, 5305, 5326, 6502, 11520]. Radiations [2446]. Radiators [1748, 9963, 7883]. Radicels [7038]. Radio [1345, 7317, 7134, 7161, 10609, 6438]. Radio-Active [1345, 7134]. Radio-Activity [7317, 7161, 10609, 6438]. Radioactivity [6710, 11453]. Radiography [2191, 10851]. Radium [2446, 2046, 5042, 5920, 5131, 6392, 6218, 7273, 6572, 6710, 7161, 8375, 8083, 7513, 8021, 9246, 9092, 8586, 9136, 8076, 5539]. Radium-Its [5539]. Rafts [2523, 10175]. Rag [11617]. Rail [1584, 1348, 1557, 2054, 1538, 2093, 1564, 2410, 1528, 1650, 2313, 3528, 3031, 2577, 4489, 5900, 5679, 6007, 5099, 5266, 5343, 6406, 6743, 7406, 7455, 9391, 9640, 8674, 9153, 9081, 10732, 9719, 9982, 10666, 10492, 9974, 11342, 11410, 263, 949, 782, 10668, 4671, 9389]. Rail-Bearing [4489]. Rail-Joint [7455]. Rail-Joints [2410]. Railbreaker [2132]. Railjoint [4264]. Railroad [1668, 2071, 1696, 2492, 1516, 3609, 3526, 2599, 3500, 3130, 3641, 4501, 3926, 4729, 4197, 4658, 4483, 4086, 4687, 3863, 3780, 5128, 5769, 5361, 5825, 5732, 5578, 5133, 5203, 5404, 5378, 6480, 6475, 7055, 7231, 6915, 6689, 7080, 8047, 8271, 8400, 8314, 8273, 8018, 7877, 8119, 8041, 7977, 8493, 7620, 7900, 9560, 8699, 8875, 9344, 9096, 9312, 9303, 8856, 9584, 9374, 9064, 9834, 10077, 10714, 10148, 10000, 10064, 9976, 10684, 10394, 9903, 10450, 9730, 11763, 11563, 11738, 11875, 1078, 1018, 177, 11414, 5443, 11064, 4902, 9804, 11065, 4644]. Railroad [4060, 10283, 2495, 3682, 3462, 4436, 4130, 3819, 4787, 5826, 7885, 9397, 8742, 8822, 8660, 8902, 10302, 10105, 10496, 10528, 9918, 11314, 11276, 11342, 1092, 7192, 8256, 1420, 7810, 4579, 9623]. Railroading [3328, 5430, 4265]. Railroads [2146, 3561, 4481, 4049, 6137, 5177, 5879, 5257, 5971, 5809, 5376, 6456, 6914, 7249, 7469, 8205, 7976, 8874, 9205, 11628, 10937, 10987, 10961, 720, 10, 834, 11176, 4031, 8966, 3882, 10132]. Rails [1343, 1390, 2724, 3889, 3981, 5949, 8312, 9571, 10707, 10295, 9834, 10058, 11448, 11075, 408, 880]. Railway [1977, 2012, 1318, 1942, 1872, 2353, 2378, 2459, 1538, 2391, 2158, 1528, 2444, 3034, 2843, 3563, 3179, 2872, 3678, 3837, 4596, 4080, 4215, 4314, 4120, 4691, 6008, 5265, 5409, 5302, 5278, 5056, 5215, 5762, 6365, 7041, 6241, 6332, 7338, 6290, 6513, 6538, 7337, 6586, 6805, 8525, 7576, 7623, 8211, 7920, 7695, 7457, 9512, 9427, 9577, 9335, 9240, 8823, 9824, 10686, 9822, 10607, 10507, 10935, 11127, 11185, 10883, 11569, 11106, 11737, 11535, 1114, 949, 1049, 304, 994, 10968, 8206, 8663, 9805, 11197, 6122, 7302, 8682, 5020, 4221, 3766, 1126]. Railway [4828, 2601, 3012, 9288, 8503, 1504, 2210, 3331, 3959, 5178, 5160, 5883, 5134, 6337, 11386, 614, 3333, 4167, 9983, 7472]. Railways [5643, 2279, 2220, 3964, 3974, 4310, 5159, 5900, 8262, 7440, 9818, 934, 3013, 3911, 8702, 2313, 2610]. Rain [5263]. Rain-Water [5263]. Rainfall [4577]. Rainier [854, 5251]. Rainmaking [107, 193]. Rainy [11857]. Raised [7822, 10247, 9751, 10869]. Raisin [2530]. Raising [1918, 1896, 1512, 2710, 4624, 5409, 5312, 6196, 7564, 9232, 10036, 9855, 11869, 6298, 10222]. Rake [7137]. Rakes [1728]. Ramapo [1454, 6905]. Rameses [4463]. Ramsay [7236, 8021, 10733]. Ranch [4246]. Rancocas [4298]. Range [2569, 5210, 7294, 8659, 11085, 643, 295]. Ranges [10561, 6147]. Rank [11130]. Rankin [3463]. Rapid [1408, 1826, 1620, 2483, 2351, 2655, 3078, 3270, 3335, 2815, 3325, 3197, 2883, 2964, 2826, 3361, 3791, 4220, 4708, 4748, 4614, 4414, 3973, 5247, 5713, 5412, 5183, 5179, 5537, 4994, 6536, 6240, 6971, 6266, 6981, 6993, 6455, 8847, 8867, 9799, 9697, 11388, 984, 1094, 7912, 4317, 2070, 2241, 5792, 6780]. Rapid-Fire [3270]. Rapidly [1911]. Rapids [3230, 2402]. Rare [10122, 7504, 3512, 8334, 8505]. Rat [176, 721]. Ratchet [2330, 8810]. Ratchet-Wrench [2330]. Rate [5088, 9596, 9874]. Rates [4773, 10872, 11215]. Ratification [5012]. Rating [9346]. Rational [10000, 8985]. Rats [6619]. Rattle [496]. Raw [5869]. Ray [3975, 6437, 10010, 2686, 1818]. Rayed [8895]. Rays [1964, 2491, 2886, 5305, 6140, 7153, 6289, 6756, 6731, 6995, 6601, 6217, 6495, 10336, 2213, 11888, 812, 2172, 5715, 7293]. Razing [4462]. Razor [10095]. Razors [1604, 10781]. Re [2108, 3789]. Re-Invention [2108, 3789]. Reached [5880]. Reaches [1981]. Reactions [845]. Reading [598].

Readings [4551, 10969]. Ready [10416, 5083]. Real [6495, 9164, 8007, 6630]. Realism [7521]. Reamer [7408, 735]. Reanimation [8105]. Reaper [2641]. Reaping [1831]. Reaping-Machine [1831]. Rear [4130, 6455, 8809, 9948, 2122, 3107, 5974]. Rear-Admiral [3107]. Rear-End [4130]. Reason [7680, 10090]. Reasoning [6874, 9142]. Rebuilding [2014, 9508, 8824, 11447]. Receiver [5100, 9076, 10138, 10817, 8968, 4324]. Receivers [6380]. Recently [1310, 1325, 1923, 1945, 1969, 1993, 2016, 2038, 2064, 2087, 2110, 2137, 2161, 2184, 2204, 2224, 2249, 2269, 2292, 2316, 2342, 2360, 2383, 2403, 2428, 2449, 2477, 2502, 1353, 1375, 1400, 1424, 1449, 1472, 1497, 1521, 1543, 1566, 1588, 1614, 1636, 1662, 1687, 1712, 1736, 1755, 1781, 1808, 1834, 1855, 1877, 1898, 2357, 3021, 2525, 2540, 3124, 3145, 3166, 3189, 3214, 3235, 3255, 3278, 3297, 3322, 3343, 3369, 3391, 3413, 3432, 3450, 3471, 3496, 3515, 3535, 3556, 3580, 3603, 3626, 3648, 3672]. Recently [2560, 2585, 2605, 2628, 2651, 2673, 2695, 2719, 2743, 2766, 2787, 2811, 2833, 2853, 2877, 2901, 2924, 2946, 2972, 2997, 3025, 3049, 3073, 3101, 2785, 2943, 3095, 4527, 3692, 3709, 4305, 4329, 4356, 4379, 4401, 4427, 4447, 4476, 4495, 4518, 4542, 4565, 4591, 4609, 4629, 4653, 4674, 4697, 4722, 4743, 4768, 4793, 4818, 4838, 4860, 4886, 3729, 3749, 3775, 3802, 3828, 3852, 3874, 3897, 3922, 3947, 3968, 3991, 4013, 4040, 4069, 4093, 4113, 4133, 4155, 4179, 4206, 4225, 4248, 4274, 4910, 4923, 5555]. Recently [5587, 5612, 5633, 5658, 5685, 5708, 5727, 5754, 5777, 5818, 5841, 5873, 5894, 5915, 5963, 5991, 6016, 6037, 6064, 6084, 6109, 6130, 6151, 4952, 4984, 5006, 5026, 5061, 5095, 5122, 5151, 5171, 5197, 5218, 5242, 5274, 5296, 5317, 5351, 5370, 5399, 5421, 5448, 5465, 5489, 5513, 5533, 6171, 6190, 6769, 6791, 6816, 6838, 6862, 6882, 6900, 6922, 6950, 6988, 7013, 7030, 7050, 7073, 7097, 7123, 7147, 7175, 7199, 7224, 7244, 7267, 7287, 7311, 7332, 6211, 6229, 6252, 6282, 6301, 6327, 6346, 6369]. Recently [6385, 6409, 6428, 6450, 6470, 6490, 6522, 6546, 6564, 6594, 6613, 6634, 6675, 6700, 6723, 6749, 7937, 7349, 7366, 7970, 7994, 8011, 8034, 8065, 8095, 8114, 8133, 8154, 8180, 8197, 8220, 8244, 8265, 8290, 8308, 8328, 8344, 8367, 8394, 8417, 8441, 8485, 8512, 7388, 7411, 7433, 7462, 7483, 7507, 7530, 7550, 7571, 7591, 7616, 7640, 7665, 7687, 7712, 7741, 7759, 7777, 7817, 7843, 7870, 7895, 7924, 7947, 9585, 8535, 8557, 9133, 9156, 9176, 9197, 9221, 9236, 9256, 9276, 9298, 9316, 9340, 9360, 9383]. Recently [9408, 9435, 9455, 9475, 9496, 9516, 9534, 9556, 9574, 9591, 9610, 9635, 9659, 8577, 8603, 8628, 8647, 8669, 8694, 8711, 8728, 8754, 8776, 8818, 8842, 8862, 8882, 8908, 8927, 8952, 8972, 8997, 9019, 9038, 9059, 9085, 9106, 9679, 9692, 10270, 10287, 10307, 10327, 10345, 10368, 10388, 10408, 10430, 10485, 10510, 10532, 10550, 10568, 10599, 10616, 10634, 10661, 10678, 10698, 10724, 10743, 10763, 10785, 9714, 9737, 9761, 9785, 9807, 9828, 9848, 9870, 9888, 9913, 9969, 9995, 10013, 10030, 10051, 10072, 10097, 10116, 10142]. Recently [10164, 10183, 10207, 10229, 10250, 10802, 10839, 11420, 11442, 11469, 11486, 11506, 11529, 11551, 11576, 11592, 11620, 11646, 11670, 11693, 11742, 11764, 11790, 11812, 11836, 11851, 11870, 11890, 11908, 11927, 10862, 10890, 10911, 10932, 10952, 11005, 11024, 11048, 11069, 11097, 11122, 11141, 11161, 11180, 11205, 11226, 11245, 11269, 11288, 11310, 11336, 11360, 11380, 11404, 2, 18, 666, 693, 711, 738, 764, 784, 816, 838, 876, 907, 928, 951, 975, 1006, 1034, 1062, 1082, 1107, 1131, 1155, 1183, 1205, 1239, 1264, 1285, 40]. Recently [73, 99, 131, 152, 180, 206, 245, 270, 298, 330, 353, 375, 402, 421, 447, 469, 499, 521, 543, 574, 601, 623, 645, 8594, 10304]. Receptacle [6587, 1303, 779]. Receptacles [10946]. Reception [1529]. Recession [7984]. Recharging [2333, 9121, 9503]. Reciprocating [6220, 8100, 7471, 9351, 10769, 10789, 11341]. Reciprocating-Cylinder [6220]. Reciprocating-Engine [10769]. Reciprocity [125]. Reckoner [7676]. Reclaiming [4234]. Reclamation [9539, 10555, 1071, 7751, 7321]. Recoil [2373, 9965]. Recommendation [4184]. Reconstruct [5129]. Reconstructed [8896, 10614]. Reconstructing [9903]. Reconstruction [2103, 1801, 2129, 3585, 3455, 2934, 3819, 3879, 5406, 5415, 6599, 6182, 10560, 11457]. Record [1767, 2486, 1890, 3285, 2548, 3926, 3860, 4457, 5910, 5358, 5795, 5043, 5732, 5278, 6662, 7231, 7307, 6851, 7362, 7838, 8896, 9160, 10084, 10452, 10705, 11896, 10991, 11129, 11372, 11046, 11676, 1217, 1276, 968, 823, 697, 5603, 1050, 7040]. Record-Breaking [6662, 11372]. Recorder [9552, 1560]. Recorders [4062, 8922]. Recording [10760, 9826, 8683, 9674, 5525, 10198, 3847]. Records [1979, 2334, 2680, 3444, 2958, 4053, 5108, 5736, 5461, 6095, 7305, 7064, 7863, 8384, 7823, 9035, 10687, 10525, 10868, 11847, 1254, 1193, 11147]. Recovering [9306]. Recovery [4550]. Recreation [11858]. Rectangular [11146]. Rectifier [9071, 10776, 1260]. Recurrence [11436]. Red [5764]. Redheugh [4263]. Redivivus [10795]. Reducer [267]. Reducing [3112, 5255, 6978, 233, 643, 295]. Reductio [11817]. Reduction [11170, 11215]. Redwood [8958, 4807, 7115]. Reef [6260, 9425]. Reel [4322, 11201]. Reeling [6858]. Reference [6957]. Refillable [1440, 9628, 9430, 11868]. Refining [3742, 3743]. Reflecting [5236, 7736, 8904, 11151]. Reflector [9955, 4245]. Reflex [4103]. Reform [1331]. Refractory [5497, 5138]. Refrigerating [12]. Refrigeration [1067, 9034, 1151]. Refrigerator [7709]. Refrigerators [3119]. Refuse [5823, 5772, 8678, 10021, 779, 10946]. Refutation [7181]. Regan [3824]. Regarding [1976, 10337]. Regas [6648]. Regelation [4967, 5528]. Regiomontanus [8475]. Region [2370, 3617]. Regions [6529]. Register [810]. Registering [3006, 8107, 7425, 10561, 1055]. Registration [2280, 7235, 6775, 7112]. Regrading [148]. Regular [5107]. Regulate [11088]. Regulating [2284]. Regulation [1843, 8937, 10211, 9856]. Regulator [3799, 5292, 6515, 8839, 11333, 806]. Regulators [4412, 9596]. Rein [1418, 5551]. Reindeer [4788, 7329]. Reinforced [8764, 8761, 8771, 8765, 10215, 11413, 10900, 10871, 9095, 8405, 11326]. Reinforcement [10986]. Reinforcing [7582]. Reinterment [5106]. Rejected [1479]. Related [8227, 11521]. Relation [1930, 6637, 8719, 11859, 11366, 4877, 4900]. Relations [10881]. Relative [2299, 3493, 3035, 3638, 6683, 10666]. Release [5340, 5382]. Reliability [4488, 4507]. Reliable [10475, 1242]. Reliance [5285, 5136, 5063, 5225, 5187, 6021, 5601]. Relic [7113]. Relics [1608, 6099, 6560, 962]. Relief [1905, 3302, 3755, 6804, 10087, 9957, 10415]. Reliefs [11260]. Relieved [7749]. Religiosa [3067]. Relining [486]. Relio [6653]. Relocation [3561]. Remains [1575, 2617, 7163, 8210, 4151, 4785]. Remarkable [1854, 1770, 1694, 2816, 3597, 4349, 4578, 4120, 5302, 5908, 8891, 9231, 9751, 10224, 10280, 10575, 3383, 5774, 6081, 6608, 7480, 5037, 2243, 11197, 2645, 4902, 9055, 917]. Remedy [2255, 9401, 202]. Remodeling [3194]. Remora [4810]. Remoras [8529]. Removable [3447, 6106]. Remover [6406]. Removing [1396, 2848, 4734, 11177, 867, 1203, 2132]. Remsen [5207]. Renaissance [7656, 965, 156]. Renard [7295]. Renovator [11089]. Reorganization [1086]. Reorganize [11769]. Reorganizing [3738]. Repair [10471, 8144, 418, 5384]. Repairing [1825, 2781, 3096, 4172, 11475, 1295, 11414]. Repairs [5519, 734]. Repeated [7646, 11305]. Repeater [3289]. Repeating [3722]. Repeats [7783]. Repel [812]. Replaced [487]. Replacer [5141]. Replacing [6125, 8969]. Report [1335, 2326, 2300, 2298, 2044, 2122, 1844, 3327, 3058, 3376, 4634, 3882, 5104, 5159, 5110, 6216, 9366, 11166, 11571, 10988, 628, 715, 1115, 3523, 2726, 1116]. Reporting [5926]. Representatives [7826]. Representing [5790]. Reproduction [1470]. Reptile [6609]. Reptiles [10950]. Republic [55, 548]. République [11828]. Requirements [11426]. Rerolling [2724]. Rescue [8490, 11541]. Rescuing [10453]. Research [9507, 7320, 550, 999, 282]. Reservation [2014]. Reservoir [6043, 5760, 6778, 660, 3129, 2845, 4185, 5354, 7581, 7496, 7961, 8766, 9002, 9647, 239]. Reservoirs [8054, 8959, 10255, 6505]. Resignation [2615, 2595]. Resilient [6338]. Resistance [2508, 1668, 1676, 3034, 5311, 5783, 9664, 8894, 8833, 10244, 11905]. Resonance [4877, 4900]. Resonator [4470, 1308]. Resources [2411, 11387, 11128, 11102, 407]. Respiration [7418]. Respirator [5621, 5828, 10091]. Respiratory [11066, 6892, 8030]. Responsibility [6868]. Responsible [10491]. Rest [11823]. Restaurant [6032]. Restoration [4663, 4974]. Restored [7772]. Restoring [871]. Restrictions [11626]. Result [4614, 10083, 858]. Results [2816, 3562, 2750, 6717, 6457, 6217, 7105, 8435, 9264, 9231, 10683, 10713, 604, 11842]. Resuscitation [4782]. Retaining [8751, 8773]. Retardation [9168]. Retirement [3107]. Retrograde [8736]. Retrospect [1316, 2531, 3695, 4913, 6176, 7353, 8332, 9287, 10807, 6]. Return [4702, 4407, 742, 306, 2887]. Returns [4975, 8448]. Retvizan [3314]. Reuleaux [10881]. Reuterdahl [11231, 11261]. Reverse [6362]. Reversible [2152, 4665, 6264]. Reversing [1870, 6813, 6322]. Reversion [1085]. Review [9670, 9683, 6198, 6757, 4899]. Revised [3999]. Revision [10681]. Revival [1527, 1883, 4253]. Revolution [821, 10739]. Revolutionizing [3477]. Revolving [4618, 5672, 7062, 8794, 10978, 2983]. Revolving-Cylinder [8794]. Rexer [8502]. Rheims [680, 670, 707]. Rheostat [7111, 7562, 736].

Rhine [8685]. Rhode [8355]. Rhodesia [8594]. Ribbon [6101]. Ribs [10137]. Ricaldoni [10439]. Rice [9291]. Richard [4934, 9307, 2706, 4099]. Rickmers [9252]. Rico [1788, 3680, 7274]. Riddle [10500]. Ride [4685]. Riding [6740]. Rieder [2156]. Rifle [2105, 1381, 5560, 5313, 9779, 10716, 10561, 280, 643, 295]. Rifle-Caliber [10716]. Rifle-Propelled [280]. Rifled [1446]. Rifles [8277]. Rigged [9704]. Right [194]. Right-Angled [194]. Righting [9483]. Rigidity [9617]. Rim [7447, 9566]. Ring [5888, 11547, 97]. Rings [853, 10540, 10925]. Rise [5014, 8892]. Risk [8041, 9578, 10666]. Risks [1794, 10254]. Ritchie [2349]. Rite [4065]. Rittenhouse [10317]. Rival [4458]. River [1800, 2083, 1561, 3404, 2688, 3106, 3078, 2893, 2723, 3240, 3181, 4635, 3806, 4214, 4612, 4784, 4623, 4253, 4028, 6029, 4972, 6045, 5792, 6142, 5982, 5519, 5376, 4994, 6235, 6358, 6269, 6178, 6915, 6679, 7195, 7261, 8071, 8168, 9391, 9261, 9204, 9418, 8867, 10936, 11944, 11293, 11351, 10845, 1300, 681, 10630, 7137, 2460, 2523, 2843, 2916, 2685, 3593, 3161, 3740, 6333, 6626, 6780, 9459, 9509, 8582, 9665, 8685, 9587, 9364, 9700, 9795, 11282, 755, 748, 747, 155, 789, 171, 8690, 5477, 8707, 11462, 1990, 8500, 9072, 555, 9886, 11457]. River [704, 10920]. Rivers [11232]. Riverside [2377, 3592]. Rivet [5279]. Riveter [6467]. Riveting [8499]. Road [3398, 4081, 4691, 5749, 5311, 6954, 6998, 9402, 9186, 9345, 9719, 11724, 11556, 11329, 596, 5111, 7584, 7379, 1671, 2054]. Roadbed [1627, 9448]. Roadbeds [3316]. Roads [1744, 4868, 4334, 7104, 6309, 9324, 9181, 10257, 11315, 11031, 11389, 185, 957, 284, 233, 3961, 5699, 10453, 7372]. Roadway [5527]. Robert [3503, 4872, 8272, 155, 1746]. Rochambeau [4026]. Roche [5090]. Rochester [3067]. Rock [4962, 6511, 9639, 10180, 8640, 10361, 5630]. Rockefeller [282]. Rocking [9912]. Rocks [11698]. Rocky [1917, 7914]. Rod [5145, 10386, 11199, 1237]. Rodney [2665]. Rods [1371, 4146, 8944, 9845]. Roentgen [3975, 10336, 11888]. Rojestvensky [7512]. Roll [10641]. Roller [4301, 9402, 8686, 9933, 10136, 9912, 11709]. Rolling [2395, 1705, 9512, 3228, 7379]. Rolling-Lift [1705]. Roman [1720, 4064, 3894, 5908, 6296, 7141, 7163, 8402, 7067, 9013]. Romance [7339, 6976]. Romans [2910]. Rome [1753, 5356, 11218, 8495, 5564, 11219, 11437]. Roof [3307, 6106]. Roofs [812]. Room [1659, 3374, 8873, 5906, 9804]. Rooms [10749]. Roosevelt [11398, 8385, 8333, 11102, 11769, 45]. Roots [11925]. Rope [1439, 6514, 9928, 11834, 9805]. Rope-Drive [9928]. Ropeway [6079]. Rosa [7500]. Rosalina [1897]. Rose [8551, 2760]. Roses [1031]. Rosin [1910]. Rotagons [199]. Rotary [2497, 2622, 3965, 4512, 6200, 8059, 8509, 9924, 1220, 10155, 11833]. Rotates [11285]. Rotating [9194, 7912]. Rotation [8504, 1779]. Rotations [4735]. Rothwell [2706]. Rotorua [7240]. Round [8745, 568, 11788, 1894]. Round-Lap [1894]. Route [2210, 2858, 4254, 4824, 4161, 4184, 8424, 10750, 9269]. Routes [4148, 6276, 7793]. Rowboat [8029]. Rowland [2705, 4510]. Royal [1975, 3376, 3242, 3262, 3079, 9947, 8597]. Royalties [3543]. Royalty [11775]. Roze [2336]. Rr [218]. Rubber [2498, 2471, 2351, 4876, 5734, 5109, 10380, 11323, 10899, 121, 3359, 2798, 939]. Rubies [417]. Rugs [3660]. Ruhmer [4532, 5314]. Ruins [4105, 6445, 4919, 8594, 9731]. Rule [5064, 11784]. Ruler [9379]. Rules [7166, 9857]. Ruling [6857]. Run [2268, 3926, 5737, 5910, 5795, 6088, 8400, 9977, 1046]. Runabout [5079, 6657, 6656, 8791, 8789, 8790, 6665]. Runner [2986, 6540, 902, 541, 1101]. Runners [637]. Running [10021, 476, 3093]. Runs [4120, 7231, 10084, 9310]. Rupture [9212]. Rural [6276, 6558, 7649, 9588]. Ruskin [1820]. Russia [1764, 2083, 2150, 6618, 6207, 6318, 7258, 6198, 6757]. Russian [1816, 1887, 2640, 2616, 2822, 3314, 3784, 4350, 4786, 6764, 6982, 6462, 6395, 6737, 6362, 6446, 6507, 7632, 8519, 7385, 7726, 8852, 8824, 1601, 6477]. Russo [7182, 6844, 7731, 7999]. Russo-Japanese [7182, 6844, 7731, 7999]. Rust [1079]. Rusting [10731].

S [3685]. S. [3685]. Saalburg [9013]. Sad [2623, 1303]. Sad-Iron [2623, 1303]. Saddle [6740, 4788]. Safe [3808, 7929, 7976, 453]. Safeguarding [1047]. Safeguards [9460]. Safes [7184]. Safety [2509, 2308, 2579, 2841, 3505, 2538, 3608, 5763, 5986, 5192, 5025, 5023, 5053, 5386, 6760, 6713, 8476, 7554, 8204, 7763, 9432, 9064, 9170, 8810, 9082, 10156, 10039, 10781, 10095, 9789, 9781, 10649, 10159, 11321, 11185, 10958, 11103, 11332, 11358, 10961, 10965, 11307, 10815, 11819, 5020, 11476, 10003]. Sahara [4502]. Sail [2828, 5187, 5137, 11371]. Sailed [616]. Sailing [2979, 4526, 3787, 5229, 5922, 9252, 10907, 2009]. Sails [2399]. Salaries [11186]. Sale [4732]. Salem [4692, 348, 10757, 10587, 11637]. Salmon [8323]. Salon [7792, 509]. Salt [1678, 4236, 5128, 6753, 6257, 11804, 7449, 3843, 9373, 2645, 4789, 5769]. Salt-Water [6257]. Salton [9509, 9415, 8690, 8707]. Salts [5306]. Salute [9562]. Salvage [2519, 3382, 3195, 10945, 1016, 3403, 4780]. Salving [9751, 11585]. Same [11844]. Sampling [6277]. Sampson [3978, 11329]. Samuel [6204, 8568]. San [4973, 8836, 1600, 1884, 1396, 2551, 4254, 9616, 8899, 8958, 8722, 8887, 8770, 10057, 7449, 1586, 9306]. Sand [3508, 4202, 8061, 10499, 11238, 10829, 9565, 3577, 9098]. Sandhurst [4335]. Sanding [9377]. Sandpaper [11689]. Sands [4268]. Sandstone [5503]. Sandy [3384, 3442, 3435, 2816, 4707, 5470, 2825, 9502]. Sanitary [3906, 5331, 8147, 11535, 671, 441, 779, 9568, 5770]. Sanitation [8226, 8422, 8573, 9648, 9166, 9669]. Santa [1102, 4319, 5185, 3702]. Santiago [3094, 3881, 613]. Santos [3221, 1938, 3248, 3419, 3631, 3620, 3250, 3408, 3163, 4262, 4711, 5813, 5832, 7129, 7016, 8127, 9419, 9325, 10286, 10613, 10426, 11756, 9050]. Santos-Dumont [3221, 3419, 3631, 3620, 3250, 3408, 3163, 4262, 4711, 5813, 5832, 7129, 7016, 8127, 9050]. Sap [1046]. Sapphires [11779, 9862]. Sargasso [9618]. Sash [2421, 3687, 10047, 10929]. Sash-Weight [2421]. Sashes [8215, 8437]. Satellites [6508]. Satisfactory [10189]. Satsuma [10130]. Saturn [6072, 853, 10610, 10540, 10925, 6118]. Sault [2774, 4529, 1630, 4100]. Saurian [10614]. Saurians [3318]. Savages [3552]. Saving [2267, 4525, 5036, 6486, 6832, 8390, 9466, 9702, 10535, 9858, 10850, 11238, 595, 441, 32, 6785, 9529, 6465, 829, 8302, 6956]. Savoie [3198]. Saw [2053, 3366, 3010, 4353, 4538, 5339, 9336, 9907, 11157, 11202, 1024, 1154, 399, 1261]. Saw-Clamp [4353]. Saw-Filing [1261]. Saw-Operating [9907]. Saw-Set [9336]. Saw-Sharpener [3366]. Saw-Tightener [5339]. Saw-Tooth [4538, 1024]. Sawdust [7498]. Sawing [10885, 11480]. Sawmills [6200]. Saws [5907, 10970]. Saxe [4267]. Saxe-Coburg-Gotha [4267]. Scaffolding [8945]. Scale [2055, 1511, 3468, 6463, 6994, 7758, 1141, 6714, 7709]. Scandal [7555]. Scarcity [4799, 4163]. Scarifier [2104]. Schebler [7092]. Schedule [7763]. Scheme [3540, 4253, 7338, 6905, 10393, 10604, 10169, 10212, 11029]. Schemes [8979]. Schenectady [11513]. Schlick [9009, 10067]. Schneider [2257, 4904, 4345, 87]. Scholarships [3856]. School [1412, 2124, 2642, 5750, 5956, 6533, 8854, 9350, 11541, 9080, 7886, 1170, 3666, 7446, 8004, 4200, 10530, 1256, 8379, 10673]. Schoolboy [11916]. Schools [11391, 6465, 1654, 9819]. Schooner [3292, 4829, 5070, 11438, 4780]. Schooners [11723]. Schultz [2416]. Schweinfurt [6339]. Science [1458, 1745, 1551, 1982, 2214, 2259, 1541, 1864, 2568, 2583, 2671, 3566, 3590, 3612, 3636, 3659, 2741, 2786, 2810, 2832, 2875, 2900, 2922, 2945, 2970, 3023, 3047, 3072, 3135, 3200, 3222, 3246, 3308, 3329, 3357, 3379, 3402, 3439, 3460, 3546, 4571, 3718, 3759, 3813, 3839, 3861, 3885, 3934, 3956, 4001, 4217, 4237, 4258, 4425, 4446, 4474, 4516, 4540, 4590, 4651, 4695, 4720, 4816, 4859, 4885, 4937, 4950, 5829, 5881, 5904, 5927, 5975, 6024, 4982, 5003, 5093, 5120, 5149, 5182, 5216, 5253, 5281, 5294, 5316, 5348, 5368, 5396]. Science [5419, 5463, 5511, 5596, 5622, 5644, 5669, 5717, 5740, 5788, 5618, 6966, 6199, 6239, 7120, 7160, 7173, 7197, 7223, 7265, 6262, 6291, 6312, 6336, 6344, 6377, 6418, 6439, 6531, 6553, 6621, 6733, 6758, 6779, 6802, 6872, 6887, 6948, 7011, 7028, 7047, 7071, 7108, 7420, 7442, 8354, 8403, 8427, 8453, 8474, 8496, 8522, 7494, 7516, 7559, 7626, 7652, 7673, 7699, 7768, 7788, 7828, 7934, 7979, 8024, 8103, 8122, 8142, 8165, 8187, 8230, 8254, 8276, 8297, 8316, 8546, 8576, 8681, 8786, 8829, 8982, 9028, 9070, 9094, 9266, 9274, 9370]. Science [9421, 9505, 9726, 9980, 10040, 10218, 10278, 10298, 10316, 10355, 10497, 10520, 10539, 10608, 10646, 10735, 10754, 10773, 10967, 11172, 11298, 11394, 11477, 11773, 11821, 11880, 11917, 11939, 11518, 11538, 11561, 11602, 11654, 11680, 11701, 11751, 11800, 1052, 682, 965, 26, 51, 609, 633, 655, 676, 702, 719, 746, 772, 795, 826, 849, 886, 916, 938, 961, 988, 1014, 1044, 1070, 1090, 1120, 1139, 1164, 1191, 85, 1216, 1273, 1294, 110, 140, 162, 190, 217, 236, 255, 279, 309, 362, 386, 412, 430, 456, 507, 529, 553, 586, 392]. Science [4415, 11916, 1935, 2026, 6602, 9368, 4751, 5413, 5207, 3175]. Sciences [3884, 2236, 1506, 2703, 3421, 4637, 5181, 7557, 8740, 11036, 9462, 1529]. Scientific [2074, 1670, 4750, 4098, 3811, 4046, 5593, 5312, 6317, 6559, 6976, 7706, 9552, 8725, 10756, 11256, 88, 17, 1169, 550, 831, 4710, 6955, 7911, 510, 4493, 10857, 4196, 5325, 3347, 4933, 5393, 5865, 6957, 7336, 8228, 10299, 10794, 10423, 10320, 9857, 9833, 9789, 10321, 10438, 11150, 11641, 11400, 10974, 11815, 11011, 11830, 82, 1172, 1198, 104, 113, 855]. Scientifically [7045]. Scientist [3907, 9523, 11233]. Scientists [1457, 3565, 10917]. Scissors [1024]. Scoop [9992]. Scooter [291, 1105, 6208]. Scope [9854, 6398]. Scotch [1144]. Scotia [62]. Scotland [1608, 3395, 5562, 2243]. Scotte [2309]. Scout [10757, 10587, 11776, 10896, 11637, 305, 348, 1136, 7858]. Scouting [11425]. Scouts [2889, 7910]. Scow [5592]. Scrap [4405, 7439]. Scraper [1960]. Scrapers [2331]. Scraping [485]. Screen [2013, 1556, 8598, 810]. Screw [1410, 2218, 2755, 4289, 4665, 3904, 4299, 7194, 8321, 8190, 9673, 9118, 9823, 1121, 11612, 973, 9124, 3009]. Screw-Eyes [2755]. Screw-Propelled [1410, 9673]. Screw-Propeller [2218, 4665]. Screw-Slotting [11612]. Screwdriver [7026]. Screws [5039, 11018, 202, 867, 1203]. Scroll [399]. Scroll-Saw [399]. Scrubber [11607]. Scrubbing [7062]. Sculpins [2356]. Sculptural [6577]. Sculptures [8813]. Sculpturing [5722]. Sea [1825, 2210, 3420, 2640, 3096, 2959, 3241, 4578, 4164, 4935, 6461, 7316, 7132, 6318, 7136, 7704, 7786, 7415, 8026, 7416, 8204, 8209, 8651, 9642, 8735, 10379, 11884, 11527, 112, 23, 1066, 1270, 7943, 9838, 5389, 9127, 1519, 6101, 5913, 6424, 6669, 9290, 8878, 8530, 11354, 8413, 10175, 962, 9509, 9618, 9415, 8989, 10258, 8707]. Sea-Going [10379]. Sea-Level [7416, 112]. Sea-Rafts [10175]. Sea-Sounding [7786]. Seacoast [611]. Seal [9866, 9816]. Sealer [2079]. Sealing [4513, 5022]. Search [1828, 3739]. Searching [9164, 10315]. Searchlight [3862, 8659, 10266]. Searchlights [4603, 55, 6135]. Seas [3350, 6917, 6496, 5654]. Seasickness [8987, 9401, 7209]. Season [10104, 11857]. Seat [2873]. Seats [10818, 11834]. Seawanhaka [8448]. Seaworthiness [145]. Second [1883, 6580, 7836, 8399, 9722, 11134, 10845, 23, 4994, 8170, 9487, 11130, 10965, 1067, 1143]. Second-Class [1883]. Secondary [5305]. Secret [6124]. Secrets [5946]. Section [2011, 1984, 5211, 7261, 10962]. Sectional [5978, 6561]. Sections [7649]. Secure [11075]. Securing [4320]. Sedgwick [3679]. See [2414, 1188, 258, 9617]. Seed [5812, 4947, 791, 7038]. Seed-Cleaner [4947]. Seedless [7830]. Seeds [8960]. Seeing [11219, 863, 11396]. Seem [8895]. Seen [4372]. Segment [11905]. Segmental [3349, 3160]. Seine [9072]. Seismic [4045]. Selden [2242, 5045, 9564, 91, 743]. Selecting [477]. Selection [7721]. Selective [2554, 8102, 8795, 4997, 9488]. Selective-Type [8795]. Selenides [3133]. Selenium [5700, 11827, 70]. Self [1747, 2623, 2938, 2898, 3202, 4195, 3887, 5022, 8174, 9867, 10136, 10561, 11139, 1021, 373, 1026, 3645, 6909, 4060, 4939]. Self-Acting [1021]. Self-Binders [9867]. Self-Cleaning [8174, 373]. Self-Closing [2938]. Self-Defense [6909]. Self-Dumping [4060]. Self-Feathering [4195]. Self-Generated [11139]. Self-Heating [2623]. Self-Launching [1747]. Self-Measuring [2898]. Self-Oiling [10136]. Self-Propelling [3202, 3645].

Self-Registering [10561]. Self-Sealing [5022]. Self-Smoking [1026]. Self-Starting [3887]. Self-Winding [4939]. Selling [9484]. Semi [3599, 3488, 4390, 11919]. Semi-Arid [11919]. Semi-Armored [3488]. Semi-Automatic [4390]. Semi-Dry [3599]. Senate [4118, 8887, 10915, 11011]. Sending [6126, 11110, 11354]. Sense [4138, 9369, 935, 8747]. Senses [5238, 5213]. Sensible [4047, 1243]. Sensitive [9601, 556]. Separate [10603]. Separation [2290, 10152]. Separator [5267, 6009]. Separators [4947]. September [2500, 3671, 4326, 5585, 6789, 7968, 9657, 10793, 11402, 1284]. Serpent [6101]. Serpents [5508, 4944, 5547]. Serpollet [1850, 2287]. Serum [2371, 10526, 7301]. Service [1430, 2497, 2060, 2598, 3862, 3909, 5302, 5107, 5824, 5782, 5783, 6290, 6558, 8100, 7471, 8823, 9877, 10496, 10714, 10132, 10684, 10020, 11483, 11944, 11187, 11459, 11819, 983, 7342, 6833, 7109, 1623, 2711, 2793, 6237, 9561, 11428, 11915, 11636, 913, 9529, 11655, 11433, 3355]. Serviceable [9124]. Session [1798]. Sessions [6871]. Set [9336, 7701]. Sets [9126, 466]. Setting [3010, 4053, 4876]. Settled [2231, 9774]. Settles [1541]. Seven [3292, 4829, 8652, 10079, 11369, 706]. Seven-Foot [10079]. Seven-Hundred-Mile [8652]. Seven-League [11369]. Seven-Masted [3292, 4829]. Seventh [9480, 9919, 10297, 1529]. Severe [9644, 11131]. Severn [11000]. Severo [4054, 3793, 4003]. Sewage [8317, 6955]. Sewerage [3465]. Sewers [8147]. Sewing [5607, 6627, 9911, 10694]. Sextants [10628]. Sfax [2030]. Shackleton [135, 10037, 11936]. Shade [7220, 4348, 11717]. Shade-Grown [4348]. Shades [565]. Shadows [6735]. Shaft [1825, 2176, 3096, 4078, 11463, 1262]. Shaft-Keyways [2176]. Shafts [2579, 4344, 1153]. Shall [5127, 11363]. Shallow [4049, 9770]. Shallow-Draft [9770]. Shallow-Tunnel [4049]. Shampooing [6165]. Shamrock [3667, 3219, 2841, 2758, 2795, 3654, 2828, 5071, 4993, 5380, 5137, 6021, 5601]. Shamrocks [6115]. Shape [6605, 9305, 10657, 414]. Shaped [8575]. Shaper [9689]. Shapes [5952]. Shapira [4778]. Shark [9569, 8055]. Sharks [1634]. Sharpener [3366, 325]. Shattered [4759]. Shaving [11948]. Shawenegan [4150]. Shear [8082]. Shearer [5833]. Shearing [8322]. Shears [6242]. Sheathing [2045, 10160]. Shed [11177]. Shedding [2467]. Sheep [5495, 5833, 10799]. Sheep-Killing [10799]. Sheep-Shearer [5833]. Sheet [7144, 9445, 725]. Shell [3442, 3435, 3792, 6982, 8489, 7526, 8189, 8642, 11942, 8084]. Shell-Digger [7526]. Shellfish [8334]. Shells [3486, 9231, 8001]. Shields [1651]. Shifting [10971, 10908]. Shikishima [1517]. Shingle [6583]. Shingles [8622]. Ship [2379, 1869, 2256, 1441, 4938, 6070, 6156, 6094, 5653, 5562, 7256, 6917, 6479, 7767, 8385, 9292, 9252, 9751, 9843, 9858, 11752, 11607, 10856, 11449, 11000, 774, 4315, 1938, 5310, 2772, 4453, 4188, 9090]. Ship-Building [2256]. Ship-Subsidy [2772]. Shipbuilders [10191]. Shipbuilding [4748, 4578, 6116, 5471, 9003, 10436, 11632, 10871, 11034, 6417]. Shipping [2863, 4773, 5429, 11356, 11045, 1894, 8585]. Ships [1410, 3632, 2581, 3054, 4387, 4800, 5117, 5866, 5040, 5405, 5608, 7821, 9144, 10580, 10581, 10582, 10583, 10585, 10624, 10067, 9895, 10642, 11166, 11238, 11364, 241, 8745, 10198, 1142]. Shipwrecks [4171]. Shipyards [4992, 9008, 4006]. Shoal [6537]. Shoals [8619]. Shock [8811, 9953, 10465]. Shock-Absorbers [8811]. Shocking [6992]. Shoe [3613, 6406, 10323, 1230]. Shoe-Plank [3613]. Shoeing [9403]. Shoes [8990, 10135, 10285, 10901, 486]. Shooting [2301]. Shoots [11925]. Shop [4941, 9403, 8527, 3046]. Shops [4727, 9044]. Shore [10105]. Short [8659]. Shot [1707]. Shot-Gun [1707]. Should [8978, 10191, 10255, 11293, 232]. Shouldering [1627]. Shovels [3469, 6076]. Show [1873, 2222, 2183, 2993, 3409, 3428, 5474, 5460, 6686, 8351, 7785, 9119, 9480, 10517, 10397, 10647, 10626, 9919, 525, 1373, 10263, 1853, 2028, 3016, 4277, 5411, 6638, 6639, 6652, 7942, 10474, 583]. Showing [9429]. Shown [2309]. Shows [3860, 9006, 10451]. Shrink [4122]. Shrinkage [1595]. Shutter [2374, 1365, 4375, 9775, 11572, 35, 8524]. Shutter-Fastener [1365]. Shutters [6316, 395]. Siberia [3564, 2735, 4788, 4008, 8337, 1534]. Siberian [4702, 1504, 2210, 5912, 6337, 5189, 1420]. Sicilian [4461]. Sicily [6932]. Side [5692, 7802, 10686, 10435, 9697, 11625, 11921, 157, 629, 590, 7342, 1320]. Side-Chair [7802]. Side-Door [5692, 9697, 7342]. Sideboard [5145]. Sideboard-Rod [5145]. Sides [194]. Sidewalk [3781]. Siege [7609]. Siemens [6604]. Sierra [3639]. Sievert [5694]. Sifter [3939, 7989]. Sigafoos [347]. Sight [8271, 7678, 9779, 11266]. Sighted [1290]. Sighting [5021]. Sights [7304]. Sign [3389, 5013, 4443]. Signal [4730, 3842, 6074, 9577, 10809, 11192, 6122, 6833, 3355]. Signaling [1919, 5021, 8874, 8658, 9335, 10064, 720, 834, 368, 251, 461, 1147, 7448, 3017]. Signals [4688, 3229, 2959, 4927, 10688, 11316, 534, 10109, 4468]. Significance [767, 10577]. Signing [5469]. Signs [2458, 8804]. Silent [8802, 612]. Silicon [4548, 9732]. Silk [2517, 3091, 9754]. Silk-Producing [2517]. Silkworm [4102]. Siloxicon [5138]. Silver [4704, 7545, 10280, 11632, 1653]. Silverware [6912]. Simbel [8382]. Similar [11949]. Simile [4604, 10219]. Simms [3986]. Simon [1169]. Simonis [10357]. Simple [1511, 1488, 1986, 1464, 2306, 2497, 2154, 1365, 1581, 1732, 3334, 2897, 2737, 3119, 3097, 3532, 3662, 2780, 3010, 3727, 3771, 4261, 3698, 5581, 5256, 5550, 6272, 7135, 6624, 7918, 8363, 7562, 7447, 8925, 9191, 8553, 8638, 10213, 10356, 9961, 11468, 11526, 178, 395, 1234, 1094, 567, 9080, 735, 689, 4324, 11545, 11785, 2291, 2314, 2963, 970, 869, 10833, 323, 10995, 10279, 11549, 10592, 420, 9527, 686, 3156]. Simple-Cylinder [10213]. Simplex [3841]. Simplified [852]. Simplifies [6688]. Simplon [2592, 3407, 4827, 7058, 8663, 7404]. Simply [656, 942, 10879]. Simpson [1746]. Simultaneous [6534]. Simultaneously [1122]. Since [2392, 3476, 7445, 9190]. Sing [8]. Singer [10301]. Single [1538, 1528, 1650, 2580, 4680, 6264, 8137, 8796, 9345, 8542, 10542, 10667, 10450, 10020, 10480, 252, 8359, 11619, 10997]. Single-Cylinder [6264, 8796]. Single-Phase [8137, 9345, 8542, 10450, 252]. Single-Plate [10667]. Single-Pull [2580]. Single-Rail [1538, 1528, 1650]. Single-Service [10020]. Singular [5326, 2624, 5652, 2960]. Sink [126]. Sinker [2439]. Sinking [4344, 4028, 4078, 11463]. Sinusoidal [10498]. Siphon [4586, 10622]. Siphons [11616, 444]. Sir [6098, 8737, 3646, 3079, 5408, 6010, 5131, 7236, 9307, 10689, 10059, 11010]. Siren [9900]. Sirup [732]. Sisal [3912]. Site [6778, 10008]. Sites [8521]. Situation [80, 617, 10668]. Six [8231, 8792, 10583, 9699, 10276, 2182]. Six-Cylinder [8231, 8792]. Six-Track [9699]. Sixteen [5922, 9123]. Sixteen-Cylinder [9123]. Sixteen-Knot [5922]. Sixth [8383]. Sixty [1080, 9113, 9676, 7109]. Sixty-Foot [9676]. Size [2693, 5479, 7252, 6605, 9073, 9417, 10056, 11844, 7981, 8141]. Sized [11933, 459, 136]. Skate [9049]. Skates [8686, 294]. Skein [3272]. Skeleton [4557, 7138, 10614, 10756]. Skeletons [1098, 10950]. Sketching [8362]. Ski [7861]. Skid [7807, 8574]. Skidding [6574, 8091, 8946]. Skimmers [416]. Skin [2467, 10246]. Skin-Shedding [2467]. Skull [7024]. Skyscrapers [11131]. Slabs [7502]. Slaby [2546, 2554, 3525]. Slack [8571]. Slag [880]. Slat [1304]. Slate [4349, 10205]. Slates [11348]. Slaughter [6092, 7900, 10018]. Slaughtering [11635, 9568]. Slaughters [5020]. Sled [4292, 3798, 8924, 11132, 541, 641, 1101]. Sleds [11666]. Sleeper [9797]. Sleeping [6623]. Sleet [6406]. Sleigh [7814, 8638, 902]. Sleigh-Runner [902]. Sleighs [146]. Slicer [5483]. Slide [11784]. Slides [1386, 6974, 7651]. Sliding [9940, 9943]. Sliding-Gear [9940, 9943]. Slime [5838]. Slip [10471]. Slipping [366]. Slotting [11612]. Slow [5537]. Sluice [3266]. Sluicing [148]. Small [1626, 2693, 2557, 6734, 6446, 9116, 9115, 8913, 10858, 3185, 11785, 7709, 7546, 1202, 11374, 661]. Smaller [2792]. Smallest [8404, 9958]. Smell [8747]. Smelling [11396]. Smelter [3310]. Smelting [1672, 4338, 7154, 9366]. Smithson [5106]. Smithsonian [7771, 1844, 2726]. Smoke [1415, 3941, 4311, 6183, 10002, 7234, 806]. Smokeless [2509, 2725, 3082, 4850, 10749, 612, 9289]. Smokes [9371]. Smoking [1026]. Smoky [10895]. Smooth [2105]. Smooth-Bore [2105]. Smugglers [9209]. Snag [8901]. Snails [8747]. Snake [3240, 11035, 11115, 7301, 10068]. Snake-Like [10068]. Snakes [5870, 5884, 9130, 11831, 5392, 10398, 5526]. Snap [2382]. Snap-Hook [2382]. Snohomish [32]. Snoqualmie [4901]. Snow [1348, 2516, 4734, 4294, 5081, 6058, 8674, 10135, 1102, 10155, 8813, 7603, 7423]. Snow-Bound [1102]. Snowfall [1917]. Snowsheds [7192]. So-Called [5102, 137, 10398]. Soap [3332, 10140, 242]. Soap-Delivery [3332]. Soar [1221]. Soaring [11848, 6376]. Societies [1412]. Society [2233, 1577, 6846, 9896, 11660, 11401, 6124, 2773, 4162, 434, 1275, 944, 10998]. Socket [2284, 13, 639]. Soda [2147, 5254]. Soda-Water [5254]. Sodding [9448]. Sodium [9768, 11392]. Soft [4385, 10035, 8571, 9429, 2310, 642]. Soft-Coal [8571, 2310]. Soil [7396, 10255, 11663]. Solar [1704, 4969, 7007, 6978, 8026, 9075, 8473, 2578, 313, 992, 11842, 8259, 11304, 8053]. Solder [571]. Soldering [11188, 10261, 7476]. Soldier [9032]. Sole [8990]. Solid [4876, 7807, 8809, 9793, 8640, 3762]. Solution [4170, 4570, 9242, 9528, 9974, 11277, 1282]. Solutions [970]. Some [2350, 1433, 2057, 1621, 1751, 2357, 2494, 1702, 3151, 3409, 3428, 3262, 2754, 2826, 2536, 3361, 3562, 3597, 4266, 4605, 3763, 4149, 4846, 3987, 4175, 4800, 5474, 5701, 5998, 6033, 5905, 5036, 5073, 5870, 5884, 5806, 6391, 7263, 6437, 6913, 6964, 6644, 8191, 7789, 8384, 7855, 8890, 8564, 9466, 8724, 9069, 8620, 9231, 8770, 9920, 9963, 10478, 10502, 10157, 11722, 11714, 11943, 11368, 11042, 11948, 11501, 11473, 11533, 391, 489, 923, 226, 367, 389, 845, 457, 9962, 11862, 2715, 7480, 4439, 10245, 8763, 4873, 8495, 6783, 3868, 482]. Some [1166, 1422, 2356, 1989, 2475, 1634, 4050, 4318, 4199, 11001, 4898, 5647, 8981, 5367, 10798, 10420, 8229, 10133, 11609, 536, 10417, 10563, 10903, 1249, 1165, 1045, 11842, 10026, 2960, 5815, 10908, 919, 7734, 5017, 10692]. Somersault [6739]. Somersaulting [10200]. Something [8104, 7694, 9285, 3318, 3848, 989]. Sondes [7650, 635]. Soo [5310, 452]. Soot [2489, 4619]. Soot-Proof [4619]. Sorter [6944]. Sorting [3618, 6824]. Sound [4749, 6759, 8145, 9097, 9294, 10109, 7958, 11113, 3426, 4363]. Sounder [11570]. Sounding [7786, 1020, 7520]. Soundings [3247]. Soundless [9289]. Sounds [3111, 7582, 4598]. Source [9224]. Sources [9549, 10692]. South [2037, 1919, 1492, 1931, 2994, 2683, 3305, 4483, 6044, 6803, 8086, 9309, 9586, 9539, 9644, 8879, 10596, 1765, 8434, 10425, 1803, 5328, 3814, 4830, 4372, 1388, 5510, 1654, 1624, 8901]. Southern [4643, 1918, 2754, 9550, 11416, 7751, 3843, 7450, 2552, 7168, 7472]. southwest [5212, 8596, 11313, 1777, 2988]. Southwestern [8966]. Space [4090]. Spaced [562]. Spain [3476, 6079, 5784]. Span [3337, 8160, 8969, 10906]. Spanish [4168, 2392, 6560, 156]. Spar [10337]. Spare [6889]. Spark [4619, 8174, 9125, 10094, 2084, 688]. Spark-Producing [10094]. Sparking [4145]. Sparklets [3762]. Sparks [10480]. Speak [10044, 1250]. Special [4277, 5050, 9735, 11391]. Specialties [7274, 6931]. Species [9164]. Specimens [2917, 1733, 11608]. Spectral [4737]. Spectrum [2046, 846, 1124]. Speech [10857]. Speed [1842, 2435, 2093, 2130, 1750, 2701, 3398, 2931, 3081, 3444, 3638, 2958, 2750, 4097, 3974, 4289, 4433, 3865, 3780, 4383, 4259, 5303, 5088, 5177, 5323, 5108, 6090, 5607, 5783, 6221, 6378, 6195, 6290, 6813, 7002, 7064, 7303, 8504, 8294, 7354, 8081, 7629, 9035, 9117, 9064, 8735, 10470, 9799, 10460, 10243, 9931, 9775, 9826, 10019, 9685, 10330, 11844, 11451, 11240, 11032, 11629, 11365, 891, 894, 492, 677, 395, 390, 1195, 8388, 9703, 9114, 4221, 8524, 7707]. Speed-Changing [492]. Speedometer [9961]. Speeds [2071, 7337]. Speedway [9122]. Speedy [1599, 8790, 8638, 9974]. Spelterini [6897]. Spencer [5973, 5602, 4492]. Sphere [9605, 8919]. Spices [4995]. Spider [2517, 10204, 5606, 7262, 350]. Spiders [3440]. Spier [5627]. Spigot [2938]. Spike [10883, 11569, 11875, 177]. Spindles [4835]. Spinning [4835]. Spiral [2006, 5670]. Spiral-Globe [2006]. Splash [1093]. Split [3334, 10732]. Spokes [11264]. Sponge [4802, 11040]. Sponges [11609]. Sponginess [9684]. Spontaneous [3038, 5524, 8375, 10256, 10171]. Spontaneously [7501]. Spontaneously-Moving [7501]. Spoons [1023]. Sport [5304, 5082, 11367]. Sportsman [11366]. Sportsmen [9119, 7942]. Spot [6072, 6118]. Spots [3646, 5796, 5767, 7378, 7888]. Spotting [9531]. Spray [11093, 1057, 663]. Spraying [11031]. Spreader [5288]. Spree [2916]. Spring [2871, 3768, 6447, 6742, 7800, 9652, 8806, 10782, 10225, 14, 96, 11545, 1025, 562, 9896]. Springfield [5313, 643]. Springless [2382]. Springs [6572, 898, 964, 734, 801, 896, 993, 8409]. Sprinkler [7255, 7523, 7987, 9651]. Sprinkling [8193]. Spur [6695]. Spy [2538]. Spy-Glass [2538]. Squadron [7346, 9885]. Square [1853, 2222, 2176, 2993, 9431, 10517, 10397, 10647, 9919, 1074, 897, 525, 256, 11371, 9704, 10263]. Square-Rigged [9704]. St [1390, 5355, 5624, 6941, 7039, 6984, 11233, 614, 10341, 6894]. St. [1460, 2338, 2331, 1854, 2217, 3655, 3488, 2888, 4105, 4776, 6075, 6164, 5698, 5426, 7044, 7128, 6259, 7065, 7252, 6912, 7087, 7000, 6855, 6878, 7043, 7325, 7139, 7336, 7001, 7345, 8168, 9067, 10335, 10441, 4468, 4709]. Stab [4715]. Stability [6374]. Stable [9744]. Stack [3310, 4462]. Stacker [3507, 7829]. Stacks [9884]. Stadium [11376]. Staff [5934]. Stage [2494, 4276, 6690, 6782, 7793, 9]. Stairway [8771]. Stairways [2220]. Stakes [2238]. Stall [4396]. Stamp [9655, 10835]. Stamp-Affixing [10835]. Stamps [11643, 10981, 369, 9484]. Stanchion [9825]. Stand [3919, 10784, 11823, 11710]. Standard [4691, 5090, 7495, 773, 940, 8407]. Standardized [8123]. Standardizing [473]. Standards [11650, 8570, 8299]. Standing [3493]. Standpoint [4137, 4138, 4214, 4523, 89]. Stands [4763, 11307]. Stanhope [4284, 6646]. Stanley [5161, 6460, 10196]. Star [2544, 5046, 7697, 9332, 920, 6735, 9728, 8169, 1124, 3438, 2731, 3642, 5068, 7211, 10005, 11043, 11631]. Starfish [6671]. Stars [8895, 11223, 2983, 8592, 10199, 11327, 30, 3509, 7734]. Start [7005, 8372, 7657]. Starting [3137, 3887, 7111, 11616, 444]. Starting-Device [3137]. Stassano [1824]. State [5868, 6776, 11496, 9619, 2686, 1867, 2063, 5738, 6775, 9042, 1159, 7085]. States [2278, 1962, 2149, 3270, 2896, 3160, 3527, 2593, 2844, 4560, 4500, 4160, 5104, 4936, 4957, 5745, 7212, 6396, 7263, 6964, 7001, 9351, 8548, 9160, 10757, 11318, 894, 862, 238, 10591, 7886, 7088, 7755, 11192, 7858, 4371, 7299, 9529, 9165, 1729, 10860, 4903, 10674, 2352, 8299, 10358, 11433, 7831, 10999, 1845, 1409, 2143, 2459, 2001, 1762, 2280, 1682, 1478, 3196, 3678, 3363, 4728, 4687, 3977, 5640, 5544, 5558, 6809, 6259, 6775, 6261, 7106, 7080, 8493, 8202, 9003, 10436, 10790, 10418, 11223, 10973, 11923, 11000, 11945, 657, 533, 220, 365, 1016, 32]. States [2348, 596, 3338, 3013, 2310, 4753, 6298]. Static [5480, 6051, 2084, 6097]. Station [1585, 2967, 3064, 3500, 4845, 4832, 3959, 3825, 4281, 6028, 6832, 6936, 8028, 8675, 8875, 8856, 10149, 10339, 11196, 11302, 1301, 357, 87, 8815, 7628, 2255, 1919, 1975, 3764, 4347, 5412, 5224, 6915, 7769, 8721, 9506, 8660, 10541, 10276, 11739, 11738, 6006, 7162, 3637]. Stationary [3952, 8675, 11092]. Stations [1383, 2881, 3973, 7782, 10959]. Statistics [5951, 9141, 10195, 10899, 11368]. Statuary [7423]. Statue [8498, 2381, 4026]. Statues [9400, 943]. Status [3437, 6116, 7677, 9284, 10516]. Statutes [4936, 4957]. Ste [2774, 4529, 4100].

Steady [10665]. Steam [3664, 1389, 2455, 1748, 2144, 1802, 1391, 2375, 1368, 2287, 1959, 3418, 2771, 2890, 3974, 4433, 4282, 3832, 3754, 4166, 3799, 4033, 4283, 4256, 4280, 4281, 4523, 4392, 6008, 6073, 5971, 5269, 5179, 5665, 5824, 5068, 5087, 5292, 7205, 6655, 6500, 6290, 6656, 7323, 6774, 6244, 6886, 7633, 8081, 7608, 7725, 7799, 8406, 8059, 8509, 7604, 7976, 9048, 9544, 8676, 8782, 9513, 8917, 8823, 8584, 10196, 9720, 9959, 9711, 10214, 10257, 9958, 10768, 9757, 11844, 10886, 10937, 11165, 358, 588, 1002, 1088, 1087, 6311, 6076, 5521, 7825]. Steam [2221, 2235, 3983, 4752, 5135, 9746]. Steam-Boiler [7604]. Steam-Driven [6500, 11844]. Steam-Turbine [1088]. Steamboat [9665, 10792, 755]. Steamboats [756, 758]. Steamer [1444, 3551, 5929, 5831, 5978, 155, 32, 5362, 7360]. Steamers [3081, 2685, 5571, 6353, 6831, 6434, 7700, 8424, 7899, 11807, 8124, 2235]. Steaming [10170]. Steamship [2390, 2299, 2170, 2933, 2710, 2566, 3903, 3910, 4406, 4047, 5303, 5415, 6807, 8921, 9751, 10543, 864, 1019, 433, 287, 186, 5209, 5476, 8144, 6379, 4753]. Steamships [2167, 2198, 1912, 3997, 4524, 7975, 425, 512, 145]. Stearns [4281, 6643]. Steeds [8172]. Steel [2181, 2329, 1605, 1679, 2199, 1436, 2369, 2456, 3269, 3656, 2621, 2915, 2724, 3292, 3823, 4829, 3889, 3981, 4500, 3972, 4230, 5104, 5948, 5950, 6091, 5953, 5949, 5498, 5527, 4990, 5803, 5951, 5946, 5954, 6680, 6309, 7131, 7903, 7929, 7976, 9414, 8611, 9205, 8912, 9667, 10275, 10058, 9954, 9723, 10254, 10492, 10191, 11079, 11342, 11389, 514, 1076, 288, 21, 79, 1160, 8831, 9356, 11704, 9686, 5384, 5827, 9684, 364, 6931]. Steel-Concrete [5803, 10254]. Steel-Making [5498]. Steel-Plate [3823, 9356]. Steels [475]. Steepest [3013]. Steering [1410, 3063, 3904, 5240, 6661, 9952, 241, 1305, 1553]. Stein [3422]. Steltje [5808]. Step [2656, 5705, 5025, 7083, 8736, 11665, 11095]. Step-Ladder [11665]. Step-Ladders [7083]. Stephenson [3614]. Stereogram [7812]. Stereophotography [10922]. Stereopticon [3943]. Stereoscope [10010, 11619]. Stereoscopic [169, 796, 11014]. Stereotyping [5852]. Sterilization [8877, 7988]. Sterilized [10123]. Sterilizer [1604]. Stern [2685, 8300]. Stern-Wheel [2685]. Stevens [5455, 4442, 6646, 11168, 11077, 760]. Still [5920, 4970, 36, 5906]. Stilts [7191]. Stimulant [7693]. Stimulate [6973]. Stitching [5860]. Stock [4812, 9512, 11013]. Stoddard [10481, 9944]. Stoker [1891, 4850]. Stokers [8447]. Stoking [2001]. Stolze [7809]. Stone [7501, 8657, 11604, 286, 1056, 642, 8532]. Stonehenge [1817]. Stones [9563, 9801]. Stop [4927, 7231, 9170, 10721, 178]. Stopcock [566]. Stopper [3571, 9847]. Stopping [1848]. Storage [1513, 1650, 2837, 2895, 5015, 6093, 5475, 5760, 7190, 6391, 6246, 8279, 9012, 9545, 10305, 10565, 11112, 11492, 1077, 383, 9071, 4491, 11688]. Storage-Battery [8279]. Store [1659, 4998]. Store-Room [1659]. Stores [11523]. Stories [250]. Storing [7305]. Storm [5186]. Storms [2679, 898, 11067]. Story [7385, 1225, 4008, 9754, 10734, 9353]. Stove [1342, 8194, 5702, 808]. Stoves [3447, 8623, 9991, 11358]. Straightening [3660, 572]. Strain [7749, 11803]. Strainer [6402, 497]. Strait [6089]. Stralau [2916]. Stranding [10523]. Strange [6805, 831, 7118, 7304, 10692]. Strasburg [341]. Strategical [6550]. Stratification [8125]. Straw [6504]. Stream [1278]. Streams [165]. Street [11429, 2285, 2410, 2458, 2942, 2967, 3064, 3389, 3031, 3812, 3936, 3906, 4522, 5018, 7255, 6332, 8262, 7469, 7523, 8193, 9651, 9293, 9337, 10154, 9822, 11410, 8428, 5439, 2601, 3012]. Street-Sprinkling [8193]. Streetcar [2642]. Streets [5691, 5127, 7131, 10087]. Strength [1930, 6314, 6683, 7648, 7851, 8209, 7578, 8888, 10573, 610, 1066]. Strengthening [2106]. Stress [57]. Stretcher [1916, 5438, 5680, 9780, 1230]. Stretching [10813]. Strike [4499]. Strikes [5499]. Striking [8060, 458, 10796, 6879, 7423]. String [5984]. Stringed [947]. Strip [11278]. Striped [8172]. Striper [2239]. Stripped [10255]. Strips [6543]. Stroke [858]. Strong [10402]. Stronger [6413]. Struck [5287]. Structural [3656, 5952, 5475, 10275]. Structure [5968, 9167, 10036, 453, 10382, 11608]. Structures [2907, 6821, 11408, 11803]. Struggle [3541]. Strung [11211]. Stud [2602, 7525]. Student [5024]. Studied [5718, 8965]. Studies [3570, 7375, 8891]. Study [3356, 4201, 4969, 6985, 6415, 6710, 7230, 9245, 10756, 7402, 7167, 9728, 6054, 11520]. Studying [431]. Stump [9987]. Stupendous [9180, 10394, 10212, 304]. Style [3358]. Suakim [6461]. Sub [5747, 1300, 170, 872]. Sub-Caliber [872]. Sub-River [1300]. Sub-Surface [170]. Sub-Target [5747]. Subjects [10857]. Submarine [2021, 1600, 2912, 3490, 2734, 4572, 3976, 4219, 4074, 4640, 3927, 4345, 5901, 6011, 4976, 6287, 6732, 6225, 6799, 6216, 6496, 6364, 6759, 7022, 8020, 8469, 7905, 7720, 8526, 9097, 8677, 10593, 10490, 10870, 10923, 11040, 11473, 863, 1054, 240, 1725, 4625, 508, 4142, 3017]. Submarines [3244, 3527, 2702, 7729, 9698, 10085, 10706, 10945, 11045, 10849, 862, 238, 4000, 9981, 11801]. Submerged [3288, 2710, 7156, 10712, 10354]. Subscriber [1459]. Subscribers [4731, 6002, 5998, 7159, 10648]. Subsidence [10223, 31]. Subsidize [3716, 10191]. Subsidy [4188, 9090, 2772]. Subsistence [8422]. Subsoil [11663]. Substitute [3134, 665, 1237, 396, 11224]. Substitutes [4511, 5734, 11614]. Substructure [6538]. Subterranean [2572, 4213]. Suburban [4868, 3909, 5162, 7395, 11187, 11428, 11915, 1135]. Subway [2815, 3325, 3039, 3361, 4220, 4414, 3973, 5166, 5412, 5594, 5167, 5540, 5156, 5431, 6536, 6681, 6258, 6971, 6993, 6780, 6992, 7105, 7056, 7953, 7782, 7491, 7477, 8312, 7583, 8377, 8184, 7534, 7719, 8977, 9571, 8717, 8912, 10149, 10791, 9879, 9699, 9975, 11145, 10826, 10959, 11740, 11388, 11427, 10986, 11625, 1080, 629, 144, 80, 8357, 4404, 4708, 5792, 6830, 7341, 6981, 6342, 7962]. Subways [7384, 11511, 11735, 8478, 8320]. Success [6712, 8349, 9304, 9321, 9240, 11390]. Successes [11473]. Successful [4486, 4164, 6013, 8188, 9479, 9325, 10313, 9929, 10069, 11829, 10876, 11895, 10972, 11636, 850, 1218, 261, 48, 3757, 6575, 11655]. Successor [11724]. Suchet [2155]. Sucking [4810]. Sucking-Fishes [4810]. Suction [8685, 922, 8052, 10008]. Suction-Gas-Propelled [8685]. Sudd [704]. Sue [11898]. Suevic [10529]. Suez [2636, 3787, 8249, 11846, 4756]. Suffer [7204]. Suffering [5156]. Suffren [1393]. Sugar [1339, 2150, 2461, 2283, 2641, 5341, 791, 11433, 9331]. Sugar-Beet [791]. Suggested [9225]. Suggestion [2458, 2700, 11294, 10946]. Suggestions [493, 724, 11476]. Suit [2409, 2242, 10649, 1291]. Sulky [2154, 1728]. Sulky-Rakes [1728]. Sulphate [344]. Sulphite [4946]. Sulphur [1722, 9581, 10028, 9583]. Sulphurous [11748]. Sultan [3796]. Sulu [10258]. Sumatra [2158, 3427, 11323]. Sumerian [8498]. Summary [7698, 8378, 9047, 10521, 10852, 11060, 11430, 11633]. Summer [4321, 7036, 7561, 9368]. Summerland [4142]. Summit [1843, 9676]. Sun [5796, 5767, 1669, 3646, 5814, 7378, 7888, 9305, 11156, 11011, 705, 8076, 11174, 10007, 10654, 11086, 11702, 11682, 11067, 7326, 9485, 11520]. Sunfish [6424]. Sunk [10256]. Sunken [1512, 8624, 9751]. Sunset [1818]. Sunstroke [10557]. Super [2492]. Super-Heated [2492]. Superheat [9065, 10213]. Superheated [2455, 6774, 8823]. Superheater [6509]. Superintendent [8527]. Superior [2714, 9671, 5213, 1072]. Superiority [112]. Superposed [1467, 3054, 3669]. Supersensitiveness [754]. Superstitions [6600, 7649, 367]. Supervision [3608]. Supplanting [4340]. Supplement [4460, 4971, 5435, 5541, 5567, 5765, 5787, 5830, 5976, 6003, 6120, 5393, 5865, 6357, 6393, 6711, 6849, 7443, 7696, 7724, 7787, 7879, 7904, 7980, 8023, 8045, 8078, 8166, 8376, 9328, 9349, 9394, 9482, 9725, 9748, 9794, 10128, 10194, 10419, 10519, 10645, 10711, 10755, 10772, 10853, 10874, 11918, 10919, 10966, 11061, 11081, 11109, 11149, 11171, 11190, 11235, 11258, 11299, 11322, 11350, 11393, 11431, 11455, 11478, 11634, 11822, 11881, 11901]. Supplemental [2873]. Supplies [9187]. Supply [1697, 5014, 2324, 1772, 1382, 1454, 3205, 4834, 5690, 5378, 5103, 6709, 6436, 7647, 8118, 7901, 8614, 8610, 8618, 10351, 10212, 10418, 10902, 11728, 164, 881, 1917, 2939, 435, 1392, 2044, 5577, 5284, 6373, 7747, 7393, 7961, 9042, 10111, 10293, 11029, 11729, 10831, 10827, 980]. Supplying [8967]. Support [1418, 5551, 6589, 7453, 8992, 9192, 1236, 567, 1262]. Supported [4420]. Supposed [4189]. Suppressing [496]. Supremacy [1381, 5946, 9875]. Surface [3097, 6551, 10339, 10255, 11802, 11663, 170, 6206, 10284]. Surfaces [7671, 10362]. Surgery [9427, 52, 2108]. Surgical [1604, 7913]. Surrey [6667]. Surrounding [6250]. Survey [8072, 615, 11328, 1406, 9323]. Surveying [727, 662]. Surveyor [893]. Susa [1815, 1370]. Susceptibility [7730]. Suspended [1538, 1528, 8960, 4167]. Suspension [3179, 3593, 7998, 11732, 11797, 11731, 491, 6940, 11940]. Suspensions [8227]. Susquehanna [11519, 9795, 10528, 11457]. Sustained [9836]. Sustaining [10505]. Swage [5308, 11202]. Swamp [10555, 1494]. Swanage [1056]. Sweden [7249]. Swedish [10392]. Sweep [1510]. Sweeper [9651]. Swift [11415]. Swimmer [9673]. Swimming [9990, 438, 7446]. Swing [11309, 11834]. Swiss [9096, 2237, 4804, 9211]. Switch [2580, 4715, 8110, 8262, 11307]. Switchboard [11094]. Switches [8397, 8743].

Switching [1919, 5549]. Switzerland [4081, 8494, 7302, 11564, 117]. Sword [6249]. Swordmaking [6293]. Sylviculture [2636]. Symbolism [4645]. Symington [759]. Sympathy [8733]. Synchronism [3461]. Synchronized [1956]. Synchronizing [10480]. Synthetic [3253]. Syntonic [4877, 4900]. Syracuse [11376]. Syria [790]. Syringe [2177]. Sysonby [10756]. System [1644, 1550, 1413, 1456, 1657, 1919, 1515, 1536, 2322, 2391, 2701, 3386, 2894, 3528, 2865, 2893, 2610, 2859, 2915, 2546, 2554, 3665, 4597, 3726, 4851, 4573, 3936, 4344, 3704, 4701, 4834, 4510, 4687, 4856, 3977, 5825, 5785, 5383, 5228, 5314, 4970, 6258, 6266, 6435, 6705, 7249, 7080, 7477, 8314, 7920, 8018, 8137, 9303, 9571, 8978, 9818, 9799, 10522, 10064, 10394, 10078, 11458, 11805, 11753, 11448, 11351, 11320, 10880, 1078, 774, 222, 834, 409, 304, 611, 599, 550, 10113, 3722, 4997, 4916, 9488, 7983, 10399, 2939, 5112, 6509, 8089]. System [4961, 7835, 432, 11138, 3465, 10674, 11090, 4615, 5793, 4998, 11806, 3397, 5565, 10541, 10897, 11324, 5310]. Systems [2459, 1696, 3633, 2838, 2796, 3678, 4622, 4436, 6332, 6455, 8765, 11796].

T [1920, 3085]. T. [2935]. Table [2151, 5191, 8436, 10741, 10446, 10445, 10653, 11498, 1142]. Table-Top [11498]. Tablet [9731]. Tabouret [1231]. Tachometric [7908]. Tack [7891]. Tacker [5680]. Tacks [6977]. Tactical [11651]. Tailed [4779]. Tailings [5886]. Tails [5549]. Take [10872, 11363]. Taken [5571, 6875]. Taking [3247, 3863]. Talamancans [5885]. Talk [11876, 8, 843]. Talking [3253, 9530, 10994, 9900]. Tall [1699, 1563, 2781, 4990, 9167, 9723, 11845, 11734]. Tallest [8380]. Tamalpais [3570]. Tame [836]. Tameness [3944, 9147]. Tammany [6617]. Tamper [5889]. Tandem [3261, 3331, 10720]. Tanganyika [2171]. Tangye [9307]. Tank [3594, 3241, 3005, 5040, 7849, 10900, 387, 462, 11175, 3530]. Tanks [1986, 7382, 122, 9503]. Tantalum [321]. Tapered [735]. Tapestry [7585, 11303, 10882]. Tar [11031]. Tarantula [11059]. Target [3174, 4124, 4160, 5747, 10705, 10561, 10575, 11375, 10265, 10811]. Tarring [9186]. Tasmanian [7773]. Tattooing [5629]. Tavritchesky [4350]. Taxation [8870]. Taxicab [11218]. Taxidermist [9607, 11286]. Taxidermy [6579, 8641]. Taximeter [10466, 532]. Taxometer [6934]. Tea [958, 7066, 7502, 10133, 8302]. Teaching [2643, 10044, 852]. Teams [7407]. Tear [6]. Technical [6484, 8980, 11697, 45, 3588, 9819]. Technically [9743]. Technology [6959]. Teeth [8661]. Tehuantepec [4787, 8822]. Telautograph [2349, 8208, 11681]. Tele [1673]. Tele-Photo [1673]. Telechirograph [5570]. Telecryptograph [6853]. Telegraph [1559, 2466, 2597, 2892, 2572, 2732, 3289, 4867, 4870, 4604, 4845, 4832, 6099, 5808, 5570, 6380, 6936, 7784, 9666, 8849, 9076, 9216, 9169, 8933, 10541, 11570, 1301, 3185, 3722, 4324, 4916, 9883, 6006, 8815, 10219, 5525, 10833, 3441]. Telegraphing [6534, 6853]. Telegraphone [2000, 2008, 5696, 5139]. Telegraphs [3153]. Telegraphy [2096, 1931, 3350, 3550, 2613, 2643, 2823, 3197, 2546, 2554, 3063, 3080, 3520, 2863, 3930, 3714, 4485, 4851, 3704, 4622, 4510, 4149, 4164, 3736, 3883, 5102, 5675, 5744, 4970, 4964, 5426, 6266, 7279, 6759, 7767, 8042, 8019, 8102, 9582, 8704, 8914, 8596, 9226, 9752, 10516, 10376, 11154, 3525, 4877, 4900, 4439, 4997, 7860, 9488, 9732, 5112, 11138, 1553, 4200]. Telepherage [3665]. Telephone [1413, 2307, 2332, 3386, 2894, 2545, 3397, 3041, 3893, 4732, 3717, 4466, 4532, 4933, 5762, 5678, 8991, 8855, 9247, 10195, 397, 4755, 10082, 6853]. Telephones [10138, 1308, 8428]. Telephonic [5259, 6962]. Telephoning [6534]. Telephony [2980, 5383, 5041, 5314, 9226, 11458, 4004, 10129, 432, 1658, 10358]. Telephot [5388]. Telephoto [893]. Telephotography [10627, 1282]. Telescope [3744, 5236, 5442, 6960, 7771, 9525, 10381, 870, 4781, 11151, 111]. Telescopes [2013, 9229, 11564]. Telescopic [1305]. Telescoping [4813, 8599]. Telethermometer [5623]. Telharmonium [9705]. Tellurian [3701]. Telpherage [9818]. Teltow [9271]. Temperance [8140]. Temperature [4805, 4213, 4412, 6681, 8937, 11088]. Temperatures [3143, 5543, 7661, 6054]. Tempered [8923]. Tempering [8457, 7394]. Temple [8825, 6535, 4709, 2003, 8382]. Temples [4974]. Templet [685]. Temporary [3181, 3755, 9506, 734]. Ten [4343]. Ten-Wheeled [4343]. Tenacity [8192]. Tender [6980]. Tendrils [3457]. Tennessee [4738, 7155, 8593]. Tennis [6539, 15]. Tension [6515, 8058, 9845, 9923, 11946, 6122, 5261, 480, 5235]. Tent [6518, 7568, 11662, 9129]. Tents [1058]. Teredo [3411]. Terminal [1919, 3500, 4570, 5224, 8856, 10149, 10154, 9903, 11739, 11302, 11738, 218, 3302, 4170, 8028, 9272, 11344, 1114, 3637]. Terminals [11737, 1158]. Termination [5791, 707]. Terms [6869]. Terrestrial [1223]. Terrific [918]. Territory [11278]. Tesla [3008, 3080]. Test [3442, 2704, 4097, 4574, 3933, 4486, 4659, 4525, 5970, 6000, 5457, 6358, 6225, 6848, 6782, 8258, 8271, 8189, 8759, 9603, 10193, 9698, 9847, 10213, 10061, 11144, 11933, 10972, 11636, 11819, 62, 136, 7983, 518, 3721, 11655, 4488, 4507, 5737, 5791, 11135]. Tested [2035, 8123, 11284]. Tester [1557, 10476]. Testing [2917, 1676, 3937, 7277, 7637, 8520, 8401, 10395, 9953, 11196, 538, 927, 941, 9080, 9286, 1274, 225, 8004, 4580, 11941, 11433, 11352, 1051, 6959]. Tests [3657, 3420, 2825, 3435, 3810, 4433, 5178, 5901, 5383, 6049, 5158, 6433, 6603, 7633, 10770, 10039, 10085, 10067, 10625, 11483, 11639, 11185, 11459, 11513, 11131, 11937, 305, 1136, 11167]. Tetrahedral [8704, 62, 10382, 11807]. Tetrahedral-Cell [62]. Texan [6670]. Texas [1518, 4456, 1578]. Textile [4573, 7556, 310, 8402, 1624, 4505]. Th [1908]. Thames [6333]. Thawing [6246, 7881, 1259]. the-Century [2390]. Theater [6414, 8426, 9901, 827, 6419, 10303]. Theatre [6620]. Their [6573, 6927, 8162, 8662, 9140, 10525, 11685, 8452, 7813, 1166, 8456, 166, 9098, 4549, 3716, 9793, 769, 464, 8337, 3984, 6181, 6671, 481, 5017]. Them [10338, 11716]. Theodore [5848]. Theories [9143, 11533, 391]. Theory [5923, 5868, 8183, 9814, 9790, 11078, 11820, 1121]. Therapeutics [10675, 10780]. There [3374, 8360]. Theresia [1417]. Thermo [3436, 8833]. Thermo-Electric [3436, 8833]. Thermometer [1372, 7907, 7425, 10758, 2355, 556]. Thermometers [1933, 4805, 8123]. Thermometric [7661]. Thermostatic [806]. Thetford [1442]. Thickness [9579]. Thill [10977]. Things [1721]. Third [1348, 2313, 3528, 2577, 5900, 5099, 5460, 6406, 6743, 8674, 10666, 9977, 9270, 1168]. Third-Rail [1348, 2313, 6406, 6743, 10666]. Thirteenth [2773]. Thirty [3330, 7183, 7068, 4492]. Thirty-Eight [7183]. Thirty-Mile [4492]. Thomas [3141, 4736, 6645, 8879, 8810, 9925, 636, 11]. Thoraco [2290]. Thoraco-Xiphopagous [2290]. Thorite [2704]. Thorium [2753]. Those [92]. Thothmes [5049]. Thoune [4081]. Thousand [4190, 8432, 7836, 7854, 8689, 9049, 8691, 7600]. Thousands [7588]. Thread [2178, 322]. Thread-Cutting [2178]. Threader [1237]. Threads [1414, 3782]. Threaten [9025]. Threatened [1336]. Threatens [9489]. Three [1443, 1411, 2914, 3593, 2593, 3446, 2785, 2943, 3095, 3797, 4484, 4119, 4691, 5070, 5864, 6394, 7900, 10580, 10581, 10585, 9792, 11281, 762, 2600, 5626, 10950, 11418, 2075, 3766, 1471]. Three-Color [5864]. Three-Masted [5070]. Three-Phase [4691, 6394, 2075, 3766]. Three-Wheel [3797]. Three-Wire [1443]. Throat [10126]. Thrower [8262]. Throwing [5606]. Thumb [751]. Thunder [4968]. Thunderstorm [5481]. Thunderstorms [8107, 1277]. Thurston [5805]. Thyridopteryx [2830]. Tiberius [4064]. Tibet [10612]. Tidal [9, 7183]. Tide [4034, 5233]. Tides [7417, 3847, 3913, 10131]. Tie [1318, 2200, 2872, 5215, 6914, 6586, 9451, 9824, 11263, 11763, 11468, 949]. Tie-Bar [9451]. Tie-Plate [2200]. Ties [5378, 8273, 11389]. Tight [69]. Tightener [5339]. Tile [9745, 9672]. Till [6814]. Tilting [6826, 7110]. Timber [1930, 4314, 5034, 7851, 7578, 11883, 881, 7773]. Time [1623, 1789, 7836, 7495, 8374, 9796, 7627, 7983, 6440, 773, 8167]. Time-Controlled [9796]. Timely [3173, 8612]. Timer [1001]. Timers [11378]. Times [5013, 7429, 9033, 9251, 9267]. Timing [6316, 6359, 7808]. Tin [4440, 4530, 1232]. Tinder [10920]. Tinfoil [2861, 396]. Tinol [10081]. Tipping [1959]. Tipping-Cart [1959]. Tipple [3964]. Tire [4876, 4172, 5085, 6658, 6889, 7685, 8946, 8574, 11934, 11502]. Tire-Repairing [4172]. Tireless [11220]. Tires [4300, 6977, 6888, 6574, 7807, 7447, 10475, 10816, 490, 481]. Tissot [11066]. Tissue [2571]. Title [4569]. Titles [4452]. To-Day [8319, 8048, 10832, 11460, 11886, 11565, 288]. To-morrow [10173]. Tobacco [5889, 7964, 9630, 9150, 10976, 2802, 4348, 10222]. Tobogganing [117]. Togo [7327, 7705]. told [7385]. Toledo [6642, 9937]. Tomato [7068, 5328]. Tomb [5049]. Tombs [1710, 3242, 4533, 7427]. Ton [2157, 3248, 3088, 2848, 4107, 5909, 5379, 5517, 5837, 7737, 8082, 8940, 9501, 9842, 10597, 147, 93, 922, 941, 9567, 4101]. Tone [1351, 1963]. Tongs [11331, 11199, 11160]. Tongue [4782]. Tongue-Traction [4782]. Toning [1386, 10277]. Tonnage [10590]. Tonneau [5075, 6649]. too [9594]. Tool [1580, 2456, 1895, 2995, 2755, 4002, 4538, 4172, 4536, 6059, 5809, 6660, 7525, 8684, 8905, 9821, 10928, 11334, 634, 256, 811, 10260]. Tool-Grinding [1580, 4002]. Toolmakers [8162]. Tools [1480, 8457, 224, 320, 10245, 4315, 4804]. Tooth [4143, 4538, 6272, 1024, 6622]. Toothed [1471]. Top [3374, 5289, 11498, 2440, 69]. Topophone [2619]. Topping [4585]. Tops [6826, 11659]. Tores [10610]. Torpedo [2396, 1493, 2023, 1769, 2244, 1346, 2912, 3490, 3491, 3492, 4212, 4233, 6011, 5386, 5794, 5667, 6068, 6390, 6205, 7022, 6444, 8272, 8189, 9203, 8548, 9118, 10576, 10066, 10379, 10589, 11411, 11419, 11217, 11651, 775, 170, 87, 63, 1274, 2807]. Torpedo-Boat [2396, 3492, 4233, 5794, 10066, 10589, 1274]. Torpedo-Boats [3491]. Torpedo-Launching [87]. Torpedoes [3063, 4345, 9752, 11032, 1553]. Total [1669, 4003, 5700, 9075, 10654, 11842, 9485]. Touch [9813, 10045]. Touch-Me-Nots [10045]. Touches [9647]. Toulon [7679]. Tour [7298, 9044, 10713, 11841, 10467, 9599, 957]. Touring [4295, 6642, 6664, 6918, 6643, 6644, 6645, 8482, 7798, 7803, 7790, 7791, 8335, 8401, 9540, 8796, 8798, 8800, 8793, 8801, 8797, 8799, 9959, 9964, 9966, 9947, 4278]. Tourists [6366]. Tourmaline [11719]. Tournament [944]. Tow [8121]. Tower [4486, 10382, 3720, 7324, 6004, 6708, 883, 458, 4963, 6245]. Towers [1561, 2551, 5703, 7954, 9765, 10206, 11104, 8683]. Towing [3362, 3241, 3005, 7849, 11175]. Town [2761, 1167, 4415, 10008]. Towns [9187]. Toxic [958]. Toy [1464, 5052, 4948, 5500, 5342, 8340, 7991, 8061, 8750, 10070, 10178, 11064]. Toys [4668, 10633, 11501, 11780, 1096]. Tracer [8688]. Trachodont [10950]. Track [1564, 2843, 3641, 5768, 5736, 6095, 6694, 6480, 8119, 8384, 9377, 10384, 10281, 10000, 9699, 10507, 11514, 89, 5038, 9650, 10687, 10771]. Track-Laying [10384]. Track-Sanding [9377]. Track-Walker [89]. Tracks [5664, 6182, 8898, 8542, 949, 2132]. Trackway [7131]. Traction [10812, 1726, 2487, 2492, 2309, 3528, 2796, 4597, 4868, 4189, 4119, 4782, 6157, 5759, 5969, 5619, 5311, 6236, 6433, 7104, 7164, 8781, 10447, 9767, 9894, 10498, 10937, 252, 1534, 8663, 6829, 10920]. Traction-Operated [10447]. Tractor [4052, 4125, 6651, 11084]. Tractor-Barges [4052]. Tracts [7546]. Trade [2034, 2275, 2300, 2280, 1478, 3176, 2885, 3006, 2657, 3540, 2635, 3653, 4231, 4775, 4161, 4678, 5640, 5231, 6775, 7459, 7826, 8074, 8738, 11108, 726, 9399, 7598, 5237, 1654, 3240, 7932]. Trade-Mark [7826, 8738, 7598, 7932]. Trade-Marks [5231]. Trades [5869, 4168]. Traffic [2255, 2025, 2862, 5162, 5205, 6847, 7469, 7395, 9180, 10477, 10087, 10825, 11389, 1158, 1135, 504, 10220]. Tragedy [4139]. Tragic [4506]. Trailing [8397]. Train [1906, 1726, 2181, 1676, 2750, 2991, 4555, 4813, 3713, 4733, 4420, 3863, 5434, 5783, 8852, 9976, 10554, 10264, 11177, 11329, 6958, 9804, 1388]. Trained [7282, 10860]. Training [2642, 6533, 9101, 11541, 11000, 7886, 11191, 8527, 4493]. Trains [1668, 1572, 3034, 2752, 3057, 4159, 4891, 4626, 4433, 3699, 5675, 6993, 7718, 7955, 7645, 10064, 11232, 629, 357, 6852, 10265, 722, 11462, 11039]. Tramp [7899]. Tramway [1490, 5565, 11259, 7278]. Tramways [2148, 2304, 6514, 8500]. Trans [2821, 5041, 10482, 1504, 6337, 1420]. Trans-Oceanic [5041]. Trans-Pacific [2821]. Trans-Siberian [1504, 6337, 1420]. Transaction [1599]. Transatlantic [3285, 2940, 3910, 4679, 4680, 5107, 5782, 6088, 6847, 8933, 10516, 10376, 10330, 11896, 11252, 11129, 10868, 433, 697, 5977]. Transcontinental [6851]. Transferable [5143]. Transformation [6578]. Transformed [3996]. Transformer [5604, 7881, 6555]. Transformers [8110, 4128]. Transit [1408, 2070, 1826, 1620, 2483, 2241, 2655, 3078, 3335, 2815, 3463, 3325, 2883, 2964, 2826, 3361, 3791, 4220, 4826, 4708, 4614, 4414, 3973, 5247, 5713, 4942, 5412, 5792, 5666, 4994, 6536, 6240, 6971, 6981, 6455, 6780, 8847, 8867, 9799, 9697, 11388, 984, 10442, 10060, 10559, 4317]. Translation [1670]. Transmission [2488, 7721, 3055, 4186, 3865, 5077, 6659, 7540, 8795, 8902, 8652, 9967, 10080, 10460, 9933, 9940, 9943, 9935, 9931, 9938, 9937, 9705, 9934, 9932, 10462, 9942, 11159, 413, 8591, 5235]. Transmissions [10459, 9930]. Transmitter [3377, 5456, 9076, 7448]. Transmitting [2959, 8239, 11266, 11154]. Transoceanic [3520, 3696]. Transplanting [1895, 5437]. Transport [2711, 2793, 2598, 9004, 10996, 11675]. Transportable [6936]. Transportation [1596, 1787, 1349, 3665, 2752, 4700, 5032, 6808, 10, 7284, 7403, 1468, 4689, 4681, 4238, 5748, 4918, 7813, 11801, 7835, 4830, 9528]. Transported [7772]. Transporter [2030, 6479]. Transporting [4707, 5507, 11356]. Transverse [1930]. Trap [9171, 9711, 9757, 10930, 176, 1002, 6158]. Trap-Lantern [6158]. Trapdoor [7262]. Trappes [2077]. Trapping [5913]. Traps [4534, 5269]. Travel [2093, 1797, 4843, 5303, 7929, 7976, 9064, 11252]. Travelers [6366]. Traveling [2220, 4107, 9008, 11243, 1141, 4804]. Travels [3422, 11775, 4372]. Tray [8507]. Tread [10498]. Treasures [3262, 5510]. Treated [7578]. Treating [3468, 121]. Treatment [4364, 2179, 5934, 9314]. Treaty [3902, 5012, 5469, 1039, 6628]. Tree [1854, 3597, 4759, 3919, 6458, 7191, 9078, 9313, 90, 3359, 1250, 9489, 11712, 11711, 11709, 11713, 11710]. Tree-Stand [3919]. Trees [1336, 2211, 1631, 3468, 4875, 5437, 7186, 9193, 1046, 727, 5214, 7860, 4007, 2552, 4199, 4196]. Trench [9111, 4106]. Trend [10234]. Trent [9149]. Treptow [2916]. Trestle [5128]. Tri [2063]. Tri-State [2063]. Trial [1456, 1762, 2094, 2118, 2614, 5739, 6607, 7303, 7042, 7106, 8002, 8735, 10876, 11641]. Trials [4097, 4640, 4383, 5071, 8355, 9416, 9685, 10768, 10896, 11637, 1012, 967, 5110, 8075]. Triangle [3761]. Triangles [194]. Tribes [9550, 4008]. Tribute [11534, 11770, 303, 45]. Trichromatic [6974]. Trick [5678, 827, 8938, 11719]. Tricks [10502, 11501, 457, 242]. Tricycle [3141, 6641]. Trillot [1435]. Trimmer [1180]. Trimming [1627]. Trinidad [3640, 2576]. Trip [3317, 3568, 4502, 4488, 4680, 4809, 6607, 6851, 7042, 8002, 8572, 10159, 1276, 776, 4492]. Triple [3120, 6645, 8404, 8905, 9118, 11802]. Triple-Cylinder [6645]. Triple-Expansion [8404]. Triple-Jawed [3120]. Triple-Screw [9118]. Triple-Surface [11802]. Tripping [5546]. Triumph [2547, 5592]. Trolley [1951, 1447, 1557, 2799, 3138, 4729, 3842, 5664, 8898, 9632, 10021, 11021, 6852, 418, 4917, 3942]. Trolley-Repair [418]. Trolleys [6870]. Trophies [5624, 10474, 10713]. Trophy [10320, 9833, 11815, 8482, 8383, 10299, 10794, 10423, 9857, 10321, 10438, 11179, 11841, 11641, 11906, 11400, 11013, 11830, 82, 1172, 1198, 104, 113]. Tropical [7274, 999]. Tropics [2412, 11877, 9672, 7539]. Trotter [1646]. Trotting [5603]. Trouble [481]. Troubles [8270, 8933]. Trousers [1581]. Truck [1728, 5290, 7797, 7756, 9337, 10784, 10483, 418, 479]. Truckee [8025]. True [3243, 6605, 7489, 9224, 767, 10577]. Truffault [4444]. Trumpet [165]. Truncated [3066]. Trunk [4109, 7104, 6870, 9205]. Trust [4773, 5429, 5045, 550, 5104]. Truth [5899, 231]. Try [1389, 9431]. Try-Cock [1389]. Try-Square [9431]. Trying [4956, 11173]. Tsushima [9597]. Tub [873]. Tube [2512, 2094, 2987, 2633, 3635, 3836, 4019, 3790, 4090, 4455, 3844, 6068, 7022, 7865, 7537, 8554, 9204, 9722, 11198, 265, 1236, 1027, 4998, 10845]. Tubes [4049, 10282, 10035, 11511]. Tubing [566]. Tubs [11092]. Tug [8500]. Tulare [9348]. Tulip [4662]. Tungsten [639]. Tungstic [3110]. Tunis [11324]. Tunnel [1408, 1445, 1671, 2408, 2241, 2592, 3335, 2883, 2964, 3395, 3407, 4218, 3791, 4254, 3786, 3815, 3737, 4197, 4827, 4049, 4158, 5900, 5879, 4994, 6915, 6435, 7058, 7468, 8186, 8847, 8698, 8583, 9111, 10416, 10339, 10169, 10282, 10035, 11514, 11251, 281, 7404, 347, 9072, 4467, 2070, 1826, 1620, 2483, 3078, 3352, 3609, 3500, 2826, 2916, 2577, 4139, 4613, 4658, 4253, 4659, 4191, 6089, 6114, 5792, 6142, 6235, 6358, 6269, 6679, 6606, 6780, 7261, 8071, 7774, 8651, 9067, 9204, 8867, 10848, 10960, 10825, 11210, 11351, 11259, 10630, 2601]. Tunnel-Boring [5879, 281]. Tunneling [2956, 5428, 9261, 9587, 9364]. Tunnels [3522, 4729, 4856, 8478, 8320, 9312, 10154, 11735, 1300, 1167, 10920, 3526, 9282, 9521, 9418, 10986]. Turbie [4032]. Turbine [2455, 2375, 3364, 2771, 3551, 2709, 3033, 2665, 3619, 4187, 4561, 4683, 3832, 4166, 3783, 4761, 4523, 5970, 5067, 5761, 4989, 5323, 5831, 5179, 5665, 5824, 5597, 6847, 6831, 7323, 6244, 6331, 6434, 7535, 7852, 8349, 7358, 7975, 8424, 8270, 7474, 9048, 9351, 8782, 9152, 9321, 8673, 8548, 8584, 9141, 10789, 10736, 10319, 10214, 10543, 11184, 11165, 11341, 1088, 333, 6311, 9837, 4684, 5362, 5521, 7809, 7360, 2235, 3983, 5135]. Turbine-Driven [7360]. Turbine-Propelled [4683, 6847, 9141]. Turbines [6353, 7102, 8100, 8358, 7899, 7471, 10235, 10683, 11002, 11318, 22, 348, 345, 191, 2237, 4752]. Turbo [9895]. Turbo-Electric [9895]. Turkestan [3422]. Turkey [4777]. Turkish [2734]. Turn [10342]. Turn-Down [10342]. Turnbuckle [10402]. Turned [9901]. Turning [2782, 5039, 9250, 11663]. Turns [5814]. Turnstile [10914]. Turntable [3273, 6808]. Turpentine [8850, 1910, 7037, 7133]. Turpentining [7373]. Turquoise [1491]. Turret [1479, 1456, 3669, 7832, 11144]. Turreted [4084]. Turrets [1467, 3054]. Turtle [3340, 9399]. Turtles [3689]. Tusk [6919]. Twelfth [3866]. Twentieth [2193, 2979, 4685, 9773, 11875]. Twenty [2889, 4524, 5732, 8164, 7793, 7854, 1270]. Twenty-Five [2889, 5732, 8164]. Twenty-Passenger [7793]. Twice [3172]. Twin [3009, 4289, 3904, 4299, 6668, 8190, 9124]. Twin-Balloon [6668]. Twin-Screw [4289, 3904, 4299, 8190, 9124, 3009]. Twine [3139, 9844].

Twins [2290, 5140]. Two [1979, 1889, 1622, 1750, 2032, 1694, 3565, 3549, 4286, 4678, 4190, 4747, 4509, 4617, 4458, 4120, 5301, 5907, 5550, 6380, 6223, 6738, 7354, 8138, 8099, 7621, 8212, 8797, 9005, 8583, 9311, 9120, 8832, 10416, 10493, 10458, 10460, 9931, 10483, 10767, 10042, 10061, 10943, 11480, 458, 5392, 5628, 2133, 4245, 483]. Two-Cycle [8797, 9120, 10483, 483]. Two-Engine [10458]. Two-Foot [4245]. Two-Gallon [8832]. Two-Hundred-Mile [10061]. Two-Passenger [10943]. Two-Speed [7354, 10460, 9931]. Tygard [6220]. Tyler [11169]. Tyne [4263]. Type [2309, 3462, 2794, 3150, 2734, 4703, 3719, 4397, 5308, 6073, 5797, 5808, 5850, 6744, 6811, 6216, 8423, 7865, 7580, 7756, 7919, 7893, 7882, 7406, 8795, 8642, 10444, 9948, 9952, 10576, 10490, 10363, 10565, 9961, 10980, 11942, 11117, 11668, 478, 1054, 86, 1220, 170, 1085, 9676, 7566, 5111, 7829, 8815, 466, 8278, 9114, 10778]. Type-Printing [5808]. Typecasting [5494]. Types [2266, 1411, 3549, 4739, 5414, 5076, 6913, 10240, 10909, 11948, 2600, 10465, 2401, 10242]. Typesetting [3548, 2553]. Typewriter [1416, 2240, 3022, 4108, 3871, 3841, 4129, 6128, 5744, 5706, 8079, 7381, 11584, 5677]. Typewriters [2558]. Typhoid [5176, 11492, 232]. Typhonoid [10778]. Typical [7790, 7791, 8335, 8765, 9120, 9939]. Tyrol [8303].

U.S.S. [1801, 11921]. Uganda [5826, 5408]. Ulm [7401]. Ultimate [6843]. Umbrella [10137]. Umbrellas [6271]. Unarmored [6827, 10586]. Unbreakable [11548]. Under-Water [11364]. Underground [2378, 2495, 3057, 4334, 5160, 6137, 165, 3911, 3012, 2444, 3304]. Underwood [2240]. Underwriters [10494]. Unearthed [4533]. Unearthing [1575]. Uneven [10385]. Unfortunate [1928]. Ungenerous [3857]. Ungentle [9007]. Unicycle [6367]. Uniformity [8700, 11650, 8233]. Unilens [9525]. Union [5566, 3819, 8897, 10496]. Unique [1702, 2060, 9430, 8872, 10462, 11259, 10880, 8434, 8505, 4828, 3012]. Unit [10480, 6852]. United [1962, 2278, 1845, 1409, 2459, 2001, 1762, 2280, 1682, 2149, 1478, 3196, 3270, 2896, 3160, 3527, 3678, 2593, 2844, 3363, 4560, 4728, 4500, 4687, 4160, 3977, 5104, 4936, 5640, 4957, 5544, 5745, 5558, 6259, 6396, 7263, 6964, 6775, 6261, 7001, 7106, 7080, 8493, 8202, 9351, 8548, 9003, 9160, 10757, 10436, 10790, 10418, 11223, 11318, 10973, 11923, 11000, 11945, 657, 533, 220, 894, 862, 238, 365, 1016, 32, 10591, 7886, 2348, 7088, 7755, 11192, 7858, 596, 3338, 3013, 4371, 7299, 9529, 9165, 2310, 1729, 10860, 4903, 10674, 4753, 8299, 10358, 11433, 6298]. United [7831, 10999]. Unites [2352]. Units [3952, 11914]. Universal [4394, 7233, 8808, 11279, 11526, 495, 10405, 11090]. Universe [8296]. Universities [5558, 9202]. University [11376, 8460, 11305, 7441, 11175]. Unjust [3857]. Unknown [10026]. Unloader [6007, 1319, 2555]. Unloading [2311, 3699, 5458, 7040, 8278, 9052]. Unparalleled [7692]. Unscientific [2127, 3060]. Unscrewing [69]. Unsinkable [7957, 9803]. Unusual [6783, 7681, 7734]. Unveiled [9841]. Unveiling [4026, 9400]. Unwary [4534]. Upbuilding [1537]. Upon [4959, 9642, 11946, 9565, 6853, 5565, 6334, 6354, 7275, 6245, 8836]. Upper [11257, 716, 635]. Upsetting [11711]. Upturned [5691]. Uralite [4441]. Urgent [10827, 10989]. Urgently [5900]. Use [2439, 1828, 1512, 2056, 3635, 4019, 4385, 3756, 5744, 5260, 6574, 7249, 9546, 9009, 8809, 8978, 10385, 9708, 10556, 11115, 11705, 11898, 11540, 10904, 10944, 10949, 10873, 11234, 11320, 922, 11111, 5934, 11784, 4602, 2472, 554, 7140, 8456, 9840, 10964, 3209, 2969, 4805, 9667, 10756, 11334, 4552, 4083, 940]. Used [4849, 7206, 7676, 9543, 10469, 11259, 11088, 11174, 9805, 7321, 5506]. Useful [3440, 5263, 8149, 10245, 10260, 6786]. Usefulness [6216]. Uses [4833, 7164, 8724, 10377, 9976, 3383, 6783, 8935, 2133, 11609, 5017, 3611, 3942, 8229]. Using [3940, 4550, 5677, 1202, 7421]. Utah [833]. Utensil [10383]. Utility [7059]. Utilization [1886, 4487, 9243, 4063, 7301]. Utilized [8302]. Utilizing [2942, 4090]. Uviol [9707].

V [11565, 2340]. Vaccine [2969, 11705]. Vacua [8046]. Vacuo [8125]. Vacuum [6105, 8872, 11214, 804, 57, 11618, 293]. Vagaries [1516, 9139]. Vagary [7726]. Valentines [9074]. Valhalla [7676]. Valley [1602, 9149, 4384, 11863]. Valtellina [2941, 4691]. Valuable [5271, 5394, 4051]. Value [4216, 1744, 6706, 8251, 7375, 7849, 8589, 9815, 10122, 10330, 11651, 820, 1087, 8935, 1519, 10424, 9208, 940]. Valueless [5087]. Valve [1535, 1748, 2756, 3112, 2557, 4586, 4716, 4056, 6166, 5142, 5646, 6243, 6550, 6200, 6202, 6273, 8059, 8509, 8130, 8665, 9551, 9170, 8644, 9924, 10979, 494, 122]. Valve-Gear [1535]. Valves [8717, 11019]. Vandalism [7488]. Vanderbilt [8170, 9230, 9270, 11603, 11580, 957, 3061, 6963, 6980, 8075, 8108, 8139, 7537, 9287]. Vanderpool [6050]. Vane [10969]. Vanishing [881]. Vapor [3703, 7541, 10878, 7300]. Vaporization [4803]. Vaporized [6776]. Vaporizer [5441, 5580]. Variability [2797]. Variable [3865, 4259, 5607, 6813, 9367, 9799, 10800, 7734]. Variable-Speed [9799]. Variag [2435, 2095]. Variations [3587, 10872]. Varied [4833]. Variety [827]. Various [3716, 3756, 4369, 3611]. Vase [9801]. Vast [2391, 2566, 3677, 3972, 7080, 9303, 1075]. Vaulx [3317]. Vegetable [4173, 7019]. Vegetable-Cutter [4173]. Vegetables [1433, 7396, 605, 10671]. Vehicle [6637, 7144, 6826, 6542, 7838, 9652, 9904, 10719, 10782, 10225, 10818, 34, 541]. Vehicles [1322, 2333, 2215, 2871, 2986, 2873, 4297, 5289, 5311, 6338, 8240, 7469, 9376, 9603, 7908]. Velocipede [1564]. Velocities [8250]. Velocity [5453, 6396]. Vending [1708, 5937, 9375, 10981]. Vending-Machine [1708]. Veneer [4323, 11283]. Venice [3980, 9508]. Venom [11035, 11115, 7301]. Venomous [5508, 4944, 5526, 5547]. Vent [4056, 5681]. Vent-Closing [4056]. Ventilated [4418, 7220]. Ventilating [4856]. Ventilation [4522, 4334, 5431, 7056, 7719, 9552, 8717, 47]. Ventilator [3365, 5505, 8993, 11333]. Venus [11879, 1125]. Verant [6999]. Verdict [4191]. Verified [11533]. Vermont [9685]. Versus [3245, 3522, 3348, 4049, 6500, 6886, 8073, 7469, 7763, 10849, 157, 252, 144, 1160, 1362]. Vertical [3118, 4370, 5960, 5032]. Very [7504, 6054]. Vessel [2979, 4703, 4526, 5229, 7211, 240, 835, 2009, 3954]. Vessels [2311, 2435, 1512, 2657, 3787, 5429, 5458, 5986, 5905, 8624, 8940, 9141, 10174, 9770, 8745, 4684, 5903, 9704, 10112, 5017]. Vest [5344]. Vestibule [5705]. Vesuvius [8708, 7519]. VI [2355]. Viaduct [2377, 2843, 3592, 797, 1253, 2129, 7472]. Viagraph [7584]. Vibration [3057, 8739, 10642, 186, 10198]. Vibrationless [9126]. Vibrations [3304, 7958, 5835, 2472]. Vibrator [10480]. Vibratory [346]. Vicinity [8680]. Vickers [6267, 8858, 10716]. Victim [5838]. Victoria [3042, 5738, 3009, 3003, 8410, 7544, 10110]. Victoria-Luise [3009]. Victorian [7474]. Victory [7692]. Vienna [4576, 4808, 10206, 10526]. Vieux [2125]. View [2289, 3243, 4047, 5280, 7489]. Viewing [8876, 6999]. Views [8770]. VII [3551, 7422, 2963]. VIII [3018]. Vila [5623]. Village [2617, 6983, 2182]. Villard [3724]. Vincennes [2666]. Vindicated [9767]. Vine [5109]. Vines [7068]. Vinne [5848]. Violet [2213, 5715]. Violin [11137, 4710, 10659, 10472]. Viper [1493, 2435, 2244, 3260]. Virag [1559]. Virchow [4366]. Virgin [7539]. Virgina [9183]. Virginia [3553, 9817, 1494, 3482]. Virus [2969]. Viscosimeter [7635]. Vise [3062, 11279, 1154, 733]. Visible [6122, 9145]. Visit [7579, 8235, 8257]. Visiting [8255, 803]. Viticulture [11082]. Vivero [6079]. Vivisection [10435]. Vixen [4289]. Vladivostock [7346]. Vocal [7694, 4598]. Vohwinkel [4167]. Voisin [591, 276]. Volcanic [4575, 3995, 5044, 9136, 606, 6078]. Volcano [4457, 8257]. Volcanoes [5720, 4050, 510]. Volga [2083]. Volt [4690, 8963, 3766]. Voltages [7721]. Volts [2591, 7444]. Voting [7215]. Voyage [7136, 9027, 673]. Vs [5278, 1480, 2002, 2725, 3635, 4405, 4787, 6045, 6157, 5691, 7471, 459, 479]. Vulcanized [5034]. Vulcanizing [2351]. Vulnerability [3669].

W [2709]. Wachusett [7961]. Wagon [1584, 1340, 3267, 3272, 4281, 7794, 7796, 9865, 10996]. Wagons [10135, 11762, 4917, 7379]. Waiter [8507]. Wales [4692]. Walker [89]. Walking [10285, 761]. Walking-Beam [761]. Wall [9101, 9777, 11111, 8773, 6873, 1445, 8530]. Walschaert [8644]. Want [3086]. Wanted [2283, 2209, 2641, 11495, 11535, 1242, 9643]. Wants [2559, 2582, 3276, 3296, 3319, 3341, 3368, 3390, 3412, 3430, 3448, 3470, 3494, 3513, 3533, 3554, 3578, 3601, 3623, 3647, 3670, 3690, 2604, 2625, 2649, 2670, 2674, 2717, 2740, 2764, 2809, 2831, 2874, 2899, 2919, 2944, 3071, 3098, 3122, 3144, 3164, 3186, 3210, 3233, 3254, 3705, 3710, 4325, 4354, 4377, 4398, 4423, 4445, 4473, 4494, 4514, 4539, 4563, 4588, 4607, 4627, 4650, 4672, 4694, 4718, 4740, 4765, 4790, 4815, 4837, 4858, 4884, 4907, 3746, 3772, 3776, 3826, 3850, 3872, 3895, 3920, 3945]. Wants [3966, 3989, 4012, 4038, 4066, 4091, 4110, 4132, 4152, 4176, 4203, 4224, 4247, 4269, 4303, 4922, 4924, 5584, 5611, 5632, 5655, 5683, 5707, 5725, 5752, 5775, 5799, 5816, 5839, 5871, 5891, 5914, 5940, 5961, 5989, 6015, 6035, 6062, 6082, 6107, 6129, 6149, 6168, 4980, 5002, 5007, 5028, 5059, 5119, 5147, 5169, 5195, 5198, 5241, 5268, 5293, 5315, 5346, 5350, 5371, 5418, 5445, 5462, 5486, 5509, 5530, 5553, 6187, 6209, 6788, 6792, 6836, 6861, 6880, 6898, 6920, 6947, 6967, 6986, 7010, 7027, 7046]. Wants [7070, 7095, 7119, 7146, 7172, 7196, 7222, 7241, 7264, 7285, 7309, 7330, 6212, 7348, 6230, 6278, 6299, 6325, 6343, 6347, 6383, 6407, 6426, 6429, 6469, 6487, 6519, 6544, 6562, 6591, 6612, 6631, 6672, 6698, 6721, 6746, 6767, 8195, 7364, 7386, 7992, 8008, 8032, 8063, 8092, 8112, 8131, 8152, 8177, 8218, 8242, 8263, 8288, 8306, 8326, 8341, 8365, 8392, 8414, 8439, 8464, 8483, 8510, 8533, 7409, 7430, 7460, 7481, 7505, 7528, 7548, 7569, 7589, 7614, 7638, 7662, 7686, 7710, 7739, 7757, 7775, 7815]. Wants [7841, 7867, 7894, 7921, 7944, 7967, 8555, 8558, 9154, 9174, 9195, 9218, 9234, 9254, 9273, 9295, 9315, 9339, 9358, 9381, 9406, 9433, 9453, 9472, 9494, 9514, 9532, 9553, 9572, 9590, 9608, 9633, 9656, 9677, 8601, 8604, 8645, 8667, 8692, 8709, 8726, 8753, 8775, 8817, 8840, 8859, 8880, 8906, 8926, 8949, 8970, 8994, 9018, 9036, 9057, 9083, 9103, 9131, 10181, 2851]. War [1458, 2435, 1794, 3476, 3908, 4728, 3986, 5031, 5206, 5496, 5327, 5278, 6651, 6936, 7717, 7861, 8590, 10174, 10018, 11817, 10403, 6440, 4413, 2392, 6044, 5277, 7182, 6844, 6207, 7345, 8339, 7596, 7731, 7999, 8048, 9190, 4756, 1765]. War-Car [3986]. War-Game [4728]. Ware [7545, 6014]. Warfare [3477, 4660, 3736, 6444, 7784, 511, 861, 9750, 5078, 6833]. Warmer [10364]. Warning [1771, 3173, 8612, 1113]. Warp [8058]. Warp-Beam [8058]. Warped [6480]. Warship [3475, 4992, 6844, 7294, 7106, 6884, 10590, 11275, 11385, 11212, 11056]. Warships [2437, 1778, 1360, 3035, 4617, 5039, 4931, 5227, 6928, 7138, 8294, 8049, 7748, 9424, 10810, 803, 8964, 5389, 1166, 8983]. Was [7772, 10423, 11605, 11346, 8690, 7909, 7701, 5442, 5287, 10256, 10921, 6875]. Wash [741]. Washer [327]. Washing [2126, 1889, 4836, 8057, 10362, 11092, 11760, 10837, 1030]. Washington [7441, 2145, 1606, 2957, 4244, 4738, 5996, 8435, 3884, 4903, 7398, 7628, 11417, 9166, 9669]. Wasps [1226]. Waste [3336, 4550, 5704, 5103, 11058, 966, 407]. Waste-Heat [3336]. Wasted [8281]. Wastefulness [10536]. Wastes [1886]. Watch [4351, 7495, 9077, 964, 672, 896, 993, 7566, 7600, 7908]. Watches [7915]. Water [2004, 5014, 809, 1958, 1749, 2324, 1392, 2512, 2395, 1772, 2335, 1382, 1454, 2044, 2308, 2492, 1986, 2147, 1349, 2094, 2756, 3205, 3223, 2533, 3530, 2967, 3064, 3594, 3332, 2633, 3635, 4849, 3836, 3726, 4019, 4397, 3799, 4365, 3790, 4834, 4529, 4321, 4455, 3844, 5117, 5704, 5563, 5690, 5577, 5987, 5254, 5284, 5263, 5834, 5103, 5292, 6373, 6972, 6461, 6744, 6753, 6257, 6246, 7081, 6886, 8282, 8017, 8118, 7901, 7397, 7747, 7865, 8241, 7393, 7358, 8361, 7537, 7961, 9600, 9025, 9541, 9595, 8554, 8610, 8618, 9042, 9117, 8839]. Water [8543, 10604, 10566, 10351, 10111, 10293, 9757, 10212, 10546, 10418, 10285, 11884, 11029, 10886, 11729, 11626, 11527, 11596, 10831, 11497, 10827, 11364, 11568, 11492, 11728, 1179, 387, 164, 660, 1100, 980, 383, 9113, 4963, 807, 7183, 1917, 2441, 2939, 1468, 3065, 7067, 7601, 8877, 7988, 7825, 2340, 462, 6379, 10877, 10878, 9187, 1176, 597, 3818, 435, 10692].

Water-Absorbing [5834]. Water-Current [8241]. Water-Curtain [7081]. Water-Feed [8839]. Water-Gages [2308]. Water-Gas [2967, 3064]. Water-Heater [3223]. Water-Level [3799]. Water-Pipe [2533]. Water-Pipes [2756]. Water-Proof [1100]. Water-Tank [3594]. Water-Tanks [1986]. Water-Tube [2094, 2633, 3635, 3836, 4019, 3790, 3844, 7865, 7537, 8554]. Water-Walking [10285]. Water-Ways [10604]. Water-Wheels [3065]. Watering [8234, 8549]. Waterproof [1632, 11667]. Waters [3587, 5598, 142, 5037, 116]. Watershed [11112]. Watertight [3953]. Waterway [382, 8281, 6123, 1874, 1073, 10341, 9817]. Waterways [3632, 2751, 7660, 9878]. Waterworks [5522]. Wave [2554, 2884, 3929, 4034, 5608, 6289, 7279, 7251, 8216, 9, 1306, 3595, 3702]. Wave-Detector [3929]. Wave-Operated [7251]. Wave-Selective [2554]. Waverley [6665]. Waves [3132, 4316, 4749, 7103, 7750, 9097, 9417, 7958, 11113, 9098]. Wax [7118, 326, 10671]. Way [1869, 5188, 5677]. Wayne [9942]. Ways [4365, 8128, 8662, 10604, 397]. Weak [488]. Wealth [3859, 8342, 9224, 8541]. Weapon [1709]. Weapons [9469, 6916]. Wear [8449, 10214, 7584]. Weather [1776, 2707, 4321, 10206, 11555, 11862, 1050, 10969, 3616, 10173, 3809]. Weatherboard [1152]. Weaving [3567, 4503]. Web [10204]. Wedsted [7500]. Weed [7660]. Weed-Infested [7660]. Weeding [1558]. Weeding-Machine [1558]. Weeds [11176]. Week [5050, 6207]. Weevils [2763]. Weigh [8296]. Weighed [9210]. Weighing [2603, 9992, 11242, 11331, 2472]. Weighing-Machine [2603]. Weight [2421, 4739, 5453, 6906, 6667, 7805, 9545, 11722]. Weights [4701, 11301]. Weird [9469]. Welch [9524]. Welder [10715]. Welding [9138, 9755, 5038, 8477, 6360]. Well [3019, 4881, 5546]. Wellcome [999]. Wellman [9148, 8924, 10086, 1053]. Wells [4536, 6800, 8364, 8130, 8600, 9139, 539, 3383, 4142, 260]. Welsbach [2050, 4165]. Wentworth [2648]. Were [10360, 751]. Wesson [9621]. West [11314, 10860, 2869, 3553, 4005, 4624, 3879, 4335, 7877, 10105, 1078, 1071, 9399, 8338]. Western [6241, 8959, 6147, 7260, 10612]. Westinghouse [3569, 3688, 5498]. Wet [599]. Wet-Grain [599]. Wetterhorn [263]. Wettin [5333]. Weybridge [10687]. Whale [1985, 4978, 8337, 7811]. Whalebacks [11354]. Whales [3093]. Wharf [1922]. Wheat [1940, 2714, 10024]. Wheel [1322, 3508, 2739, 2685, 3410, 3797, 4472, 4195, 5283, 6889, 6338, 7144, 6542, 7800, 8061, 9653, 9115, 9082, 8809, 10325, 10719, 10760, 10498, 9690, 11459, 11606, 541, 7601, 9650, 10997, 6832]. Wheel-Blocking [9082]. Wheel-Flange [9690]. Wheelbarrow [4373]. Wheeled [4343, 7586]. Wheels [6574, 7797, 8449, 8806, 9954, 11020, 11502, 366, 3065, 3867]. Where [4482, 4342, 11341, 3820, 9774, 5000, 10860, 7701]. Which [6688, 3383, 10109, 9489, 8919, 9650, 10472, 10070, 616, 9962, 11064, 4579]. Whip [13]. Whipple [4805]. Whirl [4291, 4882]. Whirlpool [2402, 8387, 3230]. Whirls [6033]. Whispering [7398]. Whistle [4192, 7732].

Whistles [11136]. Whistling [8739, 2867]. White [4550, 6072, 11781, 812, 3682, 2957, 2731, 3079, 3642, 5190, 7211, 9959, 10005, 11010, 11839, 2474, 6118]. Whitehead [8272]. Whitewashing [9312]. Whittled [2082]. Who [8636, 5287]. Whom [2953, 714]. Wichita [4581]. Wickless [2647]. Widened [10433]. Widening [3044, 11846, 3814]. Wiesbaden [5522]. Wiesen [1253]. Wilbur [823, 11676, 843, 214]. Wild [2415, 3306, 2820, 3944, 9147, 4493, 5237, 4435]. Wiley [6049]. Wilhelm [5168, 1767, 3177]. Will [3193, 2640, 1079, 7133, 9696]. William [3079, 5494, 6098, 5131, 7236, 10689, 11010, 759, 9524]. Williamsburg [5878, 6484, 11731]. Willis [3161]. Wilson [8053]. Wimshurst [11546]. Winch [11664, 1174]. Wind [2335, 5358, 5921, 5783, 6569, 8251, 8062, 7374, 7876, 6245, 1025, 562]. Wind-Driven [6569]. Winder [96]. Winding [4939]. Windmill [3290, 4531, 7216, 6275, 6322, 10022]. Windmills [1583, 10546, 7428]. Window [1556, 1514, 3687, 8215, 8437, 8993, 9445, 9491, 9232, 8598, 10179, 11080, 35, 1304, 8878]. Window-Blind [10179]. Window-Guard [1556]. Window-Sash [3687]. Windows [3365, 6686, 10362, 6669]. Wine [2261, 7085]. Wine-Making [7085]. Wing [4940, 10401]. Wings [6937, 7003, 1128]. Winning [6651, 8832, 9566, 9270, 11300, 11830, 262]. Wins [3408, 11253]. Winter [3891, 10812, 6686, 6729, 4265, 4918, 2820, 3838, 11062]. Winton [6918]. Wiped [1103]. Wire [1443, 1916, 1488, 3349, 3160, 5860, 6045, 6079, 5954, 6396, 6267, 6858, 7670, 8858, 10813, 11860, 1160, 9805, 11940]. Wire-Hanger [1488]. Wire-Reeling [6858]. Wire-Stitching [5860]. Wire-Stretcher [1916]. Wire-Wound [3160, 6396, 6267, 7670, 8858, 1160]. Wireless [2096, 1931, 3550, 2823, 3197, 2546, 2554, 3063, 3008, 3080, 3520, 2863, 4867, 3893, 3930, 3714, 4485, 4851, 4870, 4845, 4832, 3704, 4149, 4164, 3736, 3883, 5102, 5675, 5599, 5744, 5610, 6099, 4970, 5434, 4964, 5426, 6603, 7279, 6380, 6936, 7184, 7767, 8019, 8102, 9582, 8704, 8849, 8914, 9076, 9216, 9247, 8933, 8596, 9226, 10541, 9752, 10516, 10376, 11458, 11154, 11584, 11368, 11110, 1301, 998, 8968, 8379, 3185, 3525, 4755, 4877, 4900, 3722, 4324, 4439, 4997, 4916, 5259, 7448, 7860, 9031, 9326, 9488, 9732, 9883, 4004, 10129, 5112, 8815]. Wireless [5185, 432, 10833, 4200, 1553, 10358]. Wirelessly [775, 63]. Wirelessly-Controlled [775, 63]. Wires [7317, 10685]. Wisconsin [2288, 2095, 3479, 3480, 3758]. Wise [1790]. Wissahickon [10558]. Withdrawing [4536]. Within [7170, 3410]. Without [7170, 6025, 5902, 11607, 11214, 10922, 46, 1158, 322, 10690, 326, 11688, 9422, 7606, 225, 1103, 1176]. Withstand [7444]. Witness [4046]. Witnesses [727]. Witt [2464]. Wolvin [6379]. Woman [11280]. Women [1864, 5055]. Wonder [9882]. Wonderful [2534, 2985, 10028, 1072, 341, 9469, 11065]. Wonders [11542]. Wood [2007, 2389, 1908, 1792, 2265, 1652, 2738, 3311, 4342, 4921, 6570, 9581, 11525, 966, 202, 867, 1203, 4504, 7153]. Wood-Working [2738]. Wooden [2533, 3310, 9666, 8912, 9730, 10970, 8575]. Woodpeckers [5444]. Woods [2215, 9403, 8229]. Woodward [1746]. Woodworker [9173]. Word [380]. Words [9395]. Work [1226, 1351, 1984, 2241, 1361, 1431, 3384, 2840, 2638, 3589, 3407, 4577, 3809, 5166, 4941, 4960, 5417, 6641, 6572, 6917, 7058, 7913, 7608, 8321, 7838, 8725, 10606, 10706, 10111, 10413, 9723, 11213, 11857, 11152, 11865, 749, 1159, 1169, 5906, 10948, 7751, 3866, 4079, 6311, 10155, 8144, 4902, 7321, 3523, 3411, 8643, 4198, 5771, 2686, 3983, 3739, 7300, 563, 7681, 568, 11788, 7260, 9098, 10999, 10242, 201]. Workholder [4394]. Working [1958, 1470, 2738, 6008, 6422, 6918, 6197, 8404]. Workman [5880]. Workmen [3835, 4215, 5250]. Works [2952, 3248, 3266, 3688, 5545, 4946, 5537, 7495, 8074, 10062, 11303, 1076, 8235, 9215, 1605, 1679, 2199, 2456, 2733, 4393, 288, 3868]. Workshop [11544, 11610, 11677, 11686, 11708, 64, 94, 197, 223, 290, 318, 394, 419, 467, 516, 561, 638, 683, 731, 1099, 1150, 1199, 1227, 1258, 296, 11690, 11787, 66, 640, 149, 224, 320, 969, 484, 805, 1022, 865, 6179]. World [1987, 2327, 2325, 1478, 2072, 1847, 1890, 2930, 3113, 3596, 3375, 2572, 2575, 4754, 4387, 3713, 3880, 4351, 4526, 4392, 4457, 5604, 6126, 7231, 6203, 7211, 6422, 7648, 8148, 8160, 8399, 7823, 9162, 8772, 8614, 9346, 9167, 9160, 10664, 10041, 10542, 11116, 11296, 11034, 616, 610, 934, 5214, 5668, 5721, 7193, 10877, 2108, 11151, 6399, 7935, 1563, 6138, 6001, 7298, 6983, 7292, 7277, 7261, 7282, 6497, 615, 7342, 5693, 6894]. Worlds [6433, 8161, 6459]. Worm [1659, 1366]. Worst [9181]. Worsted [9839]. Would [2953, 7258, 11769]. Wound [3160, 6396, 6267, 7670, 8858, 1160]. Wrapping [9146]. Wraps [11158]. Wreck [1516, 3163, 4003, 4191, 4975, 5578, 6938, 6507, 8047, 9344, 9671, 9730, 10264, 11839]. Wrecked [1985, 4158, 9843, 11585, 6958]. Wrecking [5133, 240]. Wrecks [9560]. Wrench [1775, 2330, 1732, 2646, 2602, 3120, 4649, 6405, 8438, 9989, 9906, 11466, 570, 34, 11614]. Wright [8664, 11481, 11805, 11434, 11472, 11105, 11083, 11135, 11937, 11945, 11676, 1255, 591, 1217, 823, 306, 843, 1291, 276, 214, 6685]. Wringers [4421]. Wrinkles [2400]. Writing [3339, 7454, 8691, 4443]. Writings [9731]. Wrought [7903]. Wydts [4145]. Wyoming [2081, 3136, 4669, 10606].

X [2491, 2886, 5305, 10010, 2686]. X-Ray [10010, 2686]. X-Rays [2491, 2886, 5305]. Xiphopages [2438, 1897]. Xiphopagous [2290].

Yacht [1700, 2144, 2399, 2759, 2611, 2668, 3042, 2620, 2890, 3009, 3005, 4682, 4287, 4289, 4392, 6162, 5067, 5066, 5065, 5500, 5063, 5064, 5922, 5068, 5618, 6441, 7130, 7703, 7630, 8081, 11723, 11491, 617, 11371, 2634, 3079]. Yachting [5050, 10104]. Yachts [2914, 4683, 5597, 9126, 1268]. Yale [3644]. Yankee [1016]. Yard [1921, 4194, 4941, 9304, 11824, 8407, 1395, 2103, 7542, 8532, 10304, 3426]. Yards [1360, 3282, 2840, 3475, 6884, 11056, 4315]. Yawning [10126]. Year [1407, 1316, 2668, 2531, 3330, 3695, 4913, 6176, 6219, 8184, 7353, 8332, 7789, 10214, 9683, 10807, 11272, 4147]. Years [3252, 8101, 768, 762, 8691, 10950, 7109]. Yellow [2612, 8451]. Yerkes [4245].

Yew [1631]. Yielding [5387]. Yoke [4814]. Yolk [11412]. York [2487, 1530, 1867, 1929, 1772, 1921, 1826, 1321, 1620, 2483, 2103, 2044, 3205, 2784, 2661, 3220, 3016, 3352, 3609, 2826, 2642, 3118, 3677, 3764, 4404, 4826, 4758, 3743, 4243, 4488, 4507, 4708, 4139, 4347, 4170, 4086, 4130, 4700, 4191, 5737, 5690, 5065, 5167, 5064, 5284, 5540, 5301, 5791, 5277, 6373, 6966, 6506, 6830, 6236, 7041, 6638, 7254, 6846, 6871, 7341, 6981, 6342, 6606, 6753, 6223, 7885, 8492, 8118, 7901, 7747, 7393, 7583, 7698, 7424, 8377, 8378, 9425, 9025, 8742, 9561, 8618, 9047, 9521, 8679, 9272, 8856, 9042]. York [9742, 10103, 9877, 10351, 10111, 9699, 10132, 10521, 10061, 10264, 11428, 11915, 10963, 10897, 11496, 10852, 11060, 11430, 11633, 11294, 11856, 11011, 11636, 913, 583, 7659, 1129, 7359, 1529, 7962, 7085, 11655, 2222, 1749, 2155, 1382, 1454, 2241, 2377, 2220, 2637, 3386, 2817, 2838, 2967, 3064, 3592, 3593, 2593, 4784, 4729, 4414, 4120, 5772, 4942, 5527, 5415, 6915, 6447, 8047, 7769, 8678, 8875, 8660, 8680, 8967, 9528, 9180, 10477, 9875, 10154, 9975, 11187, 11177, 11734, 957, 409, 1135, 762]. Young [9743, 5956]. Youngest [1459]. Yucatan [7086, 7116]. Yucca [9313]. Yukon [3682, 195]. Yuma [11863, 11590].

Zanesville [4059]. Zapotecan [4533]. Zealand [7240]. Zebra [9747]. Zenobe [3085]. Zeppelin [2443, 1628, 2118, 9128, 11828, 11897, 465, 673, 458, 158, 776]. Ziegler [2727, 8490, 5735]. Zinc [1672, 9668, 3381]. Zinc-Aluminium [9668]. Zodel [2779]. Zodiacal [3087]. Zone [8226, 8573, 560]. Zoological [1321, 6966, 7263, 6223, 7089, 1129, 10297, 1733, 9355]. Zossen [4690, 5110, 3398, 5910, 4221]. Zuni [3914].

References

Russell:1909:HJa

[1] Henry Norris Russell. The heavens in January. Scientific American, 100(1):7, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

       scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
       7.pdf.

Anonymous:1909:RPIa

[2] Anonymous. Recently patented inventions. Scientific American, 100(1):12, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

       scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
       12.pdf.

Anonymous:1909:NQa

[3] Anonymous. Notes and queries. Scientific American, 100(1):12–13, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n1/pdf/scientificamerican01021909-12a.pdf.

Anonymous:1909:BNBa

[4] Anonymous. Books: New books, etc. Scientific American, 100(1):13, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n1/pdf/scientificamerican01021909-13.pdf.

Anonymous:1909:IIa

[5] Anonymous. Index of inventions. Scientific American, 100(1):13–15, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n1/pdf/scientificamerican01021909-13a.pdf.

Anonymous:1909:RTa

[6] Anonymous. Retrospect of the tear 1908. Scientific American, 100(1):2–3, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
2.pdf.

Anonymous:1909:CPP

[7] Anonymous. Carrier pigeons as photographers. Scientific American, 100(1):4, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
4.pdf.

Anonymous:1909:MPS

[8] Anonymous. Moving pictures that sing and talk. Scientific American, 100(1):5, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
5.pdf.

Anonymous:1909:TWM

[9] Anonymous. Tidal wave making on the stage. Scientific American, 100(1):8, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
8.pdf.

Anonymous:1909:MRC

[10] Anonymous. Miniature railroads for country transportation. Scientific American, 100(1):9–10, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
9.pdf.

Benjamin:1909:TCC

[11] Marcus Benjamin. Thomas Chrowder Chamberlin. Scientific American, 100(1):10, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
10.pdf.

Anonymous:1909:IRA

[12] Anonymous. Improved refrigerating apparatus. Scientific American, 100(1):11, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
11.pdf.

Anonymous:1909:WSa

[13] Anonymous. Whip socket. Scientific American, 100(1):11, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n1/pdf/scientificamerican01021909-11a.pdf.

Anonymous:1909:SD

[14] Anonymous. Spring drawer. Scientific American, 100(1):11, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n1/pdf/scientificamerican01021909-11b.pdf.

Anonymous:1909:CCP

[15] Anonymous. Combined case and post for tennis nets. Scientific American, 100(1):11, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
11c.pdf.

Anonymous:1909:IH

[16] Anonymous. An improved hinge. Scientific American, 100(1):11, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n1/pdf/scientificamerican01021909-11d.pdf.

Anonymous:1909:PIS

[17] Anonymous. Practical and instructive scientific books. Scientific American, 100(1):15, January 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n1/pdf/scientificamerican01021909-
15.pdf.

Anonymous:1909:RPIb

[18] Anonymous. Recently patented inventions. Scientific American, 100(10):192, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n10/pdf/scientificamerican03061909-
192.pdf.

Anonymous:1909:PNPa

[19] Anonymous. Patent notes: Patents. Scientific American, 100(10):192, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n10/pdf/scientificamerican03061909-192a.pdf.

Anonymous:1909:IIb

[20] Anonymous. Index of inventions. Scientific American, 100(10):192–195, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n10/pdf/scientificamerican03061909-192b.pdf.

Anonymous:1909:NS

[21] Anonymous. The new steel. Scientific American, 100(10):182, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n10/pdf/scientificamerican03061909-182.pdf.

Anonymous:1909:ABT

[22] Anonymous. American and British turbines. Scientific American, 100(10):182, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n10/pdf/scientificamerican03061909-
182a.pdf.

Anonymous:1909:SSPa

[23] Anonymous. The second sea power — America or Germany. Scientific American, 100(10):182, March 6, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n10/
pdf/scientificamerican03061909-182b.pdf.

Anonymous:1909:Ea

[24] Anonymous. Engineering. Scientific American, 100(10):183, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n10/pdf/scientificamerican03061909-183.pdf.

Anonymous:1909:Eb

[25] Anonymous. Electricity. Scientific American, 100(10):183, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n10/pdf/scientificamerican03061909-183a.pdf.

Anonymous:1909:Sa

[26] Anonymous. Science. Scientific American, 100(10):183, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n10/pdf/scientificamerican03061909-183b.pdf.

Anonymous:1909:PFCa

[27] Anonymous. Possibilities of the fusible core process. Scientific American, 100(10):184, March 6, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n10/
pdf/scientificamerican03061909-184.pdf.

Kieffer:1909:MLI

[28] H. Prime Kieffer and C. E. A movable lock for inclined canals. Scientific American, 100(10):184–185, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n10/
pdf/scientificamerican03061909-184a.pdf.

Anonymous:1909:EEB

[29] Anonymous. Explosive energy of a boiler. Scientific American, 100(10):185–186, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n10/pdf/scientificamerican03061909-
185.pdf.

Honey:1909:MES

[30] Frederic R. Honey. Morning and evening stars for 1909. Scientific American, 100(10):186–187, March 6, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n10/
pdf/scientificamerican03061909-186.pdf.

Dunlap:1909:RSN

[31] Orrin B. Dunlap. The recent subsidence of Niagara Falls. Scientific American, 100(10):187, March 6, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n10/
pdf/scientificamerican03061909-187.pdf.

Anonymous:1909:USL

[32] Anonymous. United States life-saving steamer “Snohomish”. Scientific American, 100(10):187–188, March 6, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n10/
pdf/scientificamerican03061909-187a.pdf.

Anonymous:1909:IBA

[33] Anonymous. The Isna Barrage across the Nile. Scientific American, 100(10):189–191, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n10/pdf/scientificamerican03061909-
189.pdf.

Anonymous:1909:IVW

[34] Anonymous. Improved vehicle wrench. Scientific American, 100 (10):191, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n10/pdf/scientificamerican03061909-
191.pdf.

Anonymous:1909:DOW

[35] Anonymous. Device for operating window shutter. Scientific American, 100(10):191, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n10/pdf/scientificamerican03061909-
191a.pdf.

Anonymous:1909:KS

[36] Anonymous. Kettle still. Scientific American, 100(10):191, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n10/pdf/scientificamerican03061909-191b.pdf.

Anonymous:1909:NEM

[37] Anonymous. A novel electric motor. Scientific American, 100(10):191, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n10/pdf/scientificamerican03061909-191c.pdf.

Anonymous:1909:IGH

[38] Anonymous. Improved garment hanger. Scientific American, 100 (10):191, March 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n10/pdf/scientificamerican03061909-
191d.pdf.

Anonymous:1909:Ca

[39] Anonymous. Correspondence. Scientific American, 100(11):202, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-202.pdf.

Anonymous:1909:RPIc

[40] Anonymous. Recently patented inventions. Scientific American, 100 (11):209–210, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
209.pdf.

Anonymous:1909:NQb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-210.pdf.

Anonymous:1909:BNBb

[42] Anonymous. Books: New books, etc. Scientific American, 100(11):211, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-211.pdf.

Anonymous:1909:PNPb

[43] Anonymous. Patent notes: Patents. Scientific American, 100(11):211, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-211a.pdf.

Anonymous:1909:IIc

[44] Anonymous. Index of inventions. Scientific American, 100(11):211–215, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-211b.pdf.

Anonymous:1909:TRT

[45] Anonymous. A tribute to Roosevelt’s technical judgement. Scientific American, 100(11):198, March 13, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n11/
pdf/scientificamerican03131909-198.pdf.

Anonymous:1909:CF

[46] Anonymous. A cruiser without funnels. Scientific American, 100(11):198, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
198a.pdf.

Anonymous:1909:VPC

[47] Anonymous. Ventilation of passenger coaches. Scientific American, 100(11):198, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
198b.pdf.

Anonymous:1909:SFC

[48] Anonymous. A successful friction clutch. Scientific American, 100(11):198, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
198c.pdf.

Anonymous:1909:Ec

[49] Anonymous. Engineering. Scientific American, 100(11):199, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-199.pdf.

Anonymous:1909:Ed

[50] Anonymous. Electricity. Scientific American, 100(11):199, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-199a.pdf.

Anonymous:1909:Sb

[51] Anonymous. Science. Scientific American, 100(11):199, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n11/pdf/scientificamerican03131909-199b.pdf.

Anonymous:1909:PS

[52] Anonymous. The projector in surgery. Scientific American, 100(11):200, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-200.pdf.

Anonymous:1909:GDB

[53] Anonymous. Gas-driven battleships and cruisers. Scientific American, 100(11):201, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
201.pdf.

Anonymous:1909:POP

[54] Anonymous. A peculiar optical phenomenon. Scientific American, 100(11):202, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
202a.pdf.

Anonymous:1909:DRC

[55] Anonymous. Did the “Republic” carry searchlights? Scientific American, 100(11):202, March 13, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n11/
pdf/scientificamerican03131909-202b.pdf.

Anonymous:1909:AF

[56] Anonymous. The airship of the future. Scientific American, 100(11):202, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
202c.pdf.

Anonymous:1909:SVB

[57] Anonymous. Stress in a vacuum balloon. Scientific American, 100(11):202, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
202d.pdf.

Anonymous:1909:EI

[58] Anonymous. The earthquake in Italy. Scientific American, 100(11):202, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-202e.pdf.

Anonymous:1909:CN

[59] Anonymous. Curiosities of numbers. Scientific American, 100(11):202, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-202f.pdf.

Anonymous:1909:ANa

[60] Anonymous. Aeronautical notes. Scientific American, 100(11):203, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-203.pdf.

Allen:1909:MFH

[61] William Allen. The M’Call’s ferry hydro-electric power plant. Scientific American, 100(11):203–205, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n11/
pdf/scientificamerican03131909-203a.pdf.

Anonymous:1909:TBT

[62] Anonymous. Test of the Bell tetrahedral-cell aeroplane in Nova Scotia. Scientific American, 100(11):205, March 13, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n11/pdf/scientificamerican03131909-205.pdf.

Anonymous:1909:WCT

[63] Anonymous. A wirelessly-controlled torpedo. Scientific American, 100(11):206, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
206.pdf.

Anonymous:1909:HMWa

[64] Anonymous. Handy man’s workshop. Scientific American, 100(11):207, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-207.pdf.

Thorpe:1909:CCF

[65] Edward Thorpe. A cheaply-constructed fireless cooker. Scientific American, 100(11):207, March 13, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n11/
pdf/scientificamerican03131909-207a.pdf.

Bayley:1909:FWIa

[66] I. G. Bayley. Furnishing the workshop — IV. Scientific American, 100(11):207, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
207b.pdf.

Ward:1909:EFC

[67] Frederick E. Ward. An electrical fireless cooker. Scientific American, 100(11):207–208, March 13, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n11/
pdf/scientificamerican03131909-207c.pdf.

Sweet:1909:HMB

[68] F. D. Sweet. Home-made blowpipe. Scientific American, 100(11):208, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n11/pdf/scientificamerican03131909-208.pdf.

VanDerVeer:1909:UTJ

[69] A. R. Van Der Veer. Unscrewing a tight jab top. Scientific American, 100(11):208, March 13, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n11/
pdf/scientificamerican03131909-208a.pdf.

Paulmier:1909:CSC

[70] J. Carlton Paulmier. Construction of a selenium cell. Scientific American, 100(11):208, March 13, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n11/
pdf/scientificamerican03131909-208b.pdf.

Bishop:1909:HMB

[71] Albert F. Bishop. Home-made brass furnace. Scientific American, 100(11):208, March 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n11/pdf/scientificamerican03131909-
208c.pdf.

Anonymous:1909:Cb

[72] Anonymous. Correspondence. Scientific American, 100(12):222, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n12/pdf/scientificamerican03201909-222a.pdf.

Anonymous:1909:RPId 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n12/pdf/scientificamerican03201909-
230.pdf.

Anonymous:1909:NQc

[74] Anonymous. Notes and queries. Scientific American, 100(12):230–231, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n12/pdf/scientificamerican03201909-230a.pdf.

Anonymous:1909:BNBc

[75] Anonymous. Books: New books, etc. Scientific American, 100(12):231, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n12/pdf/scientificamerican03201909-231.pdf.

Anonymous:1909:PNPc

[76] Anonymous. Patent notes: Patents. Scientific American, 100(12):231, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n12/pdf/scientificamerican03201909-231a.pdf.

Anonymous:1909:IId

[77] Anonymous. Index of inventions. Scientific American, 100(12):231–235, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n12/pdf/scientificamerican03201909-231b.pdf.

Anonymous:1909:MNG

[78] Anonymous. The menace of the noiseless gun. Scientific American, 100(12):218, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n12/pdf/scientificamerican03201909-
218.pdf.

Anonymous:1909:OAS

[79] Anonymous. Opportunity for American steel construction at Messina. Scientific American, 100(12):218, March 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n12/
pdf/scientificamerican03201909-218a.pdf.

Anonymous:1909:SSa

[80] Anonymous. The subway situation. Scientific American, 100(12):218, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n12/pdf/scientificamerican03201909-218b.pdf.

Anonymous:1909:LH

[81] Anonymous. Limitations of the hydroplane. Scientific American, 100(12):218, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n12/pdf/scientificamerican03201909-
218c.pdf.

Anonymous:1909:AAS

[82] Anonymous. Aviation in America and the Scientific American Trophy. Scientific American, 100(12):218, March 20, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n12/pdf/scientificamerican03201909-218d.pdf.

Anonymous:1909:Ee

[83] Anonymous. Engineering. Scientific American, 100(12):219, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n12/pdf/scientificamerican03201909-219.pdf.

Anonymous:1909:Ef

[84] Anonymous. Electricity. Scientific American, 100(12):219, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n12/pdf/scientificamerican03201909-219a.pdf.

Anonymous:1909:Sc

[85] Anonymous. Science. Scientific American, 100(12):219, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n12/pdf/scientificamerican03201909-219b.pdf.

Anonymous:1909:IFB

[86] Anonymous. Improved forms of the buggy-type automobile. Scientific American, 100(12):220–221, March 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n12/
pdf/scientificamerican03201909-220.pdf.

Anonymous:1909:STLa

[87] Anonymous. The Schneider torpedo-launching station. Scientific American, 100(12):222, March 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n12/
pdf/scientificamerican03201909-222.pdf.

Anonymous:1909:ASDa

[88] Anonymous. Accuracy in scientific diction. Scientific American, 100(12):222, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n12/pdf/scientificamerican03201909-
222b.pdf.

Anonymous:1909:TWS

[89] Anonymous. From the track-walker’s standpoint. Scientific American, 100(12):222–223, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n12/pdf/scientificamerican03201909-
222c.pdf.

Anonymous:1909:LTI

[90] Anonymous. The locust tree: An important field for investigation. Scientific American, 100(12):223, March 20, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n12/pdf/scientificamerican03201909-223.pdf.

Haywood:1909:WSP

[91] Charles B. Haywood. What is the selden patent? Scientific American, 100(12):223–225, March 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n12/
pdf/scientificamerican03201909-223a.pdf.

Anonymous:1909:FAI

[92] Anonymous. First aid to those injured by electricity. Scientific American, 100(12):225, March 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n12/
pdf/scientificamerican03201909-225.pdf.

Anonymous:1909:PTN

[93] Anonymous. Progress of the 20,000-ton “North Dakota”. Scientific American, 100(12):226, March 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n12/
pdf/scientificamerican03201909-226.pdf.

Anonymous:1909:HMWb

[94] Anonymous. Handy man’s workshop. Scientific American, 100 (12):227–228, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n12/pdf/scientificamerican03201909-
227.pdf.

Morgan:1909:IEE

[95] Alfred P. Morgan. An interesting experiment and its explanation. Scientific American, 100(12):228, March 20, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n12/pdf/scientificamerican03201909-228.pdf.

Brouillet:1909:HMS

[96] J. O. Brouillet. Home-made spring winder. Scientific American, 100(12):228, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n12/pdf/scientificamerican03201909-
228a.pdf.

Bayley:1909:FMB

[97] I. G. Bayley. Forging a masthead or boom ring. Scientific American, 100(12):228, March 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n12/pdf/scientificamerican03201909-
228b.pdf.

Anonymous:1909:Cc

[98] Anonymous. Correspondence. Scientific American, 100(13):243, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-243.pdf.

Anonymous:1909:RPIe

[99] Anonymous. Recently patented inventions. Scientific American, 100(13):249, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
249.pdf.

Anonymous:1909:NQd

[100] Anonymous. Notes and queries. Scientific American, 100(13):249–251, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-249a.pdf.

Anonymous:1909:BNBd

[101] Anonymous. Books: New books, etc. Scientific American, 100 (13):251–252, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
251.pdf.

Anonymous:1909:PNPd

[102] Anonymous. Patent notes: Patents. Scientific American, 100(13):252, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-252.pdf.

Anonymous:1909:IIe

[103] Anonymous. Index of inventions. Scientific American, 100(13):252–255, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-252a.pdf.

Anonymous:1909:SAFa

[104] Anonymous. The Scientific American Flying Machine Trophy for 1909. Scientific American, 100(13):238, March 27, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n13/pdf/scientificamerican03271909-238.pdf.

Anonymous:1909:HA

[105] Anonymous. The highway and the automobile. Scientific American, 100(13):238, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
238a.pdf.

Anonymous:1909:CFD

[106] Anonymous. Close of the fourth dimension contest. Scientific American, 100(13):238, March 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n13/
pdf/scientificamerican03271909-238b.pdf.

Anonymous:1909:RD

[107] Anonymous. Rainmaking by dynamite. Scientific American, 100 (13):238, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
238c.pdf.

Anonymous:1909:Aa

[108] Anonymous. Aeronautics. Scientific American, 100(13):239, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-239.pdf.

Anonymous:1909:Eg

[109] Anonymous. Electricity. Scientific American, 100(13):239, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-239a.pdf.

Anonymous:1909:Se

[110] Anonymous. Science. Scientific American, 100(13):239, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n13/pdf/scientificamerican03271909-239b.pdf.

Wood:1909:MT

[111] R. W. Wood. The Mercury telescope. Scientific American, 100 (13):240–241, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
240.pdf.

Anonymous:1909:SLS

[112] Anonymous.  Superiority of lock to sea-level canal.  Scien-

tific American, 100(13):241–242, March 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n13/
pdf/scientificamerican03271909-241.pdf.

Anonymous:1909:SAFb

[113] Anonymous. The Scientific American Flying Machine Trophy. Scientific American, 100(13):242, March 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n13/
pdf/scientificamerican03271909-242.pdf.

Anonymous:1909:GIO

[114] Anonymous. Governmental income and outgo. Scientific American, 100(13):243, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
243a.pdf.

Beasley:1909:NDD

[115] Walter L. Beasley. The Navy’s dummy drill gun. Scientific American, 100(13):244, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
244.pdf.

Willey:1909:IAW

[116] Day Allen Willey. The ice of the Arctic waters. Scientific American, 100(13):244–246, March 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n13/
pdf/scientificamerican03271909-244a.pdf.

Welch:1909:TCS

[117] Deshler Welch.  Tobogganing at Caux, Switzerland.  Scien-

tific American, 100(13):246, March 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n13/
pdf/scientificamerican03271909-246.pdf.

Anonymous:1909:PDa

[118] Anonymous. Patent department. Scientific American, 100(13):247, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-247.pdf.

Anonymous:1909:EAa

[119] Anonymous. Engine aliner. Scientific American, 100(13):247, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-247a.pdf.

Anonymous:1909:AHP

[120] Anonymous. Adjustable head for piano hammers. Scientific American, 100(13):247, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
247b.pdf.

Anonymous:1909:ATR

[121] Anonymous. Apparatus for treating rubber. Scientific American, 100(13):247, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
247c.pdf.

Anonymous:1909:VFT

[122] Anonymous. Valve for flush tanks. Scientific American, 100 (13):247–248, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
247d.pdf.

Anonymous:1909:GCP

[123] Anonymous. The German Crown Prince as an inventor. Scientific American, 100(13):248, March 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n13/
pdf/scientificamerican03271909-248.pdf.

Anonymous:1909:IOL

[124] Anonymous. Improved oar lock. Scientific American, 100(13):248, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-248a.pdf.

Anonymous:1909:RPL

[125] Anonymous. Reciprocity in patent laws. Scientific American, 100(13):248, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n13/pdf/scientificamerican03271909-
248b.pdf.

Anonymous:1909:KSC

[126] Anonymous. Kitchen sink cover. Scientific American, 100(13):248, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-248c.pdf.

Anonymous:1909:PMC

[127] Anonymous. Pneumatic milk can. Scientific American, 100(13):248, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-248d.pdf.

Anonymous:1909:OIa

[128] Anonymous. Oddities in inventions. Scientific American, 100(13):248, March 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n13/pdf/scientificamerican03271909-248e.pdf.

Anonymous:1909:Cd

[129] Anonymous. Correspondence. Scientific American, 100(14):263, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-263.pdf.

Russell:1909:HAa

[130] Henry Norris Russell. The heavens in April. Scientific American, 100(14):266, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
266.pdf.

Anonymous:1909:RPIf

[131] Anonymous. Recently patented inventions. Scientific American, 100(14):268, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
268.pdf.

Anonymous:1909:NQe

[132] Anonymous. Notes and queries. Scientific American, 100(14):268, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-268a.pdf.

Anonymous:1909:PNPe

[133] Anonymous. Patent notes: Patents. Scientific American, 100(14):268, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-268b.pdf.

Anonymous:1909:IIf

[134] Anonymous. Index of inventions. Scientific American, 100(14):268–271, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-268c.pdf.

Anonymous:1909:LSA

[135] Anonymous.  Lieut. Shackleton’s Antarctic Expedition.  Sci-

entific American, 100(14):258, April 3, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n14/
pdf/scientificamerican04031909-258.pdf.

Anonymous:1909:TFS

[136] Anonymous. To test full-sized compression members. Scientific American, 100(14):258, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
258a.pdf.

Anonymous:1909:CDC

[137] Anonymous. The so-called “Dreadnought” craze. Scientific American, 100(14):258, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
258b.pdf.

Anonymous:1909:Eh

[138] Anonymous. Engineering. Scientific American, 100(14):259, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-259.pdf.

Anonymous:1909:Ej

[139] Anonymous. Electricity. Scientific American, 100(14):259, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-259a.pdf.

Anonymous:1909:Sf

[140] Anonymous. Science. Scientific American, 100(14):259, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n14/pdf/scientificamerican04031909-259b.pdf.

Hayward:1909:EIC

[141] Charles B. Hayward. An enormous induction coil. Scientific American, 100(14):260–261, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
260.pdf.

Anonymous:1909:FLI

[142] Anonymous. Floating lights of inland waters. Scientific American, 100(14):261–262, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
261.pdf.

Anonymous:1909:ADA

[143] Anonymous. Aerial defense artillery. Scientific American, 100(14):262, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-262.pdf.

Anonymous:1909:EVC

[144] Anonymous. End versus center doors for exit on subway cars. Scientific American, 100(14):263, April 3, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n14/
pdf/scientificamerican04031909-263a.pdf.

Anonymous:1909:SPS

[145] Anonymous. Seaworthiness of passenger steamships. Scientific American, 100(14):263, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
263b.pdf.

Anonymous:1909:DMS

[146] Anonymous. Defects in motor-sleighs. Scientific American, 100(14):263, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-263c.pdf.

Anonymous:1909:ONT

[147] Anonymous. Our new 26,000-ton battleships. Scientific American, 100(14):263–265, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
263d.pdf.

Baltimore:1909:RCM

[148] J. Mayne Baltimore. Regrading a city by means of hydraulic sluicing. Scientific American, 100(14):265, April 3, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n14/pdf/scientificamerican04031909-265.pdf.

Bayley:1909:HMW

[149] I. G. Bayley. Handy man’s workshop. Scientific American, 100(14):267, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n14/pdf/scientificamerican04031909-267.pdf.

Henry:1909:HMF

[150] K. R. Henry. The handy man in the factory. Scientific American, 100(14):267, April 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n14/pdf/scientificamerican04031909-
267a.pdf.

Anonymous:1909:Ce

[151] Anonymous. Correspondence. Scientific American, 100(15):279, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-279.pdf.

Anonymous:1909:RPIg

[152] Anonymous. Recently patented inventions. Scientific American, 100(15):284, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n15/pdf/scientificamerican04101909-
284.pdf.

Anonymous:1909:PNPf

[153] Anonymous. Patent notes: Patents. Scientific American, 100(15):284, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-284a.pdf.

Anonymous:1909:IIg

[154] Anonymous. Index of inventions. Scientific American, 100(15):284–287, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-284b.pdf.

Anonymous:1909:NHR

[155] Anonymous. The new Hudson River steamer “Robert Fulton”. Scientific American, 100(15):274, April 10, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n15/
pdf/scientificamerican04101909-274.pdf.

Anonymous:1909:RSN

[156] Anonymous. The Renaissance of the Spanish Navy. Scientific American, 100(15):274, April 10, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n15/
pdf/scientificamerican04101909-274a.pdf.

Anonymous:1909:CVS

[157] Anonymous. Center versus side door cars. Scientific American, 100(15):274, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n15/pdf/scientificamerican04101909-
274b.pdf.

Anonymous:1909:RFG

[158] Anonymous. Recent flights of Germany’s first Zeppelin military airship. Scientific American, 100(15):274, April 10, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n15/pdf/scientificamerican04101909-274c.pdf.

Anonymous:1909:FDE

[159] Anonymous. The fourth dimension essays. Scientific American, 100(15):274, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n15/pdf/scientificamerican04101909-
274d.pdf.

Anonymous:1909:Ek

[160] Anonymous. Engineering. Scientific American, 100(15):275, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-275.pdf.

Anonymous:1909:El

[161] Anonymous. Electricity. Scientific American, 100(15):275, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-275a.pdf.

Anonymous:1909:Sg

[162] Anonymous. Science. Scientific American, 100(15):275, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n15/pdf/scientificamerican04101909-275b.pdf.

Anonymous:1909:LLA

[163] Anonymous. A life-line air cannon. Scientific American, 100(15):276, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-276.pdf.

Anonymous:1909:FCW

[164] Anonymous. The filtration of city water supply. Scientific American, 100(15):276, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n15/pdf/scientificamerican04101909-
276a.pdf.

Anonymous:1909:DUS

[165] Anonymous. Detecting underground streams with an ear-trumpet. Scientific American, 100(15):277, April 10, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n15/
pdf/scientificamerican04101909-277.pdf.

Sheak:1909:PTP

[166] W. Henry Sheak. Pythons and their prey. Scientific American, 100(15):277–278, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n15/pdf/scientificamerican04101909-
277a.pdf.

Anonymous:1909:CPA

[167] Anonymous. The commercial possibilities of the aeroplane. Scientific American, 100(15):279, April 10, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n15/
pdf/scientificamerican04101909-279a.pdf.

Anonymous:1909:PMa

[168] Anonymous. A problem in mechanics. Scientific American, 100(15):279, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-279b.pdf.

Anonymous:1909:SI

[169] Anonymous. Stereoscopic illusion. Scientific American, 100(15):279, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-279c.pdf.

Anonymous:1909:NTSa

[170] Anonymous. New type of sub-surface torpedo boat. Scientific American, 100(15):280, April 10, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n15/
pdf/scientificamerican04101909-280.pdf.

Anonymous:1909:QMB

[171] Anonymous. The Queensboro and Manhattan Bridges across the East River. Scientific American, 100(15):281–282, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n15/pdf/scientificamerican04101909-281.pdf.

Gradenwitz:1909:LPB

[172] Alfred Gradenwitz. A life preserver for balloonists. Scientific American, 100(15):282, April 10, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n15/
pdf/scientificamerican04101909-282.pdf.

Anonymous:1909:ANb

[173] Anonymous. Aeronautical notes. Scientific American, 100(15):282, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-282a.pdf.

Anonymous:1909:PDb

[174] Anonymous. Patent department. Scientific American, 100(15):283, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-283.pdf.

Anonymous:1909:NFD

[175] Anonymous. A novel frying device. Scientific American, 100(15):283, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-283a.pdf.

Anonymous:1909:IRT

[176] Anonymous. An improved rat trap. Scientific American, 100(15):283, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-283b.pdf.

Anonymous:1909:NRS

[177] Anonymous. A new railroad spike. Scientific American, 100(15):283, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-283c.pdf.

Anonymous:1909:SDS

[178] Anonymous. A simple drawer stop. Scientific American, 100(15):283, April 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n15/pdf/scientificamerican04101909-283d.pdf.

Anonymous:1909:Cg

[179] Anonymous. Correspondence. Scientific American, 100(16):295, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-295.pdf.

Anonymous:1909:RPIh

[180] Anonymous. Recently patented inventions. Scientific American, 100(16):301–302, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n16/pdf/scientificamerican04171909-
301.pdf.

Anonymous:1909:NQf

[181] Anonymous. Notes and queries. Scientific American, 100(16):302, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-302.pdf.

Anonymous:1909:PNPg

[182] Anonymous. Patent notes: Patents. Scientific American, 100(16):304, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-304.pdf.

Anonymous:1909:IIh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-304a.pdf.

Anonymous:1909:FAF

[184] Anonymous. The Fulton airship flight contest. Scientific American, 100(16):290, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n16/pdf/scientificamerican04171909-
290.pdf.

Anonymous:1909:BRA

[185] Anonymous. Brick roads for the automobile. Scientific American, 100(16):290, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n16/pdf/scientificamerican04171909-
290a.pdf.

Anonymous:1909:SV

[186] Anonymous. Steamship vibration. Scientific American, 100(16):290, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-290b.pdf.

Anonymous:1909:WEF

[187] Anonymous. Why not the electric fire engine? Scientific American, 100(16):290, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n16/pdf/scientificamerican04171909-
290c.pdf.

Anonymous:1909:Em

[188] Anonymous. Engineering. Scientific American, 100(16):291, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-291.pdf.

Anonymous:1909:En

[189] Anonymous. Electricity. Scientific American, 100(16):291, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-291a.pdf.

Anonymous:1909:Sh

[190] Anonymous. Science. Scientific American, 100(16):291, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n16/pdf/scientificamerican04171909-291b.pdf.

Anonymous:1909:TBNa

[191] Anonymous. The turbines of the battleship “North Dakota”. Scientific American, 100(16):292, April 17, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n16/
pdf/scientificamerican04171909-292.pdf.

Collins:1909:EA

[192] Percy Collins. The eyes of animals. Scientific American, 100 (16):293–294, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n16/pdf/scientificamerican04171909-
293.pdf.

Anonymous:1909:RE

[193] Anonymous. Rainmaking by explosion. Scientific American, 100 (16):295, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n16/pdf/scientificamerican04171909-
295a.pdf.

Anonymous:1909:FOI

[194] Anonymous. Formulæ for obtaining integral sides of right-angled triangles. Scientific American, 100(16):295, April 17, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n16/pdf/scientificamerican04171909-295b.pdf.

Raymond:1909:AYE

[195] William H. Raymond. The Alaska–Yukon Exposition. Scientific American, 100(16):296–298, April 17, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n16/
pdf/scientificamerican04171909-296.pdf.

Anonymous:1909:EGa

[196] Anonymous. An experiment with a glider. Scientific American, 100(16):298, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n16/pdf/scientificamerican04171909-
298.pdf.

Anonymous:1909:HMWc

[197] Anonymous. Handy man’s workshop. Scientific American, 100(16):299, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-299.pdf.

Ragan:1909:CFE

[198] K. Ragan. Chemical flasks from electric light globes. Scientific American, 100(16):299–300, April 17, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n16/
pdf/scientificamerican04171909-299a.pdf.

Pabody:1909:R

[199] E. F. Pabody, Jr. Rotagons. Scientific American, 100(16):300, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-300.pdf.

Heuser:1909:NHM

[200] R. V. Heuser. Novel heat motor. Scientific American, 100(16):300, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-300a.pdf.

Lurcott:1909:AJH

[201] Charles Lurcott. Another joint for handy man’s work bench. Scientific American, 100(16):300, April 17, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n16/
pdf/scientificamerican04171909-300b.pdf.

Naylor:1909:RLW

[202] G. W. Naylor. A remedy for loose wood screws. Scientific American, 100(16):300, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n16/pdf/scientificamerican04171909-
300c.pdf.

Chapman:1909:HC

[203] H. D. Chapman. Hardening copper. Scientific American, 100(16):300, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-300d.pdf.

Anonymous:1909:NBE

[204] Anonymous. New books, etc. Scientific American, 100(16):303–304, April 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n16/pdf/scientificamerican04171909-303.pdf.

Anonymous:1909:Ch

[205] Anonymous. Correspondence. Scientific American, 100(17):315, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n17/pdf/scientificamerican04241909-315.pdf.

Anonymous:1909:RPIi

[206] Anonymous. Recently patented inventions. Scientific American, 100(17):321, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
321.pdf.

Anonymous:1909:NQg

[207] Anonymous. Notes and queries. Scientific American, 100(17):321–322, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n17/pdf/scientificamerican04241909-321a.pdf.

Anonymous:1909:BNBe

[208] Anonymous. Books: New books, etc. Scientific American, 100(17):322, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n17/pdf/scientificamerican04241909-322.pdf.

Anonymous:1909:PNPh

[209] Anonymous. Patent notes: Patents. Scientific American, 100(17):322, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n17/pdf/scientificamerican04241909-322a.pdf.

Anonymous:1909:IIi

[210] Anonymous. Index of inventions. Scientific American, 100(17):322–327, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n17/pdf/scientificamerican04241909-322b.pdf.

Anonymous:1909:DC

[211] Anonymous. The “Dreadnought” competition. Scientific American, 100(17):310, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
310.pdf.

Anonymous:1909:ELT

[212] Anonymous. Economic loss through the mosquito. Scientific American, 100(17):310, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
310a.pdf.

Anonymous:1909:HFC

[213] Anonymous. The Hudson–Fulton Celebration. Scientific American, 100(17):310, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
310b.pdf.

Anonymous:1909:WWF

[214] Anonymous. Wilbur Wright’s flights in Italy. Scientific American, 100(17):310, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
310c.pdf.

Anonymous:1909:Eo

[215] Anonymous. Engineering. Scientific American, 100(17):311, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n17/pdf/scientificamerican04241909-311.pdf.

Anonymous:1909:Ep

[216] Anonymous. Electricity. Scientific American, 100(17):311, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n17/pdf/scientificamerican04241909-311a.pdf.

Anonymous:1909:Sj

[217] Anonymous. Science. Scientific American, 100(17):311, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n17/pdf/scientificamerican04241909-311b.pdf.

Hill:1909:DFO

[218] William L. Hill. Demolition of the facade of the old B. & P. Rr. Terminal, Boston. Scientific American, 100(17):312, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n17/pdf/scientificamerican04241909-312.pdf.

Springer:1909:EIC

[219] J. F. S. Springer. Experiments in ingot casting. Scientific American, 100(17):313–314, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
313.pdf.

Anonymous:1909:GIU

[220] Anonymous. Glass industry of the United States. Scientific American, 100(17):314, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
314.pdf.

Anonymous:1909:PFCb

[221] Anonymous. The pearl fishery of Ceylon. Scientific American, 100(17):316–318, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
316.pdf.

Anonymous:1909:HHM

[222] Anonymous. A house-to-house mirror system. Scientific American, 100(17):318, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
318.pdf.

Anonymous:1909:HMWd

[223] Anonymous. Handy man’s workshop. Scientific American, 100(17):319, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n17/pdf/scientificamerican04241909-319.pdf.

Bayley:1909:TW

[224] I. G. Bayley. Tools for the workshop. — I. Scientific American, 100(17):320, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
320.pdf.

Michaud:1909:MTA

[225] Gustave Michaud. Milk testing without apparatus. Scientific American, 100(17):320, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
320a.pdf.

Anonymous:1909:SIM

[226] Anonymous. Some interesting models. Scientific American, 100 (17):320, April 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n17/pdf/scientificamerican04241909-
320b.pdf.

Anonymous:1909:Ci

[227] Anonymous. Correspondence. Scientific American, 100(18):335, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n18/pdf/scientificamerican05011909-335.pdf.

Russell:1909:HMa

[228] Henry Norris Russell. The heavens in May. Scientific American, 100(18):338, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n18/pdf/scientificamerican05011909-
338.pdf.

Anonymous:1909:PNPi

[229] Anonymous. Patent notes: Patents. Scientific American, 100(18):340, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n18/pdf/scientificamerican05011909-340.pdf.

Anonymous:1909:IIj

[230] Anonymous. Index of inventions. Scientific American, 100(18):340–343, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n18/pdf/scientificamerican05011909-340a.pdf.

Anonymous:1909:TAG

[231] Anonymous. The truth about the German “Dreadnoughts”. Scientific American, 100(18):330, May 1, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n18/
pdf/scientificamerican05011909-330.pdf.

Anonymous:1909:HFS

[232] Anonymous. “house fly” should be called “typhoid fly”. Scientific American, 100(18):330, May 1, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n18/
pdf/scientificamerican05011909-330a.pdf.

Anonymous:1909:RNE

[233] Anonymous. Reducing the noise of elevated roads. Scientific American, 100(18):330, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n18/pdf/scientificamerican05011909-
330b.pdf.

Anonymous:1909:Eq

[234] Anonymous. Engineering. Scientific American, 100(18):331, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n18/pdf/scientificamerican05011909-331.pdf.

Anonymous:1909:Er

[235] Anonymous. Electricity. Scientific American, 100(18):331, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n18/pdf/scientificamerican05011909-331a.pdf.

Anonymous:1909:Sk

[236] Anonymous. Science. Scientific American, 100(18):331, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n18/pdf/scientificamerican05011909-331b.pdf.

Anonymous:1909:Ab

[237] Anonymous. The amphibicycle. Scientific American, 100(18):332, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n18/pdf/scientificamerican05011909-332.pdf.

Anonymous:1909:NCS

[238] Anonymous. A new class of submarines for the United States Navy. Scientific American, 100(18):332, May 1, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n18/pdf/scientificamerican05011909-332a.pdf.

Anonymous:1909:NCF

[239] Anonymous. The new Croton Falls Reservoir. Scientific American, 100(18):333–334, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n18/pdf/scientificamerican05011909-
333.pdf.

Anonymous:1909:SWV

[240] Anonymous. Submarine wrecking vessel. Scientific American, 100 (18):334, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n18/pdf/scientificamerican05011909-
334.pdf.

Anonymous:1909:SSb

[241] Anonymous. Steering of ships. Scientific American, 100(18):335, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n18/pdf/scientificamerican05011909-335a.pdf.

Collins:1909:TSB

[242] Percy Collins. Tricks with soap bubbles. Scientific American, 100(18):336–337, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n18/pdf/scientificamerican05011909-
336.pdf.

Perkins:1909:GIJ

[243] F. C. Perkins. The Great Ice Jam at Niagara. Scientific American, 100(18):339, May 1, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n18/pdf/scientificamerican05011909-
339.pdf.

Anonymous:1909:Cj

[244] Anonymous. Correspondence. Scientific American, 100(19):351, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-351.pdf.

Anonymous:1909:RPIj

[245] Anonymous. Recently patented inventions. Scientific American, 100(19):357–358, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n19/pdf/scientificamerican05081909-
357.pdf.

Anonymous:1909:NQh

[246] Anonymous. Notes and queries. Scientific American, 100(19):358–359, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-358.pdf.

Anonymous:1909:BNBf

[247] Anonymous. Books: New books, etc. Scientific American, 100(19):359, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-359.pdf.

Anonymous:1909:PNPj

[248] Anonymous. Patent notes: Patents. Scientific American, 100(19):359, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-359a.pdf.

Anonymous:1909:IIk

[249] Anonymous. Index of inventions. Scientific American, 100(19):359–363, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-359b.pdf.

Anonymous:1909:DBS

[250] Anonymous. Distorted bridge stories. Scientific American, 100(19):346, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-346.pdf.

Anonymous:1909:SMa

[251] Anonymous. Signaling to Mars. Scientific American, 100(19):346, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-346a.pdf.

Anonymous:1909:DCV

[252] Anonymous. Direct-current versus single-phase traction. Scientific American, 100(19):346, May 8, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n19/
pdf/scientificamerican05081909-346b.pdf.

Anonymous:1909:Ac

[253] Anonymous. Aeronautics. Scientific American, 100(19):347, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-347.pdf.

Anonymous:1909:Es

[254] Anonymous. Electricity. Scientific American, 100(19):347, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n19/pdf/scientificamerican05081909-347a.pdf.

Anonymous:1909:Sl

[255] Anonymous. Science. Scientific American, 100(19):347, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n19/pdf/scientificamerican05081909-347b.pdf.

Anonymous:1909:TBS

[256] Anonymous. A tool for boring square holes. Scientific American, 100(19):348, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n19/pdf/scientificamerican05081909-
348.pdf.

Dunlap:1909:DNI

[257] Orrin E. Dunlap. Dynamiting the Niagara ice jam. Scientific American, 100(19):349–350, May 8, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n19/
pdf/scientificamerican05081909-349.pdf.

Anonymous:1909:WDMa

[258] Anonymous. Why drunken men always see double. Scientific American, 100(19):350, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n19/pdf/scientificamerican05081909-
350.pdf.

Anonymous:1909:DH

[259] Anonymous. The discoverers of the Hudson. Scientific American, 100(19):351, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n19/pdf/scientificamerican05081909-
351a.pdf.

Anonymous:1909:WDD

[260] Anonymous. Wells and Darwin, and the Doctrine of Evolution. Scientific American, 100(19):351, May 8, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n19/
pdf/scientificamerican05081909-351b.pdf.

Anonymous:1909:RSF

[261] Anonymous. Recent successful French monoplanes. Scientific American, 100(19):352, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n19/pdf/scientificamerican05081909-
352.pdf.

Anonymous:1909:WMB

[262] Anonymous. The winning motor boats at Monaco. Scientific American, 100(19):352, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n19/pdf/scientificamerican05081909-
352a.pdf.

Anonymous:1909:EAM

[263] Anonymous. The electric aerial mono-rail up the Wetterhorn. Scientific American, 100(19):353–354, May 8, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n19/
pdf/scientificamerican05081909-353.pdf.

Anonymous:1909:PDc

[264] Anonymous. Patent department. Scientific American, 100(19):355, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-355.pdf.

Anonymous:1909:BTC

[265] Anonymous. Boiler-tube cleaner. Scientific American, 100(19):355, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-355a.pdf.

Anonymous:1909:DAM

[266] Anonymous. Delivery apparatus for mail cars. Scientific American, 100(19):355, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n19/pdf/scientificamerican05081909-
355b.pdf.

Anonymous:1909:BPR

[267] Anonymous. Back pressure reducer. Scientific American, 100 (19):355–356, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n19/pdf/scientificamerican05081909-
355c.pdf.

Anonymous:1909:OIb

[268] Anonymous. Oddities in inventions. Scientific American, 100(19):356, May 8, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n19/pdf/scientificamerican05081909-356.pdf.

Anonymous:1909:Ck

[269] Anonymous. Correspondence. Scientific American, 100(2):23, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-23.pdf.

Anonymous:1909:RPIk

[270] Anonymous. Recently patented inventions. Scientific American, 100(2):30, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
30.pdf.

Anonymous:1909:NQi

[271] Anonymous. Notes and queries. Scientific American, 100(2):30–32, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-30a.pdf.

Anonymous:1909:BNBg

[272] Anonymous. Books: New books, etc. Scientific American, 100(2):32, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-32.pdf.

Anonymous:1909:PNPk

[273] Anonymous. Patent notes: Patents. Scientific American, 100(2):32, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-32a.pdf.

Anonymous:1909:IIl

[274] Anonymous. Index of inventions. Scientific American, 100(2):32–35, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-32b.pdf.

Anonymous:1909:EDa

[275] Anonymous. The earthquake disaster. Scientific American, 100(2):18, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-18.pdf.

Anonymous:1909:WVF

[276] Anonymous. The Wright and Voisin (Farman) flying machines compared. Scientific American, 100(2):18, January 9, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n2/pdf/scientificamerican01091909-18a.pdf.

Anonymous:1909:Et

[277] Anonymous. Engineering. Scientific American, 100(2):19, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-19.pdf.

Anonymous:1909:Eu

[278] Anonymous. Electricity. Scientific American, 100(2):19, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-19a.pdf.

Anonymous:1909:Sn

[279] Anonymous. Science. Scientific American, 100(2):19, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n2/pdf/scientificamerican01091909-19b.pdf.

Anonymous:1909:NRP

[280] Anonymous. A new rifle-propelled grenade. Scientific American, 100(2):20, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
20.pdf.

Anonymous:1909:TBM

[281] Anonymous. A tunnel-boring machine. Scientific American, 100(2):20, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-20a.pdf.

Wade:1909:RIM

[282] Herbert T. Wade. The Rockefeller Institute for Medical Research. Scientific American, 100(2):21–22, January 9, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n2/pdf/scientificamerican01091909-21.pdf.

Anonymous:1909:NCM

[283] Anonymous. A novel camera for making moving pictures. Scientific American, 100(2):22, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
22.pdf.

Anonymous:1909:GRP

[284] Anonymous. Good roads and the price of food. Scientific American, 100(2):23, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
23a.pdf.

Anonymous:1909:NFP

[285] Anonymous. A novel filtration plant. Scientific American, 100(2):23, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-23b.pdf.

Anonymous:1909:CS

[286] Anonymous. Cleaning stone. Scientific American, 100(2):23, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-23c.pdf.

Anonymous:1909:MHO

[287] Anonymous. A multi-hull ocean steamship. Scientific American, 100(2):23, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
23d.pdf.

Anonymous:1909:KSW

[288] Anonymous. The Krupp Steel Works of to-day. Scientific American, 100(2):24–25, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
24.pdf.

Mitchell:1909:PBM

[289] S. A. Mitchell. The peculiar behavior of Morehouse’s Comet. Scientific American, 100(2):26, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
26.pdf.

Anonymous:1909:HMWe

[290] Anonymous. Handy man’s workshop. Scientific American, 100(2):27, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-27.pdf.

Lord:1909:HBS

[291] Frederick K. Lord. How to build a scooter. Scientific American, 100(2):28–29, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
28.pdf.

Anonymous:1909:CGa

[292] Anonymous. A cement grindstone. Scientific American, 100(2):29, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-29.pdf.

C:1909:FDH

[293] W. J. C. Further details of the home-made vacuum cleaner. Scientific American, 100(2):29, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
29a.pdf.

Handy:1909:CS

[294] L. Gessford Handy. Coasting skates. Scientific American, 100(2):29, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n2/pdf/scientificamerican01091909-29b.pdf.

Mencken:1909:RRR

[295] August Mencken. Reducing the range of a rifle. Scientific American, 100(2):29, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
29c.pdf.

Anonymous:1909:NIH

[296] Anonymous. The next issue of handy man’s workshop. Scientific American, 100(2):29, January 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n2/pdf/scientificamerican01091909-
29d.pdf.

Anonymous:1909:Cl

[297] Anonymous. Correspondence. Scientific American, 100(20):371, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-371.pdf.

Anonymous:1909:RPIl

[298] Anonymous. Recently patented inventions. Scientific American, 100(20):377, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
377.pdf.

Anonymous:1909:NQj

[299] Anonymous. Notes and queries. Scientific American, 100(20):377–378, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-377a.pdf.

Anonymous:1909:BNBh

[300] Anonymous. Books: New books, etc. Scientific American, 100 (20):378–379, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
378.pdf.

Anonymous:1909:PNPl

[301] Anonymous. Patent notes: Patents. Scientific American, 100(20):379, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-379.pdf.

Anonymous:1909:IIm

[302] Anonymous. Index of inventions. Scientific American, 100(20):379–383, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-379a.pdf.

Anonymous:1909:LDT

[303] Anonymous. A long-delayed tribute. Scientific American, 100(20):366, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-366.pdf.

Anonymous:1909:OSR

[304] Anonymous. Our stupendous railway system. Scientific American, 100(20):366, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
366a.pdf.

Anonymous:1909:CTS

[305] Anonymous. Comparative tests of the Scout cruisers. Scientific American, 100(20):366, May 15, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n20/
pdf/scientificamerican05151909-366b.pdf.

Anonymous:1909:RWB

[306] Anonymous. Return of the Wright Brothers. Scientific American, 100(20):366, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
366c.pdf.

Anonymous:1909:Ev

[307] Anonymous. Engineering. Scientific American, 100(20):367, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-367.pdf.

Anonymous:1909:Ew

[308] Anonymous. Electricity. Scientific American, 100(20):367, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-367a.pdf.

Anonymous:1909:So

[309] Anonymous. Science. Scientific American, 100(20):367, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n20/pdf/scientificamerican05151909-367b.pdf.

Anonymous:1909:CLFa

[310] Anonymous. Cork-lined fabric — a textile novelty. Scientific American, 100(20):368, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
368.pdf.

Anonymous:1909:NCP

[311] Anonymous. A new color-photography process. Scientific American, 100(20):368, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
368a.pdf.

Anonymous:1909:MPL

[312] Anonymous. The most powerful locomotives ever built. Scientific American, 100(20):369–370, May 15, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n20/
pdf/scientificamerican05151909-369.pdf.

Honey:1909:LSE

[313] Frederic R. Honey. The lunar and solar eclipses in June, 1909. Scientific American, 100(20):370, May 15, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n20/
pdf/scientificamerican05151909-370.pdf.

Anonymous:1909:NGP

[314] Anonymous. The new German photography. Scientific American, 100(20):371, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
371a.pdf.

Boyer:1909:PM

[315] Jacques Boyer. The palace of mirages. Scientific American, 100 (20):371–372, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
371b.pdf.

Homolle:1909:ED

[316] M. Homolle. The excavations at Delphi. Scientific American, 100 (20):373–374, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
373.pdf.

Anonymous:1909:GI

[317] Anonymous. Goblets of ice. Scientific American, 100(20):374, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-374.pdf.

Anonymous:1909:HMWf

[318] Anonymous. Handy man’s workshop. Scientific American, 100(20):375, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-375.pdf.

Carter:1909:HG

[319] O. D. Carter. Holder for grinders. Scientific American, 100(20):375, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-375a.pdf.

Bayley:1909:TWI

[320] I. G. Bayley. Tools for the workshop — II. Scientific American, 100(20):375, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
375b.pdf.

Morgan:1909:TD

[321] Alfred P. Morgan. A tantalum detector. Scientific American, 100 (20):375–376, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
375c.pdf.

Bergstrom:1909:TC

[322] J. A. Bergstrom. Thread cutting without a die. Scientific American, 100(20):376, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
376.pdf.

MKenzie:1909:SMC

[323] W. C. M’Kenzie. A simple method of constructing a handle. Scientific American, 100(20):376, May 15, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n20/
pdf/scientificamerican05151909-376a.pdf.

Laker:1909:MPB

[324] John W. E. Laker. Method of patching a boiler. Scientific American, 100(20):376, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
376b.pdf.

Parker:1909:PS

[325] Joseph N. Parker. Pencil sharpener. Scientific American, 100(20):376, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-376c.pdf.

Hagarty:1909:CFW

[326] Arthur E. Hagarty. Coppering flowers without wax. Scientific American, 100(20):376, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
376d.pdf.

C:1909:BPW

[327] W. J. C. Blue-print washer. Scientific American, 100(20):376, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n20/pdf/scientificamerican05151909-376e.pdf.

Anonymous:1909:PNC

[328] Anonymous. Putting on a new clutch leather. Scientific American, 100(20):376, May 15, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n20/pdf/scientificamerican05151909-
376f.pdf.

Anonymous:1909:Cm

[329] Anonymous. Correspondence. Scientific American, 100(21):391, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n21/pdf/scientificamerican05221909-391.pdf.

Anonymous:1909:RPIm

[330] Anonymous. Recently patented inventions. Scientific American, 100(21):396, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n21/pdf/scientificamerican05221909-
396.pdf.

Anonymous:1909:PNPm

[331] Anonymous. Patent notes: Patents. Scientific American, 100(21):396, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n21/pdf/scientificamerican05221909-396a.pdf.

Anonymous:1909:IIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n21/pdf/scientificamerican05221909-396b.pdf.

Anonymous:1909:TCC

[333] Anonymous. The turbine the complement of the compound engine. Scientific American, 100(21):386, May 22, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n21/pdf/scientificamerican05221909-386.pdf.

Anonymous:1909:FCL

[334] Anonymous. Futile criticisms of the lock canal. Scientific American, 100(21):386, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n21/pdf/scientificamerican05221909-
386a.pdf.

Anonymous:1909:CBA

[335] Anonymous. Conditions of the battleships after the long cruise. Scientific American, 100(21):386, May 22, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n21/
pdf/scientificamerican05221909-386b.pdf.

Anonymous:1909:OLD

[336] Anonymous. Ornamentation in lake-dwellers’ habitations. Scientific American, 100(21):386, May 22, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n21/
pdf/scientificamerican05221909-386c.pdf.

Anonymous:1909:Ex

[337] Anonymous. Engineering. Scientific American, 100(21):387, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n21/pdf/scientificamerican05221909-387.pdf.

Anonymous:1909:Ey

[338] Anonymous. Electricity. Scientific American, 100(21):387, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n21/pdf/scientificamerican05221909-387a.pdf.

Anonymous:1909:Ad

[339] Anonymous. Aeronautics. Scientific American, 100(21):387, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n21/pdf/scientificamerican05221909-387b.pdf.

Hovey:1909:KDM

[340] Horace C. Hovey. Kaemper’s discoveries in the Mammoth Cave. Scientific American, 100(21):388, May 22, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n21/
pdf/scientificamerican05221909-388.pdf.

Anonymous:1909:WMS

[341] Anonymous. A wonderful model of the Strasburg clock. Scientific American, 100(21):389–390, May 22, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n21/
pdf/scientificamerican05221909-389.pdf.

Winslow:1909:FAA

[342] B. R. Winslow. Feeding the American Army. Scientific American, 100(21):390–391, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n21/pdf/scientificamerican05221909-
390.pdf.

Anonymous:1909:RPP

[343] Anonymous. A rack and pinion problem. Scientific American, 100(21):391, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n21/pdf/scientificamerican05221909-
391a.pdf.

Anonymous:1909:SCF

[344] Anonymous. Sulphate of copper as a fungicide. Scientific American, 100(21):391, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n21/pdf/scientificamerican05221909-
391b.pdf.

Anonymous:1909:NDT

[345] Anonymous. The “North Dakota’s” turbines. Scientific American, 100(21):391, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n21/pdf/scientificamerican05221909-
391c.pdf.

Anonymous:1909:CJJa

[346] Anonymous. Col. J. J. Astor’s vibratory disintegrator. Scientific American, 100(21):392, May 22, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n21/
pdf/scientificamerican05221909-392.pdf.

Phillips:1909:STM

[347] Rowland Ashford Phillips. The Sigafoos tunnel machine. Scientific American, 100(21):393–394, May 22, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n21/
pdf/scientificamerican05221909-393.pdf.

Anonymous:1909:ITSa

[348] Anonymous. Injury to the turbines of the Scout “Salem”. Scientific American, 100(21):394, May 22, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n21/
pdf/scientificamerican05221909-394.pdf.

Holder:1909:FG

[349] Charles F. Holder. The Flying Gurnards. Scientific American, 100(21):395, May 22, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n21/pdf/scientificamerican05221909-
395.pdf.

Petrunkevitch:1909:LPS

[350] Alexander Petrunkevitch. Latrodectus — the poisonous spider. Scientific American, 100(21):395, May 22, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n21/
pdf/scientificamerican05221909-395a.pdf.

Anonymous:1909:Co

[351] Anonymous. Correspondence. Scientific American, 100(22):407, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-407.pdf.

Russell:1909:HJb

[352] Henry Norris Russell. The heavens in June. Scientific American, 100(22):410, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n22/pdf/scientificamerican05291909-
410.pdf.

Anonymous:1909:RPIn

[353] Anonymous. Recently patented inventions. Scientific American, 100(22):412, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n22/pdf/scientificamerican05291909-
412.pdf.

Anonymous:1909:PNPn

[354] Anonymous. Patent notes: Patents. Scientific American, 100(22):412, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-412a.pdf.

Anonymous:1909:IIo

[355] Anonymous. Index of inventions. Scientific American, 100(22):412–415, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-412b.pdf.

Anonymous:1909:KP

[356] Anonymous. To keep the peace. Scientific American, 100(22):402, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-402.pdf.

Anonymous:1909:SAT

[357] Anonymous. Station announcers on trains. Scientific American, 100(22):402, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n22/pdf/scientificamerican05291909-
402a.pdf.

Anonymous:1909:CSP

[358] Anonymous. Comparison of steam and producer-gas marine plants. Scientific American, 100(22):402, May 29, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n22/
pdf/scientificamerican05291909-402b.pdf.

Anonymous:1909:DAD

[359] Anonymous. Defense against dirigibles. Scientific American, 100 (22):402, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n22/pdf/scientificamerican05291909-
402c.pdf.

Anonymous:1909:Ez

[360] Anonymous. Engineering. Scientific American, 100(22):403, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-403.pdf.

Anonymous:1909:Ae

[361] Anonymous. Aeronautics. Scientific American, 100(22):403, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-403a.pdf.

Anonymous:1909:Sp

[362] Anonymous. Science. Scientific American, 100(22):403, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n22/pdf/scientificamerican05291909-403b.pdf.

Anonymous:1909:NKB

[363] Anonymous. The new Knippels Bridge at Copenhagen. Scientific American, 100(22):404, May 29, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n22/
pdf/scientificamerican05291909-404.pdf.

Springer:1909:FCS

[364] J. F. S. Springer. Fluid compression of steel ingots. Scientific American, 100(22):405–406, May 29, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n22/
pdf/scientificamerican05291909-405.pdf.

Anonymous:1909:NUS

[365] Anonymous. The new United States battleship “Mississippi”. Scientific American, 100(22):406, May 29, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n22/
pdf/scientificamerican05291909-406.pdf.

Anonymous:1909:SWC

[366] Anonymous. Slipping of wheels on curves. Scientific American, 100(22):407, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n22/pdf/scientificamerican05291909-
407a.pdf.

Anonymous:1909:SML

[367] Anonymous. Some more lunar superstitions. Scientific American, 100(22):407, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n22/pdf/scientificamerican05291909-
407b.pdf.

Anonymous:1909:SMMa

[368] Anonymous. Signaling to Mars with mirrors. Scientific American, 100(22):407, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n22/pdf/scientificamerican05291909-
407c.pdf.

Boyer:1909:MFP

[369] Jacques Boyer. The manufacture of French postage stamps. Scientific American, 100(22):407–409, May 29, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n22/
pdf/scientificamerican05291909-407d.pdf.

Anonymous:1909:PDd

[370] Anonymous. Patent department. Scientific American, 100(22):411, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-411.pdf.

Anonymous:1909:CFM

[371] Anonymous. Can-filling machine. Scientific American, 100(22):411, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-411a.pdf.

Anonymous:1909:PPM

[372] Anonymous. Poultry-picking machine. Scientific American, 100 (22):411, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n22/pdf/scientificamerican05291909-
411b.pdf.

Anonymous:1909:SCBa

[373] Anonymous. Self-cleaning brush. Scientific American, 100(22):411, May 29, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n22/pdf/scientificamerican05291909-411c.pdf.

Anonymous:1909:Cp

[374] Anonymous. Correspondence. Scientific American, 100(23):423, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n23/pdf/scientificamerican06051909-423.pdf.

Anonymous:1909:RPIo

[375] Anonymous. Recently patented inventions. Scientific American, 100(23):429–430, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
429.pdf.

Anonymous:1909:NQk

[376] Anonymous. Notes and queries. Scientific American, 100(23):430, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n23/pdf/scientificamerican06051909-430.pdf.

Anonymous:1909:BNBi

[377] Anonymous. Books: New books, etc. Scientific American, 100 (23):430–431, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
430a.pdf.

Anonymous:1909:IIp

[378] Anonymous. Index of inventions. Scientific American, 100(23):431–435, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n23/pdf/scientificamerican06051909-431.pdf.

Anonymous:1909:PNPo

[379] Anonymous. Patent notes: Patent. Scientific American, 100(23):431, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n23/pdf/scientificamerican06051909-431a.pdf.

Anonymous:1909:WAB

[380] Anonymous. A word to aeroplane builders. Scientific American, 100(23):418, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
418.pdf.

Anonymous:1909:NPa

[381] Anonymous. Naval programme for 1911. Scientific American, 100(23):418, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
418a.pdf.

Anonymous:1909:PNW

[382] Anonymous. The proposed new waterway in Canada. Scientific American, 100(23):418, June 5, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n23/
pdf/scientificamerican06051909-418b.pdf.

Anonymous:1909:SCU

[383] Anonymous. Storage of coal under water. Scientific American, 100(23):418, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
418c.pdf.

Anonymous:1909:Eba

[384] Anonymous. Engineering. Scientific American, 100(23):419, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n23/pdf/scientificamerican06051909-419.pdf.

Anonymous:1909:Ebb

[385] Anonymous. Electricity. Scientific American, 100(23):419, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n23/pdf/scientificamerican06051909-419a.pdf.

Anonymous:1909:Sq

[386] Anonymous. Science. Scientific American, 100(23):419, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n23/pdf/scientificamerican06051909-419b.pdf.

Anonymous:1909:BWT

[387] Anonymous. The bursting of a water tank and its effect. Scientific American, 100(23):420–421, June 5, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n23/
pdf/scientificamerican06051909-420.pdf.

Anonymous:1909:NFAa

[388] Anonymous. A novel French aeroplane. Scientific American, 100 (23):421, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
421.pdf.

Anonymous:1909:SNA

[389] Anonymous. Some new American aeroplanes. Scientific American, 100(23):421–422, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
421a.pdf.

Anonymous:1909:SMb

[390] Anonymous. The speed of the “Mauretania”. Scientific American, 100(23):423, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
423a.pdf.

Anonymous:1909:ASR

[391] Anonymous. Auroras: Some recent theories. Scientific American, 100(23):424, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
424.pdf.

Claudy:1909:MES

[392] C. H. Claudy. Making the eye of science. Scientific American, 100(23):425, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
425.pdf.

Shepstone:1909:EJP

[393] Harold J. Shepstone. Excavations at Jericho, Palestine. Scientific American, 100(23):426, June 5, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n23/
pdf/scientificamerican06051909-426.pdf.

Anonymous:1909:HMWg

[394] Anonymous. Handy man’s workshop. Scientific American, 100(23):427, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n23/pdf/scientificamerican06051909-427.pdf.

Anonymous:1909:SMMb

[395] Anonymous. A simple method of measuring the speed of photographic shutters. Scientific American, 100(23):427, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n23/pdf/scientificamerican06051909-427a.pdf.

Kavanaugh:1909:STL

[396] I. J. Kavanaugh. A substitute for tinfoil in Leyden jars. Scientific American, 100(23):427–428, June 5, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n23/
pdf/scientificamerican06051909-427b.pdf.

Blake:1909:WBP

[397] Thaleon Blake. Ways to brace poles for private telephone lines. Scientific American, 100(23):428, June 5, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n23/
pdf/scientificamerican06051909-428.pdf.

Graves:1909:DN

[398] W. D. Graves. The driving of a nail. Scientific American, 100(23):428, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n23/pdf/scientificamerican06051909-428a.pdf.

Parkhurst:1909:SSG

[399] W. Parkhurst and K. Parkhurst. Scroll-saw guide. Scientific American, 100(23):428, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n23/pdf/scientificamerican06051909-
428b.pdf.

Handy:1909:GA

[400] L. G. Handy. Gage for augers. Scientific American, 100(23):428, June 5, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n23/pdf/scientificamerican06051909-428c.pdf.

Anonymous:1909:Cq

[401] Anonymous. Correspondence. Scientific American, 100(24):443, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-443.pdf.

Anonymous:1909:RPIp

[402] Anonymous. Recently patented inventions. Scientific American, 100(24):448, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n24/pdf/scientificamerican06121909-
448.pdf.

Anonymous:1909:NQl

[403] Anonymous. Notes and queries. Scientific American, 100(24):448, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-448a.pdf.

Anonymous:1909:BNBj

[404] Anonymous. Books: New books, etc. Scientific American, 100(24):448, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-448b.pdf.

Anonymous:1909:IIq

[405] Anonymous. Index of inventions. Scientific American, 100(24):448–451, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-448c.pdf.

Anonymous:1909:PNPp

[406] Anonymous. Patent notes: Patents. Scientific American, 100(24):449, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-449.pdf.

Anonymous:1909:WON

[407] Anonymous. The waste of our natural resources. Scientific American, 100(24):438, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n24/pdf/scientificamerican06121909-
438.pdf.

Anonymous:1909:FBR

[408] Anonymous. Fewer broken rails. Scientific American, 100(24):438, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-438a.pdf.

Anonymous:1909:NFAb

[409] Anonymous. A new fire-alarm system for New York. Scientific American, 100(24):438, June 12, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n24/
pdf/scientificamerican06121909-438b.pdf.

Anonymous:1909:Ag

[410] Anonymous. Automobile. Scientific American, 100(24):439, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-439.pdf.

Anonymous:1909:Ebc

[411] Anonymous. Electricity. Scientific American, 100(24):439, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-439a.pdf.

Anonymous:1909:Sr

[412] Anonymous. Science. Scientific American, 100(24):439, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n24/pdf/scientificamerican06121909-439b.pdf.

Boyer:1909:BIA

[413] Jacques Boyer. Belin’s improved apparatus for the electrical transmission of pictures. Scientific American, 100(24):440, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n24/pdf/scientificamerican06121909-440.pdf.

Springer:1909:ESC

[414] J. F. Springer. Is the Earth’s shape changing. Scientific American, 100(24):441–442, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n24/pdf/scientificamerican06121909-
441.pdf.

Anonymous:1909:IAIa

[415] Anonymous. Impressions of American inventors. — II. Scientific American, 100(24):443, June 12, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n24/
pdf/scientificamerican06121909-443a.pdf.

Anonymous:1909:EHS

[416] Anonymous. Experiments with hydroplanes or skimmers. Scientific American, 100(24):444–445, June 12, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n24/
pdf/scientificamerican06121909-444.pdf.

Barton:1909:AR

[417] Victor Barton. Artificial rubies. Scientific American, 100(24):445–446, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-445.pdf.

Gradenwitz:1909:ATR

[418] Alfred Gradenwitz. An automobile trolley-repair truck. Scientific American, 100(24):446, June 12, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n24/
pdf/scientificamerican06121909-446.pdf.

Anonymous:1909:HMWh

[419] Anonymous. Handy man’s workshop. Scientific American, 100(24):447, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n24/pdf/scientificamerican06121909-447.pdf.

Ward:1909:SMC

[420] Frederick E. Ward. A simple medical coil. Scientific American, 100(24):447, June 12, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n24/pdf/scientificamerican06121909-
447a.pdf.

Anonymous:1909:RPIq 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n25/pdf/scientificamerican06191909-
465.pdf.

Anonymous:1909:NQm

[422] Anonymous. Notes and queries. Scientific American, 100(25):466, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-466.pdf.

Anonymous:1909:PNPq

[423] Anonymous. Patent notes: Patents. Scientific American, 100(25):466, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-466a.pdf.

Anonymous:1909:Ia

[424] Anonymous. Index. Scientific American, 100(25):466–471, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-466b.pdf.

Anonymous:1909:LPO

[425] Anonymous. Longer piers for ocean steamships. Scientific American, 100(25):454, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n25/pdf/scientificamerican06191909-
454.pdf.

Anonymous:1909:LTA

[426] Anonymous. Lighter than air. Scientific American, 100(25):454, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-454a.pdf.

Anonymous:1909:AFD

[427] Anonymous. The award of the fourth-dimension prize. Scientific American, 100(25):454, June 19, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n25/
pdf/scientificamerican06191909-454b.pdf.

Anonymous:1909:Ebd

[428] Anonymous. Engineering. Scientific American, 100(25):455, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-455.pdf.

Anonymous:1909:Ebe

[429] Anonymous. Electricity. Scientific American, 100(25):455, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-455a.pdf.

Anonymous:1909:St

[430] Anonymous. Science. Scientific American, 100(25):455, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n25/pdf/scientificamerican06191909-455b.pdf.

Cannon:1909:SDP

[431] William Austin Cannon. Studying the desert plant by experimental methods. Scientific American, 100(25):456, June 19, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n25/pdf/scientificamerican06191909-456.pdf.

Kendall:1909:IFS

[432] Walter Kendall. Improvements in the De Forest system of wireless telephony. Scientific American, 100(25):457, June 19, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n25/pdf/scientificamerican06191909-457.pdf.

Anonymous:1909:DTS

[433] Anonymous. Development of the transatlantic steamship. Scientific American, 100(25):458–459, June 19, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n25/
pdf/scientificamerican06191909-458.pdf.

Anonymous:1909:ASF

[434] Anonymous. The Aeronautic Society’s first Curtiss aeroplane. Scientific American, 100(25):460, June 19, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n25/
pdf/scientificamerican06191909-460.pdf.

Willey:1909:AMC

[435] Day Allen Willey. Los Angeles 200-mile conduit water supply. Scientific American, 100(25):460, June 19, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n25/
pdf/scientificamerican06191909-460a.pdf.

Anonymous:1909:MCFa

[436] Anonymous. Method of correcting faulty acoustic properties of public halls. Scientific American, 100(25):462, June 19, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n25/pdf/scientificamerican06191909-462.pdf.

Anonymous:1909:PDe

[437] Anonymous. Patent department. Scientific American, 100(25):463, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-463.pdf.

Anonymous:1909:CLB

[438] Anonymous. Combined life bail and gutter for swimming pools. Scientific American, 100(25):463, June 19, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n25/
pdf/scientificamerican06191909-463a.pdf.

Anonymous:1909:ECHa

[439] Anonymous. Expansion cutter-head for boring brass. Scientific American, 100(25):463, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n25/pdf/scientificamerican06191909-
463b.pdf.

Anonymous:1909:DGB

[440] Anonymous. Distributer for grain bins. Scientific American, 100 (25):463, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n25/pdf/scientificamerican06191909-
463c.pdf.

Anonymous:1909:SMC

[441] Anonymous. Sanitary milk and cream saving device. Scientific American, 100(25):463–464, June 19, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n25/
pdf/scientificamerican06191909-463d.pdf.

Anonymous:1909:NAL

[442] Anonymous. A new automobile lock. Scientific American, 100(25):464, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-464.pdf.

Anonymous:1909:ADD

[443] Anonymous. Apparatus for determining density. Scientific American, 100(25):464, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n25/pdf/scientificamerican06191909-
464a.pdf.

Anonymous:1909:ASDb

[444] Anonymous. Automatic starting device for siphons. Scientific American, 100(25):464, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n25/pdf/scientificamerican06191909-
464b.pdf.

Anonymous:1909:OIc

[445] Anonymous. Oddities in inventions. Scientific American, 100(25):464, June 19, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n25/pdf/scientificamerican06191909-464c.pdf.

Anonymous:1909:Cr

[446] Anonymous. Correspondence. Scientific American, 100(26):479, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-479.pdf.

Anonymous:1909:RPIr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n26/pdf/scientificamerican06261909-
484.pdf.

Anonymous:1909:NQn

[448] Anonymous. Notes and queries. Scientific American, 100(26):485, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-485.pdf.

Anonymous:1909:BNBk

[449] Anonymous. Books: New books, etc. Scientific American, 100(26):485, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-485a.pdf.

Anonymous:1909:PNPr

[450] Anonymous. Patent notes: Patents. Scientific American, 100(26):486, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-486.pdf.

Anonymous:1909:IIr

[451] Anonymous. Index of inventions. Scientific American, 100(26):486–491, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-486a.pdf.

Anonymous:1909:BSC

[452] Anonymous. Bearing of the Soo Canal accident on Panama. Scientific American, 100(26):474, June 26, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n26/
pdf/scientificamerican06261909-474.pdf.

Anonymous:1909:QBS

[453] Anonymous. Queensboro Bridge a safe structure. Scientific American, 100(26):474, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n26/pdf/scientificamerican06261909-
474a.pdf.

Anonymous:1909:Ebf

[454] Anonymous. Engineering. Scientific American, 100(26):475, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-475.pdf.

Anonymous:1909:Ebg

[455] Anonymous. Electricity. Scientific American, 100(26):475, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-475a.pdf.

Anonymous:1909:Su

[456] Anonymous. Science. Scientific American, 100(26):475, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n26/pdf/scientificamerican06261909-475b.pdf.

Anonymous:1909:STM

[457] Anonymous. Some tricks of the moving picture maker. Scientific American, 100(26):476–477, June 26, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n26/
pdf/scientificamerican06261909-476.pdf.

Anonymous:1909:MTZ

[458] Anonymous. The Metropolitan Tower and the “Zeppelin II” two striking comparisons. Scientific American, 100(26):478, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n26/pdf/scientificamerican06261909-478.pdf.

Anonymous:1909:LVM

[459] Anonymous. Large vs. moderate-sized battleships. Scientific American, 100(26):479, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n26/pdf/scientificamerican06261909-
479a.pdf.

Anonymous:1909:PBI

[460] Anonymous. Pipes in bail ingots. Scientific American, 100(26):479, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-479b.pdf.

Anonymous:1909:SMd

[461] Anonymous. Signaling to Mars. Scientific American, 100(26):479, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-479c.pdf.

Lawton:1909:FWT

[462] W. H. Lawton. Felling a water tank. Scientific American, 100(26):480, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-480.pdf.

Gradenwitz:1909:MP

[463] Alfred Gradenwitz. A motor plow. Scientific American, 100(26):480, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-480a.pdf.

Anonymous:1909:RFA

[464] Anonymous. Recent French aeroplanes and their performances. Scientific American, 100(26):481, June 26, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n26/
pdf/scientificamerican06261909-481.pdf.

Anonymous:1909:CZA

[465] Anonymous. The construction of the Zeppelin airship. Scientific American, 100(26):481, June 26, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n26/
pdf/scientificamerican06261909-481a.pdf.

Hunting:1909:MSO

[466] M. M. Hunting. A machine that sets ordinary type. Scientific American, 100(26):482, June 26, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v100/n26/
pdf/scientificamerican06261909-482.pdf.

Anonymous:1909:HMWi

[467] Anonymous. Handy man’s workshop. Scientific American, 100(26):483, June 26, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n26/pdf/scientificamerican06261909-483.pdf.

Bayley:1909:CCA

[468] I. G. Bayley. A convenient carrying arrangement for a drawing outfit. Scientific American, 100(26):483, June 26, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n26/pdf/scientificamerican06261909-483a.pdf.

Anonymous:1909:RPIs

[469] Anonymous. Recently patented inventions. Scientific American, 100(3):67, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
67.pdf.

Anonymous:1909:PNPs

[470] Anonymous. Patent notes: Patents. Scientific American, 100 (3):69, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
69.pdf.

Anonymous:1909:IIs

[471] Anonymous. Index of inventions. Scientific American, 100(3): 69–75, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
69a.pdf.

Anonymous:1909:Ah

[472] Anonymous. The 1909 automobile. Scientific American, 100(3):40, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n3/pdf/scientificamerican01161909-40.pdf.

Anonymous:1909:SAa

[473] Anonymous. Standardizing the automobile. Scientific American, 100(3):40, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
40a.pdf.

Anonymous:1909:GAI

[474] Anonymous. Growth of the automobile industry. Scientific American, 100(3):40–41, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
40b.pdf.

Anonymous:1909:ASA

[475] Anonymous. Alloy steels and automobile durability. Scientific American, 100(3):41, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
41.pdf.

Anonymous:1909:CRC

[476] Anonymous. Cost of running a car. Scientific American, 100(3):41, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n3/pdf/scientificamerican01161909-41a.pdf.

Anonymous:1909:SCa

[477] Anonymous. Selecting a car. Scientific American, 100(3):41, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n3/pdf/scientificamerican01161909-41b.pdf.

Anonymous:1909:BTA

[478] Anonymous. The buggy type automobile. Scientific American, 100(3):41, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
41c.pdf.

Rogers:1909:CTV

[479] Benjamin Rogers. The commercial truck vs the horse. Scientific American, 100(3):42–43, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
42.pdf.

Whitman:1909:HTI

[480] Roger B. Whitman. High tension ignition by magneto. Scientific American, 100(3):44–45, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
44.pdf.

Parker:1909:OLT

[481] Orrel A. Parker. Our legion tires and their trouble. Scientific American, 100(3):46, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
46.pdf.

Hanauer:1909:ALS

[482] Thomas D. Hanauer. Automobile lubrication — some elementary principles. Scientific American, 100(3):48–49, January 16, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n3/pdf/scientificamerican01161909-48.pdf.

Roberts:1909:TCA

[483] E. W. Roberts. The two-cycle automobile motor. Scientific American, 100(3):50, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
50.pdf.

Heron:1909:HMW

[484] George Heron. Handy man’s workshop. Scientific American, 100 (3):52–53, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
52.pdf.

Anonymous:1909:SCP

[485] Anonymous. Scraping carbon from the piston heads. Scientific American, 100(3):53, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
53.pdf.

Anonymous:1909:RBS

[486] Anonymous. Relining the brake shoes. Scientific American, 100 (3):53, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
53a.pdf.

Anonymous:1909:WLN

[487] Anonymous. When a lost nut cannot be replaced. Scientific American, 100(3):53, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
53b.pdf.

Anonymous:1909:GHW

[488] Anonymous. Getting home with a weak battery. Scientific American, 100(3):53, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
53c.pdf.

Anonymous:1909:SAN

[489] Anonymous. Some automobile novelties. Scientific American, 100(3):58, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
58.pdf.

Anonymous:1909:NCC

[490] Anonymous. A novel carrying case for automobile tires. Scientific American, 100(3):58, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
58a.pdf.

Anonymous:1909:NSA

[491] Anonymous. A new suspension for automobiles. Scientific American, 100(3):58, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
58b.pdf.

Anonymous:1909:NSC

[492] Anonymous. A novel speed-changing gear. Scientific American, 100(3):58, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
58c.pdf.

Anonymous:1909:ANS

[493] Anonymous. Automobile notes and suggestions. Scientific American, 100(3):60, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
60.pdf.

Anonymous:1909:CLC

[494] Anonymous. Correcting a leaky carbureter float valve. Scientific American, 100(3):66, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
66.pdf.

Anonymous:1909:GHB

[495] Anonymous. Getting home with a broken universal joint. Scientific American, 100(3):66, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
66a.pdf.

Anonymous:1909:SRB

[496] Anonymous. Suppressing rattle in brakes, mud guards, etc. Scientific American, 100(3):66, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
66b.pdf.

Anonymous:1909:NGS

[497] Anonymous. A novel gasoline strainer. Scientific American, 100 (3):66–67, January 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n3/pdf/scientificamerican01161909-
66c.pdf.

Anonymous:1909:Ct

[498] Anonymous. Correspondence. Scientific American, 100(4):83, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n4/pdf/scientificamerican01231909-83.pdf.

Anonymous:1909:RPIt

[499] Anonymous. Recently patented inventions. Scientific American, 100(4):89, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
89.pdf.

Anonymous:1909:PNPt

[500] Anonymous. Patent notes: Patents. Scientific American, 100 (4):89, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
89a.pdf.

Anonymous:1909:IIt

[501] Anonymous. Index of inventions. Scientific American, 100(4): 89–91, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
89b.pdf.

Anonymous:1909:NPb

[502] Anonymous. The Naval Programme for 1909. Scientific American, 100(4):78, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
78.pdf.

Anonymous:1909:ODM

[503] Anonymous. Our decadent merchant marine. Scientific American, 100(4):78, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
78a.pdf.

Anonymous:1909:TGC

[504] Anonymous. The traffic of a great city. Scientific American, 100(4):78, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
78b.pdf.

Anonymous:1909:Ebh

[505] Anonymous. Engineering. Scientific American, 100(4):79, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n4/pdf/scientificamerican01231909-79.pdf.

Anonymous:1909:Ebi

[506] Anonymous. Electricity. Scientific American, 100(4):79, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n4/pdf/scientificamerican01231909-79a.pdf.

Anonymous:1909:Sw

[507] Anonymous. Science. Scientific American, 100(4):79, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n4/pdf/scientificamerican01231909-79b.pdf.

Forrest:1909:SM

[508] George Forrest. The submarine mine. Scientific American, 100 (4):80–81, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
80.pdf.

Anonymous:1909:AMF

[509] Anonymous. Aeroplanes and motors at the First Paris Aeronautical Salon. Scientific American, 100(4):81–82, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n4/pdf/scientificamerican01231909-81.pdf.

Murray:1909:SAE

[510] W. J. Murray. The scientific aspect of earthquakes and volcanoes. Scientific American, 100(4):82–83, January 23, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n4/pdf/scientificamerican01231909-82.pdf.

Anonymous:1909:AWa

[511] Anonymous. Aeroplanes in warfare. Scientific American, 100 (4):83, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
83a.pdf.

Anonymous:1909:MHS

[512] Anonymous. Multi-hull steamships. Scientific American, 100(4): 83, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
83b.pdf.

Anonymous:1909:MIAa

[513] Anonymous. The meat industry of America — I. Scientific American, 100(4):84–86, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
84.pdf.

Anonymous:1909:CSC

[514] Anonymous. Concrete steel caissons. Scientific American, 100 (4):86, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
86.pdf.

Anonymous:1909:CLH

[515] Anonymous. Curious life histories. Scientific American, 100(4):87, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n4/pdf/scientificamerican01231909-87.pdf.

Anonymous:1909:HMWj

[516] Anonymous. Handy man’s workshop. Scientific American, 100 (4):88, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
88.pdf.

C:1909:MCB

[517] W. J. C. Mending a cracked bottle. Scientific American, 100(4):88, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n4/pdf/scientificamerican01231909-88a.pdf.

Laurence:1909:HTB

[518] A. C. Laurence. A hydraulic test for the boiler. Scientific American, 100(4):88, January 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n4/pdf/scientificamerican01231909-
88b.pdf.

Anonymous:1909:Cu

[519] Anonymous. Correspondence. Scientific American, 100(5):99, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n5/pdf/scientificamerican01301909-99.pdf.

Russell:1909:HF

[520] Henry Norris Russell. The heavens in February. Scientific American, 100(5):102–103, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
102.pdf.

Anonymous:1909:RPIu

[521] Anonymous. Recently patented inventions. Scientific American, 100(5):105, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
105.pdf.

Anonymous:1909:PNPu

[522] Anonymous. Patent notes: Patents. Scientific American, 100 (5):105, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
105a.pdf.

Anonymous:1909:IIu

[523] Anonymous. Index of inventions. Scientific American, 100(5): 105–107, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
105b.pdf.

Anonymous:1909:GDM

[524] Anonymous. The Gatun Dam misrepresentations. Scientific American, 100(5):94, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
94.pdf.

Anonymous:1909:NAA

[525] Anonymous. The Ninth Annual Automobile Show in Madison Square Garden. Scientific American, 100(5):94, January 30, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n5/pdf/scientificamerican01301909-94a.pdf.

Anonymous:1909:LIL

[526] Anonymous. The labor involved in liquefying helium. Scientific American, 100(5):94, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
94b.pdf.

Anonymous:1909:Ebj

[527] Anonymous. Engineering. Scientific American, 100(5):95, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n5/pdf/scientificamerican01301909-95.pdf.

Anonymous:1909:Ebk

[528] Anonymous. Electricity. Scientific American, 100(5):95, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n5/pdf/scientificamerican01301909-95a.pdf.

Anonymous:1909:Sx

[529] Anonymous. Science. Scientific American, 100(5):95, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n5/pdf/scientificamerican01301909-95b.pdf.

Anonymous:1909:BAA

[530] Anonymous. The British Army aeroplane. Scientific American, 100(5):96, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
96.pdf.

Anonymous:1909:ANc

[531] Anonymous. Aeronautical notes. Scientific American, 100(5):96, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n5/pdf/scientificamerican01301909-96a.pdf.

Whitman:1909:TF

[532] Roger B. Whitman. Taximeter frauds. Scientific American, 100 (5):97–98, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
97.pdf.

Anonymous:1909:CPU

[533] Anonymous. Clay products of the United States. Scientific American, 100(5):98–99, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
98.pdf.

Anonymous:1909:ABS

[534] Anonymous. Automatic block signals. Scientific American, 100 (5):99, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
99a.pdf.

Anonymous:1909:MIAb

[535] Anonymous. The meat industry of America — II. Scientific American, 100(5):99–102, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
99b.pdf.

Lodian:1909:SBO

[536] L. Lodian. Some breads of other lands. Scientific American, 100(5):103, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
103.pdf.

Anonymous:1909:MBH

[537] Anonymous. Milk-bottle holder. Scientific American, 100(5):104, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n5/pdf/scientificamerican01301909-104.pdf.

Anonymous:1909:DTE

[538] Anonymous. Device for testing eggs. Scientific American, 100 (5):104, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
104a.pdf.

Anonymous:1909:FOW

[539] Anonymous. Flowing oil wells by compressed air. Scientific American, 100(5):104, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
104b.pdf.

Anonymous:1909:FDD

[540] Anonymous. Finger-developing device. Scientific American, 100 (5):104, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
104c.pdf.

Anonymous:1909:VWS

[541] Anonymous. Vehicle wheel sled runner. Scientific American, 100(5):104, January 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n5/pdf/scientificamerican01301909-
104d.pdf.

Anonymous:1909:Cv

[542] Anonymous. Correspondence. Scientific American, 100(6):115, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n6/pdf/scientificamerican02061909-115.pdf.

Anonymous:1909:RPIv

[543] Anonymous. Recently patented inventions. Scientific American, 100 (6):122–123, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
122.pdf.

Anonymous:1909:NQo

[544] Anonymous. Notes and queries. Scientific American, 100(6):123, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n6/pdf/scientificamerican02061909-123.pdf.

Anonymous:1909:BNBl

[545] Anonymous. Books: New books, etc. Scientific American, 100 (6):124, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
124.pdf.

Anonymous:1909:IIv

[546] Anonymous. Index of inventions. Scientific American, 100(6): 124–127, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
124a.pdf.

Anonymous:1909:PNPv

[547] Anonymous. Patent notes: Patents. Scientific American, 100 (6):124, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
124b.pdf.

Anonymous:1909:LRD

[548] Anonymous. Lessons of the “Republic” disaster. Scientific American, 100(6):110, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
110.pdf.

Anonymous:1909:FA

[549] Anonymous. France and the aeroplane. Scientific American, 100(6):110, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
110a.pdf.

Anonymous:1909:TSA

[550] Anonymous. The “trust” system as applied to scientific research. Scientific American, 100(6):110, February 6, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n6/pdf/scientificamerican02061909-110b.pdf.

Anonymous:1909:Ebl

[551] Anonymous. Engineering. Scientific American, 100(6):111, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n6/pdf/scientificamerican02061909-111.pdf.

Anonymous:1909:Ai

[552] Anonymous. Aeronautics. Scientific American, 100(6):111, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n6/pdf/scientificamerican02061909-111a.pdf.

Anonymous:1909:Sy

[553] Anonymous. Science. Scientific American, 100(6):111, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n6/pdf/scientificamerican02061909-111b.pdf.

Mclure:1909:IUL

[554] W. Frank M’clure. The industrial use of the lifting magnet. Scientific American, 100(6):112, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
112.pdf.

Sherman:1909:ICR

[555] Edward C. Sherman. The improvement of the Charles River at Boston Mass. Scientific American, 100(6):113–114, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n6/pdf/scientificamerican02061909-113.pdf.

Mitchell:1909:ST

[556] S. A. Mitchell. A sensitive thermometer. Scientific American, 100 (6):114–115, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
114.pdf.

Anonymous:1909:AAa

[557] Anonymous. That aerolite again. Scientific American, 100(6): 115, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
115a.pdf.

Anonymous:1909:FBC

[558] Anonymous. Are filtering beds correctly constructed? Scientific American, 100(6):115–116, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
115b.pdf.

Anonymous:1909:PEH

[559] Anonymous. The physical energy of the house-fly. Scientific American, 100(6):116, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
116.pdf.

Lay:1909:PCP

[560] A. Lay. Present conditions in the Panama Canal Zone. Scientific American, 100(6):117, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
117.pdf.

Anonymous:1909:HMWk

[561] Anonymous. Handy man’s workshop. Scientific American, 100 (6):120, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
120.pdf.

Lurcott:1909:HWS

[562] Charles Lurcott. How to wind a spaced coil spring. Scientific American, 100(6):120, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
120a.pdf.

Pitkin:1909:JWB

[563] C. A. Pitkin. A joint for the work bench. Scientific American, 100(6):120, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
120b.pdf.

Fagan:1909:PKB

[564] T. Fagan. Patch for kitchen boilers. Scientific American, 100 (6):120, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
120c.pdf.

Warner:1909:AMM

[565] Fred G. Warner. Another method of making metal lamp shades. Scientific American, 100(6):120–121, February 6, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n6/pdf/scientificamerican02061909-120d.pdf.

Anonymous:1909:SGT

[566] Anonymous. Stopcock of glass tubing. Scientific American, 100 (6):121, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
121.pdf.

Anonymous:1909:SSBa

[567] Anonymous. Simple support for bicycles. Scientific American, 100(6):121, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
121a.pdf.

Searing:1909:HFR

[568] A. V. Searing, Jr. How to file round work. Scientific American, 100(6):121, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
121b.pdf.

Anonymous:1909:HMH

[569] Anonymous. How to mend a hammock or fishing net. Scientific American, 100(6):121, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
121c.pdf.

Anonymous:1909:FPW

[570] Anonymous. A fire-plug wrench. Scientific American, 100(6): 121, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
121d.pdf.

Hooper:1909:HSA

[571] William Hooper. How to solder aluminium. Scientific American, 100(6):121, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
121e.pdf.

Anonymous:1909:SBC

[572] Anonymous. Straightening buckled castings. Scientific American, 100(6):121, February 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n6/pdf/scientificamerican02061909-
121f.pdf.

Anonymous:1909:Cw

[573] Anonymous. Correspondence. Scientific American, 100(7):135, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n7/pdf/scientificamerican02131909-135.pdf.

Anonymous:1909:RPIw

[574] Anonymous. Recently patented inventions. Scientific American, 100(7):140, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
140a.pdf.

Anonymous:1909:NQp

[575] Anonymous. Notes and queries. Scientific American, 100(7):141, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n7/pdf/scientificamerican02131909-141.pdf.

Anonymous:1909:BNBm

[576] Anonymous. Books: New books, etc. Scientific American, 100 (7):141, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
141a.pdf.

Anonymous:1909:PNPw

[577] Anonymous. Patent notes: Patents. Scientific American, 100 (7):141, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
141b.pdf.

Anonymous:1909:IIw

[578] Anonymous. Index of inventions. Scientific American, 100(7): 141–143, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
141c.pdf.

Anonymous:1909:LLC

[579] Anonymous. The Lock-and-Lake Canal approved. Scientific American, 100(7):130, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
130.pdf.

Anonymous:1909:CDO

[580] Anonymous. Cuba’s debt to our engineers. Scientific American, 100(7):130, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
130a.pdf.

Anonymous:1909:CMA

[581] Anonymous. Congress and military aeronautics. Scientific American, 100(7):130, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
130b.pdf.

Anonymous:1909:HPF

[582] Anonymous. The high-pressure fire automobile. Scientific American, 100(7):130, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
130c.pdf.

Anonymous:1909:NYM

[583] Anonymous. The New York Motor Boat Show. Scientific American, 100(7):130, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
130d.pdf.

Anonymous:1909:Ebm

[584] Anonymous. Engineering. Scientific American, 100(7):131, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n7/pdf/scientificamerican02131909-131.pdf.

Anonymous:1909:Ebn

[585] Anonymous. Electricity. Scientific American, 100(7):131, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n7/pdf/scientificamerican02131909-131a.pdf.

Anonymous:1909:Sz

[586] Anonymous. Science. Scientific American, 100(7):131, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n7/pdf/scientificamerican02131909-131b.pdf.

Boyer:1909:PMP

[587] Jacques Boyer. Projecting moving pictures in the light. Scientific American, 100(7):132, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
132.pdf.

Anonymous:1909:ESC

[588] Anonymous. An Egyptian steam cultivator. Scientific American, 100(7):133, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
133.pdf.

Honey:1909:OM

[589] Frederic R. Honey. The opposition of Mars in 1909. Scientific American, 100(7):134, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
134.pdf.

Anonymous:1909:PSA

[590] Anonymous. The practical side of aerial locomotion. Scientific American, 100(7):135, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
135a.pdf.

Anonymous:1909:CWV

[591] Anonymous. Comparison of Wright and Voisin aeroplanes. Scientific American, 100(7):135, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
135b.pdf.

Anonymous:1909:MCFb

[592] Anonymous. More curious facts about numbers. Scientific American, 100(7):135, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
135c.pdf.

Anonymous:1909:ABM

[593] Anonymous. The Aurora Borealis and Morehouse’s Comet. Scientific American, 100(7):135, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
135d.pdf.

Anonymous:1909:MA

[594] Anonymous. Magnetism and the Aurora. Scientific American, 100(7):135, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
135e.pdf.

Anonymous:1909:NLS

[595] Anonymous. Novel life-saving apparatus. Scientific American, 100 (7):136, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
136.pdf.

DuPuy:1909:RBU

[596] William Atherton Du Puy. Road building by the United States in Cuba. Scientific American, 100(7):136–138, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n7/pdf/scientificamerican02131909-136a.pdf.

Reighard:1909:PAU

[597] Jacob Reighard. Photographing animals under water. Scientific American, 100(7):138–139, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
138.pdf.

Anonymous:1909:NDRa

[598] Anonymous. A new direct-reading photometer. Scientific American, 100(7):139, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
139.pdf.

Anonymous:1909:PWG

[599] Anonymous. A $250 prize for a wet-grain conveying system. Scientific American, 100(7):140, February 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v100/n7/pdf/scientificamerican02131909-
140.pdf.

Anonymous:1909:Cx

[600] Anonymous. Correspondence. Scientific American, 100(8):156, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n8/pdf/scientificamerican02201909-156a.pdf.

Anonymous:1909:RPIx

[601] Anonymous. Recently patented inventions. Scientific American, 100(8):158, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
158.pdf.

Anonymous:1909:IIx

[602] Anonymous. Index of inventions. Scientific American, 100(8): 158–163, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
158a.pdf.

Anonymous:1909:PNPx

[603] Anonymous. Patent notes: Patents. Scientific American, 100 (8):158, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
158b.pdf.

Anonymous:1909:LRB

[604] Anonymous. Lessons and results of the battleship cruise. Scientific American, 100(8):146, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
146.pdf.

Anonymous:1909:EPF

[605] Anonymous. The electric potentiality of fruit and vegetables. Scientific American, 100(8):146, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
146a.pdf.

Anonymous:1909:PVE

[606] Anonymous. Pseudo-volcanic eruptions. Scientific American, 100 (8):146, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
146b.pdf.

Anonymous:1909:Aj

[607] Anonymous. Aeronautics. Scientific American, 100(8):147, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n8/pdf/scientificamerican02201909-147.pdf.

Anonymous:1909:Ebo

[608] Anonymous. Electricity. Scientific American, 100(8):147, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n8/pdf/scientificamerican02201909-147a.pdf.

Anonymous:1909:Sba

[609] Anonymous. Science. Scientific American, 100(8):147, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n8/pdf/scientificamerican02201909-147b.pdf.

Anonymous:1909:CSN

[610] Anonymous. Comparative strength of the Navies of the world. Scientific American, 100(8):148–149, February 20, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n8/pdf/scientificamerican02201909-148.pdf.

Anonymous:1909:OSS

[611] Anonymous. Our system of seacoast fortifications. Scientific American, 100(8):149–150, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
149.pdf.

Anonymous:1909:SGC

[612] Anonymous. The silent gun — counterpart of smokeless powder. Scientific American, 100(8):150, February 20, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n8/pdf/scientificamerican02201909-150.pdf.

Anonymous:1909:SAFc

[613] Anonymous. Santiago and Atlantic Fleets compared. Scientific American, 100(8):150–151, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
150a.pdf.

Anonymous:1909:ECM

[614] Anonymous. The extension of the Chicago, Milwaukee St. Paul Railway to the Pacific Coast. Scientific American, 100(8):151–152, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n8/pdf/scientificamerican02201909-151.pdf.

Anonymous:1909:CIM

[615] Anonymous. Carnegie Institution Magnetic Survey of the World. Scientific American, 100(8):153, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
153.pdf.

Anonymous:1909:CFW

[616] Anonymous. Composition of the Fleet which sailed around the world. Scientific American, 100(8):156, February 20, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n8/pdf/scientificamerican02201909-156.pdf.

Anonymous:1909:YDO

[617] Anonymous. A yacht designer’s opinion of the Navy situation. Scientific American, 100(8):156, February 20, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n8/pdf/scientificamerican02201909-156b.pdf.

Anonymous:1909:Ak

[618] Anonymous. Alcoline. Scientific American, 100(8):156, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n8/pdf/scientificamerican02201909-156c.pdf.

Anonymous:1909:CAa

[619] Anonymous. A combination anchor. Scientific American, 100 (8):157, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
157.pdf.

Anonymous:1909:IC

[620] Anonymous. The itinerary of the cruise. Scientific American, 100(8):157, February 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n8/pdf/scientificamerican02201909-
157a.pdf.

Anonymous:1909:Cy

[621] Anonymous. Correspondence. Scientific American, 100(9):170, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n9/pdf/scientificamerican02271909-170.pdf.

Russell:1909:HMb

[622] Henry Norris Russell. The heavens in March. Scientific American, 100(9):174, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
174.pdf.

Anonymous:1909:RPIy

[623] Anonymous. Recently patented inventions. Scientific American, 100(9):176, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
176.pdf.

Anonymous:1909:NQq

[624] Anonymous. Notes and queries. Scientific American, 100(9):176, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n9/pdf/scientificamerican02271909-176a.pdf.

Anonymous:1909:BNBn

[625] Anonymous. Books: New books, etc. Scientific American, 100 (9):176, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
176b.pdf.

Anonymous:1909:IIy

[626] Anonymous. Index of inventions. Scientific American, 100(9): 176–179, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
176c.pdf.

Anonymous:1909:PNPy

[627] Anonymous. Patent notes: Patents. Scientific American, 100 (9):176, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
176d.pdf.

Anonymous:1909:ERF

[628] Anonymous. Engineers report in favor of lock canal. Scientific American, 100(9):166, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
166.pdf.

Anonymous:1909:DSD

[629] Anonymous. Double-side-door trains on the subway. Scientific American, 100(9):166, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
166a.pdf.

Anonymous:1909:CNB

[630] Anonymous. Consolidation of the Naval Bureaus. Scientific American, 100(9):166, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
166b.pdf.

Anonymous:1909:Ebp

[631] Anonymous. Engineering. Scientific American, 100(9):167, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n9/pdf/scientificamerican02271909-167.pdf.

Anonymous:1909:Ebq

[632] Anonymous. Electricity. Scientific American, 100(9):167, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n9/pdf/scientificamerican02271909-167a.pdf.

Anonymous:1909:Sbb

[633] Anonymous. Science. Scientific American, 100(9):167, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v100/n9/pdf/scientificamerican02271909-167b.pdf.

Anonymous:1909:PT

[634] Anonymous. A “putting-on” tool. Scientific American, 100(9): 168, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
168.pdf.

Fergusson:1909:EUA

[635] S. P. Fergusson. The exploration of the upper air by means of ballons sondes. Scientific American, 100(9):169–170, February 27, 1909.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n9/pdf/scientificamerican02271909-169.pdf.

Anonymous:1909:IAIb

[636] Anonymous. Impressions of American inventors — I, Thomas A. Edison. Scientific American, 100(9):171, February 27, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n9/pdf/scientificamerican02271909-171.pdf.

Anonymous:1909:MR

[637] Anonymous. Motoring on runners. Scientific American, 100(9): 172–173, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
172.pdf.

Anonymous:1909:HMWl

[638] Anonymous. Handy man’s workshop. Scientific American, 100 (9):175, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
175.pdf.

Bergstrom:1909:HMA

[639] John A. Bergstrom. Home-made adjustable socket for tungsten lamps. Scientific American, 100(9):175, February 27, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n9/pdf/scientificamerican02271909-175a.pdf.

Bayley:1909:FWIb

[640] I. Q. Bayley. Furnishing the workshop — III. Scientific American, 100(9):175, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
175b.pdf.

Clock:1909:BCS

[641] E. E. Clock. Bicycle coasting sled. Scientific American, 100 (9):175, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
175c.pdf.

Johns:1909:HDT

[642] B. A. Johns. How to drill through brick and soft stone. Scientific American, 100(9):175, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v100/n9/pdf/scientificamerican02271909-
175d.pdf.

Huggins:1909:AMR

[643] George E. Huggins. Another method of reducing the range of a Springfield rifle. Scientific American, 100(9):175, February 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v100/
n9/pdf/scientificamerican02271909-175e.pdf.

Russell:1909:HJc

[644] Henry Norris Russell. The heavens in July. Scientific American, 101(1): 12, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-12.pdf.

Anonymous:1909:RPIz

[645] Anonymous. Recently patented inventions. Scientific American, 101(1): 14, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-14.pdf.

Anonymous:1909:NQr

[646] Anonymous. Notes and queries. Scientific American, 101(1):14, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-14a.pdf.

Anonymous:1909:BNBo

[647] Anonymous. Books: New books, etc. Scientific American, 101(1):14, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-14b.pdf.

Anonymous:1909:PNPz

[648] Anonymous. Patent notes: Patents. Scientific American, 101(1):15, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-15.pdf.

Anonymous:1909:IIz

[649] Anonymous. Index of inventions. Scientific American, 101(1):15–19, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-15a.pdf.

Anonymous:1909:FFA

[650] Anonymous. Fact and fancy in aeronautics. Scientific American, 101(1): 4, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-4.pdf.

Anonymous:1909:CCC

[651] Anonymous. The Cape Cod Canal. Scientific American, 101(1):4, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-4a.pdf.

Anonymous:1909:CAb

[652] Anonymous. The Copyright Act. Scientific American, 101(1):4, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-4b.pdf.

Anonymous:1909:Ebr

[653] Anonymous. Engineering. Scientific American, 101(1):5, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n1/pdf/scientificamerican07031909-5.pdf.

Anonymous:1909:Al

[654] Anonymous. Aeronautics. Scientific American, 101(1):5, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n1/pdf/scientificamerican07031909-5a.pdf.

Anonymous:1909:Sbd

[655] Anonymous. Science. Scientific American, 101(1):5, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n1/pdf/scientificamerican07031909-5b.pdf.

Anonymous:1909:FDSa

[656] Anonymous. The fourth dimension simply explained. Scientific American, 101(1):6, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n1/pdf/scientificamerican07031909-
6.pdf.

Anonymous:1909:CIU

[657] Anonymous. The cotton industry of the United States. Scientific American, 101(1):7–8, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n1/pdf/scientificamerican07031909-
7.pdf.

Anonymous:1909:MOF

[658] Anonymous. The “Michigan” — our first “Dreadnought”. Scientific American, 101(1):9, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n1/pdf/scientificamerican07031909-
9.pdf.

Wade:1909:MID

[659] Herbert T. Wade. A museum to illustrate the development of mathematics. Scientific American, 101(1):10, July 3, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n1/pdf/scientificamerican07031909-10.pdf.

Anonymous:1909:HFW

[660] Anonymous. A huge filtered water reservoir for London. Scientific American, 101(1):11, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n1/pdf/scientificamerican07031909-
11.pdf.

Rader:1909:IMS

[661] Morris Rader. To install a motor in a small boat. Scientific American, 101(1):13, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n1/pdf/scientificamerican07031909-
13.pdf.

Mencken:1909:HMSa

[662] August Mencken. Handy man’s surveying instrument. Scientific American, 101(1):13, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n1/pdf/scientificamerican07031909-
13a.pdf.

Umoselle:1909:HMS

[663] Charles A. Umoselle. A home-made spray outfit. Scientific American, 101(1):13, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n1/pdf/scientificamerican07031909-
13b.pdf.

Bergstrom:1909:HMG

[664] J. A. Bergstrom. How to mend garden hose. Scientific American, 101(1): 13, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n1/pdf/scientificamerican07031909-13c.pdf.

Anonymous:1909:GSL

[665] Anonymous. A good substitute for leather belting. Scientific American, 101(1):13, July 3, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n1/pdf/scientificamerican07031909-
13d.pdf.

Anonymous:1909:RPIba

[666] Anonymous. Recently patented inventions. Scientific American, 101 (10):165–166, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
165.pdf.

Anonymous:1909:BNBp

[667] Anonymous. Books: New books, etc. Scientific American, 101 (10):166, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
166.pdf.

Anonymous:1909:PNPba

[668] Anonymous. Patent notes: Patents. Scientific American, 101 (10):166, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
166a.pdf.

Anonymous:1909:IIba

[669] Anonymous. Index of inventions. Scientific American, 101(10): 166–171, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
166b.pdf.

Anonymous:1909:LRA

[670] Anonymous. Lessons of the Rheims Aviation Contests. Scientific American, 101(10):154, September 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n10/
pdf/scientificamerican09041909-154.pdf.

Anonymous:1909:SDC

[671] Anonymous. A sanitary drainage canal for Buffalo. Scientific American, 101(10):154, September 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n10/
pdf/scientificamerican09041909-154a.pdf.

Anonymous:1909:WDWa

[672] Anonymous. Why do watch mainsprings break? Scientific American, 101(10):154, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
154b.pdf.

Anonymous:1909:FVZ

[673] Anonymous. The first voyage of the “Zeppelin III”. Scientific American, 101(10):154, September 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n10/
pdf/scientificamerican09041909-154c.pdf.

Anonymous:1909:Ebs

[674] Anonymous. Engineering. Scientific American, 101(10):155, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n10/pdf/scientificamerican09041909-155.pdf.

Anonymous:1909:Ebt

[675] Anonymous. Electricity. Scientific American, 101(10):155, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n10/pdf/scientificamerican09041909-155a.pdf.

Anonymous:1909:Sbe

[676] Anonymous. Science. Scientific American, 101(10):155, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n10/pdf/scientificamerican09041909-155b.pdf.

Anonymous:1909:NSI

[677] Anonymous. A new speed indicator for marine propellers. Scientific American, 101(10):156, September 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n10/
pdf/scientificamerican09041909-156.pdf.

Anonymous:1909:PHC

[678] Anonymous. Powerful hoisting and conveying machine. Scientific American, 101(10):156, September 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n10/
pdf/scientificamerican09041909-156a.pdf.

Singleton:1909:HH

[679] Esther Singleton. Hieroglyphs of the heavens. Scientific American, 101(10):157–158, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
157.pdf.

Anonymous:1909:AMR

[680] Anonymous. The Aviation Meeting at Rheims. Scientific American, 101(10):159, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
159.pdf.

Ball:1909:MRF

[681] S. Mays Ball. Measuring a river’s flow. Scientific American, 101 (10):160, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
160.pdf.

Anonymous:1909:GMM

[682] Anonymous. German Museum of Masterpieces of Science and Industry. Scientific American, 101(10):161–162, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n10/pdf/scientificamerican09041909-161.pdf.

Anonymous:1909:HMWm

[683] Anonymous. Handy man’s workshop. Scientific American, 101 (10):163, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
163.pdf.

Michaud:1909:HMC

[684] Gustave Michaud. Home-made chemical perfume. Scientific American, 101(10):163–164, September 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n10/
pdf/scientificamerican09041909-163a.pdf.

Bayley:1909:PHA

[685] I. G. Bayley. Paper hanger’s adjustable templet. Scientific American, 101(10):164, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
164.pdf.

Worts:1909:HMS

[686] George F. Worts. How to make a simple dry battery. Scientific American, 101(10):164, September 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n10/
pdf/scientificamerican09041909-164a.pdf.

Sidman:1909:HGC

[687] Charles A. Sidman. Home-grown calabash gourds. Scientific American, 101(10):164, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
164b.pdf.

Pearse:1909:HMS

[688] R. W. Pearse. The handy man’s spark plug. Scientific American, 101(10):164, September 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n10/pdf/scientificamerican09041909-
164c.pdf.

Brockman:1909:HMS

[689] Robert H. Brockman. How to make a simple electric engine. Scientific American, 101(10):164, September 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n10/
pdf/scientificamerican09041909-164d.pdf.

Anonymous:1909:Cz

[690] Anonymous. Correspondence. Scientific American, 101(11):178, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n11/pdf/scientificamerican09111909-178.pdf.

Anonymous:1909:PNPbb

[691] Anonymous. Patent notes: Patents. Scientific American, 101(11): 181–185, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
185.pdf.

Anonymous:1909:IIbb

[692] Anonymous. Index of inventions. Scientific American, 101(11): 181–186, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
185a.pdf.

Anonymous:1909:RPIbb

[693] Anonymous. Recently patented inventions. Scientific American, 101 (11):184, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
184.pdf.

Anonymous:1909:NQs

[694] Anonymous. Notes and queries. Scientific American, 101(11): 184, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
184a.pdf.

Anonymous:1909:BNBq

[695] Anonymous. Books: New books, etc. Scientific American, 101 (11):184–185, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
184b.pdf.

Anonymous:1909:DCN

[696] Anonymous. Dr. Cook and the North Pole. Scientific American, 101(11):174, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
174.pdf.

Anonymous:1909:TR

[697] Anonymous. The transatlantic record. Scientific American, 101 (11):174, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
174a.pdf.

Anonymous:1909:ELO

[698] Anonymous. Economic loss of overcrowding. Scientific American, 101(11):174, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
174b.pdf.

Anonymous:1909:EGC

[699] Anonymous. The elimination of grade crossings. Scientific American, 101(11):174, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
174c.pdf.

Anonymous:1909:Ebu

[700] Anonymous. Engineering. Scientific American, 101(11):175, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n11/pdf/scientificamerican09111909-175.pdf.

Anonymous:1909:Ebv

[701] Anonymous. Electricity. Scientific American, 101(11):175, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n11/pdf/scientificamerican09111909-175a.pdf.

Anonymous:1909:Sbf

[702] Anonymous. Science. Scientific American, 101(11):175, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n11/pdf/scientificamerican09111909-175b.pdf.

Talbot:1909:COP

[703] M. Hamilton Talbot. Counting our people by machine. Scientific American, 101(11):176, September 11, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n11/
pdf/scientificamerican09111909-176.pdf.

Willey:1909:HSC

[704] Day Allen Willey. How Sudd chokes the Nile River. Scientific American, 101(11):177–178, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n11/
pdf/scientificamerican09111909-177.pdf.

Anonymous:1909:RCE

[705] Anonymous. The recent curious eclipse of the Sun. Scientific American, 101(11):178, September 11, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n11/
pdf/scientificamerican09111909-178a.pdf.

Springer:1909:SBK

[706] J. F. Springer. The Seven Bridges of Königsberg and other puzzles. Scientific American, 101(11):179–180, September 11, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n11/pdf/scientificamerican09111909-179.pdf.

Anonymous:1909:TRA

[707] Anonymous. Termination of the Rheims Aviation Meeting. Scientific American, 101(11):180–181, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n11/
pdf/scientificamerican09111909-180.pdf.

Anonymous:1909:NCG

[708] Anonymous. A novel cantilever grain elevator. Scientific American, 101(11):182–183, September 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n11/pdf/scientificamerican09111909-
182.pdf.

Anonymous:1909:DCD

[709] Anonymous. Dr. Cook’s discovery of the North Pole. Scientific American, 101(11):183, September 11, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n11/
pdf/scientificamerican09111909-183.pdf.

Anonymous:1909:Cbb

[710] Anonymous. Correspondence. Scientific American, 101(12):195, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n12/pdf/scientificamerican09181909-195.pdf.

Anonymous:1909:RPIbc

[711] Anonymous. Recently patented inventions. Scientific American, 101 (12):200, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
200.pdf.

Anonymous:1909:PNPbc

[712] Anonymous. Patent notes: Patents. Scientific American, 101 (12):200, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
200a.pdf.

Anonymous:1909:IIbc

[713] Anonymous. Index of inventions. Scientific American, 101(12): 200–203, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
200b.pdf.

Anonymous:1909:HWH

[714] Anonymous. Honor to whom honor is due. Scientific American, 101(12):190, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
190.pdf.

Anonymous:1909:RCO

[715] Anonymous. Report of the chief of ordnance. Scientific American, 101(12):190, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
190a.pdf.

Anonymous:1909:CUA

[716] Anonymous. The composition of the upper atmosphere. Scientific American, 101(12):190, September 18, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n12/
pdf/scientificamerican09181909-190b.pdf.

Anonymous:1909:Ebw

[717] Anonymous. Engineering. Scientific American, 101(12):191, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n12/pdf/scientificamerican09181909-191.pdf.

Anonymous:1909:Ebx

[718] Anonymous. Electricity. Scientific American, 101(12):191, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n12/pdf/scientificamerican09181909-191a.pdf.

Anonymous:1909:Sbg

[719] Anonymous. Science. Scientific American, 101(12):191, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n12/pdf/scientificamerican09181909-191b.pdf.

Anonymous:1909:AEF

[720] Anonymous. An automatic electric fog-signaling apparatus for railroads. Scientific American, 101(12):192, September 18, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n12/pdf/scientificamerican09181909-192.pdf.

Anonymous:1909:RDE

[721] Anonymous. Rat destruction by electric current. Scientific American, 101(12):192, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
192a.pdf.

Gradenwitz:1909:POA

[722] Alfred Gradenwitz. Portable outlooks for artillery trains. Scientific American, 101(12):193–194, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n12/
pdf/scientificamerican09181909-193.pdf.

Anonymous:1909:HP

[723] Anonymous. A heat paradox. Scientific American, 101(12):195, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n12/pdf/scientificamerican09181909-195a.pdf.

Anonymous:1909:ESI

[724] Anonymous. Early suggestions of invading England by balloon. Scientific American, 101(12):195, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n12/
pdf/scientificamerican09181909-195b.pdf.

Anonymous:1909:SLa

[725] Anonymous. Sheet lightning. Scientific American, 101(12):195, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n12/pdf/scientificamerican09181909-195c.pdf.

Anonymous:1909:LAT

[726] Anonymous. Latin-American trade. Scientific American, 101 (12):195, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
195d.pdf.

Anonymous:1909:TWE

[727] Anonymous. Trees as witnesses in early surveying. Scientific American, 101(12):195, September 18, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n12/
pdf/scientificamerican09181909-195e.pdf.

Anonymous:1909:CPD

[728] Anonymous. Commander Peary’s discovery of the North Pole. Scientific American, 101(12):196–197, September 18, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n12/pdf/scientificamerican09181909-196.pdf.

Anonymous:1909:HCM

[729] Anonymous. How cook made his latitude observations. Scientific American, 101(12):197–198, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n12/
pdf/scientificamerican09181909-197.pdf.

Anonymous:1909:CCB

[730] Anonymous. Capt Cody’s British Army aeroplane. Scientific American, 101(12):198, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
198.pdf.

Anonymous:1909:HMWn

[731] Anonymous. Handy man’s workshop. Scientific American, 101 (12):199, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
199.pdf.

Blake:1909:CPM

[732] Thaleon Blake. The cold process of making sirup. Scientific American, 101(12):199, September 18, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n12/
pdf/scientificamerican09181909-199a.pdf.

Searing:1909:VPP

[733] A. V. Searing, Jr. Vise for polished pipe. Scientific American, 101(12):199, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
199b.pdf.

Anonymous:1909:TRB

[734] Anonymous. Temporary repairs to broken springs. Scientific American, 101(12):199, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
199c.pdf.

Bishop:1909:STR

[735] Albert F. Bishop. Simple tapered reamer. Scientific American, 101(12):199, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
199d.pdf.

Mdermott:1909:CLR

[736] F. P. M’dermott, Jr. A cheap lamp rheostat. Scientific American, 101(12):199, September 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n12/pdf/scientificamerican09181909-
199e.pdf.

Anonymous:1909:Cbc

[737] Anonymous. Correspondence. Scientific American, 101(13):215, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n13/pdf/scientificamerican09251909-215.pdf.

Anonymous:1909:RPIbd

[738] Anonymous. Recently patented inventions. Scientific American, 101 (13):229–230, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
229.pdf.

Anonymous:1909:PNPbd

[739] Anonymous. Patent notes: Patents. Scientific American, 101 (13):231, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
231.pdf.

Anonymous:1909:IIbd

[740] Anonymous. Index of inventions. Scientific American, 101(13): 231–235, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
231a.pdf.

Anonymous:1909:WA

[741] Anonymous. The “wash” of an aeroplane. Scientific American, 101(13):208, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
208.pdf.

Anonymous:1909:RHC

[742] Anonymous. The return of Halley’s Comet. Scientific American, 101(13):208, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
208a.pdf.

Anonymous:1909:SPC

[743] Anonymous. The Selden Patent Case. Scientific American, 101 (13):208, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
208b.pdf.

Anonymous:1909:Eby

[744] Anonymous. Engineering. Scientific American, 101(13):209, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n13/pdf/scientificamerican09251909-209.pdf.

Anonymous:1909:Ebz

[745] Anonymous. Electricity. Scientific American, 101(13):209, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n13/pdf/scientificamerican09251909-209a.pdf.

Anonymous:1909:Sbh

[746] Anonymous. Science. Scientific American, 101(13):209, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n13/pdf/scientificamerican09251909-209b.pdf.

Anonymous:1909:HHE

[747] Anonymous. Hudson and his exploration of the Hudson River. Scientific American, 101(13):210–211, September 25, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n13/pdf/scientificamerican09251909-210.pdf.

Anonymous:1909:FPHa

[748] Anonymous. The four periods of Hudson River history. Scientific American, 101(13):211, September 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n13/
pdf/scientificamerican09251909-211.pdf.

Anonymous:1909:FHL

[749] Anonymous. Fulton and his life work. Scientific American, 101 (13):212–214, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
212.pdf.

Anonymous:1909:WHL

[750] Anonymous. What is heat lightning. Scientific American, 101 (13):215, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
215a.pdf.

Anonymous:1909:HFT

[751] Anonymous. How finger and thumb marks were discovered through a photographic proof. Scientific American, 101(13):215, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n13/pdf/scientificamerican09251909-215b.pdf.

Anonymous:1909:EFW

[752] Anonymous. Electrical fireballs: What are they? Scientific American, 101(13):215, September 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n13/
pdf/scientificamerican09251909-215c.pdf.

Anonymous:1909:DMO

[753] Anonymous. Does municipal ownership pay. Scientific American, 101(13):215, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
215d.pdf.

Anonymous:1909:SBi

[754] Anonymous. Supersensitiveness of the blind. Scientific American, 101(13):215, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
215e.pdf.

Anonymous:1909:DHR

[755] Anonymous. Development of the Hudson River steamboat. Scientific American, 101(13):216–217, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n13/
pdf/scientificamerican09251909-216.pdf.

Anonymous:1909:ASP

[756] Anonymous. American steamboats prior to the “Clermont”. Scientific American, 101(13):218, September 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n13/
pdf/scientificamerican09251909-218.pdf.

Anonymous:1909:JF

[757] Anonymous.  John Fitch.  Scientific American, 101(13):218–

219, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
218a.pdf.

Anonymous:1909:ESP

[758] Anonymous. European steamboats prior to the “Clermont”. Scientific American, 101(13):222, September 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n13/
pdf/scientificamerican09251909-222.pdf.

Anonymous:1909:WSb

[759] Anonymous. William Symington. Scientific American, 101(13): 222, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
222a.pdf.

Anonymous:1909:JS

[760] Anonymous. John Stevens. Scientific American, 101(13):222, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n13/pdf/scientificamerican09251909-222b.pdf.

Anonymous:1909:AWB

[761] Anonymous. The American walking-beam engine. Scientific American, 101(13):223, September 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n13/pdf/scientificamerican09251909-
223.pdf.

Anonymous:1909:THY

[762] Anonymous. Three hundred years of New York. Scientific American, 101(13):224–226, September 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n13/
pdf/scientificamerican09251909-224.pdf.

Russell:1909:HO

[763] Henry Norris Russell. The heavens in October. Scientific American, 101(14):246, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n14/pdf/scientificamerican10021909-
246.pdf.

Anonymous:1909:RPIbe

[764] Anonymous. Recently patented inventions. Scientific American, 101(14):248, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n14/pdf/scientificamerican10021909-
248.pdf.

Anonymous:1909:PNPbe

[765] Anonymous. Patent notes: Patents. Scientific American, 101(14):248, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-248a.pdf.

Anonymous:1909:IIbe

[766] Anonymous. Index of inventions. Scientific American, 101(14):248–251, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-248b.pdf.

Anonymous:1909:TSH

[767] Anonymous. True significance of the Hudson–Fulton Festival. Scientific American, 101(14):238, October 2, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n14/
pdf/scientificamerican10021909-238.pdf.

Anonymous:1909:OHY

[768] Anonymous. One hundred years of the marine engine. Scientific American, 101(14):238, October 2, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n14/
pdf/scientificamerican10021909-238a.pdf.

Anonymous:1909:CII

[769] Anonymous. Combating injurious insects with the aid of their natural enemies. Scientific American, 101(14):238, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n14/pdf/scientificamerican10021909-238b.pdf.

Anonymous:1909:Eca

[770] Anonymous. Engineering. Scientific American, 101(14):239, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-239.pdf.

Anonymous:1909:Am

[771] Anonymous. Aeronautics. Scientific American, 101(14):239, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-239a.pdf.

Anonymous:1909:Sbj

[772] Anonymous. Science. Scientific American, 101(14):239, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-239b.pdf.

Stetson:1909:HWG

[773] Harlan T. Stetson. How we get standard time. Scientific American, 101(14):240, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n14/pdf/scientificamerican10021909-
240.pdf.

Anonymous:1909:CSS

[774] Anonymous. Corrugated system of ship construction. Scientific American, 101(14):241, October 2, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n14/
pdf/scientificamerican10021909-241.pdf.

Anonymous:1909:GWC

[775] Anonymous. The Gabet wirelessly-controlled automatic torpedo. Scientific American, 101(14):241–242, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n14/
pdf/scientificamerican10021909-241a.pdf.

Anonymous:1909:ZIA

[776] Anonymous. The “Zeppelin III.” airship and its trip to Berlin. Scientific American, 101(14):242–243, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n14/
pdf/scientificamerican10021909-242.pdf.

Watkins:1909:MI

[777] John Elfreth Watkins. The modern Icarus. Scientific American, 101(14):243–245, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n14/pdf/scientificamerican10021909-
243.pdf.

Anonymous:1909:PDf

[778] Anonymous. Patent department. Scientific American, 101(14):247, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-247.pdf.

Anonymous:1909:SRR

[779] Anonymous. Sanitary refuse receptacle. Scientific American, 101 (14):247, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n14/pdf/scientificamerican10021909-
247a.pdf.

Anonymous:1909:CBH

[780] Anonymous. Convenient book holder. Scientific American, 101(14):247, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-247b.pdf.

Anonymous:1909:ICH

[781] Anonymous. Improved cutter head. Scientific American, 101(14):247, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-247c.pdf.

Anonymous:1909:RF

[782] Anonymous. Rail fastener. Scientific American, 101(14):247, October 2, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n14/pdf/scientificamerican10021909-247d.pdf.

Anonymous:1909:Cbd

[783] Anonymous. Correspondence. Scientific American, 101(15):259, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-259.pdf.

Anonymous:1909:RPIbf

[784] Anonymous. Recently patented inventions. Scientific American, 101(15):265, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
265.pdf.

Anonymous:1909:NQt

[785] Anonymous. Notes and queries. Scientific American, 101(15):266, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-266.pdf.

Anonymous:1909:BNBr

[786] Anonymous. Books: New books, etc. Scientific American, 101 (15):266–267, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
266a.pdf.

Anonymous:1909:PNPbf

[787] Anonymous. Patent notes: Patents. Scientific American, 101(15):267, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-267.pdf.

Anonymous:1909:IIbf

[788] Anonymous. Index of inventions. Scientific American, 101(15):267–269, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-267a.pdf.

Anonymous:1909:PDR

[789] Anonymous. Philadelphia, the Delaware River, and John Fitch. Scientific American, 101(15):254, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-254.pdf.

Anonymous:1909:PSO

[790] Anonymous. Palestine and Syria the original cereal countries. Scientific American, 101(15):254, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-254a.pdf.

Anonymous:1909:HGS

[791] Anonymous. Home-grown sugar-beet seed. Scientific American, 101(15):254, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
254b.pdf.

Anonymous:1909:PMb

[792] Anonymous. The platinum market. Scientific American, 101(15):254, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-254c.pdf.

Anonymous:1909:Ecb

[793] Anonymous. Engineering. Scientific American, 101(15):255, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-255.pdf.

Anonymous:1909:Ao

[794] Anonymous. Aeronautics. Scientific American, 101(15):255, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-255a.pdf.

Anonymous:1909:Sbk

[795] Anonymous. Science. Scientific American, 101(15):255, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-255b.pdf.

Anonymous:1909:SMPa

[796] Anonymous. Stereoscopic moving pictures in natural colors. Scientific American, 101(15):256, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-256.pdf.

Anonymous:1909:EFV

[797] Anonymous. Erection of the fades viaduct. Scientific American, 101(15):257–258, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
257.pdf.

Anonymous:1909:PC

[798] Anonymous. The Pope and the comet. Scientific American, 101(15):259, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
259a.pdf.

Anonymous:1909:NOAa

[799] Anonymous. The number of our ancestors. Scientific American, 101(15):259, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
259b.pdf.

Anonymous:1909:NOAb

[800] Anonymous. The number of our ancestors and of our future cousins. Scientific American, 101(15):259, October 9, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n15/pdf/scientificamerican10091909-259c.pdf.

Anonymous:1909:WDMb

[801] Anonymous. Why do main springs break? Scientific American, 101(15):259, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
259d.pdf.

Mclure:1909:OAO

[802] W. Frank M’clure. Open-air orchard heating in Colorado. Scientific American, 101(15):260–261, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-260.pdf.

Anonymous:1909:VWC

[803] Anonymous. The visiting warships — a comparison. Scientific American, 101(15):261–262, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-261.pdf.

Anonymous:1909:NAP

[804] Anonymous. A novel air pump and vacuum gage. Scientific American, 101(15):262, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-262.pdf.

Swezey:1909:HMW

[805] C. L. Swezey. Handy man’s workshop. Scientific American, 101(15):263, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-263.pdf.

Johns:1909:TRS

[806] B. A. Johns. Thermostatic regulator for smoke pipes. Scientific American, 101(15):263, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-263a.pdf.

Bergstrom:1909:CPH

[807] J. A. Bergstrom. Circulating pipe for hot-water faucets. Scientific American, 101(15):263, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-263b.pdf.

L:1909:HCS

[808] H. G. L. How to connect stove pipes. Scientific American, 101(15):263, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-263c.pdf.

Allen:1909:HWC

[809] John B. Allen. Hot-water connection for kitchen boilers. Scientific American, 101(15):263–264, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-263d.pdf.

Brophy:1909:SR

[810] J. A. Brophy. A screen for the register. Scientific American, 101(15):264, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-264.pdf.

Bishop:1909:BMT

[811] A. F. Bishop. The boiler maker’s tool bag. Scientific American, 101(15):264, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
264a.pdf.

Jarman:1909:CRW

[812] A. J. Jarman. Coating roofs white to repel heat rays. Scientific American, 101(15):264, October 9, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n15/
pdf/scientificamerican10091909-264b.pdf.

Morgan:1909:MGE

[813] Alfred P. Morgan. A muffler for gas engines. Scientific American, 101(15):264, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n15/pdf/scientificamerican10091909-
264c.pdf.

Handy:1909:PCG

[814] L. G. Handy. To prevent coal gas. Scientific American, 101(15):264, October 9, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n15/pdf/scientificamerican10091909-264d.pdf.

Anonymous:1909:Cbe

[815] Anonymous. Correspondence. Scientific American, 101(16):278, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n16/pdf/scientificamerican10161909-278.pdf.

Anonymous:1909:RPIbg

[816] Anonymous. Recently patented inventions. Scientific American, 101(16):284, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
284.pdf.

Anonymous:1909:BNBs

[817] Anonymous. Books: New books, etc. Scientific American, 101 (16):285, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
285.pdf.

Anonymous:1909:PNPbg

[818] Anonymous. Patent notes: Patents. Scientific American, 101 (16):285, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
285a.pdf.

Anonymous:1909:IIbg

[819] Anonymous. Index of inventions. Scientific American, 101(16): 285–287, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
285b.pdf.

Anonymous:1909:EVH

[820] Anonymous. Educational value of the Hudson–Fulton Celebration. Scientific American, 101(16):274, October 16, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n16/pdf/scientificamerican10161909-274.pdf.

Anonymous:1909:ERa

[821] Anonymous. Evolution not revolution. Scientific American, 101 (16):274, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
274a.pdf.

Anonymous:1909:MPN

[822] Anonymous. The most powerful Naval gun. Scientific American, 101(16):274, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
274b.pdf.

Anonymous:1909:RFO

[823] Anonymous. The record flights of Orville and Wilbur Wright. Scientific American, 101(16):274, October 16, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n16/
pdf/scientificamerican10161909-274c.pdf.

Anonymous:1909:Ecc

[824] Anonymous. Engineering. Scientific American, 101(16):275, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n16/pdf/scientificamerican10161909-275.pdf.

Anonymous:1909:Ecd

[825] Anonymous. Electricity. Scientific American, 101(16):275, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n16/pdf/scientificamerican10161909-275a.pdf.

Anonymous:1909:Sbl

[826] Anonymous. Science. Scientific American, 101(16):275, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n16/pdf/scientificamerican10161909-275b.pdf.

Anonymous:1909:CTB

[827] Anonymous. Cycling trick in a Berlin variety theater. Scientific American, 101(16):276, October 16, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n16/
pdf/scientificamerican10161909-276.pdf.

Anonymous:1909:FPHb

[828] Anonymous. The first photograph of Halley’s Comet. Scientific American, 101(16):276, October 16, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n16/
pdf/scientificamerican10161909-276a.pdf.

Middleton:1909:SAL

[829] P. Harvey Middleton. Saving the American lobster. Scientific American, 101(16):277–278, October 16, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n16/
pdf/scientificamerican10161909-277.pdf.

Anonymous:1909:NOAc

[830] Anonymous. The number of our ancestors. Scientific American, 101(16):278, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
278a.pdf.

Arthur:1909:SSA

[831] James Arthur. Strange scientific anomalies at the North Pole. Scientific American, 101(16):279, October 16, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n16/
pdf/scientificamerican10161909-279.pdf.

Anonymous:1909:PCC

[832] Anonymous. Playing a ’cello by compressed air. Scientific American, 101(16):280, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
280.pdf.

Anonymous:1909:MMU

[833] Anonymous. Moving a mountain in Utah. Scientific American, 101 (16):281–282, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
281.pdf.

Anonymous:1909:NAE

[834] Anonymous. A new automatic electric signaling system for railroads. Scientific American, 101(16):282, October 16, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n16/pdf/scientificamerican10161909-282.pdf.

Baltimore:1909:VAN

[835] J. Mayne Baltimore. The vessel that accomplished the Northwest Passage. Scientific American, 101(16):283, October 16, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n16/pdf/scientificamerican10161909-283.pdf.

Holder:1909:TN

[836] Charles F. Holder. A tame nautilus. Scientific American, 101 (16):283, October 16, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n16/pdf/scientificamerican10161909-
283a.pdf.

Anonymous:1909:Cbf

[837] Anonymous. Correspondence. Scientific American, 101(17):295, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n17/pdf/scientificamerican10231909-295.pdf.

Anonymous:1909:RPIbh

[838] Anonymous. Recently patented inventions. Scientific American, 101(17):301, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
301.pdf.

Anonymous:1909:NQu

[839] Anonymous. Notes and queries. Scientific American, 101(17):301, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n17/pdf/scientificamerican10231909-301a.pdf.

Anonymous:1909:BNBt

[840] Anonymous. Books: New books, etc. Scientific American, 101 (17):301–302, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
301b.pdf.

Anonymous:1909:PNPbh

[841] Anonymous. Patent notes: Patents. Scientific American, 101 (17):302, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
302.pdf.

Anonymous:1909:IIbh

[842] Anonymous. Index of inventions. Scientific American, 101(17):302, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n17/pdf/scientificamerican10231909-302a.pdf.

Anonymous:1909:TWW

[843] Anonymous. A talk with Wilbur Wright. Scientific American, 101(17):290, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
290.pdf.

Anonymous:1909:NBD

[844] Anonymous. The new British “Dreadnoughts” and “Inflexibles”. Scientific American, 101(17):290, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-290a.pdf.

Anonymous:1909:SPR

[845] Anonymous. Some photochemical reactions. Scientific American, 101(17):290, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
290b.pdf.

Anonymous:1909:SMe

[846] Anonymous. The spectrum of Mars. Scientific American, 101 (17):290, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
290c.pdf.

Anonymous:1909:Ap

[847] Anonymous. Aeronautics. Scientific American, 101(17):291, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n17/pdf/scientificamerican10231909-291.pdf.

Anonymous:1909:Ece

[848] Anonymous. Electricity. Scientific American, 101(17):291, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n17/pdf/scientificamerican10231909-291a.pdf.

Anonymous:1909:Sbm

[849] Anonymous. Science. Scientific American, 101(17):291, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n17/pdf/scientificamerican10231909-291b.pdf.

Anonymous:1909:FGM

[850] Anonymous. The first German monoplane to make successful flights. Scientific American, 101(17):292, October 23, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n17/pdf/scientificamerican10231909-292.pdf.

Anonymous:1909:AMB

[851] Anonymous. The aviation meet at Berlin and Latham’s flight across the City. Scientific American, 101(17):292, October 23, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n17/pdf/scientificamerican10231909-292a.pdf.

Boyer:1909:SMT

[852] Jacques Boyer. A simplified method of teaching experimental physics. Scientific American, 101(17):293–294, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n17/pdf/scientificamerican10231909-293.pdf.

Honey:1909:SHR

[853] Frederic R. Honey. Saturn and his rings. Scientific American, 101(17):294, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
294.pdf.

Anonymous:1909:HMR

[854] Anonymous. The height of Mount Rainier. Scientific American, 101(17):295, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
295a.pdf.

Anonymous:1909:SAJ

[855] Anonymous. The Scientific American at Jericho. Scientific American, 101(17):295, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
295b.pdf.

Anonymous:1909:WMF

[856] Anonymous. Why not a monument to Fitch? Scientific American, 101(17):295, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
295c.pdf.

Anonymous:1909:ICC

[857] Anonymous. “Inflexible” and “Connecticut” compared. Scientific American, 101(17):295, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-295d.pdf.

Anonymous:1909:RLS

[858] Anonymous. Result of a lightning stroke. Scientific American, 101(17):295, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
295e.pdf.

Anonymous:1909:NOAd

[859] Anonymous. The number of our ancestors. Scientific American, 101(17):295, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
295f.pdf.

Anonymous:1909:HHA

[860] Anonymous. The highest human ascent. Scientific American, 101 (17):295, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
295g.pdf.

Anonymous:1909:AWc

[861] Anonymous. Aerial warfare in 1798. Scientific American, 101 (17):295, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
295h.pdf.

Anonymous:1909:LSU

[862] Anonymous. The latest submarines of the United States Navy. Scientific American, 101(17):296, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-296.pdf.

Anonymous:1909:ASE

[863] Anonymous. An all seeing eye for the submarine. Scientific American, 101(17):297, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-297.pdf.

Anonymous:1909:CJJb

[864] Anonymous. Col. John Jacob Astor’s novel steamship chair. Scientific American, 101(17):298, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-298.pdf.

Wood:1909:HMW

[865] R. W. Wood. Handy man’s workshop. Scientific American, 101 (17):299, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
299.pdf.

Ward:1909:ACE

[866] Frederick E. Ward. Alternating current experiments. Scientific American, 101(17):299–300, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-299a.pdf.

Paterson:1909:DRW

[867] James Paterson. Device for removing wood screws. Scientific American, 101(17):300, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-300.pdf.

Nichols:1909:MBM

[868] Howard W. Nichols. Mending a broken metallic filament. Scientific American, 101(17):300, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-300a.pdf.

Kaiser:1909:SFC

[869] F. A. Kaiser. A simple fruit cellar. Scientific American, 101(17):300, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n17/pdf/scientificamerican10231909-300b.pdf.

Mgahey:1909:HMP

[870] C. R. M’gahey. How to make a paper telescope barrel. Scientific American, 101(17):300, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-300c.pdf.

Hanson:1909:RDC

[871] Edward M. Hanson. Restoring a dry cell. Scientific American, 101(17):300, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
300d.pdf.

Mencken:1909:HMSb

[872] August Mencken. The handy man’s sub-caliber gun. Scientific American, 101(17):300, October 23, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n17/
pdf/scientificamerican10231909-300e.pdf.

Fairbank:1909:DET

[873] C. W. Fairbank. A device for emptying a tub. Scientific American, 101(17):300, October 23, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n17/pdf/scientificamerican10231909-
300f.pdf.

Anonymous:1909:Cbg

[874] Anonymous. Correspondence. Scientific American, 101(18):315, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-315.pdf.

Russell:1909:HN

[875] Henry Norris Russell. The heavens in November. Scientific American, 101(18):318, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
318.pdf.

Anonymous:1909:RPIbi

[876] Anonymous. Recently patented inventions. Scientific American, 101(18):320, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
320.pdf.

Anonymous:1909:BNBu

[877] Anonymous. Books: New books, etc. Scientific American, 101 (18):320–321, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
320a.pdf.

Anonymous:1909:PNPbi

[878] Anonymous. Patent notes: Patents. Scientific American, 101 (18):321, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
321.pdf.

Anonymous:1909:IIbi

[879] Anonymous. Index of inventions. Scientific American, 101(18): 321–323, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
321a.pdf.

Anonymous:1909:SCBb

[880] Anonymous. Slag a cause of broken rails. Scientific American, 101(18):310, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
310.pdf.

Anonymous:1909:OVT

[881] Anonymous. Our vanishing timber supply. Scientific American, 101(18):310, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
310a.pdf.

Anonymous:1909:DPL

[882] Anonymous. Death of Prof. Lombroso. Scientific American, 101 (18):310, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
310b.pdf.

Anonymous:1909:CLFb

[883] Anonymous. Count De Lambert’s flight over the Eiffel Tower. Scientific American, 101(18):310, October 30, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n18/
pdf/scientificamerican10301909-310c.pdf.

Anonymous:1909:Ecf

[884] Anonymous. Engineering. Scientific American, 101(18):311, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-311.pdf.

Anonymous:1909:Ecg

[885] Anonymous. Electricity. Scientific American, 101(18):311, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-311a.pdf.

Anonymous:1909:Sbn

[886] Anonymous. Science. Scientific American, 101(18):311, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-311b.pdf.

Gradenwitz:1909:BPE

[887] Alfred Gradenwitz. The Brussels Popular Electrical Laboratory. Scientific American, 101(18):312, October 30, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n18/
pdf/scientificamerican10301909-312.pdf.

Anonymous:1909:CCG

[888] Anonymous. A curious chemical growth. Scientific American, 101(18):313, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
313.pdf.

Anonymous:1909:PF

[889] Anonymous. Playing with fire. Scientific American, 101(18):313, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-313a.pdf.

Anonymous:1909:MHE

[890] Anonymous. Measuring holes by electricity. Scientific American, 101(18):313, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
313b.pdf.

Anonymous:1909:DSI

[891] Anonymous. Differential speed indicator. Scientific American, 101 (18):313, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
313c.pdf.

Anonymous:1909:APa

[892] Anonymous. Accident preventer. Scientific American, 101(18):313, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-313d.pdf.

Anonymous:1909:STa

[893] Anonymous. A surveyor’s telephoto. Scientific American, 101 (18):313, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
313e.pdf.

Anonymous:1909:HSD

[894] Anonymous. High-speed destroyers for the United States Navy. Scientific American, 101(18):314, October 30, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n18/
pdf/scientificamerican10301909-314.pdf.

Anonymous:1909:LF

[895] Anonymous. Leduc’s figures. Scientific American, 101(18):315, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-315a.pdf.

Anonymous:1909:WDWb

[896] Anonymous. Why do watch springs break? Scientific American, 101(18):315, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
315b.pdf.

Anonymous:1909:MS

[897] Anonymous. A magic square. Scientific American, 101(18):315, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-315c.pdf.

Anonymous:1909:BMS

[898] Anonymous. The breaking of main springs in storms. Scientific American, 101(18):315, October 30, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n18/
pdf/scientificamerican10301909-315d.pdf.

Anonymous:1909:OP

[899] Anonymous. An optical phenomenon. Scientific American, 101 (18):315, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
315e.pdf.

Warren:1909:HGM

[900] Frederic Blount Warren. How Germany makes forestry pay. Scientific American, 101(18):316–317, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n18/
pdf/scientificamerican10301909-316.pdf.

Anonymous:1909:PDg

[901] Anonymous. Patent department. Scientific American, 101(18):319, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n18/pdf/scientificamerican10301909-319.pdf.

Anonymous:1909:SRA

[902] Anonymous. Sleigh-runner attachment for baby carriages. Scientific American, 101(18):319, October 30, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n18/
pdf/scientificamerican10301909-319a.pdf.

Anonymous:1909:CFJ

[903] Anonymous. Combined friction and jaw clutch. Scientific American, 101(18):319, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
319b.pdf.

Anonymous:1909:PFL

[904] Anonymous. A perfected form of lock nut. Scientific American, 101(18):319, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
319c.pdf.

Anonymous:1909:OId

[905] Anonymous. Oddities in invention. Scientific American, 101 (18):319, October 30, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n18/pdf/scientificamerican10301909-
319d.pdf.

Anonymous:1909:Cbi

[906] Anonymous. Correspondence. Scientific American, 101(19):331, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n19/pdf/scientificamerican11061909-331.pdf.

Anonymous:1909:RPIbj

[907] Anonymous. Recently patented inventions. Scientific American, 101(19):337, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
337.pdf.

Anonymous:1909:BNBv

[908] Anonymous. Books: New books, etc. Scientific American, 101 (19):337–338, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
337a.pdf.

Anonymous:1909:PNPbj

[909] Anonymous. Patent notes: Patents. Scientific American, 101 (19):338, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
338.pdf.

Anonymous:1909:IIbj

[910] Anonymous. Index of inventions. Scientific American, 101(19): 338–340, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
338a.pdf.

Anonymous:1909:ACa

[911] Anonymous. The America’s Cup. Scientific American, 101(19): 326, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
326.pdf.

Anonymous:1909:DGA

[912] Anonymous. Dover’s great artificial harbor. Scientific American, 101(19):326, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
326a.pdf.

Anonymous:1909:INY

[913] Anonymous. Improvements in New York’s High-Pressure Fire Service. Scientific American, 101(19):326, November 6, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n19/pdf/scientificamerican11061909-326b.pdf.

Anonymous:1909:Eci

[914] Anonymous. Engineering. Scientific American, 101(19):327, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n19/pdf/scientificamerican11061909-327.pdf.

Anonymous:1909:Ecj

[915] Anonymous. Electricity. Scientific American, 101(19):327, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n19/pdf/scientificamerican11061909-327a.pdf.

Anonymous:1909:Sbo

[916] Anonymous. Science. Scientific American, 101(19):327, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n19/pdf/scientificamerican11061909-327b.pdf.

Perkins:1909:RMC

[917] Frank C. Perkins. A remarkable Mexican concrete conduit at Guadalajara. Scientific American, 101(19):328, November 6, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n19/pdf/scientificamerican11061909-328.pdf.

Anonymous:1909:TBE

[918] Anonymous.  Terrific boiler explosion in Milwaukee.  Scien-

tific American, 101(19):329, November 6, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n19/
pdf/scientificamerican11061909-329.pdf.

Springer:1909:SPB

[919] J. F. Springer. Some principles of ball-bearing design. Scientific American, 101(19):329–330, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n19/
pdf/scientificamerican11061909-329a.pdf.

Anonymous:1909:CSMa

[920] Anonymous. A curious star map. Scientific American, 101(19): 330, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
330.pdf.

Anonymous:1909:NOAe

[921] Anonymous. The number of our ancestors. Scientific American, 101(19):331, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
331a.pdf.

Anonymous:1909:TSD

[922] Anonymous. The 10,000-ton suction dredger “Leviathan” for use on the Mersey. Scientific American, 101(19):332, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n19/pdf/scientificamerican11061909-332.pdf.

Anonymous:1909:SCI

[923] Anonymous. Some curiosities of invention. Scientific American, 101(19):332, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
332a.pdf.

Anonymous:1909:ADb

[924] Anonymous. The attitude of diplodocus. Scientific American, 101 (19):334, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
334.pdf.

Bowdish:1909:BP

[925] B. S. Bowdish. Birds of passage. Scientific American, 101(19): 335, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
335.pdf.

Anonymous:1909:OD

[926] Anonymous. The Ophthalmo-Diaphanoscope. Scientific American, 101(19):336, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
336.pdf.

Anonymous:1909:TBP

[927] Anonymous. Testing bitumens for pavements. Scientific American, 101(19):336, November 6, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n19/pdf/scientificamerican11061909-
336a.pdf.

Anonymous:1909:RPIbk

[928] Anonymous. Recently patented inventions. Scientific American, 101(2):32, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
32.pdf.

Anonymous:1909:BNBw

[929] Anonymous. Books: New books, etc. Scientific American, 101(2):32, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n2/pdf/scientificamerican07101909-32a.pdf.

Anonymous:1909:PNPbk

[930] Anonymous. Patent notes: Patents. Scientific American, 101(2):32, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n2/pdf/scientificamerican07101909-32b.pdf.

Anonymous:1909:IIbk

[931] Anonymous. Index of inventions. Scientific American, 101(2):32, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n2/pdf/scientificamerican07101909-32c.pdf.

Anonymous:1909:ELF

[932] Anonymous. The electric locomotive of the future. Scientific American, 101(2):22, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
22.pdf.

Anonymous:1909:DOB

[933] Anonymous. Drydocks for our biggest battleships. Scientific American, 101(2):22, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
22a.pdf.

Anonymous:1909:RW

[934] Anonymous. The railways of the world. Scientific American, 101(2):22, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n2/pdf/scientificamerican07101909-22b.pdf.

Anonymous:1909:DSB

[935] Anonymous. The directive sense of bees. Scientific American, 101(2):22, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n2/pdf/scientificamerican07101909-22c.pdf.

Anonymous:1909:Eck

[936] Anonymous. Engineering. Scientific American, 101(2):23, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n2/pdf/scientificamerican07101909-23.pdf.

Anonymous:1909:Ecl

[937] Anonymous. Electricity. Scientific American, 101(2):23, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n2/pdf/scientificamerican07101909-23a.pdf.

Anonymous:1909:Sbp

[938] Anonymous. Science. Scientific American, 101(2):23, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n2/pdf/scientificamerican07101909-23b.pdf.

Kirkwood:1909:GRI

[939] J. E. Kirkwood. The Guayule rubber industry. Scientific American, 101(2):24, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
24.pdf.

Wade:1909:SEC

[940] Herbert T. Wade. The standard electrical cell — its use and its practical value. Scientific American, 101(2):25–26, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n2/pdf/scientificamerican07101909-25.pdf.

Anonymous:1909:TTM

[941] Anonymous. A 5,000-ton testing machine. Scientific American, 101(2):26, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
26.pdf.

Anonymous:1909:FDSb

[942] Anonymous. The fourth dimension simply explained. Scientific American, 101(2):27, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
27.pdf.

Smith:1909:CSM

[943] J. Hammond Smith. Carving statues by means of photography. Scientific American, 101(2):28, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
28.pdf.

Anonymous:1909:FFE

[944] Anonymous. The First Flight Exhibition and Tournament of the Aeronautic Society. Scientific American, 101(2):29, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n2/pdf/scientificamerican07101909-29.pdf.

Hodgins:1909:HP

[945] George Sherwood Hodgins. A huge planetarium. Scientific American, 101(2):30, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
30.pdf.

Anonymous:1909:PDh

[946] Anonymous. Patent department. Scientific American, 101(2):31, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n2/pdf/scientificamerican07101909-31.pdf.

Anonymous:1909:IHS

[947] Anonymous. Improved head for stringed musical instruments. Scientific American, 101(2):31, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
31a.pdf.

Anonymous:1909:PIB

[948] Anonymous. Padlock with incased bolt. Scientific American, 101(2):31, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n2/pdf/scientificamerican07101909-31b.pdf.

Anonymous:1909:MRB

[949] Anonymous. Metal rail base and cross tie for railway tracks. Scientific American, 101(2):31, July 10, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n2/pdf/scientificamerican07101909-
31c.pdf.

Anonymous:1909:Cbj

[950] Anonymous. Correspondence. Scientific American, 101(20):351, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n20/pdf/scientificamerican11131909-351.pdf.

Anonymous:1909:RPIbl

[951] Anonymous. Recently patented inventions. Scientific American, 101(20):357, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
357.pdf.

Anonymous:1909:NQv

[952] Anonymous. Notes and queries. Scientific American, 101(20): 357–358, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
357a.pdf.

Anonymous:1909:BNBx

[953] Anonymous. Books: New books, etc. Scientific American, 101 (20):358, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
358.pdf.

Anonymous:1909:PNPbl

[954] Anonymous. Patent notes: Patents. Scientific American, 101 (20):358, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
358a.pdf.

Anonymous:1909:IIbl

[955] Anonymous. Index of inventions. Scientific American, 101(20): 358–360, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
358b.pdf.

Anonymous:1909:LMD

[956] Anonymous. Local magnetic disturbances and the genesis of petroleum. Scientific American, 101(20):346, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n20/
pdf/scientificamerican11131909-346.pdf.

Anonymous:1909:FVC

[957] Anonymous. The Fifth Vanderbilt Cup Race and the Good Roads Tour from New York to Atlanta. Scientific American, 101(20):346, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n20/pdf/scientificamerican11131909-346a.pdf.

Anonymous:1909:NTC

[958] Anonymous. Non-toxic coffee and tea. Scientific American, 101 (20):346, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
346b.pdf.

Anonymous:1909:Ecn

[959] Anonymous. Engineering. Scientific American, 101(20):347, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n20/pdf/scientificamerican11131909-347.pdf.

Anonymous:1909:Eco

[960] Anonymous. Electricity. Scientific American, 101(20):347, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n20/pdf/scientificamerican11131909-347a.pdf.

Anonymous:1909:Sbq

[961] Anonymous. Science. Scientific American, 101(20):347, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n20/pdf/scientificamerican11131909-347b.pdf.

Worthington:1909:ARS

[962] Ferdinand Worthington. Archaeological relics from the sea. Scientific American, 101(20):348–349, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n20/
pdf/scientificamerican11131909-348.pdf.

Claudy:1909:OAO

[963] C. H. Claudy. Our aeronautical organization. Scientific American, 101(20):350, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
350.pdf.

Anonymous:1909:MBW

[964] Anonymous. Magnetism and the breaking of watch springs. Scientific American, 101(20):351, November 13, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n20/
pdf/scientificamerican11131909-351a.pdf.

Anonymous:1909:RCS

[965] Anonymous. The Renaissance clergy and science. Scientific American, 101(20):351, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
351b.pdf.

Anonymous:1909:MAW

[966] Anonymous. Making alcohol from wood waste. Scientific American, 101(20):352, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
352.pdf.

Anonymous:1909:TND

[967] Anonymous. Trials of the new “Dreadnought” — “North Dakota”. Scientific American, 101(20):352–354, November 13, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n20/pdf/scientificamerican11131909-352a.pdf.

Anonymous:1909:RFC

[968] Anonymous. The record flights of Count De Lambert and Henry Farman. Scientific American, 101(20):354, November 13, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n20/pdf/scientificamerican11131909-354.pdf.

Browne:1909:HMW

[969] F. Willard Browne. Handy man’s workshop. Scientific American, 101 (20):355–356, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
355.pdf.

Jarman:1909:ECM

[970] A. J. Jarman. Experimental color making with simple chemical solutions. Scientific American, 101(20):356, November 13, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n20/pdf/scientificamerican11131909-356.pdf.

Bourne:1909:IP

[971] C. S. Bourne. An interesting paradox. Scientific American, 101 (20):356, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
356a.pdf.

Brecht:1909:FC

[972] Charles Brecht. Filter for cisterns. Scientific American, 101 (20):356, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
356b.pdf.

Davis:1909:CCS

[973] Herbert S. Davis. A close-corner screw driver. Scientific American, 101(20):356, November 13, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n20/pdf/scientificamerican11131909-
356c.pdf.

Anonymous:1909:Cbk

[974] Anonymous. Correspondence. Scientific American, 101(21):371, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n21/pdf/scientificamerican11201909-371.pdf.

Anonymous:1909:RPIbm

[975] Anonymous. Recently patented inventions. Scientific American, 101(21):377, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
377.pdf.

Anonymous:1909:NQw

[976] Anonymous. Notes and queries. Scientific American, 101(21): 377, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
377a.pdf.

Anonymous:1909:PNPbm

[977] Anonymous. Patent notes: Patents. Scientific American, 101 (21):377, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
377b.pdf.

Anonymous:1909:IIbm

[978] Anonymous. Index of inventions. Scientific American, 101(21): 377–383, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
377c.pdf.

Anonymous:1909:BNBy

[979] Anonymous. Books: New books, etc. Scientific American, 101 (21):379–380, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
379.pdf.

Anonymous:1909:PCW

[980] Anonymous. Purity of Catskill Water Supply. Scientific American, 101(21):366, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
366.pdf.

Anonymous:1909:NFI

[981] Anonymous. The new fourteen-inch Army and Navy guns. Scientific American, 101(21):366, November 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n21/
pdf/scientificamerican11201909-366a.pdf.

Anonymous:1909:NBG

[982] Anonymous. The new British and German “Dreadnoughts”. Scientific American, 101(21):366, November 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n21/
pdf/scientificamerican11201909-366b.pdf.

Anonymous:1909:ALD

[983] Anonymous. Acceleration of long-distance express service. Scientific American, 101(21):366, November 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n21/
pdf/scientificamerican11201909-366c.pdf.

Anonymous:1909:LRT

[984] Anonymous. The limit of rapid transit. Scientific American, 101 (21):366, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
366d.pdf.

Anonymous:1909:OHC

[985] Anonymous. The other half of the Cullinan. Scientific American, 101(21):366, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
366e.pdf.

Anonymous:1909:Ecp

[986] Anonymous. Engineering. Scientific American, 101(21):367, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n21/pdf/scientificamerican11201909-367.pdf.

Anonymous:1909:Ecq

[987] Anonymous. Electricity. Scientific American, 101(21):367, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n21/pdf/scientificamerican11201909-367a.pdf.

Anonymous:1909:Sbr

[988] Anonymous. Science. Scientific American, 101(21):367, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n21/pdf/scientificamerican11201909-367b.pdf.

Hunting:1909:SNC

[989] M. H. Hunting. Something new in concrete block making. Scientific American, 101(21):368, November 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n21/
pdf/scientificamerican11201909-368.pdf.

Anonymous:1909:D

[990] Anonymous. A “Dreadnought” of 1863. Scientific American, 101 (21):368–369, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
368a.pdf.

Anonymous:1909:PII

[991] Anonymous. Power indicator for internal-combustion engines. Scientific American, 101(21):369–370, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n21/
pdf/scientificamerican11201909-369.pdf.

Honey:1909:SLE

[992] Frederic R. Honey. The solar and lunar eclipses of 1909. Scientific American, 101(21):370–371, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n21/
pdf/scientificamerican11201909-370.pdf.

Anonymous:1909:WWS

[993] Anonymous. Why watch springs break. Scientific American, 101 (21):371, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
371a.pdf.

Anonymous:1909:RMC

[994] Anonymous. Railway motor cars. Scientific American, 101(21): 371, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
371b.pdf.

Anonymous:1909:AED

[995] Anonymous. Another evil of deforestation. Scientific American, 101(21):371, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
371c.pdf.

Bauskett:1909:HLO

[996] Frank N. Bauskett. How lemon oil is made. Scientific American, 101(21):372, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
372.pdf.

Anonymous:1909:AAAa

[997] Anonymous. Artillery for airship attack. Scientific American, 101 (21):373, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
373.pdf.

Anonymous:1909:WCB

[998] Anonymous.  Wireless communication with balloons.  Scien-

tific American, 101(21):374, November 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n21/
pdf/scientificamerican11201909-374.pdf.

Anonymous:1909:WTR

[999] Anonymous. The Wellcome Tropical Research Laboratories. Scientific American, 101(21):375, November 20, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n21/
pdf/scientificamerican11201909-375.pdf.

Anonymous:1909:PDi

[1000] Anonymous. Patent department. Scientific American, 101(21): 376, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
376.pdf.

Anonymous:1909:CTD

[1001] Anonymous. Combined timer and distributer for internal-combustion engines. Scientific American, 101(21):376, November 20, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n21/pdf/scientificamerican11201909-376a.pdf.

Anonymous:1909:IST

[1002] Anonymous. An improved steam trap. Scientific American, 101 (21):376, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
376b.pdf.

Anonymous:1909:NGC

[1003] Anonymous. A new grain bin construction. Scientific American, 101(21):376, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
376c.pdf.

Anonymous:1909:OIe

[1004] Anonymous. Oddities in invention. Scientific American, 101 (21):376, November 20, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n21/pdf/scientificamerican11201909-
376d.pdf.

Anonymous:1909:Cbl

[1005] Anonymous. Correspondence. Scientific American, 101(22):391, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n22/pdf/scientificamerican11271909-391.pdf.

Anonymous:1909:RPIbn

[1006] Anonymous. Recently patented inventions. Scientific American, 101(22):397, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
397.pdf.

Anonymous:1909:NQx

[1007] Anonymous. Notes and queries. Scientific American, 101(22): 397, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
397a.pdf.

Anonymous:1909:BNBz

[1008] Anonymous. Books: New books, etc. Scientific American, 101 (22):398, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
398.pdf.

Anonymous:1909:PNPbn

[1009] Anonymous. Patent notes: Patents. Scientific American, 101 (22):398, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
398a.pdf.

Anonymous:1909:IIbn

[1010] Anonymous. Index of inventions. Scientific American, 101(22): 398–403, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
398b.pdf.

Anonymous:1909:PCP

[1011] Anonymous. Present conditions on the Panama Canal. Scientific American, 101(22):386, November 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n22/
pdf/scientificamerican11271909-386.pdf.

Anonymous:1909:TBNb

[1012] Anonymous. The trials of the battleship “North Dakota”. Scientific American, 101(22):386, November 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n22/
pdf/scientificamerican11271909-386a.pdf.

Anonymous:1909:Ecr

[1013] Anonymous. Electricity. Scientific American, 101(22):387, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n22/pdf/scientificamerican11271909-387.pdf.

Anonymous:1909:Sbs

[1014] Anonymous. Science. Scientific American, 101(22):387, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n22/pdf/scientificamerican11271909-387a.pdf.

Anonymous:1909:Aq

[1015] Anonymous. Aeronautics. Scientific American, 101(22):387, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n22/pdf/scientificamerican11271909-387b.pdf.

Anonymous:1909:SUS

[1016] Anonymous. Salvage of the United States cruiser “Yankee” by compressed air. Scientific American, 101(22):388, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n22/pdf/scientificamerican11271909-388.pdf.

Perkins:1909:GEC

[1017] Frank C. Perkins. A gasoline engine for corn husking. Scientific American, 101(22):388–389, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n22/
pdf/scientificamerican11271909-388a.pdf.

Anonymous:1909:GMR

[1018] Anonymous. The gyroscopic mono-railroad. Scientific American, 101 (22):389–390, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
389.pdf.

Anonymous:1909:CSE

[1019] Anonymous. Controlling steamship engines from the bridge. Scientific American, 101(22):391, November 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n22/
pdf/scientificamerican11271909-391a.pdf.

Anonymous:1909:SOA

[1020] Anonymous. Sounding the ocean of air above us. Scientific American, 101(22):392–393, November 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n22/
pdf/scientificamerican11271909-392.pdf.

Anonymous:1909:SAM

[1021] Anonymous. A self-acting mail elevator. Scientific American, 101 (22):394, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
394.pdf.

VanDerVeer:1909:HMW

[1022] A. R. Van Der Veer. Handy man’s workshop. Scientific American, 101(22):395, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
395.pdf.

Jarman:1909:PSF

[1023] A. J. Jarman. Plating spoons and forks at home. Scientific American, 101(22):395, November 27, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n22/
pdf/scientificamerican11271909-395a.pdf.

C:1909:SST

[1024] W. J. C. Scissors with a saw-tooth. Scientific American, 101 (22):395, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
395b.pdf.

Kerner:1909:HWO

[1025] Otto Kerner. How to wind an open spring. Scientific American, 101(22):395, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
395c.pdf.

Anonymous:1909:SSPb

[1026] Anonymous. The self-smoking pipe. Scientific American, 101 (22):396, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
396.pdf.

Brockmann:1909:HMK

[1027] R. H. Brockmann. Home-made knob and tube boring machine. Scientific American, 101(22):396, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n22/
pdf/scientificamerican11271909-396a.pdf.

Gorton:1909:IAI

[1028] F. R. Gorton. Inexpensive apparatus for illustrating the “hydrostatic paradox”. Scientific American, 101(22):396, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n22/pdf/scientificamerican11271909-396b.pdf.

Brouillet:1909:DDO

[1029] J. O. Brouillet. Device for drawing ovoids. Scientific American, 101(22):396, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
396c.pdf.

W:1909:DWP

[1030] W. and K. Parkhurst. Device for washing prints. Scientific American, 101(22):396, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n22/pdf/scientificamerican11271909-
396d.pdf.

Michaud:1909:BR

[1031] Gustave Michaud. Blue roses. Scientific American, 101(22):396, November 27, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n22/pdf/scientificamerican11271909-396e.pdf.

Anonymous:1909:Cbm

[1032] Anonymous. Correspondence. Scientific American, 101(23):410, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n23/pdf/scientificamerican12041909-410.pdf.

Russell:1909:HD

[1033] Henry Norris Russell. The heavens in December. Scientific American, 101(23):414, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
414.pdf.

Anonymous:1909:RPIbo

[1034] Anonymous. Recently patented inventions. Scientific American, 101(23):417, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
417.pdf.

Anonymous:1909:NQy

[1035] Anonymous. Notes and queries. Scientific American, 101(23):417, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n23/pdf/scientificamerican12041909-417a.pdf.

Anonymous:1909:BNBba

[1036] Anonymous. Books: New books, etc. Scientific American, 101 (23):418, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
418.pdf.

Anonymous:1909:PNPbo 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
418a.pdf.

Anonymous:1909:IIbo

[1038] Anonymous. Index of inventions. Scientific American, 101(23): 418–423, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
418b.pdf.

Anonymous:1909:DUG

[1039] Anonymous. A decision under the German–American Patent Treaty. Scientific American, 101(23):406, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n23/
pdf/scientificamerican12041909-406.pdf.

Anonymous:1909:CMD

[1040] Anonymous. The cherry mine disaster. Scientific American, 101 (23):406, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
406a.pdf.

Anonymous:1909:HCC

[1041] Anonymous. High and cross-country flying abroad. Scientific American, 101(23):406, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
406b.pdf.

Anonymous:1909:Ecs

[1042] Anonymous. Engineering. Scientific American, 101(23):407, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n23/pdf/scientificamerican12041909-407.pdf.

Anonymous:1909:Ect

[1043] Anonymous. Electricity. Scientific American, 101(23):407, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n23/pdf/scientificamerican12041909-407a.pdf.

Anonymous:1909:Sbt

[1044] Anonymous. Science. Scientific American, 101(23):407, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n23/pdf/scientificamerican12041909-407b.pdf.

Michaud:1909:SPD

[1045] Gustave Michaud.  Some photographic diversions.  Scientific

American, 101(23):409–410, December 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n23/
pdf/scientificamerican12041909-409.pdf.

Anonymous:1909:HDS

[1046] Anonymous. How does sap run in pine trees? Scientific American, 101(23):410, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
410a.pdf.

Anonymous:1909:SMf

[1047] Anonymous. Safeguarding mines. Scientific American, 101(23):410, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n23/pdf/scientificamerican12041909-410b.pdf.

Anonymous:1909:EME

[1048] Anonymous. An early marine engine. Scientific American, 101 (23):410, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
410c.pdf.

Anonymous:1909:NAR

[1049] Anonymous. A novel aerial railway. Scientific American, 101 (23):411, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
411.pdf.

Blake:1909:HOR

[1050] Thaleon Blake. How to observe and record the weather. Scientific American, 101(23):412, December 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n23/
pdf/scientificamerican12041909-412.pdf.

Watkins:1909:TME

[1051] John Elfreth Watkins. Testing the man-engine. Scientific American, 101(23):413, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
413.pdf.

Anonymous:1909:CSI

[1052] Anonymous. Curiosities in science and invention. Scientific American, 101(23):415, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
415.pdf.

Anonymous:1909:IDW

[1053] Anonymous. The ice drag of the Wellman Expedition. Scientific American, 101(23):415, December 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n23/
pdf/scientificamerican12041909-415a.pdf.

Anonymous:1909:ETS

[1054] Anonymous. An early type of submarine boat. Scientific American, 101(23):415, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
415b.pdf.

Anonymous:1909:COD

[1055] Anonymous. Coin-operated device for registering letters. Scientific American, 101(23):415, December 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n23/
pdf/scientificamerican12041909-415c.pdf.

Anonymous:1909:GSG

[1056] Anonymous. The Great Stone Globe at Swanage. Scientific American, 101(23):415, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
415d.pdf.

Anonymous:1909:SHF

[1057] Anonymous. Spray helmet for firemen. Scientific American, 101 (23):415, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
415e.pdf.

Smith:1909:IBT

[1058] Katherine Louise Smith. Indoor bed tents. Scientific American, 101(23):416, December 4, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n23/pdf/scientificamerican12041909-
416.pdf.

Hyde:1909:EOC

[1059] Frederic B. Hyde. An emergency oxygen cup for miners. Scientific American, 101(23):416, December 4, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n23/
pdf/scientificamerican12041909-416a.pdf.

Anonymous:1909:Cbn

[1060] Anonymous. Correspondence. Scientific American, 101(24):436, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n24/pdf/scientificamerican12111909-436.pdf.

Anonymous:1909:BNBbb

[1061] Anonymous. Books: New books, etc. Scientific American, 101 (24):451, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
451.pdf.

Anonymous:1909:RPIbp

[1062] Anonymous. Recently patented inventions. Scientific American, 101(24):452, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
452.pdf.

Anonymous:1909:IIbp

[1063] Anonymous. Index of inventions. Scientific American, 101(24): 452, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
452a.pdf.

Anonymous:1909:PNPbp

[1064] Anonymous. Patent notes: Patents. Scientific American, 101 (24):453, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
453.pdf.

Anonymous:1909:ABP

[1065] Anonymous. Agriculture the basis of prosperity. Scientific American, 101(24):426, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
426.pdf.

Anonymous:1909:SSPc

[1066] Anonymous. Sea strength of Principal Naval Powers. Scientific American, 101(24):426, December 11, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n24/
pdf/scientificamerican12111909-426a.pdf.

Anonymous:1909:SRC

[1067] Anonymous. The Second Refrigeration Congress. Scientific American, 101(24):426, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
426b.pdf.

Anonymous:1909:Ecu

[1068] Anonymous. Engineering. Scientific American, 101(24):427, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n24/pdf/scientificamerican12111909-427.pdf.

Anonymous:1909:Ecv

[1069] Anonymous. Electrical. Scientific American, 101(24):427, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n24/pdf/scientificamerican12111909-427a.pdf.

Anonymous:1909:Sbu

[1070] Anonymous. Science. Scientific American, 101(24):427, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n24/pdf/scientificamerican12111909-427b.pdf.

Anonymous:1909:RAL

[1071] Anonymous. The reclamation of the arid lands of the West. Scientific American, 101(24):428–429, December 11, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n24/pdf/scientificamerican12111909-428.pdf.

Anonymous:1909:WIM

[1072] Anonymous. The wonderful iron mines of Lake Superior — how the ore is mined and carried in bulk. Scientific American, 101 (24):430–431, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
430.pdf.

Anonymous:1909:PLG

[1073] Anonymous. The proposed Lakes-to-the-Gulf Deep Waterway. Scientific American, 101(24):433–436, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n24/
pdf/scientificamerican12111909-433.pdf.

Anonymous:1909:AMS

[1074] Anonymous. Another magic square. Scientific American, 101 (24):436, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
436a.pdf.

Anonymous:1909:AHM

[1075] Anonymous. American harvesting machinery — the birth and growth of a vast industry. Scientific American, 101(24):437–440, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n24/pdf/scientificamerican12111909-437.pdf.

Anonymous:1909:GLM

[1076] Anonymous. Gary: The largest and most modern steel works in existence. Scientific American, 101(24):441–443, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n24/pdf/scientificamerican12111909-441.pdf.

Anonymous:1909:HSO

[1077] Anonymous. The handling and storage of our huge grain crop. Scientific American, 101(24):444–445, December 11, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n24/pdf/scientificamerican12111909-444.pdf.

Anonymous:1909:CRS

[1078] Anonymous. Chicago and the railroad system of the Middle West. Scientific American, 101(24):446–447, December 11, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n24/pdf/scientificamerican12111909-446.pdf.

Anonymous:1909:IWR

[1079] Anonymous. Iron that will not rust. Scientific American, 101 (24):448, December 11, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n24/pdf/scientificamerican12111909-
448.pdf.

Anonymous:1909:CSMb

[1080] Anonymous. Chicago’s sixty miles of freight subway. Scientific American, 101(24):448, December 11, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n24/
pdf/scientificamerican12111909-448a.pdf.

Anonymous:1909:Cbo

[1081] Anonymous. Correspondence. Scientific American, 101(25):467, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n25/pdf/scientificamerican12181909-467.pdf.

Anonymous:1909:RPIbq

[1082] Anonymous. Recently patented inventions. Scientific American, 101(25):473, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
473.pdf.

Anonymous:1909:PNPbq

[1083] Anonymous. Patent notes: Patents. Scientific American, 101 (25):473, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
473a.pdf.

Anonymous:1909:IIbq

[1084] Anonymous. Index of inventions. Scientific American, 101(25): 473, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
473b.pdf.

Anonymous:1909:RTb

[1085] Anonymous. A reversion to type. Scientific American, 101(25): 462, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
462.pdf.

Anonymous:1909:NDRb

[1086] Anonymous. Navy Department reorganization. Scientific American, 101(25):462, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
462a.pdf.

Anonymous:1909:VES

[1087] Anonymous. The value of exhaust steam. Scientific American, 101(25):462, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
462b.pdf.

Anonymous:1909:STE

[1088] Anonymous. A steam-turbine electric locomotive. Scientific American, 101(25):462, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
462c.pdf.

Anonymous:1909:Ecw

[1089] Anonymous. Engineering. Scientific American, 101(25):463, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n25/pdf/scientificamerican12181909-463.pdf.

Anonymous:1909:Sbv

[1090] Anonymous. Science. Scientific American, 101(25):463, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n25/pdf/scientificamerican12181909-463a.pdf.

Anonymous:1909:Ar

[1091] Anonymous. Aeronautics. Scientific American, 101(25):463, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n25/pdf/scientificamerican12181909-463b.pdf.

Anonymous:1909:NPR

[1092] Anonymous. The new Pennsylvania Railroad electric locomotives. Scientific American, 101(25):464, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n25/
pdf/scientificamerican12181909-464.pdf.

Collins:1909:LS

[1093] Percy Collins. The life of a splash. Scientific American, 101(25): 465–466, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
465.pdf.

Anonymous:1909:SRM

[1094] Anonymous. A simple and rapid method of measuring the height of an aeroplane above the ground. Scientific American, 101(25):466, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n25/pdf/scientificamerican12181909-466.pdf.

Anonymous:1909:ECx

[1095] Anonymous. The enormous crops of 1909. Scientific American, 101(25):466, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
466a.pdf.

Boyer:1909:NAL

[1096] Jacques Boyer. The Ninth Annual Lepine Exhibition of Toys in Paris. Scientific American, 101(25):468, December 18, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n25/pdf/scientificamerican12181909-468.pdf.

Bastin:1909:CPF

[1097] S. Leonard Bastin. Carnivorous plants of the future. Scientific American, 101(25):469, December 18, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n25/
pdf/scientificamerican12181909-469.pdf.

Anonymous:1909:MAH

[1098] Anonymous. The most ancient of human skeletons. Scientific American, 101(25):470, December 18, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n25/
pdf/scientificamerican12181909-470.pdf.

Anonymous:1909:HMWo

[1099] Anonymous. Handy man’s workshop. Scientific American, 101 (25):471, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
471.pdf.

Anonymous:1909:HWP

[1100] Anonymous. How to water-proof cloth. Scientific American, 101 (25):471, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
471a.pdf.

Stone:1909:SRA

[1101] Milton Stone. Sled runner attachment for go-carts. Scientific American, 101(25):471, December 18, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n25/
pdf/scientificamerican12181909-471b.pdf.

Bergstrom:1909:SBS

[1102] John A. Bergstrom. A snow-bound Santa Claus. Scientific American, 101(25):471, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
471c.pdf.

Nichols:1909:HML

[1103] Howard M. Nichols. How to make lead pipe into iron pipe without a wiped joint. Scientific American, 101(25):471, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n25/pdf/scientificamerican12181909-471d.pdf.

Nieman:1909:MCB

[1104] Chauncey W. Nieman. A method for connecting batteries in combinations. Scientific American, 101(25):471, December 18, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n25/pdf/scientificamerican12181909-471e.pdf.

Lord:1909:PDS

[1105] Frederick K. Lord. Power-driven scooter. Scientific American, 101 (25):472, December 18, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n25/pdf/scientificamerican12181909-
472.pdf.

Anonymous:1909:Cbp

[1106] Anonymous. Correspondence. Scientific American, 101(26):487, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n26/pdf/scientificamerican12251909-487.pdf.

Anonymous:1909:RPIbr

[1107] Anonymous. Recently patented inventions. Scientific American, 101 (26):492–493, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
492.pdf.

Anonymous:1909:NQz

[1108] Anonymous. Notes and queries. Scientific American, 101(26): 493, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
493.pdf.

Anonymous:1909:BNBbc

[1109] Anonymous. Books: New books, etc. Scientific American, 101 (26):493–494, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
493a.pdf.

Anonymous:1909:PNPbr

[1110] Anonymous. Patent notes: Patents. Scientific American, 101 (26):494, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
494.pdf.

Anonymous:1909:IIbr

[1111] Anonymous. Index of inventions. Scientific American, 101(26): 494, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
494a.pdf.

Anonymous:1909:Ib

[1112] Anonymous. Index. Scientific American, 101(26):497–498, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n26/pdf/scientificamerican12251909-497.pdf.

Anonymous:1909:W

[1113] Anonymous. A warning. Scientific American, 101(26):482, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n26/pdf/scientificamerican12251909-482.pdf.

Anonymous:1909:ECRa

[1114] Anonymous. Electrification of a Chicago Railway Terminal. Scientific American, 101(26):482, December 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n26/
pdf/scientificamerican12251909-482a.pdf.

Anonymous:1909:RNB

[1115] Anonymous. Report of the Naval Bureau of Ordnance. Scientific American, 101(26):482, December 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n26/
pdf/scientificamerican12251909-482b.pdf.

Anonymous:1909:AOR

[1116] Anonymous. The Army Ordnance Report. Scientific American, 101(26):482, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
482c.pdf.

Anonymous:1909:OEP

[1117] Anonymous. The origin of European peoples. Scientific American, 101(26):482, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
482d.pdf.

Anonymous:1909:Ecy

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n26/pdf/scientificamerican12251909-483.pdf.

Anonymous:1909:Ecz

[1119] Anonymous. Electricity. Scientific American, 101(26):483, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n26/pdf/scientificamerican12251909-483a.pdf.

Anonymous:1909:Sbw

[1120] Anonymous. Science. Scientific American, 101(26):483, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n26/pdf/scientificamerican12251909-483b.pdf.

Anonymous:1909:NTSb

[1121] Anonymous. A new theory of the screw propeller. Scientific American, 101(26):484, December 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n26/
pdf/scientificamerican12251909-484.pdf.

Perkins:1909:MSF

[1122] Frank C. Perkins. A machine for simultaneously firing many blasts. Scientific American, 101(26):484, December 25, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n26/pdf/scientificamerican12251909-484a.pdf.

Anonymous:1909:PGP

[1123] Anonymous. A pollen-gathering patent dedicated to the public. Scientific American, 101(26):485, December 25, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n26/pdf/scientificamerican12251909-485.pdf.

Mitchell:1909:PSS

[1124] S. A. Mitchell.  Photographing a star spectrum.  Scientific

American, 101(26):485–486, December 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n26/
pdf/scientificamerican12251909-485a.pdf.

Honey:1909:PV

[1125] Frederic R. Honey. The phases of Venus. Scientific American, 101(26):486, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
486.pdf.

Phillips:1909:APR

[1126] Roland Ashford Phillips. An aerial passenger railway. Scientific American, 101(26):488, December 25, 1909. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v101/n26/
pdf/scientificamerican12251909-488.pdf.

Claudy:1909:HYE

[1127] C. H. Claudy. How your eyeglasses are made. Scientific American, 101(26):489–490, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
489.pdf.

Anonymous:1909:FFW

[1128] Anonymous. Flight with flapping wings. Scientific American, 101 (26):490, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n26/pdf/scientificamerican12251909-
490.pdf.

Beasley:1909:BMO

[1129] Walter L. Beasley. The baby musk ox and the Alaskan bear of the New York Zoological Gardens. Scientific American, 101(26):491, December 25, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n26/pdf/scientificamerican12251909-491.pdf.

Anonymous:1909:Cbq

[1130] Anonymous. Correspondence. Scientific American, 101(3):43, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n3/pdf/scientificamerican07171909-43a.pdf.

Anonymous:1909:RPIbs

[1131] Anonymous. Recently patented inventions. Scientific American, 101(3):49, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
49.pdf.

Anonymous:1909:NQba

[1132] Anonymous. Notes and queries. Scientific American, 101(3):50, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n3/pdf/scientificamerican07171909-50.pdf.

Anonymous:1909:BNBbd

[1133] Anonymous. Books: New books, etc. Scientific American, 101(3):50–51, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n3/pdf/scientificamerican07171909-50a.pdf.

Anonymous:1909:LD

[1134] Anonymous. Longer daylight. Scientific American, 101(3):38, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n3/pdf/scientificamerican07171909-38.pdf.

Anonymous:1909:STLb

[1135] Anonymous. Suburban traffic in London and New York. Scientific American, 101(3):38, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
38a.pdf.

Anonymous:1909:ITSb

[1136] Anonymous. Inconclusive tests of the Scout cruisers. Scientific American, 101(3):38, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
38b.pdf.

Anonymous:1909:Edb

[1137] Anonymous. Engineering. Scientific American, 101(3):39, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n3/pdf/scientificamerican07171909-39.pdf.

Anonymous:1909:Edc

[1138] Anonymous. Electricity. Scientific American, 101(3):39, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n3/pdf/scientificamerican07171909-39a.pdf.

Anonymous:1909:Sbx

[1139] Anonymous. Science. Scientific American, 101(3):39, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n3/pdf/scientificamerican07171909-39b.pdf.

Perkins:1909:EEO

[1140] Frank C. Perkins. An English electrically operated evaporative condenser plant. Scientific American, 101(3):40, July 17, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n3/pdf/scientificamerican07171909-40.pdf.

Bailey:1909:NET

[1141] Walter Bailey. A new electric traveling scale crane. Scientific American, 101(3):40–41, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
40a.pdf.

Shepstone:1909:BTS

[1142] Harold J. Shepstone. A billiard table for ships. Scientific American, 101(3):41, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
41.pdf.

Ferry:1909:SHM

[1143] F. C. Ferry. Second Honorable Mention Essay on the Fourth Dimension. Scientific American, 101(3):42–43, July 17, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n3/pdf/scientificamerican07171909-42.pdf.

Carlin:1909:NSE

[1144] William Carlin. A new Scotch elevating ferry. Scientific American, 101(3):43, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
43.pdf.

Anonymous:1909:PMD

[1145] Anonymous. The proposed Montreal drydock. Scientific American, 101(3):43, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
43b.pdf.

Anonymous:1909:NOAf

[1146] Anonymous. The number of our ancestors. Scientific American, 101(3):43, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
43c.pdf.

Anonymous:1909:SMg

[1147] Anonymous. Signaling to Mars. Scientific American, 101(3):43, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n3/pdf/scientificamerican07171909-43d.pdf.

Kieffer:1909:CCL

[1148] H. Prime Kieffer. Constructing the concrete locks of the Panama Canal. Scientific American, 101(3):44–46, July 17, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n3/pdf/scientificamerican07171909-44.pdf.

Anonymous:1909:FD

[1149] Anonymous. Foods and digestion. Scientific American, 101(3):46, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n3/pdf/scientificamerican07171909-46.pdf.

Anonymous:1909:HMWp

[1150] Anonymous. Handy man’s workshop. Scientific American, 101 (3):47–48, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
47.pdf.

Thorpe:1909:IR

[1151] Edward Thorpe. Iceless refrigeration. Scientific American, 101(3):48, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n3/pdf/scientificamerican07171909-48.pdf.

Bayley:1909:WG

[1152] I. G. Bayley. A weatherboard gage. Scientific American, 101(3):48, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n3/pdf/scientificamerican07171909-48a.pdf.

Prather:1909:QMF

[1153] Albert Prather. Quick means for finding the center of shafts. Scientific American, 101(3):48, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
48b.pdf.

Newland:1909:HMS

[1154] James G. Newland. A home-made saw vise. Scientific American, 101(3):48, July 17, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n3/pdf/scientificamerican07171909-
48c.pdf.

Anonymous:1909:RPIbt

[1155] Anonymous. Recently patented inventions. Scientific American, 101(4):65, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
65.pdf.

Anonymous:1909:BNBbe

[1156] Anonymous. Books: New books, etc. Scientific American, 101(4):65, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n4/pdf/scientificamerican07241909-65a.pdf.

Anonymous:1909:IIbs

[1157] Anonymous. Index of inventions. Scientific American, 101(4):65–67, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n4/pdf/scientificamerican07241909-65b.pdf.

Anonymous:1909:CTT

[1158] Anonymous. City terminals without traffic connection. Scientific American, 101(4):54, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
54.pdf.

Anonymous:1909:MWN

[1159] Anonymous. Magnitude of work on new State Barge Canal. Scientific American, 101(4):54, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
54a.pdf.

Anonymous:1909:WWV

[1160] Anonymous. Wire-wound versus steel guns. Scientific American, 101(4):54, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
54b.pdf.

Anonymous:1909:BPG

[1161] Anonymous. Burglar-proof glass. Scientific American, 101(4):54, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n4/pdf/scientificamerican07241909-54c.pdf.

Anonymous:1909:As

[1162] Anonymous. Aeronautics. Scientific American, 101(4):55, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n4/pdf/scientificamerican07241909-55.pdf.

Anonymous:1909:Edd

[1163] Anonymous. Electricity. Scientific American, 101(4):55, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n4/pdf/scientificamerican07241909-55a.pdf.

Anonymous:1909:Sby

[1164] Anonymous. Science. Scientific American, 101(4):55, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n4/pdf/scientificamerican07241909-55b.pdf.

Michaud:1909:SEO

[1165] Gustave Michaud. Some experiments with an orange. Scientific American, 101(4):56, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
56.pdf.

Hislam:1909:SNW

[1166] Percival A. Hislam. Some new warships and their equipment. Scientific American, 101(4):56, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
56a.pdf.

Anonymous:1909:OTH

[1167] Anonymous. Opening of the down-town Hudson tunnels. Scientific American, 101(4):57, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
57.pdf.

Smith:1909:THM

[1168] Charles Henry Smith. The Third Honorable Mention Essay in the Fourth Dimension Competition. Scientific American, 101(4):58–59, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n4/pdf/scientificamerican07241909-58.pdf.

Anonymous:1909:SWL

[1169] Anonymous. The scientific work of the Late Prof. Simon Newcomb. Scientific American, 101(4):59, July 24, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n4/pdf/scientificamerican07241909-59.pdf.

Boyer:1909:DSB

[1170] Jacques Boyer. The disinfection of school books. Scientific American, 101(4):60–61, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
60.pdf.

Perkins:1909:AF

[1171] Frank C. Perkins. An American farmobile. Scientific American, 101(4):61, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
61.pdf.

Anonymous:1909:FCA

[1172] Anonymous. The flight of the Curtiss aeroplane for the Scientific American Trophy. Scientific American, 101(4):62, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n4/pdf/scientificamerican07241909-62.pdf.

Anonymous:1909:PDj

[1173] Anonymous. Patent department. Scientific American, 101(4):63, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n4/pdf/scientificamerican07241909-63.pdf.

Anonymous:1909:PW

[1174] Anonymous. Portable winch. Scientific American, 101(4):63, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n4/pdf/scientificamerican07241909-63a.pdf.

Anonymous:1909:ADC

[1175] Anonymous. Automatic drain for compressors. Scientific American, 101(4):63, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
63b.pdf.

Phillips:1909:PW

[1176] Roland Ashford Phillips. Panning without water. Scientific American, 101(4):63–64, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
63c.pdf.

Anonymous:1909:LEM

[1177] Anonymous. Lock for electric meters. Scientific American, 101(4):64, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n4/pdf/scientificamerican07241909-64.pdf.

Anonymous:1909:EBCa

[1178] Anonymous. Expanding brush for cleansing bottles. Scientific American, 101(4):64, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n4/pdf/scientificamerican07241909-
64a.pdf.

Anonymous:1909:APW

[1179] Anonymous. Apparatus for purifying water by the injection of ozone. Scientific American, 101(4):64, July 24, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n4/pdf/scientificamerican07241909-64b.pdf.

Anonymous:1909:PHT

[1180] Anonymous. Paper hanger’s trimmer. Scientific American, 101(4):64, July 24, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n4/pdf/scientificamerican07241909-64c.pdf.

Anonymous:1909:Cbr

[1181] Anonymous. Correspondence. Scientific American, 101(5):75, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-75.pdf.

Russell:1909:HAb

[1182] Henry Norris Russell. The heavens in August. Scientific American, 101(5):78, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n5/pdf/scientificamerican07311909-
78.pdf.

Anonymous:1909:RPIbu

[1183] Anonymous. Recently patented inventions. Scientific American, 101(5):81, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n5/pdf/scientificamerican07311909-
81.pdf.

Anonymous:1909:PNPbs

[1184] Anonymous. Patent notes: Patents. Scientific American, 101(5):81, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-81a.pdf.

Anonymous:1909:IIbt

[1185] Anonymous. Index of inventions. Scientific American, 101(5):81–83, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-81b.pdf.

Anonymous:1909:GNP

[1186] Anonymous. Good news from Panama. Scientific American, 101(5):70, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-70.pdf.

Anonymous:1909:NA

[1187] Anonymous. Naval annuals of 1909. Scientific American, 101(5):70, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-70a.pdf.

Anonymous:1909:SUa

[1188] Anonymous. As others see us. Scientific American, 101(5):70, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-70b.pdf.

Anonymous:1909:Ede

[1189] Anonymous. Engineering. Scientific American, 101(5):71, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n5/pdf/scientificamerican07311909-71.pdf.

Anonymous:1909:Edf

[1190] Anonymous. Electricity. Scientific American, 101(5):71, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n5/pdf/scientificamerican07311909-71a.pdf.

Anonymous:1909:Sbz

[1191] Anonymous. Science. Scientific American, 101(5):71, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n5/pdf/scientificamerican07311909-71b.pdf.

Anonymous:1909:BN

[1192] Anonymous. Baseball at night. Scientific American, 101(5):72, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-72.pdf.

Anonymous:1909:FACa

[1193] Anonymous. The first attempt at a cross-channel flight and new French cross-country records. Scientific American, 101(5):73–74, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n5/pdf/scientificamerican07311909-73.pdf.

Anonymous:1909:HPC

[1194] Anonymous. How the peace could be kept internationally. Scientific American, 101(5):75, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n5/pdf/scientificamerican07311909-
75a.pdf.

Anonymous:1909:SMh

[1195] Anonymous. Speed of the “Mauretania”. Scientific American, 101(5):75, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n5/pdf/scientificamerican07311909-
75b.pdf.

Anonymous:1909:CEa

[1196] Anonymous. Construction of an ellipse. Scientific American, 101(5):75, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-75c.pdf.

Anonymous:1909:NPD

[1197] Anonymous. A new process for damascening, inlaying, and blending metals. Scientific American, 101(5):76–77, July 31, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n5/pdf/scientificamerican07311909-76.pdf.

Anonymous:1909:FFS

[1198] Anonymous. The first of the 1909 flights for the Scientific American Trophy. Scientific American, 101(5):77, July 31, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n5/pdf/scientificamerican07311909-77.pdf.

Anonymous:1909:HMWq

[1199] Anonymous. Handy man’s workshop. Scientific American, 101 (5):79–80, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n5/pdf/scientificamerican07311909-
79.pdf.

Bayley:1909:LEL

[1200] I. G. Bayley. A ladder extension leg. Scientific American, 101(5):80, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-80.pdf.

C:1909:ALC

[1201] W. J. C. Automatic lubricating cup. Scientific American, 101(5):80, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-80a.pdf.

Mdermott:1909:EAC

[1202] F. P. M’dermott, Jr. Experiments with alternating current, using a small direct-current motor. Scientific American, 101(5):80, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n5/pdf/scientificamerican07311909-80b.pdf.

Sinclair:1909:DRB

[1203] G. H. Sinclair. Device for removing broken wood screws. Scientific American, 101(5):80, July 31, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n5/pdf/scientificamerican07311909-
80c.pdf.

Anonymous:1909:Cbs

[1204] Anonymous. Correspondence. Scientific American, 101(6):91, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-91.pdf.

Anonymous:1909:RPIbv

[1205] Anonymous. Recently patented inventions. Scientific American, 101(6):97, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
97.pdf.

Anonymous:1909:NQbb

[1206] Anonymous. Notes and queries. Scientific American, 101(6):97–98, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-97a.pdf.

Anonymous:1909:BNBbf

[1207] Anonymous. Books: New books, etc. Scientific American, 101(6):98, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-98.pdf.

Anonymous:1909:PNPbt

[1208] Anonymous. Patent notes: Patents. Scientific American, 101(6):98, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-98a.pdf.

Anonymous:1909:IIbu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-98b.pdf.

Anonymous:1909:BCFa

[1210] Anonymous. Blériot’s channel flight. Scientific American, 101(6):86, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-86.pdf.

Anonymous:1909:GHF

[1211] Anonymous. Grandeur of the Hudson–Fulton Naval Parade. Scientific American, 101(6):86, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
86a.pdf.

Anonymous:1909:LLB

[1212] Anonymous. Leffert Lefferts Buck. Scientific American, 101(6):86, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-86b.pdf.

Anonymous:1909:AEa

[1213] Anonymous. An antique encyclopedia. Scientific American, 101(6):86, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-86c.pdf.

Anonymous:1909:Edg

[1214] Anonymous. Engineering. Scientific American, 101(6):87, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-87.pdf.

Anonymous:1909:Edh

[1215] Anonymous. Electricity. Scientific American, 101(6):87, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n6/pdf/scientificamerican08071909-87a.pdf.

Anonymous:1909:Scc

[1216] Anonymous. Science. Scientific American, 101(6):87, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n6/pdf/scientificamerican08071909-87b.pdf.

Anonymous:1909:OWR

[1217] Anonymous. Orville Wright’s record flights at Fort Myer. Scientific American, 101(6):88, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
88.pdf.

Anonymous:1909:FSC

[1218] Anonymous. The first successful cross-channel flight. Scientific American, 101(6):88, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
88a.pdf.

Warren:1909:ELM

[1219] Frederic Blount Warren. Electric lamps in the making. Scientific American, 101(6):89, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
89.pdf.

Anonymous:1909:NTR

[1220] Anonymous. A novel type of rotary conveyer. Scientific American, 101(6):90, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
90.pdf.

Anonymous:1909:HDB

[1221] Anonymous. How does a bird soar? Scientific American, 101(6):91, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-91a.pdf.

Anonymous:1909:EAb

[1222] Anonymous. Electrochemical amalgamation. Scientific American, 101(6):91, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
91b.pdf.

Anonymous:1909:TOM

[1223] Anonymous. The terrestrial origin of the Moon — a protest. Scientific American, 101(6):91, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
91c.pdf.

Anonymous:1909:CEC

[1224] Anonymous. The chronosphere — an Empire clock. Scientific American, 101(6):92, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
92.pdf.

Anonymous:1909:FS

[1225] Anonymous. The figurehead and its story. Scientific American, 101(6):92–94, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
92a.pdf.

Aaron:1909:CWM

[1226] S. F. Aaron. The cement work of the mud wasps. Scientific American, 101(6):94, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
94.pdf.

Anonymous:1909:HMWr

[1227] Anonymous. Handy man’s workshop. Scientific American, 101(6):95, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-95.pdf.

VanDerVeer:1909:FBH

[1228] A. R. Van Der Veer. Folding bracket for the hand glass. Scientific American, 101(6):95, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
95a.pdf.

Nichols:1909:ECP

[1229] Howard M. Nichols. Electric coffee pot. Scientific American, 101(6):95, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-95b.pdf.

Brachvogel:1909:ISS

[1230] John K. Brachvogel. An improvised shoe stretcher. Scientific American, 101(6):95, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
95c.pdf.

Goebel:1909:TMO

[1231] William P. Goebel. Tabouret made from an onion crate. Scientific American, 101(6):95–96, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
95d.pdf.

Mkenzie:1909:TCL

[1232] William C. M’kenzie. A tin can leader. Scientific American, 101(6):96, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-96.pdf.

Hoster:1909:CHC

[1233] T. G. Hoster. Convenient hanger for the clothes closet. Scientific American, 101(6):96, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
96a.pdf.

Anonymous:1909:SMPb

[1234] Anonymous. Simple method of pulling a cork. Scientific American, 101(6):96, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
96b.pdf.

Handy:1909:ABE

[1235] L. G. Handy. Alarm for batteries of electric gas lighters. Scientific American, 101(6):96, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
96c.pdf.

Anonymous:1909:GTS

[1236] Anonymous. Gas-tube support. Scientific American, 101(6):96, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n6/pdf/scientificamerican08071909-96d.pdf.

Anonymous:1909:SRT

[1237] Anonymous. Substitute for rod threader. Scientific American, 101(6):96, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
96e.pdf.

Johns:1909:DLM

[1238] B. A. Johns. Drop light made from an oil lamp. Scientific American, 101(6):96, August 7, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n6/pdf/scientificamerican08071909-
96f.pdf.

Anonymous:1909:RPIbw

[1239] Anonymous. Recently patented inventions. Scientific American, 101(7):117, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
117.pdf.

Anonymous:1909:PNPbu

[1240] Anonymous. Patent notes: Patents. Scientific American, 101 (7):117, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
117a.pdf.

Anonymous:1909:IIbv

[1241] Anonymous. Index of inventions. Scientific American, 101(7): 117–119, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
117b.pdf.

Anonymous:1909:WRA

[1242] Anonymous. Wanted — a reliable aeroplane motor. Scientific American, 101(7):106, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
106.pdf.

Anonymous:1909:SCCa

[1243] Anonymous. A sensible “civil calendar”. Scientific American, 101(7):106, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
106a.pdf.

Anonymous:1909:MMC

[1244] Anonymous. Medievalism in modern chemistry. Scientific American, 101(7):106, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
106b.pdf.

Anonymous:1909:KMC

[1245] Anonymous. Keeping milk by chilling it. Scientific American, 101(7):106, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
106c.pdf.

Anonymous:1909:Edi

[1246] Anonymous. Engineering. Scientific American, 101(7):107, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n7/pdf/scientificamerican08141909-107.pdf.

Anonymous:1909:Edj

[1247] Anonymous. Electricity. Scientific American, 101(7):107, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n7/pdf/scientificamerican08141909-107a.pdf.

Anonymous:1909:At

[1248] Anonymous. Aeronautics. Scientific American, 101(7):107, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n7/pdf/scientificamerican08141909-107b.pdf.

Michaud:1909:SAE

[1249] Gustave Michaud. Some acoustic experiments. Scientific American, 101(7):108, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
108.pdf.

Cooke:1909:ATS

[1250] Charles E. Cooke. Asking the tree to speak. Scientific American, 101(7):108, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
108a.pdf.

Gradenwitz:1909:MPN

[1251] Alfred Gradenwitz. Marvelous paintings of nature. Scientific American, 101(7):109, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
109.pdf.

Roberts:1909:HC

[1252] Alexander W. Roberts. Halley’s Comet. Scientific American, 101(7):110, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
110.pdf.

Anonymous:1909:WV

[1253] Anonymous. The Wiesen viaduct. Scientific American, 101(7):111, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n7/pdf/scientificamerican08141909-111.pdf.

Anonymous:1909:BCFb

[1254] Anonymous. Blériot’s Channel flight and other records in France. Scientific American, 101(7):111, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
111a.pdf.

Anonymous:1909:CGC

[1255] Anonymous. Completion of the Government contract by Orville Wright at Fort Myer. Scientific American, 101(7):111–112, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n7/pdf/scientificamerican08141909-111b.pdf.

Willey:1909:AFS

[1256] Day Allen Willey. An American forestry school. Scientific American, 101(7):113–114, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
113.pdf.

Anonymous:1909:EDk

[1257] Anonymous. An electromagnetic drill. Scientific American, 101 (7):114, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
114.pdf.

Anonymous:1909:HMWs

[1258] Anonymous. Handy man’s workshop. Scientific American, 101 (7):115, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
115.pdf.

Anonymous:1909:TD

[1259] Anonymous. Thawing dynamite. Scientific American, 101(7): 115–116, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
115a.pdf.

Ward:1909:ERC

[1260] Frederick E. Ward. An electrolytic rectifier for charging ignition batteries. Scientific American, 101(7):116, August 14, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n7/pdf/scientificamerican08141909-116.pdf.

Pitkin:1909:SFC

[1261] C. A. Pitkin. Saw-filing clamp. Scientific American, 101(7):116, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n7/pdf/scientificamerican08141909-116a.pdf.

Kerr:1909:SSW

[1262] J. Edwin Kerr. To support shaft when babbittin. Scientific American, 101(7):116, August 14, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n7/pdf/scientificamerican08141909-
116b.pdf.

Anonymous:1909:Cbt

[1263] Anonymous. Correspondence. Scientific American, 101(8):126, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n8/pdf/scientificamerican08211909-126.pdf.

Anonymous:1909:RPIbx

[1264] Anonymous. Recently patented inventions. Scientific American, 101(8):132, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
132.pdf.

Anonymous:1909:BNBbg

[1265] Anonymous. Books: New books, etc. Scientific American, 101 (8):132, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
132a.pdf.

Anonymous:1909:IIbw

[1266] Anonymous. Index of inventions. Scientific American, 101(8): 132–135, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
132b.pdf.

Anonymous:1909:PNPbv

[1267] Anonymous. Patent notes: Patents. Scientific American, 101 (8):133, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
133.pdf.

Anonymous:1909:RYG

[1268] Anonymous. Racing yachts in a gale. Scientific American, 101 (8):122, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
122.pdf.

Anonymous:1909:PNPbw

[1269] Anonymous. Peary and the North Pole. Scientific American, 101(8):122, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
122a.pdf.

Anonymous:1909:TMU

[1270] Anonymous. Twenty miles under the sea. Scientific American, 101(8):122, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
122b.pdf.

Anonymous:1909:Edl

[1271] Anonymous. Engineering. Scientific American, 101(8):123, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n8/pdf/scientificamerican08211909-123.pdf.

Anonymous:1909:Edm

[1272] Anonymous. Electricity. Scientific American, 101(8):123, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n8/pdf/scientificamerican08211909-123a.pdf.

Anonymous:1909:Scd

[1273] Anonymous. Science. Scientific American, 101(8):123, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n8/pdf/scientificamerican08211909-123b.pdf.

Eyewitness:1909:TTB

[1274] An Eyewitness. Testing a torpedo-boat defense. Scientific American, 101(8):124, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
124.pdf.

Anonymous:1909:FCC

[1275] Anonymous. The first cross-country flight of the Aeronautic Society’s biplane. Scientific American, 101(8):124, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n8/pdf/scientificamerican08211909-124a.pdf.

Anonymous:1909:RCC

[1276] Anonymous. A record cross-country motor-boat trip. Scientific American, 101(8):125–126, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
125.pdf.

Anonymous:1909:T

[1277] Anonymous. Thunderstorms. Scientific American, 101(8):126, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n8/pdf/scientificamerican08211909-126a.pdf.

Anonymous:1909:EPC

[1278] Anonymous. The effect of Polar currents on Gulf Stream plankton. Scientific American, 101(8):126, August 21, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n8/pdf/scientificamerican08211909-126b.pdf.

Anonymous:1909:EDn

[1279] Anonymous. The excavations at Delos. Scientific American, 101 (8):127, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
127.pdf.

Anonymous:1909:NAD

[1280] Anonymous. A new application of the diving bell. Scientific American, 101(8):127, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
127a.pdf.

Honey:1909:CP

[1281] Frederic R. Honey. Conjunctions of the planets. Scientific American, 101(8):130, August 21, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n8/pdf/scientificamerican08211909-
130.pdf.

Gradenwitz:1909:NSP

[1282] Alfred Gradenwitz. A novel solution of the problem of telephotography. Scientific American, 101(8):131, August 21, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n8/pdf/scientificamerican08211909-131.pdf.

Anonymous:1909:Cbu

[1283] Anonymous. Correspondence. Scientific American, 101(9):143, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n9/pdf/scientificamerican08281909-143.pdf.

Russell:1909:HS

[1284] Henry Norris Russell. The heavens in September. Scientific American, 101(9):146, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
146.pdf.

Anonymous:1909:RPIby

[1285] Anonymous. Recently patented inventions. Scientific American, 101 (9):148–149, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
148b.pdf.

Anonymous:1909:BNBbh

[1286] Anonymous. Books: New books, etc. Scientific American, 101 (9):149, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
149.pdf.

Anonymous:1909:PNPbx

[1287] Anonymous. Patent notes: Patents. Scientific American, 101 (9):149, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
149a.pdf.

Anonymous:1909:IIbx

[1288] Anonymous. Index of inventions. Scientific American, 101(9): 149–151, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
149b.pdf.

Anonymous:1909:AFG

[1289] Anonymous. Anchoring down the floor of the Gatun Locks. Scientific American, 101(9):138, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
138.pdf.

Anonymous:1909:FSP

[1290] Anonymous. Far-sighted policy of improvement. Scientific American, 101(9):138, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
138a.pdf.

Anonymous:1909:WAI

[1291] Anonymous. The Wright Aeroplane Infringement Suit. Scientific American, 101(9):138, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
138b.pdf.

Anonymous:1909:Edo

[1292] Anonymous. Engineering. Scientific American, 101(9):139, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n9/pdf/scientificamerican08281909-139.pdf.

Anonymous:1909:Edp

[1293] Anonymous. Electricity. Scientific American, 101(9):139, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n9/pdf/scientificamerican08281909-139a.pdf.

Anonymous:1909:Sce

[1294] Anonymous. Science. Scientific American, 101(9):139, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n9/pdf/scientificamerican08281909-139b.pdf.

Anonymous:1909:RCB

[1295] Anonymous. Repairing the crushed-in bow of the “Florida”. Scientific American, 101(9):140, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
140.pdf.

Anonymous:1909:ECHb

[1296] Anonymous. The Edison concrete house. Scientific American, 101 (9):141–142, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
141.pdf.

Anonymous:1909:L

[1297] Anonymous. Lightning. Scientific American, 101(9):143, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n9/pdf/scientificamerican08281909-143a.pdf.

Anonymous:1909:FACb

[1298] Anonymous. First American to cross the channel through the air. Scientific American, 101(9):143, August 28, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n9/pdf/scientificamerican08281909-143b.pdf.

Anonymous:1909:NOAg

[1299] Anonymous. The number of our ancestors. Scientific American, 101(9):143, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
143c.pdf.

Anonymous:1909:LIP

[1300] Anonymous. The Long Island Portal of the Pennsylvania sub-river tunnels. Scientific American, 101(9):143–145, August 28, 1909. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v101/
n9/pdf/scientificamerican08281909-143d.pdf.

Anonymous:1909:AWT

[1301] Anonymous. An automobile wireless telegraph station. Scientific American, 101(9):145, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
145.pdf.

Anonymous:1909:PDk

[1302] Anonymous. Patent department. Scientific American, 101(9):147, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n9/pdf/scientificamerican08281909-147.pdf.

Anonymous:1909:SIH

[1303] Anonymous. Sad-iron heater and receptacle. Scientific American, 101(9):147, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
147a.pdf.

Anonymous:1909:SAW

[1304] Anonymous. Slat adjuster for window blinds. Scientific American, 101(9):147, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
147b.pdf.

Anonymous:1909:TSG

[1305] Anonymous. Telescopic steering gear. Scientific American, 101 (9):147, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
147c.pdf.

Anonymous:1909:WGM

[1306] Anonymous. Wave and gravity motor. Scientific American, 101 (9):147, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
147d.pdf.

Anonymous:1909:PL

[1307] Anonymous. A portable lamp. Scientific American, 101(9):148, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v101/n9/pdf/scientificamerican08281909-148.pdf.

Anonymous:1909:RTc

[1308] Anonymous. Resonator for telephones. Scientific American, 101 (9):148, August 28, 1909. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v101/n9/pdf/scientificamerican08281909-
148a.pdf.

Anonymous:1900:Ca

[1309] Anonymous. Correspondence. Scientific American, 82(1):8, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-8.pdf.

Anonymous:1900:RPIa

[1310] Anonymous. Recently patented inventions. Scientific American, 82(1):12, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n1/pdf/scientificamerican01061900-
12.pdf.

Anonymous:1900:BNBa

[1311] Anonymous. Books: New books, etc. Scientific American, 82(1):12, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-12a.pdf.

Anonymous:1900:BPa

[1312] Anonymous. Business and personal. Scientific American, 82(1):12, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-12b.pdf.

Anonymous:1900:NQa

[1313] Anonymous. Notes and queries. Scientific American, 82(1):12, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-12c.pdf.

Anonymous:1900:IIa

[1314] Anonymous. Index of inventions. Scientific American, 82(1):12, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-12e.pdf.

Anonymous:1900:PNPa

[1315] Anonymous. Patent notes: Patents. Scientific American, 82(1):15, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-15.pdf.

Anonymous:1900:RY

[1316] Anonymous. Retrospect of the year 1899. Scientific American, 82(1):4–5, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n1/pdf/scientificamerican01061900-
4.pdf.

Anonymous:1900:LA

[1317] Anonymous. Landslide at Amalfi. Scientific American, 82(1):6, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-6.pdf.

Anonymous:1900:CMR

[1318] Anonymous. The Chester metallic railway tie. Scientific American, 82(1):6, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n1/pdf/scientificamerican01061900-
6a.pdf.

Fawcett:1900:MOU

[1319] Waldon Fawcett. Mechanical ore unloader. Scientific American, 82(1):7, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-7.pdf.

Mgahie:1900:ESA

[1320] Frederick H. M’gahie. The explosive side of acetylene. Scientific American, 82(1):8–9, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n1/pdf/scientificamerican01061900-
8a.pdf.

Anonymous:1900:NYZ

[1321] Anonymous. New York Zoological Park. Scientific American, 82 (1):10–11, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n1/pdf/scientificamerican01061900-
10.pdf.

Anonymous:1900:MWV

[1322] Anonymous. Motor-wheel for vehicles. Scientific American, 82(1):11, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-11.pdf.

Anonymous:1900:Ia

[1323] Anonymous. To inventors. Scientific American, 82(1):12, January 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n1/pdf/scientificamerican01061900-12d.pdf.

Anonymous:1900:Cb

[1324] Anonymous. Correspondence. Scientific American, 82(10):154, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-154.pdf.

Anonymous:1900:RPIb

[1325] Anonymous. Recently patented inventions. Scientific American, 82(10):156, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
156.pdf.

Anonymous:1900:BNBb

[1326] Anonymous. Books: New books, etc. Scientific American, 82(10):156, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-156a.pdf.

Anonymous:1900:BPb

[1327] Anonymous. Business and personal. Scientific American, 82(10):157, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-157.pdf.

Anonymous:1900:NQb

[1328] Anonymous. Notes and queries. Scientific American, 82(10):157, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-157a.pdf.

Anonymous:1900:IIb

[1329] Anonymous. Index of inventions. Scientific American, 82(10):157, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-157b.pdf.

Anonymous:1900:PNPb

[1330] Anonymous. Patent notes: Patents. Scientific American, 82(10):159, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-159.pdf.

Anonymous:1900:CR

[1331] Anonymous. Consular reform. Scientific American, 82(10):146, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-146.pdf.

Anonymous:1900:IIK

[1332] Anonymous. The Ives’ improved Kromskop. Scientific American, 82(10):146, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
146a.pdf.

Anonymous:1900:HMP

[1333] Anonymous. Hand and machine production compared. Scientific American, 82(10):146, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
146b.pdf.

Anonymous:1900:PIE

[1334] Anonymous. Protection of inventions exhibited at Paris. Scientific American, 82(10):146, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
146c.pdf.

Anonymous:1900:ARC

[1335] Anonymous. Annual Report of the Commissioner of Patents. Scientific American, 82(10):146–147, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
146d.pdf.

Anonymous:1900:CBTa

[1336] Anonymous. California big trees threatened with destruction. Scientific American, 82(10):147, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
147.pdf.

Anonymous:1900:DAO

[1337] Anonymous. Dissociation of air at ordinary pressure. Scientific American, 82(10):147, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
147a.pdf.

Anonymous:1900:HA

[1338] Anonymous. Horseflesh in America. Scientific American, 82(10):147, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-147b.pdf.

Anonymous:1900:BSI

[1339] Anonymous. The beet sugar industry. Scientific American, 82(10):147, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-147c.pdf.

Anonymous:1900:FAP

[1340] Anonymous. The first automobile patrol wagon. Scientific American, 82(10):148, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
148.pdf.

Anonymous:1900:EAA

[1341] Anonymous. An electric automobile ambulance. Scientific American, 82(10):148, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
148a.pdf.

Anonymous:1900:CHS

[1342] Anonymous. A combined heating stove and furnace. Scientific American, 82(10):149, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
149.pdf.

Anonymous:1900:APCa

[1343] Anonymous. An anchor to prevent the creeping of rails. Scientific American, 82(10):149, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
149a.pdf.

Guenaire:1900:AIB

[1344] J. Guenaire. The Alexandre III Bridge, Paris. Scientific American, 82(10):149–150, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
149b.pdf.

Anonymous:1900:RAMa

[1345] Anonymous. Radio-active matter in magnetic field. Scientific American, 82(10):150–151, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
150.pdf.

Anonymous:1900:TPN

[1346] Anonymous. Torpedo practice at Newport, R.I. Scientific American, 82(10):152, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
152.pdf.

Anonymous:1900:OLP

[1347] Anonymous. Our latest protected cruiser, the “Albany”. Scientific American, 82(10):152–153, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
152a.pdf.

Anonymous:1900:FST

[1348] Anonymous. Fighting snow on the third-rail. Scientific American, 82(10):153, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
153.pdf.

Anonymous:1900:TBW

[1349] Anonymous. Transportation of buildings by water. Scientific American, 82(10):154, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n10/pdf/scientificamerican03101900-
154a.pdf.

Michaud:1900:CG

[1350] Gustave Michaud. A chemical growth. Scientific American, 82(10):155, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-155.pdf.

Anonymous:1900:FHT

[1351] Anonymous. A feat in half-tone work. Scientific American, 82(10):155, March 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n10/pdf/scientificamerican03101900-155a.pdf.

Anonymous:1900:Cc

[1352] Anonymous. Correspondence. Scientific American, 82(11):170, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-170a.pdf.

Anonymous:1900:RPIc

[1353] Anonymous. Recently patented inventions. Scientific American, 82(11):172, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
172.pdf.

Anonymous:1900:BNBc

[1354] Anonymous. Books: New books, etc. Scientific American, 82(11):172, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-172a.pdf.

Anonymous:1900:BPc

[1355] Anonymous. Business and personal. Scientific American, 82(11):172, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-172b.pdf.

Anonymous:1900:NQc

[1356] Anonymous. Notes and queries. Scientific American, 82(11):172, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-172c.pdf.

Anonymous:1900:IIc

[1357] Anonymous. Index of inventions. Scientific American, 82(11):173, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-173a.pdf.

Anonymous:1900:PNPc

[1358] Anonymous. Patent notes: Patents. Scientific American, 82(11):175, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-175.pdf.

Anonymous:1900:LAPa

[1359] Anonymous. Liquid air promotion again. Scientific American, 82(11):162, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
162.pdf.

Anonymous:1900:CWG

[1360] Anonymous. The construction of warships at Government Navy yards. Scientific American, 82(11):162, March 17, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n11/pdf/scientificamerican03171900-162a.pdf.

Anonymous:1900:WDC

[1361] Anonymous. The work of the divison of chemistry of the department of agriculture. Scientific American, 82(11):162–163, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-162b.pdf.

Maxim:1900:FVC

[1362] Hudson Maxim. Facts versus claims for liquid air. Scientific American, 82(11):163, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
163.pdf.

Anonymous:1900:BDC

[1363] Anonymous. A billion dollar country. Scientific American, 82(11):163, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-163a.pdf.

Anonymous:1900:APH

[1364] Anonymous. The Asiatic plague at Honolulu. Scientific American, 82(11):164, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
164.pdf.

Anonymous:1900:SSF

[1365] Anonymous. A simple shutter-fastener. Scientific American, 82(11):165, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-165.pdf.

Anonymous:1900:WCD

[1366] Anonymous. A worm and chain driving-gear for elevators. Scientific American, 82(11):165, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
165a.pdf.

Anonymous:1900:ACP

[1367] Anonymous. An amateur’s camera for photographing in natural colors. Scientific American, 82(11):165–166, March 17, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n11/pdf/scientificamerican03171900-165b.pdf.

Anonymous:1900:SEL

[1368] Anonymous. Steam electric lightship for Cape Hatteras. Scientific American, 82(11):166–167, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
166.pdf.

Anonymous:1900:LLM

[1369] Anonymous. The latest of the large Mississippi bridges. Scientific American, 82(11):168, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
168.pdf.

Anonymous:1900:MJM

[1370] Anonymous. M. J. De Morgan’s excavations in the Akropolis and Palaces of Susa. Scientific American, 82(11):169–170, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-169.pdf.

Anonymous:1900:CRC

[1371] Anonymous. The Chinese rods and cords. Scientific American, 82(11):170, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
170.pdf.

Anonymous:1900:MT

[1372] Anonymous. A milligrade thermometer. Scientific American, 82 (11):170, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n11/pdf/scientificamerican03171900-
170b.pdf.

Holder:1900:PS

[1373] Charles F. Holder. The passing show. Scientific American, 82(11):171, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-171.pdf.

Anonymous:1900:Ib

[1374] Anonymous. To inventors. Scientific American, 82(11):173, March 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n11/pdf/scientificamerican03171900-173.pdf.

Anonymous:1900:RPId

[1375] Anonymous. Recently patented inventions. Scientific American, 82(12):188, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
188.pdf.

Anonymous:1900:BPd

[1376] Anonymous. Business and personal. Scientific American, 82(12):189, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n12/pdf/scientificamerican03241900-189.pdf.

Anonymous:1900:NQd

[1377] Anonymous. Notes and queries. Scientific American, 82(12):189, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n12/pdf/scientificamerican03241900-189a.pdf.

Anonymous:1900:BNBd

[1378] Anonymous. Books: New books, etc. Scientific American, 82(12):189, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n12/pdf/scientificamerican03241900-189b.pdf.

Anonymous:1900:IId

[1379] Anonymous. Index of inventions. Scientific American, 82(12):189–191, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n12/pdf/scientificamerican03241900-189c.pdf.

Anonymous:1900:PNPd

[1380] Anonymous. Patent notes: Patents. Scientific American, 82(12):191, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n12/pdf/scientificamerican03241900-191.pdf.

Anonymous:1900:SMM

[1381] Anonymous. The supremacy of the modern magazine rifle. Scientific American, 82(12):178, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
178.pdf.

Anonymous:1900:NYWa

[1382] Anonymous. New York’s water supply — present and future capacity. Scientific American, 82(12):178, March 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n12/pdf/scientificamerican03241900-178a.pdf.

Anonymous:1900:AES

[1383] Anonymous. American experiment stations at the Paris Exposition. Scientific American, 82(12):178–179, March 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n12/pdf/scientificamerican03241900-178b.pdf.

Anonymous:1900:ABN

[1384] Anonymous. An Admiralty Board for the Navy. Scientific American, 82(12):179, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
179.pdf.

Anonymous:1900:NNP

[1385] Anonymous. The new Naval Programme. Scientific American, 82(12):179, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
179a.pdf.

Anonymous:1900:TLS

[1386] Anonymous. Toning lantern slides and bromide prints by ferrocyanide of copper. Scientific American, 82(12):179, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n12/pdf/scientificamerican03241900-179b.pdf.

Anonymous:1900:DNI

[1387] Anonymous. Death of a noted inventor. Scientific American, 82(12):179, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n12/pdf/scientificamerican03241900-179c.pdf.

Shepstone:1900:PCH

[1388] H. J. Shepstone. The Princess Christian Hospital Train for South Africa. Scientific American, 82(12):180, March 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n12/pdf/scientificamerican03241900-180.pdf.

Anonymous:1900:EAT

[1389] Anonymous. An electric alarm try-cock for steam boilers. Scientific American, 82(12):181, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
181.pdf.

Anonymous:1900:CRE

[1390] Anonymous. The creeping of rails on the Eads Bridge, St. Louis. Scientific American, 82(12):181, March 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n12/pdf/scientificamerican03241900-181a.pdf.

Anonymous:1900:NFS

[1391] Anonymous. A novel form of steam engines. Scientific American, 82(12):181–182, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
181b.pdf.

Anonymous:1900:FPA

[1392] Anonymous. Filtration plant for the Albany Water Supply. Scientific American, 82(12):182, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
182.pdf.

Anonymous:1900:NFB

[1393] Anonymous. The new French battleship “Suffren” and her predecessor. Scientific American, 82(12):184, March 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n12/pdf/scientificamerican03241900-184.pdf.

Anonymous:1900:NBA

[1394] Anonymous. The new Brazilian armorclad “Marshal Deodoro”. Scientific American, 82(12):184–185, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
184a.pdf.

Anonymous:1900:ONP

[1395] Anonymous. The old and new pumping engines of Dry Dock No. 1, Brooklyn Navy Yard. Scientific American, 82(12):185–186, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n12/pdf/scientificamerican03241900-185.pdf.

Anonymous:1900:ROS

[1396] Anonymous. Removing obstructions in San Francisco Bay. Scientific American, 82(12):186–187, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
186.pdf.

Scott:1900:NAI

[1397] A. C. Scott. A new apparatus for instantaneous photo-micrography. Scientific American, 82(12):187–188, March 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v82/n12/pdf/scientificamerican03241900-
187.pdf.

Russell:1900:HAa

[1398] Henry Norris Russell. The heavens in April. Scientific American, 82(13):195, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
195.pdf.

Anonymous:1900:Cd

[1399] Anonymous. Correspondence. Scientific American, 82(13):199, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n13/pdf/scientificamerican03311900-199.pdf.

Anonymous:1900:RPIe

[1400] Anonymous. Recently patented inventions. Scientific American, 82(13):204, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
204.pdf.

Anonymous:1900:BNBe

[1401] Anonymous. Books: New books, etc. Scientific American, 82(13):204, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n13/pdf/scientificamerican03311900-204a.pdf.

Anonymous:1900:BPe

[1402] Anonymous. Business and personal. Scientific American, 82(13):205, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n13/pdf/scientificamerican03311900-205.pdf.

Anonymous:1900:NQe

[1403] Anonymous. Notes and queries. Scientific American, 82(13):205, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n13/pdf/scientificamerican03311900-205a.pdf.

Anonymous:1900:IIe

[1404] Anonymous. Index of inventions. Scientific American, 82(13):205, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n13/pdf/scientificamerican03311900-205b.pdf.

Anonymous:1900:PNPe

[1405] Anonymous. Patent notes: Patents. Scientific American, 82(13):207, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n13/pdf/scientificamerican03311900-207.pdf.

Anonymous:1900:ICS

[1406] Anonymous. The Isthmian Canal Survey. Scientific American, 82(13):194, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
194.pdf.

Anonymous:1900:DLY

[1407] Anonymous. The designs of last year’s battleships and cruisers. Scientific American, 82(13):194, March 31, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n13/pdf/scientificamerican03311900-194a.pdf.

Anonymous:1900:BGR

[1408] Anonymous. Breaking ground for the rapid transit tunnel. Scientific American, 82(13):194, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
194b.pdf.

Anonymous:1900:GNP

[1409] Anonymous. Germany’s Naval Programme — a lesson for the United States. Scientific American, 82(13):194, March 31, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n13/pdf/scientificamerican03311900-194c.pdf.

Anonymous:1900:SMSa

[1410] Anonymous. The steering of modern, screw-propelled ships. Scientific American, 82(13):194–195, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
194d.pdf.

Anonymous:1900:TTA

[1411] Anonymous. Three types of automobiles. Scientific American, 82(13):195, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
195a.pdf.

Anonymous:1900:MASa

[1412] Anonymous. Mutual Aid Societies for French school children. Scientific American, 82(13):195, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
195b.pdf.

Anonymous:1900:CTS

[1413] Anonymous. A cheap telephone system for farmers. Scientific American, 82(13):196, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
196.pdf.

Anonymous:1900:NMC

[1414] Anonymous. A new method of cutting threads. Scientific American, 82(13):196, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
196a.pdf.

Anonymous:1900:ISCa

[1415] Anonymous. An improvement in smoke consuming furnaces. Scientific American, 82(13):196, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
196b.pdf.

Anonymous:1900:CET

[1416] Anonymous. A cheap and efficient typewriter. Scientific American, 82(13):197, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
197.pdf.

Anonymous:1900:NGL

[1417] Anonymous. The North German Lloyd liner “Kaiserin Maria Theresia”. Scientific American, 82(13):197–198, March 31, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n13/pdf/scientificamerican03311900-197a.pdf.

Anonymous:1900:RSG

[1418] Anonymous. A rein support and guard for horse harness. Scientific American, 82(13):198, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
198.pdf.

Anonymous:1900:AFL

[1419] Anonymous. American freight locomotives and the engines of the “Oceanic” — a comparison of horsepower. Scientific American, 82(13):199, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
199a.pdf.

Fawcett:1900:MDT

[1420] Waldon Fawcett. The Marine Department of the Trans-Siberian Railroad. Scientific American, 82(13):200, March 31, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n13/pdf/scientificamerican03311900-200.pdf.

Anonymous:1900:PPE

[1421] Anonymous. Prof. Pictet’s experiments with liquid air. Scientific American, 82(13):201–202, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
201.pdf.

Holder:1900:SAM

[1422] C. F. Holder. Some American mummies. Scientific American, 82 (13):203–204, March 31, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n13/pdf/scientificamerican03311900-
203.pdf.

Anonymous:1900:Ce

[1423] Anonymous. Correspondence. Scientific American, 82(14):214–215, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-214a.pdf.

Anonymous:1900:RPIf

[1424] Anonymous. Recently patented inventions. Scientific American, 82(14):220, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
220.pdf.

Anonymous:1900:BPf

[1425] Anonymous. Business and personal. Scientific American, 82(14):220, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-220a.pdf.

Anonymous:1900:NQf

[1426] Anonymous. Notes and queries. Scientific American, 82(14):220, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-220b.pdf.

Anonymous:1900:IIf

[1427] Anonymous. Index of inventions. Scientific American, 82(14):220, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-220d.pdf.

Anonymous:1900:PNPf

[1428] Anonymous. Patent notes: Patents. Scientific American, 82(14):223, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-223.pdf.

Anonymous:1900:OPEa

[1429] Anonymous. Obstacles to the proposed Erie Canal improvement. Scientific American, 82(14):210, April 7, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n14/pdf/scientificamerican04071900-210.pdf.

Anonymous:1900:AGP

[1430] Anonymous. Automobilism in the German postal service. Scientific American, 82(14):210, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
210a.pdf.

Anonymous:1900:WDE

[1431] Anonymous. The work of the Division of Entomology. Scientific American, 82(14):210, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
210b.pdf.

Anonymous:1900:CACa

[1432] Anonymous. Changes in the Assistant Commissionership. Scientific American, 82(14):210–211, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
210c.pdf.

Anonymous:1900:SCV

[1433] Anonymous. Some Chinese vegetables. Scientific American, 82(14):211, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-211.pdf.

Anonymous:1900:POE

[1434] Anonymous. The Patent Office Exhibit at the Paris Exposition. Scientific American, 82(14):211, April 7, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n14/pdf/scientificamerican04071900-211a.pdf.

Anonymous:1900:MTC

[1435] Anonymous. M. Trillot on color photography. Scientific American, 82(14):211, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
211b.pdf.

Anonymous:1900:OEI

[1436] Anonymous. Our exports of iron and steel. Scientific American, 82(14):211, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
211c.pdf.

Anonymous:1900:DPP

[1437] Anonymous. Death of Prof. Pepper. Scientific American, 82(14):211, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-211d.pdf.

Anonymous:1900:PBI

[1438] Anonymous. A perfected burner for incandescent mantles. Scientific American, 82(14):212, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
212.pdf.

Anonymous:1900:RAE

[1439] Anonymous. A rope armored engine. Scientific American, 82(14):212, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-212a.pdf.

Anonymous:1900:NRB

[1440] Anonymous. A non-refillable bottle. Scientific American, 82(14):212, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-212b.pdf.

Anonymous:1900:PSP

[1441] Anonymous. A problem in ship propulsion. Scientific American, 82(14):213, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
213.pdf.

Gratacap:1900:AMT

[1442] L. P. Gratacap. The asbestos mines at Thetford, Canada. Scientific American, 82(14):213–214, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
213a.pdf.

Anonymous:1900:EEA

[1443] Anonymous. The expiration of the Edison American three-wire patent. Scientific American, 82(14):214, April 7, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n14/pdf/scientificamerican04071900-214.pdf.

Anonymous:1900:FSE

[1444] Anonymous. Ferry steamer with an elevating deck. Scientific American, 82(14):216, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
216.pdf.

Anonymous:1900:BWT

[1445] Anonymous. The Black-Wall Tunnel, London. Scientific American, 82(14):217–218, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
217.pdf.

Anonymous:1900:CBB

[1446] Anonymous. The Cullen ball-bearing rifled gun. Scientific American, 82(14):218–219, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
218.pdf.

Anonymous:1900:FTA

[1447] Anonymous. A French trolley automobile. Scientific American, 82(14):219, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n14/pdf/scientificamerican04071900-
219.pdf.

Anonymous:1900:Ic

[1448] Anonymous. To inventors. Scientific American, 82(14):220, April 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n14/pdf/scientificamerican04071900-220c.pdf.

Anonymous:1900:RPIg

[1449] Anonymous. Recently patented inventions. Scientific American, 82(15):236, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
236.pdf.

Anonymous:1900:BPg

[1450] Anonymous. Business and personal. Scientific American, 82(15):236, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n15/pdf/scientificamerican04141900-236a.pdf.

Anonymous:1900:NQg

[1451] Anonymous. Notes and queries. Scientific American, 82(15):236, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n15/pdf/scientificamerican04141900-236b.pdf.

Anonymous:1900:IIg

[1452] Anonymous. Index of inventions. Scientific American, 82(15):236, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n15/pdf/scientificamerican04141900-236c.pdf.

Anonymous:1900:PNPg

[1453] Anonymous. Patent notes: Patents. Scientific American, 82(15):239, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n15/pdf/scientificamerican04141900-239.pdf.

Anonymous:1900:NYWb

[1454] Anonymous. New York’s water supply and the Ramapo Contract. Scientific American, 82(15):226, April 14, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n15/pdf/scientificamerican04141900-226.pdf.

Anonymous:1900:PDC

[1455] Anonymous. The proposed department of commerce and industry. Scientific American, 82(15):226, April 14, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n15/pdf/scientificamerican04141900-226a.pdf.

Anonymous:1900:DTS

[1456] Anonymous. The double-turret system on trial. Scientific American, 82(15):226–227, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
226b.pdf.

Anonymous:1900:FPS

[1457] Anonymous. French prizes for scientists and inventors. Scientific American, 82(15):227, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
227.pdf.

Anonymous:1900:ASMa

[1458] Anonymous. Applied science in modern war. Scientific American, 82(15):227, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
227a.pdf.

Anonymous:1900:YS

[1459] Anonymous. The youngest subscriber. Scientific American, 82(15):227, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n15/pdf/scientificamerican04141900-227b.pdf.

Anonymous:1900:DSG

[1460] Anonymous. Death of St. George Mivart. Scientific American, 82(15):227, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
227c.pdf.

Anonymous:1900:PENa

[1461] Anonymous. Paris Exposition notes. Scientific American, 82(15):227, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n15/pdf/scientificamerican04141900-227d.pdf.

Anonymous:1900:LAPb

[1462] Anonymous. A Lister attachment for planters. Scientific American, 82(15):228, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
228.pdf.

Anonymous:1900:AMP

[1463] Anonymous. Apparatus for making perspective drawings. Scientific American, 82(15):228, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
228a.pdf.

Anonymous:1900:SIT

[1464] Anonymous. A simple ingenious toy. Scientific American, 82(15):228, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n15/pdf/scientificamerican04141900-228b.pdf.

Anonymous:1900:CAG

[1465] Anonymous. The Colt automatic gun. Scientific American, 82(15):229, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n15/pdf/scientificamerican04141900-229.pdf.

Anonymous:1900:GBCa

[1466] Anonymous. The Gordon Bennett Cup Race. Scientific American, 82(15):229, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
229a.pdf.

Anonymous:1900:STK

[1467] Anonymous. The superposed turrets of the “Kearsarge”. Scientific American, 82(15):230–231, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
230.pdf.

Fawcett:1900:WTG

[1468] Waldon Fawcett. The water transportation of grain. Scientific American, 82(15):232–233, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
232.pdf.

Mindeleff:1900:HHD

[1469] Cosmos Mindeleff. Houses and house dedication of the Navahos. Scientific American, 82(15):233–235, April 14, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n15/pdf/scientificamerican04141900-233.pdf.

Anonymous:1900:RWD

[1470] Anonymous. Reproduction of working drawings. Scientific American, 82(15):235, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
235.pdf.

Smith:1900:TTL

[1471] Hugh M. Smith. The three toothed lamprey. Scientific American, 82(15):235, April 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n15/pdf/scientificamerican04141900-
235a.pdf.

Anonymous:1900:RPIh

[1472] Anonymous. Recently patented inventions. Scientific American, 82(16):251, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
251.pdf.

Anonymous:1900:BNBf

[1473] Anonymous. Books: New books, etc. Scientific American, 82 (16):251–252, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
251a.pdf.

Anonymous:1900:BPh

[1474] Anonymous. Business and personal. Scientific American, 82(16):252, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n16/pdf/scientificamerican04211900-252.pdf.

Anonymous:1900:NQh

[1475] Anonymous. Notes and queries. Scientific American, 82(16):252, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n16/pdf/scientificamerican04211900-252a.pdf.

Anonymous:1900:IIh

[1476] Anonymous. Index of inventions. Scientific American, 82(16):252, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n16/pdf/scientificamerican04211900-252b.pdf.

Anonymous:1900:PNPh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n16/pdf/scientificamerican04211900-255.pdf.

Anonymous:1900:USW

[1478] Anonymous. United States in the world’s trade. Scientific American, 82(16):242, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
242.pdf.

Anonymous:1900:DTR

[1479] Anonymous. The double turret rejected by the Naval Board of Construction. Scientific American, 82(16):242, April 21, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n16/pdf/scientificamerican04211900-242a.pdf.

Anonymous:1900:AVE

[1480] Anonymous. American vs. English machine tools. Scientific American, 82(16):242, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
242b.pdf.

Anonymous:1900:FBA

[1481] Anonymous. Fox breeding on the Alaskan Islands. Scientific American, 82(16):242, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
242c.pdf.

Anonymous:1900:NHC

[1482] Anonymous. A new hydrocarbon cuprene. Scientific American, 82(16):243, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
243.pdf.

Anonymous:1900:PI

[1483] Anonymous. The peanut industry. Scientific American, 82(16):243, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n16/pdf/scientificamerican04211900-243a.pdf.

Anonymous:1900:CGF

[1484] Anonymous. Cotton’s great future. Scientific American, 82(16):243, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n16/pdf/scientificamerican04211900-243b.pdf.

Anonymous:1900:EPE

[1485] Anonymous. The entrance to the Paris Exposition. Scientific American, 82(16):243, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
243c.pdf.

Anonymous:1900:DFH

[1486] Anonymous. Death of Frank H. Cushing. Scientific American, 82(16):243, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
243d.pdf.

Anonymous:1900:CAM

[1487] Anonymous. A controlling apparatus for mine elevators. Scientific American, 82(16):244, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
244.pdf.

Anonymous:1900:SFW

[1488] Anonymous. A simple form of wire-hanger. Scientific American, 82(16):244, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
244a.pdf.

Anonymous:1900:CEa

[1489] Anonymous. The coming eclipse. Scientific American, 82(16):244–245, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n16/pdf/scientificamerican04211900-244b.pdf.

Anonymous:1900:CATa

[1490] Anonymous. A combined automobile and tramway omnibus. Scientific American, 82(16):245–246, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
245.pdf.

Anonymous:1900:TMP

[1491] Anonymous. The turquoise mines of Persia. Scientific American, 82(16):246–247, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
246.pdf.

Anonymous:1900:FFD

[1492] Anonymous. 700-foot floating dock at South Brooklyn. Scientific American, 82(16):248, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
248.pdf.

Anonymous:1900:ETB

[1493] Anonymous. The engines of the torpedo boat destroyer “Viper”. Scientific American, 82(16):248–249, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
248a.pdf.

Anonymous:1900:DSV

[1494] Anonymous. The Dismal Swamp of Virginia. Scientific American, 82(16):249–250, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
249.pdf.

Anonymous:1900:NACa

[1495] Anonymous. The new Assistant Commissioner of Patents. Scientific American, 82(16):250, April 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n16/pdf/scientificamerican04211900-
250.pdf.

Anonymous:1900:Cf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-267.pdf.

Anonymous:1900:RPIi

[1497] Anonymous. Recently patented inventions. Scientific American, 82(17):267, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
267a.pdf.

Anonymous:1900:BPi

[1498] Anonymous. Business and personal. Scientific American, 82(17):268, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-268.pdf.

Anonymous:1900:NQi

[1499] Anonymous. Notes and queries. Scientific American, 82(17):268, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-268a.pdf.

Anonymous:1900:BNBg

[1500] Anonymous. Books: New books, etc. Scientific American, 82(17):268, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-268b.pdf.

Anonymous:1900:IIi

[1501] Anonymous. Index of inventions. Scientific American, 82(17):268, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-268c.pdf.

Anonymous:1900:PNPi

[1502] Anonymous. Patent notes: Patents. Scientific American, 82(17):271, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-271.pdf.

Anonymous:1900:FI

[1503] Anonymous. Finality in invention. Scientific American, 82(17):258, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-258.pdf.

Anonymous:1900:PPT

[1504] Anonymous. Progress and prospects of the Trans-Siberian Railway. Scientific American, 82(17):258, April 28, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n17/pdf/scientificamerican04281900-258a.pdf.

Anonymous:1900:OPEb

[1505] Anonymous. Opening of the Paris Exposition. Scientific American, 82(17):258–259, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
258b.pdf.

Benjamin:1900:NAS

[1506] Marcus Benjamin. National Academy of Sciences. Scientific American, 82(17):259, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
259.pdf.

Anonymous:1900:AAR

[1507] Anonymous. An American automobile race. Scientific American, 82(17):259, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
259a.pdf.

Anonymous:1900:PENb

[1508] Anonymous. Paris Exposition notes. Scientific American, 82(17):259, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-259b.pdf.

Anonymous:1900:Cg

[1509] Anonymous. Cellulith. Scientific American, 82(17):259, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-259c.pdf.

Anonymous:1900:ASPa

[1510] Anonymous. Aerial sweep for the Pan-American Exposition. Scientific American, 82(17):260, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
260.pdf.

Anonymous:1900:SCC

[1511] Anonymous. A simple compass with a caliper-scale. Scientific American, 82(17):260, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
260a.pdf.

Anonymous:1900:UAR

[1512] Anonymous. The use of acetylene in raising sunken vessels. Scientific American, 82(17):261, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
261.pdf.

Anonymous:1900:NSBa

[1513] Anonymous. A new storage battery. Scientific American, 82 (17):261–262, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
261a.pdf.

Anonymous:1900:WMa

[1514] Anonymous. A window mirror. Scientific American, 82(17):262, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-262.pdf.

Anonymous:1900:HSA

[1515] Anonymous. The hydraulic system of air compression. Scientific American, 82(17):263, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
263.pdf.

Anonymous:1900:VRW

[1516] Anonymous. The vagaries of a railroad wreck. Scientific American, 82(17):264, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
264.pdf.

Anonymous:1900:NJB

[1517] Anonymous. The new Japan battleship “Shikishima”. Scientific American, 82(17):264–265, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
264a.pdf.

Anonymous:1900:FMD

[1518] Anonymous. The failure of the masonry dam at Austin, Texas. Scientific American, 82(17):265–266, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n17/pdf/scientificamerican04281900-
265.pdf.

Holder:1900:VSB

[1519] C. F. Holder. Value of sea birds. Scientific American, 82(17):266–267, April 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n17/pdf/scientificamerican04281900-266.pdf.

Russell:1900:HM

[1520] Henry Norris Russell. The heavens in May. Scientific American, 82(18):275, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
275a.pdf.

Anonymous:1900:RPIj

[1521] Anonymous. Recently patented inventions. Scientific American, 82(18):283, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
283.pdf.

Anonymous:1900:BPj

[1522] Anonymous. Business and personal. Scientific American, 82(18):284, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n18/pdf/scientificamerican05051900-284a.pdf.

Anonymous:1900:NQj

[1523] Anonymous. Notes and queries. Scientific American, 82(18):284, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n18/pdf/scientificamerican05051900-284b.pdf.

Anonymous:1900:BNBh

[1524] Anonymous. Books: New books, etc. Scientific American, 82(18):284, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n18/pdf/scientificamerican05051900-284c.pdf.

Anonymous:1900:IIj

[1525] Anonymous. Index of inventions. Scientific American, 82(18):284, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n18/pdf/scientificamerican05051900-284d.pdf.

Anonymous:1900:PNPj

[1526] Anonymous. Patent notes: Patents. Scientific American, 82(18):287, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n18/pdf/scientificamerican05051900-287.pdf.

Anonymous:1900:RAMb

[1527] Anonymous. Revival of the American merchant marine. Scientific American, 82(18):274, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
274.pdf.

Anonymous:1900:SRSa

[1528] Anonymous. The single-rail suspended railway. Scientific American, 82(18):274, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
274a.pdf.

Hovey:1900:SAR

[1529] E. O. Hovey. The Seventh Annual Reception and Exhibition of the New York Academy of Sciences. Scientific American, 82(18):274–275, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n18/pdf/scientificamerican05051900-274b.pdf.

Anonymous:1900:FDD

[1530] Anonymous. Floating dry docks for New York City. Scientific American, 82(18):275, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
275.pdf.

Anonymous:1900:PENc

[1531] Anonymous. Paris Exposition notes. Scientific American, 82(18):275, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n18/pdf/scientificamerican05051900-275b.pdf.

Anonymous:1900:DDA

[1532] Anonymous. Death of the Duke of Argyll. Scientific American, 82(18):275, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
275c.pdf.

Anonymous:1900:NFF

[1533] Anonymous. A new form of fiber-cleaning machine. Scientific American, 82(18):276, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
276.pdf.

Brown:1900:LTE

[1534] Enos Brown. Large traction engines for Siberia. Scientific American, 82(18):276, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
276a.pdf.

Anonymous:1900:IEV

[1535] Anonymous. An improved engine valve-gear. Scientific American, 82(18):276, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
276b.pdf.

Anonymous:1900:ISCb

[1536] Anonymous. The Ives system of color photography. Scientific American, 82(18):277–278, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
277.pdf.

Anonymous:1900:UAM

[1537] Anonymous. The upbuilding of the American merchant marine. Scientific American, 82(18):278–279, May 5, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n18/pdf/scientificamerican05051900-278.pdf.

Anonymous:1900:LSR

[1538] Anonymous. The Langen single-rail suspended railway. Scientific American, 82(18):280, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
280.pdf.

Anonymous:1900:EDD

[1539] Anonymous. Effective ditch-digging machine. Scientific American, 82(18):281–282, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
281.pdf.

Anonymous:1900:CPCa

[1540] Anonymous. A convenient panoramic camera. Scientific American, 82(18):282–283, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
282.pdf.

Anonymous:1900:SSG

[1541] Anonymous. Science settles a great problem. Scientific American, 82(18):284, May 5, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n18/pdf/scientificamerican05051900-
284.pdf.

Anonymous:1900:Ch

[1542] Anonymous. Correspondence. Scientific American, 82(19):299, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n19/pdf/scientificamerican05121900-299.pdf.

Anonymous:1900:RPIk

[1543] Anonymous. Recently patented inventions. Scientific American, 82(19):299, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
299a.pdf.

Anonymous:1900:BPk

[1544] Anonymous. Business and personal. Scientific American, 82(19):300, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n19/pdf/scientificamerican05121900-300.pdf.

Anonymous:1900:NQk

[1545] Anonymous. Notes and queries. Scientific American, 82(19):300, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n19/pdf/scientificamerican05121900-300a.pdf.

Anonymous:1900:IIl

[1546] Anonymous. Index of inventions. Scientific American, 82(19):300, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n19/pdf/scientificamerican05121900-300b.pdf.

Anonymous:1900:PNPk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n19/pdf/scientificamerican05121900-303.pdf.

Anonymous:1900:PNC

[1548] Anonymous. The passage of the Nicaragua Canal Bill by the House. Scientific American, 82(19):290, May 12, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n19/pdf/scientificamerican05121900-290.pdf.

Anonymous:1900:ALG

[1549] Anonymous. American locomotives in Germany. Scientific American, 82(19):290, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
290a.pdf.

Anonymous:1900:CSB

[1550] Anonymous. The cantilever system of bridge construction. Scientific American, 82(19):290, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
290b.pdf.

Anonymous:1900:SLI

[1551] Anonymous. Science for the love of it. Scientific American, 82(19):290, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n19/pdf/scientificamerican05121900-290c.pdf.

Anonymous:1900:CNG

[1552] Anonymous. The Cape Nome gold fields. Scientific American, 82(19):290, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
290d.pdf.

Talbot:1900:STM

[1553] F. A. A. Talbot. Steering torpedoes by means of wireless telegraphy. Scientific American, 82(19):291, May 12, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n19/pdf/scientificamerican05121900-291.pdf.

Anonymous:1900:PENd

[1554] Anonymous. Paris Exposition notes. Scientific American, 82(19):291, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n19/pdf/scientificamerican05121900-291a.pdf.

Anonymous:1900:AN

[1555] Anonymous. Automobile news. Scientific American, 82(19):291, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n19/pdf/scientificamerican05121900-291b.pdf.

Anonymous:1900:NWG

[1556] Anonymous. Novel window-guard and fly-screen. Scientific American, 82(19):292, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
292.pdf.

Anonymous:1900:LKE

[1557] Anonymous. Lord Kelvin’s electric trolley rail tester. Scientific American, 82(19):292, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
292a.pdf.

Anonymous:1900:NFW

[1558] Anonymous. A new form of weeding-machine. Scientific American, 82(19):292, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
292b.pdf.

Anonymous:1900:PVT

[1559] Anonymous. The Pollak and Virag telegraph. Scientific American, 82(19):293, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
293.pdf.

Coffee:1900:OAP

[1560] W. F. Coffee. The “Orograph” an automatic profile recorder. Scientific American, 82(19):293–294, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
293a.pdf.

Anonymous:1900:TAN

[1561] Anonymous. The towers and approaches of the new East River Bridge. Scientific American, 82(19):294–295, May 12, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n19/pdf/scientificamerican05121900-294.pdf.

Anonymous:1900:CMBa

[1562] Anonymous. The Children’s Museum of Brooklyn Institute. Scientific American, 82(19):296–297, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
296.pdf.

Anonymous:1900:TOB

[1563] Anonymous. Tall office building model for the World’s Fair. Scientific American, 82(19):298, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
298.pdf.

Anonymous:1900:MRT

[1564] Anonymous. The mono-rail track velocipede. Scientific American, 82(19):298, May 12, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n19/pdf/scientificamerican05121900-
298a.pdf.

Anonymous:1900:Ci

[1565] Anonymous. Correspondence. Scientific American, 82(2):22, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-22.pdf.

Anonymous:1900:RPIl

[1566] Anonymous. Recently patented inventions. Scientific American, 82(2):28, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
28.pdf.

Anonymous:1900:BNBi

[1567] Anonymous. Books: New books, etc. Scientific American, 82 (2):28, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
28a.pdf.

Anonymous:1900:BPl

[1568] Anonymous. Business and personal. Scientific American, 82(2):29, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-29.pdf.

Anonymous:1900:NQl

[1569] Anonymous. Notes and queries. Scientific American, 82(2):29, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-29a.pdf.

Anonymous:1900:IIm

[1570] Anonymous. Index of inventions. Scientific American, 82(2):29, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-29c.pdf.

Anonymous:1900:PNPl

[1571] Anonymous. Patent notes: Patents. Scientific American, 82(2):31, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-31.pdf.

Anonymous:1900:AT

[1572] Anonymous. Armored trains. Scientific American, 82(2):18, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-18.pdf.

Anonymous:1900:NC

[1573] Anonymous. The new century. Scientific American, 82(2):18, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-18a.pdf.

Anonymous:1900:EFN

[1574] Anonymous. The electric fish of the Nile. Scientific American, 82(2):18, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
18b.pdf.

Anonymous:1900:UFR

[1575] Anonymous. Unearthing fossil remains. Scientific American, 82 (2):18, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
18c.pdf.

Anonymous:1900:CP

[1576] Anonymous. Congress of Photography in 1900. Scientific American, 82(2):18, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
18d.pdf.

Anonymous:1900:GSA

[1577] Anonymous. The Geological Society of America. Scientific American, 82(2):19, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
19.pdf.

Haile:1900:PEP

[1578] Chester Haile. Powerful electric pump at Austin, Texas. Scientific American, 82(2):20, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
20.pdf.

Anonymous:1900:OED

[1579] Anonymous. Oil engine with direct-connected generator. Scientific American, 82(2):21, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
21.pdf.

Anonymous:1900:ITG

[1580] Anonymous. An ingenious tool-grinding machine. Scientific American, 82(2):21, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
21a.pdf.

Anonymous:1900:STCa

[1581] Anonymous. A simple trousers clasp. Scientific American, 82 (2):22, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
22a.pdf.

Anonymous:1900:ALL

[1582] Anonymous. An automatic lubricator for locomotives and marine engines. Scientific American, 82(2):22–23, January 13, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n2/pdf/scientificamerican01131900-22b.pdf.

Anonymous:1900:HWN

[1583] Anonymous. Homemade windmills in Nebraska. Scientific American, 82(2):24, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
24.pdf.

Anonymous:1900:BRW

[1584] Anonymous. The Bonner rail wagon. Scientific American, 82(2):25, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-25.pdf.

Anonymous:1900:HPC

[1585] Anonymous. A 70,000 horse power central station. Scientific American, 82(2):26, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
26.pdf.

Holder:1900:SJE

[1586] Chas. Fred. Holder. The San Jacinto Earthquake. Scientific American, 82(2):27, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n2/pdf/scientificamerican01131900-
27.pdf.

Anonymous:1900:Id

[1587] Anonymous. To inventors. Scientific American, 82(2):29, January 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n2/pdf/scientificamerican01131900-29b.pdf.

Anonymous:1900:RPIm

[1588] Anonymous. Recently patented inventions. Scientific American, 82(20):316, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
316.pdf.

Anonymous:1900:BNBj

[1589] Anonymous. Books: New books, etc. Scientific American, 82(20):316, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-316a.pdf.

Anonymous:1900:BPm

[1590] Anonymous. Business and personal. Scientific American, 82(20):316, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-316b.pdf.

Anonymous:1900:NQm

[1591] Anonymous. Notes and queries. Scientific American, 82(20):316, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-316c.pdf.

Anonymous:1900:IIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-317a.pdf.

Anonymous:1900:PNPm

[1593] Anonymous. Patent notes: Patents. Scientific American, 82(20):319, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-319.pdf.

Anonymous:1900:CPP

[1594] Anonymous. Capped projectiles and panic legislation. Scientific American, 82(20):306, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
306.pdf.

Anonymous:1900:SLN

[1595] Anonymous. The shrinkage of Lake Nicaragua. Scientific American, 82(20):306, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
306a.pdf.

Anonymous:1900:EFT

[1596] Anonymous. Electricity in the field of transportation — a forecast. Scientific American, 82(20):306–307, May 19, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n20/pdf/scientificamerican05191900-306b.pdf.

Anonymous:1900:NN

[1597] Anonymous. A national need. Scientific American, 82(20):307, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-307.pdf.

Benjamin:1900:MEA

[1598] Marcus Benjamin. The methods employed by the Assay Commission. Scientific American, 82(20):307, May 19, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n20/pdf/scientificamerican05191900-307a.pdf.

Anonymous:1900:SBTa

[1599] Anonymous. A speedy business transaction with China. Scientific American, 82(20):307, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
307b.pdf.

Anonymous:1900:GSB

[1600] Anonymous. Great submarine blast in San Francisco Bay. Scientific American, 82(20):308, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
308.pdf.

Field:1900:NRB

[1601] C. Field. The new Russian battleship “Petropavlovsk”. Scientific American, 82(20):308, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
308a.pdf.

Anonymous:1900:DVC

[1602] Anonymous. Death Valley, California. Scientific American, 82 (20):309–310, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
309.pdf.

Anonymous:1900:NCH

[1603] Anonymous. A new curtain-hanger. Scientific American, 82(20):310, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-310.pdf.

Anonymous:1900:SRSb

[1604] Anonymous. A sterilizer for razors and surgical instruments. Scientific American, 82(20):310–311, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
310a.pdf.

Anonymous:1900:MGAa

[1605] Anonymous. Manufacture of guns and armor at the Bethlehem Steel Works. Scientific American, 82(20):312–313, May 19, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n20/pdf/scientificamerican05191900-312.pdf.

Anonymous:1900:PNM

[1606] Anonymous. Proposed National Memorial Bridge across the Potomac at Washington. Scientific American, 82(20):313, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-313.pdf.

Anonymous:1900:AIM

[1607] Anonymous. Alexandre III Memorial Bridge, Paris. Scientific American, 82(20):313–314, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
313a.pdf.

Anonymous:1900:PCR

[1608] Anonymous. A pair of curious relics at Aberfoyle, Scotland. Scientific American, 82(20):314, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
314.pdf.

Talbot:1900:AP

[1609] F. A. A. Talbot. Aerial photography. Scientific American, 82(20):314, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-314a.pdf.

Anonymous:1900:GDM

[1610] Anonymous. Gold deposits and mining methods at Cape Nome. Scientific American, 82(20):315, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n20/pdf/scientificamerican05191900-
315.pdf.

Anonymous:1900:Ie

[1611] Anonymous. To inventors. Scientific American, 82(20):317, May 19, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n20/pdf/scientificamerican05191900-317.pdf.

Russell:1900:HJa

[1612] Henry Norris Russell. The heavens in June. Scientific American, 82(21):323, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
323.pdf.

Anonymous:1900:Cj

[1613] Anonymous. Correspondence. Scientific American, 82(21):330, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-330.pdf.

Anonymous:1900:RPIn

[1614] Anonymous. Recently patented inventions. Scientific American, 82(21):332, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
332.pdf.

Anonymous:1900:BPn

[1615] Anonymous. Business and personal. Scientific American, 82(21):332, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-332a.pdf.

Anonymous:1900:NQn

[1616] Anonymous. Notes and queries. Scientific American, 82(21):332, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-332b.pdf.

Anonymous:1900:BNBk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-332c.pdf.

Anonymous:1900:IIo

[1618] Anonymous. Index of inventions. Scientific American, 82(21):332, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-332e.pdf.

Anonymous:1900:PNPn

[1619] Anonymous. Patent notes: Patents. Scientific American, 82(21):335, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-335.pdf.

Anonymous:1900:PNYa

[1620] Anonymous. Progress of the New York Rapid Transit Tunnel. Scientific American, 82(21):322, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
322.pdf.

Anonymous:1900:SNM

[1621] Anonymous. Some needs of modern chemistry. Scientific American, 82(21):322, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
322a.pdf.

Anonymous:1900:TIB

[1622] Anonymous. Two impossible bills. Scientific American, 82(21):322, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-322b.pdf.

Anonymous:1900:PST

[1623] Anonymous. The Postal Service in the time of Queen Anne. Scientific American, 82(21):322, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
322c.pdf.

Stewart:1900:TMN

[1624] J. A. Stewart. Textile manufacturing in the New South. Scientific American, 82(21):322–323, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
322d.pdf.

Anonymous:1900:ECN

[1625] Anonymous. The emigration to Cape Nome. Scientific American, 82(21):323, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
323a.pdf.

Anonymous:1900:SAEa

[1626] Anonymous. A small automatic electric air-pump. Scientific American, 82(21):324, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
324.pdf.

Anonymous:1900:SCL

[1627] Anonymous. Shouldering car for leveling and trimming roadbed. Scientific American, 82(21):324, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
324a.pdf.

Anonymous:1900:CZA

[1628] Anonymous. Count von Zeppelin’s airship. Scientific American, 82(21):325–326, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
325.pdf.

Anonymous:1900:BON

[1629] Anonymous. Borax — old and new methods of production. Scientific American, 82(21):326–327, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
326.pdf.

Fawcett:1900:SPC

[1630] Waldon Fawcett. The Sault Power Canal. Scientific American, 82(21):328–329, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
328.pdf.

Anonymous:1900:CCY

[1631] Anonymous. Curiously cut yew trees. Scientific American, 82(21):329, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-329.pdf.

Anonymous:1900:WDM

[1632] Anonymous. A waterproof detonator for mines. Scientific American, 82(21):330, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
330a.pdf.

Anonymous:1900:MCP

[1633] Anonymous. A mechanical cotton-picker. Scientific American, 82 (21):330, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n21/pdf/scientificamerican05261900-
330b.pdf.

Holder:1900:SPS

[1634] Chas. Fred’k. Holder. Some Pacific sharks and accompanying fishes. Scientific American, 82(21):331, May 26, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n21/pdf/scientificamerican05261900-331.pdf.

Anonymous:1900:If

[1635] Anonymous. To inventors. Scientific American, 82(21):332, May 26, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n21/pdf/scientificamerican05261900-332d.pdf.

Anonymous:1900:RPIo (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
348.pdf.

Anonymous:1900:BPo

[1637] Anonymous. Business and personal. Scientific American, 82(22):348, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-348a.pdf.

Anonymous:1900:NQo

[1638] Anonymous. Notes and queries. Scientific American, 82(22):348, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-348b.pdf.

Anonymous:1900:BNBl

[1639] Anonymous. Books: New books, etc. Scientific American, 82(22):348, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-348c.pdf.

Anonymous:1900:IIp

[1640] Anonymous. Index of inventions. Scientific American, 82(22):349, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-349a.pdf.

Anonymous:1900:PNPo

[1641] Anonymous. Patent notes: Patents. Scientific American, 82(22):351, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-351.pdf.

Anonymous:1900:PHP

[1642] Anonymous. Preliminary hearing of the President’s Canal Commission. Scientific American, 82(22):338, June 2, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n22/pdf/scientificamerican06021900-338.pdf.

Anonymous:1900:MP

[1643] Anonymous. Mutually profitable. Scientific American, 82(22):338, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-338a.pdf.

Anonymous:1900:CPD

[1644] Anonymous. Congress and the pneumatic dispatch system in this City. Scientific American, 82(22):338, June 2, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n22/pdf/scientificamerican06021900-338b.pdf.

Anonymous:1900:DCP

[1645] Anonymous. Development of color-photography. Scientific American, 82(22):338, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
338c.pdf.

Anonymous:1900:ATA

[1646] Anonymous. The American trotter abroad. Scientific American, 82(22):338–339, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
338d.pdf.

Anonymous:1900:JG

[1647] Anonymous. The Japanese grasses. Scientific American, 82(22):339, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-339.pdf.

Anonymous:1900:DJG

[1648] Anonymous. Death of Jonas Gilman Clark. Scientific American, 82(22):339, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
339a.pdf.

Anonymous:1900:PENe

[1649] Anonymous. Paris Exposition notes. Scientific American, 82(22):339, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-339b.pdf.

Anonymous:1900:SRSc

[1650] Anonymous. Single-rail storage battery motor. Scientific American, 82(22):340, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
340.pdf.

Anonymous:1900:PSIa

[1651] Anonymous. Portable shields for infantry. Scientific American, 82(22):340, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
340a.pdf.

Anonymous:1900:WMb

[1652] Anonymous. Wood mosaic. Scientific American, 82(22):341, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-341.pdf.

Gamble:1900:AGS

[1653] William B. Gamble. The assaying of gold and silver ores. Scientific American, 82(22):341–342, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
341a.pdf.

Stewart:1900:CTS

[1654] J. A. Stewart. Cotton trade schools in the South. Scientific American, 82(22):342–343, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
342.pdf.

Anonymous:1900:NFCa

[1655] Anonymous. A new form of car fender. Scientific American, 82(22):343, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-343.pdf.

Anonymous:1900:PEc

[1656] Anonymous. The Paris Exposition. Scientific American, 82(22): 344–345, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
344.pdf.

Anonymous:1900:ECSa

[1657] Anonymous. Electric cab system of Paris. Scientific American, 82(22):345–346, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
345.pdf.

Wade:1900:DPI

[1658] Herbert T. Wade. Dr. Pupin’s improvements in long-distance telephony. Scientific American, 82(22):346–347, June 2, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n22/pdf/scientificamerican06021900-346.pdf.

Aaron:1900:SRB

[1659] Frank Aaron. The store-room beetle or book worm. Scientific American, 82(22):347, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n22/pdf/scientificamerican06021900-
347.pdf.

Anonymous:1900:Ig

[1660] Anonymous. To inventors. Scientific American, 82(22):349, June 2, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n22/pdf/scientificamerican06021900-349.pdf.

Anonymous:1900:Ck

[1661] Anonymous. Correspondence. Scientific American, 82(23):363, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-363b.pdf.

Anonymous:1900:RPIp (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
364.pdf.

Anonymous:1900:BPp

[1663] Anonymous. Business and personal. Scientific American, 82(23):363, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-364a.pdf.

Anonymous:1900:NQp

[1664] Anonymous. Notes and queries. Scientific American, 82(23):363, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-364b.pdf.

Anonymous:1900:BNBm

[1665] Anonymous. Books: New books, etc. Scientific American, 82(23):363, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-364c.pdf.

Anonymous:1900:IIq

[1666] Anonymous. Index of inventions. Scientific American, 82(23):363, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-364e.pdf.

Anonymous:1900:PNPp

[1667] Anonymous. Patent notes: Patents. Scientific American, 82(23):367, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-367.pdf.

Anonymous:1900:ARR

[1668] Anonymous. Atmospheric resistance to railroad trains. Scientific American, 82(23):354, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
354.pdf.

Anonymous:1900:TES

[1669] Anonymous. The total eclipse of the Sun. Scientific American, 82(23):354, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
354a.pdf.

Anonymous:1900:ST

[1670] Anonymous. Scientific translation. Scientific American, 82(23):354, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-354b.pdf.

Anonymous:1900:CLT

[1671] Anonymous. The Central London Tunnel Road. Scientific American, 82(23):354–355, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
354c.pdf.

Anonymous:1900:SPZ

[1672] Anonymous. Smelting process for zinc ores. Scientific American, 82(23):355, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
355.pdf.

Anonymous:1900:NTP

[1673] Anonymous. A new tele-photo lens. Scientific American, 82(23):355, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-355a.pdf.

Anonymous:1900:LD

[1674] Anonymous. The latest developer. Scientific American, 82(23):355, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-355b.pdf.

Anonymous:1900:IDa

[1675] Anonymous. An improvement in driers. Scientific American, 82 (23):356, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
356.pdf.

Anonymous:1900:ETT

[1676] Anonymous. Experimental train for testing atmospheric resistance. Scientific American, 82(23):356, June 9, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n23/pdf/scientificamerican06091900-356a.pdf.

Anonymous:1900:NPA

[1677] Anonymous. New plow attachment. Scientific American, 82(23):356, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-356b.pdf.

Anonymous:1900:MSa

[1678] Anonymous. The manufacture of salt. Scientific American, 82 (23):357–358, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
357.pdf.

Anonymous:1900:MGAb

[1679] Anonymous. Manufacture of guns and armor at the Bethlehem Steel Works. Scientific American, 82(23):358–359, June 9, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n23/pdf/scientificamerican06091900-358.pdf.

Webb:1900:OEM

[1680] Henry Randall Webb. Observations on the eclipse of May 28, 1900. Scientific American, 82(23):360, June 9, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n23/pdf/scientificamerican06091900-360.pdf.

Harwood:1900:MCA

[1681] W. S. Harwood. The mummification of cats in Ancient Egypt. Scientific American, 82(23):361, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
361.pdf.

Anonymous:1900:USB

[1682] Anonymous. United States battleship “Georgia” and class. Scientific American, 82(23):361–362, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
361a.pdf.

Anonymous:1900:HMN

[1683] Anonymous. An hydraulic motor of novel construction. Scientific American, 82(23):363, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
363.pdf.

Anonymous:1900:ADC

[1684] Anonymous. An attractive dwelling and conservatory for. Scientific American, 82(23):363, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n23/pdf/scientificamerican06091900-
363a.pdf.

Anonymous:1900:Ih

[1685] Anonymous. To inventors. Scientific American, 82(23):363, June 9, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n23/pdf/scientificamerican06091900-364d.pdf.

Anonymous:1900:Cl

[1686] Anonymous. Correspondence. Scientific American, 82(24):375, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-375.pdf.

Anonymous:1900:RPIq

[1687] Anonymous. Recently patented inventions. Scientific American, 82(24):380, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
380.pdf.

Anonymous:1900:BNBn

[1688] Anonymous. Books: New books, etc. Scientific American, 82(24):380, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-380a.pdf.

Anonymous:1900:BPq

[1689] Anonymous. Business and personal. Scientific American, 82(24):381, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-381.pdf.

Anonymous:1900:NQq

[1690] Anonymous. Notes and queries. Scientific American, 82(24):381, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-381a.pdf.

Anonymous:1900:IIr

[1691] Anonymous. Index of inventions. Scientific American, 82(24):381, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-381b.pdf.

Anonymous:1900:PNPq

[1692] Anonymous. Patent notes: Patents. Scientific American, 82(24):383, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-383.pdf.

Anonymous:1900:APF

[1693] Anonymous. The armor plate fiasco. Scientific American, 82(24):370, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-370.pdf.

Anonymous:1900:TRA

[1694] Anonymous. Two remarkable achievements in chemical physics. Scientific American, 82(24):370, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
370a.pdf.

Anonymous:1900:OPEc

[1695] Anonymous. Our phenomenal exports. Scientific American, 82(24):370, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-370b.pdf.

Anonymous:1900:RSA

[1696] Anonymous. The railroad systems of Asia. Scientific American, 82(24):370–371, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
370c.pdf.

Adams:1900:ESG

[1697] Alton D. Adams. Electrical supply by gas companies. Scientific American, 82(24):371, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
371.pdf.

Anonymous:1900:PENf

[1698] Anonymous. Paris Exposition notes. Scientific American, 82(24):371, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-371a.pdf.

Anonymous:1900:FTC

[1699] Anonymous. Felling tall chimneys. Scientific American, 82(24):372, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-372.pdf.

Anonymous:1900:HFY

[1700] Anonymous. The Herreshoff 70-foot yacht “Mineola”. Scientific American, 82(24):373, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
373.pdf.

Willey:1900:EHL

[1701] D. Allen Willey. Electric head light effects. Scientific American, 82(24):373–374, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
373a.pdf.

Anonymous:1900:SUA

[1702] Anonymous. Some unique attractions of the Paris Exposition. Scientific American, 82(24):374, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
374.pdf.

Anonymous:1900:AHP

[1703] Anonymous. An adjustable hoisting pulley. Scientific American, 82(24):374, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
374a.pdf.

Anonymous:1900:PSEa

[1704] Anonymous. A photograph of the solar eclipse. Scientific American, 82(24):376, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
376.pdf.

Anonymous:1900:RLB

[1705] Anonymous. Rolling-lift bridge over Fort Point Channel, Boston. Scientific American, 82(24):376, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
376a.pdf.

Anonymous:1900:MCN

[1706] Anonymous. Manufacture of carborundum at Niagara Falls. Scientific American, 82(24):377–378, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
377.pdf.

Anonymous:1900:ESG

[1707] Anonymous. Efficient shot-gun cleaner. Scientific American, 82(24):378, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n24/pdf/scientificamerican06161900-378.pdf.

Anonymous:1900:IVM

[1708] Anonymous. Improved vending-machine. Scientific American, 82 (24):378, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
378a.pdf.

Wilson:1900:CHW

[1709] Robert Wilson. A curious historic weapon. Scientific American, 82(24):379, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
379.pdf.

Anonymous:1900:ETO

[1710] Anonymous. Etruscan tombs at Orvieto. Scientific American, 82 (24):379, June 16, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n24/pdf/scientificamerican06161900-
379a.pdf.

Anonymous:1900:Cm

[1711] Anonymous. Correspondence. Scientific American, 82(25):390, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-390.pdf.

Anonymous:1900:RPIr

[1712] Anonymous. Recently patented inventions. Scientific American, 82(25):396, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
396.pdf.

Anonymous:1900:BNBo

[1713] Anonymous. Books: New books, etc. Scientific American, 82(25):396, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-396a.pdf.

Anonymous:1900:BPr

[1714] Anonymous. Business and personal. Scientific American, 82(25):396, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-396b.pdf.

Anonymous:1900:NQr

[1715] Anonymous. Notes and queries. Scientific American, 82(25):396, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-396c.pdf.

Anonymous:1900:IIs

[1716] Anonymous. Index of inventions. Scientific American, 82(25):397, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-397.pdf.

Anonymous:1900:PNPr

[1717] Anonymous. Patent notes: Patents. Scientific American, 82(25):399, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-399.pdf.

Anonymous:1900:BF

[1718] Anonymous. The battleship of the future. Scientific American, 82(25):386, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
386.pdf.

Anonymous:1900:EFL

[1719] Anonymous. Electric furnace litigation — an important patent decision. Scientific American, 82(25):386, June 23, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n25/pdf/scientificamerican06231900-386a.pdf.

Anonymous:1900:ERC

[1720] Anonymous. An experiment on the Roman Campagna with the malarial mosquito. Scientific American, 82(25):386–387, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-386b.pdf.

Varney:1900:CTC

[1721] George J. Varney. Curious things in clouds. Scientific American, 82(25):387, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
387.pdf.

Anonymous:1900:EFF

[1722] Anonymous. Experimental formation of the fluorides of sulphur. Scientific American, 82(25):387, June 23, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n25/pdf/scientificamerican06231900-387a.pdf.

Anonymous:1900:DDP

[1723] Anonymous. Death of Dr. Paul Gibier. Scientific American, 82(25):387, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-387b.pdf.

Anonymous:1900:AGE

[1724] Anonymous. An acetylene gas explosion. Scientific American, 82 (25):387, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
387c.pdf.

Bradshaw:1900:FSC

[1725] W. F. Bradshaw, Jr. The first submarine cable. Scientific American, 82(25):388, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
388.pdf.

Anonymous:1900:ATE

[1726] Anonymous. Armored traction engine and train. Scientific American, 82(25):388–389, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
388a.pdf.

Perkins:1900:ECM

[1727] Frank C. Perkins. Electrical coal-mining machines. Scientific American, 82(25):389, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
389.pdf.

Anonymous:1900:TAS

[1728] Anonymous. A truck attachment for sulky-rakes. Scientific American, 82(25):390–391, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
390a.pdf.

Larned:1900:USM

[1729] Charles W. Larned. The United States Military Academy. Scientific American, 82(25):392–393, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
392.pdf.

Anonymous:1900:FPb

[1730] Anonymous. Forest preservation. Scientific American, 82(25):393–394, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n25/pdf/scientificamerican06231900-393.pdf.

Anonymous:1900:BAB

[1731] Anonymous. The Baldwin acetylene bicycle-lamp. Scientific American, 82(25):394, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
394.pdf.

Anonymous:1900:SWN

[1732] Anonymous. A simple wrench of novel form. Scientific American, 82(25):395, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
395.pdf.

Holder:1900:DZS

[1733] Charles Frederick Holder. Diving for zoological specimens. Scientific American, 82(25):395, June 23, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n25/pdf/scientificamerican06231900-
395a.pdf.

Russell:1900:HJb

[1734] Henry Norris Russell. The heavens in July. Scientific American, 82(26):403, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
403a.pdf.

Anonymous:1900:Cn

[1735] Anonymous. Correspondence. Scientific American, 82(26):410, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n26/pdf/scientificamerican06301900-410.pdf.

Anonymous:1900:RPIs

[1736] Anonymous. Recently patented inventions. Scientific American, 82(26):410–411, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
410a.pdf.

Anonymous:1900:BPs

[1737] Anonymous. Business and personal. Scientific American, 82(26):411, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n26/pdf/scientificamerican06301900-411.pdf.

Anonymous:1900:NQs

[1738] Anonymous. Notes and queries. Scientific American, 82(26):411, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n26/pdf/scientificamerican06301900-411a.pdf.

Anonymous:1900:BNBp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n26/pdf/scientificamerican06301900-411b.pdf.

Anonymous:1900:IIt

[1740] Anonymous. Index of inventions. Scientific American, 82(26):411, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n26/pdf/scientificamerican06301900-411c.pdf.

Anonymous:1900:Ij

[1741] Anonymous. Index. Scientific American, 82(26):413–414, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n26/pdf/scientificamerican06301900-413.pdf.

Anonymous:1900:PNPs

[1742] Anonymous. Patent notes: Patents. Scientific American, 82(26):415, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n26/pdf/scientificamerican06301900-415.pdf.

Anonymous:1900:LLA

[1743] Anonymous. The latest liquid air fallacy. Scientific American, 82(26):402, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
402.pdf.

Anonymous:1900:EVGa

[1744] Anonymous. Economic value of good roads. Scientific American, 82(26):402, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
402a.pdf.

Anonymous:1900:CESa

[1745] Anonymous. Commercial expansion as a science. Scientific American, 82(26):402–403, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
402b.pdf.

Benjamin:1900:RSW

[1746] Marcus Benjamin. Robert Simpson Woodward. Scientific American, 82(26):403, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
403.pdf.

Anonymous:1900:SLF

[1747] Anonymous. A self-launching fire-ladder. Scientific American, 82 (26):404, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
404.pdf.

Anonymous:1900:IVS

[1748] Anonymous. An improved valve for steam radiators. Scientific American, 82(26):404, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
404a.pdf.

Anonymous:1900:EDW

[1749] Anonymous. Examination of the drinking water of Brooklyn, New York. Scientific American, 82(26):404–405, June 30, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n26/pdf/scientificamerican06301900-404b.pdf.

Anonymous:1900:TNHa

[1750] Anonymous. Two notable high-speed foreign cruisers. Scientific American, 82(26):405–406, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
405.pdf.

Anonymous:1900:SPP

[1751] Anonymous. Some perspectives of the Paris Exposition. Scientific American, 82(26):406–407, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
406.pdf.

Anonymous:1900:MA

[1752] Anonymous. A military automobile. Scientific American, 82(26):408, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n26/pdf/scientificamerican06301900-408.pdf.

Anonymous:1900:OPA

[1753] Anonymous. Ostia, the port of ancient rome. Scientific American, 82(26):409–410, June 30, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n26/pdf/scientificamerican06301900-
409.pdf.

Anonymous:1900:Co

[1754] Anonymous. Correspondence. Scientific American, 82(3):38, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-38.pdf.

Anonymous:1900:RPIt

[1755] Anonymous. Recently patented inventions. Scientific American, 82(3):44, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
44.pdf.

Anonymous:1900:BNBq

[1756] Anonymous. Books: New books, etc. Scientific American, 82 (3):44, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
44a.pdf.

Anonymous:1900:BPt

[1757] Anonymous. Business and personal. Scientific American, 82(3):44, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-44b.pdf.

Anonymous:1900:NQt

[1758] Anonymous. Notes and queries. Scientific American, 82(3):44–45, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-44c.pdf.

Anonymous:1900:IIu

[1759] Anonymous. Index of inventions. Scientific American, 82(3):45, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-45.pdf.

Anonymous:1900:PNPt

[1760] Anonymous. Patent notes: Patents. Scientific American, 82(3):47, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-47.pdf.

Anonymous:1900:LF

[1761] Anonymous. Legislative foolishness. Scientific American, 82(3):34, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-34.pdf.

Anonymous:1900:OTU

[1762] Anonymous. Official trial of the United States cruiser “Albany”. Scientific American, 82(3):34, January 20, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n3/pdf/scientificamerican01201900-34a.pdf.

Anonymous:1900:ICa

[1763] Anonymous. The Isthmian Canal. Scientific American, 82(3):34, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-34b.pdf.

Anonymous:1900:AGR

[1764] Anonymous. American goods in Russia. Scientific American, 82(3):34, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
34c.pdf.

Carden:1900:BBG

[1765] Godfrey L. Carden. British and Boer guns — a lesson from the South African War. Scientific American, 82(3):35, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-35.pdf.

Anonymous:1900:PP

[1766] Anonymous. Palisade preservation. Scientific American, 82(3):35, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n3/pdf/scientificamerican01201900-35a.pdf.

Anonymous:1900:KWA

[1767] Anonymous. The “Kaiser Wilhelm” again breaks her record. Scientific American, 82(3):35, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
35b.pdf.

Anonymous:1900:ME

[1768] Anonymous. Magnetism of the Earth. Scientific American, 82 (3):35, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
35c.pdf.

Anonymous:1900:TBM

[1769] Anonymous. Torpedo boat makes over 35 knots an hour. Scientific American, 82(3):35, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
35d.pdf.

Anonymous:1900:RCL

[1770] Anonymous. A remarkable collection of lamps. Scientific American, 82(3):35, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
35e.pdf.

Anonymous:1900:WI

[1771] Anonymous. A warning to the inventor. Scientific American, 82(3):35, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
35f.pdf.

Anonymous:1900:NCD

[1772] Anonymous. The new Croton Dam — New York City’s water supply. Scientific American, 82(3):36, January 20, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n3/pdf/scientificamerican01201900-36.pdf.

Anonymous:1900:GPP

[1773] Anonymous. A guard for the projecting parts of machinery. Scientific American, 82(3):37, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
37.pdf.

Anonymous:1900:PCF

[1774] Anonymous. The Pierce cushion-frame, chainless bicycle. Scientific American, 82(3):37, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
37a.pdf.

Anonymous:1900:ICP

[1775] Anonymous. An improved chain pipe-wrench. Scientific American, 82(3):37, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
37b.pdf.

Anonymous:1900:HWM

[1776] Anonymous. How a weather map is made. Scientific American, 82(3):38–39, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
38a.pdf.

James:1900:DCD

[1777] George Wharton James. Discovery of cliff dwellings in the Southwest. Scientific American, 82(3):40–41, January 20, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n3/pdf/scientificamerican01201900-40.pdf.

Anonymous:1900:AOL

[1778] Anonymous. The armament of our latest warships. Scientific American, 82(3):41–42, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
41.pdf.

Dailey:1900:EMR

[1779] Howard B. Dailey. Electro-magnetic rotation. Scientific American, 82(3):42, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
42.pdf.

Anonymous:1900:DFM

[1780] Anonymous. The Danilewsky flying machine. Scientific American, 82(3):43, January 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n3/pdf/scientificamerican01201900-
43.pdf.

Anonymous:1900:RPIu

[1781] Anonymous. Recently patented inventions. Scientific American, 82(4):60–61, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
60.pdf.

Anonymous:1900:BNBr

[1782] Anonymous. Books: New books, etc. Scientific American, 82 (4):61, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
61.pdf.

Anonymous:1900:BPu

[1783] Anonymous. Business and personal. Scientific American, 82(4):61, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-61a.pdf.

Anonymous:1900:NQu

[1784] Anonymous. Notes and queries. Scientific American, 82(4):61, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-61b.pdf.

Anonymous:1900:IIv

[1785] Anonymous. Index of inventions. Scientific American, 82(4):61, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-61c.pdf.

Anonymous:1900:PNPu

[1786] Anonymous. Patent notes: Patents. Scientific American, 82(4):63, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-63.pdf.

Anonymous:1900:NECa

[1787] Anonymous. A new era in city transportation. Scientific American, 82(4):50, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
50.pdf.

Anonymous:1900:AOP

[1788] Anonymous. The agricultural outlook in Puerto Rico. Scientific American, 82(4):50, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
50a.pdf.

Anonymous:1900:TM

[1789] Anonymous. “time is money”. Scientific American, 82(4):50, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-50b.pdf.

Anonymous:1900:WD

[1790] Anonymous. A wise decision. Scientific American, 82(4):50, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-50c.pdf.

Anonymous:1900:CCH

[1791] Anonymous. Comparative cost of horse and automobile. Scientific American, 82(4):50–51, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
50d.pdf.

Anonymous:1900:IW

[1792] Anonymous. Injection of wood. Scientific American, 82(4):51, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-51.pdf.

Anonymous:1900:PAE

[1793] Anonymous. The Pan-American Exposition of 1901. Scientific American, 82(4):51, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
51a.pdf.

Anonymous:1900:RW

[1794] Anonymous. The risks of war. Scientific American, 82(4):51, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-51b.pdf.

Anonymous:1900:DPE

[1795] Anonymous. Death of Professor Egleston. Scientific American, 82(4):51, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
51c.pdf.

Anonymous:1900:PFP

[1796] Anonymous. Possible Federal protection of birds. Scientific American, 82(4):51, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
51d.pdf.

Anonymous:1900:TP

[1797] Anonymous. Travel to Paris. Scientific American, 82(4):51, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-51e.pdf.

Anonymous:1900:CESb

[1798] Anonymous. Civil engineers in session. Scientific American, 82 (4):51, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
51f.pdf.

Anonymous:1900:FPH

[1799] Anonymous. A foot-power hammer. Scientific American, 82(4):52, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n4/pdf/scientificamerican01271900-52.pdf.

Anonymous:1900:CNE

[1800] Anonymous. Cables of the new East River Bridge. Scientific American, 82(4):52, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
52a.pdf.

Anonymous:1900:RFF

[1801] Anonymous. Reconstruction of Farragut’s flagship, the U.S.S. “Hartford”. Scientific American, 82(4):53–54, January 27, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n4/pdf/scientificamerican01271900-53.pdf.

Anonymous:1900:LSC

[1802] Anonymous. The “Locomobile” steam carriage. Scientific American, 82(4):54–55, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
54.pdf.

Mels:1900:NSA

[1803] Edgar Mels. The natives of South Africa. Scientific American, 82(4):56–57, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
56.pdf.

Anonymous:1900:DMK

[1804] Anonymous. The diamond mines of Kim-Berley. Scientific American, 82(4):57–58, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
57.pdf.

Anonymous:1900:CMBb

[1805] Anonymous. Curious mishap to a billiard ball. Scientific American, 82(4):58–59, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
58.pdf.

Anonymous:1900:MMC

[1806] Anonymous. Military motor carriage. Scientific American, 82 (4):59–60, January 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n4/pdf/scientificamerican01271900-
59.pdf.

Anonymous:1900:Cq

[1807] Anonymous. Correspondence. Scientific American, 82(5):71, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n5/pdf/scientificamerican02031900-71.pdf.

Anonymous:1900:RPIv

[1808] Anonymous. Recently patented inventions. Scientific American, 82(5):76, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
76.pdf.

Anonymous:1900:BNBs

[1809] Anonymous. Books: New books, etc. Scientific American, 82 (5):76, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
76a.pdf.

Anonymous:1900:BPv

[1810] Anonymous. Business and personal. Scientific American, 82 (5):77, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
77.pdf.

Anonymous:1900:NQv

[1811] Anonymous. Notes and queries. Scientific American, 82(5):77, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n5/pdf/scientificamerican02031900-77a.pdf.

Anonymous:1900:IIw

[1812] Anonymous. Index of inventions. Scientific American, 82(5): 77–79, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
77c.pdf.

Anonymous:1900:PNPv

[1813] Anonymous. Patent notes: Patents. Scientific American, 82 (5):79, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
79.pdf.

Anonymous:1900:ECPa

[1814] Anonymous. The Erie Canal problem. Scientific American, 82 (5):66, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
66.pdf.

Anonymous:1900:ES

[1815] Anonymous. Excavations at Susa. Scientific American, 82(5):66, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n5/pdf/scientificamerican02031900-66a.pdf.

Anonymous:1900:AMR

[1816] Anonymous. Automatic machinery and the Russian peasant. Scientific American, 82(5):66, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
66b.pdf.

Anonymous:1900:S

[1817] Anonymous. Stonehenge. Scientific American, 82(5):67, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n5/pdf/scientificamerican02031900-67.pdf.

Anonymous:1900:GRS

[1818] Anonymous. The “Green Ray” at sunset. Scientific American, 82(5):67, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
67a.pdf.

Brown:1900:NMD

[1819] Enos Brown. The new mining district at Cape Nome. Scientific American, 82(5):67, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
67b.pdf.

Anonymous:1900:JR

[1820] Anonymous. John Ruskin. Scientific American, 82(5):67, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n5/pdf/scientificamerican02031900-67c.pdf.

Anonymous:1900:Cs

[1821] Anonymous. Crops in 1899. Scientific American, 82(5):67, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n5/pdf/scientificamerican02031900-67d.pdf.

Anonymous:1900:DPH

[1822] Anonymous. Death of Prof. Hughes. Scientific American, 82 (5):67, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
67e.pdf.

Anonymous:1900:IEM

[1823] Anonymous. An improved electro-magnetic apparatus. Scientific American, 82(5):68, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
68.pdf.

Mann:1900:SP

[1824] Francis P. Mann. Stassano process. Scientific American, 82 (5):68, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
68a.pdf.

Anonymous:1900:RPS

[1825] Anonymous. Repairing a propeller shaft at sea. Scientific American, 82(5):68–69, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
68b.pdf.

Anonymous:1900:NYR

[1826] Anonymous. The New York Rapid Transit Tunnel. Scientific American, 82(5):69–70, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
69.pdf.

Holder:1900:I

[1827] Charles F. Holder. The ibis. Scientific American, 82(5):72, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n5/pdf/scientificamerican02031900-72.pdf.

Anonymous:1900:PSL

[1828] Anonymous. A portable search light for fire department use. Scientific American, 82(5):72–73, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
72a.pdf.

Anonymous:1900:NECb

[1829] Anonymous. New excavations at Carthage. Scientific American, 82(5):73, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
73.pdf.

Anonymous:1900:PPB

[1830] Anonymous. The proposed pneumatic balance locks for the Erie Canal. Scientific American, 82(5):74, February 3, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n5/pdf/scientificamerican02031900-74.pdf.

Anonymous:1900:ERM

[1831] Anonymous. An early reaping-machine. Scientific American, 82 (5):75–76, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n5/pdf/scientificamerican02031900-
75.pdf.

Anonymous:1900:Ik

[1832] Anonymous. To inventors. Scientific American, 82(5):77, February 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n5/pdf/scientificamerican02031900-77b.pdf.

Anonymous:1900:Ct

[1833] Anonymous. Correspondence. Scientific American, 82(6):86–87, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n6/pdf/scientificamerican02101900-86.pdf.

Anonymous:1900:RPIw

[1834] Anonymous. Recently patented inventions. Scientific American, 82 (6):92–93, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
92.pdf.

Anonymous:1900:BNBt

[1835] Anonymous. Books: New books, etc. Scientific American, 82 (6):93, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
93.pdf.

Anonymous:1900:BPw

[1836] Anonymous. Business and personal. Scientific American, 82(6): 93, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
93a.pdf.

Anonymous:1900:NQw

[1837] Anonymous. Notes and queries. Scientific American, 82(6):93, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n6/pdf/scientificamerican02101900-93b.pdf.

Anonymous:1900:IIx

[1838] Anonymous. Index of inventions. Scientific American, 82(6):93– 95, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
93c.pdf.

Anonymous:1900:PNPw

[1839] Anonymous. Patent notes: Patents. Scientific American, 82(6): 95, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
95.pdf.

Anonymous:1900:NCB

[1840] Anonymous. The Nicaragua Canal Bill — an affront to the nation. Scientific American, 82(6):82, February 10, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n6/pdf/scientificamerican02101900-82.pdf.

Anonymous:1900:BEH

[1841] Anonymous. The Bill for Establishing a High Court of Patents. Scientific American, 82(6):82, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
82a.pdf.

Anonymous:1900:DS

[1842] Anonymous. Distance and speed. Scientific American, 82(6): 82–83, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
82b.pdf.

Anonymous:1900:RNC

[1843] Anonymous. Regulation of the Nicaragua Canal summit level. Scientific American, 82(6):83, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
83.pdf.

Anonymous:1900:RSI

[1844] Anonymous. Report of the Smithsonian Institution. Scientific American, 82(6):83, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
83a.pdf.

Anonymous:1900:FCU

[1845] Anonymous. The foreign commerce of the United States in 1899. Scientific American, 82(6):83, February 10, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n6/pdf/scientificamerican02101900-83b.pdf.

Anonymous:1900:IAA

[1846] Anonymous. An International Association of Academies. Scientific American, 82(6):83, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
83c.pdf.

Anonymous:1900:WPP

[1847] Anonymous. The world’s production of precious metals. Scientific American, 82(6):83, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
83d.pdf.

Anonymous:1900:ADS

[1848] Anonymous. An automatic device for stopping engines. Scientific American, 82(6):84, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
84.pdf.

Hoferer:1900:NL

[1849] M. J. Hoferer and S. J. Nodes and loops. Scientific American, 82(6):84, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
84a.pdf.

Anonymous:1900:SAG

[1850] Anonymous. Serpollet acetylene generator. Scientific American, 82(6):84, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
84b.pdf.

Fawcett:1900:CC

[1851] Waldon Fawcett. Cantilever cranes. Scientific American, 82(6): 85–86, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
85.pdf.

Anonymous:1900:HCD

[1852] Anonymous. Half a century in the development of the bicycle. Scientific American, 82(6):88–89, February 10, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n6/pdf/scientificamerican02101900-88.pdf.

Anonymous:1900:AMSa

[1853] Anonymous. Automobiles at the Madison Square Garden Show. Scientific American, 82(6):89–91, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
89.pdf.

Anonymous:1900:RBTa

[1854] Anonymous. A remarkable banyan tree at St. Kitts. Scientific American, 82(6):91–92, February 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n6/pdf/scientificamerican02101900-
91.pdf.

Anonymous:1900:RPIx

[1855] Anonymous. Recently patented inventions. Scientific American, 82(7):108, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
108.pdf.

Anonymous:1900:BNBu

[1856] Anonymous. Books: New books, etc. Scientific American, 82 (7):108, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
108a.pdf.

Anonymous:1900:BPx

[1857] Anonymous. Business and personal. Scientific American, 82(7): 108, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
108b.pdf.

Anonymous:1900:NQx

[1858] Anonymous. Notes and queries. Scientific American, 82(7):108, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n7/pdf/scientificamerican02171900-108c.pdf.

Anonymous:1900:IIy

[1859] Anonymous. Index of inventions. Scientific American, 82(7): 108–111, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
108d.pdf.

Anonymous:1900:PNPx

[1860] Anonymous. Patent notes: Patents. Scientific American, 82 (7):111, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
111.pdf.

Anonymous:1900:APCb

[1861] Anonymous. The American Pacific Cable. Scientific American, 82(7):98, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
98.pdf.

Anonymous:1900:EPD

[1862] Anonymous. An extraordinary patent decision. Scientific American, 82(7):98, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
98a.pdf.

Anonymous:1900:OB

[1863] Anonymous. Our books in 1899. Scientific American, 82(7):98, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n7/pdf/scientificamerican02171900-98b.pdf.

Anonymous:1900:WS

[1864] Anonymous. Women in science. Scientific American, 82(7):98, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n7/pdf/scientificamerican02171900-98c.pdf.

Anonymous:1900:CDC

[1865] Anonymous. The Chicago Drainage Canal. Scientific American, 82(7):99, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
99.pdf.

Anonymous:1900:ECA

[1866] Anonymous. Employing the carabao for Army purposes in the Philippines. Scientific American, 82(7):99, February 17, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n7/pdf/scientificamerican02171900-99a.pdf.

Anonymous:1900:FIN

[1867] Anonymous. Factory inspection of New York State. Scientific American, 82(7):99, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
99b.pdf.

Anonymous:1900:NGP

[1868] Anonymous. The new Government Printing Office. Scientific American, 82(7):99, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
99c.pdf.

Anonymous:1900:NWC

[1869] Anonymous. A new way of cleaning ships’ bottoms. Scientific American, 82(7):100, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
100.pdf.

Anonymous:1900:NRG

[1870] Anonymous. A novel reversing gear. Scientific American, 82 (7):100, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
100a.pdf.

Damon:1900:MC

[1871] George A. Damon. Magnetic clutches. Scientific American, 82 (7):100–101, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
100b.pdf.

Anonymous:1900:FCCa

[1872] Anonymous. The four-cylinder compound locomotives of the Northern Railway of France. Scientific American, 82(7):101–104, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n7/pdf/scientificamerican02171900-101.pdf.

Anonymous:1900:AQB

[1873] Anonymous. Auto-quadricycle at the bicycle and auto mobile show. Scientific American, 82(7):104, February 17, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n7/pdf/scientificamerican02171900-104.pdf.

Stewart:1900:NIW

[1874] J. A. Stewart. The new inland waterway. Scientific American, 82 (7):105–106, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
105.pdf.

Blackford:1900:CI

[1875] Charles Minor Blackford, Jr. Curiosities of ichthyology. Scientific American, 82(7):107, February 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n7/pdf/scientificamerican02171900-
107.pdf.

Anonymous:1900:Cu

[1876] Anonymous. Correspondence. Scientific American, 82(8):124, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n8/pdf/scientificamerican02241900-124.pdf.

Anonymous:1900:RPIy

[1877] Anonymous. Recently patented inventions. Scientific American, 82 (8):124–125, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
124a.pdf.

Anonymous:1900:BNBv

[1878] Anonymous. Books: New books, etc. Scientific American, 82 (8):125, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
125.pdf.

Anonymous:1900:BPy

[1879] Anonymous. Business and personal. Scientific American, 82(8): 125, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
125a.pdf.

Anonymous:1900:NQy

[1880] Anonymous. Notes and queries. Scientific American, 82(8):125, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n8/pdf/scientificamerican02241900-125b.pdf.

Anonymous:1900:IIz

[1881] Anonymous. Index of inventions. Scientific American, 82(8): 125–127, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
125c.pdf.

Anonymous:1900:PNPy

[1882] Anonymous. Patent notes: Patents. Scientific American, 82 (8):127, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
127.pdf.

Anonymous:1900:RSC

[1883] Anonymous. Revival of the second-class mail matter question. Scientific American, 82(8):114, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
114.pdf.

Anonymous:1900:PSF

[1884] Anonymous. The problem of San Francisco Bay. Scientific American, 82(8):114, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
114a.pdf.

Anonymous:1900:PGa

[1885] Anonymous. The production of gold. Scientific American, 82 (8):114, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
114b.pdf.

Anonymous:1900:UWP

[1886] Anonymous. The utilization of wastes for power. Scientific American, 82(8):114, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
114c.pdf.

Anonymous:1900:GRP

[1887] Anonymous. A Great Russian pipe line. Scientific American, 82(8):114, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
114d.pdf.

Heilprin:1900:AIL

[1888] Angelo Heilprin. An assumed inconstancy in the level of Lake Nicaragua; a question of permanency of the Nicaragua Canal. Scientific American, 82(8):115, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
115.pdf.

Anonymous:1900:TFN

[1889] Anonymous. Two forms of a new dish-washing machine. Scientific American, 82(8):116, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
116.pdf.

Anonymous:1900:WRB

[1890] Anonymous. A world’s record in bridge building. Scientific American, 82(8):116, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
116a.pdf.

Anonymous:1900:MSF

[1891] Anonymous. A mechanical stoker for furnaces. Scientific American, 82(8):116, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
116b.pdf.

Anonymous:1900:PPA

[1892] Anonymous. The progress of practical aeronautics during 1899. Scientific American, 82(8):117, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
117.pdf.

Anonymous:1900:CPI

[1893] Anonymous. Casting pig iron by machinery. Scientific American, 82(8):118–119, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
118.pdf.

Willey:1900:RLB

[1894] D. Allen Willey. The round-lap bale method of shipping cotton. Scientific American, 82(8):120–122, February 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n8/pdf/scientificamerican02241900-120.pdf.

Anonymous:1900:TTB

[1895] Anonymous. A tool for transplanting buds. Scientific American, 82(8):122, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
122.pdf.

Anonymous:1900:IAR

[1896] Anonymous. An improved apparatus for raising liquids. Scientific American, 82(8):122, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
122a.pdf.

Anonymous:1900:XRM

[1897] Anonymous. The xiphopages, Rosalina and Maria. Scientific American, 82(8):123, February 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n8/pdf/scientificamerican02241900-
123.pdf.

Anonymous:1900:RPIz

[1898] Anonymous. Recently patented inventions. Scientific American, 82(9):140, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
140.pdf.

Anonymous:1900:BNBw

[1899] Anonymous. Books: New books, etc. Scientific American, 82(9):140, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-140a.pdf.

Anonymous:1900:BPz

[1900] Anonymous. Business and personal. Scientific American, 82(9):140, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-140b.pdf.

Anonymous:1900:NQz

[1901] Anonymous. Notes and queries. Scientific American, 82(9):140, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-140c.pdf.

Anonymous:1900:IIba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-140d.pdf.

Anonymous:1900:PNPz

[1903] Anonymous. Patent notes: Patents. Scientific American, 82(9):143, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-143.pdf.

Anonymous:1900:IAb

[1904] Anonymous. The ideal automobile. Scientific American, 82(9):130, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-130.pdf.

Anonymous:1900:PPR

[1905] Anonymous. Prospects of the Peary Relief Expedition. Scientific American, 82(9):130, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
130a.pdf.

Adams:1900:ETL

[1906] Altan D. Adams. Electric train lighting. Scientific American, 82(9):130, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-130b.pdf.

Anonymous:1900:APMa

[1907] Anonymous. Air pipes for mines, made of cloth. Scientific American, 82(9):130, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
130c.pdf.

Anonymous:1900:HTW

[1908] Anonymous. “hall i’ th’ wood,” Bolton. Scientific American, 82(9):130, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-130d.pdf.

Anonymous:1900:FPc

[1909] Anonymous. The father of photography. Scientific American, 82 (9):130–131, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
130e.pdf.

Hawkins:1900:TR

[1910] C. K. Hawkins. Turpentine and rosin. Scientific American, 82(9):131, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-131.pdf.

Anonymous:1900:ORG

[1911] Anonymous. Our rapidly growing irrigation areas. Scientific American, 82(9):131, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
131a.pdf.

Anonymous:1900:NSBb

[1912] Anonymous. New steamships building. Scientific American, 82(9):131, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-131b.pdf.

Anonymous:1900:ECPb

[1913] Anonymous. An efficient corn and potato hiller. Scientific American, 82(9):132, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
132.pdf.

Anonymous:1900:BME

[1914] Anonymous. A bicycle motor equipment. Scientific American, 82(9):132, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
132a.pdf.

Anonymous:1900:NNL

[1915] Anonymous. A new nut-lock. Scientific American, 82(9):132, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-132b.pdf.

Anonymous:1900:IWS

[1916] Anonymous. An improved wire-stretcher. Scientific American, 82 (9):133, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
133.pdf.

Crafts:1900:SWS

[1917] H. A. Crafts. The snowfall and water supply of the rocky mountains. Scientific American, 82(9):133–134, March 3, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v82/
n9/pdf/scientificamerican03031900-133a.pdf.

Anonymous:1900:BHR

[1918] Anonymous. Belgian hare raising in Southern California. Scientific American, 82(9):134–135, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
134.pdf.

Anonymous:1900:EPS

[1919] Anonymous. Electro-pneumatic switching and signaling system at the South Boston Terminal Station. Scientific American, 82(9):136–137, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v82/n9/pdf/scientificamerican03031900-136.pdf.

Anonymous:1900:NCF

[1920] Anonymous. Naval constructor Francis T. Bowles. Scientific American, 82(9):137–138, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
137.pdf.

Anonymous:1900:NYN

[1921] Anonymous. New York Navy Yard, Brooklyn. Scientific American, 82(9):138–139, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
138.pdf.

Anonymous:1900:BWP

[1922] Anonymous. Building wharf at Pango Pango. Scientific American, 82(9):139, March 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v82/n9/pdf/scientificamerican03031900-
139.pdf.

Anonymous:1900:RPIba

[1923] Anonymous. Recently patented inventions. Scientific American, 83(1): 12, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-12a.pdf.

Anonymous:1900:BPba

[1924] Anonymous. Business and personal. Scientific American, 83(1):12, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-12b.pdf.

Anonymous:1900:NQba

[1925] Anonymous. Notes and queries. Scientific American, 83(1):12, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-12c.pdf.

Anonymous:1900:IIbb

[1926] Anonymous. Index of inventions. Scientific American, 83(1):13, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-13.pdf.

Anonymous:1900:PNPba

[1927] Anonymous. Patent notes: Patents. Scientific American, 83(1):15, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-15.pdf.

Anonymous:1900:UC

[1928] Anonymous. An unfortunate compromise. Scientific American, 83(1): 4, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-4.pdf.

Anonymous:1900:GDN

[1929] Anonymous. Garbage disposal of New York City. Scientific American, 83(1):4, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n1/pdf/scientificamerican07071900-
4a.pdf.

Anonymous:1900:RBTb

[1930] Anonymous. Relation between transverse and crushing strength of timber. Scientific American, 83(1):4, July 7, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n1/pdf/scientificamerican07071900-4b.pdf.

Anonymous:1900:WTS

[1931] Anonymous. Wireless telegraphy in South Africa. Scientific American, 83(1):4, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n1/pdf/scientificamerican07071900-
4c.pdf.

Anonymous:1900:RDC

[1932] Anonymous. Recent discoveries in Corinth. Scientific American, 83(1): 4, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-4d.pdf.

Anonymous:1900:QT

[1933] Anonymous. Quartz thermometers. Scientific American, 83(1):4–5, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-4e.pdf.

Anonymous:1900:MGD

[1934] Anonymous. Muir Glacier destroyed. Scientific American, 83(1):5, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-5.pdf.

Anonymous:1900:AAA

[1935] Anonymous. American Association for the Advancement of Science. Scientific American, 83(1):5, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n1/pdf/scientificamerican07071900-
5a.pdf.

Anonymous:1900:PENg

[1936] Anonymous. Paris Exposition notes. Scientific American, 83(1):5, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-5b.pdf.

Fawcett:1900:DM

[1937] Waldon Fawcett. Dredges on the Mississippi. Scientific American, 83(1): 6, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-6.pdf.

Mann:1900:NAS

[1938] Francis P. Mann. The new air ship of M. De Santos Dumont. Scientific American, 83(1):7, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n1/pdf/scientificamerican07071900-
7.pdf.

Howard:1900:DBM

[1939] L. O. Howard. The differences between malarial and non-malarial mosquitoes. Scientific American, 83(1):8–9, July 7, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n1/pdf/scientificamerican07071900-8.pdf.

Anonymous:1900:WCC

[1940] Anonymous. The wheat crop of California. Scientific American, 83(1): 9, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-9.pdf.

Colvin:1900:GMG

[1941] Wilber Colvin. Gold mining in Georgia. Scientific American, 83(1):10– 11, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n1/pdf/scientificamerican07071900-10.pdf.

Anonymous:1900:ERD

[1942] Anonymous. The Elevated Railway Drawbridge, Boston. Scientific American, 83(1):11–12, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n1/pdf/scientificamerican07071900-
11.pdf.

Anonymous:1900:ADD

[1943] Anonymous. An air-directing device for electric fans. Scientific American, 83(1):12, July 7, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n1/pdf/scientificamerican07071900-
12.pdf.

Anonymous:1900:Cv

[1944] Anonymous. Correspondence. Scientific American, 83(10):150– 151, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
150.pdf.

Anonymous:1900:RPIbb

[1945] Anonymous. Recently patented inventions. Scientific American, 83(10):156, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
156.pdf.

Anonymous:1900:BNBx

[1946] Anonymous. Books: New books, etc. Scientific American, 83 (10):157, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
157.pdf.

Anonymous:1900:BPbb

[1947] Anonymous. Business and personal. Scientific American, 83 (10):157, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
157a.pdf.

Anonymous:1900:NQbb

[1948] Anonymous. Notes and queries. Scientific American, 83(10): 157, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
157b.pdf.

Anonymous:1900:IIbc

[1949] Anonymous. Index of inventions. Scientific American, 83(10): 157, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
157d.pdf.

Anonymous:1900:PNPbb

[1950] Anonymous. Patent notes: Patents. Scientific American, 83 (10):159, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
159.pdf.

Anonymous:1900:CTa

[1951] Anonymous. From cable to trolley. Scientific American, 83(10): 146, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
146.pdf.

Anonymous:1900:ABE

[1952] Anonymous. American and British Engineering Competition. Scientific American, 83(10):146, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
146a.pdf.

Anonymous:1900:CIM

[1953] Anonymous. Crystalline iodide of mercury. Scientific American, 83(10):146, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
146b.pdf.

Anonymous:1900:EMM

[1954] Anonymous. Experiments of M. Moissan. Scientific American, 83 (10):146–147, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
146c.pdf.

Anonymous:1900:PEE

[1955] Anonymous. Paris Exposition — Electrical Congress. Scientific American, 83(10):147, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
147.pdf.

Anonymous:1900:MLD

[1956] Anonymous. Marconi’s latest developments — synchronized messages. Scientific American, 83(10):147, September 8, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n10/pdf/scientificamerican09081900-147a.pdf.

Anonymous:1900:ENL

[1957] Anonymous. Explosion of a naphtha launch. Scientific American, 83(10):147, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
147b.pdf.

Anonymous:1900:AWU

[1958] Anonymous. Accumulators working under water. Scientific American, 83(10):147, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
147c.pdf.

Anonymous:1900:STCb

[1959] Anonymous. Steam tipping-cart. Scientific American, 83(10): 148, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
148.pdf.

Anonymous:1900:IS

[1960] Anonymous. An improved scraper. Scientific American, 83(10): 148, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
148a.pdf.

Fawcett:1900:II

[1961] Waldon Fawcett. Irrigation in Idaho. Scientific American, 83 (10):149, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
149.pdf.

Anonymous:1900:PEU

[1962] Anonymous. Paris Exposition — United States Mining and Metallurgical Exhibit. Scientific American, 83(10):152–153, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n10/pdf/scientificamerican09081900-152.pdf.

Anonymous:1900:MHT

[1963] Anonymous. The making of a half-tone engraving. Scientific American, 83(10):153–154, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
153.pdf.

Anonymous:1900:ECRa

[1964] Anonymous. Experiments on the cathode rays. Scientific American, 83(10):154, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
154.pdf.

Anonymous:1900:LFA

[1965] Anonymous. The last of a famous “America” Cup challenger. Scientific American, 83(10):155, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
155.pdf.

Anonymous:1900:FCBa

[1966] Anonymous. The first-class battleship “Alabama”. Scientific American, 83(10):155–156, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n10/pdf/scientificamerican09081900-
155a.pdf.

Anonymous:1900:Il

[1967] Anonymous. To inventors. Scientific American, 83(10):157, September 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n10/pdf/scientificamerican09081900-157c.pdf.

Anonymous:1900:Cw

[1968] Anonymous. Correspondence. Scientific American, 83(11):167, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n11/pdf/scientificamerican09151900-167.pdf.

Anonymous:1900:RPIbc

[1969] Anonymous. Recently patented inventions. Scientific American, 83 (11):172–173, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
172.pdf.

Anonymous:1900:BPbc

[1970] Anonymous. Business and personal. Scientific American, 83 (11):173, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
173.pdf.

Anonymous:1900:NQbc

[1971] Anonymous. Notes and queries. Scientific American, 83(11): 173, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
173a.pdf.

Anonymous:1900:IIbd

[1972] Anonymous. Index of inventions. Scientific American, 83(11): 173–175, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
173b.pdf.

Anonymous:1900:PNPbc

[1973] Anonymous. Patent notes: Patents. Scientific American, 83 (11):175, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
175.pdf.

Anonymous:1900:AON

[1974] Anonymous. The armament of our new battleships and cruisers. Scientific American, 83(11):162, September 15, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n11/pdf/scientificamerican09151900-162.pdf.

Anonymous:1900:PAP

[1975] Anonymous. Proposed abandonment of Port Royal Naval Station. Scientific American, 83(11):162, September 15, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n11/pdf/scientificamerican09151900-162a.pdf.

Anonymous:1900:CFR

[1976] Anonymous. Curious facts regarding mosquitoes. Scientific American, 83(11):162–163, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
162b.pdf.

Anonymous:1900:AER

[1977] Anonymous. Another electric railway for London. Scientific American, 83(11):163, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
163.pdf.

Anonymous:1900:MIE

[1978] Anonymous. Malarial infection on the East Coast of Africa. Scientific American, 83(11):163, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
163a.pdf.

Anonymous:1900:DBT

[1979] Anonymous. The “Deutschland” breaks two records. Scientific American, 83(11):163, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
163b.pdf.

Anonymous:1900:GAA

[1980] Anonymous. The Great Ararat ascended. Scientific American, 83 (11):163, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
163c.pdf.

Anonymous:1900:DAE

[1981] Anonymous. The Duke of Abruzzi’s Expedition reaches the highest altitude. Scientific American, 83(11):163, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n11/pdf/scientificamerican09151900-163d.pdf.

Anonymous:1900:SNa

[1982] Anonymous. Science notes. Scientific American, 83(11):163, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n11/pdf/scientificamerican09151900-163e.pdf.

Anonymous:1900:AAG

[1983] Anonymous. An automatic acetylene-generator. Scientific American, 83(11):164, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
164.pdf.

Anonymous:1900:PEM

[1984] Anonymous. Paris Exposition — models in Hungarian section, illustrating engineering work on Lower Danube. Scientific American, 83 (11):164–165, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
164a.pdf.

Anonymous:1900:PBW

[1985] Anonymous. Pilot boat wrecked by a whale. Scientific American, 83(11):165, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
165.pdf.

Anonymous:1900:SIL

[1986] Anonymous. A simple indicator for locomobile water-tanks. Scientific American, 83(11):165, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
165a.pdf.

Anonymous:1900:GBCb

[1987] Anonymous. The greatest business concern in the world. Scientific American, 83(11):166, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
166.pdf.

Anonymous:1900:PEL

[1988] Anonymous. Paris Exposition — large engine. Scientific American, 83(11):168, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
168.pdf.

Holder:1900:SLL

[1989] Charles F. Holder. Some living lamps. Scientific American, 83 (11):168–169, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
168a.pdf.

Mvey:1900:LIG

[1990] G. H. P. M’vey. Laying a 24-inch gas main across the Harlem River. Scientific American, 83(11):169–170, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n11/pdf/scientificamerican09151900-169.pdf.

Michaud:1900:CON

[1991] Gustave Michaud. The climate of our new possessions. Scientific American, 83(11):171–172, September 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n11/pdf/scientificamerican09151900-
171.pdf.

Anonymous:1900:Cx

[1992] Anonymous. Correspondence. Scientific American, 83(12):182, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n12/pdf/scientificamerican09221900-182.pdf.

Anonymous:1900:RPIbd

[1993] Anonymous. Recently patented inventions. Scientific American, 83 (12):189, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
189.pdf.

Anonymous:1900:BPbd

[1994] Anonymous. Business and personal. Scientific American, 83 (12):189, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
189a.pdf.

Anonymous:1900:NQbd

[1995] Anonymous. Notes and queries. Scientific American, 83(12): 189, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
189b.pdf.

Anonymous:1900:BNBy

[1996] Anonymous. Books: New books, etc. Scientific American, 83 (12):189, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
189c.pdf.

Anonymous:1900:IIbe

[1997] Anonymous. Index of inventions. Scientific American, 83(12): 190, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
190.pdf.

Anonymous:1900:PNPbd

[1998] Anonymous. Patent notes: Patents. Scientific American, 83 (12):191, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
191.pdf.

Anonymous:1900:AAE

[1999] Anonymous. Awards to American exhibitors at Paris. Scientific American, 83(12):178, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
178.pdf.

Anonymous:1900:PTa

[2000] Anonymous. Poulsen’s telegraphone. Scientific American, 83(12): 178, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
178a.pdf.

Anonymous:1900:MSU

[2001] Anonymous. Mechanical stoking in the United States. Scientific American, 83(12):178, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
178b.pdf.

Anonymous:1900:EVGb

[2002] Anonymous. English vs. German Atlantic liners. Scientific American, 83(12):179, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
179.pdf.

Anonymous:1900:TLN

[2003] Anonymous. The Temple Library at Nippur. Scientific American, 83(12):179, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
179a.pdf.

Adams:1900:DWP

[2004] Altan D. Adams. Distant water powers. Scientific American, 83 (12):179, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
179b.pdf.

Anonymous:1900:EAC

[2005] Anonymous. Exports of American coal. Scientific American, 83 (12):179, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
179c.pdf.

Anonymous:1900:SGL

[2006] Anonymous. The spiral-globe lamp. Scientific American, 83 (12):180, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
180.pdf.

Anonymous:1900:AOW

[2007] Anonymous. An automatic oval wood dish machine. Scientific American, 83(12):180, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
180a.pdf.

Anonymous:1900:PTb

[2008] Anonymous. Poulsen telegraphone. Scientific American, 83(12): 181, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
181.pdf.

Fawcett:1900:NEA

[2009] Waldon Fawcett. The new era of the American sailing vessel. Scientific American, 83(12):182–183, September 22, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n12/pdf/scientificamerican09221900-182a.pdf.

Anonymous:1900:PEAa

[2010] Anonymous. Paris Exposition — Austrian Palace. Scientific American, 83(12):184, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
184.pdf.

Anonymous:1900:MASb

[2011] Anonymous. Models in Austrian section of the Paris Exposition. Scientific American, 83(12):185, September 22, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n12/pdf/scientificamerican09221900-185.pdf.

Anonymous:1900:CICa

[2012] Anonymous. Coney Island’s centrifugal railway. Scientific American, 83(12):185–186, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
185a.pdf.

Anonymous:1900:CST

[2013] Anonymous. Color screen for telescopes. Scientific American, 83 (12):186, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
186.pdf.

Anonymous:1900:RNR

[2014] Anonymous. Rebuilding Niagara’s reservation bridges. Scientific American, 83(12):187–188, September 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n12/pdf/scientificamerican09221900-
187.pdf.

Russell:1900:HO

[2015] Henry Norris Russell. The heavens in October. Scientific American, 83(13):195, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
195.pdf.

Anonymous:1900:RPIbe

[2016] Anonymous. Recently patented inventions. Scientific American, 83 (13):204, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
204.pdf.

Anonymous:1900:BPbe

[2017] Anonymous. Business and personal. Scientific American, 83 (13):205, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
205.pdf.

Anonymous:1900:NQbe

[2018] Anonymous. Notes and queries. Scientific American, 83(13): 205, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
205a.pdf.

Anonymous:1900:IIbf

[2019] Anonymous. Index of inventions. Scientific American, 83(13): 205, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
205b.pdf.

Anonymous:1900:PNPbe

[2020] Anonymous. Patent notes: Patents. Scientific American, 83 (13):207, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
207.pdf.

Anonymous:1900:AIK

[2021] Anonymous. Another “invention” — a 70 knot submarine boat? Scientific American, 83(13):194, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
194.pdf.

Anonymous:1900:GD

[2022] Anonymous. The Galveston Disaster. Scientific American, 83 (13):194, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
194a.pdf.

Anonymous:1900:NTD

[2023] Anonymous. New torpedo destroyers for Holland. Scientific American, 83(13):194, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
194b.pdf.

Anonymous:1900:HLC

[2024] Anonymous. Hand-labor in Chinese mines. Scientific American, 83(13):194, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
194c.pdf.

Anonymous:1900:TEP

[2025] Anonymous. Traffic in European ports. Scientific American, 83 (13):194–195, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
194d.pdf.

Anonymous:1900:BAA

[2026] Anonymous. The British Association for the Advancement of Science. Scientific American, 83(13):195, September 29, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n13/pdf/scientificamerican09291900-195a.pdf.

Anonymous:1900:IPC

[2027] Anonymous. The International Physical Congress. Scientific American, 83(13):195, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
195b.pdf.

Anonymous:1900:CAS

[2028] Anonymous. The Chicago Automobile Show. Scientific American, 83(13):195, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
195c.pdf.

Anonymous:1900:FEE

[2029] Anonymous. The “Fortis” electric exerciser. Scientific American, 83(13):196, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
196.pdf.

Anonymous:1900:PTS

[2030] Anonymous. A phosphate transporter at Sfax. Scientific American, 83(13):196–197, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
196a.pdf.

Anonymous:1900:DO

[2031] Anonymous. The damage to the “Oregon”. Scientific American, 83(13):197, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
197.pdf.

Anonymous:1900:TNHb

[2032] Anonymous. Two novel harrows. Scientific American, 83(13): 197, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
197a.pdf.

Anonymous:1900:MPE

[2033] Anonymous. The Mareorama at the Paris Exposition. Scientific American, 83(13):198, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
198.pdf.

Anonymous:1900:JTM

[2034] Anonymous. A Japanese trade mark. Scientific American, 83(13): 198–200, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
198a.pdf.

Kime:1900:PQL

[2035] J. W. Kime. Penetrative quality of light as tested by photography. Scientific American, 83(13):200, September 29, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n13/pdf/scientificamerican09291900-200.pdf.

Anonymous:1900:NAI

[2036] Anonymous. The new Army 16-inch gun. Scientific American, 83 (13):201–202, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
201.pdf.

Anonymous:1900:ANS

[2037] Anonymous. Aborigines of the Nilgiris in South India. Scientific American, 83(13):203–204, September 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n13/pdf/scientificamerican09291900-
203.pdf.

Anonymous:1900:RPIbf

[2038] Anonymous. Recently patented inventions. Scientific American, 83(14):220, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
220.pdf.

Anonymous:1900:BNBz

[2039] Anonymous. Books: New books, etc. Scientific American, 83(14):220, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-220a.pdf.

Anonymous:1900:BPbf

[2040] Anonymous. Business and personal. Scientific American, 83(14):220, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-220b.pdf.

Anonymous:1900:NQbf

[2041] Anonymous. Notes and queries. Scientific American, 83(14):220, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-220c.pdf.

Anonymous:1900:IIbg

[2042] Anonymous. Index of inventions. Scientific American, 83(14):220, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-220d.pdf.

Anonymous:1900:PNPbf

[2043] Anonymous. Patent notes: Patents. Scientific American, 83(14):223, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-223.pdf.

Anonymous:1900:RMA

[2044] Anonymous. The report of the Merchants’ Association on the New York Water Supply. Scientific American, 83(14):210, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-210.pdf.

Anonymous:1900:GAC

[2045] Anonymous. Galvanic action of copper sheathing. Scientific American, 83(14):210, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
210a.pdf.

Anonymous:1900:SR

[2046] Anonymous. Spectrum of radium. Scientific American, 83(14):210, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-210b.pdf.

Anonymous:1900:PO

[2047] Anonymous. Pallado-oxalates. Scientific American, 83(14):210, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-210c.pdf.

Anonymous:1900:MO

[2048] Anonymous. The mastery of the ocean. Scientific American, 83(14):211, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-211.pdf.

Anonymous:1900:EA

[2049] Anonymous. Experiments on amalgams. Scientific American, 83 (14):211, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
211a.pdf.

Anonymous:1900:OWB

[2050] Anonymous. Operation of Welsbach burners. Scientific American, 83(14):211, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
211b.pdf.

Anonymous:1900:CBG

[2051] Anonymous. A cactus in the Botanical Garden at Berlin. Scientific American, 83(14):211, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
211c.pdf.

Anonymous:1900:PEAb

[2052] Anonymous. Paris Exposition Awards. Scientific American, 83(14):211, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-211d.pdf.

Anonymous:1900:NGS

[2053] Anonymous. A novel grooving-saw and dado-head. Scientific American, 83(14):212, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
212.pdf.

Anonymous:1900:LMR

[2054] Anonymous. The Langen Mono-Rail Road at the Paris Exposition. Scientific American, 83(14):212, October 6, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n14/pdf/scientificamerican10061900-212a.pdf.

Anonymous:1900:CPCb

[2055] Anonymous. A cotton per cent scale. Scientific American, 83(14):212, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n14/pdf/scientificamerican10061900-212b.pdf.

Anonymous:1900:UFB

[2056] Anonymous. The use of flexible bridles on kites. Scientific American, 83(14):213, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
213.pdf.

Anonymous:1900:SME

[2057] Anonymous. Some metallurgical exhibits at the Paris Exposition. Scientific American, 83(14):213, October 6, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n14/pdf/scientificamerican10061900-213a.pdf.

Anonymous:1900:MIG

[2058] Anonymous. Manufacture of illuminating gas. Scientific American, 83(14):214–215, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
214.pdf.

Anonymous:1900:AAP

[2059] Anonymous. Aleuts of the Alaska Peninsula. Scientific American, 83(14):216, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
216.pdf.

Anonymous:1900:UPP

[2060] Anonymous. A unique pigeon postal service. Scientific American, 83(14):217, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
217.pdf.

Anonymous:1900:LMD

[2061] Anonymous. The Langston mooring device. Scientific American, 83(14):217–218, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
217a.pdf.

Kellogg:1900:CAE

[2062] Arthur T. Kellogg. Chicago Automobile Exhibition and race meet. Scientific American, 83(14):218–219, October 6, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n14/pdf/scientificamerican10061900-218.pdf.

Anonymous:1900:TSF

[2063] Anonymous. The Tri-State Fair Automobile Races. Scientific American, 83(14):219, October 6, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n14/pdf/scientificamerican10061900-
219.pdf.

Anonymous:1900:RPIbg

[2064] Anonymous. Recently patented inventions. Scientific American, 83(15):236, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
236.pdf.

Anonymous:1900:BPbg

[2065] Anonymous. Business and personal. Scientific American, 83 (15):237, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
237.pdf.

Anonymous:1900:NQbg

[2066] Anonymous. Notes and queries. Scientific American, 83(15):237, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n15/pdf/scientificamerican10131900-237a.pdf.

Anonymous:1900:BNBba

[2067] Anonymous. Books: New books, etc. Scientific American, 83 (15):237, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
237b.pdf.

Anonymous:1900:IIbh

[2068] Anonymous. Index of inventions. Scientific American, 83(15):237, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n15/pdf/scientificamerican10131900-237c.pdf.

Anonymous:1900:PNPbg

[2069] Anonymous. Patent notes: Patents. Scientific American, 83 (15):239, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
239.pdf.

Anonymous:1900:ERT

[2070] Anonymous. Extension of the Rapid Transit Tunnel to Brooklyn. Scientific American, 83(15):226, October 13, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n15/pdf/scientificamerican10131900-226.pdf.

Anonymous:1900:RS

[2071] Anonymous. Railroad speeds in 1899. Scientific American, 83 (15):226, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
226a.pdf.

Anonymous:1900:WNP

[2072] Anonymous. The world’s Naval progress for 1900. Scientific American, 83(15):226, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
226b.pdf.

Adams:1900:CBF

[2073] Alton D. Adams. Culm banks and fuel gas. Scientific American, 83(15):227, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
227.pdf.

Anonymous:1900:NPS

[2074] Anonymous. The Nobel Prizes for scientific discoveries. Scientific American, 83(15):227, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
227a.pdf.

Perkins:1900:HPE

[2075] Frank C. Perkins. High power engines and three-phase generators at Paris. Scientific American, 83(15):228, October 13, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n15/pdf/scientificamerican10131900-228.pdf.

Maginnis:1900:PAH

[2076] Owen B. Maginnis. The practical application of the hydrostatic level in building construction. Scientific American, 83(15):229, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n15/pdf/scientificamerican10131900-229.pdf.

Anonymous:1900:FMO

[2077] Anonymous. The French Meteorological Observatory at Trappes. Scientific American, 83(15):229–230, October 13, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n15/pdf/scientificamerican10131900-229a.pdf.

Anonymous:1900:CGM

[2078] Anonymous. A cable-grip for mining-cars. Scientific American, 83(15):230, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
230.pdf.

Anonymous:1900:CEM

[2079] Anonymous. A convenient envelop moistener and sealer. Scientific American, 83(15):230, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
230a.pdf.

Anonymous:1900:EPF

[2080] Anonymous. Electrical palace and fountain at the Paris Exposition. Scientific American, 83(15):231, October 13, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n15/pdf/scientificamerican10131900-231.pdf.

Anonymous:1900:LMW

[2081] Anonymous. Launch of the monitor “Wyoming”. Scientific American, 83(15):232, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
232.pdf.

Anonymous:1900:IWF

[2082] Anonymous. An ingeniously whittled fan. Scientific American, 83(15):233, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
233.pdf.

Anonymous:1900:EDV

[2083] Anonymous. Electrical dredge for the Volga River, Russia. Scientific American, 83(15):233–234, October 13, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n15/pdf/scientificamerican10131900-
233a.pdf.

Dailey:1900:ASD

[2084] Howard B. Dailey. Acceleration of the spark discharge in static machines. Scientific American, 83(15):234, October 13, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n15/pdf/scientificamerican10131900-234.pdf.

Anonymous:1900:BCE

[2085] Anonymous. “Broadway Chambers” — exhibits at the Paris Exposition. Scientific American, 83(15):235–236, October 13, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n15/pdf/scientificamerican10131900-235.pdf.

Anonymous:1900:Cy

[2086] Anonymous. Correspondence. Scientific American, 83(16):246–247, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n16/pdf/scientificamerican10201900-246a.pdf.

Anonymous:1900:RPIbh

[2087] Anonymous. Recently patented inventions. Scientific American, 83(16):252, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
252.pdf.

Anonymous:1900:BNBbb

[2088] Anonymous. Books: New books, etc. Scientific American, 83 (16):252, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
252a.pdf.

Anonymous:1900:BPbh

[2089] Anonymous. Business and personal. Scientific American, 83 (16):253, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
253.pdf.

Anonymous:1900:NQbh

[2090] Anonymous. Notes and queries. Scientific American, 83(16):253, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n16/pdf/scientificamerican10201900-253a.pdf.

Anonymous:1900:IIbi

[2091] Anonymous. Index of inventions. Scientific American, 83(16):253, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n16/pdf/scientificamerican10201900-253b.pdf.

Anonymous:1900:PNPbh

[2092] Anonymous. Patent notes: Patents. Scientific American, 83 (16):255, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
255.pdf.

Anonymous:1900:MRH

[2093] Anonymous. The mono-rail for high-speed travel. Scientific American, 83(16):242, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
242.pdf.

Anonymous:1900:WTB

[2094] Anonymous. The water-tube boiler on trial. Scientific American, 83(16):242, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
242a.pdf.

Anonymous:1900:WV

[2095] Anonymous. The “Wisconsin” and the “Variag”. Scientific American, 83(16):242, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
242b.pdf.

Anonymous:1900:LDW

[2096] Anonymous. The latest developments of wireless telegraphy. Scientific American, 83(16):242–243, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
242c.pdf.

Anonymous:1900:CCCa

[2097] Anonymous. Calcium carbide at the Congress of Applied Chemistry. Scientific American, 83(16):243, October 20, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n16/pdf/scientificamerican10201900-243.pdf.

Anonymous:1900:CACb

[2098] Anonymous. The Congress of Applied Chemistry at Paris. Scientific American, 83(16):243, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
243a.pdf.

Anonymous:1900:BE

[2099] Anonymous. Borax in Europe. Scientific American, 83(16):243, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n16/pdf/scientificamerican10201900-243b.pdf.

Anonymous:1900:NMF

[2100] Anonymous. A new military folding bicycle. Scientific American, 83(16):244, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
244.pdf.

Anonymous:1900:CLD

[2101] Anonymous. A collapsible lathe-dog. Scientific American, 83 (16):244, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
244a.pdf.

Anonymous:1900:PEB

[2102] Anonymous. Paris Exposition — building of the German Marine. Scientific American, 83(16):245, October 20, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n16/pdf/scientificamerican10201900-245.pdf.

Anonymous:1900:RCD

[2103] Anonymous. Reconstruction in concrete of Dry Dock No. 2, New York Navy Yard. Scientific American, 83(16):246, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n16/pdf/scientificamerican10201900-246.pdf.

Anonymous:1900:MSb

[2104] Anonymous. A mechanical scarifier. Scientific American, 83 (16):248, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
248.pdf.

Anonymous:1900:IRI

[2105] Anonymous. The 16-inch rifle and 20-inch smooth-bore compared. Scientific American, 83(16):248, October 20, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n16/pdf/scientificamerican10201900-248a.pdf.

Anonymous:1900:SCB

[2106] Anonymous. Strengthening the cantilever bridge at Niagara. Scientific American, 83(16):249–250, October 20, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n16/pdf/scientificamerican10201900-249.pdf.

Aaron:1900:PM

[2107] S. Frank Aaron. The pantry moth. Scientific American, 83(16):250, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n16/pdf/scientificamerican10201900-250.pdf.

Thomas:1900:RIW

[2108] R. E. Thomas. Re-invention in the world of surgery. Scientific American, 83(16):251, October 20, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n16/pdf/scientificamerican10201900-
251.pdf.

Anonymous:1900:Cz

[2109] Anonymous. Correspondence. Scientific American, 83(17):266, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n17/pdf/scientificamerican10271900-266.pdf.

Anonymous:1900:RPIbi

[2110] Anonymous. Recently patented inventions. Scientific American, 83(17):268, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
268.pdf.

Anonymous:1900:BPbi

[2111] Anonymous. Business and personal. Scientific American, 83 (17):268, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
268a.pdf.

Anonymous:1900:NQbi

[2112] Anonymous. Notes and queries. Scientific American, 83(17):268, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n17/pdf/scientificamerican10271900-268b.pdf.

Anonymous:1900:BNBbc

[2113] Anonymous. Books: New books, etc. Scientific American, 83 (17):268–269, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
268c.pdf.

Anonymous:1900:IIbj

[2114] Anonymous. Index of inventions. Scientific American, 83(17):269, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n17/pdf/scientificamerican10271900-269a.pdf.

Anonymous:1900:PNPbi

[2115] Anonymous. Patent notes: Patents. Scientific American, 83 (17):271, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
271.pdf.

Anonymous:1900:APMb

[2116] Anonymous. To the armor-plate makers. Scientific American, 83(17):258, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
258.pdf.

Anonymous:1900:ACC

[2117] Anonymous. An America Cup Contest in 1901. Scientific American, 83(17):258, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
258a.pdf.

Anonymous:1900:ZAT

[2118] Anonymous. Zeppelin’s airship on trial. Scientific American, 83 (17):258, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
258b.pdf.

Anonymous:1900:NLL

[2119] Anonymous. The new liner “La Lorraine”. Scientific American, 83(17):258, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
258c.pdf.

Anonymous:1900:LFG

[2120] Anonymous. The local forecaster on the Galveston Hurricane. Scientific American, 83(17):258, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
258d.pdf.

Anonymous:1900:PIN

[2121] Anonymous. “The progress of invention in the Nineteenth Century”. Scientific American, 83(17):259, October 27, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n17/pdf/scientificamerican10271900-259.pdf.

Anonymous:1900:RRA

[2122] Anonymous. Report of Rear Admiral Farquhar on the Newport Naval Maneuvers. Scientific American, 83(17):259, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n17/pdf/scientificamerican10271900-259a.pdf.

Anonymous:1900:ICb

[2123] Anonymous. Indigo in Cambodia. Scientific American, 83(17):259, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n17/pdf/scientificamerican10271900-259b.pdf.

Anonymous:1900:SGE

[2124] Anonymous. The school gardens of Europe. Scientific American, 83(17):259–260, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
259c.pdf.

Anonymous:1900:VPE

[2125] Anonymous. “Vieux Paris” at the Exposition. Scientific American, 83(17):260, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
260.pdf.

Anonymous:1900:ICWa

[2126] Anonymous. An improved clothes-washing machine. Scientific American, 83(17):260, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
260a.pdf.

Headland:1900:UC

[2127] Isaac Taylor Headland. The unscientific Chinese. Scientific American, 83(17):261, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
261.pdf.

Anonymous:1900:GGP

[2128] Anonymous. Gas-generating plant at the Paris Exposition. Scientific American, 83(17):261–262, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
261a.pdf.

Anonymous:1900:RKV

[2129] Anonymous. The reconstruction of the Kinzua Viaduct. Scientific American, 83(17):262, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
262.pdf.

Anonymous:1900:PMD

[2130] Anonymous. Photographic method of determining speed of automobiles. Scientific American, 83(17):262–263, October 27, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n17/pdf/scientificamerican10271900-262a.pdf.

Anonymous:1900:EFA

[2131] Anonymous. The electric fire automobiles of the Paris Fire Department. Scientific American, 83(17):264, October 27, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n17/pdf/scientificamerican10271900-264.pdf.

Stark:1900:RRO

[2132] George Stark. Railbreaker for removing old tracks. Scientific American, 83(17):265, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n17/pdf/scientificamerican10271900-
265.pdf.

Howard:1900:TIU

[2133] L. O. Howard. Two interesting uses of insects by natives in Natal. Scientific American, 83(17):267, October 27, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n17/pdf/scientificamerican10271900-267.pdf.

Anonymous:1900:Im

[2134] Anonymous. To inventors. Scientific American, 83(17):269, October 27, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n17/pdf/scientificamerican10271900-269.pdf.

Russell:1900:HN

[2135] Henry Norris Russell. The heavens in November. Scientific American, 83(18):275, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
275.pdf.

Anonymous:1900:Cba

[2136] Anonymous. Correspondence. Scientific American, 83(18):282, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n18/pdf/scientificamerican11031900-282.pdf.

Anonymous:1900:RPIbj

[2137] Anonymous. Recently patented inventions. Scientific American, 83(18):284, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
284.pdf.

Anonymous:1900:BPbj

[2138] Anonymous. Business and personal. Scientific American, 83 (18):284, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
284a.pdf.

Anonymous:1900:NQbj

[2139] Anonymous. Notes and queries. Scientific American, 83(18): 284, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
284b.pdf.

Anonymous:1900:BNBbd

[2140] Anonymous. Books: New books, etc. Scientific American, 83 (18):284, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
284c.pdf.

Anonymous:1900:IIbk

[2141] Anonymous. Index of inventions. Scientific American, 83(18): 284–286, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
284e.pdf.

Anonymous:1900:PNPbj

[2142] Anonymous. Patent notes: Patents. Scientific American, 83 (18):287, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
287.pdf.

Anonymous:1900:IES

[2143] Anonymous. Irrigation in the Eastern States. Scientific American, 83(18):274, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
274.pdf.

Anonymous:1900:KSY

[2144] Anonymous. A 40-knot steam yacht. Scientific American, 83 (18):274, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
274a.pdf.

Anonymous:1900:NMB

[2145] Anonymous. The National Memorial Bridge at Washington. Scientific American, 83(18):274, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
274b.pdf.

Anonymous:1900:ORC

[2146] Anonymous. Our railroads at the close of the century. Scientific American, 83(18):274, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
274c.pdf.

Anonymous:1900:SWF

[2147] Anonymous. Soda water fountain in Great Britain. Scientific American, 83(18):274–275, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
274d.pdf.

Anonymous:1900:CTPa

[2148] Anonymous. Congress of Tramways at Paris. Scientific American, 83(18):275, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
275a.pdf.

Anonymous:1900:USN

[2149] Anonymous. The United States National Museum. Scientific American, 83(18):275, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
275b.pdf.

Anonymous:1900:HRC

[2150] Anonymous. How Russia corners sugar. Scientific American, 83 (18):275, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
275c.pdf.

Anonymous:1900:NOT

[2151] Anonymous. A new operating table for horses. Scientific American, 83(18):276, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
276.pdf.

Anonymous:1900:REN

[2152] Anonymous. A reversible engine of novel form. Scientific American, 83(18):277, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
277.pdf.

Anonymous:1900:ELP

[2153] Anonymous. Electric locomotive at the Paris Exhibition. Scientific American, 83(18):277, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
277a.pdf.

Anonymous:1900:SSAa

[2154] Anonymous. A simple sulky attachment for harrows. Scientific American, 83(18):277, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
277b.pdf.

Anonymous:1900:FCCb

[2155] Anonymous. The French cruisers “Cecille” and “Suchet” at New York. Scientific American, 83(18):278–279, November 3, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n18/pdf/scientificamerican11031900-278.pdf.

Anonymous:1900:REE

[2156] Anonymous. The Rieder electro-engraving apparatus. Scientific American, 83(18):280, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
280.pdf.

Anonymous:1900:CTF

[2157] Anonymous. The construction of a 150-ton flywheel. Scientific American, 83(18):280, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
280a.pdf.

Anonymous:1900:RRI

[2158] Anonymous. A rack railway on the Island of Sumatra. Scientific American, 83(18):281, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
281.pdf.

Anonymous:1900:AEP

[2159] Anonymous. The aerostatic exhibits at Paris. Scientific American, 83(18):282–283, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n18/pdf/scientificamerican11031900-
282a.pdf.

Anonymous:1900:In

[2160] Anonymous. To inventors. Scientific American, 83(18):284, November 3, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n18/pdf/scientificamerican11031900-284d.pdf.

Anonymous:1900:RPIbk

[2161] Anonymous. Recently patented inventions. Scientific American, 83(19):300, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
300.pdf.

Anonymous:1900:BPbk

[2162] Anonymous. Business and personal. Scientific American, 83 (19):300, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
300a.pdf.

Anonymous:1900:NQbk

[2163] Anonymous. Notes and queries. Scientific American, 83(19): 300, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
300b.pdf.

Anonymous:1900:BNBbe

[2164] Anonymous. Books: New books, etc. Scientific American, 83 (19):300, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
300c.pdf.

Anonymous:1900:IIbl

[2165] Anonymous. Index of inventions. Scientific American, 83(19): 300–303, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
300d.pdf.

Anonymous:1900:PNPbk

[2166] Anonymous. Patent notes: Patents. Scientific American, 83 (19):302, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
302.pdf.

Anonymous:1900:ASPb

[2167] Anonymous. Atlantic steamships — present and future. Scientific American, 83(19):289, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
289.pdf.

Anonymous:1900:CGA

[2168] Anonymous. That centrifugal gun again. Scientific American, 83 (19):290, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
290.pdf.

Anonymous:1900:AEM

[2169] Anonymous. American engineering methods for the development of India. Scientific American, 83(19):290, November 10, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n19/pdf/scientificamerican11101900-290a.pdf.

Anonymous:1900:SCA

[2170] Anonymous. Steamship competition on the Atlantic. Scientific American, 83(19):290, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
290b.pdf.

Anonymous:1900:LTE

[2171] Anonymous. Lake Tanganyika explorations. Scientific American, 83 (19):290–291, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
290c.pdf.

Villard:1900:CR

[2172] P. Villard. Cathode rays. Scientific American, 83(19):291, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n19/pdf/scientificamerican11101900-291.pdf.

Anonymous:1900:PGb

[2173] Anonymous. Purification of gases. Scientific American, 83(19): 291, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
291a.pdf.

Anonymous:1900:DMM

[2174] Anonymous. Death of Max Müller. Scientific American, 83 (19):291, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
291b.pdf.

Anonymous:1900:OE

[2175] Anonymous. Oysters in Europe. Scientific American, 83(19): 291–292, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
291c.pdf.

Anonymous:1900:SMSb

[2176] Anonymous. A square for marking shaft-keyways. Scientific American, 83(19):292, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
292.pdf.

Anonymous:1900:HS

[2177] Anonymous. The hydrometer syringe. Scientific American, 83 (19):292, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
292a.pdf.

Anonymous:1900:PTC

[2178] Anonymous. A portable thread-cutting machine. Scientific American, 83(19):292, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
292b.pdf.

Brush:1900:CTP

[2179] Edward Hale Brush. The color treatment of the Pan-American Exposition. Scientific American, 83(19):293–294, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n19/pdf/scientificamerican11101900-293.pdf.

Fawcett:1900:PAP

[2180] Waldon Fawcett. The production of American pottery. Scientific American, 83(19):296–297, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
296.pdf.

Anonymous:1900:ATS

[2181] Anonymous. An accident to a train of steel cars. Scientific American, 83(19):297–298, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
297.pdf.

Brush:1900:SNV

[2182] Edward Hale Brush. Six Nations village at the Pan-American Exposition. Scientific American, 83(19):298, November 10, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n19/pdf/scientificamerican11101900-298.pdf.

Anonymous:1900:PSEb

[2183] Anonymous. The puppet show at the Exposition. Scientific American, 83(19):299, November 10, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n19/pdf/scientificamerican11101900-
299.pdf.

Anonymous:1900:RPIbl

[2184] Anonymous. Recently patented inventions. Scientific American, 83(2):28, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n2/pdf/scientificamerican07141900-
28.pdf.

Anonymous:1900:BNBbf

[2185] Anonymous. Books: New books, etc. Scientific American, 83(2):28–29, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-28a.pdf.

Anonymous:1900:BPbl

[2186] Anonymous. Business and personal. Scientific American, 83(2):29, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-29.pdf.

Anonymous:1900:NQbl

[2187] Anonymous. Notes and queries. Scientific American, 83(2):29, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-29a.pdf.

Anonymous:1900:IIbm

[2188] Anonymous. Index of inventions. Scientific American, 83(2):29, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-29c.pdf.

Anonymous:1900:PNPbl

[2189] Anonymous. Patent notes: Patents. Scientific American, 83(2):31, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-31.pdf.

Anonymous:1900:FPD

[2190] Anonymous. Fire-proof docks — the moral of a great calamity. Scientific American, 83(2):18, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n2/pdf/scientificamerican07141900-
18.pdf.

Anonymous:1900:IR

[2191] Anonymous. An improvement in radiography. Scientific American, 83(2):18, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n2/pdf/scientificamerican07141900-
18a.pdf.

Anonymous:1900:CCCb

[2192] Anonymous. Curious chemical compound. Scientific American, 83(2):18, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n2/pdf/scientificamerican07141900-
18b.pdf.

Adams:1900:TCE

[2193] Alton D. Adams. Twentieth Century energy. Scientific American, 83(2):19, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n2/pdf/scientificamerican07141900-
19.pdf.

Anonymous:1900:PENh

[2194] Anonymous. Paris Exposition notes. Scientific American, 83(2):19, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-19a.pdf.

Anonymous:1900:NNF

[2195] Anonymous. A novel nail forming machine. Scientific American, 83(2):20, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n2/pdf/scientificamerican07141900-
20.pdf.

Anonymous:1900:E

[2196] Anonymous. Electroengraving. Scientific American, 83(2):20, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-20a.pdf.

Anonymous:1900:ACA

[2197] Anonymous. An automatic coupling for air-brakes. Scientific American, 83(2):20–21, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n2/pdf/scientificamerican07141900-
20b.pdf.

Anonymous:1900:BNG

[2198] Anonymous. The burning of the North German Lloyd Steamships and Docks, Hoboken, N.J. Scientific American, 83(2):21–23, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-21.pdf.

Anonymous:1900:MGAc

[2199] Anonymous. The manufacture of guns and armor at the Bethlehem Steel Works. Scientific American, 83(2):24–26, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-24.pdf.

Anonymous:1900:CTPb

[2200] Anonymous. A compensation tie-plate. Scientific American, 83(2):26, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-26.pdf.

Holder:1900:CC

[2201] Charles Frederick Holder. A California cycleway. Scientific American, 83(2):27, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n2/pdf/scientificamerican07141900-
27.pdf.

Anonymous:1900:Io

[2202] Anonymous. To inventors. Scientific American, 83(2):29, July 14, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n2/pdf/scientificamerican07141900-29b.pdf.

Anonymous:1900:Cbc

[2203] Anonymous. Correspondence. Scientific American, 83(20):311, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n20/pdf/scientificamerican11171900-311.pdf.

Anonymous:1900:RPIbm

[2204] Anonymous. Recently patented inventions. Scientific American, 83(20):316, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
316.pdf.

Anonymous:1900:BPbm

[2205] Anonymous. Business and personal. Scientific American, 83 (20):316, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
316a.pdf.

Anonymous:1900:NQbm

[2206] Anonymous. Notes and queries. Scientific American, 83(20): 316, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
316b.pdf.

Anonymous:1900:IIbn

[2207] Anonymous. Index of inventions. Scientific American, 83(20): 316–319, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
316d.pdf.

Anonymous:1900:PNPbm

[2208] Anonymous. Patent notes: Patents. Scientific American, 83 (20):318, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
319.pdf.

Anonymous:1900:WME

[2209] Anonymous. Wanted — a marine engine governor. Scientific American, 83(20):306, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
306.pdf.

Anonymous:1900:SRC

[2210] Anonymous. The sea route as a competitor of the Siberian Railway. Scientific American, 83(20):306, November 17, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n20/pdf/scientificamerican11171900-306a.pdf.

Anonymous:1900:CBTb

[2211] Anonymous. California’s big trees. Scientific American, 83(20): 306, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
306b.pdf.

Anonymous:1900:HMA

[2212] Anonymous. Home-made apparatus from home materials. Scientific American, 83(20):307, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
307.pdf.

Bichat:1900:AEP

[2213] ?. Bichat and ?. Swingedauw. Actino-electric phenomena produced by violet rays. Scientific American, 83(20):307, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n20/pdf/scientificamerican11171900-307a.pdf.

Anonymous:1900:SNb

[2214] Anonymous. Science notes. Scientific American, 83(20):307, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n20/pdf/scientificamerican11171900-307b.pdf.

Anonymous:1900:WEV

[2215] Anonymous. The woods electric vehicles. Scientific American, 83 (20):308, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
308.pdf.

Perkins:1900:EML

[2216] Frank C. Perkins. Electric mining locomotives. Scientific American, 83(20):309, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
309.pdf.

Anonymous:1900:SLPa

[2217] Anonymous. The St. Lawrence Power Company’s plant at Massena, N.Y. Scientific American, 83(20):310, November 17, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n20/pdf/scientificamerican11171900-310.pdf.

Anonymous:1900:SPN

[2218] Anonymous. A screw-propeller of novel form. Scientific American, 83(20):311, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
311a.pdf.

Anonymous:1900:NB

[2219] Anonymous. The new battleships. Scientific American, 83(20): 312–313, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
312.pdf.

Anonymous:1900:TSE

[2220] Anonymous. Traveling stairways for the Elevated Railways, New York. Scientific American, 83(20):313, November 17, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n20/pdf/scientificamerican11171900-313.pdf.

Hopkins:1900:HMS

[2221] George M. Hopkins. A home-made steam engine. Scientific American, 83(20):314, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n20/pdf/scientificamerican11171900-
314.pdf.

Anonymous:1900:ASMb

[2222] Anonymous. The Automobile Show at Madison Square Garden, New York. Scientific American, 83(20):314–315, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n20/pdf/scientificamerican11171900-314a.pdf.

Anonymous:1900:Ip

[2223] Anonymous. To inventors. Scientific American, 83(20):316, November 17, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n20/pdf/scientificamerican11171900-316c.pdf.

Anonymous:1900:RPIbn

[2224] Anonymous. Recently patented inventions. Scientific American, 83(21):332, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
332.pdf.

Anonymous:1900:BPbn

[2225] Anonymous. Business and personal. Scientific American, 83 (21):332, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
332a.pdf.

Anonymous:1900:NQbn

[2226] Anonymous. Notes and queries. Scientific American, 83(21): 332, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
332b.pdf.

Anonymous:1900:BNBbg

[2227] Anonymous. Books: New books, etc. Scientific American, 83 (21):332, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
332c.pdf.

Anonymous:1900:IIbo

[2228] Anonymous. Index of inventions. Scientific American, 83(21): 332–334, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
332d.pdf.

Anonymous:1900:PNPbn

[2229] Anonymous. Patent notes: Patents. Scientific American, 83 (21):335, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
335.pdf.

Anonymous:1900:OLC

[2230] Anonymous. Object lessons in centrifugal force. Scientific American, 83(21):322, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
322.pdf.

Anonymous:1900:ACF

[2231] Anonymous. Armor controversy finally settled. Scientific American, 83(21):322, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
322a.pdf.

Anonymous:1900:BPAa

[2232] Anonymous. Bogus patent attorneys in Great Britain. Scientific American, 83(21):322, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
322b.pdf.

Anonymous:1900:AMSb

[2233] Anonymous. The Annual Meeting of the Society of Naval Architects and Marine Engineers. Scientific American, 83(21):322, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n21/pdf/scientificamerican11241900-322c.pdf.

Anonymous:1900:CPE

[2234] Anonymous. Close of the Paris Exposition. Scientific American, 83(21):322, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
322d.pdf.

Melville:1900:STL

[2235] George W. Melville. The steam turbine in large merchant steamers. Scientific American, 83(21):323, November 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n21/pdf/scientificamerican11241900-323.pdf.

Hale:1900:NAS

[2236] William H. Hale. National Academy of Sciences, Providence Meeting. Scientific American, 83(21):323, November 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n21/pdf/scientificamerican11241900-323a.pdf.

Perkins:1900:ST

[2237] Frank C. Perkins. Swiss turbines. Scientific American, 83(21): 324, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
324.pdf.

Anonymous:1900:HCP

[2238] Anonymous. A holder for car-platform stakes. Scientific American, 83(21):325, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
325.pdf.

Anonymous:1900:PSIb

[2239] Anonymous. A paint-striper of improved form. Scientific American, 83(21):325, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
325a.pdf.

Anonymous:1900:UT

[2240] Anonymous. The Underwood typewriter. Scientific American, 83 (21):325–326, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
325b.pdf.

Anonymous:1900:PWR

[2241] Anonymous. Progress of work on the Rapid Transit Tunnel, New York. Scientific American, 83(21):326, November 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n21/pdf/scientificamerican11241900-326.pdf.

Anonymous:1900:SPSa

[2242] Anonymous. The selden patent suit. Scientific American, 83 (21):327, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
327.pdf.

Field:1900:RGF

[2243] C. Field. Remarkable gravestone at Forres, Scotland. Scientific American, 83(21):328, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
328.pdf.

Anonymous:1900:TBV

[2244] Anonymous. The torpedo boat “Viper” in dry dock. Scientific American, 83(21):328, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
328a.pdf.

Brush:1900:PPA

[2245] Edward Hale Brush. Progress of the Pan-American Exposition. Scientific American, 83(21):329–330, November 24, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n21/pdf/scientificamerican11241900-329.pdf.

Hopkins:1900:IEC

[2246] George M. Hopkins. II. the electric chime. Scientific American, 83(21):331, November 24, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n21/pdf/scientificamerican11241900-
331.pdf.

Russell:1900:HD

[2247] Henry Norris Russell. The heavens in December. Scientific American, 83(22):339, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
339.pdf.

Anonymous:1900:Cbd

[2248] Anonymous. Correspondence. Scientific American, 83(22):342, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-342.pdf.

Anonymous:1900:RPIbo

[2249] Anonymous. Recently patented inventions. Scientific American, 83(22):347, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
347.pdf.

Anonymous:1900:BNBbh

[2250] Anonymous. Books: New books, etc. Scientific American, 83 (22):347, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
347a.pdf.

Anonymous:1900:BPbo

[2251] Anonymous. Business and personal. Scientific American, 83 (22):348, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
348.pdf.

Anonymous:1900:NQbo

[2252] Anonymous. Notes and queries. Scientific American, 83(22):348, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-348a.pdf.

Anonymous:1900:IIbp

[2253] Anonymous. Index of inventions. Scientific American, 83(22):348, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-348b.pdf.

Anonymous:1900:PNPbo 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
350.pdf.

Anonymous:1900:CTG

[2255] Anonymous. Congestion of traffic at the Grand Central Station and its remedy. Scientific American, 83(22):338, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-338.pdf.

Anonymous:1900:PCSa

[2256] Anonymous. Pacific Coast ship-building. Scientific American, 83 (22):338, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
338a.pdf.

Anonymous:1900:CEK

[2257] Anonymous. The comparative efficiency of the Krupp, Armstrong, and Schneider–Canet guns. Scientific American, 83(22):338, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-338b.pdf.

Right:1900:C

[2258] Aug Right. Coherers. Scientific American, 83(22):339, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-339a.pdf.

Anonymous:1900:SNc

[2259] Anonymous. Science notes. Scientific American, 83(22):339, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-339b.pdf.

Anonymous:1900:AL

[2260] Anonymous. Artificial lightning. Scientific American, 83(22):340, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-340.pdf.

Holder:1900:FFW

[2261] Charles Frederick Holder. Fighting fire with wine. Scientific American, 83(22):340, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
340a.pdf.

Anonymous:1900:NBH

[2262] Anonymous. A new bicycle handle-bar. Scientific American, 83 (22):341, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
341.pdf.

Fawcett:1900:HAC

[2263] Waldon Fawcett. Handling anthracite coal. Scientific American, 83 (22):341–342, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
341a.pdf.

Anonymous:1900:LAA

[2264] Anonymous. Liquid air automobile. Scientific American, 83(22): 343, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
343.pdf.

Anonymous:1900:MPW

[2265] Anonymous. The making of pictures in wood. Scientific American, 83(22):344, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
344.pdf.

Anonymous:1900:NACb

[2266] Anonymous. The new armored cruisers of the “California” and “Maryland” types. Scientific American, 83(22):344–345, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n22/pdf/scientificamerican12011900-344a.pdf.

Anonymous:1900:PPL

[2267] Anonymous. The Pollak Prize for Life-Saving Devices. Scientific American, 83(22):345–346, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
345.pdf.

Anonymous:1900:ECRb

[2268] Anonymous. Electric clock run by earth plates. Scientific American, 83(22):346–347, December 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n22/pdf/scientificamerican12011900-
346.pdf.

Anonymous:1900:RPIbp

[2269] Anonymous. Recently patented inventions. Scientific American, 83 (23):363–364, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
363.pdf.

Anonymous:1900:BPbp

[2270] Anonymous. Business and personal. Scientific American, 83 (23):364, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
364.pdf.

Anonymous:1900:NQbp

[2271] Anonymous. Notes and queries. Scientific American, 83(23):364, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n23/pdf/scientificamerican12081900-364a.pdf.

Anonymous:1900:BNBbi

[2272] Anonymous. Books: New books, etc. Scientific American, 83 (23):364, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
364b.pdf.

Anonymous:1900:IIbq

[2273] Anonymous. Index of inventions. Scientific American, 83(23):364, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n23/pdf/scientificamerican12081900-364c.pdf.

Anonymous:1900:PNPbp

[2274] Anonymous. Patent notes: Patents. Scientific American, 83 (23):367, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
367.pdf.

Anonymous:1900:PTM

[2275] Anonymous. The Patent and Trade Mark Commission agree on a bill. Scientific American, 83(23):354, December 8, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n23/pdf/scientificamerican12081900-354.pdf.

Anonymous:1900:FDC

[2276] Anonymous. Fireproof dock construction. Scientific American, 83 (23):354, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
354a.pdf.

Anonymous:1900:CICb

[2277] Anonymous. Congress and the Isthmian Canal. Scientific American, 83(23):354, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
354b.pdf.

Anonymous:1900:CCU

[2278] Anonymous. The chief constructor of the United States Navy. Scientific American, 83(23):354–355, December 8, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n23/pdf/scientificamerican12081900-354c.pdf.

Anonymous:1900:ICR

[2279] Anonymous. International Congress of Railways at Paris. Scientific American, 83(23):355, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
355.pdf.

Anonymous:1900:RUS

[2280] Anonymous. Registration of United States patents and trade marks in Cuba. Scientific American, 83(23):355, December 8, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n23/pdf/scientificamerican12081900-355a.pdf.

Anonymous:1900:LP

[2281] Anonymous. Letters from Peary. Scientific American, 83(23):355, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n23/pdf/scientificamerican12081900-355b.pdf.

Anonymous:1900:FDL

[2282] Anonymous. Failure of the display of Leonids. Scientific American, 83(23):355, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
355c.pdf.

Anonymous:1900:SIW

[2283] Anonymous. Sugar inventions wanted. Scientific American, 83 (23):355, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
355d.pdf.

Anonymous:1900:BIR

[2284] Anonymous. The brunt improved regulating incandescent-lamp socket. Scientific American, 83(23):356, December 8, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n23/pdf/scientificamerican12081900-356.pdf.

Anonymous:1900:NEM

[2285] Anonymous. A novel electro-magnetic brake for street cars. Scientific American, 83(23):356, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
356a.pdf.

Anonymous:1900:ACB

[2286] Anonymous. American coal briquettes. Scientific American, 83 (23):356–357, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
356b.pdf.

Anonymous:1900:SSAb

[2287] Anonymous. The Serpollet steam automobile. Scientific American, 83(23):357–358, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
357.pdf.

Anonymous:1900:FCBb

[2288] Anonymous. The first-class battleship “Wisconsin”. Scientific American, 83(23):358–359, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
358.pdf.

Anonymous:1900:BEV

[2289] Anonymous. Bird’s-eye view of the Pan-American Exposition. Scientific American, 83(23):361, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
361.pdf.

Anonymous:1900:SBTb

[2290] Anonymous. Separation of the Brazilian thoraco-xiphopagous twins. Scientific American, 83(23):361–362, December 8, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n23/pdf/scientificamerican12081900-361a.pdf.

Hopkins:1900:ISEa

[2291] George M. Hopkins. III. simple electric motor. Scientific American, 83(23):362–363, December 8, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n23/pdf/scientificamerican12081900-
362.pdf.

Anonymous:1900:RPIbq

[2292] Anonymous. Recently patented inventions. Scientific American, 83(24):379, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
379.pdf.

Anonymous:1900:BNBbj

[2293] Anonymous. Books: New books, etc. Scientific American, 83 (24):379, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
379a.pdf.

Anonymous:1900:BPbq

[2294] Anonymous. Business and personal. Scientific American, 83 (24):380, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
380.pdf.

Anonymous:1900:NQbq

[2295] Anonymous. Notes and queries. Scientific American, 83(24): 380, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
380a.pdf.

Anonymous:1900:IIbr

[2296] Anonymous. Index of inventions. Scientific American, 83(24): 380–382, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
380c.pdf.

Anonymous:1900:PNPbq

[2297] Anonymous. Patent notes: Patents. Scientific American, 83 (24):382, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
382.pdf.

Anonymous:1900:RICa

[2298] Anonymous. Report of the Isthmian Canal Commission. Scientific American, 83(24):370, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
370.pdf.

Anonymous:1900:RCS

[2299] Anonymous. The relative cost of steamship construction in Europe and America. Scientific American, 83(24):370, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n24/pdf/scientificamerican12151900-370a.pdf.

Anonymous:1900:RCP

[2300] Anonymous. Report of the Commission on Patents and Trade Marks. Scientific American, 83(24):370–371, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n24/pdf/scientificamerican12151900-370b.pdf.

Anonymous:1900:SC

[2301] Anonymous. Shooting at the clouds. Scientific American, 83 (24):371, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
371.pdf.

Anonymous:1900:AFM

[2302] Anonymous. Alcohol as a fuel for motor carriages. Scientific American, 83(24):371, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
371a.pdf.

Anonymous:1900:DCA

[2303] Anonymous. The direct combination of alumina and oxide of calcium in the electric furnace. Scientific American, 83(24):371, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n24/pdf/scientificamerican12151900-371b.pdf.

Anonymous:1900:ETL

[2304] Anonymous. Electric tramways in London. Scientific American, 83(24):371, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
371c.pdf.

Anonymous:1900:CKF

[2305] Anonymous. Curious knot formation. Scientific American, 83 (24):372, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
372.pdf.

Anonymous:1900:SPA

[2306] Anonymous. A simple portable acetylene house-lamp. Scientific American, 83(24):372, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
372a.pdf.

Anonymous:1900:NTM

[2307] Anonymous. New telephone meter. Scientific American, 83(24): 372, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
372b.pdf.

Anonymous:1900:SAW

[2308] Anonymous. A safety attachment for water-gages. Scientific American, 83(24):373, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
373.pdf.

Anonymous:1900:STE

[2309] Anonymous. Scotte traction engine — type shown at the Paris Exposition. Scientific American, 83(24):373–374, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n24/pdf/scientificamerican12151900-373a.pdf.

Hale:1900:MSC

[2310] Henry Hale. Modern soft-coal mining and handling in the United States. Scientific American, 83(24):374–375, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n24/pdf/scientificamerican12151900-374.pdf.

Anonymous:1900:EFC

[2311] Anonymous. Electric freight conveyers for loading and unloading vessels. Scientific American, 83(24):376, December 15, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n24/pdf/scientificamerican12151900-376.pdf.

Anonymous:1900:NIN

[2312] Anonymous. The new 12-inch Naval gun. Scientific American, 83 (24):376–377, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
376a.pdf.

Anonymous:1900:TRE

[2313] Anonymous. The third-rail equipment of the Manhattan Elevated Railways. Scientific American, 83(24):377, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n24/pdf/scientificamerican12151900-377.pdf.

Hopkins:1900:ISEb

[2314] George M. Hopkins. IV. simple electric motor. Scientific American, 83(24):378, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n24/pdf/scientificamerican12151900-
378.pdf.

Anonymous:1900:Iq

[2315] Anonymous. To inventors. Scientific American, 83(24):380, December 15, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n24/pdf/scientificamerican12151900-380b.pdf.

Anonymous:1900:RPIbr

[2316] Anonymous. Recently patented inventions. Scientific American, 83(25):396, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
396.pdf.

Anonymous:1900:BPbr

[2317] Anonymous. Business and personal. Scientific American, 83 (25):396, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
396a.pdf.

Anonymous:1900:NQbr

[2318] Anonymous. Notes and queries. Scientific American, 83(25): 396, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
396b.pdf.

Anonymous:1900:BNBbk

[2319] Anonymous. Books: New books, etc. Scientific American, 83 (25):396, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
396c.pdf.

Anonymous:1900:IIbs

[2320] Anonymous. Index of inventions. Scientific American, 83(25): 396–399, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
396e.pdf.

Anonymous:1900:PNPbr

[2321] Anonymous. Patent notes: Patents. Scientific American, 83 (25):399, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
399.pdf.

Anonymous:1900:MSC

[2322] Anonymous. The metric system in Congress. Scientific American, 83(25):386, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
386.pdf.

Anonymous:1900:BAC

[2323] Anonymous. Boston and the “America” Cup. Scientific American, 83(25):386, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
386a.pdf.

Anonymous:1900:FOW

[2324] Anonymous. Filtration of our water supply. Scientific American, 83(25):386, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
386b.pdf.

Anonymous:1900:IEC

[2325] Anonymous. Imports and exports of the countries of the world compared. Scientific American, 83(25):386, December 22, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n25/pdf/scientificamerican12221900-386c.pdf.

Anonymous:1900:RCB

[2326] Anonymous. Report of the Chief of the Bureau of Ordnance. Scientific American, 83(25):386, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
386d.pdf.

Anonymous:1900:GIO

[2327] Anonymous. The greatest iron ore mine in the world. Scientific American, 83(25):386–387, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
386e.pdf.

Blackford:1900:ABP

[2328] Charles Minor Blackford, Jr. Animals that bloom and plants that eat meat. Scientific American, 83(25):387, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n25/pdf/scientificamerican12221900-387.pdf.

Anonymous:1900:ISI

[2329] Anonymous. The Iron and Steel Institute at Paris. Scientific American, 83(25):387, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
387a.pdf.

Anonymous:1900:NRW

[2330] Anonymous. A new ratchet-wrench. Scientific American, 83 (25):388, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
388.pdf.

Anonymous:1900:PSS

[2331] Anonymous. Power scrapers on the St. Lawrence Power Company’s Canal. Scientific American, 83(25):388, December 22, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n25/pdf/scientificamerican12221900-388a.pdf.

Anonymous:1900:PEP

[2332] Anonymous. The Plecher electro-pneumatic telephone. Scientific American, 83(25):388, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
388b.pdf.

Anonymous:1900:RME

[2333] Anonymous. A recharging motor for electric vehicles. Scientific American, 83(25):389, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
389.pdf.

Anonymous:1900:MEP

[2334] Anonymous. The manufacture of Edison phonograph records. Scientific American, 83(25):390–391, December 22, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n25/pdf/scientificamerican12221900-390.pdf.

Anonymous:1900:NWW

[2335] Anonymous. A novel wind or water motor. Scientific American, 83(25):391, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
391.pdf.

Anonymous:1900:RDA

[2336] Anonymous. The roze dirigible airship. Scientific American, 83 (25):392, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
392.pdf.

Robinson:1900:DD

[2337] W. E. Robinson. Disappearing demon. Scientific American, 83 (25):393, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
393.pdf.

Anonymous:1900:PCSb

[2338] Anonymous. The protected cruisers of the “St. Louis” class. Scientific American, 83(25):393–394, December 22, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n25/pdf/scientificamerican12221900-393a.pdf.

Holder:1900:AB

[2339] Charles F. Holder. Armored burrowers. Scientific American, 83 (25):394, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
394.pdf.

Hopkins:1900:VIW

[2340] George M. Hopkins. V. an inexpensive water motor. Scientific American, 83(25):395, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n25/pdf/scientificamerican12221900-
395.pdf.

Anonymous:1900:It

[2341] Anonymous. To inventors. Scientific American, 83(25):396, December 22, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n25/pdf/scientificamerican12221900-396d.pdf.

Anonymous:1900:RPIbs

[2342] Anonymous. Recently patented inventions. Scientific American, 83(26):413, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
413.pdf.

Anonymous:1900:BNBbl

[2343] Anonymous. Books: New books, etc. Scientific American, 83(26): 413–414, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
413a.pdf.

Anonymous:1900:BPbs

[2344] Anonymous. Business and personal. Scientific American, 83 (26):414, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
414.pdf.

Anonymous:1900:NQbs

[2345] Anonymous. Notes and queries. Scientific American, 83(26): 414, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
414a.pdf.

Anonymous:1900:IIbt

[2346] Anonymous. Index of inventions. Scientific American, 83(26): 414–419, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
414c.pdf.

Anonymous:1900:PNPbs

[2347] Anonymous. Patent notes: Patents. Scientific American, 83 (26):419, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
419.pdf.

Byrn:1900:CPU

[2348] Edward W. Byrn. A century of progress in the United States. Scientific American, 83(26):402–403, December 29, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n26/pdf/scientificamerican12291900-402.pdf.

Anonymous:1900:RT

[2349] Anonymous. Ritchie’s telautograph. Scientific American, 83(26): 404, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
404.pdf.

Anonymous:1900:SAEb

[2350] Anonymous. Some automobiles exhibited at the Paris Exposition. Scientific American, 83(26):405, December 29, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n26/pdf/scientificamerican12291900-405.pdf.

Anonymous:1900:RMV

[2351] Anonymous. A rapid method of vulcanizing rubber. Scientific American, 83(26):406, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
406.pdf.

Strauss:1900:USN

[2352] Jos. Strauss. Unites States Naval Proving Ground, Indian Head. Scientific American, 83(26):407, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
407.pdf.

Anonymous:1900:HGR

[2353] Anonymous. The Halford gradient railway. Scientific American, 83(26):408, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
408.pdf.

Willey:1900:LN

[2354] D. A. Willey. Logging in the Northwest. Scientific American, 83 (26):409–410, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
409.pdf.

Hopkins:1900:VMT

[2355] George M. Hopkins. VI. A metallic thermometer. Scientific American, 83(26):411, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
411.pdf.

Holder:1900:SCS

[2356] Charles Frederick Holder. Some curious sculpins. Scientific American, 83(26):411–412, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
411a.pdf.

Anonymous:1900:SRPa

[2357] Anonymous. Some recently patented novelties. Scientific American, 83(26):412, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n26/pdf/scientificamerican12291900-
412.pdf.

Anonymous:1900:Iu

[2358] Anonymous. To inventors. Scientific American, 83(26):414, December 29, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n26/pdf/scientificamerican12291900-414b.pdf.

Anonymous:1900:Cbe

[2359] Anonymous. Correspondence. Scientific American, 83(3):44, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-44.pdf.

Anonymous:1900:RPIbt

[2360] Anonymous. Recently patented inventions. Scientific American, 83(3):44–45, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
44a.pdf.

Anonymous:1900:BNBbm

[2361] Anonymous. Books: New books, etc. Scientific American, 83(3):45, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-45.pdf.

Anonymous:1900:BPbt

[2362] Anonymous. Business and personal. Scientific American, 83(3):45, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-45a.pdf.

Anonymous:1900:NQbt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-45b.pdf.

Anonymous:1900:IIbu

[2364] Anonymous. Index of inventions. Scientific American, 83(3):45, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-45c.pdf.

Anonymous:1900:PNPbt

[2365] Anonymous. Patent notes: Patents. Scientific American, 83(3):47, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-47.pdf.

Anonymous:1900:AECa

[2366] Anonymous. American engineering competition. Scientific American, 83(3):34, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
34.pdf.

Anonymous:1900:OBL

[2367] Anonymous. Our battleships in the light of the “Belleisle” experiment. Scientific American, 83(3):34, July 21, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n3/pdf/scientificamerican07211900-34a.pdf.

Anonymous:1900:D

[2368] Anonymous. The “Deutschland”. Scientific American, 83(3):34, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-34b.pdf.

Anonymous:1900:PSC

[2369] Anonymous. Pressed steel cars. Scientific American, 83(3):35, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-35.pdf.

Anonymous:1900:FCR

[2370] Anonymous. French Congo region. Scientific American, 83(3):35, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-35a.pdf.

Anonymous:1900:AAS

[2371] Anonymous. Anti-alcohol serum in France. Scientific American, 83(3):35, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
35b.pdf.

Anonymous:1900:PENi

[2372] Anonymous. Paris Exposition notes. Scientific American, 83(3):35, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-35c.pdf.

Anonymous:1900:CGG

[2373] Anonymous. Canet gun with gas-recoil-check. Scientific American, 83(3):36, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
36.pdf.

Anonymous:1900:FS

[2374] Anonymous. A fireproof shutter. Scientific American, 83(3):36, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-36a.pdf.

Anonymous:1900:NST

[2375] Anonymous. A new steam turbine. Scientific American, 83(3):36, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-36b.pdf.

Anonymous:1900:CCa

[2376] Anonymous. China and the Chinese. Scientific American, 83(3):37–38, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-37.pdf.

Anonymous:1900:RDV

[2377] Anonymous. Riverside Drive Viaduct, New York. Scientific American, 83(3):38–39, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
38.pdf.

Anonymous:1900:LNU

[2378] Anonymous. London’s new underground electric railway. Scientific American, 83(3):40, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
40.pdf.

Anonymous:1900:ACD

[2379] Anonymous. Attack on the coast-defense ship “Belleisle”. Scientific American, 83(3):41–42, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
41.pdf.

Anonymous:1900:IMG

[2380] Anonymous. An improved marking-gage. Scientific American, 83(3):42, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n3/pdf/scientificamerican07211900-42.pdf.

Anonymous:1900:SLPb

[2381] Anonymous. The Statue of Lafayette at Paris. Scientific American, 83(3):43, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
43.pdf.

Anonymous:1900:SHS

[2382] Anonymous. A springless harness snap-hook. Scientific American, 83(3):43, July 21, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n3/pdf/scientificamerican07211900-
43a.pdf.

Anonymous:1900:RPIbu

[2383] Anonymous. Recently patented inventions. Scientific American, 83(4):60–61, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n4/pdf/scientificamerican07281900-
60.pdf.

Anonymous:1900:BNBbn

[2384] Anonymous. Books: New books, etc. Scientific American, 83(4):61, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-61.pdf.

Anonymous:1900:BPbu

[2385] Anonymous. Business and personal. Scientific American, 83(4):61, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-61a.pdf.

Anonymous:1900:NQbu

[2386] Anonymous. Notes and queries. Scientific American, 83(4):61, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-61b.pdf.

Anonymous:1900:IIbv

[2387] Anonymous. Index of inventions. Scientific American, 83(4):61, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-61c.pdf.

Anonymous:1900:PNPbu

[2388] Anonymous. Patent notes: Patents. Scientific American, 83(4):63, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-63.pdf.

Anonymous:1900:FW

[2389] Anonymous. Fireproofing wood. Scientific American, 83(4):49, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-49.pdf.

Anonymous:1900:ECSb

[2390] Anonymous. The end-of the-century steamship. Scientific American, 83(4):50, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n4/pdf/scientificamerican07281900-
50.pdf.

Anonymous:1900:OVR

[2391] Anonymous. Our vast railway system. Scientific American, 83(4):50, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-50a.pdf.

Anonymous:1900:ION

[2392] Anonymous. Increase of our Navy since the Spanish War. Scientific American, 83(4):50–51, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n4/pdf/scientificamerican07281900-
50b.pdf.

Russell:1900:HAb

[2393] Henry Norris Russell. The heavens in August. Scientific American, 83(4):51, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n4/pdf/scientificamerican07281900-
51.pdf.

Anonymous:1900:IER

[2394] Anonymous. An International Exhibition and Race-Meet at Chicago. Scientific American, 83(4):51, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n4/pdf/scientificamerican07281900-
51a.pdf.

Anonymous:1900:LWR

[2395] Anonymous. A large water rolling mill. Scientific American, 83(4):52, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-52.pdf.

Anonymous:1900:ATB

[2396] Anonymous. Accident to the torpedo-boat “Dupont”. Scientific American, 83(4):52, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n4/pdf/scientificamerican07281900-
52a.pdf.

Trowbridge:1900:EPJ

[2397] John Trowbridge. The electrical plant of the Jefferson Physical Laboratory. Scientific American, 83(4):53–54, July 28, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n4/pdf/scientificamerican07281900-53.pdf.

Anonymous:1900:ICc

[2398] Anonymous. An ingenious calculator. Scientific American, 83(4):54, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-54.pdf.

Anonymous:1900:YSI

[2399] Anonymous. A yacht that sails itself. Scientific American, 83(4):56, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-56.pdf.

Anonymous:1900:LW

[2400] Anonymous. Locomobile wrinkles. Scientific American, 83(4):57, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n4/pdf/scientificamerican07281900-57.pdf.

Mason:1900:TAI

[2401] Otis T. Mason. Types of American Indian basketry. Scientific American, 83(4):57–58, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n4/pdf/scientificamerican07281900-
57a.pdf.

Anonymous:1900:TNW

[2402] Anonymous. Through Niagara Whirlpool Rapids in a boat. Scientific American, 83(4):59–60, July 28, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n4/pdf/scientificamerican07281900-
59.pdf.

Anonymous:1900:RPIbv

[2403] Anonymous. Recently patented inventions. Scientific American, 83(5):76, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
76.pdf.

Anonymous:1900:BPbv

[2404] Anonymous. Business and personal. Scientific American, 83(5):77, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-77.pdf.

Anonymous:1900:NQbv

[2405] Anonymous. Notes and queries. Scientific American, 83(5):77, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-77a.pdf.

Anonymous:1900:IIbw

[2406] Anonymous. Index of inventions. Scientific American, 83(5):77–79, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-77c.pdf.

Anonymous:1900:PNPbv

[2407] Anonymous. Patent notes: Patents. Scientific American, 83(5):79, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-79.pdf.

Anonymous:1900:EEP

[2408] Anonymous. Electrical equipment of the park avenue tunnel. Scientific American, 83(5):66, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
66.pdf.

Anonymous:1900:IPS

[2409] Anonymous. An important patent suit. Scientific American, 83(5):66, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-66a.pdf.

Anonymous:1900:RJS

[2410] Anonymous. Rail-joints and street noises. Scientific American, 83(5):66, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-66b.pdf.

Anonymous:1900:RICb

[2411] Anonymous. Resources and improvements in China. Scientific American, 83(5):66–67, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
66c.pdf.

Anonymous:1900:NAT

[2412] Anonymous. The new agriculture of the tropics. Scientific American, 83(5):67, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
67.pdf.

Anonymous:1900:Iv

[2413] Anonymous. Indican. Scientific American, 83(5):67, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-67a.pdf.

Anonymous:1900:SU

[2414] Anonymous. As others see us. Scientific American, 83(5):67, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-67b.pdf.

Anonymous:1900:PWA

[2415] Anonymous. The preservation of wild animals of Africa. Scientific American, 83(5):67, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
67c.pdf.

Anonymous:1900:QSC

[2416] Anonymous. The Queen–Schultz chronoscope. Scientific American, 83(5):68, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
68.pdf.

Anonymous:1900:APN

[2417] Anonymous. Artillery practice of the National Guard. Scientific American, 83(5):68–69, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
68a.pdf.

Anonymous:1900:HAF

[2418] Anonymous. The Hone automatic flood-gate. Scientific American, 83(5):69, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
69.pdf.

Anonymous:1900:PPP

[2419] Anonymous. A perforator for printing presses. Scientific American, 83(5):69, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
69a.pdf.

Anonymous:1900:IPa

[2420] Anonymous. An improved packing. Scientific American, 83(5):69–70, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-69b.pdf.

Anonymous:1900:CFS

[2421] Anonymous. A convenient form of sash-weight. Scientific American, 83(5):70, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
70.pdf.

Sharp:1900:GO

[2422] Benjamin Sharp. The guardian of the oyster. Scientific American, 83(5):71, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
71.pdf.

Anonymous:1900:HAL

[2423] Anonymous. The Hamburg–American liner “Deutschland”. Scientific American, 83(5):72, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n5/pdf/scientificamerican08041900-
72.pdf.

Mann:1900:PGP

[2424] F. P. Mann. The power generating plant at the Paris Exposition. Scientific American, 83(5):73–74, August 4, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n5/pdf/scientificamerican08041900-73.pdf.

Anonymous:1900:CFC

[2425] Anonymous. Clever feat with the camera. Scientific American, 83(5):75, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-75.pdf.

Anonymous:1900:MB

[2426] Anonymous. A musical bicycle. Scientific American, 83(5):75–76, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-75a.pdf.

Anonymous:1900:Ix

[2427] Anonymous. To inventors. Scientific American, 83(5):77, August 4, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n5/pdf/scientificamerican08041900-77b.pdf.

Anonymous:1900:RPIbw

[2428] Anonymous. Recently patented inventions. Scientific American, 83(6):93, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
93.pdf.

Anonymous:1900:BNBbo

[2429] Anonymous. Books: New books, etc. Scientific American, 83(6):93, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n6/pdf/scientificamerican08111900-93a.pdf.

Anonymous:1900:BPbw

[2430] Anonymous. Business and personal. Scientific American, 83(6):93, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n6/pdf/scientificamerican08111900-93b.pdf.

Anonymous:1900:NQbw

[2431] Anonymous. Notes and queries. Scientific American, 83(6):93, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n6/pdf/scientificamerican08111900-93c.pdf.

Anonymous:1900:IIbx

[2432] Anonymous. Index of inventions. Scientific American, 83(6):93–95, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n6/pdf/scientificamerican08111900-93e.pdf.

Anonymous:1900:PNPbw

[2433] Anonymous. Patent notes: Patents. Scientific American, 83(6):95, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n6/pdf/scientificamerican08111900-95.pdf.

Anonymous:1900:PAG

[2434] Anonymous. The protection of American game. Scientific American, 83(6):82, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
82.pdf.

Anonymous:1900:HSW

[2435] Anonymous. High speed in war vessels — the “Variag” and the “Viper”. Scientific American, 83(6):82, August 11, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n6/pdf/scientificamerican08111900-82a.pdf.

Anonymous:1900:AECb

[2436] Anonymous. American engineering competition. Scientific American, 83(6):82–83, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
82b.pdf.

Adams:1900:PFW

[2437] Alton D. Adams. Petroleum fuel for warships. Scientific American, 83(6):83, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
83.pdf.

Anonymous:1900:OX

[2438] Anonymous. Operation on the xiphopages. Scientific American, 83(6):83, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
83a.pdf.

Anonymous:1900:NES

[2439] Anonymous. A novel engine for die-sinker’s use. Scientific American, 83(6):84, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
84.pdf.

Dailey:1900:ETM

[2440] Howard B. Dailey. Electrical top and motor. Scientific American, 83(6):84, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
84a.pdf.

Crafts:1900:WMM

[2441] H. A. Crafts. Water measurement and manipulation in Colorado. Scientific American, 83(6):85, August 11, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n6/pdf/scientificamerican08111900-85.pdf.

Anonymous:1900:PEd

[2442] Anonymous. The Paris Exposition. Scientific American, 83(6): 86–87, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
86.pdf.

Anonymous:1900:ACZ

[2443] Anonymous. The ascension of Count Zeppelin’s airship. Scientific American, 83(6):88, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
88.pdf.

Anonymous:1900:UER

[2444] Anonymous. Underground Electric Railway, London — methods of construction. Scientific American, 83(6):89–90, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n6/pdf/scientificamerican08111900-89.pdf.

Willey:1900:PFA

[2445] D. Allen Willey. The Petrified Forest of Arizona. Scientific American, 83(6):91, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n6/pdf/scientificamerican08111900-
91.pdf.

Anonymous:1900:RER

[2446] Anonymous. Recent experiments with the radiations of polonium and radium. Scientific American, 83(6):92, August 11, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n6/pdf/scientificamerican08111900-92.pdf.

Anonymous:1900:Iy

[2447] Anonymous. To inventors. Scientific American, 83(6):93, August 11, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n6/pdf/scientificamerican08111900-93d.pdf.

Anonymous:1900:Cbf

[2448] Anonymous. Correspondence. Scientific American, 83(7):102, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n7/pdf/scientificamerican08181900-102.pdf.

Anonymous:1900:RPIbx

[2449] Anonymous. Recently patented inventions. Scientific American, 83 (7):108–109, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
108.pdf.

Anonymous:1900:BNBbp

[2450] Anonymous. Books: New books, etc. Scientific American, 83 (7):109, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
109.pdf.

Anonymous:1900:BPbx

[2451] Anonymous. Business and personal. Scientific American, 83(7):109, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n7/pdf/scientificamerican08181900-109a.pdf.

Anonymous:1900:NQbx

[2452] Anonymous. Notes and queries. Scientific American, 83(7):109, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n7/pdf/scientificamerican08181900-109b.pdf.

Anonymous:1900:IIby

[2453] Anonymous. Index of inventions. Scientific American, 83(7): 109–111, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
109d.pdf.

Anonymous:1900:PNPbx

[2454] Anonymous. Patent notes: Patents. Scientific American, 83(7):111, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n7/pdf/scientificamerican08181900-111.pdf.

Anonymous:1900:ESSa

[2455] Anonymous. Economies of superheated steam in the steam turbine. Scientific American, 83(7):98, August 18, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n7/pdf/scientificamerican08181900-98.pdf.

Anonymous:1900:TSB

[2456] Anonymous. Tool steel at the Bethlehem Steel Works. Scientific American, 83(7):98, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
98a.pdf.

Anonymous:1900:CAP

[2457] Anonymous. Chinese artillery and the present crisis. Scientific American, 83(7):98, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
98b.pdf.

Anonymous:1900:SSS

[2458] Anonymous. A suggestion as to street signs. Scientific American, 83(7):98, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
98c.pdf.

Anonymous:1900:LRS

[2459] Anonymous. The leading railway systems of the United States. Scientific American, 83(7):98, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
98d.pdf.

Anonymous:1900:EMR

[2460] Anonymous. The entrance to the Mississippi River. Scientific American, 83(7):98–99, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
98e.pdf.

Anonymous:1900:ISP

[2461] Anonymous. Irrigating the sugar plantations of Hawaii. Scientific American, 83(7):99, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
99.pdf.

Anonymous:1900:NFCb

[2462] Anonymous. New form of cellulose. Scientific American, 83(7):99, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n7/pdf/scientificamerican08181900-99a.pdf.

Anonymous:1900:IMC

[2463] Anonymous. International Mining Congress. Scientific American, 83(7):99, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
99b.pdf.

Anonymous:1900:DWC

[2464] Anonymous. Death of De Witt Clinton Haskin. Scientific American, 83(7):99, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
99c.pdf.

Anonymous:1900:Cbh

[2465] Anonymous. Caviar. Scientific American, 83(7):99, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n7/pdf/scientificamerican08181900-99d.pdf.

Anonymous:1900:TLG

[2466] Anonymous. Telegraph line in Germany. Scientific American, 83 (7):99, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
99e.pdf.

Anonymous:1900:CCS

[2467] Anonymous. A curious case of skin-shedding. Scientific American, 83(7):100, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
100.pdf.

Marsh:1900:MMC

[2468] Alonzo R. Marsh. The Marsh motor cycle. Scientific American, 83(7):100–101, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
100a.pdf.

Anonymous:1900:NFE

[2469] Anonymous. New form of electric automobile. Scientific American, 83(7):101, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
101.pdf.

Anonymous:1900:MPL

[2470] Anonymous. The most powerful locomotive ever built. Scientific American, 83(7):101, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
101a.pdf.

Anonymous:1900:MMR

[2471] Anonymous. The manufacture of mechanical rubber goods. Scientific American, 83(7):102–103, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
102a.pdf.

McGarvey:1900:UHV

[2472] Edward McGarvey. The use of harmonic vibrations in weighing. Scientific American, 83(7):104–105, August 18, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n7/pdf/scientificamerican08181900-104.pdf.

Anonymous:1900:PHE

[2473] Anonymous. Palace of Hungary at the Exposition. Scientific American, 83(7):105–106, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
105.pdf.

Mindeleff:1900:WIN

[2474] Cosmos Mindeleff. White Indians of New Mexico. Scientific American, 83(7):107, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
107.pdf.

Holder:1900:SPJ

[2475] Charles F. Holder. Some Pacific jellyfishes. Scientific American, 83(7):107–108, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n7/pdf/scientificamerican08181900-
107a.pdf.

Anonymous:1900:Iz

[2476] Anonymous. To inventors. Scientific American, 83(7):109, August 18, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n7/pdf/scientificamerican08181900-109c.pdf.

Anonymous:1900:RPIby

[2477] Anonymous. Recently patented inventions. Scientific American, 83 (8):124–125, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
124.pdf.

Anonymous:1900:BPby

[2478] Anonymous. Business and personal. Scientific American, 83(8):125, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n8/pdf/scientificamerican08251900-125.pdf.

Anonymous:1900:NQby

[2479] Anonymous. Notes and queries. Scientific American, 83(8):125, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n8/pdf/scientificamerican08251900-125a.pdf.

Anonymous:1900:BNBbq

[2480] Anonymous. Books: New books, etc. Scientific American, 83 (8):125, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
125b.pdf.

Anonymous:1900:IIbz

[2481] Anonymous. Index of inventions. Scientific American, 83(8): 125–127, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
125d.pdf.

Anonymous:1900:PNPby

[2482] Anonymous. Patent notes: Patents. Scientific American, 83(8):127, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n8/pdf/scientificamerican08251900-127.pdf.

Anonymous:1900:PNYb

[2483] Anonymous. Progress of the New York Rapid Transit Tunnel. Scientific American, 83(8):114, August 25, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v83/
n8/pdf/scientificamerican08251900-114.pdf.

Anonymous:1900:HLL

[2484] Anonymous. High-lift locks for the Erie Canal. Scientific American, 83(8):114, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
114a.pdf.

Anonymous:1900:RAP

[2485] Anonymous. Recent armor plate bids. Scientific American, 83 (8):114, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
114b.pdf.

Anonymous:1900:LAR

[2486] Anonymous. Lowering of the Atlantic record. Scientific American, 83(8):114, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
114c.pdf.

Anonymous:1900:CATb

[2487] Anonymous. Compressed air traction in New York City. Scientific American, 83(8):114–115, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
114d.pdf.

Adams:1900:LET

[2488] Alton D. Adams. Limits of electric transmission. Scientific American, 83(8):115, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
115.pdf.

Anonymous:1900:CSa

[2489] Anonymous. The chemistry of soot. Scientific American, 83(8):115, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n8/pdf/scientificamerican08251900-115a.pdf.

Anonymous:1900:PENj

[2490] Anonymous. Paris Exposition notes. Scientific American, 83 (8):115, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
115b.pdf.

Anonymous:1900:EXR

[2491] Anonymous. Experiments with X-rays in electrostatic field. Scientific American, 83(8):115, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
115c.pdf.

Anonymous:1900:SHW

[2492] Anonymous. Super-heated water motor for railroad traction. Scientific American, 83(8):116–117, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
116.pdf.

Fawcett:1900:PCC

[2493] Waldon Fawcett. Problems of the Chinese campaign. Scientific American, 83(8):117–118, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
117.pdf.

Anonymous:1900:SSE

[2494] Anonymous. Some stage effects in “Ben Hur”. Scientific American, 83(8):119, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
119.pdf.

Anonymous:1900:PMU

[2495] Anonymous. The Paris Metropolitan Underground Railroad. Scientific American, 83(8):120–l21, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
120.pdf.

Anonymous:1900:HDC

[2496] Anonymous. The Howell disappearing carriage. Scientific American, 83(8):121–122, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
121.pdf.

Anonymous:1900:SRPb

[2497] Anonymous. A simple rotary pump for light service. Scientific American, 83(8):122, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
122.pdf.

Anonymous:1900:GCC

[2498] Anonymous. Gathering and curing crude rubber. Scientific American, 83(8):123–124, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n8/pdf/scientificamerican08251900-
123.pdf.

Anonymous:1900:Iba

[2499] Anonymous. To inventors. Scientific American, 83(8):125, August 25, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n8/pdf/scientificamerican08251900-125c.pdf.

Russell:1900:HS

[2500] Henry Norris Russell. The heavens in September. Scientific American, 83(9):131, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
131.pdf.

Anonymous:1900:Cbi

[2501] Anonymous. Correspondence. Scientific American, 83(9):134, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n9/pdf/scientificamerican09011900-134.pdf.

Anonymous:1900:RPIbz

[2502] Anonymous. Recently patented inventions. Scientific American, 83(9):140, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
140.pdf.

Anonymous:1900:BNBbr

[2503] Anonymous. Books: New books, etc. Scientific American, 83 (9):140, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
140a.pdf.

Anonymous:1900:BPbz

[2504] Anonymous. Business and personal. Scientific American, 83(9): 140, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
140b.pdf.

Anonymous:1900:NQbz

[2505] Anonymous. Notes and queries. Scientific American, 83(9):140– 141, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
140c.pdf.

Anonymous:1900:IIca

[2506] Anonymous. Index of inventions. Scientific American, 83(9): 141–143, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
141a.pdf.

Anonymous:1900:PNPbz

[2507] Anonymous. Patent notes: Patents. Scientific American, 83 (9):143, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
143.pdf.

Anonymous:1900:ARM

[2508] Anonymous. Air resistance to moving bodies. Scientific American, 83(9):130, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
130.pdf.

Anonymous:1900:ESSb

[2509] Anonymous. The element of safety in smokeless powder. Scientific American, 83(9):130, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
130a.pdf.

Anonymous:1900:NGB

[2510] Anonymous. A new gas battery. Scientific American, 83(9): 130, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
130b.pdf.

Anonymous:1900:ICd

[2511] Anonymous. An interesting clock. Scientific American, 83(9): 130, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
130c.pdf.

Anonymous:1900:ICWb

[2512] Anonymous. Inquiry concerning water tube boilers. Scientific American, 83(9):130, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
130d.pdf.

Anonymous:1900:PC

[2513] Anonymous. Polyglot China. Scientific American, 83(9):131, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n9/pdf/scientificamerican09011900-131a.pdf.

Anonymous:1900:DPJ

[2514] Anonymous. The death of Prof. J. E. Keeler. Scientific American, 83(9):131, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
131b.pdf.

Anonymous:1900:APE

[2515] Anonymous. Awards at the Paris Exposition. Scientific American, 83(9):131, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
131c.pdf.

Anonymous:1900:LSP

[2516] Anonymous. Locomotive snow plows in Europe. Scientific American, 83(9):132, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
132.pdf.

Anonymous:1900:SPSb

[2517] Anonymous. The silk-producing spider of Madagascar. Scientific American, 83(9):133, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
133.pdf.

Anonymous:1900:FCBc

[2518] Anonymous. First class battleship “Kentucky”. Scientific American, 83(9):134–135, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
134a.pdf.

Anonymous:1900:SO

[2519] Anonymous. Salvage of the “Oregon”. Scientific American, 83 (9):136, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
136.pdf.

Anonymous:1900:BC

[2520] Anonymous. The Brooks Comet. Scientific American, 83(9): 136, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
136a.pdf.

Anonymous:1900:PEI

[2521] Anonymous. Paris Exposition — interior of the Petit Palais. Scientific American, 83(9):137, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
137.pdf.

Anonymous:1900:NLE

[2522] Anonymous. Nernst lamp at the Exposition. Scientific American, 83(9):138–139, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
138.pdf.

Anonymous:1900:LRC

[2523] Anonymous. Lumber rafts on the Columbia River. Scientific American, 83(9):139, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v83/n9/pdf/scientificamerican09011900-
139.pdf.

Anonymous:1900:Ibb

[2524] Anonymous. To inventors. Scientific American, 83(9):141, September 1, 1900. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v83/n9/pdf/scientificamerican09011900-141.pdf.

Anonymous:1901:RPIa

[2525] Anonymous. Recently patented inventions. Scientific American, 84(1):12, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n1/pdf/scientificamerican01051901-
12.pdf.

Anonymous:1901:BPa

[2526] Anonymous. Business and personal. Scientific American, 84(1):13, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n1/pdf/scientificamerican01051901-13.pdf.

Anonymous:1901:NQa

[2527] Anonymous. Notes and queries. Scientific American, 84(1):13, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n1/pdf/scientificamerican01051901-13a.pdf.

Anonymous:1901:IIa

[2528] Anonymous. Index of inventions. Scientific American, 84(1):13, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n1/pdf/scientificamerican01051901-13b.pdf.

Anonymous:1901:PNPa

[2529] Anonymous. Patent notes: Patents. Scientific American, 84(1):15, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n1/pdf/scientificamerican01051901-15.pdf.

Anonymous:1901:CRC

[2530] Anonymous. California raisin culture. Scientific American, 84(1):3, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n1/pdf/scientificamerican01051901-3.pdf.

Anonymous:1901:RY

[2531] Anonymous. Retrospect of the year 1900. Scientific American, 84(1):4–5, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n1/pdf/scientificamerican01051901-
4.pdf.

Anonymous:1901:PFA

[2532] Anonymous. A pyrotechnic fire-alarm. Scientific American, 84(1):6, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n1/pdf/scientificamerican01051901-6.pdf.

Anonymous:1901:MCW

[2533] Anonymous. Method of constructing wooden water-pipe. Scientific American, 84(1):6, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n1/pdf/scientificamerican01051901-
6a.pdf.

Anonymous:1901:WC

[2534] Anonymous. A wonderful clock. Scientific American, 84(1):7, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n1/pdf/scientificamerican01051901-7.pdf.

Peckham:1901:NIC

[2535] W. C. Peckham. A new induction coil. Scientific American, 84(1):7, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n1/pdf/scientificamerican01051901-7a.pdf.

Anonymous:1901:SIE

[2536] Anonymous. Some interesting experiments in dirigible airships. Scientific American, 84(1):10, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n1/pdf/scientificamerican01051901-
10.pdf.

Anonymous:1901:FAC

[2537] Anonymous. The French armored cruiser “Montcalm”. Scientific American, 84(1):10, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n1/pdf/scientificamerican01051901-
10a.pdf.

Anonymous:1901:SMSa

[2538] Anonymous. A safety military spy-glass. Scientific American, 84 (1):10–11, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n1/pdf/scientificamerican01051901-
10b.pdf.

Benjamin:1901:BAB

[2539] Marcus Benjamin. The bolometer or actinic balance. Scientific American, 84(1):11–12, January 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n1/pdf/scientificamerican01051901-
11.pdf.

Anonymous:1901:RPIb

[2540] Anonymous. Recently patented inventions. Scientific American, 84(10):155, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
155.pdf.

Anonymous:1901:NQb

[2541] Anonymous. Notes and queries. Scientific American, 84(10):155–156, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n10/pdf/scientificamerican03091901-155b.pdf.

Anonymous:1901:IIb

[2542] Anonymous. Index of inventions. Scientific American, 84(10):156, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n10/pdf/scientificamerican03091901-156.pdf.

Anonymous:1901:PNPb

[2543] Anonymous. Patent notes: Patents. Scientific American, 84(10):159, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n10/pdf/scientificamerican03091901-159.pdf.

Anonymous:1901:NSP

[2544] Anonymous. The new star in Perseus. Scientific American, 84(10):146, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n10/pdf/scientificamerican03091901-146.pdf.

Anonymous:1901:ITP

[2545] Anonymous. An important telephone patent decision. Scientific American, 84(10):146, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
146a.pdf.

Anonymous:1901:SSWa

[2546] Anonymous. The Slaby system of wireless duplex telegraphy. Scientific American, 84(10):146–147, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
146b.pdf.

Anonymous:1901:TIPa

[2547] Anonymous. A triumph of ingenuity and a patent with a history. Scientific American, 84(10):147, March 9, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n10/pdf/scientificamerican03091901-147.pdf.

Anonymous:1901:NPPa

[2548] Anonymous. A new permanent phonograph record. Scientific American, 84(10):147, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
147a.pdf.

Anonymous:1901:PMP

[2549] Anonymous. The Pollok Memorial Prize. Scientific American, 84(10):147, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
147b.pdf.

Anonymous:1901:IGA

[2550] Anonymous. An improved gasoline automobile. Scientific American, 84(10):148, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
148.pdf.

Anonymous:1901:SGT

[2551] Anonymous. San Gimignano and its towers. Scientific American, 84(10):148, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
148a.pdf.

Holder:1901:TPS

[2552] Charles P. Holder. Trees and plants of Southern California. Scientific American, 84(10):149, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
149.pdf.

Anonymous:1901:PTM

[2553] Anonymous. The Paige typesetting machine. Scientific American, 84(10):150–151, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
150.pdf.

Anonymous:1901:SSWb

[2554] Anonymous. The Slaby system of wave-selective wireless duplex telegraphy. Scientific American, 84(10):152–153, March 9, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n10/pdf/scientificamerican03091901-152.pdf.

MClure:1901:AOU

[2555] W. Frank M’Clure. Automatic ore unloader. Scientific American, 84(10):153–154, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
153.pdf.

Anonymous:1901:PLBa

[2556] Anonymous. A pail for live bait. Scientific American, 84(10):154, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n10/pdf/scientificamerican03091901-154.pdf.

Anonymous:1901:VMS

[2557] Anonymous. A valve motion for small high-pressure pumps. Scientific American, 84(10):154, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
154a.pdf.

Anonymous:1901:NBC

[2558] Anonymous. A new book or copy holder for typewriters. Scientific American, 84(10):154, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
154b.pdf.

Anonymous:1901:BPWa

[2559] Anonymous. Business and personal wants. Scientific American, 84(10):155, March 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n10/pdf/scientificamerican03091901-
155a.pdf.

Anonymous:1901:RPIc

[2560] Anonymous. Recently patented inventions. Scientific American, 84(11):171, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
171.pdf.

Anonymous:1901:BNBa

[2561] Anonymous. Books: New books, etc. Scientific American, 84(11):171, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n11/pdf/scientificamerican03161901-171a.pdf.

Anonymous:1901:NQc

[2562] Anonymous. Notes and queries. Scientific American, 84(11):172–173, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n11/pdf/scientificamerican03161901-172.pdf.

Anonymous:1901:IIc

[2563] Anonymous. Index of inventions. Scientific American, 84(11):173, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n11/pdf/scientificamerican03161901-173.pdf.

Anonymous:1901:PNPc

[2564] Anonymous. Patent notes: Patents. Scientific American, 84(11):175, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n11/pdf/scientificamerican03161901-175.pdf.

Anonymous:1901:PNM

[2565] Anonymous. The proposed Naval Memorial Arch. Scientific American, 84(11):162, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
162.pdf.

Anonymous:1901:VIB

[2566] Anonymous. Vast and increasing bulk of the modern steamship. Scientific American, 84(11):162, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
162a.pdf.

Anonymous:1901:DAB

[2567] Anonymous. Decision against the Berliner microphone patent. Scientific American, 84(11):162, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
162b.pdf.

Anonymous:1901:CSa

[2568] Anonymous. Curiosity and science. Scientific American, 84(11): 162–163, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
162c.pdf.

Anonymous:1901:ECR

[2569] Anonymous. Eros, the celestial range finder. Scientific American, 84(11):163, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
163.pdf.

Anonymous:1901:AK

[2570] Anonymous. The automobile in the Klondike. Scientific American, 84(11):163, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
163a.pdf.

Anonymous:1901:EAT

[2571] Anonymous. Experiments with animal tissue. Scientific American, 84(11):163, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
163b.pdf.

Anonymous:1901:LSTa

[2572] Anonymous. The longest subterranean telegraph cable in the world. Scientific American, 84(11):163, March 16, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n11/pdf/scientificamerican03161901-163c.pdf.

Anonymous:1901:MOF

[2573] Anonymous. The manufacture of ostrich feathers. Scientific American, 84(11):164, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
164.pdf.

Anonymous:1901:OBF

[2574] Anonymous. The Old Boonton Forge. Scientific American, 84(11):165, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n11/pdf/scientificamerican03161901-165.pdf.

Anonymous:1901:MPE

[2575] Anonymous. Most powerful express locomotive in the world. Scientific American, 84(11):165–166, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
165a.pdf.

Anonymous:1901:PLA

[2576] Anonymous. The Pitch Lake and the asphalt industry of Trinidad. Scientific American, 84(11):166–167, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
166.pdf.

Anonymous:1901:TRB

[2577] Anonymous. Third rail in the Baltimore Belt Line Tunnel. Scientific American, 84(11):168–169, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
168.pdf.

Holder:1901:SM

[2578] Charles F. Holder. Solar motors. Scientific American, 84(11):169–170, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n11/pdf/scientificamerican03161901-169.pdf.

Anonymous:1901:FSC

[2579] Anonymous. The Fleming safety collar for shafts. Scientific American, 84(11):170, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
170.pdf.

Anonymous:1901:ESPa

[2580] Anonymous. An efficient single-pull ceiling-switch. Scientific American, 84(11):170, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
170a.pdf.

Anonymous:1901:FBC

[2581] Anonymous. A folding bunk for cars and ships. Scientific American, 84(11):170, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
170b.pdf.

Anonymous:1901:BPWb

[2582] Anonymous. Business and personal wants. Scientific American, 84(11):171, March 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n11/pdf/scientificamerican03161901-
171b.pdf.

Anonymous:1901:SNa

[2583] Anonymous. Science notes. Scientific American, 84(12):179, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n12/pdf/scientificamerican03231901-179c.pdf.

Anonymous:1901:Ca

[2584] Anonymous. Correspondence. Scientific American, 84(12):183, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n12/pdf/scientificamerican03231901-183.pdf.

Anonymous:1901:RPId 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
187.pdf.

Anonymous:1901:BNBb

[2586] Anonymous. Books: New books, etc. Scientific American, 84(12):187, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n12/pdf/scientificamerican03231901-187a.pdf.

Anonymous:1901:NQd

[2587] Anonymous. Notes and queries. Scientific American, 84(12):188, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n12/pdf/scientificamerican03231901-188.pdf.

Anonymous:1901:IId

[2588] Anonymous. Index of inventions. Scientific American, 84(12):188, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n12/pdf/scientificamerican03231901-188a.pdf.

Anonymous:1901:PNPd

[2589] Anonymous. Patent notes: Patents. Scientific American, 84(12):191, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n12/pdf/scientificamerican03231901-191.pdf.

Anonymous:1901:AEP

[2590] Anonymous. American engineering progress. Scientific American, 84(12):178, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
178.pdf.

Anonymous:1901:CVM

[2591] Anonymous. Current of 3,000 volts at the motors. Scientific American, 84(12):178, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
178a.pdf.

Anonymous:1901:CST

[2592] Anonymous. Construction of the simplon tunnel. Scientific American, 84(12):178–179, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
178b.pdf.

Anonymous:1901:TIPb

[2593] Anonymous. Three important patent decisions by the United States Court of Appeals at New York. Scientific American, 84(12):179, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n12/pdf/scientificamerican03231901-179.pdf.

Anonymous:1901:GE

[2594] Anonymous. The Glasgow Exhibition of 1901. Scientific American, 84(12):179, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
179a.pdf.

Anonymous:1901:RCDa

[2595] Anonymous. Resignation of Commissioner Duell. Scientific American, 84(12):179, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
179b.pdf.

Anonymous:1901:PPC

[2596] Anonymous. Production of phosphorus and chlorides of carbon by means of electric furnaces. Scientific American, 84(12):180, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n12/pdf/scientificamerican03231901-180.pdf.

Anonymous:1901:BPPa

[2597] Anonymous. The Burry page-printing telegraph. Scientific American, 84(12):181–182, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
181.pdf.

Anonymous:1901:TSP

[2598] Anonymous. Transport service to the Philippines — I. Scientific American, 84(12):182–183, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
182.pdf.

Anonymous:1901:PRA

[2599] Anonymous. A peculiar railroad accident. Scientific American, 84(12):184, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
184.pdf.

Beard:1901:TCT

[2600] J. Carter Beard. Three characteristic types of American dinosaurs. Scientific American, 84(12):184–185, March 23, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n12/pdf/scientificamerican03231901-184a.pdf.

Stewart:1901:NSR

[2601] J. A. Stewart. The New Street Railway Tunnel under Boston Harbor. Scientific American, 84(12):185–186, March 23, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n12/pdf/scientificamerican03231901-185.pdf.

Anonymous:1901:NPS

[2602] Anonymous. A novel pipe and stud wrench. Scientific American, 84(12):186, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
186.pdf.

Anonymous:1901:AWM

[2603] Anonymous. An automatic weighing-machine for liquids. Scientific American, 84(12):186, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
186a.pdf.

Anonymous:1901:BPWc

[2604] Anonymous. Business and personal wants. Scientific American, 84(12):187, March 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n12/pdf/scientificamerican03231901-
187b.pdf.

Anonymous:1901:RPIe

[2605] Anonymous. Recently patented inventions. Scientific American, 84 (13):203–204, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
203.pdf.

Anonymous:1901:NQe

[2606] Anonymous. Notes and queries. Scientific American, 84(13):204, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-204a.pdf.

Anonymous:1901:BNBc

[2607] Anonymous. Books: New books, etc. Scientific American, 84(13):204, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-204b.pdf.

Anonymous:1901:IIe

[2608] Anonymous. Index of inventions. Scientific American, 84(13):204, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-204c.pdf.

Anonymous:1901:PNPe

[2609] Anonymous. Patent notes: Patents. Scientific American, 84(13):207, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-207.pdf.

Anonymous:1901:GSL

[2610] Anonymous. The Ganz system and the London Electric Railways. Scientific American, 84(13):194, March 30, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n13/pdf/scientificamerican03301901-194.pdf.

Anonymous:1901:CYD

[2611] Anonymous. Cup-yacht designing. Scientific American, 84(13):194, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-194a.pdf.

Anonymous:1901:YF

[2612] Anonymous. Yellow fever. Scientific American, 84(13):194, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-194b.pdf.

Anonymous:1901:NMA

[2613] Anonymous. New method of aerial telegraphy. Scientific American, 84(13):194–195, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
194c.pdf.

Anonymous:1901:BTI

[2614] Anonymous. Builders’ trial of the “Illinois”. Scientific American, 84(13):195, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
195.pdf.

Anonymous:1901:POR

[2615] Anonymous. The Patent Office — resignation of the Commissioner of Patents. Scientific American, 84(13):195, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-195a.pdf.

Anonymous:1901:RAM

[2616] Anonymous. Russian Army and Marine. Scientific American, 84(13):195, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
195b.pdf.

Anonymous:1901:ROI

[2617] Anonymous. The remains of an old Indian village. Scientific American, 84(13):195, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
195c.pdf.

Anonymous:1901:MMM

[2618] Anonymous. The mutoscope and machinery in motion. Scientific American, 84(13):196, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
196.pdf.

Anonymous:1901:T

[2619] Anonymous. The topophone. Scientific American, 84(13):197–198, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-197.pdf.

Anonymous:1901:IRY

[2620] Anonymous. The international racing-yacht “Independence”. Scientific American, 84(13):198–199, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
198.pdf.

Anonymous:1901:NMM

[2621] Anonymous. A new method of manufacturing steel pipe. Scientific American, 84(13):200–201, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
200.pdf.

Anonymous:1901:PRE

[2622] Anonymous. Porter rotary engine. Scientific American, 84(13):201, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-201.pdf.

Anonymous:1901:NSH

[2623] Anonymous. A novel self-heating sad-iron. Scientific American, 84(13):202, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
202.pdf.

Holder:1901:SIP

[2624] Charles F. Holder. A singular industry in the poisonous insects of California. Scientific American, 84(13):202–203, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n13/pdf/scientificamerican03301901-202a.pdf.

Anonymous:1901:BPWd

[2625] Anonymous. Business and personal wants. Scientific American, 84(13):204, March 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n13/pdf/scientificamerican03301901-
204.pdf.

Russell:1901:HA

[2626] Henry Norris Russell. The heavens in April. Scientific American, 84(14):211, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
211.pdf.

Anonymous:1901:Cb

[2627] Anonymous. Correspondence. Scientific American, 84(14):215, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n14/pdf/scientificamerican04061901-215.pdf.

Anonymous:1901:RPIf

[2628] Anonymous. Recently patented inventions. Scientific American, 84(14):219, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
219.pdf.

Anonymous:1901:NQf

[2629] Anonymous. Notes and queries. Scientific American, 84(14):219, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n14/pdf/scientificamerican04061901-219b.pdf.

Anonymous:1901:IIf

[2630] Anonymous. Index of inventions. Scientific American, 84(14):219–220, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n14/pdf/scientificamerican04061901-219d.pdf.

Anonymous:1901:PNPf

[2631] Anonymous. Patent notes: Patents. Scientific American, 84(14):223, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n14/pdf/scientificamerican04061901-223.pdf.

Anonymous:1901:PNE

[2632] Anonymous. Prospects of the new Erie Canal. Scientific American, 84(14):210, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
210.pdf.

Anonymous:1901:WTB

[2633] Anonymous. Water-tube boilers in the British Navy. Scientific American, 84(14):210, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
210a.pdf.

Anonymous:1901:LFH

[2634] Anonymous. Longitudinal framing for the Herreshoff Cup-Yacht. Scientific American, 84(14):210, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
210b.pdf.

Anonymous:1901:OAT

[2635] Anonymous. Our advancing trade. Scientific American, 84(14):210, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n14/pdf/scientificamerican04061901-210c.pdf.

Anonymous:1901:SSC

[2636] Anonymous. Sylviculture and the Suez Canal. Scientific American, 84(14):210–211, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
210d.pdf.

Anonymous:1901:ANY

[2637] Anonymous. Automobiles in New York. Scientific American, 84 (14):211, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
211a.pdf.

Anonymous:1901:MDH

[2638] Anonymous. Method of detecting hypo in photographic work. Scientific American, 84(14):211, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
211b.pdf.

Anonymous:1901:EIF

[2639] Anonymous. Electrically illuminated fountains. Scientific American, 84(14):212, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
212.pdf.

Anonymous:1901:NRBa

[2640] Anonymous. The new Russian battleship “Czarevitch” as she will appear when completed for sea. Scientific American, 84(14):212–213, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n14/pdf/scientificamerican04061901-212a.pdf.

Anonymous:1901:SCR

[2641] Anonymous. A sugar cane reaper wanted in Hawaii. Scientific American, 84(14):213–214, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
213.pdf.

Anonymous:1901:TSN

[2642] Anonymous. Training school for New York Streetcar Motormen. Scientific American, 84(14):214, April 6, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n14/pdf/scientificamerican04061901-214.pdf.

Anonymous:1901:OIT

[2643] Anonymous. The Omnigraph — an instrument for teaching telegraphy. Scientific American, 84(14):216, April 6, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n14/pdf/scientificamerican04061901-216.pdf.

Anonymous:1901:HEC

[2644] Anonymous. Hydraulic electric-car brake. Scientific American, 84 (14):216–217, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
216a.pdf.

Holder:1901:RSD

[2645] Charles F. Holder. A remarkable salt deposit. Scientific American, 84(14):217, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
217.pdf.

Anonymous:1901:IPW

[2646] Anonymous. An improved pipe-wrench. Scientific American, 84 (14):218, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
218.pdf.

Anonymous:1901:NWO

[2647] Anonymous. A new wickless oil burner. Scientific American, 84(14):218, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n14/pdf/scientificamerican04061901-218a.pdf.

Anonymous:1901:WCGa

[2648] Anonymous. The Wentworth cotton gin. Scientific American, 84 (14):218, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
218b.pdf.

Anonymous:1901:BPWe

[2649] Anonymous. Business and personal wants. Scientific American, 84(14):219, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n14/pdf/scientificamerican04061901-
219a.pdf.

Anonymous:1901:Ia

[2650] Anonymous. To inventors. Scientific American, 84(14):219, April 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n14/pdf/scientificamerican04061901-219c.pdf.

Anonymous:1901:RPIg

[2651] Anonymous. Recently patented inventions. Scientific American, 84(15):235, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
235.pdf.

Anonymous:1901:NQg

[2652] Anonymous. Notes and queries. Scientific American, 84(15):235–236, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n15/pdf/scientificamerican04131901-235b.pdf.

Anonymous:1901:IIg

[2653] Anonymous. Index of inventions. Scientific American, 84(15):236, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n15/pdf/scientificamerican04131901-236.pdf.

Anonymous:1901:PNPg

[2654] Anonymous. Patent notes: Patents. Scientific American, 84(15):239, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n15/pdf/scientificamerican04131901-239.pdf.

Anonymous:1901:BRTa

[2655] Anonymous. The Brooklyn Rapid Transit Extension. Scientific American, 84(15):226, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
226.pdf.

Anonymous:1901:SB

[2656] Anonymous. A step backward. Scientific American, 84(15):226, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n15/pdf/scientificamerican04131901-226a.pdf.

Anonymous:1901:LVA

[2657] Anonymous. Lake vessels for the Atlantic trade. Scientific American, 84(15):226, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
226b.pdf.

Proctor:1901:GEE

[2658] Mary Proctor. Government Eclipse Expedition. Scientific American, 84(15):226–227, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
226c.pdf.

Anonymous:1901:EPM

[2659] Anonymous. Electrical protection of the mails. Scientific American, 84(15):227, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
227.pdf.

Anonymous:1901:NCP

[2660] Anonymous. The new Commissioner of Patents. Scientific American, 84(15):227, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
227a.pdf.

Anonymous:1901:NYBa

[2661] Anonymous. The New York Botanical Garden Museum. Scientific American, 84(15):227, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
227b.pdf.

Anonymous:1901:LPM

[2662] Anonymous. Launch of the Pacific mail liner “Korea”. Scientific American, 84(15):228, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
228.pdf.

Anonymous:1901:IBE

[2663] Anonymous. An interesting boiler explosion. Scientific American, 84(15):228–229, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
228a.pdf.

Weeks:1901:RDN

[2664] Arthur B. Weeks. Recent developments at the Niagara Falls Power Plant. Scientific American, 84(15):229, April 13, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n15/pdf/scientificamerican04131901-229.pdf.

Anonymous:1901:RHT

[2665] Anonymous. Rodney Hunt Turbine Plant at the Great Northern Paper Company’s mill. Scientific American, 84(15):230–231, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n15/pdf/scientificamerican04131901-230.pdf.

Anonymous:1901:BV

[2666] Anonymous. Balloons at Vincennes. Scientific American, 84(15):232, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n15/pdf/scientificamerican04131901-232.pdf.

Anonymous:1901:NPD

[2667] Anonymous. A new pile driver. Scientific American, 84(15):232–233, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n15/pdf/scientificamerican04131901-232a.pdf.

Anonymous:1901:FYD

[2668] Anonymous. Fifteen years’ development of the 90-foot racing yacht. Scientific American, 84(15):233–234, April 13, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n15/pdf/scientificamerican04131901-233.pdf.

Anonymous:1901:IL

[2669] Anonymous. An improvement in lubricators. Scientific American, 84(15):234, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
234.pdf.

Anonymous:1901:BPWf

[2670] Anonymous. Business and personal wants. Scientific American, 84(15):235, April 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n15/pdf/scientificamerican04131901-
235a.pdf.

Anonymous:1901:SNb

[2671] Anonymous. Science notes. Scientific American, 84(16):243, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n16/pdf/scientificamerican04201901-243a.pdf.

Anonymous:1901:Cc

[2672] Anonymous. Correspondence. Scientific American, 84(16):247, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n16/pdf/scientificamerican04201901-247.pdf.

Anonymous:1901:RPIh

[2673] Anonymous. Recently patented inventions. Scientific American, 84(16):251, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
251.pdf.

Anonymous:1901:BPWg

[2674] Anonymous. Business and personal wants. Scientific American, 84(16):251–252, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
251a.pdf.

Anonymous:1901:NQh

[2675] Anonymous. Notes and queries. Scientific American, 84(16):252, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n16/pdf/scientificamerican04201901-252.pdf.

Anonymous:1901:IIh

[2676] Anonymous. Index of inventions. Scientific American, 84(16):252, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n16/pdf/scientificamerican04201901-252a.pdf.

Anonymous:1901:PNPh

[2677] Anonymous. Patent notes: Patents. Scientific American, 84(16):253, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n16/pdf/scientificamerican04201901-253.pdf.

Anonymous:1901:ECM

[2678] Anonymous. The Erie Canal muddle. Scientific American, 84(16):242, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n16/pdf/scientificamerican04201901-242.pdf.

Anonymous:1901:ACD

[2679] Anonymous. Application of the coherer to detection of storms. Scientific American, 84(16):242, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
242a.pdf.

Anonymous:1901:IPD

[2680] Anonymous. Improved process of duplicating phonograph records. Scientific American, 84(16):242, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
242b.pdf.

Anonymous:1901:NPG

[2681] Anonymous. New products in the glass industry. Scientific American, 84(16):242, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
242c.pdf.

Anonymous:1901:NCC

[2682] Anonymous. New chemical compounds. Scientific American, 84 (16):242–243, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
242d.pdf.

Anonymous:1901:LCF

[2683] Anonymous. Local conditions favorable to the cotton industry of the South. Scientific American, 84(16):243, April 20, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n16/pdf/scientificamerican04201901-243.pdf.

Anonymous:1901:MGP

[2684] Anonymous. A moving grazing pen. Scientific American, 84(16):244, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n16/pdf/scientificamerican04201901-244.pdf.

Anonymous:1901:SWS

[2685] Anonymous. Stern-wheel steamers on the Mississippi River. Scientific American, 84(16):244, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
244a.pdf.

Peckham:1901:PSX

[2686] William C. Peckham. The present state of X-ray work. Scientific American, 84(16):245, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
245.pdf.

Dailey:1901:EG

[2687] Howard B. Dailey. Electrical gyroscopes. Scientific American, 84 (16):245–246, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
245a.pdf.

Anonymous:1901:CNEa

[2688] Anonymous. Construction of the new East River Bridge. Scientific American, 84(16):246–247, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
246.pdf.

Fawcett:1901:MBO

[2689] Waldon Fawcett. Modern British ordnance. Scientific American, 84(16):248–249, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
248.pdf.

Collins:1901:MLP

[2690] J. H. Collins. Making large plaster casts. Scientific American, 84(16):249–250, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
249.pdf.

Anonymous:1901:NHD

[2691] Anonymous. A novel haystacker and derrick. Scientific American, 84(16):250, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
250.pdf.

Anonymous:1901:PII

[2692] Anonymous. A pen with increased ink capacity. Scientific American, 84(16):250, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
250a.pdf.

Anonymous:1901:CIA

[2693] Anonymous. A completely inclosed arc lamp of small size. Scientific American, 84(16):250, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
250b.pdf.

Anonymous:1901:MCBa

[2694] Anonymous. A machine for cleaning blackboard erasers. Scientific American, 84(16):250, April 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n16/pdf/scientificamerican04201901-
250c.pdf.

Anonymous:1901:RPIi

[2695] Anonymous. Recently patented inventions. Scientific American, 84(17):267–268, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
267.pdf.

Anonymous:1901:BNBd

[2696] Anonymous. Books: New books, etc. Scientific American, 84(17):268, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n17/pdf/scientificamerican04271901-268.pdf.

Anonymous:1901:NQi

[2697] Anonymous. Notes and queries. Scientific American, 84(17):268, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n17/pdf/scientificamerican04271901-268b.pdf.

Anonymous:1901:IIi

[2698] Anonymous. Index of inventions. Scientific American, 84(17):268, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n17/pdf/scientificamerican04271901-268c.pdf.

Anonymous:1901:PNPi

[2699] Anonymous. Patent notes: Patents. Scientific American, 84(17):271, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n17/pdf/scientificamerican04271901-271.pdf.

Anonymous:1901:LFS

[2700] Anonymous. The locomotive of the future — a suggestion. Scientific American, 84(17):258, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
258.pdf.

Anonymous:1901:BHH

[2701] Anonymous. Behr on his high-speed monorail system. Scientific American, 84(17):258, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
258a.pdf.

Anonymous:1901:SBN

[2702] Anonymous. Submarines for the British Navy. Scientific American, 84(17):258–259, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
258b.pdf.

Benjamin:1901:NASa

[2703] Marcus Benjamin. The National Academy of Sciences. Scientific American, 84(17):259, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
259.pdf.

Anonymous:1901:OBT

[2704] Anonymous. The Ordnance Board’s test of thorite. Scientific American, 84(17):259, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
259a.pdf.

Anonymous:1901:DPH

[2705] Anonymous. Death of Prof. Henry A. Rowland. Scientific American, 84(17):259, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
259b.pdf.

Anonymous:1901:DRP

[2706] Anonymous. Death of Richard P. Rothwell. Scientific American, 84(17):259, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
259c.pdf.

Anonymous:1901:NWC

[2707] Anonymous. A new weather cannon. Scientific American, 84(17):260, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n17/pdf/scientificamerican04271901-260.pdf.

Stout:1901:HMG

[2708] Will. B. Stout. How to make a gramophone. Scientific American, 84(17):260, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
260a.pdf.

Anonymous:1901:PKW

[2709] Anonymous. Parsons 1,000-k. w. turbine and alternator. Scientific American, 84(17):260–261, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
260b.pdf.

Anonymous:1901:RSS

[2710] Anonymous. Raising a submerged steamship. Scientific American, 84(17):261, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
261.pdf.

Anonymous:1901:ATS

[2711] Anonymous. Army Transport Service in the Philippines — II. Scientific American, 84(17):262–263, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
262.pdf.

Anonymous:1901:IP

[2712] Anonymous. An improved plow. Scientific American, 84(17):263, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n17/pdf/scientificamerican04271901-263.pdf.

Anonymous:1901:ABF

[2713] Anonymous. An automatic boiler-feeder. Scientific American, 84 (17):263, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
263a.pdf.

Anonymous:1901:WHL

[2714] Anonymous. A wheat hospital on lake superior. Scientific American, 84(17):264, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
264.pdf.

Beard:1901:SNF

[2715] J. Carter Beard. Some new features in ant life. Scientific American, 84(17):265–266, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
265.pdf.

Anonymous:1901:CHL

[2716] Anonymous. The Cooper Hewitt lamp. Scientific American, 84 (17):266–267, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
266.pdf.

Anonymous:1901:BPWh

[2717] Anonymous. Business and personal wants. Scientific American, 84(17):268, April 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n17/pdf/scientificamerican04271901-
268a.pdf.

Russell:1901:HMa

[2718] Henry Norris Russell. The heavens in May. Scientific American, 84(18):279, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
279a.pdf.

Anonymous:1901:RPIj

[2719] Anonymous. Recently patented inventions. Scientific American, 84(18):283, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
283.pdf.

Anonymous:1901:NQj

[2720] Anonymous. Notes and queries. Scientific American, 84(18):283–284, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n18/pdf/scientificamerican05041901-283b.pdf.

Anonymous:1901:IIj

[2721] Anonymous. Index of inventions. Scientific American, 84(18):284, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n18/pdf/scientificamerican05041901-284.pdf.

Anonymous:1901:PNPj

[2722] Anonymous. Patent notes: Patents. Scientific American, 84(18):287, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n18/pdf/scientificamerican05041901-287.pdf.

Anonymous:1901:NRBb

[2723] Anonymous. The North River Bridge Bill. Scientific American, 84(18):274, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
274.pdf.

Anonymous:1901:ROS

[2724] Anonymous. Rerolling old steel rails. Scientific American, 84(18):274, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n18/pdf/scientificamerican05041901-274a.pdf.

Anonymous:1901:NCV

[2725] Anonymous. Nitro cellulose vs nitroglycerine smokeless powders. Scientific American, 84(18):274, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
274b.pdf.

Anonymous:1901:SR

[2726] Anonymous. The Smithsonian Report. Scientific American, 84 (18):274–275, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
274c.pdf.

Anonymous:1901:BZE

[2727] Anonymous. The Baldwin–Ziegler Expedition to Franz Josef Land. Scientific American, 84(18):275, May 4, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n18/pdf/scientificamerican05041901-275.pdf.

Anonymous:1901:NAC

[2728] Anonymous. The new Assistant Commissioner of Patents. Scientific American, 84(18):275, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
275a.pdf.

Anonymous:1901:OPAa

[2729] Anonymous. Opening of the Pan-American Exposition. Scientific American, 84(18):275, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
275b.pdf.

Anonymous:1901:PMC

[2730] Anonymous. A perpetual motion clock. Scientific American, 84(18):276, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n18/pdf/scientificamerican05041901-276.pdf.

Anonymous:1901:LWS

[2731] Anonymous. Launch of the White Star liner “Celtic”. Scientific American, 84(18):276–277, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
276a.pdf.

Anonymous:1901:MPP

[2732] Anonymous. The Murray page-printing telegraph. Scientific American, 84(18):277–278, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
277.pdf.

Anonymous:1901:NIW

[2733] Anonymous. The Nile Irrigation Works. Scientific American, 84 (18):279, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
279.pdf.

Anonymous:1901:TSB

[2734] Anonymous. Turkish submarine boat of the Nordenfeldt type. Scientific American, 84(18):280, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
280.pdf.

Anonymous:1901:SGA

[2735] Anonymous. Siberia in the Glacial Age. Scientific American, 84(18):280–281, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
280a.pdf.

Anonymous:1901:HFM

[2736] Anonymous. Hofman’s flying machine. Scientific American, 84(18):281, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n18/pdf/scientificamerican05041901-281.pdf.

Anonymous:1901:SCMa

[2737] Anonymous. A simple car mover. Scientific American, 84(18):282, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n18/pdf/scientificamerican05041901-282.pdf.

Anonymous:1901:CMW

[2738] Anonymous. Combination metal and wood-working machine. Scientific American, 84(18):282, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
282a.pdf.

Anonymous:1901:NFW

[2739] Anonymous. A novel fifth-wheel. Scientific American, 84(18):282, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n18/pdf/scientificamerican05041901-282b.pdf.

Anonymous:1901:BPWi

[2740] Anonymous. Business and personal wants. Scientific American, 84(18):283, May 4, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n18/pdf/scientificamerican05041901-
283a.pdf.

Anonymous:1901:SNc

[2741] Anonymous. Science notes. Scientific American, 84(19):291, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n19/pdf/scientificamerican05111901-291a.pdf.

Anonymous:1901:Cd

[2742] Anonymous. Correspondence. Scientific American, 84(19):294, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n19/pdf/scientificamerican05111901-294.pdf.

Anonymous:1901:RPIk

[2743] Anonymous. Recently patented inventions. Scientific American, 84(19):299, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
299.pdf.

Anonymous:1901:NQk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n19/pdf/scientificamerican05111901-299b.pdf.

Anonymous:1901:BNBe

[2745] Anonymous. Books: New books, etc. Scientific American, 84(19):300, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n19/pdf/scientificamerican05111901-300.pdf.

Anonymous:1901:IIk

[2746] Anonymous. Index of inventions. Scientific American, 84(19):300, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n19/pdf/scientificamerican05111901-300a.pdf.

Anonymous:1901:PNPk

[2747] Anonymous. Patent notes: Patents. Scientific American, 84(19):303, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n19/pdf/scientificamerican05111901-303.pdf.

Anonymous:1901:CDC

[2748] Anonymous. Chicago Drainage Canal complications. Scientific American, 84(19):290, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
290.pdf.

Anonymous:1901:LAB

[2749] Anonymous. Liquid air as a blasting agent. Scientific American, 84(19):290, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
290a.pdf.

Anonymous:1901:TBR

[2750] Anonymous. Train brake results of high speed. Scientific American, 84(19):290, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
290b.pdf.

Anonymous:1901:WCC

[2751] Anonymous. Waterways and canals of Canada. Scientific American, 84(19):290, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
290c.pdf.

Anonymous:1901:TTL

[2752] Anonymous. Transportation of trains on Lake Baikal. Scientific American, 84(19):290, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
290d.pdf.

Anonymous:1901:TCa

[2753] Anonymous. Thorium compounds. Scientific American, 84(19): 290–291, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
290e.pdf.

Anonymous:1901:SFF

[2754] Anonymous. Some figures of the first cost and operation of Southern cotton mills. Scientific American, 84(19):291, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n19/pdf/scientificamerican05111901-291.pdf.

Anonymous:1901:TDS

[2755] Anonymous. A tool for driving screw-eyes. Scientific American, 84(19):292, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
292.pdf.

Anonymous:1901:AFV

[2756] Anonymous. Automatic frost valve for water-pipes. Scientific American, 84(19):292, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
292a.pdf.

Anonymous:1901:NIE

[2757] Anonymous. A new igniter for explosion-engines. Scientific American, 84(19):292, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
292b.pdf.

Anonymous:1901:LSI

[2758] Anonymous. The launch of “Shamrock II”. Scientific American, 84(19):293, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
293.pdf.

Anonymous:1901:CYC

[2759] Anonymous. Construction of the yacht “Constitution”. Scientific American, 84(19):295, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
295.pdf.

Simonsen:1901:HMI

[2760] Parker Simonsen. How to make an iron rose. Scientific American, 84(19):296, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
296.pdf.

Anonymous:1901:DT

[2761] Anonymous. A disappearing town. Scientific American, 84(19):296–297, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n19/pdf/scientificamerican05111901-296a.pdf.

Mindeleff:1901:CDP

[2762] Cosmos Mindeleff. A cliff-dwelling park in Colorado. Scientific American, 84(19):297–298, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
297.pdf.

Aaron:1901:BW

[2763] S Frank Aaron. The bean weevils. Scientific American, 84(19):298, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n19/pdf/scientificamerican05111901-298.pdf.

Anonymous:1901:BPWj

[2764] Anonymous. Business and personal wants. Scientific American, 84(19):299, May 11, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n19/pdf/scientificamerican05111901-
299a.pdf.

Anonymous:1901:Ce

[2765] Anonymous. Correspondence. Scientific American, 84(2):26–27, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n2/pdf/scientificamerican01121901-26.pdf.

Anonymous:1901:RPIl

[2766] Anonymous. Recently patented inventions. Scientific American, 84(2):28, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
28.pdf.

Anonymous:1901:BPb

[2767] Anonymous. Business and personal. Scientific American, 84(2):28, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n2/pdf/scientificamerican01121901-28a.pdf.

Anonymous:1901:NQl

[2768] Anonymous. Notes and queries. Scientific American, 84(2):28, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n2/pdf/scientificamerican01121901-28b.pdf.

Anonymous:1901:BNBf (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
28c.pdf.

Anonymous:1901:IIl

[2770] Anonymous. Index of inventions. Scientific American, 84(2):28, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n2/pdf/scientificamerican01121901-28d.pdf.

Anonymous:1901:EST

[2771] Anonymous. The efficiency of the steam turbine. Scientific American, 84(2):18, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
18.pdf.

Anonymous:1901:SSBa

[2772] Anonymous. The Ship-Subsidy Bill. Scientific American, 84(2):18, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n2/pdf/scientificamerican01121901-18a.pdf.

Hovey:1901:TAM

[2773] Edmund O. Hovey. The Thirteenth Annual Meeting of the Geological Society of America. Scientific American, 84(2):19, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n2/pdf/scientificamerican01121901-19.pdf.

Anonymous:1901:PISa

[2774] Anonymous. Proposed improvements at the Sault Ste. Marie Canal. Scientific American, 84(2):19, January 12, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n2/pdf/scientificamerican01121901-19a.pdf.

Anonymous:1901:MPL

[2775] Anonymous. Moving picture litigation. Scientific American, 84 (2):19, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
19b.pdf.

Anonymous:1901:DLA

[2776] Anonymous. Death of Lord Armstrong. Scientific American, 84 (2):19, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
19c.pdf.

Anonymous:1901:NQFa

[2777] Anonymous. A new quick-firing gun. Scientific American, 84(2):20, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n2/pdf/scientificamerican01121901-20.pdf.

Anonymous:1901:IHIa

[2778] Anonymous. Irrigation in the Hawaiian Islands. Scientific American, 84(2):20–21, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
20a.pdf.

Anonymous:1901:ZFC

[2779] Anonymous. The Zodel flexible coupling. Scientific American, 84(2):21, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
21.pdf.

Anonymous:1901:SMB

[2780] Anonymous. A simple method of broaching brasses and bearings. Scientific American, 84(2):21, January 12, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n2/pdf/scientificamerican01121901-21a.pdf.

Anonymous:1901:BRTb

[2781] Anonymous. The building and repairing of tall chimneys. Scientific American, 84(2):21–22, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
21b.pdf.

Anonymous:1901:LLT

[2782] Anonymous. A large lathe for turning granite columns. Scientific American, 84(2):22–23, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
22.pdf.

Willey:1901:CEB

[2783] D. A. Willey. Curious effects of boiler explosions. Scientific American, 84(2):24–25, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
24.pdf.

Anonymous:1901:DNF

[2784] Anonymous. Dredges for the new 40-foot channels of New York Harbor. Scientific American, 84(2):25–26, January 12, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n2/pdf/scientificamerican01121901-25.pdf.

Anonymous:1901:TRPa

[2785] Anonymous. Three recently patented novelties. Scientific American, 84(2):27, January 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n2/pdf/scientificamerican01121901-
27.pdf.

Anonymous:1901:SNd

[2786] Anonymous. Science notes. Scientific American, 84(20):307, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-307b.pdf.

Anonymous:1901:RPIm

[2787] Anonymous. Recently patented inventions. Scientific American, 84(20):315, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
315.pdf.

Anonymous:1901:BNBg

[2788] Anonymous. Books: New books, etc. Scientific American, 84(20):315, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-315a.pdf.

Anonymous:1901:NQm

[2789] Anonymous. Notes and queries. Scientific American, 84(20):316, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-316a.pdf.

Anonymous:1901:IIm

[2790] Anonymous. Index of inventions. Scientific American, 84(20):316, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-316b.pdf.

Anonymous:1901:PNPl

[2791] Anonymous. Patent notes: Patents. Scientific American, 84(20):319, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-319.pdf.

Anonymous:1901:EBS

[2792] Anonymous. The existence of bodies smaller than atoms — the corpuscular hypothesis. Scientific American, 84(20):306, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-306.pdf.

Anonymous:1901:PAA

[2793] Anonymous. Proposed abolition of the Army Transport Service on the Atlantic. Scientific American, 84(20):306, May 18, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n20/pdf/scientificamerican05181901-306a.pdf.

Anonymous:1901:NTC

[2794] Anonymous. A new type of coast defense gun. Scientific American, 84(20):306, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
306b.pdf.

Anonymous:1901:SCa

[2795] Anonymous. “Shamrock” and “Constitution”. Scientific American, 84(20):306, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
306c.pdf.

Anonymous:1901:NET

[2796] Anonymous. New electric traction systems in Europe. Scientific American, 84(20):306–307, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
306d.pdf.

Anonymous:1901:VLE

[2797] Anonymous. Variability in light of eros. Scientific American, 84(20):307, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-307.pdf.

Brown:1901:CRM

[2798] Enos Brown. Cultivation of rubber in Mexico. Scientific American, 84(20):307, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
307a.pdf.

Anonymous:1901:ATP

[2799] Anonymous. An attachment for trolley poles. Scientific American, 84(20):308, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
308.pdf.

Davison:1901:CC

[2800] Ralph C. Davison. A concrete chimney. Scientific American, 84(20):308, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-308a.pdf.

Anonymous:1901:NIC

[2801] Anonymous. A new ice-conveying apparatus. Scientific American, 84(20):308, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
308b.pdf.

Fawcett:1901:TI

[2802] Waldon Fawcett. The tobacco industry. Scientific American, 84 (20):309–310, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
309.pdf.

Anonymous:1901:NGB

[2803] Anonymous. New gas battery. Scientific American, 84(20):310, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n20/pdf/scientificamerican05181901-310.pdf.

Anonymous:1901:MPG

[2804] Anonymous. Manufacture of plate glass. Scientific American, 84 (20):311, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
311.pdf.

Mayer:1901:MBI

[2805] Alfred G. Mayer. The mastodon at the Brooklyn Institute Museum. Scientific American, 84(20):312, May 18, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n20/pdf/scientificamerican05181901-312.pdf.

Anonymous:1901:LCa

[2806] Anonymous. Launch of the “Constitution”. Scientific American, 84(20):312–313, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
312a.pdf.

Gathmann:1901:GIT

[2807] Emil Gathmann. Gathmann 18-inch torpedo gun. Scientific American, 84(20):313–314, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
313.pdf.

Anonymous:1901:MP

[2808] Anonymous. Malaria and its prevention. Scientific American, 84 (20):314, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
314.pdf.

Anonymous:1901:BPWk

[2809] Anonymous. Business and personal wants. Scientific American, 84(20):316, May 18, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n20/pdf/scientificamerican05181901-
316.pdf.

Anonymous:1901:SNe

[2810] Anonymous. Science notes. Scientific American, 84(21):323, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n21/pdf/scientificamerican05251901-323b.pdf.

Anonymous:1901:RPIn

[2811] Anonymous. Recently patented inventions. Scientific American, 84(21):331, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
331.pdf.

Anonymous:1901:NQn

[2812] Anonymous. Notes and queries. Scientific American, 84(21):331–332, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n21/pdf/scientificamerican05251901-331b.pdf.

Anonymous:1901:IIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n21/pdf/scientificamerican05251901-332.pdf.

Anonymous:1901:PNPm

[2814] Anonymous. Patent notes: Patents. Scientific American, 84(21):335, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n21/pdf/scientificamerican05251901-335.pdf.

Anonymous:1901:PGR

[2815] Anonymous. Pipe galleries for the rapid transit subway. Scientific American, 84(21):322, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
322.pdf.

Anonymous:1901:RRS

[2816] Anonymous. Remarkable results at sandy hook. Scientific American, 84(21):322, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
322a.pdf.

Anonymous:1901:GLN

[2817] Anonymous. The growth of London and New York. Scientific American, 84(21):322, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
322b.pdf.

Anonymous:1901:FIC

[2818] Anonymous. Forestry in India and Canada. Scientific American, 84(21):322, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
322c.pdf.

Anonymous:1901:NPPb

[2819] Anonymous. New process for plaster of Paris. Scientific American, 84(21):322–323, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
322d.pdf.

Holder:1901:WAW

[2820] Charles Frederick Holder. Wild animals in winter. Scientific American, 84(21):323, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
323.pdf.

Anonymous:1901:NTP

[2821] Anonymous. New trans-Pacific cable. Scientific American, 84(21):323, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n21/pdf/scientificamerican05251901-323a.pdf.

Anonymous:1901:RBE

[2822] Anonymous. Russian bayonet exercises. Scientific American, 84 (21):324, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
324.pdf.

Anonymous:1901:PDA

[2823] Anonymous. The Popoff–Ducretet apparatus for wireless telegraphy. Scientific American, 84(21):324, May 25, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n21/pdf/scientificamerican05251901-324a.pdf.

Anonymous:1901:HBJ

[2824] Anonymous. A hydraulic bicycle joint. Scientific American, 84 (21):324–325, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
324b.pdf.

Anonymous:1901:GTM

[2825] Anonymous. Government tests of Maximite at Sandy Hook. Scientific American, 84(21):325–326, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
325.pdf.

Anonymous:1901:SFN

[2826] Anonymous. Some features of the New York Rapid Transit Tunnel. Scientific American, 84(21):327, May 25, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n21/pdf/scientificamerican05251901-327.pdf.

Anonymous:1901:IF

[2827] Anonymous. An ice fountain. Scientific American, 84(21):328, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n21/pdf/scientificamerican05251901-328.pdf.

Anonymous:1901:SUS

[2828] Anonymous. The “Shamrock” under sail. Scientific American, 84(21):328, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
328a.pdf.

Watson:1901:AP

[2829] W. F. Watson. The amateur in photomicrography. Scientific American, 84(21):329, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
329.pdf.

Moore:1901:IIT

[2830] N. Hudson Moore. An interesting insect — Thyridopteryx Ephemeræformis. Scientific American, 84(21):330, May 25, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n21/pdf/scientificamerican05251901-330.pdf.

Anonymous:1901:BPWl

[2831] Anonymous. Business and personal wants. Scientific American, 84(21):331, May 25, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n21/pdf/scientificamerican05251901-
331a.pdf.

Anonymous:1901:SNf

[2832] Anonymous. Science notes. Scientific American, 84(22):339, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n22/pdf/scientificamerican06011901-339c.pdf.

Anonymous:1901:RPIo

[2833] Anonymous. Recently patented inventions. Scientific American, 84(22):347, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
347.pdf.

Anonymous:1901:NQo

[2834] Anonymous. Notes and queries. Scientific American, 84(22):347, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n22/pdf/scientificamerican06011901-347b.pdf.

Anonymous:1901:IIo

[2835] Anonymous. Index of inventions. Scientific American, 84(22):347–351, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n22/pdf/scientificamerican06011901-347c.pdf.

Anonymous:1901:PNPn

[2836] Anonymous. Patent notes: Patents. Scientific American, 84(22):351, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n22/pdf/scientificamerican06011901-351.pdf.

Anonymous:1901:NESa

[2837] Anonymous. The new Edison storage battery. Scientific American, 84(22):338, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
338.pdf.

Anonymous:1901:LCS

[2838] Anonymous. Last of the cable systems in New York. Scientific American, 84(22):338, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
338a.pdf.

Anonymous:1901:OPAb

[2839] Anonymous. Opening of the Pan-American Exposition. Scientific American, 84(22):338–339, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
338b.pdf.

Anonymous:1901:HCW

[2840] Anonymous. High-class work at our Navy yards. Scientific American, 84(22):339, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
339.pdf.

Anonymous:1901:FSS

[2841] Anonymous. The “factor of safety” in “Shamrock”. Scientific American, 84(22):339, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
339a.pdf.

Anonymous:1901:HJa

[2842] Anonymous. The heavens in June. Scientific American, 84(22):339, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n22/pdf/scientificamerican06011901-339b.pdf.

Anonymous:1901:DTR

[2843] Anonymous. Double-track railway viaduct over the Des Moines River. Scientific American, 84(22):340, June 1, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n22/pdf/scientificamerican06011901-340.pdf.

Anonymous:1901:USE

[2844] Anonymous. The United States Experimental Model Basin. Scientific American, 84(22):341, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
341.pdf.

Anonymous:1901:JPR

[2845] Anonymous. Jerome Park Reservoir. Scientific American, 84 (22):342–343, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
342.pdf.

Brown:1901:PNI

[2846] Enos Brown. Pine needles industry in Oregon. Scientific American, 84(22):344, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
344.pdf.

Anonymous:1901:AFM

[2847] Anonymous. Alcohol as fuel for motor carriages. Scientific American, 84(22):344, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
344a.pdf.

Anonymous:1901:RDT

[2848] Anonymous. Removing the disabled 65-ton gun of the “Kearsarge”. Scientific American, 84(22):345–346, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
345.pdf.

Anonymous:1901:EMC

[2849] Anonymous. The experiments of m.curie. Scientific American, 84(22):346, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
346.pdf.

Anonymous:1901:PMM

[2850] Anonymous. A portable mortising machine. Scientific American, 84(22):346, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
346a.pdf.

Anonymous:1901:BPWm

[2851] Anonymous. Business and personal wants. Scientific American, 84(22):347, June 1, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n22/pdf/scientificamerican06011901-
347a.pdf.

Anonymous:1901:Cf

[2852] Anonymous. Correspondence. Scientific American, 84(23):358, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n23/pdf/scientificamerican06081901-358.pdf.

Anonymous:1901:RPIp

[2853] Anonymous. Recently patented inventions. Scientific American, 84(23):363, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
363.pdf.

Anonymous:1901:NQp

[2854] Anonymous. Notes and queries. Scientific American, 84(23):364, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n23/pdf/scientificamerican06081901-364.pdf.

Anonymous:1901:BNBh

[2855] Anonymous. Books: New books, etc. Scientific American, 84(23):364, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n23/pdf/scientificamerican06081901-364a.pdf.

Anonymous:1901:IIp

[2856] Anonymous. Index of inventions. Scientific American, 84(23):364, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n23/pdf/scientificamerican06081901-364b.pdf.

Anonymous:1901:PNPo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n23/pdf/scientificamerican06081901-367.pdf.

Anonymous:1901:CRG

[2858] Anonymous. The Canadian route from the Great Lakes. Scientific American, 84(23):354, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
354.pdf.

Anonymous:1901:HSAa

[2859] Anonymous. Hydraulic system of air compression. Scientific American, 84(23):354, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
354a.pdf.

Anonymous:1901:CEK

[2860] Anonymous. Comparative efficiency of Krupp and other highpower guns. Scientific American, 84(23):354, June 8, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n23/pdf/scientificamerican06081901-354b.pdf.

Anonymous:1901:GTI

[2861] Anonymous. Growth of the tinfoil industry. Scientific American, 84(23):354–355, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
354c.pdf.

Anonymous:1901:MPD

[2862] Anonymous. Methods proposed for dealing with London’s congested traffic. Scientific American, 84(23):355, June 8, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n23/pdf/scientificamerican06081901-355.pdf.

Anonymous:1901:WTP

[2863] Anonymous. Wireless telegraphy for the prevention of shipping disasters. Scientific American, 84(23):355, June 8, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n23/pdf/scientificamerican06081901-355a.pdf.

Fawcett:1901:CMA

[2864] Waldon Fawcett. Copper mining in America. Scientific American, 84(23):356–357, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
356.pdf.

Anonymous:1901:EAK

[2865] Anonymous. Electric automobile — Krieger system. Scientific American, 84(23):357, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
357.pdf.

Anonymous:1901:NFMa

[2866] Anonymous. A new flying machine. Scientific American, 84(23):357, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n23/pdf/scientificamerican06081901-357a.pdf.

Daniel:1901:WA

[2867] Frank C. Daniel. A whistling arc. Scientific American, 84(23):358, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n23/pdf/scientificamerican06081901-358a.pdf.

Anonymous:1901:LBO

[2868] Anonymous. Launch of the battleship “Ohio”. Scientific American, 84(23):359, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
359.pdf.

Anonymous:1901:CCW

[2869] Anonymous. The cultivation of cocoa in the West India Islands. Scientific American, 84(23):360, June 8, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n23/pdf/scientificamerican06081901-360.pdf.

Anonymous:1901:PAE

[2870] Anonymous. The Pan-American Exposition. Scientific American, 84(23):360–362, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
360a.pdf.

Anonymous:1901:PSV

[2871] Anonymous. A pneumatic spring for vehicles. Scientific American, 84(23):362, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
362.pdf.

Anonymous:1901:NMR

[2872] Anonymous. A new metal railway tie. Scientific American, 84(23):362, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n23/pdf/scientificamerican06081901-362a.pdf.

Anonymous:1901:SSV

[2873] Anonymous. A supplemental seat for vehicles. Scientific American, 84(23):362, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
362b.pdf.

Anonymous:1901:BPWn

[2874] Anonymous. Business and personal wants. Scientific American, 84(23):363–364, June 8, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n23/pdf/scientificamerican06081901-
363a.pdf.

Anonymous:1901:SNg

[2875] Anonymous. Science notes. Scientific American, 84(24):371, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n24/pdf/scientificamerican06151901-371c.pdf.

Anonymous:1901:Cg

[2876] Anonymous. Correspondence. Scientific American, 84(24):375, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n24/pdf/scientificamerican06151901-375.pdf.

Anonymous:1901:RPIq

[2877] Anonymous. Recently patented inventions. Scientific American, 84(24):379, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
379.pdf.

Anonymous:1901:NQq

[2878] Anonymous. Notes and queries. Scientific American, 84(24):380, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n24/pdf/scientificamerican06151901-380a.pdf.

Anonymous:1901:IIq

[2879] Anonymous. Index of inventions. Scientific American, 84(24):380, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n24/pdf/scientificamerican06151901-380b.pdf.

Anonymous:1901:PNPp

[2880] Anonymous. Patent notes: Patents. Scientific American, 84(24):383, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n24/pdf/scientificamerican06151901-383.pdf.

Anonymous:1901:GPS

[2881] Anonymous. The greatest of power stations. Scientific American, 84(24):370, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
370.pdf.

Anonymous:1901:CEa

[2882] Anonymous. A costly equipment. Scientific American, 84(24):370, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n24/pdf/scientificamerican06151901-370a.pdf.

Anonymous:1901:RTTa

[2883] Anonymous. Rapid transit tunnel to Brooklyn assured. Scientific American, 84(24):370, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
370b.pdf.

Anonymous:1901:WLG

[2884] Anonymous. The wave line of the “Georgia”. Scientific American, 84(24):370, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
370c.pdf.

Anonymous:1901:GFT

[2885] Anonymous. Germany’s foreign trade in machinery. Scientific American, 84(24):370, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
370d.pdf.

Anonymous:1901:LAX

[2886] Anonymous. Laws of absorption of X-rays for different bodies. Scientific American, 84(24):370–371, June 15, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n24/pdf/scientificamerican06151901-370e.pdf.

Proctor:1901:ERE

[2887] M. Proctor. Expected return of Encke’s Comet. Scientific American, 84(24):371, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
371.pdf.

Anonymous:1901:SLEa

[2888] Anonymous. The St. Louis Exposition of 1903. Scientific American, 84(24):371, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
371a.pdf.

Anonymous:1901:TFK

[2889] Anonymous. Twenty-five knot Naval Scouts. Scientific American, 84(24):371, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
371b.pdf.

Anonymous:1901:SYL

[2890] Anonymous. The steam yacht “Lysistrata”. Scientific American, 84(24):372, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
372.pdf.

Anonymous:1901:ECF

[2891] Anonymous. An efficient car-fender. Scientific American, 84(24):372, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n24/pdf/scientificamerican06151901-372a.pdf.

Anonymous:1901:ENF

[2892] Anonymous. The electrograph — a new facsimile telegraph. Scientific American, 84(24):373–374, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
373.pdf.

Anonymous:1901:FSN

[2893] Anonymous. Floor system of the new East River Bridge. Scientific American, 84(24):374, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
374.pdf.

Anonymous:1901:CTS

[2894] Anonymous. A Chinese telephone system. Scientific American, 84(24):376, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
376.pdf.

Anonymous:1901:NESb

[2895] Anonymous. The new Edison storage battery. Scientific American, 84(24):376, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
376a.pdf.

Anonymous:1901:NFD

[2896] Anonymous. The new floating drydock for the United States Navy. Scientific American, 84(24):377–378, June 15, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n24/pdf/scientificamerican06151901-377.pdf.

Anonymous:1901:SBC

[2897] Anonymous. A simple brick-carrying device. Scientific American, 84(24):378, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
378.pdf.

Anonymous:1901:SMP

[2898] Anonymous. A self-measuring pump for liquids. Scientific American, 84(24):378, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
378a.pdf.

Anonymous:1901:BPWo

[2899] Anonymous. Business and personal wants. Scientific American, 84(24):380, June 15, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n24/pdf/scientificamerican06151901-
380.pdf.

Anonymous:1901:SNh

[2900] Anonymous. Science notes. Scientific American, 84(25):387, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n25/pdf/scientificamerican06221901-387b.pdf.

Anonymous:1901:RPIr

[2901] Anonymous. Recently patented inventions. Scientific American, 84(25):395, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
395.pdf.

Anonymous:1901:BNBi

[2902] Anonymous. Books: New books, etc. Scientific American, 84 (25):395–396, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
395a.pdf.

Anonymous:1901:NQr

[2903] Anonymous. Notes and queries. Scientific American, 84(25):396–397, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n25/pdf/scientificamerican06221901-396a.pdf.

Anonymous:1901:IIr

[2904] Anonymous. Index of inventions. Scientific American, 84(25):397, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n25/pdf/scientificamerican06221901-397a.pdf.

Anonymous:1901:PNPq

[2905] Anonymous. Patent notes: Patents. Scientific American, 84(25):399, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n25/pdf/scientificamerican06221901-399.pdf.

Anonymous:1901:ALA

[2906] Anonymous. American locomotives abroad. Scientific American, 84(25):386, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
386.pdf.

Anonymous:1901:PISb

[2907] Anonymous. Protection of iron structures. Scientific American, 84(25):386, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
386a.pdf.

Anonymous:1901:OFB

[2908] Anonymous. Our fastest battleship. Scientific American, 84(25):386, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n25/pdf/scientificamerican06221901-386b.pdf.

Witz:1901:GE

[2909] Aime Witz. Gas engines. Scientific American, 84(25):387, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n25/pdf/scientificamerican06221901-387.pdf.

Anonymous:1901:CAR

[2910] Anonymous. The clothing of the Ancient Romans. Scientific American, 84(25):387, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
387a.pdf.

Anonymous:1901:APHa

[2911] Anonymous. An adjustable power-hammer. Scientific American, 84(25):388, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
388.pdf.

Anonymous:1901:LSTb

[2912] Anonymous. Launch of the submarine torpedo boat “Fulton”. Scientific American, 84(25):388, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
388a.pdf.

Guenaire:1901:FMB

[2913] J. Guenaire. The form of mercury bath — experiments at the Observatory of Paris. Scientific American, 84(25):388–389, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n25/pdf/scientificamerican06221901-388b.pdf.

Anonymous:1901:TCY

[2914] Anonymous. The three cup-yachts of 1901. Scientific American, 84(25):389–390, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
389.pdf.

Anonymous:1901:PSS

[2915] Anonymous. Pressed-steel system of car construction. Scientific American, 84(25):391, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
391.pdf.

Anonymous:1901:STT

[2916] Anonymous. The Stralau–Treptow Tunnel under the River Spree. Scientific American, 84(25):392, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
392.pdf.

Adams:1901:MTS

[2917] Frank D. Adams. Method of testing specimens of marble. Scientific American, 84(25):393, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
393.pdf.

Aaron:1901:LHI

[2918] S. Frank Aaron. The life history of an insect parasite. Scientific American, 84(25):394, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
394.pdf.

Anonymous:1901:BPWp

[2919] Anonymous. Business and personal wants. Scientific American, 84(25):396, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n25/pdf/scientificamerican06221901-
396.pdf.

Anonymous:1901:Ib

[2920] Anonymous. To inventors. Scientific American, 84(25):397, June 22, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n25/pdf/scientificamerican06221901-397.pdf.

Anonymous:1901:HJb

[2921] Anonymous. The heavens in July. Scientific American, 84(26):403, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-403.pdf.

Anonymous:1901:SNi

[2922] Anonymous. Science notes. Scientific American, 84(26):403, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-403b.pdf.

Anonymous:1901:Ch

[2923] Anonymous. Correspondence. Scientific American, 84(26):407, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-407.pdf.

Anonymous:1901:RPIs

[2924] Anonymous. Recently patented inventions. Scientific American, 84(26):411, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
411.pdf.

Anonymous:1901:BNBj

[2925] Anonymous. Books: New books, etc. Scientific American, 84(26):411, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-411a.pdf.

Anonymous:1901:NQs

[2926] Anonymous. Notes and queries. Scientific American, 84(26):412, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-412.pdf.

Anonymous:1901:IIs

[2927] Anonymous. Index of inventions. Scientific American, 84(26):412, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-412a.pdf.

Anonymous:1901:PNPr

[2928] Anonymous. Patent notes: Patents. Scientific American, 84(26):415, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-415.pdf.

Anonymous:1901:Ic

[2929] Anonymous. Index. Scientific American, 84(26):417–418, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-417.pdf.

Anonymous:1901:BWa

[2930] Anonymous. Biggest in the world. Scientific American, 84(26):402, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-402.pdf.

Anonymous:1901:ESPb

[2931] Anonymous. Extraordinary speed in the Paris–Bordeaux Automobile Race. Scientific American, 84(26):402, June 29, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n26/pdf/scientificamerican06291901-402a.pdf.

Anonymous:1901:LRF

[2932] Anonymous. The lesson of the recent ferryboat disaster. Scientific American, 84(26):402, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
402b.pdf.

Anonymous:1901:NCL

[2933] Anonymous. The new Clyde Line steamship “Apache”. Scientific American, 84(26):402, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
402c.pdf.

Anonymous:1901:RGC

[2934] Anonymous. Reconstruction of German cruisers. Scientific American, 84(26):402–403, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
402d.pdf.

Anonymous:1901:DCC

[2935] Anonymous. Death of T. C. Clarke. Scientific American, 84(26):403, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-403a.pdf.

Anonymous:1901:NHM

[2936] Anonymous. A novel house mail-box. Scientific American, 84(26):404, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-404.pdf.

Anonymous:1901:CDA

[2937] Anonymous. Curious docking accident. Scientific American, 84(26):404, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n26/pdf/scientificamerican06291901-404a.pdf.

Anonymous:1901:SCS

[2938] Anonymous. A self-closing spigot or faucet. Scientific American, 84(26):404, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
404b.pdf.

Embree:1901:WSS

[2939] F. B. Embree. The water supply system of Chicago. Scientific American, 84(26):405–406, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
405.pdf.

Anonymous:1901:PTL

[2940] Anonymous. Provisioning a transatlantic liner. Scientific American, 84(26):406, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
406.pdf.

Anonymous:1901:GEL

[2941] Anonymous. The Ganz electric line of Valtellina. Scientific American, 84(26):408, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
408.pdf.

Anonymous:1901:MUO

[2942] Anonymous. Method of utilizing old street cars. Scientific American, 84(26):409–410, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
409.pdf.

Anonymous:1901:TRPb

[2943] Anonymous. Three recently patented novelties. Scientific American, 84(26):410, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
410.pdf.

Anonymous:1901:BPWq

[2944] Anonymous. Business and personal wants. Scientific American, 84(26):411–412, June 29, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n26/pdf/scientificamerican06291901-
411b.pdf.

Anonymous:1901:SNj

[2945] Anonymous. Science notes. Scientific American, 84(3):35, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-35a.pdf.

Anonymous:1901:RPIt

[2946] Anonymous. Recently patented inventions. Scientific American, 84(3):43–44, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
43.pdf.

Anonymous:1901:BPc

[2947] Anonymous. Business and personal. Scientific American, 84(3):44, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-44.pdf.

Anonymous:1901:NQt

[2948] Anonymous. Notes and queries. Scientific American, 84(3):44, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-44a.pdf.

Anonymous:1901:BNBk

[2949] Anonymous. Books: New books, etc. Scientific American, 84 (3):44–45, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
44b.pdf.

Anonymous:1901:IIt

[2950] Anonymous. Index of inventions. Scientific American, 84(3):45, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-45.pdf.

Anonymous:1901:PNPs

[2951] Anonymous. Patent notes: Patents. Scientific American, 84(3):46, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-46.pdf.

Anonymous:1901:BWb

[2952] Anonymous. A brake that works. Scientific American, 84(3):34, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-34.pdf.

Anonymous:1901:WGW

[2953] Anonymous. Whom the gods would destroy they first make mad. Scientific American, 84(3):34, January 19, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n3/pdf/scientificamerican01191901-34a.pdf.

Anonymous:1901:MMa

[2954] Anonymous. That message from Mars. Scientific American, 84 (3):34, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
34b.pdf.

Anonymous:1901:CCI

[2955] Anonymous. Continental criticism of the 16-inch gun. Scientific American, 84(3):34, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
34c.pdf.

Anonymous:1901:TCb

[2956] Anonymous. The tunneling craze. Scientific American, 84(3):34, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-34d.pdf.

Anonymous:1901:EWHa

[2957] Anonymous. The enlargement of the White House, Washington. Scientific American, 84(3):34, January 19, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n3/pdf/scientificamerican01191901-34e.pdf.

Anonymous:1901:SRB

[2958] Anonymous. Speed records of the bicycle for 1900. Scientific American, 84(3):34, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
34f.pdf.

Anonymous:1901:TSTa

[2959] Anonymous. Transmitting signals through the sea. Scientific American, 84(3):34–35, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
34g.pdf.

Rice:1901:SSF

[2960] George D. Rice. Some of the singular foods of the Filipinos. Scientific American, 84(3):35, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
35.pdf.

Anonymous:1901:OHC

[2961] Anonymous. An overhead hoisting and carrying device. Scientific American, 84(3):36, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
36.pdf.

Anonymous:1901:PLP

[2962] Anonymous. Panhard & Levassor petroleum locomotive. Scientific American, 84(3):36, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
36a.pdf.

Hopkins:1901:VSH

[2963] George M. Hopkins. VII. simple hygroscope. Scientific American, 84(3):36, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
36b.pdf.

Anonymous:1901:RTTb

[2964] Anonymous. The rapid transit tunnel. Scientific American, 84 (3):37, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
37.pdf.

Anonymous:1901:Ma

[2965] Anonymous. The Mirograph. Scientific American, 84(3):37–38, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-37a.pdf.

Anonymous:1901:IG

[2966] Anonymous. An interesting group. Scientific American, 84(3):38, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-38.pdf.

Anonymous:1901:MWGa

[2967] Anonymous. Manufacture of water-gas at the Forty-Fourth Street Station New York — I. Scientific American, 84(3):39, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n3/pdf/scientificamerican01191901-39.pdf.

Fawcett:1901:ACF

[2968] Waldon Fawcett. American car ferries. Scientific American, 84 (3):40–41, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
40.pdf.

Anonymous:1901:VVP

[2969] Anonymous. Vaccine virus — its preparation and its use. Scientific American, 84(3):41–42, January 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n3/pdf/scientificamerican01191901-
41.pdf.

Anonymous:1901:SNk

[2970] Anonymous. Science notes. Scientific American, 84(4):51, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-51b.pdf.

Anonymous:1901:Ci

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-54.pdf.

Anonymous:1901:RPIu

[2972] Anonymous. Recently patented inventions. Scientific American, 84(4):59, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
59.pdf.

Anonymous:1901:BPd

[2973] Anonymous. Business and personal. Scientific American, 84(4):59, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-59a.pdf.

Anonymous:1901:NQu

[2974] Anonymous. Notes and queries. Scientific American, 84(4):59–60, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-59b.pdf.

Anonymous:1901:BNBl

[2975] Anonymous. Books: New books, etc. Scientific American, 84 (4):60, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
60.pdf.

Anonymous:1901:IIu

[2976] Anonymous. Index of inventions. Scientific American, 84(4):60, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-60a.pdf.

Anonymous:1901:PNPt

[2977] Anonymous. Patent notes: Patents. Scientific American, 84(4):62, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-62.pdf.

Anonymous:1901:EFD

[2978] Anonymous. Effect of forced draught on coal consumption. Scientific American, 84(4):50, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
50.pdf.

Anonymous:1901:TCS

[2979] Anonymous. The Twentieth Century sailing vessel. Scientific American, 84(4):50, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
50a.pdf.

Anonymous:1901:PLD

[2980] Anonymous. Pupin’s long distance telephony. Scientific American, 84(4):50, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
50b.pdf.

Anonymous:1901:KAa

[2981] Anonymous. The Kress airship. Scientific American, 84(4):50, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-50c.pdf.

Anonymous:1901:ICP

[2982] Anonymous. The international convention for the protection of industrial property. Scientific American, 84(4):50–51, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-50d.pdf.

Holden:1901:ORD

[2983] Edward S. Holden. Orbits of revolving double stars. Scientific American, 84(4):51, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
51.pdf.

Anonymous:1901:PBCa

[2984] Anonymous. A prize for beer-cooling mixtures. Scientific American, 84(4):51, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
51a.pdf.

Anonymous:1901:WPC

[2985] Anonymous. A wonderful piece of clockwork. Scientific American, 84(4):52, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
52.pdf.

Anonymous:1901:RV

[2986] Anonymous. A runner for vehicles. Scientific American, 84(4):52, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n4/pdf/scientificamerican01261901-52a.pdf.

Anonymous:1901:BPPb

[2987] Anonymous. The Burton plastic pressure tube expander. Scientific American, 84(4):52, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
52b.pdf.

Willey:1901:IFS

[2988] D. Allen Willey. Irrigation farming in the Southwest. Scientific American, 84(4):53–54, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
53.pdf.

Perkins:1901:GEL

[2989] Frank C. Perkins. German express locomotive with auxiliary driving axle. Scientific American, 84(4):55, January 26, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n4/pdf/scientificamerican01261901-55.pdf.

Anonymous:1901:EMH

[2990] Anonymous. An eight-mile house moving. Scientific American, 84(4):56, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
56.pdf.

Anonymous:1901:TLC

[2991] Anonymous. Train lighting from the car axle. Scientific American, 84(4):56–57, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
56a.pdf.

Anonymous:1901:RGA

[2992] Anonymous. Recent gasoline automobiles. Scientific American, 84(4):57, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
57.pdf.

Anonymous:1901:ABS

[2993] Anonymous. The Annual Bicycle Show at Madison Square Garden. Scientific American, 84(4):57–58, January 26, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n4/pdf/scientificamerican01261901-57a.pdf.

Anonymous:1901:EES

[2994] Anonymous. Electrical engineers in South Africa. Scientific American, 84(4):58, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
58.pdf.

Anonymous:1901:CMCa

[2995] Anonymous. A combined miner’s candlestick and tool. Scientific American, 84(4):58, January 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n4/pdf/scientificamerican01261901-
58a.pdf.

Russell:1901:HF

[2996] Henry Norris Russell. The heavens in February. Scientific American, 84(5):67, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
67.pdf.

Anonymous:1901:RPIv

[2997] Anonymous. Recently patented inventions. Scientific American, 84(5):76, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
76.pdf.

Anonymous:1901:BPe

[2998] Anonymous. Business and personal. Scientific American, 84 (5):76, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
76a.pdf.

Anonymous:1901:NQv

[2999] Anonymous. Notes and queries. Scientific American, 84(5):76– 77, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
76b.pdf.

Anonymous:1901:BNBm

[3000] Anonymous. Books: New books, etc. Scientific American, 84 (5):77, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
77.pdf.

Anonymous:1901:IIv

[3001] Anonymous. Index of inventions. Scientific American, 84(5):77, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n5/pdf/scientificamerican02021901-77a.pdf.

Anonymous:1901:PNPu

[3002] Anonymous. Patent notes: Patents. Scientific American, 84 (5):79, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
79.pdf.

Anonymous:1901:V

[3003] Anonymous. Victoria. Scientific American, 84(5):66, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n5/pdf/scientificamerican02021901-66.pdf.

Anonymous:1901:NB

[3004] Anonymous. The Naval Bill. Scientific American, 84(5):66, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n5/pdf/scientificamerican02021901-66a.pdf.

Anonymous:1901:YRT

[3005] Anonymous. Yacht racing and the towing tank. Scientific American, 84(5):66, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
66b.pdf.

Anonymous:1901:IRT

[3006] Anonymous. The importance of registering trade marks. Scientific American, 84(5):66, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
66c.pdf.

Anonymous:1901:EPP

[3007] Anonymous. Electrolytic preparation of phosphorus. Scientific American, 84(5):66, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
66d.pdf.

Anonymous:1901:TWL

[3008] Anonymous. Tesla’s wireless light. Scientific American, 84(5):67, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n5/pdf/scientificamerican02021901-67a.pdf.

Anonymous:1901:TSC

[3009] Anonymous. The twin-screw cruising yacht “Prinzessin Victoria-Luise”. Scientific American, 84(5):67, February 2, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n5/pdf/scientificamerican02021901-67b.pdf.

Anonymous:1901:SSS

[3010] Anonymous. A simple saw setting and gumming machine. Scientific American, 84(5):68, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
68.pdf.

Holder:1901:FBC

[3011] Charles Frederick Holder. Famous basaltic columns. Scientific American, 84(5):68, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
68a.pdf.

Stewart:1901:UEU

[3012] J. A. Stewart. Unique elevated-underground street railway in Boston. Scientific American, 84(5):69–70, February 2, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n5/pdf/scientificamerican02021901-69.pdf.

Fawcett:1901:SRU

[3013] Waldon Fawcett. The steepest railways in the United States. Scientific American, 84(5):70–71, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
70.pdf.

Anonymous:1901:DPE

[3014] Anonymous. Death of Prof. Elisha Gray. Scientific American, 84(5):72, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
72.pdf.

Anonymous:1901:NFMb

[3015] Anonymous. A new flying machine. Scientific American, 84 (5):72, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
72a.pdf.

Anonymous:1901:NYCa

[3016] Anonymous. The New York Cycle and Automobile Show. Scientific American, 84(5):73, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
73.pdf.

Mundy:1901:SS

[3017] Arthur J. Mundy. Submarine signaling. Scientific American, 84 (5):73–74, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
73a.pdf.

Hopkins:1901:VMB

[3018] George.M. Hopkins. VIII Mercurial barometer. Scientific American, 84(5):74, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
74.pdf.

Anonymous:1901:BOW

[3019] Anonymous. The Beaumont oil well. Scientific American, 84 (5):74–75, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
74a.pdf.

Anonymous:1901:PPF

[3020] Anonymous. Press for preparing fomentations. Scientific American, 84(5):75, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
75.pdf.

Anonymous:1901:NIR

[3021] Anonymous. Novel inventions recently patented. Scientific American, 84(5):75, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
75a.pdf.

Anonymous:1901:PMT

[3022] Anonymous. Picture making on the typewriter. Scientific American, 84(5):75, February 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n5/pdf/scientificamerican02021901-
75b.pdf.

Anonymous:1901:SNl

[3023] Anonymous. Science notes. Scientific American, 84(6):83, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n6/pdf/scientificamerican02091901-83a.pdf.

Anonymous:1901:Cj

[3024] Anonymous. Correspondence. Scientific American, 84(6):87, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n6/pdf/scientificamerican02091901-87.pdf.

Anonymous:1901:RPIw

[3025] Anonymous. Recently patented inventions. Scientific American, 84(6):91, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
91.pdf.

Anonymous:1901:BNBn

[3026] Anonymous. Books: New books, etc. Scientific American, 84 (6):91–92, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
91a.pdf.

Anonymous:1901:BPf

[3027] Anonymous. Business and personal. Scientific American, 84 (6):92, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
92.pdf.

Anonymous:1901:NQw

[3028] Anonymous. Notes and queries. Scientific American, 84(6):92, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n6/pdf/scientificamerican02091901-92a.pdf.

Anonymous:1901:IIw

[3029] Anonymous. Index of inventions. Scientific American, 84(6):92, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n6/pdf/scientificamerican02091901-92b.pdf.

Anonymous:1901:PNPv

[3030] Anonymous. Patent notes: Patents. Scientific American, 84 (6):95, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
95.pdf.

Anonymous:1901:RJS

[3031] Anonymous. Rail joints and street noises. Scientific American, 84(6):82, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
82.pdf.

Anonymous:1901:PPH

[3032] Anonymous. Pound-per-horsepower economy. Scientific American, 84(6):82, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
82a.pdf.

Anonymous:1901:PTP

[3033] Anonymous. Progress of turbine propulsion. Scientific American, 84(6):82, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
82b.pdf.

Anonymous:1901:ARR

[3034] Anonymous. Air resistance of railway trains. Scientific American, 84(6):82, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
82c.pdf.

Anonymous:1901:RFE

[3035] Anonymous. Relative fighting efficiency of warships. Scientific American, 84(6):82, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
82d.pdf.

Anonymous:1901:CCP

[3036] Anonymous. Comparative cost of prime movers. Scientific American, 84(6):82–83, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
82e.pdf.

Anonymous:1901:EAa

[3037] Anonymous. Efficiency in action. Scientific American, 84(6):83, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n6/pdf/scientificamerican02091901-83.pdf.

Anonymous:1901:SBI

[3038] Anonymous. The spontaneous breaking of an inkwell. Scientific American, 84(6):84, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
84.pdf.

Anonymous:1901:SCB

[3039] Anonymous. The subway of the City of Berlin. Scientific American, 84(6):84, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
84a.pdf.

Anonymous:1901:LALa

[3040] Anonymous. A lens attachment for lamps. Scientific American, 84(6):84, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
84b.pdf.

Anonymous:1901:TAC

[3041] Anonymous. The telephone auto-commutator. Scientific American, 84(6):85–86, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
85.pdf.

Anonymous:1901:HAY

[3042] Anonymous. The Hamburg American yacht “Prinzessin Victoria Luise”. Scientific American, 84(6):86, February 9, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n6/pdf/scientificamerican02091901-86.pdf.

Anonymous:1901:WCM

[3043] Anonymous. What can be made of orange peel. Scientific American, 84(6):88, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
88.pdf.

Anonymous:1901:PWL

[3044] Anonymous. The proposed widening of London Bridge. Scientific American, 84(6):88, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
88a.pdf.

Fawcett:1901:CC

[3045] Waldon Fawcett. The canals of Canada. Scientific American, 84 (6):89–90, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
89.pdf.

Anonymous:1901:OCS

[3046] Anonymous. The Old Cutler’s Shop at the Exposition. Scientific American, 84(6):90, February 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n6/pdf/scientificamerican02091901-
90.pdf.

Anonymous:1901:SNm

[3047] Anonymous. Science notes. Scientific American, 84(7):99, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n7/pdf/scientificamerican02161901-99a.pdf.

Anonymous:1901:Ck

[3048] Anonymous. Correspondence. Scientific American, 84(7):103, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n7/pdf/scientificamerican02161901-103.pdf.

Anonymous:1901:RPIx

[3049] Anonymous. Recently patented inventions. Scientific American, 84 (7):107–108, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
107.pdf.

Anonymous:1901:BNBo

[3050] Anonymous. Books: New books, etc. Scientific American, 84 (7):108, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
108.pdf.

Anonymous:1901:NQx

[3051] Anonymous. Notes and queries. Scientific American, 84(7):108– 109, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
108b.pdf.

Anonymous:1901:IIx

[3052] Anonymous. Index of inventions. Scientific American, 84(7):109, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n7/pdf/scientificamerican02161901-109.pdf.

Anonymous:1901:PNPw

[3053] Anonymous. Patent notes: Patents. Scientific American, 84 (7):111, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
111.pdf.

Anonymous:1901:STO

[3054] Anonymous. Superposed turrets for our latest battle-ships. Scientific American, 84(7):98, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
98.pdf.

Anonymous:1901:ATL

[3055] Anonymous. An aluminium transmission line. Scientific American, 84(7):98, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
98a.pdf.

Anonymous:1901:PBCb

[3056] Anonymous. A point in boiler construction. Scientific American, 84(7):98, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
98b.pdf.

Anonymous:1901:VUT

[3057] Anonymous. Vibration from underground trains. Scientific American, 84(7):98, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
98c.pdf.

Anonymous:1901:ARC

[3058] Anonymous. The Annual Report of the Commissioner of Patents. Scientific American, 84(7):98, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
98d.pdf.

Anonymous:1901:ELD

[3059] Anonymous. England’s long-deferred decadence? Scientific American, 84(7):98–99, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
98e.pdf.

Headland:1901:UCC

[3060] Isaac Taylor Headland. The unscientific character of Chinese play. Scientific American, 84(7):99, February 16, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n7/pdf/scientificamerican02161901-99.pdf.

Anonymous:1901:VLB

[3061] Anonymous. Vanderbilt locomotive boiler. Scientific American, 84(7):100, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
100.pdf.

Anonymous:1901:CPB

[3062] Anonymous. A combination pipe and bench vise. Scientific American, 84(7):101, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
101.pdf.

Anonymous:1901:STW

[3063] Anonymous. Steering torpedoes by wireless telegraphy. Scientific American, 84(7):101–102, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
101a.pdf.

Anonymous:1901:MWGb

[3064] Anonymous. Manufacture of water-gas at the Forty Fourth Street Station, New York — II. Scientific American, 84(7):102, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n7/pdf/scientificamerican02161901-102.pdf.

Fawcett:1901:IWW

[3065] Waldon Fawcett. Ingenious water-wheels. Scientific American, 84 (7):104–105, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
104.pdf.

Anonymous:1901:TM

[3066] Anonymous. The truncated man. Scientific American, 84(7):105, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n7/pdf/scientificamerican02161901-105.pdf.

Moore:1901:MRR

[3067] N. Hudson Moore. The Mantis Religiosa in Rochester, N.Y. Scientific American, 84(7):105–106, February 16, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n7/pdf/scientificamerican02161901-105a.pdf.

Anonymous:1901:NH

[3068] Anonymous. A novel hinge. Scientific American, 84(7):106, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n7/pdf/scientificamerican02161901-106.pdf.

Anonymous:1901:NFI

[3069] Anonymous. A new form of insulator. Scientific American, 84 (7):106, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
106a.pdf.

Anonymous:1901:IDE

[3070] Anonymous. An instrument for drawing ellipses. Scientific American, 84(7):106–107, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
106b.pdf.

Anonymous:1901:BPWr

[3071] Anonymous. Business and personal wants. Scientific American, 84(7):108, February 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n7/pdf/scientificamerican02161901-
108a.pdf.

Anonymous:1901:SNn

[3072] Anonymous. Science notes. Scientific American, 84(8):115, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n8/pdf/scientificamerican02231901-115d.pdf.

Anonymous:1901:RPIy

[3073] Anonymous. Recently patented inventions. Scientific American, 84(8):123, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
123.pdf.

Anonymous:1901:BNBp

[3074] Anonymous. Books: New books, etc. Scientific American, 84 (8):123, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
123a.pdf.

Anonymous:1901:NQy

[3075] Anonymous. Notes and queries. Scientific American, 84(8):124, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n8/pdf/scientificamerican02231901-124.pdf.

Anonymous:1901:IIy

[3076] Anonymous. Index of inventions. Scientific American, 84(8):124, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n8/pdf/scientificamerican02231901-124a.pdf.

Anonymous:1901:PNPx

[3077] Anonymous. Patent notes: Patents. Scientific American, 84 (8):127, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
127.pdf.

Anonymous:1901:ERR

[3078] Anonymous. East River Rapid Transit Tunnel. Scientific American, 84(8):114, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
114.pdf.

Anonymous:1901:SWW

[3079] Anonymous. Sir William White and the new Royal Yacht. Scientific American, 84(8):114, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
114a.pdf.

Anonymous:1901:TWTa

[3080] Anonymous. Tesla’s wireless telegraphy. Scientific American, 84 (8):114, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
114b.pdf.

Anonymous:1901:HSP

[3081] Anonymous. High-speed passenger steamers on the Hudson. Scientific American, 84(8):114, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
114c.pdf.

Anonymous:1901:SPG

[3082] Anonymous. Smokeless powder and gun erosion. Scientific American, 84(8):114, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
114d.pdf.

Anonymous:1901:PIL

[3083] Anonymous. Prizes for inventions a legitimate encouragement. Scientific American, 84(8):114–115, February 23, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n8/pdf/scientificamerican02231901-114e.pdf.

Anonymous:1901:YAB

[3084] Anonymous. Are you about to build? Scientific American, 84 (8):115, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
115.pdf.

Anonymous:1901:ZG

[3085] Anonymous. Zenobe T. Gramme. Scientific American, 84(8): 115, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
115a.pdf.

Anonymous:1901:OBP

[3086] Anonymous. Our business and personal want column. Scientific American, 84(8):115, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
115b.pdf.

Anonymous:1901:ZL

[3087] Anonymous. Zodiacal light. Scientific American, 84(8):115, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n8/pdf/scientificamerican02231901-115c.pdf.

Anonymous:1901:PNT

[3088] Anonymous. Powerful nine-ton dipper dredge. Scientific American, 84(8):116, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
116.pdf.

Anonymous:1901:PIB

[3089] Anonymous. Petroleum as an illuminant for buoys. Scientific American, 84(8):117, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
117.pdf.

Anonymous:1901:OME

[3090] Anonymous. The orchestra-man of the Exposition. Scientific American, 84(8):117–118, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
117a.pdf.

Anonymous:1901:SM

[3091] Anonymous. Silk manufacture. Scientific American, 84(8):118– 119, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
118.pdf.

Anonymous:1901:DA

[3092] Anonymous. De Dietrich automobiles. Scientific American, 84 (8):120, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
120.pdf.

Holder:1901:RW

[3093] Charles Frederick Holder. Running down whales. Scientific American, 84(8):121, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
121.pdf.

Anonymous:1901:GNB

[3094] Anonymous. A glimpse of the Naval Battle of Santiago. Scientific American, 84(8):121–122, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
121a.pdf.

Anonymous:1901:TRPc

[3095] Anonymous. Three recently patented novelties. Scientific American, 84(8):122, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
122.pdf.

Anonymous:1901:RBS

[3096] Anonymous. Repairing a broken shaft at sea. Scientific American, 84(8):122, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
122a.pdf.

Anonymous:1901:SEC

[3097] Anonymous. A simple and efficient clipper grinder and surface grinder. Scientific American, 84(8):122, February 23, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v84/
n8/pdf/scientificamerican02231901-122b.pdf.

Anonymous:1901:BPWs

[3098] Anonymous. Business and personal wants. Scientific American, 84(8):123, February 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n8/pdf/scientificamerican02231901-
123b.pdf.

Russell:1901:HMb

[3099] Henry Norris Russell. The heavens in March, 1901. Scientific American, 84(9):131, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
131a.pdf.

Anonymous:1901:Cl

[3100] Anonymous. Correspondence. Scientific American, 84(9):135–136, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-135.pdf.

Anonymous:1901:RPIz

[3101] Anonymous. Recently patented inventions. Scientific American, 84(9):139, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
139.pdf.

Anonymous:1901:BNBq

[3102] Anonymous. Books: New books, etc. Scientific American, 84(9):139, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-139a.pdf.

Anonymous:1901:NQz

[3103] Anonymous. Notes and queries. Scientific American, 84(9):140, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-140.pdf.

Anonymous:1901:IIz

[3104] Anonymous. Index of inventions. Scientific American, 84(9):140, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-140a.pdf.

Anonymous:1901:PNPy

[3105] Anonymous. Patent notes: Patents. Scientific American, 84(9):143, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-143.pdf.

Anonymous:1901:CNEb

[3106] Anonymous. Cables for the new East River Bridge. Scientific American, 84(9):130, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
130.pdf.

Anonymous:1901:RRA

[3107] Anonymous. The retirement of Rear-Admiral Hichborn. Scientific American, 84(9):130, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
130a.pdf.

Anonymous:1901:ACa

[3108] Anonymous. Aeronautical Congress. Scientific American, 84(9):130, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-130b.pdf.

Anonymous:1901:LP

[3109] Anonymous. Lotus poison. Scientific American, 84(9):130, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-130c.pdf.

Anonymous:1901:ETA

[3110] Anonymous. Estimation of tungstic acid in its ores. Scientific American, 84(9):131, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
131.pdf.

Anonymous:1901:MSE

[3111] Anonymous. Musical sounds from the electric arc. Scientific American, 84(9):131, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
131b.pdf.

Anonymous:1901:ARV

[3112] Anonymous. An automatic reducing valve of novel form. Scientific American, 84(9):132, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
132.pdf.

Anonymous:1901:LCW

[3113] Anonymous. The largest camera in the world. Scientific American, 84(9):132, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
132a.pdf.

Anonymous:1901:POB

[3114] Anonymous. The potter of “Old Berry” at the Exposition. Scientific American, 84(9):133, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
133.pdf.

Anonymous:1901:PCa

[3115] Anonymous. Pod corn. Scientific American, 84(9):133, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-133a.pdf.

Clay:1901:EN

[3116] Albert T. Clay. Excavations at Nippur. Scientific American, 84 (9):133–134, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
133b.pdf.

Anonymous:1901:KAb

[3117] Anonymous. The Kress aeroplane. Scientific American, 84(9):137, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v84/n9/pdf/scientificamerican03021901-137.pdf.

Anonymous:1901:VGN

[3118] Anonymous. The vertical growth of New York City. Scientific American, 84(9):137–138, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
137a.pdf.

Anonymous:1901:SDL

[3119] Anonymous. A simple door latch for refrigerators. Scientific American, 84(9):138, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
138.pdf.

Anonymous:1901:TJP

[3120] Anonymous. A triple-jawed pipe-wrench. Scientific American, 84 (9):138, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
138a.pdf.

Anonymous:1901:ICM

[3121] Anonymous. An improvement in capsuling-machines. Scientific American, 84(9):138, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
138b.pdf.

Anonymous:1901:BPWt

[3122] Anonymous. Business and personal wants. Scientific American, 84(9):139, March 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v84/n9/pdf/scientificamerican03021901-
139b.pdf.

Anonymous:1901:Cm

[3123] Anonymous. Correspondence. Scientific American, 85(1):7, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-7.pdf.

Anonymous:1901:RPIba

[3124] Anonymous. Recently patented inventions. Scientific American, 85(1): 11, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-11.pdf.

Anonymous:1901:BNBr

[3125] Anonymous. Books: New books, etc. Scientific American, 85(1):11, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-11a.pdf.

Anonymous:1901:NQba

[3126] Anonymous. Notes and queries. Scientific American, 85(1):12, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-12.pdf.

Anonymous:1901:IIba

[3127] Anonymous. Index of inventions. Scientific American, 85(1):13, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-13.pdf.

Anonymous:1901:PNPz

[3128] Anonymous. Patent notes: Patents. Scientific American, 85(1):15, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-15.pdf.

Anonymous:1901:HEM

[3129] Anonymous. Handling excavated material at Jerome Park Reservoir. Scientific American, 85(1):1, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n1/pdf/scientificamerican07061901-
1.pdf.

Anonymous:1901:RCM

[3130] Anonymous. Railroad connections with Manhattan Island. Scientific American, 85(1):2, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n1/pdf/scientificamerican07061901-
2.pdf.

Anonymous:1901:ALE

[3131] Anonymous. American locomotives in England. Scientific American, 85(1):2, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n1/pdf/scientificamerican07061901-
2a.pdf.

Anonymous:1901:HOW

[3132] Anonymous. Height of ocean waves. Scientific American, 85(1):2, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-2b.pdf.

Anonymous:1901:SCc

[3133] Anonymous. Selenides of cobalt. Scientific American, 85(1):2–3, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-2c.pdf.

Adams:1901:CSA

[3134] Alton D. Adams. Coke as a substitute for anthracite coal. Scientific American, 85(1):3, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n1/pdf/scientificamerican07061901-
3.pdf.

Anonymous:1901:SNo

[3135] Anonymous. Science notes. Scientific American, 85(1):3, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-3a.pdf.

Anonymous:1901:WFQ

[3136] Anonymous. A Wyoming fossil quarry. Scientific American, 85(1):4, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-4.pdf.

Anonymous:1901:SDH

[3137] Anonymous. A starting-device for hydrocarbon motors. Scientific American, 85(1):5, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n1/pdf/scientificamerican07061901-
5.pdf.

Anonymous:1901:TPC

[3138] Anonymous. A trolley palace car. Scientific American, 85(1):5, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-5a.pdf.

Anonymous:1901:NTH

[3139] Anonymous. A novel twine holder and cutter. Scientific American, 85(1):5, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n1/pdf/scientificamerican07061901-
5b.pdf.

Holder:1901:PME

[3140] Charles F. Holder. Pyramids of Del Mar — erosion on the Pacific Coast. Scientific American, 85(1):8, July 6, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n1/pdf/scientificamerican07061901-8.pdf.

Anonymous:1901:TAT

[3141] Anonymous. The Thomas auto-tricycle. Scientific American, 85(1):9, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n1/pdf/scientificamerican07061901-9.pdf.

Anonymous:1901:LBG

[3142] Anonymous. The lance-boats of the German cavalry. Scientific American, 85(1):9–10, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n1/pdf/scientificamerican07061901-
9a.pdf.

Anonymous:1901:NIM

[3143] Anonymous. A new instrument for the measurement of high temperatures. Scientific American, 85(1):10, July 6, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n1/pdf/scientificamerican07061901-10.pdf.

Anonymous:1901:BPWu

[3144] Anonymous. Business and personal wants. Scientific American, 85(1):11–12, July 6, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n1/pdf/scientificamerican07061901-
11b.pdf.

Anonymous:1901:RPIbb

[3145] Anonymous. Recently patented inventions. Scientific American, 85(10):156, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
156.pdf.

Anonymous:1901:BNBs

[3146] Anonymous. Books: New books, etc. Scientific American, 85 (10):156, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
156a.pdf.

Anonymous:1901:NQbb

[3147] Anonymous. Notes and queries. Scientific American, 85(10): 157, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
157.pdf.

Anonymous:1901:IIbb

[3148] Anonymous. Index of inventions. Scientific American, 85(10): 157, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
157a.pdf.

Anonymous:1901:PNPba

[3149] Anonymous. Patent notes: Patents. Scientific American, 85 (10):157, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
157b.pdf.

Anonymous:1901:PTG

[3150] Anonymous. A promising type of gun. Scientific American, 85 (10):146, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
146.pdf.

Anonymous:1901:SCMb

[3151] Anonymous. Some chemical mysteries. Scientific American, 85 (10):146, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
146a.pdf.

Anonymous:1901:DP

[3152] Anonymous. The design of propellers. Scientific American, 85 (10):146, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
146b.pdf.

Anonymous:1901:MTB

[3153] Anonymous. Marconi telegraphs on the British Fleet. Scientific American, 85(10):146–147, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
146c.pdf.

Hale:1901:PAM

[3154] William H. Hale. Pan-American Meeting of the American Institute of Electrical Engineers. Scientific American, 85(10):147, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n10/pdf/scientificamerican09071901-147.pdf.

Chamberlain:1901:EM

[3155] John Chamberlain. Extermination of the mosquito. Scientific American, 85(10):147, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
147a.pdf.

Wright:1901:SHM

[3156] Budd A. Wright. A simple home-made lathe. Scientific American, 85(10):148, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
148.pdf.

Anonymous:1901:LALb

[3157] Anonymous. A laboratory apparatus for liquefying gases. Scientific American, 85(10):148, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
148a.pdf.

Anonymous:1901:MMb

[3158] Anonymous. The manufacture of matches. Scientific American, 85 (10):149–150, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
149.pdf.

Adams:1901:DCE

[3159] Alton D. Adams. Distribution of combined electrical energy. Scientific American, 85(10):150–151, September 7, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n10/pdf/scientificamerican09071901-150.pdf.

Anonymous:1901:NIS

[3160] Anonymous. New 5-inch segmental wire-wound gun for the United States Army. Scientific American, 85(10):152–153, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n10/pdf/scientificamerican09071901-152.pdf.

Anonymous:1901:WAB

[3161] Anonymous. The Willis Avenue Bridge over the Harlem River. Scientific American, 85(10):153–154, September 7, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n10/pdf/scientificamerican09071901-153.pdf.

Benjamin:1901:HMM

[3162] Marcus Benjamin. How money is made. Scientific American, 85 (10):154–155, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
154.pdf.

Anonymous:1901:WSD

[3163] Anonymous. The wreck of the Santos-Dumont balloon. Scientific American, 85(10):155, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
155.pdf.

Anonymous:1901:BPWv

[3164] Anonymous. Business and personal wants. Scientific American, 85(10):156, September 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n10/pdf/scientificamerican09071901-
156b.pdf.

Anonymous:1901:Cn

[3165] Anonymous. Correspondence. Scientific American, 85(11):166, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n11/pdf/scientificamerican09141901-166.pdf.

Anonymous:1901:RPIbc

[3166] Anonymous. Recently patented inventions. Scientific American, 85 (11):172, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
172.pdf.

Anonymous:1901:BNBt

[3167] Anonymous. Books: New books, etc. Scientific American, 85 (11):172, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
172a.pdf.

Anonymous:1901:NQbc

[3168] Anonymous. Notes and queries. Scientific American, 85(11): 173, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
173.pdf.

Anonymous:1901:IIbc

[3169] Anonymous. Index of inventions. Scientific American, 85(11): 173, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
173b.pdf.

Anonymous:1901:PNPbb

[3170] Anonymous. Patent notes: Patents. Scientific American, 85 (11):174, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
174.pdf.

Anonymous:1901:AP

[3171] Anonymous. The attempt on the President. Scientific American, 85(11):162, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
162.pdf.

Anonymous:1901:CTC

[3172] Anonymous. “Columbia” — twice a Cup defender. Scientific American, 85(11):162, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
162a.pdf.

Anonymous:1901:TW

[3173] Anonymous. A timely warning. Scientific American, 85(11): 162, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
162b.pdf.

Anonymous:1901:TPA

[3174] Anonymous. Target practice at Aldershot. Scientific American, 85 (11):162, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
162c.pdf.

Hovey:1901:AAA

[3175] E. O. Hovey. The American Association for the Advancement of Science. Scientific American, 85(11):163, September 14, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n11/pdf/scientificamerican09141901-163.pdf.

Anonymous:1901:BCT

[3176] Anonymous. The best class trade paper. Scientific American, 85 (11):163, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
163a.pdf.

Anonymous:1901:NAL

[3177] Anonymous. The new Atlantic liner “Kronprinz Wilhelm”. Scientific American, 85(11):164, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
164.pdf.

Anonymous:1901:ICL

[3178] Anonymous. An improved chalk-line holder. Scientific American, 85(11):165, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
165.pdf.

Anonymous:1901:LSR

[3179] Anonymous. The Loschwitz suspension railway. Scientific American, 85(11):165, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
165a.pdf.

Anonymous:1901:GP

[3180] Anonymous. A grease-pump. Scientific American, 85(11):165, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n11/pdf/scientificamerican09141901-165b.pdf.

Anonymous:1901:TFN

[3181] Anonymous. The temporary footways of the new East River Bridge. Scientific American, 85(11):166–167, September 14, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n11/pdf/scientificamerican09141901-166a.pdf.

Anonymous:1901:FHG

[3182] Anonymous. The five-horned giraffe. Scientific American, 85 (11):168, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
168.pdf.

Anonymous:1901:AMa

[3183] Anonymous. The ascent of the Matterhorn. Scientific American, 85 (11):168–169, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
168a.pdf.

Anonymous:1901:NAR

[3184] Anonymous. The Newport Automobile Races. Scientific American, 85 (11):169–170, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
169.pdf.

Collins:1901:HCE

[3185] A. Frederick Collins. How to construct an efficient wireless telegraph apparatus at a small cost. Scientific American, 85(11):170–171, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n11/pdf/scientificamerican09141901-170.pdf.

Anonymous:1901:BPWw

[3186] Anonymous. Business and personal wants. Scientific American, 85 (11):172, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n11/pdf/scientificamerican09141901-
172b.pdf.

Anonymous:1901:Id

[3187] Anonymous. To inventors. Scientific American, 85(11):173, September 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n11/pdf/scientificamerican09141901-173a.pdf.

Anonymous:1901:Co

[3188] Anonymous. Correspondence. Scientific American, 85(12):183, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n12/pdf/scientificamerican09211901-183.pdf.

Anonymous:1901:RPIbd

[3189] Anonymous. Recently patented inventions. Scientific American, 85 (12):188, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
188.pdf.

Anonymous:1901:NQbd

[3190] Anonymous. Notes and queries. Scientific American, 85(12): 189, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
189.pdf.

Anonymous:1901:IIbd

[3191] Anonymous. Index of inventions. Scientific American, 85(12): 189, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
189b.pdf.

Anonymous:1901:PNPbc

[3192] Anonymous. Patent notes: Patents. Scientific American, 85 (12):190, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
190.pdf.

Anonymous:1901:GWD

[3193] Anonymous. “God’s will be done — not ours”. Scientific American, 85(12):178, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
178.pdf.

Anonymous:1901:PRC

[3194] Anonymous. Proposed remodeling of the Croton Dam. Scientific American, 85(12):178, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
178a.pdf.

Anonymous:1901:GSF

[3195] Anonymous. A great salvage feat. Scientific American, 85(12): 178, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
178b.pdf.

Anonymous:1901:CGU

[3196] Anonymous. Canaigre growing in the United States. Scientific American, 85(12):178, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
178c.pdf.

Anonymous:1901:RDW

[3197] Anonymous. Rapid development of wireless telegraphy. Scientific American, 85(12):178–179, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
178d.pdf.

Anonymous:1901:NFL

[3198] Anonymous. The new French liner “La Savoie”. Scientific American, 85(12):179, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
179.pdf.

Anonymous:1901:GCD

[3199] Anonymous. Germany’s commercial depression — its causes and effects. Scientific American, 85(12):179, September 21, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n12/pdf/scientificamerican09211901-179a.pdf.

Anonymous:1901:SNp

[3200] Anonymous. Science notes. Scientific American, 85(12):179, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n12/pdf/scientificamerican09211901-179b.pdf.

Anonymous:1901:NDA

[3201] Anonymous. A new driving axle for automobiles. Scientific American, 85(12):180, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
180.pdf.

Anonymous:1901:SPA

[3202] Anonymous. Self-propelling automatic grab bucket. Scientific American, 85(12):180–181, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
180a.pdf.

Anonymous:1901:HBL

[3203] Anonymous. House boat “Loudoun”. Scientific American, 85 (12):181, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
181.pdf.

Anonymous:1901:AEB

[3204] Anonymous. An adjustable electric book lamp. Scientific American, 85(12):181–182, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
181a.pdf.

Anonymous:1901:CDW

[3205] Anonymous. The Croton Dam for the water supply of New York City. Scientific American, 85(12):182, September 21, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n12/pdf/scientificamerican09211901-182.pdf.

Anonymous:1901:IHIb

[3206] Anonymous. Improved hydrographic instruments. Scientific American, 85(12):184, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
184.pdf.

Anonymous:1901:MCa

[3207] Anonymous. The manufacture of caramels. Scientific American, 85 (12):185–186, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
185.pdf.

Holder:1901:CMD

[3208] Charles Frederick Holder. A curious means of defense. Scientific American, 85(12):186–187, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
186.pdf.

Anonymous:1901:IUO

[3209] Anonymous. Increasing use of oil on highways. Scientific American, 85(12):187, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
187.pdf.

Anonymous:1901:BPWx

[3210] Anonymous. Business and personal wants. Scientific American, 85 (12):188, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n12/pdf/scientificamerican09211901-
188a.pdf.

Anonymous:1901:Ie

[3211] Anonymous. To inventors. Scientific American, 85(12):189, September 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n12/pdf/scientificamerican09211901-189a.pdf.

Russell:1901:HO

[3212] Henry Norris Russell. The heavens in October. Scientific American, 85(13):195, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
195.pdf.

Anonymous:1901:Cp

[3213] Anonymous. Correspondence. Scientific American, 85(13):198– 199, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
198a.pdf.

Anonymous:1901:RPIbe

[3214] Anonymous. Recently patented inventions. Scientific American, 85 (13):204, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
204.pdf.

Anonymous:1901:NQbe

[3215] Anonymous. Notes and queries. Scientific American, 85(13): 204, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
204b.pdf.

Anonymous:1901:IIbe

[3216] Anonymous. Index of inventions. Scientific American, 85(13): 205, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
205a.pdf.

Anonymous:1901:PNPbd

[3217] Anonymous. Patent notes: Patents. Scientific American, 85 (13):207, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
207.pdf.

Anonymous:1901:PAD

[3218] Anonymous. Peary’s Arctic discoveries. Scientific American, 85 (13):194, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
194.pdf.

Anonymous:1901:CSb

[3219] Anonymous. “Columbia” and “Shamrock”. Scientific American, 85 (13):194, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
194a.pdf.

Anonymous:1901:NYBb

[3220] Anonymous. New York–Buffalo Endurance Race. Scientific American, 85(13):194, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
194b.pdf.

Anonymous:1901:MSDa

[3221] Anonymous. The M. Santos-Dumont Balloon No. 6 fails. Scientific American, 85(13):195, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
195a.pdf.

Anonymous:1901:SNq

[3222] Anonymous. Science notes. Scientific American, 85(13):195, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n13/pdf/scientificamerican09281901-195b.pdf.

Anonymous:1901:EWHb

[3223] Anonymous. An electrical water-heater. Scientific American, 85 (13):196, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
196.pdf.

Anonymous:1901:IFD

[3224] Anonymous. An improved fertilizer-distributer. Scientific American, 85(13):196, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
196a.pdf.

Anonymous:1901:HIO

[3225] Anonymous. Harbor improvement, Oakland, Cal. Scientific American, 85(13):196–197, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
196b.pdf.

Anonymous:1901:CLE

[3226] Anonymous. Curious locomotive explosion. Scientific American, 85 (13):197, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
197.pdf.

Anonymous:1901:LCb

[3227] Anonymous. A locomotive curiosity. Scientific American, 85(13): 197–198, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
197a.pdf.

Fawcett:1901:RLB

[3228] Waldon Fawcett. Rolling lift bridges. Scientific American, 85 (13):198, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
198.pdf.

Anonymous:1901:FFS

[3229] Anonymous. Foster’s fog signals. Scientific American, 85(13): 200, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
200.pdf.

Dunlap:1901:FAP

[3230] Orrin E. Dunlap. Foolhardy attempts at passing the Whirlpool Rapids of Niagara. Scientific American, 85(13):201–202, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n13/pdf/scientificamerican09281901-201.pdf.

Anonymous:1901:DBL

[3231] Anonymous. A destructive bolt of lightning. Scientific American, 85(13):202, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
202.pdf.

Anonymous:1901:NMC

[3232] Anonymous. A new motor cycle. Scientific American, 85(13): 203, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
203.pdf.

Anonymous:1901:BPWy

[3233] Anonymous. Business and personal wants. Scientific American, 85 (13):204, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n13/pdf/scientificamerican09281901-
204a.pdf.

Anonymous:1901:Ih

[3234] Anonymous. To inventors. Scientific American, 85(13):205, September 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n13/pdf/scientificamerican09281901-205.pdf.

Anonymous:1901:RPIbf

[3235] Anonymous. Recently patented inventions. Scientific American, 85(14):220, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
220.pdf.

Anonymous:1901:BNBu

[3236] Anonymous. Books: New books, etc. Scientific American, 85(14):220, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n14/pdf/scientificamerican10051901-220b.pdf.

Anonymous:1901:NQbf

[3237] Anonymous. Notes and queries. Scientific American, 85(14):220, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n14/pdf/scientificamerican10051901-220c.pdf.

Anonymous:1901:IIbf

[3238] Anonymous. Index of inventions. Scientific American, 85(14):220, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n14/pdf/scientificamerican10051901-220d.pdf.

Anonymous:1901:PNPbe

[3239] Anonymous. Patent notes: Patents. Scientific American, 85(14):223, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n14/pdf/scientificamerican10051901-223.pdf.

Anonymous:1901:TDC

[3240] Anonymous. Trade Dollar Consolidated Mining Company’s Electric Plant, Snake River, Idaho. Scientific American, 85(14):209, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n14/pdf/scientificamerican10051901-209.pdf.

Anonymous:1901:TTD

[3241] Anonymous. The towing tank and the deep sea. Scientific American, 85(14):210, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
210.pdf.

Anonymous:1901:RTA

[3242] Anonymous. The Royal tombs at Abydos. Scientific American, 85(14):210, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
210a.pdf.

Anonymous:1901:TPV

[3243] Anonymous. The true point of view. Scientific American, 85(14):210, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n14/pdf/scientificamerican10051901-210b.pdf.

Anonymous:1901:HSB

[3244] Anonymous. The Holland submarines for the British Navy. Scientific American, 85(14):210–211, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
210c.pdf.

Anonymous:1901:BVC

[3245] Anonymous. Belleville versus cylindrical boilers. Scientific American, 85(14):211, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
211.pdf.

Anonymous:1901:SNr

[3246] Anonymous. Science notes. Scientific American, 85(14):211, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n14/pdf/scientificamerican10051901-211a.pdf.

Anonymous:1901:FKT

[3247] Anonymous. The “Fool Killer” taking soundings. Scientific American, 85(14):211–212, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
211b.pdf.

Anonymous:1901:ETF

[3248] Anonymous. Eighty-ton floating crane for the Santos harbor works. Scientific American, 85(14):212, October 5, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n14/pdf/scientificamerican10051901-212.pdf.

Anonymous:1901:EGF

[3249] Anonymous. An electrical glass-furnace. Scientific American, 85 (14):212–213, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
212a.pdf.

Anonymous:1901:SDN

[3250] Anonymous. The “Santos-Dumont No. 6”. Scientific American, 85 (14):213–214, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
213.pdf.

Colvin:1901:EGM

[3251] Wilber Colvin. Electricity in gold mining. Scientific American, 85 (14):216–217, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
216.pdf.

Anonymous:1901:FYA

[3252] Anonymous. Fifty years of “America” Cup Contests. Scientific American, 85(14):217–218, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
217.pdf.

Anonymous:1901:STM

[3253] Anonymous. Synthetic talking machine. Scientific American, 85 (14):219, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
219.pdf.

Anonymous:1901:BPWz

[3254] Anonymous. Business and personal wants. Scientific American, 85(14):220, October 5, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n14/pdf/scientificamerican10051901-
220a.pdf.

Anonymous:1901:RPIbg

[3255] Anonymous. Recently patented inventions. Scientific American, 85(15):236, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
236.pdf.

Anonymous:1901:BNBv

[3256] Anonymous. Books: New books, etc. Scientific American, 85 (15):236, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
236a.pdf.

Anonymous:1901:NQbg

[3257] Anonymous. Notes and queries. Scientific American, 85(15):237, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n15/pdf/scientificamerican10121901-237.pdf.

Anonymous:1901:IIbg

[3258] Anonymous. Index of inventions. Scientific American, 85(15):237, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n15/pdf/scientificamerican10121901-237a.pdf.

Anonymous:1901:PNPbf

[3259] Anonymous. Patent notes: Patents. Scientific American, 85 (15):239, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
239.pdf.

Anonymous:1901:LVC

[3260] Anonymous. The loss of the “Viper” and the “Cobra”. Scientific American, 85(15):226, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
226.pdf.

Anonymous:1901:TCLa

[3261] Anonymous. The tandem compound locomotive. Scientific American, 85(15):226, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
226a.pdf.

Anonymous:1901:SER

[3262] Anonymous. Some of the English Royal treasures. Scientific American, 85(15):226–227, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
226b.pdf.

Anonymous:1901:EPI

[3263] Anonymous. Extensive project for irrigating Egypt. Scientific American, 85(15):227, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
227.pdf.

Anonymous:1901:IGF

[3264] Anonymous. Inventing games and forms of entertainment. Scientific American, 85(15):227, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
227a.pdf.

Anonymous:1901:EAb

[3265] Anonymous. The Ezekiel airship. Scientific American, 85(15):228, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n15/pdf/scientificamerican10121901-228.pdf.

Anonymous:1901:GSG

[3266] Anonymous. Great sluice gates for the Nile irrigation works. Scientific American, 85(15):228, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
228a.pdf.

Anonymous:1901:AFW

[3267] Anonymous. An automobile forage wagon. Scientific American, 85(15):229, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
229.pdf.

Anonymous:1901:QAD

[3268] Anonymous. Queen Alexandra’s dairy. Scientific American, 85 (15):229–230, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
229a.pdf.

Anonymous:1901:JS

[3269] Anonymous. Jupiter steel. Scientific American, 85(15):230–231, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n15/pdf/scientificamerican10121901-230.pdf.

Anonymous:1901:NCR

[3270] Anonymous. The new 50-caliber rapid-fire guns of the United States Navy. Scientific American, 85(15):232, October 12, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n15/pdf/scientificamerican10121901-232.pdf.

Anonymous:1901:CAC

[3271] Anonymous. The 1901 contest for the “America” Cup. Scientific American, 85(15):232–235, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
232a.pdf.

Anonymous:1901:SWA

[3272] Anonymous. A skein for wagon axles. Scientific American, 85 (15):235, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
235.pdf.

Anonymous:1901:THC

[3273] Anonymous. Turntable for hand-cars. Scientific American, 85 (15):235, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
235a.pdf.

Anonymous:1901:NF

[3274] Anonymous. A novel filter. Scientific American, 85(15):235, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n15/pdf/scientificamerican10121901-235b.pdf.

Anonymous:1901:FPC

[3275] Anonymous. A fireproof pipe covering. Scientific American, 85 (15):235, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
235c.pdf.

Anonymous:1901:BPWba

[3276] Anonymous. Business and personal wants. Scientific American, 85(15):236, October 12, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n15/pdf/scientificamerican10121901-
236b.pdf.

Anonymous:1901:Cq

[3277] Anonymous. Correspondence. Scientific American, 85(16):247, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n16/pdf/scientificamerican10191901-247.pdf.

Anonymous:1901:RPIbh

[3278] Anonymous. Recently patented inventions. Scientific American, 85(16):251, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
251.pdf.

Anonymous:1901:BNBw

[3279] Anonymous. Books: New books, etc. Scientific American, 85 (16):251, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
251a.pdf.

Anonymous:1901:NQbh

[3280] Anonymous. Notes and queries. Scientific American, 85(16): 252–253, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
252a.pdf.

Anonymous:1901:IIbh

[3281] Anonymous. Index of inventions. Scientific American, 85(16):253, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n16/pdf/scientificamerican10191901-253.pdf.

Anonymous:1901:DNC

[3282] Anonymous. Delayed Naval construction in private yards. Scientific American, 85(16):242, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
242.pdf.

Anonymous:1901:AAC

[3283] Anonymous. The aftermath of the “America” Cup Races. Scientific American, 85(16):242, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
242a.pdf.

Anonymous:1901:DBB

[3284] Anonymous. Deterioration of the Brooklyn Bridge. Scientific American, 85(16):242, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
242b.pdf.

Anonymous:1901:LTR

[3285] Anonymous. The latest transatlantic record breaker. Scientific American, 85(16):242–243, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
242c.pdf.

Irwin:1901:DBC

[3286] William Gilbert Irwin. The development of the bituminous coal industry. Scientific American, 85(16):243, October 19, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n16/pdf/scientificamerican10191901-243.pdf.

Anonymous:1901:MCc

[3287] Anonymous. The manufacture of cordage. Scientific American, 85 (16):244–245, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
244.pdf.

Anonymous:1901:NSE

[3288] Anonymous. A new submerged electric motor and propeller. Scientific American, 85(16):245, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
245.pdf.

Anonymous:1901:NATa

[3289] Anonymous. A new automatic telegraph repeater. Scientific American, 85(16):245–246, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
245a.pdf.

Anonymous:1901:CW

[3290] Anonymous. A curious windmill. Scientific American, 85(16):246, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n16/pdf/scientificamerican10191901-246.pdf.

Anonymous:1901:BMN

[3291] Anonymous. Burial monuments of North Africa. Scientific American, 85(16):248, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
248.pdf.

Anonymous:1901:SMSb

[3292] Anonymous. A seven-masted steel schooner. Scientific American, 85(16):248–249, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
248a.pdf.

Anonymous:1901:CEP

[3293] Anonymous. A curious example of petrifaction. Scientific American, 85(16):249, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
249.pdf.

Scott:1901:PEA

[3294] A. C. Scott. Photographing the electric arc. Scientific American, 85(16):249–250, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
249a.pdf.

Anonymous:1901:ONDa

[3295] Anonymous. The okapi — a newly-discovered animal. Scientific American, 85(16):250, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
250.pdf.

Anonymous:1901:BPWbb

[3296] Anonymous. Business and personal wants. Scientific American, 85(16):252, October 19, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n16/pdf/scientificamerican10191901-
252.pdf.

Anonymous:1901:RPIbi

[3297] Anonymous. Recently patented inventions. Scientific American, 85(17):268, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
268.pdf.

Anonymous:1901:BNBx

[3298] Anonymous. Books: New books, etc. Scientific American, 85 (17):268, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
268a.pdf.

Anonymous:1901:NQbi

[3299] Anonymous. Notes and queries. Scientific American, 85(17):269, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n17/pdf/scientificamerican10261901-269.pdf.

Anonymous:1901:IIbi

[3300] Anonymous. Index of inventions. Scientific American, 85(17):269, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n17/pdf/scientificamerican10261901-269a.pdf.

Anonymous:1901:PNPbg

[3301] Anonymous. Patent notes: Patents. Scientific American, 85 (17):271, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
271.pdf.

Anonymous:1901:PRB

[3302] Anonymous. Proposed relief of the Brooklyn Bridge Terminal. Scientific American, 85(17):258, October 26, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n17/pdf/scientificamerican10261901-258.pdf.

Anonymous:1901:NE

[3303] Anonymous. Naval estimates. Scientific American, 85(17):258, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n17/pdf/scientificamerican10261901-258a.pdf.

Anonymous:1901:VLU

[3304] Anonymous. Vibrations of London Underground. Scientific American, 85(17):258, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
258b.pdf.

Anonymous:1901:MCS

[3305] Anonymous. Mining conditions in South Africa. Scientific American, 85(17):258–259, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
258c.pdf.

Anonymous:1901:FWA

[3306] Anonymous. Feeding the wild animals. Scientific American, 85 (17):259, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
259.pdf.

Anonymous:1901:RGP

[3307] Anonymous. Roof gardens on private houses. Scientific American, 85(17):259, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
259a.pdf.

Anonymous:1901:SNs

[3308] Anonymous. Science notes. Scientific American, 85(17):259, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n17/pdf/scientificamerican10261901-259b.pdf.

Anonymous:1901:CPM

[3309] Anonymous. A clever piece of molding. Scientific American, 85 (17):260, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
260.pdf.

Anonymous:1901:WSS

[3310] Anonymous. A wooden smelter stack. Scientific American, 85 (17):260, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
260a.pdf.

Anonymous:1901:LWP

[3311] Anonymous. A large wood pulley. Scientific American, 85(17):260, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n17/pdf/scientificamerican10261901-260b.pdf.

Anonymous:1901:ELM

[3312] Anonymous. Electric lifting magnets. Scientific American, 85 (17):261, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
261.pdf.

Anonymous:1901:NEC

[3313] Anonymous. A night excursion car. Scientific American, 85(17): 261–262, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
261a.pdf.

Anonymous:1901:RBR

[3314] Anonymous. The Russian battleship “Retvizan”. Scientific American, 85(17):263, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
263.pdf.

Anonymous:1901:FMF

[3315] Anonymous. French military folding bicycle. Scientific American, 85(17):264, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
264.pdf.

Fawcett:1901:OR

[3316] Waldon Fawcett. Oiled roadbeds. Scientific American, 85(17): 264–265, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
264a.pdf.

Anonymous:1901:CVB

[3317] Anonymous. Count De La Vaulx’s balloon trip across the Mediterranean. Scientific American, 85(17):265–266, October 26, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n17/pdf/scientificamerican10261901-265.pdf.

Beard:1901:SAA

[3318] J. Carter Beard. Something about ancient American saurians. Scientific American, 85(17):267, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
267.pdf.

Anonymous:1901:BPWbc

[3319] Anonymous. Business and personal wants. Scientific American, 85(17):268, October 26, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n17/pdf/scientificamerican10261901-
268b.pdf.

Russell:1901:HN

[3320] Henry Norris Russell. The heavens in November. Scientific American, 85(18):275, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
275.pdf.

Anonymous:1901:Cr

[3321] Anonymous. Correspondence. Scientific American, 85(18):283, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n18/pdf/scientificamerican11021901-283.pdf.

Anonymous:1901:RPIbj

[3322] Anonymous. Recently patented inventions. Scientific American, 85(18):284, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
284.pdf.

Anonymous:1901:NQbj

[3323] Anonymous. Notes and queries. Scientific American, 85(18): 284, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
284b.pdf.

Anonymous:1901:IIbj

[3324] Anonymous. Index of inventions. Scientific American, 85(18): 284, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
284c.pdf.

Anonymous:1901:RCR

[3325] Anonymous. Rapid construction of the rapid transit subway. Scientific American, 85(18):274, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
274.pdf.

Anonymous:1901:PBB

[3326] Anonymous. Painting of big bridges. Scientific American, 85 (18):274, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
274a.pdf.

Anonymous:1901:ARB

[3327] Anonymous. Annual report of the Bureau of Ordnance. Scientific American, 85(18):274, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
274b.pdf.

Anonymous:1901:IAE

[3328] Anonymous. Important application of electric power in railroading. Scientific American, 85(18):274–275, November 2, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n18/pdf/scientificamerican11021901-274c.pdf.

Anonymous:1901:SNt

[3329] Anonymous. Science notes. Scientific American, 85(18):275, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n18/pdf/scientificamerican11021901-275a.pdf.

Anonymous:1901:TYG

[3330] Anonymous. Thirty years’ growth of the locomotive. Scientific American, 85(18):276, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
276.pdf.

Anonymous:1901:TCLb

[3331] Anonymous. Tandem compound locomotive for the Northern Pacific Railway. Scientific American, 85(18):276–277, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n18/pdf/scientificamerican11021901-276a.pdf.

Anonymous:1901:SDA

[3332] Anonymous. A soap-delivery attachment for water pipes. Scientific American, 85(18):277, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
277.pdf.

Brown:1901:CMR

[3333] Enos Brown. A California Marine Railway. Scientific American, 85(18):277, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
277a.pdf.

Anonymous:1901:HMS

[3334] Anonymous. How to make a simple split pulley. Scientific American, 85(18):277, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
277b.pdf.

Anonymous:1901:PCR

[3335] Anonymous. Progress of construction of the rapid transit tunnel. Scientific American, 85(18):278–279, November 2, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n18/pdf/scientificamerican11021901-278.pdf.

Anonymous:1901:WHA

[3336] Anonymous. Waste-heat auxiliary engine. Scientific American, 85 (18):279, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
279.pdf.

Anonymous:1901:ECS

[3337] Anonymous. The electric cable span at Carquinez. Scientific American, 85(18):280–281, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
280.pdf.

Fawcett:1901:BKU

[3338] Waldon Fawcett. Bee-keeping in the United States. Scientific American, 85(18):281–282, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
281.pdf.

Anonymous:1901:IWM

[3339] Anonymous. An ingenious writing machine. Scientific American, 85(18):283, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
283a.pdf.

Frederick:1901:LT

[3340] M. C. Frederick. The leather turtle. Scientific American, 85 (18):283, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
283b.pdf.

Anonymous:1901:BPWbd

[3341] Anonymous. Business and personal wants. Scientific American, 85(18):284, November 2, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n18/pdf/scientificamerican11021901-
284a.pdf.

Anonymous:1901:LNa

[3342] Anonymous. Legal notes. Scientific American, 85(19):300, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n19/pdf/scientificamerican11091901-300.pdf.

Anonymous:1901:RPIbk

[3343] Anonymous. Recently patented inventions. Scientific American, 85(19):301, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
301.pdf.

Anonymous:1901:IIbk

[3344] Anonymous. Index of inventions. Scientific American, 85(19): 301–302, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
301a.pdf.

Anonymous:1901:NQbk

[3345] Anonymous. Notes and queries. Scientific American, 85(19): 302, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
302a.pdf.

Anonymous:1901:BNBy

[3346] Anonymous. Books: New books, etc. Scientific American, 85 (19):305, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
305.pdf.

Anonymous:1901:NDS

[3347] Anonymous. New department of the Scientific American. Scientific American, 85(19):290, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
290.pdf.

Anonymous:1901:FVN

[3348] Anonymous. Forced versus natural draft. Scientific American, 85 (19):290, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
290a.pdf.

Anonymous:1901:AIB

[3349] Anonymous. Accident to the 10-inch brown segmental wire gun. Scientific American, 85(19):290, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
290b.pdf.

Anonymous:1901:MTH

[3350] Anonymous. Marconi telegraphy on the high seas. Scientific American, 85(19):290, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
290c.pdf.

Anonymous:1901:GEP

[3351] Anonymous. Gas engine plants. Scientific American, 85(19): 290, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
290d.pdf.

Anonymous:1901:NYCb

[3352] Anonymous. The New York Central Tunnel. Scientific American, 85(19):290, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
290e.pdf.

Anonymous:1901:BAM

[3353] Anonymous. Balloon and automobile match. Scientific American, 85(19):290, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
290f.pdf.

Anonymous:1901:RAA

[3354] Anonymous. Recent automobile accidents. Scientific American, 85 (19):290–291, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
290g.pdf.

Walsh:1901:NCS

[3355] George E. Walsh. The new Coast Signal Service. Scientific American, 85(19):291, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
291.pdf.

Anonymous:1901:HSAb

[3356] Anonymous. How to study autumn leaves. Scientific American, 85(19):291, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
291a.pdf.

Anonymous:1901:SNu

[3357] Anonymous. Science notes. Scientific American, 85(19):291, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n19/pdf/scientificamerican11091901-291b.pdf.

Wisby:1901:SA

[3358] Hrolf Wisby. Style in automobiles. Scientific American, 85(19): 292–293, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
292.pdf.

Brown:1901:ACR

[3359] Enos Brown. The artificial cultivation of the rubber tree for industrial purposes. Scientific American, 85(19):293, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n19/pdf/scientificamerican11091901-293.pdf.

Shepstone:1901:BNB

[3360] Harold J. Shepstone. Building the new Beachy Head lighthouse. Scientific American, 85(19):294–295, November 9, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n19/pdf/scientificamerican11091901-294.pdf.

Anonymous:1901:SMC

[3361] Anonymous. Some methods of construction of the rapid transit subway. Scientific American, 85(19):297, November 9, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n19/pdf/scientificamerican11091901-297.pdf.

Anonymous:1901:TNN

[3362] Anonymous. Towing the new Naval drydock to Algiers, La. Scientific American, 85(19):297, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
297a.pdf.

Anonymous:1901:USG

[3363] Anonymous. United States and German patent practice. Scientific American, 85(19):298, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
298.pdf.

Anonymous:1901:CJJ

[3364] Anonymous. Col. J. J. Astor’s marine turbine. Scientific American, 85(19):298–299, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
298a.pdf.

Anonymous:1901:AVW

[3365] Anonymous. An adjustable ventilator for windows. Scientific American, 85(19):299, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
299.pdf.

Anonymous:1901:ISS

[3366] Anonymous. An improved saw-sharpener. Scientific American, 85 (19):299, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
299a.pdf.

Anonymous:1901:DCI

[3367] Anonymous. Devices curious and interesting. Scientific American, 85(19):299–300, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
299b.pdf.

Anonymous:1901:BPWbe

[3368] Anonymous. Business and personal wants. Scientific American, 85(19):302, November 9, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n19/pdf/scientificamerican11091901-
302.pdf.

Anonymous:1901:RPIbl

[3369] Anonymous. Recently patented inventions. Scientific American, 85(2):27, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
27.pdf.

Anonymous:1901:BNBz

[3370] Anonymous. Books: New books, etc. Scientific American, 85(2):28, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-28.pdf.

Anonymous:1901:NQbl

[3371] Anonymous. Notes and queries. Scientific American, 85(2):28–29, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-28b.pdf.

Anonymous:1901:IIbl

[3372] Anonymous. Index of inventions. Scientific American, 85(2):29, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-29.pdf.

Anonymous:1901:PNPbh

[3373] Anonymous. Patent notes: Patents. Scientific American, 85(2):31, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-31.pdf.

Anonymous:1901:TRT

[3374] Anonymous. There is room at the top. Scientific American, 85(2):18, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-18.pdf.

Anonymous:1901:LNW

[3375] Anonymous. The leading Navies of the world compared. Scientific American, 85(2):18, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
18a.pdf.

Anonymous:1901:RRB

[3376] Anonymous. Report of Royal British Observatory for 1900. Scientific American, 85(2):18–19, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
18b.pdf.

Anonymous:1901:BTP

[3377] Anonymous. The Berliner transmitter patent held to be invalid. Scientific American, 85(2):19, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
19.pdf.

Anonymous:1901:PBA

[3378] Anonymous. Paris–Berlin Automobile Race. Scientific American, 85(2):19, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
19a.pdf.

Anonymous:1901:SNv

[3379] Anonymous. Science notes. Scientific American, 85(2):19, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-19b.pdf.

Anonymous:1901:DPJ

[3380] Anonymous. Death of Prof. John Fiske. Scientific American, 85(2):19, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-19c.pdf.

Wohlford:1901:ZMK

[3381] Harry L. Wohlford. Zinc miring in Kansas. Scientific American, 85(2):20, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
20.pdf.

Anonymous:1901:CSF

[3382] Anonymous. A clever salvage feat. Scientific American, 85(2):20–21, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-20a.pdf.

Barbour:1901:FRA

[3383] Erwin Hinckley Barbour. A few remarkable artesian wells and the uses to which they are put. Scientific American, 85(2):21, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n2/pdf/scientificamerican07131901-21.pdf.

Anonymous:1901:AWO

[3384] Anonymous. Admirable work of the ordnance board at sandy hook. Scientific American, 85(2):22, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
22.pdf.

Anonymous:1901:IEP

[3385] Anonymous. Interesting exhibits at the Pan-American. Scientific American, 85(2):23, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
23.pdf.

Anonymous:1901:CBT

[3386] Anonymous. The common-battery telephone system of the city of New York. Scientific American, 85(2):24–25, July 13, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n2/pdf/scientificamerican07131901-24.pdf.

Anonymous:1901:IDCa

[3387] Anonymous. An improved double-cylinder force-pump. Scientific American, 85(2):26, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
26.pdf.

Anonymous:1901:MEC

[3388] Anonymous. A miniature electric carriage at the Pan-American Exposition. Scientific American, 85(2):26, July 13, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n2/pdf/scientificamerican07131901-26a.pdf.

Anonymous:1901:NSS

[3389] Anonymous. A new street sign for lamp-posts. Scientific American, 85(2):26–27, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
26b.pdf.

Anonymous:1901:BPWbf

[3390] Anonymous. Business and personal wants. Scientific American, 85(2):28, July 13, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n2/pdf/scientificamerican07131901-
28a.pdf.

Anonymous:1901:RPIbm

[3391] Anonymous. Recently patented inventions. Scientific American, 85(20):318, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
318.pdf.

Anonymous:1901:IIbm

[3392] Anonymous. Index of inventions. Scientific American, 85(20): 318, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
318b.pdf.

Anonymous:1901:NQbm

[3393] Anonymous. Notes and queries. Scientific American, 85(20): 321, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
321.pdf.

Anonymous:1901:PIN

[3394] Anonymous. Proposed increase of the Navy. Scientific American, 85(20):308, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
308.pdf.

Anonymous:1901:TBS

[3395] Anonymous. The tunnel between Scotland and Ireland. Scientific American, 85(20):308, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
308a.pdf.

Anonymous:1901:PLBb

[3396] Anonymous. The problem of the locomotive boiler. Scientific American, 85(20):308, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
308b.pdf.

Anonymous:1901:NLT

[3397] Anonymous. New London Telephone System. Scientific American, 85(20):308, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
308c.pdf.

Anonymous:1901:BZH

[3398] Anonymous. The Berlin–Zossen high speed electric road. Scientific American, 85(20):308–309, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
308d.pdf.

Anonymous:1901:PEA

[3399] Anonymous. Paris Exhibition of alcohol-consuming devices. Scientific American, 85(20):309, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
309.pdf.

Anonymous:1901:EPA

[3400] Anonymous. End of the Pan-American. Scientific American, 85 (20):309, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
309a.pdf.

Anonymous:1901:PCb

[3401] Anonymous. Prizes covering over. Scientific American, 85(20): 309, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
309b.pdf.

Anonymous:1901:SNw

[3402] Anonymous. Science notes. Scientific American, 85(20):309– 310, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
309c.pdf.

Fawcett:1901:NSO

[3403] Waldon Fawcett. A novel salvage operation. Scientific American, 85(20):310, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
310.pdf.

Anonymous:1901:BAN

[3404] Anonymous. Building the approaches of the new East River Bridge. Scientific American, 85(20):311, November 16, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n20/pdf/scientificamerican11161901-311.pdf.

Anonymous:1901:FFC

[3405] Anonymous. French first-class battleship “Charlemagne”. Scientific American, 85(20):311–312, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
311a.pdf.

Hopkins:1901:FMI

[3406] George M. Hopkins. A few meteorological instruments. Scientific American, 85(20):312, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
312.pdf.

Anonymous:1901:WST

[3407] Anonymous. Work on the simplon tunnel. Scientific American, 85(20):313, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
313.pdf.

Anonymous:1901:SDW

[3408] Anonymous. Santos-Dumont wins the Deutsch Prize. Scientific American, 85(20):314, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
314.pdf.

Anonymous:1901:SEAa

[3409] Anonymous. Some exhibits at the automobile show. Scientific American, 85(20):314–316, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
314a.pdf.

Anonymous:1901:WWW

[3410] Anonymous. The wheel within wheel. Scientific American, 85 (20):316, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
316.pdf.

Mcurdy:1901:TW

[3411] J. G. M’curdy. The teredo and its work. Scientific American, 85 (20):317, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
317.pdf.

Anonymous:1901:BPWbg

[3412] Anonymous. Business and personal wants. Scientific American, 85(20):318, November 16, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n20/pdf/scientificamerican11161901-
318a.pdf.

Anonymous:1901:RPIbn

[3413] Anonymous. Recently patented inventions. Scientific American, 85(21):334, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
334.pdf.

Anonymous:1901:IIbn

[3414] Anonymous. Index of inventions. Scientific American, 85(21): 335, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
335.pdf.

Anonymous:1901:BNBba

[3415] Anonymous. Books: New books, etc. Scientific American, 85 (21):335, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
335a.pdf.

Anonymous:1901:NQbn

[3416] Anonymous. Notes and queries. Scientific American, 85(21): 336, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
336.pdf.

Anonymous:1901:ONN

[3417] Anonymous. Our new Navy. Scientific American, 85(21):324, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n21/pdf/scientificamerican11231901-324.pdf.

Anonymous:1901:CCS

[3418] Anonymous. A compound condensing steam carriage. Scientific American, 85(21):324, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
324a.pdf.

Anonymous:1901:MSDb

[3419] Anonymous. M. Santos-Dumont’s plans to cross from Nice to Corsica. Scientific American, 85(21):324, November 23, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n21/pdf/scientificamerican11231901-324b.pdf.

Anonymous:1901:FTS

[3420] Anonymous. Further tests at sea of the Belleville Boiler. Scientific American, 85(21):324–325, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
324c.pdf.

Benjamin:1901:NASb

[3421] Marcus Benjamin. The National Academy of Sciences. Scientific American, 85(21):325, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
325.pdf.

Anonymous:1901:DST

[3422] Anonymous. Dr. Stein’s travels in Chinese Turkestan. Scientific American, 85(21):325, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
325a.pdf.

Anonymous:1901:GPE

[3423] Anonymous. German port of Emden. Scientific American, 85 (21):325, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
325b.pdf.

Moore:1901:CLP

[3424] Frederick Moore. Cable laying in the Philippines. Scientific American, 85(21):326, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
326.pdf.

Dunlap:1901:NCA

[3425] Orrin E. Dunlap. New concrete arch bridges at Niagara. Scientific American, 85(21):327–328, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
327.pdf.

Wyckoff:1901:PSN

[3426] A. B. Wyckoff. The Puget Sound Navy Yard. Scientific American, 85(21):328–329, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
328.pdf.

Anonymous:1901:LOC

[3427] Anonymous. The Lick Observatory—Crocker Eclipse Expedition to Sumatra. Scientific American, 85(21):331, November 23, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n21/pdf/scientificamerican11231901-331.pdf.

Anonymous:1901:SEAb

[3428] Anonymous. Some exhibits at the automobile show — II. Scientific American, 85(21):331–332, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
331a.pdf.

Carson:1901:O

[3429] A. C. Carson. The opossum. Scientific American, 85(21):332– 333, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
332.pdf.

Anonymous:1901:BPWbh

[3430] Anonymous. Business and personal wants. Scientific American, 85(21):334, November 23, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n21/pdf/scientificamerican11231901-
334a.pdf.

Anonymous:1901:LNb

[3431] Anonymous. Legal notes. Scientific American, 85(22):350, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n22/pdf/scientificamerican11301901-350a.pdf.

Anonymous:1901:RPIbo

[3432] Anonymous. Recently patented inventions. Scientific American, 85(22):351, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
351.pdf.

Anonymous:1901:IIbo

[3433] Anonymous. Index of inventions. Scientific American, 85(22): 351–352, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
351a.pdf.

Anonymous:1901:NQbo

[3434] Anonymous. Notes and queries. Scientific American, 85(22): 353, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
353.pdf.

Anonymous:1901:RHE

[3435] Anonymous. The recent high-explosive shell tests at sandy hook. Scientific American, 85(22):340, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
340.pdf.

Anonymous:1901:NTE

[3436] Anonymous. A new thermo-electric couple. Scientific American, 85(22):340, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
340a.pdf.

Anonymous:1901:GSN

[3437] Anonymous. The growth and status of natural gas. Scientific American, 85(22):340–341, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
340b.pdf.

Russell:1901:HDN

[3438] Henry Norris Russell. The heavens in December and The New Star in Perseus. Scientific American, 85(22):341, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n22/pdf/scientificamerican11301901-341.pdf.

Anonymous:1901:SNx

[3439] Anonymous. Science notes. Scientific American, 85(22):341, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n22/pdf/scientificamerican11301901-341a.pdf.

Aaron:1901:US

[3440] S. Frank Aaron. Useful spiders. Scientific American, 85(22): 342, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
342.pdf.

Anonymous:1901:CCT

[3441] Anonymous. The Cape to Cairo Telegraph. Scientific American, 85(22):343, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
343.pdf.

Anonymous:1901:ICT

[3442] Anonymous. Important competitive test of high-explosive shell at sandy hook. Scientific American, 85(22):344–345, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n22/pdf/scientificamerican11301901-344.pdf.

Anonymous:1901:OLD

[3443] Anonymous. Our latest design for battleships. Scientific American, 85(22):345–346, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
345.pdf.

Anonymous:1901:NAS

[3444] Anonymous. New automobile speed records. Scientific American, 85(22):347, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
347.pdf.

Anonymous:1901:PDb

[3445] Anonymous. Patent department. Scientific American, 85(22): 348–349, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
348.pdf.

Anonymous:1901:TNE

[3446] Anonymous. Three new electrical devices. Scientific American, 85 (22):349, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
349.pdf.

Anonymous:1901:RFG

[3447] Anonymous. Removable firebox and grate for cooking stoves. Scientific American, 85(22):350, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
350.pdf.

Anonymous:1901:BPWbi

[3448] Anonymous. Business and personal wants. Scientific American, 85(22):352, November 30, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n22/pdf/scientificamerican11301901-
352.pdf.

Anonymous:1901:Ct

[3449] Anonymous. Correspondence. Scientific American, 85(23):363, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n23/pdf/scientificamerican12071901-363.pdf.

Anonymous:1901:RPIbp

[3450] Anonymous. Recently patented inventions. Scientific American, 85(23):367, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
367.pdf.

Anonymous:1901:IIbp

[3451] Anonymous. Index of inventions. Scientific American, 85(23):367, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n23/pdf/scientificamerican12071901-367b.pdf.

Anonymous:1901:PNPbi 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
370.pdf.

Anonymous:1901:BNBbb

[3453] Anonymous. Books: New books, etc. Scientific American, 85 (23):370, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
370a.pdf.

Anonymous:1901:NQbp

[3454] Anonymous. Notes and queries. Scientific American, 85(23):371, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n23/pdf/scientificamerican12071901-371.pdf.

Anonymous:1901:RCDb

[3455] Anonymous. Reconstruction of the Croton Dam. Scientific American, 85(23):358, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
358.pdf.

Anonymous:1901:IDCb

[3456] Anonymous. Irrigation of the Delta of the Colorado. Scientific American, 85(23):358, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
358a.pdf.

Anonymous:1901:CGT

[3457] Anonymous. The conversion of grape tendrils into fruit clusters. Scientific American, 85(23):358–359, December 7, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n23/pdf/scientificamerican12071901-358b.pdf.

Anonymous:1901:CDP

[3458] Anonymous. Cave drawings of the Palæolithic Epoch. Scientific American, 85(23):359, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
359.pdf.

Proctor:1901:NNP

[3459] Mary Proctor. Nebula in Nova Persei. Scientific American, 85 (23):359, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
359a.pdf.

Anonymous:1901:SNy

[3460] Anonymous. Science notes. Scientific American, 85(23):359, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n23/pdf/scientificamerican12071901-359b.pdf.

Woodbury:1901:NSI

[3461] Fred P. Woodbury. A new synchronism indicator for alternators. Scientific American, 85(23):360, December 7, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n23/pdf/scientificamerican12071901-360.pdf.

Anonymous:1901:NATb

[3462] Anonymous. New Atlantic type of passenger locomotive on the Pennsylvania Railroad. Scientific American, 85(23):360, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n23/pdf/scientificamerican12071901-360a.pdf.

Anonymous:1901:RBT

[3463] Anonymous. Rankin bridge for the transit of molten iron. Scientific American, 85(23):361, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
361.pdf.

Anonymous:1901:MOC

[3464] Anonymous. The manufacture of oil-cloth. Scientific American, 85(23):362, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
362.pdf.

Moore:1901:NDS

[3465] Frederick Moore. The new drainage and sewerage system of New Orleans. Scientific American, 85(23):364–365, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n23/pdf/scientificamerican12071901-364.pdf.

Anonymous:1901:PFL

[3466] Anonymous. Portraiture by flash light. Scientific American, 85 (23):365, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
365.pdf.

Anonymous:1901:PDc

[3467] Anonymous. Patent department. Scientific American, 85(23):366, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n23/pdf/scientificamerican12071901-366.pdf.

Anonymous:1901:TFT

[3468] Anonymous. Treating fruit-trees for scale. Scientific American, 85(23):366, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
366a.pdf.

Anonymous:1901:IDS

[3469] Anonymous. Improvement in dredge shovels. Scientific American, 85(23):366, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
366b.pdf.

Anonymous:1901:BPWbj

[3470] Anonymous. Business and personal wants. Scientific American, 85(23):367, December 7, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n23/pdf/scientificamerican12071901-
367a.pdf.

Anonymous:1901:RPIbq

[3471] Anonymous. Recently patented inventions. Scientific American, 85(24):394, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
394.pdf.

Anonymous:1901:IIbq

[3472] Anonymous. Index of inventions. Scientific American, 85(24): 394, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
394a.pdf.

Anonymous:1901:NQbq

[3473] Anonymous. Notes and queries. Scientific American, 85(24): 397, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
397.pdf.

Anonymous:1901:PPN

[3474] Anonymous. Peace and a powerful Navy. Scientific American, 85 (24):374, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
374.pdf.

Anonymous:1901:WCGb

[3475] Anonymous. Warship construction at the Government Navy yards. Scientific American, 85(24):374, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
374a.pdf.

Anonymous:1901:ONDb

[3476] Anonymous. Our Naval development since the war with Spain. Scientific American, 85(24):374, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
374b.pdf.

Anonymous:1901:RNW

[3477] Anonymous. “Revolutionizing” Naval warfare. Scientific American, 85(24):374–375, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
374c.pdf.

Anonymous:1901:B

[3478] Anonymous. Battleships. Scientific American, 85(24):375–376, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n24/pdf/scientificamerican12141901-375.pdf.

Anonymous:1901:ACF

[3479] Anonymous. “Alabama” Class — First-Class battleship “Wisconsin”. Scientific American, 85(24):376–378, December 14, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n24/pdf/scientificamerican12141901-376.pdf.

Anonymous:1901:EBW

[3480] Anonymous. Engines of the battleship “Wisconsin”. Scientific American, 85(24):378, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
378.pdf.

Anonymous:1901:MCF

[3481] Anonymous. “Maine” class — first-class battleship “Maine”. Scientific American, 85(24):378, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
378a.pdf.

Anonymous:1901:VCF

[3482] Anonymous. “Virginia” class — first-class battleship “Georgia”. Scientific American, 85(24):378–381, December 14, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n24/pdf/scientificamerican12141901-378b.pdf.

Anonymous:1901:Mb

[3483] Anonymous. Monitors. Scientific American, 85(24):381, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n24/pdf/scientificamerican12141901-381.pdf.

Anonymous:1901:GA

[3484] Anonymous. Guns and armor. Scientific American, 85(24):382– 383, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
382.pdf.

Anonymous:1901:GMB

[3485] Anonymous. Gun-mounts and breech mechanisms. Scientific American, 85(24):383, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
383.pdf.

Anonymous:1901:APHb

[3486] Anonymous. Armor plate and high-explosive shells. Scientific American, 85(24):384–385, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
384.pdf.

Anonymous:1901:ACb

[3487] Anonymous. Armored cruisers. Scientific American, 85(24): 385, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
385.pdf.

Anonymous:1901:SAC

[3488] Anonymous. The semi-armored cruisers of the “St. Louis” class. Scientific American, 85(24):386, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
386.pdf.

Anonymous:1901:PCc

[3489] Anonymous. Protected cruisers. Scientific American, 85(24):386– 387, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
386a.pdf.

Anonymous:1901:STB

[3490] Anonymous. Submarine torpedo boats. Scientific American, 85 (24):387–388, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
387.pdf.

Anonymous:1901:TBDa

[3491] Anonymous. Torpedo-boats and destroyers. Scientific American, 85(24):388, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
388.pdf.

Anonymous:1901:TBDb

[3492] Anonymous. Torpedo-boat destroyer “Perry”. Scientific American, 85 (24):388–390, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
388a.pdf.

Anonymous:1901:ORS

[3493] Anonymous. Our relative standing among the Naval Powers. Scientific American, 85(24):390–394, December 14, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n24/pdf/scientificamerican12141901-390.pdf.

Anonymous:1901:BPWbk

[3494] Anonymous. Business and personal wants. Scientific American, 85(24):400, December 14, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n24/pdf/scientificamerican12141901-
400.pdf.

Anonymous:1901:Cu

[3495] Anonymous. Correspondence. Scientific American, 85(25):410– 411, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
410a.pdf.

Anonymous:1901:RPIbr

[3496] Anonymous. Recently patented inventions. Scientific American, 85(25):416, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
416.pdf.

Anonymous:1901:IIbr

[3497] Anonymous. Index of inventions. Scientific American, 85(25): 416–417, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
416b.pdf.

Anonymous:1901:BNBbc

[3498] Anonymous. Books: New books, etc. Scientific American, 85 (25):418, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
418.pdf.

Anonymous:1901:NQbr

[3499] Anonymous. Notes and queries. Scientific American, 85(25): 419, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
419.pdf.

Anonymous:1901:PRT

[3500] Anonymous. Pennsylvania Railroad Terminal Station and Tunnel. Scientific American, 85(25):406, December 21, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n25/pdf/scientificamerican12211901-406.pdf.

Anonymous:1901:NPa

[3501] Anonymous. Nicaragua or Panama. Scientific American, 85(25): 406, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
406a.pdf.

Anonymous:1901:IAD

[3502] Anonymous. An interesting automobile damage case. Scientific American, 85(25):407, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
407.pdf.

Anonymous:1901:RFM

[3503] Anonymous. The Robert Fulton Memorial. Scientific American, 85(25):407, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
407a.pdf.

Anonymous:1901:IKO

[3504] Anonymous. The international kerosene oil motor. Scientific American, 85(25):408, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
408.pdf.

Anonymous:1901:SDE

[3505] Anonymous. Safety device for engines and condensers. Scientific American, 85(25):409, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
409.pdf.

Anonymous:1901:IMP

[3506] Anonymous. Iron making on the Pacific Coast. Scientific American, 85(25):409–410, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
409a.pdf.

Anonymous:1901:IHS

[3507] Anonymous. An improved hay stacker and builders’ derrick. Scientific American, 85(25):410, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
410.pdf.

Anonymous:1901:GSW

[3508] Anonymous. Giant sand wheel for the Calumet and Hecla Mining Company. Scientific American, 85(25):411, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n25/pdf/scientificamerican12211901-411.pdf.

Proctor:1901:MHS

[3509] Mary Proctor. Measuring the heat of the stars. Scientific American, 85(25):412–413, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
412.pdf.

Wisby:1901:EMC

[3510] Hrolf Wisby. The evolution of the motor cycle. Scientific American, 85(25):413–414, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
413.pdf.

Anonymous:1901:HF

[3511] Anonymous. Hypnosis in frogs. Scientific American, 85(25): 414–415, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
414.pdf.

Holder:1901:RF

[3512] C. F. Holder. A rare fish. Scientific American, 85(25):415, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n25/pdf/scientificamerican12211901-415.pdf.

Anonymous:1901:BPWbl

[3513] Anonymous. Business and personal wants. Scientific American, 85(25):416, December 21, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n25/pdf/scientificamerican12211901-
416a.pdf.

Russell:1901:HJ

[3514] Henry Norris Russell. The heavens in January, 1902. Scientific American, 85(26):423, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
423.pdf.

Anonymous:1901:RPIbs

[3515] Anonymous. Recently patented inventions. Scientific American, 85 (26):432–433, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
432.pdf.

Anonymous:1901:IIbs

[3516] Anonymous. Index of inventions. Scientific American, 85(26): 434, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
434.pdf.

Anonymous:1901:Ij

[3517] Anonymous. Index. Scientific American, 85(26):435–436, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n26/pdf/scientificamerican12281901-435.pdf.

Anonymous:1901:BNBbd

[3518] Anonymous. Books: New books. Scientific American, 85(26): 437, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
437.pdf.

Anonymous:1901:NQbs

[3519] Anonymous. Notes and queries. Scientific American, 85(26): 438, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
438.pdf.

Anonymous:1901:TWTb

[3520] Anonymous. Transoceanic wireless telegraphy. Scientific American, 85(26):422, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
422.pdf.

Anonymous:1901:ANN

[3521] Anonymous. Appreciation of the Naval number. Scientific American, 85(26):422, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
422a.pdf.

Anonymous:1901:BVT

[3522] Anonymous. Bridges versus tunnels. Scientific American, 85 (26):422, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
422b.pdf.

Loewy:1901:ARW

[3523] M. Loewy. Annual report of the work of the Observatory of Paris for 1900. Scientific American, 85(26):422–423, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n26/pdf/scientificamerican12281901-422c.pdf.

Anonymous:1901:PP

[3524] Anonymous. A pendulum propeller. Scientific American, 85 (26):424, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
424.pdf.

Collins:1901:SAP

[3525] A. Frederick Collins. The Slaby–Arco portable field equipment for wireless telegraphy. Scientific American, 85(26):425–426, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n26/pdf/scientificamerican12281901-425.pdf.

Anonymous:1901:PLI

[3526] Anonymous. The Pennsylvania–Long Island Connecting Railroad Tunnels. Scientific American, 85(26):426–427, December 28, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n26/pdf/scientificamerican12281901-426.pdf.

Anonymous:1901:NSU

[3527] Anonymous. New submarines for the United States Navy. Scientific American, 85(26):427, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
427.pdf.

Anonymous:1901:CTT

[3528] Anonymous. The combined third and traction rail system of haulage. Scientific American, 85(26):428, December 28, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n26/pdf/scientificamerican12281901-428.pdf.

Anonymous:1901:AAP

[3529] Anonymous. Alcohol automobiles at the Paris Alcohol Exhibition. Scientific American, 85(26):428, December 28, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n26/pdf/scientificamerican12281901-428a.pdf.

Anonymous:1901:MLG

[3530] Anonymous. Method of lighting the Great Water Tank at the Nouveau Cirque, Paris. Scientific American, 85(26):429, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n26/pdf/scientificamerican12281901-429.pdf.

Anonymous:1901:PDd

[3531] Anonymous. Patent department. Scientific American, 85(26): 430, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
430.pdf.

Anonymous:1901:SII

[3532] Anonymous. Simple interesting inventions. Scientific American, 85(26):431, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
431.pdf.

Anonymous:1901:BPWbm

[3533] Anonymous. Business and personal wants. Scientific American, 85(26):433, December 28, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n26/pdf/scientificamerican12281901-
433.pdf.

Anonymous:1901:Cv

[3534] Anonymous. Correspondence. Scientific American, 85(3):38, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n3/pdf/scientificamerican07201901-38.pdf.

Anonymous:1901:RPIbt

[3535] Anonymous. Recently patented inventions. Scientific American, 85(3):44, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
44.pdf.

Anonymous:1901:BNBbe

[3536] Anonymous. Books: New books, etc. Scientific American, 85(3):44, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n3/pdf/scientificamerican07201901-44a.pdf.

Anonymous:1901:NQbt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n3/pdf/scientificamerican07201901-45a.pdf.

Anonymous:1901:IIbt

[3538] Anonymous. Index of inventions. Scientific American, 85(3):45, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n3/pdf/scientificamerican07201901-45b.pdf.

Anonymous:1901:PNPbj

[3539] Anonymous. Patent notes: Patents. Scientific American, 85(3):47, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n3/pdf/scientificamerican07201901-47.pdf.

Anonymous:1901:NST

[3540] Anonymous. A new scheme of trade mark blackmail. Scientific American, 85(3):34, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
34.pdf.

Anonymous:1901:SCCa

[3541] Anonymous. The struggle of the cup champions. Scientific American, 85(3):34, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
34a.pdf.

Anonymous:1901:NPb

[3542] Anonymous. A Naval problem. Scientific American, 85(3):34–35, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n3/pdf/scientificamerican07201901-34b.pdf.

Anonymous:1901:DAP

[3543] Anonymous. Defective assignments of patents and royalties. Scientific American, 85(3):35, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
35.pdf.

Anonymous:1901:IAM

[3544] Anonymous. Introducing American methods in England. Scientific American, 85(3):35, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
35a.pdf.

Anonymous:1901:MCBb

[3545] Anonymous. Manufacture of celluloid beads. Scientific American, 85(3):35, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
35b.pdf.

Anonymous:1901:SNz

[3546] Anonymous. Science notes. Scientific American, 85(3):35, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n3/pdf/scientificamerican07201901-35c.pdf.

Ruhmer:1901:P

[3547] Ernst Ruhmer. The “photographophone”. Scientific American, 85(3):36, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
36.pdf.

Anonymous:1901:NTM

[3548] Anonymous. A new typesetting machine. Scientific American, 85(3):37–38, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
37.pdf.

Anonymous:1901:TLT

[3549] Anonymous. Two of the latest types of battleships. Scientific American, 85(3):39, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
39.pdf.

Anonymous:1901:MWT

[3550] Anonymous. Marconi’s wireless telegraphy automobile for military purposes. Scientific American, 85(3):40, July 20, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n3/pdf/scientificamerican07201901-40.pdf.

Anonymous:1901:FPT

[3551] Anonymous. The first passenger turbine steamer “Edward VII”. Scientific American, 85(3):40–41, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
40a.pdf.

Fawcett:1901:MPA

[3552] Waldon Fawcett. The most primitive American savages. Scientific American, 85(3):41–42, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
41.pdf.

Anonymous:1901:WVF

[3553] Anonymous. The West Virginia Flood. Scientific American, 85(3):43, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n3/pdf/scientificamerican07201901-43.pdf.

Anonymous:1901:BPWbn

[3554] Anonymous. Business and personal wants. Scientific American, 85(3):45, July 20, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n3/pdf/scientificamerican07201901-
45.pdf.

Larkin:1901:HA

[3555] Edgar L. Larkin. The heavens in August. Scientific American, 85(4):51, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-51.pdf.

Anonymous:1901:RPIbu 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
59.pdf.

Anonymous:1901:BNBbf

[3557] Anonymous. Books: New books, etc. Scientific American, 85(4):59–60, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-59a.pdf.

Anonymous:1901:NQbu

[3558] Anonymous. Notes and queries. Scientific American, 85(4):60, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-60a.pdf.

Anonymous:1901:IIbu

[3559] Anonymous. Index of inventions. Scientific American, 85(4):60, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-60b.pdf.

Anonymous:1901:PNPbk

[3560] Anonymous. Patent notes: Patents. Scientific American, 85(4):62, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-62.pdf.

Anonymous:1901:ROR

[3561] Anonymous. The relocation of our railroads. Scientific American, 85(4):50, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
50.pdf.

Anonymous:1901:SRA

[3562] Anonymous. Some results of the Albany Filtration Plant. Scientific American, 85(4):50, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
50a.pdf.

Anonymous:1901:EAR

[3563] Anonymous. An epitome of American railway methods. Scientific American, 85(4):50, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
50b.pdf.

Anonymous:1901:FPS

[3564] Anonymous. The future possibilities of Siberia. Scientific American, 85(4):50–51, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
50c.pdf.

Anonymous:1901:TEC

[3565] Anonymous. Two eminent California scientists. Scientific American, 85(4):51, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
51a.pdf.

Anonymous:1901:SNba

[3566] Anonymous. Science notes. Scientific American, 85(4):51, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-51b.pdf.

Anonymous:1901:NWD

[3567] Anonymous. New weaving device. Scientific American, 85(4):51, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-51c.pdf.

Anonymous:1901:FTD

[3568] Anonymous. The fastest trip of the “Deutschland”. Scientific American, 85(4):51, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
51d.pdf.

Anonymous:1901:WEB

[3569] Anonymous. The Westinghouse electric brake and heater. Scientific American, 85(4):52, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
52.pdf.

Anonymous:1901:FSM

[3570] Anonymous. Fog studies on Mount Tamalpais, California. Scientific American, 85(4):52–53, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
52a.pdf.

Anonymous:1901:NCS

[3571] Anonymous. A novel chain-stopper. Scientific American, 85(4):53, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-53.pdf.

Anonymous:1901:IBa

[3572] Anonymous. An improved blowpipe. Scientific American, 85(4):53, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-53a.pdf.

Hall:1901:EHC

[3573] Leon M. Hall. Electric hoists on the Comstock. Scientific American, 85(4):53–54, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
53b.pdf.

Anonymous:1901:MBG

[3574] Anonymous. The manufacture of blown glass-lamp chimneys. Scientific American, 85(4):54–55, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
54.pdf.

Anonymous:1901:PBM

[3575] Anonymous. The Paris–Berlin Motor Carriage Race. Scientific American, 85(4):56–57, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
56.pdf.

Anonymous:1901:CAa

[3576] Anonymous. The conquest of the air. Scientific American, 85(4):57–58, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-57.pdf.

Milhenny:1901:NSB

[3577] James M’ilhenny. A natural sand blast. Scientific American, 85(4):58, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n4/pdf/scientificamerican07271901-58.pdf.

Anonymous:1901:BPWbo

[3578] Anonymous. Business and personal wants. Scientific American, 85(4):60, July 27, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n4/pdf/scientificamerican07271901-
60.pdf.

Anonymous:1901:Cx

[3579] Anonymous. Correspondence. Scientific American, 85(5):70–71, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-70a.pdf.

Anonymous:1901:RPIbv

[3580] Anonymous. Recently patented inventions. Scientific American, 85(5):76, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
76.pdf.

Anonymous:1901:BNBbg

[3581] Anonymous. Books: New books, etc. Scientific American, 85(5):76, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-76a.pdf.

Anonymous:1901:NQbv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-77.pdf.

Anonymous:1901:IIbv

[3583] Anonymous. Index of inventions. Scientific American, 85(5):77, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-77a.pdf.

Anonymous:1901:PNPbl

[3584] Anonymous. Patent notes: Patents. Scientific American, 85(5):78, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-78.pdf.

Anonymous:1901:FRB

[3585] Anonymous. Future reconstruction of the Brooklyn Bridge. Scientific American, 85(5):66, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
66.pdf.

Anonymous:1901:RDB

[3586] Anonymous. Recent developments in the British Navy. Scientific American, 85(5):66, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
66a.pdf.

Anonymous:1901:EMD

[3587] Anonymous. Electrical method for determining the variations of mineral waters. Scientific American, 85(5):66–67, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n5/pdf/scientificamerican08031901-66b.pdf.

Pellat:1901:NPT

[3588] H. Pellat. National Physico-Technical Laboratories. Scientific American, 85(5):67, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
67.pdf.

Anonymous:1901:WEE

[3589] Anonymous. The work of the Egyptian Exploration Fund. Scientific American, 85(5):67, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
67a.pdf.

Anonymous:1901:SNbb

[3590] Anonymous. Science notes. Scientific American, 85(5):67, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-67b.pdf.

Anonymous:1901:BIP

[3591] Anonymous. Baby incubators at the Pan-American Exposition. Scientific American, 85(5):68, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
68.pdf.

Anonymous:1901:NRV

[3592] Anonymous. The new Riverside Viaduct, New York. Scientific American, 85(5):69–70, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
69.pdf.

Anonymous:1901:TGS

[3593] Anonymous. The three great suspension bridges across the East River, New York. Scientific American, 85(5):70, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n5/pdf/scientificamerican08031901-70.pdf.

Anonymous:1901:NKW

[3594] Anonymous. A new kind of water-tank. Scientific American, 85(5):71, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-71.pdf.

Holder:1901:WM

[3595] Charles F. Holder. A wave motor. Scientific American, 85(5):72, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-72.pdf.

Anonymous:1901:LFD

[3596] Anonymous. The largest floating dock in the world. Scientific American, 85(5):72–73, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
72a.pdf.

Anonymous:1901:SRT

[3597] Anonymous. Some remarkable tree growths. Scientific American, 85(5):73, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
73.pdf.

Anonymous:1901:NFMc

[3598] Anonymous. Nemethy’s flying-machine. Scientific American, 85 (5):73–74, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
73a.pdf.

Anonymous:1901:NSD

[3599] Anonymous. A new semi-dry battery. Scientific American, 85(5):74, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n5/pdf/scientificamerican08031901-74.pdf.

Anonymous:1901:FBB

[3600] Anonymous. The failure of the Brooklyn Bridge. Scientific American, 85(5):74–76, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
74a.pdf.

Anonymous:1901:BPWbp

[3601] Anonymous. Business and personal wants. Scientific American, 85(5):76–77, August 3, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n5/pdf/scientificamerican08031901-
76b.pdf.

Anonymous:1901:Cy

[3602] Anonymous. Correspondence. Scientific American, 85(6):87, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n6/pdf/scientificamerican08101901-87.pdf.

Anonymous:1901:RPIbw

[3603] Anonymous. Recently patented inventions. Scientific American, 85(6):92, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
92.pdf.

Anonymous:1901:BNBbh

[3604] Anonymous. Books: New books, etc. Scientific American, 85(6):92, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n6/pdf/scientificamerican08101901-92a.pdf.

Anonymous:1901:NQbw

[3605] Anonymous. Notes and queries. Scientific American, 85(6):92–93, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n6/pdf/scientificamerican08101901-92c.pdf.

Anonymous:1901:IIbw

[3606] Anonymous. Index of inventions. Scientific American, 85(6):93, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n6/pdf/scientificamerican08101901-93a.pdf.

Anonymous:1901:PNPbm

[3607] Anonymous. Patent notes: Patents. Scientific American, 85(6):94, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n6/pdf/scientificamerican08101901-94.pdf.

Anonymous:1901:SSBb

[3608] Anonymous. Supervision and safety of the Brooklyn Bridge. Scientific American, 85(6):82, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
82.pdf.

Anonymous:1901:NYCc

[3609] Anonymous. The New York Central Railroad Tunnel nuisance. Scientific American, 85(6):82, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
82a.pdf.

Anonymous:1901:DOB

[3610] Anonymous. Divergent opinions on battleship design. Scientific American, 85(6):82–83, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
82b.pdf.

Walsh:1901:VUP

[3611] George E. Walsh. Various uses of paper. Scientific American, 85(6):83, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
83.pdf.

Anonymous:1901:SNbc

[3612] Anonymous. Science notes. Scientific American, 85(6):83, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n6/pdf/scientificamerican08101901-83a.pdf.

Anonymous:1901:CBC

[3613] Anonymous. A combined box-cap and shoe-plank holder. Scientific American, 85(6):84, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
84.pdf.

Anonymous:1901:SL

[3614] Anonymous. Stephenson’s locomotive. Scientific American, 85 (6):84, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
84a.pdf.

Anonymous:1901:NCO

[3615] Anonymous. A novel can-opener. Scientific American, 85(6):84, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n6/pdf/scientificamerican08101901-84b.pdf.

Talbot:1901:ECC

[3616] Frederick A. Talbot. The electrical causes of changes in the weather. Scientific American, 85(6):85–86, August 10, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n6/pdf/scientificamerican08101901-85.pdf.

Anonymous:1901:CMCb

[3617] Anonymous. Coke making in the Connellsville region. Scientific American, 85(6):86, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
86.pdf.

Anonymous:1901:ESE

[3618] Anonymous. The examination and sorting of eggs. Scientific American, 85(6):88, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
88.pdf.

Anonymous:1901:TEK

[3619] Anonymous. The turbine engines of the “King Edward”. Scientific American, 85(6):88–89, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
88a.pdf.

Anonymous:1901:SDB

[3620] Anonymous. The Santos-Dumont balloon. Scientific American, 85(6):89–90, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
89.pdf.

Roberts:1901:NFB

[3621] Edward Coleridge Roberts. A new form of barometer. Scientific American, 85(6):90, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
90.pdf.

Anonymous:1901:NBD

[3622] Anonymous. The new battleship design controversy. Scientific American, 85(6):90–91, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
90a.pdf.

Anonymous:1901:BPWbq

[3623] Anonymous. Business and personal wants. Scientific American, 85(6):92, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n6/pdf/scientificamerican08101901-
92b.pdf.

Anonymous:1901:Ik

[3624] Anonymous. To inventors. Scientific American, 85(6):93, August 10, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n6/pdf/scientificamerican08101901-93.pdf.

Anonymous:1901:Cz

[3625] Anonymous. Correspondence. Scientific American, 85(7):102, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n7/pdf/scientificamerican08171901-102a.pdf.

Anonymous:1901:RPIbx

[3626] Anonymous. Recently patented inventions. Scientific American, 85(7):108, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
108.pdf.

Anonymous:1901:NQbx

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n7/pdf/scientificamerican08171901-109.pdf.

Anonymous:1901:BNBbi

[3628] Anonymous. Books: New books, etc. Scientific American, 85 (7):109, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
109a.pdf.

Anonymous:1901:IIbx

[3629] Anonymous. Index of inventions. Scientific American, 85(7):109, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n7/pdf/scientificamerican08171901-109b.pdf.

Anonymous:1901:PNPbn

[3630] Anonymous. Patent notes: Patents. Scientific American, 85(7):110, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n7/pdf/scientificamerican08171901-110.pdf.

Anonymous:1901:MSDc

[3631] Anonymous. Mishap to the Santos-Dumont airship. Scientific American, 85(7):98, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
98.pdf.

Anonymous:1901:BSD

[3632] Anonymous. Big ships and deep waterways. Scientific American, 85(7):98, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
98a.pdf.

Anonymous:1901:HHS

[3633] Anonymous. Humidity and heating systems. Scientific American, 85(7):98–99, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
98b.pdf.

Anonymous:1901:PHP

[3634] Anonymous. The proposed history of the Patent Office. Scientific American, 85(7):99, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
99.pdf.

Anonymous:1901:WTV

[3635] Anonymous. Water-tube vs. fire-tube boilers for Naval use. Scientific American, 85(7):99, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
99a.pdf.

Anonymous:1901:SNbd

[3636] Anonymous. Science notes. Scientific American, 85(7):99–100, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n7/pdf/scientificamerican08171901-99b.pdf.

Weeks:1901:TSB

[3637] Arthur B. Weeks. In the Terminal Station at Buffalo. Scientific American, 85(7):100, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
100.pdf.

Anonymous:1901:RSI

[3638] Anonymous. Relative speed indicator. Scientific American, 85 (7):100–101, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
100a.pdf.

Holder:1901:BSM

[3639] Charles F. Holder. Burning the Sierra Madres. Scientific American, 85(7):101, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
101.pdf.

Anonymous:1901:EII

[3640] Anonymous. East Indians in the Island of Trinidad. Scientific American, 85(7):101–102, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
101a.pdf.

Anonymous:1901:RTN

[3641] Anonymous. Railroad track nut-lock. Scientific American, 85 (7):102, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
102.pdf.

Anonymous:1901:WSL

[3642] Anonymous. The White Star liner “Celtic”. Scientific American, 85(7):103, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
103.pdf.

Willey:1901:EGB

[3643] Day Allen Willey. The erection of the Gokteik Bridge. Scientific American, 85(7):104, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
104.pdf.

Anonymous:1901:YLD

[3644] Anonymous. Yale’s large dinosaur. Scientific American, 85(7):105, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n7/pdf/scientificamerican08171901-105.pdf.

Fawcett:1901:SPF

[3645] Waldon Fawcett. Self-propelling fire engines. Scientific American, 85(7):105–107, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
105a.pdf.

Anonymous:1901:SNLa

[3646] Anonymous. Sir Norman Lockyer on Sun spots. Scientific American, 85(7):107–108, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
107.pdf.

Anonymous:1901:BPWbr

[3647] Anonymous. Business and personal wants. Scientific American, 85(7):108, August 17, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n7/pdf/scientificamerican08171901-
108a.pdf.

Anonymous:1901:RPIby

[3648] Anonymous. Recently patented inventions. Scientific American, 85(8):124, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
124.pdf.

Anonymous:1901:BNBbj

[3649] Anonymous. Books: New books, etc. Scientific American, 85 (8):124, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
124a.pdf.

Anonymous:1901:NQby

[3650] Anonymous. Notes and queries. Scientific American, 85(8):125, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n8/pdf/scientificamerican08241901-125.pdf.

Anonymous:1901:IIby

[3651] Anonymous. Index of inventions. Scientific American, 85(8): 126–127, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
126.pdf.

Anonymous:1901:PNPbo

[3652] Anonymous. Patent notes: Patents. Scientific American, 85(8):126, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n8/pdf/scientificamerican08241901-126a.pdf.

Anonymous:1901:THE

[3653] Anonymous. Trade and higher education. Scientific American, 85(8):114, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
114.pdf.

Anonymous:1901:SIC

[3654] Anonymous. “Shamrock II” and “Columbia” compared. Scientific American, 85(8):114, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
114a.pdf.

Anonymous:1901:DSP

[3655] Anonymous. Danger to St. Paul’s Cathedral. Scientific American, 85(8):114, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
114b.pdf.

Anonymous:1901:MSS

[3656] Anonymous. Modern structural steel. Scientific American, 85 (8):114–115, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
114c.pdf.

Anonymous:1901:AAF

[3657] Anonymous. American adulterated fabrics and tests. Scientific American, 85(8):115, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
115.pdf.

Anonymous:1901:BND

[3658] Anonymous. Baron Nordenskjöld dead. Scientific American, 85 (8):115, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
115a.pdf.

Anonymous:1901:SNbe

[3659] Anonymous. Science notes. Scientific American, 85(8):115, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n8/pdf/scientificamerican08241901-115b.pdf.

Anonymous:1901:DSR

[3660] Anonymous. A device for straightening rugs. Scientific American, 85(8):116, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
116.pdf.

Anonymous:1901:FLH

[3661] Anonymous. Freight locomotive for heavy grades. Scientific American, 85(8):116, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
116a.pdf.

Anonymous:1901:SLEb

[3662] Anonymous. A simple locomotive engine indicator. Scientific American, 85(8):116, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
116b.pdf.

Anonymous:1901:MGE

[3663] Anonymous. Modern grain elevators. Scientific American, 85 (8):117–118, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
117.pdf.

Adams:1901:SHC

[3664] Alton D. Adams. Steam heat with condensing engines. Scientific American, 85(8):118–119, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
118.pdf.

Anonymous:1901:TSTb

[3665] Anonymous. The telepherage system of transportation. Scientific American, 85(8):120, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
120.pdf.

Fawcett:1901:FDD

[3666] Waldon Fawcett. A fire department drill school. Scientific American, 85(8):120–121, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
120a.pdf.

Anonymous:1901:ECA

[3667] Anonymous. The Eleventh Challenger for the America Cup — “Shamrock II”. Scientific American, 85(8):122, August 24, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n8/pdf/scientificamerican08241901-122.pdf.

Anonymous:1901:NLF

[3668] Anonymous. A new luminous fountain. Scientific American, 85 (8):122–123, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
122a.pdf.

Anonymous:1901:VST

[3669] Anonymous. Vulnerability of the superposed turret. Scientific American, 85(8):123, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
123.pdf.

Anonymous:1901:BPWbs

[3670] Anonymous. Business and personal wants. Scientific American, 85(8):124, August 24, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n8/pdf/scientificamerican08241901-
124b.pdf.

Russell:1901:HS

[3671] Henry Norris Russell. The heavens in September. Scientific American, 85(9):134–135, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
134.pdf.

Anonymous:1901:RPIbz

[3672] Anonymous. Recently patented inventions. Scientific American, 85(9):140, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
140.pdf.

Anonymous:1901:NQbz

[3673] Anonymous. Notes and queries. Scientific American, 85(9):141, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n9/pdf/scientificamerican08311901-141a.pdf.

Anonymous:1901:IIbz

[3674] Anonymous. Index of inventions. Scientific American, 85(9):141, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n9/pdf/scientificamerican08311901-141b.pdf.

Anonymous:1901:PNPbp

[3675] Anonymous. Patent notes: Patents. Scientific American, 85(9):143, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n9/pdf/scientificamerican08311901-143.pdf.

Anonymous:1901:GIC

[3676] Anonymous. Germany’s industrial crisis. Scientific American, 85(9):130, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
130.pdf.

Anonymous:1901:VMI

[3677] Anonymous. Vast municipal improvements in New York City. Scientific American, 85(9):130, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
130a.pdf.

Anonymous:1901:RSU

[3678] Anonymous. Railway systems of the United States. Scientific American, 85(9):130–131, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
130b.pdf.

Benjamin:1901:CSM

[3679] Marcus Benjamin. Charles Sedgwick Minot. Scientific American, 85(9):131, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
131.pdf.

Anonymous:1901:EPR

[3680] Anonymous. Exports to Porto Rico. Scientific American, 85(9):131, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n9/pdf/scientificamerican08311901-131a.pdf.

Anonymous:1901:APN

[3681] Anonymous. Advantages of paper negatives. Scientific American, 85(9):131, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
131b.pdf.

Anonymous:1901:CBW

[3682] Anonymous. Cantilever bridge on the White Pass and Yukon Railroad. Scientific American, 85(9):132, August 31, 1901. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v85/
n9/pdf/scientificamerican08311901-132.pdf.

Anonymous:1901:AMb

[3683] Anonymous. An automobile mower. Scientific American, 85(9):132, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n9/pdf/scientificamerican08311901-132a.pdf.

Anonymous:1901:PME

[3684] Anonymous. Palermo–Monreale Electric Incline. Scientific American, 85(9):133–134, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
133.pdf.

Anonymous:1901:LI

[3685] Anonymous. Lengthening of the S. S. “Iroquois”. Scientific American, 85(9):136–137, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
136.pdf.

Anonymous:1901:NLH

[3686] Anonymous. A novel lubricator and heater. Scientific American, 85(9):138, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
138.pdf.

Anonymous:1901:BWS

[3687] Anonymous. The bogue window-sash. Scientific American, 85 (9):138, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
138a.pdf.

Anonymous:1901:WWE

[3688] Anonymous. Works of the Westinghouse Electric and Manufacturing Company. Scientific American, 85(9):138–139, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v85/n9/pdf/scientificamerican08311901-138b.pdf.

Holder:1901:TG

[3689] Charles F. Holder. The turtles of Galapagos. Scientific American, 85(9):139–140, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
139.pdf.

Anonymous:1901:BPWbt

[3690] Anonymous. Business and personal wants. Scientific American, 85(9):141, August 31, 1901. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v85/n9/pdf/scientificamerican08311901-
141.pdf.

Anonymous:1902:Ca

[3691] Anonymous. Correspondence. Scientific American, 86(1):7, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n1/pdf/scientificamerican01041902-7a.pdf.

Anonymous:1902:RPIa

[3692] Anonymous. Recently patented inventions. Scientific American, 86(1):11, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
11.pdf.

Anonymous:1902:IIa

[3693] Anonymous. Index of inventions. Scientific American, 86(1):12, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n1/pdf/scientificamerican01041902-12a.pdf.

Anonymous:1902:NQa

[3694] Anonymous. Notes and queries. Scientific American, 86(1):14–15, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n1/pdf/scientificamerican01041902-14.pdf.

Anonymous:1902:RY

[3695] Anonymous. Retrospect of the year 1901. Scientific American, 86(1):2–3, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
2.pdf.

Anonymous:1902:MTE

[3696] Anonymous. The Marconi transoceanic experiments. Scientific American, 86(1):4, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
4.pdf.

Gratacap:1902:GJD

[3697] L. P. Gratacap. The Great Jurassic Dinosaur. Scientific American, 86(1):5–6, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
5.pdf.

Anonymous:1902:SMN

[3698] Anonymous. A simple mechanical nail-puller. Scientific American, 86(1):6, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
6.pdf.

Anonymous:1902:MUP

[3699] Anonymous. Method of unloading passengers from moving trains. Scientific American, 86(1):6, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
6a.pdf.

Anonymous:1902:BMM

[3700] Anonymous. Bread-making by machinery. Scientific American, 86(1):7, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n1/pdf/scientificamerican01041902-7.pdf.

Anonymous:1902:CT

[3701] Anonymous. The Craddock Tellurian. Scientific American, 86(1):8, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n1/pdf/scientificamerican01041902-8.pdf.

Penniman:1902:SCW

[3702] H. W. H. Penniman. The Santa Cruz wave motor. Scientific American, 86(1):8, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
8a.pdf.

Anonymous:1902:EVG

[3703] Anonymous. An ether-vapor generator. Scientific American, 86(1):9–10, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n1/pdf/scientificamerican01041902-9.pdf.

Anonymous:1902:NSW

[3704] Anonymous. A new system of wireless telegraphy. Scientific American, 86(1):10, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
10.pdf.

Anonymous:1902:BPWa

[3705] Anonymous. Business and personal wants. Scientific American, 86(1):12, January 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n1/pdf/scientificamerican01041902-
12.pdf.

Russell:1902:HMa

[3706] Henry Norris Russell. The heavens in March. Scientific American, 86(10):167, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
167.pdf.

Anonymous:1902:Cb

[3707] Anonymous. Correspondence. Scientific American, 86(10):171, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n10/pdf/scientificamerican03081902-171.pdf.

Anonymous:1902:LNa

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n10/pdf/scientificamerican03081902-175.pdf.

Anonymous:1902:RPIb

[3709] Anonymous. Recently patented inventions. Scientific American, 86(10):176, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
176.pdf.

Anonymous:1902:BPWb

[3710] Anonymous. Business and personal wants. Scientific American, 86(10):176, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
176a.pdf.

Anonymous:1902:IIb

[3711] Anonymous. Index of inventions. Scientific American, 86(10):176–178, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n10/pdf/scientificamerican03081902-176b.pdf.

Anonymous:1902:NQb

[3712] Anonymous. Notes and queries. Scientific American, 86(10):179, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n10/pdf/scientificamerican03081902-179.pdf.

Anonymous:1902:FLD

[3713] Anonymous. The fastest long distance train in the world. Scientific American, 86(10):166, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
166.pdf.

Anonymous:1902:CWTa

[3714] Anonymous. Cost of wireless telegraphy. Scientific American, 86(10):166, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
166a.pdf.

Anonymous:1902:HPP

[3715] Anonymous. A 70,000 horse power plant. Scientific American, 86(10):166, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
166b.pdf.

Anonymous:1902:HVC

[3716] Anonymous. How various countries subsidize their mercantile marine. Scientific American, 86(10):166–167, March 8, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n10/pdf/scientificamerican03081902-166c.pdf.

Anonymous:1902:ETEa

[3717] Anonymous. Earth telephone experiments of M. Ducretet. Scientific American, 86(10):167, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
167a.pdf.

Anonymous:1902:SNa

[3718] Anonymous. Science notes. Scientific American, 86(10):167, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n10/pdf/scientificamerican03081902-167b.pdf.

Anonymous:1902:NTM

[3719] Anonymous. A new type of motor car. Scientific American, 86(10):168, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n10/pdf/scientificamerican03081902-168.pdf.

Crandall:1902:PFE

[3720] William S. Crandall. A pneumatic fire escape and tower. Scientific American, 86(10):168, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
168a.pdf.

Moore:1902:TNF

[3721] Frederick Moore. Test of the new floating drydock at Algiers, La. Scientific American, 86(10):169, March 8, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n10/pdf/scientificamerican03081902-169.pdf.

Collins:1902:GRW

[3722] A. Frederick Collins. The Guarini repeating wireless telegraph system. Scientific American, 86(10):169–170, March 8, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n10/pdf/scientificamerican03081902-169a.pdf.

Fawcett:1902:OHF

[3723] Waldon Fawcett. Ore handling at furnaces. Scientific American, 86(10):172–173, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
172.pdf.

Anonymous:1902:VFM

[3724] Anonymous. The Villard flying machine. Scientific American, 86(10):173, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
173.pdf.

Anonymous:1902:NLP

[3725] Anonymous. A new level and plumb. Scientific American, 86(10):174, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n10/pdf/scientificamerican03081902-174.pdf.

Anonymous:1902:EHW

[3726] Anonymous. Electrical hot-water heating system. Scientific American, 86(10):174, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
174a.pdf.

Anonymous:1902:SIIa

[3727] Anonymous. Simple and interesting inventions. Scientific American, 86(10):174–175, March 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n10/pdf/scientificamerican03081902-
174b.pdf.

Anonymous:1902:LNb

[3728] Anonymous. Legal notes. Scientific American, 86(11):190, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n11/pdf/scientificamerican03151902-190.pdf.

Anonymous:1902:RPIc

[3729] Anonymous. Recently patented inventions. Scientific American, 86 (11):192–193, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
192.pdf.

Anonymous:1902:IIc

[3730] Anonymous. Index of inventions. Scientific American, 86(11):193–195, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n11/pdf/scientificamerican03151902-193a.pdf.

Anonymous:1902:BNBa

[3731] Anonymous. Books: New books, etc. Scientific American, 86 (11):195–196, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
195.pdf.

Anonymous:1902:PNPa

[3732] Anonymous. Patent notes: Patents. Scientific American, 86(11):197, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n11/pdf/scientificamerican03151902-197.pdf.

Anonymous:1902:NQc

[3733] Anonymous. Notes and queries. Scientific American, 86(11):197, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n11/pdf/scientificamerican03151902-197a.pdf.

Anonymous:1902:NCL

[3734] Anonymous. The new Cunard liners. Scientific American, 86(11):182, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n11/pdf/scientificamerican03151902-182.pdf.

Anonymous:1902:PAH

[3735] Anonymous. The Park Avenue Hotel Fire. Scientific American, 86(11):182, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
182a.pdf.

Anonymous:1902:WTN

[3736] Anonymous. Wireless telegraphy in Naval warfare. Scientific American, 86(11):182, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
182b.pdf.

Anonymous:1902:PIP

[3737] Anonymous. A pledge to improve the park avenue tunnel. Scientific American, 86(11):182–183, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
182c.pdf.

Anonymous:1902:BNP

[3738] Anonymous. The British Naval Program for 1903 — reorganizing the fleet. Scientific American, 86(11):183, March 15, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n11/pdf/scientificamerican03151902-183.pdf.

Perkins:1902:SLP

[3739] Frank C. Perkins. The search light in photographic work. Scientific American, 86(11):184, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
184.pdf.

Anonymous:1902:ARB

[3740] Anonymous. The Atbara River bridge. Scientific American, 86 (11):184–185, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
184a.pdf.

Anonymous:1902:MHB

[3741] Anonymous. The Mergenthaler–Horton basket-making machinery. Scientific American, 86(11):185–186, March 15, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n11/pdf/scientificamerican03151902-185.pdf.

Anonymous:1902:ERC

[3742] Anonymous. Electrolytic refining of copper. Scientific American, 86(11):187, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
187.pdf.

Anonymous:1902:NYAa

[3743] Anonymous. The New York Assay Office — melting and refining. Scientific American, 86(11):188–189, March 15, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n11/pdf/scientificamerican03151902-188.pdf.

Anonymous:1902:NET

[3744] Anonymous. A new equatorial telescope for Oxford. Scientific American, 86(11):189–190, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
189.pdf.

Anonymous:1902:BFA

[3745] Anonymous. “Bataclan” — the French Army high-jumping horse. Scientific American, 86(11):191, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
191.pdf.

Anonymous:1902:BPWc

[3746] Anonymous. Business and personal wants. Scientific American, 86(11):193, March 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n11/pdf/scientificamerican03151902-
193.pdf.

Anonymous:1902:Cc

[3747] Anonymous. Correspondence. Scientific American, 86(12):204, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-204.pdf.

Anonymous:1902:LNc

[3748] Anonymous. Legal notes. Scientific American, 86(12):210, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-210.pdf.

Anonymous:1902:RPId

[3749] Anonymous. Recently patented inventions. Scientific American, 86(12):211, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
211.pdf.

Anonymous:1902:IId

[3750] Anonymous. Index of inventions. Scientific American, 86(12):212–213, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-212.pdf.

Anonymous:1902:BNBb

[3751] Anonymous. Books: New books, etc. Scientific American, 86(12):213, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-213.pdf.

Anonymous:1902:NQd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-213a.pdf.

Anonymous:1902:PNPb

[3753] Anonymous. Patent notes: Patents. Scientific American, 86(12):215, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-215.pdf.

Anonymous:1902:NES

[3754] Anonymous. The new era for the steam engine. Scientific American, 86(12):200, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
200.pdf.

Anonymous:1902:TRBa

[3755] Anonymous. Temporary relief at the Brooklyn Bridge. Scientific American, 86(12):200, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
200a.pdf.

Anonymous:1902:UVM

[3756] Anonymous. Use of various motors in aerostatics. Scientific American, 86(12):200–201, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
200b.pdf.

Irvine:1902:ISI

[3757] Ansley Irvine. Incomes of successful inventors. Scientific American, 86(12):201, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
201.pdf.

Anonymous:1902:LMW

[3758] Anonymous. Lead mining in Wisconsin. Scientific American, 86(12):201, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
201a.pdf.

Anonymous:1902:SNb

[3759] Anonymous. Science notes. Scientific American, 86(12):201, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-201b.pdf.

Anonymous:1902:MID

[3760] Anonymous. The Motsinger ignition dynamo. Scientific American, 86(12):202, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
202.pdf.

DesJardins:1902:CT

[3761] B. M. Des Jardins. The computing triangle. Scientific American, 86(12):202–203, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
202a.pdf.

Wood:1902:SCA

[3762] R. W. Wood. Solid carbonic acid from “sparklets”. Scientific American, 86(12):203, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
203.pdf.

Anonymous:1902:SDAa

[3763] Anonymous. Some disastrous acetylene gas explosions. Scientific American, 86(12):203, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
203a.pdf.

Anonymous:1902:ENY

[3764] Anonymous. The engines of the New York Edison Power Station. Scientific American, 86(12):205, March 22, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n12/pdf/scientificamerican03221902-205.pdf.

Anonymous:1902:CI

[3765] Anonymous. The cranberry industry. Scientific American, 86 (12):206–207, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
206.pdf.

Perkins:1902:TPV

[3766] Frank C. Perkins. Three-phase 10,000-volt railway at Gross-Lichterfelde. Scientific American, 86(12):207, March 22, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n12/pdf/scientificamerican03221902-207.pdf.

Anonymous:1902:PDa

[3767] Anonymous. Patent department. Scientific American, 86(12):208, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-208.pdf.

Anonymous:1902:NSB

[3768] Anonymous. A new spring block. Scientific American, 86(12):208, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-208a.pdf.

Anonymous:1902:DMH

[3769] Anonymous. Double-moldboard hilling-up plow. Scientific American, 86(12):208, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
208b.pdf.

Anonymous:1902:HBa

[3770] Anonymous. Hydrocarbon burner. Scientific American, 86(12):208, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n12/pdf/scientificamerican03221902-208c.pdf.

Anonymous:1902:SIIb

[3771] Anonymous. Simple and interesting inventions. Scientific American, 86(12):209, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
209.pdf.

Anonymous:1902:BPWd

[3772] Anonymous. Business and personal wants. Scientific American, 86 (12):211–212, March 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n12/pdf/scientificamerican03221902-
211a.pdf.

Russell:1902:HAa

[3773] Henry Norris Russell. The heavens in April, 1902. Scientific American, 86(13):219, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
219.pdf.

Anonymous:1902:Cd

[3774] Anonymous. Correspondence. Scientific American, 86(13):222, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-222.pdf.

Anonymous:1902:RPIe

[3775] Anonymous. Recently patented inventions. Scientific American, 86(13):228, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
228.pdf.

Anonymous:1902:BPWe

[3776] Anonymous. Business and personal wants. Scientific American, 86(13):228, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
228a.pdf.

Anonymous:1902:IIe

[3777] Anonymous. Index of inventions. Scientific American, 86(13):228–230, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-228b.pdf.

Anonymous:1902:BNBc

[3778] Anonymous. Books: New books, etc. Scientific American, 86(13):230, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-230.pdf.

Anonymous:1902:NQe

[3779] Anonymous. Notes and queries. Scientific American, 86(13):231, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-231.pdf.

Anonymous:1902:SER

[3780] Anonymous. Speed on the elevated railroad curves. Scientific American, 86(13):218, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
218.pdf.

Anonymous:1902:MSB

[3781] Anonymous. Moving sidewalk for the Brooklyn Bridge. Scientific American, 86(13):218, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
218a.pdf.

Anonymous:1902:FPTa

[3782] Anonymous. Faulty pipe threads. Scientific American, 86(13):218, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-218b.pdf.

Anonymous:1902:NTD

[3783] Anonymous. A novel turbine destroyer. Scientific American, 86(13):218, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-218c.pdf.

Anonymous:1902:APR

[3784] Anonymous. Armor plate for the Russian Government. Scientific American, 86(13):218, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
218d.pdf.

Anonymous:1902:FIH

[3785] Anonymous. Fire inspection of hotels. Scientific American, 86(13):218, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-218e.pdf.

Anonymous:1902:OPT

[3786] Anonymous. Objections to the Pennsylvania Tunnel Bill. Scientific American, 86(13):218, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
218f.pdf.

Anonymous:1902:SVSa

[3787] Anonymous. Sailing vessels in the Suez Canal. Scientific American, 86(13):218, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
218g.pdf.

Asher:1902:EPD

[3788] James Asher. Electricity and power direct from heat. Scientific American, 86(13):219, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
219a.pdf.

Townsend:1902:CIR

[3789] Irving U. Townsend. A curious instance of re-invention. Scientific American, 86(13):220, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
220.pdf.

Anonymous:1902:OBW

[3790] Anonymous. An oil-burning water-tube locomotive boiler. Scientific American, 86(13):220, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
220a.pdf.

Anonymous:1902:CPAa

[3791] Anonymous. Cave-in at the park avenue rapid transit tunnel. Scientific American, 86(13):221, March 29, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n13/pdf/scientificamerican03291902-221.pdf.

Anonymous:1902:SPCa

[3792] Anonymous. Shell practice in city armories. Scientific American, 86(13):222–223, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
222a.pdf.

Anonymous:1902:SAa

[3793] Anonymous. The Severo airship. Scientific American, 86(13):224, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-224.pdf.

Anonymous:1902:ADa

[3794] Anonymous. Automobile department. Scientific American, 86(13):225, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-225.pdf.

Anonymous:1902:FFa

[3795] Anonymous. Flexibility of the frame. Scientific American, 86(13):225, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-225a.pdf.

Anonymous:1902:ACBa

[3796] Anonymous. Automobile coupé built for the Sultan of Morocco. Scientific American, 86(13):226, March 29, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n13/pdf/scientificamerican03291902-226.pdf.

Anonymous:1902:NTWa

[3797] Anonymous. A novel three-wheel automobile. Scientific American, 86(13):226–227, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
226a.pdf.

Anonymous:1902:MS

[3798] Anonymous. A motor sled. Scientific American, 86(13):227, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-227.pdf.

Anonymous:1902:NWL

[3799] Anonymous. A new water-level regulator for steam boilers. Scientific American, 86(13):227, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n13/pdf/scientificamerican03291902-
227a.pdf.

Anonymous:1902:ANa

[3800] Anonymous. Automobile novelties. Scientific American, 86(13):227, March 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n13/pdf/scientificamerican03291902-227b.pdf.

Anonymous:1902:LNd

[3801] Anonymous. Legal notes. Scientific American, 86(14):244, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-244.pdf.

Anonymous:1902:RPIf

[3802] Anonymous. Recently patented inventions. Scientific American, 86(14):245, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
245.pdf.

Anonymous:1902:IIf

[3803] Anonymous. Index of inventions. Scientific American, 86(14):246, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-246.pdf.

Anonymous:1902:NQf

[3804] Anonymous. Notes and queries. Scientific American, 86(14):249, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-249.pdf.

Anonymous:1902:RFL

[3805] Anonymous. Recent fire losses in America. Scientific American, 86(14):234, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
234.pdf.

Anonymous:1902:ERBa

[3806] Anonymous. East River Bridge No. 3. Scientific American, 86(14):234, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-234a.pdf.

Anonymous:1902:HD

[3807] Anonymous. A healthy decadence. Scientific American, 86(14):234, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-234b.pdf.

Anonymous:1902:SE

[3808] Anonymous. Safe explosives. Scientific American, 86(14):234, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-234c.pdf.

Anonymous:1902:WWB

[3809] Anonymous. The work of the Weather Bureau in 1901. Scientific American, 86(14):234, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
234d.pdf.

Anonymous:1902:CAE

[3810] Anonymous. Coming automobile endurance tests. Scientific American, 86(14):234–235, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
234e.pdf.

Anonymous:1902:SNc

[3811] Anonymous. Scientific nomenclature. Scientific American, 86(14):235, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-235.pdf.

Anonymous:1902:BHS

[3812] Anonymous. A brief history of street gas. Scientific American, 86(14):235, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
235a.pdf.

Anonymous:1902:SNd

[3813] Anonymous. Science notes. Scientific American, 86(14):235, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-235b.pdf.

Moore:1902:SWH

[3814] Frederick Moore. The South widening her field of agriculture. Scientific American, 86(14):236, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
236.pdf.

Anonymous:1902:PAT

[3815] Anonymous. The park avenue tunnel cave-in. Scientific American, 86(14):237, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
237.pdf.

Shepstone:1902:NBF

[3816] H. J. Shepstone. The new Bermuda Floating Dock. Scientific American, 86(14):237–238, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
237a.pdf.

Anonymous:1902:GPB

[3817] Anonymous. The Glass Palace in “Beauty and the Beast”. Scientific American, 86(14):238–239, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
238.pdf.

Trowbridge:1902:LAB

[3818] John Trowbridge. Lightning above and below water. Scientific American, 86(14):239, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
239.pdf.

Anonymous:1902:RUP

[3819] Anonymous. Reconstruction of the Union Pacific Railroad. Scientific American, 86(14):240, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
240.pdf.

Beard:1902:WIF

[3820] J. Carter Beard. Where instinct fails. Scientific American, 86 (14):241–242, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
241.pdf.

Anonymous:1902:PDb

[3821] Anonymous. Patent department. Scientific American, 86(14):242, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-242.pdf.

Anonymous:1902:CBC

[3822] Anonymous. The chainless-bicycle coaster-brake gear-changer. Scientific American, 86(14):243, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
243.pdf.

Anonymous:1902:HSP

[3823] Anonymous. The Hartford steel-plate exhauster. Scientific American, 86(14):243, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
243a.pdf.

Anonymous:1902:RLB

[3824] Anonymous. The Regan locomotive boiler. Scientific American, 86(14):243, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
243b.pdf.

Anonymous:1902:SI

[3825] Anonymous. Station indicator. Scientific American, 86(14):243–244, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n14/pdf/scientificamerican04051902-243c.pdf.

Anonymous:1902:BPWf

[3826] Anonymous. Business and personal wants. Scientific American, 86(14):245, April 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n14/pdf/scientificamerican04051902-
245a.pdf.

Anonymous:1902:Ce

[3827] Anonymous. Correspondence. Scientific American, 86(15):257, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n15/pdf/scientificamerican04121902-257.pdf.

Anonymous:1902:RPIg

[3828] Anonymous. Recently patented inventions. Scientific American, 86(15):263, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
263.pdf.

Anonymous:1902:IIg

[3829] Anonymous. Index of inventions. Scientific American, 86(15):263, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n15/pdf/scientificamerican04121902-263b.pdf.

Anonymous:1902:PNPc

[3830] Anonymous. Patent notes: Patents. Scientific American, 86(15):267, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n15/pdf/scientificamerican04121902-267.pdf.

Anonymous:1902:NQg

[3831] Anonymous. Notes and queries. Scientific American, 86(15):267, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n15/pdf/scientificamerican04121902-267a.pdf.

Anonymous:1902:LST

[3832] Anonymous. The latest steam turbine plant. Scientific American, 86(15):252, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
252.pdf.

Anonymous:1902:LLE

[3833] Anonymous. The largest locomotive engine in Great Britain. Scientific American, 86(15):252, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
252a.pdf.

Anonymous:1902:NMB

[3834] Anonymous. A novel method of bridge erection. Scientific American, 86(15):252, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
252b.pdf.

Anonymous:1902:AME

[3835] Anonymous. American methods for English workmen. Scientific American, 86(15):252, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
252c.pdf.

Anonymous:1902:CWTb

[3836] Anonymous. A comparison of water-tube boilers. Scientific American, 86(15):252, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
252d.pdf.

Anonymous:1902:FRE

[3837] Anonymous. French railway enterprise in Abyssinia. Scientific American, 86(15):252–253, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
252e.pdf.

Walsh:1902:EWF

[3838] George E. Walsh. The expansion of winter farming. Scientific American, 86(15):253, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
253.pdf.

Anonymous:1902:SNe

[3839] Anonymous. Science notes. Scientific American, 86(15):253, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n15/pdf/scientificamerican04121902-253a.pdf.

Fawcett:1902:BL

[3840] Waldon Fawcett. Broadside launchings. Scientific American, 86 (15):254, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
254.pdf.

Anonymous:1902:ISTa

[3841] Anonymous. The improved simplex typewriter. Scientific American, 86(15):255, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
255.pdf.

Anonymous:1902:NSTa

[3842] Anonymous. Night signal for trolley lines. Scientific American, 86(15):255, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
255a.pdf.

Clapp:1902:SPN

[3843] A. W. Clapp. Southern Pacific new line across Great Salt Lake. Scientific American, 86(15):256, April 12, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n15/pdf/scientificamerican04121902-256.pdf.

Anonymous:1902:WTL

[3844] Anonymous. A water-tube locomotive boiler. Scientific American, 86(15):256, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
256a.pdf.

Schafer:1902:CFB

[3845] L. C. Schafer. Capsizing of the French bark “Asie”. Scientific American, 86(15):258, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
258.pdf.

Anonymous:1902:OLH

[3846] Anonymous. An object lesson in heavy ordnance. Scientific American, 86(15):258, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
258a.pdf.

Willey:1902:IRT

[3847] Day Allen Willey. Indicating and recording the tides. Scientific American, 86(15):259–260, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
259.pdf.

Frederick:1902:SAP

[3848] M. C. Frederick. Something about pelicans. Scientific American, 86(15):261, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
261.pdf.

Anonymous:1902:FPE

[3849] Anonymous. The future of photography: An early prophecy by Arago. Scientific American, 86(15):262, April 12, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n15/pdf/scientificamerican04121902-262.pdf.

Anonymous:1902:BPWg

[3850] Anonymous. Business and personal wants. Scientific American, 86(15):263, April 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n15/pdf/scientificamerican04121902-
263a.pdf.

Anonymous:1902:LNe

[3851] Anonymous. Legal notes. Scientific American, 86(16):280, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n16/pdf/scientificamerican04191902-280.pdf.

Anonymous:1902:RPIh

[3852] Anonymous. Recently patented inventions. Scientific American, 86(16):281, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
281.pdf.

Anonymous:1902:IIh

[3853] Anonymous. Index of inventions. Scientific American, 86(16):281, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n16/pdf/scientificamerican04191902-281b.pdf.

Anonymous:1902:PNPd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n16/pdf/scientificamerican04191902-285.pdf.

Anonymous:1902:NQh

[3855] Anonymous. Notes and queries. Scientific American, 86(16):285, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n16/pdf/scientificamerican04191902-285a.pdf.

Anonymous:1902:ASOa

[3856] Anonymous. The American scholarships at Oxford. Scientific American, 86(16):270, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
270.pdf.

Anonymous:1902:UU

[3857] Anonymous. Unjust and ungenerous. Scientific American, 86(16):270, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n16/pdf/scientificamerican04191902-270a.pdf.

Anonymous:1902:AAa

[3858] Anonymous. The airship and the aeroplane. Scientific American, 86(16):270, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
270b.pdf.

Anonymous:1902:MWI

[3859] Anonymous. Men of wealth as inventors. Scientific American, 86 (16):270–271, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
270c.pdf.

Anonymous:1902:RAA

[3860] Anonymous. Record of American auto shows. Scientific American, 86(16):271, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
271.pdf.

Anonymous:1902:SNf

[3861] Anonymous. Science notes. Scientific American, 86(16):271, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n16/pdf/scientificamerican04191902-271a.pdf.

Anonymous:1902:ESF

[3862] Anonymous. Electric searchlight for field service. Scientific American, 86(16):272, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
272.pdf.

Anonymous:1902:RTT

[3863] Anonymous. Railroad train taking its own photograph. Scientific American, 86(16):272, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
272a.pdf.

Anonymous:1902:HMD

[3864] Anonymous. Home-made device for decorating glass. Scientific American, 86(16):273, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
273.pdf.

Anonymous:1902:NVS

[3865] Anonymous. A new variable speed transmission. Scientific American, 86(16):273–274, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
273a.pdf.

Byrn:1902:MWT

[3866] Edward W. Byrn. The mechanical work of the twelfth census. Scientific American, 86(16):275, April 19, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n16/pdf/scientificamerican04191902-275.pdf.

vonBlon:1902:BW

[3867] John L. von Blon. A boat on wheels. Scientific American, 86(16):276, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n16/pdf/scientificamerican04191902-276.pdf.

Geissel:1902:SIF

[3868] Henry L. Geissel. Some interesting features of the Krupp Works at Essen. Scientific American, 86(16):276–277, April 19, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n16/pdf/scientificamerican04191902-276a.pdf.

Muckermann:1902:IG

[3869] Herman Muckermann. An insect geometrician. Scientific American, 86(16):278, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
278.pdf.

Anonymous:1902:PDc

[3870] Anonymous. Patent department. Scientific American, 86(16):279, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n16/pdf/scientificamerican04191902-279.pdf.

Anonymous:1902:HLT

[3871] Anonymous. The Hanson–Lee typewriter. Scientific American, 86 (16):279–280, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
279a.pdf.

Anonymous:1902:BPWh

[3872] Anonymous. Business and personal wants. Scientific American, 86(16):281, April 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n16/pdf/scientificamerican04191902-
281a.pdf.

Anonymous:1902:Cf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-293.pdf.

Anonymous:1902:RPIi

[3874] Anonymous. Recently patented inventions. Scientific American, 86(17):298–299, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
298.pdf.

Anonymous:1902:IIi

[3875] Anonymous. Index of inventions. Scientific American, 86(17):299, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-299a.pdf.

Anonymous:1902:BNBd

[3876] Anonymous. Books: New books, etc. Scientific American, 86(17):301, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-301.pdf.

Anonymous:1902:PNPe

[3877] Anonymous. Patent notes: Patents. Scientific American, 86(17):303, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-303.pdf.

Anonymous:1902:NQi

[3878] Anonymous. Notes and queries. Scientific American, 86(17):303, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-303a.pdf.

Anonymous:1902:RWP

[3879] Anonymous. Reconstruction of West Point Academy. Scientific American, 86(17):288, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
288.pdf.

Anonymous:1902:FW

[3880] Anonymous. The Fleets of the world. Scientific American, 86(17):288, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-288a.pdf.

Anonymous:1902:CAM

[3881] Anonymous. The cost of ammunition at Manila and Santiago. Scientific American, 86(17):288, April 26, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n17/pdf/scientificamerican04261902-288b.pdf.

Anonymous:1902:RPC

[3882] Anonymous. Report of the Prussian Commission on American Railroads. Scientific American, 86(17):288, April 26, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n17/pdf/scientificamerican04261902-288c.pdf.

Anonymous:1902:WTP

[3883] Anonymous. Wireless telegraphy and the promoter. Scientific American, 86(17):288–289, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
288d.pdf.

Benjamin:1902:RMN

[3884] Marcus Benjamin. Recent meeting of the National Academy of Sciences in Washington. Scientific American, 86(17):289, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-289.pdf.

Anonymous:1902:SNg

[3885] Anonymous. Science notes. Scientific American, 86(17):289, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-289a.pdf.

Anonymous:1902:IEC

[3886] Anonymous. The improvement of the Erie Canal. Scientific American, 86(17):290–291, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
290.pdf.

Anonymous:1902:SSG

[3887] Anonymous. A self-starting gasoline motor of novel design. Scientific American, 86(17):291, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
291.pdf.

Anonymous:1902:GEFa

[3888] Anonymous. A gas-engine friction clutch. Scientific American, 86 (17):291–292, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
291a.pdf.

Anonymous:1902:MSRa

[3889] Anonymous. The manufacture of steel rails. Scientific American, 86(17):292, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
292.pdf.

Anonymous:1902:CEJ

[3890] Anonymous. A canal excavator for Java. Scientific American, 86(17):294, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
294.pdf.

Aaron:1902:IW

[3891] S. Frank Aaron. Insects in winter. Scientific American, 86(17):295, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n17/pdf/scientificamerican04261902-295.pdf.

Anonymous:1902:BIC

[3892] Anonymous. Breaking up 15-inch cast-iron guns. Scientific American, 86(17):295–296, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
295a.pdf.

Anonymous:1902:AWTa

[3893] Anonymous. An acetylene wireless telephone apparatus. Scientific American, 86(17):296, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
296.pdf.

Anonymous:1902:RFE

[3894] Anonymous. Roman Forum excavations. Scientific American, 86 (17):297, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
297.pdf.

Anonymous:1902:BPWi

[3895] Anonymous. Business and personal wants. Scientific American, 86(17):299, April 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n17/pdf/scientificamerican04261902-
299.pdf.

Russell:1902:HMb

[3896] Henry Norris Russell. The heavens in May. Scientific American, 86(18):307, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
307a.pdf.

Anonymous:1902:RPIj

[3897] Anonymous. Recently patented inventions. Scientific American, 86(18):316, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
316.pdf.

Anonymous:1902:IIj

[3898] Anonymous. Index of inventions. Scientific American, 86(18):317, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-317a.pdf.

Anonymous:1902:BNBe

[3899] Anonymous. Books: New books, etc. Scientific American, 86(18):318, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-318.pdf.

Anonymous:1902:NQj

[3900] Anonymous. Notes and queries. Scientific American, 86(18):320, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-320.pdf.

Anonymous:1902:PNPf

[3901] Anonymous. Patent notes: Patents. Scientific American, 86(18):321, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-321.pdf.

Anonymous:1902:PCT

[3902] Anonymous. The Panama Canal Treaty. Scientific American, 86(18):306, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
306.pdf.

Anonymous:1902:GSC

[3903] Anonymous. Great steamship combine. Scientific American, 86(18):306, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-306a.pdf.

Anonymous:1902:STS

[3904] Anonymous. Steering with twin-screw engines. Scientific American, 86(18):306, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
306b.pdf.

Anonymous:1902:CAA

[3905] Anonymous. Combining the airship and the aeroplane. Scientific American, 86(18):306, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
306c.pdf.

Anonymous:1902:SCS

[3906] Anonymous. Sanitary condition of street cars. Scientific American, 86(18):306–307, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
306d.pdf.

Anonymous:1902:LKE

[3907] Anonymous. Lord Kelvin — England’s greatest living scientist. Scientific American, 86(18):307, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
307.pdf.

Anonymous:1902:NAU

[3908] Anonymous. New airship under construction for the British War Office. Scientific American, 86(18):308, May 3, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n18/pdf/scientificamerican05031902-308.pdf.

Anonymous:1902:PLSa

[3909] Anonymous. Powerful locomotive for suburban service. Scientific American, 86(18):308–309, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
308a.pdf.

Anonymous:1902:GTS

[3910] Anonymous. The growth of the transatlantic steamship. Scientific American, 86(18):309–310, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
309.pdf.

Perkins:1902:UER

[3911] Frank C. Perkins. Underground electric railways abroad. Scientific American, 86(18):311, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
311.pdf.

Hamilton:1902:SIB

[3912] S. Harbert Hamilton. The sisal industry in the Bahamas. Scientific American, 86(18):312, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
312.pdf.

Willey:1902:MPT

[3913] Day Allen Willey. Mechanism for predicting the tides. Scientific American, 86(18):312–313, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
312a.pdf.

Anonymous:1902:ZAC

[3914] Anonymous. Zuni, an ancient city of the Pueblos. Scientific American, 86(18):313–314, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
313.pdf.

Anonymous:1902:PDd

[3915] Anonymous. Patent department. Scientific American, 86(18):314–315, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-314.pdf.

Anonymous:1902:ACBb

[3916] Anonymous. Adjustable claw-bar. Scientific American, 86(18):315, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-315.pdf.

Anonymous:1902:OIa

[3917] Anonymous. Oddities in inventions. Scientific American, 86(18):315, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-315a.pdf.

Anonymous:1902:ADG

[3918] Anonymous. Automatic drive-gate. Scientific American, 86(18):315, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-315b.pdf.

Anonymous:1902:TS

[3919] Anonymous. Tree-stand. Scientific American, 86(18):315, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n18/pdf/scientificamerican05031902-315c.pdf.

Anonymous:1902:BPWj

[3920] Anonymous. Business and personal wants. Scientific American, 86(18):317, May 3, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n18/pdf/scientificamerican05031902-
317.pdf.

Anonymous:1902:Cg

[3921] Anonymous. Correspondence. Scientific American, 86(19):329–330, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-329a.pdf.

Anonymous:1902:RPIk

[3922] Anonymous. Recently patented inventions. Scientific American, 86(19):334, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
334.pdf.

Anonymous:1902:IIk

[3923] Anonymous. Index of inventions. Scientific American, 86(19):334, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-334b.pdf.

Anonymous:1902:PNPg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-339.pdf.

Anonymous:1902:NQk

[3925] Anonymous. Notes and queries. Scientific American, 86(19):339, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-339a.pdf.

Anonymous:1902:FAR

[3926] Anonymous. Fastest authentic railroad run on record. Scientific American, 86(19):324, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
324.pdf.

Anonymous:1902:PSNa

[3927] Anonymous. Perils of submarine navigation. Scientific American, 86(19):324, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
324a.pdf.

Anonymous:1902:GGA

[3928] Anonymous. The Gathmann gun again. Scientific American, 86 (19):324, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
324b.pdf.

Anonymous:1902:FEW

[3929] Anonymous. Fessenden’s electric wave-detector. Scientific American, 86(19):324, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
324c.pdf.

Anonymous:1902:AWTb

[3930] Anonymous. Attuned wireless telegraphy. Scientific American, 86 (19):324, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
324d.pdf.

Anonymous:1902:NAB

[3931] Anonymous. The Naval Appropriation Bill. Scientific American, 86(19):324–325, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
324e.pdf.

Anonymous:1902:ABL

[3932] Anonymous. American and British locomotives in Egypt. Scientific American, 86(19):325, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
325.pdf.

Anonymous:1902:RAE

[3933] Anonymous. The recent automobile endurance test. Scientific American, 86(19):325, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
325a.pdf.

Anonymous:1902:SNh

[3934] Anonymous. Science notes. Scientific American, 86(19):325, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-325b.pdf.

Anonymous:1902:NJB

[3935] Anonymous. The new Japanese battleship “Mikasa”. Scientific American, 86(19):326–327, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
326.pdf.

Anonymous:1902:NSSa

[3936] Anonymous. New system of street cleaning. Scientific American, 86(19):327, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
327.pdf.

Anonymous:1902:THE

[3937] Anonymous. The testing of highspeed engines. Scientific American, 86(19):327–328, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
327a.pdf.

Anonymous:1902:EMC

[3938] Anonymous. Experiments of M. Curie. Scientific American, 86(19):328, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-328.pdf.

Anonymous:1902:ASa

[3939] Anonymous. Ash-sifter. Scientific American, 86(19):328, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-328a.pdf.

Anonymous:1902:NMU

[3940] Anonymous. A new means of using compressed air in the manufacture of glassware. Scientific American, 86(19):329, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-329.pdf.

Anonymous:1902:LSD

[3941] Anonymous. London smoke deposits. Scientific American, 86(19):330, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-330.pdf.

Willey:1902:NUT

[3942] Day Allen Willey. Novel uses for the trolley current. Scientific American, 86(19):330–332, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
330a.pdf.

Anonymous:1902:NSL

[3943] Anonymous. A new stereopticon light. Scientific American, 86(19):332, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n19/pdf/scientificamerican05101902-332.pdf.

Holder:1902:TWA

[3944] Charles F. Holder. Tameness of wild animals. Scientific American, 86(19):333, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
333.pdf.

Anonymous:1902:BPWk

[3945] Anonymous. Business and personal wants. Scientific American, 86(19):334, May 10, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n19/pdf/scientificamerican05101902-
334a.pdf.

Anonymous:1902:Ch

[3946] Anonymous. Correspondence. Scientific American, 86(2):22–23, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-22a.pdf.

Anonymous:1902:RPIl

[3947] Anonymous. Recently patented inventions. Scientific American, 86(2):28, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
28.pdf.

Anonymous:1902:IIl

[3948] Anonymous. Index of inventions. Scientific American, 86(2):28, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-28b.pdf.

Anonymous:1902:PNPh

[3949] Anonymous. Patent notes: Patents. Scientific American, 86(2):31, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-31.pdf.

Anonymous:1902:NQl

[3950] Anonymous. Notes and queries. Scientific American, 86(2):31, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-31a.pdf.

Anonymous:1902:AMPa

[3951] Anonymous. The alcohol motor and its possibilities. Scientific American, 86(2):18, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
18.pdf.

Anonymous:1902:LSM

[3952] Anonymous. Large stationary and marine engine units. Scientific American, 86(2):18, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
18a.pdf.

Anonymous:1902:ACW

[3953] Anonymous. Automatic closing of watertight bulkheads. Scientific American, 86(2):18, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
18b.pdf.

Irvine:1902:FIV

[3954] Ansley Irvine. The first iron vessel in Great Britain. Scientific American, 86(2):18–19, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
18c.pdf.

Benjamin:1902:CK

[3955] Marcus Benjamin. Clarence king. Scientific American, 86(2):19, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-19.pdf.

Anonymous:1902:SNi

[3956] Anonymous. Science notes. Scientific American, 86(2):19, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-19a.pdf.

Anonymous:1902:HL

[3957] Anonymous. A hardware locomotive. Scientific American, 86 (2):20, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
20.pdf.

Anonymous:1902:LBM

[3958] Anonymous. The launch of the battleship “Missouri”. Scientific American, 86(2):20, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
20a.pdf.

Anonymous:1902:OPS

[3959] Anonymous. Opening the power station of the Manhattan Elevated Railway. Scientific American, 86(2):21–22, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-21.pdf.

Baker:1902:CCA

[3960] Thomas R. Baker. A contrivance for collecting atmospheric dust for microscopic examination. Scientific American, 86(2):22, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-22.pdf.

Fawcett:1902:MRM

[3961] Waldon Fawcett. Making roads by machinery. Scientific American, 86(2):24–25, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
24.pdf.

Anonymous:1902:NL

[3962] Anonymous. The Nernst lamp. Scientific American, 86(2):25–26, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-25.pdf.

Anonymous:1902:PDe

[3963] Anonymous. Patent department. Scientific American, 86(2):26–27, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-26.pdf.

Anonymous:1902:ACT

[3964] Anonymous. An automatic car tipple for mine railways. Scientific American, 86(2):27, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
27.pdf.

Anonymous:1902:NREa

[3965] Anonymous. A new rotary engine. Scientific American, 86(2):27, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n2/pdf/scientificamerican01111902-27a.pdf.

Anonymous:1902:BPWl

[3966] Anonymous. Business and personal wants. Scientific American, 86(2):28, January 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n2/pdf/scientificamerican01111902-
28a.pdf.

Anonymous:1902:LNf

[3967] Anonymous. Legal notes. Scientific American, 86(20):352, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-352.pdf.

Anonymous:1902:RPIm

[3968] Anonymous. Recently patented inventions. Scientific American, 86(20):354, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
354.pdf.

Anonymous:1902:IIm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-355a.pdf.

Anonymous:1902:PNPi

[3970] Anonymous. Patent notes: Patents. Scientific American, 86(20):357, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-357.pdf.

Anonymous:1902:NQm

[3971] Anonymous. Notes and queries. Scientific American, 86(20):357, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-357a.pdf.

Anonymous:1902:OVS

[3972] Anonymous. Our vast steel industry. Scientific American, 86(20):342, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-342.pdf.

Anonymous:1902:SRTa

[3973] Anonymous. The stations of the rapid transit subway. Scientific American, 86(20):342, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
342a.pdf.

Anonymous:1902:ESH

[3974] Anonymous. Electricity or steam for high-speed railways. Scientific American, 86(20):342, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
342b.pdf.

Anonymous:1902:RRB

[3975] Anonymous. Roentgen ray burns. Scientific American, 86(20):342, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-342c.pdf.

Anonymous:1902:GAS

[3976] Anonymous. Guarding against the submarine. Scientific American, 86(20):342–343, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
342d.pdf.

Anonymous:1902:USM

[3977] Anonymous. The United States and the metric system. Scientific American, 86(20):343, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
343.pdf.

Anonymous:1902:ASb

[3978] Anonymous. Admiral Sampson. Scientific American, 86(20):343, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-343a.pdf.

Anonymous:1902:Pa

[3979] Anonymous. The photorama. Scientific American, 86(20):344–345, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-344.pdf.

Anonymous:1902:VLI

[3980] Anonymous. The Venice of Long Island. Scientific American, 86 (20):345, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
345.pdf.

Anonymous:1902:MSRb

[3981] Anonymous. Manufacture of steel rails — II. Scientific American, 86(20):346–347, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
346.pdf.

Anonymous:1902:NH

[3982] Anonymous. A new heater. Scientific American, 86(20):347, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-347.pdf.

Perkins:1902:IST

[3983] Frank C. Perkins. The introduction of the steam turbine for light and power work. Scientific American, 86(20):348–349, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-348.pdf.

Holder:1902:BTF

[3984] Charles Frederick Holder. Buttes and their formation. Scientific American, 86(20):349–350, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
349.pdf.

Anonymous:1902:ADb

[3985] Anonymous. Automobile department. Scientific American, 86(20):350, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-350.pdf.

Anonymous:1902:SAW

[3986] Anonymous. The Simms armored war-car. Scientific American, 86(20):350–351, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
350a.pdf.

Anonymous:1902:SMA

[3987] Anonymous. Some modern automobile accumulators. Scientific American, 86(20):351–352, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
351.pdf.

Beasley:1902:MCA

[3988] Walter L. Beasley. Monster crustaceans at the American Museum of Natural History. Scientific American, 86(20):353, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n20/pdf/scientificamerican05171902-353.pdf.

Anonymous:1902:BPWm

[3989] Anonymous. Business and personal wants. Scientific American, 86(20):355, May 17, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n20/pdf/scientificamerican05171902-
355.pdf.

Anonymous:1902:Cj

[3990] Anonymous. Correspondence. Scientific American, 86(21):364, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-364.pdf.

Anonymous:1902:RPIn

[3991] Anonymous. Recently patented inventions. Scientific American, 86(21):370, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
370.pdf.

Anonymous:1902:IIn

[3992] Anonymous. Index of inventions. Scientific American, 86(21):370, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-370b.pdf.

Anonymous:1902:BNBf

[3993] Anonymous. Books: New books, etc. Scientific American, 86(21):372, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-372.pdf.

Anonymous:1902:NQn

[3994] Anonymous. Notes and queries. Scientific American, 86(21):373, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-373.pdf.

Anonymous:1902:VEI

[3995] Anonymous. Volcanic eruptions and the Isthmian Canal. Scientific American, 86(21):360, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
360.pdf.

Anonymous:1902:HT

[3996] Anonymous. Havana transformed. Scientific American, 86(21):360, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-360a.pdf.

Anonymous:1902:LFSa

[3997] Anonymous. Liquid fuel for steamships. Scientific American, 86 (21):360, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
360b.pdf.

Anonymous:1902:FQ

[3998] Anonymous. Fusion of quartz. Scientific American, 86(21):360–361, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-360c.pdf.

Anonymous:1902:NRD

[3999] Anonymous. A new revised design patent law. Scientific American, 86(21):361, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
361.pdf.

Ashley:1902:BS

[4000] Ernest Ashley. The British submarines. Scientific American, 86(21):361, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-361a.pdf.

Anonymous:1902:SNj

[4001] Anonymous. Science notes. Scientific American, 86(21):361, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-361b.pdf.

Anonymous:1902:NTG

[4002] Anonymous. A novel tool-grinding attachment. Scientific American, 86(21):362, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
362.pdf.

Anonymous:1902:TWS

[4003] Anonymous. Total wreck of the Severo airship. Scientific American, 86(21):362–363, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
362a.pdf.

Fawcett:1902:LAW

[4004] Waldon Fawcett. The latest advance in wireless telephony. Scientific American, 86(21):363, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
363.pdf.

Anonymous:1902:DWI

[4005] Anonymous. The disaster in the West Indies and its explanation. Scientific American, 86(21):365, May 24, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n21/pdf/scientificamerican05241902-365.pdf.

Fawoett:1902:HMS

[4006] Waldon Fawoett. Handling material in shipyards. Scientific American, 86(21):366–368, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
366.pdf.

Headland:1902:QCT

[4007] Isaac Taylor Headland. Queer Chinese trees. Scientific American, 86(21):368, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
368.pdf.

Beasley:1902:SFG

[4008] Walter L. Beasley. Story of the fishskin garments of the Amur tribes of East Siberia. Scientific American, 86(21):368, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-368a.pdf.

Anonymous:1902:PDf

[4009] Anonymous. Patent department. Scientific American, 86(21):369, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-369.pdf.

Anonymous:1902:LA

[4010] Anonymous. Lathe attachment. Scientific American, 86(21):369, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-369a.pdf.

Anonymous:1902:IICa

[4011] Anonymous. Illuminated Indian club. Scientific American, 86(21):369, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n21/pdf/scientificamerican05241902-369b.pdf.

Anonymous:1902:BPWn

[4012] Anonymous. Business and personal wants. Scientific American, 86(21):370, May 24, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n21/pdf/scientificamerican05241902-
370a.pdf.

Anonymous:1902:RPIo

[4013] Anonymous. Recently patented inventions. Scientific American, 86(22):386, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
386.pdf.

Anonymous:1902:IIo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-386c.pdf.

Anonymous:1902:BNBg

[4015] Anonymous. Books: New books, etc. Scientific American, 86(22):388, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-388.pdf.

Anonymous:1902:PNPj

[4016] Anonymous. Patent notes: Patents. Scientific American, 86(22):389, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-389.pdf.

Anonymous:1902:NQo

[4017] Anonymous. Notes and queries. Scientific American, 86(22):389, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-389a.pdf.

Anonymous:1902:FMC

[4018] Anonymous. Fireproof mail cars. Scientific American, 86(22):376, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-376.pdf.

Anonymous:1902:EUW

[4019] Anonymous. Early use of the water-tube locomotive boiler. Scientific American, 86(22):376, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
376a.pdf.

Anonymous:1902:DPI

[4020] Anonymous. Dangers of peddled ice cream. Scientific American, 86(22):376, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
376b.pdf.

Anonymous:1902:EOF

[4021] Anonymous. The expansion of our foreign commerce. Scientific American, 86(22):376, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
376c.pdf.

Anonymous:1902:IDP

[4022] Anonymous. The imperfect definition of the photographic image on dry plates. Scientific American, 86(22):376–377, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-376d.pdf.

Russell:1902:HJa

[4023] Henry Norris Russell. The heavens in June. Scientific American, 86(22):377, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
377.pdf.

Adams:1902:IGM

[4024] Alton D. Adams. Inspection of gas meters. Scientific American, 86(22):377, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
377a.pdf.

Anonymous:1902:PMO

[4025] Anonymous. Progressive-motion ore-conveyer. Scientific American, 86(22):378, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
378.pdf.

Byrn:1902:URS

[4026] Edward W. Byrn. The unveiling of the Rochambeau Statue. Scientific American, 86(22):378–379, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
378a.pdf.

Anonymous:1902:HMB

[4027] Anonymous. Hygienic mechanical bread-making. Scientific American, 86(22):379–380, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
379.pdf.

Anonymous:1902:SCE

[4028] Anonymous. Sinking the caissons for East River Bridge No. 3. Scientific American, 86(22):381, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
381.pdf.

Anonymous:1902:Pb

[4029] Anonymous. Pompeii. Scientific American, 86(22):382, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-382.pdf.

Frederick:1902:CP

[4030] M. C. Frederick. Cynips Polycera. Scientific American, 86(22):382–383, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-382a.pdf.

Fawcett:1902:RC

[4031] Waldon Fawcett. Railroads in Cuba. Scientific American, 86 (22):383–384, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
383.pdf.

Anonymous:1902:NTRa

[4032] Anonymous. The Nice–Turbie Race. Scientific American, 86(22):384, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-384.pdf.

Anonymous:1902:OCS

[4033] Anonymous. An oscillating compound steam engine of novel design. Scientific American, 86(22):384–385, May 31, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n22/pdf/scientificamerican05311902-384a.pdf.

Anonymous:1902:WTM

[4034] Anonymous. Wave and tide motor. Scientific American, 86(22):385, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-385.pdf.

Anonymous:1902:LH

[4035] Anonymous. Lamp-hanger. Scientific American, 86(22):385, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-385a.pdf.

Anonymous:1902:FCP

[4036] Anonymous. Friction clutch pulley. Scientific American, 86(22):385, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-385b.pdf.

Anonymous:1902:NB

[4037] Anonymous. The nation’s baskets. Scientific American, 86(22):386, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n22/pdf/scientificamerican05311902-386a.pdf.

Anonymous:1902:BPWo

[4038] Anonymous. Business and personal wants. Scientific American, 86(22):386, May 31, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n22/pdf/scientificamerican05311902-
386b.pdf.

Anonymous:1902:Ck

[4039] Anonymous. Correspondence. Scientific American, 86(23):397, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n23/pdf/scientificamerican06071902-397b.pdf.

Anonymous:1902:RPIp (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
402.pdf.

Anonymous:1902:IIp

[4041] Anonymous. Index of inventions. Scientific American, 86(23):402, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n23/pdf/scientificamerican06071902-402b.pdf.

Anonymous:1902:BNBh

[4042] Anonymous. Books: New books, etc. Scientific American, 86(23):406, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n23/pdf/scientificamerican06071902-406.pdf.

Anonymous:1902:PNPk

[4043] Anonymous. Patent notes: Patents. Scientific American, 86(23):407, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n23/pdf/scientificamerican06071902-407.pdf.

Anonymous:1902:NQp

[4044] Anonymous. Notes and queries. Scientific American, 86(23):407, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n23/pdf/scientificamerican06071902-407a.pdf.

Anonymous:1902:SDI

[4045] Anonymous. Seismic disturbances and the Isthmian Canal. Scientific American, 86(23):392, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
392.pdf.

Anonymous:1902:SWP

[4046] Anonymous. A scientific witness of Pelée’s eruption. Scientific American, 86(23):392, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
392a.pdf.

Anonymous:1902:SVSb

[4047] Anonymous. Sensible view of the steamship combine. Scientific American, 86(23):392, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
392b.pdf.

Anonymous:1902:HNB

[4048] Anonymous. The House Naval Bill. Scientific American, 86(23):392, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n23/pdf/scientificamerican06071902-392c.pdf.

Anonymous:1902:TVS

[4049] Anonymous. Tubes versus shallow-tunnel railroads. Scientific American, 86(23):392, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
392d.pdf.

Holder:1902:SAE

[4050] Charles F. Holder. Some active and extinct volcanoes. Scientific American, 86(23):393, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
393.pdf.

Walsh:1902:VM

[4051] George E. Walsh. Valuable minerals. Scientific American, 86(23):393, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n23/pdf/scientificamerican06071902-393a.pdf.

Anonymous:1902:TBN

[4052] Anonymous. Tractor-barges on the Nivernais Canal. Scientific American, 86(23):394, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
394.pdf.

Anonymous:1902:SHG

[4053] Anonymous. Setting the heavy granite columns on the hall of records. Scientific American, 86(23):395, June 7, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n23/pdf/scientificamerican06071902-395.pdf.

Anonymous:1902:SACa

[4054] Anonymous. The Severo airship catastrophe. Scientific American, 86(23):395–396, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
395a.pdf.

Anonymous:1902:BALa

[4055] Anonymous. Building of American locomotives — I. Scientific American, 86(23):396–397, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
396.pdf.

Anonymous:1902:VCV

[4056] Anonymous. Vent-closing valve for pumps. Scientific American, 86(23):397, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
397.pdf.

Anonymous:1902:FIB

[4057] Anonymous. Folding ice-box. Scientific American, 86(23):397, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n23/pdf/scientificamerican06071902-397a.pdf.

Anonymous:1902:HME

[4058] Anonymous. Hose made entirely of metal. Scientific American, 86(23):398, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
398.pdf.

Anonymous:1902:CBZ

[4059] Anonymous. Curious bridge at Zanesville, Ohio. Scientific American, 86(23):398, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
398a.pdf.

Willey:1902:SDC

[4060] Day Allen Willey. Self-dumping cars in railroad construction. Scientific American, 86(23):399–400, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
399.pdf.

Anonymous:1902:NFL

[4061] Anonymous. New form of Lippmann electrometer. Scientific American, 86(23):400, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
400.pdf.

Anonymous:1902:ERA

[4062] Anonymous. Earthquake recorders in America. Scientific American, 86(23):400, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
400a.pdf.

Holder:1902:UPG

[4063] Charles F. Holder. The utilization of Pampas grass. Scientific American, 86(23):400–401, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
400b.pdf.

Anonymous:1902:ABH

[4064] Anonymous. An antique bronze head of the Roman Emperor Tiberius. Scientific American, 86(23):401, June 7, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n23/pdf/scientificamerican06071902-401.pdf.

Brush:1902:AIR

[4065] Edward Hale Brush. An ancient Iroquois rite. Scientific American, 86(23):401, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
401a.pdf.

Anonymous:1902:BPWp

[4066] Anonymous. Business and personal wants. Scientific American, 86(23):402, June 7, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n23/pdf/scientificamerican06071902-
402a.pdf.

Anonymous:1902:Cl

[4067] Anonymous. Correspondence. Scientific American, 86(24):414–415, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-414a.pdf.

Anonymous:1902:LNg

[4068] Anonymous. Legal notes. Scientific American, 86(24):421, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-421.pdf.

Anonymous:1902:RPIq

[4069] Anonymous. Recently patented inventions. Scientific American, 86(24):422, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
422.pdf.

Anonymous:1902:IIq

[4070] Anonymous. Index of inventions. Scientific American, 86(24):422, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-422b.pdf.

Anonymous:1902:BNBi

[4071] Anonymous. Books: New books. Scientific American, 86(24):424, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-424.pdf.

Anonymous:1902:PNPl

[4072] Anonymous. Patent notes: Patents. Scientific American, 86(24):425, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-425.pdf.

Anonymous:1902:NQq

[4073] Anonymous. Notes and queries. Scientific American, 86(24):425, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-425a.pdf.

Anonymous:1902:NBC

[4074] Anonymous. Naval Bureau Chiefs and the submarine boat. Scientific American, 86(24):410, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
410.pdf.

Anonymous:1902:DIM

[4075] Anonymous. A daring investigation of Mont Pelée. Scientific American, 86(24):410, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
410a.pdf.

Anonymous:1902:CPAb

[4076] Anonymous. The commercial prospects of Africa. Scientific American, 86(24):410–411, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
410b.pdf.

Benjamin:1902:CIP

[4077] Marcus Benjamin. Chemical indexing by the Patent Office. Scientific American, 86(24):411, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
411.pdf.

Anonymous:1902:SSF

[4078] Anonymous. Sinking a shaft by freezing. Scientific American, 86(24):411, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
411a.pdf.

Clapp:1902:PWG

[4079] A. W. Clapp. Preliminary work on the $3,000,000 Government contract in the Philippines. Scientific American, 86(24):411, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-411b.pdf.

Anonymous:1902:MLR

[4080] Anonymous. Moving a large railway bridge. Scientific American, 86(24):412, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
412.pdf.

Anonymous:1902:ERBb

[4081] Anonymous. The electric road of Berthoud–Thoune, Switzerland. Scientific American, 86(24):413–414, June 14, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n24/pdf/scientificamerican06141902-413.pdf.

Brooks:1902:BC

[4082] William R. Brooks. The Brooks Comet of 1902. Scientific American, 86(24):414, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
414.pdf.

Perkins:1902:MUE

[4083] Frank C. Perkins. The modern use of electricity in printing. Scientific American, 86(24):415, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
415.pdf.

Anonymous:1902:EBD

[4084] Anonymous. Early British double-turreted monitors. Scientific American, 86(24):416, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
416.pdf.

Mclure:1902:BC

[4085] W. Frank M’clure. Bee-culture. Scientific American, 86(24):416–418, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-416a.pdf.

Anonymous:1902:RFH

[4086] Anonymous. Railroad freight handling in New York City. Scientific American, 86(24):418–419, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
418.pdf.

Anonymous:1902:BER

[4087] Anonymous. The Baker electric racing automobile. Scientific American, 86(24):419, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
419.pdf.

Anonymous:1902:PDg

[4088] Anonymous. Patent department. Scientific American, 86(24):420, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-420.pdf.

Anonymous:1902:Pc

[4089] Anonymous. Planter. Scientific American, 86(24):420, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n24/pdf/scientificamerican06141902-420a.pdf.

Anonymous:1902:UTS

[4090] Anonymous. Utilizing the tube space in bicycle frames. Scientific American, 86(24):420–421, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
420b.pdf.

Anonymous:1902:BPWq

[4091] Anonymous. Business and personal wants. Scientific American, 86(24):422, June 14, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n24/pdf/scientificamerican06141902-
422a.pdf.

Anonymous:1902:Cm

[4092] Anonymous. Correspondence. Scientific American, 86(25):433, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n25/pdf/scientificamerican06211902-433a.pdf.

Anonymous:1902:RPIr

[4093] Anonymous. Recently patented inventions. Scientific American, 86(25):437, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n25/pdf/scientificamerican06211902-
437.pdf.

Anonymous:1902:IIr

[4094] Anonymous. Index of inventions. Scientific American, 86(25):437–440, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n25/pdf/scientificamerican06211902-437b.pdf.

Anonymous:1902:BNBj

[4095] Anonymous. Books: New books. Scientific American, 86(25):440–441, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n25/pdf/scientificamerican06211902-440.pdf.

Anonymous:1902:NQr

[4096] Anonymous. Notes and queries. Scientific American, 86(25):441, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n25/pdf/scientificamerican06211902-441.pdf.

Anonymous:1902:ARAa

[4097] Anonymous. Analysis of the recent automobile endurance test and speed trials. Scientific American, 86(25):428, June 21, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n25/pdf/scientificamerican06211902-428.pdf.

Anonymous:1902:PAS

[4098] Anonymous. Practical application of scientific education in Germany. Scientific American, 86(25):428, June 21, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n25/pdf/scientificamerican06211902-428a.pdf.

Anonymous:1902:RMM

[4099] Anonymous. Richard L. Maddox, M.D. Scientific American, 86 (25):428–429, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n25/pdf/scientificamerican06211902-
428b.pdf.

Irwin:1902:DSS

[4100] William Gilbert Irwin. The development of the Sault Ste. Marie Canal. Scientific American, 86(25):429, June 21, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n25/pdf/scientificamerican06211902-429.pdf.

Willet:1902:MFT

[4101] Day Allen Willet. Making fifty-ton anchor chains. Scientific American, 86(25):429, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n25/pdf/scientificamerican06211902-
429a.pdf.

Fawcett:1902:SCA

[4102] Waldon Fawcett. Silkworm culture in America. Scientific American, 86(25):430–431, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n25/pdf/scientificamerican06211902-
430.pdf.

Beard:1902:RAP

[4103] J. Carter Beard. Reflex action of plants as compared with the instincts of insects. Scientific American, 86(25):431–432, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n25/pdf/scientificamerican06211902-431.pdf.

Anonymous:1902:MPa

[4104] Anonymous. Mixing pump. Scientific American, 86(25):432, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n25/pdf/scientificamerican06211902-432.pdf.

Anonymous:1902:RSPa

[4105] Anonymous. The ruins of St. Pierre. Scientific American, 86(25):433, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n25/pdf/scientificamerican06211902-433.pdf.

Jones:1902:ETD

[4106] George J. Jones. Extraordinary trench digging. Scientific American, 86(25):434, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n25/pdf/scientificamerican06211902-
434.pdf.

Anonymous:1902:TTE

[4107] Anonymous. A 20-ton traveling electric crane. Scientific American, 86(25):435, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n25/pdf/scientificamerican06211902-
435.pdf.

Anonymous:1902:EBT

[4108] Anonymous. The Elliott book-typewriter. Scientific American, 86(25):436, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n25/pdf/scientificamerican06211902-
436.pdf.

Anonymous:1902:NFT

[4109] Anonymous. A new form of trunk. Scientific American, 86(25):436, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n25/pdf/scientificamerican06211902-436a.pdf.

Anonymous:1902:BPWr

[4110] Anonymous. Business and personal wants. Scientific American, 86(25):437, June 21, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n25/pdf/scientificamerican06211902-
437a.pdf.

Russell:1902:HJb

[4111] Henry Norris Russell. The heavens in July. Scientific American, 86(26):445, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
445.pdf.

Anonymous:1902:Cn

[4112] Anonymous. Correspondence. Scientific American, 86(26):449, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-449.pdf.

Anonymous:1902:RPIs

[4113] Anonymous. Recently patented inventions. Scientific American, 86(26):454, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
454.pdf.

Anonymous:1902:IIs

[4114] Anonymous. Index of inventions. Scientific American, 86(26):455, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-455.pdf.

Anonymous:1902:Ia

[4115] Anonymous. Index. Scientific American, 86(26):457–458, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-457.pdf.

Anonymous:1902:NQs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-459.pdf.

Anonymous:1902:PNPm

[4117] Anonymous. Patent notes: Patents. Scientific American, 86(26):459, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-459a.pdf.

Anonymous:1902:SPCb

[4118] Anonymous. The Senate Panama Canal Bill. Scientific American, 86(26):444, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
444.pdf.

Anonymous:1902:TPT

[4119] Anonymous. Three phase traction in America. Scientific American, 86(26):444, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
444a.pdf.

Anonymous:1902:TRR

[4120] Anonymous. Two remarkable railway runs from New York to Chicago. Scientific American, 86(26):444, June 28, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n26/pdf/scientificamerican06281902-444b.pdf.

Anonymous:1902:GI

[4121] Anonymous. Government irrigation. Scientific American, 86(26): 444–445, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
444c.pdf.

King:1902:FSF

[4122] C. D. King. Forced and shrink fits. Scientific American, 86(26):445, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-445a.pdf.

Anonymous:1902:EN

[4123] Anonymous. Engineering notes. Scientific American, 86(26):445, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-445b.pdf.

Anonymous:1902:CET

[4124] Anonymous. Chevallier electric target. Scientific American, 86(26):446, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-446.pdf.

Anonymous:1902:TLO

[4125] Anonymous. The tractor in lumbering operations. Scientific American, 86(26):447–448, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
447.pdf.

Anonymous:1902:FFb

[4126] Anonymous. Fiji and the Fijians. Scientific American, 86(26):448–449, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-448.pdf.

Willey:1902:PIC

[4127] Day Allen Willey. The packing industry of Chicago. Scientific American, 86(26):450–451, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
450.pdf.

Collins:1902:GCA

[4128] A. Frederick Collins. The Grisson continuous-alternating current transformers. Scientific American, 86(26):452, June 28, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n26/pdf/scientificamerican06281902-452.pdf.

Anonymous:1902:NTK

[4129] Anonymous. New typewriter keyboard. Scientific American, 86 (26):453, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
453.pdf.

Anonymous:1902:RRE

[4130] Anonymous. A recent rear-end collision on the New York Central Railroad. Scientific American, 86(26):453, June 28, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n26/pdf/scientificamerican06281902-453a.pdf.

Anonymous:1902:Co

[4131] Anonymous. Crusher. Scientific American, 86(26):453, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n26/pdf/scientificamerican06281902-453b.pdf.

Anonymous:1902:BPWs

[4132] Anonymous. Business and personal wants. Scientific American, 86(26):454, June 28, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n26/pdf/scientificamerican06281902-
454a.pdf.

Anonymous:1902:RPIt

[4133] Anonymous. Recently patented inventions. Scientific American, 86(3):44, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
44.pdf.

Anonymous:1902:IIt

[4134] Anonymous. Index of inventions. Scientific American, 86(3):45, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n3/pdf/scientificamerican01181902-45.pdf.

Anonymous:1902:NQt

[4135] Anonymous. Notes and queries. Scientific American, 86(3):47, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n3/pdf/scientificamerican01181902-47.pdf.

Anonymous:1902:ICH

[4136] Anonymous. The Isthmian Canal — historical. Scientific American, 86(3):34, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
34.pdf.

Anonymous:1902:CES

[4137] Anonymous. The canal from the engineering standpoint. Scientific American, 86(3):34, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
34a.pdf.

Anonymous:1902:CSC

[4138] Anonymous. The canal from the standpoint of common sense. Scientific American, 86(3):34, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
34b.pdf.

Anonymous:1902:NYC

[4139] Anonymous. The New York Central Tunnel tragedy. Scientific American, 86(3):34–35, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
34c.pdf.

Anonymous:1902:PGM

[4140] Anonymous. The perfecting of the gasoline motor. Scientific American, 86(3):35, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
35.pdf.

Anonymous:1902:EOP

[4141] Anonymous. Electro optical phenomena. Scientific American, 86 (3):35, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
35a.pdf.

Kempton:1902:DSO

[4142] Dwight Kempton. The drilling of submarine oil wells as performed at Summerland, Cal. Scientific American, 86(3):36, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n3/pdf/scientificamerican01181902-36.pdf.

Anonymous:1902:NHT

[4143] Anonymous. A novel harrow-tooth. Scientific American, 86(3):37, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n3/pdf/scientificamerican01181902-37.pdf.

Anonymous:1902:AAP

[4144] Anonymous. Alcohol automobiles at the Paris Alcohol Exhibition. Scientific American, 86(3):37, January 18, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n3/pdf/scientificamerican01181902-37a.pdf.

Anonymous:1902:WEC

[4145] Anonymous. The Wydts electro-catalytic sparking plug. Scientific American, 86(3):37–38, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
37b.pdf.

Anonymous:1902:LDO

[4146] Anonymous. A lever-device for oil-rods. Scientific American, 86 (3):38, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
38.pdf.

Thompson:1902:YK

[4147] James Harold Thompson. The year of 1901 in the Klondike. Scientific American, 86(3):38, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
38a.pdf.

Anonymous:1902:CPN

[4148] Anonymous. A comparison of the Panama and Nicaragua Canal routes. Scientific American, 86(3):39, January 18, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n3/pdf/scientificamerican01181902-39.pdf.

Anonymous:1902:SEW

[4149] Anonymous. Some experiments with wireless telegraphy. Scientific American, 86(3):40–41, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
40.pdf.

Perkins:1902:SFP

[4150] Frank C. Perkins. Shawenegan Falls Power Plant. Scientific American, 86(3):41–42, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
41.pdf.

Lucas:1902:CEM

[4151] Frederic A. Lucas. Constructing an extinct monster from fossil remains. Scientific American, 86(3):43, January 18, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n3/pdf/scientificamerican01181902-43.pdf.

Anonymous:1902:BPWt

[4152] Anonymous. Business and personal wants. Scientific American, 86(3):44–45, January 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n3/pdf/scientificamerican01181902-
44a.pdf.

Anonymous:1902:Cp

[4153] Anonymous. Correspondence. Scientific American, 86(4):54–55, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n4/pdf/scientificamerican01251902-54.pdf.

Anonymous:1902:LNh

[4154] Anonymous. Legal notes. Scientific American, 86(4):61, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n4/pdf/scientificamerican01251902-61.pdf.

Anonymous:1902:RPIu

[4155] Anonymous. Recently patented inventions. Scientific American, 86(4):62, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
62.pdf.

Anonymous:1902:IIu

[4156] Anonymous. Index of inventions. Scientific American, 86(4):62, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n4/pdf/scientificamerican01251902-62b.pdf.

Anonymous:1902:NQu

[4157] Anonymous. Notes and queries. Scientific American, 86(4):65, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n4/pdf/scientificamerican01251902-65.pdf.

Anonymous:1902:WNH

[4158] Anonymous. The wrecked new haven car in the tunnel accident. Scientific American, 86(4):50, January 25, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n4/pdf/scientificamerican01251902-50.pdf.

Anonymous:1902:EET

[4159] Anonymous. The electric elevated trains. Scientific American, 86(4):50, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
50a.pdf.

Anonymous:1902:TPU

[4160] Anonymous. Target practice in the United States Navy. Scientific American, 86(4):50, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
50b.pdf.

Anonymous:1902:NTRb

[4161] Anonymous. A new trade route to Persia. Scientific American, 86(4):50–51, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
50c.pdf.

Hovey:1902:AMG

[4162] Edmund O. Hovey. The Annual Meeting of the Geological Society of America. Scientific American, 86(4):51, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n4/pdf/scientificamerican01251902-51.pdf.

Anonymous:1902:SOM

[4163] Anonymous. A scarcity of officers in the merchant marine. Scientific American, 86(4):51, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
51a.pdf.

Anonymous:1902:SWT

[4164] Anonymous. Successful wireless telegraphy at sea. Scientific American, 86(4):51, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
51b.pdf.

Anonymous:1902:WGB

[4165] Anonymous. The Welsbach gasoline burner. Scientific American, 86(4):52, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
52.pdf.

Anonymous:1902:NSTb

[4166] Anonymous. Novel steam turbine and dynamo combination. Scientific American, 86(4):52, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
52a.pdf.

Dolezalek:1902:LSR

[4167] Dolezalek. The Langen Suspended Railway of Bar-Men-Vohwinkel-El-Berfeld. Scientific American, 86(4):53–54, January 25, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n4/pdf/scientificamerican01251902-53.pdf.

Mills:1902:TCO

[4168] E. T. Mills. Trades and crafts in the Old Spanish Missions. Scientific American, 86(4):56–57, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
56.pdf.

Anonymous:1902:FFP

[4169] Anonymous. Feeding a 26-foot python. Scientific American, 86 (4):57, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
57.pdf.

Anonymous:1902:PSNb

[4170] Anonymous. Proposed solution of the New York Central Terminal problem. Scientific American, 86(4):58, January 25, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n4/pdf/scientificamerican01251902-58.pdf.

Anonymous:1902:DPS

[4171] Anonymous. Device for preventing shipwrecks. Scientific American, 86(4):59, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
59.pdf.

Anonymous:1902:TRT

[4172] Anonymous. Tire-repairing tool. Scientific American, 86(4):59, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n4/pdf/scientificamerican01251902-59a.pdf.

Anonymous:1902:MVC

[4173] Anonymous. A mechanical vegetable-cutter. Scientific American, 86(4):59, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
59b.pdf.

Anonymous:1902:INL

[4174] Anonymous. An improved nut-lock. Scientific American, 86(4):59, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n4/pdf/scientificamerican01251902-59c.pdf.

Anonymous:1902:SNIa

[4175] Anonymous. Some new interesting inventions. Scientific American, 86(4):60, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
60.pdf.

Anonymous:1902:BPWu

[4176] Anonymous. Business and personal wants. Scientific American, 86(4):62, January 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n4/pdf/scientificamerican01251902-
62a.pdf.

Anonymous:1902:Cq

[4177] Anonymous. Correspondence. Scientific American, 86(5):72, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n5/pdf/scientificamerican02011902-72a.pdf.

Russell:1902:HF

[4178] Henry Norris Russell. The heavens in February. Scientific American, 86(5):73, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
73a.pdf.

Anonymous:1902:RPIv 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
77.pdf.

Anonymous:1902:IIv

[4180] Anonymous. Index of inventions. Scientific American, 86(5):78, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n5/pdf/scientificamerican02011902-78a.pdf.

Anonymous:1902:BNBk

[4181] Anonymous. Books: New books, etc. Scientific American, 86 (5):80, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
80.pdf.

Anonymous:1902:PNPn

[4182] Anonymous. Patent notes: Patents. Scientific American, 86 (5):81, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
81.pdf.

Anonymous:1902:NQv

[4183] Anonymous. Notes and queries. Scientific American, 86(5):81, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n5/pdf/scientificamerican02011902-81a.pdf.

Anonymous:1902:RPR

[4184] Anonymous. Recommendation of the Panama route. Scientific American, 86(5):68, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
68.pdf.

Anonymous:1902:FEJ

[4185] Anonymous. The faulty embankments of the Jerome Park Reservoir. Scientific American, 86(5):68, February 1, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n5/pdf/scientificamerican02011902-68a.pdf.

Anonymous:1902:LDT

[4186] Anonymous. Long distance transmission. Scientific American, 86(5):68, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
68b.pdf.

Anonymous:1902:AMT

[4187] Anonymous. Advance of the marine turbine. Scientific American, 86(5):68, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
68c.pdf.

Anonymous:1902:SSB

[4188] Anonymous. The Ship Subsidy Bill. Scientific American, 86 (5):68–69, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
68d.pdf.

Anonymous:1902:SDE

[4189] Anonymous. The supposed dangers of electric traction. Scientific American, 86(5):69, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
69.pdf.

Anonymous:1902:THT

[4190] Anonymous. Two hundred thousand dollars in prizes for airships. Scientific American, 86(5):69, February 1, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n5/pdf/scientificamerican02011902-69a.pdf.

Anonymous:1902:VNY

[4191] Anonymous. Verdict in New York Central Tunnel wreck inquest. Scientific American, 86(5):69, February 1, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n5/pdf/scientificamerican02011902-69b.pdf.

Anonymous:1902:FAW

[4192] Anonymous. A fire alarm whistle blown by gas. Scientific American, 86(5):70, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
70.pdf.

Anonymous:1902:LJE

[4193] Anonymous. Lifting jack for electric cars. Scientific American, 86(5):70, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
70a.pdf.

Anonymous:1902:GCY

[4194] Anonymous. The Great Clearing Yard at Chicago. Scientific American, 86(5):70–71, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
70b.pdf.

Anonymous:1902:SFP

[4195] Anonymous. A self-feathering paddle-wheel. Scientific American, 86(5):71, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
71.pdf.

Willey:1902:SMM

[4196] Day Allen Willey. Scientific methods of moving trees. Scientific American, 86(5):71–72, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
71a.pdf.

Anonymous:1902:PRTa

[4197] Anonymous. The Pennsylvania Railroad Tunnel Bridge. Scientific American, 86(5):72, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
72.pdf.

Moore:1902:WIC

[4198] Frederick Moore. Work of the Isthmian Canal engineers. Scientific American, 86(5):73, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
73.pdf.

Holder:1902:STF

[4199] Charles F. Holder. Some trees and forests of California. Scientific American, 86(5):74, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
74.pdf.

Shepstone:1902:MSW

[4200] H. J. Shepstone. The Marconi School of Wireless Telegraphy. Scientific American, 86(5):74–75, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
74a.pdf.

Anonymous:1902:ESMa

[4201] Anonymous. The experimental study of the motion of fluids. Scientific American, 86(5):75–76, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
75.pdf.

Anonymous:1902:MBS

[4202] Anonymous. Modeling in beach sand. Scientific American, 86 (5):76, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
76.pdf.

Anonymous:1902:BPWv

[4203] Anonymous. Business and personal wants. Scientific American, 86(5):78, February 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n5/pdf/scientificamerican02011902-
78.pdf.

Anonymous:1902:Cr

[4204] Anonymous. Correspondence. Scientific American, 86(6):88–89, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n6/pdf/scientificamerican02081902-88.pdf.

Anonymous:1902:LNi

[4205] Anonymous. Legal notes. Scientific American, 86(6):94, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n6/pdf/scientificamerican02081902-94.pdf.

Anonymous:1902:RPIw

[4206] Anonymous. Recently patented inventions. Scientific American, 86(6):95, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
95.pdf.

Anonymous:1902:BNBl

[4207] Anonymous. Books: New books, etc. Scientific American, 86 (6):95, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
95a.pdf.

Anonymous:1902:IIw

[4208] Anonymous. Index of inventions. Scientific American, 86(6):96, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n6/pdf/scientificamerican02081902-96a.pdf.

Anonymous:1902:NQw

[4209] Anonymous. Notes and queries. Scientific American, 86(6):97, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n6/pdf/scientificamerican02081902-97.pdf.

Anonymous:1902:PNPo

[4210] Anonymous. Patent notes: Patents. Scientific American, 86 (6):99, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
99.pdf.

Anonymous:1902:LMH

[4211] Anonymous. Lessons of the Murray Hill explosion. Scientific American, 86(6):84, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
84.pdf.

Anonymous:1902:FTB

[4212] Anonymous. Frailty of torpedo boat destroyers. Scientific American, 86(6):84, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
84a.pdf.

Anonymous:1902:STb

[4213] Anonymous. Subterranean temperature. Scientific American, 86 (6):84, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
84b.pdf.

Anonymous:1902:ERBc

[4214] Anonymous. The East River Bridge from the standpoint of art. Scientific American, 86(6):84, February 8, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n6/pdf/scientificamerican02081902-84c.pdf.

Anonymous:1902:RWD

[4215] Anonymous. Railway workmen and disbursements. Scientific American, 86(6):84–85, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
84d.pdf.

Adams:1902:PVN

[4216] Alton D. Adams. Practical value of Nernst lamps. Scientific American, 86(6):85, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
85.pdf.

Anonymous:1902:SNk

[4217] Anonymous. Science notes. Scientific American, 86(6):85, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n6/pdf/scientificamerican02081902-85a.pdf.

Anonymous:1902:CAT

[4218] Anonymous. Construction of the aspen tunnel. Scientific American, 86(6):86, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
86.pdf.

Anonymous:1902:MSC

[4219] Anonymous. The manufacture of submarine cables. Scientific American, 86(6):87–88, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
87.pdf.

Anonymous:1902:EDR

[4220] Anonymous. Explosion of dynamite at the rapid transit subway. Scientific American, 86(6):90–91, February 8, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n6/pdf/scientificamerican02081902-90.pdf.

Perkins:1902:HSG

[4221] Frank C. Perkins. High-speed German railway at Zossen. Scientific American, 86(6):91–92, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
91.pdf.

Anonymous:1902:PPIa

[4222] Anonymous. Profit in patented inventions. Scientific American, 86(6):92–93, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
92.pdf.

Anonymous:1902:INI

[4223] Anonymous. Interesting new inventions. Scientific American, 86 (6):93–94, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
93.pdf.

Anonymous:1902:BPWw

[4224] Anonymous. Business and personal wants. Scientific American, 86(6):96, February 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n6/pdf/scientificamerican02081902-
96.pdf.

Anonymous:1902:RPIx

[4225] Anonymous. Recently patented inventions. Scientific American, 86(7):112, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
112.pdf.

Anonymous:1902:BNBm

[4226] Anonymous. Books: New books, etc. Scientific American, 86 (7):112, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
112a.pdf.

Anonymous:1902:IIx

[4227] Anonymous. Index of inventions. Scientific American, 86(7):112, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n7/pdf/scientificamerican02151902-112c.pdf.

Anonymous:1902:PNPp

[4228] Anonymous. Patent notes: Patents. Scientific American, 86 (7):115, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
115.pdf.

Anonymous:1902:NQx

[4229] Anonymous. Notes and queries. Scientific American, 86(7):115, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n7/pdf/scientificamerican02151902-115a.pdf.

Anonymous:1902:PSP

[4230] Anonymous. Possibilities of a steel panic. Scientific American, 86(7):102, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
102.pdf.

Anonymous:1902:GET

[4231] Anonymous. Germany’s export trade in her hour of need. Scientific American, 86(7):102, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
102a.pdf.

Anonymous:1902:DDB

[4232] Anonymous. Double-decking the Brooklyn Bridge. Scientific American, 86(7):102, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
102b.pdf.

Anonymous:1902:TBF

[4233] Anonymous. The torpedo-boat fiasco. Scientific American, 86 (7):102, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
102c.pdf.

Anonymous:1902:ROC

[4234] Anonymous. Reclaiming our coast lands. Scientific American, 86 (7):102–103, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
102d.pdf.

Anonymous:1902:NPE

[4235] Anonymous. A new Polar Expedition. Scientific American, 86 (7):103, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
103.pdf.

Anonymous:1902:NSG

[4236] Anonymous. New salt of glucinium. Scientific American, 86 (7):103, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
103a.pdf.

Anonymous:1902:SNl

[4237] Anonymous. Science notes. Scientific American, 86(7):103, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n7/pdf/scientificamerican02151902-103b.pdf.

Fawcett:1902:MTC

[4238] Waldon Fawcett. Mail transportation in cities. Scientific American, 86(7):104, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
104.pdf.

Hopkins:1902:HME

[4239] George M. Hopkins. Home-made electric night lamp. Scientific American, 86(7):105, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
105.pdf.

Anonymous:1902:CHG

[4240] Anonymous. Curious horn growth. Scientific American, 86(7): 105, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
105a.pdf.

Anonymous:1902:PCL

[4241] Anonymous. Portable coal-loading machine. Scientific American, 86 (7):105–106, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
105b.pdf.

Simonsen:1902:HMC

[4242] Parker S. Simonsen. How to make a cheap hydrometer. Scientific American, 86(7):106, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
106.pdf.

Anonymous:1902:NYAb

[4243] Anonymous. The New York Assay Office. Scientific American, 86(7):107, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
107.pdf.

Anonymous:1902:IW

[4244] Anonymous. The improvement of Washington. Scientific American, 86(7):108–109, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
108.pdf.

Proctor:1902:YOT

[4245] Mary Proctor. The Yerkes Observatory two-foot reflector. Scientific American, 86(7):109–110, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
109.pdf.

Holder:1902:PR

[4246] Charles F. Holder. A pigeon ranch. Scientific American, 86 (7):111, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
111.pdf.

Anonymous:1902:BPWx

[4247] Anonymous. Business and personal wants. Scientific American, 86(7):112, February 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n7/pdf/scientificamerican02151902-
112b.pdf.

Anonymous:1902:RPIy

[4248] Anonymous. Recently patented inventions. Scientific American, 86(8):128, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
128.pdf.

Anonymous:1902:IIy

[4249] Anonymous. Index of inventions. Scientific American, 86(8):128, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n8/pdf/scientificamerican02221902-128b.pdf.

Anonymous:1902:BNBn

[4250] Anonymous. Books: New books, etc. Scientific American, 86 (8):130, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
130.pdf.

Anonymous:1902:PNPq

[4251] Anonymous. Patent notes: Patents. Scientific American, 86 (8):131, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
131.pdf.

Anonymous:1902:NQy

[4252] Anonymous. Notes and queries. Scientific American, 86(8):131, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n8/pdf/scientificamerican02221902-131a.pdf.

Anonymous:1902:ROH

[4253] Anonymous. Revival of the Old Hudson River Tunnel scheme. Scientific American, 86(8):118, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
118.pdf.

Anonymous:1902:DSB

[4254] Anonymous. The Darien or San Blas Tunnel Canal route. Scientific American, 86(8):118, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
118a.pdf.

Anonymous:1902:NDD

[4255] Anonymous. Naval development during the next decade. Scientific American, 86(8):118–119, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
118b.pdf.

Anonymous:1902:SBI

[4256] Anonymous. Steam boiler inspection. Scientific American, 86 (8):119, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
119.pdf.

Anonymous:1902:OFA

[4257] Anonymous. Our forthcoming automobile and outing number. Scientific American, 86(8):119, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
119a.pdf.

Anonymous:1902:SNm

[4258] Anonymous. Science notes. Scientific American, 86(8):119, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n8/pdf/scientificamerican02221902-119b.pdf.

Anonymous:1902:VSG

[4259] Anonymous. A variable speed gear for pumps. Scientific American, 86(8):120, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
120.pdf.

Fawcett:1902:PEH

[4260] Waldon Fawcett. Photographing by electric headlight. Scientific American, 86(8):120, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
120a.pdf.

Anonymous:1902:SME

[4261] Anonymous. A simple method of etching pictures. Scientific American, 86(8):121, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
121.pdf.

Anonymous:1902:SDAb

[4262] Anonymous. The Santos-Dumont Airship No. 7. Scientific American, 86(8):121, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
121a.pdf.

Anonymous:1902:NRBa

[4263] Anonymous. The new Redheugh Bridge at Newcastle-On-Tyne. Scientific American, 86(8):122, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
122.pdf.

Anonymous:1902:IR

[4264] Anonymous. Improved railjoint. Scientific American, 86(8):122– 123, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
122a.pdf.

Brown:1902:WRA

[4265] Enos Brown. Winter railroading in Alaska. Scientific American, 86 (8):124–125, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
124.pdf.

Anonymous:1902:SAE

[4266] Anonymous. Some aeronautical experiments. Scientific American, 86(8):125–126, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
125.pdf.

Anonymous:1902:DSC

[4267] Anonymous. The Duke of Saxe-Coburg-Gotha’s collection of Nefs. Scientific American, 86(8):126–127, February 22, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n8/pdf/scientificamerican02221902-126.pdf.

Gratacap:1902:SCC

[4268] L. P. Gratacap. The sands of Cape Cod. Scientific American, 86(8):127, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
127.pdf.

Anonymous:1902:BPWy

[4269] Anonymous. Business and personal wants. Scientific American, 86(8):128, February 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n8/pdf/scientificamerican02221902-
128a.pdf.

Anonymous:1902:IIz

[4270] Anonymous. Index of inventions. Scientific American, 86(9):156, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-156.pdf.

Anonymous:1902:BNBo

[4271] Anonymous. Books: New books, etc. Scientific American, 86(9):158, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-158.pdf.

Anonymous:1902:NQz

[4272] Anonymous. Notes and queries. Scientific American, 86(9):159, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-159.pdf.

Anonymous:1902:PNPr

[4273] Anonymous. Patent notes: Patents. Scientific American, 86(9):161, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-161.pdf.

Anonymous:1902:RPIz

[4274] Anonymous. Recently patented inventions. Scientific American, 86(9):161, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
161a.pdf.

Anonymous:1902:Ib

[4275] Anonymous. Introductory. Scientific American, 86(9):134, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-134.pdf.

Anonymous:1902:ASc

[4276] Anonymous. The automobile on the stage. Scientific American, 86(9):134, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
134a.pdf.

Anonymous:1902:SFC

[4277] Anonymous. Special features at the Chicago Show. Scientific American, 86(9):134, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
134b.pdf.

Wisby:1902:PAT

[4278] Hrolf Wisby. A practical automobile touring outfit. Scientific American, 86(9):134–135, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
134c.pdf.

Anonymous:1902:ANb

[4279] Anonymous. Automobile notes. Scientific American, 86(9):135, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-135.pdf.

Anonymous:1902:SCa

[4280] Anonymous. Steam carriages. Scientific American, 86(9):136–137, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-136.pdf.

Anonymous:1902:SSS

[4281] Anonymous. The Stearns steam station wagon. Scientific American, 86(9):137, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
137.pdf.

Anonymous:1902:LSC

[4282] Anonymous. The lane steam carriage. Scientific American, 86(9):138, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-138.pdf.

Anonymous:1902:OSA

[4283] Anonymous. The overman steam automobile. Scientific American, 86(9):139, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
139.pdf.

Anonymous:1902:GQS

[4284] Anonymous. The Grout Queen Stanhope. Scientific American, 86(9):139, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
139a.pdf.

Anonymous:1902:GA

[4285] Anonymous. Gasoline automobiles. Scientific American, 86(9):140, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-140.pdf.

Anonymous:1902:KTF

[4286] Anonymous. The Knox two and four-passenger carriage. Scientific American, 86(9):140–141, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
140a.pdf.

Anonymous:1902:EYM

[4287] Anonymous. The Emperor’s yacht “Meteor III”. Scientific American, 86(9):141, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
141.pdf.

Anonymous:1902:GAM

[4288] Anonymous. Gasoline automobiles — 1902 models. Scientific American, 86(9):141–143, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
141a.pdf.

Anonymous:1902:HST

[4289] Anonymous. High speed twin-screw yacht “Vixen”. Scientific American, 86(9):143, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
143.pdf.

Anonymous:1902:MA

[4290] Anonymous. The marine automobile. Scientific American, 86 (9):143–144, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
143a.pdf.

Anonymous:1902:CW

[4291] Anonymous. The cycle whirl. Scientific American, 86(9):144, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-144.pdf.

Anonymous:1902:ASd

[4292] Anonymous. An automobile sled. Scientific American, 86(9):145, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-145.pdf.

Anonymous:1902:BDC

[4293] Anonymous. A bicycle-driven catamaran. Scientific American, 86 (9):145, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
145a.pdf.

Anonymous:1902:SB

[4294] Anonymous. A snow bicycle. Scientific American, 86(9):145, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-145b.pdf.

Anonymous:1902:PGT

[4295] Anonymous. The Packard gasoline touring car. Scientific American, 86(9):146–147, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
146.pdf.

Anonymous:1902:DGC

[4296] Anonymous. The Duryea gasoline carriage. Scientific American, 86(9):147, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
147.pdf.

Anonymous:1902:EV

[4297] Anonymous. Electric vehicles. Scientific American, 86(9):148, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-148.pdf.

Anonymous:1902:HRa

[4298] Anonymous. The houseboat “Rancocas”. Scientific American, 86 (9):148, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
148a.pdf.

Anonymous:1902:TSH

[4299] Anonymous. Twin-screw houseboat designed for the late Pierre Lorillard. Scientific American, 86(9):149, March 1, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v86/
n9/pdf/scientificamerican03011902-149.pdf.

Anonymous:1902:AT

[4300] Anonymous. Automobile tires. Scientific American, 86(9):150, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-150.pdf.

Anonymous:1902:ARAb

[4301] Anonymous. Automobile roller axle bearings. Scientific American, 86(9):152, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
152.pdf.

Anonymous:1902:ANc

[4302] Anonymous. Automobile novelties. Scientific American, 86(9):152–155, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v86/n9/pdf/scientificamerican03011902-152a.pdf.

Anonymous:1902:BPWz

[4303] Anonymous. Business and personal wants. Scientific American, 86(9):155, March 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v86/n9/pdf/scientificamerican03011902-
155.pdf.

Anonymous:1902:Cs

[4304] Anonymous. Correspondence. Scientific American, 87(1):6–7, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-6a.pdf.

Anonymous:1902:RPIba

[4305] Anonymous. Recently patented inventions. Scientific American, 87(1): 12, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-12.pdf.

Anonymous:1902:IIba

[4306] Anonymous. Index of inventions. Scientific American, 87(1):12, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-12b.pdf.

Anonymous:1902:PNPs

[4307] Anonymous. Patent notes: Patents. Scientific American, 87(1):15, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-15.pdf.

Anonymous:1902:NQba

[4308] Anonymous. Notes and queries. Scientific American, 87(1):15, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-15a.pdf.

Anonymous:1902:PCP

[4309] Anonymous. Prospective completion of the Panama Interoceanic Canal. Scientific American, 87(1):2, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
2.pdf.

Anonymous:1902:POA

[4310] Anonymous. A Prussian opinion of American railways. Scientific American, 87(1):2, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
2a.pdf.

Anonymous:1902:NLS

[4311] Anonymous. The need of a locomotive smoke consumer. Scientific American, 87(1):2, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
2b.pdf.

Hale:1902:AIE

[4312] William H. Hale. The American Institute of Electrical Engineers — Great Barrington Meeting. Scientific American, 87(1):2–3, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-2c.pdf.

Benjamin:1902:AH

[4313] Marcus Benjamin. Asaph Hall. Scientific American, 87(1):3, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-3.pdf.

Anonymous:1902:TRBb

[4314] Anonymous. Timber railway bridges in Australia. Scientific American, 87(1):3–4, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
3a.pdf.

Fawcett:1902:PTS

[4315] Waldon Fawcett. Pneumatic tools in ship yards. Scientific American, 87(1):4–6, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
4.pdf.

Anonymous:1902:EEM

[4316] Anonymous. Experiments with electro-magnetic waves on mont blanc. Scientific American, 87(1):6, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
6.pdf.

Willey:1902:RTM

[4317] Day Allen Willey. Rapid transit of material. Scientific American, 87(1):8–9, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
8.pdf.

Holder:1902:SCL

[4318] Charles F. Holder. Some California lizards. Scientific American, 87(1): 9, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-9.pdf.

Frederick:1902:SBB

[4319] M. C. Frederick. Santa Barbara’s big grapevine. Scientific American, 87(1):9–10, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
9a.pdf.

Anonymous:1902:DSD

[4320] Anonymous. Device for securing door-knobs. Scientific American, 87(1):10, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
10.pdf.

Anonymous:1902:SWW

[4321] Anonymous. A summer weather water boiler. Scientific American, 87(1):10, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n1/pdf/scientificamerican07051902-
10a.pdf.

Anonymous:1902:HRb

[4322] Anonymous. Harvester reel. Scientific American, 87(1):10–11, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-10b.pdf.

Anonymous:1902:IVB

[4323] Anonymous. Improved veneer basket. Scientific American, 87(1):11, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-11.pdf.

Collins:1902:SEW

[4324] A. Frederick Collins. A simple and efficient wireless telegraph receiver. Scientific American, 87(1):11, July 5, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n1/pdf/scientificamerican07051902-11a.pdf.

Anonymous:1902:BPWba

[4325] Anonymous. Business and personal wants. Scientific American, 87(1): 12, July 5, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n1/pdf/scientificamerican07051902-12a.pdf.

Russell:1902:HS

[4326] Henry Norris Russell. The heavens in September, 1902. Scientific American, 87(10):149, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
149.pdf.

Hopkins:1902:C

[4327] George M. Hopkins. Correspondence. Scientific American, 87(10): 152–153, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
152a.pdf.

Anonymous:1902:LNj

[4328] Anonymous. Legal notes. Scientific American, 87(10):158, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n10/pdf/scientificamerican09061902-158.pdf.

Anonymous:1902:RPIbb

[4329] Anonymous. Recently patented inventions. Scientific American, 87(10):159, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
159.pdf.

Anonymous:1902:IIbb

[4330] Anonymous. Index of inventions. Scientific American, 87(10): 160, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
160.pdf.

Anonymous:1902:PNPt

[4331] Anonymous. Patent notes: Patents. Scientific American, 87 (10):162, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
162.pdf.

Anonymous:1902:BNBp

[4332] Anonymous. Books: New books, etc. Scientific American, 87 (10):162, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
162a.pdf.

Anonymous:1902:NQbb

[4333] Anonymous. Notes and queries. Scientific American, 87(10): 163, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
163.pdf.

Anonymous:1902:VUR

[4334] Anonymous. Ventilation of underground roads. Scientific American, 87(10):148, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
148.pdf.

Anonymous:1902:WPM

[4335] Anonymous. West Point methods for Sandhurst. Scientific American, 87(10):148, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
148a.pdf.

Anonymous:1902:PBC

[4336] Anonymous. A point in boiler construction. Scientific American, 87(10):148, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
148b.pdf.

Anonymous:1902:OFB

[4337] Anonymous. Our fastest battleship. Scientific American, 87(10): 148, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
148c.pdf.

Anonymous:1902:EGS

[4338] Anonymous. Electric glass smelting. Scientific American, 87 (10):148, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
148d.pdf.

Anonymous:1902:EIC

[4339] Anonymous. The emerald industry of Colombia. Scientific American, 87(10):148–149, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
148e.pdf.

Anonymous:1902:SACb

[4340] Anonymous. The supplanting of agriculture by chemistry. Scientific American, 87(10):149, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
149a.pdf.

Anonymous:1902:AGH

[4341] Anonymous. Another great hydraulic electric power plant. Scientific American, 87(10):149, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
149b.pdf.

Anonymous:1902:WWE

[4342] Anonymous. Where wood engraving originated. Scientific American, 87(10):149, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
149c.pdf.

Anonymous:1902:NTWb

[4343] Anonymous. New ten-wheeled passenger engine. Scientific American, 87(10):150, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
150.pdf.

Anonymous:1902:NSSb

[4344] Anonymous. A new system of sinking shafts. Scientific American, 87(10):150–151, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
150a.pdf.

Anonymous:1902:SCL

[4345] Anonymous. Schneider–Canet launching apparatus for submarine torpedoes. Scientific American, 87(10):151, September 6, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n10/pdf/scientificamerican09061902-151.pdf.

Anonymous:1902:HPMa

[4346] Anonymous. Hydraulic process for manufacturing gun-cotton charges. Scientific American, 87(10):151–152, September 6, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n10/pdf/scientificamerican09061902-151a.pdf.

Anonymous:1902:NYE

[4347] Anonymous. The New York Edison Power Station. Scientific American, 87(10):152, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
152.pdf.

Fawcett:1902:SGT

[4348] Waldon Fawcett. Shade-grown tobacco. Scientific American, 87 (10):154–155, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
154.pdf.

Anonymous:1902:RBO

[4349] Anonymous. Remarkable blasting operations in slate quarries. Scientific American, 87(10):155, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
155.pdf.

Anonymous:1902:NRBb

[4350] Anonymous. The new Russian battleship “Kniaz Potemkin Tavritchesky”. Scientific American, 87(10):156, September 6, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n10/pdf/scientificamerican09061902-156.pdf.

Anonymous:1902:LWW

[4351] Anonymous. Largest watch in the world. Scientific American, 87 (10):156, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
156a.pdf.

Anonymous:1902:OIb

[4352] Anonymous. Oddities in inventions. Scientific American, 87 (10):157, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
157.pdf.

Anonymous:1902:NSCa

[4353] Anonymous. A new saw-clamp. Scientific American, 87(10): 157–158, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
157a.pdf.

Anonymous:1902:BPWbb

[4354] Anonymous. Business and personal wants. Scientific American, 87(10):159, September 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n10/pdf/scientificamerican09061902-
159a.pdf.

Anonymous:1902:Cu

[4355] Anonymous. Correspondence. Scientific American, 87(11):170– 171, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
170a.pdf.

Anonymous:1902:RPIbc

[4356] Anonymous. Recently patented inventions. Scientific American, 87 (11):176, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
176.pdf.

Anonymous:1902:IIbc

[4357] Anonymous. Index of inventions. Scientific American, 87(11): 176, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
176b.pdf.

Anonymous:1902:NQbc

[4358] Anonymous. Notes and queries. Scientific American, 87(11): 177, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
177.pdf.

Anonymous:1902:BNBq

[4359] Anonymous. Books: New books, etc. Scientific American, 87 (11):179, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
179.pdf.

Anonymous:1902:PNPu

[4360] Anonymous. Patent notes: Patents. Scientific American, 87 (11):180, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
180.pdf.

Anonymous:1902:M

[4361] Anonymous. The motorman. Scientific American, 87(11):166, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n11/pdf/scientificamerican09131902-166.pdf.

Anonymous:1902:ACR

[4362] Anonymous. “America” Cup Races in 1903. Scientific American, 87(11):166, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
166a.pdf.

Anonymous:1902:DLI

[4363] Anonymous. The defenses of Long Island Sound. Scientific American, 87(11):166, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
166b.pdf.

Anonymous:1902:ETL

[4364] Anonymous. Electrical treatment for lead poisoning. Scientific American, 87(11):166–167, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
166c.pdf.

Anonymous:1902:NWM

[4365] Anonymous. New ways of measuring water. Scientific American, 87(11):167, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
167.pdf.

Anonymous:1902:DPV

[4366] Anonymous. The death of Prof. Virchow. Scientific American, 87 (11):167, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
167a.pdf.

Anonymous:1902:MG

[4367] Anonymous. Malleable glass. Scientific American, 87(11):168, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n11/pdf/scientificamerican09131902-168.pdf.

Anonymous:1902:LCM

[4368] Anonymous. The Loomis carbureter and muffler. Scientific American, 87(11):169, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
169.pdf.

Anonymous:1902:VAG

[4369] Anonymous. Various applications of the gasoline motor. Scientific American, 87(11):169–170, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
169a.pdf.

Anonymous:1902:HVB

[4370] Anonymous. Headlight with a vertical beam. Scientific American, 87(11):170, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
170.pdf.

Fawcett:1902:EUS

[4371] Waldon Fawcett. The enlargement of the United States Capitol. Scientific American, 87(11):172, September 13, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n11/pdf/scientificamerican09131902-172.pdf.

Rost:1902:QCS

[4372] E. C. Rost. Queer craft seen on my travels in the Philippines, China and South America. Scientific American, 87(11):173, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n11/pdf/scientificamerican09131902-173.pdf.

Anonymous:1902:AW

[4373] Anonymous. Adjustable wheelbarrow. Scientific American, 87 (11):174, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
174.pdf.

Anonymous:1902:ACF

[4374] Anonymous. Automatic chemical fire extinguisher. Scientific American, 87(11):174, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
174a.pdf.

Anonymous:1902:SAb

[4375] Anonymous. A shutter attachment. Scientific American, 87(11): 174–175, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
174b.pdf.

MCurdy:1902:BCN

[4376] James G. M’Curdy. Boring clams of the Northwest Coast. Scientific American, 87(11):175, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
175.pdf.

Anonymous:1902:BPWbc

[4377] Anonymous. Business and personal wants. Scientific American, 87 (11):176, September 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n11/pdf/scientificamerican09131902-
176a.pdf.

Anonymous:1902:Cv

[4378] Anonymous. Correspondence. Scientific American, 87(12):187, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n12/pdf/scientificamerican09201902-187.pdf.

Anonymous:1902:RPIbd

[4379] Anonymous. Recently patented inventions. Scientific American, 87 (12):192, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
192.pdf.

Anonymous:1902:IIbd

[4380] Anonymous. Index of inventions. Scientific American, 87(12): 192, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
192b.pdf.

Anonymous:1902:NQbd

[4381] Anonymous. Notes and queries. Scientific American, 87(12): 195, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
195.pdf.

Anonymous:1902:PNPv

[4382] Anonymous. Patent notes: Patents. Scientific American, 87 (12):196, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
196.pdf.

Anonymous:1902:STA

[4383] Anonymous. The speed trials of the “Arrow”. Scientific American, 87(12):182, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
182.pdf.

Anonymous:1902:INV

[4384] Anonymous. The irrigation of the Nile Valley. Scientific American, 87(12):182, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
182a.pdf.

Anonymous:1902:SCD

[4385] Anonymous. Soft coal for domestic use. Scientific American, 87 (12):182, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
182b.pdf.

Anonymous:1902:DI

[4386] Anonymous. Distribution of industries. Scientific American, 87 (12):182–183, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
182c.pdf.

Anonymous:1902:AWF

[4387] Anonymous. “All the world’s fighting ships”. Scientific American, 87(12):183, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
183.pdf.

Anonymous:1902:IMC

[4388] Anonymous. International Mining Congress — Butte Meeting. Scientific American, 87(12):183, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
183a.pdf.

Athens:1902:EC

[4389] Paul Elsner Athens. Excavations at Corinth. Scientific American, 87(12):184, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
184.pdf.

Anonymous:1902:SAP

[4390] Anonymous. The semi-automatic piano-player. Scientific American, 87(12):185–186, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
185.pdf.

Smith:1902:IDO

[4391] Edmund S. Smith. An improved drying oven. Scientific American, 87(12):186, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
186.pdf.

Anonymous:1902:SYA

[4392] Anonymous. The steam yacht “Arrow” — the fastest boat in the world. Scientific American, 87(12):188–189, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n12/pdf/scientificamerican09201902-188.pdf.

Anonymous:1902:NIW

[4393] Anonymous. The Nile Irrigation Works. Scientific American, 87 (12):189–191, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
189.pdf.

Anonymous:1902:UW

[4394] Anonymous. Universal workholder. Scientific American, 87(12): 191, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
191.pdf.

Anonymous:1902:IK

[4395] Anonymous. An improved knife. Scientific American, 87(12): 191, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
191a.pdf.

Anonymous:1902:SHC

[4396] Anonymous. A stall for horses or cattle. Scientific American, 87 (12):191, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
191b.pdf.

Anonymous:1902:NTWc

[4397] Anonymous. A new type of water cooler. Scientific American, 87 (12):191, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
191c.pdf.

Anonymous:1902:BPWbd

[4398] Anonymous. Business and personal wants. Scientific American, 87 (12):192, September 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n12/pdf/scientificamerican09201902-
192a.pdf.

Anonymous:1902:Cx

[4399] Anonymous. Correspondence. Scientific American, 87(13):203, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n13/pdf/scientificamerican09271902-203.pdf.

Anonymous:1902:LNk

[4400] Anonymous. Legal notes. Scientific American, 87(13):208, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n13/pdf/scientificamerican09271902-208.pdf.

Anonymous:1902:RPIbe

[4401] Anonymous. Recently patented inventions. Scientific American, 87 (13):209, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
209.pdf.

Anonymous:1902:BNBr

[4402] Anonymous. Books: New books, etc. Scientific American, 87 (13):210, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
210a.pdf.

Anonymous:1902:NQbe

[4403] Anonymous. Notes and queries. Scientific American, 87(13): 212, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
212.pdf.

Anonymous:1902:EPN

[4404] Anonymous. Excellent progress on the New York Subway. Scientific American, 87(13):198, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
198.pdf.

Anonymous:1902:LLV

[4405] Anonymous. Long life vs the scrap heap. Scientific American, 87 (13):198, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
198a.pdf.

Anonymous:1902:LAS

[4406] Anonymous. The latest Atlantic steamship. Scientific American, 87(13):198, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
198b.pdf.

Anonymous:1902:RLP

[4407] Anonymous. The return of Lieut. Peary. Scientific American, 87 (13):198, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
198c.pdf.

Russell:1902:HO

[4408] Henry Norris Russell. The heavens in October, 1902. Scientific American, 87(13):199, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
199.pdf.

Gehlen:1902:FMB

[4409] A. Gehlen. The first map bearing the name of America. Scientific American, 87(13):199, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
199a.pdf.

Anonymous:1902:PEB

[4410] Anonymous. Portable engine boiler explosion. Scientific American, 87(13):200, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
200.pdf.

Anonymous:1902:MPE

[4411] Anonymous. Most powerful express locomotive in Great Britain. Scientific American, 87(13):200, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
200a.pdf.

Anonymous:1902:THR

[4412] Anonymous. Temperature and humidity regulators. Scientific American, 87(13):201, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
201.pdf.

Willey:1902:WA

[4413] Day Allan Willey. The war automobile. Scientific American, 87 (13):201–202, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
201a.pdf.

Anonymous:1902:RTS

[4414] Anonymous. The rapid transit subway, New York. Scientific American, 87(13):202–203, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
202.pdf.

Holder:1902:TMS

[4415] Charles F. Holder. Town making a science. Scientific American, 87 (13):204–205, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
204.pdf.

Anonymous:1902:JMP

[4416] Anonymous. The Jaubert method of producing oxygen gas. Scientific American, 87(13):205–206, September 27, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n13/pdf/scientificamerican09271902-205.pdf.

Anonymous:1902:PDh

[4417] Anonymous. Patent department. Scientific American, 87(13): 206, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
206.pdf.

Anonymous:1902:V

[4418] Anonymous. Ventilated hat. Scientific American, 87(13):207, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n13/pdf/scientificamerican09271902-207.pdf.

Anonymous:1902:GJ

[4419] Anonymous. Gurgless jug. Scientific American, 87(13):207, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n13/pdf/scientificamerican09271902-207a.pdf.

Anonymous:1902:MSTa

[4420] Anonymous. Magnetically-supported train. Scientific American, 87 (13):207, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
207b.pdf.

Anonymous:1902:MW

[4421] Anonymous. Mop-wringers. Scientific American, 87(13):207, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n13/pdf/scientificamerican09271902-207c.pdf.

Anonymous:1902:CHB

[4422] Anonymous. Combined horn and bell. Scientific American, 87(13): 207–208, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
207d.pdf.

Anonymous:1902:BPWbe

[4423] Anonymous. Business and personal wants. Scientific American, 87 (13):209, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
209a.pdf.

Anonymous:1902:IIbe

[4424] Anonymous. Index of inventions. Scientific American, 87(13): 210, September 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n13/pdf/scientificamerican09271902-
210.pdf.

Anonymous:1902:SNn

[4425] Anonymous. Science notes. Scientific American, 87(14):217, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-217a.pdf.

Anonymous:1902:Cy

[4426] Anonymous. Correspondence. Scientific American, 87(14):221, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-221a.pdf.

Anonymous:1902:RPIbf

[4427] Anonymous. Recently patented inventions. Scientific American, 87(14):227, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
227.pdf.

Anonymous:1902:IIbf

[4428] Anonymous. Index of inventions. Scientific American, 87(14):227, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-227b.pdf.

Anonymous:1902:BNBs

[4429] Anonymous. Books: New books, etc. Scientific American, 87(14):229, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-229.pdf.

Anonymous:1902:PNPw

[4430] Anonymous. Patent notes: Patents. Scientific American, 87(14):230, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-230.pdf.

Anonymous:1902:NQbf

[4431] Anonymous. Notes and queries. Scientific American, 87(14):231, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-231.pdf.

Anonymous:1902:FOM

[4432] Anonymous. Fuel oil in the merchant marine. Scientific American, 87(14):216, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
216.pdf.

Anonymous:1902:ISTb

[4433] Anonymous. Important speed tests of steam and electric trains. Scientific American, 87(14):216, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
216a.pdf.

Anonymous:1902:CCC

[4434] Anonymous. The collapse of the Campanile — the cause. Scientific American, 87(14):216–217, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
216b.pdf.

Walsh:1902:POW

[4435] George Ethelbert Walsh. Protecting our wild animals. Scientific American, 87(14):217, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
217.pdf.

Anonymous:1902:PMR

[4436] Anonymous. Paris Metropolitan Railroad — new systems. Scientific American, 87(14):218, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
218.pdf.

Anonymous:1902:EC

[4437] Anonymous. Electric call. Scientific American, 87(14):219, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-219.pdf.

Fawcett:1902:MPD

[4438] Waldon Fawcett. Modern pile drivers. Scientific American, 87 (14):219–220, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
219a.pdf.

Collins:1902:SND

[4439] A. Frederick Collins. Some new detectors for wireless telegraphy. Scientific American, 87(14):220–221, October 4, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n14/pdf/scientificamerican10041902-220.pdf.

Anonymous:1902:MTPa

[4440] Anonymous. The manufacture of tin plate — I. Scientific American, 87(14):221, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
221.pdf.

Anonymous:1902:MNF

[4441] Anonymous. The manufacture of the new fireproof material Uralite. Scientific American, 87(14):222–223, October 4, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n14/pdf/scientificamerican10041902-222.pdf.

Anonymous:1902:SAd

[4442] Anonymous. Stevens’ airship. Scientific American, 87(14):223–224, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-223.pdf.

Rothermel:1902:PMS

[4443] John G. Rothermel. Prehistoric Mexican sign writing. Scientific American, 87(14):225–226, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
225.pdf.

Anonymous:1902:TR

[4444] Anonymous. The Truffault racer. Scientific American, 87(14):226, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n14/pdf/scientificamerican10041902-226.pdf.

Anonymous:1902:BPWbf

[4445] Anonymous. Business and personal wants. Scientific American, 87(14):227, October 4, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n14/pdf/scientificamerican10041902-
227a.pdf.

Anonymous:1902:SNo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n15/pdf/scientificamerican10111902-235c.pdf.

Anonymous:1902:RPIbg

[4447] Anonymous. Recently patented inventions. Scientific American, 87(15):244, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
244.pdf.

Anonymous:1902:IIbg

[4448] Anonymous. Index of inventions. Scientific American, 87(15): 244–245, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
244b.pdf.

Anonymous:1902:BNBt

[4449] Anonymous. Books: New books, etc. Scientific American, 87 (15):245–246, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
245.pdf.

Anonymous:1902:NQbg

[4450] Anonymous. Notes and queries. Scientific American, 87(15):246, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n15/pdf/scientificamerican10111902-246.pdf.

Anonymous:1902:PNPx

[4451] Anonymous. Patent notes: Patents. Scientific American, 87 (15):250, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
250.pdf.

Anonymous:1902:TPCa

[4452] Anonymous. Titles to the Panama Canal. Scientific American, 87(15):234, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
234.pdf.

Anonymous:1902:BGS

[4453] Anonymous. The British Government and the Ship Combine. Scientific American, 87(15):234, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
234a.pdf.

Anonymous:1902:FDN

[4454] Anonymous. Further development of Niagara Falls Power. Scientific American, 87(15):234, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
234b.pdf.

Anonymous:1902:WTB

[4455] Anonymous. The water tube boiler problem. Scientific American, 87(15):234, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
234c.pdf.

Anonymous:1902:POT

[4456] Anonymous. Power from oil in Texas. Scientific American, 87 (15):234–235, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
234d.pdf.

Anonymous:1902:WVR

[4457] Anonymous. The world’s volcano record for 1902. Scientific American, 87(15):235, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
235.pdf.

Anonymous:1902:TRA

[4458] Anonymous. Two rival airship ascents. Scientific American, 87 (15):235, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
235a.pdf.

Anonymous:1902:CBN

[4459] Anonymous. Colliers for the British Navy. Scientific American, 87(15):235, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
235b.pdf.

Anonymous:1902:CSa

[4460] Anonymous. The current supplement. Scientific American, 87 (15):235, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
235d.pdf.

Anonymous:1902:SCb

[4461] Anonymous. A Sicilian cyclone. Scientific American, 87(15):235, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n15/pdf/scientificamerican10111902-235e.pdf.

Mitchell:1902:RBS

[4462] George D. Mitchell. Razing a big stack with dynamite. Scientific American, 87(15):236, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
236.pdf.

Field:1902:RC

[4463] Alberta Field. The Rameses Colossus. Scientific American, 87 (15):236, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
236a.pdf.

Hall:1902:CEL

[4464] Alfred E. Hall. Curious effect of lightning on a brick chimney. Scientific American, 87(15):236, October 11, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n15/pdf/scientificamerican10111902-236b.pdf.

Anonymous:1902:EMP

[4465] Anonymous. The electrical manufacture of peat-fuel. Scientific American, 87(15):237–238, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
237.pdf.

Anonymous:1902:FAT

[4466] Anonymous. Faller automatic telephone exchange. Scientific American, 87(15):238–239, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
238.pdf.

Riffle:1902:BBT

[4467] Franklin Riffle. The big bend tunnel for power development. Scientific American, 87(15):239, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
239.pdf.

Fyfe:1902:ESC

[4468] H. C. Fyfe. Experiments at St. Catherine’s with fog signals. Scientific American, 87(15):240, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
240.pdf.

Holder:1902:CPA

[4469] Charles F. Holder. The Chinese press in America. Scientific American, 87(15):241, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
241.pdf.

Anonymous:1902:IPR

[4470] Anonymous. An improved piano resonator. Scientific American, 87(15):242, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
242.pdf.

Anonymous:1902:EDO

[4471] Anonymous. An electric detector of ores. Scientific American, 87(15):243, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
243.pdf.

Anonymous:1902:PPIb

[4472] Anonymous. Puncture-proof inflatable wheel. Scientific American, 87(15):243, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
243a.pdf.

Anonymous:1902:BPWbg

[4473] Anonymous. Business and personal wants. Scientific American, 87(15):244, October 11, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n15/pdf/scientificamerican10111902-
244a.pdf.

Anonymous:1902:SNp

[4474] Anonymous. Science notes. Scientific American, 87(16):253, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n16/pdf/scientificamerican10181902-253c.pdf.

Anonymous:1902:Cz

[4475] Anonymous. Correspondence. Scientific American, 87(16):257, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n16/pdf/scientificamerican10181902-257a.pdf.

Anonymous:1902:RPIbh

[4476] Anonymous. Recently patented inventions. Scientific American, 87(16):262, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
262.pdf.

Anonymous:1902:IIbh

[4477] Anonymous. Index of inventions. Scientific American, 87(16):263, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n16/pdf/scientificamerican10181902-263.pdf.

Anonymous:1902:NQbh

[4478] Anonymous. Notes and queries. Scientific American, 87(16):265, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n16/pdf/scientificamerican10181902-265.pdf.

Anonymous:1902:PNPy

[4479] Anonymous. Patent notes: Patents. Scientific American, 87 (16):266, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
266.pdf.

Anonymous:1902:PNO

[4480] Anonymous. The pressing need of our Navy. Scientific American, 87(16):252, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
252.pdf.

Anonymous:1902:BCA

[4481] Anonymous. British criticism of American railroads. Scientific American, 87(16):252, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
252a.pdf.

Anonymous:1902:WOM

[4482] Anonymous. Where our manufactures go. Scientific American, 87(16):252, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
252b.pdf.

Anonymous:1902:RES

[4483] Anonymous. Railroad expansion in South Africa. Scientific American, 87(16):252–253, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
252c.pdf.

Anonymous:1902:THA

[4484] Anonymous. The three hundredth anniversary of Oxford’s Bodleian Library. Scientific American, 87(16):253, October 18, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n16/pdf/scientificamerican10181902-253.pdf.

Anonymous:1902:EWT

[4485] Anonymous. An example of wireless telegraphy’s efficiency. Scientific American, 87(16):253, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
253a.pdf.

Anonymous:1902:STTa

[4486] Anonymous. Successful test of a tower elevator. Scientific American, 87(16):253, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
253b.pdf.

Anonymous:1902:CUA

[4487] Anonymous. The commercial utilization of atmospheric elements. Scientific American, 87(16):254–255, October 18, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n16/pdf/scientificamerican10181902-254.pdf.

Anonymous:1902:NYBa

[4488] Anonymous. New York–Boston Automobile Reliability Test — the first day’s trip. Scientific American, 87(16):255–256, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n16/pdf/scientificamerican10181902-255.pdf.

Anonymous:1902:IRB

[4489] Anonymous. Improved rail-bearing plate. Scientific American, 87(16):256, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
256.pdf.

Anonymous:1902:LBA

[4490] Anonymous. The latest British armored cruiser “Good Hope”. Scientific American, 87(16):257, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
257.pdf.

Fawcett:1902:CS

[4491] Waldon Fawcett. Coal storage. Scientific American, 87(16):258– 259, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
258.pdf.

Ward:1902:SAT

[4492] William Edward Ward. The Spencer airship and its thirty-mile trip. Scientific American, 87(16):259–260, October 18, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n16/pdf/scientificamerican10181902-259.pdf.

Shepstone:1902:STW

[4493] Harold J. Shepstone. The scientific training of wild animals. Scientific American, 87(16):260–261, October 18, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n16/pdf/scientificamerican10181902-260.pdf.

Anonymous:1902:BPWbh

[4494] Anonymous. Business and personal wants. Scientific American, 87(16):262, October 18, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n16/pdf/scientificamerican10181902-
262a.pdf.

Anonymous:1902:RPIbi

[4495] Anonymous. Recently patented inventions. Scientific American, 87(17):279, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
279.pdf.

Anonymous:1902:NQbi

[4496] Anonymous. Notes and queries. Scientific American, 87(17):279, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n17/pdf/scientificamerican10251902-279b.pdf.

Anonymous:1902:IIbi

[4497] Anonymous. Index of inventions. Scientific American, 87(17):279, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n17/pdf/scientificamerican10251902-279c.pdf.

Anonymous:1902:BNBu

[4498] Anonymous. Books: New books, etc. Scientific American, 87 (17):283, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
283.pdf.

Anonymous:1902:ACS

[4499] Anonymous. The aftermath of the coal strike. Scientific American, 87(17):268, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
268.pdf.

Anonymous:1902:OUS

[4500] Anonymous. Output of the United States Steel Corporation. Scientific American, 87(17):268, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
268a.pdf.

Anonymous:1902:EEG

[4501] Anonymous. Electric equipment of a great railroad. Scientific American, 87(17):268, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
268b.pdf.

Anonymous:1902:BTA

[4502] Anonymous. Balloon trip across the Sahara. Scientific American, 87(17):268, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
268c.pdf.

Anonymous:1902:RIW

[4503] Anonymous. Recent improvements in weaving. Scientific American, 87(17):268–269, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
268d.pdf.

Walsh:1902:MWN

[4504] George E. Walsh. Maine’s wood novelty mills. Scientific American, 87(17):269, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
269.pdf.

VonBrennerberg:1902:TN

[4505] William Von Brennerberg. A textile novelty. Scientific American, 87(17):269, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
269a.pdf.

Anonymous:1902:BAT

[4506] Anonymous. The De Bradsky airship and the tragic end of its inventor. Scientific American, 87(17):270–271, October 25, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n17/pdf/scientificamerican10251902-270.pdf.

Anonymous:1902:NYBb

[4507] Anonymous. New York–Boston Automobile Reliability Test. Scientific American, 87(17):271–272, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
271.pdf.

Ionescu:1902:KCB

[4508] Nicolae Ionescu. King Charles I Bridge across the Danube. Scientific American, 87(17):272–273, October 25, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n17/pdf/scientificamerican10251902-272.pdf.

Anonymous:1902:TNI

[4509] Anonymous. Two new incandescent gas burners. Scientific American, 87(17):273, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
273.pdf.

Anonymous:1902:RMS

[4510] Anonymous. The Rowland multiplex system of page-printing telegraphy. Scientific American, 87(17):274–276, October 25, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n17/pdf/scientificamerican10251902-274.pdf.

Anonymous:1902:SCHa

[4511] Anonymous. Substitutes for coal in heating and cooking. Scientific American, 87(17):276–277, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
276.pdf.

Anonymous:1902:NREb

[4512] Anonymous. A novel rotary engine. Scientific American, 87 (17):278, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
278.pdf.

Anonymous:1902:ESMb

[4513] Anonymous. Envelope-sealing machine. Scientific American, 87 (17):278, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
278a.pdf.

Anonymous:1902:BPWbi

[4514] Anonymous. Business and personal wants. Scientific American, 87(17):279, October 25, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n17/pdf/scientificamerican10251902-
279a.pdf.

Russell:1902:HN

[4515] Henry Norris Russell. The heavens in November. Scientific American, 87(18):287, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
287.pdf.

Anonymous:1902:SNq

[4516] Anonymous. Science notes. Scientific American, 87(18):287, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n18/pdf/scientificamerican11011902-287b.pdf.

Anonymous:1902:Cba

[4517] Anonymous. Correspondence. Scientific American, 87(18):291, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n18/pdf/scientificamerican11011902-291.pdf.

Anonymous:1902:RPIbj

[4518] Anonymous. Recently patented inventions. Scientific American, 87(18):296, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
296.pdf.

Anonymous:1902:NQbj

[4519] Anonymous. Notes and queries. Scientific American, 87(18): 296, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
296b.pdf.

Anonymous:1902:IIbj

[4520] Anonymous. Index of inventions. Scientific American, 87(18): 297, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
297.pdf.

Anonymous:1902:BNBv

[4521] Anonymous. Books: New books, etc. Scientific American, 87 (18):300, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
300.pdf.

Anonymous:1902:SV

[4522] Anonymous. Street ventilation. Scientific American, 87(18): 286, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
286.pdf.

Anonymous:1902:STC

[4523] Anonymous. The steam turbine from the commercial standpoint. Scientific American, 87(18):286, November 1, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n18/pdf/scientificamerican11011902-286a.pdf.

Anonymous:1902:TFK

[4524] Anonymous. The twenty five knot Cunard steamships. Scientific American, 87(18):286–287, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
286b.pdf.

Anonymous:1902:TNL

[4525] Anonymous. Test of a Norwegian life-saving boat. Scientific American, 87(18):287, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
287a.pdf.

Anonymous:1902:PWL

[4526] Anonymous. The “Preussen,” the world’s largest sailing vessel. Scientific American, 87(18):288, November 1, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n18/pdf/scientificamerican11011902-288.pdf.

Anonymous:1902:HMR

[4527] Anonymous. A huge meteorite recently discovered. Scientific American, 87(18):288, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
288a.pdf.

Anonymous:1902:AFE

[4528] Anonymous. An ancient fire engine. Scientific American, 87 (18):288, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
288b.pdf.

Anonymous:1902:SSM

[4529] Anonymous. The Sault Ste. Marie Water Power Canal. Scientific American, 87(18):289–290, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
289.pdf.

Anonymous:1902:MTPb

[4530] Anonymous. Manufacture of tin plate — II. Scientific American, 87(18):290, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
290.pdf.

Anonymous:1902:PW

[4531] Anonymous. A portable windmill. Scientific American, 87(18): 292, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
292.pdf.

Anonymous:1902:ROT

[4532] Anonymous. Ruhmer’s optic telephone. Scientific American, 87 (18):292, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
292a.pdf.

Beasley:1902:ZTU

[4533] Walter L. Beasley. The Zapotecan tombs unearthed. Scientific American, 87(18):293, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
293.pdf.

Anonymous:1902:TUI

[4534] Anonymous. Traps for the unwary inventor. Scientific American, 87(18):294, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
294.pdf.

Anonymous:1902:IDC

[4535] Anonymous. Improved drawing compasses. Scientific American, 87(18):295, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
295.pdf.

Anonymous:1902:TWC

[4536] Anonymous. Tool for withdrawing casings from wells. Scientific American, 87(18):295, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
295a.pdf.

Anonymous:1902:OB

[4537] Anonymous. Oil burner. Scientific American, 87(18):295, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n18/pdf/scientificamerican11011902-295b.pdf.

Anonymous:1902:STTb

[4538] Anonymous. Saw-tooth tool. Scientific American, 87(18):295, November 1, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n18/pdf/scientificamerican11011902-295c.pdf.

Anonymous:1902:BPWbj 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n18/pdf/scientificamerican11011902-
296a.pdf.

Anonymous:1902:SNr

[4540] Anonymous. Science notes. Scientific American, 87(19):305, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n19/pdf/scientificamerican11081902-305b.pdf.

Anonymous:1902:LNl

[4541] Anonymous. Legal notes. Scientific American, 87(19):314, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n19/pdf/scientificamerican11081902-314.pdf.

Anonymous:1902:RPIbk

[4542] Anonymous. Recently patented inventions. Scientific American, 87(19):315, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
315.pdf.

Anonymous:1902:NQbk

[4543] Anonymous. Notes and queries. Scientific American, 87(19): 315–316, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
315b.pdf.

Anonymous:1902:IIbk

[4544] Anonymous. Index of inventions. Scientific American, 87(19): 316, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
316.pdf.

Anonymous:1902:BNBw

[4545] Anonymous. Books: New books, etc. Scientific American, 87 (19):319, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
319.pdf.

Anonymous:1902:ECQ

[4546] Anonymous. The Erie Canal question. Scientific American, 87 (19):304, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
304.pdf.

Anonymous:1902:LFN

[4547] Anonymous. Liquid fuel for Naval purposes. Scientific American, 87(19):304, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
304a.pdf.

Anonymous:1902:NHS

[4548] Anonymous. A new hydride of silicon. Scientific American, 87 (19):304, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
304b.pdf.

Anonymous:1902:FPTb

[4549] Anonymous. Fruit parasites and their destruction. Scientific American, 87(19):304–305, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
304c.pdf.

Anonymous:1902:RWU

[4550] Anonymous. Recovery of waste in using white metals. Scientific American, 87(19):305, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
305.pdf.

Grimshaw:1902:BR

[4551] Robert Grimshaw. Barometer readings. Scientific American, 87 (19):305, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
305a.pdf.

Holder:1902:CUE

[4552] Charles F. Holder. A curious use for elk horn. Scientific American, 87(19):306, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
306.pdf.

Anonymous:1902:ELB

[4553] Anonymous. Explosion of a locomotive boiler. Scientific American, 87(19):306, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
306a.pdf.

Fawcett:1902:H

[4554] Waldon Fawcett. Hathamite. Scientific American, 87(19):307– 308, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
307.pdf.

Anonymous:1902:ATC

[4555] Anonymous. The automatic train controller. Scientific American, 87 (19):308–309, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
308.pdf.

Anonymous:1902:IG

[4556] Anonymous. The influence of grades. Scientific American, 87 (19):309, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
309.pdf.

Anonymous:1902:LS

[4557] Anonymous. The Lansing skeleton. Scientific American, 87(19): 310, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
310.pdf.

Anonymous:1902:PEF

[4558] Anonymous. The Pantheon experiments with Foucault’s Pendulum. Scientific American, 87(19):310–311, November 8, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n19/pdf/scientificamerican11081902-310a.pdf.

Brown:1902:DG

[4559] Enos Brown. Dredging for gold. Scientific American, 87(19): 311–312, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
311.pdf.

Anonymous:1902:NCD

[4560] Anonymous. The new cover design for United States Letters Patent. Scientific American, 87(19):312, November 8, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n19/pdf/scientificamerican11081902-312.pdf.

Anonymous:1902:CJJ

[4561] Anonymous. Col. J. J. Astor dedicates his marine turbine patents to the public. Scientific American, 87(19):312, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n19/pdf/scientificamerican11081902-312a.pdf.

Anonymous:1902:OIc

[4562] Anonymous. Oddities in inventions. Scientific American, 87 (19):313, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
313.pdf.

Anonymous:1902:BPWbk

[4563] Anonymous. Business and personal wants. Scientific American, 87(19):315, November 8, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n19/pdf/scientificamerican11081902-
315a.pdf.

Anonymous:1902:Cbb

[4564] Anonymous. Correspondence. Scientific American, 87(2):23, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-23.pdf.

Anonymous:1902:RPIbl

[4565] Anonymous. Recently patented inventions. Scientific American, 87(2):28, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
28.pdf.

Anonymous:1902:IIbl

[4566] Anonymous. Index of inventions. Scientific American, 87(2):29, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-29.pdf.

Anonymous:1902:PNPz

[4567] Anonymous. Patent notes: Patents. Scientific American, 87(2):31, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-31.pdf.

Anonymous:1902:NQbl

[4568] Anonymous. Notes and queries. Scientific American, 87(2):31, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-31a.pdf.

Anonymous:1902:TPCb

[4569] Anonymous. Title to the Panama Canal. Scientific American, 87(2):18, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-18.pdf.

Anonymous:1902:SBT

[4570] Anonymous. Solution of the bridge terminal problem. Scientific American, 87(2):18, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
18a.pdf.

Anonymous:1902:SAf

[4571] Anonymous. Science in America. Scientific American, 87(2):18, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-18b.pdf.

Anonymous:1902:BSa

[4572] Anonymous. The “blindness” of the submarine. Scientific American, 87(2):18–19, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
18c.pdf.

Anonymous:1902:MSTb

[4573] Anonymous. The metric system in the textile industry. Scientific American, 87(2):19, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
19.pdf.

Anonymous:1902:FAA

[4574] Anonymous. The French alcohol automobile endurance test. Scientific American, 87(2):19, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
19a.pdf.

Anonymous:1902:AMPb

[4575] Anonymous. Analysis of Mont Pelée’s volcanic dust. Scientific American, 87(2):19, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
19b.pdf.

Anonymous:1902:PVAa

[4576] Anonymous. The Paris–Vienna Automobile Race. Scientific American, 87(2):19, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
19c.pdf.

Anonymous:1902:MRB

[4577] Anonymous. Measuring the rainfall — the British Rainfall Organization and its work. Scientific American, 87(2):20–21, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-20.pdf.

Anonymous:1902:RSF

[4578] Anonymous. A remarkable shipbuilding feat, 13,000 feet above the sea level. Scientific American, 87(2):21–22, July 12, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n2/pdf/scientificamerican07121902-21.pdf.

Rost:1902:FOR

[4579] E. C. Rost. The famous Oroya Railroad of Peru which climbs higher than any other on the globe. Scientific American, 87(2):22, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-22.pdf.

Shepstone:1902:TNB

[4580] Harold J. Shepstone. Testing the new Bermuda Dock. Scientific American, 87(2):24, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
24.pdf.

Draper:1902:GHW

[4581] William R. Draper. Grass houses of the Wichita Indians. Scientific American, 87(2):25, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
25.pdf.

Holder:1902:AN

[4582] Charles F. Holder. Armored nests. Scientific American, 87(2):25, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-25a.pdf.

Parker:1902:EMD

[4583] W. E. Parker. An experimental motor and dynamo. Scientific American, 87(2):26, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
26.pdf.

Anonymous:1902:PPa

[4584] Anonymous. A perplexing puzzle. Scientific American, 87(2):26–27, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-26a.pdf.

Anonymous:1902:BTD

[4585] Anonymous. Beet-topping device. Scientific American, 87(2):27, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-27.pdf.

Anonymous:1902:ASOb

[4586] Anonymous. An automatic siphon overflow valve. Scientific American, 87(2):27, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
27a.pdf.

Anonymous:1902:FDH

[4587] Anonymous. Fastening device for horses. Scientific American, 87(2):27, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-27b.pdf.

Anonymous:1902:BPWbl

[4588] Anonymous. Business and personal wants. Scientific American, 87(2):28, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n2/pdf/scientificamerican07121902-
28a.pdf.

Anonymous:1902:BNa

[4589] Anonymous. Book notices. Scientific American, 87(2):30, July 12, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n2/pdf/scientificamerican07121902-30.pdf.

Anonymous:1902:SNs

[4590] Anonymous. Science notes. Scientific American, 87(20):323, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n20/pdf/scientificamerican11151902-323a.pdf.

Anonymous:1902:RPIbm

[4591] Anonymous. Recently patented inventions. Scientific American, 87 (20):331–332, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
331a.pdf.

Anonymous:1902:NQbm

[4592] Anonymous. Notes and queries. Scientific American, 87(20): 332, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
332a.pdf.

Anonymous:1902:IIbm

[4593] Anonymous. Index of inventions. Scientific American, 87(20): 332, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
332b.pdf.

Anonymous:1902:PNPba

[4594] Anonymous. Patent notes: Patents. Scientific American, 87 (20):335, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
335.pdf.

Anonymous:1902:BNBx

[4595] Anonymous. Books: New books, etc. Scientific American, 87 (20):335, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
335a.pdf.

Anonymous:1902:MAE

[4596] Anonymous. Municipal architecture and the elevated railway. Scientific American, 87(20):322, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
322.pdf.

Anonymous:1902:ASE

[4597] Anonymous. The Arnold system of electrical traction. Scientific American, 87(20):322–323, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
322a.pdf.

Holder:1902:VSF

[4598] Charles Frederick Holder. Vocal sounds of fishes. Scientific American, 87(20):323, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
323.pdf.

Harris:1902:PCD

[4599] Thomas C. Harris. A punch centering device. Scientific American, 87(20):324, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
324.pdf.

Anonymous:1902:IPM

[4600] Anonymous. The inclined plane of the Morris Canal. Scientific American, 87(20):324, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
324a.pdf.

Anonymous:1902:MPL

[4601] Anonymous. Making a photographic lens. Scientific American, 87 (20):324–327, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
324b.pdf.

McCurdy:1902:UPE

[4602] James G. McCurdy. Use of photographs as evidence. Scientific American, 87(20):328, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
328.pdf.

Anonymous:1902:ASAa

[4603] Anonymous. Acetylene searchlights for automobiles. Scientific American, 87(20):328–329, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
328a.pdf.

Anonymous:1902:IEF

[4604] Anonymous. The improved electrograph — a fac-simile telegraph. Scientific American, 87(20):329–330, November 15, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n20/pdf/scientificamerican11151902-329.pdf.

Anonymous:1902:SCHb

[4605] Anonymous. Some curious homes. Scientific American, 87(20): 330, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
330.pdf.

Anonymous:1902:BAD

[4606] Anonymous. De Bradsky airship disaster. Scientific American, 87 (20):331, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
331.pdf.

Anonymous:1902:BPWbm

[4607] Anonymous. Business and personal wants. Scientific American, 87(20):332, November 15, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n20/pdf/scientificamerican11151902-
332.pdf.

Anonymous:1902:Cbd

[4608] Anonymous. Correspondence. Scientific American, 87(21):343, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n21/pdf/scientificamerican11221902-343.pdf.

Anonymous:1902:RPIbn

[4609] Anonymous. Recently patented inventions. Scientific American, 87(21):348, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
348.pdf.

Anonymous:1902:NQbn

[4610] Anonymous. Notes and queries. Scientific American, 87(21): 348, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
348b.pdf.

Anonymous:1902:IIbn

[4611] Anonymous. Index of inventions. Scientific American, 87(21): 348, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
348c.pdf.

Anonymous:1902:LER

[4612] Anonymous. The lesson of the East River Bridge Fire. Scientific American, 87(21):338, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
338.pdf.

Anonymous:1902:PGB

[4613] Anonymous. Pipe galleries for the Broadway Tunnel. Scientific American, 87(21):338, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
338a.pdf.

Anonymous:1902:RIR

[4614] Anonymous. A result of improved rapid transit. Scientific American, 87(21):338, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
338b.pdf.

Walsh:1902:NAL

[4615] George Ethelbert Walsh. The new American lighthouse system. Scientific American, 87(21):338–339, November 22, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n21/pdf/scientificamerican11221902-338c.pdf.

Weir:1902:PCC

[4616] James Weir, Jr. The physiology of certain colors. Scientific American, 87(21):339, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
339.pdf.

Anonymous:1902:TNW

[4617] Anonymous. Two new warships for Chile. Scientific American, 87(21):339, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
339a.pdf.

Anonymous:1902:RCC

[4618] Anonymous. Revolving chimney cap. Scientific American, 87 (21):340, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
340.pdf.

Anonymous:1902:NSP

[4619] Anonymous. A new soot-proof spark plug for gasoline motors. Scientific American, 87(21):340, November 22, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n21/pdf/scientificamerican11221902-340a.pdf.

Anonymous:1902:AAb

[4620] Anonymous. An automobile ambulance. Scientific American, 87 (21):340, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
340b.pdf.

Anonymous:1902:HBb

[4621] Anonymous. Horseshoer’s box. Scientific American, 87(21):340, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n21/pdf/scientificamerican11221902-340c.pdf.

Anonymous:1902:PBA

[4622] Anonymous. Popp–Branly aerial telegraphy systems. Scientific American, 87(21):341–342, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
341.pdf.

Anonymous:1902:NER

[4623] Anonymous. The new East River Bridge Fire. Scientific American, 87(21):344–346, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
344.pdf.

Anonymous:1902:RPA

[4624] Anonymous. The raising of peas and asparagus in the West. Scientific American, 87(21):346, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
346.pdf.

Fawcett:1902:SBP

[4625] Waldon Fawcett. The submarine boat “Protector”. Scientific American, 87(21):346–347, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
346a.pdf.

Anonymous:1902:HCT

[4626] Anonymous. Head-on collision of trains in Los Angeles, Cal. Scientific American, 87(21):347, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
347.pdf.

Anonymous:1902:BPWbn

[4627] Anonymous. Business and personal wants. Scientific American, 87(21):348, November 22, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n21/pdf/scientificamerican11221902-
348a.pdf.

Russell:1902:HD

[4628] Henry Norris Russell. The heavens in December. Scientific American, 87(22):355, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
355.pdf.

Anonymous:1902:RPIbo

[4629] Anonymous. Recently patented inventions. Scientific American, 87(22):365, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
365.pdf.

Anonymous:1902:NQbo

[4630] Anonymous. Notes and queries. Scientific American, 87(22): 366, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
366.pdf.

Anonymous:1902:IIbo

[4631] Anonymous. Index of inventions. Scientific American, 87(22): 366, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
366a.pdf.

Anonymous:1902:PNPbb

[4632] Anonymous. Patent notes: Patents. Scientific American, 87 (22):369, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
369.pdf.

Anonymous:1902:BNBy

[4633] Anonymous. Books: New books, etc. Scientific American, 87 (22):369, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
369a.pdf.

Anonymous:1902:RCBa

[4634] Anonymous. Report of the Chief of the Bureau of Ordnance. Scientific American, 87(22):354, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
354.pdf.

Anonymous:1902:CNE

[4635] Anonymous. Criticism of the new East River Bridge. Scientific American, 87(22):354, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
354a.pdf.

Anonymous:1902:HPMb

[4636] Anonymous. A 200 horse power motor propelled boat. Scientific American, 87(22):354–355, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
354b.pdf.

Benjamin:1902:NAS

[4637] Marcus Benjamin. The National Academy of Sciences. Scientific American, 87(22):355, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
355a.pdf.

Fawcett:1902:AHI

[4638] Waldon Fawcett. The American hemp industry. Scientific American, 87(22):356, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
356.pdf.

Anonymous:1902:MLB

[4639] Anonymous. A machine for laying bricks. Scientific American, 87 (22):357, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
357.pdf.

Anonymous:1902:OTO

[4640] Anonymous. Official trials of our submarine boats. Scientific American, 87(22):357–358, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
357a.pdf.

Anonymous:1902:NBH

[4641] Anonymous. The new Beachy Head lighthouse. Scientific American, 87(22):358–359, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
358.pdf.

Anonymous:1902:NE

[4642] Anonymous. Natural effigies. Scientific American, 87(22):360, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n22/pdf/scientificamerican11291902-360.pdf.

Campbell:1902:DME

[4643] W. W. Campbell. The D. O. Mills Expedition to the Southern Hemisphere. Scientific American, 87(22):360–361, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n22/pdf/scientificamerican11291902-360a.pdf.

Willey:1902:PRF

[4644] Day Allen Willey. Protecting a railroad from flood currents. Scientific American, 87(22):361–362, November 29, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n22/pdf/scientificamerican11291902-361.pdf.

Beasley:1902:SHI

[4645] Walter L. Beasley. The symbolism of the Huichol Indians. Scientific American, 87(22):362–363, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
362.pdf.

Anonymous:1902:PDi

[4646] Anonymous. Patent department. Scientific American, 87(22): 363, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
363.pdf.

Anonymous:1902:OId

[4647] Anonymous. Oddities in inventions. Scientific American, 87 (22):364, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
364.pdf.

Anonymous:1902:PF

[4648] Anonymous. Printing frame. Scientific American, 87(22):364, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n22/pdf/scientificamerican11291902-364a.pdf.

Anonymous:1902:W

[4649] Anonymous. Wrench. Scientific American, 87(22):364, November 29, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n22/pdf/scientificamerican11291902-364b.pdf.

Anonymous:1902:BPWbo 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n22/pdf/scientificamerican11291902-
365a.pdf.

Anonymous:1902:SNt

[4651] Anonymous. Science notes. Scientific American, 87(23):373, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n23/pdf/scientificamerican12061902-373b.pdf.

Anonymous:1902:Cbe

[4652] Anonymous. Correspondence. Scientific American, 87(23):376–377, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n23/pdf/scientificamerican12061902-376a.pdf.

Anonymous:1902:RPIbp

[4653] Anonymous. Recently patented inventions. Scientific American, 87(23):381, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
381.pdf.

Anonymous:1902:NQbp

[4654] Anonymous. Notes and queries. Scientific American, 87(23):381, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n23/pdf/scientificamerican12061902-381b.pdf.

Anonymous:1902:IIbp

[4655] Anonymous. Index of inventions. Scientific American, 87(23):382, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n23/pdf/scientificamerican12061902-382.pdf.

Anonymous:1902:PNPbc

[4656] Anonymous. Patent notes: Patents. Scientific American, 87 (23):385, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
385.pdf.

Anonymous:1902:BNBz

[4657] Anonymous. Books: New books, etc. Scientific American, 87 (23):385, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
385a.pdf.

Anonymous:1902:PRTb

[4658] Anonymous. The Pennsylvania Railroad Tunnel. Scientific American, 87(23):372, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
372.pdf.

Anonymous:1902:TMC

[4659] Anonymous. Test of the maximum capacity of the Niagara Falls Tunnel. Scientific American, 87(23):372, December 6, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n23/pdf/scientificamerican12061902-372a.pdf.

Anonymous:1902:GNW

[4660] Anonymous. The gun in Naval warfare. Scientific American, 87 (23):372, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
372b.pdf.

Anonymous:1902:DFK

[4661] Anonymous. The death of Friedrich Krupp. Scientific American, 87 (23):372–373, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
372c.pdf.

Townsend:1902:APT

[4662] Irving U. Townsend. An American parallel to the tulip craze in Holland. Scientific American, 87(23):373, December 6, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n23/pdf/scientificamerican12061902-373.pdf.

Anonymous:1902:RP

[4663] Anonymous. Restoration of the Parthenon. Scientific American, 87(23):373, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
373a.pdf.

Anonymous:1902:NFH

[4664] Anonymous. Non-freezing hydrant. Scientific American, 87(23): 374, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
374.pdf.

Anonymous:1902:RSPb

[4665] Anonymous. Reversible screw-propeller. Scientific American, 87 (23):374, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
374a.pdf.

Anonymous:1902:NME

[4666] Anonymous. The Navipendulum method of experiment. Scientific American, 87(23):374–375, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
374b.pdf.

Dunlap:1902:NPC

[4667] Orrin E. Dunlap. The new plant of the Canadian Niagara Falls Company. Scientific American, 87(23):375–376, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n23/pdf/scientificamerican12061902-375.pdf.

Anonymous:1902:MTD

[4668] Anonymous. The manufacture of toys and dolls. Scientific American, 87(23):376, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
376.pdf.

Anonymous:1902:NMWa

[4669] Anonymous. The new monitor “Wyoming”. Scientific American, 87(23):378, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
378.pdf.

Dawley:1902:GE

[4670] Thomas R. Dawley, Jr. Guatemala’s earthquakes. Scientific American, 87(23):379–380, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
379.pdf.

Willey:1902:EMR

[4671] Day Allen Willey. An experimental mono-rail line. Scientific American, 87(23):380, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
380.pdf.

Anonymous:1902:BPWbp

[4672] Anonymous. Business and personal wants. Scientific American, 87(23):381, December 6, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n23/pdf/scientificamerican12061902-
381a.pdf.

Anonymous:1902:NQbq

[4673] Anonymous. Notes and queries. Scientific American, 87(24): 419–426, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
419.pdf.

Anonymous:1902:RPIbq

[4674] Anonymous. Recently patented inventions. Scientific American, 87 (24):427–428, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
427.pdf.

Anonymous:1902:BNBba

[4675] Anonymous. Books: New books, etc. Scientific American, 87(24): 429–430, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
429.pdf.

Anonymous:1902:IIbq

[4676] Anonymous. Index of inventions. Scientific American, 87(24): 430–434, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
430.pdf.

Anonymous:1902:AMM

[4677] Anonymous. The American merchant marine. Scientific American, 87(24):388–389, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
388.pdf.

Anonymous:1902:TGF

[4678] Anonymous. The two great freighters for the Pacific trade. Scientific American, 87(24):389, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
389.pdf.

Anonymous:1902:LFT

[4679] Anonymous. The latest of the fast transatlantic liners. Scientific American, 87(24):389–393, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
389a.pdf.

Anonymous:1902:PLSb

[4680] Anonymous. Provisioning a liner for a single transatlantic trip. Scientific American, 87(24):393, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
393.pdf.

Fawcett:1902:DTG

[4681] Waldon Fawcett. The development of transportation on the Great Lakes. Scientific American, 87(24):394–395, December 13, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n24/pdf/scientificamerican12131902-394.pdf.

Anonymous:1902:DAY

[4682] Anonymous. The development of the auxiliary yacht. Scientific American, 87(24):395–396, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
395.pdf.

Anonymous:1902:ETP

[4683] Anonymous. English turbine-propelled yachts. Scientific American, 87(24):396, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
396.pdf.

Fyfe:1902:ATP

[4684] Heheert C. Fyfe. The application of turbine propulsion to passenger vessels. Scientific American, 87(24):396–397, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n24/pdf/scientificamerican12131902-396a.pdf.

Anonymous:1902:RCBb

[4685] Anonymous. A ride to Chicago and back on the locomotive of the Twentieth Century Limited. Scientific American, 87(24):398–399, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n24/pdf/scientificamerican12131902-398.pdf.

Walker:1902:NHA

[4686] Herberd T. Walker. Notes on the history of the American locomotive. Scientific American, 87(24):399–402, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n24/pdf/scientificamerican12131902-399.pdf.

Anonymous:1902:RSU

[4687] Anonymous. The railroad system of the United States. Scientific American, 87(24):402–404, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
402.pdf.

Adams:1902:BIS

[4688] B. B. Adams. Block and interlocking signals. Scientific American, 87 (24):404–408, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
404.pdf.

Fawcett:1902:CAT

[4689] Waldon Fawcett. Compressed air as a transportation agent. Scientific American, 87(24):408–409, December 13, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n24/pdf/scientificamerican12131902-408.pdf.

Anonymous:1902:VAC

[4690] Anonymous. A 10,000-volt alternating current locomotive for the Berlin–Zossen Experiments. Scientific American, 87(24):410– 411, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
410.pdf.

Anonymous:1902:VRF

[4691] Anonymous. The Valtellina railway — the first standard three-phase road. Scientific American, 87(24):412, December 13, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n24/pdf/scientificamerican12131902-412.pdf.

Anonymous:1902:HSH

[4692] Anonymous. The Hon. Salem H. Wales. Scientific American, 87 (24):412–416, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
412a.pdf.

Anonymous:1902:AB

[4693] Anonymous. Air brakes. Scientific American, 87(24):416–418, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n24/pdf/scientificamerican12131902-416.pdf.

Anonymous:1902:BPWbq

[4694] Anonymous. Business and personal wants. Scientific American, 87 (24):435–438, December 13, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n24/pdf/scientificamerican12131902-
435.pdf.

Anonymous:1902:SNu

[4695] Anonymous. Science notes. Scientific American, 87(25):441, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n25/pdf/scientificamerican12201902-441a.pdf.

Anonymous:1902:Cbf

[4696] Anonymous. Correspondence. Scientific American, 87(25):445, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n25/pdf/scientificamerican12201902-445.pdf.

Anonymous:1902:RPIbr

[4697] Anonymous. Recently patented inventions. Scientific American, 87(25):451, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
451.pdf.

Anonymous:1902:NQbr

[4698] Anonymous. Notes and queries. Scientific American, 87(25): 452, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
452a.pdf.

Anonymous:1902:IIbr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
452b.pdf.

Anonymous:1902:TPN

[4700] Anonymous. The transportation problem in New York City. Scientific American, 87(25):440, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
440.pdf.

Anonymous:1902:PDS

[4701] Anonymous. A proposed duodecimal system of weights and measures. Scientific American, 87(25):440, December 20, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n25/pdf/scientificamerican12201902-440a.pdf.

Anonymous:1902:RJS

[4702] Anonymous. Return of the Jesup Siberian Expedition. Scientific American, 87(25):440, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
440b.pdf.

Anonymous:1902:NTA

[4703] Anonymous. A new type of auxiliary vessel for the British Navy. Scientific American, 87(25):440–441, December 20, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n25/pdf/scientificamerican12201902-440c.pdf.

Anonymous:1902:HMG

[4704] Anonymous. The hall-marking of gold and silver in England. Scientific American, 87(25):441, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
441.pdf.

Anonymous:1902:PCM

[4705] Anonymous. A pneumatic coating machine. Scientific American, 87(25):442, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
442.pdf.

Brown:1902:DLB

[4706] Enos Brown. Dredger for levee building. Scientific American, 87 (25):442, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
442a.pdf.

Anonymous:1902:TIG

[4707] Anonymous. Transporting the 16-inch gun to sandy hook. Scientific American, 87(25):443, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
443.pdf.

Anonymous:1902:NYBc

[4708] Anonymous. New York and Brooklyn Rapid Transit Subway. Scientific American, 87(25):444, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
444.pdf.

Shepstone:1902:TMS

[4709] Harold J. Shepstone. Temple models for the St. Louis Exposition. Scientific American, 87(25):446–447, December 20, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n25/pdf/scientificamerican12201902-446.pdf.

Fyfe:1902:SV

[4710] H. C. Fyfe. A “scientific” violin. Scientific American, 87(25): 447, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
447.pdf.

Anonymous:1902:SDN

[4711] Anonymous. The “Santos-Dumont No. 9”. Scientific American, 87 (25):447–448, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
447a.pdf.

Beasley:1902:EPP

[4712] Walter L. Beasley. The Egyptian paint palettes. Scientific American, 87(25):448, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
448.pdf.

Anonymous:1902:PDj

[4713] Anonymous. Patent department. Scientific American, 87(25): 449, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
449.pdf.

Anonymous:1902:PH

[4714] Anonymous. Pivoted hub. Scientific American, 87(25):449, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n25/pdf/scientificamerican12201902-449a.pdf.

Anonymous:1902:SS

[4715] Anonymous. The “stab” switch. Scientific American, 87(25): 449, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
449b.pdf.

Anonymous:1902:LV

[4716] Anonymous. Lock valve. Scientific American, 87(25):449–450, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n25/pdf/scientificamerican12201902-449c.pdf.

Anonymous:1902:LNm

[4717] Anonymous. Legal notes. Scientific American, 87(25):450, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n25/pdf/scientificamerican12201902-450.pdf.

Anonymous:1902:BPWbr

[4718] Anonymous. Business and personal wants. Scientific American, 87(25):452, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n25/pdf/scientificamerican12201902-
452.pdf.

Anonymous:1902:Pe

[4719] Anonymous. Patents. Scientific American, 87(25):455, December 20, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n25/pdf/scientificamerican12201902-455.pdf.

Anonymous:1902:SNv

[4720] Anonymous. Science notes. Scientific American, 87(26):459– 460, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
459c.pdf.

Anonymous:1902:Cbg

[4721] Anonymous. Correspondence. Scientific American, 87(26):462, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n26/pdf/scientificamerican12271902-462.pdf.

Anonymous:1902:RPIbs

[4722] Anonymous. Recently patented inventions. Scientific American, 87(26):468, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
468.pdf.

Anonymous:1902:NQbs

[4723] Anonymous. Notes and queries. Scientific American, 87(26): 469, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
469a.pdf.

Anonymous:1902:IIbs 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
469b.pdf.

Anonymous:1902:Ic

[4725] Anonymous. Index. Scientific American, 87(26):473–474, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n26/pdf/scientificamerican12271902-473.pdf.

Anonymous:1902:PNPbd

[4726] Anonymous. Patent notes: Patents. Scientific American, 87 (26):475, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
475.pdf.

Anonymous:1902:BALb

[4727] Anonymous. Building American locomotives in British shops. Scientific American, 87(26):458, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
458.pdf.

Anonymous:1902:NWG

[4728] Anonymous. Naval war-game between the United States and Germany. Scientific American, 87(26):458, December 27, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n26/pdf/scientificamerican12271902-458a.pdf.

Anonymous:1902:NRT

[4729] Anonymous. The new railroad and trolley tunnels entering New York. Scientific American, 87(26):458, December 27, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n26/pdf/scientificamerican12271902-458b.pdf.

Anonymous:1902:ASCa

[4730] Anonymous. The Army Signal Corps. Scientific American, 87 (26):459, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
459.pdf.

Anonymous:1902:OS

[4731] Anonymous. To our subscribers. Scientific American, 87(26): 459, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
459a.pdf.

Anonymous:1902:ESP

[4732] Anonymous. European sale of Pupin’s telephone patents. Scientific American, 87(26):459, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
459b.pdf.

Anonymous:1902:GAT

[4733] Anonymous. A German ambulance train. Scientific American, 87 (26):460, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
460.pdf.

Anonymous:1902:MRD

[4734] Anonymous. Machine for removing and disposing of snow. Scientific American, 87(26):461, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
461.pdf.

Dailey:1902:EMR

[4735] Howard B. Dailey. Electro-magnetic rotations. Scientific American, 87(26):461–462, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
461a.pdf.

Anonymous:1902:TAE

[4736] Anonymous. Thomas Alva Edison. Scientific American, 87(26): 463, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
463.pdf.

Hopkins:1902:SP

[4737] George M. Hopkins. Spectral photography. Scientific American, 87(26):464, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
464.pdf.

Anonymous:1902:NAC

[4738] Anonymous. The new armored cruisers “Tennessee” and “Washington”. Scientific American, 87(26):464–466, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n26/pdf/scientificamerican12271902-464a.pdf.

Anonymous:1902:LTF

[4739] Anonymous. Leading types of 1902 French light-weight automobiles. Scientific American, 87(26):466, December 27, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n26/pdf/scientificamerican12271902-466.pdf.

Anonymous:1902:BPWbs

[4740] Anonymous. Business and personal wants. Scientific American, 87(26):469, December 27, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n26/pdf/scientificamerican12271902-
469.pdf.

Anonymous:1902:Cbh

[4741] Anonymous. Correspondence. Scientific American, 87(3):39, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-39.pdf.

Anonymous:1902:LNn

[4742] Anonymous. Legal notes. Scientific American, 87(3):43, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-43.pdf.

Anonymous:1902:RPIbt

[4743] Anonymous. Recently patented inventions. Scientific American, 87(3):44, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
44.pdf.

Anonymous:1902:IIbt

[4744] Anonymous. Index of inventions. Scientific American, 87(3):44, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-44b.pdf.

Anonymous:1902:PNPbe

[4745] Anonymous. Patent notes: Patents. Scientific American, 87(3):47, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-47.pdf.

Anonymous:1902:NQbt

[4746] Anonymous. Notes and queries. Scientific American, 87(3):47, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-47a.pdf.

Anonymous:1902:TIP

[4747] Anonymous. The two important problems of the Panama Canal. Scientific American, 87(3):34, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
34.pdf.

Anonymous:1902:RDS

[4748] Anonymous. The rapid development of the shipbuilding industry of Japan. Scientific American, 87(3):34, July 19, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n3/pdf/scientificamerican07191902-34a.pdf.

Anonymous:1902:PSW

[4749] Anonymous. Photography of sound waves. Scientific American, 87(3):34, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
34b.pdf.

Anonymous:1902:ESD

[4750] Anonymous. Exhibition of scientific devices in Paris. Scientific American, 87(3):34–35, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
34c.pdf.

Benjamin:1902:AAA

[4751] Marcus Benjamin. American Association for the Advancement of Science. Scientific American, 87(3):35, July 19, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n3/pdf/scientificamerican07191902-35.pdf.

Perkins:1902:MST

[4752] Frank C. Perkins. Marine steam turbines. Scientific American, 87(3):36, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-36.pdf.

Perkins:1902:FOB

[4753] Frank C. Perkins. First oil-burning steamship built in the United States. Scientific American, 87(3):36, July 19, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n3/pdf/scientificamerican07191902-36a.pdf.

Anonymous:1902:AHL

[4754] Anonymous. Another “heaviest” locomotive in the world. Scientific American, 87(3):37, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
37.pdf.

Collins:1902:CWT

[4755] A. Frederick Collins. The Collins wireless telephone. Scientific American, 87(3):37–38, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
37a.pdf.

Byrne:1902:TSC

[4756] H. H. Byrne. Through the Suez Canal on a Man-of-War. Scientific American, 87(3):38–39, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
38.pdf.

Jones:1902:MPF

[4757] George J. Jones. Milk in powdered form. Scientific American, 87(3):40–41, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
40.pdf.

Anonymous:1902:NCN

[4758] Anonymous. A new curiosity at the New York Aquarium. Scientific American, 87(3):41, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
41.pdf.

Anonymous:1902:PLSc

[4759] Anonymous. A peculiar lightning-shattered tree. Scientific American, 87(3):41, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
41a.pdf.

Anonymous:1902:PDk

[4760] Anonymous. Patent department. Scientific American, 87(3):42, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-42.pdf.

Anonymous:1902:NTE

[4761] Anonymous. New turbine engine. Scientific American, 87(3):42, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-42a.pdf.

Anonymous:1902:EL

[4762] Anonymous. Electric lamp. Scientific American, 87(3):42, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-42b.pdf.

Anonymous:1902:BSb

[4763] Anonymous. Bicycle stands. Scientific American, 87(3):42, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-42c.pdf.

Anonymous:1902:PPb

[4764] Anonymous. Pipe puller. Scientific American, 87(3):42–43, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-42d.pdf.

Anonymous:1902:BPWbt

[4765] Anonymous. Business and personal wants. Scientific American, 87(3):44, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n3/pdf/scientificamerican07191902-
44a.pdf.

Anonymous:1902:BNc

[4766] Anonymous. Book notices. Scientific American, 87(3):46, July 19, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n3/pdf/scientificamerican07191902-46.pdf.

Anonymous:1902:Cbi

[4767] Anonymous. Correspondence. Scientific American, 87(4):54, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n4/pdf/scientificamerican07261902-54.pdf.

Anonymous:1902:RPIbu

[4768] Anonymous. Recently patented inventions. Scientific American, 87(4):60, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
60.pdf.

Anonymous:1902:IIbu

[4769] Anonymous. Index of inventions. Scientific American, 87(4):61, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n4/pdf/scientificamerican07261902-61.pdf.

Anonymous:1902:BNBbb

[4770] Anonymous. Books: New books, etc. Scientific American, 87(4):62, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n4/pdf/scientificamerican07261902-62.pdf.

Anonymous:1902:PNPbf

[4771] Anonymous. Patent notes: Patents. Scientific American, 87(4):62, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n4/pdf/scientificamerican07261902-62a.pdf.

Anonymous:1902:NQbu

[4772] Anonymous. Notes and queries. Scientific American, 87(4):63, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n4/pdf/scientificamerican07261902-63.pdf.

Anonymous:1902:STH

[4773] Anonymous. The shipping trust and higher rates. Scientific American, 87(4):50, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
50.pdf.

Anonymous:1902:NFE

[4774] Anonymous. The need for fireproof electric cars. Scientific American, 87(4):50, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
50a.pdf.

Anonymous:1902:NCP

[4775] Anonymous. New Cuban Patent and Trade Mark Law. Scientific American, 87(4):50–51, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
50b.pdf.

Anonymous:1902:SLA

[4776] Anonymous. St. Louis airship prizes. Scientific American, 87(4):51, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n4/pdf/scientificamerican07261902-51.pdf.

Anonymous:1902:MIT

[4777] Anonymous. The Meerschaum industry of Turkey. Scientific American, 87(4):51, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
51a.pdf.

Anonymous:1902:SPFa

[4778] Anonymous. Shapira the philological forger. Scientific American, 87(4):51, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
51b.pdf.

Beasley:1902:LTB

[4779] Walter L. Beasley. A long-tailed breed of Japanese fowls. Scientific American, 87(4):52, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
52.pdf.

MCurdy:1902:SSM

[4780] James G. M’Curdy. Salvage of the schooner “Minnie A. Caine”. Scientific American, 87(4):52, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
52a.pdf.

Proctor:1902:PGT

[4781] Mary Proctor. Plans for a great telescope. Scientific American, 87(4):52–53, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
52b.pdf.

Anonymous:1902:TTR

[4782] Anonymous. Tongue-traction for resuscitation of the asphyxiated. Scientific American, 87(4):53, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
53.pdf.

Stonebridge:1902:MCA

[4783] G. E. Stonebridge. The militia and the Colt automatic gun. Scientific American, 87(4):53–54, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
53a.pdf.

Anonymous:1902:MCN

[4784] Anonymous. Making the cables on the new East River Bridge, New York. Scientific American, 87(4):55, July 26, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n4/pdf/scientificamerican07261902-55.pdf.

Rost:1902:RPF

[4785] E. C. Rost. The remains of Pizarro, founder of Peru and conqueror of the Ancient Incas. Scientific American, 87(4):56, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n4/pdf/scientificamerican07261902-56.pdf.

Anonymous:1902:RAC

[4786] Anonymous. The Russian armored cruiser “Gromoboi”. Scientific American, 87(4):56, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
56a.pdf.

Anonymous:1902:TRV

[4787] Anonymous. The Tehuantepec Railroad vs the Isthmian Canal. Scientific American, 87(4):57–58, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
57.pdf.

Beasley:1902:PRS

[4788] Walter L. Beasley. Primitive reindeer saddle and cradle from Northeastern Siberia. Scientific American, 87(4):58, July 26, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n4/pdf/scientificamerican07261902-58.pdf.

Jones:1902:FGS

[4789] F. O. Jones. The formation and geology of the salt deposit. Scientific American, 87(4):59, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
59.pdf.

Anonymous:1902:BPWbu

[4790] Anonymous. Business and personal wants. Scientific American, 87(4):60, July 26, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n4/pdf/scientificamerican07261902-
60a.pdf.

Russell:1902:HAb

[4791] Henry Norris Russell. The heavens in August. Scientific American, 87(5):67, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
67.pdf.

Anonymous:1902:LNo

[4792] Anonymous. Legal notes. Scientific American, 87(5):75, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-75.pdf.

Anonymous:1902:RPIbv

[4793] Anonymous. Recently patented inventions. Scientific American, 87(5):76–77, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
76.pdf.

Anonymous:1902:IIbv

[4794] Anonymous. Index of inventions. Scientific American, 87(5):77, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-77a.pdf.

Anonymous:1902:BNBbc

[4795] Anonymous. Books: New books, etc. Scientific American, 87(5):78, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-78.pdf.

Anonymous:1902:NQbv

[4796] Anonymous. Notes and queries. Scientific American, 87(5):79, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-79.pdf.

Anonymous:1902:PNPbg

[4797] Anonymous. Patent notes: Patents. Scientific American, 87(5):81, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-81.pdf.

Anonymous:1902:CAb

[4798] Anonymous. Commercial Alaska. Scientific American, 87(5):66, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-66.pdf.

Anonymous:1902:SH

[4799] Anonymous. The scarcity of horses. Scientific American, 87(5):66, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-66a.pdf.

Anonymous:1902:SOS

[4800] Anonymous. Some old ships. Scientific American, 87(5):66, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-66b.pdf.

Anonymous:1902:NFR

[4801] Anonymous. A new form of radiation. Scientific American, 87(5):66–67, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-66c.pdf.

Anonymous:1902:SFL

[4802] Anonymous. Sponge fishing in the Levant. Scientific American, 87(5):67, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
67a.pdf.

Anonymous:1902:VE

[4803] Anonymous. Vaporization experiments. Scientific American, 87(5):67, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-67b.pdf.

Perkins:1902:STC

[4804] F. C. Perkins. Swiss traveling crane tools. Scientific American, 87(5):68, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-68.pdf.

Anonymous:1902:NWT

[4805] Anonymous. The new Whipple temperature indicator for use with platinum thermometers. Scientific American, 87(5):69, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-69.pdf.

Fawcett:1902:AL

[4806] Waldon Fawcett. An automatic lightship. Scientific American, 87(5):69–70, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
69a.pdf.

Brown:1902:RLP

[4807] Enos Brown. A redwood lumber plant. Scientific American, 87 (5):70–71, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
70.pdf.

Anonymous:1902:PVAb

[4808] Anonymous. The Paris–Vienna Automobile Race. Scientific American, 87(5):72–73, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
72.pdf.

Anonymous:1902:TAA

[4809] Anonymous. A trip across the Atlantic in a kerosene boat. Scientific American, 87(5):73, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
73.pdf.

Geare:1902:RSF

[4810] R. I. Geare. The remora or sucking-fishes. Scientific American, 87(5):73, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-73a.pdf.

Anonymous:1902:PDl

[4811] Anonymous. Patent department. Scientific American, 87(5):74, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-74.pdf.

Anonymous:1902:SPFb

[4812] Anonymous. Stock and poultry-feeding device. Scientific American, 87(5):74, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
74a.pdf.

Anonymous:1902:ATT

[4813] Anonymous. Anti-telescoping train guard. Scientific American, 87(5):74, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
74b.pdf.

Anonymous:1902:NYAc

[4814] Anonymous. Neck-yoke attachment. Scientific American, 87(5):74–75, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n5/pdf/scientificamerican08021902-74c.pdf.

Anonymous:1902:BPWbv

[4815] Anonymous. Business and personal wants. Scientific American, 87(5):77, August 2, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n5/pdf/scientificamerican08021902-
77.pdf.

Anonymous:1902:SNw

[4816] Anonymous. Science notes. Scientific American, 87(6):85, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-85a.pdf.

Anonymous:1902:Cbj

[4817] Anonymous. Correspondence. Scientific American, 87(6):89, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-89.pdf.

Anonymous:1902:RPIbw

[4818] Anonymous. Recently patented inventions. Scientific American, 87(6):94, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
94.pdf.

Anonymous:1902:IIbw

[4819] Anonymous. Index of inventions. Scientific American, 87(6):95, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-95.pdf.

Anonymous:1902:BNBbd

[4820] Anonymous. Books: New books, etc. Scientific American, 87(6):95, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-95a.pdf.

Anonymous:1902:PNPbh

[4821] Anonymous. Patent notes: Patents. Scientific American, 87(6):97, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-97.pdf.

Anonymous:1902:NQbw

[4822] Anonymous. Notes and queries. Scientific American, 87(6):97, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-97a.pdf.

Anonymous:1902:GEFb

[4823] Anonymous. A great engineering feat completed. Scientific American, 87(6):84, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
84.pdf.

Anonymous:1902:HRE

[4824] Anonymous. The Halifax route to Europe. Scientific American, 87(6):84, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
84a.pdf.

Anonymous:1902:MD

[4825] Anonymous. The Marconi disclaimer. Scientific American, 87(6):84, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-84b.pdf.

Anonymous:1902:FTF

[4826] Anonymous. Future transit facilities of New York City. Scientific American, 87(6):84–85, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
84c.pdf.

Anonymous:1902:STd

[4827] Anonymous. Simplon tunnel. Scientific American, 87(6):85, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-85.pdf.

Price:1902:UR

[4828] C. E. Price. A unique railway. Scientific American, 87(6):86, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-86.pdf.

Anonymous:1902:LFSb

[4829] Anonymous. Launch of the first seven-masted steel schooner. Scientific American, 87(6):87, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
87.pdf.

Rost:1902:HAT

[4830] C. E. Rost. Hints for Americans — transportation in South America. Scientific American, 87(6):87–88, August 9, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n6/pdf/scientificamerican08091902-87a.pdf.

Fyfe:1902:GFD

[4831] Robert C. Fyfe. The Great Floating Docks of Bermuda and Algiers, La. Scientific American, 87(6):88–89, August 9, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n6/pdf/scientificamerican08091902-88.pdf.

Anonymous:1902:NMWb

[4832] Anonymous. The new Marconi Wireless Telegraph Station at Cape Breton. Scientific American, 87(6):90–91, August 9, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n6/pdf/scientificamerican08091902-90.pdf.

Anonymous:1902:VMU

[4833] Anonymous. Varied modern uses of the automobile. Scientific American, 87(6):91–92, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
91.pdf.

Anonymous:1902:PWS

[4834] Anonymous. A pneumatic water supply system. Scientific American, 87(6):93, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
93.pdf.

Anonymous:1902:ISS

[4835] Anonymous. Improvements in spinning spindles. Scientific American, 87(6):93, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
93a.pdf.

Anonymous:1902:BWB

[4836] Anonymous. Bottle-washing brush. Scientific American, 87(6):93, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n6/pdf/scientificamerican08091902-93b.pdf.

Anonymous:1902:BPWbw

[4837] Anonymous. Business and personal wants. Scientific American, 87(6):94, August 9, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n6/pdf/scientificamerican08091902-
94a.pdf.

Anonymous:1902:RPIbx

[4838] Anonymous. Recently patented inventions. Scientific American, 87(7):110, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
110.pdf.

Anonymous:1902:IIbx

[4839] Anonymous. Index of inventions. Scientific American, 87(7):111, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n7/pdf/scientificamerican08161902-111.pdf.

Anonymous:1902:BNBbe

[4840] Anonymous. Books: New books, etc. Scientific American, 87 (7):112, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
112.pdf.

Anonymous:1902:NQbx

[4841] Anonymous. Notes and queries. Scientific American, 87(7):113, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n7/pdf/scientificamerican08161902-113.pdf.

Anonymous:1902:PNPbi

[4842] Anonymous. Patent notes: Patents. Scientific American, 87(7):114, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n7/pdf/scientificamerican08161902-114.pdf.

Anonymous:1902:NAF

[4843] Anonymous. Needless annoyances in ferryboat travel. Scientific American, 87(7):100, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
100.pdf.

Anonymous:1902:NMA

[4844] Anonymous. New method of armor plate making. Scientific American, 87(7):100, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
100a.pdf.

Anonymous:1902:MOW

[4845] Anonymous. Mid ocean wireless telegraph station. Scientific American, 87(7):100, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
100b.pdf.

Anonymous:1902:SFG

[4846] Anonymous. Some further gunnery experiments with the “Belleisle”. Scientific American, 87(7):100–101, August 16, 1902. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
100c.pdf.

Anonymous:1902:HOF

[4847] Anonymous. How ostrich feathers are culled. Scientific American, 87(7):101, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
101.pdf.

Anonymous:1902:MPb

[4848] Anonymous. Military photography. Scientific American, 87(7):101, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n7/pdf/scientificamerican08161902-101a.pdf.

Anonymous:1902:AWU

[4849] Anonymous. The amount of water used in irrigation. Scientific American, 87(7):101, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
101b.pdf.

Anonymous:1902:SAS

[4850] Anonymous. A smokeless automatic stoker. Scientific American, 87(7):102, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
102.pdf.

Anonymous:1902:FSW

[4851] Anonymous. The De Forest system of wireless telegraphy. Scientific American, 87(7):102–103, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
102a.pdf.

Yeigh:1902:GIM

[4852] Frank Yeigh. A group of Indian masks. Scientific American, 87 (7):103–105, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
103.pdf.

Rost:1902:MPE

[4853] E. C. Rost. Mont Pelée in eruption. Scientific American, 87 (7):106–108, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
106.pdf.

Anonymous:1902:PDm

[4854] Anonymous. Patent department. Scientific American, 87(7):108– 109, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
108.pdf.

Anonymous:1902:IGD

[4855] Anonymous. Improved grain door. Scientific American, 87(7):109, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n7/pdf/scientificamerican08161902-109.pdf.

Anonymous:1902:SVT

[4856] Anonymous. System for ventilating tunnels. Scientific American, 87(7):109, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
109a.pdf.

Anonymous:1902:HBc

[4857] Anonymous. Heater and boiler. Scientific American, 87(7):109, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n7/pdf/scientificamerican08161902-109b.pdf.

Anonymous:1902:BPWbx

[4858] Anonymous. Business and personal wants. Scientific American, 87(7):110, August 16, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n7/pdf/scientificamerican08161902-
110a.pdf.

Anonymous:1902:SNx

[4859] Anonymous. Science notes. Scientific American, 87(8):117, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n8/pdf/scientificamerican08231902-117b.pdf.

Anonymous:1902:RPIby

[4860] Anonymous. Recently patented inventions. Scientific American, 87(8):126, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
126.pdf.

Anonymous:1902:IIby

[4861] Anonymous. Index of inventions. Scientific American, 87(8):127, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n8/pdf/scientificamerican08231902-127.pdf.

Anonymous:1902:BNBbf

[4862] Anonymous. Books: New books, etc. Scientific American, 87 (8):128, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
128.pdf.

Anonymous:1902:NQby

[4863] Anonymous. Notes and queries. Scientific American, 87(8):129, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n8/pdf/scientificamerican08231902-129.pdf.

Anonymous:1902:PNPbj

[4864] Anonymous. Patent notes: Patents. Scientific American, 87(8):130, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n8/pdf/scientificamerican08231902-130.pdf.

Anonymous:1902:DCF

[4865] Anonymous. Dynamite cruiser fiasco. Scientific American, 87 (8):116, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
116.pdf.

Anonymous:1902:IBB

[4866] Anonymous. The inspection of the Brooklyn Bridge. Scientific American, 87(8):116, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
116a.pdf.

Anonymous:1902:AOW

[4867] Anonymous. The Armstrong Orling wireless telegraph. Scientific American, 87(8):116, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
116b.pdf.

Anonymous:1902:ETEb

[4868] Anonymous. Electric traction on English suburban roads. Scientific American, 87(8):116, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
116c.pdf.

Anonymous:1902:GB

[4869] Anonymous. The greatest of battleships. Scientific American, 87(8):116, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
116d.pdf.

Anonymous:1902:FWT

[4870] Anonymous. Fessenden wireless telegraph patents allowed. Scientific American, 87(8):116, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
116e.pdf.

Walsh:1902:PAP

[4871] George Ethelbert Walsh. The passing of American plagues. Scientific American, 87(8):117, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
117.pdf.

Anonymous:1902:RBM

[4872] Anonymous. Robert Bach M’Master. Scientific American, 87 (8):117, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
117a.pdf.

Dawley:1902:SRA

[4873] Thomas R. Dawley. Some recent archaeological discoveries in Mexico City. Scientific American, 87(8):118–119, August 23, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n8/pdf/scientificamerican08231902-118.pdf.

Anonymous:1902:IPP

[4874] Anonymous. Improved picture projecting apparatus. Scientific American, 87(8):119, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
119.pdf.

Anonymous:1902:MLT

[4875] Anonymous. Moving large trees. Scientific American, 87(8):119, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n8/pdf/scientificamerican08231902-119a.pdf.

Anonymous:1902:SRTb

[4876] Anonymous. Solid rubber tire setting machine. Scientific American, 87(8):119–120, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
119b.pdf.

Collins:1902:ERRa

[4877] A. Frederick Collins. Electrical resonance and its relation to syntonic wireless telegraphy — I. Scientific American, 87(8):120–121, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n8/pdf/scientificamerican08231902-120.pdf.

Anonymous:1902:BALc

[4878] Anonymous. The building of American locomotives — II. Scientific American, 87(8):121, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
121.pdf.

Willey:1902:HGC

[4879] Day Allen Willey. How golf clubs are made in large quantities. Scientific American, 87(8):122–123, August 23, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n8/pdf/scientificamerican08231902-122.pdf.

Anonymous:1902:BCL

[4880] Anonymous. The battleships “Connecticut” and “Louisiana”. Scientific American, 87(8):123–124, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
123.pdf.

Anonymous:1902:PAW

[4881] Anonymous. Power from an artesian well. Scientific American, 87(8):124, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
124.pdf.

Anonymous:1902:MPR

[4882] Anonymous. Motor-paced race on a cycle whirl. Scientific American, 87(8):125, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
125.pdf.

Anonymous:1902:BOG

[4883] Anonymous. Burning oil gusher in Louisiana. Scientific American, 87(8):125, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
125a.pdf.

Anonymous:1902:BPWby

[4884] Anonymous. Business and personal wants. Scientific American, 87(8):126, August 23, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n8/pdf/scientificamerican08231902-
126a.pdf.

Anonymous:1902:SNy

[4885] Anonymous. Science notes. Scientific American, 87(9):133, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n9/pdf/scientificamerican08301902-133c.pdf.

Anonymous:1902:RPIbz

[4886] Anonymous. Recently patented inventions. Scientific American, 87(9):142, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
142.pdf.

Anonymous:1902:IIbz

[4887] Anonymous. Index of inventions. Scientific American, 87(9):142, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n9/pdf/scientificamerican08301902-142b.pdf.

Anonymous:1902:BNBbg

[4888] Anonymous. Books: New books, etc. Scientific American, 87 (9):144, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
144.pdf.

Anonymous:1902:PNPbk

[4889] Anonymous. Patent notes: Patents. Scientific American, 87(9):144, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n9/pdf/scientificamerican08301902-144a.pdf.

Anonymous:1902:NQbz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n9/pdf/scientificamerican08301902-145.pdf.

Anonymous:1902:ETF

[4891] Anonymous. Express trains of the future. Scientific American, 87(9):132, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
132.pdf.

Anonymous:1902:LLF

[4892] Anonymous. Light on the liquid fuel question. Scientific American, 87(9):132, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
132a.pdf.

Anonymous:1902:RBE

[4893] Anonymous. The recent “Belleisle” experiments. Scientific American, 87(9):132–133, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
132b.pdf.

Anonymous:1902:GMH

[4894] Anonymous. George M. Hopkins. Scientific American, 87(9):133, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n9/pdf/scientificamerican08301902-133.pdf.

Anonymous:1902:LPC

[4895] Anonymous. The laying of a Pacific cable. Scientific American, 87(9):133, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
133a.pdf.

Anonymous:1902:EH

[4896] Anonymous. Electrified houses. Scientific American, 87(9):133, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n9/pdf/scientificamerican08301902-133b.pdf.

Anonymous:1902:MCH

[4897] Anonymous. A mammoth combined harvester. Scientific American, 87(9):134, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
134.pdf.

Hopkins:1902:SMN

[4898] George M. Hopkins. Some modifications of the normal photograph. Scientific American, 87(9):135, August 30, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n9/pdf/scientificamerican08301902-135.pdf.

Anonymous:1902:CNR

[4899] Anonymous. The Coronation Naval Review. Scientific American, 87(9):135–136, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
135a.pdf.

Collins:1902:ERRb

[4900] A. Frederick Collins. Electrical resonance and its relation to syntonic wireless telegraphy — II. Scientific American, 87(9):136–137, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v87/n9/pdf/scientificamerican08301902-136.pdf.

Brown:1902:SFP

[4901] Enos Brown. The Snoqualmie Falls Power Plant. Scientific American, 87(9):137, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
137.pdf.

Holder:1902:REF

[4902] Charles F. Holder. Remarkable engineering feats in railroad work. Scientific American, 87(9):138, August 30, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n9/pdf/scientificamerican08301902-138.pdf.

Moore:1902:USN

[4903] Frederick Moore. The United States Naval Observatory at Washington. Scientific American, 87(9):139–140, August 30, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n9/pdf/scientificamerican08301902-139.pdf.

Anonymous:1902:NSCb

[4904] Anonymous. A new Schneider–Canet Naval gun. Scientific American, 87(9):140, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
140.pdf.

Anonymous:1902:AC

[4905] Anonymous. The Ardennes Circuit. Scientific American, 87(9): 140–141, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
140a.pdf.

Anonymous:1902:FDE

[4906] Anonymous. Formation of the diamond by the electric furnace. Scientific American, 87(9):141, August 30, 1902. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v87/
n9/pdf/scientificamerican08301902-141.pdf.

Anonymous:1902:BPWbz

[4907] Anonymous. Business and personal wants. Scientific American, 87(9):142, August 30, 1902. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v87/n9/pdf/scientificamerican08301902-
142a.pdf.

Anonymous:1903:Ca

[4908] Anonymous. Correspondence. Scientific American, 88(1):7, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n1/pdf/scientificamerican01031903-7.pdf.

Russell:1903:HJa

[4909] Henry Norris Russell. The heavens in January. Scientific American, 88(1):7, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
7a.pdf.

Anonymous:1903:RPIa

[4910] Anonymous. Recently patented inventions. Scientific American, 88(1):12, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
12.pdf.

Anonymous:1903:IIa

[4911] Anonymous. Index of inventions. Scientific American, 88(1):12–15, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n1/pdf/scientificamerican01031903-12b.pdf.

Anonymous:1903:PNPa

[4912] Anonymous. Patent notes: Patents. Scientific American, 88(1):16, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n1/pdf/scientificamerican01031903-16.pdf.

Anonymous:1903:RY

[4913] Anonymous. Retrospect of the year 1902. Scientific American, 88(1):2–3, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
2.pdf.

Townsend:1903:BCK

[4914] Irving U. Townsend. The Barber–Colman knotter. Scientific American, 88(1):4, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
4.pdf.

Anonymous:1903:BD

[4915] Anonymous. Babbitting device. Scientific American, 88(1):5, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n1/pdf/scientificamerican01031903-5.pdf.

Collins:1903:FWT

[4916] A. Frederick Collins. The Fessenden wireless telegraph system. Scientific American, 88(1):5–6, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
5a.pdf.

Perkins:1903:ETW

[4917] Frank C. Perkins. Electric trolley wagons and omnibuses. Scientific American, 88(1):6, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
6.pdf.

Fawcett:1903:WOT

[4918] Waldon Fawcett. Winter obstacles to transportation. Scientific American, 88(1):8–9, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
8.pdf.

Brown:1903:ARP

[4919] Enos Brown. The ancient ruins of Palenke. Scientific American, 88(1):9–10, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
9.pdf.

Anonymous:1903:CMG

[4920] Anonymous. A curious mushroom growth. Scientific American, 88(1):10, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
10.pdf.

Anonymous:1903:WPP

[4921] Anonymous. Wood paving in Paris. Scientific American, 88(1):11, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n1/pdf/scientificamerican01031903-11.pdf.

Anonymous:1903:BPWa

[4922] Anonymous. Business and personal wants. Scientific American, 88(1):12, January 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n1/pdf/scientificamerican01031903-
12a.pdf.

Anonymous:1903:RPIb

[4923] Anonymous. Recently patented inventions. Scientific American, 88(10):178, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
178.pdf.

Anonymous:1903:BPWb

[4924] Anonymous. Business and personal wants. Scientific American, 88(10):178, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
178a.pdf.

Anonymous:1903:NQa

[4925] Anonymous. Notes and queries. Scientific American, 88(10):178–179, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-178b.pdf.

Anonymous:1903:IIb

[4926] Anonymous. Index of inventions. Scientific American, 88(10):179–183, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-179.pdf.

Anonymous:1903:BSA

[4927] Anonymous. Block signals and the automatic stop. Scientific American, 88(10):168, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
168.pdf.

Anonymous:1903:DA

[4928] Anonymous. Diverging again. Scientific American, 88(10):168, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-168a.pdf.

Anonymous:1903:EFB

[4929] Anonymous. The effect of fire on bridge cables. Scientific American, 88(10):168, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
168b.pdf.

Anonymous:1903:NCa

[4930] Anonymous. The new Cunarders. Scientific American, 88(10):168, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-168c.pdf.

Anonymous:1903:DCW

[4931] Anonymous. Delay in the construction of warships. Scientific American, 88(10):168, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
168d.pdf.

Anonymous:1903:EIF

[4932] Anonymous. Extraordinary immigration figures. Scientific American, 88(10):168, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
168e.pdf.

Anonymous:1903:SAI

[4933] Anonymous. The Scientific American and the introduction of the telephone into Europe. Scientific American, 88(10):168, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-168f.pdf.

Anonymous:1903:DDR

[4934] Anonymous. The death of Dr. Richard J. Gatling. Scientific American, 88(10):169, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
169.pdf.

Anonymous:1903:DFS

[4935] Anonymous. To detect fire at sea. Scientific American, 88(10):169, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-169a.pdf.

Anonymous:1903:AUS

[4936] Anonymous. Amendment to the United States Patent Statutes. Scientific American, 88(10):169, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
169b.pdf.

Anonymous:1903:SNa

[4937] Anonymous. Science notes. Scientific American, 88(10):169, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-169c.pdf.

Anonymous:1903:NDS

[4938] Anonymous. A new departure in ship logs. Scientific American, 88(10):170, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
170.pdf.

Willey:1903:SWC

[4939] Day Allen Willey. Self-winding clock. Scientific American, 88(10):171, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-171.pdf.

Collins:1903:ABW

[4940] T. Byard Collins. The action of a bird’s wing and its bearing on the problem of mechanical flight. Scientific American, 88(10):171–172, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-171a.pdf.

Anonymous:1903:ELY

[4941] Anonymous. Electric locomotives for yard and shop work. Scientific American, 88(10):172, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
172.pdf.

Anonymous:1903:ETF

[4942] Anonymous. Extension of transit facilities in New York. Scientific American, 88(10):173, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
173.pdf.

Fyfe:1903:LHC

[4943] Herbert C. Fyfe. The latest Himalaya-Climbing Expedition. Scientific American, 88(10):174–176, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
174.pdf.

Geare:1903:VSIa

[4944] Randolph I. Geare. Venomous serpents — IV. Scientific American, 88(10):176, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
176.pdf.

Anonymous:1903:AH

[4945] Anonymous. Artificial horizon. Scientific American, 88(10):177, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-177.pdf.

Anonymous:1903:NAC

[4946] Anonymous. A new acid chamber for sulphite works. Scientific American, 88(10):177, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
177a.pdf.

Anonymous:1903:SCG

[4947] Anonymous. Seed-cleaner for grain separators. Scientific American, 88(10):177, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n10/pdf/scientificamerican03071903-
177b.pdf.

Anonymous:1903:MT

[4948] Anonymous. Mechanical toy. Scientific American, 88(10):177, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-177c.pdf.

Anonymous:1903:NBEa

[4949] Anonymous. New books, etc. Scientific American, 88(10):183–184, March 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n10/pdf/scientificamerican03071903-183.pdf.

Anonymous:1903:SNb

[4950] Anonymous. Science notes. Scientific American, 88(11):187–188, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-187b.pdf.

Anonymous:1903:Cb

[4951] Anonymous. Correspondence. Scientific American, 88(11):191, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-191.pdf.

Anonymous:1903:RPIc

[4952] Anonymous. Recently patented inventions. Scientific American, 88(11):196, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
196.pdf.

Anonymous:1903:NQb

[4953] Anonymous. Notes and queries. Scientific American, 88(11):196, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-196c.pdf.

Anonymous:1903:IIc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-196d.pdf.

Anonymous:1903:PNPb

[4955] Anonymous. Patent notes: Patents. Scientific American, 88(11):199, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-199.pdf.

Anonymous:1903:TPQ

[4956] Anonymous. Trying to put a quart into a pint cup. Scientific American, 88(11):186, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
186.pdf.

Anonymous:1903:EAU

[4957] Anonymous. Enactment of the amendment of United States Patent Statutes. Scientific American, 88(11):186, March 14, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n11/pdf/scientificamerican03141903-186a.pdf.

Anonymous:1903:OAC

[4958] Anonymous. Our agricultural colleges. Scientific American, 88 (11):186–187, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
186b.pdf.

Anonymous:1903:AGU

[4959] Anonymous. Action of gelatine upon glass. Scientific American, 88(11):187, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
187.pdf.

Anonymous:1903:NEM

[4960] Anonymous. New electromagnet for magneto-optic work. Scientific American, 88(11):187, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
187a.pdf.

Jones:1903:NPF

[4961] H. D. Jones. The new Philadelphia filtration system. Scientific American, 88(11):188–189, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
188.pdf.

Anonymous:1903:LER

[4962] Anonymous. The Locke electric rock drill. Scientific American, 88(11):189, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
189.pdf.

Bennet:1903:GWT

[4963] H. J. Bennet. The Gorter Water Tower. Scientific American, 88 (11):189–190, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
189a.pdf.

Anonymous:1903:WTB

[4964] Anonymous. Wireless telegraphy between Guadeloupe and Martinique. Scientific American, 88(11):190, March 14, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n11/pdf/scientificamerican03141903-190.pdf.

Anonymous:1903:MDI

[4965] Anonymous. A motor-driven ice-boat. Scientific American, 88(11):190, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-190a.pdf.

Anonymous:1903:MBM

[4966] Anonymous. Mechanical brick-making. Scientific American, 88(11):190, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-190b.pdf.

Anonymous:1903:CCRa

[4967] Anonymous. On the curious case of regelation. Scientific American, 88(11):191, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
191a.pdf.

Anonymous:1903:CTA

[4968] Anonymous. The cause of thunder again. Scientific American, 88(11):191, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
191b.pdf.

Anonymous:1903:MPEa

[4969] Anonymous. The Mills Peruvian Expedition for the study of solar motion. Scientific American, 88(11):191, March 14, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n11/pdf/scientificamerican03141903-191c.pdf.

Anonymous:1903:SAW

[4970] Anonymous. Still another wireless telegraphy system. Scientific American, 88(11):191, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
191d.pdf.

Anonymous:1903:CSa

[4971] Anonymous. The current supplement. Scientific American, 88 (11):191–192, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
191e.pdf.

Anonymous:1903:CFNa

[4972] Anonymous. Completion of the floor of the new East River Bridge. Scientific American, 88(11):192–193, March 14, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n11/pdf/scientificamerican03141903-192.pdf.

Anonymous:1903:HPD

[4973] Anonymous. The Hunter’s Point Dry Dock, San Francisco. Scientific American, 88(11):193–194, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
193.pdf.

Anonymous:1903:RFP

[4974] Anonymous. The restoration of the foundations of the philæ temples. Scientific American, 88(11):194–195, March 14, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n11/pdf/scientificamerican03141903-194.pdf.

Anonymous:1903:LWR

[4975] Anonymous. Lloyd’s wreck returns. Scientific American, 88(11):195, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-195.pdf.

Anonymous:1903:NFS

[4976] Anonymous. New French submarine boat. Scientific American, 88(11):195, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
195a.pdf.

Anonymous:1903:AFN

[4977] Anonymous. An armchair formed by natural growth. Scientific American, 88(11):195, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
195b.pdf.

Beasley:1903:LFW

[4978] Walter L. Beasley. The largest finback whale. Scientific American, 88(11):195–196, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
195c.pdf.

Anonymous:1903:MI

[4979] Anonymous. Miscellaneous invention. Scientific American, 88(11):196, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-196a.pdf.

Anonymous:1903:BPWc

[4980] Anonymous. Business and personal wants. Scientific American, 88(11):196, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n11/pdf/scientificamerican03141903-
196b.pdf.

Anonymous:1903:NBEb

[4981] Anonymous. New books, etc. Scientific American, 88(11):199–200, March 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n11/pdf/scientificamerican03141903-199a.pdf.

Anonymous:1903:SNc

[4982] Anonymous. Science notes. Scientific American, 88(12):203, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n12/pdf/scientificamerican03211903-203b.pdf.

Anonymous:1903:Cc

[4983] Anonymous. Correspondence. Scientific American, 88(12):207, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n12/pdf/scientificamerican03211903-207.pdf.

Anonymous:1903:RPId

[4984] Anonymous. Recently patented inventions. Scientific American, 88(12):212, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
212.pdf.

Anonymous:1903:NQc

[4985] Anonymous. Notes and queries. Scientific American, 88(12):213–214, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n12/pdf/scientificamerican03211903-213.pdf.

Anonymous:1903:IId

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n12/pdf/scientificamerican03211903-214.pdf.

Anonymous:1903:PNPc

[4987] Anonymous. Patent notes: Patents. Scientific American, 88(12):217, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n12/pdf/scientificamerican03211903-217.pdf.

Anonymous:1903:BNBa

[4988] Anonymous. Books: New books, etc. Scientific American, 88(12):217, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n12/pdf/scientificamerican03211903-217a.pdf.

Anonymous:1903:HPT

[4989] Anonymous. A 75,000 horse power turbine plant. Scientific American, 88(12):202, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
202.pdf.

Anonymous:1903:PSF

[4990] Anonymous. Permanence of steel framework in tall buildings. Scientific American, 88(12):202, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
202a.pdf.

Anonymous:1903:CCD

[4991] Anonymous. The changes in the Croton Dam. Scientific American, 88(12):202, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
202b.pdf.

Anonymous:1903:WCP

[4992] Anonymous. Warship construction in private and Government shipyards. Scientific American, 88(12):202, March 21, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n12/pdf/scientificamerican03211903-202c.pdf.

Anonymous:1903:SI

[4993] Anonymous. “Shamrock II”. Scientific American, 88(12):202–203, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n12/pdf/scientificamerican03211903-202d.pdf.

Anonymous:1903:SER

[4994] Anonymous. The Second East River Rapid Transit Tunnel to Brooklyn. Scientific American, 88(12):203, March 21, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n12/pdf/scientificamerican03211903-203.pdf.

Atkinson:1903:SPO

[4995] A. S. Atkinson. Spices — pure and otherwise. Scientific American, 88(12):203, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
203a.pdf.

Anonymous:1903:MGE

[4996] Anonymous. The Montreal grain elevator. Scientific American, 88(12):204, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
204.pdf.

Collins:1903:BSS

[4997] A. Frederick Collins. Bull’s selective system of wireless telegraphy. Scientific American, 88(12):205–206, March 21, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n12/pdf/scientificamerican03211903-205.pdf.

Willey:1903:PTS

[4998] Day Allen Willey. The pneumatic tube system of a modern department store. Scientific American, 88(12):206–207, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n12/pdf/scientificamerican03211903-206.pdf.

Anonymous:1903:CEB

[4999] Anonymous. A curious elliptical bridge. Scientific American, 88(12):208, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
208.pdf.

Holder:1903:WOG

[5000] Charles F. Holder. Where olives grow. Scientific American, 88 (12):209–210, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
209.pdf.

Beasley:1903:CJC

[5001] Walter L. Beasley. Curious jade carvings from Mexico. Scientific American, 88(12):211, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
211.pdf.

Anonymous:1903:BPWd

[5002] Anonymous. Business and personal wants. Scientific American, 88(12):212, March 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n12/pdf/scientificamerican03211903-
212a.pdf.

Anonymous:1903:SNd

[5003] Anonymous. Science notes. Scientific American, 88(13):221, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-221b.pdf.

Anonymous:1903:Cd

[5004] Anonymous. Correspondence. Scientific American, 88(13):225, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-225.pdf.

Anonymous:1903:LNa

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-229.pdf.

Anonymous:1903:RPIe

[5006] Anonymous. Recently patented inventions. Scientific American, 88(13):231, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n13/pdf/scientificamerican03281903-
231.pdf.

Anonymous:1903:BPWe

[5007] Anonymous. Business and personal wants. Scientific American, 88(13):231, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n13/pdf/scientificamerican03281903-
231a.pdf.

Anonymous:1903:NQd

[5008] Anonymous. Notes and queries. Scientific American, 88(13):232, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-232.pdf.

Anonymous:1903:IIe

[5009] Anonymous. Index of inventions. Scientific American, 88(13):232–235, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-232a.pdf.

Anonymous:1903:PNPd

[5010] Anonymous. Patent notes: Patents. Scientific American, 88(13):235, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-235.pdf.

Anonymous:1903:BNBb

[5011] Anonymous. Books: New books, etc. Scientific American, 88 (13):235–236, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n13/pdf/scientificamerican03281903-
235a.pdf.

Anonymous:1903:RPC

[5012] Anonymous. Ratification of the Panama Canal Treaty. Scientific American, 88(13):220, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n13/pdf/scientificamerican03281903-
220.pdf.

Anonymous:1903:ST

[5013] Anonymous. A sign of the times. Scientific American, 88(13):220, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-220a.pdf.

Adams:1903:RWP

[5014] Alton D. Adams. Rise of water power in electrical supply. Scientific American, 88(13):220, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n13/pdf/scientificamerican03281903-
220b.pdf.

Anonymous:1903:LES

[5015] Anonymous. The latest Edison storage battery patent. Scientific American, 88(13):221, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n13/pdf/scientificamerican03281903-
221.pdf.

Russell:1903:HAa

[5016] Henry Norris Russell. The heavens in April. Scientific American, 88(13):221, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n13/pdf/scientificamerican03281903-
221a.pdf.

Trowbridge:1903:VMQ

[5017] John Trowbridge. Vessels made from quartz and some of their uses. Scientific American, 88(13):222, March 28, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n13/pdf/scientificamerican03281903-222.pdf.

Anonymous:1903:MSC

[5018] Anonymous. A mechanical street cleaner. Scientific American, 88 (13):222–223, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n13/pdf/scientificamerican03281903-
222a.pdf.

Perkins:1903:EMM

[5019] Frank C. Perkins. Electricity in modern mining drainage installations. Scientific American, 88(13):223–224, March 28, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n13/pdf/scientificamerican03281903-223.pdf.

Moore:1903:ARS

[5020] Frederick Moore. American railway slaughters and British railway safety. Scientific American, 88(13):226–227, March 28, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n13/pdf/scientificamerican03281903-226.pdf.

Anonymous:1903:NSG

[5021] Anonymous. A new sighting gear for Naval guns and gunnery signaling apparatus. Scientific American, 88(13):227–228, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-227.pdf.

Anonymous:1903:SSM

[5022] Anonymous. Self-sealing moth bag. Scientific American, 88(13):230, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-230.pdf.

Anonymous:1903:SEa

[5023] Anonymous. Safety envelop. Scientific American, 88(13):230, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-230a.pdf.

Anonymous:1903:SCa

[5024] Anonymous. Student’s chair. Scientific American, 88(13):230, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-230b.pdf.

Anonymous:1903:SCSa

[5025] Anonymous. Safety car step. Scientific American, 88(13):230, March 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n13/pdf/scientificamerican03281903-230c.pdf.

Anonymous:1903:RPIf

[5026] Anonymous. Recently patented inventions. Scientific American, 88(14):248, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
248.pdf.

Anonymous:1903:MNM

[5027] Anonymous. Miscellaneous notes: Miscellaneous. Scientific American, 88(14):248, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
248b.pdf.

Anonymous:1903:BPWf

[5028] Anonymous. Business and personal wants. Scientific American, 88(14):248, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
248c.pdf.

Anonymous:1903:IIf

[5029] Anonymous. Index of inventions. Scientific American, 88(14):248–252, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-248d.pdf.

Anonymous:1903:BAR

[5030] Anonymous. The blood of all races. Scientific American, 88(14):238, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-238.pdf.

Anonymous:1903:FLG

[5031] Anonymous. Final lessons of the German American war game. Scientific American, 88(14):238, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
238a.pdf.

Anonymous:1903:VT

[5032] Anonymous. Vertical transportation. Scientific American, 88 (14):238–239, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
238b.pdf.

Anonymous:1903:BBCa

[5033] Anonymous. The British Battleship Construction Programme for 1903. Scientific American, 88(14):239, April 4, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n14/pdf/scientificamerican04041903-239.pdf.

Anonymous:1903:VTE

[5034] Anonymous. Vulcanized timber in England. Scientific American, 88(14):239, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
239a.pdf.

Anonymous:1903:Ce

[5035] Anonymous. “callitypy”. Scientific American, 88(14):239–240, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-239b.pdf.

Anonymous:1903:SMA

[5036] Anonymous. Some modern appliances for life saving at fires. Scientific American, 88(14):240–241, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
240.pdf.

Dunlap:1903:RDN

[5037] Orrin E. Dunlap. Remarkable diversion of Niagara’s waters. Scientific American, 88(14):241–242, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
241.pdf.

Fawcett:1903:ETW

[5038] Waldon Fawcett. Electric track welding. Scientific American, 88 (14):242, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
242.pdf.

Anonymous:1903:ATS

[5039] Anonymous. The advantages of in-turning screws on warships. Scientific American, 88(14):243, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
243.pdf.

Anonymous:1903:NLT

[5040] Anonymous. “Narragansett” — the largest of tank ships. Scientific American, 88(14):243, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
243a.pdf.

Anonymous:1903:PJT

[5041] Anonymous. Pupin and Jentsch on trans-oceanic telephony. Scientific American, 88(14):243, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
243b.pdf.

Anonymous:1903:FCE

[5042] Anonymous. Further Curie experiments with radium. Scientific American, 88(14):243, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
243c.pdf.

Anonymous:1903:NBA

[5043] Anonymous. News from the British Antarctic Expedition — a new record. Scientific American, 88(14):243, April 4, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n14/pdf/scientificamerican04041903-243d.pdf.

Anonymous:1903:VDF

[5044] Anonymous. Volcanic dust falls in Georgia. Scientific American, 88(14):243, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
243e.pdf.

Anonymous:1903:SPA

[5045] Anonymous. The selden patent and the automobile trust. Scientific American, 88(14):243, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
243f.pdf.

Anonymous:1903:NSa

[5046] Anonymous. A new star. Scientific American, 88(14):243, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-243g.pdf.

Anonymous:1903:NBH

[5047] Anonymous. The new Buffalo Harbor Breakwater. Scientific American, 88(14):244–245, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
244.pdf.

Watson:1903:PEN

[5048] W. F. Watson. Photographic experiment with nature’s lenses. Scientific American, 88(14):245–246, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
245.pdf.

Anonymous:1903:DTT

[5049] Anonymous. Discovery of the Tomb of Thothmes IV. Scientific American, 88(14):246, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
246.pdf.

Anonymous:1903:NWS

[5050] Anonymous. Next week’s special automobile and yachting number. Scientific American, 88(14):246, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
246a.pdf.

Anonymous:1903:PDa

[5051] Anonymous. Patent department. Scientific American, 88(14):247, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-247.pdf.

Anonymous:1903:AT

[5052] Anonymous. Automatic toy. Scientific American, 88(14):247, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-247a.pdf.

Anonymous:1903:SHF

[5053] Anonymous. Safety heel for footwear. Scientific American, 88(14):247, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-247b.pdf.

Anonymous:1903:DME

[5054] Anonymous. Diaphragm meter and exerciser. Scientific American, 88(14):247, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n14/pdf/scientificamerican04041903-
247c.pdf.

Anonymous:1903:WI

[5055] Anonymous. Women inventors. Scientific American, 88(14):247–248, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-247d.pdf.

Anonymous:1903:RI

[5056] Anonymous. Railway improvements. Scientific American, 88(14):248, April 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n14/pdf/scientificamerican04041903-248a.pdf.

Anonymous:1903:PNPe

[5057] Anonymous. Patent notes: Patent. Scientific American, 88(15):275, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-275.pdf.

Anonymous:1903:NQe

[5058] Anonymous. Notes and queries. Scientific American, 88(15):279–280, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-279.pdf.

Anonymous:1903:BPWg

[5059] Anonymous. Business and personal wants. Scientific American, 88(15):281, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
281.pdf.

Anonymous:1903:BNBc

[5060] Anonymous. Books: New books, etc. Scientific American, 88(15):282, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-282.pdf.

Anonymous:1903:RPIg

[5061] Anonymous. Recently patented inventions. Scientific American, 88(15):282–287, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
282a.pdf.

Anonymous:1903:IIg

[5062] Anonymous. Index of inventions. Scientific American, 88(15):287–292, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-287.pdf.

Anonymous:1903:NCY

[5063] Anonymous. The new Cup yacht “Reliance”. Scientific American, 88(15):254, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
254.pdf.

Anonymous:1903:NMR

[5064] Anonymous. The new measurement rule of the New York Yacht Club. Scientific American, 88(15):254–255, April 11, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n15/pdf/scientificamerican04111903-254a.pdf.

Anonymous:1903:HNY

[5065] Anonymous. The home of the New York Yacht Club. Scientific American, 88(15):255–257, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
255.pdf.

Anonymous:1903:GFY

[5066] Anonymous. A group of famous yacht models. Scientific American, 88(15):257–258, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
257.pdf.

Anonymous:1903:ETY

[5067] Anonymous. The “Emerald” turbine yacht. Scientific American, 88(15):258, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
258.pdf.

Anonymous:1903:SYN

[5068] Anonymous. The steam yacht “North Star”. Scientific American, 88(15):258–259, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
258a.pdf.

Anonymous:1903:GLC

[5069] Anonymous. A gasoline launch for cruising. Scientific American, 88(15):259, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
259.pdf.

Anonymous:1903:NTM

[5070] Anonymous. New three-masted auxiliary schooner. Scientific American, 88(15):259–260, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
259a.pdf.

Anonymous:1903:LETa

[5071] Anonymous. The launch and early trials of “Shamrock III”. Scientific American, 88(15):260–261, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
260.pdf.

Anonymous:1903:FAP

[5072] Anonymous. The foreign automobiles of prominent Americans. Scientific American, 88(15):261, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
261.pdf.

Anonymous:1903:SMb

[5073] Anonymous. Some 1903 models. Scientific American, 88(15):262–263, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-262.pdf.

Anonymous:1903:LGC

[5074] Anonymous. The locomobile gasoline car. Scientific American, 88 (15):264–265, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
264.pdf.

Anonymous:1903:CGTa

[5075] Anonymous. The Columbia gasoline tonneau. Scientific American, 88(15):265–266, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
265.pdf.

Anonymous:1903:LTM

[5076] Anonymous. The latest types of motor bicycles. Scientific American, 88(15):266–267, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
266.pdf.

Anonymous:1903:NTDa

[5077] Anonymous. A novel transmission device for gasoline automobiles. Scientific American, 88(15):267, April 11, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n15/pdf/scientificamerican04111903-267.pdf.

Fyfe:1903:AW

[5078] Herbert C. Fyfe. The automobile in warfare. Scientific American, 88(15):268–272, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
268.pdf.

Anonymous:1903:CGR

[5079] Anonymous. The Cadillac gasoline runabout. Scientific American, 88(15):272, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
272.pdf.

Anonymous:1903:AAC

[5080] Anonymous. Automobile air cushions. Scientific American, 88 (15):272–273, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
272a.pdf.

Anonymous:1903:NSL

[5081] Anonymous. A novel snow locomotive. Scientific American, 88(15):274, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-274.pdf.

Willey:1903:NS

[5082] Day Allen Willey. A new sport. Scientific American, 88(15):276, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-276.pdf.

Hopkins:1903:AHP

[5083] N. Monroe Hopkins. The automobile and its horse power: a ready method of determining the latter experimentally. Scientific American, 88(15):276–277, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
276a.pdf.

Anonymous:1903:AN

[5084] Anonymous. Automobile novelties. Scientific American, 88(15):278, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-278.pdf.

Anonymous:1903:NDT

[5085] Anonymous. A new detachable tire. Scientific American, 88(15):278, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-278a.pdf.

Anonymous:1903:DPG

[5086] Anonymous. A double piston gasoline motor. Scientific American, 88(15):278, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
278b.pdf.

Anonymous:1903:VAS

[5087] Anonymous. A valueless automobile steam engine. Scientific American, 88(15):278, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
278c.pdf.

Anonymous:1903:DIS

[5088] Anonymous. A device for indicating the speed rate of automobiles. Scientific American, 88(15):278, April 11, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n15/pdf/scientificamerican04111903-278d.pdf.

Anonymous:1903:LBAa

[5089] Anonymous. The lamp that beats Aladdin’s. Scientific American, 88(15):280, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
280.pdf.

Anonymous:1903:RNS

[5090] Anonymous. The Roche new standard pole plug. Scientific American, 88(15):280, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n15/pdf/scientificamerican04111903-
280a.pdf.

Anonymous:1903:HA

[5091] Anonymous. A home attraction. Scientific American, 88(15):280, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-280b.pdf.

Anonymous:1903:EA

[5092] Anonymous. An efficient automobile. Scientific American, 88(15):280, April 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n15/pdf/scientificamerican04111903-280c.pdf.

Anonymous:1903:SNe

[5093] Anonymous. Science notes. Scientific American, 88(16):295, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n16/pdf/scientificamerican04181903-295f.pdf.

Anonymous:1903:Cf

[5094] Anonymous. Correspondence. Scientific American, 88(16):299, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n16/pdf/scientificamerican04181903-299.pdf.

Anonymous:1903:RPIh

[5095] Anonymous. Recently patented inventions. Scientific American, 88(16):303, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
303a.pdf.

Anonymous:1903:NQf

[5096] Anonymous. Notes and queries. Scientific American, 88(16):304, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n16/pdf/scientificamerican04181903-304a.pdf.

Anonymous:1903:IIh

[5097] Anonymous. Index of inventions. Scientific American, 88(16):304–307, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n16/pdf/scientificamerican04181903-304b.pdf.

Anonymous:1903:PNPf

[5098] Anonymous. Patent notes: Patents. Scientific American, 88(16):307, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n16/pdf/scientificamerican04181903-307.pdf.

Anonymous:1903:PTR

[5099] Anonymous. Protect the third rail. Scientific American, 88(16):294, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n16/pdf/scientificamerican04181903-294.pdf.

Anonymous:1903:LMR

[5100] Anonymous. The Lodge–Muirhead receiver. Scientific American, 88(16):294, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
294a.pdf.

Anonymous:1903:FGA

[5101] Anonymous. Fatal gun accident in the Navy. Scientific American, 88(16):294, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
294b.pdf.

Anonymous:1903:CDW

[5102] Anonymous. The so-called dangers of wireless telegraphy. Scientific American, 88(16):294, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
294c.pdf.

Anonymous:1903:WCW

[5103] Anonymous. Waste of city water supply. Scientific American, 88(16):294, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
294d.pdf.

Anonymous:1903:ARU

[5104] Anonymous. Annual report of the United States Steel Trust. Scientific American, 88(16):294–295, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
294e.pdf.

Anonymous:1903:AGI

[5105] Anonymous. The accession of Germany to the International Convention. Scientific American, 88(16):295, April 18, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n16/pdf/scientificamerican04181903-295.pdf.

Anonymous:1903:RJS

[5106] Anonymous. The reinterment of James Smithson in America. Scientific American, 88(16):295, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
295a.pdf.

Anonymous:1903:RTM

[5107] Anonymous. A regular transatlantic Marconi service. Scientific American, 88(16):295, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
295b.pdf.

Anonymous:1903:NAAa

[5108] Anonymous. New American automobile speed records. Scientific American, 88(16):295, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
295c.pdf.

Anonymous:1903:RVH

[5109] Anonymous. Rubber vine in Honduras. Scientific American, 88(16):295, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n16/pdf/scientificamerican04181903-295d.pdf.

Anonymous:1903:RBZ

[5110] Anonymous. Report of the Berlin Zossen Trials. Scientific American, 88(16):295, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
295e.pdf.

Fyfe:1903:PNT

[5111] Herbert C. Fyfe. The Pedrail — a new type of road locomotive. Scientific American, 88(16):296, April 18, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n16/pdf/scientificamerican04181903-296.pdf.

Fyfe:1903:LMS

[5112] H. C. Fyfe. The Lodge–Muirhead system of wireless telegraphy. Scientific American, 88(16):297–298, April 18, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n16/pdf/scientificamerican04181903-297.pdf.

Anonymous:1903:NAB

[5113] Anonymous. The new American-built liner “Minnesota”. Scientific American, 88(16):298, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
298.pdf.

Anonymous:1903:MPEb

[5114] Anonymous. The most powerful express passenger locomotive. Scientific American, 88(16):300, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
300.pdf.

Collins:1903:NEF

[5115] A. Frederick Collins. A new electrical fire engine. Scientific American, 88(16):300–301, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
300a.pdf.

Anonymous:1903:JDA

[5116] Anonymous. The Jaquet–Droz androids. Scientific American, 88 (16):301–302, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
301.pdf.

Anonymous:1903:ADB

[5117] Anonymous. Apparatus for discharging bilge water from ships. Scientific American, 88(16):302–303, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
302.pdf.

Inkersley:1903:CPB

[5118] Arthur Inkersley. A century plant in bloom. Scientific American, 88(16):303, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
303.pdf.

Anonymous:1903:BPWh

[5119] Anonymous. Business and personal wants. Scientific American, 88(16):304, April 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n16/pdf/scientificamerican04181903-
304.pdf.

Anonymous:1903:SNf

[5120] Anonymous. Science notes. Scientific American, 88(17):311, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-311b.pdf.

Anonymous:1903:Cg

[5121] Anonymous. Correspondence. Scientific American, 88(17):315, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-315.pdf.

Anonymous:1903:RPIi

[5122] Anonymous. Recently patented inventions. Scientific American, 88(17):321, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
321.pdf.

Anonymous:1903:NQg

[5123] Anonymous. Notes and queries. Scientific American, 88(17):322, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-322.pdf.

Anonymous:1903:BNBd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-323.pdf.

Anonymous:1903:IIi

[5125] Anonymous. Index of inventions. Scientific American, 88(17):323, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-323a.pdf.

Anonymous:1903:PNPg

[5126] Anonymous. Patent notes: Patents. Scientific American, 88(17):327, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-327.pdf.

Anonymous:1903:SWD

[5127] Anonymous. Shall we have double-decked streets? Scientific American, 88(17):310, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
310.pdf.

Anonymous:1903:CGS

[5128] Anonymous. Collapse of the Great Salt Lake Railroad Trestle. Scientific American, 88(17):310, April 25, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n17/pdf/scientificamerican04251903-310a.pdf.

Anonymous:1903:DIP

[5129] Anonymous. Does it pay to reconstruct battleships? Scientific American, 88(17):310, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
310b.pdf.

Anonymous:1903:EML

[5130] Anonymous. The extermination of the mosquito on Long Island. Scientific American, 88(17):310–311, April 25, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n17/pdf/scientificamerican04251903-310c.pdf.

Anonymous:1903:SWC

[5131] Anonymous. Sir William Crookes on radium. Scientific American, 88(17):311, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
311.pdf.

Anonymous:1903:EC

[5132] Anonymous. “Egyptianized” clay. Scientific American, 88(17):311, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-311a.pdf.

Anonymous:1903:PRW

[5133] Anonymous. Powerful railroad wrecking cranes. Scientific American, 88(17):312, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
312.pdf.

Anonymous:1903:PPF

[5134] Anonymous. Past, present, and future of the Chicago, Aurora, and Elgin Railway. Scientific American, 88(17):312, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-312a.pdf.

Perkins:1903:MDS

[5135] Frank C. Perkins. Modern development of the steam turbine. Scientific American, 88(17):313–314, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
313.pdf.

Anonymous:1903:LR

[5136] Anonymous. Launch of the “Reliance”. Scientific American, 88 (17):314–315, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
314.pdf.

Anonymous:1903:SIU

[5137] Anonymous. “Shamrock III.” under sail. Scientific American, 88(17):316, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
316.pdf.

Dunlap:1903:SRF

[5138] Orrin E. Dunlap. “siloxicon” — a refractory furnace lining. Scientific American, 88(17):316–317, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
316a.pdf.

Fyfe:1903:TBP

[5139] Herbert C. Fyfe. The telegraphone and the British Post Office. Scientific American, 88(17):317–318, April 25, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n17/pdf/scientificamerican04251903-317.pdf.

Anonymous:1903:KT

[5140] Anonymous. The Korean twins. Scientific American, 88(17):318, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-318.pdf.

Anonymous:1903:CR

[5141] Anonymous. Car replacer. Scientific American, 88(17):319, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-319.pdf.

Anonymous:1903:MVE

[5142] Anonymous. Mixing valve for explosion-engines. Scientific American, 88(17):319, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
319a.pdf.

Anonymous:1903:TJB

[5143] Anonymous. Transferable job bank. Scientific American, 88(17):319, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-319b.pdf.

Anonymous:1903:FPa

[5144] Anonymous. Fuse plug. Scientific American, 88(17):319, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-319c.pdf.

Anonymous:1903:SRF

[5145] Anonymous. Sideboard-rod fastening. Scientific American, 88(17):319, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-319d.pdf.

Anonymous:1903:LNb

[5146] Anonymous. Legal notes. Scientific American, 88(17):320, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n17/pdf/scientificamerican04251903-320.pdf.

Anonymous:1903:BPWi

[5147] Anonymous. Business and personal wants. Scientific American, 88(17):321, April 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n17/pdf/scientificamerican04251903-
321a.pdf.

Russell:1903:HMa

[5148] Henry Norris Russell. The heavens in May. Scientific American, 88(18):331, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
331.pdf.

Anonymous:1903:SNg

[5149] Anonymous. Science notes. Scientific American, 88(18):331, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n18/pdf/scientificamerican05021903-331a.pdf.

Anonymous:1903:Ch

[5150] Anonymous. Correspondence. Scientific American, 88(18):335, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n18/pdf/scientificamerican05021903-335.pdf.

Anonymous:1903:RPIj

[5151] Anonymous. Recently patented inventions. Scientific American, 88(18):341–342, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
341.pdf.

Anonymous:1903:NQh

[5152] Anonymous. Notes and queries. Scientific American, 88(18):342–343, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n18/pdf/scientificamerican05021903-342a.pdf.

Anonymous:1903:BNBe

[5153] Anonymous. Books: New books, etc. Scientific American, 88(18):343, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n18/pdf/scientificamerican05021903-343.pdf.

Anonymous:1903:IIj

[5154] Anonymous. Index of inventions. Scientific American, 88(18):343–347, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n18/pdf/scientificamerican05021903-343a.pdf.

Anonymous:1903:PNPh

[5155] Anonymous. Patent notes: Patents. Scientific American, 88(18):347, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n18/pdf/scientificamerican05021903-347.pdf.

Anonymous:1903:SCSb

[5156] Anonymous. The subway contractor and the suffering public. Scientific American, 88(18):330, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
330.pdf.

Anonymous:1903:FII

[5157] Anonymous. Failure of the “Iowa’s” 12-inch gun. Scientific American, 88(18):330, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
330a.pdf.

Anonymous:1903:TOF

[5158] Anonymous. Tests of oil fuel on locomotives. Scientific American, 88(18):330, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
330b.pdf.

Anonymous:1903:ERA

[5159] Anonymous. English report on the American railways. Scientific American, 88(18):330, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
330c.pdf.

Anonymous:1903:ELU

[5160] Anonymous. Electrification of the London “Underground” Railway. Scientific American, 88(18):330–331, May 2, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n18/pdf/scientificamerican05021903-330d.pdf.

Anonymous:1903:SAb

[5161] Anonymous. The Stanley airship. Scientific American, 88(18):332, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n18/pdf/scientificamerican05021903-332.pdf.

Anonymous:1903:PEE

[5162] Anonymous. Powerful English engine for suburban traffic. Scientific American, 88(18):332–333, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
332a.pdf.

Fyfe:1903:OFE

[5163] Herbert C. Fyfe. Ore finding by electricity. Scientific American, 88(18):333–334, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
333.pdf.

Benjamin:1903:AGB

[5164] Marcus Benjamin. Alexander Graham Bell. Scientific American, 88(18):334, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
334.pdf.

Anonymous:1903:EFP

[5165] Anonymous. Experiments in food preservatives conducted by the department of agriculture. Scientific American, 88(18):336–338, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n18/pdf/scientificamerican05021903-336.pdf.

Anonymous:1903:CWS

[5166] Anonymous. Condition of the work on the subway. Scientific American, 88(18):338, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
338.pdf.

Anonymous:1903:MNY

[5167] Anonymous. Motors for the New York City subway. Scientific American, 88(18):339, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
339.pdf.

Anonymous:1903:LAL

[5168] Anonymous. The latest Atlantic liner, “Kaiser Wilhelm II”. Scientific American, 88(18):339–340, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
339a.pdf.

Anonymous:1903:BPWj

[5169] Anonymous. Business and personal wants. Scientific American, 88(18):342, May 2, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n18/pdf/scientificamerican05021903-
342.pdf.

Anonymous:1903:Ci

[5170] Anonymous. Correspondence. Scientific American, 88(19):355, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-355.pdf.

Anonymous:1903:RPIk

[5171] Anonymous. Recently patented inventions. Scientific American, 88(19):361, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
361.pdf.

Anonymous:1903:NQi

[5172] Anonymous. Notes and queries. Scientific American, 88(19):362–363, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-362a.pdf.

Anonymous:1903:BNBf

[5173] Anonymous. Books: New books, etc. Scientific American, 88(19):363, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-363.pdf.

Anonymous:1903:IIk

[5174] Anonymous. Index of inventions. Scientific American, 88(19):364, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-364.pdf.

Anonymous:1903:PNPi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-367.pdf.

Anonymous:1903:FPT

[5176] Anonymous. Filtration plants and typhoid fever. Scientific American, 88(19):350, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
350.pdf.

Anonymous:1903:HSA

[5177] Anonymous. High speed on American railroads. Scientific American, 88(19):350, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
350a.pdf.

Anonymous:1903:ATG

[5178] Anonymous. Acceleration tests on the Great Eastern Railway. Scientific American, 88(19):350, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
350b.pdf.

Anonymous:1903:RDS

[5179] Anonymous. Rapid development of the steam turbine. Scientific American, 88(19):350, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
350c.pdf.

Anonymous:1903:NED

[5180] Anonymous. New electrical discoveries. Scientific American, 88 (19):350–351, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
350d.pdf.

Benjamin:1903:APP

[5181] Marcus Benjamin. Abstracts of papers presented at the National Academy of Sciences. Scientific American, 88(19):351, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-351.pdf.

Anonymous:1903:SNh

[5182] Anonymous. Science notes. Scientific American, 88(19):351, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-351a.pdf.

Anonymous:1903:RCB

[5183] Anonymous. Rapid construction of the battleship “Louisiana”. Scientific American, 88(19):352, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
352.pdf.

Anonymous:1903:LL

[5184] Anonymous. “Looping the loop” in 1846. Scientific American, 88(19):352–353, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
352a.pdf.

Holder:1903:SCW

[5185] Charles F. Holder. Santa Catalina’s wireless newspaper. Scientific American, 88(19):353, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
353.pdf.

Anonymous:1903:DSO

[5186] Anonymous. Dust storm observations. Scientific American, 88(19):353, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-353a.pdf.

Anonymous:1903:RUS

[5187] Anonymous. The “Reliance” under sail. Scientific American, 88 (19):354, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
354.pdf.

Anonymous:1903:ACAa

[5188] Anonymous. An automatic centrifugal amusement-way. Scientific American, 88(19):356–357, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
356.pdf.

Beasley:1903:SCJ

[5189] Walter L. Beasley. Siberian collection of the Jesup North Pacific Expedition. Scientific American, 88(19):357–358, May 9, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n19/pdf/scientificamerican05091903-357.pdf.

Anonymous:1903:WBI

[5190] Anonymous. The White blackberry “Iceberg”. Scientific American, 88(19):358, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
358.pdf.

Anonymous:1903:CGTb

[5191] Anonymous. A convenient game table. Scientific American, 88(19):359, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-359.pdf.

Anonymous:1903:SCE

[5192] Anonymous. A safety catch for elevators. Scientific American, 88(19):359, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
359a.pdf.

Anonymous:1903:DDB

[5193] Anonymous. Daylight developing box. Scientific American, 88(19):359, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n19/pdf/scientificamerican05091903-359b.pdf.

Anonymous:1903:OIa

[5194] Anonymous. Oddities in inventions. Scientific American, 88 (19):359–360, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
359c.pdf.

Anonymous:1903:BPWk

[5195] Anonymous. Business and personal wants. Scientific American, 88(19):362, May 9, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n19/pdf/scientificamerican05091903-
362.pdf.

Anonymous:1903:Cj

[5196] Anonymous. Correspondence. Scientific American, 88(2):22–23, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n2/pdf/scientificamerican01101903-22.pdf.

Anonymous:1903:RPIl

[5197] Anonymous. Recently patented inventions. Scientific American, 88(2):28, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
28.pdf.

Anonymous:1903:BPWl

[5198] Anonymous. Business and personal wants. Scientific American, 88(2):28, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
28a.pdf.

Anonymous:1903:NQj

[5199] Anonymous. Notes and queries. Scientific American, 88(2):28, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n2/pdf/scientificamerican01101903-28b.pdf.

Anonymous:1903:BNBg

[5200] Anonymous. Books: New books, etc. Scientific American, 88 (2):28, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
28c.pdf.

Anonymous:1903:IIl

[5201] Anonymous. Index of inventions. Scientific American, 88(2):28–33, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n2/pdf/scientificamerican01101903-28d.pdf.

Anonymous:1903:PNPj

[5202] Anonymous. Patent notes: Patents. Scientific American, 88(2):33, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n2/pdf/scientificamerican01101903-33.pdf.

Anonymous:1903:RAH

[5203] Anonymous. Railroad accidents here and abroad. Scientific American, 88(2):18, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
18.pdf.

Anonymous:1903:PIO

[5204] Anonymous. Proposed increase of our Navy. Scientific American, 88(2):18, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
18a.pdf.

Anonymous:1903:TCM

[5205] Anonymous. Traffic congestion in Manhattan. Scientific American, 88(2):18–19, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
18b.pdf.

Anonymous:1903:GAWa

[5206] Anonymous. The German–American War Game. Scientific American, 88(2):19, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
19.pdf.

Benjamin:1903:PIR

[5207] Marcus Benjamin. Prof. Ira Remsen, President of the American Association for the Advancement of Science. Scientific American, 88(2):19, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n2/pdf/scientificamerican01101903-19a.pdf.

Anonymous:1903:LDA

[5208] Anonymous. The Lebaudy dirigible airship. Scientific American, 88(2):20, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
20.pdf.

Brown:1903:DSP

[5209] Enos Brown. Destruction of steamship “Progresso”. Scientific American, 88(2):21, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
21.pdf.

Anonymous:1903:NBO

[5210] Anonymous. The new British one-man folding range finder. Scientific American, 88(2):21–22, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
21a.pdf.

Anonymous:1903:OFS

[5211] Anonymous. Opening of the first section of the Pacific cable. Scientific American, 88(2):24–25, January 10, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n2/pdf/scientificamerican01101903-24.pdf.

Willey:1903:OFS

[5212] Day Allen Willey. Oil fires in the southwest. Scientific American, 88(2):25–26, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
25.pdf.

Beard:1903:SLA

[5213] J. Carter Beard. Are the senses of the lower animals superior to ours? Scientific American, 88(2):26–27, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n2/pdf/scientificamerican01101903-26.pdf.

Beasley:1903:OLT

[5214] Walter L. Beasley. One of the largest trees in the world. Scientific American, 88(2):27, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
27.pdf.

Anonymous:1903:RTF

[5215] Anonymous. Railway tie and fastening. Scientific American, 88 (2):27, January 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n2/pdf/scientificamerican01101903-
27a.pdf.

Anonymous:1903:SNi

[5216] Anonymous. Science notes. Scientific American, 88(20):371, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-371b.pdf.

Anonymous:1903:Ck

[5217] Anonymous. Correspondence. Scientific American, 88(20):374–375, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-374.pdf.

Anonymous:1903:RPIm

[5218] Anonymous. Recently patented inventions. Scientific American, 88(20):380, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
380a.pdf.

Anonymous:1903:NQk

[5219] Anonymous. Notes and queries. Scientific American, 88(20):381, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-381a.pdf.

Anonymous:1903:BNBh

[5220] Anonymous. Books: New books, etc. Scientific American, 88(20):382, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-382.pdf.

Anonymous:1903:IIm

[5221] Anonymous. Index of inventions. Scientific American, 88(20):382, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-382a.pdf.

Anonymous:1903:PNPk

[5222] Anonymous. Patent notes: Patents. Scientific American, 88 (20):385–386, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
385.pdf.

Anonymous:1903:WNF

[5223] Anonymous. We need fewer doctors. Scientific American, 88(20):370, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-370.pdf.

Anonymous:1903:PBB

[5224] Anonymous. The proposed Brooklyn Bridge Terminal Station. Scientific American, 88(20):370, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
370a.pdf.

Anonymous:1903:RCa

[5225] Anonymous. “Reliance” and “Columbia”. Scientific American, 88 (20):370, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
370b.pdf.

Anonymous:1903:OEE

[5226] Anonymous. Our enormous exports. Scientific American, 88(20):370, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-370c.pdf.

Anonymous:1903:WAW

[5227] Anonymous. What to abolish from warships. Scientific American, 88(20):370, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
370d.pdf.

Anonymous:1903:MSa

[5228] Anonymous. The metric system. Scientific American, 88(20):370, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-370e.pdf.

Anonymous:1903:FIS

[5229] Anonymous. The first iron sailing vessel. Scientific American, 88(20):370–371, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
370f.pdf.

Watson:1903:CD

[5230] Egbert P. Watson. Chimney draft. Scientific American, 88(20):371, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-371.pdf.

Anonymous:1903:TMP

[5231] Anonymous. Trade-marks in the Philippines. Scientific American, 88(20):371, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
371a.pdf.

Anonymous:1903:ACAb

[5232] Anonymous. Air cooling apparatus. Scientific American, 88(20):372, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-372.pdf.

Anonymous:1903:BBO

[5233] Anonymous. A bell-buoy operated by tide motor. Scientific American, 88(20):372, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
372a.pdf.

Anonymous:1903:DM

[5234] Anonymous. Dovetailing machine. Scientific American, 88(20):372–373, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n20/pdf/scientificamerican05161903-372b.pdf.

Wright:1903:LDH

[5235] Hamilton Wright. Long-distance high-tension transmission of power in California. Scientific American, 88(20):373–374, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n20/pdf/scientificamerican05161903-373.pdf.

Anonymous:1903:CRT

[5236] Anonymous. The Crossley reflecting telescope and its new mounting. Scientific American, 88(20):376, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
376.pdf.

Shepstone:1903:TWA

[5237] Harold J. Shepstone. The trade in wild animals. Scientific American, 88(20):376–378, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
376a.pdf.

Beard:1903:LAO

[5238] J. Carter Beard. Have the lower animals other senses than ours? Scientific American, 88(20):378–379, May 16, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n20/pdf/scientificamerican05161903-378.pdf.

Guarini:1903:AEC

[5239] Emile Guarini. Artemiev’s electrical coat of mail. Scientific American, 88(20):379, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
379.pdf.

Anonymous:1903:NES

[5240] Anonymous. A new electrical steering gear. Scientific American, 88(20):380, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
380.pdf.

Anonymous:1903:BPWm

[5241] Anonymous. Business and personal wants. Scientific American, 88(20):381, May 16, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n20/pdf/scientificamerican05161903-
381.pdf.

Anonymous:1903:RPIn

[5242] Anonymous. Recently patented inventions. Scientific American, 88(21):399, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
399.pdf.

Anonymous:1903:NQl

[5243] Anonymous. Notes and queries. Scientific American, 88(21):400–401, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-400a.pdf.

Anonymous:1903:NBEc

[5244] Anonymous. New books, etc. Scientific American, 88(21):401, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-401.pdf.

Anonymous:1903:PNPl

[5245] Anonymous. Patent notes: Patent. Scientific American, 88(21):404– 406, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1903:AEB

[5246] Anonymous. The artistic element in bridge design. Scientific American, 88(21):388, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
388.pdf.

Anonymous:1903:ART

[5247] Anonymous. American rapid transit in London. Scientific American, 88(21):388, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
388a.pdf.

Anonymous:1903:NCb

[5248] Anonymous. The new Cunarders. Scientific American, 88(21):388, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-388b.pdf.

Anonymous:1903:FCB

[5249] Anonymous. Flies as carriers of bacteria. Scientific American, 88(21):388, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
388c.pdf.

Anonymous:1903:BWA

[5250] Anonymous. British workmen on American industry. Scientific American, 88(21):388–389, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
388d.pdf.

Mayo:1903:EPP

[5251] Earl Mayo. Electric power plant below Mount Rainier. Scientific American, 88(21):389, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
389.pdf.

Anonymous:1903:CPG

[5252] Anonymous. Carrier pigeons for the German navy. Scientific American, 88(21):389, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
389a.pdf.

Anonymous:1903:SNj

[5253] Anonymous. Science notes. Scientific American, 88(21):389, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-389b.pdf.

Anonymous:1903:NSW

[5254] Anonymous. A new soda-water machine. Scientific American, 88 (21):390, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
390.pdf.

Anonymous:1903:EOR

[5255] Anonymous. An electric ore-reducing apparatus. Scientific American, 88(21):390, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
390a.pdf.

Anonymous:1903:SKE

[5256] Anonymous. A simple kerosene engine. Scientific American, 88(21):391, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-391.pdf.

Anonymous:1903:MCB

[5257] Anonymous. Motor coaches for British railroads. Scientific American, 88(21):391–392, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
391a.pdf.

Mclure:1903:CIP

[5258] W. Frank M’clure. The coal industry of Pennsylvania. Scientific American, 88(21):392–393, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
392.pdf.

Collins:1903:WTC

[5259] T. Byard Collins. Wireless telephonic communication between moving ferry boats. Scientific American, 88(21):394–395, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-394.pdf.

Anonymous:1903:LBAb

[5260] Anonymous. A land boat for armory use. Scientific American, 88(21):395–396, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
395.pdf.

Guarini:1903:GHT

[5261] Emile Guarini. Guilleminot high-tension and high-frequency coils. Scientific American, 88(21):396, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
396.pdf.

Anonymous:1903:PDb

[5262] Anonymous. Patent department. Scientific American, 88(21):397, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-397.pdf.

Anonymous:1903:UAR

[5263] Anonymous. Useful attachment for rain-water leaders. Scientific American, 88(21):397, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
397a.pdf.

Anonymous:1903:HAL

[5264] Anonymous. Heater attachment for lamps and gas burners. Scientific American, 88(21):397, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
397b.pdf.

Anonymous:1903:GRC

[5265] Anonymous. Gage for railway car axles. Scientific American, 88 (21):397–398, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
397c.pdf.

Anonymous:1903:RCc

[5266] Anonymous. Rail connection. Scientific American, 88(21):398, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-398.pdf.

Anonymous:1903:CSb

[5267] Anonymous. Centrifugal separator. Scientific American, 88(21):398, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-398a.pdf.

Anonymous:1903:BPWn

[5268] Anonymous. Business and personal wants. Scientific American, 88(21):399–400, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n21/pdf/scientificamerican05231903-
399a.pdf.

Anonymous:1903:PST

[5269] Anonymous. Popular steam traps. Scientific American, 88(21):400, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-400.pdf.

Anonymous:1903:IIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n21/pdf/scientificamerican05231903-401a.pdf.

Anonymous:1903:VBa

[5271] Anonymous. Valuable books! Scientific American, 88(21):403–404, May 23, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Russell:1903:HJb

[5272] Henry Norris Russell. The heavens in June. Scientific American, 88(22):409, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
409.pdf.

Anonymous:1903:Cl

[5273] Anonymous. Correspondence. Scientific American, 88(22):413, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-413.pdf.

Anonymous:1903:RPIo

[5274] Anonymous. Recently patented inventions. Scientific American, 88(22):418, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
418.pdf.

Anonymous:1903:IIo

[5275] Anonymous. Index of inventions. Scientific American, 88(22):418, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-418b.pdf.

Anonymous:1903:PNPm

[5276] Anonymous. Patent notes: Patents. Scientific American, 88(22):419, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-419.pdf.

Anonymous:1903:WDN

[5277] Anonymous. The War Department and New York City docks. Scientific American, 88(22):408, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
408.pdf.

Anonymous:1903:RHV

[5278] Anonymous. Railway horrors vs. the “horrors of war” an appalling record. Scientific American, 88(22):408, May 30, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n22/pdf/scientificamerican05301903-408a.pdf.

Anonymous:1903:BRH

[5279] Anonymous. Boiler rivet heads. Scientific American, 88(22):408, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-408b.pdf.

Anonymous:1903:AOV

[5280] Anonymous. An American officer’s view of German industries. Scientific American, 88(22):408–409, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
408c.pdf.

Anonymous:1903:SNk

[5281] Anonymous. Science notes. Scientific American, 88(22):409, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-409a.pdf.

Fyfe:1903:IAM

[5282] Herbert C. Fyfe. The Ivel agricultural motor. Scientific American, 88(22):410, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
410.pdf.

Anonymous:1903:NAF

[5283] Anonymous. New automatic fender and wheel guard. Scientific American, 88(22):410–411, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
410a.pdf.

Anonymous:1903:PGC

[5284] Anonymous. The problem of the Great Croton Dam — New York City Water Supply. Scientific American, 88(22):411–412, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-411.pdf.

Anonymous:1903:FRC

[5285] Anonymous. First race of the Cup Defender “Reliance”. Scientific American, 88(22):414, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
414.pdf.

Boyer:1903:BFF

[5286] Jacques Boyer. Butterfly farms in France. Scientific American, 88(22):415–416, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
415.pdf.

Anonymous:1903:MWW

[5287] Anonymous. A man who was struck by lightning and lives. Scientific American, 88(22):416, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
416.pdf.

Anonymous:1903:MSb

[5288] Anonymous. Manure spreader. Scientific American, 88(22):417, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-417.pdf.

Anonymous:1903:FTV

[5289] Anonymous. Folding top for vehicles. Scientific American, 88(22):417, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-417a.pdf.

Anonymous:1903:ET

[5290] Anonymous. Elevating truck. Scientific American, 88(22):417, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n22/pdf/scientificamerican05301903-417b.pdf.

Anonymous:1903:OGC

[5291] Anonymous. An opera-glass condiment holder. Scientific American, 88(22):417, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
417c.pdf.

Anonymous:1903:WRS

[5292] Anonymous. Water regulator for steam boilers. Scientific American, 88(22):417, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
417d.pdf.

Anonymous:1903:BPWo

[5293] Anonymous. Business and personal wants. Scientific American, 88(22):418, May 30, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n22/pdf/scientificamerican05301903-
418a.pdf.

Anonymous:1903:SNl

[5294] Anonymous. Science notes. Scientific American, 88(23):425, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n23/pdf/scientificamerican06061903-425b.pdf.

Anonymous:1903:LNc

[5295] Anonymous. Legal notes. Scientific American, 88(23):434, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n23/pdf/scientificamerican06061903-434.pdf.

Anonymous:1903:RPIp

[5296] Anonymous. Recently patented inventions. Scientific American, 88(23):435–436, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
435.pdf.

Anonymous:1903:NQm

[5297] Anonymous. Notes and queries. Scientific American, 88(23):436, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n23/pdf/scientificamerican06061903-436a.pdf.

Anonymous:1903:BNBi

[5298] Anonymous. Books: New books, etc. Scientific American, 88(23):436, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n23/pdf/scientificamerican06061903-436b.pdf.

Anonymous:1903:IIp

[5299] Anonymous. Index of inventions. Scientific American, 88(23):436, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n23/pdf/scientificamerican06061903-436c.pdf.

Anonymous:1903:PNPn

[5300] Anonymous. Patent notes: Patents. Scientific American, 88(23):440, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n23/pdf/scientificamerican06061903-440.pdf.

Anonymous:1903:THF

[5301] Anonymous. The two hundred and fiftieth anniversary of New York City. Scientific American, 88(23):424, June 6, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n23/pdf/scientificamerican06061903-424.pdf.

Anonymous:1903:REE

[5302] Anonymous. Remarkable efficiency of elevated railway electric service. Scientific American, 88(23):424, June 6, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n23/pdf/scientificamerican06061903-424a.pdf.

Anonymous:1903:CHS

[5303] Anonymous. Cost of high-speed steamship travel. Scientific American, 88(23):424, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
424b.pdf.

Anonymous:1903:ANS

[5304] Anonymous. The abuse of a noble sport. Scientific American, 88(23):424–425, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
424c.pdf.

Anonymous:1903:PGS

[5305] Anonymous. Prof. Goodspeed on secondary radiation induced by X-rays. Scientific American, 88(23):425, June 6, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n23/pdf/scientificamerican06061903-425.pdf.

Anonymous:1903:AME

[5306] Anonymous. Alkali manufacture by electricity from natural salts. Scientific American, 88(23):425, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
425a.pdf.

Anonymous:1903:OFO

[5307] Anonymous. The Ostergren fuel oil engine. Scientific American, 88(23):426, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
426.pdf.

Anonymous:1903:ITSa

[5308] Anonymous. An improved type of swage block. Scientific American, 88(23):427, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
427.pdf.

Anonymous:1903:MMG

[5309] Anonymous. A method of making glass models of mines. Scientific American, 88(23):427, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
427a.pdf.

Willey:1903:SSC

[5310] Day Allen Willey. The “Soo” Ship Canal System — the fiftieth anniversary of its commencement. Scientific American, 88(23):427–429, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n23/pdf/scientificamerican06061903-427b.pdf.

Anonymous:1903:NAD

[5311] Anonymous. A new apparatus for determining the resistance of road vehicles to traction. Scientific American, 88(23):430–431, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n23/pdf/scientificamerican06061903-430.pdf.

Anonymous:1903:SPR

[5312] Anonymous. Scientific poultry raising. Scientific American, 88 (23):431–432, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
431.pdf.

Anonymous:1903:NSMa

[5313] Anonymous. The new Springfield magazine rifle. Scientific American, 88(23):432, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
432.pdf.

Anonymous:1903:RSL

[5314] Anonymous. Ruhmer’s system of light-telephony. Scientific American, 88(23):433, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
433.pdf.

Anonymous:1903:BPWp

[5315] Anonymous. Business and personal wants. Scientific American, 88(23):436, June 6, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n23/pdf/scientificamerican06061903-
436.pdf.

Anonymous:1903:SNm

[5316] Anonymous. Science notes. Scientific American, 88(24):445, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-445c.pdf.

Anonymous:1903:RPIq

[5317] Anonymous. Recently patented inventions. Scientific American, 88(24):455–456, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
455.pdf.

Anonymous:1903:NQn

[5318] Anonymous. Notes and queries. Scientific American, 88(24):456–457, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-456a.pdf.

Anonymous:1903:BNBj

[5319] Anonymous. Books: New books, etc. Scientific American, 88(24):457, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-457.pdf.

Anonymous:1903:IIq

[5320] Anonymous. Index of inventions. Scientific American, 88(24):457, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-457a.pdf.

Anonymous:1903:PNPo

[5321] Anonymous. Patent notes: Patents. Scientific American, 88(24):459, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-459a.pdf.

Anonymous:1903:MACa

[5322] Anonymous. Mosely on American competition. Scientific American, 88(24):444, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
444.pdf.

Anonymous:1903:HST

[5323] Anonymous. The high-speed turbine as an air compressor. Scientific American, 88(24):444, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
444a.pdf.

Anonymous:1903:NNP

[5324] Anonymous. A new Niagara Power Canal. Scientific American, 88(24):444, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
444b.pdf.

Anonymous:1903:SEB

[5325] Anonymous. Scientific Expedition to the Bahama Islands. Scientific American, 88(24):444, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
444c.pdf.

Anonymous:1903:SRPa

[5326] Anonymous. On a singular radiation phenomenon. Scientific American, 88(24):444–445, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
444d.pdf.

Anonymous:1903:OWK

[5327] Anonymous. Observation war kites. Scientific American, 88(24):445, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-445.pdf.

Mitchell:1903:TCS

[5328] Guy E. Mitchell. Tomato culture in the South. Scientific American, 88(24):445, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
445a.pdf.

Anonymous:1903:KC

[5329] Anonymous. A kite competition. Scientific American, 88(24):445, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-445b.pdf.

Fawcett:1903:NAL

[5330] Waldon Fawcett. A novel automobile lawn mower. Scientific American, 88(24):446, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
446.pdf.

Anonymous:1903:NSP

[5331] Anonymous. A new and sanitary process of pouring brass in casting. Scientific American, 88(24):446–447, June 13, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n24/pdf/scientificamerican06131903-446a.pdf.

Anonymous:1903:NIE

[5332] Anonymous. New instruments for enabling the deaf to hear. Scientific American, 88(24):447–448, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
447.pdf.

Jane:1903:CGB

[5333] Fred T. Jane. A comparison of the German battleship “Wettin” with the “Maine”. Scientific American, 88(24):448–449, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-448.pdf.

Anonymous:1903:NMCa

[5334] Anonymous. The new motor chair. Scientific American, 88(24):450, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-450.pdf.

Anonymous:1903:DFAa

[5335] Anonymous. A daring feat in the amusement line. Scientific American, 88(24):450–451, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
450a.pdf.

Beasley:1903:EH

[5336] Walter L. Beasley. Evolution of the horse. Scientific American, 88(24):451–452, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
451.pdf.

Anonymous:1903:PDc

[5337] Anonymous. Patent department. Scientific American, 88(24):452–453, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-452.pdf.

Anonymous:1903:NPM

[5338] Anonymous. New prepayment meter. Scientific American, 88(24):453, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-453.pdf.

Anonymous:1903:DST

[5339] Anonymous. A durable saw-tightener. Scientific American, 88(24):453, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-453a.pdf.

Anonymous:1903:CCRb

[5340] Anonymous. Cage and cable release. Scientific American, 88(24):453, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-453b.pdf.

Anonymous:1903:SCb

[5341] Anonymous. Sugar cabinet. Scientific American, 88(24):453–454, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-453c.pdf.

Anonymous:1903:PTH

[5342] Anonymous. A peculiar toy hoop. Scientific American, 88(24):454, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-454.pdf.

Anonymous:1903:RF

[5343] Anonymous. Rail fastening. Scientific American, 88(24):454, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-454a.pdf.

Anonymous:1903:VPP

[5344] Anonymous. A vest pocket punching bag. Scientific American, 88(24):454, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
454b.pdf.

Anonymous:1903:IOB

[5345] Anonymous. An improved oil burner. Scientific American, 88(24):454, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n24/pdf/scientificamerican06131903-454c.pdf.

Anonymous:1903:BPWq

[5346] Anonymous. Business and personal wants. Scientific American, 88(24):456, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
456.pdf.

Hiscox:1903:GGO

[5347] Gardner D. Hiscox. Gas gasoline and oil engines. Scientific American, 88(24):459, June 13, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n24/pdf/scientificamerican06131903-
459.pdf.

Anonymous:1903:SNn

[5348] Anonymous. Science notes. Scientific American, 88(25):465, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n25/pdf/scientificamerican06201903-465c.pdf.

Anonymous:1903:Cm

[5349] Anonymous. Correspondence. Scientific American, 88(25):469, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n25/pdf/scientificamerican06201903-469a.pdf.

Anonymous:1903:BPWr

[5350] Anonymous. Business and personal wants. Scientific American, 88(25):474, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
474.pdf.

Anonymous:1903:RPIr

[5351] Anonymous. Recently patented inventions. Scientific American, 88(25):474, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
474a.pdf.

Anonymous:1903:IIr

[5352] Anonymous. Index of inventions. Scientific American, 88(25):474–477, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n25/pdf/scientificamerican06201903-474b.pdf.

Anonymous:1903:PNPp

[5353] Anonymous. Patent notes: Patents. Scientific American, 88(25):476, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n25/pdf/scientificamerican06201903-476.pdf.

Anonymous:1903:MAJ

[5354] Anonymous. The Merchants’ Association and the Jerome Park Reservoir. Scientific American, 88(25):464, June 20, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n25/pdf/scientificamerican06201903-464.pdf.

Anonymous:1903:CDC

[5355] Anonymous. Chicago Drainage Canal and the City of St. Louis. Scientific American, 88(25):464, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
464a.pdf.

Anonymous:1903:OMR

[5356] Anonymous. The oldest map of rome. Scientific American, 88 (25):464–465, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
464b.pdf.

Anonymous:1903:MPC

[5357] Anonymous. A mud-puddle commune; or, the beginnings of mind. Scientific American, 88(25):465, June 20, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n25/pdf/scientificamerican06201903-465.pdf.

Anonymous:1903:HWR

[5358] Anonymous. Highest wind record. Scientific American, 88(25):465, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n25/pdf/scientificamerican06201903-465a.pdf.

Anonymous:1903:FML

[5359] Anonymous. Failure of the monolith lathe. Scientific American, 88(25):465, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
465b.pdf.

Anonymous:1903:OCC

[5360] Anonymous. Orange culture in California. Scientific American, 88(25):466, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
466.pdf.

Anonymous:1903:FEE

[5361] Anonymous. French express engine for an English railroad. Scientific American, 88(25):467, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
467.pdf.

Fyfe:1903:FCC

[5362] H. C. Fyfe. The first cross-channel turbine steamer. Scientific American, 88(25):467–468, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
467a.pdf.

Anonymous:1903:POA

[5363] Anonymous. A portable oscillograph for alternating currents. Scientific American, 88(25):468–469, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
468.pdf.

MClure:1903:GOH

[5364] W. Frank M’Clure. Gigantic ore-handling machinery. Scientific American, 88(25):469, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
469.pdf.

Anonymous:1903:AOA

[5365] Anonymous. The Atlantic Ocean and the ‘America’ Cup. Scientific American, 88(25):470, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
470.pdf.

Anonymous:1903:PMAa

[5366] Anonymous. The Paris–Madrid Automobile Race. Scientific American, 88(25):470–472, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
470a.pdf.

Kime:1903:SEA

[5367] J. W. Kime. Some experiments with actinic light. Scientific American, 88(25):472–473, June 20, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n25/pdf/scientificamerican06201903-
472.pdf.

Anonymous:1903:SNo

[5368] Anonymous. Science notes. Scientific American, 88(26):481, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-481a.pdf.

Anonymous:1903:LNd

[5369] Anonymous. Legal notes. Scientific American, 88(26):490, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-490.pdf.

Anonymous:1903:RPIs

[5370] Anonymous. Recently patented inventions. Scientific American, 88(26):491, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
491.pdf.

Anonymous:1903:BPWs

[5371] Anonymous. Business and personal wants. Scientific American, 88(26):491, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
491a.pdf.

Anonymous:1903:NQo

[5372] Anonymous. Notes and queries. Scientific American, 88(26):491, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-491b.pdf.

Anonymous:1903:IIs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-492.pdf.

Anonymous:1903:PNPq

[5374] Anonymous. Patent notes: Patents. Scientific American, 88(26):492, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-492a.pdf.

Anonymous:1903:I

[5375] Anonymous. Index. Scientific American, 88(26):495–497, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-495.pdf.

Anonymous:1903:RNE

[5376] Anonymous. The railroads and the new East River Bridge. Scientific American, 88(26):480, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
480.pdf.

Anonymous:1903:PA

[5377] Anonymous. Painting by the acre. Scientific American, 88(26):480, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-480a.pdf.

Anonymous:1903:RTO

[5378] Anonymous. Railroad ties and our forest supply. Scientific American, 88(26):480, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
480b.pdf.

Anonymous:1903:NTB

[5379] Anonymous. The new 13,000-ton battleships. Scientific American, 88(26):480, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
480c.pdf.

Anonymous:1903:SID

[5380] Anonymous. “Shamrock III” in drydock. Scientific American, 88 (26):480–481, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
480d.pdf.

Russell:1903:HJc

[5381] Henry Norris Russell. The heavens in July. Scientific American, 88(26):481, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
481.pdf.

Anonymous:1903:DCCa

[5382] Anonymous. Dumping car with convenient lock and release. Scientific American, 88(26):482, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
482.pdf.

Anonymous:1903:GTP

[5383] Anonymous. German tests of the Pupin system of longdistance telephony. Scientific American, 88(26):482–483, June 27, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n26/pdf/scientificamerican06271903-482a.pdf.

Jones:1903:SCR

[5384] George J. Jones. A steel caboose and repair car. Scientific American, 88(26):483, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
483.pdf.

Perkins:1903:EHC

[5385] Frank C. Perkins. Electric haulage on canals. Scientific American, 88(26):483–485, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
483a.pdf.

Anonymous:1903:SMT

[5386] Anonymous. A safety magazine torpedo cane. Scientific American, 88(26):485, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
485.pdf.

Anonymous:1903:AYC

[5387] Anonymous. Anchor with yielding chain connection. Scientific American, 88(26):485, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
485a.pdf.

Gradenwitz:1903:TNA

[5388] A. Gradenwitz. The “telephot,” a novel apparatus for photographing at great distances. Scientific American, 88(26):486, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-486.pdf.

Fyfe:1903:CWS

[5389] Herbert C. Fyfe. Coaling warships at sea. Scientific American, 88(26):486–488, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
486a.pdf.

Anonymous:1903:HLT

[5390] Anonymous. Hailstones larger than hens’ eggs. Scientific American, 88(26):488, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
488.pdf.

Holder:1903:IFA

[5391] Charles F. Holder. Interesting facts about pelicans. Scientific American, 88(26):489, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
489.pdf.

Bryan:1903:BBT

[5392] W. T. Bryan. A battle between two deadly snakes. Scientific American, 88(26):489, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
489a.pdf.

Anonymous:1903:SASa

[5393] Anonymous. Scientific American supplement catalogue. Scientific American, 88(26):493, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
493.pdf.

Anonymous:1903:VBb

[5394] Anonymous. Valuable books. Scientific American, 88(26):494, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n26/pdf/scientificamerican06271903-494.pdf.

Anonymous:1903:PPP

[5395] Anonymous. Practical pointers for patentees. Scientific American, 88(26):494, June 27, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n26/pdf/scientificamerican06271903-
494a.pdf.

Anonymous:1903:SNp

[5396] Anonymous. Science notes. Scientific American, 88(3):37, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-37a.pdf.

Anonymous:1903:Cn

[5397] Anonymous. Correspondence. Scientific American, 88(3):41, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-41.pdf.

Anonymous:1903:LNe

[5398] Anonymous. Legal notes. Scientific American, 88(3):44, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-44.pdf.

Anonymous:1903:RPIt

[5399] Anonymous. Recently patented inventions. Scientific American, 88(3):46–47, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
46.pdf.

Anonymous:1903:NQp

[5400] Anonymous. Notes and queries. Scientific American, 88(3):47–48, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-47a.pdf.

Anonymous:1903:BNBk

[5401] Anonymous. Books: New books, etc. Scientific American, 88 (3):48, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
48.pdf.

Anonymous:1903:IIt

[5402] Anonymous. Index of inventions. Scientific American, 88(3):48, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-48a.pdf.

Anonymous:1903:PNPr

[5403] Anonymous. Patent notes: Patents. Scientific American, 88(3):52, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-52.pdf.

Anonymous:1903:RB

[5404] Anonymous. Railroad building in 1902. Scientific American, 88 (3):36, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
36.pdf.

Anonymous:1903:NSN

[5405] Anonymous. New ships for the Navy. Scientific American, 88 (3):36, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
36a.pdf.

Anonymous:1903:REC

[5406] Anonymous. Reconstruction of the Erie Canal. Scientific American, 88(3):36, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
36b.pdf.

Anonymous:1903:CFNb

[5407] Anonymous. The coal fields of Natal. Scientific American, 88 (3):36, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
36c.pdf.

Anonymous:1903:SHJ

[5408] Anonymous. Sir Harry Johnston’s explorations in Uganda and Central Africa. Scientific American, 88(3):37, January 17, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n3/pdf/scientificamerican01171903-37.pdf.

Anonymous:1903:RBR

[5409] Anonymous. The raising of a big railway bridge. Scientific American, 88(3):37–38, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
37b.pdf.

Anonymous:1903:EFE

[5410] Anonymous. European fire engines. Scientific American, 88(3):38, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-38.pdf.

Anonymous:1903:PASa

[5411] Anonymous. The Paris Automobile Show. Scientific American, 88(3):39, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
39.pdf.

Anonymous:1903:GCS

[5412] Anonymous. Grand Central Station improvements and connection with rapid transit subway. Scientific American, 88(3):39–40, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-39a.pdf.

Benjamin:1903:AAA

[5413] Marcus Benjamin. American Association for the Advancement of Science. Scientific American, 88(3):40–41, January 17, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n3/pdf/scientificamerican01171903-40.pdf.

Anonymous:1903:LTA

[5414] Anonymous. Leading types of American racing automobiles. Scientific American, 88(3):42–43, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
42.pdf.

Anonymous:1903:RSN

[5415] Anonymous. Reconstruction of the steamship “New York”. Scientific American, 88(3):43–44, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
43.pdf.

Anonymous:1903:PDd

[5416] Anonymous. Patent department. Scientific American, 88(3):45, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n3/pdf/scientificamerican01171903-45.pdf.

Anonymous:1903:PPD

[5417] Anonymous. Portable power device for farm work. Scientific American, 88(3):45–46, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
45a.pdf.

Anonymous:1903:BPWt

[5418] Anonymous. Business and personal wants. Scientific American, 88(3):47, January 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n3/pdf/scientificamerican01171903-
47.pdf.

Anonymous:1903:SNq

[5419] Anonymous. Science notes. Scientific American, 88(4):55, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n4/pdf/scientificamerican01241903-55b.pdf.

Anonymous:1903:Co

[5420] Anonymous. Correspondence. Scientific American, 88(4):59, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n4/pdf/scientificamerican01241903-59a.pdf.

Anonymous:1903:RPIu

[5421] Anonymous. Recently patented inventions. Scientific American, 88(4):64, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
64.pdf.

Anonymous:1903:IIu

[5422] Anonymous. Index of inventions. Scientific American, 88(4):65, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n4/pdf/scientificamerican01241903-65.pdf.

Anonymous:1903:PNPs

[5423] Anonymous. Patent notes: Patents. Scientific American, 88(4):67, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n4/pdf/scientificamerican01241903-67.pdf.

Anonymous:1903:BNBl

[5424] Anonymous. Books: New books, etc. Scientific American, 88 (4):68, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
68.pdf.

Anonymous:1903:NQq

[5425] Anonymous. Notes and queries. Scientific American, 88(4):68, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n4/pdf/scientificamerican01241903-68a.pdf.

Anonymous:1903:WTS

[5426] Anonymous. Wireless telegraphy and the “St. Louis”. Scientific American, 88(4):54, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
54.pdf.

Anonymous:1903:MCL

[5427] Anonymous. Magnitude of commerce on the lakes. Scientific American, 88(4):54, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
54a.pdf.

Anonymous:1903:NMTa

[5428] Anonymous. New methods of tunneling. Scientific American, 88(4):54, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
54b.pdf.

Anonymous:1903:DNV

[5429] Anonymous. A dozen new vessels for the shipping trust. Scientific American, 88(4):54, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
54c.pdf.

Anonymous:1903:HR

[5430] Anonymous. The humors of railroading. Scientific American, 88(4):54, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
54d.pdf.

Anonymous:1903:SV

[5431] Anonymous. Subway ventilation. Scientific American, 88(4):54, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n4/pdf/scientificamerican01241903-54e.pdf.

Anonymous:1903:CON

[5432] Anonymous. Colonists for our new public lands. Scientific American, 88(4):54–55, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
54f.pdf.

Anonymous:1903:MOD

[5433] Anonymous. Morse’s old diary and its prophecy. Scientific American, 88(4):55, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
55.pdf.

Anonymous:1903:WMM

[5434] Anonymous. Wireless messages to a moving train. Scientific American, 88(4):55, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
55a.pdf.

Anonymous:1903:CSc

[5435] Anonymous. The current supplement. Scientific American, 88 (4):55, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
55c.pdf.

Anonymous:1903:IPF

[5436] Anonymous. Improved paper file. Scientific American, 88(4):56, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n4/pdf/scientificamerican01241903-56.pdf.

Anonymous:1903:NMTb

[5437] Anonymous. A new method of transplanting large trees. Scientific American, 88(4):56, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
56a.pdf.

Anonymous:1903:BS

[5438] Anonymous. Belt stretcher. Scientific American, 88(4):56, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n4/pdf/scientificamerican01241903-56b.pdf.

Partridge:1903:ASC

[5439] W. E. Partridge. An automobile street car. Scientific American, 88(4):57–58, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
57.pdf.

James:1903:IP

[5440] George Wharton James. Indian physiognomies. Scientific American, 88(4):58–59, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
58.pdf.

Anonymous:1903:IVG

[5441] Anonymous. An improved vaporizer for gasoline engines. Scientific American, 88(4):59, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
59.pdf.

Anonymous:1903:HTW

[5442] Anonymous. How a telescope was made at the Jesuit College, Montreal. Scientific American, 88(4):60, January 24, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n4/pdf/scientificamerican01241903-60.pdf.

Davies:1903:RDC

[5443] I. W. Davies. Railroad development in Cuba. Scientific American, 88(4):61–63, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
61.pdf.

Frederick:1903:QCC

[5444] M. C. Frederick. Queer caprice of California woodpeckers. Scientific American, 88(4):63, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
63.pdf.

Anonymous:1903:BPWu

[5445] Anonymous. Business and personal wants. Scientific American, 88(4):64, January 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n4/pdf/scientificamerican01241903-
64a.pdf.

Russell:1903:HF

[5446] Henry Norris Russell. The heavens in February. Scientific American, 88(5):76, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
76.pdf.

Anonymous:1903:Cp

[5447] Anonymous. Correspondence. Scientific American, 88(5):77, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n5/pdf/scientificamerican01311903-77.pdf.

Anonymous:1903:RPIv

[5448] Anonymous. Recently patented inventions. Scientific American, 88(5):82–83, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
82.pdf.

Anonymous:1903:IIv

[5449] Anonymous. Index of inventions. Scientific American, 88(5):83, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n5/pdf/scientificamerican01311903-83a.pdf.

Anonymous:1903:BNBm

[5450] Anonymous. Books: New books, etc. Scientific American, 88 (5):85, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
85.pdf.

Anonymous:1903:NQr

[5451] Anonymous. Notes and queries. Scientific American, 88(5):86, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n5/pdf/scientificamerican01311903-86.pdf.

Anonymous:1903:PNPt

[5452] Anonymous. Patent notes: Patents. Scientific American, 88(5):87, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n5/pdf/scientificamerican01311903-87.pdf.

Anonymous:1903:WVH

[5453] Anonymous. Weight and velocity in heavy guns. Scientific American, 88(5):72, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
72.pdf.

Anonymous:1903:AA

[5454] Anonymous. The American automobile. Scientific American, 88 (5):72, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
72a.pdf.

Anonymous:1903:ASH

[5455] Anonymous. Abram Stevens Hewitt. Scientific American, 88(5):73, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n5/pdf/scientificamerican01311903-73.pdf.

Anonymous:1903:RBT

[5456] Anonymous. The recent Berliner transmitter patent decision. Scientific American, 88(5):73, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
73a.pdf.

Anonymous:1903:TNI

[5457] Anonymous. Test of the new 16-inch gun. Scientific American, 88(5):75, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
75.pdf.

Anonymous:1903:NDG

[5458] Anonymous. A new derrick or “Grasshopper” elevator for unloading grain from vessels. Scientific American, 88(5):75–76, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n5/pdf/scientificamerican01311903-75a.pdf.

Anonymous:1903:LHE

[5459] Anonymous. The landing of the Honolulu end of the Pacific cable. Scientific American, 88(5):78–80, January 31, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n5/pdf/scientificamerican01311903-78.pdf.

Anonymous:1903:TAA

[5460] Anonymous. The third annual automobile show. Scientific American, 88(5):80, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
80.pdf.

Anonymous:1903:PHR

[5461] Anonymous. The passing of the hall of records. Scientific American, 88(5):81–82, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
81.pdf.

Anonymous:1903:BPWv

[5462] Anonymous. Business and personal wants. Scientific American, 88(5):83, January 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n5/pdf/scientificamerican01311903-
83.pdf.

Anonymous:1903:SNr

[5463] Anonymous. Science notes. Scientific American, 88(6):91, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n6/pdf/scientificamerican02071903-91a.pdf.

Anonymous:1903:Cq

[5464] Anonymous. Correspondence. Scientific American, 88(6):95, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n6/pdf/scientificamerican02071903-95.pdf.

Anonymous:1903:RPIw

[5465] Anonymous. Recently patented inventions. Scientific American, 88 (6):100–101, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
100.pdf.

Anonymous:1903:BNBn

[5466] Anonymous. Books: New books, etc. Scientific American, 88 (6):101, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
101.pdf.

Anonymous:1903:NQs

[5467] Anonymous. Notes and queries. Scientific American, 88(6):102, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n6/pdf/scientificamerican02071903-102.pdf.

Anonymous:1903:IIw

[5468] Anonymous. Index of inventions. Scientific American, 88(6):104, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n6/pdf/scientificamerican02071903-104.pdf.

Anonymous:1903:SPC

[5469] Anonymous. The signing of the Panama Canal Treaty. Scientific American, 88(6):90, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
90.pdf.

Anonymous:1903:EIG

[5470] Anonymous. Explosion of a 12-inch gun at sandy hook. Scientific American, 88(6):90, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
90a.pdf.

Anonymous:1903:SD

[5471] Anonymous. Shipbuilding during 1902. Scientific American, 88 (6):90, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
90b.pdf.

Anonymous:1903:ACCa

[5472] Anonymous. “America” Cup Contest. Scientific American, 88 (6):90, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
90c.pdf.

Anonymous:1903:DD

[5473] Anonymous. The diseases of dust. Scientific American, 88(6): 90–91, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
90d.pdf.

Anonymous:1903:SAA

[5474] Anonymous. Some afterthoughts on the automobile show. Scientific American, 88(6):91, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
91.pdf.

Anonymous:1903:SDE

[5475] Anonymous. Structural details of the Edison storage battery. Scientific American, 88(6):92, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
92.pdf.

Brown:1903:OFS

[5476] Enos Brown. Oil fuel on the steamship “Mariposa”. Scientific American, 88(6):92, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
92a.pdf.

Dunlap:1903:ICN

[5477] Orrin E. Dunlap. The ice conditions of Niagara River. Scientific American, 88(6):93, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
93.pdf.

Anonymous:1903:IM

[5478] Anonymous. An impression of Marconi. Scientific American, 88 (6):94, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
94.pdf.

Anonymous:1903:OES

[5479] Anonymous. An operative engine the size of a dime. Scientific American, 88(6):94, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
94a.pdf.

Anonymous:1903:CHL

[5480] Anonymous. The Cooper Hewitt lamp and static converter. Scientific American, 88(6):96–97, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
96.pdf.

Anonymous:1903:CKT

[5481] Anonymous. Cows killed in thunderstorm. Scientific American, 88(6):98, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
98.pdf.

Anonymous:1903:PDe

[5482] Anonymous. Patent department. Scientific American, 88(6):99, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n6/pdf/scientificamerican02071903-99.pdf.

Anonymous:1903:NSb

[5483] Anonymous. A new slicer. Scientific American, 88(6):99, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n6/pdf/scientificamerican02071903-99a.pdf.

Anonymous:1903:NPO

[5484] Anonymous. A new pattern for overalls. Scientific American, 88(6):99, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
99b.pdf.

Anonymous:1903:OIb

[5485] Anonymous. Oddities in inventions. Scientific American, 88(6): 99–100, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
99c.pdf.

Anonymous:1903:BPWw

[5486] Anonymous. Business and personal wants. Scientific American, 88(6):101, February 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n6/pdf/scientificamerican02071903-
101a.pdf.

Anonymous:1903:Ct

[5487] Anonymous. Correspondence. Scientific American, 88(7):115, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n7/pdf/scientificamerican02141903-115.pdf.

Anonymous:1903:LNf

[5488] Anonymous. Legal notes. Scientific American, 88(7):119, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n7/pdf/scientificamerican02141903-119.pdf.

Anonymous:1903:RPIx

[5489] Anonymous. Recently patented inventions. Scientific American, 88(7):120, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
120.pdf.

Anonymous:1903:NQt

[5490] Anonymous. Notes and queries. Scientific American, 88(7):121, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n7/pdf/scientificamerican02141903-121a.pdf.

Anonymous:1903:BNBo

[5491] Anonymous. Books: New books, etc. Scientific American, 88 (7):123, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
123.pdf.

Anonymous:1903:IIx

[5492] Anonymous. Index of inventions. Scientific American, 88(7):124, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n7/pdf/scientificamerican02141903-124.pdf.

Anonymous:1903:PNPu

[5493] Anonymous. Patent notes: Patents. Scientific American, 88 (7):126, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
126.pdf.

Anonymous:1903:IDW

[5494] Anonymous. The inventions of Dr. William Church — the first patented typecasting and composing machine. Scientific American, 88(7):109–110, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
109.pdf.

Anonymous:1903:HMT

[5495] Anonymous. “How much, then, is a man better than a sheep?”. Scientific American, 88(7):110, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
110.pdf.

Anonymous:1903:GAWb

[5496] Anonymous. German–American War Game. Scientific American, 88(7):110, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
110a.pdf.

Anonymous:1903:RPN

[5497] Anonymous. A refractory piston needed. Scientific American, 88 (7):110–111, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
110b.pdf.

Anonymous:1903:MWA

[5498] Anonymous. Mr. Westinghouse on American methods in steelmaking. Scientific American, 88(7):111, February 14, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n7/pdf/scientificamerican02141903-111.pdf.

Dunlap:1903:LSN

[5499] Orrin E. Dunlap. Lightning strikes the Niagara Power Plant. Scientific American, 88(7):111, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
111a.pdf.

Anonymous:1903:NATa

[5500] Anonymous. A novel aerial toy yacht. Scientific American, 88 (7):112, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
112.pdf.

Anonymous:1903:KAC

[5501] Anonymous. The Krebs automatic carbureter for gasoline engines. Scientific American, 88(7):112, February 14, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n7/pdf/scientificamerican02141903-112a.pdf.

Anonymous:1903:ODB

[5502] Anonymous. An old double-barreled cannon. Scientific American, 88(7):113, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
113.pdf.

MClure:1903:SQO

[5503] W. Frank M’Clure. The sandstone quarries of Ohio. Scientific American, 88(7):113–114, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
113a.pdf.

Anonymous:1903:NOB

[5504] Anonymous. A new oil burner. Scientific American, 88(7):114, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n7/pdf/scientificamerican02141903-114.pdf.

Anonymous:1903:V

[5505] Anonymous. Ventilator. Scientific American, 88(7):114, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n7/pdf/scientificamerican02141903-114a.pdf.

Anonymous:1903:CUJ

[5506] Anonymous. A cave used as a jail. Scientific American, 88(7): 116–117, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
116.pdf.

Anonymous:1903:TLA

[5507] Anonymous. Transporting lumber across deep gorges. Scientific American, 88(7):117, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
117.pdf.

Geare:1903:VS

[5508] Randolph I. Geare. Venomous serpents — I. Scientific American, 88(7):118, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
118.pdf.

Anonymous:1903:BPWx

[5509] Anonymous. Business and personal wants. Scientific American, 88(7):121, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
121.pdf.

StArmont:1903:SAH

[5510] Jack St. Armont. South America and her treasures. Scientific American, 88(7):122, February 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n7/pdf/scientificamerican02141903-
122.pdf.

Anonymous:1903:SNs

[5511] Anonymous. Science notes. Scientific American, 88(8):131, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n8/pdf/scientificamerican02211903-131a.pdf.

Anonymous:1903:Cu

[5512] Anonymous. Correspondence. Scientific American, 88(8):135, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n8/pdf/scientificamerican02211903-135.pdf.

Anonymous:1903:RPIy

[5513] Anonymous. Recently patented inventions. Scientific American, 88(8):140, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
140.pdf.

Anonymous:1903:NQu

[5514] Anonymous. Notes and queries. Scientific American, 88(8):140– 141, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
140b.pdf.

Anonymous:1903:IIy

[5515] Anonymous. Index of inventions. Scientific American, 88(8):141, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n8/pdf/scientificamerican02211903-141.pdf.

Anonymous:1903:PNPv

[5516] Anonymous. Patent notes: Patents. Scientific American, 88 (8):143, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
143.pdf.

Anonymous:1903:TB

[5517] Anonymous. An 18,000 ton battleship. Scientific American, 88 (8):130, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
130.pdf.

Anonymous:1903:NCI

[5518] Anonymous. The new coke industry. Scientific American, 88 (8):130, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
130a.pdf.

Anonymous:1903:PRE

[5519] Anonymous. Proposed repairs to the East River Bridge cables. Scientific American, 88(8):130, February 21, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n8/pdf/scientificamerican02211903-130b.pdf.

Anonymous:1903:HGC

[5520] Anonymous. A Harvard grant from the Carnegie Institution. Scientific American, 88(8):130–131, February 21, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n8/pdf/scientificamerican02211903-130c.pdf.

Gleason:1903:ST

[5521] H. M. Gleason. The steam turbine. Scientific American, 88 (8):131, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
131.pdf.

Anonymous:1903:OWW

[5522] Anonymous. The Ozone Waterworks at Wiesbaden and Paderborn. Scientific American, 88(8):132–133, February 21, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n8/pdf/scientificamerican02211903-132.pdf.

Anonymous:1903:NCBa

[5523] Anonymous. The new Chilian battleship “Libertad”. Scientific American, 88(8):133–134, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
133.pdf.

Anonymous:1903:SBM

[5524] Anonymous. The spontaneous bending of marble. Scientific American, 88(8):134, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
134.pdf.

Hopkins:1903:RTA

[5525] M. Hopkins. Recording telegraph for amateurs. Scientific American, 88(8):136, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
136.pdf.

Geare:1903:VSIb

[5526] Randolph I. Geare. Venomous snakes — II. Scientific American, 88(8):137, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
137.pdf.

Anonymous:1903:NYS

[5527] Anonymous. New York’s steel roadway. Scientific American, 88 (8):138, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
138.pdf.

Anonymous:1903:CCRc

[5528] Anonymous. A curious case of regelation. Scientific American, 88(8):138, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
138a.pdf.

Brown:1903:ONC

[5529] Enos Brown. The orange in Northern California. Scientific American, 88(8):138–139, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
138b.pdf.

Anonymous:1903:BPWy

[5530] Anonymous. Business and personal wants. Scientific American, 88(8):140, February 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n8/pdf/scientificamerican02211903-
140a.pdf.

Russell:1903:HMb

[5531] Henry Norris Russell. The heavens in March. Scientific American, 88(9):153, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
153.pdf.

Anonymous:1903:LNg

[5532] Anonymous. Legal notes. Scientific American, 88(9):155–156, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n9/pdf/scientificamerican02281903-155.pdf.

Anonymous:1903:RPIz

[5533] Anonymous. Recently patented inventions. Scientific American, 88(9):160, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
160.pdf.

Anonymous:1903:NQv

[5534] Anonymous. Notes and queries. Scientific American, 88(9):161, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n9/pdf/scientificamerican02281903-161a.pdf.

Anonymous:1903:IIz

[5535] Anonymous. Index of inventions. Scientific American, 88(9): 161–164, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
161b.pdf.

Anonymous:1903:BNBp

[5536] Anonymous. Books: New books, etc. Scientific American, 88 (9):165, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
165.pdf.

Anonymous:1903:RSC

[5537] Anonymous. Rapid and slow construction of public works. Scientific American, 88(9):148, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
148.pdf.

Anonymous:1903:CHI

[5538] Anonymous. The Cooper Hewitt interrupter. Scientific American, 88(9):148, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
148a.pdf.

Kanolt:1903:REP

[5539] C. W. Kanolt. Radium-its extraordinary properties. Scientific American, 88(9):149, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
149.pdf.

Anonymous:1903:PSE

[5540] Anonymous. Proposed subway and elevated extension in New York City. Scientific American, 88(9):149, February 28, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n9/pdf/scientificamerican02281903-149a.pdf.

Anonymous:1903:CSd

[5541] Anonymous. The current supplement. Scientific American, 88 (9):149, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
149b.pdf.

Anonymous:1903:NMCb

[5542] Anonymous. New method for cutting asphalt pavement. Scientific American, 88(9):150, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
150.pdf.

Anonymous:1903:PLT

[5543] Anonymous. The production of low temperatures. Scientific American, 88(9):150, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
150a.pdf.

Anonymous:1903:IHU

[5544] Anonymous. The injured hull of the United States cruiser “Brooklyn”. Scientific American, 88(9):151, February 28, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n9/pdf/scientificamerican02281903-151.pdf.

Anonymous:1903:CAD

[5545] Anonymous. The construction of the Asyut Dam; Nile irrigation works. Scientific American, 88(9):152, February 28, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v88/
n9/pdf/scientificamerican02281903-152.pdf.

Bond:1903:THW

[5546] Rev. Lewis Bond. Tripping hoist for well buckets. Scientific American, 88(9):153, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
153a.pdf.

Geare:1903:VSIc

[5547] Randolph I. Geare. Venomous serpents. — III. Scientific American, 88(9):154, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
154.pdf.

Anonymous:1903:PDf

[5548] Anonymous. Patent department. Scientific American, 88(9):158, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n9/pdf/scientificamerican02281903-158.pdf.

Anonymous:1903:DPS

[5549] Anonymous. Device for preventing the switching of cows’ tails. Scientific American, 88(9):158, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
158a.pdf.

Anonymous:1903:TSF

[5550] Anonymous. Two simple fastening devices. Scientific American, 88(9):158, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
158b.pdf.

Anonymous:1903:RS

[5551] Anonymous. Rein support. Scientific American, 88(9):158, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v88/n9/pdf/scientificamerican02281903-158c.pdf.

Anonymous:1903:NIM

[5552] Anonymous. A new insulating material. Scientific American, 88 (9):158–159, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
158d.pdf.

Anonymous:1903:BPWz

[5553] Anonymous. Business and personal wants. Scientific American, 88(9):161, February 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v88/n9/pdf/scientificamerican02281903-
161.pdf.

Anonymous:1903:Cv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-8a.pdf.

Anonymous:1903:RPIba

[5555] Anonymous. Recently patented inventions. Scientific American, 89(1): 15, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-15.pdf.

Anonymous:1903:NQw

[5556] Anonymous. Notes and queries. Scientific American, 89(1):15–16, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-15b.pdf.

Anonymous:1903:IIba

[5557] Anonymous. Index of inventions. Scientific American, 89(1):16, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-16.pdf.

Anonymous:1903:UUS

[5558] Anonymous. Universities in the United States and Great Britain. Scientific American, 89(1):4, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
4.pdf.

Anonymous:1903:AE

[5559] Anonymous. The alderman and the engineer. Scientific American, 89(1):4, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
4a.pdf.

Anonymous:1903:NAR

[5560] Anonymous. The new Army rifle. Scientific American, 89(1):4, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-4b.pdf.

Anonymous:1903:GOF

[5561] Anonymous. The growth of our foreign commerce. Scientific American, 89(1):4, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
4c.pdf.

Anonymous:1903:SCT

[5562] Anonymous. A ship canal through Scotland. Scientific American, 89(1): 4, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-4d.pdf.

Anonymous:1903:DWH

[5563] Anonymous. Destroying the water hyacinth by a new chemical process. Scientific American, 89(1):4–5, July 4, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n1/pdf/scientificamerican07041903-4e.pdf.

Anonymous:1903:MR

[5564] Anonymous. Marconi in Rome. Scientific American, 89(1):5, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-5.pdf.

Anonymous:1903:ETO

[5565] Anonymous. An electric tramway operated upon the Lorain Contact Box System. Scientific American, 89(1):5, July 4, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n1/pdf/scientificamerican07041903-5a.pdf.

Anonymous:1903:UEB

[5566] Anonymous. The union engineering building. Scientific American, 89(1):5, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
5b.pdf.

Anonymous:1903:CSe

[5567] Anonymous. The current supplement. Scientific American, 89(1):5, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-5c.pdf.

Anonymous:1903:BCa

[5568] Anonymous. Borelly’s Comet. Scientific American, 89(1):5, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-5d.pdf.

Anonymous:1903:AMC

[5569] Anonymous. Alexander Melville Clark. Scientific American, 89(1):5, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-5e.pdf.

Anonymous:1903:TNF

[5570] Anonymous. The telechirograph — a new facsimile telegraph. Scientific American, 89(1):6, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
6.pdf.

Anonymous:1903:HLS

[5571] Anonymous. How large steamers are divided and taken through the Canadian Canal. Scientific American, 89(1):7, July 4, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n1/pdf/scientificamerican07041903-7.pdf.

Anonymous:1903:PFM

[5572] Anonymous. Production of fluorine in Moissan’s electrolytic apparatus. Scientific American, 89(1):7–8, July 4, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n1/pdf/scientificamerican07041903-7a.pdf.

Willey:1903:AMH

[5573] Day Allen Willey. Automatic machinery for handling coal. Scientific American, 89(1):8, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
8.pdf.

Field:1903:QDA

[5574] C. Field. A quaint diving apparatus of the eighteenth century. Scientific American, 89(1):9, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
9.pdf.

Anonymous:1903:HMP

[5575] Anonymous. How to make a paper cane. Scientific American, 89(1):9, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-9a.pdf.

Anonymous:1903:ACCb

[5576] Anonymous. American cars to compete in the Gordon–Bennett Cup Race. Scientific American, 89(1):10, July 4, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n1/pdf/scientificamerican07041903-10.pdf.

Anonymous:1903:GGC

[5577] Anonymous. The Great Goose Creek Dam for the Denver Water Supply. Scientific American, 89(1):11–12, July 4, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n1/pdf/scientificamerican07041903-11.pdf.

Anonymous:1903:PCR

[5578] Anonymous. Pullman cars in a railroad wreck. Scientific American, 89(1):12, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
12.pdf.

Anonymous:1903:PDg

[5579] Anonymous. Patent department. Scientific American, 89(1):13, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-13.pdf.

Anonymous:1903:VIC

[5580] Anonymous. Vaporizer for internal combustion engines. Scientific American, 89(1):14, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n1/pdf/scientificamerican07041903-
14.pdf.

Anonymous:1903:SCF

[5581] Anonymous. A simple chimney flue cleaner. Scientific American, 89(1): 14, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-14a.pdf.

Anonymous:1903:EMB

[5582] Anonymous. Engine for motor bicycles. Scientific American, 89(1):14, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-14b.pdf.

Anonymous:1903:OIc

[5583] Anonymous. Oddities in inventions. Scientific American, 89(1):14, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-14c.pdf.

Anonymous:1903:BPWba

[5584] Anonymous. Business and personal wants. Scientific American, 89(1): 15, July 4, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n1/pdf/scientificamerican07041903-15a.pdf.

Russell:1903:HS

[5585] Henry Norris Russell. The heavens in September. Scientific American, 89(10):163, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
163.pdf.

Anonymous:1903:LNh

[5586] Anonymous. Legal notes. Scientific American, 89(10):174, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n10/pdf/scientificamerican09051903-174.pdf.

Anonymous:1903:RPIbb

[5587] Anonymous. Recently patented inventions. Scientific American, 89(10):175, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
175.pdf.

Anonymous:1903:BNBq

[5588] Anonymous. Books: New books, etc. Scientific American, 89 (10):175, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
175c.pdf.

Anonymous:1903:NQx

[5589] Anonymous. Notes and queries. Scientific American, 89(10): 176, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
176.pdf.

Anonymous:1903:IIbb

[5590] Anonymous. Index of inventions. Scientific American, 89(10): 176–177, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
176a.pdf.

Anonymous:1903:PNPw

[5591] Anonymous. Patent notes: Patents. Scientific American, 89 (10):177, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
177.pdf.

Anonymous:1903:TS

[5592] Anonymous. The triumph of the scow. Scientific American, 89 (10):162, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
162.pdf.

Anonymous:1903:SHA

[5593] Anonymous. “scientific hooliganism” again. Scientific American, 89(10):162, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
162a.pdf.

Anonymous:1903:GSP

[5594] Anonymous. The great subway power plant. Scientific American, 89(10):162, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
162b.pdf.

Anonymous:1903:EME

[5595] Anonymous. Electricity on the Miami and Erie Canal. Scientific American, 89(10):162–163, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
162c.pdf.

Anonymous:1903:SNt

[5596] Anonymous. Science notes. Scientific American, 89(10):163, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n10/pdf/scientificamerican09051903-163a.pdf.

Anonymous:1903:TYI

[5597] Anonymous. Turbine yachts at the international races. Scientific American, 89(10):164, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
164.pdf.

Anonymous:1903:MCM

[5598] Anonymous. Manufacture of carbonated mineral waters. Scientific American, 89(10):165–166, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
165.pdf.

Anonymous:1903:GWF

[5599] Anonymous. Guarini’s wireless fire alarm apparatus. Scientific American, 89(10):166, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
166.pdf.

Anonymous:1903:HEM

[5600] Anonymous. How the Earth’s movements are noted. Scientific American, 89(10):167, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
167.pdf.

Anonymous:1903:SRR

[5601] Anonymous. The “Shamrock”–“Reliance” races of 1903. Scientific American, 89(10):168–169, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
168.pdf.

Talbot:1903:SA

[5602] Frederick A. Talbot. The Spencer airship for 1903. Scientific American, 89(10):169–170, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
169.pdf.

Balch:1903:NRH

[5603] S. W. Balch. A new record in horse trotting. Scientific American, 89(10):170, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
170.pdf.

Anonymous:1903:LTW

[5604] Anonymous. The largest transformer in the world. Scientific American, 89(10):171, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
171.pdf.

Anonymous:1903:KNG

[5605] Anonymous. Kunzite — a new gem. Scientific American, 89 (10):171, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
171a.pdf.

Hutchinson:1903:BTS

[5606] Charles E. Hutchinson. A bolas-throwing spider. Scientific American, 89(10):172, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
172.pdf.

Anonymous:1903:VSG

[5607] Anonymous. Variable speed gear for motor-driven sewing machines. Scientific American, 89(10):173, September 5, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n10/pdf/scientificamerican09051903-173.pdf.

Anonymous:1903:WMP

[5608] Anonymous. Wave motor for pumping ships. Scientific American, 89(10):173, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
173a.pdf.

Anonymous:1903:OId

[5609] Anonymous. Oddities in invention. Scientific American, 89(10): 173–174, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
173b.pdf.

Anonymous:1903:MW

[5610] Anonymous.  Marconi wireless.  Scientific American, 89(10):

175, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
175a.pdf.

Anonymous:1903:BPWbb

[5611] Anonymous. Business and personal wants. Scientific American, 89(10):175, September 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n10/pdf/scientificamerican09051903-
175b.pdf.

Anonymous:1903:RPIbc

[5612] Anonymous. Recently patented inventions. Scientific American, 89 (11):192, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
192.pdf.

Anonymous:1903:BNBr

[5613] Anonymous. Books: New books, etc. Scientific American, 89 (11):192, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
192b.pdf.

Anonymous:1903:NQy

[5614] Anonymous. Notes and queries. Scientific American, 89(11): 193, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
193.pdf.

Anonymous:1903:IIbc

[5615] Anonymous. Index of inventions. Scientific American, 89(11): 193–194, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
193a.pdf.

Anonymous:1903:AGH

[5616] Anonymous. Another great hydraulic-electric power plant. Scientific American, 89(11):182, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
182.pdf.

Anonymous:1903:LFLa

[5617] Anonymous. The last of a famous liner. Scientific American, 89 (11):182, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
182a.pdf.

Anonymous:1903:YDE

[5618] Anonymous. Is yacht designing an exact science. Scientific American, 89(11):182, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
182b.pdf.

Anonymous:1903:LETb

[5619] Anonymous. Limits of electric traction with direct current. Scientific American, 89(11):182–183, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
182c.pdf.

Anonymous:1903:ABB

[5620] Anonymous. The agreement between the British Government and the Cunard and Morgan Combine Mercantile Fleets. Scientific American, 89(11):183, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
183.pdf.

Anonymous:1903:MIA

[5621] Anonymous. More information about the prize for a dust-arresting respirator. Scientific American, 89(11):183, September 12, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n11/pdf/scientificamerican09121903-183a.pdf.

Anonymous:1903:SNu

[5622] Anonymous. Science notes. Scientific American, 89(11):183, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n11/pdf/scientificamerican09121903-183b.pdf.

Anonymous:1903:ETJ

[5623] Anonymous. Electro-telethermometer of Juan Vila Forns. Scientific American, 89(11):184, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
184.pdf.

Anonymous:1903:EAT

[5624] Anonymous. English artillery trophies at Mont St. Michel. Scientific American, 89(11):184, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
184a.pdf.

Gradenwitz:1903:NMM

[5625] A. Gradenwitz. A novel measuring machine. Scientific American, 89(11):184, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
184b.pdf.

Bosch:1903:PCT

[5626] A. Bosch. Photography in colors — the three color method. Scientific American, 89(11):185–186, September 12, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n11/pdf/scientificamerican09121903-185.pdf.

Main:1903:SFD

[5627] Walter H. Main. Spier falls dam and power plant. Scientific American, 89(11):186–187, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
186.pdf.

Guarini:1903:EAA

[5628] Emile Guarini. Electricity as an aid to agriculture — a description of two model electrical farms. Scientific American, 89(11):188–189, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n11/pdf/scientificamerican09121903-188.pdf.

Geare:1903:T

[5629] Randolph I. Geare. Tattooing. Scientific American, 89(11):189– 190, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
189.pdf.

Byers:1903:ERO

[5630] Charles Alma Byers. Elephant Rock and other bowlders. Scientific American, 89(11):191, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
191.pdf.

Lodian:1903:CP

[5631] L. Lodian. “coal-porcelain”. Scientific American, 89(11):191, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n11/pdf/scientificamerican09121903-191a.pdf.

Anonymous:1903:BPWbc

[5632] Anonymous. Business and personal wants. Scientific American, 89 (11):192, September 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n11/pdf/scientificamerican09121903-
192a.pdf.

Anonymous:1903:RPIbd

[5633] Anonymous. Recently patented inventions. Scientific American, 89 (12):208, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
208.pdf.

Anonymous:1903:NQz

[5634] Anonymous. Notes and queries. Scientific American, 89(12): 208–209, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
208b.pdf.

Anonymous:1903:IIbd

[5635] Anonymous. Index of inventions. Scientific American, 89(12): 209, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
209.pdf.

Anonymous:1903:PNPx

[5636] Anonymous. Patent notes: Patents. Scientific American, 89 (12):211, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
211.pdf.

Anonymous:1903:GEM

[5637] Anonymous. Great economy of the motor cycle. Scientific American, 89(12):198, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
198.pdf.

Anonymous:1903:EBC

[5638] Anonymous. Eye-bar cables. Scientific American, 89(12):198, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n12/pdf/scientificamerican09191903-198a.pdf.

Anonymous:1903:IBB

[5639] Anonymous. Intermediate batteries on battleships. Scientific American, 89(12):198, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
198b.pdf.

Anonymous:1903:DTP

[5640] Anonymous. Distribution of trade at the ports of the United States. Scientific American, 89(12):198, September 19, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n12/pdf/scientificamerican09191903-198c.pdf.

Anonymous:1903:CCHa

[5641] Anonymous. The cause and cure of hay fever. Scientific American, 89(12):198, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
198d.pdf.

Anonymous:1903:AI

[5642] Anonymous. Alcohol as an illuminant. Scientific American, 89 (12):198, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
198e.pdf.

Adams:1903:EER

[5643] Alton D. Adams. Expansion of electric railways in Massachusetts. Scientific American, 89(12):199, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
199.pdf.

Anonymous:1903:SNv

[5644] Anonymous. Science notes. Scientific American, 89(12):199, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n12/pdf/scientificamerican09191903-199a.pdf.

Anonymous:1903:FPb

[5645] Anonymous. A floating piston. Scientific American, 89(12): 200, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
200.pdf.

Anonymous:1903:NFV

[5646] Anonymous. A non-freezing valve. Scientific American, 89(12): 200, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
200a.pdf.

Jones:1903:SIM

[5647] H. D. Jones. Some interesting models. Scientific American, 89 (12):200–201, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
200b.pdf.

MClure:1903:APH

[5648] W. Frank M’Clure. The automobile as a plow horse. Scientific American, 89(12):201, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
201.pdf.

Dodge:1903:PMM

[5649] Charles Richard Dodge. The Pulque and Mescal of Mexico. Scientific American, 89(12):201–202, September 19, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n12/pdf/scientificamerican09191903-201a.pdf.

Anonymous:1903:BIB

[5650] Anonymous. Blackwell’s Island Bridge. Scientific American, 89(12): 202–203, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
202.pdf.

Anonymous:1903:EMD

[5651] Anonymous. Experiments with motor-driven aeroplanes. Scientific American, 89(12):204, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
204.pdf.

Holder:1903:SP

[5652] C. F. Holder. A singular profession. Scientific American, 89(12): 205–206, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
205.pdf.

Anonymous:1903:SB

[5653] Anonymous. A ship brake. Scientific American, 89(12):206, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n12/pdf/scientificamerican09191903-206.pdf.

Guarini:1903:SF

[5654] Emile Guarini. The seas of fog. Scientific American, 89(12): 207, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
207.pdf.

Anonymous:1903:BPWbd

[5655] Anonymous. Business and personal wants. Scientific American, 89 (12):208, September 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n12/pdf/scientificamerican09191903-
208a.pdf.

Anonymous:1903:Cw

[5656] Anonymous. Correspondence. Scientific American, 89(13):218– 219, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
218a.pdf.

Anonymous:1903:LNi

[5657] Anonymous. Legal notes. Scientific American, 89(13):225, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n13/pdf/scientificamerican09261903-225.pdf.

Anonymous:1903:RPIbe

[5658] Anonymous. Recently patented inventions. Scientific American, 89 (13):226–227, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
226.pdf.

Anonymous:1903:NQba

[5659] Anonymous. Notes and queries. Scientific American, 89(13): 227–228, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
227a.pdf.

Anonymous:1903:BNBs

[5660] Anonymous. Books: New books, etc. Scientific American, 89 (13):228, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
228.pdf.

Anonymous:1903:IIbe

[5661] Anonymous. Index of inventions. Scientific American, 89(13): 228–229, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
228a.pdf.

Anonymous:1903:PNPy

[5662] Anonymous. Patent notes: Patents. Scientific American, 89 (13):231, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
231.pdf.

Anonymous:1903:LLU

[5663] Anonymous. Life is longer under modern conditions. Scientific American, 89(13):214, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
214.pdf.

Anonymous:1903:ETT

[5664] Anonymous. Elevated and trolley tracks on the new bridges. Scientific American, 89(13):214, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
214a.pdf.

Anonymous:1903:STOa

[5665] Anonymous. Steam turbine for ocean liners. Scientific American, 89(13):214, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
214b.pdf.

Anonymous:1903:MACb

[5666] Anonymous. Merchants’ Association on city transit. Scientific American, 89(13):214, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
214c.pdf.

Anonymous:1903:TEB

[5667] Anonymous. Torpedo experiments with the “Belleisle” by the British Admiralty. Scientific American, 89(13):214–215, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n13/pdf/scientificamerican09261903-214d.pdf.

Cline:1903:EWF

[5668] V. E. Cline. Eucalypts and the world’s fuel. Scientific American, 89(13):215, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
215.pdf.

Anonymous:1903:SNw

[5669] Anonymous. Science notes. Scientific American, 89(13):215, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n13/pdf/scientificamerican09261903-215a.pdf.

Anonymous:1903:NSBa

[5670] Anonymous. A novel spiral bridge. Scientific American, 89(13): 216, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
216.pdf.

Anonymous:1903:BA

[5671] Anonymous. The Berliner aeroplane. Scientific American, 89 (13):216, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
216a.pdf.

Anonymous:1903:RLL

[5672] Anonymous. Revolving lanterns for lighthouses. Scientific American, 89(13):217–218, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
217.pdf.

Anonymous:1903:BCP

[5673] Anonymous. The Bessemer copper process. Scientific American, 89(13):218, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
218.pdf.

Anonymous:1903:Hb

[5674] Anonymous. The hyposcope. Scientific American, 89(13):220, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n13/pdf/scientificamerican09261903-220.pdf.

Anonymous:1903:CTW

[5675] Anonymous. Communication with trains by wireless telegraphy. Scientific American, 89(13):220–221, September 26, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n13/pdf/scientificamerican09261903-220a.pdf.

Willoughby:1903:PNN

[5676] Charles C. Willoughby. Piedras Negras, a newly discovered prehistoric city in Guatemala. Scientific American, 89(13):221–222, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n13/pdf/scientificamerican09261903-221.pdf.

Backes:1903:CNW

[5677] Jacob Backes. Callitypy — a new way of using the typewriter. Scientific American, 89(13):222, September 26, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n13/pdf/scientificamerican09261903-222.pdf.

Anonymous:1903:TT

[5678] Anonymous. A trick with the telephone. Scientific American, 89 (13):223, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
223.pdf.

Anonymous:1903:IRJ

[5679] Anonymous. Improved rail joint. Scientific American, 89(13): 223, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
223a.pdf.

Anonymous:1903:CCS

[5680] Anonymous. Combined carpet stretcher and tacker. Scientific American, 89(13):223, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
223b.pdf.

Anonymous:1903:FDV

[5681] Anonymous. Fastening device for vent pipes. Scientific American, 89(13):223, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
223c.pdf.

Anonymous:1903:OIe

[5682] Anonymous. Oddities in invention. Scientific American, 89(13): 224, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
224.pdf.

Anonymous:1903:BPWbe

[5683] Anonymous. Business and personal wants. Scientific American, 89 (13):227, September 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n13/pdf/scientificamerican09261903-
227.pdf.

Russell:1903:HO

[5684] Henry Norris Russell. The heavens in October. Scientific American, 89(14):235, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
235.pdf.

Anonymous:1903:RPIbf

[5685] Anonymous. Recently patented inventions. Scientific American, 89(14):246, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
246.pdf.

Anonymous:1903:NQbb

[5686] Anonymous. Notes and queries. Scientific American, 89(14):246–247, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n14/pdf/scientificamerican10031903-246c.pdf.

Anonymous:1903:BNBt

[5687] Anonymous. Books: New books, etc. Scientific American, 89(14):247, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n14/pdf/scientificamerican10031903-247.pdf.

Anonymous:1903:IIbf

[5688] Anonymous. Index of inventions. Scientific American, 89(14):247, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n14/pdf/scientificamerican10031903-247a.pdf.

Anonymous:1903:PNPz

[5689] Anonymous. Patent notes: Patents. Scientific American, 89(14):250, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n14/pdf/scientificamerican10031903-250.pdf.

Anonymous:1903:FWS

[5690] Anonymous. The future water supply of New York City. Scientific American, 89(14):234, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
234.pdf.

Anonymous:1903:PGV

[5691] Anonymous. Pipe galleries vs. upturned streets. Scientific American, 89(14):234, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
234a.pdf.

Anonymous:1903:LSD

[5692] Anonymous. At last, the side-door passenger car. Scientific American, 89(14):234, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
234b.pdf.

Carden:1903:HPW

[5693] Golfrey L. Carden. The 60,000 horsepower plant at the World’s Fair. Scientific American, 89(14):234–235, October 3, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n14/pdf/scientificamerican10031903-234c.pdf.

Anonymous:1903:SPMa

[5694] Anonymous. The Sievert process of mechanical glass blowing. Scientific American, 89(14):236, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
236.pdf.

Anonymous:1903:LGP

[5695] Anonymous. London’s gasoline-propelled cabs. Scientific American, 89(14):237, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
237.pdf.

Anonymous:1903:NT

[5696] Anonymous. The new telegraphone. Scientific American, 89 (14):237–238, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
237a.pdf.

Perkins:1903:EFD

[5697] Frank C. Perkins. European floating and dock cranes. Scientific American, 89(14):239, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
239.pdf.

Anonymous:1903:MCS

[5698] Anonymous. Moving the columns of St. John the Divine. Scientific American, 89(14):240, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
240.pdf.

Fawcett:1903:LPG

[5699] Waldon Fawcett. Latest progress in good roads construction. Scientific American, 89(14):241–242, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
241.pdf.

Anonymous:1903:NTE

[5700] Anonymous. A nearly total eclipse of the Moon observed by means of photoelectric selenium cells. Scientific American, 89(14):242, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n14/pdf/scientificamerican10031903-242.pdf.

Anonymous:1903:SCO

[5701] Anonymous. Some curious oyster growths. Scientific American, 89(14):243, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
243.pdf.

Jones:1903:SAF

[5702] George J. Jones. Stove with an adjustable firebox. Scientific American, 89(14):244–245, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
244.pdf.

Anonymous:1903:ICCa

[5703] Anonymous. Illuminated cross for church towers. Scientific American, 89(14):245, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
245.pdf.

Anonymous:1903:DPW

[5704] Anonymous. Device for preventing waste of water. Scientific American, 89(14):245, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
245a.pdf.

Anonymous:1903:FSV

[5705] Anonymous. Folding step for vestibule cars. Scientific American, 89(14):245, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
245b.pdf.

Anonymous:1903:LPT

[5706] Anonymous. A low-priced typewriter. Scientific American, 89(14):246, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n14/pdf/scientificamerican10031903-246a.pdf.

Anonymous:1903:BPWbf

[5707] Anonymous. Business and personal wants. Scientific American, 89(14):246, October 3, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n14/pdf/scientificamerican10031903-
246b.pdf.

Anonymous:1903:RPIbg

[5708] Anonymous. Recently patented inventions. Scientific American, 89(15):264, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
264.pdf.

Anonymous:1903:NQbc

[5709] Anonymous. Notes and queries. Scientific American, 89(15):264, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n15/pdf/scientificamerican10101903-264b.pdf.

Anonymous:1903:BNBu

[5710] Anonymous. Books: New books, etc. Scientific American, 89 (15):265, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
265.pdf.

Anonymous:1903:IIbg

[5711] Anonymous. Index of inventions. Scientific American, 89(15):265, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n15/pdf/scientificamerican10101903-265a.pdf.

Anonymous:1903:PNPba

[5712] Anonymous. Patent notes: Patents. Scientific American, 89 (15):266, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
266.pdf.

Anonymous:1903:DDC

[5713] Anonymous. The double-deck car for rapid transit. Scientific American, 89(15):254, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
254.pdf.

Anonymous:1903:KHC

[5714] Anonymous. Keeping the health of a city. Scientific American, 89 (15):254–255, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
254a.pdf.

Weir:1903:CCE

[5715] James Weir. Certain curious effects of the violet rays. Scientific American, 89(15):255, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
255.pdf.

Anonymous:1903:FNF

[5716] Anonymous. Fruit and nuts as food. Scientific American, 89 (15):255, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
255a.pdf.

Anonymous:1903:SNx

[5717] Anonymous. Science notes. Scientific American, 89(15):255, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n15/pdf/scientificamerican10101903-255b.pdf.

Guarini:1903:FBM

[5718] Emile Guarini. The flight of birds mechanically studied. Scientific American, 89(15):256–257, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
256.pdf.

Anonymous:1903:BBCb

[5719] Anonymous. The Barney brick conveyer. Scientific American, 89 (15):257–258, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
257.pdf.

Bosch:1903:PDM

[5720] A Bosch. The petroleum districts and mud-volcanoes of Java. Scientific American, 89(15):259, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
259.pdf.

Fullerton:1903:LHC

[5721] A. W. Fullerton. The largest hydraulic canal locks in the world. Scientific American, 89(15):260, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
260.pdf.

Anonymous:1903:ISM

[5722] Anonymous. An ingenious sculpturing machine. Scientific American, 89(15):260–262, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
260a.pdf.

Anonymous:1903:NA

[5723] Anonymous. A new aeroplane. Scientific American, 89(15):262, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n15/pdf/scientificamerican10101903-262.pdf.

Holder:1903:IG

[5724] Charles F. Holder. Indian granaries. Scientific American, 89 (15):263, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
263.pdf.

Anonymous:1903:BPWbg

[5725] Anonymous. Business and personal wants. Scientific American, 89(15):264, October 10, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n15/pdf/scientificamerican10101903-
264a.pdf.

Anonymous:1903:Cx

[5726] Anonymous. Correspondence. Scientific American, 89(16):275, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n16/pdf/scientificamerican10171903-275.pdf.

Anonymous:1903:RPIbh

[5727] Anonymous. Recently patented inventions. Scientific American, 89(16):280, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
280.pdf.

Anonymous:1903:NQbd

[5728] Anonymous. Notes and queries. Scientific American, 89(16):280, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n16/pdf/scientificamerican10171903-280b.pdf.

Anonymous:1903:BNBv

[5729] Anonymous. Books: New books, etc. Scientific American, 89 (16):280, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
280c.pdf.

Anonymous:1903:IIbh

[5730] Anonymous. Index of inventions. Scientific American, 89(16): 280–281, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
280d.pdf.

Anonymous:1903:PNPbb

[5731] Anonymous. Patent notes: Patents. Scientific American, 89 (16):281, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
281.pdf.

Anonymous:1903:OHT

[5732] Anonymous. One hundred and twenty-five miles an hour — a new railroad record. Scientific American, 89(16):270, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n16/pdf/scientificamerican10171903-270.pdf.

Anonymous:1903:DI

[5733] Anonymous. Dangers in irrigation. Scientific American, 89(16):270, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n16/pdf/scientificamerican10171903-270a.pdf.

Anonymous:1903:IRS

[5734] Anonymous. India rubber and its substitutes. Scientific American, 89(16):270, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
270b.pdf.

Anonymous:1903:NZE

[5735] Anonymous. News from the Ziegler Expedition. Scientific American, 89(16):270, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
270c.pdf.

Anonymous:1903:NATb

[5736] Anonymous. New automobile track records. Scientific American, 89(16):271, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
271.pdf.

Anonymous:1903:FDR

[5737] Anonymous. First day’s run of the New York–Pittsburgh Automobile Endurance Test. Scientific American, 89(16):271, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n16/pdf/scientificamerican10171903-271a.pdf.

Anonymous:1903:QVI

[5738] Anonymous. Queen Victoria’s Ivory Chair of State. Scientific American, 89(16):271, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
271b.pdf.

Anonymous:1903:BMB

[5739] Anonymous. Battleship “Missouri’s” builders’ trial. Scientific American, 89(16):271, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
271c.pdf.

Anonymous:1903:SNy

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n16/pdf/scientificamerican10171903-271d.pdf.

Anonymous:1903:FLA

[5741] Anonymous. The failure of Langley’s aerodrome. Scientific American, 89(16):272, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
272.pdf.

Anonymous:1903:EHa

[5742] Anonymous. Electric heater. Scientific American, 89(16):273, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n16/pdf/scientificamerican10171903-273.pdf.

Anonymous:1903:MDC

[5743] Anonymous. Machine for drying cocoa beans. Scientific American, 89(16):273, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
273a.pdf.

Anonymous:1903:KTU

[5744] Anonymous. The Kamm typewriter for use with wireless telegraphy. Scientific American, 89(16):273–274, October 17, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n16/pdf/scientificamerican10171903-273b.pdf.

Anonymous:1903:LBU

[5745] Anonymous. The latest battleships for the United States Navy. Scientific American, 89(16):274, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
274.pdf.

Anonymous:1903:CCa

[5746] Anonymous. Chimney cap. Scientific American, 89(16):275, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n16/pdf/scientificamerican10171903-275a.pdf.

Anonymous:1903:STG

[5747] Anonymous. The sub-target gun machine. Scientific American, 89(16):276, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
276.pdf.

Fawcett:1903:TL

[5748] Waldon Fawcett. The transportation of lumber. Scientific American, 89(16):276–278, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
276a.pdf.

Anonymous:1903:ERL

[5749] Anonymous. An early road locomotive. Scientific American, 89 (16):278, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
278.pdf.

Anonymous:1903:CSF

[5750] Anonymous. A children’s school farm in the heart of a great city. Scientific American, 89(16):279, October 17, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n16/pdf/scientificamerican10171903-279.pdf.

Anonymous:1903:LAG

[5751] Anonymous. The latest applications of the gasoline automobile motor. Scientific American, 89(16):279, October 17, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n16/pdf/scientificamerican10171903-279a.pdf.

Anonymous:1903:BPWbh

[5752] Anonymous. Business and personal wants. Scientific American, 89(16):280, October 17, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n16/pdf/scientificamerican10171903-
280a.pdf.

Anonymous:1903:Cy

[5753] Anonymous. Correspondence. Scientific American, 89(17):291, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n17/pdf/scientificamerican10241903-291.pdf.

Anonymous:1903:RPIbi

[5754] Anonymous. Recently patented inventions. Scientific American, 89 (17):296–297, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
296.pdf.

Anonymous:1903:NQbe

[5755] Anonymous. Notes and queries. Scientific American, 89(17):297, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n17/pdf/scientificamerican10241903-297a.pdf.

Anonymous:1903:BNBw

[5756] Anonymous. Books: New books, etc. Scientific American, 89 (17):297, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
297b.pdf.

Anonymous:1903:IIbi

[5757] Anonymous. Index of inventions. Scientific American, 89(17): 297–298, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
297c.pdf.

Anonymous:1903:PNPbc

[5758] Anonymous. Patent notes: Patents. Scientific American, 89 (17):298, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
298.pdf.

Anonymous:1903:EET

[5759] Anonymous. The economy of electric traction. Scientific American, 89(17):286, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
286.pdf.

Anonymous:1903:SRPb

[5760] Anonymous. The storage reservoir as a preventive of floods. Scientific American, 89(17):286, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
286a.pdf.

Anonymous:1903:FTA

[5761] Anonymous. The first turbine Atlantic liner. Scientific American, 89(17):286, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
286b.pdf.

Anonymous:1903:TR

[5762] Anonymous. The telephone and the railway. Scientific American, 89(17):286–287, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
286c.pdf.

Anonymous:1903:BQE

[5763] Anonymous. The British quest for an eighteenpenny safety lamp. Scientific American, 89(17):287, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
287.pdf.

Anonymous:1903:NGR

[5764] Anonymous. New German red phosphorus match. Scientific American, 89(17):287, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
287a.pdf.

Anonymous:1903:CSg

[5765] Anonymous. The current supplement. Scientific American, 89 (17):287, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
287b.pdf.

Anonymous:1903:AFP

[5766] Anonymous. Aluminium in flowering plants. Scientific American, 89(17):287, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
287c.pdf.

Anonymous:1903:SSP

[5767] Anonymous. Sun spots photographed. Scientific American, 89 (17):287, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
287d.pdf.

Anonymous:1903:FT

[5768] Anonymous. A flooded track. Scientific American, 89(17):288, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n17/pdf/scientificamerican10241903-288.pdf.

Anonymous:1903:CRA

[5769] Anonymous. Construction of the railroad across Salt Lake. Scientific American, 89(17):288, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
288a.pdf.

Willey:1903:SHC

[5770] Day Allen Willey. Sanitary house cleaning. Scientific American, 89 (17):289–290, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
289.pdf.

Moore:1903:WFC

[5771] Frederick Moore. Work of the Fisheries Commission. Scientific American, 89(17):290–291, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
290.pdf.

Anonymous:1903:DNY

[5772] Anonymous. The disposal of New York’s refuse. Scientific American, 89(17):292–294, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
292.pdf.

Anonymous:1903:HEL

[5773] Anonymous. A heavy electric locomotive. Scientific American, 89(17):294, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
294.pdf.

Beasley:1903:RCG

[5774] Walter L. Beasley. Remarkable ceremonial garments. Scientific American, 89(17):295, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
295.pdf.

Anonymous:1903:BPWbi

[5775] Anonymous. Business and personal wants. Scientific American, 89(17):297, October 24, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n17/pdf/scientificamerican10241903-
297.pdf.

Anonymous:1903:LNj

[5776] Anonymous. Legal notes. Scientific American, 89(18):312, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n18/pdf/scientificamerican10311903-312.pdf.

Anonymous:1903:RPIbj

[5777] Anonymous. Recently patented inventions. Scientific American, 89 (18):313–314, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
313.pdf.

Anonymous:1903:NQbf

[5778] Anonymous. Notes and queries. Scientific American, 89(18): 314–315, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
314a.pdf.

Anonymous:1903:BNBx

[5779] Anonymous. Books: New books, etc. Scientific American, 89 (18):315–316, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
315.pdf.

Anonymous:1903:IIbj

[5780] Anonymous. Index of inventions. Scientific American, 89(18): 316–317, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
316.pdf.

Anonymous:1903:PNPbd

[5781] Anonymous. Patent notes: Patents. Scientific American, 89 (18):318, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
318.pdf.

Anonymous:1903:TCSa

[5782] Anonymous. A transatlantic clipper service. Scientific American, 89(18):302, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
302.pdf.

Anonymous:1903:WRH

[5783] Anonymous. Wind resistance in high-speed train service. Scientific American, 89(18):302, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
302a.pdf.

Anonymous:1903:PDAa

[5784] Anonymous. Prehistoric drawings of animals, etc., in the Cave of Altamira, Spain. Scientific American, 89(18):302–303, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n18/pdf/scientificamerican10311903-302b.pdf.

Anonymous:1903:GSE

[5785] Anonymous. The Ganz system of electric canal haulage. Scientific American, 89(18):303, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
303.pdf.

Anonymous:1903:HMA

[5786] Anonymous. “home mechanics for amateurs”. Scientific American, 89(18):303, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
303a.pdf.

Anonymous:1903:CSh

[5787] Anonymous. The current supplement. Scientific American, 89 (18):303, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
303b.pdf.

Anonymous:1903:SNz

[5788] Anonymous. Science notes. Scientific American, 89(18):303, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n18/pdf/scientificamerican10311903-303c.pdf.

Anonymous:1903:ALL

[5789] Anonymous. Another “largest locomotive”. Scientific American, 89(18):304, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
304.pdf.

Harris:1903:MFA

[5790] R. A. Harris. A machine for facilitating the analysis of curves representing periodic phenomena. Scientific American, 89(18):304–305, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n18/pdf/scientificamerican10311903-304a.pdf.

Anonymous:1903:TNY

[5791] Anonymous. Termination of the New York–Pittsburgh Automobile Endurance Test. Scientific American, 89(18):305–307, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n18/pdf/scientificamerican10311903-305.pdf.

Anonymous:1903:HRTa

[5792] Anonymous. Harlem River Tunnel of the Rapid Transit Subway. Scientific American, 89(18):307, October 31, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n18/pdf/scientificamerican10311903-307.pdf.

Willey:1903:BSG

[5793] Day Allen Willey. Baltimore’s system of garbage disposal. Scientific American, 89(18):308, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
308.pdf.

Anonymous:1903:TBD

[5794] Anonymous. The torpedo-boat destroyer “Perry”. Scientific American, 89(18):309, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
309.pdf.

Anonymous:1903:LDR

[5795] Anonymous. The long-distance record run of an American electric automobile. Scientific American, 89(18):310, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n18/pdf/scientificamerican10311903-310.pdf.

Anonymous:1903:SS

[5796] Anonymous. Sun spots. Scientific American, 89(18):310, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n18/pdf/scientificamerican10311903-310a.pdf.

Anonymous:1903:NTH

[5797] Anonymous. A new type of hollow grate-bar. Scientific American, 89(18):311, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
311.pdf.

Anonymous:1903:OIf

[5798] Anonymous. Oddities in invention. Scientific American, 89(18): 311–312, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
311a.pdf.

Anonymous:1903:BPWbj

[5799] Anonymous. Business and personal wants. Scientific American, 89(18):314, October 31, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n18/pdf/scientificamerican10311903-
314.pdf.

Russell:1903:HN

[5800] Henry Norris Russell. The heavens in November. Scientific American, 89(19):323, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
323a.pdf.

Anonymous:1903:IIbk

[5801] Anonymous. Index of inventions. Scientific American, 89(19): 332, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
332a.pdf.

Anonymous:1903:NCc

[5802] Anonymous. The new Cunarders. Scientific American, 89(19): 322, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
322.pdf.

Anonymous:1903:SCC

[5803] Anonymous. Steel-concrete construction. Scientific American, 89 (19):322, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
322a.pdf.

Anonymous:1903:NFG

[5804] Anonymous. New field gun for the British Army. Scientific American, 89(19):322, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
322b.pdf.

Anonymous:1903:DPR

[5805] Anonymous. The death of Prof. R. H. Thurston. Scientific American, 89(19):323, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
323.pdf.

Anonymous:1903:SPE

[5806] Anonymous. Some powerful English express locomotives. Scientific American, 89(19):324, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
324.pdf.

Anonymous:1903:HL

[5807] Anonymous. The handling of locomotives. Scientific American, 89(19):324, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
324a.pdf.

Anonymous:1903:STP

[5808] Anonymous. Steltje’s type-printing telegraph. Scientific American, 89(19):325, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
325.pdf.

Anonymous:1903:PPT

[5809] Anonymous. A portable pneumatic tool outfit for railroads. Scientific American, 89(19):325–326, November 7, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n19/pdf/scientificamerican11071903-325a.pdf.

Guarini:1903:ROP

[5810] Emile Guarini. Recent observations of the Planet Mercury. Scientific American, 89(19):326–327, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
326.pdf.

Inkersley:1903:DGA

[5811] Arthur Inkersley. Dr. Greth’s airship. Scientific American, 89 (19):327, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
327.pdf.

Anonymous:1903:COS

[5812] Anonymous. A California onion-seed farm. Scientific American, 89 (19):328–329, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
328.pdf.

Anonymous:1903:NSDa

[5813] Anonymous. The new Santos-Dumont airship. Scientific American, 89(19):329, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
329.pdf.

Anonymous:1903:HTS

[5814] Anonymous. A house that turns with the Sun. Scientific American, 89(19):330, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
330.pdf.

Shepstone:1903:SNH

[5815] Harold J. Shepstone. Some new hybrids. Scientific American, 89 (19):331–332, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
331.pdf.

Anonymous:1903:BPWbk

[5816] Anonymous. Business and personal wants. Scientific American, 89(19):332, November 7, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n19/pdf/scientificamerican11071903-
332.pdf.

Anonymous:1903:Cz

[5817] Anonymous. Correspondence. Scientific American, 89(2):27, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-27.pdf.

Anonymous:1903:RPIbk

[5818] Anonymous. Recently patented inventions. Scientific American, 89(2):32–33, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
32.pdf.

Anonymous:1903:NQbg

[5819] Anonymous. Notes and queries. Scientific American, 89(2):34–35, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-34.pdf.

Anonymous:1903:BNBy

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-35.pdf.

Anonymous:1903:IIbl

[5821] Anonymous. Index of inventions. Scientific American, 89(2):36, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-36.pdf.

Anonymous:1903:PNPbe

[5822] Anonymous. Patent notes: Patents. Scientific American, 89(2):37, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-37.pdf.

Anonymous:1903:DCR

[5823] Anonymous. Destruction of city refuse. Scientific American, 89(2):22, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-22.pdf.

Anonymous:1903:STOb

[5824] Anonymous. The steam turbine for ocean service. Scientific American, 89(2):22, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
22a.pdf.

Anonymous:1903:GOR

[5825] Anonymous. Growth of our railroad system. Scientific American, 89(2):22, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
22b.pdf.

Anonymous:1903:PUR

[5826] Anonymous. Progress of the Uganda Railroad. Scientific American, 89(2):22–23, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
22c.pdf.

Marriott:1903:NS

[5827] Crittenden Marriott. Nickel-steel. Scientific American, 89(2):23, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-23.pdf.

Anonymous:1903:POR

[5828] Anonymous. A prize offered for a respirator. Scientific American, 89(2):23, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
23a.pdf.

Anonymous:1903:SNba

[5829] Anonymous. Science notes. Scientific American, 89(2):23, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-23b.pdf.

Anonymous:1903:CSi

[5830] Anonymous. The current supplement. Scientific American, 89(2):23, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n2/pdf/scientificamerican07111903-23c.pdf.

Anonymous:1903:NCD

[5831] Anonymous. The new Calais–Dover turbine steamer. Scientific American, 89(2):24, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
24.pdf.

Anonymous:1903:NSDb

[5832] Anonymous. The new Santos-Dumont airships. Scientific American, 89(2):24–25, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
24a.pdf.

Anonymous:1903:MSS

[5833] Anonymous. A mechanical sheep-shearer. Scientific American, 89(2):25–26, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
25.pdf.

Anonymous:1903:WAP

[5834] Anonymous. The water-absorbing properties of plants. Scientific American, 89(2):26, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
26.pdf.

Herr:1903:IDD

[5835] Shirl Herr. An instrument for detecting delicate acoustic vibrations. Scientific American, 89(2):26, July 11, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n2/pdf/scientificamerican07111903-26a.pdf.

Anonymous:1903:PHP

[5836] Anonymous. Philadelphia’s high-pressure fire pipe line. Scientific American, 89(2):28, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
28.pdf.

Anonymous:1903:TFD

[5837] Anonymous. A 6,000-ton floating depositing pontoon dock. Scientific American, 89(2):29–30, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
29.pdf.

Beard:1903:ASM

[5838] J. Carter Beard. The amœba — a slime monster and its victim. Scientific American, 89(2):31, July 11, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n2/pdf/scientificamerican07111903-31.pdf.

Anonymous:1903:BPWbl

[5839] Anonymous. Business and personal wants. Scientific American, 89(2):33–34, July 11, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n2/pdf/scientificamerican07111903-
33.pdf.

Anonymous:1903:Cba

[5840] Anonymous. Correspondence. Scientific American, 89(20):348, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n20/pdf/scientificamerican11141903-348.pdf.

Anonymous:1903:RPIbl

[5841] Anonymous. Recently patented inventions. Scientific American, 89(20):353, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
353.pdf.

Anonymous:1903:NQbh

[5842] Anonymous. Notes and queries. Scientific American, 89(20): 357, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
357.pdf.

Anonymous:1903:PNPbf

[5843] Anonymous. Patent notes: Patents. Scientific American, 89 (20):358, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
358.pdf.

Anonymous:1903:BNBz

[5844] Anonymous. Books: New books, etc. Scientific American, 89(20): 358–359, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
358a.pdf.

Anonymous:1903:IIbm

[5845] Anonymous. Index of inventions. Scientific American, 89(20): 359, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
359.pdf.

Anonymous:1903:MPM

[5846] Anonymous. Modern printing methods. Scientific American, 89 (20):338, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
338.pdf.

Anonymous:1903:HNP

[5847] Anonymous. How a newspaper is produced. Scientific American, 89 (20):338–339, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
338a.pdf.

Anonymous:1903:MTV

[5848] Anonymous. A morning with Theodore L. De Vinne. Scientific American, 89(20):339, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
339.pdf.

Anonymous:1903:MPP

[5849] Anonymous. The manufacture of paper and paper pulp. Scientific American, 89(20):340, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
340.pdf.

Anonymous:1903:TF

[5850] Anonymous. Type founding. Scientific American, 89(20):340– 341, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
340a.pdf.

Anonymous:1903:A

[5851] Anonymous. The autoplate. Scientific American, 89(20):341, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n20/pdf/scientificamerican11141903-341.pdf.

Anonymous:1903:SEc

[5852] Anonymous. Stereotyping and electrotyping. Scientific American, 89(20):341, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
341a.pdf.

Anonymous:1903:MC

[5853] Anonymous. Machine composition. Scientific American, 89(20): 341–342, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
341b.pdf.

Anonymous:1903:Ga

[5854] Anonymous. The graphotype. Scientific American, 89(20):342– 343, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
342.pdf.

Anonymous:1903:IMP

[5855] Anonymous. The invention of the modern press. Scientific American, 89(20):343, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
343.pdf.

Anonymous:1903:NP

[5856] Anonymous. The newspaper press. Scientific American, 89(20): 343, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
343a.pdf.

Anonymous:1903:MBP

[5857] Anonymous. Magazine and book presses. Scientific American, 89 (20):343–344, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
343b.pdf.

Anonymous:1903:AJP

[5858] Anonymous. Automatic jobbing press. Scientific American, 89 (20):344, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
344.pdf.

Anonymous:1903:MPD

[5859] Anonymous. Mechanical “printer’s devils”. Scientific American, 89(20):345, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
345.pdf.

Anonymous:1903:CFW

[5860] Anonymous. Combination folding and wire-stitching machine. Scientific American, 89(20):346, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
346.pdf.

Anonymous:1903:BCM

[5861] Anonymous. Book-covering machine for applying paper covers to books, pamphlets, and magazines. Scientific American, 89(20):346, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n20/pdf/scientificamerican11141903-346a.pdf.

Anonymous:1903:B

[5862] Anonymous. Bookbinding. Scientific American, 89(20):346–348, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n20/pdf/scientificamerican11141903-346b.pdf.

Guarini:1903:D

[5863] Emile Guarini. The dupligraph. Scientific American, 89(20): 349, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
349.pdf.

Anonymous:1903:TCP

[5864] Anonymous. Three-color process. Scientific American, 89(20): 349, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
349a.pdf.

Anonymous:1903:SASb

[5865] Anonymous. The Scientific American and the supplement. Scientific American, 89(20):350, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
350.pdf.

Anonymous:1903:DCN

[5866] Anonymous. Delay in the construction of Naval ships. Scientific American, 89(20):350, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
350a.pdf.

Anonymous:1903:DPM

[5867] Anonymous. Death of Prof. Mommsen. Scientific American, 89 (20):350, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
350b.pdf.

Anonymous:1903:TCSb

[5868] Anonymous. On the theory of the critical state. Scientific American, 89(20):350, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
350c.pdf.

Anonymous:1903:GCR

[5869] Anonymous. A graphical comparison of raw and finished products in the printing and paper-making trades. Scientific American, 89 (20):351, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
351.pdf.

Anonymous:1903:SOCa

[5870] Anonymous. Some of our common snakes — I. Scientific American, 89(20):352–353, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
352.pdf.

Anonymous:1903:BPWbm

[5871] Anonymous. Business and personal wants. Scientific American, 89(20):356, November 14, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n20/pdf/scientificamerican11141903-
356.pdf.

Anonymous:1903:Cbb

[5872] Anonymous. Correspondence. Scientific American, 89(21):371, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n21/pdf/scientificamerican11211903-371.pdf.

Anonymous:1903:RPIbm

[5873] Anonymous. Recently patented inventions. Scientific American, 89(21):376, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
376.pdf.

Anonymous:1903:BNBba

[5874] Anonymous. Books: New books, etc. Scientific American, 89 (21):376, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
376b.pdf.

Anonymous:1903:IIbn

[5875] Anonymous. Index of inventions. Scientific American, 89(21): 376–377, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
376c.pdf.

Anonymous:1903:PNPbg

[5876] Anonymous. Patent notes: Patents. Scientific American, 89 (21):379, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
379.pdf.

Anonymous:1903:GDH

[5877] Anonymous. Garbage disposal and the harbor channels. Scientific American, 89(21):366, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
366.pdf.

Anonymous:1903:WBF

[5878] Anonymous. The Williamsburg Bridge fiasco. Scientific American, 89(21):366, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
366a.pdf.

Anonymous:1903:ITB

[5879] Anonymous. Improvements in tunnel-boring practice in London for deep-level railroads. Scientific American, 89(21):366–367, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n21/pdf/scientificamerican11211903-366b.pdf.

Anonymous:1903:WEH

[5880] Anonymous. The Workman Expedition to the Himalayas — an altitude of 23,394 feet reached. Scientific American, 89(21):367, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n21/pdf/scientificamerican11211903-367.pdf.

Anonymous:1903:SNbb

[5881] Anonymous. Science notes. Scientific American, 89(21):367, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n21/pdf/scientificamerican11211903-367a.pdf.

Collins:1903:AC

[5882] A. Frederick Collins. Artificial camphor. Scientific American, 89 (21):368, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
368.pdf.

Anonymous:1903:JR

[5883] Anonymous. The Jungfrau Railway. Scientific American, 89(21): 369–370, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
369.pdf.

Anonymous:1903:SOCb

[5884] Anonymous. Some of our common snakes — II. Scientific American, 89(21):372, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
372.pdf.

Hayward:1903:TFP

[5885] C. D. Hayward. The Talamancans — a forgotten people of Panama. Scientific American, 89(21):373–374, November 21, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n21/pdf/scientificamerican11211903-373.pdf.

Stovall:1903:HEH

[5886] Dennis H. Stovall. Hydraulic elevators for hoisting placer tailings. Scientific American, 89(21):374, November 21, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n21/pdf/scientificamerican11211903-374.pdf.

Anonymous:1903:PDh

[5887] Anonymous. Patent department. Scientific American, 89(21): 375, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
375.pdf.

Anonymous:1903:IPR

[5888] Anonymous. Improved packing ring. Scientific American, 89 (21):375, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
375a.pdf.

Anonymous:1903:TTC

[5889] Anonymous. Tobacco tamper and cover for tobacco pipe. Scientific American, 89(21):375, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
375b.pdf.

Anonymous:1903:ICCb

[5890] Anonymous. An improved construction for coherers. Scientific American, 89(21):375, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
375c.pdf.

Anonymous:1903:BPWbn

[5891] Anonymous. Business and personal wants. Scientific American, 89(21):376, November 21, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n21/pdf/scientificamerican11211903-
376a.pdf.

Anonymous:1903:Cbc

[5892] Anonymous. Correspondence. Scientific American, 89(22):387, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n22/pdf/scientificamerican11281903-387.pdf.

Anonymous:1903:LNk

[5893] Anonymous. Legal notes. Scientific American, 89(22):392, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n22/pdf/scientificamerican11281903-392.pdf.

Anonymous:1903:RPIbn 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
393.pdf.

Anonymous:1903:NQbi

[5895] Anonymous. Notes and queries. Scientific American, 89(22): 395, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
395a.pdf.

Anonymous:1903:BNBbb

[5896] Anonymous. Books: New books, etc. Scientific American, 89 (22):395, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
395b.pdf.

Anonymous:1903:IIbo

[5897] Anonymous. Index of inventions. Scientific American, 89(22): 395–396, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
395c.pdf.

Anonymous:1903:PNPbh

[5898] Anonymous. Patent notes: Patents. Scientific American, 89 (22):399, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
399.pdf.

Anonymous:1903:TAB

[5899] Anonymous. The truth about the Belleville Boiler. Scientific American, 89(22):382, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
382.pdf.

Anonymous:1903:FMC

[5900] Anonymous. Fireproof metallic cars and a protected third rail urgently needed on electric elevated and tunnel railways. Scientific American, 89(22):382, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
382a.pdf.

Anonymous:1903:GTL

[5901] Anonymous. Government tests of the lake submarine boat. Scientific American, 89(22):382, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
382b.pdf.

Anonymous:1903:PPL

[5902] Anonymous. Photographic prints without light. Scientific American, 89(22):382–383, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
382c.pdf.

Hoff:1903:NVN

[5903] Arthur Bainbridge Hoff. Naming the vessels of the Navy. Scientific American, 89(22):383, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
383.pdf.

Anonymous:1903:SNbc

[5904] Anonymous. Science notes. Scientific American, 89(22):383, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n22/pdf/scientificamerican11281903-383a.pdf.

Anonymous:1903:SLG

[5905] Anonymous. Some large glass vessels. Scientific American, 89 (22):384, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
384.pdf.

Baker:1903:SAC

[5906] Thomas R. Baker. A still for amateur and class-room work. Scientific American, 89(22):384, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
384a.pdf.

Anonymous:1903:TNH

[5907] Anonymous. Two new hack saws. Scientific American, 89(22): 384–385, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
384b.pdf.

Anonymous:1903:REG

[5908] Anonymous. A remarkable early Græco–Roman chariot. Scientific American, 89(22):385–386, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
385.pdf.

Anonymous:1903:HTN

[5909] Anonymous. Hundred-ton Naval floating crane. Scientific American, 89(22):386–387, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
386.pdf.

Anonymous:1903:GAB

[5910] Anonymous. A graphic account of the Berlin–Zossen record run. Scientific American, 89(22):388–389, November 28, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n22/pdf/scientificamerican11281903-388.pdf.

Fyfe:1903:FLC

[5911] Herbert C. Fyfe. The Finsen light-cure in England. Scientific American, 89(22):389–390, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
389.pdf.

Anonymous:1903:SMc

[5912] Anonymous. A Siberian mammoth. Scientific American, 89(22): 390, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
390.pdf.

Holder:1903:TBG

[5913] Charles F. Holder. Trapping big game of the sea. Scientific American, 89(22):391, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
391.pdf.

Anonymous:1903:BPWbo

[5914] Anonymous. Business and personal wants. Scientific American, 89(22):395, November 28, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n22/pdf/scientificamerican11281903-
395.pdf.

Anonymous:1903:RPIbo

[5915] Anonymous. Recently patented inventions. Scientific American, 89(23):414, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
414.pdf.

Anonymous:1903:NQbj

[5916] Anonymous. Notes and queries. Scientific American, 89(23):415, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n23/pdf/scientificamerican12051903-415.pdf.

Anonymous:1903:BNBbc

[5917] Anonymous. Books: New books, etc. Scientific American, 89 (23):415, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
415a.pdf.

Anonymous:1903:IIbp

[5918] Anonymous. Index of inventions. Scientific American, 89(23): 415–419, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
415b.pdf.

Anonymous:1903:PNPbi

[5919] Anonymous. Patent notes: Patents. Scientific American, 89 (23):419, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
419.pdf.

Anonymous:1903:SAR

[5920] Anonymous. Still another radium puzzle. Scientific American, 89(23):402, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
402.pdf.

Anonymous:1903:WBD

[5921] Anonymous. Wind bracing during erection. Scientific American, 89(23):402, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
402a.pdf.

Anonymous:1903:SKS

[5922] Anonymous. The sixteen-knot sailing yacht. Scientific American, 89(23):402, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
402b.pdf.

Anonymous:1903:ETA

[5923] Anonymous. An electromagnetic theory of the Aurora Borealis. Scientific American, 89(23):402, December 5, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n23/pdf/scientificamerican12051903-402c.pdf.

Anonymous:1903:AHE

[5924] Anonymous. Arsenic in hens’ eggs. Scientific American, 89(23): 402–403, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
402d.pdf.

Russell:1903:HD

[5925] Henry Norris Russell. The heavens in December. Scientific American, 89(23):403, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
403.pdf.

Anonymous:1903:RE

[5926] Anonymous. Reporting by electrophone. Scientific American, 89 (23):403, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
403a.pdf.

Anonymous:1903:SNbd

[5927] Anonymous. Science notes. Scientific American, 89(23):403, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n23/pdf/scientificamerican12051903-403b.pdf.

Anonymous:1903:MPL

[5928] Anonymous. A motor “Pullman” — Lord Anglesey’s new Mors. Scientific American, 89(23):404, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
404.pdf.

Anonymous:1903:GGO

[5929] Anonymous. The “Grangesberg” — a giant ore-carrying steamer. Scientific American, 89(23):404–405, December 5, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n23/pdf/scientificamerican12051903-404a.pdf.

Anonymous:1903:PED

[5930] Anonymous. Photography by electric discharges. Scientific American, 89(23):405, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
405.pdf.

Anonymous:1903:EBP

[5931] Anonymous. Electric blue-print machinery. Scientific American, 89 (23):405–407, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
405a.pdf.

Anonymous:1903:OMP

[5932] Anonymous. The Obelisk of Mont Pelé. Scientific American, 89 (23):407, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
407.pdf.

Beasley:1903:PCK

[5933] Walter L. Beasley. Picturesque Chinese kites. Scientific American, 89(23):408, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
408.pdf.

Crawford:1903:SUT

[5934] J. S. Crawford. Staff — its use and treatment. Scientific American, 89(23):408–410, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
408a.pdf.

Anonymous:1903:LA

[5935] Anonymous. The Lebaudy airship. Scientific American, 89(23):411, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n23/pdf/scientificamerican12051903-411.pdf.

Anonymous:1903:PDi

[5936] Anonymous. Patent department. Scientific American, 89(23):412, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n23/pdf/scientificamerican12051903-412.pdf.

Anonymous:1903:DVC

[5937] Anonymous. Display and vending case. Scientific American, 89 (23):412, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
412a.pdf.

Anonymous:1903:CMa

[5938] Anonymous. Centrifugal machine. Scientific American, 89(23):412, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n23/pdf/scientificamerican12051903-412b.pdf.

Anonymous:1903:OIg

[5939] Anonymous. Oddities in inventions. Scientific American, 89 (23):413, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
413.pdf.

Anonymous:1903:BPWbp

[5940] Anonymous. Business and personal wants. Scientific American, 89(23):414, December 5, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n23/pdf/scientificamerican12051903-
414a.pdf.

Anonymous:1903:NQbk

[5941] Anonymous. Notes and queries. Scientific American, 89(24): 449, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
449.pdf.

Anonymous:1903:BNBbd

[5942] Anonymous. Books: New books, etc. Scientific American, 89 (24):450, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
450.pdf.

Anonymous:1903:IIbq

[5943] Anonymous. Index of inventions. Scientific American, 89(24): 450–454, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
450a.pdf.

Anonymous:1903:PNPbj

[5944] Anonymous. Patent notes: Patents. Scientific American, 89 (24):453, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
453.pdf.

Anonymous:1903:Hc

[5945] Anonymous. Historical. Scientific American, 89(24):422, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n24/pdf/scientificamerican12121903-422.pdf.

Anonymous:1903:SOS

[5946] Anonymous. Secrets of our supremacy in iron and steel. Scientific American, 89(24):422–423, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
422a.pdf.

Meigs:1903:RDG

[5947] John F. Meigs. Recent developments in guns and armor. Scientific American, 89(24):423, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
423.pdf.

Anonymous:1903:FOS

[5948] Anonymous. The future of our steel industry. Scientific American, 89(24):423, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
423a.pdf.

Anonymous:1903:MSR

[5949] Anonymous. The manufacture of steel rails. Scientific American, 89 (24):424–427, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
424.pdf.

Anonymous:1903:MGSa

[5950] Anonymous. Manufacture of gun steel and armor plate. Scientific American, 89(24):427–430, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
427.pdf.

Anonymous:1903:SOI

[5951] Anonymous. Statistics of our iron and steel production. Scientific American, 89(24):430–431, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
430.pdf.

Anonymous:1903:MBB

[5952] Anonymous. Manufacture of bridge and building structural shapes. Scientific American, 89(24):432–433, December 12, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n24/pdf/scientificamerican12121903-432.pdf.

Anonymous:1903:MSP

[5953] Anonymous. Manufacture of steel pipe. Scientific American, 89 (24):434–435, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
434.pdf.

Anonymous:1903:SWN

[5954] Anonymous. Steel wire and nail making. Scientific American, 89 (24):436–438, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
436.pdf.

Anonymous:1903:CMb

[5955] Anonymous. Chain making. Scientific American, 89(24):439– 440, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
439.pdf.

Anonymous:1903:YAN

[5956] Anonymous. “Young America” — a nautical preparatory school. Scientific American, 89(24):441, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
441.pdf.

Anonymous:1903:PGA

[5957] Anonymous. Pompeiian gladiatorial arms. Scientific American, 89 (24):442–443, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
442.pdf.

Fisher:1903:LAD

[5958] Walter K. Fisher. The Laysan albatross and its dance. Scientific American, 89(24):444, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
444.pdf.

Anonymous:1903:MB

[5959] Anonymous. Marceau’s “Butterfly”. Scientific American, 89(24): 446, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
446.pdf.

Anonymous:1903:VHM

[5960] Anonymous. A vertical house moving. Scientific American, 89 (24):446–447, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
446a.pdf.

Anonymous:1903:BPWbq

[5961] Anonymous. Business and personal wants. Scientific American, 89(24):448, December 12, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n24/pdf/scientificamerican12121903-
448.pdf.

Anonymous:1903:Cbd

[5962] Anonymous. Correspondence. Scientific American, 89(25):463, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n25/pdf/scientificamerican12191903-463.pdf.

Anonymous:1903:RPIbp

[5963] Anonymous. Recently patented inventions. Scientific American, 89 (25):470–471, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
470.pdf.

Anonymous:1903:NQbl

[5964] Anonymous. Notes and queries. Scientific American, 89(25): 472, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
472.pdf.

Anonymous:1903:BNBbe

[5965] Anonymous. Books: New books, etc. Scientific American, 89(25): 472–473, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
472a.pdf.

Anonymous:1903:IIbr

[5966] Anonymous. Index of inventions. Scientific American, 89(25): 473, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
473.pdf.

Anonymous:1903:PNPbk

[5967] Anonymous. Patent notes: Patents. Scientific American, 89 (25):475, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
475.pdf.

Anonymous:1903:MSd

[5968] Anonymous. A monumental structure. Scientific American, 89 (25):458, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
458.pdf.

Anonymous:1903:IDE

[5969] Anonymous. Important development in electric traction. Scientific American, 89(25):458, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
458a.pdf.

Anonymous:1903:CCT

[5970] Anonymous. The Cunard Commission turbine test. Scientific American, 89(25):458, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
458b.pdf.

Anonymous:1903:NSMb

[5971] Anonymous. A new steam motor car for railroads. Scientific American, 89(25):458, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
458c.pdf.

Anonymous:1903:NCAa

[5972] Anonymous. The next Congress of Americanists. Scientific American, 89(25):458, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
458d.pdf.

Anonymous:1903:DHS

[5973] Anonymous. The death of Herbert Spencer. Scientific American, 89(25):459, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
459.pdf.

Anonymous:1903:DRA

[5974] Anonymous. Death of Rear Admiral Bancroft Gherardi. Scientific American, 89(25):459, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
459a.pdf.

Anonymous:1903:SNbe

[5975] Anonymous. Science notes. Scientific American, 89(25):459, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n25/pdf/scientificamerican12191903-459b.pdf.

Anonymous:1903:CSj

[5976] Anonymous. The current supplement. Scientific American, 89 (25):459, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
459c.pdf.

Shepstone:1903:LNT

[5977] Harold J. Shepstone. The launch of the new transatlantic liner “Baltic”. Scientific American, 89(25):460, December 19, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n25/pdf/scientificamerican12191903-460.pdf.

Anonymous:1903:SSO

[5978] Anonymous. A sectional steamer for our colonial possessions. Scientific American, 89(25):460, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
460a.pdf.

Trowbridge:1903:NL

[5979] John Trowbridge. The noise of lightning. Scientific American, 89 (25):461, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
461.pdf.

Anonymous:1903:OP

[5980] Anonymous. An optical pyrometer. Scientific American, 89(25): 461–462, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
461a.pdf.

Anonymous:1903:OC

[5981] Anonymous. An odd casting. Scientific American, 89(25):462– 463, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
462.pdf.

Anonymous:1903:ONE

[5982] Anonymous. The opening of the new East River Bridge. Scientific American, 89(25):464, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
464.pdf.

Bolce:1903:ODL

[5983] Harold J. Bolce. Ornithological discoveries on Laysan. Scientific American, 89(25):465–466, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
465.pdf.

Anonymous:1903:CSBa

[5984] Anonymous. California string beans. Scientific American, 89 (25):466, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
466.pdf.

Anonymous:1903:PDj

[5985] Anonymous. Patent department. Scientific American, 89(25): 467, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
467.pdf.

Anonymous:1903:SAD

[5986] Anonymous. Safety alarm device for marine vessels. Scientific American, 89(25):467, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
467a.pdf.

Anonymous:1903:IFW

[5987] Anonymous. Improved feed-water filter. Scientific American, 89 (25):467–468, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
467b.pdf.

Anonymous:1903:OIh

[5988] Anonymous. Oddities in inventions. Scientific American, 89(25): 468–469, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
468.pdf.

Anonymous:1903:BPWbr

[5989] Anonymous. Business and personal wants. Scientific American, 89(25):471, December 19, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n25/pdf/scientificamerican12191903-
471.pdf.

Anonymous:1903:Cbe

[5990] Anonymous. Correspondence. Scientific American, 89(26):483– 484, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
483.pdf.

Anonymous:1903:RPIbq

[5991] Anonymous. Recently patented inventions. Scientific American, 89(26):489, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
489.pdf.

Anonymous:1903:NQbm

[5992] Anonymous. Notes and queries. Scientific American, 89(26): 489, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
489b.pdf.

Anonymous:1903:BNBbf

[5993] Anonymous. Books: New books, etc. Scientific American, 89 (26):489, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
489c.pdf.

Anonymous:1903:IIbs

[5994] Anonymous. Index of inventions. Scientific American, 89(26): 489–492, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
489d.pdf.

Anonymous:1903:PNPbl

[5995] Anonymous. Patent notes: Patents. Scientific American, 89 (26):491, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
491.pdf.

Anonymous:1903:PNPbm

[5996] Anonymous. Proposed new Patent Office Building at Washington. Scientific American, 89(26):478, December 26, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n26/pdf/scientificamerican12261903-478.pdf.

Anonymous:1903:BGE

[5997] Anonymous. Brunel and the “Great Eastern”. Scientific American, 89(26):478, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
478a.pdf.

Anonymous:1903:SES

[5998] Anonymous. Some early subscribers heard from. Scientific American, 89(26):478, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
478b.pdf.

Anonymous:1903:CME

[5999] Anonymous. A convention on mosquito extermination. Scientific American, 89(26):478–479, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
478c.pdf.

Anonymous:1903:ETE

[6000] Anonymous. An English test of the Edison battery. Scientific American, 89(26):479, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
479.pdf.

Anonymous:1903:FWF

[6001] Anonymous. Final World’s Fair dividend. Scientific American, 89 (26):479, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
479a.pdf.

Anonymous:1903:OS

[6002] Anonymous. To our subscribers. Scientific American, 89(26): 479, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
479b.pdf.

Anonymous:1903:CSk

[6003] Anonymous. The current supplement. Scientific American, 89 (26):479, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
479c.pdf.

Anonymous:1903:ETD

[6004] Anonymous. Eiffel Tower to be dismantled. Scientific American, 89(26):479, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
479d.pdf.

Anonymous:1903:DNP

[6005] Anonymous. Distribution of the Nobel Prizes. Scientific American, 89(26):479, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
479e.pdf.

Gradenwitz:1903:BPT

[6006] Alfred Gradenwitz. The Berlin Printing-Telegraph Central Station. Scientific American, 89(26):480, December 26, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n26/pdf/scientificamerican12261903-480.pdf.

Anonymous:1903:PRU

[6007] Anonymous. A pneumatic rail unloader. Scientific American, 89 (26):481, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
481.pdf.

Anonymous:1903:FBS

[6008] Anonymous. The first British steam railway to adopt electrical working. Scientific American, 89(26):481–482, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n26/pdf/scientificamerican12261903-481a.pdf.

Anonymous:1903:ICSa

[6009] Anonymous. Improved cream separator. Scientific American, 89 (26):482, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
482.pdf.

Anonymous:1903:SOJ

[6010] Anonymous. Sir Oliver Joseph Lodge, F.R.S. Scientific American, 89 (26):482–483, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
482a.pdf.

Anonymous:1903:LST

[6011] Anonymous. The lake submarine torpedo boat “Protector”. Scientific American, 89(26):485–486, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
485.pdf.

Anonymous:1903:LFLb

[6012] Anonymous. The latest form of “looping the loop”. Scientific American, 89(26):486, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
486.pdf.

Anonymous:1903:SEM

[6013] Anonymous. A successful experiment with a motor-drlven aeroplane. Scientific American, 89(26):486, December 26, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n26/pdf/scientificamerican12261903-486a.pdf.

Geare:1903:PW

[6014] Randolph I. Geare. Pewter ware. Scientific American, 89(26): 487–489, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
487.pdf.

Anonymous:1903:BPWbs

[6015] Anonymous. Business and personal wants. Scientific American, 89(26):489, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n26/pdf/scientificamerican12261903-
489a.pdf.

Anonymous:1903:RPIbr

[6016] Anonymous. Recently patented inventions. Scientific American, 89(3):52, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
52.pdf.

Anonymous:1903:NQbn

[6017] Anonymous. Notes and queries. Scientific American, 89(3):52, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n3/pdf/scientificamerican07181903-52c.pdf.

Anonymous:1903:IIbt

[6018] Anonymous. Index of inventions. Scientific American, 89(3):52, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n3/pdf/scientificamerican07181903-52d.pdf.

Anonymous:1903:PNPbn

[6019] Anonymous. Patent notes: Patents. Scientific American, 89(3):54, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n3/pdf/scientificamerican07181903-54.pdf.

Anonymous:1903:DFAb

[6020] Anonymous. Dam failures not the act of god. Scientific American, 89(3):42, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
42.pdf.

Anonymous:1903:SR

[6021] Anonymous. “Shamrock” and “Reliance”. Scientific American, 89(3):42, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
42a.pdf.

Anonymous:1903:BPF

[6022] Anonymous. The book paper of the future. Scientific American, 89(3):42–43, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
42b.pdf.

Thompson:1903:PCA

[6023] J. H. Thompson. Present conditions in Alaska. Scientific American, 89(3):43, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
43.pdf.

Anonymous:1903:SNbf

[6024] Anonymous. Science notes. Scientific American, 89(3):43, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n3/pdf/scientificamerican07181903-43a.pdf.

Anonymous:1903:CDD

[6025] Anonymous. Core drilling without diamonds. Scientific American, 89(3):44, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
44.pdf.

Anonymous:1903:EHb

[6026] Anonymous. An electric hammer. Scientific American, 89(3):44–45, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n3/pdf/scientificamerican07181903-44a.pdf.

Jones:1903:EBP

[6027] George J. Jones. Electric blue-print making. Scientific American, 89(3):45–46, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
45.pdf.

Anonymous:1903:FAC

[6028] Anonymous. The first Atlantic cable station in America. Scientific American, 89(3):46, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
46.pdf.

Anonymous:1903:ANE

[6029] Anonymous. Approach to the new East River Bridge. Scientific American, 89(3):46–47, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
46a.pdf.

Anonymous:1903:NLA

[6030] Anonymous. The new liner “Arabic”. Scientific American, 89(3):48, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n3/pdf/scientificamerican07181903-48.pdf.

Anonymous:1903:Gb

[6031] Anonymous. The “Gleniffer”. Scientific American, 89(3):48–50, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n3/pdf/scientificamerican07181903-48a.pdf.

Anonymous:1903:AR

[6032] Anonymous. The automatic restaurant. Scientific American, 89 (3):50–51, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
50.pdf.

Anonymous:1903:SFC

[6033] Anonymous. Some freak cycle-whirls and loop-the-loops. Scientific American, 89(3):51, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
51.pdf.

Anonymous:1903:ILM

[6034] Anonymous. The ideal lawn mower grinder. Scientific American, 89(3):52, July 18, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
52a.pdf.

Anonymous:1903:BPWbt 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n3/pdf/scientificamerican07181903-
52b.pdf.

Anonymous:1903:LNl

[6036] Anonymous. Legal notes. Scientific American, 89(4):68, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-68.pdf.

Anonymous:1903:RPIbs

[6037] Anonymous. Recently patented inventions. Scientific American, 89(4):69–70, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
69.pdf.

Anonymous:1903:NQbo

[6038] Anonymous. Notes and queries. Scientific American, 89(4):70–71, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-70a.pdf.

Anonymous:1903:BNBbg

[6039] Anonymous. Books: New books, etc. Scientific American, 89(4):71–72, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-71.pdf.

Anonymous:1903:IIbu

[6040] Anonymous. Index of inventions. Scientific American, 89(4):72–73, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-72.pdf.

Anonymous:1903:PNPbo

[6041] Anonymous. Patent notes: Patents. Scientific American, 89(4):73, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-73.pdf.

Anonymous:1903:PAB

[6042] Anonymous. Public appropriations and the board of aldermen. Scientific American, 89(4):58, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
58.pdf.

Anonymous:1903:MAN

[6043] Anonymous. The Merchants’ Association and the new reservoir. Scientific American, 89(4):58, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
58a.pdf.

Anonymous:1903:ICSb

[6044] Anonymous. Industrial chemistry in the South during the Civil War. Scientific American, 89(4):58, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
58b.pdf.

Anonymous:1903:CVW

[6045] Anonymous. Chain vs. wire cables in the new East River Bridge. Scientific American, 89(4):59, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
59.pdf.

Anonymous:1903:CCHb

[6046] Anonymous. Combs and cigar holders made of milk. Scientific American, 89(4):59, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
59a.pdf.

Anonymous:1903:AC

[6047] Anonymous. The “Akouphone” — a correction. Scientific American, 89(4):59, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
59b.pdf.

Anonymous:1903:DCCb

[6048] Anonymous. Death of Charles C. Martin. Scientific American, 89(4):59, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-59c.pdf.

Anonymous:1903:NTDb

[6049] Anonymous. New tests by Dr. Wiley. Scientific American, 89(4):59, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-59d.pdf.

Anonymous:1903:DEV

[6050] Anonymous. The death of Eugene Vanderpool. Scientific American, 89(4):59, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
59e.pdf.

Anonymous:1903:ISE

[6051] Anonymous. An improved static electric machine. Scientific American, 89(4):60, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
60.pdf.

Anonymous:1903:FCD

[6052] Anonymous. French commerce-destroyer “Guichen”. Scientific American, 89(4):60, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
60a.pdf.

Anonymous:1903:GBA

[6053] Anonymous. The Gordon Bennett Automobile Cup Race. Scientific American, 89(4):61–63, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
61.pdf.

Trowbridge:1903:SGM

[6054] John Trowbridge. The study of gases and metals at very high temperatures. Scientific American, 89(4):63–64, July 25, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n4/pdf/scientificamerican07251903-63.pdf.

Anonymous:1903:PLP

[6055] Anonymous. Progress of the Louisiana Purchase Exposition. Scientific American, 89(4):65, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
65.pdf.

Anonymous:1903:CRM

[6056] Anonymous. From cruiser to racing machine. Scientific American, 89(4):65–66, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
65a.pdf.

Anonymous:1903:PDk

[6057] Anonymous. Patent department. Scientific American, 89(4):66, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-66.pdf.

Anonymous:1903:SGL

[6058] Anonymous. Snow guard for lumbermen’s overshoes. Scientific American, 89(4):66–67, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
66a.pdf.

Anonymous:1903:MTH

[6059] Anonymous. Magnetic tool holder. Scientific American, 89(4):67, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-67.pdf.

Anonymous:1903:AFD

[6060] Anonymous. Automobile frame and driving axle gear. Scientific American, 89(4):67, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
67a.pdf.

Anonymous:1903:OIi

[6061] Anonymous. Oddities in inventions. Scientific American, 89(4):67, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n4/pdf/scientificamerican07251903-67b.pdf.

Anonymous:1903:BPWbu

[6062] Anonymous. Business and personal wants. Scientific American, 89(4):70, July 25, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n4/pdf/scientificamerican07251903-
70.pdf.

Russell:1903:HAb

[6063] Henry Norris Russell. The heavens in August, 1903. Scientific American, 89(5):79, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
79.pdf.

Anonymous:1903:RPIbt

[6064] Anonymous. Recently patented inventions. Scientific American, 89(5):88, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
88.pdf.

Anonymous:1903:BNBbh

[6065] Anonymous. Books: New books, etc. Scientific American, 89(5):88, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n5/pdf/scientificamerican08011903-88b.pdf.

Anonymous:1903:IIbv

[6066] Anonymous. Index of inventions. Scientific American, 89(5):88–89, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n5/pdf/scientificamerican08011903-88c.pdf.

Anonymous:1903:PNPbp

[6067] Anonymous. Patent notes: Patents. Scientific American, 89(5):90, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n5/pdf/scientificamerican08011903-90.pdf.

Anonymous:1903:TTO

[6068] Anonymous. The torpedo tube and our new battleships. Scientific American, 89(5):78, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
78.pdf.

Anonymous:1903:ED

[6069] Anonymous. Electrochemical discoveries. Scientific American, 89(5):78, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n5/pdf/scientificamerican08011903-78a.pdf.

Anonymous:1903:NSC

[6070] Anonymous. Nordenskiold’s ship crushed in the ice pack. Scientific American, 89(5):78–79, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
78b.pdf.

Anonymous:1903:BCb

[6071] Anonymous. Borelly’s Comet. Scientific American, 89(5):79, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n5/pdf/scientificamerican08011903-79a.pdf.

Anonymous:1903:WSS

[6072] Anonymous. White spot on Saturn. Scientific American, 89(5):79, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n5/pdf/scientificamerican08011903-79b.pdf.

Anonymous:1903:ITSb

[6073] Anonymous. An improved type of steam nozzle. Scientific American, 89(5):80, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
80.pdf.

Anonymous:1903:EFS

[6074] Anonymous. The Edem fog signal. Scientific American, 89(5):80, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n5/pdf/scientificamerican08011903-80a.pdf.

Anonymous:1903:ACS

[6075] Anonymous. Auto chairs for the St. Louis Fair. Scientific American, 89(5):81, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
81.pdf.

Fawcett:1903:ISS

[6076] Waldon Fawcett. Improvements in steam shovels. Scientific American, 89(5):81–82, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
81a.pdf.

Anonymous:1903:PPM

[6077] Anonymous. The projections of the Planet Mars. Scientific American, 89(5):82–83, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
82.pdf.

Winterton:1903:VEG

[6078] J. Winterton. The volcanic eruptions in Guatemala. Scientific American, 89(5):84, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
84.pdf.

Anonymous:1903:OCW

[6079] Anonymous. The ore-carrying wire ropeway at Vivero, Spain. Scientific American, 89(5):84–86, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
84a.pdf.

Anonymous:1903:PMAb

[6080] Anonymous. A prism for measuring the actual dip of the horizon. Scientific American, 89(5):86, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
86.pdf.

Beasley:1903:RFD

[6081] Walter L. Beasley. A remarkable fossil discovery. Scientific American, 89(5):87, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
87.pdf.

Anonymous:1903:BPWbv

[6082] Anonymous. Business and personal wants. Scientific American, 89(5):88, August 1, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n5/pdf/scientificamerican08011903-
88a.pdf.

Anonymous:1903:LNm

[6083] Anonymous. Legal notes. Scientific American, 89(6):104, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n6/pdf/scientificamerican08081903-104.pdf.

Anonymous:1903:RPIbu

[6084] Anonymous. Recently patented inventions. Scientific American, 89(6):105–106, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
105.pdf.

Anonymous:1903:NQbp

[6085] Anonymous. Notes and queries. Scientific American, 89(6):106–108, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n6/pdf/scientificamerican08081903-106a.pdf.

Anonymous:1903:BNBbi

[6086] Anonymous. Books: New books, etc. Scientific American, 89 (6):108–109, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
108.pdf.

Anonymous:1903:IIbw

[6087] Anonymous. Index of inventions. Scientific American, 89(6):109, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n6/pdf/scientificamerican08081903-109.pdf.

Anonymous:1903:TRK

[6088] Anonymous. The transatlantic run of the “Kearsarge”. Scientific American, 89(6):94, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
94.pdf.

Anonymous:1903:BST

[6089] Anonymous. The Behring Strait Tunnel again. Scientific American, 89(6):94, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
94a.pdf.

Anonymous:1903:SAc

[6090] Anonymous. The speed of the automobile. Scientific American, 89(6):94, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
94b.pdf.

Anonymous:1903:MPI

[6091] Anonymous. Marvelous production of iron and steel. Scientific American, 89(6):94, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
94c.pdf.

Anonymous:1903:SG

[6092] Anonymous. The slaughter goes on. Scientific American, 89(6):94, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n6/pdf/scientificamerican08081903-94d.pdf.

Anonymous:1903:MNA

[6093] Anonymous. More news about Edison’s storage battery. Scientific American, 89(6):94–95, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
94e.pdf.

Anonymous:1903:PFC

[6094] Anonymous. Proposed Forth–Clyde Ship Canal. Scientific American, 89(6):95, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
95.pdf.

Anonymous:1903:RAT

[6095] Anonymous. Recent automobile track records. Scientific American, 89(6):96–97, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
96.pdf.

Anonymous:1903:MLM

[6096] Anonymous. A multiface Linotype machine. Scientific American, 89(6):97, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
97.pdf.

Harrison:1903:ESD

[6097] Walton Harrison. Eccentricities of the static discharge. Scientific American, 89(6):98, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
98.pdf.

Anonymous:1903:PSW

[6098] Anonymous. Prof. Sir William Crookes, F.R.S. Scientific American, 89(6):99, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
99.pdf.

Anonymous:1903:MWT

[6099] Anonymous. Marconi wireless telegraph relics. Scientific American, 89(6):100, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
100.pdf.

Anonymous:1903:MBR

[6100] Anonymous. Motor boat races in Ireland for the Harmsworth Cup. Scientific American, 89(6):101, August 8, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n6/pdf/scientificamerican08081903-101.pdf.

Holder:1903:RFS

[6101] Charles F. Holder. The ribbon fish and the sea serpent. Scientific American, 89(6):101–102, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
101a.pdf.

Anonymous:1903:PDl

[6102] Anonymous. Patent department. Scientific American, 89(6):102, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n6/pdf/scientificamerican08081903-102.pdf.

Anonymous:1903:FPAa

[6103] Anonymous. A fountain painting apparatus. Scientific American, 89(6):102–103, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
102a.pdf.

Anonymous:1903:NPE

[6104] Anonymous. A novel picture exhibitor. Scientific American, 89(6):103, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n6/pdf/scientificamerican08081903-103.pdf.

Anonymous:1903:VPB

[6105] Anonymous. Vacuum pump for bulbs of electric lamps. Scientific American, 89(6):103, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
103a.pdf.

Anonymous:1903:FCR

[6106] Anonymous. Freight car with removable roof. Scientific American, 89(6):103, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
103b.pdf.

Anonymous:1903:BPWbw

[6107] Anonymous. Business and personal wants. Scientific American, 89(6):106, August 8, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n6/pdf/scientificamerican08081903-
106.pdf.

Anonymous:1903:Cbf

[6108] Anonymous. Correspondence. Scientific American, 89(7):119, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n7/pdf/scientificamerican08151903-119.pdf.

Anonymous:1903:RPIbv

[6109] Anonymous. Recently patented inventions. Scientific American, 89(7):124, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
124.pdf.

Anonymous:1903:NQbq

[6110] Anonymous. Notes and queries. Scientific American, 89(7):124– 125, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
124c.pdf.

Anonymous:1903:BNBbj (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
125.pdf.

Anonymous:1903:IIbx

[6112] Anonymous. Index of inventions. Scientific American, 89(7): 125–126, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
125a.pdf.

Anonymous:1903:PNPbq

[6113] Anonymous. Patent notes: Patents. Scientific American, 89(7):127, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n7/pdf/scientificamerican08151903-127.pdf.

Anonymous:1903:BTC

[6114] Anonymous. Broadway Tunnel controversy. Scientific American, 89(7):114, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
114.pdf.

Anonymous:1903:DS

[6115] Anonymous. A day with the “Shamrocks”. Scientific American, 89(7):114–115, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
114a.pdf.

Anonymous:1903:PSA

[6116] Anonymous. Present status of American shipbuilding. Scientific American, 89(7):115, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
115.pdf.

Anonymous:1903:NIG

[6117] Anonymous. The Ninth International Geological Congress. Scientific American, 89(7):115, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
115a.pdf.

Proctor:1903:PBO

[6118] Mary Proctor. Prof. Barnard’s observations of the White Spot on Saturn. Scientific American, 89(7):115, August 15, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n7/pdf/scientificamerican08151903-115b.pdf.

Anonymous:1903:NGD

[6119] Anonymous. A new gold district. Scientific American, 89(7):115, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n7/pdf/scientificamerican08151903-115c.pdf.

Anonymous:1903:CSl

[6120] Anonymous. The current supplement. Scientific American, 89 (7):115, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
115d.pdf.

Massa:1903:IDM

[6121] F. N. Massa. An instrument for describing mathematical curves. Scientific American, 89(7):116, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
116.pdf.

Guarini:1903:AVR

[6122] Emile Guarini. An audible and visible railway signal for grade crossings, operated by a high tension current. Scientific American, 89(7):117, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n7/pdf/scientificamerican08151903-117.pdf.

Irwin:1903:WIO

[6123] William Gilbert Irwin. Waterway improvement on the Ohio. Scientific American, 89(7):117–118, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
117a.pdf.

Beasley:1903:SCS

[6124] Walter L. Beasley. The secret cannibal society of the Kwakiutl. Scientific American, 89(7):120–122, August 15, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n7/pdf/scientificamerican08151903-120.pdf.

Anonymous:1903:ADR

[6125] Anonymous. An automatic device for replacing bowling alley pins. Scientific American, 89(7):122, August 15, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n7/pdf/scientificamerican08151903-122.pdf.

Anonymous:1903:SPMb

[6126] Anonymous. The sending of the President’s message around the world. Scientific American, 89(7):122–123, August 15, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n7/pdf/scientificamerican08151903-122a.pdf.

Anonymous:1903:NAAb

[6127] Anonymous. A novel amusement apparatus. Scientific American, 89(7):123, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
123.pdf.

Anonymous:1903:ELP

[6128] Anonymous. An efficient low-priced typewriter. Scientific American, 89(7):124, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
124a.pdf.

Anonymous:1903:BPWbx

[6129] Anonymous. Business and personal wants. Scientific American, 89(7):124, August 15, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n7/pdf/scientificamerican08151903-
124b.pdf.

Anonymous:1903:RPIbw

[6130] Anonymous. Recently patented inventions. Scientific American, 89(8):140, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
140.pdf.

Anonymous:1903:NQbr

[6131] Anonymous. Notes and queries. Scientific American, 89(8):141, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n8/pdf/scientificamerican08221903-141.pdf.

Anonymous:1903:BNBbk

[6132] Anonymous. Books: New books, etc. Scientific American, 89 (8):141, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
141a.pdf.

Anonymous:1903:IIby

[6133] Anonymous. Index of inventions. Scientific American, 89(8): 141–142, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
141b.pdf.

Anonymous:1903:PNPbr

[6134] Anonymous. Patent notes: Patents. Scientific American, 89(8):143, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n8/pdf/scientificamerican08221903-143.pdf.

Perkins:1903:MS

[6135] Frank C. Perkins. Modern searchlights. Scientific American, 89 (8):129, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
129.pdf.

Anonymous:1903:NCe

[6136] Anonymous. The new chemistry. Scientific American, 89(8):130, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n8/pdf/scientificamerican08221903-130.pdf.

Anonymous:1903:FPU

[6137] Anonymous. Fire peril on underground railroads. Scientific American, 89(8):130, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
130a.pdf.

Anonymous:1903:FGE

[6138] Anonymous. Foreign gas engines at the World’s Fair. Scientific American, 89(8):130–131, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
130b.pdf.

Watson:1903:MBP

[6139] Egbert P. Watson. The mandatory blue print. Scientific American, 89(8):131, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
131.pdf.

Anonymous:1903:RDM

[6140] Anonymous. N-rays discovered by M. Blondlot. Scientific American, 89(8):131, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
131a.pdf.

Anonymous:1903:LAE

[6141] Anonymous. Langley’s aerodrome experiments. Scientific American, 89(8):131, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
131b.pdf.

Anonymous:1903:HRTb

[6142] Anonymous. The Hudson River Tunnel. Scientific American, 89 (8):132–133, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
132.pdf.

Anonymous:1903:BDB

[6143] Anonymous. The beaver as a dam-builder. Scientific American, 89(8):133–134, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
133.pdf.

Proctor:1903:BC

[6144] Mary Proctor. Borelly’s Comet. Scientific American, 89(8):135, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n8/pdf/scientificamerican08221903-135.pdf.

Campbell:1903:OBC

[6145] W. W. Campbell. Observations of Borelly’s Comet at Lick Observatory. Scientific American, 89(8):135, August 22, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n8/pdf/scientificamerican08221903-135a.pdf.

Anonymous:1903:MLP

[6146] Anonymous. The making of a lead pencil. Scientific American, 89(8):137, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
137.pdf.

Willey:1903:AWC

[6147] Day Allen Willey. Alfalfa on western cattle ranges. Scientific American, 89(8):138, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
138.pdf.

Beard:1903:FP

[6148] J. Carter Beard. Flowers of prey. Scientific American, 89(8):139, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n8/pdf/scientificamerican08221903-139.pdf.

Anonymous:1903:BPWby

[6149] Anonymous. Business and personal wants. Scientific American, 89(8):140, August 22, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n8/pdf/scientificamerican08221903-
140a.pdf.

Anonymous:1903:Cbg

[6150] Anonymous. Correspondence. Scientific American, 89(9):151, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n9/pdf/scientificamerican08291903-151.pdf.

Anonymous:1903:RPIbx

[6151] Anonymous. Recently patented inventions. Scientific American, 89(9):156, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
156.pdf.

Anonymous:1903:NQbs

[6152] Anonymous. Notes and queries. Scientific American, 89(9):156, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n9/pdf/scientificamerican08291903-156b.pdf.

Anonymous:1903:IIbz

[6153] Anonymous. Index of inventions. Scientific American, 89(9): 156–157, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
156c.pdf.

Anonymous:1903:PNPbs

[6154] Anonymous. Patent notes: Patents. Scientific American, 89(9):159, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v89/n9/pdf/scientificamerican08291903-159.pdf.

Anonymous:1903:CPC

[6155] Anonymous. Colombia and the Panama Canal. Scientific American, 89(9):146, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
146.pdf.

Anonymous:1903:PASb

[6156] Anonymous. A problem for the air-ship builder. Scientific American, 89(9):146, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
146a.pdf.

Anonymous:1903:DCV

[6157] Anonymous. Direct current vs. induction motor for electric traction. Scientific American, 89(9):146, August 29, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n9/pdf/scientificamerican08291903-146b.pdf.

Bolce:1903:TLI

[6158] Harold Bolce. The trap-lantern as an insect exterminator. Scientific American, 89(9):147, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
147.pdf.

Hovey:1903:BMM

[6159] Horace C. Hovey. Balloon measurements of Mammoth Cave’s height. Scientific American, 89(9):147, August 29, 1903. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v89/
n9/pdf/scientificamerican08291903-147a.pdf.

Anonymous:1903:CHB

[6160] Anonymous. Concrete house building. Scientific American, 89 (9):148–149, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
148.pdf.

Anonymous:1903:APF

[6161] Anonymous. Automatic paper-feeders for printing machinery. Scientific American, 89(9):149–150, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
149.pdf.

Anonymous:1903:DFR

[6162] Anonymous. Development of the 90-foot racing yacht. Scientific American, 89(9):152–154, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
152.pdf.

Anonymous:1903:IBM

[6163] Anonymous. Improved breech mechanism for heavy guns. Scientific American, 89(9):154, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
154.pdf.

Anonymous:1903:LMS

[6164] Anonymous. Light motor for a St. Louis Contest airship. Scientific American, 89(9):154–155, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
154a.pdf.

Anonymous:1903:MGSb

[6165] Anonymous. Machine for grooming and shampooing horses. Scientific American, 89(9):155, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
155.pdf.

Anonymous:1903:IVM

[6166] Anonymous. Improved valve mechanism. Scientific American, 89(9):155, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
155a.pdf.

Anonymous:1903:CGC

[6167] Anonymous. Cable grip or clutch device. Scientific American, 89(9):155, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
155b.pdf.

Anonymous:1903:BPWbz

[6168] Anonymous. Business and personal wants. Scientific American, 89(9):156, August 29, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v89/n9/pdf/scientificamerican08291903-
156a.pdf.

Anonymous:1904:Ca

[6169] Anonymous. Correspondence. Scientific American, 90(1):8, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-8.pdf.

Russell:1904:HJa

[6170] Henry Norris Russell. The heavens in January. Scientific American, 90(1):9, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n1/pdf/scientificamerican01021904-
9.pdf.

Anonymous:1904:RPIa

[6171] Anonymous. Recently patented inventions. Scientific American, 90(1):14–15, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n1/pdf/scientificamerican01021904-
14.pdf.

Anonymous:1904:NQa

[6172] Anonymous. Notes and queries. Scientific American, 90(1):15–16, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-15a.pdf.

Anonymous:1904:BNBa

[6173] Anonymous. Books: New books, etc. Scientific American, 90(1):16, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-16.pdf.

Anonymous:1904:IIa

[6174] Anonymous. Index of inventions. Scientific American, 90(1):16–19, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-16a.pdf.

Anonymous:1904:PNPa

[6175] Anonymous. Patent notes: Patents. Scientific American, 90(1):18, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-18.pdf.

Anonymous:1904:RY

[6176] Anonymous. Retrospect of the year 1903. Scientific American, 90(1):4–5, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n1/pdf/scientificamerican01021904-
4.pdf.

Anonymous:1904:H

[6177] Anonymous. The Hydrobion. Scientific American, 90(1):6, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-6.pdf.

Anonymous:1904:IER

[6178] Anonymous. The illumination of the East River Bridge. Scientific American, 90(1):6, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n1/pdf/scientificamerican01021904-
6a.pdf.

Anonymous:1904:DFD

[6179] Anonymous. The Durban Floating Dock and Workshop. Scientific American, 90(1):7–8, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n1/pdf/scientificamerican01021904-
7.pdf.

Anonymous:1904:AD

[6180] Anonymous. An albino deer. Scientific American, 90(1):10, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-10.pdf.

Holder:1904:MTC

[6181] Charles F. Holder. Meteorites and their collectors. Scientific American, 90(1):10, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n1/pdf/scientificamerican01021904-
10a.pdf.

Anonymous:1904:RLT

[6182] Anonymous. Reconstruction of the Lackawanna tracks through Newark. Scientific American, 90(1):10–13, January 2, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n1/pdf/scientificamerican01021904-10b.pdf.

Anonymous:1904:SC

[6183] Anonymous. Smoke consumer. Scientific American, 90(1):13, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-13.pdf.

Anonymous:1904:MPCa

[6184] Anonymous. Metallic pipe coupling. Scientific American, 90(1):13, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-13a.pdf.

Anonymous:1904:LMH

[6185] Anonymous. A long-maned horse. Scientific American, 90(1):13, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-13b.pdf.

Anonymous:1904:IDC

[6186] Anonymous. Improved dumping car. Scientific American, 90(1):13, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n1/pdf/scientificamerican01021904-13c.pdf.

Anonymous:1904:BPWa

[6187] Anonymous. Business and personal wants. Scientific American, 90(1):15, January 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n1/pdf/scientificamerican01021904-
15.pdf.

Anonymous:1904:Cb

[6188] Anonymous. Correspondence. Scientific American, 90(10):195, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-195.pdf.

Russell:1904:HMa

[6189] Henry Norris Russell. The heavens in March. Scientific American, 90(10):195, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
195a.pdf.

Anonymous:1904:RPIb

[6190] Anonymous. Recently patented inventions. Scientific American, 90(10):202, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
202.pdf.

Anonymous:1904:NQb

[6191] Anonymous. Notes and queries. Scientific American, 90(10):202, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-202b.pdf.

Anonymous:1904:BNBb

[6192] Anonymous. Books: New books, etc. Scientific American, 90(10):202, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-202c.pdf.

Anonymous:1904:IIb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-203.pdf.

Anonymous:1904:PNPb

[6194] Anonymous. Patent notes: Patents. Scientific American, 90(10):206, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-206.pdf.

Anonymous:1904:HSM

[6195] Anonymous. The high-speed motor-boat model. Scientific American, 90(10):190, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
190.pdf.

Anonymous:1904:RCP

[6196] Anonymous. Raising carrier pigeons. Scientific American, 90(10):190, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-190a.pdf.

Anonymous:1904:WCC

[6197] Anonymous. On the working of combined coherers. Scientific American, 90(10):190–191, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
190b.pdf.

Hurd:1904:FFEa

[6198] Archibald S. Hurd. Fleets in the Far East — an English review of the position of Russia and Japan. Scientific American, 90(10):191, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-191.pdf.

Anonymous:1904:SNa

[6199] Anonymous. Science notes. Scientific American, 90(10):191, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-191a.pdf.

Anonymous:1904:RVS

[6200] Anonymous. Rotary valve for sawmills. Scientific American, 90(10):192, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-192.pdf.

Anonymous:1904:ABG

[6201] Anonymous. The Adam–Boudin globular boat. Scientific American, 90(10):192, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
192a.pdf.

Anonymous:1904:VCH

[6202] Anonymous. Valve for control of hydraulic elevators. Scientific American, 90(10):192, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
192b.pdf.

Anonymous:1904:LPW

[6203] Anonymous. The largest photograph in the world. Scientific American, 90(10):193–194, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
193.pdf.

Anonymous:1904:SPL

[6204] Anonymous. Samuel Pierpont Langley. Scientific American, 90(10):194, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-194.pdf.

Anonymous:1904:MT

[6205] Anonymous. The modern torpedo. Scientific American, 90(10):196, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-196.pdf.

Phillips:1904:JHS

[6206] T. E. R. Phillips. Jupiter and his surface currents. Scientific American, 90(10):197, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
197.pdf.

Anonymous:1904:RLF

[6207] Anonymous. Russia’s losses in the first week of the War. Scientific American, 90(10):200, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
200.pdf.

Anonymous:1904:S

[6208] Anonymous. The “Scooter”. Scientific American, 90(10):201, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n10/pdf/scientificamerican03051904-201.pdf.

Anonymous:1904:BPWb

[6209] Anonymous. Business and personal wants. Scientific American, 90(10):202, March 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n10/pdf/scientificamerican03051904-
202a.pdf.

Anonymous:1904:Cc

[6210] Anonymous. Correspondence. Scientific American, 90(11):215, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n11/pdf/scientificamerican03121904-215.pdf.

Anonymous:1904:RPIc

[6211] Anonymous. Recently patented inventions. Scientific American, 90 (11):219–220, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
219a.pdf.

Anonymous:1904:BPWc

[6212] Anonymous. Business and personal wants. Scientific American, 90(11):220, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
220.pdf.

Anonymous:1904:IIc

[6213] Anonymous. Index of inventions. Scientific American, 90(11):220–223, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n11/pdf/scientificamerican03121904-220a.pdf.

Anonymous:1904:PNPc

[6214] Anonymous. Patent notes: Patents. Scientific American, 90(11):220, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n11/pdf/scientificamerican03121904-220b.pdf.

Anonymous:1904:DAH

[6215] Anonymous. The Darlington apartment house disaster. Scientific American, 90(11):210, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
210.pdf.

Anonymous:1904:RABa

[6216] Anonymous. Report of the Army Board as to the usefulness of the lake type of submarine boat for coast defense. Scientific American, 90(11):210–211, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
210a.pdf.

Anonymous:1904:RRE

[6217] Anonymous. Results of recent experiments with N-rays. Scientific American, 90(11):211, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
211.pdf.

Anonymous:1904:EEG

[6218] Anonymous. Examination of the emanations given off by radium. Scientific American, 90(11):211, March 12, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n11/pdf/scientificamerican03121904-211a.pdf.

Anonymous:1904:YBE

[6219] Anonymous. A year’s battleships in England. Scientific American, 90(11):211, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
211b.pdf.

Anonymous:1904:TRC

[6220] Anonymous. The Tygard reciprocating-cylinder double-action gasoline motor. Scientific American, 90(11):212, March 12, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n11/pdf/scientificamerican03121904-212.pdf.

Anonymous:1904:DHS

[6221] Anonymous. The development of the high-speed launch or automobile boat. Scientific American, 90(11):213–214, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n11/pdf/scientificamerican03121904-213.pdf.

Anonymous:1904:JD

[6222] Anonymous. The Japanese destroyers. Scientific American, 90 (11):214–215, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
214.pdf.

Anonymous:1904:TIA

[6223] Anonymous. Two interesting animals at the New York Zoological Park. Scientific American, 90(11):216, March 12, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n11/pdf/scientificamerican03121904-216.pdf.

Willey:1904:LBF

[6224] Day Allen Willey. The lesson of the Baltimore Fire. Scientific American, 90(11):216–218, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
216a.pdf.

Anonymous:1904:OTL

[6225] Anonymous. Official test of the lake submarine boat “Protector”. Scientific American, 90(11):218, March 12, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n11/pdf/scientificamerican03121904-218.pdf.

Jones:1904:QHP

[6226] Helen Lukens Jones. Queer hiding place for bees. Scientific American, 90(11):219, March 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n11/pdf/scientificamerican03121904-
219.pdf.

Anonymous:1904:Cd

[6227] Anonymous. Correspondence. Scientific American, 90(12):231, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-231.pdf.

Anonymous:1904:LNa

[6228] Anonymous. Legal notes. Scientific American, 90(12):237, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-237.pdf.

Anonymous:1904:RPId

[6229] Anonymous. Recently patented inventions. Scientific American, 90 (12):238–239, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
238.pdf.

Anonymous:1904:BPWd

[6230] Anonymous. Business and personal wants. Scientific American, 90(12):239, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
239.pdf.

Anonymous:1904:NQc

[6231] Anonymous. Notes and queries. Scientific American, 90(12):239, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-239a.pdf.

Anonymous:1904:BNBc

[6232] Anonymous. Books: New books, etc. Scientific American, 90 (12):239–240, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
239b.pdf.

Anonymous:1904:IId

[6233] Anonymous. Index of inventions. Scientific American, 90(12):240–243, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-240.pdf.

Anonymous:1904:PNPd

[6234] Anonymous. Patent notes: Patents. Scientific American, 90(12):240, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-240a.pdf.

Anonymous:1904:CNR

[6235] Anonymous. The completion of the North River Tunnel. Scientific American, 90(12):226, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
226.pdf.

Anonymous:1904:DCT

[6236] Anonymous. Direct-current traction on the New York Central. Scientific American, 90(12):226, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
226a.pdf.

Anonymous:1904:MPH

[6237] Anonymous. The motive power for High-Pressure Fire Service. Scientific American, 90(12):226–227, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
226b.pdf.

Anonymous:1904:BNP

[6238] Anonymous. The British Naval Programme for 1904. Scientific American, 90(12):227, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
227.pdf.

Anonymous:1904:SNb

[6239] Anonymous. Science notes. Scientific American, 90(12):227, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-227a.pdf.

Anonymous:1904:DDC

[6240] Anonymous. A double-deck car for rapid transit. Scientific American, 90(12):228, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
228.pdf.

Anonymous:1904:GCG

[6241] Anonymous. De Glehn compound for the Great Western Railway, England. Scientific American, 90(12):228–229, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-228a.pdf.

Anonymous:1904:NPS

[6242] Anonymous. New pruning shears. Scientific American, 90(12):229, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-229.pdf.

Anonymous:1904:IGV

[6243] Anonymous. An improved grain valve. Scientific American, 90(12):229, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n12/pdf/scientificamerican03191904-229a.pdf.

Anonymous:1904:STH

[6244] Anonymous. Steam turbine of 11,000 horse-power. Scientific American, 90(12):229–230, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
229b.pdf.

Guarini:1904:EUP

[6245] Emile Guarini. Experiments upon the pressure of wind at the Eiffel Tower. Scientific American, 90(12):230, March 19, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n12/pdf/scientificamerican03191904-230.pdf.

Anonymous:1904:TFW

[6246] Anonymous. Thawing out frozen water pipes by electricity from storage batteries. Scientific American, 90(12):231, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n12/pdf/scientificamerican03191904-231a.pdf.

Anonymous:1904:EH

[6247] Anonymous. Electricity in the household. Scientific American, 90(12):232, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
232.pdf.

Anonymous:1904:OKM

[6248] Anonymous. Our knowledge of the Moon. Scientific American, 90 (12):232–234, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
232a.pdf.

Tilden:1904:JSE

[6249] G. H. Tilden. The Japanese sword exercise. — I. Scientific American, 90(12):235–236, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
235.pdf.

Bushnell:1904:CSM

[6250] D. I. Bushnell, Jr. The Cahokia and surrounding mound groups. Scientific American, 90(12):236, March 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n12/pdf/scientificamerican03191904-
236.pdf.

Anonymous:1904:Ce

[6251] Anonymous. Correspondence. Scientific American, 90(13):251, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-251.pdf.

Anonymous:1904:RPIe

[6252] Anonymous. Recently patented inventions. Scientific American, 90(13):257, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
257.pdf.

Anonymous:1904:NQd

[6253] Anonymous. Notes and queries. Scientific American, 90(13):257–258, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-257b.pdf.

Anonymous:1904:BNBd

[6254] Anonymous. Books: New books, etc. Scientific American, 90(13):258, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-258.pdf.

Anonymous:1904:IIe

[6255] Anonymous. Index of inventions. Scientific American, 90(13):258–263, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-258a.pdf.

Anonymous:1904:PNPe

[6256] Anonymous. Patent notes: Patents. Scientific American, 90(13):260, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-260.pdf.

Anonymous:1904:SWM

[6257] Anonymous. Salt-water mains for fire protection. Scientific American, 90(13):246, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
246.pdf.

Anonymous:1904:FES

[6258] Anonymous. Future extension of the subway system. Scientific American, 90(13):246, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
246a.pdf.

Anonymous:1904:FMU

[6259] Anonymous. A floral map of the United States at the St. Louis Exposition. Scientific American, 90(13):246–247, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n13/pdf/scientificamerican03261904-246b.pdf.

Anonymous:1904:BCR

[6260] Anonymous. Borings on a coral reef. Scientific American, 90(13):247, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-247.pdf.

Anonymous:1904:US

[6261] Anonymous. The United States in 1903. Scientific American, 90(13):247, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
247a.pdf.

Anonymous:1904:SNc

[6262] Anonymous. Science notes. Scientific American, 90(13):247, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-247b.pdf.

Anonymous:1904:CP

[6263] Anonymous. Concrete piling. Scientific American, 90(13):248, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-248.pdf.

Anonymous:1904:SCR

[6264] Anonymous. A single-cylinder, reversible, four-cycle gasoline engine. Scientific American, 90(13):249, March 26, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n13/pdf/scientificamerican03261904-249.pdf.

Anonymous:1904:NGMa

[6265] Anonymous. A new gasoline motor for airships. Scientific American, 90(13):249, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
249a.pdf.

Anonymous:1904:NSRa

[6266] Anonymous. A new system of rapid telegraphy. Scientific American, 90(13):249–250, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
249b.pdf.

Anonymous:1904:NVM

[6267] Anonymous. The new Vickers–Maxim 9.2-inch wire-wound gun. Scientific American, 90(13):250, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
250.pdf.

Anonymous:1904:APEa

[6268] Anonymous. The Abruzzi Polar Expedition — I. Scientific American, 90(13):252, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
252.pdf.

Anonymous:1904:HRT

[6269] Anonymous. The Hudson River Tunnel. Scientific American, 90 (13):253–254, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
253.pdf.

Anonymous:1904:PDa

[6270] Anonymous. Patent department. Scientific American, 90(13):255, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-255.pdf.

Anonymous:1904:CLU

[6271] Anonymous. Combination lock for umbrellas. Scientific American, 90(13):255, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
255a.pdf.

Anonymous:1904:HTS

[6272] Anonymous. Harrow tooth and simple fastening device. Scientific American, 90(13):255, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
255b.pdf.

Anonymous:1904:VML

[6273] Anonymous. Valve mechanism for locomotives. Scientific American, 90(13):255, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
255c.pdf.

Anonymous:1904:BD

[6274] Anonymous. Bending device. Scientific American, 90(13):256, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-256.pdf.

Anonymous:1904:IW

[6275] Anonymous. Improved windmill. Scientific American, 90(13):256, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-256a.pdf.

Anonymous:1904:MBRa

[6276] Anonymous. Mail box for rural free delivery routes. Scientific American, 90(13):256, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
256b.pdf.

Anonymous:1904:OSM

[6277] Anonymous. Ore-sampling machine. Scientific American, 90(13):256, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n13/pdf/scientificamerican03261904-256c.pdf.

Anonymous:1904:BPWe

[6278] Anonymous. Business and personal wants. Scientific American, 90(13):257, March 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n13/pdf/scientificamerican03261904-
257a.pdf.

Russell:1904:HAa

[6279] Henry Norris Russell. The heavens in April. Scientific American, 90(14):267, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
267.pdf.

Anonymous:1904:Cf

[6280] Anonymous. Correspondence. Scientific American, 90(14):271, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n14/pdf/scientificamerican04021904-271.pdf.

Anonymous:1904:LNb

[6281] Anonymous. Legal notes. Scientific American, 90(14):276, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n14/pdf/scientificamerican04021904-276.pdf.

Anonymous:1904:RPIf (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
277.pdf.

Anonymous:1904:NQe

[6283] Anonymous. Notes and queries. Scientific American, 90(14):278, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n14/pdf/scientificamerican04021904-278.pdf.

Anonymous:1904:BNBe

[6284] Anonymous. Books: New books, etc. Scientific American, 90(14):279, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n14/pdf/scientificamerican04021904-279.pdf.

Anonymous:1904:IIf

[6285] Anonymous. Index of inventions. Scientific American, 90(14):279–283, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n14/pdf/scientificamerican04021904-279a.pdf.

Anonymous:1904:PNPf

[6286] Anonymous. Patent notes: Patents. Scientific American, 90(14):283, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n14/pdf/scientificamerican04021904-283.pdf.

Anonymous:1904:BS

[6287] Anonymous. The blindness of the submarine. Scientific American, 90(14):266, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
266.pdf.

Anonymous:1904:HEP

[6288] Anonymous. High-explosive projectiles. Scientific American, 90 (14):266, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
266a.pdf.

Anonymous:1904:DWL

[6289] Anonymous. Dispersion and wave length of N-rays. Scientific American, 90(14):266–267, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
266b.pdf.

Anonymous:1904:HSS

[6290] Anonymous. High-speed steam railway service. Scientific American, 90(14):267, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
267a.pdf.

Anonymous:1904:SNd

[6291] Anonymous. Science notes. Scientific American, 90(14):267, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n14/pdf/scientificamerican04021904-267b.pdf.

Anonymous:1904:PND

[6292] Anonymous. The Perspectartigraph — a novel drawing instrument. Scientific American, 90(14):268, April 2, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n14/pdf/scientificamerican04021904-268.pdf.

Tilden:1904:JSI

[6293] G. H. Tilden. Japanese swordmaking. — II. Scientific American, 90(14):269–270, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
269.pdf.

Anonymous:1904:BMA

[6294] Anonymous. The battleship “Mikasa” in action. Scientific American, 90(14):270, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
270.pdf.

Anonymous:1904:APEb

[6295] Anonymous. The Abruzzi Polar Expedition — II. Scientific American, 90(14):272, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
272.pdf.

Anonymous:1904:ERF

[6296] Anonymous. Excavations in the Roman Forum. Scientific American, 90(14):273, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
273.pdf.

Anonymous:1904:ABI

[6297] Anonymous. Accident to the battleship “Illinois”. Scientific American, 90(14):274, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
274.pdf.

Willey:1904:ROU

[6298] Day Allen Willey. Raising ostriches in the United States. Scientific American, 90(14):275, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
275.pdf.

Anonymous:1904:BPWf

[6299] Anonymous. Business and personal wants. Scientific American, 90(14):277, April 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n14/pdf/scientificamerican04021904-
277a.pdf.

Anonymous:1904:Cg

[6300] Anonymous. Correspondence. Scientific American, 90(15):291, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-291.pdf.

Anonymous:1904:RPIg

[6301] Anonymous. Recently patented inventions. Scientific American, 90(15):297, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
297.pdf.

Anonymous:1904:NQf

[6302] Anonymous. Notes and queries. Scientific American, 90(15):297–299, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-297b.pdf.

Anonymous:1904:BNBf

[6303] Anonymous. Books: New books, etc. Scientific American, 90(15):299, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-299.pdf.

Anonymous:1904:IIg

[6304] Anonymous. Index of inventions. Scientific American, 90(15):299–303, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-299a.pdf.

Anonymous:1904:PNPg

[6305] Anonymous. Patent notes: Patents. Scientific American, 90(15):303, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-303.pdf.

Anonymous:1904:FPA

[6306] Anonymous. That fire peril again. Scientific American, 90(15):286, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-286.pdf.

Anonymous:1904:BPA

[6307] Anonymous. The blocking of the Port Arthur entrance. Scientific American, 90(15):286, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
286a.pdf.

Anonymous:1904:DPC

[6308] Anonymous. Departure of the Panama Canal Commission. Scientific American, 90(15):286, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
286b.pdf.

Anonymous:1904:SRC

[6309] Anonymous. Steel roads for country districts. Scientific American, 90(15):286, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
286c.pdf.

Anonymous:1904:CIC

[6310] Anonymous. Cast-iron columns in the Darlington Hotel collapse. Scientific American, 90(15):286–287, April 9, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n15/pdf/scientificamerican04091904-286d.pdf.

Dinger:1904:STF

[6311] Lieut. H. C. Dinger. The steam turbine and its field in marine work. Scientific American, 90(15):287, April 9, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n15/pdf/scientificamerican04091904-287.pdf.

Anonymous:1904:SNe

[6312] Anonymous. Science notes. Scientific American, 90(15):287, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-287a.pdf.

Anonymous:1904:AOL

[6313] Anonymous. Automatic ore loader. Scientific American, 90(15):288, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-288.pdf.

Anonymous:1904:PSM

[6314] Anonymous. Pulling strength of men and animals. Scientific American, 90(15):289–290, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
289.pdf.

Anonymous:1904:NEP

[6315] Anonymous. New electrolytic phenomenon. Scientific American, 90(15):290, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
290.pdf.

Anonymous:1904:MTP

[6316] Anonymous. Method of timing photographic shutters. Scientific American, 90(15):291, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
291a.pdf.

Anonymous:1904:MBCa

[6317] Anonymous. Malting and brewing conducted on scientific principles. — I. malting. Scientific American, 90(15):292–293, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-292.pdf.

Anonymous:1904:RLH

[6318] Anonymous. Russia’s last hope on the sea. Scientific American, 90(15):293–294, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
293.pdf.

Anonymous:1904:IBF

[6319] Anonymous. An interesting balancing feat. Scientific American, 90(15):294, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
294.pdf.

Anonymous:1904:PDb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-295.pdf.

Anonymous:1904:IC

[6321] Anonymous. Improved calipers. Scientific American, 90(15):295, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-295a.pdf.

Anonymous:1904:WRAa

[6322] Anonymous. Windmill with reversing attachment. Scientific American, 90(15):295, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
295b.pdf.

Anonymous:1904:CAB

[6323] Anonymous. Convenient attachment for bird cages. Scientific American, 90(15):295, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
295c.pdf.

Anonymous:1904:OIa

[6324] Anonymous. Oddities in inventions. Scientific American, 90(15):296, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n15/pdf/scientificamerican04091904-296.pdf.

Anonymous:1904:BPWg

[6325] Anonymous. Business and personal wants. Scientific American, 90(15):297, April 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n15/pdf/scientificamerican04091904-
297a.pdf.

Anonymous:1904:Ch

[6326] Anonymous. Correspondence. Scientific American, 90(16):311, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n16/pdf/scientificamerican04161904-311.pdf.

Anonymous:1904:RPIh

[6327] Anonymous. Recently patented inventions. Scientific American, 90(16):316, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
316.pdf.

Anonymous:1904:NQg

[6328] Anonymous. Notes and queries. Scientific American, 90(16):316, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n16/pdf/scientificamerican04161904-316b.pdf.

Anonymous:1904:IIh

[6329] Anonymous. Index of inventions. Scientific American, 90(16):316–319, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n16/pdf/scientificamerican04161904-316c.pdf.

Anonymous:1904:PNPh

[6330] Anonymous. Patent notes: Patents. Scientific American, 90(16):319, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n16/pdf/scientificamerican04161904-319.pdf.

Anonymous:1904:TEN

[6331] Anonymous. Turbine engines for the new Cunarders. Scientific American, 90(16):306, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
306.pdf.

Anonymous:1904:GOS

[6332] Anonymous. The growth of our street railway systems. Scientific American, 90(16):306, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
306a.pdf.

Anonymous:1904:BRT

[6333] Anonymous. A barrage on the River Thames. Scientific American, 90(16):306, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
306b.pdf.

Anonymous:1904:ALU

[6334] Anonymous. Action of light upon the formation of accumulators. Scientific American, 90(16):306, April 16, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n16/pdf/scientificamerican04161904-306c.pdf.

Maury:1904:LC

[6335] A. C. Maury. Light in crystals. Scientific American, 90(16):307, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n16/pdf/scientificamerican04161904-307.pdf.

Anonymous:1904:SNf

[6336] Anonymous. Science notes. Scientific American, 90(16):307, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n16/pdf/scientificamerican04161904-307a.pdf.

Anonymous:1904:MAT

[6337] Anonymous. Military aspects of the Trans-Siberian Railway. Scientific American, 90(16):308–310, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
308.pdf.

Anonymous:1904:RWV

[6338] Anonymous. Resilient wheel for vehicles. Scientific American, 90(16):310, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
310.pdf.

Steens:1904:CMD

[6339] A. Steens. The cylindrical movable dams at Schweinfurt, Germany. Scientific American, 90(16):312, April 16, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n16/pdf/scientificamerican04161904-312.pdf.

Anonymous:1904:BAa

[6340] Anonymous. The beedle airship. Scientific American, 90(16):313, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n16/pdf/scientificamerican04161904-313.pdf.

Guarini:1904:POI

[6341] Emile Guarini. A portable oxygen-inhaling apparatus. Scientific American, 90(16):314, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
314.pdf.

Anonymous:1904:PBN

[6342] Anonymous. A peculiar blaze in the New York Subway. Scientific American, 90(16):315, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
315.pdf.

Anonymous:1904:BPWh

[6343] Anonymous. Business and personal wants. Scientific American, 90(16):316, April 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n16/pdf/scientificamerican04161904-
316a.pdf.

Anonymous:1904:SNg

[6344] Anonymous. Science notes. Scientific American, 90(17):323, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-323b.pdf.

Anonymous:1904:LNc

[6345] Anonymous. Legal notes. Scientific American, 90(17):333, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-333.pdf.

Anonymous:1904:RPIi 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
334.pdf.

Anonymous:1904:BPWi

[6347] Anonymous. Business and personal wants. Scientific American, 90(17):334, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
334a.pdf.

Anonymous:1904:NQh

[6348] Anonymous. Notes and queries. Scientific American, 90(17):334–335, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-334b.pdf.

Anonymous:1904:BNBg

[6349] Anonymous. Books: New books, etc. Scientific American, 90(17):335, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-335.pdf.

Anonymous:1904:IIi

[6350] Anonymous. Index of inventions. Scientific American, 90(17):335–339, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-335a.pdf.

Anonymous:1904:PNPi

[6351] Anonymous. Patent notes: Patents. Scientific American, 90(17):337, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-337.pdf.

Anonymous:1904:BMD

[6352] Anonymous. The battleship “Missouri” disaster. Scientific American, 90(17):322, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
322.pdf.

Anonymous:1904:ADT

[6353] Anonymous. The advantages and drawbacks of turbines for ocean steamers. Scientific American, 90(17):322, April 23, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n17/pdf/scientificamerican04231904-322a.pdf.

Anonymous:1904:EML

[6354] Anonymous. Experiments made at the laboratory in the catacombs; effect of darkness upon animals, etc. Scientific American, 90(17):322–323, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-322b.pdf.

Macartney:1904:CDH

[6355] M. H. H. Macartney. The caves and dene-holes at Chiselhurst. Scientific American, 90(17):323, April 23, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n17/pdf/scientificamerican04231904-323.pdf.

Anonymous:1904:NPMa

[6356] Anonymous. New process of manufacturing ozone. Scientific American, 90(17):323, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
323a.pdf.

Anonymous:1904:CSa

[6357] Anonymous. The current supplement. Scientific American, 90(17):323, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-323c.pdf.

Anonymous:1904:DTP

[6358] Anonymous. Driving a test pile for the Hudson River Tunnel. Scientific American, 90(17):324, April 23, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n17/pdf/scientificamerican04231904-324.pdf.

Anonymous:1904:NECa

[6359] Anonymous. New electric chronometer for timing automobiles. Scientific American, 90(17):325, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
325.pdf.

Perkins:1904:NEA

[6360] Frank C. Perkins. A new English aluminium welding machine. Scientific American, 90(17):325–326, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
325a.pdf.

Comstock:1904:IP

[6361] Daniel F. Comstock. An interesting paradox. Scientific American, 90(17):326, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
326.pdf.

Anonymous:1904:RRP

[6362] Anonymous. The Russian reverse at Port Arthur. Scientific American, 90(17):327, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
327.pdf.

Jones:1904:GPO

[6363] Helen Lukens Jones. The greatest of pipe organs. Scientific American, 90(17):328–329, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
328.pdf.

Anonymous:1904:SMa

[6364] Anonymous. The submarine mine. Scientific American, 90(17):330, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-330.pdf.

Anonymous:1904:CRB

[6365] Anonymous. Concrete railway bridges. Scientific American, 90(17):331, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n17/pdf/scientificamerican04231904-331.pdf.

Gerald:1904:FET

[6366] W. G. Fitz Gerald. A fire escape for tourists and travelers. Scientific American, 90(17):332, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
332.pdf.

Guarini:1904:NIU

[6367] Emile Guarini. A novel Italian unicycle. Scientific American, 90(17):332, April 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n17/pdf/scientificamerican04231904-
332a.pdf.

Anonymous:1904:Ci

[6368] Anonymous. Correspondence. Scientific American, 90(18):347, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n18/pdf/scientificamerican04301904-347.pdf.

Anonymous:1904:RPIj

[6369] Anonymous. Recently patented inventions. Scientific American, 90(18):351, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n18/pdf/scientificamerican04301904-
351.pdf.

Anonymous:1904:NQi

[6370] Anonymous. Notes and queries. Scientific American, 90(18):351–352, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n18/pdf/scientificamerican04301904-351b.pdf.

Anonymous:1904:IIj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n18/pdf/scientificamerican04301904-352.pdf.

Anonymous:1904:PNPj

[6372] Anonymous. Patent notes: Patents. Scientific American, 90(18):353, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n18/pdf/scientificamerican04301904-353.pdf.

Anonymous:1904:ACN

[6373] Anonymous. The Aqueduct Commissioners and the New York City Water Supply. Scientific American, 90(18):342, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n18/pdf/scientificamerican04301904-342.pdf.

Anonymous:1904:LBB

[6374] Anonymous. Longitudinal bulkheads and battleship stability. Scientific American, 90(18):342, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n18/pdf/scientificamerican04301904-
342a.pdf.

Anonymous:1904:PPA

[6375] Anonymous. The preparation of pure argon. Scientific American, 90(18):342–343, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n18/pdf/scientificamerican04301904-
342b.pdf.

Serviss:1904:SF

[6376] Garrett P. Serviss, Jr. Soaring flight. Scientific American, 90(18):343, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n18/pdf/scientificamerican04301904-343.pdf.

Anonymous:1904:SNh

[6377] Anonymous. Science notes. Scientific American, 90(18):343, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n18/pdf/scientificamerican04301904-343a.pdf.

Anonymous:1904:GHS

[6378] Anonymous. A German high-speed locomotive. Scientific American, 90(18):344, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n18/pdf/scientificamerican04301904-
344.pdf.

Mclure:1904:WLF

[6379] W. Frank M’clure. The “Wolvin,” the largest fresh-water steamship afloat. Scientific American, 90(18):345–346, April 30, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n18/pdf/scientificamerican04301904-345.pdf.

Anonymous:1904:THM

[6380] Anonymous. Two home made wireless telegraph receivers. Scientific American, 90(18):347, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n18/pdf/scientificamerican04301904-
347a.pdf.

Anonymous:1904:MIC

[6381] Anonymous. The manufacture of incised, or cut glass. Scientific American, 90(18):348–350, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n18/pdf/scientificamerican04301904-
348.pdf.

Anonymous:1904:AMB

[6382] Anonymous. The automobile and motor-boat races at Nice and Monte Carlo. Scientific American, 90(18):350–351, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n18/pdf/scientificamerican04301904-350.pdf.

Anonymous:1904:BPWj

[6383] Anonymous. Business and personal wants. Scientific American, 90(18):351, April 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n18/pdf/scientificamerican04301904-
351a.pdf.

Russell:1904:HMb

[6384] Henry Norris Russell. The heavens in May. Scientific American, 90(19):359, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
359.pdf.

Anonymous:1904:RPIk

[6385] Anonymous. Recently patented inventions. Scientific American, 90(19):370–372, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
370.pdf.

Anonymous:1904:NQj

[6386] Anonymous. Notes and queries. Scientific American, 90(19):372–374, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-372a.pdf.

Anonymous:1904:BNBh

[6387] Anonymous. Books: New books, etc. Scientific American, 90 (19):375–376, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
375.pdf.

Anonymous:1904:IIk

[6388] Anonymous. Index of inventions. Scientific American, 90(19):376–379, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-376.pdf.

Anonymous:1904:CLP

[6389] Anonymous. Commemoration of the Louisiana Purchase. Scientific American, 90(19):358, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
358.pdf.

Anonymous:1904:BTB

[6390] Anonymous. Battleship or torpedo boat. Scientific American, 90(19):358, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
358a.pdf.

Anonymous:1904:SIF

[6391] Anonymous. Some interesting facts about the development of the Edison storage battery. Scientific American, 90(19):358–359, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-358b.pdf.

Anonymous:1904:ARM

[6392] Anonymous. Action of radium on metals. Scientific American, 90(19):359, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
359a.pdf.

Anonymous:1904:CSb

[6393] Anonymous. The current supplement. Scientific American, 90(19):359, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-359b.pdf.

Anonymous:1904:ETPa

[6394] Anonymous. Experiments with three-phase electric arcs. Scientific American, 90(19):360, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
360.pdf.

Anonymous:1904:RIB

[6395] Anonymous. The Russian ice-breaker “Ermack”. Scientific American, 90(19):360, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
360a.pdf.

Anonymous:1904:HVI

[6396] Anonymous. High-velocity 6-inch wire-wound gun for the United States Army. Scientific American, 90(19):361–362, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n19/pdf/scientificamerican05071904-361.pdf.

Anonymous:1904:GON

[6397] Anonymous. Growth of our Navy. Scientific American, 90(19):363, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-363.pdf.

Williams:1904:LPE

[6398] Samuel Williams. The Louisiana Purchase Exposition — its scope and its purpose. Scientific American, 90(19):364–365, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-364.pdf.

Wetmore:1904:LCW

[6399] Claude H. Wetmore. The largest clock in the world. Scientific American, 90(19):365–366, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
365.pdf.

Anonymous:1904:PDc

[6400] Anonymous. Patent department. Scientific American, 90(19):367, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-367.pdf.

Anonymous:1904:NECb

[6401] Anonymous. A new electric clock. Scientific American, 90(19):367, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-367a.pdf.

Anonymous:1904:SL

[6402] Anonymous. Strainer for locomotives. Scientific American, 90(19):367, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-367b.pdf.

Anonymous:1904:HLP

[6403] Anonymous. The hydraulic letter press. Scientific American, 90 (19):367–368, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
367c.pdf.

Anonymous:1904:NMB

[6404] Anonymous. A novel mail-bag catcher. Scientific American, 90(19):368, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-368.pdf.

Anonymous:1904:NW

[6405] Anonymous. A new wrench. Scientific American, 90(19):368, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n19/pdf/scientificamerican05071904-368a.pdf.

Anonymous:1904:CTR

[6406] Anonymous. A combined third-rail contact shoe and sleet remover. Scientific American, 90(19):368, May 7, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n19/pdf/scientificamerican05071904-368b.pdf.

Anonymous:1904:BPWk

[6407] Anonymous. Business and personal wants. Scientific American, 90(19):372, May 7, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n19/pdf/scientificamerican05071904-
372.pdf.

Anonymous:1904:Cj

[6408] Anonymous. Correspondence. Scientific American, 90(2):27, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-27.pdf.

Anonymous:1904:RPIl 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
32.pdf.

Anonymous:1904:NQk

[6410] Anonymous. Notes and queries. Scientific American, 90(2):32, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-32b.pdf.

Anonymous:1904:IIl

[6411] Anonymous. Index of inventions. Scientific American, 90(2):32–35, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-32c.pdf.

Anonymous:1904:PNPk

[6412] Anonymous. Patent notes: Patents. Scientific American, 90(2):34, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-34.pdf.

Anonymous:1904:CSPa

[6413] Anonymous. A call for stronger passenger cars. Scientific American, 90(2):22, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
22.pdf.

Anonymous:1904:CTF

[6414] Anonymous. Can the theater fire be prevented? Scientific American, 90(2):22, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
22a.pdf.

Anonymous:1904:NSB

[6415] Anonymous. A new study of bird life. Scientific American, 90(2):22–23, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-22b.pdf.

Anonymous:1904:DJM

[6416] Anonymous. The death of Jeremiah M. Allen. Scientific American, 90(2):23, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
23.pdf.

Fyfe:1904:NBS

[6417] H. C. Fyfe. The new British Shipbuilding Programme. Scientific American, 90(2):23, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
23a.pdf.

Anonymous:1904:SNi

[6418] Anonymous. Science notes. Scientific American, 90(2):23, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-23b.pdf.

Mclure:1904:FT

[6419] W. Frank M’clure. A floating theater. Scientific American, 90(2):24, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-24.pdf.

Guarini:1904:NCA

[6420] Emile Guarini. A new chloroforming apparatus. Scientific American, 90(2):24–25, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
24a.pdf.

Anonymous:1904:EME

[6421] Anonymous. Electricity on the Miami and Erie Canal. Scientific American, 90(2):25–26, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
25.pdf.

Anonymous:1904:OWL

[6422] Anonymous. The oldest working locomotive in the world. Scientific American, 90(2):28, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
28.pdf.

Anonymous:1904:MA

[6423] Anonymous. The manufacture of automobiles. Scientific American, 90(2):29–30, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
29.pdf.

Holder:1904:DSS

[6424] Charles F. Holder. Deep-sea sunfish. Scientific American, 90(2):30–31, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-30.pdf.

Burti:1904:HPE

[6425] V. Burti. How a python eats. Scientific American, 90(2):31, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n2/pdf/scientificamerican01091904-31.pdf.

Anonymous:1904:BPWl

[6426] Anonymous. Business and personal wants. Scientific American, 90(2):32, January 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n2/pdf/scientificamerican01091904-
32a.pdf.

Anonymous:1904:Ck

[6427] Anonymous. Correspondence. Scientific American, 90(20):387, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n20/pdf/scientificamerican05141904-387.pdf.

Anonymous:1904:RPIm

[6428] Anonymous. Recently patented inventions. Scientific American, 90(20):391–392, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
391a.pdf.

Anonymous:1904:BPWm

[6429] Anonymous. Business and personal wants. Scientific American, 90(20):392, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
392.pdf.

Anonymous:1904:NQl

[6430] Anonymous. Notes and queries. Scientific American, 90(20):392, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n20/pdf/scientificamerican05141904-392a.pdf.

Anonymous:1904:IIm

[6431] Anonymous. Index of inventions. Scientific American, 90(20):392–395, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n20/pdf/scientificamerican05141904-392b.pdf.

Anonymous:1904:PNPl

[6432] Anonymous. Patent notes: Patents. Scientific American, 90(20):395, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n20/pdf/scientificamerican05141904-395.pdf.

Anonymous:1904:ETTa

[6433] Anonymous. Electric traction tests at the worlds fair. Scientific American, 90(20):382, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
382.pdf.

Anonymous:1904:TOS

[6434] Anonymous. Turbine ocean steamers. Scientific American, 90(20):382, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n20/pdf/scientificamerican05141904-382a.pdf.

Anonymous:1904:NSTa

[6435] Anonymous. New system of tunnel construction. Scientific American, 90(20):382, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
382b.pdf.

Anonymous:1904:OAI

[6436] Anonymous. Our available iron-ore supply. Scientific American, 90(20):382, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
382c.pdf.

Anonymous:1904:SNR

[6437] Anonymous. On some novel N-ray phenomena. Scientific American, 90(20):382–383, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
382d.pdf.

Metzenbaum:1904:IRA

[6438] M. Metzenbaum. Induced radio-activity and aluminium. Scientific American, 90(20):383, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
383.pdf.

Anonymous:1904:SNj

[6439] Anonymous. Science notes. Scientific American, 90(20):383, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n20/pdf/scientificamerican05141904-383a.pdf.

Forster:1904:NMI

[6440] Edward J. Forster. New method of inflating balloons in war time. Scientific American, 90(20):384, May 14, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n20/pdf/scientificamerican05141904-384.pdf.

Anonymous:1904:CYI

[6441] Anonymous. The challenging yacht “Ingomar”. Scientific American, 90(20):385, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
385.pdf.

Anonymous:1904:AML

[6442] Anonymous. Automatic match lighter operated by alarm clock. Scientific American, 90(20):385–386, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
385a.pdf.

Tristan:1904:IMA

[6443] J. Fid Tristan. An improved Michaud apparatus. Scientific American, 90(20):386, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
386.pdf.

Anonymous:1904:TBM

[6444] Anonymous. The torpedo boat in modern warfare. Scientific American, 90(20):387, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
387a.pdf.

Beasley:1904:RAI

[6445] Walter L. Beasley. Ruins of the Ancient Inca Empire. Scientific American, 90(20):388–390, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
388.pdf.

Anonymous:1904:SARa

[6446] Anonymous. The small arms of the Russian and Japanese Armies. Scientific American, 90(20):390–391, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
390.pdf.

Anonymous:1904:NYSa

[6447] Anonymous. New York’s spring automobile parade. Scientific American, 90(20):391, May 14, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n20/pdf/scientificamerican05141904-
391.pdf.

Anonymous:1904:Cl

[6448] Anonymous. Correspondence. Scientific American, 90(21):403, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-403.pdf.

Anonymous:1904:LNd

[6449] Anonymous. Legal notes. Scientific American, 90(21):409, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-409.pdf.

Anonymous:1904:RPIn

[6450] Anonymous. Recently patented inventions. Scientific American, 90(21):410, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
410.pdf.

Anonymous:1904:NQm

[6451] Anonymous. Notes and queries. Scientific American, 90(21):410–411, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-410b.pdf.

Anonymous:1904:BNBi

[6452] Anonymous. Books: New books, etc. Scientific American, 90(21):411, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-411.pdf.

Anonymous:1904:IIn

[6453] Anonymous. Index of inventions. Scientific American, 90(21):411–415, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-411a.pdf.

Anonymous:1904:PNPm

[6454] Anonymous. Patent notes: Patents. Scientific American, 90(21):415, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-415.pdf.

Anonymous:1904:RCC

[6455] Anonymous. Rear collisions on city rapid transit systems. Scientific American, 90(21):398, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
398.pdf.

Anonymous:1904:ENG

[6456] Anonymous. Economy of narrow-gage railroads. Scientific American, 90(21):398, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
398a.pdf.

Anonymous:1904:RGA

[6457] Anonymous. Results of the German Antarctic Expedition. Scientific American, 90(21):398, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
398b.pdf.

Anonymous:1904:CT

[6458] Anonymous. The catalpa tree. Scientific American, 90(21):398–399, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-398c.pdf.

Love:1904:IWF

[6459] Robertus Love. Illumination of the worlds fair. Scientific American, 90(21):399, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
399.pdf.

Anonymous:1904:DHM

[6460] Anonymous. Death of Henry M. Stanley. Scientific American, 90(21):399, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
399a.pdf.

Anonymous:1904:DSW

[6461] Anonymous. The distillation of sea water at Suakim. Scientific American, 90(21):400, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
400.pdf.

Anonymous:1904:RAG

[6462] Anonymous. The Russian Army and its guns. Scientific American, 90(21):400–402, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
400a.pdf.

Anonymous:1904:MMC

[6463] Anonymous. The manufacture of mens clothing on a large scale. Scientific American, 90(21):404–406, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
404.pdf.

Anonymous:1904:IED

[6464] Anonymous. Interesting experiments to determine the correct lines for gasoline motor boats. Scientific American, 90(21):406–407, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-406.pdf.

Fitz-Gerald:1904:LSA

[6465] William G. Fitz-Gerald. Life-saving apparatus in British schools. Scientific American, 90(21):407, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
407.pdf.

Anonymous:1904:PDd

[6466] Anonymous. Patent department. Scientific American, 90(21):408, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-408.pdf.

Anonymous:1904:NRB

[6467] Anonymous. New riveter and brace. Scientific American, 90(21):408, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n21/pdf/scientificamerican05211904-408a.pdf.

Anonymous:1904:ECL

[6468] Anonymous. Express call for large office buildings. Scientific American, 90(21):408, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
408b.pdf.

Anonymous:1904:BPWn

[6469] Anonymous. Business and personal wants. Scientific American, 90(21):410, May 21, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n21/pdf/scientificamerican05211904-
410a.pdf.

Anonymous:1904:RPIo

[6470] Anonymous. Recently patented inventions. Scientific American, 90(22):427, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
427.pdf.

Anonymous:1904:NQn

[6471] Anonymous. Notes and queries. Scientific American, 90(22):427, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n22/pdf/scientificamerican05281904-427b.pdf.

Anonymous:1904:BNBj

[6472] Anonymous. Books: New books, etc. Scientific American, 90 (22):427–428, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
427c.pdf.

Anonymous:1904:IIo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n22/pdf/scientificamerican05281904-428.pdf.

Anonymous:1904:PNPn

[6474] Anonymous. Patent notes: Patents. Scientific American, 90(22):431, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n22/pdf/scientificamerican05281904-431.pdf.

Anonymous:1904:IRFa

[6475] Anonymous. Increasing railroad fatalities. Scientific American, 90 (22):418, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
418.pdf.

Anonymous:1904:JDP

[6476] Anonymous. The Japanese disaster at Port Arthur. Scientific American, 90(22):418, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
418a.pdf.

Jane:1904:RNG

[6477] Fred T. Jane. Russian Naval guns. Scientific American, 90(22):418–419, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n22/pdf/scientificamerican05281904-418b.pdf.

Horwill:1904:ME

[6478] Herbert W. Horwill. The mother of exhibitions. Scientific American, 90(22):419, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
419.pdf.

Anonymous:1904:TBAa

[6479] Anonymous. Transporter bridge across the Manchester Ship Canal. Scientific American, 90(22):420, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
420.pdf.

Anonymous:1904:BCW

[6480] Anonymous. A bad case of warped railroad track. Scientific American, 90(22):421, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
421.pdf.

Guarini:1904:C

[6481] Emile Guarini. The cyclograph. Scientific American, 90(22):421–422, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n22/pdf/scientificamerican05281904-421a.pdf.

Kunz:1904:FDJ

[6482] George F. Kunz. Facsimile of Dr. Joseph Priestley’s invoice on coming to America in 1794. Scientific American, 90(22):422, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n22/pdf/scientificamerican05281904-422.pdf.

Anonymous:1904:EOC

[6483] Anonymous. An electrically-operated clock. Scientific American, 90(22):423, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
423.pdf.

Anonymous:1904:CWB

[6484] Anonymous. The casting of the Williamsburg Bridge entablatures — a notable technical achievement. Scientific American, 90(22):424–425, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n22/pdf/scientificamerican05281904-424.pdf.

Anonymous:1904:JAG

[6485] Anonymous. The Japanese Army and its guns. Scientific American, 90(22):425–426, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
425.pdf.

Anonymous:1904:CDL

[6486] Anonymous. Capt. Doenvig’s life-saving globe. Scientific American, 90(22):426, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
426.pdf.

Anonymous:1904:BPWo

[6487] Anonymous. Business and personal wants. Scientific American, 90(22):427, May 28, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n22/pdf/scientificamerican05281904-
427a.pdf.

Russell:1904:HJb

[6488] Henry Norris Russell. The heavens in June. Scientific American, 90(23):435, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
435.pdf.

Anonymous:1904:Cm

[6489] Anonymous. Correspondence. Scientific American, 90(23):439, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-439.pdf.

Anonymous:1904:RPIp

[6490] Anonymous. Recently patented inventions. Scientific American, 90(23):446–447, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
446.pdf.

Anonymous:1904:NQo

[6491] Anonymous. Notes and queries. Scientific American, 90(23):447, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-447a.pdf.

Anonymous:1904:BNBk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-447b.pdf.

Anonymous:1904:IIp

[6493] Anonymous. Index of inventions. Scientific American, 90(23):447–451, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-447c.pdf.

Anonymous:1904:PNPo

[6494] Anonymous. Patent notes: Patents. Scientific American, 90(23):451, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-451.pdf.

Anonymous:1904:RTR

[6495] Anonymous. The N-rays: Are they real or illusory? Scientific American, 90(23):434, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
434.pdf.

Anonymous:1904:SMH

[6496] Anonymous. Submarine mines on the high seas. Scientific American, 90(23):434, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
434a.pdf.

Anonymous:1904:WCA

[6497] Anonymous. A World’s Congress of Academies. Scientific American, 90(23):434–435, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
434b.pdf.

Anonymous:1904:CHH

[6498] Anonymous. Charles H. Haswell’s ninety-fifth birthday. Scientific American, 90(23):435, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
435a.pdf.

Perkins:1904:BAE

[6499] Frank C. Perkins. Belgian and American electric lifting magnets. Scientific American, 90(23):436, June 4, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n23/pdf/scientificamerican06041904-436.pdf.

Anonymous:1904:EOV

[6500] Anonymous. Electrically-operated versus steam-driven mine hoists. Scientific American, 90(23):437, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
437.pdf.

Anonymous:1904:JND

[6501] Anonymous. The Japanese Naval disasters. Scientific American, 90(23):438, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
438.pdf.

Anonymous:1904:NFR

[6502] Anonymous. New form of radiation. Scientific American, 90(23):438, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-438a.pdf.

Anonymous:1904:EP

[6503] Anonymous. Electrochemical phonograph. Scientific American, 90(23):438, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
438b.pdf.

Mclure:1904:MSH

[6504] W. Frank M’clure. The manufacture of straw hats. Scientific American, 90(23):440–441, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
440.pdf.

Crafts:1904:PRC

[6505] H. A. Crafts. The petroleum reservoirs of California. Scientific American, 90(23):441, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
441.pdf.

Anonymous:1904:ANY

[6506] Anonymous. An arcade for New York City. Scientific American, 90(23):442, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
442.pdf.

Anonymous:1904:WRC

[6507] Anonymous. Wreck of the Russian cruiser “Bogatyr”. Scientific American, 90(23):442, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
442a.pdf.

Anonymous:1904:PJS

[6508] Anonymous. Photography of Jupiter’s satellites. Scientific American, 90(23):443, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
443.pdf.

Gradenwitz:1904:PSS

[6509] Alfred Gradenwitz. The Pielock superheater system. Scientific American, 90(23):443, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
443a.pdf.

Anonymous:1904:PDe

[6510] Anonymous. Patent department. Scientific American, 90(23):444, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-444.pdf.

Anonymous:1904:RDA

[6511] Anonymous. Rock drill attachment. Scientific American, 90(23):444, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-444a.pdf.

Anonymous:1904:NFC

[6512] Anonymous. A novel fly-catcher. Scientific American, 90(23):444, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-444b.pdf.

Anonymous:1904:IRC

[6513] Anonymous. Improved railway chair. Scientific American, 90(23):444, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-444c.pdf.

Anonymous:1904:CDR

[6514] Anonymous. Compensating device for rope tramways. Scientific American, 90(23):444–445, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
444d.pdf.

Anonymous:1904:TRO

[6515] Anonymous. A tension regulator for overhead cables. Scientific American, 90(23):445, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
445.pdf.

Anonymous:1904:CRa

[6516] Anonymous. A clothing rack. Scientific American, 90(23):445, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n23/pdf/scientificamerican06041904-445a.pdf.

Anonymous:1904:DDN

[6517] Anonymous. Device for distributing noxious fumes. Scientific American, 90(23):445, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
445b.pdf.

Anonymous:1904:IPT

[6518] Anonymous. An improved portable tent. Scientific American, 90(23):445, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
445c.pdf.

Anonymous:1904:BPWp

[6519] Anonymous. Business and personal wants. Scientific American, 90(23):447, June 4, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n23/pdf/scientificamerican06041904-
447.pdf.

Anonymous:1904:Cn

[6520] Anonymous. Correspondence. Scientific American, 90(24):459, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-459.pdf.

Anonymous:1904:LNe

[6521] Anonymous. Legal notes. Scientific American, 90(24):465, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-465.pdf.

Anonymous:1904:RPIq

[6522] Anonymous. Recently patented inventions. Scientific American, 90(24):466, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
466.pdf.

Anonymous:1904:NQp

[6523] Anonymous. Notes and queries. Scientific American, 90(24):467, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-467.pdf.

Anonymous:1904:BNBl

[6524] Anonymous. Books: New books, etc. Scientific American, 90(24):467, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-467a.pdf.

Anonymous:1904:IIq

[6525] Anonymous. Index of inventions. Scientific American, 90(24):467–471, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-467b.pdf.

Anonymous:1904:PNPp

[6526] Anonymous. Patent notes: Patents. Scientific American, 90(24):471, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-471.pdf.

Anonymous:1904:GAM

[6527] Anonymous. Growth of American merchant marine. Scientific American, 90(24):454, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
454.pdf.

Anonymous:1904:EKJ

[6528] Anonymous. Electricity in Korea and Japan. Scientific American, 90(24):454, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
454a.pdf.

Anonymous:1904:PR

[6529] Anonymous. The Polar regions. Scientific American, 90(24):454–455, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-454b.pdf.

Anonymous:1904:CCC

[6530] Anonymous. Calcium carbide and its commercial development. Scientific American, 90(24):455, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
455.pdf.

Anonymous:1904:SNk

[6531] Anonymous. Science notes. Scientific American, 90(24):455, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-455a.pdf.

Anonymous:1904:FAA

[6532] Anonymous. The first American auto-boat contest. Scientific American, 90(24):456, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
456.pdf.

Anonymous:1904:BNS

[6533] Anonymous. The British Naval School for the Training of Divers. Scientific American, 90(24):457–458, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
457.pdf.

Anonymous:1904:STTa

[6534] Anonymous. Simultaneous telegraphing and telephoning. Scientific American, 90(24):459, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
459a.pdf.

Jones:1904:ABT

[6535] Harry Dillon Jones. An American Buddhist temple. Scientific American, 90(24):460, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
460.pdf.

Anonymous:1904:CRT

[6536] Anonymous. Completing the rapid transit subway. Scientific American, 90(24):461–462, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
461.pdf.

Anonymous:1904:PLC

[6537] Anonymous. Proposed lighthouse for Cape Hatteras Diamond Shoal. Scientific American, 90(24):462, June 11, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n24/pdf/scientificamerican06111904-462.pdf.

Anonymous:1904:IRS

[6538] Anonymous. Improved railway substructure. Scientific American, 90(24):463, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
463.pdf.

Anonymous:1904:NTRa

[6539] Anonymous. A new tennis racket. Scientific American, 90(24):464, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-464.pdf.

Anonymous:1904:RABb

[6540] Anonymous. Runner attachment for bicycles. Scientific American, 90(24):464, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
464a.pdf.

Anonymous:1904:PBC

[6541] Anonymous. Pipe or bar cutter. Scientific American, 90(24):464, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-464b.pdf.

Anonymous:1904:VW

[6542] Anonymous. Vehicle wheel. Scientific American, 90(24):464, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n24/pdf/scientificamerican06111904-464c.pdf.

Anonymous:1904:MMG

[6543] Anonymous. Machine for moistening gummed paper strips. Scientific American, 90(24):464, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
464d.pdf.

Anonymous:1904:BPWq

[6544] Anonymous. Business and personal wants. Scientific American, 90(24):466, June 11, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n24/pdf/scientificamerican06111904-
466a.pdf.

Anonymous:1904:Co

[6545] Anonymous. Correspondence. Scientific American, 90(25):479, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-479.pdf.

Anonymous:1904:RPIr

[6546] Anonymous. Recently patented inventions. Scientific American, 90(25):483, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n25/pdf/scientificamerican06181904-
483a.pdf.

Anonymous:1904:NQq

[6547] Anonymous. Notes and queries. Scientific American, 90(25):484, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-484.pdf.

Anonymous:1904:IIr

[6548] Anonymous. Index of inventions. Scientific American, 90(25):484–487, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-484a.pdf.

Anonymous:1904:PNPq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-487.pdf.

Anonymous:1904:MSV

[6550] Anonymous. Military and strategical valve of oil fuel. Scientific American, 90(25):474, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n25/pdf/scientificamerican06181904-
474.pdf.

Anonymous:1904:SIO

[6551] Anonymous. Surface indications of oil. Scientific American, 90(25):474, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-474a.pdf.

McNeely:1904:FPB

[6552] R. W. McNeely. Fine performance of the battleship “Kentucky”. Scientific American, 90(25):474–475, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v90/n25/pdf/scientificamerican06181904-
474b.pdf.

Anonymous:1904:SNl

[6553] Anonymous. Science notes. Scientific American, 90(25):475, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-475.pdf.

Anonymous:1904:DM

[6554] Anonymous. The dredger “Marquess”. Scientific American, 90(25):476, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-476.pdf.

Guarini:1904:BAE

[6555] Emile Guarini. The Berthier actino-electric transformer. Scientific American, 90(25):476, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n25/pdf/scientificamerican06181904-
476a.pdf.

Anonymous:1904:NAPa

[6556] Anonymous. New apparatus for photographing the background of the eye. Scientific American, 90(25):477, June 18, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n25/pdf/scientificamerican06181904-477.pdf.

Anonymous:1904:ACL

[6557] Anonymous. Automatic control of locomotives. Scientific American, 90(25):477–478, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n25/pdf/scientificamerican06181904-
477a.pdf.

Anonymous:1904:MBRb

[6558] Anonymous. Mail box for rural delivery service. Scientific American, 90(25):478, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n25/pdf/scientificamerican06181904-
478.pdf.

Anonymous:1904:MBCb

[6559] Anonymous. Malting and brewing conducted on scientific principles — II brewing. Scientific American, 90(25):480–481, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-480.pdf.

Anonymous:1904:RSAa

[6560] Anonymous. Relics of the Spanish Armada. Scientific American, 90(25):482–483, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n25/pdf/scientificamerican06181904-
482.pdf.

Anonymous:1904:SB

[6561] Anonymous. A sectional bookcase. Scientific American, 90(25):483, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n25/pdf/scientificamerican06181904-483.pdf.

Anonymous:1904:BPWr

[6562] Anonymous. Business and personal wants. Scientific American, 90(25):483, June 18, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n25/pdf/scientificamerican06181904-
483b.pdf.

Anonymous:1904:Cq

[6563] Anonymous. Correspondence. Scientific American, 90(26):495, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-495.pdf.

Anonymous:1904:RPIs

[6564] Anonymous. Recently patented inventions. Scientific American, 90(26):501, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
501.pdf.

Anonymous:1904:NQr

[6565] Anonymous. Notes and queries. Scientific American, 90(26):501, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-501b.pdf.

Anonymous:1904:IIs

[6566] Anonymous. Index of inventions. Scientific American, 90(26):501–507, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-501c.pdf.

Anonymous:1904:PNPr

[6567] Anonymous. Patent notes: Patents. Scientific American, 90(26):507, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-507.pdf.

Anonymous:1904:EPOa

[6568] Anonymous. Edison and the Patent Office. Scientific American, 90(26):490, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
490.pdf.

Anonymous:1904:WDG

[6569] Anonymous. Wind-driven generators for farming. Scientific American, 90(26):490, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
490a.pdf.

Anonymous:1904:WAI

[6570] Anonymous. Wood alcohol for industrial purposes. Scientific American, 90(26):490–491, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
490b.pdf.

Anonymous:1904:IARa

[6571] Anonymous. The international automobile race. Scientific American, 90(26):491, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
491.pdf.

Anonymous:1904:MCW

[6572] Anonymous. M. Curie’s work on radium emanations from mineral springs. Scientific American, 90(26):491, June 25, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n26/pdf/scientificamerican06251904-491a.pdf.

Anonymous:1904:CTH

[6573] Anonymous. Cameos and their history. Scientific American, 90(26):491, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-491b.pdf.

Anonymous:1904:NSTb

[6574] Anonymous. Non skidding tires and protective bands for use on the wheels of automobiles. Scientific American, 90(26):492, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-492.pdf.

King:1904:SIE

[6575] Charles R. King. Successful Italian Express Compound Locomotive. Scientific American, 90(26):492–493, June 25, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n26/pdf/scientificamerican06251904-492a.pdf.

Anonymous:1904:LDL

[6576] Anonymous. Looping the double loop. Scientific American, 90(26):493, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-493.pdf.

Gradenwitz:1904:APP

[6577] Alfred Gradenwitz. The autoplastic photographic process of producing sculptural portraits. Scientific American, 90(26):493–494, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-493a.pdf.

Ginther:1904:TGC

[6578] C. M. Ginther. The transformation of the Grand Cañon into a National Park. Scientific American, 90(26):494, June 25, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n26/pdf/scientificamerican06251904-494.pdf.

Beasley:1904:MAC

[6579] Walter L. Beasley. Modeling animals in clay the passing of taxidermy. Scientific American, 90(26):496, June 25, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n26/pdf/scientificamerican06251904-496.pdf.

Anonymous:1904:SMB

[6580] Anonymous. The second motor-boat race. Scientific American, 90(26):497, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
497.pdf.

Anonymous:1904:BAR

[6581] Anonymous. The Boston Automobile Race. Scientific American, 90(26):498, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
498.pdf.

Anonymous:1904:PDf

[6582] Anonymous. Patent department. Scientific American, 90(26):499, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-499.pdf.

Anonymous:1904:SG

[6583] Anonymous. Shingle gage. Scientific American, 90(26):499, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-499a.pdf.

Anonymous:1904:IB

[6584] Anonymous. Improved buckle. Scientific American, 90(26):499, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-499b.pdf.

Anonymous:1904:LPO

[6585] Anonymous. Loose pulley oiler. Scientific American, 90(26):499, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-499c.pdf.

Anonymous:1904:RTF

[6586] Anonymous. Railway tie and fastening. Scientific American, 90(26):499, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-499d.pdf.

Anonymous:1904:CBM

[6587] Anonymous. Cigarette box with match receptacle. Scientific American, 90(26):500, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
500.pdf.

Anonymous:1904:AEF

[6588] Anonymous. Automatic electric fire alarm. Scientific American, 90(26):500, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
500a.pdf.

Anonymous:1904:ISB

[6589] Anonymous. Improved support for bicycle handle bars. Scientific American, 90(26):500, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
500b.pdf.

Anonymous:1904:DA

[6590] Anonymous. Dredge attachment. Scientific American, 90(26):500, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n26/pdf/scientificamerican06251904-500c.pdf.

Anonymous:1904:BPWs

[6591] Anonymous. Business and personal wants. Scientific American, 90(26):501, June 25, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n26/pdf/scientificamerican06251904-
501a.pdf.

Anonymous:1904:Cu

[6592] Anonymous. Correspondence. Scientific American, 90(3):43, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-43.pdf.

Anonymous:1904:LNf

[6593] Anonymous. Legal notes. Scientific American, 90(3):48, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-48.pdf.

Anonymous:1904:RPIt

[6594] Anonymous. Recently patented inventions. Scientific American, 90(3):50–51, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n3/pdf/scientificamerican01161904-
50.pdf.

Anonymous:1904:NQs

[6595] Anonymous. Notes and queries. Scientific American, 90(3):51–52, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-51a.pdf.

Anonymous:1904:BNBm

[6596] Anonymous. Books: New books, etc. Scientific American, 90 (3):52–53, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n3/pdf/scientificamerican01161904-
52.pdf.

Anonymous:1904:IIt

[6597] Anonymous. Index of inventions. Scientific American, 90(3):53–55, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-53.pdf.

Anonymous:1904:PNPs

[6598] Anonymous. Patent notes: Patents. Scientific American, 90(3):56, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-56.pdf.

Anonymous:1904:PRB

[6599] Anonymous. Proposed reconstruction of the Brooklyn Bridge. Scientific American, 90(3):38, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n3/pdf/scientificamerican01161904-
38.pdf.

Anonymous:1904:LS

[6600] Anonymous. Lunar superstitions. Scientific American, 90(3):38, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-38a.pdf.

Anonymous:1904:PMI

[6601] Anonymous. The problem of the mutual influence of cathode rays. Scientific American, 90(3):38, January 16, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n3/pdf/scientificamerican01161904-38b.pdf.

Anonymous:1904:AAA

[6602] Anonymous. The American Association for the Advancement of Science. Scientific American, 90(3):39, January 16, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n3/pdf/scientificamerican01161904-39.pdf.

Anonymous:1904:FWT

[6603] Anonymous. The De Forest wireless tests across the Irish Channel. Scientific American, 90(3):40, January 16, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n3/pdf/scientificamerican01161904-40.pdf.

Anonymous:1904:SHN

[6604] Anonymous. The Siemens & Halske new electric hand-drill. Scientific American, 90(3):41, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n3/pdf/scientificamerican01161904-
41.pdf.

Anonymous:1904:OAD

[6605] Anonymous. The Orthodiagraph. an apparatus for determining the true shape and size of internal organs. Scientific American, 90(3):41–42, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-41a.pdf.

Anonymous:1904:PTN

[6606] Anonymous. The Pennsylvania Tunnel at New York City. Scientific American, 90(3):42, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n3/pdf/scientificamerican01161904-
42.pdf.

Anonymous:1904:BHA

[6607] Anonymous. The Barton 150-horsepower airship’s forth-coming trial trip. Scientific American, 90(3):44–46, January 16, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n3/pdf/scientificamerican01161904-44.pdf.

Beasley:1904:RDB

[6608] Walter L. Beasley. Remarkable discoveries by Bandelier of Inca civilization in Peru. Scientific American, 90(3):46–47, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-46.pdf.

Spaid:1904:HR

[6609] A. R. M. Spaid. A harmless reptile. Scientific American, 90(3):47, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-47.pdf.

Anonymous:1904:PDg

[6610] Anonymous. Patent department. Scientific American, 90(3):49, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-49.pdf.

Anonymous:1904:OIb

[6611] Anonymous. Oddities in inventions. Scientific American, 90(3):49, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n3/pdf/scientificamerican01161904-49a.pdf.

Anonymous:1904:BPWt

[6612] Anonymous. Business and personal wants. Scientific American, 90(3):51, January 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n3/pdf/scientificamerican01161904-
51.pdf.

Anonymous:1904:RPIu 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
68.pdf.

Anonymous:1904:NQt

[6614] Anonymous. Notes and queries. Scientific American, 90(4):68, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n4/pdf/scientificamerican01231904-68b.pdf.

Anonymous:1904:IIu

[6615] Anonymous. Index of inventions. Scientific American, 90(4):68–71, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n4/pdf/scientificamerican01231904-68c.pdf.

Anonymous:1904:PNPt

[6616] Anonymous. Patent notes: Patents. Scientific American, 90(4):71, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n4/pdf/scientificamerican01231904-71.pdf.

Anonymous:1904:TNB

[6617] Anonymous. Tammany and the new bridge. Scientific American, 90(4):58, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
58.pdf.

Anonymous:1904:NRJ

[6618] Anonymous. The Navies of Russia and Japan compared. Scientific American, 90(4):58, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
58a.pdf.

Anonymous:1904:DDD

[6619] Anonymous. Discovery of Dr. Danysz, of Pasteur Institute, for exterminating rats. Scientific American, 90(4):58–59, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n4/pdf/scientificamerican01231904-58b.pdf.

Lautenschlaeger:1904:TFP

[6620] Carl Lautenschlaeger. The theatre fire and its prevention in Germany. Scientific American, 90(4):59, January 23, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n4/pdf/scientificamerican01231904-59.pdf.

Anonymous:1904:SNm

[6621] Anonymous. Science notes. Scientific American, 90(4):59, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n4/pdf/scientificamerican01231904-59a.pdf.

Mclure:1904:GMT

[6622] W. Frank M’clure. A great mammoth’s tooth. Scientific American, 90(4):60, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
60.pdf.

Jones:1904:ESC

[6623] George J. Jones. Electric sleeping cars. Scientific American, 90 (4):60–61, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
60a.pdf.

Anonymous:1904:SLB

[6624] Anonymous. A simple laboratory blowpipe apparatus. Scientific American, 90(4):61, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
61.pdf.

Anonymous:1904:NJA

[6625] Anonymous. New Japanese armored cruisers “Kasaga” and “Niasin”. Scientific American, 90(4):61–62, January 23, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n4/pdf/scientificamerican01231904-61a.pdf.

Anonymous:1904:MBA

[6626] Anonymous. The Manhattan Bridge across the East River. Scientific American, 90(4):62–63, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
62.pdf.

Anonymous:1904:ISM

[6627] Anonymous. The invention of the sewing machine. Scientific American, 90(4):65–66, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
65.pdf.

Rost:1904:PHI

[6628] E. C. Rost. The phonograph and how it figured in the Treaty of Jolo. Scientific American, 90(4):66, January 23, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n4/pdf/scientificamerican01231904-66.pdf.

Anonymous:1904:BH

[6629] Anonymous. A bottle house. Scientific American, 90(4):67, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n4/pdf/scientificamerican01231904-67.pdf.

Inkersley:1904:RLC

[6630] Arthur Inkersley. A real log cabin. Scientific American, 90(4):67, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n4/pdf/scientificamerican01231904-67a.pdf.

Anonymous:1904:BPWu

[6631] Anonymous. Business and personal wants. Scientific American, 90(4):68, January 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n4/pdf/scientificamerican01231904-
68a.pdf.

Anonymous:1904:BNBn 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
100a.pdf.

Anonymous:1904:IIv

[6633] Anonymous. Index of inventions. Scientific American, 90(5): 101–102, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
101.pdf.

Anonymous:1904:RPIv

[6634] Anonymous. Recently patented inventions. Scientific American, 90 (5):102–104, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
102.pdf.

Anonymous:1904:NQu

[6635] Anonymous. Notes and queries. Scientific American, 90(5):105– 109, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
105.pdf.

Anonymous:1904:MAF

[6636] Anonymous. Mushroom automobile firms. Scientific American, 90(5):74, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
74.pdf.

Anonymous:1904:RAR

[6637] Anonymous. The racing automobile and its relation to the development of the pleasure vehicle. Scientific American, 90(5):74, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n5/pdf/scientificamerican01301904-74a.pdf.

Anonymous:1904:FAN

[6638] Anonymous. The Fourth Annual New York Automobile Show. Scientific American, 90(5):74–75, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
74b.pdf.

Anonymous:1904:ICA

[6639] Anonymous. Improvements and changes in automobile construction as notes at the Paris Show. Scientific American, 90(5):75, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n5/pdf/scientificamerican01301904-75.pdf.

Anonymous:1904:IBA

[6640] Anonymous. An improved buckboard automobile. Scientific American, 90(5):76, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
76.pdf.

Anonymous:1904:ETPb

[6641] Anonymous. An electric tricycle for postal work. Scientific American, 90(5):76, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
76a.pdf.

Anonymous:1904:HPTa

[6642] Anonymous. The 24 horsepower Pope–Toledo touring car. Scientific American, 90(5):76–77, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
76b.pdf.

Anonymous:1904:SGT

[6643] Anonymous. The Stearns gasoline touring car. Scientific American, 90(5):77, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
77.pdf.

Anonymous:1904:STC

[6644] Anonymous. Some 1904 touring cars. Scientific American, 90 (5):79, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
79.pdf.

Anonymous:1904:TTC

[6645] Anonymous. The Thomas triple-cylinder touring car. Scientific American, 90(5):80–81, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
80.pdf.

Anonymous:1904:SDG

[6646] Anonymous. The Stevens–Duryea gasoline Stanhope. Scientific American, 90(5):81–82, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
81.pdf.

Anonymous:1904:KDC

[6647] Anonymous. The Knox double-cylinder cars with fan-cooled motors. Scientific American, 90(5):82–83, January 30, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n5/pdf/scientificamerican01301904-82.pdf.

Anonymous:1904:RAC

[6648] Anonymous. The Regas air-cooled automobile. Scientific American, 90(5):83, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
83.pdf.

Anonymous:1904:NFT

[6649] Anonymous. The new Franklin tonneau. Scientific American, 90(5):83, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
83a.pdf.

Anonymous:1904:MAC

[6650] Anonymous. The Mitchell air-cooled cars. Scientific American, 90(5):83, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
83b.pdf.

Anonymous:1904:PWMa

[6651] Anonymous. The prize-winning military tractor of the British War Office. Scientific American, 90(5):83, January 30, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n5/pdf/scientificamerican01301904-83c.pdf.

Anonymous:1904:PAS

[6652] Anonymous. The Paris Automobile Show. Scientific American, 90(5):84, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
84.pdf.

Anonymous:1904:IARb

[6653] Anonymous. An interesting automobile relio. Scientific American, 90(5):86, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
86.pdf.

Anonymous:1904:IAM

[6654] Anonymous. The Ivel agricultural motor. Scientific American, 90 (5):86–87, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
86a.pdf.

Anonymous:1904:CSLa

[6655] Anonymous. The Coldwell steam lawn mower. Scientific American, 90(5):87, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
87.pdf.

Anonymous:1904:NSRb

[6656] Anonymous. A new steam runabout. Scientific American, 90(5):87, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n5/pdf/scientificamerican01301904-87a.pdf.

Anonymous:1904:NAR

[6657] Anonymous. The new autocar runabout. Scientific American, 90(5):87, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
87b.pdf.

Anonymous:1904:NDT

[6658] Anonymous. A new detachable tire. Scientific American, 90(5):88, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n5/pdf/scientificamerican01301904-88.pdf.

Anonymous:1904:CPG

[6659] Anonymous. The Cadillac planetary gear transmission. Scientific American, 90(5):88, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
88a.pdf.

Anonymous:1904:CTA

[6660] Anonymous. A combination tool for automobiles. Scientific American, 90(5):88–90, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
88b.pdf.

Anonymous:1904:BLS

[6661] Anonymous. The brown lipe steering device. Scientific American, 90(5):90, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
90.pdf.

Anonymous:1904:FRB

[6662] Anonymous. The Ford record-breaking racer. Scientific American, 90(5):90, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
90a.pdf.

Anonymous:1904:GFP

[6663] Anonymous. Glass-front protector for Phaetons. Scientific American, 90(5):90, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
90b.pdf.

Anonymous:1904:NCTa

[6664] Anonymous. The new Columbia touring car. Scientific American, 90(5):92, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
92.pdf.

Anonymous:1904:WER

[6665] Anonymous. The Waverley Electric Runabout with Edison battery. Scientific American, 90(5):92, January 30, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n5/pdf/scientificamerican01301904-92a.pdf.

Anonymous:1904:GHC

[6666] Anonymous. A gasoline Hansom cab. Scientific American, 90 (5):92, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
92b.pdf.

Anonymous:1904:NLW

[6667] Anonymous. A new light-weight electric surrey. Scientific American, 90(5):92, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
92c.pdf.

Anonymous:1904:TBAb

[6668] Anonymous. A twin-balloon airship. Scientific American, 90(5):94, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n5/pdf/scientificamerican01301904-94.pdf.

Holder:1904:WS

[6669] Charles F. Holder. The windows of the sea. Scientific American, 90(5):94, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
94a.pdf.

Willey:1904:NTO

[6670] Day Allen Willey. New Texan oil deposits. Scientific American, 90(5):96–98, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
96.pdf.

Mclure:1904:STI

[6671] W. Frank M’clure. Starfish and their injuries. Scientific American, 90(5):98, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
98.pdf.

Anonymous:1904:BPWv

[6672] Anonymous. Business and personal wants. Scientific American, 90(5):100, January 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n5/pdf/scientificamerican01301904-
100.pdf.

Anonymous:1904:Cv

[6673] Anonymous. Correspondence. Scientific American, 90(6):119, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n6/pdf/scientificamerican02061904-119.pdf.

Russell:1904:HF

[6674] Henry Norris Russell. The heavens in February. Scientific American, 90(6):119, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
119a.pdf.

Anonymous:1904:RPIw

[6675] Anonymous. Recently patented inventions. Scientific American, 90(6):124, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
124.pdf.

Anonymous:1904:NQv

[6676] Anonymous. Notes and queries. Scientific American, 90(6):124, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n6/pdf/scientificamerican02061904-124b.pdf.

Anonymous:1904:IIw

[6677] Anonymous. Index of inventions. Scientific American, 90(6): 124–127, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
124c.pdf.

Anonymous:1904:PNPu

[6678] Anonymous. Patent notes: Patents. Scientific American, 90(6): 126, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
126.pdf.

Anonymous:1904:NCF

[6679] Anonymous. Near completion of the first Hudson River Tunnel. Scientific American, 90(6):114, February 6, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n6/pdf/scientificamerican02061904-114.pdf.

Anonymous:1904:CAS

[6680] Anonymous. Corrosion of architectural steel. Scientific American, 90(6):114, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
114a.pdf.

Anonymous:1904:ETS

[6681] Anonymous. Equable temperature in the subway. Scientific American, 90(6):114, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
114b.pdf.

Anonymous:1904:IFE

[6682] Anonymous. An Imperial Floating Exhibition. Scientific American, 90(6):114, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
114c.pdf.

Anonymous:1904:RSN

[6683] Anonymous. Relative strength of the Naval Powers. Scientific American, 90(6):114, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
114d.pdf.

Anonymous:1904:GCI

[6684] Anonymous. The grants of the Carnegie Institution. Scientific American, 90(6):114–115, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
114e.pdf.

Benjamin:1904:CDW

[6685] Marcus Benjamin. Carroll Davidson Wright. Scientific American, 90(6):115, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
115.pdf.

Anonymous:1904:PFS

[6686] Anonymous. Preventing frost on show windows in winter. Scientific American, 90(6):115, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
115a.pdf.

Anonymous:1904:DFE

[6687] Anonymous. Dispersion of fogs by electricity. Scientific American, 90(6):115, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
115b.pdf.

Anonymous:1904:IMW

[6688] Anonymous. An ingenious machine which greatly simplifies the process of printing. Scientific American, 90(6):116, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n6/pdf/scientificamerican02061904-116.pdf.

Anonymous:1904:PCR

[6689] Anonymous. Pullman cars in railroad accidents. Scientific American, 90(6):117, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
117.pdf.

Anonymous:1904:NSM

[6690] Anonymous. The new stage of the Metropolitan Opera House. Scientific American, 90(6):117–118, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
117a.pdf.

Collins:1904:MAF

[6691] A. Frederick Collins. Making asbestos fire curtains and fireproof fabrics. Scientific American, 90(6):120, February 6, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n6/pdf/scientificamerican02061904-120.pdf.

Beasley:1904:AJ

[6692] Walter L. Beasley. The Aborigines of Japan. Scientific American, 90(6):121–122, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
121.pdf.

Anonymous:1904:CDC

[6693] Anonymous. Cooking device for camping purposes. Scientific American, 90(6):123, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
123.pdf.

Anonymous:1904:ATG

[6694] Anonymous. Automatic track gage. Scientific American, 90(6): 123, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
123a.pdf.

Anonymous:1904:SAL

[6695] Anonymous. Spur attachment for leggings. Scientific American, 90(6):123, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
123b.pdf.

Anonymous:1904:HCH

[6696] Anonymous. A handy coin holder. Scientific American, 90(6): 123, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
123c.pdf.

Anonymous:1904:LJ

[6697] Anonymous. Lifting jack. Scientific American, 90(6):123, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n6/pdf/scientificamerican02061904-123d.pdf.

Anonymous:1904:BPWw

[6698] Anonymous. Business and personal wants. Scientific American, 90(6):124, February 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n6/pdf/scientificamerican02061904-
124a.pdf.

Anonymous:1904:LNg

[6699] Anonymous. Legal notes. Scientific American, 90(7):140, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n7/pdf/scientificamerican02131904-140.pdf.

Anonymous:1904:RPIx

[6700] Anonymous. Recently patented inventions. Scientific American, 90 (7):141–142, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
141.pdf.

Anonymous:1904:NQw

[6701] Anonymous. Notes and queries. Scientific American, 90(7):142– 143, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
142a.pdf.

Anonymous:1904:BNBo

[6702] Anonymous. Books: New books, etc. Scientific American, 90 (7):143, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
143.pdf.

Anonymous:1904:IIx

[6703] Anonymous. Index of inventions. Scientific American, 90(7): 143–147, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
143a.pdf.

Anonymous:1904:PNPv

[6704] Anonymous. Patent notes: Patents. Scientific American, 90 (7):146, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
146.pdf.

Anonymous:1904:OSN

[6705] Anonymous. Our system of Naval comparison. Scientific American, 90(7):130, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
130.pdf.

Anonymous:1904:VCP

[6706] Anonymous. Value of correct propeller design. Scientific American, 90(7):130, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
130a.pdf.

Anonymous:1904:PAPa

[6707] Anonymous. Precautions against poisoning the queen bee. Scientific American, 90(7):130, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
130b.pdf.

Anonymous:1904:ET

[6708] Anonymous. The Eiffel Tower. Scientific American, 90(7):130– 131, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
130c.pdf.

Anonymous:1904:FSJ

[6709] Anonymous. The fuel supply of the Japanese Navy. Scientific American, 90(7):131, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
131.pdf.

Anonymous:1904:PSR

[6710] Anonymous. Progress in the study of radium and radioactivity. Scientific American, 90(7):131, February 13, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n7/pdf/scientificamerican02131904-131a.pdf.

Anonymous:1904:CSc

[6711] Anonymous. The current supplement. Scientific American, 90 (7):131, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
131b.pdf.

Anonymous:1904:HS

[6712] Anonymous. The hydroscope and its success. Scientific American, 90(7):131, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
131c.pdf.

Anonymous:1904:SLE

[6713] Anonymous. Safety locomotive for explosives factory. Scientific American, 90(7):132, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
132.pdf.

Jones:1904:BSD

[6714] George J. Jones. Boiler scale detection. Scientific American, 90 (7):132, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
132a.pdf.

Anonymous:1904:PPB

[6715] Anonymous. A practical primary battery. Scientific American, 90(7):132, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
132b.pdf.

Anonymous:1904:JN

[6716] Anonymous. The Japanese Navy. Scientific American, 90(7): 134–135, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
134.pdf.

Anonymous:1904:RFA

[6717] Anonymous. Results of the Florida Automobile Race Meet. Scientific American, 90(7):136, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
136.pdf.

Blackiston:1904:GEF

[6718] G. P. Blackiston. A great engineering feat. Scientific American, 90(7):138, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
138.pdf.

Holder:1904:NM

[6719] Charles F. Holder. Natural monuments. Scientific American, 90 (7):139, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
139.pdf.

Anonymous:1904:BG

[6720] Anonymous. The blue gnu. Scientific American, 90(7):139, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n7/pdf/scientificamerican02131904-139a.pdf.

Anonymous:1904:BPWx

[6721] Anonymous. Business and personal wants. Scientific American, 90(7):142, February 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n7/pdf/scientificamerican02131904-
142.pdf.

Anonymous:1904:Cw

[6722] Anonymous. Correspondence. Scientific American, 90(8):160, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n8/pdf/scientificamerican02201904-160.pdf.

Anonymous:1904:RPIy 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
163.pdf.

Anonymous:1904:NQx

[6724] Anonymous. Notes and queries. Scientific American, 90(8):164, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n8/pdf/scientificamerican02201904-164.pdf.

Anonymous:1904:BNBp

[6725] Anonymous. Books: New books, etc. Scientific American, 90 (8):164, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
164a.pdf.

Anonymous:1904:IIy

[6726] Anonymous. Index of inventions. Scientific American, 90(8): 164–167, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
164b.pdf.

Anonymous:1904:PNPw

[6727] Anonymous. Patent notes: Patents. Scientific American, 90 (8):168, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
168.pdf.

Anonymous:1904:MBG

[6728] Anonymous. The man behind the gun. Scientific American, 90 (8):150, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
150.pdf.

Anonymous:1904:WI

[6729] Anonymous. Winter insects. Scientific American, 90(8):150, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n8/pdf/scientificamerican02201904-150a.pdf.

Anonymous:1904:FPF

[6730] Anonymous. Forcing plants by fire. Scientific American, 90(8): 150, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
150b.pdf.

Anonymous:1904:NKR

[6731] Anonymous. New kind of rays emitted from the brain and nerve centers. Scientific American, 90(8):151, February 20, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n8/pdf/scientificamerican02201904-151.pdf.

Anonymous:1904:ESC

[6732] Anonymous. Extensive submarine construction programme for Great Britain. Scientific American, 90(8):151, February 20, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v90/
n8/pdf/scientificamerican02201904-151a.pdf.

Anonymous:1904:SNn

[6733] Anonymous. Science notes. Scientific American, 90(8):151, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n8/pdf/scientificamerican02201904-151b.pdf.

Anonymous:1904:MBS

[6734] Anonymous. Magnetic brake for small electric motors. Scientific American, 90(8):152, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
152.pdf.

Guarini:1904:PSS

[6735] Emile Guarini. Photography of star shadows. Scientific American, 90(8):152–154, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
152a.pdf.

Anonymous:1904:GBF

[6736] Anonymous. The Great Baltimore Fire. Scientific American, 90 (8):154–155, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
154.pdf.

Anonymous:1904:RN

[6737] Anonymous. The Russian Navy. Scientific American, 90(8): 157–159, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
157.pdf.

Anonymous:1904:TNM

[6738] Anonymous. Two new motor bicycles. Scientific American, 90 (8):159, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
159.pdf.

Anonymous:1904:SMC

[6739] Anonymous. Somersault monocycle course. Scientific American, 90 (8):159–160, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
159a.pdf.

Anonymous:1904:CPR

[6740] Anonymous. Combined pack and riding saddle. Scientific American, 90(8):161, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
161.pdf.

Anonymous:1904:LA

[6741] Anonymous. Land anchor. Scientific American, 90(8):161, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n8/pdf/scientificamerican02201904-161a.pdf.

Anonymous:1904:SDA

[6742] Anonymous. Spring draft attachment. Scientific American, 90 (8):161, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
161b.pdf.

Anonymous:1904:NTRb

[6743] Anonymous. A novel third-rail protector. Scientific American, 90(8):161, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
161c.pdf.

Anonymous:1904:DTW

[6744] Anonymous. The disk type of water meter. Scientific American, 90 (8):161–162, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
161d.pdf.

Anonymous:1904:OIc

[6745] Anonymous. Oddities in inventions. Scientific American, 90(8): 162, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
162.pdf.

Anonymous:1904:BPWy

[6746] Anonymous. Business and personal wants. Scientific American, 90(8):163, February 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n8/pdf/scientificamerican02201904-
163a.pdf.

Anonymous:1904:Cx

[6747] Anonymous. Correspondence. Scientific American, 90(9):175, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n9/pdf/scientificamerican02271904-175.pdf.

Anonymous:1904:LNh

[6748] Anonymous. Legal notes. Scientific American, 90(9):182, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n9/pdf/scientificamerican02271904-182.pdf.

Anonymous:1904:RPIz

[6749] Anonymous. Recently patented inventions. Scientific American, 90(9):183, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
183.pdf.

Anonymous:1904:NQy

[6750] Anonymous. Notes and queries. Scientific American, 90(9):183– 184, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
183b.pdf.

Anonymous:1904:IIz

[6751] Anonymous. Index of inventions. Scientific American, 90(9): 184–187, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
184.pdf.

Anonymous:1904:PNPx

[6752] Anonymous. Patent notes: Patents. Scientific American, 90 (9):187, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
187.pdf.

Anonymous:1904:SWF

[6753] Anonymous. Salt water fire protection for New York City. Scientific American, 90(9):170, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
170.pdf.

Anonymous:1904:PNA

[6754] Anonymous. The proposed Naval Appropriation. Scientific American, 90(9):170, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
170a.pdf.

Anonymous:1904:EM

[6755] Anonymous. Electricity in mines. Scientific American, 90(9): 170, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
170b.pdf.

Anonymous:1904:ERH

[6756] Anonymous. Emission of N-rays by the human body, especially the muscles and nerves. Scientific American, 90(9):170–171, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n9/pdf/scientificamerican02271904-170c.pdf.

Hurd:1904:FFEb

[6757] Archibald S. Hurd. Fleets in the Far East — an English review of the position of Russia and Japan. Scientific American, 90(9):171, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n9/pdf/scientificamerican02271904-171.pdf.

Anonymous:1904:SNo

[6758] Anonymous. Science notes. Scientific American, 90(9):171, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n9/pdf/scientificamerican02271904-171a.pdf.

Anonymous:1904:SST

[6759] Anonymous. Submarine sound telegraphy. Scientific American, 90(9):172, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
172.pdf.

Anonymous:1904:SHP

[6760] Anonymous. Safety hook provided with counting mechanism. Scientific American, 90(9):173, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
173.pdf.

Anonymous:1904:PCA

[6761] Anonymous. A powerful compressed air clamp. Scientific American, 90(9):173, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
173a.pdf.

Anonymous:1904:NAF

[6762] Anonymous. Novel appliances for fighting fire. Scientific American, 90(9):173–174, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
173b.pdf.

Anonymous:1904:MFM

[6763] Anonymous. Modern flour milling. Scientific American, 90(9): 176–178, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
176.pdf.

Anonymous:1904:DRC

[6764] Anonymous. The disabled Russian cruiser “Novik”. Scientific American, 90(9):178, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
178.pdf.

Aaron:1904:POP

[6765] S. Frank Aaron. The parasite of the oak pruner. Scientific American, 90(9):179, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
179.pdf.

Field:1904:EA

[6766] C. Field. Early armorclads. Scientific American, 90(9):180–181, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v90/n9/pdf/scientificamerican02271904-180.pdf.

Anonymous:1904:BPWz

[6767] Anonymous. Business and personal wants. Scientific American, 90(9):183, February 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v90/n9/pdf/scientificamerican02271904-
183a.pdf.

Anonymous:1904:Cy

[6768] Anonymous. Correspondence. Scientific American, 91(1):9, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-9.pdf.

Anonymous:1904:RPIba

[6769] Anonymous. Recently patented inventions. Scientific American, 91(1):14–15, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n1/pdf/scientificamerican07021904-
14.pdf.

Anonymous:1904:NQz

[6770] Anonymous. Notes and queries. Scientific American, 91(1):15, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-15a.pdf.

Anonymous:1904:BNBq

[6771] Anonymous. Books: New books, etc. Scientific American, 91(1):15–16, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-15b.pdf.

Anonymous:1904:IIba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-16.pdf.

Anonymous:1904:PNPy

[6773] Anonymous. Patent notes: Patents. Scientific American, 91(1):19, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-19.pdf.

Anonymous:1904:SS

[6774] Anonymous. Superheated steam. Scientific American, 91(1):4, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-4.pdf.

Anonymous:1904:SRU

[6775] Anonymous. State registration of United States trade marks. Scientific American, 91(1):4, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n1/pdf/scientificamerican07021904-
4a.pdf.

Anonymous:1904:SCV

[6776] Anonymous. On the state of the carbon vaporized in glow lamps. Scientific American, 91(1):4–5, July 2, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n1/pdf/scientificamerican07021904-4b.pdf.

Russell:1904:HJc

[6777] Henry Norris Russell. The heavens in July. Scientific American, 91(1): 5, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-5.pdf.

Anonymous:1904:PRS

[6778] Anonymous. The purchase of a reservoir site. Scientific American, 91(1):5, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n1/pdf/scientificamerican07021904-
5a.pdf.

Anonymous:1904:SNp

[6779] Anonymous. Science notes. Scientific American, 91(1):5, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-5b.pdf.

Anonymous:1904:RTS

[6780] Anonymous. The Rapid Transit Subway Tunnel under the Harlem River. Scientific American, 91(1):6, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n1/pdf/scientificamerican07021904-
6.pdf.

Anonymous:1904:FIA

[6781] Anonymous. The fifth international automobile cup race. Scientific American, 91(1):7, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n1/pdf/scientificamerican07021904-
7.pdf.

Anonymous:1904:TFS

[6782] Anonymous. Test of fireproof stage properties in London. Scientific American, 91(1):7–8, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n1/pdf/scientificamerican07021904-
7a.pdf.

Fitzgerald:1904:SUU

[6783] W. G. Fitzgerald. Some unusual uses for fire-engines. Scientific American, 91(1):10, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n1/pdf/scientificamerican07021904-
10.pdf.

Anonymous:1904:JNG

[6784] Anonymous. Japanese Naval guns. Scientific American, 91(1):11–12, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-11.pdf.

Collins:1904:NLS

[6785] A. Frederick Collins. A new life-saving gun. Scientific American, 91(1): 12, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-12.pdf.

Weeks:1904:UIN

[6786] Arthur B. Weeks. A useful instrument in the Niagara Falls Power House. Scientific American, 91(1):13, July 2, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n1/pdf/scientificamerican07021904-13.pdf.

Anonymous:1904:IAC

[6787] Anonymous. Insulating apparatus for canned goods to prevent ptomaine poisoning. Scientific American, 91(1):13, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n1/pdf/scientificamerican07021904-13a.pdf.

Anonymous:1904:BPWba

[6788] Anonymous. Business and personal wants. Scientific American, 91(1): 15, July 2, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n1/pdf/scientificamerican07021904-15.pdf.

Russell:1904:HS

[6789] Henry Norris Russell. The heavens in September. Scientific American, 91(10):155, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
155.pdf.

Anonymous:1904:Cz

[6790] Anonymous. Correspondence. Scientific American, 91(10):158, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n10/pdf/scientificamerican09031904-158a.pdf.

Anonymous:1904:RPIbb

[6791] Anonymous. Recently patented inventions. Scientific American, 91 (10):164–165, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
164.pdf.

Anonymous:1904:BPWbb

[6792] Anonymous. Business and personal wants. Scientific American, 91(10):165, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
165.pdf.

Anonymous:1904:NQba

[6793] Anonymous. Notes and queries. Scientific American, 91(10): 165, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
165a.pdf.

Anonymous:1904:BNBr

[6794] Anonymous. Books: New books, etc. Scientific American, 91 (10):168, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
168.pdf.

Anonymous:1904:IIbb

[6795] Anonymous. Index of inventions. Scientific American, 91(10): 168–171, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
168a.pdf.

Anonymous:1904:P

[6796] Anonymous. Patents. Scientific American, 91(10):171, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n10/pdf/scientificamerican09031904-171.pdf.

Anonymous:1904:NCL

[6797] Anonymous. A novel canal lift at Foxton. Scientific American, 91(10):153, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
153.pdf.

Anonymous:1904:MBF

[6798] Anonymous. The Manhattan Bridge fiasco. Scientific American, 91(10):154, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
154.pdf.

Anonymous:1904:PS

[6799] Anonymous. Perils of the submarine. Scientific American, 91 (10):154, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
154a.pdf.

Anonymous:1904:EPOb

[6800] Anonymous. Electric power for oil wells. Scientific American, 91 (10):154, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
154b.pdf.

Anonymous:1904:PIO

[6801] Anonymous. The production of iron ores in 1903. Scientific American, 91(10):154–155, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
154c.pdf.

Anonymous:1904:SNq

[6802] Anonymous. Science notes. Scientific American, 91(10):155, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n10/pdf/scientificamerican09031904-155a.pdf.

Anonymous:1904:LACa

[6803] Anonymous. Launch of the armored cruiser “South Dakota”. Scientific American, 91(10):156, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
156.pdf.

Anonymous:1904:NMM

[6804] Anonymous. A novel method of making relief maps. Scientific American, 91(10):157, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
157.pdf.

Anonymous:1904:SRA

[6805] Anonymous. A strange railway accident. Scientific American, 91 (10):157–158, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
157a.pdf.

Anonymous:1904:DPN

[6806] Anonymous. Direct photography in natural colors on paper. Scientific American, 91(10):158, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
158.pdf.

Anonymous:1904:MLSa

[6807] Anonymous. The “Minnesota” the largest steamship ever built in America. Scientific American, 91(10):160, September 3, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n10/pdf/scientificamerican09031904-160.pdf.

Anonymous:1904:EET

[6808] Anonymous. Express engine and turntable in the transportation building. Scientific American, 91(10):160–161, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n10/pdf/scientificamerican09031904-160a.pdf.

Anonymous:1904:FHC

[6809] Anonymous. Festival Hall and the Colonnade of States — the architectural masterpiece of the Louisiana Purchase Exposition. Scientific American, 91(10):161, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
161.pdf.

Collins:1904:DFE

[6810] A. Frederick Collins. Dispelling fog by electricity — an experimental proof. Scientific American, 91(10):162–163, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n10/pdf/scientificamerican09031904-162.pdf.

Anonymous:1904:QLT

[6811] Anonymous. Quoin for locking type forms. Scientific American, 91(10):163, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
163.pdf.

Anonymous:1904:IMB

[6812] Anonymous. The International Motor-Boat Contest “Napier II”. Scientific American, 91(10):163, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
163a.pdf.

Anonymous:1904:IVS

[6813] Anonymous. Improved variable speed and reversing gear. Scientific American, 91(10):163, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
163b.pdf.

Anonymous:1904:GTB

[6814] Anonymous. Georgia Till Basket and Paducah Factory. Scientific American, 91(10):166–167, September 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n10/pdf/scientificamerican09031904-
166.pdf.

Anonymous:1904:Cba

[6815] Anonymous. Correspondence. Scientific American, 91(11):179, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n11/pdf/scientificamerican09101904-179.pdf.

Anonymous:1904:RPIbc

[6816] Anonymous. Recently patented inventions. Scientific American, 91 (11):184, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
184.pdf.

Anonymous:1904:NQbb

[6817] Anonymous. Notes and queries. Scientific American, 91(11): 184–185, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
184b.pdf.

Anonymous:1904:BNBs

[6818] Anonymous. Books: New books, etc. Scientific American, 91 (11):185, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
185.pdf.

Anonymous:1904:IIbc

[6819] Anonymous. Index of inventions. Scientific American, 91(11): 185–187, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
185a.pdf.

Anonymous:1904:PNPz

[6820] Anonymous. Patent notes: Patents. Scientific American, 91 (11):187, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
187.pdf.

Anonymous:1904:ACP

[6821] Anonymous. The art commission and public structures. Scientific American, 91(11):174, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
174.pdf.

Anonymous:1904:AIB

[6822] Anonymous. The agricultural importance of bacteria. Scientific American, 91(11):174, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
174a.pdf.

Anonymous:1904:PPF

[6823] Anonymous. Possibilities of peat as fuel. Scientific American, 91 (11):174–175, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
174b.pdf.

Bickerton:1904:SA

[6824] A. W. Bickerton. The sorting of atoms. Scientific American, 91 (11):175, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
175.pdf.

Anonymous:1904:ANa

[6825] Anonymous. Automobile notes. Scientific American, 91(11): 175, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
175a.pdf.

Anonymous:1904:TAV

[6826] Anonymous. Tilting attachment for vehicle tops. Scientific American, 91(11):176, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
176.pdf.

Anonymous:1904:UCG

[6827] Anonymous. The unarmored composite gunboat “Dubuque”. Scientific American, 91(11):176, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
176a.pdf.

Anonymous:1904:PPc

[6828] Anonymous. Pendulum power. Scientific American, 91(11): 176, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
176b.pdf.

Perkins:1904:BTD

[6829] Frank C. Perkins. The buckeye traction ditcher. Scientific American, 91(11):177–178, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
177.pdf.

Anonymous:1904:CNY

[6830] Anonymous. The completion of the New York Subway. Scientific American, 91(11):178, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
178.pdf.

Anonymous:1904:NCTb

[6831] Anonymous. The new Cunard turbine steamers. Scientific American, 91(11):179, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
179a.pdf.

Anonymous:1904:GLS

[6832] Anonymous. Government life-saving station and the Ferris Wheel. Scientific American, 91(11):180–181, September 10, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n11/pdf/scientificamerican09101904-180.pdf.

Sullivan:1904:SSM

[6833] M. C. Sullivan. Signal service in modern warfare. Scientific American, 91(11):182–183, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
182.pdf.

Anonymous:1904:NGMb

[6834] Anonymous. A novel gliding machine. Scientific American, 91 (11):183, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
183.pdf.

Baltimore:1904:PBF

[6835] J. Mayne Baltimore. Prof. Botts’s flying machine. Scientific American, 91(11):183, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
183a.pdf.

Anonymous:1904:BPWbc

[6836] Anonymous. Business and personal wants. Scientific American, 91 (11):184, September 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n11/pdf/scientificamerican09101904-
184a.pdf.

Anonymous:1904:LNi

[6837] Anonymous. Legal notes. Scientific American, 91(12):202, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n12/pdf/scientificamerican09171904-202.pdf.

Anonymous:1904:RPIbd

[6838] Anonymous. Recently patented inventions. Scientific American, 91 (12):203, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
203.pdf.

Anonymous:1904:NQbc

[6839] Anonymous. Notes and queries. Scientific American, 91(12): 203–204, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
203b.pdf.

Anonymous:1904:BNBt

[6840] Anonymous. Books: New books, etc. Scientific American, 91 (12):204, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
204.pdf.

Anonymous:1904:IIbd

[6841] Anonymous. Index of inventions. Scientific American, 91(12): 204–206, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
204a.pdf.

Anonymous:1904:PNPba

[6842] Anonymous. Patent notes: Patents. Scientific American, 91 (12):206, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
206.pdf.

Anonymous:1904:UDG

[6843] Anonymous. The ultimate development of the gas engine. Scientific American, 91(12):190, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
190.pdf.

Anonymous:1904:EWM

[6844] Anonymous. Endurance of warship machinery in the Russo-Japanese War. Scientific American, 91(12):190, September 17, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n12/pdf/scientificamerican09171904-190a.pdf.

Anonymous:1904:PC

[6845] Anonymous. Production of coal in 1903. Scientific American, 91 (12):190, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
190b.pdf.

Anonymous:1904:NYM

[6846] Anonymous. New York Meeting of the Society of Chemical Industry. Scientific American, 91(12):191, September 17, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n12/pdf/scientificamerican09171904-191.pdf.

Anonymous:1904:LFT

[6847] Anonymous. Launch of the first turbine-propelled liner for transatlantic traffic. Scientific American, 91(12):191, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n12/pdf/scientificamerican09171904-191a.pdf.

Anonymous:1904:TDG

[6848] Anonymous. A test of the Doenvig globe. Scientific American, 91(12):191, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
191b.pdf.

Anonymous:1904:CSd

[6849] Anonymous. The current supplement. Scientific American, 91 (12):191, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
191c.pdf.

Anonymous:1904:JAA

[6850] Anonymous. The Jungner alkaline accumulator. Scientific American, 91(12):192, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
192.pdf.

Anonymous:1904:RTA

[6851] Anonymous. A record transcontinental automobile trip. Scientific American, 91(12):192, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
192a.pdf.

Bennett:1904:THM

[6852] Herbert I. Bennett. A trolley for heavy multiple-unit electric trains. Scientific American, 91(12):193, September 17, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n12/pdf/scientificamerican09171904-193.pdf.

Guarini:1904:MTT

[6853] Emile Guarini. The Malcotti telecryptograph for telegraphing upon telephone lines. Scientific American, 91(12):193–194, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n12/pdf/scientificamerican09171904-193a.pdf.

Hayward:1904:CCC

[6854] C. B. Hayward. Coffee and coffee culture. Scientific American, 91 (12):194–195, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
194.pdf.

Anonymous:1904:MLSb

[6855] Anonymous. Modern locomotives at the St. Louis Exposition. Scientific American, 91(12):195–196, September 17, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n12/pdf/scientificamerican09171904-195.pdf.

Wallach:1904:DCH

[6856] Isabel R. Wallach. The dino collection of historic armor. Scientific American, 91(12):196–199, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
196.pdf.

Anonymous:1904:RP

[6857] Anonymous. Ruling pen. Scientific American, 91(12):200, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n12/pdf/scientificamerican09171904-200.pdf.

Anonymous:1904:WRAb

[6858] Anonymous. Wire-reeling apparatus. Scientific American, 91 (12):200, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
200a.pdf.

Anonymous:1904:PDh

[6859] Anonymous. Pump diaphragm. Scientific American, 91(12): 200, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
200b.pdf.

Anonymous:1904:OId

[6860] Anonymous. Oddities in invention. Scientific American, 91(12): 200–201, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
200c.pdf.

Anonymous:1904:BPWbd

[6861] Anonymous. Business and personal wants. Scientific American, 91 (12):203, September 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n12/pdf/scientificamerican09171904-
203a.pdf.

Anonymous:1904:RPIbe

[6862] Anonymous. Recently patented inventions. Scientific American, 91 (13):220, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
220.pdf.

Anonymous:1904:NQbd

[6863] Anonymous. Notes and queries. Scientific American, 91(13): 220, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
220b.pdf.

Anonymous:1904:BNBu

[6864] Anonymous. Books: New books, etc. Scientific American, 91 (13):220, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
220c.pdf.

Anonymous:1904:IIbe

[6865] Anonymous. Index of inventions. Scientific American, 91(13): 220–222, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
220d.pdf.

Anonymous:1904:PNPbb

[6866] Anonymous. Patent notes: Patents. Scientific American, 91 (13):222, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
222.pdf.

Anonymous:1904:FHB

[6867] Anonymous. The forlorn hope of the Baltic Fleet. Scientific American, 91(13):210, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
210.pdf.

Anonymous:1904:GR

[6868] Anonymous. A grave responsibility. Scientific American, 91(13): 210, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
210a.pdf.

Anonymous:1904:PAPb

[6869] Anonymous. Proposed amendment permitting the extension of the terms of patents. Scientific American, 91(13):210, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n13/pdf/scientificamerican09241904-210b.pdf.

Anonymous:1904:TLT

[6870] Anonymous. Trunk lines and trolleys in the East. Scientific American, 91(13):210–211, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
210c.pdf.

Anonymous:1904:NYSb

[6871] Anonymous. The New York Sessions of the Eighth International Geographic Congress. Scientific American, 91(13):211, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n13/pdf/scientificamerican09241904-211.pdf.

Anonymous:1904:SNs

[6872] Anonymous. Science notes. Scientific American, 91(13):211, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n13/pdf/scientificamerican09241904-211a.pdf.

MClure:1904:MWP

[6873] W. Frank M’Clure. The manufacture of wall paper. Scientific American, 91(13):212, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
212.pdf.

Anonymous:1904:RH

[6874] Anonymous. A reasoning horse. Scientific American, 91(13): 213, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
213.pdf.

Rigge:1904:WWP

[6875] William F. Rigge. When was the photograph taken? Scientific American, 91(13):213–214, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
213a.pdf.

Anonymous:1904:CNC

[6876] Anonymous. Completing the new Croton Dam. Scientific American, 91(13):214–215, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
214.pdf.

Guarrini:1904:LI

[6877] Emile Guarrini. The “Lebaudy II”. Scientific American, 91(13): 216–218, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
216.pdf.

Anonymous:1904:NDS

[6878] Anonymous. Native dwellings at the St. Louis Exposition. Scientific American, 91(13):218, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
218.pdf.

Fulda:1904:SEI

[6879] O. Fulda. Striking examples of insect mimicry. Scientific American, 91(13):219, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
219.pdf.

Anonymous:1904:BPWbe

[6880] Anonymous. Business and personal wants. Scientific American, 91 (13):220, September 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n13/pdf/scientificamerican09241904-
220a.pdf.

Russell:1904:HO

[6881] Henry Norris Russell. The heavens in October. Scientific American, 91(14):227, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
227.pdf.

Anonymous:1904:RPIbf

[6882] Anonymous. Recently patented inventions. Scientific American, 91(14):235, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
235a.pdf.

Anonymous:1904:IIbf

[6883] Anonymous. Index of inventions. Scientific American, 91(14):237–239, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n14/pdf/scientificamerican10011904-237a.pdf.

Anonymous:1904:WCG

[6884] Anonymous. Warship construction in Government yards. Scientific American, 91(14):226, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
226.pdf.

Anonymous:1904:CCM

[6885] Anonymous. The challenge cup motor-boat contest. Scientific American, 91(14):226, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
226a.pdf.

Anonymous:1904:WVS

[6886] Anonymous. Water versus steam power. Scientific American, 91 (14):226–227, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
226b.pdf.

Anonymous:1904:SNt

[6887] Anonymous. Science notes. Scientific American, 91(14):227, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n14/pdf/scientificamerican10011904-227a.pdf.

Anonymous:1904:NAPb

[6888] Anonymous. A novel automatic pump for inflating automobile tires. Scientific American, 91(14):228, October 1, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n14/pdf/scientificamerican10011904-228.pdf.

Anonymous:1904:NSW

[6889] Anonymous. Novel spare wheel and tire device for automobiles. Scientific American, 91(14):228, October 1, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n14/pdf/scientificamerican10011904-228a.pdf.

Anonymous:1904:MAM

[6890] Anonymous. A machine for automatically making incandescent gas mantles. Scientific American, 91(14):228–229, October 1, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n14/pdf/scientificamerican10011904-228b.pdf.

Anonymous:1904:DBP

[6891] Anonymous. Domestic bath plumbing. Scientific American, 91 (14):229–230, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
229.pdf.

Guarini:1904:NRA

[6892] Emile Guarini. A novel respiratory apparatus. Scientific American, 91(14):230–231, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
230.pdf.

Anonymous:1904:LCM

[6893] Anonymous. Launching of the “Connecticut” and “Milwaukee”. Scientific American, 91(14):231, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
231.pdf.

Walker:1904:LHW

[6894] Herbert T. Walker. Locomotive history at the World’s Fair, St. Louis. Scientific American, 91(14):232–233, October 1, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n14/pdf/scientificamerican10011904-232.pdf.

Beasley:1904:APL

[6895] Walter L. Beasley. Artificial propagation of lobsters. Scientific American, 91(14):233–234, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
233.pdf.

Anonymous:1904:AB

[6896] Anonymous. The auto-bolide. Scientific American, 91(14):234, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n14/pdf/scientificamerican10011904-234.pdf.

Anonymous:1904:SEP

[6897] Anonymous. Spelterini’s Egyptian pictures. Scientific American, 91(14):235, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
235.pdf.

Anonymous:1904:BPWbf

[6898] Anonymous. Business and personal wants. Scientific American, 91(14):237, October 1, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n14/pdf/scientificamerican10011904-
237.pdf.

Anonymous:1904:Cbb

[6899] Anonymous. Correspondence. Scientific American, 91(15):247, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-247.pdf.

Anonymous:1904:RPIbg

[6900] Anonymous. Recently patented inventions. Scientific American, 91(15):252, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
252.pdf.

Anonymous:1904:NQbe

[6901] Anonymous. Notes and queries. Scientific American, 91(15):252, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-252b.pdf.

Anonymous:1904:BNBv

[6902] Anonymous. Books: New books, etc. Scientific American, 91(15):252, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-252c.pdf.

Anonymous:1904:IIbg

[6903] Anonymous. Index of inventions. Scientific American, 91(15):253–255, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-253.pdf.

Anonymous:1904:PNPbc

[6904] Anonymous. Patent notes: Patents. Scientific American, 91(15):255, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-255.pdf.

Anonymous:1904:RSAb

[6905] Anonymous. That Ramapo scheme again. Scientific American, 91(15):242, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
242.pdf.

Anonymous:1904:LWC

[6906] Anonymous. The launching weight of the “Connecticut”. Scientific American, 91(15):242, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
242a.pdf.

Anonymous:1904:MPCb

[6907] Anonymous. The menace of the Pullman car. Scientific American, 91(15):242, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
242b.pdf.

Anonymous:1904:CFa

[6908] Anonymous. Curious foods. Scientific American, 91(15):242–243, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-242c.pdf.

Ingersoll:1904:OSD

[6909] Ernest Ingersoll. Oddities in self-defense. Scientific American, 91(15):243, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
243.pdf.

Anonymous:1904:ANb

[6910] Anonymous. Automobile notes. Scientific American, 91(15):243, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-243a.pdf.

Guarini:1904:CB

[6911] Emile Guarini. The camera book. Scientific American, 91(15):244, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-244.pdf.

Anonymous:1904:HPE

[6912] Anonymous. Hydraulic press for embossing silverware at the St. Louis Fair. Scientific American, 91(15):244, October 8, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n15/pdf/scientificamerican10081904-244a.pdf.

Anonymous:1904:SNTa

[6913] Anonymous. Some new types of lifeboats. Scientific American, 91 (15):245–246, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
245.pdf.

Anonymous:1904:MCT

[6914] Anonymous. Metal cross tie for railroads. Scientific American, 91(15):246, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
246.pdf.

Anonymous:1904:MPR

[6915] Anonymous. Model of the Pennsylvania Railroad Hudson River Tunnel Station at New York. Scientific American, 91(15):246, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n15/pdf/scientificamerican10081904-246a.pdf.

Wallach:1904:AWD

[6916] Isabel R. Wallach. Armor and weapons of the De Dino Collection — II. Scientific American, 91(15):248, October 8, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n15/pdf/scientificamerican10081904-248.pdf.

Anonymous:1904:PNS

[6917] Anonymous. Peary’s new ship for work in Arctic seas. Scientific American, 91(15):249–250, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
249.pdf.

Anonymous:1904:PWMb

[6918] Anonymous. Perfect working model of a Winton touring car. Scientific American, 91(15):251, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
251.pdf.

Beasley:1904:LNT

[6919] Walter L. Beasley. The largest narwhal tusk. Scientific American, 91(15):251, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
251a.pdf.

Anonymous:1904:BPWbg

[6920] Anonymous. Business and personal wants. Scientific American, 91(15):252, October 8, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n15/pdf/scientificamerican10081904-
252a.pdf.

Anonymous:1904:Cbc

[6921] Anonymous. Correspondence. Scientific American, 91(16):263, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n16/pdf/scientificamerican10151904-263.pdf.

Anonymous:1904:RPIbh

[6922] Anonymous. Recently patented inventions. Scientific American, 91(16):270, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
270.pdf.

Anonymous:1904:NQbf

[6923] Anonymous. Notes and queries. Scientific American, 91(16):271, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n16/pdf/scientificamerican10151904-271a.pdf.

Anonymous:1904:BNBw

[6924] Anonymous. Books: New books, etc. Scientific American, 91 (16):271, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
271b.pdf.

Anonymous:1904:IIbh

[6925] Anonymous. Index of inventions. Scientific American, 91(16): 271–273, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
271c.pdf.

Anonymous:1904:PNPbd 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
274.pdf.

Anonymous:1904:EFT

[6927] Anonymous. Elevator fatalities and their prevention. Scientific American, 91(16):258, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
258.pdf.

Anonymous:1904:BGF

[6928] Anonymous. Big guns for future warships. Scientific American, 91(16):258, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
258a.pdf.

Evans:1904:EM

[6929] T. H. Evans. Experiments on the mosquito. Scientific American, 91(16):258, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
258b.pdf.

Anonymous:1904:ANBa

[6930] Anonymous. Arsenic needed for the body and found in different foods. Scientific American, 91(16):258–259, October 15, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n16/pdf/scientificamerican10151904-258c.pdf.

Walsh:1904:ASS

[6931] George E. Walsh. Automobile steel specialties. Scientific American, 91(16):259, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
259.pdf.

Anonymous:1904:MBE

[6932] Anonymous. Municipal bakery experiments in Sicily. Scientific American, 91(16):259, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
259a.pdf.

Anonymous:1904:ANc

[6933] Anonymous. Automobile notes. Scientific American, 91(16):259, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n16/pdf/scientificamerican10151904-259b.pdf.

Anonymous:1904:T

[6934] Anonymous. The taxometer. Scientific American, 91(16):260, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n16/pdf/scientificamerican10151904-260.pdf.

Anonymous:1904:DH

[6935] Anonymous. The diving horse. Scientific American, 91(16):260, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n16/pdf/scientificamerican10151904-260a.pdf.

Anonymous:1904:TWT

[6936] Anonymous. Transportable wireless telegraph station for war purposes. Scientific American, 91(16):260–261, October 15, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n16/pdf/scientificamerican10151904-260b.pdf.

Aaion:1904:IW

[6937] S. Frank Aaion. Insect wings. Scientific American, 91(16):261–262, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n16/pdf/scientificamerican10151904-261.pdf.

Anonymous:1904:DAW

[6938] Anonymous. Dastardly attempt to wreck the “Connecticut”. Scientific American, 91(16):262, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
262.pdf.

Anonymous:1904:CPP

[6939] Anonymous. The “cire-perdue” process of bronze casting. Scientific American, 91(16):264–265, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
264.pdf.

Hayward:1904:NSB

[6940] Charles B. Hayward. A novel suspension bridge. Scientific American, 91(16):265, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
265.pdf.

Anonymous:1904:EBH

[6941] Anonymous. Exhibits in the Boiler House and the Palace of Machinery, St. Louis. Scientific American, 91(16):265–267, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n16/pdf/scientificamerican10151904-265a.pdf.

Anonymous:1904:PDi

[6942] Anonymous. Patent department. Scientific American, 91(16):268, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n16/pdf/scientificamerican10151904-268.pdf.

Anonymous:1904:PPD

[6943] Anonymous. A portable pneumatic duster. Scientific American, 91(16):268, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
268a.pdf.

Anonymous:1904:FPS

[6944] Anonymous. Fruit and potato sorter. Scientific American, 91 (16):268, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
268b.pdf.

Anonymous:1904:DME

[6945] Anonymous. Device for muffling the exhaust from engines. Scientific American, 91(16):268, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
268c.pdf.

Anonymous:1904:OIe

[6946] Anonymous. Oddities in invention. Scientific American, 91(16): 268–269, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
268d.pdf.

Anonymous:1904:BPWbh

[6947] Anonymous. Business and personal wants. Scientific American, 91(16):271, October 15, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n16/pdf/scientificamerican10151904-
271.pdf.

Anonymous:1904:SNu

[6948] Anonymous. Science notes. Scientific American, 91(17):279, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n17/pdf/scientificamerican10221904-279b.pdf.

Anonymous:1904:Cbd

[6949] Anonymous. Correspondence. Scientific American, 91(17):283, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n17/pdf/scientificamerican10221904-283.pdf.

Anonymous:1904:RPIbi

[6950] Anonymous. Recently patented inventions. Scientific American, 91(17):288, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
288.pdf.

Anonymous:1904:NQbg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n17/pdf/scientificamerican10221904-288b.pdf.

Anonymous:1904:BNBx

[6952] Anonymous. Books: New books, etc. Scientific American, 91 (17):288, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
288c.pdf.

Anonymous:1904:IIbi

[6953] Anonymous. Index of inventions. Scientific American, 91(17): 288–290, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
288d.pdf.

Anonymous:1904:ARR

[6954] Anonymous. Automobile road racing. Scientific American, 91 (17):278, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
278.pdf.

Holder:1904:SDS

[6955] Charles F. Holder. Scientific disposition of sewage. Scientific American, 91(17):278–279, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
278a.pdf.

Walsh:1904:LSM

[6956] George E. Walsh. A life-saving museum. Scientific American, 91(17):279, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
279.pdf.

Anonymous:1904:SARb

[6957] Anonymous. The Scientific American reference book. Scientific American, 91(17):279, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
279a.pdf.

Radcliffe:1904:TWC

[6958] W. L. Radcliffe. Train wrecked by collision with dynamite. Scientific American, 91(17):280, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
280.pdf.

Willey:1904:TMB

[6959] Day Allen Willey. Testing machines at the Boston Institute of Technology. Scientific American, 91(17):280–281, October 22, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n17/pdf/scientificamerican10221904-280a.pdf.

Anonymous:1904:FPT

[6960] Anonymous. The Fery pyrometric telescope. Scientific American, 91(17):281, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
281.pdf.

Anonymous:1904:NII

[6961] Anonymous. A novel interrupter for induction coils. Scientific American, 91(17):281–282, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
281a.pdf.

Perkins:1904:NET

[6962] Frank C. Perkins. A new English telephonic fire alarm. Scientific American, 91(17):282–283, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
282.pdf.

Anonymous:1904:FCV

[6963] Anonymous. First competition for the Vanderbilt International Automobile Cup. Scientific American, 91(17):283–284, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n17/pdf/scientificamerican10221904-283a.pdf.

Anonymous:1904:SRL

[6964] Anonymous. Some recent launchings in the United States Navy. Scientific American, 91(17):285–287, October 22, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n17/pdf/scientificamerican10221904-285.pdf.

Priestley:1904:PL

[6965] Herbert I. Priestley. A peculiar lily. Scientific American, 91(17):287, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n17/pdf/scientificamerican10221904-287.pdf.

Anonymous:1904:ANS

[6966] Anonymous. An animal new to science at the New York Zoological Park. Scientific American, 91(17):287, October 22, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n17/pdf/scientificamerican10221904-287a.pdf.

Anonymous:1904:BPWbi

[6967] Anonymous. Business and personal wants. Scientific American, 91(17):288, October 22, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n17/pdf/scientificamerican10221904-
288a.pdf.

Anonymous:1904:Cbe

[6968] Anonymous. Correspondence. Scientific American, 91(18):299, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n18/pdf/scientificamerican10291904-299.pdf.

Anonymous:1904:IIbj

[6969] Anonymous. Index of inventions. Scientific American, 91(18): 304–306, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
304a.pdf.

Anonymous:1904:PNPbe

[6970] Anonymous. Patent notes: Patents. Scientific American, 91 (18):307, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
307.pdf.

Anonymous:1904:IRT

[6971] Anonymous. Inauguration of the rapid transit subway. Scientific American, 91(18):294, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
294.pdf.

Anonymous:1904:CEW

[6972] Anonymous. Cost of electric water power plants. Scientific American, 91(18):294, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
294a.pdf.

Anonymous:1904:SCD

[6973] Anonymous. To stimulate chemical discovery. Scientific American, 91(18):294, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
294b.pdf.

Anonymous:1904:NPMb

[6974] Anonymous. A new process of making trichromatic lantern slides. Scientific American, 91(18):295, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
295.pdf.

Russell:1904:HN

[6975] Henry Norris Russell. The heavens in November. Scientific American, 91(18):295, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
295a.pdf.

Anonymous:1904:RSAc

[6976] Anonymous. Romance of a scientific age. Scientific American, 91(18):295, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
295b.pdf.

Anonymous:1904:DPT

[6977] Anonymous. Device for pulling tacks and nails from pneumatic tires. Scientific American, 91(18):296, October 29, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n18/pdf/scientificamerican10291904-296.pdf.

Anonymous:1904:SRF

[6978] Anonymous. A solar reducing furnace. Scientific American, 91 (18):296, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
296a.pdf.

Anonymous:1904:ILK

[6979] Anonymous. Improved lime kiln. Scientific American, 91(18):296, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n18/pdf/scientificamerican10291904-296b.pdf.

Anonymous:1904:PLV

[6980] Anonymous. Passenger locomotive with Vanderbilt tender. Scientific American, 91(18):297, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
297.pdf.

Anonymous:1904:ONY

[6981] Anonymous. Opening of the New York Rapid Transit Subway. Scientific American, 91(18):297–298, October 29, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n18/pdf/scientificamerican10291904-297a.pdf.

Anonymous:1904:ESF

[6982] Anonymous. Effect of shell fire on the Russian cruiser “Askold”. Scientific American, 91(18):300–301, October 29, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n18/pdf/scientificamerican10291904-300.pdf.

Anonymous:1904:EVW

[6983] Anonymous. Eskimo village at the World’s Fair. Scientific American, 91(18):301–302, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
301.pdf.

Anonymous:1904:PAC

[6984] Anonymous. Preparations for the Airship Contest at St. Louis. Scientific American, 91(18):302, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
302.pdf.

Anonymous:1904:MBN

[6985] Anonymous. The mounting of butterflies for nature study. Scientific American, 91(18):303, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
303.pdf.

Anonymous:1904:BPWbj

[6986] Anonymous. Business and personal wants. Scientific American, 91(18):304, October 29, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n18/pdf/scientificamerican10291904-
304.pdf.

Anonymous:1904:Cbf

[6987] Anonymous. Correspondence. Scientific American, 91(19):314, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n19/pdf/scientificamerican11051904-314.pdf.

Anonymous:1904:RPIbj

[6988] Anonymous. Recently patented inventions. Scientific American, 91(19):321, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
321.pdf.

Anonymous:1904:NQbh

[6989] Anonymous. Notes and queries. Scientific American, 91(19): 322, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
322a.pdf.

Anonymous:1904:BNBy

[6990] Anonymous. Books: New books, etc. Scientific American, 91 (19):322, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
322b.pdf.

Anonymous:1904:IIbk

[6991] Anonymous. Index of inventions. Scientific American, 91(19): 322–327, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
322c.pdf.

Anonymous:1904:SDS

[6992] Anonymous. The shocking disfigurement of the subway. Scientific American, 91(19):310, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
310.pdf.

Anonymous:1904:RAS

[6993] Anonymous. Rapid acceleration of the subway trains. Scientific American, 91(19):310, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
310a.pdf.

Anonymous:1904:SDEa

[6994] Anonymous. Scale divisions and errors of graduation. Scientific American, 91(19):310, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
310b.pdf.

Anonymous:1904:NMO

[6995] Anonymous. A new method of observing the N-rays. Scientific American, 91(19):310, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
310c.pdf.

Bellet:1904:PPM

[6996] Daniel Bellet. The personalities of Professor and Madame Curie. Scientific American, 91(19):311, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
311.pdf.

Anonymous:1904:BAF

[6997] Anonymous. Benbow’s airship flight. Scientific American, 91 (19):311, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
311a.pdf.

Anonymous:1904:OHF

[6998] Anonymous. One hundred and fifty miles an hour on electric road. Scientific American, 91(19):311, November 5, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n19/pdf/scientificamerican11051904-311b.pdf.

Guarini:1904:VND

[6999] Emile Guarini. The Verant: A new device for viewing photographs. Scientific American, 91(19):312, November 5, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n19/pdf/scientificamerican11051904-312.pdf.

Anonymous:1904:MCF

[7000] Anonymous. Model of a continuous-feed gas producer at the St. Louis Fair. Scientific American, 91(19):312, November 5, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n19/pdf/scientificamerican11051904-312a.pdf.

Anonymous:1904:USM

[7001] Anonymous. The United States Mint at the St. Louis Exposition. Scientific American, 91(19):313, November 5, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n19/pdf/scientificamerican11051904-313.pdf.

Anonymous:1904:NSI

[7002] Anonymous. A novel speed indicator for automobiles. Scientific American, 91(19):313–314, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
313a.pdf.

Aaron:1904:MWI

[7003] S. Frank Aaron. The musical wings of insects. Scientific American, 91(19):315–316, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
315.pdf.

Willey:1904:CA

[7004] Day Allen Willey. The Carrara of America. Scientific American, 91 (19):317–318, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
317.pdf.

Anonymous:1904:SBF

[7005] Anonymous. The start of the Baltic Fleet. Scientific American, 91(19):318, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
318.pdf.

Anonymous:1904:PDj

[7006] Anonymous. Patent department. Scientific American, 91(19): 319, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
319.pdf.

Anonymous:1904:SMd

[7007] Anonymous. Solar motor. Scientific American, 91(19):319, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n19/pdf/scientificamerican11051904-319a.pdf.

Anonymous:1904:ABH

[7008] Anonymous. Attachment for bolt-heading machines. Scientific American, 91(19):319, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
319b.pdf.

Anonymous:1904:OIf

[7009] Anonymous. Oddities in invention. Scientific American, 91(19): 319–320, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
319c.pdf.

Anonymous:1904:BPWbk

[7010] Anonymous. Business and personal wants. Scientific American, 91(19):322, November 5, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n19/pdf/scientificamerican11051904-
322.pdf.

Anonymous:1904:SNv

[7011] Anonymous. Science notes. Scientific American, 91(2):23, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-23b.pdf.

Anonymous:1904:Cbg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-27.pdf.

Anonymous:1904:RPIbk

[7013] Anonymous. Recently patented inventions. Scientific American, 91(2): 32, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-32.pdf.

Anonymous:1904:IIbl

[7014] Anonymous. Index of inventions. Scientific American, 91(2):32–35, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-32c.pdf.

Anonymous:1904:PNPbf

[7015] Anonymous. Patent notes: Patents. Scientific American, 91(2):35, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-35.pdf.

Anonymous:1904:SDEb

[7016] Anonymous. Santos-Dumont and the Exposition authorities. Scientific American, 91(2):22, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n2/pdf/scientificamerican07091904-
22.pdf.

Anonymous:1904:MG

[7017] Anonymous. Machine grinding. Scientific American, 91(2):22, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-22a.pdf.

Anonymous:1904:PME

[7018] Anonymous. The progress of mosquito extermination. Scientific American, 91(2):22–23, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n2/pdf/scientificamerican07091904-
22b.pdf.

Anonymous:1904:VB

[7019] Anonymous. A vegetable butter. Scientific American, 91(2):23, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-23.pdf.

Anonymous:1904:HFF

[7020] Anonymous. Historic forest fires. Scientific American, 91(2):23, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-23a.pdf.

Anonymous:1904:PCF

[7021] Anonymous. The preparation of cereal foods. Scientific American, 91(2):24–27, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n2/pdf/scientificamerican07091904-
24.pdf.

Anonymous:1904:STTb

[7022] Anonymous. The submarine torpedo tube. Scientific American, 91(2): 28, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-28.pdf.

Anonymous:1904:GBC

[7023] Anonymous. The Gordon Bennett Cup Race. Scientific American, 91(2):29–30, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n2/pdf/scientificamerican07091904-
29.pdf.

Wraith:1904:CFH

[7024] J. C. Wraith. A curious formation of horse’s skull. Scientific American, 91(2):31, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n2/pdf/scientificamerican07091904-
31.pdf.

Gradenwitz:1904:NNF

[7025] Alfred Gradenwitz. The new Nemethy flying-machine and the principle of its construction. Scientific American, 91(2):31, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-31a.pdf.

Anonymous:1904:ISa

[7026] Anonymous. Improved screwdriver. Scientific American, 91(2):32, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-32a.pdf.

Anonymous:1904:BPWbl

[7027] Anonymous. Business and personal wants. Scientific American, 91(2): 32, July 9, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n2/pdf/scientificamerican07091904-32b.pdf.

Anonymous:1904:SNw

[7028] Anonymous. Science notes. Scientific American, 91(20):331, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n20/pdf/scientificamerican11121904-331a.pdf.

Anonymous:1904:Cbh

[7029] Anonymous. Correspondence. Scientific American, 91(20):335, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n20/pdf/scientificamerican11121904-335.pdf.

Anonymous:1904:RPIbl

[7030] Anonymous. Recently patented inventions. Scientific American, 91 (20):341–342, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
341.pdf.

Anonymous:1904:NQbi

[7031] Anonymous. Notes and queries. Scientific American, 91(20): 342–343, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
342a.pdf.

Anonymous:1904:BNBz

[7032] Anonymous. Books: New books, etc. Scientific American, 91 (20):343, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
343.pdf.

Anonymous:1904:IIbm

[7033] Anonymous. Index of inventions. Scientific American, 91(20): 343–347, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
343a.pdf.

Anonymous:1904:PNPbg

[7034] Anonymous. Patent notes: Patents. Scientific American, 91 (20):347, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
347.pdf.

Anonymous:1904:FAI

[7035] Anonymous. Free alcohol for industrial purposes. Scientific American, 91(20):330, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
330.pdf.

Anonymous:1904:ISc

[7036] Anonymous. Indian summer. Scientific American, 91(20):330, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n20/pdf/scientificamerican11121904-330a.pdf.

Reuter:1904:TD

[7037] L. Reuter. Turpentine derivatives. Scientific American, 91(20): 331, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
331.pdf.

Guarini:1904:PMS

[7038] Emile Guarini. Penetration of mercury by seed radicels. Scientific American, 91(20):332, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
332.pdf.

Anonymous:1904:ECC

[7039] Anonymous. Erecting the columns for the Choir of the Cathedral of St. John the Divine. Scientific American, 91(20):332, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n20/pdf/scientificamerican11121904-332a.pdf.

Mclure:1904:ROU

[7040] W. Frank M’clure. A record in ore unloading. Scientific American, 91(20):333, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
333.pdf.

Anonymous:1904:ERL

[7041] Anonymous. Electric railway locomotives for the New York Central. Scientific American, 91(20):333–334, November 12, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n20/pdf/scientificamerican11121904-333a.pdf.

Anonymous:1904:TTA

[7042] Anonymous. The trial trip of the armored cruiser “Colorado”. Scientific American, 91(20):336–337, November 12, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n20/pdf/scientificamerican11121904-336.pdf.

Anonymous:1904:NSL

[7043] Anonymous. Notable St. Louis airships. Scientific American, 91 (20):337–338, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
337.pdf.

Anonymous:1904:CDS

[7044] Anonymous. The cliff dwellers at the St. Louis Exposition. Scientific American, 91(20):339, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
339.pdf.

Willey:1904:LDS

[7045] Day Allen Willey. A locomotive duel scientifically conducted. Scientific American, 91(20):340, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
340.pdf.

Anonymous:1904:BPWbm

[7046] Anonymous. Business and personal wants. Scientific American, 91(20):342, November 12, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n20/pdf/scientificamerican11121904-
342.pdf.

Anonymous:1904:SNx

[7047] Anonymous. Science notes. Scientific American, 91(21):351, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n21/pdf/scientificamerican11191904-351a.pdf.

Anonymous:1904:Cbi

[7048] Anonymous. Correspondence. Scientific American, 91(21):355, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n21/pdf/scientificamerican11191904-355.pdf.

Anonymous:1904:LNj

[7049] Anonymous. Legal notes. Scientific American, 91(21):361, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n21/pdf/scientificamerican11191904-361.pdf.

Anonymous:1904:RPIbm 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
362.pdf.

Anonymous:1904:NQbj

[7051] Anonymous. Notes and queries. Scientific American, 91(21): 363, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
363a.pdf.

Anonymous:1904:BNBba

[7052] Anonymous. Books: New books, etc. Scientific American, 91 (21):363, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
363b.pdf.

Anonymous:1904:IIbn

[7053] Anonymous. Index of inventions. Scientific American, 91(21): 363–367, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
363c.pdf.

Anonymous:1904:PNPbh

[7054] Anonymous. Patent notes: Patents. Scientific American, 91 (21):367, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
367.pdf.

Anonymous:1904:IRFb

[7055] Anonymous. Increasing railroad fatalities. Scientific American, 91 (21):350, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
350.pdf.

Anonymous:1904:VS

[7056] Anonymous. The ventilation of the subway. Scientific American, 91(21):350, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
350a.pdf.

Anonymous:1904:MDP

[7057] Anonymous. The mining debris problem in California. Scientific American, 91(21):350, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
350b.pdf.

Anonymous:1904:WST

[7058] Anonymous. Work on the simplon tunnel. Scientific American, 91 (21):350–351, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
350c.pdf.

Wilcox:1904:AUI

[7059] Marrion Wilcox. The automobile and its utility in the industrial development of Latin America. Scientific American, 91(21):351, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n21/pdf/scientificamerican11191904-351.pdf.

Collins:1904:CI

[7060] A. Frederick Collins. A carboy inclinator. Scientific American, 91 (21):352, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
352.pdf.

Anonymous:1904:ICG

[7061] Anonymous. An improved cattle guard. Scientific American, 91 (21):352, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
352a.pdf.

Anonymous:1904:RSM

[7062] Anonymous. Revolving scrubbing machine. Scientific American, 91(21):352, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
352b.pdf.

Anonymous:1904:MCP

[7063] Anonymous. The Miethe color projector. Scientific American, 91 (21):353, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
353.pdf.

Anonymous:1904:RAMa

[7064] Anonymous. Recent automobile and motor-boat speed records. Scientific American, 91(21):353–354, November 19, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n21/pdf/scientificamerican11191904-353a.pdf.

Anonymous:1904:GNP

[7065] Anonymous. The German National Pavilion at St. Louis. Scientific American, 91(21):356–357, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
356.pdf.

Johnson:1904:CTC

[7066] Charles C. Johnson. The cultivation of tea in Ceylon. Scientific American, 91(21):357–358, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
357.pdf.

Galpin:1904:WOR

[7067] F. W. Galpin. The water organ — a Roman keyboard instrument. Scientific American, 91(21):358–360, November 19, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n21/pdf/scientificamerican11191904-358.pdf.

Jones:1904:TVT

[7068] H. L. Jones. Tomato vines thirty feet long. Scientific American, 91(21):360, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
360.pdf.

Anonymous:1904:CPC

[7069] Anonymous. A curious piece of casting. Scientific American, 91 (21):360, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
360a.pdf.

Anonymous:1904:BPWbn

[7070] Anonymous. Business and personal wants. Scientific American, 91(21):363, November 19, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n21/pdf/scientificamerican11191904-
363.pdf.

Anonymous:1904:SNy

[7071] Anonymous. Science notes. Scientific American, 91(22):371, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n22/pdf/scientificamerican11261904-371b.pdf.

Anonymous:1904:Cbj

[7072] Anonymous. Correspondence. Scientific American, 91(22):375, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n22/pdf/scientificamerican11261904-375.pdf.

Anonymous:1904:RPIbn

[7073] Anonymous. Recently patented inventions. Scientific American, 91(22):382, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
382.pdf.

Anonymous:1904:NQbk

[7074] Anonymous. Notes and queries. Scientific American, 91(22): 383, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
383.pdf.

Anonymous:1904:BNBbb

[7075] Anonymous. Books: New books, etc. Scientific American, 91 (22):383, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
383a.pdf.

Anonymous:1904:IIbo

[7076] Anonymous. Index of inventions. Scientific American, 91(22): 383–386, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
383b.pdf.

Anonymous:1904:PNPbi

[7077] Anonymous. Patent notes: Patents. Scientific American, 91 (22):387, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
387.pdf.

Anonymous:1904:PA

[7078] Anonymous. Port Arthur. Scientific American, 91(22):370, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n22/pdf/scientificamerican11261904-370.pdf.

Anonymous:1904:HJP

[7079] Anonymous. Hydraulic jet propulsion. Scientific American, 91 (22):370, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
370a.pdf.

Anonymous:1904:VRS

[7080] Anonymous. The vast railroad system of the United States. Scientific American, 91(22):370–371, November 26, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n22/pdf/scientificamerican11261904-370b.pdf.

Anonymous:1904:WCF

[7081] Anonymous. Water-curtain fire protection. Scientific American, 91 (22):371, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
371.pdf.

Anonymous:1904:LFM

[7082] Anonymous. Loads on factory motors. Scientific American, 91 (22):371, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
371a.pdf.

Anonymous:1904:ASL

[7083] Anonymous. Automatic step-ladders. Scientific American, 91 (22):372, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
372.pdf.

Anonymous:1904:GMP

[7084] Anonymous. Gasoline-motor-propelled fire engine. Scientific American, 91(22):372, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
372a.pdf.

Steele:1904:WMI

[7085] John S. Steele. The wine-making industry of New York State. Scientific American, 91(22):373–374, November 26, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n22/pdf/scientificamerican11261904-373.pdf.

Geare:1904:ARYa

[7086] Randolph I. Geare. The ancient races of Yucatan and Mexico. — I. Scientific American, 91(22):375, November 26, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n22/pdf/scientificamerican11261904-375a.pdf.

Anonymous:1904:JSL

[7087] Anonymous. Japan at the St. Louis Fair. Scientific American, 91(22):376, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
376.pdf.

Claudy:1904:NAC

[7088] C. H. Claudy. The new asteroid camera at the United States Naval Observatory. Scientific American, 91(22):377–378, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n22/pdf/scientificamerican11261904-377.pdf.

Beard:1904:NAZ

[7089] J. Carter Beard. New antelopes at the N.Y. Zoological Park. Scientific American, 91(22):379, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
379.pdf.

Anonymous:1904:PDk

[7090] Anonymous. Patent departement. Scientific American, 91(22): 380, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
380.pdf.

Anonymous:1904:BM

[7091] Anonymous. Blow-off mechanism. Scientific American, 91(22): 380, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
380a.pdf.

Anonymous:1904:SAC

[7092] Anonymous. The Schebler automatic carbureter. Scientific American, 91(22):380, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
380b.pdf.

Anonymous:1904:ACH

[7093] Anonymous. An adjustable candle holder. Scientific American, 91 (22):380, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
380c.pdf.

Anonymous:1904:OIg

[7094] Anonymous. Oddities in invention. Scientific American, 91(22): 381, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
381.pdf.

Anonymous:1904:BPWbo

[7095] Anonymous. Business and personal wants. Scientific American, 91(22):382, November 26, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n22/pdf/scientificamerican11261904-
382a.pdf.

Anonymous:1904:Cbk

[7096] Anonymous. Correspondence. Scientific American, 91(23):395, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n23/pdf/scientificamerican12031904-395.pdf.

Anonymous:1904:RPIbo

[7097] Anonymous. Recently patented inventions. Scientific American, 91(23):400, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
400.pdf.

Anonymous:1904:NQbl

[7098] Anonymous. Notes and queries. Scientific American, 91(23):400, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n23/pdf/scientificamerican12031904-400b.pdf.

Anonymous:1904:IIbp 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
400c.pdf.

Anonymous:1904:PNPbj

[7100] Anonymous. Patent notes: Patents. Scientific American, 91 (23):403, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
403.pdf.

Anonymous:1904:PPe

[7101] Anonymous. Production and population. Scientific American, 91 (23):390, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
390.pdf.

Anonymous:1904:GT

[7102] Anonymous. Gas turbines. Scientific American, 91(23):390, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n23/pdf/scientificamerican12031904-390a.pdf.

Anonymous:1904:POW

[7103] Anonymous. Power of ocean waves. Scientific American, 91 (23):390, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
390b.pdf.

Anonymous:1904:ETTb

[7104] Anonymous. Electric traction for trunk roads. Scientific American, 91(23):390, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
390c.pdf.

Anonymous:1904:SR

[7105] Anonymous. Subway results. Scientific American, 91(23):390, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n23/pdf/scientificamerican12031904-390d.pdf.

Anonymous:1904:USW

[7106] Anonymous. A United States warship on trial. Scientific American, 91(23):390–391, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
390e.pdf.

Reed:1904:HD

[7107] W. M. Reed. The heavens in December. Scientific American, 91(23):391, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
391.pdf.

Anonymous:1904:SNz

[7108] Anonymous. Science notes. Scientific American, 91(23):391, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n23/pdf/scientificamerican12031904-391a.pdf.

Walker:1904:LSS

[7109] Herbert T. Walker. A locomotive in service sixty years. Scientific American, 91(23):392, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
392.pdf.

Anonymous:1904:TLC

[7110] Anonymous. A tilting lock-chamber. Scientific American, 91 (23):392, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
392a.pdf.

Anonymous:1904:NFL

[7111] Anonymous. New form of liquid rheostat for starting electric motors. Scientific American, 91(23):393, December 3, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n23/pdf/scientificamerican12031904-393.pdf.

Guarini:1904:RMP

[7112] Emile Guarini. The registration of meteorological phenomena in Lapland. Scientific American, 91(23):393–394, December 3, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n23/pdf/scientificamerican12031904-393a.pdf.

Anonymous:1904:IR

[7113] Anonymous. An interesting relic. Scientific American, 91(23):394, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n23/pdf/scientificamerican12031904-394.pdf.

Bellet:1904:HB

[7114] Daniel Bellet. Henri Becquerel. Scientific American, 91(23):395, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n23/pdf/scientificamerican12031904-395a.pdf.

Crafts:1904:CRM

[7115] H. A. Crafts. The California redwood in modern engineering. Scientific American, 91(23):396–398, December 3, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n23/pdf/scientificamerican12031904-396.pdf.

Geare:1904:ARYb

[7116] Randolph I. Geare. The ancient races of Yucatan and Mexico. — II. Scientific American, 91(23):398, December 3, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n23/pdf/scientificamerican12031904-398.pdf.

Anonymous:1904:Cbl

[7117] Anonymous. Comet “1904 a”. Scientific American, 91(23):398, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n23/pdf/scientificamerican12031904-398a.pdf.

Beasley:1904:WFC

[7118] Walter L. Beasley. Wax farming in China — a strange industry. Scientific American, 91(23):399, December 3, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n23/pdf/scientificamerican12031904-399.pdf.

Anonymous:1904:BPWbp

[7119] Anonymous. Business and personal wants. Scientific American, 91(23):400, December 3, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n23/pdf/scientificamerican12031904-
400a.pdf.

Anonymous:1904:SNba

[7120] Anonymous. Science notes. Scientific American, 91(24):407– 408, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
407b.pdf.

Anonymous:1904:Cbn

[7121] Anonymous. Correspondence. Scientific American, 91(24):411– 412, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
411.pdf.

Anonymous:1904:LNk

[7122] Anonymous. Legal notes. Scientific American, 91(24):417–418, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n24/pdf/scientificamerican12101904-417.pdf.

Anonymous:1904:RPIbp

[7123] Anonymous. Recently patented inventions. Scientific American, 91 (24):418–419, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
418.pdf.

Anonymous:1904:NQbm

[7124] Anonymous. Notes and queries. Scientific American, 91(24): 419–420, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
419a.pdf.

Anonymous:1904:BNBbc

[7125] Anonymous. Books: New books, etc. Scientific American, 91 (24):420, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
420.pdf.

Anonymous:1904:IIbq

[7126] Anonymous. Index of inventions. Scientific American, 91(24): 420–423, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
420a.pdf.

Anonymous:1904:PNPbk

[7127] Anonymous. Patent notes: Patents. Scientific American, 91 (24):423, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
423.pdf.

Anonymous:1904:CSLb

[7128] Anonymous. Close of the St. Louis Fair. Scientific American, 91 (24):406, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
406.pdf.

Anonymous:1904:BSD

[7129] Anonymous. Baldwin and Santos-Dumont. Scientific American, 91(24):406, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
406a.pdf.

Anonymous:1904:DFY

[7130] Anonymous. Death of a famous yacht designer. Scientific American, 91(24):406, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
406b.pdf.

Anonymous:1904:STS

[7131] Anonymous. Steel trackway on streets. Scientific American, 91 (24):406, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
406c.pdf.

Anonymous:1904:MOS

[7132] Anonymous. Meteorological observatories at sea. Scientific American, 91(24):406–407, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
406d.pdf.

Smoot:1904:WLD

[7133] Thomas Arthur Smoot. Will “lightwood” displace the long-leaf pine in turpentine distillation? Scientific American, 91(24):407, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n24/pdf/scientificamerican12101904-407.pdf.

Anonymous:1904:RAMb

[7134] Anonymous. Radio-active minerals. Scientific American, 91(24): 407, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
407a.pdf.

Anonymous:1904:SBA

[7135] Anonymous. A simple burglar alarm. Scientific American, 91 (24):408, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
408.pdf.

Anonymous:1904:SVL

[7136] Anonymous. A sea voyage on land. Scientific American, 91(24): 408–409, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
408a.pdf.

Peltier:1904:RRC

[7137] J. Peltier. River rake for channel deepening. Scientific American, 91(24):409, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
409.pdf.

Anonymous:1904:SMW

[7138] Anonymous. Skeleton models of warships. Scientific American, 91 (24):409–411, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
409a.pdf.

Anonymous:1904:RES

[7139] Anonymous. The Racial Exhibit at the St. Louis Fair. Scientific American, 91(24):412–413, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
412.pdf.

Perkins:1904:UED

[7140] Frank C. Perkins. The use of electricity in driving coal-conveying machinery. Scientific American, 91(24):413–415, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n24/pdf/scientificamerican12101904-413.pdf.

Anonymous:1904:RGD

[7141] Anonymous. The Roman galleys discovered in Lake Nemi. Scientific American, 91(24):415–416, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
415.pdf.

Anonymous:1904:PDl

[7142] Anonymous. Patent department. Scientific American, 91(24): 416, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
416.pdf.

Anonymous:1904:IEM

[7143] Anonymous. Improved exercising machine. Scientific American, 91(24):416, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
416a.pdf.

Anonymous:1904:SMV

[7144] Anonymous. Sheet metal vehicle wheel. Scientific American, 91 (24):416, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
416b.pdf.

Anonymous:1904:MM

[7145] Anonymous. Milking machine. Scientific American, 91(24):416– 417, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
416c.pdf.

Anonymous:1904:BPWbq

[7146] Anonymous. Business and personal wants. Scientific American, 91(24):419, December 10, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n24/pdf/scientificamerican12101904-
419.pdf.

Anonymous:1904:RPIbq

[7147] Anonymous. Recently patented inventions. Scientific American, 91 (25):444–445, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
444.pdf.

Anonymous:1904:NQbn

[7148] Anonymous. Notes and queries. Scientific American, 91(25): 445–446, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
445a.pdf.

Anonymous:1904:BNBbd

[7149] Anonymous. Books: New books, etc. Scientific American, 91(25): 446–454, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
446.pdf.

Anonymous:1904:IIbr

[7150] Anonymous. Index of inventions. Scientific American, 91(25): 450–455, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
450.pdf.

Anonymous:1904:PNPbl

[7151] Anonymous. Patent notes: Patents. Scientific American, 91 (25):455, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
455.pdf.

Anonymous:1904:PCP

[7152] Anonymous. Panama Canal problems. Scientific American, 91 (25):426, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
426.pdf.

Anonymous:1904:BPWbr

[7153] Anonymous. Blondlot and Prof. Wood on the N-rays. Scientific American, 91(25):426, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
426a.pdf.

Anonymous:1904:ESI

[7154] Anonymous. The electric smelting of iron ore. Scientific American, 91(25):426, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
426b.pdf.

Anonymous:1904:LACb

[7155] Anonymous. Launch of the armored cruiser “Tennessee”. Scientific American, 91(25):426, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
426c.pdf.

Anonymous:1904:NAS

[7156] Anonymous. News about the submerged coal experiments. Scientific American, 91(25):426–427, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
426d.pdf.

Anonymous:1904:AEG

[7157] Anonymous. “American estates and gardens”. Scientific American, 91(25):427, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
427.pdf.

Anonymous:1904:NBP

[7158] Anonymous. The new British Patent Act. Scientific American, 91(25):427, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
427a.pdf.

Anonymous:1904:OS

[7159] Anonymous. To our subscribers. Scientific American, 91(25): 427, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
427b.pdf.

Anonymous:1904:SNbb

[7160] Anonymous. Science notes. Scientific American, 91(25):427, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n25/pdf/scientificamerican12171904-427c.pdf.

Anonymous:1904:RRA

[7161] Anonymous. Radium and radio-activity. Scientific American, 91 (25):428, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
428.pdf.

Shepstone:1904:GLP

[7162] Harold J. Shepstone. The Great London Power Station. Scientific American, 91(25):429–430, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
429.pdf.

Anonymous:1904:RRN

[7163] Anonymous. Roman remains in North Africa. Scientific American, 91(25):430–431, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
430.pdf.

Anonymous:1904:HPTb

[7164] Anonymous. The high-powered traction engine and its many uses. Scientific American, 91(25):433–434, December 17, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n25/pdf/scientificamerican12171904-433.pdf.

Ramakers:1904:PIE

[7165] L. Ramakers. Photography of the interior of the eye. Scientific American, 91(25):435, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
435.pdf.

Anonymous:1904:AFM

[7166] Anonymous. Aeroplanes in France and M. Archdeacon’s apparatus. — rules for concourses. Scientific American, 91(25):435–436, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n25/pdf/scientificamerican12171904-435a.pdf.

Guarini:1904:BIS

[7167] Emile Guarini. The bioscope — an instrument for the study of insect life. Scientific American, 91(25):437–438, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n25/pdf/scientificamerican12171904-437.pdf.

Willey:1904:ISC

[7168] Day Allen Willey. Irrigation in Southern California. Scientific American, 91(25):439, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
439.pdf.

Anonymous:1904:BIM

[7169] Anonymous. Battleships “Idaho” and “Mississippi”. Scientific American, 91(25):440, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
440.pdf.

Anonymous:1904:MBW

[7170] Anonymous. The modern battleship within and without. Scientific American, 91(25):440–442, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
440a.pdf.

Galpin:1904:HWM

[7171] P. W. Galpin. The how and why of the modern keyboard. — I. Scientific American, 91(25):442–443, December 17, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n25/pdf/scientificamerican12171904-442.pdf.

Anonymous:1904:BPWbs

[7172] Anonymous. Business and personal wants. Scientific American, 91(25):445, December 17, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n25/pdf/scientificamerican12171904-
445.pdf.

Anonymous:1904:SNbc

[7173] Anonymous. Science notes. Scientific American, 91(26):459, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n26/pdf/scientificamerican12241904-459b.pdf.

Anonymous:1904:Cbo

[7174] Anonymous. Correspondence. Scientific American, 91(26):463, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n26/pdf/scientificamerican12241904-463.pdf.

Anonymous:1904:RPIbr

[7175] Anonymous. Recently patented inventions. Scientific American, 91(26):468, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
468.pdf.

Anonymous:1904:NQbo

[7176] Anonymous. Notes and queries. Scientific American, 91(26): 468, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
468b.pdf.

Anonymous:1904:BNBbe

[7177] Anonymous. Books: New books, etc. Scientific American, 91 (26):468, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
468c.pdf.

Anonymous:1904:IIbs

[7178] Anonymous. Index of inventions. Scientific American, 91(26): 468–470, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
468d.pdf.

Anonymous:1904:PNPbm 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
471.pdf.

Anonymous:1904:GPI

[7180] Anonymous. A great public improvement. Scientific American, 91(26):458, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
458.pdf.

Anonymous:1904:CRc

[7181] Anonymous. A canard and its refutation. Scientific American, 91 (26):458, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
458a.pdf.

Anonymous:1904:CRJ

[7182] Anonymous. The crisis of the Russo-Japanese War. Scientific American, 91(26):458–459, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
458b.pdf.

Cleveland:1904:TWP

[7183] W. S. Cleveland. Tidal water power-means of obtaining a constant head of about thirty-eight feet. Scientific American, 91(26):459, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n26/pdf/scientificamerican12241904-459.pdf.

Anonymous:1904:WBA

[7184] Anonymous. Wireless burglar alarm for safes. Scientific American, 91(26):459, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
459a.pdf.

Anonymous:1904:PPg

[7185] Anonymous. Pneumatic pump. Scientific American, 91(26): 460, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
460.pdf.

Anonymous:1904:EAF

[7186] Anonymous. Electrical apparatus for felling trees. Scientific American, 91(26):460, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
460a.pdf.

Anonymous:1904:DHC

[7187] Anonymous. Destructive head-on collision. Scientific American, 91(26):460, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
460b.pdf.

Beasley:1904:NHP

[7188] Walter L. Beasley. A new home for the Peary Meteorite. Scientific American, 91(26):461, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
461.pdf.

Anonymous:1904:ABD

[7189] Anonymous. The Aerial Bridge at Duluth. Scientific American, 91 (26):461–462, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
461a.pdf.

Anonymous:1904:HCS

[7190] Anonymous. The Haschke compound storage battery. Scientific American, 91(26):462, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
462.pdf.

Anonymous:1904:TS

[7191] Anonymous. A tree on stilts. Scientific American, 91(26):464, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n26/pdf/scientificamerican12241904-464.pdf.

Bennett:1904:FPS

[7192] H. I. Bennett. Fire protection for snowsheds on the Central Pacific Railroad. Scientific American, 91(26):464, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n26/pdf/scientificamerican12241904-464a.pdf.

Johnson:1904:LGW

[7193] Oliver Johnson. The largest geyser in the world. Scientific American, 91(26):465–466, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
465.pdf.

Anonymous:1904:NAA

[7194] Anonymous. A novel adaptation of the Archimedean screw. Scientific American, 91(26):467, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
467.pdf.

Anonymous:1904:RF

[7195] Anonymous. A river of fish. Scientific American, 91(26):467, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n26/pdf/scientificamerican12241904-467a.pdf.

Anonymous:1904:BPWbt

[7196] Anonymous. Business and personal wants. Scientific American, 91(26):468, December 24, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n26/pdf/scientificamerican12241904-
468a.pdf.

Anonymous:1904:SNbd

[7197] Anonymous. Science notes. Scientific American, 91(3):39, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-39a.pdf.

Anonymous:1904:Cbp

[7198] Anonymous. Correspondence. Scientific American, 91(3):43, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-43.pdf.

Anonymous:1904:RPIbs

[7199] Anonymous. Recently patented inventions. Scientific American, 91(3):49–50, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n3/pdf/scientificamerican07161904-
49.pdf.

Anonymous:1904:NQbp

[7200] Anonymous. Notes and queries. Scientific American, 91(3):51, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-51.pdf.

Anonymous:1904:BNBbf

[7201] Anonymous. Books: New books, etc. Scientific American, 91(3):51–52, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-51a.pdf.

Anonymous:1904:IIbt

[7202] Anonymous. Index of inventions. Scientific American, 91(3):52–55, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-52.pdf.

Anonymous:1904:PNPbn

[7203] Anonymous. Patent notes: Patents. Scientific American, 91(3):55, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-55.pdf.

Anonymous:1904:HLM

[7204] Anonymous. How long must we suffer. Scientific American, 91(3):38, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-38.pdf.

Anonymous:1904:CSBa

[7205] Anonymous. Corrosion of steam boilers. Scientific American, 91(3):38, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-38a.pdf.

Anonymous:1904:NMG

[7206] Anonymous. New method of gold extraction from pyrites used in Italy. Scientific American, 91(3):38, July 16, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n3/pdf/scientificamerican07161904-38b.pdf.

Anonymous:1904:BF

[7207] Anonymous. Borax, and food. Scientific American, 91(3):39, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-39.pdf.

Anonymous:1904:NAC

[7208] Anonymous. A novel air compressor. Scientific American, 91(3):40, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-40.pdf.

Gradenwitz:1904:APS

[7209] Alfred Gradenwitz. An apparatus for preventing seasickness. Scientific American, 91(3):40, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n3/pdf/scientificamerican07161904-
40a.pdf.

Anonymous:1904:PNM

[7210] Anonymous. The plans for the new Manhattan Bridge. Scientific American, 91(3):41–42, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n3/pdf/scientificamerican07161904-
41.pdf.

Anonymous:1904:NWS

[7211] Anonymous. The new White Star liner “Baltic” — the largest vessel in the world. Scientific American, 91(3):44–46, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-44.pdf.

Anonymous:1904:CCS

[7212] Anonymous. Cascades and colonnade of states at the Exposition. Scientific American, 91(3):46, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n3/pdf/scientificamerican07161904-
46.pdf.

Briggs:1904:LP

[7213] N. R. Briggs. Lensless photography. Scientific American, 91(3):46, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-46a.pdf.

Anonymous:1904:NAPc

[7214] Anonymous. The new Australian Patent Act. Scientific American, 91(3):47, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n3/pdf/scientificamerican07161904-
47.pdf.

Anonymous:1904:VM

[7215] Anonymous. Voting machine. Scientific American, 91(3):47, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-47a.pdf.

Anonymous:1904:CWa

[7216] Anonymous. A compensating windmill. Scientific American, 91 (3):47–48, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n3/pdf/scientificamerican07161904-
47b.pdf.

Anonymous:1904:IG

[7217] Anonymous. Improved gage. Scientific American, 91(3):48, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-48.pdf.

Anonymous:1904:HPC

[7218] Anonymous. Hose or pipe coupling. Scientific American, 91(3):48, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-48a.pdf.

Anonymous:1904:GLH

[7219] Anonymous. Grain lifter for headers. Scientific American, 91(3):48, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-48b.pdf.

Anonymous:1904:VES

[7220] Anonymous. Ventilated eye-shade. Scientific American, 91(3):48, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-48c.pdf.

Anonymous:1904:CCa

[7221] Anonymous. Chain clamp. Scientific American, 91(3):48, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n3/pdf/scientificamerican07161904-48d.pdf.

Anonymous:1904:BPWbu

[7222] Anonymous. Business and personal wants. Scientific American, 91(3):50, July 16, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n3/pdf/scientificamerican07161904-
50.pdf.

Anonymous:1904:SNbe

[7223] Anonymous. Science notes. Scientific American, 91(4):59, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n4/pdf/scientificamerican07231904-59b.pdf.

Anonymous:1904:RPIbt

[7224] Anonymous. Recently patented inventions. Scientific American, 91(4):67–68, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
67a.pdf.

Anonymous:1904:NQbq

[7225] Anonymous. Notes and queries. Scientific American, 91(4):68, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n4/pdf/scientificamerican07231904-68a.pdf.

Anonymous:1904:IIbu

[7226] Anonymous. Index of inventions. Scientific American, 91(4):68–71, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n4/pdf/scientificamerican07231904-68b.pdf.

Anonymous:1904:PNPbo

[7227] Anonymous. Patent notes: Patents. Scientific American, 91(4):71, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n4/pdf/scientificamerican07231904-71.pdf.

Anonymous:1904:WWI

[7228] Anonymous. What we inherit with the Panama Canal. Scientific American, 91(4):58, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
58.pdf.

Anonymous:1904:CEH

[7229] Anonymous. Cost of electric heating. Scientific American, 91(4):58–59, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n4/pdf/scientificamerican07231904-58a.pdf.

Anonymous:1904:SFC

[7230] Anonymous. Study of forest conditions. Scientific American, 91(4):59, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n4/pdf/scientificamerican07231904-59.pdf.

Anonymous:1904:LDN

[7231] Anonymous. Long-distance non-stop railroad runs in Great Britain — a new world’s record. Scientific American, 91(4):59, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n4/pdf/scientificamerican07231904-59a.pdf.

Anonymous:1904:ACA

[7232] Anonymous. Automatic coupling for air-brake hose. Scientific American, 91(4):60, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
60.pdf.

Anonymous:1904:IAU

[7233] Anonymous. An ingenious adjustable universal coupling. Scientific American, 91(4):60, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
60a.pdf.

Fitzgerald:1904:BFM

[7234] W. G. Fitzgerald. The British fireman’s mask and smoke jacket. Scientific American, 91(4):61, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
61.pdf.

Anonymous:1904:RBI

[7235] Anonymous. Registration balloons in Italy. Scientific American, 91(4):61–62, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
61a.pdf.

Anonymous:1904:SWR

[7236] Anonymous. Sir William Ramsay. Scientific American, 91(4):62–63, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n4/pdf/scientificamerican07231904-62.pdf.

Glasgow:1904:NFB

[7237] C. Field Glasgow and Lieutenant-Colonel. The new French battleship “Démocratie”. Scientific American, 91(4):64, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n4/pdf/scientificamerican07231904-64.pdf.

Anonymous:1904:GPE

[7238] Anonymous. The Government Philippine Exposition. Scientific American, 91(4):64–66, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
64a.pdf.

Beasley:1904:FQN

[7239] Walter L. Beasley. The flamingo and its queer nest. Scientific American, 91(4):66–67, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
66.pdf.

Warnock:1904:GRG

[7240] James A. Warnock. The Great Rotorua Geyser of New Zealand. Scientific American, 91(4):67, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
67.pdf.

Anonymous:1904:BPWbv

[7241] Anonymous. Business and personal wants. Scientific American, 91(4):68, July 23, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n4/pdf/scientificamerican07231904-
68.pdf.

Russell:1904:HAb

[7242] Henry Norris Russell. The heavens in August. Scientific American, 91(5):75, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
75.pdf.

Anonymous:1904:Cbq

[7243] Anonymous. Correspondence. Scientific American, 91(5):78, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n5/pdf/scientificamerican07301904-78.pdf.

Anonymous:1904:RPIbu

[7244] Anonymous. Recently patented inventions. Scientific American, 91(5):84, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
84.pdf.

Anonymous:1904:NQbr

[7245] Anonymous. Notes and queries. Scientific American, 91(5):85, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n5/pdf/scientificamerican07301904-85.pdf.

Anonymous:1904:BNBbg

[7246] Anonymous. Books: New books, etc. Scientific American, 91(5):85, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n5/pdf/scientificamerican07301904-85a.pdf.

Anonymous:1904:IIbv

[7247] Anonymous. Index of inventions. Scientific American, 91(5):85–87, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n5/pdf/scientificamerican07301904-85b.pdf.

Anonymous:1904:PNPbp

[7248] Anonymous. Patent notes: Patents. Scientific American, 91(5):87, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n5/pdf/scientificamerican07301904-87.pdf.

Anonymous:1904:RSU

[7249] Anonymous. Railroads in Sweden to use electric system. Scientific American, 91(5):74, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
74.pdf.

Anonymous:1904:REC

[7250] Anonymous. Recent excavations at Carthage. Scientific American, 91(5):74, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
74a.pdf.

Anonymous:1904:WOC

[7251] Anonymous. Wave-operated clocks at Paris. Scientific American, 91(5):74, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
74b.pdf.

Anonymous:1904:GSS

[7252] Anonymous. The great size of the St. Louis Exposition. Scientific American, 91(5):74, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
74c.pdf.

Anonymous:1904:NCP

[7253] Anonymous. Need of cotton-picking machines. Scientific American, 91(5):74–75, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
74d.pdf.

Anonymous:1904:NYC

[7254] Anonymous. New York Central Express compound locomotive. Scientific American, 91(5):76, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
76.pdf.

Anonymous:1904:ASS

[7255] Anonymous. Automobile street sprinkler. Scientific American, 91(5):77, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n5/pdf/scientificamerican07301904-77.pdf.

Anonymous:1904:PMSa

[7256] Anonymous. Planing machine for ships’ decks. Scientific American, 91(5):77, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
77a.pdf.

Anonymous:1904:BDL

[7257] Anonymous. The Berger disinfecting lamp. Scientific American, 91(5):78, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
78a.pdf.

Anonymous:1904:WLP

[7258] Anonymous. What the loss of the Port Arthur Fleet would mean to Russia. Scientific American, 91(5):79, July 30, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n5/pdf/scientificamerican07301904-79.pdf.

Dailey:1904:EII

[7259] Howard B. Dailey. Electrostatic illuminations: Interesting experiments for the induction machine. Scientific American, 91(5):80, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n5/pdf/scientificamerican07301904-80.pdf.

Willey:1904:WWC

[7260] Day Allen Willey. The work of a western cyclone. Scientific American, 91(5):81–82, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
81.pdf.

Anonymous:1904:SHR

[7261] Anonymous. A section of the Hudson River Tunnel at the World’s Fair. Scientific American, 91(5):82, July 30, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n5/pdf/scientificamerican07301904-82.pdf.

Hutchinson:1904:TS

[7262] C. E. Hutchinson. A trapdoor spider. Scientific American, 91(5):83, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n5/pdf/scientificamerican07301904-83.pdf.

Anonymous:1904:SNAa

[7263] Anonymous. Some new additions to the United States Zoological Park. Scientific American, 91(5):83, July 30, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n5/pdf/scientificamerican07301904-83a.pdf.

Anonymous:1904:BPWbw

[7264] Anonymous. Business and personal wants. Scientific American, 91(5):84–85, July 30, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n5/pdf/scientificamerican07301904-
84a.pdf.

Anonymous:1904:SNbf

[7265] Anonymous. Science notes. Scientific American, 91(6):91, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n6/pdf/scientificamerican08061904-91b.pdf.

Anonymous:1904:Cbr

[7266] Anonymous. Correspondence. Scientific American, 91(6):95, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n6/pdf/scientificamerican08061904-95.pdf.

Anonymous:1904:RPIbv

[7267] Anonymous. Recently patented inventions. Scientific American, 91(6):100, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
100.pdf.

Anonymous:1904:NQbs

[7268] Anonymous. Notes and queries. Scientific American, 91(6):100, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n6/pdf/scientificamerican08061904-100b.pdf.

Anonymous:1904:IIbw

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n6/pdf/scientificamerican08061904-100c.pdf.

Anonymous:1904:PNPbq

[7270] Anonymous. Patent notes: Patents. Scientific American, 91(6):103, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n6/pdf/scientificamerican08061904-103.pdf.

Anonymous:1904:EEF

[7271] Anonymous. The Exposition as an educational force. Scientific American, 91(6):90, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
90.pdf.

Anonymous:1904:ONC

[7272] Anonymous. Osmon, a new combustible from peat. Scientific American, 91(6):90, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
90a.pdf.

Anonymous:1904:MCE

[7273] Anonymous. M. Curie’s experiments with radium emanations. Scientific American, 91(6):90, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
90b.pdf.

Anonymous:1904:TSP

[7274] Anonymous. Tropical specialties for Porto Rico. Scientific American, 91(6):90–91, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
90c.pdf.

Anonymous:1904:IMU

[7275] Anonymous. The influence of milking upon the quantity and quality of milk. Scientific American, 91(6):91, August 6, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n6/pdf/scientificamerican08061904-91.pdf.

Maury:1904:PCA

[7276] A. C. Maury. Phosphorescent crystallization of arsenious acid. Scientific American, 91(6):91, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
91a.pdf.

Anonymous:1904:LTP

[7277] Anonymous. The locomotive-testing plant at the World’s Fair. Scientific American, 91(6):92, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
92.pdf.

Plummer:1904:CCT

[7278] John Plummer. Counterweighted cable tramway. Scientific American, 91(6):93, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
93.pdf.

Anonymous:1904:NWM

[7279] Anonymous. A novel wave meter for wireless telegraphy. Scientific American, 91(6):93–94, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
93a.pdf.

Anonymous:1904:HC

[7280] Anonymous. A handless clock. Scientific American, 91(6):94, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n6/pdf/scientificamerican08061904-94.pdf.

Baden-Powell:1904:AE

[7281] Major B. Baden-Powell. Aeroplane experiments. Scientific American, 91(6):96, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
96.pdf.

Anonymous:1904:TAW

[7282] Anonymous. Trained animals at the World’s Fair. Scientific American, 91(6):97–98, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
97.pdf.

Anonymous:1904:BAb

[7283] Anonymous. The bee as an artisan. Scientific American, 91(6):98, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n6/pdf/scientificamerican08061904-98.pdf.

Chambers:1904:TON

[7284] R. E. Chambers and Rev. Transportation old and new. Scientific American, 91(6):99, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
99.pdf.

Anonymous:1904:BPWbx

[7285] Anonymous. Business and personal wants. Scientific American, 91(6):100, August 6, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n6/pdf/scientificamerican08061904-
100a.pdf.

Anonymous:1904:Cbs

[7286] Anonymous. Correspondence. Scientific American, 91(7):111, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n7/pdf/scientificamerican08131904-111.pdf.

Anonymous:1904:RPIbw

[7287] Anonymous. Recently patented inventions. Scientific American, 91(7):116, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
116.pdf.

Anonymous:1904:NQbt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n7/pdf/scientificamerican08131904-117.pdf.

Anonymous:1904:BNBbh

[7289] Anonymous. Books: New books, etc. Scientific American, 91 (7):117, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
117a.pdf.

Anonymous:1904:IIbx

[7290] Anonymous. Index of inventions. Scientific American, 91(7): 117–119, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
117b.pdf.

Anonymous:1904:PNPbr

[7291] Anonymous. Patent notes: Patents. Scientific American, 91(7):119, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n7/pdf/scientificamerican08131904-119.pdf.

Anonymous:1904:FLF

[7292] Anonymous. Four leading foreign exhibits at the World’s Fair. Scientific American, 91(7):106, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
106.pdf.

Anonymous:1904:PPR

[7293] Anonymous. The penetrating power of “N_{1}-Rays”. Scientific American, 91(7):106–107, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
106a.pdf.

Anonymous:1904:PRF

[7294] Anonymous. The position of the range finder on a warship. Scientific American, 91(7):107, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
107.pdf.

Anonymous:1904:RNB

[7295] Anonymous. Renard’s new boiler for airships. Scientific American, 91(7):107, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
107a.pdf.

Anonymous:1904:NC

[7296] Anonymous. A new Cunarder. Scientific American, 91(7):107, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n7/pdf/scientificamerican08131904-107b.pdf.

Anonymous:1904:NGC

[7297] Anonymous. The National Geographic Congress. Scientific American, 91(7):107, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
107c.pdf.

Anonymous:1904:ATW

[7298] Anonymous. The automobile tour to the World’s Fair. Scientific American, 91(7):107, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
107d.pdf.

Fawcett:1904:USM

[7299] Waldon Fawcett. The United States Marine Corps. Scientific American, 91(7):108, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
108.pdf.

Perkins:1904:MVL

[7300] Frank C. Perkins. The mercury vapor lamp for photographic work. Scientific American, 91(7):108, August 13, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n7/pdf/scientificamerican08131904-108a.pdf.

Shaw:1904:USV

[7301] A. F. Shaw. The utilization of snake venom in preparing antibubonic serum. Scientific American, 91(7):109–110, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n7/pdf/scientificamerican08131904-109.pdf.

Guarini:1904:MPR

[7302] Emile Guarini. The Mount Pilatus Railway, Switzerland. Scientific American, 91(7):110–111, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
110.pdf.

Anonymous:1904:STB

[7303] Anonymous. Speed trial of the battleship “Ohio”. Scientific American, 91(7):112, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
112.pdf.

Inkersley:1904:SSF

[7304] Arthur Inkersley. Strange sights in the Far North. Scientific American, 91(7):112–114, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
112a.pdf.

Anonymous:1904:CSPb

[7305] Anonymous. Cabinet for storing phonograph records. Scientific American, 91(7):114, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
114.pdf.

Anonymous:1904:MB

[7306] Anonymous. Medical battery. Scientific American, 91(7):114, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n7/pdf/scientificamerican08131904-114a.pdf.

Anonymous:1904:MAR

[7307] Anonymous. A marvelous automobile record — 1031/2miles an hour. Scientific American, 91(7):115, August 13, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n7/pdf/scientificamerican08131904-115.pdf.

Anonymous:1904:OKI

[7308] Anonymous. The oldest known inhabitant of the globe. Scientific American, 91(7):115, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
115a.pdf.

Anonymous:1904:BPWby

[7309] Anonymous. Business and personal wants. Scientific American, 91(7):116, August 13, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n7/pdf/scientificamerican08131904-
116a.pdf.

Anonymous:1904:Cbt

[7310] Anonymous. Correspondence. Scientific American, 91(8):126, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n8/pdf/scientificamerican08201904-126a.pdf.

Anonymous:1904:RPIbx

[7311] Anonymous. Recently patented inventions. Scientific American, 91(8):132, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
132.pdf.

Anonymous:1904:NQbu

[7312] Anonymous. Notes and queries. Scientific American, 91(8):132, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n8/pdf/scientificamerican08201904-132b.pdf.

Anonymous:1904:BNBbi (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
132c.pdf.

Anonymous:1904:IIby

[7314] Anonymous. Index of inventions. Scientific American, 91(8): 132–135, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
132d.pdf.

Anonymous:1904:PNPbs

[7315] Anonymous. Patent notes: Patents. Scientific American, 91(8):135, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n8/pdf/scientificamerican08201904-135.pdf.

Anonymous:1904:JDL

[7316] Anonymous. Japanese daring on land and caution on the sea. Scientific American, 91(8):122, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
122.pdf.

Anonymous:1904:RAI

[7317] Anonymous. Radio-activity induced by the emanations of incandescent metal wires. Scientific American, 91(8):122, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n8/pdf/scientificamerican08201904-122a.pdf.

Anonymous:1904:DBF

[7318] Anonymous. Deductions from the Baltimore Fire. Scientific American, 91(8):122–123, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
122b.pdf.

Anonymous:1904:EA

[7319] Anonymous. Electricity in agriculture. Scientific American, 91 (8):123, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
123.pdf.

Anonymous:1904:PEA

[7320] Anonymous. A plea for the endowment of astronomical research. Scientific American, 91(8):123, August 20, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n8/pdf/scientificamerican08201904-123a.pdf.

Lawson:1904:MDU

[7321] William A. Lawson. Mammoth dredger used for reclamation work in California. Scientific American, 91(8):124, August 20, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n8/pdf/scientificamerican08201904-124.pdf.

Guarini:1904:IIM

[7322] Emile Guarini. An instrument for indicating mean astronomical noon. Scientific American, 91(8):125, August 20, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n8/pdf/scientificamerican08201904-125.pdf.

Anonymous:1904:NSTc

[7323] Anonymous. New steam turbine. Scientific American, 91(8):125, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n8/pdf/scientificamerican08201904-125a.pdf.

Perkins:1904:NGT

[7324] Frank C. Perkins. A new German tower electric crane. Scientific American, 91(8):125–126, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
125b.pdf.

Anonymous:1904:PMSb

[7325] Anonymous. Pumping machinery at the St. Louis Fair. Scientific American, 91(8):126, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
126.pdf.

Proctor:1904:ES

[7326] Mary Proctor. The 1905 eclipse of the Sun. Scientific American, 91(8):127, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
127.pdf.

Anonymous:1904:ATF

[7327] Anonymous. Admiral Togo’s flagship “Mikasa”. Scientific American, 91(8):128–130, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
128.pdf.

Anonymous:1904:I

[7328] Anonymous. The iguana. Scientific American, 91(8):130, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n8/pdf/scientificamerican08201904-130.pdf.

Geare:1904:RIA

[7329] R. I. Geare. The reindeer industry in Alaska. Scientific American, 91(8):131, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
131.pdf.

Anonymous:1904:BPWbz

[7330] Anonymous. Business and personal wants. Scientific American, 91(8):132, August 20, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n8/pdf/scientificamerican08201904-
132a.pdf.

Anonymous:1904:Cbu

[7331] Anonymous. Correspondence. Scientific American, 91(9):143, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n9/pdf/scientificamerican08271904-143.pdf.

Anonymous:1904:RPIby

[7332] Anonymous. Recently patented inventions. Scientific American, 91(9):148, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
148.pdf.

Anonymous:1904:NQbv

[7333] Anonymous. Notes and queries. Scientific American, 91(9):148, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n9/pdf/scientificamerican08271904-148b.pdf.

Anonymous:1904:IIbz

[7334] Anonymous. Index of inventions. Scientific American, 91(9): 148–151, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
148c.pdf.

Anonymous:1904:PNPbt

[7335] Anonymous. Patent notes: Patents. Scientific American, 91(9):151, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n9/pdf/scientificamerican08271904-151.pdf.

Anonymous:1904:SAS

[7336] Anonymous. The Scientific American and the St. Louis Fair. Scientific American, 91(9):138, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
138.pdf.

Anonymous:1904:RSH

[7337] Anonymous. Railway speeds here and in Europe. Scientific American, 91(9):138, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
138a.pdf.

Anonymous:1904:GRS

[7338] Anonymous. A great railway scheme. Scientific American, 91 (9):138, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
138b.pdf.

Anonymous:1904:RL

[7339] Anonymous. The romance of light. Scientific American, 91(9): 138–139, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
138c.pdf.

Anonymous:1904:GMM

[7340] Anonymous. Glass models of microscopic forms of life — an interesting art. Scientific American, 91(9):140–141, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n9/pdf/scientificamerican08271904-140.pdf.

Anonymous:1904:NYSc

[7341] Anonymous. The New York Subway instruction car. Scientific American, 91(9):141, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
141.pdf.

Byers:1904:SDC

[7342] Charles Alma Byers. Side-door coaches for World’s Fair service. Scientific American, 91(9):141–142, August 27, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n9/pdf/scientificamerican08271904-141a.pdf.

Anonymous:1904:ICM

[7343] Anonymous. Ice-cutting machine. Scientific American, 91(9):142, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v91/n9/pdf/scientificamerican08271904-142.pdf.

Anonymous:1904:DPP

[7344] Anonymous. The direct production of photographs in colors by a new method. Scientific American, 91(9):142, August 27, 1904. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v91/
n9/pdf/scientificamerican08271904-142a.pdf.

Anonymous:1904:WDE

[7345] Anonymous. War Department Exhibit at the St. Louis Fair. Scientific American, 91(9):144–145, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
144.pdf.

Anonymous:1904:DVS

[7346] Anonymous. The defeat of the Vladivostock squadron. Scientific American, 91(9):146–147, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
146.pdf.

Baltimore:1904:NBA

[7347] J. Mayne Baltimore. The new Baldwin airship. Scientific American, 91(9):147–148, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
147.pdf.

Anonymous:1904:BPWca

[7348] Anonymous. Business and personal wants. Scientific American, 91(9):148, August 27, 1904. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v91/n9/pdf/scientificamerican08271904-
148a.pdf.

Anonymous:1905:RPIa

[7349] Anonymous. Recently patented inventions. Scientific American, 92(1):12, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
12.pdf.

Anonymous:1905:NQa

[7350] Anonymous. Notes and queries. Scientific American, 92(1):12, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n1/pdf/scientificamerican01071905-12b.pdf.

Anonymous:1905:IIa

[7351] Anonymous. Index of inventions. Scientific American, 92(1):12–15, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n1/pdf/scientificamerican01071905-12c.pdf.

Anonymous:1905:PNPa

[7352] Anonymous. Patent notes: Patents. Scientific American, 92(1):15, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n1/pdf/scientificamerican01071905-15.pdf.

Anonymous:1905:RYa

[7353] Anonymous. Retrospect of the year 1904. Scientific American, 92(1):2–3, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
2.pdf.

Anonymous:1905:GTS

[7354] Anonymous. A gearless two-speed bicycle. Scientific American, 92(1):4, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n1/pdf/scientificamerican01071905-4.pdf.

Anonymous:1905:BCG

[7355] Anonymous. A bad case of gun erosion. Scientific American, 92(1):4, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n1/pdf/scientificamerican01071905-4a.pdf.

Anonymous:1905:NPLa

[7356] Anonymous. New photographic lamp. Scientific American, 92(1):4, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n1/pdf/scientificamerican01071905-4b.pdf.

Mclure:1905:MM

[7357] W. Frank M’clure. The manufacture of macaroni. Scientific American, 92(1):5, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
5.pdf.

Anonymous:1905:LWT

[7358] Anonymous. The largest water turbine in existence. Scientific American, 92(1):6, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
6.pdf.

Haupt:1905:NYH

[7359] Lewis M. Haupt. The New York Harbor entrance. Scientific American, 92(1):7, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
7.pdf.

Guarini:1905:TDS

[7360] Emile Guarini. The turbine-driven steamer “Manxman”. Scientific American, 92(1):8, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
8.pdf.

Brown:1905:PC

[7361] Enos Brown. The prune and its cultivation. Scientific American, 92(1):9–10, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
9.pdf.

Anonymous:1905:LLR

[7362] Anonymous. The largest lobster on record. Scientific American, 92(1):10, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
10.pdf.

Beard:1905:QC

[7363] J. Carter Beard. The queerest of creatures — I. Scientific American, 92(1):11, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
11.pdf.

Anonymous:1905:BPWa

[7364] Anonymous. Business and personal wants. Scientific American, 92(1):12, January 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n1/pdf/scientificamerican01071905-
12a.pdf.

Anonymous:1905:Ca

[7365] Anonymous. Correspondence. Scientific American, 92(10):203, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-203.pdf.

Anonymous:1905:RPIb

[7366] Anonymous. Recently patented inventions. Scientific American, 92(10):209, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
209.pdf.

Anonymous:1905:NQb

[7367] Anonymous. Notes and queries. Scientific American, 92(10):210, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-210.pdf.

Anonymous:1905:BNBa

[7368] Anonymous. Books: New books, etc. Scientific American, 92(10):210, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-210a.pdf.

Anonymous:1905:IIb

[7369] Anonymous. Index of inventions. Scientific American, 92(10):210–215, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-210b.pdf.

Anonymous:1905:PNPb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-215.pdf.

Anonymous:1905:PAM

[7371] Anonymous. A Port Arthur mystery explained. Scientific American, 92(10):198, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
198.pdf.

Anonymous:1905:PML

[7372] Anonymous. Proposed Manhattan loop for Brooklyn Elevated Roads. Scientific American, 92(10):198, March 11, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n10/pdf/scientificamerican03111905-198a.pdf.

Anonymous:1905:NET

[7373] Anonymous. New experiments in turpentining. Scientific American, 92(10):198, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
198b.pdf.

Anonymous:1905:WPBa

[7374] Anonymous. Wind pressure on bridges. Scientific American, 92(10):198, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-198c.pdf.

Anonymous:1905:SFV

[7375] Anonymous. Studies of the food value of fruit. Scientific American, 92(10):198–199, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
198d.pdf.

Erichsen:1905:DD

[7376] Hugo Erichsen. The dangers of dining. Scientific American, 92(10):199, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-199.pdf.

Anonymous:1905:NCF

[7377] Anonymous. New compound formed in electric furnace. Scientific American, 92(10):199, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
199a.pdf.

Anonymous:1905:NSS

[7378] Anonymous. New Sun spots. Scientific American, 92(10):199, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-199b.pdf.

Weeks:1905:RRH

[7379] Arthur B. Weeks. A rolling road for horses and wagons. Scientific American, 92(10):200, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
200.pdf.

Anonymous:1905:CED

[7380] Anonymous. Curious effects of damage by ice. Scientific American, 92(10):200, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
200a.pdf.

Anonymous:1905:TM

[7381] Anonymous. A typewriter a minute. Scientific American, 92 (10):201–202, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
201.pdf.

Anonymous:1905:MIT

[7382] Anonymous. A magnetic indicator for tanks. Scientific American, 92(10):203, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
203a.pdf.

Anonymous:1905:NCLa

[7383] Anonymous. The new Cunard liner “Caronia”. Scientific American, 92(10):204–205, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
204.pdf.

Anonymous:1905:CFS

[7384] Anonymous. Chicago’s freight subways. Scientific American, 92 (10):205–206, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
205.pdf.

Anonymous:1905:SORa

[7385] Anonymous. Story of the operations of the Russian Fleet at Pt, Arthur as told by one of the commanders. Scientific American, 92(10):206–208, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n10/pdf/scientificamerican03111905-206.pdf.

Anonymous:1905:BPWb

[7386] Anonymous. Business and personal wants. Scientific American, 92(10):209, March 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n10/pdf/scientificamerican03111905-
209a.pdf.

Anonymous:1905:Cb

[7387] Anonymous. Correspondence. Scientific American, 92(11):223, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n11/pdf/scientificamerican03181905-223.pdf.

Anonymous:1905:RPIc

[7388] Anonymous. Recently patented inventions. Scientific American, 92(11):229, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
229.pdf.

Anonymous:1905:NQc

[7389] Anonymous. Notes and queries. Scientific American, 92(11):230, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n11/pdf/scientificamerican03181905-230.pdf.

Anonymous:1905:BNBb

[7390] Anonymous. Books: New books, etc. Scientific American, 92 (11):230–231, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
230a.pdf.

Anonymous:1905:IIc

[7391] Anonymous. Index of inventions. Scientific American, 92(11):231–235, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n11/pdf/scientificamerican03181905-231.pdf.

Anonymous:1905:PNPc

[7392] Anonymous. Patent notes: Patents. Scientific American, 92(11):235, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n11/pdf/scientificamerican03181905-235.pdf.

Anonymous:1905:LEN

[7393] Anonymous. Lake Erie and the New York Water Supply. Scientific American, 92(11):218, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
218.pdf.

Anonymous:1905:TB

[7394] Anonymous. The tempering of bronze. Scientific American, 92(11):218, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n11/pdf/scientificamerican03181905-218a.pdf.

Anonymous:1905:STL

[7395] Anonymous. Suburban traffic in London. Scientific American, 92(11):218, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
218b.pdf.

Anonymous:1905:BSI

[7396] Anonymous. Bacterial soil inoculation for vegetables. Scientific American, 92(11):218–219, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
218c.pdf.

Anonymous:1905:IDP

[7397] Anonymous. An important discovery in the purification of contaminated water. Scientific American, 92(11):219, March 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n11/pdf/scientificamerican03181905-219.pdf.

Sabine:1905:WGC

[7398] Wallace C. Sabine. The Whispering Gallery in the Capitol at Washington. Scientific American, 92(11):219, March 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n11/pdf/scientificamerican03181905-219a.pdf.

Anonymous:1905:FAC

[7399] Anonymous. The French aeroplane contest. Scientific American, 92(11):220, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
220.pdf.

Anonymous:1905:IAA

[7400] Anonymous. Inoculating animals against disease. Scientific American, 92(11):221, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
221.pdf.

Brassler:1905:ACC

[7401] Charles A. Brassler. The astronomical clock of the City Hall of Ulm. Scientific American, 92(11):221–222, March 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n11/pdf/scientificamerican03181905-221a.pdf.

Curtiss:1905:SBB

[7402] Edward Whitehead Curtiss. A study of the Brooklyn Bridge problem. Scientific American, 92(11):222–223, March 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n11/pdf/scientificamerican03181905-222.pdf.

Collins:1905:TME

[7403] A. Frederick Collins. Transportation of materials by electricity. Scientific American, 92(11):224–226, March 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n11/pdf/scientificamerican03181905-224.pdf.

King:1905:CST

[7404] Charles R. King. The completion of the simplon tunnel. Scientific American, 92(11):226, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
226.pdf.

Anonymous:1905:PAA

[7405] Anonymous. Port Arthur after the capitulation. Scientific American, 92(11):227–228, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
227.pdf.

Anonymous:1905:NTR

[7406] Anonymous. A new type of rail joint. Scientific American, 92(11):228, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n11/pdf/scientificamerican03181905-228.pdf.

Anonymous:1905:IBH

[7407] Anonymous. Improved backing harness for teams. Scientific American, 92(11):228, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
228a.pdf.

Anonymous:1905:REB

[7408] Anonymous. Reamer for enlarging blast holes. Scientific American, 92(11):228, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
228b.pdf.

Anonymous:1905:BPWc

[7409] Anonymous. Business and personal wants. Scientific American, 92(11):229, March 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n11/pdf/scientificamerican03181905-
229a.pdf.

Anonymous:1905:Cc

[7410] Anonymous. Correspondence. Scientific American, 92(12):243, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n12/pdf/scientificamerican03251905-243b.pdf.

Anonymous:1905:RPId

[7411] Anonymous. Recently patented inventions. Scientific American, 92(12):248, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
248.pdf.

Anonymous:1905:NQd

[7412] Anonymous. Notes and queries. Scientific American, 92(12):248–250, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n12/pdf/scientificamerican03251905-248b.pdf.

Anonymous:1905:BNBc

[7413] Anonymous. Books: New books, etc. Scientific American, 92 (12):250–251, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
250.pdf.

Anonymous:1905:IId

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n12/pdf/scientificamerican03251905-251.pdf.

Anonymous:1905:JLS

[7415] Anonymous. Japan’s losses on the sea. Scientific American, 92(12):238, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n12/pdf/scientificamerican03251905-238.pdf.

Anonymous:1905:SLC

[7416] Anonymous. A sea-level canal at Panama. Scientific American, 92(12):238, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
238a.pdf.

Anonymous:1905:TCP

[7417] Anonymous. The tides at Colon and Panama. Scientific American, 92(12):238–239, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
238b.pdf.

Anonymous:1905:EAR

[7418] Anonymous. An electrical analogue for respiration. Scientific American, 92(12):239, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
239.pdf.

Anonymous:1905:BNP

[7419] Anonymous. The British Naval Programme for 1905. Scientific American, 92(12):239, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
239a.pdf.

Anonymous:1905:SNa

[7420] Anonymous. Science notes. Scientific American, 92(12):239, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n12/pdf/scientificamerican03251905-239b.pdf.

Gradenwitz:1905:NPU

[7421] Alfred Gradenwitz. A novel process for using blue light as an anæsthetic. Scientific American, 92(12):240, March 25, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n12/pdf/scientificamerican03251905-240.pdf.

Anonymous:1905:NBB

[7422] Anonymous. The new British battleship “King Edward VII”. Scientific American, 92(12):240, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
240a.pdf.

Willey:1905:SSS

[7423] Day Allen Willey. Striking snow statuary. Scientific American, 92(12):241, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
241.pdf.

Anonymous:1905:NYH

[7424] Anonymous. The New York Hippodrome. Scientific American, 92 (12):241–242, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
241a.pdf.

Anonymous:1905:RT

[7425] Anonymous. Registering thermometer. Scientific American, 92(12):243, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n12/pdf/scientificamerican03251905-243.pdf.

Anonymous:1905:NCLb

[7426] Anonymous. A new cadmium lamp. Scientific American, 92(12):243, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n12/pdf/scientificamerican03251905-243a.pdf.

Anonymous:1905:TMMa

[7427] Anonymous. The Tombs at Mukden of the Ming Dynasty. Scientific American, 92(12):244–245, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
244.pdf.

Hayward:1905:PGW

[7428] Charles B. Hayward. Powerful German windmills. Scientific American, 92(12):245–246, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
245.pdf.

Anonymous:1905:MFA

[7429] Anonymous. Multi-firearms of ancient times. Scientific American, 92(12):246–247, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
246.pdf.

Anonymous:1905:BPWd

[7430] Anonymous. Business and personal wants. Scientific American, 92(12):248, March 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n12/pdf/scientificamerican03251905-
248a.pdf.

Russell:1905:HAa

[7431] Henry Norris Russell. The heavens in April. Scientific American, 92(13):259, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
259.pdf.

Anonymous:1905:Cd

[7432] Anonymous. Correspondence. Scientific American, 92(13):262–263, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-262.pdf.

Anonymous:1905:RPIe (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
269.pdf.

Anonymous:1905:NQe

[7434] Anonymous. Notes and queries. Scientific American, 92(13):269–270, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-269b.pdf.

Anonymous:1905:BNBd

[7435] Anonymous. Books: New books, etc. Scientific American, 92 (13):270–271, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
270.pdf.

Anonymous:1905:IIe

[7436] Anonymous. Index of inventions. Scientific American, 92(13):271–275, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-271.pdf.

Anonymous:1905:PNPd

[7437] Anonymous. Patent notes: Patents. Scientific American, 92(13):275, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-275.pdf.

Anonymous:1905:CMC

[7438] Anonymous. Casualties — military and civil. Scientific American, 92(13):258, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
258.pdf.

Anonymous:1905:CSH

[7439] Anonymous. “Consigned to the scrap heap”. Scientific American, 92(13):258, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
258a.pdf.

Anonymous:1905:ERP

[7440] Anonymous. Electric railways and population. Scientific American, 92(13):258, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
258b.pdf.

Hopkins:1905:GWU

[7441] N. Monroe Hopkins. The George Washington University. Scientific American, 92(13):258–259, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
258c.pdf.

Anonymous:1905:SNb

[7442] Anonymous. Science notes. Scientific American, 92(13):259, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-259a.pdf.

Anonymous:1905:CSa

[7443] Anonymous. The current supplement. Scientific American, 92(13):259, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-259b.pdf.

Anonymous:1905:CWV

[7444] Anonymous. Cables that withstand 90,000 volts. Scientific American, 92(13):260, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
260.pdf.

Anonymous:1905:PASa

[7445] Anonymous. The progress of aviation since 1891. Scientific American, 92(13):260, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
260a.pdf.

Gerald:1905:SSC

[7446] W. G. Fitz Gerald. A swimming school for cavalry horses. Scientific American, 92(13):261, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
261.pdf.

Anonymous:1905:SRD

[7447] Anonymous. A simple rim for detachable tires. Scientific American, 92(13):261–262, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
261a.pdf.

Collins:1905:HMW

[7448] A. Frederick Collins. A home-made wireless signaling transmitter. Scientific American, 92(13):263, April 1, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n13/pdf/scientificamerican04011905-263.pdf.

Brown:1905:SMS

[7449] Enos Brown. Salt making on San Francisco Bay. Scientific American, 92(13):264–266, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
264.pdf.

Guarini:1905:BDS

[7450] Emile Guarini. The Barossa Dam, Southern Australia. Scientific American, 92(13):266, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
266.pdf.

Anonymous:1905:NBI

[7451] Anonymous. The new Blackwell’s Island Bridge. Scientific American, 92(13):266, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
266a.pdf.

Anonymous:1905:PDa

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-267.pdf.

Anonymous:1905:LS

[7453] Anonymous. Lineman’s support. Scientific American, 92(13):267, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-267a.pdf.

Anonymous:1905:ICF

[7454] Anonymous. An ink-controlled fountain drawing and writing pen. Scientific American, 92(13):267, April 1, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n13/pdf/scientificamerican04011905-267b.pdf.

Anonymous:1905:RJC

[7455] Anonymous. Rail-joint connection. Scientific American, 92(13):267, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-267c.pdf.

Anonymous:1905:HC

[7456] Anonymous. Hose coupling. Scientific American, 92(13):267–268, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-267d.pdf.

Anonymous:1905:RG

[7457] Anonymous. Railway gage. Scientific American, 92(13):268, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-268.pdf.

Anonymous:1905:OIa

[7458] Anonymous. Oddities in invention. Scientific American, 92(13):268, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-268a.pdf.

Anonymous:1905:NTMa

[7459] Anonymous. The new trade mark law. Scientific American, 92(13):268, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n13/pdf/scientificamerican04011905-268b.pdf.

Anonymous:1905:BPWe

[7460] Anonymous. Business and personal wants. Scientific American, 92(13):269, April 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n13/pdf/scientificamerican04011905-
269a.pdf.

Anonymous:1905:Ce

[7461] Anonymous. Correspondence. Scientific American, 92(14):286, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n14/pdf/scientificamerican04081905-286.pdf.

Anonymous:1905:RPIf

[7462] Anonymous. Recently patented inventions. Scientific American, 92(14):289, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
289.pdf.

Anonymous:1905:NQf

[7463] Anonymous. Notes and queries. Scientific American, 92(14):289–290, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n14/pdf/scientificamerican04081905-289b.pdf.

Anonymous:1905:BNBe

[7464] Anonymous. Books: New books, etc. Scientific American, 92 (14):290–291, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
290.pdf.

Anonymous:1905:IIf

[7465] Anonymous. Index of inventions. Scientific American, 92(14):291–295, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n14/pdf/scientificamerican04081905-291.pdf.

Anonymous:1905:PNPe

[7466] Anonymous. Patent notes: Patents. Scientific American, 92(14):295, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n14/pdf/scientificamerican04081905-295.pdf.

Anonymous:1905:ORG

[7467] Anonymous. The ocean race for the German Emperor’s Cup. Scientific American, 92(14):278, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
278.pdf.

Anonymous:1905:DDT

[7468] Anonymous. The dangers and difficulties of tunnel boring. Scientific American, 92(14):278, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
278a.pdf.

Anonymous:1905:ESR

[7469] Anonymous. Electric street railroads versus motor-propelled vehicles for passenger traffic. Scientific American, 92(14):278–279, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n14/pdf/scientificamerican04081905-278b.pdf.

Erichsen:1905:AL

[7470] Hugo Erichsen. An antitoxin for laziness. Scientific American, 92(14):279, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
279.pdf.

Anonymous:1905:TVR

[7471] Anonymous. Turbines vs. reciprocating engines in marine service. Scientific American, 92(14):279, April 8, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n14/pdf/scientificamerican04081905-279a.pdf.

Willey:1905:LPV

[7472] Day Allen Willey. The lofty Pecos Viaduct on the Southern Pacific Railway. Scientific American, 92(14):280, April 8, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n14/pdf/scientificamerican04081905-280.pdf.

Anonymous:1905:LB

[7473] Anonymous. The largest of bears. Scientific American, 92(14):280–281, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n14/pdf/scientificamerican04081905-280a.pdf.

Anonymous:1905:VFT

[7474] Anonymous. The “Victorian” the first turbine Atlantic liner. Scientific American, 92(14):281–282, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
281.pdf.

Chandler:1905:MC

[7475] Edward F. Chandler. Miniature cameras. Scientific American, 92(14):282, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
282.pdf.

Weeks:1905:HME

[7476] Arthur B. Weeks. How to make an electric soldering iron. Scientific American, 92(14):283, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
283.pdf.

Anonymous:1905:DDS

[7477] Anonymous. The double-deck subway system. Scientific American, 92(14):283–284, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
283a.pdf.

Anonymous:1905:EGE

[7478] Anonymous. The entries for the German Emperor’s Ocean Cup Race. Scientific American, 92(14):285, April 8, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n14/pdf/scientificamerican04081905-285.pdf.

Anonymous:1905:GPD

[7479] Anonymous. The great premier diamond. Scientific American, 92(14):285–286, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
285a.pdf.

Beasley:1905:SRO

[7480] Walter L. Beasley. Some remarkable old Chinese bronzes. Scientific American, 92(14):287–288, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
287.pdf.

Anonymous:1905:BPWf

[7481] Anonymous. Business and personal wants. Scientific American, 92(14):289, April 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n14/pdf/scientificamerican04081905-
289a.pdf.

Anonymous:1905:Cf

[7482] Anonymous. Correspondence. Scientific American, 92(15):303, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-303.pdf.

Anonymous:1905:RPIg

[7483] Anonymous. Recently patented inventions. Scientific American, 92(15):309–310, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
309.pdf.

Anonymous:1905:NQg

[7484] Anonymous. Notes and queries. Scientific American, 92(15):310–311, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-310a.pdf.

Anonymous:1905:BNBf

[7485] Anonymous. Books: New books, etc. Scientific American, 92 (15):311–312, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
311.pdf.

Anonymous:1905:IIg

[7486] Anonymous. Index of inventions. Scientific American, 92(15):312–315, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-312.pdf.

Anonymous:1905:PNPf

[7487] Anonymous. Patent notes: Patents. Scientific American, 92(15):315, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-315.pdf.

Anonymous:1905:VNF

[7488] Anonymous. Vandalism at Niagara Falls. Scientific American, 92(15):298, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
298.pdf.

Anonymous:1905:TPV

[7489] Anonymous. The true point of view. Scientific American, 92(15):298, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-298a.pdf.

Anonymous:1905:FFM

[7490] Anonymous. The fastest 40-foot motor boat. Scientific American, 92(15):298, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
298b.pdf.

Anonymous:1905:CPS

[7491] Anonymous. Complete the present subway plan. Scientific American, 92(15):298, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
298c.pdf.

Ostwald:1905:CPP

[7492] W. Ostwald. The chemistry of painting and the preservation of canvases. Scientific American, 92(15):298–299, April 15, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n15/pdf/scientificamerican04151905-298d.pdf.

Adams:1905:NPG

[7493] Alton D. Adams. Niagara Power at Goat Island. Scientific American, 92(15):299, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
299.pdf.

Anonymous:1905:SNc

[7494] Anonymous. Science notes. Scientific American, 92(15):299, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-299a.pdf.

Anonymous:1905:STM

[7495] Anonymous. Standard time at a modern watch works. Scientific American, 92(15):300–301, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
300.pdf.

Anonymous:1905:ONCa

[7496] Anonymous. The opening of the new Croton Reservoir. Scientific American, 92(15):302, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
302.pdf.

Anonymous:1905:LFB

[7497] Anonymous. Liquid fuel burner. Scientific American, 92(15):303, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-303a.pdf.

Anonymous:1905:AS

[7498] Anonymous. Alcohol from sawdust. Scientific American, 92(15):304, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-304.pdf.

Anonymous:1905:CHa

[7499] Anonymous. A column of horseshoes. Scientific American, 92(15):305, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-305.pdf.

Anonymous:1905:GRW

[7500] Anonymous. The giantess Rosa Wedsted. Scientific American, 92(15):305, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
305a.pdf.

Anonymous:1905:SMS

[7501] Anonymous. A spontaneously-moving stone ball. Scientific American, 92(15):305–307, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
305b.pdf.

Lodian:1905:CTS

[7502] L. Lodian. Concentrated tea slabs. Scientific American, 92(15):308, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n15/pdf/scientificamerican04151905-308.pdf.

Anonymous:1905:CDL

[7503] Anonymous. A curiously distorted locomotive. Scientific American, 92(15):308, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
308a.pdf.

Baltimore:1905:VRB

[7504] J. Mayne Baltimore. A very rare bronze casting. Scientific American, 92(15):308, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
308b.pdf.

Anonymous:1905:BPWg

[7505] Anonymous. Business and personal wants. Scientific American, 92(15):310, April 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n15/pdf/scientificamerican04151905-
310.pdf.

Anonymous:1905:Cg

[7506] Anonymous. Correspondence. Scientific American, 92(16):322, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-322.pdf.

Anonymous:1905:RPIh

[7507] Anonymous. Recently patented inventions. Scientific American, 92(16):328, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
328.pdf.

Anonymous:1905:NQh

[7508] Anonymous. Notes and queries. Scientific American, 92(16):329–330, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-329a.pdf.

Anonymous:1905:IIh

[7509] Anonymous. Index of inventions. Scientific American, 92(16):330–335, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-330.pdf.

Anonymous:1905:PNPg

[7510] Anonymous. Patent notes: Patents. Scientific American, 92(16):335, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-335.pdf.

Anonymous:1905:MLDa

[7511] Anonymous. Maximum load due to crowding. Scientific American, 92(16):318, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
318.pdf.

Anonymous:1905:R

[7512] Anonymous. Rojestvensky. Scientific American, 92(16):318, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-318a.pdf.

Anonymous:1905:RM

[7513] Anonymous. Radium on metals. Scientific American, 92(16):318, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-318b.pdf.

DelMar:1905:OHMa

[7514] Alex. Del Mar. Our heritage of the mechanical arts. Scientific American, 92(16):319, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
319.pdf.

Anonymous:1905:NFA

[7515] Anonymous. Novelties in fire alarms. Scientific American, 92(16):319, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-319a.pdf.

Anonymous:1905:SNd

[7516] Anonymous. Science notes. Scientific American, 92(16):319, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-319b.pdf.

Anonymous:1905:MLDb

[7517] Anonymous. Maximum load due to a crowd of people. Scientific American, 92(16):320, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
320.pdf.

Anonymous:1905:DCD

[7518] Anonymous. Disappearing coast-defense guns. Scientific American, 92(16):320, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
320a.pdf.

Willey:1905:VA

[7519] Day Allen Willey. The Vesuvius of America. Scientific American, 92(16):321–322, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
321.pdf.

Gradenwitz:1905:LSA

[7520] Alfred Gradenwitz. A leadless sounding apparatus. Scientific American, 92(16):323, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
323.pdf.

Fitz-Gerald:1905:RMM

[7521] W. G. Fitz-Gerald. Realism in military maneuvers. Scientific American, 92(16):324, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
324.pdf.

Brown:1905:CAB

[7522] Enos Brown. The Concrete Amphitheater at Berkeley, Cal. Scientific American, 92(16):324–326, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
324a.pdf.

Anonymous:1905:ESS

[7523] Anonymous. An electric street sprinkler. Scientific American, 92(16):326, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
326.pdf.

Anonymous:1905:PDb

[7524] Anonymous. Patent department. Scientific American, 92(16):327, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-327.pdf.

Anonymous:1905:IBS

[7525] Anonymous. Improved belt-stud tool. Scientific American, 92(16):327, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-327a.pdf.

Anonymous:1905:ISD

[7526] Anonymous. Improved shell-digger. Scientific American, 92(16):327, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-327b.pdf.

Anonymous:1905:OIb

[7527] Anonymous. Oddities in inventions. Scientific American, 92(16):327, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n16/pdf/scientificamerican04221905-327c.pdf.

Anonymous:1905:BPWh

[7528] Anonymous. Business and personal wants. Scientific American, 92(16):329, April 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n16/pdf/scientificamerican04221905-
329.pdf.

Russell:1905:HMa

[7529] Henry Norris Russell. The heavens in May. Scientific American, 92(17):339, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
339a.pdf.

Anonymous:1905:RPIi

[7530] Anonymous. Recently patented inventions. Scientific American, 92(17):349, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
349.pdf.

Anonymous:1905:NQi

[7531] Anonymous. Notes and queries. Scientific American, 92(17):350–351, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n17/pdf/scientificamerican04291905-350.pdf.

Anonymous:1905:IIi

[7532] Anonymous. Index of inventions. Scientific American, 92(17):351–355, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n17/pdf/scientificamerican04291905-351.pdf.

Anonymous:1905:PNPh

[7533] Anonymous. Patent notes: Patents. Scientific American, 92(17):355, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n17/pdf/scientificamerican04291905-355.pdf.

Anonymous:1905:SCC

[7534] Anonymous. Subway competition and cheap fares. Scientific American, 92(17):338, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
338.pdf.

Anonymous:1905:ATA

[7535] Anonymous. Another turbine Atlantic liner. Scientific American, 92(17):338, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
338a.pdf.

Anonymous:1905:CLC

[7536] Anonymous. “Connecticut”–“Louisiana” contest. Scientific American, 92(17):338, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
338b.pdf.

Anonymous:1905:VWT

[7537] Anonymous. The Vanderbilt water-tube boiler for locomotives. Scientific American, 92(17):338, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
338c.pdf.

DelMar:1905:OHMb

[7538] Alex. Del Mar. Our heritage of the mechanical arts — II. Scientific American, 92(17):339, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
339.pdf.

Lodian:1905:VLL

[7539] L. Lodian. The virgin lace (lace-bark) of the tropics. Scientific American, 92(17):340, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
340.pdf.

Anonymous:1905:EDT

[7540] Anonymous. An efficient disk transmission for automobiles. Scientific American, 92(17):340, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
340a.pdf.

Collins:1905:CHM

[7541] A. Frederick Collins. The Cooper Hewitt mercury vapor converter. Scientific American, 92(17):341, April 29, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n17/pdf/scientificamerican04291905-341.pdf.

Anonymous:1905:NMDa

[7542] Anonymous. New Masonry Drydock at the Boston Navy Yard. Scientific American, 92(17):341–342, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
341a.pdf.

Anonymous:1905:HHC

[7543] Anonymous. How hydrographic charts are made. Scientific American, 92(17):343, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
343.pdf.

Anonymous:1905:HVF

[7544] Anonymous. Harnessing the Victoria Falls for electrical power. Scientific American, 92(17):344–345, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
344.pdf.

Anonymous:1905:PSW

[7545] Anonymous. Pompeiian silver ware. Scientific American, 92(17):346, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n17/pdf/scientificamerican04291905-346.pdf.

Mcurdy:1905:NML

[7546] James G. M’curdy. Novel method of logging small tracts. Scientific American, 92(17):347–348, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
347.pdf.

Holder:1905:PLF

[7547] Charles F. Holder. Photographing leaping fishes. Scientific American, 92(17):348, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
348.pdf.

Anonymous:1905:BPWi

[7548] Anonymous. Business and personal wants. Scientific American, 92(17):349–350, April 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n17/pdf/scientificamerican04291905-
349a.pdf.

Anonymous:1905:Ci

[7549] Anonymous. Correspondence. Scientific American, 92(18):363, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n18/pdf/scientificamerican05061905-363.pdf.

Anonymous:1905:RPIj

[7550] Anonymous. Recently patented inventions. Scientific American, 92(18):369, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
369.pdf.

Anonymous:1905:NQj

[7551] Anonymous. Notes and queries. Scientific American, 92(18):369–370, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n18/pdf/scientificamerican05061905-369b.pdf.

Anonymous:1905:IIj

[7552] Anonymous. Index of inventions. Scientific American, 92(18):370–375, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n18/pdf/scientificamerican05061905-370.pdf.

Anonymous:1905:PNPi

[7553] Anonymous. Patent notes: Patents. Scientific American, 92(18):375, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n18/pdf/scientificamerican05061905-375.pdf.

Anonymous:1905:SGA

[7554] Anonymous. Safety of the Great Assouan Dam. Scientific American, 92(18):358, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
358.pdf.

Anonymous:1905:MBS

[7555] Anonymous. The Manhattan Bridge scandal. Scientific American, 92(18):358, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
358a.pdf.

Anonymous:1905:TFP

[7556] Anonymous. Textile fabrics of paper. Scientific American, 92 (18):358–359, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
358b.pdf.

Benjamin:1905:NAS

[7557] Marcus Benjamin. The National Academy of Sciences. Scientific American, 92(18):359, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
359.pdf.

DelMar:1905:OHMc

[7558] Alex. Del Mar. Our heritage of the mechanical arts — III. Scientific American, 92(18):359, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
359a.pdf.

Anonymous:1905:SNe

[7559] Anonymous. Science notes. Scientific American, 92(18):359, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n18/pdf/scientificamerican05061905-359b.pdf.

Anonymous:1905:NBD

[7560] Anonymous. A novel balanced door. Scientific American, 92(18):360, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n18/pdf/scientificamerican05061905-360.pdf.

Gignilliat:1905:MOH

[7561] L. R. Gignilliat. Manning oars and halyards as a summer outing. Scientific American, 92(18):360–362, May 6, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n18/pdf/scientificamerican05061905-360a.pdf.

Anonymous:1905:SR

[7562] Anonymous. A simple rheostat. Scientific American, 92(18):362, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n18/pdf/scientificamerican05061905-362.pdf.

Willey:1905:LCE

[7563] Day Allen Willey. The Lewis and Clark Exposition. Scientific American, 92(18):364–366, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
364.pdf.

Anonymous:1905:RAA

[7564] Anonymous. Raising an aeroplane with an automobile. Scientific American, 92(18):366, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
366.pdf.

Anonymous:1905:FBA

[7565] Anonymous. Finding of the body of Admiral Paul Jones in Paris. Scientific American, 92(18):367, May 6, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n18/pdf/scientificamerican05061905-367.pdf.

Collins:1905:NTW

[7566] A. Frederick Collins. A new type of watch demagnetizer. Scientific American, 92(18):368, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
368.pdf.

Anonymous:1905:IMF

[7567] Anonymous. Improvement in magazine firearms. Scientific American, 92(18):368, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
368a.pdf.

Anonymous:1905:TF

[7568] Anonymous. Tent fastener. Scientific American, 92(18):368, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n18/pdf/scientificamerican05061905-368b.pdf.

Anonymous:1905:BPWj

[7569] Anonymous. Business and personal wants. Scientific American, 92(18):369, May 6, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n18/pdf/scientificamerican05061905-
369a.pdf.

Anonymous:1905:Cj

[7570] Anonymous. Correspondence. Scientific American, 92(19):383, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n19/pdf/scientificamerican05131905-383.pdf.

Anonymous:1905:RPIk

[7571] Anonymous. Recently patented inventions. Scientific American, 92(19):389, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
389.pdf.

Anonymous:1905:NQk

[7572] Anonymous. Notes and queries. Scientific American, 92(19):389, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n19/pdf/scientificamerican05131905-389b.pdf.

Anonymous:1905:BNBg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n19/pdf/scientificamerican05131905-390.pdf.

Anonymous:1905:IIk

[7574] Anonymous. Index of inventions. Scientific American, 92(19):390–395, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n19/pdf/scientificamerican05131905-390a.pdf.

Anonymous:1905:PNPj

[7575] Anonymous. Patent notes: Patents. Scientific American, 92(19):395, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n19/pdf/scientificamerican05131905-395.pdf.

Anonymous:1905:IRCa

[7576] Anonymous. The International Railway Congress. Scientific American, 92(19):378, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
378.pdf.

Anonymous:1905:EMP

[7577] Anonymous. The economics of the moving platform. Scientific American, 92(19):378, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
378a.pdf.

Anonymous:1905:STT

[7578] Anonymous. The strength of timber treated with preservatives. Scientific American, 92(19):378–379, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
378b.pdf.

Anonymous:1905:VIC

[7579] Anonymous. A visit to the interior of the Crater of Mont Pelée and a partial ascent of the new “dome”. Scientific American, 92(19):379, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n19/pdf/scientificamerican05131905-379.pdf.

Anonymous:1905:NTCa

[7580] Anonymous. A new type of cow-milking machine. Scientific American, 92(19):380, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
380.pdf.

Anonymous:1905:CJP

[7581] Anonymous. Completing the Jerome Park Reservoir. Scientific American, 92(19):381–382, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
381.pdf.

Anonymous:1905:ARG

[7582] Anonymous. The auxetophone for reinforcing gramophone sounds. Scientific American, 92(19):382, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
382.pdf.

Anonymous:1905:MPS

[7583] Anonymous. Moving platform subway for New York City. Scientific American, 92(19):382–383, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
382a.pdf.

Lavell:1905:VIM

[7584] Charles Lavell. The Viagraph: An instrument for measuring road wear. Scientific American, 92(19):384, May 13, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n19/pdf/scientificamerican05131905-384.pdf.

Anonymous:1905:MGT

[7585] Anonymous. The making of a Gobelin tapestry. Scientific American, 92(19):384–386, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
384a.pdf.

Anonymous:1905:NFW

[7586] Anonymous. A novel five-wheeled automobile. Scientific American, 92(19):386, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
386.pdf.

Anonymous:1905:BCMa

[7587] Anonymous. Bee culture for medicinal purposes. Scientific American, 92(19):387, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
387.pdf.

Byers:1905:PFC

[7588] Charles Alma Byers. A petrified forest covering thousands of acres. Scientific American, 92(19):388, May 13, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n19/pdf/scientificamerican05131905-388.pdf.

Anonymous:1905:BPWk

[7589] Anonymous. Business and personal wants. Scientific American, 92(19):389, May 13, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n19/pdf/scientificamerican05131905-
389a.pdf.

Anonymous:1905:Ck

[7590] Anonymous. Correspondence. Scientific American, 92(2):27, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n2/pdf/scientificamerican01141905-27.pdf.

Anonymous:1905:RPIl

[7591] Anonymous. Recently patented inventions. Scientific American, 92(2):30–31, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
30.pdf.

Anonymous:1905:NQl

[7592] Anonymous. Notes and queries. Scientific American, 92(2):31, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n2/pdf/scientificamerican01141905-31a.pdf.

Anonymous:1905:BNBh

[7593] Anonymous. Books: New books, etc. Scientific American, 92 (2):31, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
31b.pdf.

Anonymous:1905:IIl

[7594] Anonymous. Index of inventions. Scientific American, 92(2):31–35, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n2/pdf/scientificamerican01141905-31c.pdf.

Anonymous:1905:PNPk

[7595] Anonymous. Patent notes: Patents. Scientific American, 92(2):35, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n2/pdf/scientificamerican01141905-35.pdf.

Anonymous:1905:LJW

[7596] Anonymous. Lessons of the Japanese War. Scientific American, 92(2):18, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
18.pdf.

Anonymous:1905:PFA

[7597] Anonymous. Performance of French and American locomotives compared. Scientific American, 92(2):18, January 14, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n2/pdf/scientificamerican01141905-18a.pdf.

Greeley:1905:PNT

[7598] P. Greeley. The proposed new trade-mark law. Scientific American, 92(2):19, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
19.pdf.

Anonymous:1905:ENa

[7599] Anonymous. Engineering notes. Scientific American, 92(2):19, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n2/pdf/scientificamerican01141905-19a.pdf.

Etzel:1905:FTD

[7600] William Etzel. A four-thousand-dollar watch. Scientific American, 92(2):20, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
20.pdf.

Gradenwitz:1905:NWW

[7601] Alfred Gradenwitz. A novel water wheel. Scientific American, 92(2):20, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
20a.pdf.

Collins:1905:GLB

[7602] A. Frederick Collins. The Gros life-belt. Scientific American, 92(2):21, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
21.pdf.

Willey:1905:SC

[7603] Day Allen Willey. Snow crystals. Scientific American, 92(2):21–22, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n2/pdf/scientificamerican01141905-21a.pdf.

Anonymous:1905:PSB

[7604] Anonymous. Packing for steam-boiler manholes. Scientific American, 92(2):22, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
22.pdf.

Anonymous:1905:AIA

[7605] Anonymous. An annual Italian Automobile Cup Race. Scientific American, 92(2):22, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
22a.pdf.

MacDonald:1905:DMI

[7606] J. A. MacDonald. Determining the Meridian without instruments. Scientific American, 92(2):23, January 14, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n2/pdf/scientificamerican01141905-23.pdf.

Anonymous:1905:WBB

[7607] Anonymous. Why barrels are broken by ice expansion. Scientific American, 92(2):23, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
23a.pdf.

Anonymous:1905:IPS

[7608] Anonymous. An interesting powerful steam dredger for harbor work. Scientific American, 92(2):24, January 14, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n2/pdf/scientificamerican01141905-24.pdf.

Barry:1905:ISC

[7609] Richard Barry. The investment, siege, and capture of Port Arthur. Scientific American, 92(2):24–27, January 14, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n2/pdf/scientificamerican01141905-24a.pdf.

Anonymous:1905:PDc

[7610] Anonymous. Patent department. Scientific American, 92(2):28, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n2/pdf/scientificamerican01141905-28.pdf.

Anonymous:1905:IGH

[7611] Anonymous. Improved garment hanger. Scientific American, 92 (2):28, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
28a.pdf.

Anonymous:1905:NAAa

[7612] Anonymous. Novel attachment for automobiles. Scientific American, 92(2):28, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
28b.pdf.

Anonymous:1905:OIc

[7613] Anonymous. Oddities in invention. Scientific American, 92(2): 28–29, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
28c.pdf.

Anonymous:1905:BPWl

[7614] Anonymous. Business and personal wants. Scientific American, 92(2):31, January 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n2/pdf/scientificamerican01141905-
31.pdf.

Anonymous:1905:Cl

[7615] Anonymous. Correspondence. Scientific American, 92(20):403, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-403.pdf.

Anonymous:1905:RPIm

[7616] Anonymous. Recently patented inventions. Scientific American, 92(20):409–410, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
409.pdf.

Anonymous:1905:NQm

[7617] Anonymous. Notes and queries. Scientific American, 92(20):410–411, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-410a.pdf.

Anonymous:1905:BNBi

[7618] Anonymous. Books: New books, etc. Scientific American, 92(20):412, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-412.pdf.

Anonymous:1905:IIm

[7619] Anonymous. Index of inventions. Scientific American, 92(20):412–415, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-412a.pdf.

Anonymous:1905:RRH

[7620] Anonymous. The recent railroad horror. Scientific American, 92 (20):398, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
398.pdf.

Anonymous:1905:TGO

[7621] Anonymous. Two great ocean contests. Scientific American, 92(20):398, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-398a.pdf.

Anonymous:1905:CCDa

[7622] Anonymous. Complete the coast defenses. Scientific American, 92(20):398, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
398b.pdf.

Anonymous:1905:IRCb

[7623] Anonymous. The International Railway Congress. Scientific American, 92(20):398–399, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
398c.pdf.

Anonymous:1905:EPM

[7624] Anonymous. An electric process for manufacturing peat. Scientific American, 92(20):399, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
399.pdf.

Anonymous:1905:AC

[7625] Anonymous. Artificial cotton. Scientific American, 92(20):399, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-399a.pdf.

Anonymous:1905:SNf

[7626] Anonymous. Science notes. Scientific American, 92(20):399, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-399b.pdf.

Brassler:1905:ACA

[7627] Charles A. Brassler. An astronomical clock with an automatic adjustment of time and church calendar. Scientific American, 92(20):400, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-400.pdf.

Willey:1905:CFN

[7628] Day Allen Willey. Concrete foundations of the new passenger station, Washington, D.C. Scientific American, 92(20):401–402, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-401.pdf.

Anonymous:1905:MHS

[7629] Anonymous. The modern high-speed locomotive. Scientific American, 92(20):402, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
402.pdf.

Anonymous:1905:GMY

[7630] Anonymous. The great 3,000-mile yacht race. Scientific American, 92(20):402–403, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
402a.pdf.

Anonymous:1905:MA

[7631] Anonymous. The Montgomery aeroplane. Scientific American, 92 (20):404–406, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
404.pdf.

Anonymous:1905:DRM

[7632] Anonymous. Destroying a Russian mine. Scientific American, 92 (20):406, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
406.pdf.

Anonymous:1905:CTS

[7633] Anonymous. Comparative tests of steam and electric locomotives. Scientific American, 92(20):407, May 20, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n20/pdf/scientificamerican05201905-407.pdf.

Anonymous:1905:PDd

[7634] Anonymous. Patent and department. Scientific American, 92(20):408, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n20/pdf/scientificamerican05201905-408.pdf.

Anonymous:1905:IV

[7635] Anonymous. An improved viscosimeter. Scientific American, 92 (20):408, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
408a.pdf.

Anonymous:1905:IFEa

[7636] Anonymous. An improved fire extinguisher. Scientific American, 92(20):408, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
408b.pdf.

Anonymous:1905:ATF

[7637] Anonymous. Apparatus for testing flashing points of oils. Scientific American, 92(20):408, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
408c.pdf.

Anonymous:1905:BPWm

[7638] Anonymous. Business and personal wants. Scientific American, 92(20):410, May 20, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n20/pdf/scientificamerican05201905-
410.pdf.

Anonymous:1905:Cm

[7639] Anonymous. Correspondence. Scientific American, 92(21):423, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-423.pdf.

Anonymous:1905:RPIn

[7640] Anonymous. Recently patented inventions. Scientific American, 92(21):428, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n21/pdf/scientificamerican05271905-
428.pdf.

Anonymous:1905:NQn

[7641] Anonymous. Notes and queries. Scientific American, 92(21):430–431, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-430.pdf.

Anonymous:1905:BNBj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-431.pdf.

Anonymous:1905:IIn

[7643] Anonymous. Index of inventions. Scientific American, 92(21):431–435, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-431a.pdf.

Anonymous:1905:PNPl

[7644] Anonymous. Patent notes: Patents. Scientific American, 92(21):435, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-435.pdf.

Anonymous:1905:HEF

[7645] Anonymous. High explosives on freight trains. Scientific American, 92(21):418, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n21/pdf/scientificamerican05271905-
418.pdf.

Anonymous:1905:ORQ

[7646] Anonymous. An oft-repeated question. Scientific American, 92(21):418, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-418a.pdf.

Anonymous:1905:CSF

[7647] Anonymous. Coal supply of the future. Scientific American, 92(21):418, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-418b.pdf.

Anonymous:1905:BSN

[7648] Anonymous. Battleship strength of the Navies of the world. Scientific American, 92(21):418, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n21/pdf/scientificamerican05271905-
418c.pdf.

Anonymous:1905:SPR

[7649] Anonymous. Superstitions that prevail in rural sections. Scientific American, 92(21):418–419, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n21/pdf/scientificamerican05271905-
418d.pdf.

Anonymous:1905:FOB

[7650] Anonymous. The first observations with “ballons-sondes” in America. Scientific American, 92(21):419, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v92/n21/pdf/scientificamerican05271905-
419.pdf.

Honeking:1905:MPP

[7651] J. A. Honeking. A method for putting printed matter on finished lantern slides. Scientific American, 92(21):419, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-419a.pdf.

Anonymous:1905:SNg

[7652] Anonymous. Science notes. Scientific American, 92(21):419, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-419b.pdf.

Anonymous:1905:HL

[7653] Anonymous. A historical locomotive. Scientific American, 92(21):420, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-420.pdf.

Geare:1905:C

[7654] Randolph I. Geare. The cobra. Scientific American, 92(21):420, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-420a.pdf.

Guarini:1905:OH

[7655] Emile Guarini. Oxygen at home. Scientific American, 92(21):421, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-421.pdf.

Anonymous:1905:RB

[7656] Anonymous. The Renaissance of the bayonet. Scientific American, 92(21):421–422, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n21/pdf/scientificamerican05271905-
421a.pdf.

Anonymous:1905:SORb

[7657] Anonymous. Start of the ocean race. Scientific American, 92(21):422, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n21/pdf/scientificamerican05271905-422.pdf.

Anonymous:1905:ARCa

[7658] Anonymous. American racing cars for the international cup race. Scientific American, 92(21):424, May 27, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n21/pdf/scientificamerican05271905-424.pdf.

Beasley:1905:NYA

[7659] Walter L. Beasley. The New York Aquarium. Scientific American, 92(21):425–426, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n21/pdf/scientificamerican05271905-
425.pdf.

Anonymous:1905:NPLb

[7660] Anonymous. A novel power launch for clearing weed-infested waterways. Scientific American, 92(21):426, May 27, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n21/pdf/scientificamerican05271905-426.pdf.

Ramakers:1905:MTA

[7661] L. Ramakers. Modern thermometric apparatus for high temperatures. Scientific American, 92(21):427, May 27, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n21/pdf/scientificamerican05271905-427.pdf.

Anonymous:1905:BPWn

[7662] Anonymous. Business and personal wants. Scientific American, 92(21):428, May 27, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n21/pdf/scientificamerican05271905-
428a.pdf.

Russell:1905:HJa

[7663] Henry Norris Russell. The heavens in June. Scientific American, 92(22):439, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
439.pdf.

Anonymous:1905:Cn

[7664] Anonymous. Correspondence. Scientific American, 92(22):443, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-443.pdf.

Anonymous:1905:RPIo

[7665] Anonymous. Recently patented inventions. Scientific American, 92(22):449–450, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
449.pdf.

Anonymous:1905:NQo

[7666] Anonymous. Notes and queries. Scientific American, 92(22):451–452, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-451.pdf.

Anonymous:1905:IIo

[7667] Anonymous. Index of inventions. Scientific American, 92(22):452–455, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-452.pdf.

Anonymous:1905:PNPm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-455.pdf.

Anonymous:1905:APP

[7669] Anonymous. Alternative plans for the Panama Canal. Scientific American, 92(22):438, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
438.pdf.

Anonymous:1905:AWW

[7670] Anonymous. An attack on the wire-wound gun. Scientific American, 92(22):438, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
438a.pdf.

Anonymous:1905:PSP

[7671] Anonymous. The preparation of surfaces for painting. Scientific American, 92(22):438–439, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
438b.pdf.

Anonymous:1905:EB

[7672] Anonymous. Electricity and bread. Scientific American, 92(22):439, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-439a.pdf.

Anonymous:1905:SNh

[7673] Anonymous. Science notes. Scientific American, 92(22):439–440, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-439b.pdf.

Anonymous:1905:EMA

[7674] Anonymous. Electric mail automobiles. Scientific American, 92 (22):440, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
440.pdf.

Guarini:1905:GC

[7675] Emile Guarini. The graphic chronometer. Scientific American, 92(22):441, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
441.pdf.

Anonymous:1905:EDR

[7676] Anonymous. Electric dead reckoner used on “Valhalla” in the ocean race. Scientific American, 92(22):441–442, June 3, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n22/pdf/scientificamerican06031905-441a.pdf.

Anonymous:1905:HPS

[7677] Anonymous. History and present status of the Panama Canal. Scientific American, 92(22):442–443, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
442.pdf.

Anonymous:1905:GGS

[7678] Anonymous. The Grubb gun-sight. Scientific American, 92(22):444, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-444.pdf.

Anonymous:1905:ATM

[7679] Anonymous. The Algiers–Toulon Motor-Boat Race. Scientific American, 92(22):446–447, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
446.pdf.

Briggs:1905:DAR

[7680] Nelson R. Briggs. Do animals reason? Scientific American, 92(22):447, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-447.pdf.

Russell:1905:UIC

[7681] Edgar Russell. An unusual incident of cable work. Scientific American, 92(22):447, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
447a.pdf.

Anonymous:1905:PDe

[7682] Anonymous. Patent department. Scientific American, 92(22):448, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-448.pdf.

Anonymous:1905:IDC

[7683] Anonymous. Improved dumping car. Scientific American, 92(22):448, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-448a.pdf.

Anonymous:1905:IDM

[7684] Anonymous. An improved ditching machine. Scientific American, 92(22):448, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
448b.pdf.

Anonymous:1905:PPT

[7685] Anonymous. A puncture-proof tire. Scientific American, 92(22):448, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n22/pdf/scientificamerican06031905-448c.pdf.

Anonymous:1905:BPWo

[7686] Anonymous. Business and personal wants. Scientific American, 92(22):450, June 3, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n22/pdf/scientificamerican06031905-
450.pdf.

Anonymous:1905:RPIp

[7687] Anonymous. Recently patented inventions. Scientific American, 92(23):469, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
469.pdf.

Anonymous:1905:NQp

[7688] Anonymous. Notes and queries. Scientific American, 92(23):470, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n23/pdf/scientificamerican06101905-470.pdf.

Anonymous:1905:BNBk

[7689] Anonymous. Books: New books, etc. Scientific American, 92 (23):470–471, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
470a.pdf.

Anonymous:1905:IIp

[7690] Anonymous. Index of inventions. Scientific American, 92(23):471–475, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n23/pdf/scientificamerican06101905-471.pdf.

Anonymous:1905:PNPn

[7691] Anonymous. Patent notes: Patents. Scientific American, 92(23):475, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n23/pdf/scientificamerican06101905-475.pdf.

Anonymous:1905:UV

[7692] Anonymous. An unparalleled victory. Scientific American, 92(23):458, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n23/pdf/scientificamerican06101905-458.pdf.

Anonymous:1905:ESP

[7693] Anonymous. Electricity as a stimulant to plant growth. Scientific American, 92(23):458, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
458a.pdf.

Anonymous:1905:SNAa

[7694] Anonymous. Something new about the vocal mechanism. Scientific American, 92(23):458–459, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
458b.pdf.

Anonymous:1905:RAL

[7695] Anonymous. The railway around Lake Baikal. Scientific American, 92(23):459, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
459.pdf.

Anonymous:1905:CSb

[7696] Anonymous. The current supplement. Scientific American, 92(23):459, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n23/pdf/scientificamerican06101905-459a.pdf.

Anonymous:1905:NS

[7697] Anonymous. A new star. Scientific American, 92(23):459, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n23/pdf/scientificamerican06101905-459b.pdf.

Anonymous:1905:MSNa

[7698] Anonymous. Meteorological Summary, New York, N.Y., May, 1905. Scientific American, 92(23):459, June 10, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n23/pdf/scientificamerican06101905-459c.pdf.

Anonymous:1905:SNi

[7699] Anonymous. Science notes. Scientific American, 92(23):459, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n23/pdf/scientificamerican06101905-459d.pdf.

Anonymous:1905:NAC

[7700] Anonymous. A new apparatus for coaling steamers. Scientific American, 92(23):460, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
460.pdf.

Rigge:1905:WWC

[7701] William F. Rigge. Where was the camera set up? Scientific American, 92(23):461, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
461.pdf.

Perkins:1905:EAF

[7702] Frank C. Perkins. An electrical aerial ferry. Scientific American, 92(23):461–462, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
461a.pdf.

Anonymous:1905:FOY

[7703] Anonymous. The finish of the ocean yacht race. Scientific American, 92(23):462, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
462.pdf.

Anonymous:1905:BSJ

[7704] Anonymous. The Battle of the Sea of Japan. Scientific American, 92(23):463, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
463.pdf.

Anonymous:1905:ATH

[7705] Anonymous. Admiral Togo and his Flagship. Scientific American, 92(23):464–465, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
464.pdf.

Anonymous:1905:MSD

[7706] Anonymous. Modern scientific dairy methods. Scientific American, 92(23):465–466, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
465.pdf.

Walker:1905:OHS

[7707] Herbert T. Walker. An old high-speed locomotive. Scientific American, 92(23):467–468, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
467.pdf.

Anonymous:1905:NMP

[7708] Anonymous. A novel milk pail. Scientific American, 92(23):468, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n23/pdf/scientificamerican06101905-468.pdf.

Jones:1905:IER

[7709] George J. Jones. An iceless electric refrigerator on a small scale. Scientific American, 92(23):468, June 10, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n23/pdf/scientificamerican06101905-468a.pdf.

Anonymous:1905:BPWp

[7710] Anonymous. Business and personal wants. Scientific American, 92(23):469, June 10, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n23/pdf/scientificamerican06101905-
469a.pdf.

Anonymous:1905:Co

[7711] Anonymous. Correspondence. Scientific American, 92(24):482, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-482a.pdf.

Anonymous:1905:RPIq

[7712] Anonymous. Recently patented inventions. Scientific American, 92(24):489, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
489.pdf.

Anonymous:1905:NQq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-490.pdf.

Anonymous:1905:BNBl

[7714] Anonymous. Books: New books, etc. Scientific American, 92(24):491, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-491.pdf.

Anonymous:1905:IIq

[7715] Anonymous. Index of inventions. Scientific American, 92(24):491–495, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-491a.pdf.

Anonymous:1905:PNPo

[7716] Anonymous. Patent notes: Patents. Scientific American, 92(24):495, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-495.pdf.

Anonymous:1905:EWN

[7717] Anonymous. The effect of the war on Naval construction. Scientific American, 92(24):478, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
478.pdf.

Anonymous:1905:FLDa

[7718] Anonymous. Fast long-distance trains. Scientific American, 92(24):478, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-478a.pdf.

Anonymous:1905:VS

[7719] Anonymous. Ventilation of the subway. Scientific American, 92(24):478, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-478b.pdf.

Anonymous:1905:SBD

[7720] Anonymous. The submarine boat disasters. Scientific American, 92(24):478–479, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
478c.pdf.

Adams:1905:STV

[7721] Alton D. Adams. Selection of transmission voltages. Scientific American, 92(24):479, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
479.pdf.

Anonymous:1905:NMM

[7722] Anonymous. New method of milk analysis by centrifugal apparatus. Scientific American, 92(24):479, June 17, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n24/pdf/scientificamerican06171905-479a.pdf.

Anonymous:1905:EIC

[7723] Anonymous. An explanation of ice caves. Scientific American, 92(24):479, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
479b.pdf.

Anonymous:1905:CSc

[7724] Anonymous. The current supplement. Scientific American, 92(24):479, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-479c.pdf.

Anonymous:1905:LHS

[7725] Anonymous. Loss of heat in steam. Scientific American, 92(24):479, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-479d.pdf.

Anonymous:1905:VRP

[7726] Anonymous. A vagary of the Russian Press Censor. Scientific American, 92(24):479, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
479e.pdf.

Anonymous:1905:AAF

[7727] Anonymous. The Alvares aeroplane flying machine. Scientific American, 92(24):480, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
480.pdf.

Brown:1905:CLA

[7728] Enos Brown. The creation of life by artificial means. Scientific American, 92(24):480, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
480a.pdf.

Anonymous:1905:NSB

[7729] Anonymous. The new submarines for the British Navy. Scientific American, 92(24):481, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
481.pdf.

Gradenwitz:1905:PMN

[7730] Alfred Gradenwitz. On a process of measuring nervous susceptibility. Scientific American, 92(24):481–482, June 17, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n24/pdf/scientificamerican06171905-481a.pdf.

Anonymous:1905:MER

[7731] Anonymous. Mines and entanglements in the Russo-Japanese War. Scientific American, 92(24):482, June 17, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n24/pdf/scientificamerican06171905-482.pdf.

Anonymous:1905:PBW

[7732] Anonymous. A power brake and whistle for automobiles. Scientific American, 92(24):483, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
483.pdf.

Claudy:1905:MPL

[7733] C. H. Claudy. A model photographic laboratory. Scientific American, 92(24):483–485, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
483a.pdf.

Thompson:1905:SUD

[7734] G. A. Thompson. Some unusual discoveries of variable stars at Harvard Observatory. Scientific American, 92(24):485–486, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-485.pdf.

Gradenwitz:1905:CM

[7735] Alfred Gradenwitz. The capillary motor. Scientific American, 92 (24):486, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
486.pdf.

Anonymous:1905:NRL

[7736] Anonymous. A novel reflecting lantern. Scientific American, 92(24):487, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-487.pdf.

Anonymous:1905:TMMb

[7737] Anonymous. A 750-ton metal mixer. Scientific American, 92(24):488, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-488.pdf.

Anonymous:1905:BI

[7738] Anonymous. A beehive incubator. Scientific American, 92(24):488, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n24/pdf/scientificamerican06171905-488a.pdf.

Anonymous:1905:BPWq

[7739] Anonymous. Business and personal wants. Scientific American, 92(24):489, June 17, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n24/pdf/scientificamerican06171905-
489a.pdf.

Anonymous:1905:Cp

[7740] Anonymous. Correspondence. Scientific American, 92(25):502–503, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n25/pdf/scientificamerican06241905-502.pdf.

Anonymous:1905:RPIr

[7741] Anonymous. Recently patented inventions. Scientific American, 92(25):507, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
507.pdf.

Anonymous:1905:NQr

[7742] Anonymous. Notes and queries. Scientific American, 92(25):507, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n25/pdf/scientificamerican06241905-507b.pdf.

Anonymous:1905:BNBm

[7743] Anonymous. Books: New books, etc. Scientific American, 92(25):509, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n25/pdf/scientificamerican06241905-509.pdf.

Anonymous:1905:IIr

[7744] Anonymous. Index of inventions. Scientific American, 92(25):509–512, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n25/pdf/scientificamerican06241905-509a.pdf.

Anonymous:1905:I

[7745] Anonymous. Index. Scientific American, 92(25):513–514, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n25/pdf/scientificamerican06241905-513.pdf.

Anonymous:1905:PNPp

[7746] Anonymous. Patent notes: Patents. Scientific American, 92(25):515, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n25/pdf/scientificamerican06241905-515.pdf.

Anonymous:1905:INY

[7747] Anonymous. Increasing the New York Water Supply. Scientific American, 92(25):498, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
498.pdf.

Anonymous:1905:WF

[7748] Anonymous. The warships of the future. Scientific American, 92(25):498, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
498a.pdf.

Anonymous:1905:ESH

[7749] Anonymous. Eye-strain and how it can be relieved. Scientific American, 92(25):498–499, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
498b.pdf.

Anonymous:1905:EWL

[7750] Anonymous. Electric waves and liquids. Scientific American, 92 (25):499, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
499.pdf.

Blanchard:1905:RWS

[7751] C. J. Blanchard. Reclamation work in Southern California. Scientific American, 92(25):499, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
499a.pdf.

Anonymous:1905:ENb

[7752] Anonymous. Engineering notes. Scientific American, 92(25):499, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n25/pdf/scientificamerican06241905-499b.pdf.

Anonymous:1905:DM

[7753] Anonymous. The development of a mollusk. Scientific American, 92(25):500, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
500.pdf.

Hayward:1905:GA

[7754] Charles B. Hayward. The Gillespie aeroplane. Scientific American, 92(25):501, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
501.pdf.

Claudy:1905:USN

[7755] C. H. Claudy. The United States Naval Observatory Eclipse Expedition. Scientific American, 92(25):503–504, June 24, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n25/pdf/scientificamerican06241905-503.pdf.

Anonymous:1905:NTEa

[7756] Anonymous. New type of electric truck. Scientific American, 92 (25):505–506, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
505.pdf.

Anonymous:1905:BPWr

[7757] Anonymous. Business and personal wants. Scientific American, 92(25):507, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
507a.pdf.

Anonymous:1905:PAL

[7758] Anonymous. The production of automobiles on a lapge scale. Scientific American, 92(25):508–509, June 24, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n25/pdf/scientificamerican06241905-
508.pdf.

Anonymous:1905:RPIs

[7759] Anonymous. Recently patented inventions. Scientific American, 92(3):47, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
47.pdf.

Anonymous:1905:NQs

[7760] Anonymous. Notes and queries. Scientific American, 92(3):47, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n3/pdf/scientificamerican01211905-47b.pdf.

Anonymous:1905:IIs

[7761] Anonymous. Index of inventions. Scientific American, 92(3):47–51, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n3/pdf/scientificamerican01211905-47c.pdf.

Anonymous:1905:PNPq

[7762] Anonymous. Patent notes: Patents. Scientific American, 92(3):51, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n3/pdf/scientificamerican01211905-51.pdf.

Anonymous:1905:SVS

[7763] Anonymous. Schedule versus safety. Scientific American, 92(3):38, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n3/pdf/scientificamerican01211905-38.pdf.

Anonymous:1905:AEF

[7764] Anonymous. The American exhibit of foreign automobiles. Scientific American, 92(3):38, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
38a.pdf.

Anonymous:1905:MB

[7765] Anonymous. A monumental building. Scientific American, 92 (3):38–39, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
38b.pdf.

Walsh:1905:LB

[7766] George E. Walsh. The life of a barrel. Scientific American, 92(3):39, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
39.pdf.

Anonymous:1905:DSB

[7767] Anonymous. Determining a ship’s bearing by wireless telegraphy. Scientific American, 92(3):39, January 21, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n3/pdf/scientificamerican01211905-39a.pdf.

Anonymous:1905:SNj

[7768] Anonymous. Science notes. Scientific American, 92(3):39, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n3/pdf/scientificamerican01211905-39b.pdf.

Anonymous:1905:NGC

[7769] Anonymous. The new Grand Central Station, New York. Scientific American, 92(3):40, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
40.pdf.

Mclure:1905:CFM

[7770] W. Frank M’clure. A chicken-feeding machine. Scientific American, 92(3):41, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
41.pdf.

Anonymous:1905:NCH

[7771] Anonymous. New cælostat and horizontal telescope of the Astrophysical Observatory of the Smithsonian Institution. Scientific American, 92(3):41–42, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
41a.pdf.

Anonymous:1905:BHG

[7772] Anonymous. The brontosaur how a giant prehistoric animal was discovered, transported and restored. Scientific American, 92(3):42–43, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n3/pdf/scientificamerican01211905-42.pdf.

Shepstone:1905:TBG

[7773] Harold J. Shepstone. The Tasmanian blue gum — an ideal timber for harbor building. Scientific American, 92(3):43–44, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n3/pdf/scientificamerican01211905-43.pdf.

Anonymous:1905:CEB

[7774] Anonymous. Completion of the East Boston Tunnel. Scientific American, 92(3):44–46, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
44.pdf.

Anonymous:1905:BPWs

[7775] Anonymous. Business and personal wants. Scientific American, 92(3):47, January 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n3/pdf/scientificamerican01211905-
47a.pdf.

Serviss:1905:HF

[7776] Garrett P. Serviss. The heavens in February. Scientific American, 92(4):55, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
55.pdf.

Anonymous:1905:RPIt

[7777] Anonymous. Recently patented inventions. Scientific American, 92(4):90, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
90.pdf.

Anonymous:1905:NQt

[7778] Anonymous. Notes and queries. Scientific American, 92(4):90–91, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n4/pdf/scientificamerican01281905-90a.pdf.

Anonymous:1905:BNBn

[7779] Anonymous. Books: New books, etc. Scientific American, 92 (4):91–92, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
91.pdf.

Anonymous:1905:IIt

[7780] Anonymous. Index of inventions. Scientific American, 92(4):92–95, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n4/pdf/scientificamerican01281905-92.pdf.

Anonymous:1905:OAP

[7781] Anonymous. Our assets at Panama. Scientific American, 92(4):54, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n4/pdf/scientificamerican01281905-54.pdf.

Anonymous:1905:AFS

[7782] Anonymous. Asphalt flooring at subway stations. Scientific American, 92(4):54, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
54a.pdf.

Anonymous:1905:HRI

[7783] Anonymous. History repeats itself. Scientific American, 92(4):54, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n4/pdf/scientificamerican01281905-54b.pdf.

Anonymous:1905:WT

[7784] Anonymous. Warfare by telegraph. Scientific American, 92(4):54, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n4/pdf/scientificamerican01281905-54c.pdf.

Anonymous:1905:FAA

[7785] Anonymous. The fifth annual automobile show. Scientific American, 92(4):54, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
54d.pdf.

Anonymous:1905:EAS

[7786] Anonymous. An electric-acoustic sea-sounding apparatus. Scientific American, 92(4):55, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
55a.pdf.

Anonymous:1905:CSd

[7787] Anonymous. The current supplement. Scientific American, 92 (4):55, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
55b.pdf.

Anonymous:1905:SNk

[7788] Anonymous. Science notes. Scientific American, 92(4):55, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n4/pdf/scientificamerican01281905-55c.pdf.

Anonymous:1905:SLA

[7789] Anonymous. Some leading automobiles of the present year. Scientific American, 92(4):56, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
56.pdf.

Anonymous:1905:TAF

[7790] Anonymous. A typical American four-cylinder gasoline touring car motor. Scientific American, 92(4):57, January 28, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n4/pdf/scientificamerican01281905-57.pdf.

Anonymous:1905:TATa

[7791] Anonymous. Typical American touring cars for 1905. Scientific American, 92(4):58–59, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
58.pdf.

Anonymous:1905:FAI

[7792] Anonymous. Foreign automobiles at the importers’ salon. Scientific American, 92(4):60–61, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
60.pdf.

Anonymous:1905:TPA

[7793] Anonymous. Twenty-passenger auto-stage for long-distance routes. Scientific American, 92(4):62, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
62.pdf.

Anonymous:1905:GPP

[7794] Anonymous. A gasoline police patrol wagon. Scientific American, 92(4):62, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
62a.pdf.

Anonymous:1905:BMO

[7795] Anonymous. The brush motor omnibus. Scientific American, 92(4):62, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
62b.pdf.

Anonymous:1905:NOD

[7796] Anonymous. The new Oldsmobile delivery wagon. Scientific American, 92(4):62–63, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
62c.pdf.

Anonymous:1905:GTD

[7797] Anonymous. A gasoline truck driven by all four wheels. Scientific American, 92(4):63, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
63.pdf.

Anonymous:1905:NTCb

[7798] Anonymous. The Northern touring car. Scientific American, 92 (4):64, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
64.pdf.

Anonymous:1905:NGSa

[7799] Anonymous. The new grout steam automobile. Scientific American, 92(4):64, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
64a.pdf.

Anonymous:1905:NSW

[7800] Anonymous. A new spring wheel for automobiles. Scientific American, 92(4):64, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
64b.pdf.

Anonymous:1905:FCM

[7801] Anonymous. A four-cylinder motor bicycle. Scientific American, 92(4):65, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
65.pdf.

Anonymous:1905:NSCa

[7802] Anonymous. A novel side-chair attachment for motor bicycles. Scientific American, 92(4):65, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
65a.pdf.

Anonymous:1905:NTCc

[7803] Anonymous. The National touring car. Scientific American, 92 (4):66–67, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
66.pdf.

Anonymous:1905:NAAb

[7804] Anonymous. A new American automobile. Scientific American, 92(4):68, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
68.pdf.

Anonymous:1905:PLW

[7805] Anonymous. A powerful light-weight gasoline motor. Scientific American, 92(4):70, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
70.pdf.

Anonymous:1905:IAb

[7806] Anonymous. Ignition accumulators. Scientific American, 92(4):70, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n4/pdf/scientificamerican01281905-70a.pdf.

Anonymous:1905:SNS

[7807] Anonymous. Solid and non-skid tires for automobiles. Scientific American, 92(4):72, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
72.pdf.

Anonymous:1905:TAR

[7808] Anonymous. Timing automobile races. Scientific American, 92 (4):72, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
72a.pdf.

Gradenwitz:1905:SGT

[7809] Alfred Gradenwitz. The Stolze gas turbine. Scientific American, 92(4):74, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
74.pdf.

Page:1905:OPR

[7810] Edward H. L. Page. The Old Portage Railroad in Pennsylvania. Scientific American, 92(4):74–76, January 28, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n4/pdf/scientificamerican01281905-74a.pdf.

Inkersley:1905:PWB

[7811] Arthur Inkersley. A pavement of whales’ bones. Scientific American, 92(4):76, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
76.pdf.

Anonymous:1905:PSa

[7812] Anonymous. The parallax stereogram. Scientific American, 92 (4):76, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
76a.pdf.

Hayward:1905:BTC

[7813] Charles B. Hayward. Bananas — their culture and transportation. Scientific American, 92(4):78, January 28, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n4/pdf/scientificamerican01281905-78.pdf.

Anonymous:1905:AMS

[7814] Anonymous. An American motor sleigh. Scientific American, 92(4):84, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
84.pdf.

Anonymous:1905:BPWt

[7815] Anonymous. Business and personal wants. Scientific American, 92(4):88, January 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n4/pdf/scientificamerican01281905-
88.pdf.

Anonymous:1905:Cq

[7816] Anonymous. Correspondence. Scientific American, 92(5):103, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n5/pdf/scientificamerican02041905-103a.pdf.

Anonymous:1905:RPIu

[7817] Anonymous. Recently patented inventions. Scientific American, 92 (5):109–110, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
109.pdf.

Anonymous:1905:NQu

[7818] Anonymous. Notes and queries. Scientific American, 92(5):110, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n5/pdf/scientificamerican02041905-110a.pdf.

Anonymous:1905:BNBo

[7819] Anonymous. Books: New books, etc. Scientific American, 92 (5):110, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
110b.pdf.

Anonymous:1905:IIu

[7820] Anonymous. Index of inventions. Scientific American, 92(5): 110–115, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
110c.pdf.

Anonymous:1905:GDL

[7821] Anonymous. Gas-driven locomotives and ships? Scientific American, 92(5):98, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
98.pdf.

Anonymous:1905:CPA

[7822] Anonymous. Can the Port Arthur Fleet be raised? Scientific American, 92(5):98, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
98a.pdf.

Anonymous:1905:WAR

[7823] Anonymous. World’s automobile records in Florida. Scientific American, 92(5):98, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
98b.pdf.

Anonymous:1905:OE

[7824] Anonymous. Our exports. Scientific American, 92(5):98–99, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n5/pdf/scientificamerican02041905-98c.pdf.

Haswell:1905:FWS

[7825] Charles H. Haswell. Foaming of the water in a steam boiler and its effect. Scientific American, 92(5):99, February 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n5/pdf/scientificamerican02041905-99.pdf.

Anonymous:1905:PTM

[7826] Anonymous. Passage of the trade-mark law by the House of Representatives. Scientific American, 92(5):99, February 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n5/pdf/scientificamerican02041905-99a.pdf.

Anonymous:1905:EEK

[7827] Anonymous. Eclipse Expedition from Kirkwood Observatory, Bloomington, Ind. Scientific American, 92(5):99, February 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n5/pdf/scientificamerican02041905-99b.pdf.

Anonymous:1905:SNl

[7828] Anonymous. Science notes. Scientific American, 92(5):99, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n5/pdf/scientificamerican02041905-99c.pdf.

Gradenwitz:1905:NTH

[7829] Alfred Gradenwitz. A novel type of hay stacker. Scientific American, 92(5):100, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
100.pdf.

Collins:1905:SCB

[7830] A. Frederick Collins. A seedless, coreless, bloomless apple. Scientific American, 92(5):100, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
100a.pdf.

Willey:1905:NCC

[7831] Day Allen Willey. The new concrete chapel of the United States Naval Academy, Annapolis. Scientific American, 92(5):101–102, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n5/pdf/scientificamerican02041905-101.pdf.

Anonymous:1905:TBB

[7832] Anonymous. The turret and barbette of a battleship. Scientific American, 92(5):103, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
103.pdf.

Anonymous:1905:ONCb

[7833] Anonymous. Opening of the Niagara Canadian Power Company’s plant. — I. Scientific American, 92(5):104–105, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n5/pdf/scientificamerican02041905-104.pdf.

Anonymous:1905:AAR

[7834] Anonymous. The Annual Automobile Race Meet at Ormond, Florida. Scientific American, 92(5):105–106, February 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n5/pdf/scientificamerican02041905-105.pdf.

Jones:1905:NSE

[7835] George J. Jones. New system of egg transportation. Scientific American, 92(5):106, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
106.pdf.

Anonymous:1905:HBE

[7836] Anonymous. Hopkins “band electro-chronograph,” carable of dividing a second of time into one hundred thousand and one million parts. Scientific American, 92(5):107, February 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n5/pdf/scientificamerican02041905-107.pdf.

Anonymous:1905:IFEb

[7837] Anonymous. An improved fire escape. Scientific American, 92 (5):108, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
108.pdf.

Anonymous:1905:VAK

[7838] Anonymous. Vehicle attachment for keeping record of driver’s work. Scientific American, 92(5):108, February 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n5/pdf/scientificamerican02041905-108a.pdf.

Anonymous:1905:ADAa

[7839] Anonymous. An automatic door alarm. Scientific American, 92 (5):108, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
108b.pdf.

Anonymous:1905:OCM

[7840] Anonymous. Oil can with means for controlling delivery of oil. Scientific American, 92(5):108, February 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n5/pdf/scientificamerican02041905-108c.pdf.

Anonymous:1905:BPWu

[7841] Anonymous. Business and personal wants. Scientific American, 92(5):110, February 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n5/pdf/scientificamerican02041905-
110.pdf.

Anonymous:1905:Cr

[7842] Anonymous. Correspondence. Scientific American, 92(6):122, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n6/pdf/scientificamerican02111905-122.pdf.

Anonymous:1905:RPIv

[7843] Anonymous. Recently patented inventions. Scientific American, 92(6):129, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
129.pdf.

Anonymous:1905:NQv

[7844] Anonymous. Notes and queries. Scientific American, 92(6):130, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n6/pdf/scientificamerican02111905-130a.pdf.

Anonymous:1905:BNBp 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
130b.pdf.

Anonymous:1905:IIv

[7846] Anonymous. Index of inventions. Scientific American, 92(6): 130–135, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
130c.pdf.

Anonymous:1905:PNPr

[7847] Anonymous. Patent notes: Patents. Scientific American, 92 (6):135, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
135.pdf.

Anonymous:1905:CCDb

[7848] Anonymous. Closing of the Croton Dam. Scientific American, 92(6):118, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
118.pdf.

Anonymous:1905:VTT

[7849] Anonymous. Value of the towing tank. Scientific American, 92 (6):118, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
118a.pdf.

Anonymous:1905:GIF

[7850] Anonymous. Growing interest in forest preservation. Scientific American, 92(6):118, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
118b.pdf.

Anonymous:1905:SPT

[7851] Anonymous. Strength of preserved timber. Scientific American, 92(6):118, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
118c.pdf.

Anonymous:1905:DTC

[7852] Anonymous. A design for turbine cruisers. Scientific American, 92(6):118, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
118d.pdf.

Anonymous:1905:ODH

[7853] Anonymous. Opening of Dinosaur Hall of the American Museum of Natural History. Scientific American, 92(6):118–119, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n6/pdf/scientificamerican02111905-118e.pdf.

Anonymous:1905:TTD

[7854] Anonymous. Twenty thousand dollars for new international race. Scientific American, 92(6):119, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
119.pdf.

Anonymous:1905:SNCa

[7855] Anonymous. Some new Cooper Hewitt inventions. Scientific American, 92(6):119, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
119a.pdf.

Anonymous:1905:PEC

[7856] Anonymous. Prizes for essays on cement. Scientific American, 92(6):119, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
119b.pdf.

Anonymous:1905:CEK

[7857] Anonymous. A compact and economical kerosene engine. Scientific American, 92(6):120, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
120.pdf.

Dinger:1905:SCU

[7858] H. C. Dinger. Scout Cruisers for the United States Navy. Scientific American, 92(6):120–121, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
120a.pdf.

Anonymous:1905:FDE

[7859] Anonymous. Fog dispersion by electricity. Scientific American, 92 (6):121–122, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
121.pdf.

Collins:1905:TAW

[7860] A. Frederick Collins. Trees as antennæ in wireless telegraphy. Scientific American, 92(6):123, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
123.pdf.

Anonymous:1905:SPW

[7861] Anonymous. The ski in peace and war. Scientific American, 92 (6):124–126, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
124.pdf.

Anonymous:1905:OPC

[7862] Anonymous. Ontario Power Company’s power plant at Niagara Falls — II. Scientific American, 92(6):126, February 11, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n6/pdf/scientificamerican02111905-126.pdf.

Anonymous:1905:NAR

[7863] Anonymous. New automobile records at the Florida Races. Scientific American, 92(6):127, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
127.pdf.

Anonymous:1905:PDf

[7864] Anonymous. Patent department. Scientific American, 92(6):128, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n6/pdf/scientificamerican02111905-128.pdf.

Anonymous:1905:ITW

[7865] Anonymous. An improved type of water-tube boiler. Scientific American, 92(6):128, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
128a.pdf.

Anonymous:1905:OId

[7866] Anonymous. Oddities in inventions. Scientific American, 92(6): 128, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
128b.pdf.

Anonymous:1905:BPWv

[7867] Anonymous. Business and personal wants. Scientific American, 92(6):130, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
130.pdf.

Anonymous:1905:IIGa

[7868] Anonymous. An improved incandescent gas lamp. Scientific American, 92(6):131, February 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n6/pdf/scientificamerican02111905-
131.pdf.

Anonymous:1905:Ct

[7869] Anonymous. Correspondence. Scientific American, 92(7):143, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n7/pdf/scientificamerican02181905-143.pdf.

Anonymous:1905:RPIw

[7870] Anonymous. Recently patented inventions. Scientific American, 92 (7):149–150, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
149.pdf.

Anonymous:1905:NQw

[7871] Anonymous. Notes and queries. Scientific American, 92(7):150– 151, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
150a.pdf.

Anonymous:1905:BNBq

[7872] Anonymous. Books: New books, etc. Scientific American, 92 (7):151, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
151.pdf.

Anonymous:1905:IIw

[7873] Anonymous. Index of inventions. Scientific American, 92(7): 152–155, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
152.pdf.

Anonymous:1905:PNPs

[7874] Anonymous. Patent notes: Patents. Scientific American, 92 (7):155, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
155.pdf.

Anonymous:1905:ON

[7875] Anonymous. The old and the new. Scientific American, 92 (7):138, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
138.pdf.

Anonymous:1905:WPBb

[7876] Anonymous. Wind pressure on bridges and buildings. Scientific American, 92(7):138, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
138a.pdf.

Anonymous:1905:PRK

[7877] Anonymous. The proposed railroad to Key West. Scientific American, 92(7):139, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
139.pdf.

Meynell:1905:GEF

[7878] G. Meynell. Gas engines for factory power. Scientific American, 92(7):139, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
139a.pdf.

Anonymous:1905:CSe

[7879] Anonymous. The current supplement. Scientific American, 92 (7):139, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
139b.pdf.

Anonymous:1905:PDP

[7880] Anonymous. A patent dedicated to the public. Scientific American, 92(7):139, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
139c.pdf.

Anonymous:1905:TOT

[7881] Anonymous. Transformer outfits for thawing pipes. Scientific American, 92(7):140, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
140.pdf.

Anonymous:1905:NTL

[7882] Anonymous. A new type of lifeboat. Scientific American, 92 (7):140, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
140a.pdf.

Guarini:1905:ER

[7883] Emile Guarini. Electrical radiators. Scientific American, 92(7): 141, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
141.pdf.

Anonymous:1905:AA

[7884] Anonymous. An ambulance automobile. Scientific American, 92 (7):141–142, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
141a.pdf.

Anonymous:1905:EEN

[7885] Anonymous. Electrical equipment of the New York Central Railroad. Scientific American, 92(7):142, February 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n7/pdf/scientificamerican02181905-142.pdf.

Beasley:1905:USN

[7886] Walter L. Beasley. The United States Naval Training School at Newport. Scientific American, 92(7):144, February 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n7/pdf/scientificamerican02181905-144.pdf.

Anonymous:1905:OFH

[7887] Anonymous. The origin and form of hoarfrost. Scientific American, 92(7):145–146, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
145.pdf.

Anonymous:1905:WWK

[7888] Anonymous. What we know about Sun spots. Scientific American, 92(7):147, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
147.pdf.

Anonymous:1905:AFLa

[7889] Anonymous. Another form of looping the loop. Scientific American, 92(7):147, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
147a.pdf.

Anonymous:1905:PDg

[7890] Anonymous. Patent department. Scientific American, 92(7):148, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n7/pdf/scientificamerican02181905-148.pdf.

Anonymous:1905:NMT

[7891] Anonymous. A novel magazine tack hammer. Scientific American, 92(7):148, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
148a.pdf.

Anonymous:1905:IAc

[7892] Anonymous. An improved amalgamator. Scientific American, 92 (7):148, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
148b.pdf.

Anonymous:1905:NTG

[7893] Anonymous. A new type of grate. Scientific American, 92(7): 148, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
148c.pdf.

Anonymous:1905:BPWw

[7894] Anonymous. Business and personal wants. Scientific American, 92(7):150, February 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n7/pdf/scientificamerican02181905-
150.pdf.

Anonymous:1905:RPIx

[7895] Anonymous. Recently patented inventions. Scientific American, 92(8):169, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
169.pdf.

Anonymous:1905:NQx

[7896] Anonymous. Notes and queries. Scientific American, 92(8):169, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n8/pdf/scientificamerican02251905-169b.pdf.

Anonymous:1905:IIx

[7897] Anonymous. Index of inventions. Scientific American, 92(8): 169–175, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
169c.pdf.

Anonymous:1905:PNPt

[7898] Anonymous. Patent notes: Patents. Scientific American, 92 (8):175, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
175.pdf.

Anonymous:1905:TTSa

[7899] Anonymous. Turbines for tramp steamers. Scientific American, 92(8):158, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
158.pdf.

Anonymous:1905:TMR

[7900] Anonymous. Three months of railroad slaughter. Scientific American, 92(8):158, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
158a.pdf.

Anonymous:1905:FWS

[7901] Anonymous. Future water supply of New York City. Scientific American, 92(8):158–159, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
158b.pdf.

Brown:1905:HAF

[7902] Enos Brown. The Heaton Airship Failure. Scientific American, 92(8):159, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
159.pdf.

Anonymous:1905:MHD

[7903] Anonymous. The Moore–Heskett direct process for the manufacture of wrought iron and steel. Scientific American, 92(8):159, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n8/pdf/scientificamerican02251905-159a.pdf.

Anonymous:1905:CSg

[7904] Anonymous. The current supplement. Scientific American, 92 (8):159, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
159b.pdf.

Anonymous:1905:ES

[7905] Anonymous. Explosion of a submarine. Scientific American, 92 (8):159, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
159c.pdf.

Anonymous:1905:CAE

[7906] Anonymous. The Charcot Antarctic Expedition. Scientific American, 92(8):159, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
159d.pdf.

Anonymous:1905:HMM

[7907] Anonymous. How to make a mercury thermometer. Scientific American, 92(8):160, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
160.pdf.

Guarini:1905:TWB

[7908] Emile Guarini. A tachometric watch for bicycles and other vehicles. Scientific American, 92(8):160, February 25, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n8/pdf/scientificamerican02251905-160a.pdf.

Inkersley:1905:BDH

[7909] Arthur Inkersley. A blast of dynamite and how it was photographed. Scientific American, 92(8):161, February 25, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n8/pdf/scientificamerican02251905-161.pdf.

Anonymous:1905:NKB

[7910] Anonymous. New 25-knot British Scouts. Scientific American, 92 (8):161–162, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
161a.pdf.

Kelvin:1905:MSE

[7911] Lord Kelvin. My scientific education. Scientific American, 92 (8):162, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
162.pdf.

Hopkins:1905:ECR

[7912] N. Monroe Hopkins. An easily-constructed rotating anode equipment for rapid electro-analysis. Scientific American, 92(8):163, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n8/pdf/scientificamerican02251905-163.pdf.

Anonymous:1905:HEM

[7913] Anonymous. A huge electro-magnet for optical and surgical work. Scientific American, 92(8):163, February 25, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n8/pdf/scientificamerican02251905-163a.pdf.

Willey:1905:RMA

[7914] Day Allen Willey. Rocky mountain avalanches. Scientific American, 92(8):164, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
164.pdf.

Brassler:1905:AWF

[7915] Charles A. Brassler. Antique watches of the famous Marfels Collection. Scientific American, 92(8):165–166, February 25, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n8/pdf/scientificamerican02251905-165.pdf.

Beard:1905:QCI

[7916] J. Carter Beard. The queerest of creatures — II. Scientific American, 92(8):167, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
167.pdf.

Anonymous:1905:IHR

[7917] Anonymous. An improved hay rack. Scientific American, 92 (8):168, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
168.pdf.

Anonymous:1905:SCF

[7918] Anonymous. A simple car fender. Scientific American, 92(8):168, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n8/pdf/scientificamerican02251905-168a.pdf.

Anonymous:1905:NTEb

[7919] Anonymous. A new type of electric coupé. Scientific American, 92(8):168, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
168b.pdf.

Anonymous:1905:NGR

[7920] Anonymous. A novel gravity railway system. Scientific American, 92(8):168, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
168c.pdf.

Anonymous:1905:BPWx

[7921] Anonymous. Business and personal wants. Scientific American, 92(8):169, February 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n8/pdf/scientificamerican02251905-
169a.pdf.

Russell:1905:HMb

[7922] Henry Norris Russell. The heavens in March. Scientific American, 92(9):179, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
179.pdf.

Anonymous:1905:Cu

[7923] Anonymous. Correspondence. Scientific American, 92(9):183, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-183.pdf.

Anonymous:1905:RPIy (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
189.pdf.

Anonymous:1905:NQy

[7925] Anonymous. Notes and queries. Scientific American, 92(9):190, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-190a.pdf.

Anonymous:1905:BNBr

[7926] Anonymous. Books: New books, etc. Scientific American, 92(9):190, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-190b.pdf.

Anonymous:1905:IIy

[7927] Anonymous. Index of inventions. Scientific American, 92(9):190–195, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-190c.pdf.

Anonymous:1905:PNPu

[7928] Anonymous. Patent notes: Patents. Scientific American, 92(9):195, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-195.pdf.

Anonymous:1905:SCSa

[7929] Anonymous. Steel cars and safe travel. Scientific American, 92(9):178, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-178.pdf.

Anonymous:1905:MMEa

[7930] Anonymous. The military method of educating boys. Scientific American, 92(9):178, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
178a.pdf.

Anonymous:1905:FAQ

[7931] Anonymous. The free-alcohol question. Scientific American, 92(9):178, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-178b.pdf.

Anonymous:1905:NTMb

[7932] Anonymous. New Trade-Mark Law of April 1, 1905. Scientific American, 92(9):178–179, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
178c.pdf.

Putnam:1905:GB

[7933] B. L. Putnam. Grandfather’s barometer. Scientific American, 92 (9):179, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
179a.pdf.

Anonymous:1905:SNm

[7934] Anonymous. Science notes. Scientific American, 92(9):179, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-179b.pdf.

Wright:1905:LFW

[7935] H. E. Wright. The largest ferryboat in the world. Scientific American, 92(9):180, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
180.pdf.

Anonymous:1905:CMPa

[7936] Anonymous. The Capitaine marine producer-gas engine. Scientific American, 92(9):180, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
180a.pdf.

Anonymous:1905:NMR

[7937] Anonymous. A new meteorite recently placed in the American Museum of Natural History. Scientific American, 92(9):181, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-181.pdf.

Willey:1905:MMG

[7938] Day Allen Willey. Modern methods of gold prospecting and mining. Scientific American, 92(9):181–182, March 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v92/
n9/pdf/scientificamerican03041905-181a.pdf.

Gignilliat:1905:EBM

[7939] R. L. Gignilliat. Education of boys by the military method. Scientific American, 92(9):184, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
184.pdf.

Anonymous:1905:NMI

[7940] Anonymous. A novel motor ice boat. Scientific American, 92(9):186, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-186.pdf.

Anonymous:1905:HA

[7941] Anonymous. The Heaton airship. Scientific American, 92(9):186, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v92/n9/pdf/scientificamerican03041905-186a.pdf.

Anonymous:1905:MBE

[7942] Anonymous. Motor boat engines at the Sportsmen’s Show. Scientific American, 92(9):186–187, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
186b.pdf.

Christison:1905:CFD

[7943] Sanderson Christison. Curious fishes of the deep sea. Scientific American, 92(9):188, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
188.pdf.

Anonymous:1905:BPWy

[7944] Anonymous. Business and personal wants. Scientific American, 92(9):190, March 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v92/n9/pdf/scientificamerican03041905-
190.pdf.

Russell:1905:HJb

[7945] Henry Norris Russell. The heavens in July. Scientific American, 93(1): 5, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-5.pdf.

Anonymous:1905:Cv

[7946] Anonymous. Correspondence. Scientific American, 93(1):8–9, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-8.pdf.

Anonymous:1905:RPIz

[7947] Anonymous. Recently patented inventions. Scientific American, 93(1): 14, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-14.pdf.

Anonymous:1905:NQz

[7948] Anonymous. Notes and queries. Scientific American, 93(1):14, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-14b.pdf.

Anonymous:1905:BNBs

[7949] Anonymous. Books: New books, etc. Scientific American, 93(1):15–16, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-15.pdf.

Anonymous:1905:IIz

[7950] Anonymous. Index of inventions. Scientific American, 93(1):16–19, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-16.pdf.

Anonymous:1905:PNPv

[7951] Anonymous. Patent notes: Patents. Scientific American, 93(1):19, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-19.pdf.

Anonymous:1905:AHG

[7952] Anonymous. American homes and gardens — the new magazine. Scientific American, 93(1):4, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n1/pdf/scientificamerican07011905-
4.pdf.

Anonymous:1905:ACS

[7953] Anonymous. Atmospheric conditions in the subway. Scientific American, 93(1):4, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n1/pdf/scientificamerican07011905-
4a.pdf.

Anonymous:1905:CTa

[7954] Anonymous. A city of towers. Scientific American, 93(1):4, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-4b.pdf.

Anonymous:1905:FTDa

[7955] Anonymous. Are fast trains dangerous? Scientific American, 93(1):4–5, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-4c.pdf.

Anonymous:1905:ENc

[7956] Anonymous. Engineering notes. Scientific American, 93(1):5, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-5a.pdf.

Geare:1905:EUL

[7957] Randolph I. Geare. The Englehardt unsinkable lifeboat. Scientific American, 93(1):6, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n1/pdf/scientificamerican07011905-
6.pdf.

Guarini:1905:AIV

[7958] Emile Guarini. An apparatus for indicating the vibrations of sound waves. Scientific American, 93(1):6–7, July 1, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n1/pdf/scientificamerican07011905-6a.pdf.

Anonymous:1905:BMA

[7959] Anonymous. The bird as a model for the aeroplane. Scientific American, 93(1):7–8, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n1/pdf/scientificamerican07011905-
7.pdf.

Anonymous:1905:EF

[7960] Anonymous. Exploring for fossils. Scientific American, 93(1):9–10, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-9.pdf.

Anonymous:1905:WRB

[7961] Anonymous. The Wachusett Reservoir for Boston Water Supply. Scientific American, 93(1):11–12, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n1/pdf/scientificamerican07011905-
11.pdf.

Perkins:1905:PNY

[7962] Frank C. Perkins. Photographing the New York Subway. Scientific American, 93(1):12, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n1/pdf/scientificamerican07011905-
12.pdf.

Anonymous:1905:EPF

[7963] Anonymous. An efficient portable fire escape. Scientific American, 93(1):13, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n1/pdf/scientificamerican07011905-
13.pdf.

Anonymous:1905:NTP

[7964] Anonymous. A novel tobacco pipe. Scientific American, 93(1):13, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-13a.pdf.

Anonymous:1905:IFC

[7965] Anonymous. Improved folding coop. Scientific American, 93(1):13, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-13b.pdf.

Anonymous:1905:IC

[7966] Anonymous. An improved carbureter. Scientific American, 93(1):13, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-13c.pdf.

Anonymous:1905:BPWz

[7967] Anonymous. Business and personal wants. Scientific American, 93(1): 14, July 1, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n1/pdf/scientificamerican07011905-14a.pdf.

Russell:1905:HS

[7968] Henry Norris Russell. The heavens in September. Scientific American, 93(10):175, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
175.pdf.

Anonymous:1905:Cw

[7969] Anonymous. Correspondence. Scientific American, 93(10):176, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n10/pdf/scientificamerican09021905-179.pdf.

Anonymous:1905:RPIba

[7970] Anonymous. Recently patented inventions. Scientific American, 93(10):185, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
185.pdf.

Anonymous:1905:NQba

[7971] Anonymous. Notes and queries. Scientific American, 93(10): 185–186, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
185b.pdf.

Anonymous:1905:BNBt

[7972] Anonymous. Books: New books, etc. Scientific American, 93 (10):186, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
186.pdf.

Anonymous:1905:IIba

[7973] Anonymous. Index of inventions. Scientific American, 93(10): 186–191, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
186a.pdf.

Anonymous:1905:PNPw

[7974] Anonymous. Patent notes: Patents. Scientific American, 93 (10):191, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
191.pdf.

Anonymous:1905:MPT

[7975] Anonymous. Maneuvering power of turbine steamships. Scientific American, 93(10):174, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
174.pdf.

Anonymous:1905:SCSb

[7976] Anonymous. Steel cars and safe travel on steam railroads. Scientific American, 93(10):174, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
174a.pdf.

Anonymous:1905:RHE

[7977] Anonymous. Railroad housekeeping economies. Scientific American, 93(10):174–175, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
174b.pdf.

Anonymous:1905:LG

[7978] Anonymous. Lumber grades. Scientific American, 93(10):175, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n10/pdf/scientificamerican09021905-175a.pdf.

Anonymous:1905:SNn

[7979] Anonymous. Science notes. Scientific American, 93(10):175, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n10/pdf/scientificamerican09021905-175b.pdf.

Anonymous:1905:CSi

[7980] Anonymous. The current supplement. Scientific American, 93 (10):175, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
175c.pdf.

Mansfield:1905:NSP

[7981] Charles Monroe Mansfield. Natural-size photography. Scientific American, 93(10):176, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
176.pdf.

Anonymous:1905:GD

[7982] Anonymous. A giant dredge. Scientific American, 93(10):176, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n10/pdf/scientificamerican09021905-176a.pdf.

Crafts:1905:TTG

[7983] H. A. Crafts. Time test of a great irrigating system. Scientific American, 93(10):177, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
177.pdf.

Adams:1905:RNF

[7984] Alton D. Adams. Recession of Niagara Fails. Scientific American, 93(10):178, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
178.pdf.

Field:1905:GG

[7985] C. Field. Grenades and grenadiers. Scientific American, 93(10): 179–180, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
179a.pdf.

Anonymous:1905:KAE

[7986] Anonymous. Knabenshue’s airship and its exploits. Scientific American, 93(10):181, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
181.pdf.

Anonymous:1905:MSC

[7987] Anonymous. Motor sprinkler at Cologne. Scientific American, 93 (10):182, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
182.pdf.

Guarini:1905:SW

[7988] Emile Guarini. Sterilization of water. Scientific American, 93(10): 182–184, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
182a.pdf.

Anonymous:1905:IAS

[7989] Anonymous. An improved ash sifter. Scientific American, 93 (10):184, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
184.pdf.

Anonymous:1905:PE

[7990] Anonymous. A pneumatic eraser. Scientific American, 93(10): 184, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
184a.pdf.

Anonymous:1905:NT

[7991] Anonymous. A new toy. Scientific American, 93(10):184, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n10/pdf/scientificamerican09021905-184b.pdf.

Anonymous:1905:BPWba

[7992] Anonymous. Business and personal wants. Scientific American, 93(10):185, September 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n10/pdf/scientificamerican09021905-
185a.pdf.

Anonymous:1905:Cx

[7993] Anonymous. Correspondence. Scientific American, 93(11):198, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n11/pdf/scientificamerican09091905-198.pdf.

Anonymous:1905:RPIbb

[7994] Anonymous. Recently patented inventions. Scientific American, 93 (11):205–206, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
205.pdf.

Anonymous:1905:NQbb

[7995] Anonymous. Notes and queries. Scientific American, 93(11): 206, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
206a.pdf.

Anonymous:1905:IIbb

[7996] Anonymous. Index of inventions. Scientific American, 93(11): 207–211, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
207a.pdf.

Anonymous:1905:PNPx

[7997] Anonymous. Patent notes: Patents. Scientific American, 93 (11):211, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
211.pdf.

Anonymous:1905:BMS

[7998] Anonymous. The Budapest and Manhattan Suspension Bridges. Scientific American, 93(11):194, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
194.pdf.

Anonymous:1905:NLR

[7999] Anonymous. Naval losses of the Russo-Japanese War. Scientific American, 93(11):194, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
194a.pdf.

Anonymous:1905:PFE

[8000] Anonymous. Peace in the Far East. Scientific American, 93(11): 194–195, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
194b.pdf.

Geare:1905:PPS

[8001] Randolph I. Geare. Pearls and pearl shells. Scientific American, 93(11):195, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
195.pdf.

Anonymous:1905:TTB

[8002] Anonymous. Trial trip of the Barton airship. Scientific American, 93(11):196–197, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
196.pdf.

Anonymous:1905:EBB

[8003] Anonymous. The Elizabeth Bridge at Budapest. Scientific American, 93(11):197–198, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
197.pdf.

Ramakers:1905:LTM

[8004] L. Ramakers. The Laboratory for the Testing of Materials of the Charlottenburg Polytechnic School of Berlin. Scientific American, 93(11):199, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
199.pdf.

Owens:1905:OI

[8005] I. B. Owens. The oyster industry. Scientific American, 93(11): 200–201, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
200.pdf.

Anonymous:1905:FC

[8006] Anonymous. A fireless cooker. Scientific American, 93(11):204, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n11/pdf/scientificamerican09091905-204.pdf.

Collins:1905:APR

[8007] A. Frederick Collins. Artificial production of real diamonds. Scientific American, 93(11):204–205, September 9, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n11/pdf/scientificamerican09091905-204a.pdf.

Anonymous:1905:BPWbb

[8008] Anonymous. Business and personal wants. Scientific American, 93(11):206, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n11/pdf/scientificamerican09091905-
206.pdf.

Anonymous:1905:NBE

[8009] Anonymous. New books, etc. Scientific American, 93(11):207, September 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n11/pdf/scientificamerican09091905-207.pdf.

Anonymous:1905:Cy

[8010] Anonymous. Correspondence. Scientific American, 93(12):219, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n12/pdf/scientificamerican09161905-219.pdf.

Anonymous:1905:RPIbc

[8011] Anonymous. Recently patented inventions. Scientific American, 93 (12):225, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
225.pdf.

Anonymous:1905:NQbc

[8012] Anonymous. Notes and queries. Scientific American, 93(12): 225–226, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
225b.pdf.

Anonymous:1905:BNBu

[8013] Anonymous. Books: New books, etc. Scientific American, 93(12): 226–227, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
226.pdf.

Anonymous:1905:IIbc

[8014] Anonymous. Index of inventions. Scientific American, 93(12): 227–231, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
227.pdf.

Anonymous:1905:PNPy

[8015] Anonymous. Patent notes: Patents. Scientific American, 93 (12):231, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
231.pdf.

Anonymous:1905:PPC

[8016] Anonymous. Planning the Panama Canal. Scientific American, 93 (12):214, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
214.pdf.

Anonymous:1905:FCW

[8017] Anonymous. Filtration for Croton water. Scientific American, 93 (12):214, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
214a.pdf.

Anonymous:1905:ORS

[8018] Anonymous. Our railroad system. Scientific American, 93(12): 214, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
214b.pdf.

Anonymous:1905:PWT

[8019] Anonymous. The progress of wireless telegraphy in Great Britain. Scientific American, 93(12):214–215, September 16, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n12/pdf/scientificamerican09161905-214c.pdf.

Anonymous:1905:BSB

[8020] Anonymous. The buoyancy of submarine boats. Scientific American, 93(12):215, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
215.pdf.

Anonymous:1905:RRB

[8021] Anonymous. Ramsay, radium, and Burke. Scientific American, 93 (12):215, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
215a.pdf.

Anonymous:1905:LA

[8022] Anonymous. Lost arts. Scientific American, 93(12):215, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n12/pdf/scientificamerican09161905-215b.pdf.

Anonymous:1905:CSj

[8023] Anonymous. The current supplement. Scientific American, 93 (12):215, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
215c.pdf.

Anonymous:1905:SNo

[8024] Anonymous. Science notes. Scientific American, 93(12):215, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n12/pdf/scientificamerican09161905-215d.pdf.

Bennett:1905:TCI

[8025] Herbert I. Bennett. The Truckee–Carson Irrigation Project. Scientific American, 93(12):216–218, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
216.pdf.

Anonymous:1905:OSE

[8026] Anonymous. Observation of solar eclipse at sea. Scientific American, 93(12):218–219, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
218.pdf.

Brown:1905:LBP

[8027] Enos Brown. Luther Burbank and plant breeding. Scientific American, 93(12):220–221, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
220.pdf.

Anonymous:1905:GCS

[8028] Anonymous. The Grand Central Station Terminal improvements. Scientific American, 93(12):222–223, September 16, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n12/pdf/scientificamerican09161905-222.pdf.

Sawyer:1905:AR

[8029] E. O. Sawyer. An aerial rowboat. Scientific American, 93(12): 223, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
223.pdf.

Inkersley:1905:RAF

[8030] Arthur Inkersley. A respiratory apparatus for firemen. Scientific American, 93(12):224, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
224.pdf.

Anonymous:1905:MPE

[8031] Anonymous. The most powerful express locomotive in Great Britain. Scientific American, 93(12):224, September 16, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n12/pdf/scientificamerican09161905-224a.pdf.

Anonymous:1905:BPWbc

[8032] Anonymous. Business and personal wants. Scientific American, 93 (12):225, September 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n12/pdf/scientificamerican09161905-
225a.pdf.

Anonymous:1905:Cz

[8033] Anonymous. Correspondence. Scientific American, 93(13):239, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n13/pdf/scientificamerican09231905-239a.pdf.

Anonymous:1905:RPIbd

[8034] Anonymous. Recently patented inventions. Scientific American, 93 (13):246, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
246.pdf.

Anonymous:1905:NQbd

[8035] Anonymous. Notes and queries. Scientific American, 93(13): 246–247, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
246b.pdf.

Anonymous:1905:BNBv

[8036] Anonymous. Books: New books, etc. Scientific American, 93 (13):247, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
247.pdf.

Anonymous:1905:IIbd

[8037] Anonymous. Index of inventions. Scientific American, 93(13): 247–251, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
247a.pdf.

Anonymous:1905:PNPz

[8038] Anonymous. Patent notes: Patents. Scientific American, 93 (13):251, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
251.pdf.

Anonymous:1905:BA

[8039] Anonymous. The bicycle and the automobile. Scientific American, 93(13):234, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
234.pdf.

Anonymous:1905:LFM

[8040] Anonymous. Loss of the flagship “Mikasa”. Scientific American, 93(13):234, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
234a.pdf.

Anonymous:1905:RDE

[8041] Anonymous. Risk of derailment on elevated railroad curves. Scientific American, 93(13):234, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
234b.pdf.

Anonymous:1905:IITa

[8042] Anonymous. An important innovation in telegraphy. Scientific American, 93(13):234–235, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
234c.pdf.

Tait:1905:EHG

[8043] George M. S. Tait. The effect of hydrogen on gas engine compression. Scientific American, 93(13):235, September 23, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n13/pdf/scientificamerican09231905-235.pdf.

Anonymous:1905:P

[8044] Anonymous. The phonocard. Scientific American, 93(13):235, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n13/pdf/scientificamerican09231905-235a.pdf.

Anonymous:1905:CSk

[8045] Anonymous. The current supplement. Scientific American, 93 (13):235, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
235b.pdf.

Anonymous:1905:PHV

[8046] Anonymous. Producing high vacua. Scientific American, 93(13): 235, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
235c.pdf.

Anonymous:1905:ERW

[8047] Anonymous. The elevated railroad wreck in New York. Scientific American, 93(13):236, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
236.pdf.

Anonymous:1905:ONG

[8048] Anonymous. Our Naval guns in the Civil War and to-day. Scientific American, 93(13):237, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
237.pdf.

Anonymous:1905:ECB

[8049] Anonymous. Electrical control of bulkhead doors on warships. Scientific American, 93(13):237–238, September 23, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n13/pdf/scientificamerican09231905-237a.pdf.

Anonymous:1905:FPG

[8050] Anonymous. The first producer-gas boat. Scientific American, 93 (13):238–239, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
238.pdf.

Anonymous:1905:GBB

[8051] Anonymous. A Government-built battleship. Scientific American, 93(13):239, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
239.pdf.

Willey:1905:HSD

[8052] Day Allen Willey. Hydraulic suction dredge on the Mississippi. Scientific American, 93(13):240, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
240.pdf.

Maye:1905:NSO

[8053] M. Benedict Maye. The new Solar Observatory at Mount Wilson. Scientific American, 93(13):240–242, September 23, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n13/pdf/scientificamerican09231905-240a.pdf.

Anonymous:1905:ARCb

[8054] Anonymous. Ancient reservoirs at Carthage. Scientific American, 93 (13):242–243, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
242.pdf.

Holder:1905:IS

[8055] Charles Frederick Holder. An interesting shark. Scientific American, 93(13):243, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
243.pdf.

Anonymous:1905:PDh

[8056] Anonymous. Patent department. Scientific American, 93(13): 244, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
244.pdf.

Anonymous:1905:NCW

[8057] Anonymous. A novel carriage-washing apparatus. Scientific American, 93(13):244, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
244a.pdf.

Anonymous:1905:NWB

[8058] Anonymous. New warp-beam tension devices for looms. Scientific American, 93(13):244–245, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
244b.pdf.

Anonymous:1905:NRVa

[8059] Anonymous. New rotary valve for steam engines. Scientific American, 93(13):245, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
245.pdf.

Anonymous:1905:SIAa

[8060] Anonymous. A striking illusion apparatus. Scientific American, 93 (13):245, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
245a.pdf.

Anonymous:1905:SWT

[8061] Anonymous. Sand wheel toy. Scientific American, 93(13):245, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n13/pdf/scientificamerican09231905-245b.pdf.

Anonymous:1905:NWI

[8062] Anonymous. A new wind instrument. Scientific American, 93 (13):245, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
245c.pdf.

Anonymous:1905:BPWbd

[8063] Anonymous. Business and personal wants. Scientific American, 93 (13):246, September 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n13/pdf/scientificamerican09231905-
246a.pdf.

Anonymous:1905:Cba

[8064] Anonymous. Correspondence. Scientific American, 93(14):259, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n14/pdf/scientificamerican09301905-259.pdf.

Anonymous:1905:RPIbe

[8065] Anonymous. Recently patented inventions. Scientific American, 93 (14):266, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
266.pdf.

Anonymous:1905:NQbe

[8066] Anonymous. Notes and queries. Scientific American, 93(14): 266–267, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
266b.pdf.

Anonymous:1905:BNBw

[8067] Anonymous. Books: New books, etc. Scientific American, 93(14): 267–268, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
267.pdf.

Anonymous:1905:IIbe

[8068] Anonymous. Index of inventions. Scientific American, 93(14): 268–271, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
268.pdf.

Anonymous:1905:PNPba

[8069] Anonymous. Patent notes: Patents. Scientific American, 93 (14):271, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
271.pdf.

Anonymous:1905:BEB

[8070] Anonymous. Building the Erie Barge Canal. Scientific American, 93(14):254, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
254.pdf.

Anonymous:1905:ANR

[8071] Anonymous. Another North River Tunnel. Scientific American, 93 (14):254, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
254a.pdf.

Anonymous:1905:MSNb

[8072] Anonymous. Magnetic survey of the North Pacific Ocean. Scientific American, 93(14):254, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
254b.pdf.

Anonymous:1905:CVC

[8073] Anonymous. Chains versus cables in the Manhattan Bridge. Scientific American, 93(14):254, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
254c.pdf.

Anonymous:1905:TMC

[8074] Anonymous. Trade marks on copyrighted works. Scientific American, 93(14):254–255, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
254d.pdf.

Anonymous:1905:VCEa

[8075] Anonymous. Vanderbilt Cup Elimination Trials. Scientific American, 93(14):255, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
255.pdf.

Darwin:1905:RCS

[8076] G. H. Darwin. Is radium the cause of the Sun’s heat and light? Scientific American, 93(14):255, September 30, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n14/pdf/scientificamerican09301905-255a.pdf.

Anonymous:1905:PEA

[8077] Anonymous. Patents for exported articles — practical advice to manufacturers. Scientific American, 93(14):255, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n14/pdf/scientificamerican09301905-255b.pdf.

Anonymous:1905:CSl

[8078] Anonymous. The current supplement. Scientific American, 93 (14):255, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
255c.pdf.

Anonymous:1905:ET

[8079] Anonymous. An early typewriter. Scientific American, 93(14): 256, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
256.pdf.

Anonymous:1905:PMM

[8080] Anonymous. Pie-making by machinery. Scientific American, 93 (14):256, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
256a.pdf.

Anonymous:1905:HSS

[8081] Anonymous. High-speed steam yacht and power-boat races. Scientific American, 93(14):257–258, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
257.pdf.

Anonymous:1905:TSLa

[8082] Anonymous. 180-ton shear legs for the British Naval Dockyard. Scientific American, 93(14):258, September 30, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n14/pdf/scientificamerican09301905-258.pdf.

Anonymous:1905:HIR

[8083] Anonymous. Heliotropism induced by radium. Scientific American, 93(14):259, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
259a.pdf.

Crafts:1905:ECC

[8084] H. A. Crafts. Evolution of the California clam-shell dredger. Scientific American, 93(14):260, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
260.pdf.

Anonymous:1905:MF

[8085] Anonymous. The making of a felt hat. Scientific American, 93 (14):260–262, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
260a.pdf.

Anonymous:1905:CRA

[8086] Anonymous. Combined rack and adhesion locomotive for South Africa. Scientific American, 93(14):262, September 30, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n14/pdf/scientificamerican09301905-262.pdf.

Tingley:1905:LAC

[8087] Charles Edward Tingley. Littlefield and the artificial creation of life. Scientific American, 93(14):263, September 30, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n14/pdf/scientificamerican09301905-263.pdf.

Ramakers:1905:NMI

[8088] L. Ramakers. A new method of identifying criminals. Scientific American, 93(14):264, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
264.pdf.

Gradenwitz:1905:NPS

[8089] Alfred Gradenwitz. A new projection system for lectures. Scientific American, 93(14):265, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
265.pdf.

Guarini:1905:NAA

[8090] Emile Guarini. A new apparatus for automatically making and breaking an electric circuit. Scientific American, 93(14):265, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n14/pdf/scientificamerican09301905-265a.pdf.

Anonymous:1905:NDP

[8091] Anonymous. A novel device for preventing automobiles from skidding. Scientific American, 93(14):265, September 30, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n14/pdf/scientificamerican09301905-265b.pdf.

Anonymous:1905:BPWbe

[8092] Anonymous. Business and personal wants. Scientific American, 93 (14):266, September 30, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n14/pdf/scientificamerican09301905-
266a.pdf.

Russell:1905:HO

[8093] Henry Norris Russell. The heavens in October. Scientific American, 93(15):275, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
275.pdf.

Anonymous:1905:Cbb

[8094] Anonymous. Correspondence. Scientific American, 93(15):279, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-279.pdf.

Anonymous:1905:RPIbf

[8095] Anonymous. Recently patented inventions. Scientific American, 93(15):285, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
285.pdf.

Anonymous:1905:NQbf

[8096] Anonymous. Notes and queries. Scientific American, 93(15):286–287, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-286.pdf.

Anonymous:1905:IIbf

[8097] Anonymous. Index of inventions. Scientific American, 93(15):287–291, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-287.pdf.

Anonymous:1905:PNPbb

[8098] Anonymous. Patent notes: Patents. Scientific American, 93(15):291, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-291.pdf.

Anonymous:1905:TEC

[8099] Anonymous. Two engineers in the cab. Scientific American, 93(15):274, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-274.pdf.

Anonymous:1905:CTR

[8100] Anonymous. Comparison of turbines and reciprocating marine engines in service. Scientific American, 93(15):274, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-274a.pdf.

Anonymous:1905:PCF

[8101] Anonymous. Panama Canal in four years. Scientific American, 93(15):274, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
274b.pdf.

Anonymous:1905:RAS

[8102] Anonymous. Recent advances in selective wireless telegraphy. Scientific American, 93(15):274–275, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
274c.pdf.

Anonymous:1905:SNp

[8103] Anonymous. Science notes. Scientific American, 93(15):275, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-275a.pdf.

Anonymous:1905:SAAa

[8104] Anonymous. Something about alligators. Scientific American, 93(15):276, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
276.pdf.

Gradenwitz:1905:NPR

[8105] Alfred Gradenwitz. A novel process of reanimation. Scientific American, 93(15):276–277, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
276a.pdf.

Mclure:1905:BM

[8106] W. Frank M’clure. Brick making. Scientific American, 93(15):277–278, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-277.pdf.

Anonymous:1905:AOA

[8107] Anonymous. An apparatus for observing and automatically registering thunderstorms. Scientific American, 93(15):278, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-278.pdf.

Anonymous:1905:VCEb

[8108] Anonymous. The Vanderbilt Cup Elimination Race. Scientific American, 93(15):280–282, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
280.pdf.

Bellet:1905:GBN

[8109] Daniel Bellet. Gliding boats — the navigation of the future. Scientific American, 93(15):283–284, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
283.pdf.

Anonymous:1905:ASM

[8110] Anonymous. An automatic switch for making and breaking the primary circuit of transformers. Scientific American, 93(15):284, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n15/pdf/scientificamerican10071905-284.pdf.

Anonymous:1905:CBE

[8111] Anonymous. Circuit-breaker for explosive engines. Scientific American, 93(15):284, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
284a.pdf.

Anonymous:1905:BPWbf

[8112] Anonymous. Business and personal wants. Scientific American, 93(15):285, October 7, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n15/pdf/scientificamerican10071905-
285a.pdf.

Anonymous:1905:Cbc

[8113] Anonymous. Correspondence. Scientific American, 93(16):298, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n16/pdf/scientificamerican10141905-298.pdf.

Anonymous:1905:RPIbg

[8114] Anonymous. Recently patented inventions. Scientific American, 93 (16):305–306, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
305.pdf.

Anonymous:1905:NQbg

[8115] Anonymous. Notes and queries. Scientific American, 93(16): 306–307, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
306a.pdf.

Anonymous:1905:IIbg

[8116] Anonymous. Index of inventions. Scientific American, 93(16): 307–311, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
307.pdf.

Anonymous:1905:PNPbc

[8117] Anonymous. Patent notes: Patents. Scientific American, 93 (16):311, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
311.pdf.

Anonymous:1905:FFW

[8118] Anonymous. Filtration and the future water supply for New York City. Scientific American, 93(16):294, October 14, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n16/pdf/scientificamerican10141905-294.pdf.

Anonymous:1905:PRT

[8119] Anonymous. Present railroad track inefficient. Scientific American, 93(16):294, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
294a.pdf.

Anonymous:1905:NFQ

[8120] Anonymous. The need for Federal quarantine control. Scientific American, 93(16):294, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
294b.pdf.

Spears:1905:ETB

[8121] John R. Spears. Evolution of the tow barge. Scientific American, 93(16):295, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
295.pdf.

Anonymous:1905:SNq

[8122] Anonymous. Science notes. Scientific American, 93(16):295, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n16/pdf/scientificamerican10141905-295a.pdf.

Wade:1905:HCT

[8123] Herbert T. Wade. How clinical thermometers are standardized and tested. Scientific American, 93(16):296, October 14, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n16/pdf/scientificamerican10141905-296.pdf.

Mclure:1905:EHG

[8124] W. Frank M’clure. The “Elbert H. Gary” — the greatest of orecarrying steamers. Scientific American, 93(16):297–298, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n16/pdf/scientificamerican10141905-297.pdf.

Dailey:1905:SVP

[8125] Howard B. Dailey. Stratification in vacuo: Its production with the influence machine. Scientific American, 93(16):299, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n16/pdf/scientificamerican10141905-299.pdf.

Anonymous:1905:APA

[8126] Anonymous. Aftermath of Port Arthur. Scientific American, 93 (16):300–301, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
300.pdf.

Anonymous:1905:SDN

[8127] Anonymous. The Santos-Dumont “No. 14”. Scientific American, 93(16):302, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
302.pdf.

Anonymous:1905:HWB

[8128] Anonymous. A hundred ways of breaking your neck. Scientific American, 93(16):302–304, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
302a.pdf.

White:1905:APE

[8129] W. P. White. Air pumps for experimental purposes. Scientific American, 93(16):304, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
304.pdf.

Anonymous:1905:VOW

[8130] Anonymous. Valve for oil wells. Scientific American, 93(16):304, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n16/pdf/scientificamerican10141905-304a.pdf.

Anonymous:1905:BPWbg

[8131] Anonymous. Business and personal wants. Scientific American, 93(16):306, October 14, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n16/pdf/scientificamerican10141905-
306.pdf.

Anonymous:1905:Cbd

[8132] Anonymous. Correspondence. Scientific American, 93(17):318–319, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n17/pdf/scientificamerican10211905-318.pdf.

Anonymous:1905:RPIbh

[8133] Anonymous. Recently patented inventions. Scientific American, 93(17):325, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
325.pdf.

Anonymous:1905:NQbh

[8134] Anonymous. Notes and queries. Scientific American, 93(17): 325–326, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
325b.pdf.

Anonymous:1905:IIbh

[8135] Anonymous. Index of inventions. Scientific American, 93(17): 326–331, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
326.pdf.

Anonymous:1905:PNPbd 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
331a.pdf.

Anonymous:1905:SPS

[8137] Anonymous. The single-phase system in America. Scientific American, 93(17):314, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
314.pdf.

Anonymous:1905:OTG

[8138] Anonymous. Our two great canal projects. Scientific American, 93(17):314, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
314a.pdf.

Anonymous:1905:VCR

[8139] Anonymous. Vanderbilt Cup Race. Scientific American, 93(17):314, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n17/pdf/scientificamerican10211905-314b.pdf.

Hodgins:1905:TM

[8140] George S. Hodgins. A temperance match. Scientific American, 93(17):315, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
315.pdf.

Wallin:1905:SID

[8141] J. E. Wallace Wallin. The size illusion of the depressed letter P. Scientific American, 93(17):315, October 21, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n17/pdf/scientificamerican10211905-315a.pdf.

Anonymous:1905:SNr

[8142] Anonymous. Science notes. Scientific American, 93(17):315, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n17/pdf/scientificamerican10211905-315b.pdf.

Anonymous:1905:DFM

[8143] Anonymous. The Dufaux flying machine. Scientific American, 93 (17):316–317, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
316.pdf.

Gradenwitz:1905:IRW

[8144] Alfred Gradenwitz. Interesting repair work on a German steamship. Scientific American, 93(17):317–318, October 21, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n17/pdf/scientificamerican10211905-317.pdf.

Anonymous:1905:NPD

[8145] Anonymous. A novel process for disengaging forces by means of sound. Scientific American, 93(17):319, October 21, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n17/pdf/scientificamerican10211905-319.pdf.

Anonymous:1905:EPDa

[8146] Anonymous. Electric power developments at Niagara Falls — II. Scientific American, 93(17):320–321, October 21, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n17/pdf/scientificamerican10211905-320.pdf.

Anonymous:1905:NMC

[8147] Anonymous. A new method of constructing sanitary sewers. Scientific American, 93(17):322, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
322.pdf.

Anonymous:1905:GIW

[8148] Anonymous. Giants of the insect world. Scientific American, 93 (17):323–324, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
323.pdf.

Anonymous:1905:UCG

[8149] Anonymous. A useful chain grip for automobiles. Scientific American, 93(17):324, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
324.pdf.

Gradenwitz:1905:BHL

[8150] Alfred Gradenwitz. A binocular head lens. Scientific American, 93(17):324, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
324a.pdf.

Anonymous:1905:AFLb

[8151] Anonymous. Anchor and fastener for lightning conductors. Scientific American, 93(17):324, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
324b.pdf.

Anonymous:1905:BPWbh

[8152] Anonymous. Business and personal wants. Scientific American, 93(17):325, October 21, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n17/pdf/scientificamerican10211905-
325a.pdf.

Anonymous:1905:Cbf

[8153] Anonymous. Correspondence. Scientific American, 93(18):339, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n18/pdf/scientificamerican10281905-339.pdf.

Anonymous:1905:RPIbi

[8154] Anonymous. Recently patented inventions. Scientific American, 93(18):345, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
345.pdf.

Anonymous:1905:NQbi

[8155] Anonymous. Notes and queries. Scientific American, 93(18):346, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n18/pdf/scientificamerican10281905-346.pdf.

Anonymous:1905:BNBx

[8156] Anonymous. Books: New books, etc. Scientific American, 93 (18):347–348, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
347.pdf.

Anonymous:1905:IIbi

[8157] Anonymous. Index of inventions. Scientific American, 93(18): 348–351, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
348.pdf.

Anonymous:1905:PNPbe

[8158] Anonymous. Patent notes: Patents. Scientific American, 93 (18):351, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
351.pdf.

Anonymous:1905:LLC

[8159] Anonymous. A lesson from the Lewis and Clark Fair. Scientific American, 93(18):334, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
334.pdf.

Anonymous:1905:LSB

[8160] Anonymous. Long-span bridges of the world. Scientific American, 93(18):334, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
334a.pdf.

Anonymous:1905:LOW

[8161] Anonymous. Life on other worlds. Scientific American, 93(18): 334–335, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
334b.pdf.

Anonymous:1905:FTT

[8162] Anonymous. The first toolmakers and their methods. Scientific American, 93(18):335, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
335.pdf.

Anonymous:1905:LCa

[8163] Anonymous. Literature for convalescents. Scientific American, 93(18):335, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
335a.pdf.

Anonymous:1905:TFE

[8164] Anonymous. Twenty-five electric locomotives for N.Y., N.H., and H.R.R. Co. Scientific American, 93(18):335, October 28, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n18/pdf/scientificamerican10281905-335b.pdf.

Anonymous:1905:SNt

[8165] Anonymous. Science notes. Scientific American, 93(18):335, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n18/pdf/scientificamerican10281905-335c.pdf.

Anonymous:1905:CSm

[8166] Anonymous. The current supplement. Scientific American, 93 (18):335, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
335d.pdf.

Willey:1905:TD

[8167] Day Allen Willey. “The time of day”. Scientific American, 93 (18):336–337, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
336.pdf.

Anonymous:1905:SLR

[8168] Anonymous. The St. Lawrence River Bridge, Quebec. Scientific American, 93(18):337–338, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
337.pdf.

Larkin:1905:MDS

[8169] Edgar L. Larkin. Measuring the distance of a star. Scientific American, 93(18):338–339, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
338.pdf.

Anonymous:1905:SAAb

[8170] Anonymous. The Second Annual Automobile Race for the Vanderbilt Cup. Scientific American, 93(18):339–341, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n18/pdf/scientificamerican10281905-339a.pdf.

Anonymous:1905:ADAb

[8171] Anonymous. Archæological discoveries at antinoë. Scientific American, 93(18):342, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
342.pdf.

Beard:1905:SS

[8172] J. Carter Beard. Striped steeds. Scientific American, 93(18): 342–343, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
342a.pdf.

Anonymous:1905:PDi

[8173] Anonymous. Patent department. Scientific American, 93(18):344, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n18/pdf/scientificamerican10281905-344.pdf.

Anonymous:1905:NSCb

[8174] Anonymous. A new self-cleaning spark plug. Scientific American, 93(18):344, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
344a.pdf.

Anonymous:1905:HCH

[8175] Anonymous. Hanger for cable heads. Scientific American, 93 (18):344, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
344b.pdf.

Anonymous:1905:OIe

[8176] Anonymous. Oddities in inventions. Scientific American, 93 (18):344, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
344c.pdf.

Anonymous:1905:BPWbi

[8177] Anonymous. Business and personal wants. Scientific American, 93(18):345, October 28, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n18/pdf/scientificamerican10281905-
345a.pdf.

Russell:1905:HN

[8178] Henry Norris Russell. The heavens in November. Scientific American, 93(19):355, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
355.pdf.

Anonymous:1905:Cbg

[8179] Anonymous. Correspondence. Scientific American, 93(19):359, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n19/pdf/scientificamerican11041905-359.pdf.

Anonymous:1905:RPIbj

[8180] Anonymous. Recently patented inventions. Scientific American, 93(19):366, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
366.pdf.

Anonymous:1905:IIbj

[8181] Anonymous. Index of inventions. Scientific American, 93(19): 366–371, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
366a.pdf.

Anonymous:1905:PNPbf

[8182] Anonymous. Patent notes: Patents. Scientific American, 93 (19):371, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
371.pdf.

Anonymous:1905:DT

[8183] Anonymous. A discredited theory. Scientific American, 93(19): 354, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
354.pdf.

Anonymous:1905:OYO

[8184] Anonymous. One year’s operation of the subway. Scientific American, 93(19):354, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
354a.pdf.

Anonymous:1905:MHM

[8185] Anonymous. Mr. Hildenbrand and the Manhattan Bridge problem. Scientific American, 93(19):354–355, November 4, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n19/pdf/scientificamerican11041905-354b.pdf.

Anonymous:1905:TBA

[8186] Anonymous. Tunnel boring in Ancient Palestine. Scientific American, 93(19):355, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
355a.pdf.

Anonymous:1905:SNu

[8187] Anonymous. Science notes. Scientific American, 93(19):355, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n19/pdf/scientificamerican11041905-355b.pdf.

Anonymous:1905:LFS

[8188] Anonymous. Ludlow’s first successful flight. Scientific American, 93(19):356, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
356.pdf.

Anonymous:1905:TTSb

[8189] Anonymous. Test of a torpedo shell. Scientific American, 93 (19):357–358, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
357.pdf.

Anonymous:1905:TSLb

[8190] Anonymous. The twin-screw liner “Amerika”. Scientific American, 93(19):360–361, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
360.pdf.

Anonymous:1905:SCA

[8191] Anonymous. Some cities of Ancient Egypt. Scientific American, 93 (19):362–363, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
362.pdf.

Fielde:1905:TLA

[8192] Adele M. Fielde. Tenacity of life in ants. Scientific American, 93 (19):363–364, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
363.pdf.

Anonymous:1905:SSC

[8193] Anonymous. A street-sprinkling curb. Scientific American, 93 (19):364, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
364.pdf.

Anonymous:1905:OGS

[8194] Anonymous. An odorless gas stove. Scientific American, 93 (19):364, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
364a.pdf.

Anonymous:1905:BAP

[8195] Anonymous. Business ana personal wants. Scientific American, 93(19):365, November 4, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n19/pdf/scientificamerican11041905-
365.pdf.

Anonymous:1905:Cbh

[8196] Anonymous. Correspondence. Scientific American, 93(2):27, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-27.pdf.

Anonymous:1905:RPIbk

[8197] Anonymous. Recently patented inventions. Scientific American, 93(2):33–34, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
33.pdf.

Anonymous:1905:NQbj

[8198] Anonymous. Notes and queries. Scientific American, 93(2):34–35, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-34a.pdf.

Anonymous:1905:BNBy

[8199] Anonymous. Books: New books, etc. Scientific American, 93(2):35, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-35.pdf.

Anonymous:1905:IIbk

[8200] Anonymous. Index of inventions. Scientific American, 93(2):35–39, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-35a.pdf.

Anonymous:1905:PNPbg

[8201] Anonymous. Patent notes: Patents. Scientific American, 93(2):39, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-39.pdf.

Anonymous:1905:USC

[8202] Anonymous. The United States and the Command of the Pacific. Scientific American, 93(2):22, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
22.pdf.

Anonymous:1905:CPL

[8203] Anonymous. The canal from Pittsburgh to the Lakes. Scientific American, 93(2):22, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
22a.pdf.

Anonymous:1905:SS

[8204] Anonymous. Safety on the sea. Scientific American, 93(2):22, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-22b.pdf.

Anonymous:1905:NER

[8205] Anonymous. New electric railroads. Scientific American, 93(2):22, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-22c.pdf.

Eddy:1905:ERD

[8206] Horace T. Eddy. Electric railway developments. Scientific American, 93(2):23, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
23.pdf.

Irish:1905:FFM

[8207] W. E. Irish. Feathered flying machines. Scientific American, 93(2):24, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-24.pdf.

Gradenwitz:1905:IGT

[8208] Alfred Gradenwitz. The improved Gruhn telautograph. Scientific American, 93(2):25, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
25.pdf.

Anonymous:1905:SSN

[8209] Anonymous. Sea strength of the Naval Powers. Scientific American, 93(2):26, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
26.pdf.

Anonymous:1905:ACR

[8210] Anonymous. Ancient Chaldean remains found in recent excavations. Scientific American, 93(2):28, July 8, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n2/pdf/scientificamerican07081905-28.pdf.

Anonymous:1905:LFR

[8211] Anonymous. The “leap-frog” railway. Scientific American, 93(2):29–30, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-29.pdf.

Anonymous:1905:TPI

[8212] Anonymous. Two pairs of interlocked deer horns. Scientific American, 93(2):31, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
31.pdf.

Guarini:1905:FCJ

[8213] Emile Guarini. Folding canoes and jointed oars. Scientific American, 93(2):31, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
31a.pdf.

Anonymous:1905:PDj

[8214] Anonymous. Patent department. Scientific American, 93(2):32, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-32.pdf.

Anonymous:1905:GFA

[8215] Anonymous. Glass-framing attachment for window sashes. Scientific American, 93(2):32, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n2/pdf/scientificamerican07081905-
32a.pdf.

Anonymous:1905:NWM

[8216] Anonymous. A novel wave motor. Scientific American, 93(2):32, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-32b.pdf.

Anonymous:1905:CPC

[8217] Anonymous. A convenient pipe coupling. Scientific American, 93(2):32, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-32c.pdf.

Anonymous:1905:BPWbj

[8218] Anonymous. Business and personal wants. Scientific American, 93(2): 34, July 8, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n2/pdf/scientificamerican07081905-34.pdf.

Anonymous:1905:Cbi

[8219] Anonymous. Correspondence. Scientific American, 93(20):379, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n20/pdf/scientificamerican11111905-379.pdf.

Anonymous:1905:RPIbl

[8220] Anonymous. Recently patented inventions. Scientific American, 93 (20):385–386, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
385.pdf.

Anonymous:1905:NQbk

[8221] Anonymous. Notes and queries. Scientific American, 93(20): 386, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
386a.pdf.

Anonymous:1905:BNBz

[8222] Anonymous. Books: New books, etc. Scientific American, 93 (20):386, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
386b.pdf.

Anonymous:1905:IIbl

[8223] Anonymous. Index of inventions. Scientific American, 93(20): 386–391, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
386c.pdf.

Anonymous:1905:PNPbh

[8224] Anonymous. Patent notes: Patents. Scientific American, 93 (20):391, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
391.pdf.

Anonymous:1905:AFE

[8225] Anonymous. An American Floating Exposition. Scientific American, 93(20):374, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
374.pdf.

Anonymous:1905:SPC

[8226] Anonymous. Sanitation of the Panama Canal zone. Scientific American, 93(20):374, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
374a.pdf.

Anonymous:1905:CSR

[8227] Anonymous. Colloids, suspensions, and related phenomena. Scientific American, 93(20):374–375, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
374b.pdf.

Anonymous:1905:SAB

[8228] Anonymous. “The Scientific American boy”. Scientific American, 93(20):375, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
375.pdf.

Lodian:1905:SCU

[8229] L. Lodian. Some curious uses of American woods. Scientific American, 93(20):375, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
375a.pdf.

Anonymous:1905:SNv

[8230] Anonymous. Science notes. Scientific American, 93(20):375, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n20/pdf/scientificamerican11111905-375b.pdf.

Anonymous:1905:SCH

[8231] Anonymous. A six-cylinder 400-horse-power racing gasoline power boat. Scientific American, 93(20):376, November 11, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n20/pdf/scientificamerican11111905-376.pdf.

Winslow:1905:M

[8232] D. R. Winslow. Microphotography. Scientific American, 93(20): 376–378, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
376a.pdf.

Gradenwitz:1905:UOI

[8233] Alfred Gradenwitz. Uniformity of organic and inorganic bodies. Scientific American, 93(20):378, November 11, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n20/pdf/scientificamerican11111905-378.pdf.

Anonymous:1905:AAW

[8234] Anonymous. Automatic apparatus for watering plants. Scientific American, 93(20):379, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
379a.pdf.

Boyer:1905:VFG

[8235] Jacques Boyer. A visit to a French gas works. Scientific American, 93(20):380–382, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
380.pdf.

Dunlap:1905:CEF

[8236] Orrin E. Dunlap. Curious engineering feat at Niagara. Scientific American, 93(20):382, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
382.pdf.

Beard:1905:DA

[8237] J. Carter Beard. Double animals. Scientific American, 93(20): 383, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
383.pdf.

Anonymous:1905:PDk

[8238] Anonymous. Patent department. Scientific American, 93(20): 384, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
384.pdf.

Anonymous:1905:TD

[8239] Anonymous. Transmitting device. Scientific American, 93(20): 384, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
384a.pdf.

Anonymous:1905:BAV

[8240] Anonymous. Balance adjuster for vehicles. Scientific American, 93(20):384, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
384b.pdf.

Anonymous:1905:IWC

[8241] Anonymous. An improved water-current motor. Scientific American, 93(20):384–385, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
384c.pdf.

Anonymous:1905:BPWbk

[8242] Anonymous. Business and personal wants. Scientific American, 93(20):386, November 11, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n20/pdf/scientificamerican11111905-
386.pdf.

Anonymous:1905:Cbj

[8243] Anonymous. Correspondence. Scientific American, 93(21):399, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n21/pdf/scientificamerican11181905-399.pdf.

Anonymous:1905:RPIbm

[8244] Anonymous. Recently patented inventions. Scientific American, 93(21):405, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
405.pdf.

Anonymous:1905:NQbl

[8245] Anonymous. Notes and queries. Scientific American, 93(21): 405, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
405b.pdf.

Anonymous:1905:IIbm

[8246] Anonymous. Index of inventions. Scientific American, 93(21): 406–411, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
406.pdf.

Anonymous:1905:PNPbi

[8247] Anonymous. Patent notes: Patents. Scientific American, 93 (21):411, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
411.pdf.

Anonymous:1905:OBP

[8248] Anonymous. The offer to build the Panama Canal by contract. Scientific American, 93(21):394, November 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n21/pdf/scientificamerican11181905-394.pdf.

Anonymous:1905:SCE

[8249] Anonymous. The Suez Canal explosion. Scientific American, 93 (21):394, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
394a.pdf.

Anonymous:1905:HVG

[8250] Anonymous. High velocities and gun erosion. Scientific American, 93(21):394, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
394b.pdf.

Anonymous:1905:EVW

[8251] Anonymous. The electrical value of wind power. Scientific American, 93(21):394–395, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
394c.pdf.

Anonymous:1905:ANP

[8252] Anonymous. Another North Pole Balloon Expedition. Scientific American, 93(21):395, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
395.pdf.

Anonymous:1905:PCa

[8253] Anonymous. The prevention of “colds”. Scientific American, 93 (21):395, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
395a.pdf.

Anonymous:1905:SNw

[8254] Anonymous. Science notes. Scientific American, 93(21):395, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n21/pdf/scientificamerican11181905-395b.pdf.

Anonymous:1905:VBF

[8255] Anonymous. The visiting British fleet. Scientific American, 93 (21):396, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
396.pdf.

Collins:1905:LIR

[8256] A. Frederick Collins. The Long Island Railroad electrified. Scientific American, 93(21):397–398, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
397.pdf.

Rudin:1905:VVP

[8257] Albert Rudin. A visit to the Volcano Poas. Scientific American, 93(21):400, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
400.pdf.

Anonymous:1905:AET

[8258] Anonymous. An automobile economy test. Scientific American, 93(21):401, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
401.pdf.

Ohalloran:1905:MPS

[8259] Rose O’halloran. The maximum of the present solar cycle. Scientific American, 93(21):402–403, November 18, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n21/pdf/scientificamerican11181905-402.pdf.

Anonymous:1905:CHb

[8260] Anonymous. Captive hippopotami. Scientific American, 93(21): 403, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
403.pdf.

Anonymous:1905:ABP

[8261] Anonymous. An automatic bicycle pump. Scientific American, 93 (21):404, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
404.pdf.

Anonymous:1905:AST

[8262] Anonymous. Automatic switch thrower for street railways. Scientific American, 93(21):404, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
404a.pdf.

Anonymous:1905:BPWbl

[8263] Anonymous. Business and personal wants. Scientific American, 93(21):405, November 18, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n21/pdf/scientificamerican11181905-
405a.pdf.

Anonymous:1905:Cbk

[8264] Anonymous. Correspondence. Scientific American, 93(22):419, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n22/pdf/scientificamerican11251905-419.pdf.

Anonymous:1905:RPIbn

[8265] Anonymous. Recently patented inventions. Scientific American, 93 (22):425–426, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
425.pdf.

Anonymous:1905:NQbm

[8266] Anonymous. Notes and queries. Scientific American, 93(22): 426, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
426a.pdf.

Anonymous:1905:BNBba

[8267] Anonymous. Books: New books, etc. Scientific American, 93 (22):426, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
426b.pdf.

Anonymous:1905:IIbn

[8268] Anonymous. Index of inventions. Scientific American, 93(22): 426–431, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
426c.pdf.

Anonymous:1905:PNPbj

[8269] Anonymous. Patent notes: Patents. Scientific American, 93 (22):431, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
431.pdf.

Anonymous:1905:TT

[8270] Anonymous. Turbine troubles. Scientific American, 93(22):414, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n22/pdf/scientificamerican11251905-414.pdf.

Anonymous:1905:EST

[8271] Anonymous. Eye sight test for railroad men. Scientific American, 93(22):414, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
414a.pdf.

Anonymous:1905:RWT

[8272] Anonymous. Robert Whitehead and the torpedo. Scientific American, 93(22):414, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
414b.pdf.

Anonymous:1905:MPR

[8273] Anonymous. Modern preservation of railroad ties. Scientific American, 93(22):414–415, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
414c.pdf.

Allis:1905:FP

[8274] Dexter W. Allis. Fun with the phonograph. Scientific American, 93(22):415, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
415.pdf.

Anonymous:1905:FHB

[8275] Anonymous. Frederic H. Betts. Scientific American, 93(22): 415, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
415a.pdf.

Anonymous:1905:SNx

[8276] Anonymous. Science notes. Scientific American, 93(22):415, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n22/pdf/scientificamerican11251905-415b.pdf.

Gradenwitz:1905:ACC

[8277] Alfred Gradenwitz. An automatic cartridge-counter for magazine rifles. Scientific American, 93(22):416, November 25, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n22/pdf/scientificamerican11251905-416.pdf.

Mclure:1905:NTO

[8278] W. Frank M’clure. A new type of ore-unloading bucket. Scientific American, 93(22):416, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
416a.pdf.

Anonymous:1905:SBL

[8279] Anonymous. A storage-battery locomotive. Scientific American, 93 (22):417–418, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
417.pdf.

Gradenwitz:1905:NM

[8280] Alfred Gradenwitz. A new megaphone. Scientific American, 93 (22):418, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
418.pdf.

Brown:1905:OWW

[8281] G. L. M. Brown. The Orinoco — a wasted waterway. Scientific American, 93(22):420–422, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
420.pdf.

Anonymous:1905:EEW

[8282] Anonymous. Extraordinary effect of water erosion caused by the Assouan Dam. Scientific American, 93(22):422, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n22/pdf/scientificamerican11251905-422.pdf.

Dunlap:1905:FCC

[8283] Orrin E. Dunlap. The felling of the concrete column at Niagara. Scientific American, 93(22):423, November 25, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n22/pdf/scientificamerican11251905-423.pdf.

Anonymous:1905:PDl

[8284] Anonymous. Patent department. Scientific American, 93(22): 424, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
424.pdf.

Anonymous:1905:IBC

[8285] Anonymous. An improved bottle closure. Scientific American, 93 (22):424, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
424a.pdf.

Anonymous:1905:QMD

[8286] Anonymous. Quartz mill for dry-grinding of ore. Scientific American, 93(22):424, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
424b.pdf.

Anonymous:1905:MMP

[8287] Anonymous. Machine for molding plastic materials. Scientific American, 93(22):424, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
424c.pdf.

Anonymous:1905:BPWbm

[8288] Anonymous. Business and personal wants. Scientific American, 93(22):426, November 25, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n22/pdf/scientificamerican11251905-
426.pdf.

Anonymous:1905:Cbl

[8289] Anonymous. Correspondence. Scientific American, 93(23):439, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n23/pdf/scientificamerican12021905-439.pdf.

Anonymous:1905:RPIbo

[8290] Anonymous. Recently patented inventions. Scientific American, 93(23):444, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
444.pdf.

Anonymous:1905:IIbo

[8291] Anonymous. Index of inventions. Scientific American, 93(23): 444–449, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
444b.pdf.

Anonymous:1905:PNPbk

[8292] Anonymous. Patent notes: Patents. Scientific American, 93 (23):451, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
451.pdf.

Anonymous:1905:FL

[8293] Anonymous. “Festina Lente”. Scientific American, 93(23):434, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n23/pdf/scientificamerican12021905-434.pdf.

Anonymous:1905:APL

[8294] Anonymous. Admiral Prince Louis of Battenberg on speed in warships. Scientific American, 93(23):434, December 2, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n23/pdf/scientificamerican12021905-434a.pdf.

Anonymous:1905:EMC

[8295] Anonymous. Electric mining in California. Scientific American, 93 (23):434–435, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
434b.pdf.

Larkin:1905:HMD

[8296] Edgar L. Larkin. How much does the universe weigh? Scientific American, 93(23):435, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
435.pdf.

Anonymous:1905:SNy

[8297] Anonymous. Science notes. Scientific American, 93(23):435, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n23/pdf/scientificamerican12021905-435a.pdf.

Mels:1905:PC

[8298] Edgar Mels. The photobiography of a chimney. Scientific American, 93(23):436, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
436.pdf.

Wade:1905:USN

[8299] Herbert T. Wade. United States National Bureau of Standards. Scientific American, 93(23):437, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
437.pdf.

Anonymous:1905:SFB

[8300] Anonymous. Stern frames and brackets for the new Cunarders. Scientific American, 93(23):438, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
438.pdf.

Anonymous:1905:NDA

[8301] Anonymous. A nailless device for attaching horseshoes. Scientific American, 93(23):438, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
438a.pdf.

Smith:1905:LSD

[8302] C. Smith. Labor-saving devices utilized in the manufacture of tea. Scientific American, 93(23):439–440, December 2, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n23/pdf/scientificamerican12021905-439a.pdf.

Fitz-Gerald:1905:EPT

[8303] W. G. Fitz-Gerald. The earth pyramids of the Tyrol. Scientific American, 93(23):440–441, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
440.pdf.

Beard:1905:CA

[8304] J. Carter Beard. Composite animals. Scientific American, 93 (23):442, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
442.pdf.

Brassler:1905:PMP

[8305] Charles A. Brassler. The Pergamon Marbles in the Pergamon Museum of Berlin. Scientific American, 93(23):442–444, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n23/pdf/scientificamerican12021905-442a.pdf.

Anonymous:1905:BPWbn

[8306] Anonymous. Business and personal wants. Scientific American, 93(23):444, December 2, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n23/pdf/scientificamerican12021905-
444a.pdf.

Anonymous:1905:HD

[8307] Anonymous. The heavens in December. Scientific American, 93 (24):454–455, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
454c.pdf.

Anonymous:1905:RPIbp

[8308] Anonymous. Recently patented inventions. Scientific American, 93(24):465, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
465.pdf.

Anonymous:1905:NQbn

[8309] Anonymous. Notes and queries. Scientific American, 93(24):465, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n24/pdf/scientificamerican12091905-465b.pdf.

Anonymous:1905:IIbp

[8310] Anonymous. Index of inventions. Scientific American, 93(24): 465–471, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
465c.pdf.

Anonymous:1905:PNPbl

[8311] Anonymous. Patent notes: Patents. Scientific American, 93 (24):471, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
471.pdf.

Anonymous:1905:GRS

[8312] Anonymous. Guard rails in the subway. Scientific American, 93 (24):454, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
454.pdf.

Anonymous:1905:ADR

[8313] Anonymous. An all-day race between battleships. Scientific American, 93(24):454, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
454a.pdf.

Anonymous:1905:GOR

[8314] Anonymous. The growth of our railroad system. Scientific American, 93(24):454, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
454b.pdf.

Anonymous:1905:MOB

[8315] Anonymous. The mysteries of the ocean bed. Scientific American, 93(24):455, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
455.pdf.

Anonymous:1905:SNz

[8316] Anonymous. Science notes. Scientific American, 93(24):455, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n24/pdf/scientificamerican12091905-455a.pdf.

Glyndon:1905:BPS

[8317] Albert Glyndon. The bacterial purification of sewage. Scientific American, 93(24):456, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
456.pdf.

Anonymous:1905:NGM

[8318] Anonymous. The new gasoline motor car. Scientific American, 93 (24):457–458, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
457.pdf.

Anonymous:1905:FID

[8319] Anonymous. The flax industry of to-day. Scientific American, 93 (24):458–459, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
458.pdf.

Anonymous:1905:PNJ

[8320] Anonymous. Progress of the New Jersey Tunnels and Subways. Scientific American, 93(24):459–461, December 9, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n24/pdf/scientificamerican12091905-459.pdf.

Anonymous:1905:LPS

[8321] Anonymous. The lifting power of a screw propeller for aeronautical work. Scientific American, 93(24):461, December 9, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n24/pdf/scientificamerican12091905-461.pdf.

Anonymous:1905:NMDb

[8322] Anonymous. A novel method of dog-shearing. Scientific American, 93(24):462, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
462.pdf.

Willey:1905:SFN

[8323] Day Allen Willey. The salmon fisheries of the Northwest. Scientific American, 93(24):462–464, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
462a.pdf.

Anonymous:1905:DGB

[8324] Anonymous. Depth gage for brace bits. Scientific American, 93 (24):464, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
464.pdf.

Anonymous:1905:IE

[8325] Anonymous. An improved excavator. Scientific American, 93 (24):464, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
464a.pdf.

Anonymous:1905:BPWbo

[8326] Anonymous. Business and personal wants. Scientific American, 93(24):465, December 9, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n24/pdf/scientificamerican12091905-
465a.pdf.

Anonymous:1905:Cbm

[8327] Anonymous. Correspondence. Scientific American, 93(25):479, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n25/pdf/scientificamerican12161905-479.pdf.

Anonymous:1905:RPIbq

[8328] Anonymous. Recently patented inventions. Scientific American, 93 (25):490–491, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
490.pdf.

Anonymous:1905:NQbo

[8329] Anonymous. Notes and queries. Scientific American, 93(25): 491, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
491a.pdf.

Anonymous:1905:BNBbb

[8330] Anonymous. Books: New books, etc. Scientific American, 93(25): 492–494, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
492.pdf.

Anonymous:1905:IIbq

[8331] Anonymous. Index of inventions. Scientific American, 93(25): 494–499, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
494.pdf.

Anonymous:1905:RYb

[8332] Anonymous. Retrospect of the year 1905. Scientific American, 93 (25):474–475, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
474.pdf.

Anonymous:1905:RID

[8333] Anonymous. The Roosevelt Irrigation Dam in Arizona. Scientific American, 93(25):476–477, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
476.pdf.

Inkersley:1905:RS

[8334] Arthur Inkersley. Rare shellfish. Scientific American, 93(25): 478, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
478.pdf.

Anonymous:1905:TATb

[8335] Anonymous. A typical American touring car. Scientific American, 93(25):480, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
480.pdf.

Wilson:1905:NAF

[8336] E. Wilson. Natural and artificial flight. Scientific American, 93 (25):481, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
481.pdf.

Beasley:1905:KST

[8337] Walter L. Beasley. The Koryaks of Siberia and their Great Whale Festival. Scientific American, 93(25):482–483, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n25/pdf/scientificamerican12161905-482.pdf.

Willey:1905:OFW

[8338] Day Allen Willey. The oil fields of the West. Scientific American, 93 (25):484–485, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
484.pdf.

Anonymous:1905:JNA

[8339] Anonymous. The Japanese Navy after the War. Scientific American, 93(25):485–488, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
485.pdf.

Anonymous:1905:NPT

[8340] Anonymous. Novelties at the Parisian Toy Exposition. Scientific American, 93(25):489, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
489.pdf.

Anonymous:1905:BPWbp

[8341] Anonymous. Business and personal wants. Scientific American, 93(25):491, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
491.pdf.

Anonymous:1905:WPP

[8342] Anonymous. The wealth patents — patents of nations. Scientific American, 93(25):499, December 16, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n25/pdf/scientificamerican12161905-
499.pdf.

Anonymous:1905:Cbn

[8343] Anonymous. Correspondence. Scientific American, 93(26):507, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n26/pdf/scientificamerican12231905-507.pdf.

Anonymous:1905:RPIbr

[8344] Anonymous. Recently patented inventions. Scientific American, 93 (26):513–514, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
513.pdf.

Anonymous:1905:NQbp

[8345] Anonymous. Notes and queries. Scientific American, 93(26): 514, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
514a.pdf.

Anonymous:1905:BNBbc

[8346] Anonymous. Books: New books, etc. Scientific American, 93 (26):514, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
514b.pdf.

Anonymous:1905:IIbr

[8347] Anonymous. Index of inventions. Scientific American, 93(26): 514–519, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
514c.pdf.

Anonymous:1905:PNPbm

[8348] Anonymous. Patent notes: Patents. Scientific American, 93 (26):519, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
519.pdf.

Anonymous:1905:LSM

[8349] Anonymous. Latest success of the marine turbine. Scientific American, 93(26):502, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
502.pdf.

Anonymous:1905:QNP

[8350] Anonymous. The quest of the Northwest Passage. Scientific American, 93(26):502–503, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
502a.pdf.

Anonymous:1905:ESG

[8351] Anonymous. The electrical show at the garden. Scientific American, 93(26):503, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
503.pdf.

Anonymous:1905:DEA

[8352] Anonymous. The death of Edward Atkinson. Scientific American, 93(26):503, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
503a.pdf.

Anonymous:1905:PO

[8353] Anonymous. The Patent Office. Scientific American, 93(26): 503, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
503b.pdf.

Anonymous:1905:SNba

[8354] Anonymous. Science notes. Scientific American, 93(26):503, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n26/pdf/scientificamerican12231905-503c.pdf.

Anonymous:1905:OTB

[8355] Anonymous. Official trials of the battleship “Rhode Island”. Scientific American, 93(26):504, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
504.pdf.

Chapman:1905:GPH

[8356] William B. Chapman. The gas producer for heating processes. Scientific American, 93(26):504, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
504a.pdf.

Stewart:1905:NPS

[8357] J. A. Stewart. The new Philadelphia subway. Scientific American, 93(26):505–506, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
505.pdf.

Anonymous:1905:TC

[8358] Anonymous. The turbines of the “Carmania”. Scientific American, 93(26):508, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
508.pdf.

Beard:1905:SMA

[8359] J. Carter Beard. Single and multiple animals. Scientific American, 93(26):510, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
510.pdf.

Anonymous:1905:TLM

[8360] Anonymous. Is there life on the Moon? Scientific American, 93(26):511, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
511.pdf.

Anonymous:1905:PWH

[8361] Anonymous. A portable water heater. Scientific American, 93 (26):512, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
512.pdf.

Anonymous:1905:NSD

[8362] Anonymous. A novel sketching device. Scientific American, 93 (26):512, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
512a.pdf.

Anonymous:1905:SMB

[8363] Anonymous. Simple mail-box delivery device. Scientific American, 93(26):512, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
512b.pdf.

Anonymous:1905:PMO

[8364] Anonymous. Pulling machine for oil wells. Scientific American, 93(26):512, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
512c.pdf.

Anonymous:1905:BPWbq

[8365] Anonymous. Business and personal wants. Scientific American, 93(26):514, December 23, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n26/pdf/scientificamerican12231905-
514.pdf.

Anonymous:1905:Cbo

[8366] Anonymous. Correspondence. Scientific American, 93(3):46–47, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-46a.pdf.

Anonymous:1905:RPIbs

[8367] Anonymous. Recently patented inventions. Scientific American, 93(3):53, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
53.pdf.

Anonymous:1905:NQbq

[8368] Anonymous. Notes and queries. Scientific American, 93(3):54–55, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-54.pdf.

Anonymous:1905:BNBbd

[8369] Anonymous. Books: New books, etc. Scientific American, 93(3):55, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-55.pdf.

Anonymous:1905:IIbs

[8370] Anonymous. Index of inventions. Scientific American, 93(3):56–59, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-56.pdf.

Anonymous:1905:PNPbn

[8371] Anonymous. Patent notes: Patents. Scientific American, 93(3):59, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-59.pdf.

Anonymous:1905:PSc

[8372] Anonymous. A poor start. Scientific American, 93(3):42, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-42.pdf.

Anonymous:1905:PNPbo

[8373] Anonymous. Peary and the North Pole. Scientific American, 93(3):42, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-42a.pdf.

Anonymous:1905:TAP

[8374] Anonymous. Time for another Principia. Scientific American, 93(3):42–43, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
42b.pdf.

Anonymous:1905:BOA

[8375] Anonymous. Burke’s own account of the spontaneous action of radium on gelatin media. Scientific American, 93(3):43, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-43.pdf.

Anonymous:1905:CSn

[8376] Anonymous. The current supplement. Scientific American, 93(3):43, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-43a.pdf.

Anonymous:1905:OEN

[8377] Anonymous. Opening of extensions of the New York City subway. Scientific American, 93(3):43, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
43b.pdf.

Anonymous:1905:OMS

[8378] Anonymous. Official Meteorological Summary, New York, N.Y., June, 1905. Scientific American, 93(3):43, July 15, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n3/pdf/scientificamerican07151905-43c.pdf.

Beasley:1905:NWS

[8379] Walter L. Beasley. The Navy Wireless School. Scientific American, 93(3):44, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
44.pdf.

Anonymous:1905:TCC

[8380] Anonymous. Tallest concrete chimney ever constructed. Scientific American, 93(3):44–45, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
44a.pdf.

Ludlow:1905:NA

[8381] Israel Ludlow. A new aeroplane. Scientific American, 93(3):45–46, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-45.pdf.

Anonymous:1905:TAS

[8382] Anonymous. The Temple of Abu Simbel. Scientific American, 93(3):46, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-46.pdf.

Anonymous:1905:SIAb

[8383] Anonymous. The Sixth International Automobile Race for the Bennett Trophy. Scientific American, 93(3):47, July 15, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n3/pdf/scientificamerican07151905-47.pdf.

Anonymous:1905:SNAb

[8384] Anonymous. Some new automobile track records. Scientific American, 93(3):47, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
47a.pdf.

Anonymous:1905:PASb

[8385] Anonymous. Peary’s Arctic ship, the “Roosevelt”. Scientific American, 93(3):47–48, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
47b.pdf.

Durant:1905:EOG

[8386] Edward Durant. An electrically-operated gyroscope. Scientific American, 93(3):50, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
50.pdf.

Anonymous:1905:WI

[8387] Anonymous. The “whirlpool” illusion. Scientific American, 93(3):50, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-50a.pdf.

Gradenwitz:1905:IMH

[8388] Alfred Gradenwitz. Improved methods in high-speed chronophotography. Scientific American, 93(3):50–51, July 15, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n3/pdf/scientificamerican07151905-50b.pdf.

Anonymous:1905:PDm

[8389] Anonymous. Patent department. Scientific American, 93(3):52, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-52.pdf.

Anonymous:1905:GSD

[8390] Anonymous. A gas-saving device. Scientific American, 93(3):52, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-52a.pdf.

Anonymous:1905:OIf

[8391] Anonymous. Oddities in inventions. Scientific American, 93(3):52, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n3/pdf/scientificamerican07151905-52b.pdf.

Anonymous:1905:BPWbr

[8392] Anonymous. Business and personal wants. Scientific American, 93(3):53, July 15, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n3/pdf/scientificamerican07151905-
53a.pdf.

Anonymous:1905:Cbp

[8393] Anonymous. Correspondence. Scientific American, 93(4):67, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n4/pdf/scientificamerican07221905-67.pdf.

Anonymous:1905:RPIbt

[8394] Anonymous. Recently patented inventions. Scientific American, 93(4):72, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
72a.pdf.

Anonymous:1905:IIbt

[8395] Anonymous. Index of inventions. Scientific American, 93(4):72–75, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n4/pdf/scientificamerican07221905-72b.pdf.

Anonymous:1905:PNPbp

[8396] Anonymous. Patent notes: Patents. Scientific American, 93(4):75, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n4/pdf/scientificamerican07221905-75.pdf.

Anonymous:1905:FTS

[8397] Anonymous. Facing and trailing switches. Scientific American, 93(4):62, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n4/pdf/scientificamerican07221905-62.pdf.

Anonymous:1905:BB

[8398] Anonymous. Below the belt. Scientific American, 93(4):62, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n4/pdf/scientificamerican07221905-62a.pdf.

Anonymous:1905:SCW

[8399] Anonymous. The second city in the world. Scientific American, 93(4):62, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
62b.pdf.

Anonymous:1905:FLDb

[8400] Anonymous. Fastest long-distance railroad run. Scientific American, 93(4):62, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
62c.pdf.

Anonymous:1905:TTC

[8401] Anonymous. Testing the touring cars. Scientific American, 93(4):62, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n4/pdf/scientificamerican07221905-62d.pdf.

DelMar:1905:RTM

[8402] Alex Del Mar. Roman textile manufactures. Scientific American, 93(4):63, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
63.pdf.

Anonymous:1905:SNbb

[8403] Anonymous. Science notes. Scientific American, 93(4):63, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n4/pdf/scientificamerican07221905-63a.pdf.

Anonymous:1905:SWM

[8404] Anonymous. The smallest working model of a triple-expansion engine. Scientific American, 93(4):64, July 22, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n4/pdf/scientificamerican07221905-64.pdf.

Guarini:1905:LRC

[8405] Emile Guarini. A lighthouse of reinforced concrete. Scientific American, 93(4):64, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
64a.pdf.

Anonymous:1905:NGSb

[8406] Anonymous. A novel gyrating steam engine. Scientific American, 93(4):65, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
65.pdf.

Wade:1905:NBY

[8407] Herbert T. Wade. The new British Yard Standard. Scientific American, 93(4):65–66, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
65a.pdf.

Anonymous:1905:MCL

[8408] Anonymous. The mammoth Cunard Liners. Scientific American, 93(4):66, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
66.pdf.

Fuller:1905:AFG

[8409] Myron I. Fuller. Artificial fountain and geyser springs. Scientific American, 93(4):67, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
67a.pdf.

Anonymous:1905:CVF

[8410] Anonymous. Completion of Victoria Falls Bridge. Scientific American, 93(4):68–69, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
68.pdf.

Zahm:1905:AFO

[8411] A. E. Zahm. Atmospheric friction as an obstacle in locomotion. Scientific American, 93(4):70–71, July 22, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n4/pdf/scientificamerican07221905-70.pdf.

Claudy:1905:MFM

[8412] C. H. Claudy. The Moro fire maker. Scientific American, 93(4):71, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n4/pdf/scientificamerican07221905-71.pdf.

Shepstone:1905:MMS

[8413] Harold J. Shepstone. Mounting a monster sea elephant. Scientific American, 93(4):71, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
71a.pdf.

Anonymous:1905:BPWbs

[8414] Anonymous. Business and personal wants. Scientific American, 93(4):72, July 22, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n4/pdf/scientificamerican07221905-
72.pdf.

Russell:1905:HAb

[8415] Henry Norris Russell. The heavens in August. Scientific American, 93(5):79, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
79.pdf.

Anonymous:1905:Cbq

[8416] Anonymous. Correspondence. Scientific American, 93(5):83, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-83.pdf.

Anonymous:1905:RPIbu

[8417] Anonymous. Recently patented inventions. Scientific American, 93(5):89–90, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
89.pdf.

Anonymous:1905:NQbr

[8418] Anonymous. Notes and queries. Scientific American, 93(5):90–91, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-90a.pdf.

Anonymous:1905:BNBbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-91.pdf.

Anonymous:1905:IIbu

[8420] Anonymous. Index of inventions. Scientific American, 93(5):92–95, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-92.pdf.

Anonymous:1905:PNPbq

[8421] Anonymous. Patent notes: Patents. Scientific American, 93(5):95, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-95.pdf.

Anonymous:1905:SSP

[8422] Anonymous. Sanitation and subsistence at Panama. Scientific American, 93(5):78, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
78.pdf.

Anonymous:1905:FLT

[8423] Anonymous. The first and last of its type. Scientific American, 93(5):78, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-78a.pdf.

Anonymous:1905:TSC

[8424] Anonymous. The turbine steamers on the Canadian route. Scientific American, 93(5):78, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
78b.pdf.

Anonymous:1905:FGP

[8425] Anonymous. The future of the gas-producer engine. Scientific American, 93(5):78, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
78c.pdf.

Anonymous:1905:NKF

[8426] Anonymous. A new kind of fireproof theater. Scientific American, 93(5):79, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
79a.pdf.

Anonymous:1905:SNbc

[8427] Anonymous. Science notes. Scientific American, 93(5):79, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-79b.pdf.

Jones:1905:TES

[8428] George J. Jones. Telephones on every street corner. Scientific American, 93(5):80, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
80.pdf.

Anonymous:1905:POJ

[8429] Anonymous. The Primitive Observatory of Jeypore. Scientific American, 93(5):80, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
80a.pdf.

Anonymous:1905:EH

[8430] Anonymous. The evolution of the horse. Scientific American, 93(5):81–82, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
81.pdf.

Gernsback:1905:NI

[8431] Huck Gernsback. A new interrupter. Scientific American, 93(5):82, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-82.pdf.

Anonymous:1905:FTDb

[8432] Anonymous. Five thousand degrees of heat. Scientific American, 93(5):83–84, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
83a.pdf.

Fernald:1905:MFP

[8433] James G. Fernald. A modern filtration plant. Scientific American, 93(5):85–86, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
85.pdf.

Coleman:1905:ULS

[8434] P. C. Coleman. A unique locomotive for South Africa. Scientific American, 93(5):86–87, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
86.pdf.

Anonymous:1905:RHC

[8435] Anonymous. Results of the hill-climbing contest at Mount Washington. Scientific American, 93(5):87, July 29, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n5/pdf/scientificamerican07291905-87.pdf.

Anonymous:1905:NGT

[8436] Anonymous. A new game of table billiards. Scientific American, 93(5):88, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
88.pdf.

Anonymous:1905:NMH

[8437] Anonymous. A new method of hanging window sashes. Scientific American, 93(5):88, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
88a.pdf.

Anonymous:1905:IPW

[8438] Anonymous. An improved pipe wrench. Scientific American, 93(5):88, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n5/pdf/scientificamerican07291905-88b.pdf.

Anonymous:1905:BPWbt

[8439] Anonymous. Business and personal wants. Scientific American, 93(5):90, July 29, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n5/pdf/scientificamerican07291905-
90.pdf.

Anonymous:1905:Cbr

[8440] Anonymous. Correspondence. Scientific American, 93(6):103, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-103.pdf.

Anonymous:1905:RPIbv

[8441] Anonymous. Recently patented inventions. Scientific American, 93(6):109–110, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
109.pdf.

Anonymous:1905:NQbs

[8442] Anonymous. Notes and queries. Scientific American, 93(6):110–111, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-110a.pdf.

Anonymous:1905:BNBbf

[8443] Anonymous. Books: New books, etc. Scientific American, 93(6):111, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-111.pdf.

Anonymous:1905:IIbv

[8444] Anonymous. Index of inventions. Scientific American, 93(6):111–115, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-111a.pdf.

Anonymous:1905:PNPbr

[8445] Anonymous. Patent notes: Patents. Scientific American, 93(6):115, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-115.pdf.

Anonymous:1905:LBD

[8446] Anonymous. The lesson of the “Bennington” disaster. Scientific American, 93(6):98, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
98.pdf.

Anonymous:1905:MSL

[8447] Anonymous. Mechanical stokers on locomotives. Scientific American, 93(6):98, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
98a.pdf.

Anonymous:1905:SCR

[8448] Anonymous. Seawanhaka Cup returns to America. Scientific American, 93(6):98, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
98b.pdf.

Anonymous:1905:WCW

[8449] Anonymous. Wear of car wheels on curves. Scientific American, 93(6):98, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
98c.pdf.

Anonymous:1905:PAF

[8450] Anonymous. Protection against fire damp. Scientific American, 93(6):98–99, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
98d.pdf.

Anonymous:1905:YFP

[8451] Anonymous. Yellow fever prevented by mosquito extermination. Scientific American, 93(6):99, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
99.pdf.

Banks:1905:HAB

[8452] Edgar James Banks. How the Ancient Babylonians drained their city. Scientific American, 93(6):99, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
99a.pdf.

Anonymous:1905:SNbd

[8453] Anonymous. Science notes. Scientific American, 93(6):99, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-99b.pdf.

Clure:1905:IOC

[8454] W. Frank Mc Clure. The improvement of the Ohio Canals. Scientific American, 93(6):100, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
100.pdf.

Dunlap:1905:IME

[8455] Orrin E. Dunlap. Ice-making by electricity. Scientific American, 93(6):100–101, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
100a.pdf.

Powell:1905:HCB

[8456] L. B. Powell. Hollow concrete building blocks: Their manufacture and use. Scientific American, 93(6):101–102, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-101.pdf.

Anonymous:1905:ETT

[8457] Anonymous. Electric tempering of tools. Scientific American, 93(6):103, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-103a.pdf.

Anonymous:1905:LHC

[8458] Anonymous. The life history of a coral. Scientific American, 93(6):104, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n6/pdf/scientificamerican08051905-104.pdf.

Anonymous:1905:NJB

[8459] Anonymous. The new Japanese battleships. Scientific American, 93(6):105–106, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
105.pdf.

Brown:1905:FUC

[8460] Enos Brown. The future University of California. Scientific American, 93(6):106, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
106.pdf.

Anonymous:1905:CMPb

[8461] Anonymous. Canals of Mars photographed. Scientific American, 93(6):107, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
107.pdf.

Anonymous:1905:PCb

[8462] Anonymous. The Pygmies of the Congo. Scientific American, 93 (6):107–108, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
107a.pdf.

Anonymous:1905:AFLc

[8463] Anonymous. America’s first long-distance motor-boat race. Scientific American, 93(6):108–109, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
108.pdf.

Anonymous:1905:BPWbu

[8464] Anonymous. Business and personal wants. Scientific American, 93(6):110, August 5, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n6/pdf/scientificamerican08051905-
110.pdf.

Anonymous:1905:Cbs

[8465] Anonymous. Correspondence. Scientific American, 93(7):123, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n7/pdf/scientificamerican08121905-123a.pdf.

Anonymous:1905:IIbw

[8466] Anonymous. Index of inventions. Scientific American, 93(7): 128–131, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
128a.pdf.

Anonymous:1905:PNPbs

[8467] Anonymous. Patent notes: Patents. Scientific American, 93(7):130, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n7/pdf/scientificamerican08121905-130.pdf.

Anonymous:1905:EPDb

[8468] Anonymous. The electric power developments at Niagara Falls. Scientific American, 93(7):118, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
118.pdf.

Anonymous:1905:CAS

[8469] Anonymous. The cause of accidents to submarine boats. Scientific American, 93(7):118, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
118a.pdf.

Anonymous:1905:EP

[8470] Anonymous. An excellent precedent. Scientific American, 93 (7):118, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
118b.pdf.

Anonymous:1905:BCMb

[8471] Anonymous. Blood corpuscles on Mont Blanc. Scientific American, 93(7):118–119, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
118c.pdf.

Anonymous:1905:PNPbt

[8472] Anonymous. A prize for a non-poisonous diamond cutter’s composition. Scientific American, 93(7):119, August 12, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n7/pdf/scientificamerican08121905-119.pdf.

Davidson:1905:PSC

[8473] J. W. Davidson. Photographing the solar corona. Scientific American, 93(7):119, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
119a.pdf.

Anonymous:1905:SNbe

[8474] Anonymous. Science notes. Scientific American, 93(7):119, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n7/pdf/scientificamerican08121905-119b.pdf.

Brassler:1905:AR

[8475] Charles A. Brassler. The astrolabe of regiomontanus. Scientific American, 93(7):120, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
120.pdf.

Anonymous:1905:SBA

[8476] Anonymous. A safety buffer for automobiles. Scientific American, 93(7):120–121, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
120a.pdf.

Gradenwitz:1905:EPH

[8477] Alfred Gradenwitz. The electrolytic production of hydrogen and oxygen for welding purposes. Scientific American, 93(7):121, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n7/pdf/scientificamerican08121905-121.pdf.

Anonymous:1905:NJT

[8478] Anonymous. The New Jersey Tunnels and Subways. Scientific American, 93(7):122–123, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
122.pdf.

Anonymous:1905:NOC

[8479] Anonymous. A novel oil can. Scientific American, 93(7):123, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n7/pdf/scientificamerican08121905-123.pdf.

Anonymous:1905:BD

[8480] Anonymous. The “Bennington” disaster. Scientific American, 93 (7):124–126, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
124.pdf.

Anonymous:1905:EPDc

[8481] Anonymous. Electric power developments at Niagara Falls — I. Scientific American, 93(7):126, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
126.pdf.

Anonymous:1905:AGT

[8482] Anonymous. Award of the Glidden Trophy for touring cars. Scientific American, 93(7):127, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
127.pdf.

Anonymous:1905:BPWbv

[8483] Anonymous. Business and personal wants. Scientific American, 93(7):128, August 12, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n7/pdf/scientificamerican08121905-
128.pdf.

Anonymous:1905:Cbt

[8484] Anonymous. Correspondence. Scientific American, 93(8):139, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n8/pdf/scientificamerican08191905-139.pdf.

Anonymous:1905:RPIbw

[8485] Anonymous. Recently patented inventions. Scientific American, 93(8):146, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
146.pdf.

Anonymous:1905:NQbt

[8486] Anonymous. Notes and queries. Scientific American, 93(8):146– 147, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
146b.pdf.

Anonymous:1905:IIbx

[8487] Anonymous. Index of inventions. Scientific American, 93(8): 147–151, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
147.pdf.

Anonymous:1905:PNPbu

[8488] Anonymous. Patent notes: Patents. Scientific American, 93(8):151, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n8/pdf/scientificamerican08191905-151.pdf.

Anonymous:1905:FIS

[8489] Anonymous. Failure of the Isham shell. Scientific American, 93(8):134, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
134.pdf.

Anonymous:1905:RFZ

[8490] Anonymous. Rescue of the Fiala–Ziegler Expedition. Scientific American, 93(8):134, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
134a.pdf.

Anonymous:1905:BMB

[8491] Anonymous. Bids for the Manhattan Bridge. Scientific American, 93(8):134, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
134b.pdf.

Anonymous:1905:ELN

[8492] Anonymous. Electric locomotives for the New York Central. Scientific American, 93(8):134, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
134c.pdf.

Anonymous:1905:ROA

[8493] Anonymous. Railroad and other accidents in the United States. Scientific American, 93(8):134, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
134d.pdf.

Anonymous:1905:BHS

[8494] Anonymous. Bombs for hail in Switzerland. Scientific American, 93(8):134–135, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
134e.pdf.

DelMar:1905:IAR

[8495] Alex. Del Mar. Inventions of ancient rome: Some forerunners of modern ingenuity. Scientific American, 93(8):135, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n8/pdf/scientificamerican08191905-135.pdf.

Anonymous:1905:SNbf

[8496] Anonymous. Science notes. Scientific American, 93(8):135, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n8/pdf/scientificamerican08191905-135a.pdf.

Brown:1905:HMC

[8497] Enos Brown. Hydraulic mining with centrifugal pumps. Scientific American, 93(8):136, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
136.pdf.

Banks:1905:SSK

[8498] Edgar James Banks. The statue of the Sumerian King David. Scientific American, 93(8):137, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
137.pdf.

Gradenwitz:1905:HPR

[8499] Alfred Gradenwitz. A hand-pressure riveting machine. Scientific American, 93(8):137, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
137a.pdf.

Payne:1905:ATT

[8500] Henry Mace Payne. Aerial tramways in the Tug River coal fields. Scientific American, 93(8):138, August 19, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n8/pdf/scientificamerican08191905-138.pdf.

Anonymous:1905:MMEb

[8501] Anonymous. Measuring the mechanical equivalent of heat. Scientific American, 93(8):140, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
140.pdf.

Anonymous:1905:RAM

[8502] Anonymous. The Rexer automatic machine-gun. Scientific American, 93(8):140–141, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
140a.pdf.

Willey:1905:BRC

[8503] Day Allen Willey. The building of a railway car. Scientific American, 93(8):141–142, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
141.pdf.

Anonymous:1905:AME

[8504] Anonymous. An apparatus for measuring the Earth’s speed of rotation. Scientific American, 93(8):142–143, August 19, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n8/pdf/scientificamerican08191905-142.pdf.

Jones:1905:UCR

[8505] Harry Dillon Jones. A unique collection of rare birds. Scientific American, 93(8):143–144, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
143.pdf.

Anonymous:1905:PDn

[8506] Anonymous. Patent department. Scientific American, 93(8):145, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n8/pdf/scientificamerican08191905-145.pdf.

Anonymous:1905:IWT

[8507] Anonymous. An improved waiter’s tray. Scientific American, 93(8):145, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
145a.pdf.

Anonymous:1905:DA

[8508] Anonymous. Drying apparatus. Scientific American, 93(8):145, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n8/pdf/scientificamerican08191905-145b.pdf.

Anonymous:1905:NRVb

[8509] Anonymous. New rotary valve for steam engines. Scientific American, 93(8):145, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
145c.pdf.

Anonymous:1905:BPWbw

[8510] Anonymous. Business and personal wants. Scientific American, 93(8):146, August 19, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n8/pdf/scientificamerican08191905-
146a.pdf.

Anonymous:1905:Cbu

[8511] Anonymous. Correspondence. Scientific American, 93(9):159, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n9/pdf/scientificamerican08261905-159.pdf.

Anonymous:1905:RPIbx

[8512] Anonymous. Recently patented inventions. Scientific American, 93 (9):165–166, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
165.pdf.

Anonymous:1905:NQbu

[8513] Anonymous. Notes and queries. Scientific American, 93(9):167, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n9/pdf/scientificamerican08261905-167.pdf.

Anonymous:1905:BNBbg

[8514] Anonymous. Books: New books, etc. Scientific American, 93 (9):167, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
167a.pdf.

Anonymous:1905:IIby

[8515] Anonymous. Index of inventions. Scientific American, 93(9): 167–171, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
167b.pdf.

Anonymous:1905:PNPbv

[8516] Anonymous. Patent notes: Patents. Scientific American, 93(9):171, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n9/pdf/scientificamerican08261905-171.pdf.

Anonymous:1905:DBA

[8517] Anonymous. Delay in building the Ambrose Channel. Scientific American, 93(9):154, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
154.pdf.

Anonymous:1905:OPN

[8518] Anonymous. An old problem in a new form. Scientific American, 93(9):154, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
154a.pdf.

Anonymous:1905:RAH

[8519] Anonymous. Russian Army hydrogen balloons. Scientific American, 93(9):154, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
154b.pdf.

Anonymous:1905:GNC

[8520] Anonymous. The Government’s new coal-testing plant. Scientific American, 93(9):154–155, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
154c.pdf.

Adams:1905:PSA

[8521] Alton D. Adams. Power sites about Niagara Falls. Scientific American, 93(9):155, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
155.pdf.

Anonymous:1905:SNbg

[8522] Anonymous. Science notes. Scientific American, 93(9):155, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n9/pdf/scientificamerican08261905-155a.pdf.

Anonymous:1905:PG

[8523] Anonymous. Procession of the giants. Scientific American, 93 (9):156, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
156.pdf.

Tracy:1905:DSP

[8524] J. Irving Tracy. Determining the speed of a photographic shutter. Scientific American, 93(9):156, August 26, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n9/pdf/scientificamerican08261905-156a.pdf.

Anonymous:1905:AGR

[8525] Anonymous. American gasoline railway motor cars. Scientific American, 93(9):157–158, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
157.pdf.

Anonymous:1905:SM

[8526] Anonymous. The submarine mine. Scientific American, 93(9):160, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n9/pdf/scientificamerican08261905-160.pdf.

Johnson:1905:TMS

[8527] Charles C. Johnson. The training of the modern shop superintendent. Scientific American, 93(9):161–162, August 26, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n9/pdf/scientificamerican08261905-161.pdf.

Anonymous:1905:LCb

[8528] Anonymous. The latest in cycles. Scientific American, 93(9):162, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v93/n9/pdf/scientificamerican08261905-162.pdf.

Holder:1905:R

[8529] Charles Frederick Holder. The remoras. Scientific American, 93 (9):162–163, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
162a.pdf.

Hudson:1905:GSW

[8530] W. S. Hudson. The Great Sea Wall at Galveston. Scientific American, 93(9):163–164, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
163.pdf.

Anonymous:1905:CBAa

[8531] Anonymous. A curious bridge accident. Scientific American, 93 (9):164, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
164.pdf.

Anonymous:1905:OSD

[8532] Anonymous. Opening of the Stone Drydock at the Boston Navy Yard. Scientific American, 93(9):164–165, August 26, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v93/
n9/pdf/scientificamerican08261905-164a.pdf.

Anonymous:1905:BPWbx

[8533] Anonymous. Business and personal wants. Scientific American, 93(9):166, August 26, 1905. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v93/n9/pdf/scientificamerican08261905-
166.pdf.

Anonymous:1906:Ca

[8534] Anonymous. Correspondence. Scientific American, 94(1):9, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-9.pdf.

Anonymous:1906:RPIa

[8535] Anonymous. Recently patented inventions. Scientific American, 94(1):14, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n1/pdf/scientificamerican01061906-
14.pdf.

Anonymous:1906:NQa

[8536] Anonymous. Notes and queries. Scientific American, 94(1):15, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-15.pdf.

Anonymous:1906:BNBa

[8537] Anonymous. Books: New books, etc. Scientific American, 94(1):15, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-15a.pdf.

Anonymous:1906:IIa

[8538] Anonymous. Index of inventions. Scientific American, 94(1):15–19, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-15b.pdf.

Anonymous:1906:PNPa

[8539] Anonymous. Patent notes: Patents. Scientific American, 94(1):19, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-19.pdf.

Anonymous:1906:POa

[8540] Anonymous. The passing of the “Oregon”. Scientific American, 94(1):4, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-4.pdf.

Anonymous:1906:WOF

[8541] Anonymous. Wealth of our farms. Scientific American, 94(1):4, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-4a.pdf.

Anonymous:1906:SPLa

[8542] Anonymous. Single-phase locomotives on direct-current tracks. Scientific American, 94(1):4, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n1/pdf/scientificamerican01061906-
4b.pdf.

Anonymous:1906:WPH

[8543] Anonymous. Water power at high pressure. Scientific American, 94(1):4–5, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n1/pdf/scientificamerican01061906-
4c.pdf.

Anonymous:1906:DCP

[8544] Anonymous. The development of the Capitaine producer gas marine engine. Scientific American, 94(1):5, January 6, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n1/pdf/scientificamerican01061906-5.pdf.

Gilbert:1906:HP

[8545] W. R. Gilbert. Hybridization of plants. Scientific American, 94(1):5, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-5a.pdf.

Anonymous:1906:SNa

[8546] Anonymous. Science notes. Scientific American, 94(1):5, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-5b.pdf.

Boyer:1906:NDQ

[8547] Jacques Boyer. A new determination of the Quito Arc of the Meridian by the French Geodetic Commission of Ecuador. Scientific American, 94(1):6, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n1/pdf/scientificamerican01061906-
6.pdf.

Anonymous:1906:NTT

[8548] Anonymous. The new turbine torpedo of the United States Navy. Scientific American, 94(1):7–8, January 6, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n1/pdf/scientificamerican01061906-7.pdf.

Anonymous:1906:AWP

[8549] Anonymous. Automatic watering pot. Scientific American, 94(1):8, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-8.pdf.

Willey:1906:CD

[8550] Day Allen Willey. The Cavité Drydock. Scientific American, 94 (1):10–11, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n1/pdf/scientificamerican01061906-
10.pdf.

Aaron:1906:LFR

[8551] S. Frank Aaron. A little friend of the rose. Scientific American, 94(1):12, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-12.pdf.

Anonymous:1906:PDa

[8552] Anonymous. Patent department. Scientific American, 94(1):13, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-13.pdf.

Anonymous:1906:SDC

[8553] Anonymous. A simple door check. Scientific American, 94(1):13, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-13a.pdf.

Anonymous:1906:IWT

[8554] Anonymous. Improved water-tube boiler. Scientific American, 94(1):13, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n1/pdf/scientificamerican01061906-13b.pdf.

Anonymous:1906:BPWa

[8555] Anonymous. Business and personal wants. Scientific American, 94(1):14, January 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n1/pdf/scientificamerican01061906-
14a.pdf.

Anonymous:1906:Cb

[8556] Anonymous. Correspondence. Scientific American, 94(10):211, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-211.pdf.

Anonymous:1906:RPIb

[8557] Anonymous. Recently patented inventions. Scientific American, 94 (10):216–217, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
216.pdf.

Anonymous:1906:BPWb

[8558] Anonymous. Business and personal wants. Scientific American, 94(10):217, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
217.pdf.

Anonymous:1906:NQb

[8559] Anonymous. Notes and queries. Scientific American, 94(10):217, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-217a.pdf.

Anonymous:1906:BNBb

[8560] Anonymous. Books: New books, etc. Scientific American, 94(10):218, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-218.pdf.

Anonymous:1906:IIb

[8561] Anonymous. Index of inventions. Scientific American, 94(10):218–223, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-218a.pdf.

Anonymous:1906:PNPb

[8562] Anonymous. Patent notes: Patents. Scientific American, 94(10):223, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-223.pdf.

Anonymous:1906:CF

[8563] Anonymous. A change of front. Scientific American, 94(10):206, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-206.pdf.

Anonymous:1906:SLP

[8564] Anonymous. Some lessons from the past at Panama. Scientific American, 94(10):206, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
206a.pdf.

Anonymous:1906:GDK

[8565] Anonymous. Gatun Dam, the key to the Canal problem. Scientific American, 94(10):206–207, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
206b.pdf.

Anonymous:1906:MAE

[8566] Anonymous. Mikkelsen’s Arctic Expedition. Scientific American, 94(10):207, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
207.pdf.

Anonymous:1906:MF

[8567] Anonymous. The moth and the flame. Scientific American, 94(10):207, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-207a.pdf.

Anonymous:1906:SPLb

[8568] Anonymous. Samuel Pierpont Langley. Scientific American, 94(10):207, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-207b.pdf.

Anonymous:1906:NBB

[8569] Anonymous. The new British battleship “Dreadnought”. Scientific American, 94(10):208–209, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
208.pdf.

Wade:1906:SL

[8570] Herbert T. Wade. Standards of length. Scientific American, 94(10):209, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-209.pdf.

E:1906:MBS

[8571] Thaleon Blake C. E. A method of burning slack coal in a common soft-coal furnace. Scientific American, 94(10):210, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n10/pdf/scientificamerican03101906-210.pdf.

Anonymous:1906:TTP

[8572] Anonymous. A trip through the Panama Canal in 1915. Scientific American, 94(10):212–214, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
212.pdf.

Anonymous:1906:SPCa

[8573] Anonymous. Sanitation of the Panama Canal Zone. Scientific American, 94(10):214–215, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
214.pdf.

Anonymous:1906:NDN

[8574] Anonymous. A new detachable non-skid tire protector. Scientific American, 94(10):215, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
215.pdf.

Mason:1906:ASW

[8575] F. C. Mason. Arrowheads shaped with a wooden implement. Scientific American, 94(10):215, March 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n10/pdf/scientificamerican03101906-
215a.pdf.

Anonymous:1906:SNb

[8576] Anonymous. Science notes. Scientific American, 94(11):227, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n11/pdf/scientificamerican03171906-227c.pdf.

Anonymous:1906:RPIc

[8577] Anonymous. Recently patented inventions. Scientific American, 94(11):237, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
237.pdf.

Anonymous:1906:NQc

[8578] Anonymous. Notes and queries. Scientific American, 94(11):238, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n11/pdf/scientificamerican03171906-238.pdf.

Anonymous:1906:BNBc

[8579] Anonymous. Books: New books, etc. Scientific American, 94(11):239, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n11/pdf/scientificamerican03171906-239.pdf.

Anonymous:1906:IIc

[8580] Anonymous. Index of inventions. Scientific American, 94(11):239–243, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n11/pdf/scientificamerican03171906-239a.pdf.

Anonymous:1906:PNPc

[8581] Anonymous. Patent notes: Patents. Scientific American, 94(11):243, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n11/pdf/scientificamerican03171906-243.pdf.

Anonymous:1906:DNR

[8582] Anonymous. Diversion of Niagara River. Scientific American, 94(11):226, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
226.pdf.

Anonymous:1906:TIT

[8583] Anonymous. Two important tunnel proposals. Scientific American, 94(11):226, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
226a.pdf.

Anonymous:1906:STE

[8584] Anonymous. Steam turbine economy. Scientific American, 94(11):226, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n11/pdf/scientificamerican03171906-226b.pdf.

Anonymous:1906:SBN

[8585] Anonymous. The Shipping Bill in a nutshell. Scientific American, 94(11):226, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
226c.pdf.

Anonymous:1906:LFR

[8586] Anonymous. Is lead a form of radium? Scientific American, 94 (11):226–227, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
226d.pdf.

Anonymous:1906:BLE

[8587] Anonymous. Blue light and electricity as anæsthetics. Scientific American, 94(11):227, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
227.pdf.

Anonymous:1906:APG

[8588] Anonymous. The advantages of producer gas for large power plants. Scientific American, 94(11):227, March 17, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n11/pdf/scientificamerican03171906-227a.pdf.

Anonymous:1906:VCC

[8589] Anonymous. Value of commercial cultures for legumes. Scientific American, 94(11):227, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
227b.pdf.

Anonymous:1906:NAW

[8590] Anonymous. New armored war automobiles. Scientific American, 94(11):228, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
228.pdf.

Dunlap:1906:ATL

[8591] Orrin E. Dunlap. Aluminium transmission line across Niagara Gorge. Scientific American, 94(11):229, March 17, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n11/pdf/scientificamerican03171906-229.pdf.

Honey:1906:MES

[8592] F. R. Honey. Morning and evening stars in 1906. Scientific American, 94(11):230, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
230.pdf.

Anonymous:1906:NACa

[8593] Anonymous. The new armored cruiser “Tennessee”. Scientific American, 94(11):230–231, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
230a.pdf.

Geare:1906:RDR

[8594] Randolph I. Geare. Recently discovered ruins in Rhodesia. Scientific American, 94(11):231–232, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
231.pdf.

Anonymous:1906:GGL

[8595] Anonymous. The Government’s gasoline lifeboats. Scientific American, 94(11):233–234, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
233.pdf.

Anonymous:1906:WTS

[8596] Anonymous. Wireless telegraphy in Southwest Africa. Scientific American, 94(11):234, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
234.pdf.

Ramakers:1906:RMN

[8597] L. Ramakers. The Royal Museum of Natural History, Brussels. Scientific American, 94(11):234–235, March 17, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n11/pdf/scientificamerican03171906-234a.pdf.

Anonymous:1906:IWS

[8598] Anonymous. An improved window screen. Scientific American, 94(11):236, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
236.pdf.

Anonymous:1906:TCD

[8599] Anonymous. Telescoping a car of dynamite. Scientific American, 94(11):236, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
236a.pdf.

Anonymous:1906:EPDa

[8600] Anonymous. Ejector pump for driven wells. Scientific American, 94(11):236, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
236b.pdf.

Anonymous:1906:BPWc

[8601] Anonymous. Business and personal wants. Scientific American, 94(11):237, March 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n11/pdf/scientificamerican03171906-
237a.pdf.

Anonymous:1906:Cc

[8602] Anonymous. Correspondence. Scientific American, 94(12):251, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n12/pdf/scientificamerican03241906-251.pdf.

Anonymous:1906:RPId

[8603] Anonymous. Recently patented inventions. Scientific American, 94(12):257, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
257.pdf.

Anonymous:1906:BPWd

[8604] Anonymous. Business and personal wants. Scientific American, 94(12):257, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
257a.pdf.

Anonymous:1906:NQd

[8605] Anonymous. Notes and queries. Scientific American, 94(12):257–258, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n12/pdf/scientificamerican03241906-257b.pdf.

Anonymous:1906:BNBd

[8606] Anonymous. Books: New books, etc. Scientific American, 94 (12):258–259, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
258.pdf.

Anonymous:1906:IId

[8607] Anonymous. Index of inventions. Scientific American, 94(12):259–263, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n12/pdf/scientificamerican03241906-259.pdf.

Anonymous:1906:PNPd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n12/pdf/scientificamerican03241906-263.pdf.

Anonymous:1906:WN

[8609] Anonymous. “What’s in a name?”. Scientific American, 94(12):246, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n12/pdf/scientificamerican03241906-246.pdf.

Anonymous:1906:NWSa

[8610] Anonymous. New water supply legislation. Scientific American, 94(12):246, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
246a.pdf.

Anonymous:1906:SMC

[8611] Anonymous. Steel manufacturers and the chain bridge. Scientific American, 94(12):246, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
246b.pdf.

Anonymous:1906:TWa

[8612] Anonymous. A timely warning. Scientific American, 94(12):246, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n12/pdf/scientificamerican03241906-246c.pdf.

Anonymous:1906:DLF

[8613] Anonymous. The duration of lightning flashes. Scientific American, 94(12):246–247, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
246d.pdf.

Anonymous:1906:IEW

[8614] Anonymous. The impending exhaustion of the world’s iron supply. Scientific American, 94(12):247, March 24, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n12/pdf/scientificamerican03241906-247.pdf.

Anonymous:1906:EGE

[8615] Anonymous. The efficiency of the gas engine. Scientific American, 94(12):247–248, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
247a.pdf.

Anonymous:1906:EPDb

[8616] Anonymous. Electric power developments at Niagara Falls. — III. Scientific American, 94(12):248–249, March 24, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n12/pdf/scientificamerican03241906-248.pdf.

Anonymous:1906:MIG

[8617] Anonymous. Mounting the 12-inch guns of the “Connecticut”. Scientific American, 94(12):249–250, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
249.pdf.

Anonymous:1906:NWSb

[8618] Anonymous. The new water supply for New York City. Scientific American, 94(12):250–251, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
250.pdf.

Anonymous:1906:DSL

[8619] Anonymous. The diamond shoals lighthouse. Scientific American, 94(12):251–252, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
251a.pdf.

Anonymous:1906:SRF

[8620] Anonymous. Some recent foreign flying machines. Scientific American, 94(12):252–254, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
252.pdf.

Perkins:1906:NFC

[8621] Frank C. Perkins. A new French compound locomotive. Scientific American, 94(12):254, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
254.pdf.

Willey:1906:HSM

[8622] Day Allen Willey. How shingles are made. Scientific American, 94(12):254–256, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
254a.pdf.

Anonymous:1906:ABS

[8623] Anonymous. Air box for stoves. Scientific American, 94(12):256, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n12/pdf/scientificamerican03241906-256.pdf.

Anonymous:1906:AMS

[8624] Anonymous. Apparatus for marking sunken vessels. Scientific American, 94(12):256, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
256a.pdf.

Anonymous:1906:OIa

[8625] Anonymous. Oddities in inventions. Scientific American, 94 (12):256–257, March 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n12/pdf/scientificamerican03241906-
256b.pdf.

Russell:1906:HAa

[8626] Henry Norris Russell. The heavens in April, 1906. Scientific American, 94(13):267, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
267.pdf.

Anonymous:1906:Cd

[8627] Anonymous. Correspondence. Scientific American, 94(13):271, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n13/pdf/scientificamerican03311906-271.pdf.

Anonymous:1906:RPIe

[8628] Anonymous. Recently patented inventions. Scientific American, 94(13):276, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
276.pdf.

Anonymous:1906:BNBe

[8629] Anonymous. Books: New books, etc. Scientific American, 94(13):276, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n13/pdf/scientificamerican03311906-276b.pdf.

Anonymous:1906:IIe

[8630] Anonymous. Index of inventions. Scientific American, 94(13):276–279, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n13/pdf/scientificamerican03311906-276c.pdf.

Anonymous:1906:PNPe

[8631] Anonymous. Patent notes: Patents. Scientific American, 94(13):279, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n13/pdf/scientificamerican03311906-279.pdf.

Anonymous:1906:EDC

[8632] Anonymous. Earthquakes and dam construction at Panama. Scientific American, 94(13):266, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
266.pdf.

Anonymous:1906:QMP

[8633] Anonymous. A quart measure in a pint pot. Scientific American, 94(13):266, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
266a.pdf.

Anonymous:1906:WAE

[8634] Anonymous. What are atoms, electrons, and ions? Scientific American, 94(13):266, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
266b.pdf.

Anonymous:1906:PBE

[8635] Anonymous. Proposed Bill for the Extension of Patents. Scientific American, 94(13):267, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
267a.pdf.

Anonymous:1906:PWE

[8636] Anonymous. People who have eaten books. Scientific American, 94(13):267, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
267b.pdf.

Anonymous:1906:EM

[8637] Anonymous. Electrical music. Scientific American, 94(13):268–269, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n13/pdf/scientificamerican03311906-268.pdf.

Anonymous:1906:SSM

[8638] Anonymous. A simple and speedy motor sleigh. Scientific American, 94(13):269–270, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
269.pdf.

Anonymous:1906:NCa

[8639] Anonymous. The new Cunarders. Scientific American, 94(13):270, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n13/pdf/scientificamerican03311906-270.pdf.

Ramakers:1906:BCS

[8640] L. Ramakers. Boring out columns in solid rock. Scientific American, 94(13):272, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
272.pdf.

Boyer:1906:AFT

[8641] Jacques Boyer. Artistic French taxidermy. Scientific American, 94 (13):272–274, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
272a.pdf.

Anonymous:1906:NTA

[8642] Anonymous. New type of armor-piercing shell. Scientific American, 94(13):274, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
274.pdf.

Metcalf:1906:PBF

[8643] C. Metcalf. A puzzling bit of foundry work. Scientific American, 94(13):274–275, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
274a.pdf.

Anonymous:1906:WVA

[8644] Anonymous. The Walschaert valve on American locomotives. Scientific American, 94(13):275, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
275.pdf.

Anonymous:1906:BPWe

[8645] Anonymous. Business and personal wants. Scientific American, 94(13):276, March 31, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n13/pdf/scientificamerican03311906-
276a.pdf.

Anonymous:1906:Ce

[8646] Anonymous. Correspondence. Scientific American, 94(14):287, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n14/pdf/scientificamerican04071906-287.pdf.

Anonymous:1906:RPIf

[8647] Anonymous. Recently patented inventions. Scientific American, 94(14):293, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
293.pdf.

Anonymous:1906:NQe

[8648] Anonymous. Notes and queries. Scientific American, 94(14):294, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n14/pdf/scientificamerican04071906-294.pdf.

Anonymous:1906:IIf

[8649] Anonymous. Index of inventions. Scientific American, 94(14):294–299, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n14/pdf/scientificamerican04071906-294a.pdf.

Anonymous:1906:PNPf

[8650] Anonymous. Patent notes: Patents. Scientific American, 94(14):299, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n14/pdf/scientificamerican04071906-299.pdf.

Anonymous:1906:BST

[8651] Anonymous. The Behring Sea Tunnel. Scientific American, 94(14):282, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n14/pdf/scientificamerican04071906-282.pdf.

Anonymous:1906:SHM

[8652] Anonymous. Seven-hundred-mile electric transmission. Scientific American, 94(14):282, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
282a.pdf.

Anonymous:1906:CAM

[8653] Anonymous. Cheap alcohol for motors. Scientific American, 94(14):282, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n14/pdf/scientificamerican04071906-282b.pdf.

Anonymous:1906:PAA

[8654] Anonymous. A practical application of the aversion of eels from light. Scientific American, 94(14):282, April 7, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n14/pdf/scientificamerican04071906-282c.pdf.

Anonymous:1906:CLM

[8655] Anonymous. The chances of living in a modern battle. Scientific American, 94(14):282, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
282d.pdf.

Anonymous:1906:OLM

[8656] Anonymous. An open letter to the Mayor. Scientific American, 94(14):283, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
283.pdf.

Anonymous:1906:MPS

[8657] Anonymous. The modern philosopher’s stone. Scientific American, 94(14):283, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
283a.pdf.

Anonymous:1906:NASa

[8658] Anonymous. A new Army signaling apparatus. Scientific American, 94(14):283, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
283b.pdf.

Anonymous:1906:CLSa

[8659] Anonymous. A combined long and short range automobile searchlight. Scientific American, 94(14):284, April 7, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n14/pdf/scientificamerican04071906-284.pdf.

Anonymous:1906:PRE

[8660] Anonymous. Pennsylvania Railroad’s extension to New York and Long Island — the Long Island City Power Station — I. Scientific American, 94(14):285–287, April 7, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n14/pdf/scientificamerican04071906-285.pdf.

Jones:1906:HAT

[8661] Harry Dillon Jones. How artificial teeth are made. Scientific American, 94(14):288, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
288.pdf.

Anonymous:1906:ATW

[8662] Anonymous. Ants and their ways. Scientific American, 94(14):289–290, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n14/pdf/scientificamerican04071906-289.pdf.

Gradenwitz:1906:ETS

[8663] Alfred Gradenwitz. Electric traction on the simplon railway. Scientific American, 94(14):290, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
290.pdf.

Anonymous:1906:WAP

[8664] Anonymous. The Wright aeroplane and its performances. Scientific American, 94(14):291–292, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
291.pdf.

Anonymous:1906:EVL

[8665] Anonymous. An exhaust valve for locomotives. Scientific American, 94(14):292, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
292.pdf.

Anonymous:1906:CBMa

[8666] Anonymous. Cabinet for bread making. Scientific American, 94 (14):292, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
292a.pdf.

Anonymous:1906:BPWf

[8667] Anonymous. Business and personal wants. Scientific American, 94(14):293–294, April 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n14/pdf/scientificamerican04071906-
293a.pdf.

Anonymous:1906:Cg

[8668] Anonymous. Correspondence. Scientific American, 94(15):307, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-307.pdf.

Anonymous:1906:RPIg

[8669] Anonymous. Recently patented inventions. Scientific American, 94(15):313, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
313.pdf.

Anonymous:1906:NQf

[8670] Anonymous. Notes and queries. Scientific American, 94(15):313, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-313b.pdf.

Anonymous:1906:IIg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-313c.pdf.

Anonymous:1906:PNPg

[8672] Anonymous. Patent notes: Patents. Scientific American, 94(15):319, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-319.pdf.

Anonymous:1906:MLT

[8673] Anonymous. More light on turbine economy. Scientific American, 94(15):302, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
302.pdf.

Anonymous:1906:EST

[8674] Anonymous. Effects of snow on the third rail. Scientific American, 94(15):302, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
302a.pdf.

Anonymous:1906:CFS

[8675] Anonymous. Combined floating and stationary power station. Scientific American, 94(15):302, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
302b.pdf.

Anonymous:1906:CHE

[8676] Anonymous. Comparison of heavy electric and steam locomotive performance. Scientific American, 94(15):302, April 14, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n15/pdf/scientificamerican04141906-302c.pdf.

Anonymous:1906:OSF

[8677] Anonymous. The organization of the submarine flotilla of the British Fleet. Scientific American, 94(15):302, April 14, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n15/pdf/scientificamerican04141906-302d.pdf.

Anonymous:1906:ENY

[8678] Anonymous. The economics of New York’s refuse destructor. Scientific American, 94(15):303, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
303.pdf.

Anonymous:1906:PNYa

[8679] Anonymous. Proposed New York, Brockton, and Boston Canal. Scientific American, 94(15):303, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
303a.pdf.

Anonymous:1906:RBA

[8680] Anonymous. Recent balloon ascensions in the vicinity of New York. Scientific American, 94(15):303, April 14, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n15/pdf/scientificamerican04141906-303b.pdf.

Anonymous:1906:SNc

[8681] Anonymous. Science notes. Scientific American, 94(15):303, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-303c.pdf.

Mclure:1906:RMC

[8682] W. Frank M’clure. A railway mail catcher. Scientific American, 94(15):304, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
304.pdf.

Gradenwitz:1906:ARO

[8683] Alfred Gradenwitz. An apparatus for recording the oscillations of lighthouse towers. Scientific American, 94(15):304, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-304a.pdf.

Anonymous:1906:PDC

[8684] Anonymous. A pneumatic deck-calking tool. Scientific American, 94(15):304, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
304b.pdf.

Anonymous:1906:HPSa

[8685] Anonymous. 100-horse-power suction-gas-propelled boat on the River Rhine. Scientific American, 94(15):305–306, April 14, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n15/pdf/scientificamerican04141906-305.pdf.

Anonymous:1906:MRS

[8686] Anonymous. Motor roller skates. Scientific American, 94(15):306, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-306.pdf.

Anonymous:1906:IEI

[8687] Anonymous. An improved engine indicator. Scientific American, 94(15):307, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
307a.pdf.

Anonymous:1906:NPT

[8688] Anonymous. A novel profile tracer. Scientific American, 94(15):308, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-308.pdf.

Anonymous:1906:HEE

[8689] Anonymous. How Easter eggs are made by the thousand. Scientific American, 94(15):308–309, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
308a.pdf.

Day:1906:ISB

[8690] Allen Day. The inundation of the Salton Basin by the Colorado River and how it was caused. Scientific American, 94(15):310–312, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n15/pdf/scientificamerican04141906-310.pdf.

Banks:1906:LWF

[8691] Edgar James Banks. Letter writing five thousand years ago. Scientific American, 94(15):312, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
312.pdf.

Anonymous:1906:BPWg

[8692] Anonymous. Business and personal wants. Scientific American, 94(15):313, April 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n15/pdf/scientificamerican04141906-
313a.pdf.

Anonymous:1906:Ch

[8693] Anonymous. Correspondence. Scientific American, 94(16):327, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n16/pdf/scientificamerican04211906-327.pdf.

Anonymous:1906:RPIh

[8694] Anonymous. Recently patented inventions. Scientific American, 94(16):332–334, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
332a.pdf.

Anonymous:1906:IIh

[8695] Anonymous. Index of inventions. Scientific American, 94(16):334–339, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n16/pdf/scientificamerican04211906-334.pdf.

Anonymous:1906:PNPh

[8696] Anonymous. Patent notes: Patents. Scientific American, 94(16):339, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n16/pdf/scientificamerican04211906-339.pdf.

Anonymous:1906:IPN

[8697] Anonymous. International protection of Niagara Falls. Scientific American, 94(16):322, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
322.pdf.

Anonymous:1906:ECT

[8698] Anonymous. The English channel tunnel. Scientific American, 94(16):322, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
322a.pdf.

Anonymous:1906:EEG

[8699] Anonymous. Electric equipment of a great railroad. Scientific American, 94(16):322, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
322b.pdf.

Anonymous:1906:UB

[8700] Anonymous. Uniformity in battleships. Scientific American, 94(16):322, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n16/pdf/scientificamerican04211906-322c.pdf.

Anonymous:1906:DAN

[8701] Anonymous. The development of aerial navigation for military purposes in Europe. Scientific American, 94(16):322–323, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n16/pdf/scientificamerican04211906-322d.pdf.

Yates:1906:CRI

[8702] L. H. Yates. Canals and railways in India. Scientific American, 94(16):323, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
323.pdf.

Anonymous:1906:ME

[8703] Anonymous. Mosquito extermination. Scientific American, 94(16):323, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n16/pdf/scientificamerican04211906-323a.pdf.

Anonymous:1906:BTK

[8704] Anonymous. The bell tetrahedral kite in wireless telegraphy. Scientific American, 94(16):324, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
324.pdf.

Anonymous:1906:APL

[8705] Anonymous. Armor plate — the last answer of the banker to the burglar. Scientific American, 94(16):325–326, April 21, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n16/pdf/scientificamerican04211906-325.pdf.

Anonymous:1906:IFA

[8706] Anonymous. An improved form of automatic carbureter. Scientific American, 94(16):326, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
326.pdf.

James:1906:OCR

[8707] George Wharton James. The overflow of the Colorado River and the Salton Sea — II. Scientific American, 94(16):328–329, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n16/pdf/scientificamerican04211906-328.pdf.

Anonymous:1906:VPP

[8708] Anonymous. Vesuvius: Past and present. Scientific American, 94 (16):330–331, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
330.pdf.

Anonymous:1906:BPWh

[8709] Anonymous. Business and personal wants. Scientific American, 94(16):332, April 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n16/pdf/scientificamerican04211906-
332.pdf.

Russell:1906:HMa

[8710] Henry Norris Russell. The heavens in May. Scientific American, 94(17):343, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
343.pdf.

Anonymous:1906:RPIi

[8711] Anonymous. Recently patented inventions. Scientific American, 94(17):352–354, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
352.pdf.

Anonymous:1906:NQg

[8712] Anonymous. Notes and queries. Scientific American, 94(17):354–355, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n17/pdf/scientificamerican04281906-354a.pdf.

Anonymous:1906:BNBf

[8713] Anonymous. Books: New books, etc. Scientific American, 94(17):355, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n17/pdf/scientificamerican04281906-355.pdf.

Anonymous:1906:IIi

[8714] Anonymous. Index of inventions. Scientific American, 94(17):355–359, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n17/pdf/scientificamerican04281906-355a.pdf.

Anonymous:1906:Pa

[8715] Anonymous. Patents. Scientific American, 94(17):359, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n17/pdf/scientificamerican04281906-359.pdf.

Anonymous:1906:EPC

[8716] Anonymous. Earthquake-proof construction. Scientific American, 94(17):342, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
342.pdf.

Anonymous:1906:SVA

[8717] Anonymous. Subway ventilation by air valves. Scientific American, 94(17):342, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
342a.pdf.

Anonymous:1906:KD

[8718] Anonymous. The “Kearsarge” disaster. Scientific American, 94(17):342, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n17/pdf/scientificamerican04281906-342b.pdf.

Anonymous:1906:RBP

[8719] Anonymous. The relation between pain and inflammation. Scientific American, 94(17):343, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
343a.pdf.

Anonymous:1906:FEP

[8720] Anonymous. Feeling the Earth’s pulse. Scientific American, 94 (17):344–345, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
344.pdf.

Anonymous:1906:EEL

[8721] Anonymous. Electrical equipment of the Long Island City Power Station. Scientific American, 94(17):345–346, April 28, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n17/pdf/scientificamerican04281906-345.pdf.

Anonymous:1906:SFC

[8722] Anonymous. San Francisco and its catastrophe. Scientific American, 94(17):347, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
347.pdf.

Anonymous:1906:MMP

[8723] Anonymous. Meerschaum and its manufacture into pipes. Scientific American, 94(17):348–349, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
348.pdf.

Anonymous:1906:SNU

[8724] Anonymous. Some novel uses of electricity. Scientific American, 94(17):349–350, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
349.pdf.

Anonymous:1906:FSW

[8725] Anonymous. Franklin’s scientific work. Scientific American, 94 (17):350–351, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
350.pdf.

Anonymous:1906:BPWi

[8726] Anonymous. Business and personal wants. Scientific American, 94(17):354, April 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n17/pdf/scientificamerican04281906-
354.pdf.

Anonymous:1906:Ci

[8727] Anonymous. Correspondence. Scientific American, 94(18):367, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-367.pdf.

Anonymous:1906:RPIj

[8728] Anonymous. Recently patented inventions. Scientific American, 94(18):373, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
373.pdf.

Anonymous:1906:NQh

[8729] Anonymous. Notes and queries. Scientific American, 94(18):374, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-374.pdf.

Anonymous:1906:BNBg

[8730] Anonymous. Books: New books, etc. Scientific American, 94 (18):374–375, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
374a.pdf.

Anonymous:1906:IIj

[8731] Anonymous. Index of inventions. Scientific American, 94(18):375–379, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-375.pdf.

Anonymous:1906:PNPi

[8732] Anonymous. Patent notes: Patents. Scientific American, 94(18):379, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-379.pdf.

Anonymous:1906:CSD

[8733] Anonymous. Courage and sympathy in disaster. Scientific American, 94(18):362, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
362.pdf.

Anonymous:1906:OP

[8734] Anonymous. An opinion from Panama. Scientific American, 94(18):362, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-362a.pdf.

Anonymous:1906:TSS

[8735] Anonymous. Trial speed and sea speed. Scientific American, 94(18):362, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-362b.pdf.

Anonymous:1906:RS

[8736] Anonymous. A retrograde step. Scientific American, 94(18):362, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-362c.pdf.

Anonymous:1906:PSGa

[8737] Anonymous. Prof. Sir George H. Darwin. Scientific American, 94(18):363, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
363.pdf.

Anonymous:1906:PAT

[8738] Anonymous. The proposed amendment to the trade-mark law. Scientific American, 94(18):363, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
363a.pdf.

Anonymous:1906:WVD

[8739] Anonymous. The whistling vibration of a drop. Scientific American, 94(18):363, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
363b.pdf.

Benjamin:1906:NAS

[8740] Marcus Benjamin. The National Academy of Sciences. Scientific American, 94(18):363, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
363c.pdf.

Willey:1906:ACL

[8741] Day Allen Willey. The automatic conveyer in lumbering. Scientific American, 94(18):364–365, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
364.pdf.

Anonymous:1906:ELN

[8742] Anonymous. The electric locomotives of New York, New Haven and Hartford Railroad. Scientific American, 94(18):365, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-365.pdf.

Anonymous:1906:NMH

[8743] Anonymous. A new method of heating switches. Scientific American, 94(18):365–366, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
365a.pdf.

Anonymous:1906:HHM

[8744] Anonymous. The Hendrick Hudson Memorial Bridge. Scientific American, 94(18):366, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
366.pdf.

Field:1906:CNA

[8745] C. Field. Curiosities of Naval architecture — round ships and globular vessels. Scientific American, 94(18):368–369, May 5, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n18/pdf/scientificamerican05051906-368.pdf.

Forstall:1906:HCE

[8746] A. W. Forstall. How to construct an experimental polyphase motor. Scientific American, 94(18):370–371, May 5, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n18/pdf/scientificamerican05051906-370.pdf.

Gradenwitz:1906:SSS

[8747] Alfred Gradenwitz. The sense of smell in snails. Scientific American, 94(18):371, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
371.pdf.

Dale:1906:MPC

[8748] William Dale. A machine for picking cotton. Scientific American, 94(18):371–372, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
371a.pdf.

Anonymous:1906:CDI

[8749] Anonymous. A convenient drafting instrument. Scientific American, 94(18):372, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
372.pdf.

Anonymous:1906:NT

[8750] Anonymous. A novel toy. Scientific American, 94(18):372, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-372a.pdf.

Anonymous:1906:RDB

[8751] Anonymous. Retaining device for barrels. Scientific American, 94(18):372, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
372b.pdf.

Anonymous:1906:BP

[8752] Anonymous. Book purse. Scientific American, 94(18):372, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n18/pdf/scientificamerican05051906-372c.pdf.

Anonymous:1906:BPWj

[8753] Anonymous. Business and personal wants. Scientific American, 94(18):373, May 5, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n18/pdf/scientificamerican05051906-
373a.pdf.

Anonymous:1906:RPIk

[8754] Anonymous. Recently patented inventions. Scientific American, 94(19):398, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
398.pdf.

Anonymous:1906:NQi

[8755] Anonymous. Notes and queries. Scientific American, 94(19):400, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n19/pdf/scientificamerican05121906-400.pdf.

Anonymous:1906:BNBh

[8756] Anonymous. Books: New books, etc. Scientific American, 94(19):400, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n19/pdf/scientificamerican05121906-400a.pdf.

Anonymous:1906:IIk

[8757] Anonymous. Index of inventions. Scientific American, 94(19):401–406, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n19/pdf/scientificamerican05121906-401.pdf.

Anonymous:1906:PNPj

[8758] Anonymous. Patent notes: Patents. Scientific American, 94(19):405, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n19/pdf/scientificamerican05121906-405.pdf.

Anonymous:1906:FET

[8759] Anonymous. A fuel efficiency test of automobiles. Scientific American, 94(19):382, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
382.pdf.

Anonymous:1906:PMD

[8760] Anonymous. Prevention and mastery of disease. Scientific American, 94(19):382, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
382a.pdf.

Anonymous:1906:FQR

[8761] Anonymous. Fireproof qualities of reinforced concrete. Scientific American, 94(19):382, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
382b.pdf.

Anonymous:1906:GCM

[8762] Anonymous. The Government as a cement manufacturer. Scientific American, 94(19):382–383, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
382c.pdf.

Crafts:1906:SFG

[8763] H. A. Crafts. Some features of the great earthquake. Scientific American, 94(19):383, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
383.pdf.

Anonymous:1906:ALR

[8764] Anonymous. The advantages and limitations of reinforced concrete. Scientific American, 94(19):383, May 12, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n19/pdf/scientificamerican05121906-383a.pdf.

Anonymous:1906:TSRa

[8765] Anonymous. Typical systems of reinforced concrete construction. Scientific American, 94(19):384–386, May 12, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n19/pdf/scientificamerican05121906-384.pdf.

Anonymous:1906:CJP

[8766] Anonymous. Concreting the Jerome Park Reservoir. Scientific American, 94(19):386–387, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
386.pdf.

Anonymous:1906:IFC

[8767] Anonymous. The impinging flame in cement burning. Scientific American, 94(19):387, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
387.pdf.

Anonymous:1906:CMM

[8768] Anonymous. Concrete mixing machinery. Scientific American, 94 (19):388–389, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
388.pdf.

Anonymous:1906:CBMb

[8769] Anonymous. Concrete block machines. Scientific American, 94 (19):390–391, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
390.pdf.

Anonymous:1906:SVL

[8770] Anonymous. Some views and lessons of the San Francisco Disaster. Scientific American, 94(19):392–394, May 12, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n19/pdf/scientificamerican05121906-392.pdf.

Anonymous:1906:SRC

[8771] Anonymous. A stairway of reinforced concrete. Scientific American, 94(19):394, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
394.pdf.

Anonymous:1906:IBL

[8772] Anonymous. The Ingalls Building — the largest concrete office building in the world. Scientific American, 94(19):394–395, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n19/pdf/scientificamerican05121906-394a.pdf.

Dunlap:1906:GCR

[8773] Orrin E. Dunlap. A great concrete retaining wall. Scientific American, 94(19):395–396, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
395.pdf.

Anonymous:1906:APC

[8774] Anonymous. The artistic possibilities of cement. Scientific American, 94(19):396, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
396.pdf.

Anonymous:1906:BPWk

[8775] Anonymous. Business and personal wants. Scientific American, 94(19):398, May 12, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n19/pdf/scientificamerican05121906-
398a.pdf.

Anonymous:1906:RPIl

[8776] Anonymous. Recently patented inventions. Scientific American, 94(2):48, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
48.pdf.

Anonymous:1906:NQj

[8777] Anonymous. Notes and queries. Scientific American, 94(2):48, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n2/pdf/scientificamerican01131906-48a.pdf.

Anonymous:1906:BNBi

[8778] Anonymous. Books: New books, etc. Scientific American, 94 (2):48, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
48b.pdf.

Anonymous:1906:IIl

[8779] Anonymous. Index of inventions. Scientific American, 94(2):58–63, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n2/pdf/scientificamerican01131906-58.pdf.

Anonymous:1906:PNPk

[8780] Anonymous. Patent notes: Patents. Scientific American, 94(2):63, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n2/pdf/scientificamerican01131906-63.pdf.

Anonymous:1906:FDC

[8781] Anonymous. Future of direct-current traction. Scientific American, 94(2):22, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
22.pdf.

Anonymous:1906:DST

[8782] Anonymous. Designing the steam turbine. Scientific American, 94(2):22, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
22a.pdf.

Anonymous:1906:A

[8783] Anonymous. The automobile in 1906. Scientific American, 94 (2):22–23, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
22b.pdf.

Anonymous:1906:PEA

[8784] Anonymous. Progress with the electric automobile. Scientific American, 94(2):23, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
23.pdf.

Smith:1906:IDM

[8785] K. L. Smith. Increasing demand for the Mississippi pearl. Scientific American, 94(2):23, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
23a.pdf.

Anonymous:1906:SNd

[8786] Anonymous. Science notes. Scientific American, 94(2):23, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n2/pdf/scientificamerican01131906-23b.pdf.

Anonymous:1906:CCA

[8787] Anonymous. Crossing the continent by automobile. Scientific American, 94(2):24–26, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
24.pdf.

Anonymous:1906:MEB

[8788] Anonymous. A motor equipment for bicycles. Scientific American, 94(2):26, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
26.pdf.

Anonymous:1906:LFP

[8789] Anonymous. A light four-passenger runabout. Scientific American, 94(2):26, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
26a.pdf.

Anonymous:1906:LSR

[8790] Anonymous. A light and speedy runabout. Scientific American, 94(2):26, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
26b.pdf.

Anonymous:1906:FCA

[8791] Anonymous. A four-cylinder air-cooled runabout. Scientific American, 94(2):27, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
27.pdf.

Anonymous:1906:FMS

[8792] Anonymous. The Frayer–Miller six-cylinder air-cooled motor. Scientific American, 94(2):28, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
28.pdf.

Anonymous:1906:MAC

[8793] Anonymous. The Marmon air-cooled touring car. Scientific American, 94(2):28, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
28a.pdf.

Anonymous:1906:RCM

[8794] Anonymous. A revolving-cylinder motor car. Scientific American, 94(2):29, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
29.pdf.

Anonymous:1906:NFC

[8795] Anonymous. A new four-cylinder air-cooled motor and selectivetype transmission. Scientific American, 94(2):29, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n2/pdf/scientificamerican01131906-29a.pdf.

Anonymous:1906:CSC

[8796] Anonymous. The Cadillac single-cylinder light touring car. Scientific American, 94(2):30, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
30.pdf.

Anonymous:1906:OTC

[8797] Anonymous. The Oldsmobile two-cycle and four-cycle touring cars. Scientific American, 94(2):30, January 13, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n2/pdf/scientificamerican01131906-30a.pdf.

Anonymous:1906:DLT

[8798] Anonymous. A detachable limousine touring car. Scientific American, 94(2):31, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
31.pdf.

Anonymous:1906:PHP

[8799] Anonymous. The Peerless 30-horse-power touring car. Scientific American, 94(2):31, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
31a.pdf.

Anonymous:1906:HPA

[8800] Anonymous. High-powered American touring cars. Scientific American, 94(2):32, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
32.pdf.

Anonymous:1906:NCT

[8801] Anonymous. A novel 4-cylinder touring car. Scientific American, 94(2):34, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
34.pdf.

Anonymous:1906:SNAa

[8802] Anonymous. A silent non-adjustable ball bearing. Scientific American, 94(2):34, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
34a.pdf.

Anonymous:1906:ANI

[8803] Anonymous. Automobile novelties and improvements. Scientific American, 94(2):36, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
36.pdf.

Anonymous:1906:DSA

[8804] Anonymous. Danger signs for automobilists. Scientific American, 94(2):36, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
36a.pdf.

Anonymous:1906:ELC

[8805] Anonymous. Electric light for closed carriages. Scientific American, 94(2):36, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
36b.pdf.

Anonymous:1906:NSW

[8806] Anonymous. New spring wheels. Scientific American, 94(2):36, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n2/pdf/scientificamerican01131906-36c.pdf.

Anonymous:1906:DIO

[8807] Anonymous. A dashboard ignition outfit. Scientific American, 94(2):38, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
38.pdf.

Anonymous:1906:IUJ

[8808] Anonymous. An improved universal joint. Scientific American, 94(2):38, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
38a.pdf.

Anonymous:1906:WCU

[8809] Anonymous. Wheel clutch for use with a solid rear axle. Scientific American, 94(2):38, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
38b.pdf.

Anonymous:1906:TSRb

[8810] Anonymous. The Thomas safety ratchet device. Scientific American, 94(2):38–40, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
38c.pdf.

Anonymous:1906:ASA

[8811] Anonymous. Automobile shock-absorbers. Scientific American, 94(2):40, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
40.pdf.

Anonymous:1906:FAC

[8812] Anonymous. The Franklin air-cooled cars. Scientific American, 94(2):40, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
40a.pdf.

Gradenwitz:1906:SS

[8813] Alfred Gradenwitz. Snow sculptures. Scientific American, 94 (2):42, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
42.pdf.

Guarini:1906:EH

[8814] Emile Guarini. Electricity in hospitals. Scientific American, 94 (2):44–46, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
44.pdf.

Gradenwitz:1906:NTP

[8815] Alfred Gradenwitz. A new type of portable wireless telegraph station. Scientific American, 94(2):46, January 13, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n2/pdf/scientificamerican01131906-46.pdf.

Anonymous:1906:CME

[8816] Anonymous. A clever mechanical and electrical automaton. Scientific American, 94(2):46, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
46a.pdf.

Anonymous:1906:BPWl

[8817] Anonymous. Business and personal wants. Scientific American, 94(2):49, January 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n2/pdf/scientificamerican01131906-
49.pdf.

Anonymous:1906:RPIm

[8818] Anonymous. Recently patented inventions. Scientific American, 94(20):421, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
421.pdf.

Anonymous:1906:NQk

[8819] Anonymous. Notes and queries. Scientific American, 94(20):422, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n20/pdf/scientificamerican05191906-422.pdf.

Anonymous:1906:IIm

[8820] Anonymous. Index of inventions. Scientific American, 94(20):422–427, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n20/pdf/scientificamerican05191906-422a.pdf.

Anonymous:1906:PNPl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n20/pdf/scientificamerican05191906-427.pdf.

Anonymous:1906:PCT

[8822] Anonymous. Panama Canal and the Tehuantepec Railroad. Scientific American, 94(20):410, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
410.pdf.

Anonymous:1906:SSRa

[8823] Anonymous. Superheated steam in railway service. Scientific American, 94(20):410, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
410a.pdf.

Anonymous:1906:RRN

[8824] Anonymous. Rebuilding the Russian Navy. Scientific American, 94(20):410, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
410b.pdf.

Anonymous:1906:DAH

[8825] Anonymous. The discovery of the Ancient Hebrew Temple of Onias in Egypt. Scientific American, 94(20):410, May 19, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n20/pdf/scientificamerican05191906-410c.pdf.

Anonymous:1906:EAPa

[8826] Anonymous. Exports of agricultural products. Scientific American, 94(20):411, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
411.pdf.

Anonymous:1906:EC

[8827] Anonymous. Electricity for canals. Scientific American, 94(20):411, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n20/pdf/scientificamerican05191906-411a.pdf.

Anonymous:1906:LKC

[8828] Anonymous. Lord Kelvin’s conception of an atom. Scientific American, 94(20):411, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
411b.pdf.

Anonymous:1906:SNe

[8829] Anonymous. Science notes. Scientific American, 94(20):411, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n20/pdf/scientificamerican05191906-411c.pdf.

Anonymous:1906:GL

[8830] Anonymous. Gasoline locomotives. Scientific American, 94(20):412, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n20/pdf/scientificamerican05191906-412.pdf.

Collins:1906:NME

[8831] A. Frederick Collins. A new machine for engraving steel plates. Scientific American, 94(20):412, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
412a.pdf.

Anonymous:1906:WCA

[8832] Anonymous. The winning cars in the Automobile Club of America’s two-gallon fuel efficiency contest. Scientific American, 94(20):414, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n20/pdf/scientificamerican05191906-414.pdf.

Anonymous:1906:TEP

[8833] Anonymous. A thermo-electric pyrometer of low resistance and compensator. Scientific American, 94(20):415, May 19, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n20/pdf/scientificamerican05191906-415.pdf.

Anonymous:1906:APMa

[8834] Anonymous. The art of piano making — I. Scientific American, 94(20):416–418, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
416.pdf.

Larkin:1906:EO

[8835] Edgar L. Larkin. Earthquake observations. Scientific American, 94(20):418, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
418.pdf.

Inkersley:1906:EEF

[8836] Arthur Inkersley. Effects of the earthquake and fire upon the City of San Francisco and its buildings. Scientific American, 94(20):418–420, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n20/pdf/scientificamerican05191906-418a.pdf.

Anonymous:1906:IAC

[8837] Anonymous. An improved adjustable clutch. Scientific American, 94(20):420, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
420.pdf.

Anonymous:1906:IHO

[8838] Anonymous. An improved hand-operated meat press. Scientific American, 94(20):420, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
420a.pdf.

Anonymous:1906:WFR

[8839] Anonymous. Water-feed regulator with automatic alarm. Scientific American, 94(20):420, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
420b.pdf.

Anonymous:1906:BPWm

[8840] Anonymous. Business and personal wants. Scientific American, 94(20):421, May 19, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n20/pdf/scientificamerican05191906-
421a.pdf.

Anonymous:1906:Cj

[8841] Anonymous. Correspondence. Scientific American, 94(21):435, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n21/pdf/scientificamerican05261906-435.pdf.

Anonymous:1906:RPIn

[8842] Anonymous. Recently patented inventions. Scientific American, 94(21):441, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
441.pdf.

Anonymous:1906:NQl

[8843] Anonymous. Notes and queries. Scientific American, 94(21):442, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n21/pdf/scientificamerican05261906-442.pdf.

Anonymous:1906:BNBj

[8844] Anonymous. Books: New books, etc. Scientific American, 94 (21):442–443, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
442a.pdf.

Anonymous:1906:IIn

[8845] Anonymous. Index of inventions. Scientific American, 94(21):443–447, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n21/pdf/scientificamerican05261906-443.pdf.

Anonymous:1906:PNPm

[8846] Anonymous. Patent notes: Patents. Scientific American, 94(21):447, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n21/pdf/scientificamerican05261906-447.pdf.

Anonymous:1906:CRT

[8847] Anonymous. Conditions of the rapid transit tunnel. Scientific American, 94(21):430, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
430.pdf.

Anonymous:1906:EAL

[8848] Anonymous. Efficiency of the American locomotive. Scientific American, 94(21):430, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
430a.pdf.

Anonymous:1906:DWTa

[8849] Anonymous. Directed wireless telegraph messages. Scientific American, 94(21):430–431, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
430b.pdf.

Anonymous:1906:MT

[8850] Anonymous. The manufacture of turpentine. Scientific American, 94(21):431, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
431.pdf.

Anonymous:1906:NCP

[8851] Anonymous. A new and cheap process for generating hydrogen. Scientific American, 94(21):431, May 26, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n21/pdf/scientificamerican05261906-431a.pdf.

Anonymous:1906:RGE

[8852] Anonymous. A Russian gasoline–electric train. Scientific American, 94(21):431, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
431b.pdf.

Anonymous:1906:APMb

[8853] Anonymous. The art of piano making — II. Scientific American, 94(21):432–435, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
432.pdf.

Anonymous:1906:AS

[8854] Anonymous. An automobile school. Scientific American, 94(21): 435–436, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
435a.pdf.

Anonymous:1906:LAT

[8855] Anonymous. The Lorimer automatic telephone. Scientific American, 94(21):437–438, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
437.pdf.

Anonymous:1906:PRT

[8856] Anonymous. Pennsylvania Railroad’s Terminal Station, New York City. Scientific American, 94(21):438–439, May 26, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n21/pdf/scientificamerican05261906-438.pdf.

Anonymous:1906:LGI

[8857] Anonymous. The Laocoön Group as it ought to be. Scientific American, 94(21):439, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
439.pdf.

Anonymous:1906:NVM

[8858] Anonymous. The new Vickers–Maxim 12-inch breech-loading wirewound gun. Scientific American, 94(21):440, May 26, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n21/pdf/scientificamerican05261906-440.pdf.

Anonymous:1906:BPWn

[8859] Anonymous. Business and personal wants. Scientific American, 94(21):441, May 26, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n21/pdf/scientificamerican05261906-
441a.pdf.

Russell:1906:HJa

[8860] Henry Norris Russell. The heavens in June. Scientific American, 94(22):451, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
451.pdf.

Anonymous:1906:Ck

[8861] Anonymous. Correspondence. Scientific American, 94(22):454–455, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n22/pdf/scientificamerican06021906-454.pdf.

Anonymous:1906:RPIo

[8862] Anonymous. Recently patented inventions. Scientific American, 94(22):461, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
461.pdf.

Anonymous:1906:NQm

[8863] Anonymous. Notes and queries. Scientific American, 94(22):461, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n22/pdf/scientificamerican06021906-461b.pdf.

Anonymous:1906:BNBk

[8864] Anonymous. Books: New books, etc. Scientific American, 94(22):461, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n22/pdf/scientificamerican06021906-461c.pdf.

Anonymous:1906:IIo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n22/pdf/scientificamerican06021906-461d.pdf.

Anonymous:1906:PNPn

[8866] Anonymous. Patent notes: Patents. Scientific American, 94(22):466, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n22/pdf/scientificamerican06021906-466.pdf.

Anonymous:1906:RTE

[8867] Anonymous. The rapid transit East River Tunnel defects. Scientific American, 94(22):450, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
450.pdf.

Anonymous:1906:PCPa

[8868] Anonymous. Present conditions at Panama. Scientific American, 94(22):450, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
450a.pdf.

Anonymous:1906:REP

[8869] Anonymous. Recent experiments in the preservation of meat. Scientific American, 94(22):450, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
450b.pdf.

Anonymous:1906:DAF

[8870] Anonymous. Denatured alcohol freed from taxation. Scientific American, 94(22):450, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
450c.pdf.

Anonymous:1906:NCb

[8871] Anonymous. A new compound. Scientific American, 94(22):451, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n22/pdf/scientificamerican06021906-451a.pdf.

Anonymous:1906:UVA

[8872] Anonymous. A unique vacuum apparatus. Scientific American, 94(22):451, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
451b.pdf.

Anonymous:1906:LRB

[8873] Anonymous. Lecture-room boomerang. Scientific American, 94 (22):451, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
451c.pdf.

Anonymous:1906:BFS

[8874] Anonymous. The Brierley fog-signaling apparatus for railroads. Scientific American, 94(22):452–453, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
452.pdf.

Anonymous:1906:EPR

[8875] Anonymous. The excavation for the Pennsylvania Railroad Station, New York. Scientific American, 94(22):453–454, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n22/pdf/scientificamerican06021906-453.pdf.

Anonymous:1906:AVD

[8876] Anonymous. Apparatus for viewing diffraction color photographs. Scientific American, 94(22):455, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
455.pdf.

Gradenwitz:1906:WS

[8877] Alfred Gradenwitz. Water sterilization. Scientific American, 94 (22):456–458, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
456.pdf.

Holder:1906:TWS

[8878] Charles Frederick Holder. Through a window of the sea. Scientific American, 94(22):458–459, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
458.pdf.

Anonymous:1906:TEA

[8879] Anonymous. Thomas A. Edison automobiling to the South. Scientific American, 94(22):460, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
460.pdf.

Anonymous:1906:BPWo

[8880] Anonymous. Business and personal wants. Scientific American, 94(22):461, June 2, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n22/pdf/scientificamerican06021906-
461a.pdf.

Anonymous:1906:Cl

[8881] Anonymous. Correspondence. Scientific American, 94(23):474–475, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-474.pdf.

Anonymous:1906:RPIp

[8882] Anonymous. Recently patented inventions. Scientific American, 94(23):481–482, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
481.pdf.

Anonymous:1906:NQn

[8883] Anonymous. Notes and queries. Scientific American, 94(23):482–483, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-482a.pdf.

Anonymous:1906:BNBl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-483.pdf.

Anonymous:1906:IIp

[8885] Anonymous. Index of inventions. Scientific American, 94(23):483–487, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-483a.pdf.

Anonymous:1906:PNPo

[8886] Anonymous. Patent notes: Patents. Scientific American, 94(23):487, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-487.pdf.

Anonymous:1906:SFE

[8887] Anonymous. The San Francisco Earthquake and the Senate Canal Committee. Scientific American, 94(23):470, June 9, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n23/pdf/scientificamerican06091906-470.pdf.

Anonymous:1906:SJN

[8888] Anonymous. Strength of the Japanese Navy. Scientific American, 94(23):470, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
470a.pdf.

Anonymous:1906:DA

[8889] Anonymous. Denaturized alcohol. Scientific American, 94(23):470, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-470b.pdf.

Anonymous:1906:SFA

[8890] Anonymous. Some facts about Portland cement. Scientific American, 94(23):470–471, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
470c.pdf.

Anonymous:1906:ERS

[8891] Anonymous. Ehrlich’s remarkable studies of cancer in mice. Scientific American, 94(23):471, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
471.pdf.

Anonymous:1906:CRA

[8892] Anonymous. Concrete: Its rise and its applications. Scientific American, 94(23):471, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
471a.pdf.

Anonymous:1906:FL

[8893] Anonymous. Fluid lenses. Scientific American, 94(23):471, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-471b.pdf.

Anonymous:1906:RHB

[8894] Anonymous. Resistance of the human body to automobile accidents. Scientific American, 94(23):471, June 9, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n23/pdf/scientificamerican06091906-471c.pdf.

Anonymous:1906:WDSa

[8895] Anonymous. Why do stars seem rayed? Scientific American, 94(23):471, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
471d.pdf.

Anonymous:1906:RCRa

[8896] Anonymous. The reconstructed Christie racer and its record at the automobile carnival. Scientific American, 94(23):472, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-472.pdf.

Anonymous:1906:UPM

[8897] Anonymous. Union Pacific Motor Car No. 7. Scientific American, 94(23):472–473, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
472a.pdf.

Anonymous:1906:ETT

[8898] Anonymous. The earthquake and the tracks of a trolley line. Scientific American, 94(23):473, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
473.pdf.

Anonymous:1906:FSF

[8899] Anonymous. Fireproofing at San Francisco. Scientific American, 94(23):473–474, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
473a.pdf.

Guarini:1906:MCG

[8900] Emile Guarini. Mechanical compression of the ground in the construction of foundations. Scientific American, 94(23):476–477, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-476.pdf.

Willey:1906:SBS

[8901] Day Allen Willey. The snag boats of the South. Scientific American, 94(23):478, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
478.pdf.

Anonymous:1906:PTL

[8902] Anonymous. Power transmission line of the Long Island Railroad. Scientific American, 94(23):478–479, June 9, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n23/pdf/scientificamerican06091906-478a.pdf.

Anonymous:1906:PDb

[8903] Anonymous. Patent department. Scientific American, 94(23):480, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n23/pdf/scientificamerican06091906-480.pdf.

Anonymous:1906:NRL

[8904] Anonymous. A novel reflecting lamp globe. Scientific American, 94(23):480, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
480a.pdf.

Anonymous:1906:TGB

[8905] Anonymous. Triple gage, butt gage, and tool holder. Scientific American, 94(23):480, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
480b.pdf.

Anonymous:1906:BPWp

[8906] Anonymous. Business and personal wants. Scientific American, 94(23):482, June 9, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n23/pdf/scientificamerican06091906-
482.pdf.

Anonymous:1906:Cm

[8907] Anonymous. Correspondence. Scientific American, 94(24):495, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n24/pdf/scientificamerican06161906-495.pdf.

Anonymous:1906:RPIq

[8908] Anonymous. Recently patented inventions. Scientific American, 94(24):501, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
501.pdf.

Anonymous:1906:NQo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n24/pdf/scientificamerican06161906-501c.pdf.

Anonymous:1906:IIq

[8910] Anonymous. Index of inventions. Scientific American, 94(24):502–507, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n24/pdf/scientificamerican06161906-502.pdf.

Anonymous:1906:PNPp

[8911] Anonymous. Patent notes: Patents. Scientific American, 94(24):507, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n24/pdf/scientificamerican06161906-507.pdf.

Anonymous:1906:SWC

[8912] Anonymous. Steel and wooden cars in the subway fire. Scientific American, 94(24):490, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
490.pdf.

Anonymous:1906:ORS

[8913] Anonymous. Ocean races for small boats. Scientific American, 94(24):490, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
490a.pdf.

Anonymous:1906:DWTb

[8914] Anonymous. A decade of wireless telegraphy. Scientific American, 94(24):490–491, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
490b.pdf.

Dodge:1906:NPF

[8915] Charles Richards Dodge. The need of a pure-food law. Scientific American, 94(24):491, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
491.pdf.

Anonymous:1906:BNP

[8916] Anonymous. The British Naval Programme for 1906. Scientific American, 94(24):491, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
491a.pdf.

Anonymous:1906:SJI

[8917] Anonymous. Steam jacketing investigations. Scientific American, 94(24):491, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
491b.pdf.

Anonymous:1906:DMN

[8918] Anonymous. Dried milk — a new process. Scientific American, 94(24):492, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
492.pdf.

Gradenwitz:1906:MSI

[8919] Alfred Gradenwitz. The magic sphere — an illusion apparatus in which curious optical effects are produced by centrifugal force. Scientific American, 94(24):493, June 16, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n24/pdf/scientificamerican06161906-493.pdf.

Anonymous:1906:HAB

[8920] Anonymous. The hot-air boiler. Scientific American, 94(24):494, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n24/pdf/scientificamerican06161906-494.pdf.

Anonymous:1906:LLS

[8921] Anonymous. The latest and largest steamship. Scientific American, 94(24):496–498, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
496.pdf.

Anonymous:1906:EER

[8922] Anonymous. European earthquake recorders. Scientific American, 94(24):498–499, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
498.pdf.

Bresler:1906:TDM

[8923] Arthur F. Bresler. Tempered ductile metals. Scientific American, 94(24):500, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
500.pdf.

Anonymous:1906:WMB

[8924] Anonymous. Wellman’s motor bicycle sled. Scientific American, 94(24):500, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
500a.pdf.

Anonymous:1906:SAM

[8925] Anonymous. A simple adding machine. Scientific American, 94(24):501, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n24/pdf/scientificamerican06161906-501a.pdf.

Anonymous:1906:BPWq

[8926] Anonymous. Business and personal wants. Scientific American, 94(24):501, June 16, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n24/pdf/scientificamerican06161906-
501b.pdf.

Anonymous:1906:RPIr

[8927] Anonymous. Recently patented inventions. Scientific American, 94(25):521, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
521.pdf.

Anonymous:1906:NQp

[8928] Anonymous. Notes and queries. Scientific American, 94(25):521–522, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n25/pdf/scientificamerican06231906-521b.pdf.

Anonymous:1906:BNBm

[8929] Anonymous. Books: New books, etc. Scientific American, 94 (25):522–523, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
522.pdf.

Anonymous:1906:IIr

[8930] Anonymous. Index of inventions. Scientific American, 94(25):523–527, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n25/pdf/scientificamerican06231906-523.pdf.

Anonymous:1906:PNPq

[8931] Anonymous. Patent notes: Patents. Scientific American, 94(25):527, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n25/pdf/scientificamerican06231906-527.pdf.

Anonymous:1906:PCA

[8932] Anonymous. Present condition of the Assouan Dam. Scientific American, 94(25):510, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
510.pdf.

Anonymous:1906:TWT

[8933] Anonymous. Transatlantic wireless telegraph troubles. Scientific American, 94(25):510, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
510a.pdf.

Anonymous:1906:IIBa

[8934] Anonymous. The influence of increased barometric pressure on the human body. Scientific American, 94(25):510–511, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n25/pdf/scientificamerican06231906-510b.pdf.

Hale:1906:IUV

[8935] Henry Hale. The industrial uses and value of alcohol. Scientific American, 94(25):511, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
511.pdf.

Anonymous:1906:ANa

[8936] Anonymous. Automobile notes. Scientific American, 94(25):511–512, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n25/pdf/scientificamerican06231906-511a.pdf.

Anonymous:1906:ARA

[8937] Anonymous. Artificial regulation of atmospheric humidity and temperature. Scientific American, 94(25):512–513, June 23, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n25/pdf/scientificamerican06231906-512.pdf.

Michaud:1906:CT

[8938] Gustave Michaud. A chemical trick. Scientific American, 94(25):513, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n25/pdf/scientificamerican06231906-513.pdf.

Prindle:1906:AI

[8939] Edwin J. Prindle. The art of inventing. Scientific American, 94 (25):513–515, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
513a.pdf.

Anonymous:1906:OHT

[8940] Anonymous. A one-hundred-ton derrick crane for fitting out vessels. Scientific American, 94(25):515, June 23, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n25/pdf/scientificamerican06231906-515.pdf.

Anonymous:1906:LBE

[8941] Anonymous. A locomotive boiler explosion. Scientific American, 94(25):515–516, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
515a.pdf.

Boyer:1906:FCM

[8942] Jacques Boyer. French candle making. Scientific American, 94 (25):516–518, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
516.pdf.

Anonymous:1906:PDd

[8943] Anonymous. Patent department. Scientific American, 94(25):518, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n25/pdf/scientificamerican06231906-518.pdf.

Anonymous:1906:DMP

[8944] Anonymous. Device for moving pump piston rods. Scientific American, 94(25):518, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
518a.pdf.

Anonymous:1906:IS

[8945] Anonymous. Improved scaffolding. Scientific American, 94(25): 518–519, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
518b.pdf.

Anonymous:1906:INS

[8946] Anonymous. An improved non-skidding tire. Scientific American, 94(25):519, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
519.pdf.

Anonymous:1906:NIP

[8947] Anonymous. A new insulator pin. Scientific American, 94(25):519, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n25/pdf/scientificamerican06231906-519a.pdf.

Anonymous:1906:OIb

[8948] Anonymous. Oddities in invention. Scientific American, 94(25): 519–521, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
519b.pdf.

Anonymous:1906:BPWr

[8949] Anonymous. Business and personal wants. Scientific American, 94(25):521, June 23, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n25/pdf/scientificamerican06231906-
521a.pdf.

Russell:1906:HJb

[8950] Henry Norris Russell. The heavens in July. Scientific American, 94(26):531, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
531.pdf.

Anonymous:1906:Cn

[8951] Anonymous. Correspondence. Scientific American, 94(26):535, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n26/pdf/scientificamerican06301906-535.pdf.

Anonymous:1906:RPIs

[8952] Anonymous. Recently patented inventions. Scientific American, 94(26):540, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
540.pdf.

Anonymous:1906:NQq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n26/pdf/scientificamerican06301906-540b.pdf.

Anonymous:1906:IIs

[8954] Anonymous. Index of inventions. Scientific American, 94(26):540–544, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n26/pdf/scientificamerican06301906-540c.pdf.

Anonymous:1906:I

[8955] Anonymous. Index. Scientific American, 94(26):545–546, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n26/pdf/scientificamerican06301906-545.pdf.

Anonymous:1906:PNPr

[8956] Anonymous. Patent notes: Patents. Scientific American, 94(26):547, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n26/pdf/scientificamerican06301906-547.pdf.

Anonymous:1906:LCP

[8957] Anonymous. A lock canal for Panama. Scientific American, 94(26):530, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n26/pdf/scientificamerican06301906-530.pdf.

Anonymous:1906:RSF

[8958] Anonymous. Redwood in the San Francisco Fire. Scientific American, 94(26):530, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
530a.pdf.

Anonymous:1906:CWI

[8959] Anonymous. Comparison of western irrigation reservoirs and the new Croton Dam. Scientific American, 94(26):530, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n26/pdf/scientificamerican06301906-530b.pdf.

Anonymous:1906:SSA

[8960] Anonymous. Seeds and suspended animation. Scientific American, 94(26):530, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
530c.pdf.

Anonymous:1906:OCP

[8961] Anonymous. Orientation of carrier pigeons. Scientific American, 94(26):531, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
531a.pdf.

Anonymous:1906:IAa

[8962] Anonymous. The “Italia” airship. Scientific American, 94(26):532, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n26/pdf/scientificamerican06301906-532.pdf.

Gradenwitz:1906:LVC

[8963] Alfred Gradenwitz. The laying of a 10,000-volt cable. Scientific American, 94(26):532–533, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
532a.pdf.

Beasley:1906:MFO

[8964] Walter L. Beasley. Making the flags of our warships. Scientific American, 94(26):533–534, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
533.pdf.

Wiles:1906:HGS

[8965] Charles Wiles. How glaciers are studied. Scientific American, 94 (26):534–535, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
534.pdf.

Willey:1906:OFS

[8966] Day Allen Willey. Oil fuel on Southwestern railroads. Scientific American, 94(26):536, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
536.pdf.

Anonymous:1906:SNYa

[8967] Anonymous. Supplying New York with bluefish. Scientific American, 94(26):537–538, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
537.pdf.

Beardsley:1906:EWR

[8968] D. L. Beardsley. An electrolytic wireless receiver. Scientific American, 94(26):538, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
538.pdf.

Anonymous:1906:RBD

[8969] Anonymous. Replacing the Broadway Drawbridge with a new span. Scientific American, 94(26):539, June 30, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n26/pdf/scientificamerican06301906-539.pdf.

Anonymous:1906:BPWs

[8970] Anonymous. Business and personal wants. Scientific American, 94(26):540, June 30, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n26/pdf/scientificamerican06301906-
540a.pdf.

Anonymous:1906:Co

[8971] Anonymous. Correspondence. Scientific American, 94(3):71, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n3/pdf/scientificamerican01201906-71a.pdf.

Anonymous:1906:RPIt

[8972] Anonymous. Recently patented inventions. Scientific American, 94(3):77–78, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
77.pdf.

Anonymous:1906:NQr

[8973] Anonymous. Notes and queries. Scientific American, 94(3):78–79, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n3/pdf/scientificamerican01201906-78a.pdf.

Anonymous:1906:BNBn

[8974] Anonymous. Books: New books, etc. Scientific American, 94 (3):79, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
79.pdf.

Anonymous:1906:IIt

[8975] Anonymous. Index of inventions. Scientific American, 94(3):79–83, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n3/pdf/scientificamerican01201906-79a.pdf.

Anonymous:1906:PNPs

[8976] Anonymous. Patent notes: Patents. Scientific American, 94(3):83, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n3/pdf/scientificamerican01201906-83.pdf.

Anonymous:1906:PGN

[8977] Anonymous. Pipe galleries in the new subway. Scientific American, 94(3):66, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
66.pdf.

Anonymous:1906:WSU

[8978] Anonymous. We should use the metric system. Scientific American, 94(3):66, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
66a.pdf.

Anonymous:1906:ASPa

[8979] Anonymous. Alternative schemes for the Panama Canal. Scientific American, 94(3):66, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
66b.pdf.

Anonymous:1906:ETI

[8980] Anonymous. Electro-technical industries. Scientific American, 94 (3):66–67, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
66c.pdf.

Keane:1906:SCE

[8981] Arthur H. J. Keane. Some costly eggs. Scientific American, 94(3):67, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
67.pdf.

Anonymous:1906:SNf

[8982] Anonymous. Science notes. Scientific American, 94(3):67, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n3/pdf/scientificamerican01201906-67a.pdf.

Willey:1906:FDC

[8983] Day Allen Willey. Floating depot for coaling warships. Scientific American, 94(3):68, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
68.pdf.

Anonymous:1906:PEP

[8984] Anonymous. Proposed excavation of the Panama Canal by floating dredges. Scientific American, 94(3):68–70, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n3/pdf/scientificamerican01201906-68a.pdf.

Howell:1906:RMC

[8985] S. M. Howell. A rational method of cooling gas-engine cylinders. Scientific American, 94(3):70–71, January 20, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n3/pdf/scientificamerican01201906-70.pdf.

Dodge:1906:BF

[8986] Charles Richards Dodge. Bromelia fiber. Scientific American, 94(3):71, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
71.pdf.

Anonymous:1906:CDPa

[8987] Anonymous. A chair designed for the prevention of seasickness. Scientific American, 94(3):72, January 20, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n3/pdf/scientificamerican01201906-72.pdf.

Anonymous:1906:NCM

[8988] Anonymous. A new composing machine. Scientific American, 94(3):73, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
73.pdf.

Anonymous:1906:HPSb

[8989] Anonymous. Historical photograph from the Sea of Japan. Scientific American, 94(3):74–76, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
74.pdf.

Anonymous:1906:ICS

[8990] Anonymous. Improved cushion sole for shoes. Scientific American, 94(3):76, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
76.pdf.

Anonymous:1906:IFT

[8991] Anonymous. An improved form of telephone instrument. Scientific American, 94(3):76, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
76a.pdf.

Anonymous:1906:CLSb

[8992] Anonymous. Clothes-line support. Scientific American, 94(3):76, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n3/pdf/scientificamerican01201906-76b.pdf.

Anonymous:1906:AWV

[8993] Anonymous. Adjustable window ventilator. Scientific American, 94(3):76, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
76c.pdf.

Anonymous:1906:BPWt

[8994] Anonymous. Business and personal wants. Scientific American, 94(3):78, January 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n3/pdf/scientificamerican01201906-
78.pdf.

Russell:1906:HF

[8995] Henry Norris Russell. The heavens in February. Scientific American, 94(4):87, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
87.pdf.

Anonymous:1906:Cp

[8996] Anonymous. Correspondence. Scientific American, 94(4):91, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n4/pdf/scientificamerican01271906-91.pdf.

Anonymous:1906:RPIu

[8997] Anonymous. Recently patented inventions. Scientific American, 94(4):97, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
97.pdf.

Anonymous:1906:NQs

[8998] Anonymous. Notes and queries. Scientific American, 94(4):97–98, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n4/pdf/scientificamerican01271906-97b.pdf.

Anonymous:1906:BNBo

[8999] Anonymous. Books: New books, etc. Scientific American, 94 (4):98–99, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
98.pdf.

Anonymous:1906:IIu

[9000] Anonymous. Index of inventions. Scientific American, 94(4):99– 103, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
99.pdf.

Anonymous:1906:PNPt

[9001] Anonymous. Patent notes: Patents. Scientific American, 94(4): 103, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
103.pdf.

Anonymous:1906:FCC

[9002] Anonymous. Final completion of the Croton Reservoir. Scientific American, 94(4):86, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
86.pdf.

Anonymous:1906:SUS

[9003] Anonymous. Shipbuilding in the United States. Scientific American, 94(4):86, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
86a.pdf.

Anonymous:1906:ATM

[9004] Anonymous. Army transport and the Merchant Marine. Scientific American, 94(4):86, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
86b.pdf.

Anonymous:1906:TCA

[9005] Anonymous. Two centuries after Franklin. Scientific American, 94(4):86–87, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
86c.pdf.

Anonymous:1906:INA

[9006] Anonymous. Improvements noted at the automobile shows. Scientific American, 94(4):87, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
87a.pdf.

Anonymous:1906:UAB

[9007] Anonymous. The ungentle art of burglary. Scientific American, 94(4):88–89, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
88.pdf.

Anonymous:1906:NIO

[9008] Anonymous. A new and interesting overhead traveling gear for expediting construction in shipyards. Scientific American, 94(4):90, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n4/pdf/scientificamerican01271906-90.pdf.

Anonymous:1906:SMG

[9009] Anonymous. Schlick’s marine gyroscope in use. Scientific American, 94(4):90, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
90a.pdf.

Collins:1906:HPD

[9010] A. Frederick Collins. High-potential discharges. Scientific American, 94(4):92–93, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
92.pdf.

Anonymous:1906:ACA

[9011] Anonymous. The Aero Club of America’s exhibit of aeronautical apparatus. Scientific American, 94(4):93–94, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n4/pdf/scientificamerican01271906-93.pdf.

Anonymous:1906:CIE

[9012] Anonymous. Changes and improvements in the Edison nickel-iron storage cell. Scientific American, 94(4):94, January 27, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n4/pdf/scientificamerican01271906-94.pdf.

Schneider:1906:ROS

[9013] Rudolf Schneider. The “Roman Ordnance” at the Saalburg. Scientific American, 94(4):95–96, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
95.pdf.

Anonymous:1906:ICB

[9014] Anonymous. Improved carrier for barrels. Scientific American, 94(4):96, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
96.pdf.

Anonymous:1906:NGA

[9015] Anonymous. A novel gun attachment. Scientific American, 94 (4):96, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
96a.pdf.

Anonymous:1906:CHH

[9016] Anonymous. Corn harvester and husker. Scientific American, 94(4):96, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
96b.pdf.

Anonymous:1906:OIc

[9017] Anonymous. Oddities in inventions. Scientific American, 94(4):96, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n4/pdf/scientificamerican01271906-96c.pdf.

Anonymous:1906:BPWu

[9018] Anonymous. Business and personal wants. Scientific American, 94(4):97, January 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n4/pdf/scientificamerican01271906-
97a.pdf.

Anonymous:1906:RPIv

[9019] Anonymous. Recently patented inventions. Scientific American, 94(5):116, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
116.pdf.

Anonymous:1906:NQt

[9020] Anonymous. Notes and queries. Scientific American, 94(5):117– 118, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
117.pdf.

Anonymous:1906:BNBp

[9021] Anonymous. Books: New books, etc. Scientific American, 94 (5):118, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
118.pdf.

Anonymous:1906:IIv

[9022] Anonymous. Index of inventions. Scientific American, 94(5): 118–123, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
118a.pdf.

Anonymous:1906:PNPu

[9023] Anonymous. Patent notes: Patents. Scientific American, 94(5): 123, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
123.pdf.

Anonymous:1906:CLL

[9024] Anonymous. A canal by lake and locks. Scientific American, 94(5):106, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
106.pdf.

Anonymous:1906:DWF

[9025] Anonymous. Does a water famine threaten New York City? Scientific American, 94(5):106, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
106a.pdf.

Anonymous:1906:NCAa

[9026] Anonymous. New conceptions in astronomy. Scientific American, 94(5):106–107, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
106b.pdf.

Anonymous:1906:VD

[9027] Anonymous. The voyage of the “Discovery”. Scientific American, 94(5):107, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
107.pdf.

Anonymous:1906:SNg

[9028] Anonymous. Science notes. Scientific American, 94(5):107, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n5/pdf/scientificamerican02031906-107a.pdf.

Gradenwitz:1906:NMD

[9029] Alfred Gradenwitz. Novel methods of determining the amount of atmospheric dust. Scientific American, 94(5):108, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n5/pdf/scientificamerican02031906-108.pdf.

Anonymous:1906:BPL

[9030] Anonymous. The Bates Plan for lock and lake canal at Panama — II. Scientific American, 94(5):109–110, February 3, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n5/pdf/scientificamerican02031906-109.pdf.

Collins:1906:BNM

[9031] A. Frederick Collins. Braun’s new method of directing wireless messages. Scientific American, 94(5):110–111, February 3, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n5/pdf/scientificamerican02031906-110.pdf.

Gerald:1906:SMI

[9032] William G. Fitz Gerald. The soldier–mountaineers of Italy. Scientific American, 94(5):111, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
111.pdf.

Anonymous:1906:GCP

[9033] Anonymous. Giant creatures of prehistoric times. Scientific American, 94(5):113–114, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
113.pdf.

Gradenwitz:1906:REM

[9034] Alfred Gradenwitz. Refrigeration with electric motive power. Scientific American, 94(5):114, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
114.pdf.

Anonymous:1906:NASb

[9035] Anonymous. New automobile speed records in Florida. Scientific American, 94(5):115, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
115.pdf.

Anonymous:1906:BPWv

[9036] Anonymous. Business and personal wants. Scientific American, 94(5):116, February 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n5/pdf/scientificamerican02031906-
116a.pdf.

Anonymous:1906:Cq

[9037] Anonymous. Correspondence. Scientific American, 94(6):131, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n6/pdf/scientificamerican02101906-131.pdf.

Anonymous:1906:RPIw

[9038] Anonymous. Recently patented inventions. Scientific American, 94 (6):136–137, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
136a.pdf.

Anonymous:1906:NQu

[9039] Anonymous. Notes and queries. Scientific American, 94(6):137, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n6/pdf/scientificamerican02101906-137a.pdf.

Anonymous:1906:IIw

[9040] Anonymous. Index of inventions. Scientific American, 94(6): 138–143, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
138.pdf.

Anonymous:1906:PNPv

[9041] Anonymous. Patent notes: Patents. Scientific American, 94 (6):143, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
143.pdf.

Anonymous:1906:SCN

[9042] Anonymous. State Commission on New York Water Supply. Scientific American, 94(6):126, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
126.pdf.

Anonymous:1906:DMB

[9043] Anonymous. The delay of the Manhattan Bridge. Scientific American, 94(6):126, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
126a.pdf.

Anonymous:1906:TS

[9044] Anonymous. On a tour of the shops. Scientific American, 94 (6):126–127, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
126b.pdf.

Adams:1906:PDL

[9045] Alton D. Adams. Proposed dam for lake erie. Scientific American, 94(6):127, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
127.pdf.

Anonymous:1906:PED

[9046] Anonymous. Prize for electric device. Scientific American, 94 (6):127, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
127a.pdf.

Anonymous:1906:OMS

[9047] Anonymous. Official Meteorological Summary, New York, N.Y., January, 1906. Scientific American, 94(6):127, February 10, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n6/pdf/scientificamerican02101906-127b.pdf.

Anonymous:1906:AMP

[9048] Anonymous. An American modification of the Parsons steam turbine. Scientific American, 94(6):128, February 10, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n6/pdf/scientificamerican02101906-128.pdf.

Anonymous:1906:MSN

[9049] Anonymous. The motor skate — a new thousand league boot. Scientific American, 94(6):129, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
129.pdf.

Ramakers:1906:HSD

[9050] L. Ramakers. The Helicoptère Santos-Dumont’s latest flying machine. Scientific American, 94(6):129–130, February 10, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n6/pdf/scientificamerican02101906-129a.pdf.

Dailey:1906:CHG

[9051] H. B. Dailey. A convenient hydrogen generator. Scientific American, 94(6):130, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
130.pdf.

Willey:1906:OUM

[9052] Day Allen Willey. Ore-unloading by machinery. Scientific American, 94(6):131–132, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
131a.pdf.

Anonymous:1906:FRO

[9053] Anonymous. Final races at the Ormond Automobile Meet. Scientific American, 94(6):133–134, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
133.pdf.

Gradenwitz:1906:MGL

[9054] Alfred Gradenwitz. Moving a German lighthouse. Scientific American, 94(6):134, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
134.pdf.

Holder:1906:RPA

[9055] Charles Frederick Holder. Remarkable phosphorescent animals. Scientific American, 94(6):135–136, February 10, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n6/pdf/scientificamerican02101906-135.pdf.

Anonymous:1906:IAb

[9056] Anonymous. An improved annunciator. Scientific American, 94 (6):136, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
136.pdf.

Anonymous:1906:BPWw

[9057] Anonymous. Business and personal wants. Scientific American, 94(6):137, February 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n6/pdf/scientificamerican02101906-
137.pdf.

Anonymous:1906:Cr

[9058] Anonymous. Correspondence. Scientific American, 94(7):151, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n7/pdf/scientificamerican02171906-151.pdf.

Anonymous:1906:RPIx

[9059] Anonymous. Recently patented inventions. Scientific American, 94(7):156, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
156b.pdf.

Anonymous:1906:NQv

[9060] Anonymous. Notes and queries. Scientific American, 94(7):158– 159, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
158a.pdf.

Anonymous:1906:BNBq

[9061] Anonymous. Books: New books, etc. Scientific American, 94 (7):159–160, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
159.pdf.

Anonymous:1906:IIx

[9062] Anonymous. Index of inventions. Scientific American, 94(7): 160–163, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
160.pdf.

Anonymous:1906:PNPw

[9063] Anonymous. Patent notes: Patents. Scientific American, 94 (7):163, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
163.pdf.

Anonymous:1906:SSRb

[9064] Anonymous. Speed and safety in railroad travel. Scientific American, 94(7):146, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
146.pdf.

Anonymous:1906:SBP

[9065] Anonymous. Superheat and boiler pressures. Scientific American, 94(7):146, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
146a.pdf.

Anonymous:1906:FPC

[9066] Anonymous. Foreshore protection with chain-cable groynes. Scientific American, 94(7):146, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
146b.pdf.

Anonymous:1906:ESC

[9067] Anonymous. Electrifying the St. Clair Tunnel. Scientific American, 94(7):146, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
146c.pdf.

Anonymous:1906:EFG

[9068] Anonymous. The electric fusion of glass. Scientific American, 94 (7):146–147, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
146d.pdf.

Anonymous:1906:SRD

[9069] Anonymous. Some recent discoveries in oceanography. Scientific American, 94(7):147, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
147.pdf.

Anonymous:1906:SNh

[9070] Anonymous. Science notes. Scientific American, 94(7):147, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n7/pdf/scientificamerican02171906-147a.pdf.

Collins:1906:MAR

[9071] A. Frederick Collins. Mercury arc rectifier for charging storage batteries. Scientific American, 94(7):148–149, February 17, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n7/pdf/scientificamerican02171906-148.pdf.

Ramakers:1906:TUS

[9072] L. Ramakers. The tunnel under the Seine River. Scientific American, 94(7):149–150, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
149.pdf.

Anonymous:1906:SM

[9073] Anonymous. The size of molecules. Scientific American, 94(7): 150, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
150.pdf.

Anonymous:1906:MVM

[9074] Anonymous. Making Valentines by the million. Scientific American, 94(7):152–153, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
152.pdf.

Anonymous:1906:TSE

[9075] Anonymous. The total solar eclipse of 1905. Scientific American, 94(7):153–154, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
153.pdf.

Anonymous:1906:HBW

[9076] Anonymous. The human body as a wireless telegraph transmitter and receiver. Scientific American, 94(7):154, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n7/pdf/scientificamerican02171906-154.pdf.

Anonymous:1906:NWD

[9077] Anonymous. A novelty in watch dials. Scientific American, 94 (7):155, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
155.pdf.

Anonymous:1906:CAT

[9078] Anonymous. A curious arched tree. Scientific American, 94 (7):155, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
155a.pdf.

Anonymous:1906:AC

[9079] Anonymous. The “auto-carte”. Scientific American, 94(7):155, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n7/pdf/scientificamerican02171906-155b.pdf.

Baker:1906:SDS

[9080] Thomas R. Baker. A simple dynamometer for school gymnasium testing. Scientific American, 94(7):155, February 17, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n7/pdf/scientificamerican02171906-155c.pdf.

Anonymous:1906:IRJ

[9081] Anonymous. An improved rail joint. Scientific American, 94 (7):156, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
156.pdf.

Anonymous:1906:WBS

[9082] Anonymous. Wheel-blocking safety device for automobiles. Scientific American, 94(7):156, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
156a.pdf.

Anonymous:1906:BPWx

[9083] Anonymous. Business and personal wants. Scientific American, 94(7):158, February 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n7/pdf/scientificamerican02171906-
158.pdf.

Anonymous:1906:Cs

[9084] Anonymous. Correspondence. Scientific American, 94(8):171, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n8/pdf/scientificamerican02241906-171.pdf.

Anonymous:1906:RPIy

[9085] Anonymous. Recently patented inventions. Scientific American, 94(8):176, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
176b.pdf.

Anonymous:1906:BNBr

[9086] Anonymous. Books: New books, etc. Scientific American, 94 (8):178, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
178a.pdf.

Anonymous:1906:NQw

[9087] Anonymous. Notes and queries. Scientific American, 94(8):178, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n8/pdf/scientificamerican02241906-178b.pdf.

Anonymous:1906:IIy

[9088] Anonymous. Index of inventions. Scientific American, 94(8): 178–183, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
178c.pdf.

Anonymous:1906:PNPx

[9089] Anonymous. Patent notes: Patents. Scientific American, 94 (8):183, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
183.pdf.

Anonymous:1906:SSB

[9090] Anonymous. The Ship Subsidy Bill. Scientific American, 94 (8):166, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
166.pdf.

Anonymous:1906:D

[9091] Anonymous. The “Dreadnought”. Scientific American, 94(8): 166, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
166a.pdf.

Anonymous:1906:EHR

[9092] Anonymous. Evolution of heat from radium. Scientific American, 94(8):166, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
166b.pdf.

DeVries:1906:EE

[9093] Hugo De Vries. Experimental evolution. Scientific American, 94 (8):167, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
167.pdf.

Anonymous:1906:SNi

[9094] Anonymous. Science notes. Scientific American, 94(8):167, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n8/pdf/scientificamerican02241906-167a.pdf.

Crafts:1906:RCP

[9095] H. A. Crafts. Reinforced concrete on the Pacific Coast. Scientific American, 94(8):168–169, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
168.pdf.

Anonymous:1906:GCR

[9096] Anonymous. A gravity cable railroad in the Swiss Alps. Scientific American, 94(8):169–170, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
169.pdf.

Anonymous:1906:DSM

[9097] Anonymous. Detonation of submarine mines by means of sound waves. Scientific American, 94(8):170, February 24, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n8/pdf/scientificamerican02241906-170.pdf.

Willey:1906:SWT

[9098] Day Allen Willey. Sand waves and their work. Scientific American, 94(8):172–174, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
172.pdf.

Luxton:1906:IA

[9099] G. Luxton. An ice automobile. Scientific American, 94(8):174, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n8/pdf/scientificamerican02241906-174.pdf.

Anonymous:1906:INM

[9100] Anonymous. The Inhalatorium, a new medical institution. Scientific American, 94(8):175, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
175.pdf.

Anonymous:1906:TWb

[9101] Anonymous. Training wall. Scientific American, 94(8):176, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n8/pdf/scientificamerican02241906-176.pdf.

Anonymous:1906:CCL

[9102] Anonymous. Combination chair and life preserver. Scientific American, 94(8):176, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
176a.pdf.

Anonymous:1906:BPWy

[9103] Anonymous. Business and personal wants. Scientific American, 94(8):178, February 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n8/pdf/scientificamerican02241906-
178.pdf.

Russell:1906:HMb

[9104] Henry Norris Russell. The heavens in March. Scientific American, 94(9):187, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
187.pdf.

Anonymous:1906:Ct

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-191.pdf.

Anonymous:1906:RPIz

[9106] Anonymous. Recently patented inventions. Scientific American, 94(9):198–199, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
198.pdf.

Anonymous:1906:NQx

[9107] Anonymous. Notes and queries. Scientific American, 94(9):199–200, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-199a.pdf.

Anonymous:1906:IIz

[9108] Anonymous. Index of inventions. Scientific American, 94(9):200–203, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-200.pdf.

Anonymous:1906:PNPy

[9109] Anonymous. Patent notes: Patents. Scientific American, 94(9):203, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-203.pdf.

Anonymous:1906:HLL

[9110] Anonymous. For a high-level lock canal. Scientific American, 94(9):186, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-186.pdf.

Anonymous:1906:TMT

[9111] Anonymous. The trench method of tunnel construction. Scientific American, 94(9):186, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
186a.pdf.

Anonymous:1906:DG

[9112] Anonymous. Distillation of gold. Scientific American, 94(9):186, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-186b.pdf.

Archdeacon:1906:SMH

[9113] Ernest Archdeacon. Sixty miles an hour on the water. Scientific American, 94(9):187, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
187a.pdf.

Meacham:1906:HBL

[9114] William M. Meacham. Hydroplane boats: Latest type of high-speed craft. Scientific American, 94(9):188–189, March 3, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n9/pdf/scientificamerican03031906-188.pdf.

Anonymous:1906:MPW

[9115] Anonymous. A motor paddle wheel for small boats. Scientific American, 94(9):189, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
189.pdf.

Anonymous:1906:MMD

[9116] Anonymous. The “Motogodille” a motor device for propelling small boats. Scientific American, 94(9):189–190, March 3, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n9/pdf/scientificamerican03031906-189a.pdf.

Anonymous:1906:SEL

[9117] Anonymous. Speed and energy loss of projectiles in water. Scientific American, 94(9):190, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
190.pdf.

Anonymous:1906:TSM

[9118] Anonymous. A triple-screw motor torpedo boat. Scientific American, 94(9):190, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
190a.pdf.

Anonymous:1906:LPB

[9119] Anonymous. Light power boats at the sportsmen’s show. Scientific American, 94(9):191, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
191a.pdf.

Anonymous:1906:TTC

[9120] Anonymous. A typical two-cycle marine motor. Scientific American, 94(9):192, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
192.pdf.

Anonymous:1906:ELR

[9121] Anonymous. Electric launch with recharging equipment. Scientific American, 94(9):192, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
192a.pdf.

Anonymous:1906:NSM

[9122] Anonymous. The new “Speedway” marine engine. Scientific American, 94(9):192, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
192b.pdf.

Anonymous:1906:SCMa

[9123] Anonymous. A sixteen-cylinder motor-boat engine of 180 horsepower. Scientific American, 94(9):193, March 3, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n9/pdf/scientificamerican03031906-193.pdf.

Fay:1906:STS

[9124] Thomas J. Fay and E. E. A serviceable twin-screw motor boat. Scientific American, 94(9):193–194, March 3, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v94/
n9/pdf/scientificamerican03031906-193a.pdf.

Anonymous:1906:ISC

[9125] Anonymous. Improvements in spark coil construction. Scientific American, 94(9):194, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
194.pdf.

Anonymous:1906:VEL

[9126] Anonymous. Vibrationless electric light sets for yachts. Scientific American, 94(9):194, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
194a.pdf.

Gradenwitz:1906:ASB

[9127] Alfred Gradenwitz. Artificial sea baths. Scientific American, 94(9):195, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-195.pdf.

Anonymous:1906:CZD

[9128] Anonymous. Count Von Zeppelin’s dirigible airship. Scientific American, 94(9):195–196, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
195a.pdf.

Guarini:1906:NT

[9129] Emile Guarini. A novel tent. Scientific American, 94(9):196, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-196.pdf.

Anonymous:1906:HSF

[9130] Anonymous. How snakes feed. Scientific American, 94(9):196, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v94/n9/pdf/scientificamerican03031906-196a.pdf.

Anonymous:1906:BPWz

[9131] Anonymous. Business and personal wants. Scientific American, 94(9):199, March 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v94/n9/pdf/scientificamerican03031906-
199.pdf.

Anonymous:1906:Cu

[9132] Anonymous. Correspondence. Scientific American, 95(1):9, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-9.pdf.

Anonymous:1906:RPIba

[9133] Anonymous. Recently patented inventions. Scientific American, 95(1): 14, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-14a.pdf.

Anonymous:1906:IIbb

[9134] Anonymous. Index of inventions. Scientific American, 95(1):14–19, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-14b.pdf.

Anonymous:1906:PNPz

[9135] Anonymous. Patent notes: Patents. Scientific American, 95(1):19, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-19.pdf.

Anonymous:1906:VAR

[9136] Anonymous. Volcanic activity and radium. Scientific American, 95(1): 4, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-4.pdf.

Anonymous:1906:CLO

[9137] Anonymous. The capture of a live okapi. Scientific American, 95(1):4, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-4a.pdf.

Anonymous:1906:NBW

[9138] Anonymous. A new blowpipe welding process. Scientific American, 95(1):4, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
4b.pdf.

Anonymous:1906:VW

[9139] Anonymous. The vagaries of wells. Scientific American, 95(1):4, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-4c.pdf.

Anonymous:1906:NIT

[9140] Anonymous. Nerve impulses and their propagation. Scientific American, 95(1):4, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
4d.pdf.

Anonymous:1906:STP

[9141] Anonymous. Statistics of turbine-propelled vessels, May 1, 1906. Scientific American, 95(1):5, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
5.pdf.

Anonymous:1906:CAR

[9142] Anonymous. A case of alleged reasoning in a dog. Scientific American, 95(1):5, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
5a.pdf.

Anonymous:1906:JET

[9143] Anonymous. Japanese earthquake theories. Scientific American, 95(1): 5, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-5b.pdf.

Anonymous:1906:HED

[9144] Anonymous. How England disposes of old ships. Scientific American, 95(1):5, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
5c.pdf.

Honey:1906:VH

[9145] Frederick R. Honey. The visible heavens. Scientific American, 95(1):5, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-5d.pdf.

Collins:1906:EMC

[9146] A. Frederick Collins. An electro-mechanical coin counting and wrapping machine. Scientific American, 95(1):6, July 7, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n1/pdf/scientificamerican07071906-6.pdf.

Holder:1906:TWA

[9147] Charles Frederick Holder. The tameness of wild animals. Scientific American, 95(1):6, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
6a.pdf.

Anonymous:1906:WAH

[9148] Anonymous. Wellman’s airship for his North Polar Expedition. Scientific American, 95(1):7–8, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
7.pdf.

Anonymous:1906:PLL

[9149] Anonymous. The Peterborough Lift-Lock of the Trent Valley Canal. Scientific American, 95(1):8, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
8.pdf.

Anonymous:1906:TLC

[9150] Anonymous. From the tobacco leaf to the cigar. Scientific American, 95(1):10–13, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n1/pdf/scientificamerican07071906-
10.pdf.

Anonymous:1906:MPa

[9151] Anonymous. A mechanical prizefighter. Scientific American, 95(1):13, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-13.pdf.

Anonymous:1906:ITa

[9152] Anonymous. An improved turbine. Scientific American, 95(1):13, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-13a.pdf.

Anonymous:1906:IRB

[9153] Anonymous. Improved rail bond. Scientific American, 95(1):13, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-13b.pdf.

Anonymous:1906:BPWba

[9154] Anonymous. Business and personal wants. Scientific American, 95(1): 14, July 7, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n1/pdf/scientificamerican07071906-14.pdf.

Anonymous:1906:Cv

[9155] Anonymous. Correspondence. Scientific American, 95(10):175, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n10/pdf/scientificamerican09081906-175.pdf.

Anonymous:1906:RPIbb

[9156] Anonymous. Recently patented inventions. Scientific American, 95(10):180, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
180a.pdf.

Anonymous:1906:IIbc

[9157] Anonymous. Index of inventions. Scientific American, 95(10): 180–182, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
180b.pdf.

Anonymous:1906:PNPba

[9158] Anonymous. Patent notes: Patents. Scientific American, 95 (10):183, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
183.pdf.

Anonymous:1906:LPM

[9159] Anonymous. Lack of progress on the Manhattan Bridge. Scientific American, 95(10):170, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
170.pdf.

Anonymous:1906:WRU

[9160] Anonymous. World’s record for United States Army gun. Scientific American, 95(10):170, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
170a.pdf.

Anonymous:1906:NCAb

[9161] Anonymous. The next challenger for “America’s” Cup. Scientific American, 95(10):170, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
170b.pdf.

Anonymous:1906:GGM

[9162] Anonymous. The greatest gem mine in the world. Scientific American, 95(10):170, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
170c.pdf.

Anonymous:1906:ACB

[9163] Anonymous. Adulteration of cereal breakfast foods. Scientific American, 95(10):170, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
170d.pdf.

Anonymous:1906:SRO

[9164] Anonymous. Searching for the real origin of species. Scientific American, 95(10):171, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
171.pdf.

Fawcett:1906:MAU

[9165] Waldon Fawcett. A motor ambulance for the United States Army. Scientific American, 95(10):172, September 8, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n10/pdf/scientificamerican09081906-172.pdf.

Winslow:1906:SCWa

[9166] Ben Winslow. Sanitation of the City of Washington. — II. Scientific American, 95(10):173–174, September 8, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n10/pdf/scientificamerican09081906-173.pdf.

Anonymous:1906:OBF

[9167] Anonymous. An office building 612 feet tall — the loftiest masonry structure in the world. Scientific American, 95(10):174, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n10/pdf/scientificamerican09081906-174.pdf.

Bastin:1906:MPR

[9168] S. Leonard Bastin. The marvels of plant retardation. Scientific American, 95(10):176–177, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
176.pdf.

Anonymous:1906:MAP

[9169] Anonymous. The Murray automatic page-printing telegraph — its history and its progress. Scientific American, 95(10):178, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n10/pdf/scientificamerican09081906-178.pdf.

Anonymous:1906:SSV

[9170] Anonymous. Safety stop valve. Scientific American, 95(10): 179, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
179.pdf.

Anonymous:1906:IAT

[9171] Anonymous. An improved animal trap. Scientific American, 95 (10):179, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
179a.pdf.

Anonymous:1906:PHG

[9172] Anonymous. A portable head gate. Scientific American, 95 (10):179, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
179b.pdf.

Anonymous:1906:WC

[9173] Anonymous. Woodworker’s clamp. Scientific American, 95(10): 179, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
179c.pdf.

Anonymous:1906:BPWbb

[9174] Anonymous. Business and personal wants. Scientific American, 95(10):180, September 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n10/pdf/scientificamerican09081906-
180.pdf.

Anonymous:1906:Cw

[9175] Anonymous. Correspondence. Scientific American, 95(11):191, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n11/pdf/scientificamerican09151906-191.pdf.

Anonymous:1906:RPIbc

[9176] Anonymous. Recently patented inventions. Scientific American, 95 (11):197, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
197.pdf.

Anonymous:1906:NQy

[9177] Anonymous. Notes and queries. Scientific American, 95(11): 198–199, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
198.pdf.

Anonymous:1906:IIbd

[9178] Anonymous. Index of inventions. Scientific American, 95(11): 199–202, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
199.pdf.

Anonymous:1906:PNPbb

[9179] Anonymous. Patent notes: Patents. Scientific American, 95 (11):203, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
203.pdf.

Anonymous:1906:STG

[9180] Anonymous. Stupendous traffic of Greater New York. Scientific American, 95(11):186, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
186.pdf.

Anonymous:1906:WEG

[9181] Anonymous. The worst enemy of the good roads movement. Scientific American, 95(11):186, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
186a.pdf.

Anonymous:1906:ELP

[9182] Anonymous. Erosion — the last problem for the gunmaker. Scientific American, 95(11):186, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
186b.pdf.

Gerard:1906:VP

[9183] W. R. Gerard. The “Virgina” potato. Scientific American, 95 (11):187, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
187.pdf.

Anonymous:1906:NA

[9184] Anonymous. A new airship. Scientific American, 95(11):187, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n11/pdf/scientificamerican09151906-187a.pdf.

Anonymous:1906:PPa

[9185] Anonymous. Photograph of the “Pennsylvania”. Scientific American, 95(11):187, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
187b.pdf.

Anonymous:1906:RTF

[9186] Anonymous. Road tarring in France. Scientific American, 95(11): 188–190, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
188.pdf.

Parker:1906:HWS

[9187] Thomas Parker. Hot-water supplies to towns. Scientific American, 95(11):190, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
190.pdf.

Willey:1906:GM

[9188] D. A. Willey. Gila monsters. Scientific American, 95(11):192, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n11/pdf/scientificamerican09151906-192.pdf.

Gradenwitz:1906:MBC

[9189] Alfred Gradenwitz. Modern bell casting. Scientific American, 95 (11):192–193, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
192a.pdf.

Anonymous:1906:GON

[9190] Anonymous. The growth of our Navy since the War. Scientific American, 95(11):194–195, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
194.pdf.

Anonymous:1906:SCMb

[9191] Anonymous. A simple canceling machine. Scientific American, 95 (11):196, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
196.pdf.

Anonymous:1906:SFH

[9192] Anonymous. Support for fire hose. Scientific American, 95(11): 196, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
196a.pdf.

Anonymous:1906:IPC

[9193] Anonymous. Implement for plowing or cultivating around trees. Scientific American, 95(11):196, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
196b.pdf.

Anonymous:1906:MAB

[9194] Anonymous. Means for automatically balancing rotating masses. Scientific American, 95(11):196, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
196c.pdf.

Anonymous:1906:BPWbc

[9195] Anonymous. Business and personal wants. Scientific American, 95 (11):197, September 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n11/pdf/scientificamerican09151906-
197a.pdf.

Anonymous:1906:Cx

[9196] Anonymous. Correspondence. Scientific American, 95(12):211, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n12/pdf/scientificamerican09221906-211.pdf.

Anonymous:1906:RPIbd

[9197] Anonymous. Recently patented inventions. Scientific American, 95 (12):218, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
218.pdf.

Anonymous:1906:NQz

[9198] Anonymous. Notes and queries. Scientific American, 95(12): 218–219, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
218b.pdf.

Anonymous:1906:BNBs

[9199] Anonymous. Books: New books, etc. Scientific American, 95 (12):219, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
219.pdf.

Anonymous:1906:IIbe

[9200] Anonymous. Index of inventions. Scientific American, 95(12): 219–223, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
219a.pdf.

Anonymous:1906:PNPbc

[9201] Anonymous. Patent notes: Patents. Scientific American, 95 (12):223, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
223.pdf.

Anonymous:1906:GAU

[9202] Anonymous. Growth of American universities. Scientific American, 95(12):206, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
206.pdf.

Anonymous:1906:GET

[9203] Anonymous. Gasoline engines and the torpedo boat. Scientific American, 95(12):206, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
206a.pdf.

Anonymous:1906:NTH

[9204] Anonymous. North tube of Hudson River Tunnel completed. Scientific American, 95(12):206, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
206b.pdf.

Anonymous:1906:SPCb

[9205] Anonymous. Steel passenger cars for trunk railroads. Scientific American, 95(12):206, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
206c.pdf.

Anonymous:1906:JD

[9206] Anonymous. James Dredge. Scientific American, 95(12):206– 207, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
206d.pdf.

Cleveland:1906:ACA

[9207] W. F. Cleveland. The advantages of crank axles for locomotives. Scientific American, 95(12):207, September 22, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n12/pdf/scientificamerican09221906-207.pdf.

Sprague:1906:INV

[9208] Elizabeth C. Sprague. Investigating the nutritive value of meat. Scientific American, 95(12):207, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
207a.pdf.

Anonymous:1906:ASPb

[9209] Anonymous. The alcohol smugglers of Paris. Scientific American, 95 (12):208–209, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
208.pdf.

Honey:1906:HPW

[9210] Frederick R. Honey. How a planet is weighed. Scientific American, 95(12):210, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
210.pdf.

Willey:1906:PSA

[9211] Day Allen Willey. The pioneers of the Swiss Army. Scientific American, 95(12):212, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
212.pdf.

Graves:1906:PR

[9212] W. D. Graves. A peculiar rupture. Scientific American, 95 (12):213, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
213.pdf.

Anonymous:1906:MMC

[9213] Anonymous. A mantel made of cigar boxes. Scientific American, 95(12):213, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
213a.pdf.

Miners:1906:CAG

[9214] W. F. Miners. A curious accident to a grain elevator. Scientific American, 95(12):213, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
213b.pdf.

Lewis:1906:MWC

[9215] Lawrence Lewis. The Minnequa Works of the Colorado Fuel and Iron Company. Scientific American, 95(12):214–216, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n12/pdf/scientificamerican09221906-214.pdf.

Anonymous:1906:LMP

[9216] Anonymous. The Lodge–Muirhead portable wireless telegraph plant for military purposes. Scientific American, 95(12):216, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n12/pdf/scientificamerican09221906-216.pdf.

Anonymous:1906:MPb

[9217] Anonymous. The Mathushek piano. Scientific American, 95 (12):217, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
217.pdf.

Anonymous:1906:BPWbd

[9218] Anonymous. Business and personal wants. Scientific American, 95 (12):218, September 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n12/pdf/scientificamerican09221906-
218a.pdf.

Russell:1906:HO

[9219] Henry Norris Russell. The heavens in October. Scientific American, 95(13):227, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
227.pdf.

Anonymous:1906:Cy

[9220] Anonymous. Correspondence. Scientific American, 95(13):231, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n13/pdf/scientificamerican09291906-231.pdf.

Anonymous:1906:RPIbe

[9221] Anonymous. Recently patented inventions. Scientific American, 95 (13):236, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
236a.pdf.

Anonymous:1906:IIbf

[9222] Anonymous. Index of inventions. Scientific American, 95(13): 236–239, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
236b.pdf.

Anonymous:1906:PNPbd

[9223] Anonymous. Patent notes: Patents. Scientific American, 95 (13):239, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
239.pdf.

Anonymous:1906:ATS

[9224] Anonymous. Agriculture the true source of our wealth. Scientific American, 95(13):226, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
226.pdf.

Anonymous:1906:SCG

[9225] Anonymous. A suggested cure of gun erosion. Scientific American, 95(13):226, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
226a.pdf.

Anonymous:1906:WTT

[9226] Anonymous. Wireless telegraphy and telephony in the Japanese Navy. Scientific American, 95(13):226, September 29, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n13/pdf/scientificamerican09291906-226b.pdf.

Anonymous:1906:EQB

[9227] Anonymous. The erection of the Quebec Bridge. Scientific American, 95(13):228, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
228.pdf.

Coar:1906:IOG

[9228] Charles H. Coar. An improved ozone generator. Scientific American, 95(13):229–230, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
229.pdf.

Gruppen:1906:HME

[9229] Carl L. Gruppen. A home-made equatorial mounting for telescopes. Scientific American, 95(13):230, September 29, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n13/pdf/scientificamerican09291906-230.pdf.

Anonymous:1906:MER

[9230] Anonymous. Machines in the Elimination Race for the Vanderbilt Cup. Scientific American, 95(13):232–234, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n13/pdf/scientificamerican09291906-232.pdf.

Anonymous:1906:SRR

[9231] Anonymous. Some remarkable results with armor-piercing shells. Scientific American, 95(13):234, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
234.pdf.

Anonymous:1906:IWR

[9232] Anonymous. An improved window raising and locking device. Scientific American, 95(13):235, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
235.pdf.

Anonymous:1906:ICH

[9233] Anonymous. An improved cutter head. Scientific American, 95 (13):235, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
235a.pdf.

Anonymous:1906:BPWbe

[9234] Anonymous. Business and personal wants. Scientific American, 95 (13):236, September 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n13/pdf/scientificamerican09291906-
236.pdf.

Anonymous:1906:Cz

[9235] Anonymous. Correspondence. Scientific American, 95(14):247, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n14/pdf/scientificamerican10061906-247.pdf.

Anonymous:1906:RPIbf

[9236] Anonymous. Recently patented inventions. Scientific American, 95(14):252, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
252a.pdf.

Anonymous:1906:IIbg

[9237] Anonymous. Index of inventions. Scientific American, 95(14):252–255, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n14/pdf/scientificamerican10061906-252b.pdf.

Anonymous:1906:PNPbe

[9238] Anonymous. Patent notes: Patents. Scientific American, 95(14):255, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n14/pdf/scientificamerican10061906-255.pdf.

Anonymous:1906:HPN

[9239] Anonymous. Huge proportions of the new Government irrigation dams. Scientific American, 95(14):242, October 6, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n14/pdf/scientificamerican10061906-242.pdf.

Anonymous:1906:SIR

[9240] Anonymous. Success of the independent railway motor car. Scientific American, 95(14):242, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
242a.pdf.

Anonymous:1906:CGa

[9241] Anonymous. The caliber of a gun. Scientific American, 95(14):242, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n14/pdf/scientificamerican10061906-242b.pdf.

Anonymous:1906:NPS

[9242] Anonymous. A new problem and its solution. Scientific American, 95(14):242, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
242c.pdf.

Anonymous:1906:NPC

[9243] Anonymous. A new process for the commercial utilization of atmospheric nitrogen. Scientific American, 95(14):242–243, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n14/pdf/scientificamerican10061906-242d.pdf.

Anonymous:1906:CDPb

[9244] Anonymous. Carbon dioxide pockets in France. Scientific American, 95(14):243, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
243.pdf.

Anonymous:1906:HSI

[9245] Anonymous. How to study industrial chemistry. Scientific American, 95(14):243, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
243a.pdf.

Anonymous:1906:ARO

[9246] Anonymous. Action of radium on organisms. Scientific American, 95(14):243, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
243b.pdf.

Anonymous:1906:NWT

[9247] Anonymous. A new wireless telephone. Scientific American, 95(14):243, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n14/pdf/scientificamerican10061906-243c.pdf.

Dunlap:1906:NHP

[9248] Orrin E. Dunlap. A new 30,000-horse-power plant at Niagara Falls. Scientific American, 95(14):244–245, October 6, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n14/pdf/scientificamerican10061906-244.pdf.

Anonymous:1906:AAP

[9249] Anonymous. Archdeacon’s air-propelled motor bicycle. Scientific American, 95(14):245–246, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
245.pdf.

Anonymous:1906:ADT

[9250] Anonymous. Automatic device for turning off the gas. Scientific American, 95(14):246, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
246.pdf.

Geare:1906:MCMa

[9251] Randolph I. Geare. Masks of classic and modern times, — I. Scientific American, 95(14):248–250, October 6, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n14/pdf/scientificamerican10061906-248.pdf.

Anonymous:1906:RCRb

[9252] Anonymous. The “R. C. Rickmers” — the largest sailing ship afloat. Scientific American, 95(14):250, October 6, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n14/pdf/scientificamerican10061906-250.pdf.

Alger:1906:BEP

[9253] M. Alger. The Birkeland–Eyde process and the artificial production of nitrates from the atmosphere. Scientific American, 95(14):250–251, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n14/pdf/scientificamerican10061906-250a.pdf.

Anonymous:1906:BPWbf

[9254] Anonymous. Business and personal wants. Scientific American, 95(14):252, October 6, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n14/pdf/scientificamerican10061906-
252.pdf.

Anonymous:1906:Cba

[9255] Anonymous. Correspondence. Scientific American, 95(15):263, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n15/pdf/scientificamerican10131906-263.pdf.

Anonymous:1906:RPIbg

[9256] Anonymous. Recently patented inventions. Scientific American, 95(15):269, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
269.pdf.

Anonymous:1906:NQba

[9257] Anonymous. Notes and queries. Scientific American, 95(15):270, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n15/pdf/scientificamerican10131906-270a.pdf.

Anonymous:1906:IIbh

[9258] Anonymous. Index of inventions. Scientific American, 95(15): 270–275, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
270b.pdf.

Anonymous:1906:PNPbf

[9259] Anonymous. Patent notes: Patents. Scientific American, 95 (15):275, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
275.pdf.

Anonymous:1906:FF

[9260] Anonymous. A forecast and its fulfillment. Scientific American, 95(15):258, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
258.pdf.

Anonymous:1906:NMR

[9261] Anonymous. New method of river tunneling. Scientific American, 95(15):258, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
258a.pdf.

Anonymous:1906:GPA

[9262] Anonymous. Growing prosperity in agriculture. Scientific American, 95(15):258, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
258b.pdf.

Anonymous:1906:WIC

[9263] Anonymous. What is the interior condition of the Earth? Scientific American, 95(15):258–259, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
258c.pdf.

Anonymous:1906:RFI

[9264] Anonymous. Results of the first international balloon race. Scientific American, 95(15):259, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
259.pdf.

Anonymous:1906:LGB

[9265] Anonymous. Levavasseur’s gliding boat. Scientific American, 95 (15):259, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
259a.pdf.

Anonymous:1906:SNj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n15/pdf/scientificamerican10131906-259b.pdf.

Geare:1906:MCMb

[9267] Randolph I. Geare. Masks of classic and modern times — II. Scientific American, 95(15):260–261, October 13, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n15/pdf/scientificamerican10131906-260.pdf.

Turner:1906:HM

[9268] Arthur B. Turner. The harvest Moon. Scientific American, 95 (15):262, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
262.pdf.

Yates:1906:NRI

[9269] L. H. Yates. The new route to Ireland. Scientific American, 95(15):262, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
262a.pdf.

Anonymous:1906:WFM

[9270] Anonymous. The winning foreign machines in the Third International Race for the Vanderbilt Cup. Scientific American, 95(15):263–265, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n15/pdf/scientificamerican10131906-263a.pdf.

Anonymous:1906:TC

[9271] Anonymous. The Teltow Canal. Scientific American, 95(15): 266–267, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
266.pdf.

Anonymous:1906:PNYb

[9272] Anonymous. Progress of the New York Central Terminal improvements. Scientific American, 95(15):267–268, October 13, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n15/pdf/scientificamerican10131906-267.pdf.

Anonymous:1906:BPWbg

[9273] Anonymous. Business and personal wants. Scientific American, 95(15):270, October 13, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n15/pdf/scientificamerican10131906-
270.pdf.

Anonymous:1906:SNk

[9274] Anonymous. Science notes. Scientific American, 95(16):279, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n16/pdf/scientificamerican10201906-279a.pdf.

Anonymous:1906:Cbb

[9275] Anonymous. Correspondence. Scientific American, 95(16):283, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n16/pdf/scientificamerican10201906-283.pdf.

Anonymous:1906:RPIbh

[9276] Anonymous. Recently patented inventions. Scientific American, 95 (16):289–290, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
289.pdf.

Anonymous:1906:NQbb

[9277] Anonymous. Notes and queries. Scientific American, 95(16):290, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n16/pdf/scientificamerican10201906-290a.pdf.

Anonymous:1906:BNBt

[9278] Anonymous. Books: New books, etc. Scientific American, 95 (16):290, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
290b.pdf.

Anonymous:1906:IIbi

[9279] Anonymous. Index of inventions. Scientific American, 95(16): 291–295, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
291.pdf.

Anonymous:1906:PNPbg

[9280] Anonymous. Patent notes: Patents. Scientific American, 95 (16):295, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
295.pdf.

Anonymous:1906:KB

[9281] Anonymous. A 211/2-knot battleship. Scientific American, 95 (16):278, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
278.pdf.

Anonymous:1906:CPD

[9282] Anonymous. Completion of the Pennsylvania Double Tunnels. Scientific American, 95(16):278, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
278a.pdf.

Anonymous:1906:PCB

[9283] Anonymous. Panama Canal to be built by contract. Scientific American, 95(16):278, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
278b.pdf.

Anonymous:1906:SLB

[9284] Anonymous. The status of the liquid barretter. Scientific American, 95(16):278–279, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
278c.pdf.

Anonymous:1906:SAC

[9285] Anonymous. Something about cereal breakfast foods. Scientific American, 95(16):279, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
279.pdf.

Erichsen:1906:NTP

[9286] Hugo Erichsen. The need and the testing of pure drugs. Scientific American, 95(16):280–281, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
280.pdf.

Anonymous:1906:RVC

[9287] Anonymous. A retrospect of the Vanderbilt Cup Race. Scientific American, 95(16):281–282, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
281.pdf.

Wheeler:1906:CPR

[9288] Kittredge Wheeler. The Canadian Pacific Railway Company’s irrigation project. Scientific American, 95(16):284–285, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n16/pdf/scientificamerican10201906-284.pdf.

Billman:1906:EPS

[9289] Obed C. Billman. An electric powderless, smokeless, flashless, and soundless gun. Scientific American, 95(16):286, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n16/pdf/scientificamerican10201906-286.pdf.

Holder:1906:PDF

[9290] Charles Frederick Holder. Photographing a devil fish — the chameleon of the sea. Scientific American, 95(16):286–288, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n16/pdf/scientificamerican10201906-286a.pdf.

Anonymous:1906:AMD

[9291] Anonymous. Apparatus for mixing different grades of rice. Scientific American, 95(16):288, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
288.pdf.

Anonymous:1906:NMP

[9292] Anonymous. A new method of propelling a ship. Scientific American, 95(16):288, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
288a.pdf.

Anonymous:1906:MFC

[9293] Anonymous. Means for facilitating the cleansing of street catch basins. Scientific American, 95(16):288, October 20, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n16/pdf/scientificamerican10201906-288b.pdf.

Anonymous:1906:SDP

[9294] Anonymous. Sound deflector for pianos. Scientific American, 95 (16):288, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
288c.pdf.

Anonymous:1906:BPWbh

[9295] Anonymous. Business and personal wants. Scientific American, 95(16):290, October 20, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n16/pdf/scientificamerican10201906-
290.pdf.

Russell:1906:HN

[9296] Henry Norris Russell. The heavens in November. Scientific American, 95(17):299, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
299.pdf.

Anonymous:1906:Cbc

[9297] Anonymous. Correspondence. Scientific American, 95(17):302–303, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n17/pdf/scientificamerican10271906-302.pdf.

Anonymous:1906:RPIbi

[9298] Anonymous. Recently patented inventions. Scientific American, 95(17):309, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
309.pdf.

Anonymous:1906:NQbc

[9299] Anonymous. Notes and queries. Scientific American, 95(17): 309–310, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
309b.pdf.

Anonymous:1906:BNBu

[9300] Anonymous. Books: New books, etc. Scientific American, 95 (17):310, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
310.pdf.

Anonymous:1906:IIbj

[9301] Anonymous. Index of inventions. Scientific American, 95(17): 310–315, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
310a.pdf.

Anonymous:1906:PNPbh

[9302] Anonymous. Patent notes: Patents. Scientific American, 95 (17):315, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
315.pdf.

Anonymous:1906:OVR

[9303] Anonymous. Our vast railroad system. Scientific American, 95 (17):298, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
298.pdf.

Anonymous:1906:BNY

[9304] Anonymous. A brilliant Navy yard success. Scientific American, 95(17):298, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
298a.pdf.

Anonymous:1906:SS

[9305] Anonymous. The shape of the Sun. Scientific American, 95(17):298, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n17/pdf/scientificamerican10271906-298b.pdf.

Inkersley:1906:RMM

[9306] Arthur Inkersley. Recovering metals melted in the San Francisco Fire. Scientific American, 95(17):299, October 27, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n17/pdf/scientificamerican10271906-299a.pdf.

Anonymous:1906:SRT

[9307] Anonymous. Sir Richard Tangye. Scientific American, 95(17):299, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n17/pdf/scientificamerican10271906-299b.pdf.

Anonymous:1906:FIB

[9308] Anonymous. The first international balloon race. Scientific American, 95(17):300–301, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
300.pdf.

Anonymous:1906:BDS

[9309] Anonymous. The battleships “Dreadnought” and “South Carolina”. Scientific American, 95(17):303–306, October 27, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n17/pdf/scientificamerican10271906-303.pdf.

Jenkins:1906:MRL

[9310] C. Francis Jenkins. A motor that runs by lightning. Scientific American, 95(17):306, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
306.pdf.

Anonymous:1906:TRF

[9311] Anonymous. Two recent fire-fighting inventions. Scientific American, 95(17):307–308, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
307.pdf.

Anonymous:1906:MCW

[9312] Anonymous. Motor car for whitewashing railroad tunnels. Scientific American, 95(17):308, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
308.pdf.

Anonymous:1906:HAY

[9313] Anonymous. A huge arched yucca tree. Scientific American, 95 (17):308, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
308a.pdf.

Perkins:1906:NEM

[9314] Frank C. Perkins. A novel electrical medical treatment. Scientific American, 95(17):308, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
308b.pdf.

Anonymous:1906:BPWbi

[9315] Anonymous. Business and personal wants. Scientific American, 95(17):309, October 27, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n17/pdf/scientificamerican10271906-
309a.pdf.

Anonymous:1906:RPIbj

[9316] Anonymous. Recently patented inventions. Scientific American, 95(18):329, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
329.pdf.

Anonymous:1906:NQbd

[9317] Anonymous. Notes and queries. Scientific American, 95(18): 329–330, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
329b.pdf.

Anonymous:1906:BNBv

[9318] Anonymous. Books: New books, etc. Scientific American, 95 (18):330, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
330.pdf.

Anonymous:1906:IIbk

[9319] Anonymous. Index of inventions. Scientific American, 95(18): 330–335, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
330a.pdf.

Anonymous:1906:PNPbi

[9320] Anonymous. Patent notes: Patents. Scientific American, 95 (18):335, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
335.pdf.

Anonymous:1906:LMT

[9321] Anonymous. The latest marine turbine success. Scientific American, 95(18):318, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
318.pdf.

Anonymous:1906:CBd

[9322] Anonymous. “Caliber” and “Bore”. Scientific American, 95(18): 318, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
318a.pdf.

Anonymous:1906:CIM

[9323] Anonymous. The Carnegie Institution Magnetic Survey. Scientific American, 95(18):318, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
318b.pdf.

Anonymous:1906:RA

[9324] Anonymous. Roads for automobiles. Scientific American, 95 (18):318, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
318c.pdf.

Anonymous:1906:SFS

[9325] Anonymous. Successful flight of Santos Dumont’s aeroplane. Scientific American, 95(18):318–319, November 3, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n18/pdf/scientificamerican11031906-318d.pdf.

Collins:1906:IWC

[9326] A. Frederick Collins. The international wireless conference. Scientific American, 95(18):319, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
319.pdf.

Anonymous:1906:ANb

[9327] Anonymous. Aeronautical notes. Scientific American, 95(18): 319, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
319a.pdf.

Anonymous:1906:CSa

[9328] Anonymous. The current supplement. Scientific American, 95 (18):319, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
319b.pdf.

Anonymous:1906:MIC

[9329] Anonymous. Manufacture of iron in China. Scientific American, 95(18):319, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
319c.pdf.

Anonymous:1906:KCL

[9330] Anonymous. The 251/4-knot Cunard liner “Mauretania”. Scientific American, 95(18):320, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
320.pdf.

Willey:1906:SMC

[9331] Day Allen Willey. Sugar making in Cuba. Scientific American, 95 (18):321–322, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
321.pdf.

Anonymous:1906:CSb

[9332] Anonymous. The creation of a star. Scientific American, 95(18): 323–324, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
323.pdf.

Beasley:1906:INMa

[9333] Walter L. Beasley. The Iona Naval Magazine — I. Scientific American, 95(18):325–326, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
325.pdf.

Anonymous:1906:PDe

[9334] Anonymous. Patent department. Scientific American, 95(18): 327, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
327.pdf.

Anonymous:1906:RSD

[9335] Anonymous. Railway signaling device. Scientific American, 95 (18):327, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
327a.pdf.

Anonymous:1906:ISS

[9336] Anonymous. An improved saw-set. Scientific American, 95(18): 327–328, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
327b.pdf.

Anonymous:1906:SCT

[9337] Anonymous. Street cleaner’s truck. Scientific American, 95(18): 328, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
328.pdf.

Anonymous:1906:NCL

[9338] Anonymous. A novel cigar lighter. Scientific American, 95(18): 328–329, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
328a.pdf.

Anonymous:1906:BPWbj

[9339] Anonymous. Business and personal wants. Scientific American, 95(18):329, November 3, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n18/pdf/scientificamerican11031906-
329a.pdf.

Anonymous:1906:RPIbk

[9340] Anonymous. Recently patented inventions. Scientific American, 95(19):350, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
350.pdf.

Anonymous:1906:NQbe

[9341] Anonymous. Notes and queries. Scientific American, 95(19): 351, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
351a.pdf.

Anonymous:1906:IIbl

[9342] Anonymous. Index of inventions. Scientific American, 95(19): 351–355, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
351b.pdf.

Anonymous:1906:PNPbj

[9343] Anonymous. Patent notes: Patents. Scientific American, 95 (19):355, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
355.pdf.

Anonymous:1906:ERW

[9344] Anonymous. The electric railroad wreck at Atlantic City. Scientific American, 95(19):338, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
338.pdf.

Anonymous:1906:SPE

[9345] Anonymous. Single-phase electric road between Baltimore and the National Capital. Scientific American, 95(19):338, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n19/pdf/scientificamerican11101906-338a.pdf.

Anonymous:1906:NRW

[9346] Anonymous. New rating of the world’s Fleets. Scientific American, 95(19):338–339, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
338b.pdf.

Anonymous:1906:FJ

[9347] Anonymous. A French “jungle”. Scientific American, 95(19): 339, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
339.pdf.

Anonymous:1906:TLB

[9348] Anonymous. Tulare Lake Basin again filling up. Scientific American, 95(19):339, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
339a.pdf.

Anonymous:1906:CSe

[9349] Anonymous. The current supplement. Scientific American, 95 (19):339, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
339b.pdf.

Anonymous:1906:NDS

[9350] Anonymous. The new dueling school in Paris. Scientific American, 95(19):340, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
340.pdf.

Anonymous:1906:CTR

[9351] Anonymous. Comparison of a turbine and a reciprocating engine for the United States Navy. Scientific American, 95(19):340, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n19/pdf/scientificamerican11101906-340a.pdf.

Gradenwitz:1906:PDA

[9352] Alfred Gradenwitz. The Parseval dirigible airship. Scientific American, 95(19):341–342, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
341.pdf.

Perrin:1906:SDF

[9353] William Henry Perrin. The story of the discovery of the first aniline dye. Scientific American, 95(19):342–343, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n19/pdf/scientificamerican11101906-342.pdf.

Beasley:1906:INMb

[9354] Walter L. Beasley. The Iona Naval Magazine — II. Scientific American, 95(19):343–344, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
343.pdf.

Shepstone:1906:CHN

[9355] Harold J. Shepstone. Carl Hagenbeck’s novel zoological park. Scientific American, 95(19):345–346, November 10, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n19/pdf/scientificamerican11101906-345.pdf.

Collins:1906:NSP

[9356] A. Frederick Collins. A new steel-plate printing and embossing machine. Scientific American, 95(19):347, November 10, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n19/pdf/scientificamerican11101906-347.pdf.

Marcosson:1906:AMB

[9357] Isaac F. Marcosson. An aid to modern business. Scientific American, 95(19):348–349, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
348.pdf.

Anonymous:1906:BPWbk

[9358] Anonymous. Business and personal wants. Scientific American, 95(19):351, November 10, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n19/pdf/scientificamerican11101906-
351.pdf.

Anonymous:1906:Cbe

[9359] Anonymous. Correspondence. Scientific American, 95(2):27, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-27.pdf.

Anonymous:1906:RPIbl

[9360] Anonymous. Recently patented inventions. Scientific American, 95(2):33–34, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
33.pdf.

Anonymous:1906:NQbf

[9361] Anonymous. Notes and queries. Scientific American, 95(2):34–36, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-34a.pdf.

Anonymous:1906:BNBw

[9362] Anonymous. Books: New books, etc. Scientific American, 95(2):36–39, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-36.pdf.

Anonymous:1906:PNPbk

[9363] Anonymous. Patent notes: Patents. Scientific American, 95(2):39–40, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-39.pdf.

Anonymous:1906:TFM

[9364] Anonymous. Tunneling by the freezing method under the East River. Scientific American, 95(2):22, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
22.pdf.

Anonymous:1906:GAP

[9365] Anonymous. The Government as an armor-plate manufacturer. Scientific American, 95(2):22, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
22a.pdf.

Anonymous:1906:CCRa

[9366] Anonymous. The Canadian Commission’s Report on the Electric Smelting of Iron Ores. Scientific American, 95(2):22, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-22b.pdf.

Anonymous:1906:NAV

[9367] Anonymous. A new Algol variable. Scientific American, 95(2):22–23, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-22c.pdf.

Bale:1906:SMA

[9368] William H. Bale. Summer Meeting of the American Association for the Advancement of Science. Scientific American, 95(2):23, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-23.pdf.

Anonymous:1906:SCA

[9369] Anonymous. The sense of color in animals. Scientific American, 95(2):23, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
23a.pdf.

Anonymous:1906:SNl

[9370] Anonymous. Science notes. Scientific American, 95(2):23, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-23b.pdf.

Claudy:1906:MSC

[9371] C. H. Claudy. A machine that smokes cigars. Scientific American, 95(2):24, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
24.pdf.

Anonymous:1906:GLN

[9372] Anonymous. The galleys of lake nemi. Scientific American, 95(2):25–26, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-25.pdf.

Gerald:1906:CS

[9373] William G. Fitz Gerald. A city of salt. Scientific American, 95(2):28, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-28.pdf.

Anonymous:1906:RCF

[9374] Anonymous. A railroad collision as a form of amusement. Scientific American, 95(2):30, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
30.pdf.

Anonymous:1906:MCV

[9375] Anonymous. The Meredith cigar-vending machine. Scientific American, 95(2):31, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
31.pdf.

Anonymous:1906:BMV

[9376] Anonymous. Brake mechanism for vehicles. Scientific American, 95(2):31, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
31a.pdf.

Anonymous:1906:ITSa

[9377] Anonymous. An improved track-sanding device. Scientific American, 95(2):31, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
31b.pdf.

Anonymous:1906:FC

[9378] Anonymous. Folding crib. Scientific American, 95(2):31, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-31c.pdf.

Anonymous:1906:IPR

[9379] Anonymous. An improved parallel ruler. Scientific American, 95(2):32, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-32.pdf.

Anonymous:1906:OId

[9380] Anonymous. Oddities in inventions. Scientific American, 95(2):32, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n2/pdf/scientificamerican07141906-32a.pdf.

Anonymous:1906:BPWbl

[9381] Anonymous. Business and personal wants. Scientific American, 95(2):34, July 14, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n2/pdf/scientificamerican07141906-
34.pdf.

Anonymous:1906:Cbf

[9382] Anonymous. Correspondence. Scientific American, 95(20):363, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n20/pdf/scientificamerican11171906-363.pdf.

Anonymous:1906:RPIbm

[9383] Anonymous. Recently patented inventions. Scientific American, 95(20):369, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
369.pdf.

Anonymous:1906:NQbg

[9384] Anonymous. Notes and queries. Scientific American, 95(20): 369–370, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
369b.pdf.

Anonymous:1906:BNBx

[9385] Anonymous. Books: New books, etc. Scientific American, 95 (20):370, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
370.pdf.

Anonymous:1906:IIbm

[9386] Anonymous. Index of inventions. Scientific American, 95(20): 370–374, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
370a.pdf.

Anonymous:1906:PNPbl

[9387] Anonymous. Patent notes: Patents. Scientific American, 95 (20):375, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
375.pdf.

Anonymous:1906:DP

[9388] Anonymous. Drawbridge protection. Scientific American, 95(20): 358, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
358.pdf.

Anonymous:1906:BMR

[9389] Anonymous. The Behr Mono-Rail in Brooklyn. Scientific American, 95(20):358, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
358a.pdf.

Anonymous:1906:EIU

[9390] Anonymous. Exports and imports under the new grouping. Scientific American, 95(20):358, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
358b.pdf.

Anonymous:1906:CCRb

[9391] Anonymous. Cape to Cairo by rail, river, and lake. Scientific American, 95(20):358–359, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
358c.pdf.

Anonymous:1906:PFN

[9392] Anonymous. Peary’s “Farthest North”. Scientific American, 95 (20):359, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
359.pdf.

Anonymous:1906:BF

[9393] Anonymous. Breakfast foods. Scientific American, 95(20):359, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n20/pdf/scientificamerican11171906-359a.pdf.

Anonymous:1906:CSf

[9394] Anonymous. The current supplement. Scientific American, 95 (20):359, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
359b.pdf.

Anonymous:1906:NWL

[9395] Anonymous. The number of words and letters in the bible. Scientific American, 95(20):359, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
359c.pdf.

Shepstone:1906:PBG

[9396] Harold J. Shepstone. A prehistoric burial ground. Scientific American, 95(20):360–361, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
360.pdf.

Anonymous:1906:CCRc

[9397] Anonymous. The Cape to Cairo Railroad. Scientific American, 95 (20):361–362, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
361.pdf.

Anonymous:1906:PAF

[9398] Anonymous. Progress of aeronautics in France. Scientific American, 95(20):363–364, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
363a.pdf.

Gerald:1906:TTW

[9399] W. G. Fitz Gerald. The turtle trade of the West Indies. Scientific American, 95(20):365–366, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
365.pdf.

Jones:1906:USE

[9400] George J. Jones. Unveiling statues by electricity. Scientific American, 95(20):366, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
366.pdf.

Anonymous:1906:NRS

[9401] Anonymous. A new remedy for seasickness. Scientific American, 95(20):367, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
367.pdf.

Anonymous:1906:GPR

[9402] Anonymous. A gasoline-propelled road roller. Scientific American, 95(20):367, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
367a.pdf.

Anonymous:1906:OSS

[9403] Anonymous. An ox-shoeing shop in the woods. Scientific American, 95(20):367, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
367b.pdf.

Anonymous:1906:NDH

[9404] Anonymous. Novel door hinge and check. Scientific American, 95 (20):368, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
368.pdf.

Anonymous:1906:IMM

[9405] Anonymous. An improved milking machine. Scientific American, 95(20):368, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
368a.pdf.

Anonymous:1906:BPWbm

[9406] Anonymous. Business and personal wants. Scientific American, 95(20):369, November 17, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n20/pdf/scientificamerican11171906-
369a.pdf.

Anonymous:1906:Cbg

[9407] Anonymous. Correspondence. Scientific American, 95(21):382, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n21/pdf/scientificamerican11241906-382a.pdf.

Anonymous:1906:RPIbn

[9408] Anonymous. Recently patented inventions. Scientific American, 95(21):389, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
389.pdf.

Anonymous:1906:NQbh

[9409] Anonymous. Notes and queries. Scientific American, 95(21): 390, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
390a.pdf.

Anonymous:1906:BNBy

[9410] Anonymous. Books: New books, etc. Scientific American, 95 (21):390, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
390b.pdf.

Anonymous:1906:IIbn

[9411] Anonymous. Index of inventions. Scientific American, 95(21): 390–395, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
390c.pdf.

Anonymous:1906:PNPbm

[9412] Anonymous. Patent notes: Patents. Scientific American, 95 (21):395, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
395.pdf.

Anonymous:1906:CPO

[9413] Anonymous. The congestion at the Patent Office. Scientific American, 95(21):378, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
378.pdf.

Anonymous:1906:ECS

[9414] Anonymous. Extraordinary conditions in the steel industry. Scientific American, 95(21):378, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
378a.pdf.

Anonymous:1906:GSS

[9415] Anonymous. Growth of the Salton Sea arrested. Scientific American, 95(21):378, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
378b.pdf.

Anonymous:1906:GTD

[9416] Anonymous. Gun trials of the “Dreadnought”. Scientific American, 95(21):378, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
378c.pdf.

Anonymous:1906:SOW

[9417] Anonymous. The size of ocean waves. Scientific American, 95 (21):378, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
378d.pdf.

Anonymous:1906:PPE

[9418] Anonymous. Progress of the Pennsylvania East River Tunnels. Scientific American, 95(21):378, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
378e.pdf.

Anonymous:1906:SDL

[9419] Anonymous. Santos Dumont’s latest flight. Scientific American, 95 (21):378–379, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
378f.pdf.

Gaut:1906:WDB

[9420] Helen Lukens Gaut. What do the birds eat? Scientific American, 95(21):379, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
379.pdf.

Anonymous:1906:SNm

[9421] Anonymous. Science notes. Scientific American, 95(21):379, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n21/pdf/scientificamerican11241906-379a.pdf.

Jones:1906:CF

[9422] George J. Jones. Cooking without fire. Scientific American, 95 (21):380, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
380.pdf.

Michaud:1906:CCE

[9423] Gustave Michaud. A curious chemical experiment. Scientific American, 95(21):380, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
380a.pdf.

Anonymous:1906:NACb

[9424] Anonymous. A new apparatus for the coaling of warships. Scientific American, 95(21):381, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
381.pdf.

Anonymous:1906:BRN

[9425] Anonymous. Blasting out a reef in New York Harbor. Scientific American, 95(21):382, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
382.pdf.

Honey:1906:OM

[9426] Frederic R. Honey. The opposition of Mars in 1907. Scientific American, 95(21):383, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
383.pdf.

Anonymous:1906:RAS

[9427] Anonymous. Railway accidents and surgery. Scientific American, 95 (21):383–385, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
383a.pdf.

Willey:1906:ERH

[9428] Day Allen Willey. Effects of the recent hurricane at Mobile. Scientific American, 95(21):385–386, November 24, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n21/pdf/scientificamerican11241906-385.pdf.

Gradenwitz:1906:SCS

[9429] Alfred Gradenwitz. Soft crystals showing apparent life. Scientific American, 95(21):387, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
387.pdf.

Anonymous:1906:UNR

[9430] Anonymous. A unique non-refillable bottle. Scientific American, 95(21):388, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
388.pdf.

Anonymous:1906:ITSb

[9431] Anonymous. An improved try-square. Scientific American, 95 (21):388, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
388a.pdf.

Anonymous:1906:SCL

[9432] Anonymous. Safety can lifter. Scientific American, 95(21):388, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n21/pdf/scientificamerican11241906-388b.pdf.

Anonymous:1906:BPWbn

[9433] Anonymous. Business and personal wants. Scientific American, 95(21):390, November 24, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n21/pdf/scientificamerican11241906-
390.pdf.

Russell:1906:HD

[9434] Henry Norris Russell. The heavens in December. Scientific American, 95(22):399, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
399.pdf.

Anonymous:1906:RPIbo

[9435] Anonymous. Recently patented inventions. Scientific American, 95(22):409, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
409.pdf.

Anonymous:1906:NQbi

[9436] Anonymous. Notes and queries. Scientific American, 95(22): 409–410, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
409b.pdf.

Anonymous:1906:BNBz

[9437] Anonymous. Books: New books, etc. Scientific American, 95 (22):410–411, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
410.pdf.

Anonymous:1906:IIbo

[9438] Anonymous. Index of inventions. Scientific American, 95(22): 411–415, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
411.pdf.

Anonymous:1906:PNPbn

[9439] Anonymous. Patent notes: Patents. Scientific American, 95 (22):415, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
415.pdf.

Anonymous:1906:PPb

[9440] Anonymous. The President at Panama. Scientific American, 95 (22):398, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
398.pdf.

Anonymous:1906:JLL

[9441] Anonymous. Japan launches a larger “Dreadnought”. Scientific American, 95(22):398, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
398a.pdf.

Anonymous:1906:RAE

[9442] Anonymous. Recent aeronautic experiments in France. Scientific American, 95(22):398, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
398b.pdf.

Anonymous:1906:EAPb

[9443] Anonymous. English aeroplane prizes. Scientific American, 95 (22):398, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
398c.pdf.

Anonymous:1906:PI

[9444] Anonymous. Peat in Italy. Scientific American, 95(22):399, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n22/pdf/scientificamerican12011906-399a.pdf.

Anonymous:1906:FMC

[9445] Anonymous. The first machine for the commercial production of window glass by the sheet process. Scientific American, 95(22):400, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n22/pdf/scientificamerican12011906-400.pdf.

Anonymous:1906:ICP

[9446] Anonymous. The interior construction of a piano player. Scientific American, 95(22):401–403, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
401.pdf.

Anonymous:1906:PCPb

[9447] Anonymous. Present conditions at Panama. Scientific American, 95 (22):403–405, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
403.pdf.

Anonymous:1906:BRS

[9448] Anonymous. Beautifying the roadbed by sodding. Scientific American, 95(22):406, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
406.pdf.

Anonymous:1906:AE

[9449] Anonymous. The ancestors of the elephant. Scientific American, 95 (22):406–407, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
406a.pdf.

Anonymous:1906:PDf

[9450] Anonymous. Patent department. Scientific American, 95(22):408, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n22/pdf/scientificamerican12011906-408.pdf.

Anonymous:1906:ITB

[9451] Anonymous. An improved tie-bar. Scientific American, 95(22):408, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n22/pdf/scientificamerican12011906-408a.pdf.

Anonymous:1906:PCD

[9452] Anonymous. Pollen-collecting device. Scientific American, 95 (22):408, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
408b.pdf.

Anonymous:1906:BPWbo

[9453] Anonymous. Business and personal wants. Scientific American, 95(22):409, December 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n22/pdf/scientificamerican12011906-
409a.pdf.

Anonymous:1906:Cbh

[9454] Anonymous. Correspondence. Scientific American, 95(23):426, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n23/pdf/scientificamerican12081906-426.pdf.

Anonymous:1906:RPIbp

[9455] Anonymous. Recently patented inventions. Scientific American, 95(23):433, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
433.pdf.

Anonymous:1906:NQbj

[9456] Anonymous. Notes and queries. Scientific American, 95(23):433, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n23/pdf/scientificamerican12081906-433b.pdf.

Anonymous:1906:IIbp

[9457] Anonymous. Index of inventions. Scientific American, 95(23): 434–439, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
434.pdf.

Anonymous:1906:PNPbo

[9458] Anonymous. Patent notes: Patents. Scientific American, 95 (23):439, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
439.pdf.

Anonymous:1906:BHR

[9459] Anonymous. Bridging the Hudson River. Scientific American, 95 (23):418, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
418.pdf.

Anonymous:1906:SPCc

[9460] Anonymous. Safeguards for the Panama Canal locks. Scientific American, 95(23):418, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
418a.pdf.

Anonymous:1906:AFA

[9461] Anonymous. Airships in the French Army. Scientific American, 95 (23):418–419, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
418b.pdf.

Hale:1906:NAS

[9462] William H. Hale. National Academy of Sciences. Scientific American, 95(23):419, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
419.pdf.

Christy:1906:PAN

[9463] H. R. Christy. Points about needles. Scientific American, 95 (23):419, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
419a.pdf.

Boyer:1906:PIF

[9464] Jacques Boyer. The pharmaceutical industry of France. Scientific American, 95(23):420–423, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
420.pdf.

Anonymous:1906:PDDa

[9465] Anonymous. The De Pluvy diving dress. Scientific American, 95(23):423, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
423.pdf.

Anonymous:1906:SNEa

[9466] Anonymous. Some novel electric labor-saving devices. Scientific American, 95(23):424, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
424.pdf.

Willey:1906:PKA

[9467] Day Allen Willey. A pocket-knife artist. Scientific American, 95 (23):425, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
425.pdf.

Anonymous:1906:OMM

[9468] Anonymous. Oceanographic Museum of Monaco. Scientific American, 95(23):427, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
427.pdf.

Field:1906:WWW

[9469] C. Field. Weapons weird and wonderful. Scientific American, 95 (23):428–430, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
428.pdf.

Gradenwitz:1906:MAE

[9470] Alfred Gradenwitz. The microphotoscope — an apparatus for examining minute maps. Scientific American, 95(23):430, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n23/pdf/scientificamerican12081906-430.pdf.

Bowdish:1906:BOI

[9471] B. S. Bowdish. Birds of an ocean island. Scientific American, 95 (23):431–432, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
431.pdf.

Anonymous:1906:BPWbp

[9472] Anonymous. Business and personal wants. Scientific American, 95(23):433, December 8, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n23/pdf/scientificamerican12081906-
433a.pdf.

Anonymous:1906:PHC

[9473] Anonymous. The progress of the hollow concrete block industry. Scientific American, 95(23):433, December 8, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n23/pdf/scientificamerican12081906-433c.pdf.

Anonymous:1906:Cbi

[9474] Anonymous. Correspondence. Scientific American, 95(24):446– 447, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
446a.pdf.

Anonymous:1906:RPIbq

[9475] Anonymous. Recently patented inventions. Scientific American, 95 (24):453–454, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
453.pdf.

Anonymous:1906:NQbk

[9476] Anonymous. Notes and queries. Scientific American, 95(24): 454, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
454a.pdf.

Anonymous:1906:IIbq

[9477] Anonymous. Index of inventions. Scientific American, 95(24): 454–458, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
454b.pdf.

Anonymous:1906:PNPbp

[9478] Anonymous. Patent notes: Patents. Scientific American, 95 (24):459, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
459.pdf.

Anonymous:1906:GFS

[9479] Anonymous. Genesis of the first successful aeroplane. Scientific American, 95(24):442, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
442.pdf.

Anonymous:1906:SAS

[9480] Anonymous. The seventh annual show of the automobile club of America. Scientific American, 95(24):442–443, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n24/pdf/scientificamerican12151906-442a.pdf.

Anonymous:1906:POb

[9481] Anonymous. The Patent Office in 1906. Scientific American, 95 (24):443, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
443.pdf.

Anonymous:1906:CSg

[9482] Anonymous. The current supplement. Scientific American, 95 (24):443, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
443a.pdf.

Scott:1906:RCL

[9483] Stuart Stevens Scott. Righting a capsized lighthouse caisson. Scientific American, 95(24):444, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
444.pdf.

Grimshaw:1906:AAS

[9484] Robert Grimshaw. Automatic apparatus for selling postage stamps, post cards, and newspapers. Scientific American, 95(24):445– 446, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
445.pdf.

Proctor:1906:TES

[9485] Mary Proctor. Total eclipse of the Sun in January, 1907. Scientific American, 95(24):446, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
446.pdf.

Anonymous:1906:MPK

[9486] Anonymous. The motive power of a 25 1/4 knots liner. Scientific American, 95(24):447, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
447.pdf.

Anonymous:1906:SAE

[9487] Anonymous. The Second Annual Exhibition of the Aero Club of America. Scientific American, 95(24):447–449, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n24/pdf/scientificamerican12151906-447a.pdf.

Collins:1906:PSS

[9488] A. Frederick Collins. The Poulsen selective system of wireless telegraphy. Scientific American, 95(24):450, December 15, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n24/pdf/scientificamerican12151906-450.pdf.

Copp:1906:DWT

[9489] G. G. Copp. A disease which threatens the American chestnut tree. Scientific American, 95(24):451, December 15, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n24/pdf/scientificamerican12151906-451.pdf.

Anonymous:1906:PDg

[9490] Anonymous. Patent department. Scientific American, 95(24): 452, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
452.pdf.

Anonymous:1906:GMW

[9491] Anonymous. Gage for marking window casements. Scientific American, 95(24):452, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
452a.pdf.

Anonymous:1906:CAPa

[9492] Anonymous. A collapsible ash-pan. Scientific American, 95(24): 452, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
452b.pdf.

Anonymous:1906:IHC

[9493] Anonymous. An improved hand camera. Scientific American, 95 (24):452–453, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
452c.pdf.

Anonymous:1906:BPWbq

[9494] Anonymous. Business and personal wants. Scientific American, 95(24):454, December 15, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n24/pdf/scientificamerican12151906-
454.pdf.

Anonymous:1906:Cbj

[9495] Anonymous. Correspondence. Scientific American, 95(25):467, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n25/pdf/scientificamerican12221906-467.pdf.

Anonymous:1906:RPIbr

[9496] Anonymous. Recently patented inventions. Scientific American, 95(25):473, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
473.pdf.

Anonymous:1906:NQbl

[9497] Anonymous. Notes and queries. Scientific American, 95(25): 473–474, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
473b.pdf.

Anonymous:1906:BNBba

[9498] Anonymous. Books: New books, etc. Scientific American, 95 (25):475, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
475.pdf.

Anonymous:1906:IIbr

[9499] Anonymous. Index of inventions. Scientific American, 95(25): 475–479, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
475a.pdf.

Anonymous:1906:PNPbq

[9500] Anonymous. Patent notes: Patents. Scientific American, 95(25): 476–479, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
476.pdf.

Anonymous:1906:POT

[9501] Anonymous. Plans of our 20,000-ton battleship. Scientific American, 95(25):462, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
462.pdf.

Anonymous:1906:SHC

[9502] Anonymous. Sandy Hook coast defenses. Scientific American, 95 (25):462–463, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
462a.pdf.

Bolling:1906:NRO

[9503] Randolph Bolling. Notes on recharging oxygen gas tanks. Scientific American, 95(25):463, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
463.pdf.

Anonymous:1906:LDO

[9504] Anonymous. Long-distance ocean race for motor boats. Scientific American, 95(25):463, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
463a.pdf.

Anonymous:1906:SNn

[9505] Anonymous. Science notes. Scientific American, 95(25):463, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n25/pdf/scientificamerican12221906-463b.pdf.

Anonymous:1906:ONG

[9506] Anonymous. Opening of the new Grand Central Temporary Station. Scientific American, 95(25):463, December 22, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n25/pdf/scientificamerican12221906-463c.pdf.

Albrecht:1906:GMR

[9507] Herrmann Albrecht. The German marine research boat “Planet”. Scientific American, 95(25):464–465, December 22, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n25/pdf/scientificamerican12221906-464.pdf.

Anonymous:1906:RCV

[9508] Anonymous. The rebuilding of the Campanile at Venice. Scientific American, 95(25):465–466, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
465.pdf.

Anonymous:1906:CCRd

[9509] Anonymous. Controlling the Colorado River and Salton Sea. Scientific American, 95(25):467–469, December 22, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n25/pdf/scientificamerican12221906-467a.pdf.

Anonymous:1906:RAAa

[9510] Anonymous. Recent airship and aeroplane experiments in Europe. Scientific American, 95(25):470–472, December 22, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n25/pdf/scientificamerican12221906-470.pdf.

Anonymous:1906:NHP

[9511] Anonymous. A new hand-propelled automobile for children. Scientific American, 95(25):472, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
472.pdf.

Anonymous:1906:NFR

[9512] Anonymous. A new form of railway rolling stock construction. Scientific American, 95(25):472, December 22, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n25/pdf/scientificamerican12221906-472a.pdf.

Anonymous:1906:LSE

[9513] Anonymous. Lubricator for steam engine cylinders. Scientific American, 95(25):472, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
472b.pdf.

Anonymous:1906:BPWbr

[9514] Anonymous. Business and personal wants. Scientific American, 95(25):473, December 22, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n25/pdf/scientificamerican12221906-
473a.pdf.

Anonymous:1906:Cbk

[9515] Anonymous. Correspondence. Scientific American, 95(26):487, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n26/pdf/scientificamerican12291906-487.pdf.

Anonymous:1906:RPIbs

[9516] Anonymous. Recently patented inventions. Scientific American, 95(26):492, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
492.pdf.

Anonymous:1906:NQbm

[9517] Anonymous. Notes and queries. Scientific American, 95(26): 493, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
493a.pdf.

Anonymous:1906:BNBbb

[9518] Anonymous. Books: New books, etc. Scientific American, 95 (26):493, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
493b.pdf.

Anonymous:1906:IIbs

[9519] Anonymous. Index of inventions. Scientific American, 95(26): 493–498, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
493c.pdf.

Anonymous:1906:PNPbr

[9520] Anonymous. Patent notes: Patents. Scientific American, 95 (26):499, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
499.pdf.

Anonymous:1906:PCN

[9521] Anonymous. Present condition of the New York Tunnels. Scientific American, 95(26):482, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
482.pdf.

Anonymous:1906:PMP

[9522] Anonymous. The President’s message on the Panama Canal. Scientific American, 95(26):482, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
482a.pdf.

Anonymous:1906:CSh

[9523] Anonymous. Catechism of a scientist. Scientific American, 95 (26):483, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
483.pdf.

Benjamin:1906:WHW

[9524] Marous Benjamin. William Henry Welch. Scientific American, 95 (26):483, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
483a.pdf.

Anonymous:1906:UNF

[9525] Anonymous. The “Unilens” a novel form of telescope. Scientific American, 95(26):484, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
484.pdf.

Boyer:1906:NLL

[9526] Jacques Boyer. The new Lusol lamp. Scientific American, 95(26): 484–485, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
484a.pdf.

Woodruff:1906:SED

[9527] Paul H. Woodruff. A simple experimental dynamo. Scientific American, 95(26):485–486, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
485.pdf.

Anonymous:1906:SNYb

[9528] Anonymous. The solution of New York’s transportation problem. Scientific American, 95(26):486, December 29, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n26/pdf/scientificamerican12291906-486.pdf.

Fawcett:1906:AUS

[9529] Waldon Fawcett. The apparatus of the United States Life-Saving Service. Scientific American, 95(26):488–489, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n26/pdf/scientificamerican12291906-488.pdf.

Anonymous:1906:IPT

[9530] Anonymous. Increasing the power of a talking machine by means of compressed air. Scientific American, 95(26):490, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n26/pdf/scientificamerican12291906-490.pdf.

Collins:1906:ESN

[9531] Percy Collins. Eye-spotting in nature. Scientific American, 95 (26):490–492, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
490a.pdf.

Anonymous:1906:BPWbs

[9532] Anonymous. Business and personal wants. Scientific American, 95(26):493, December 29, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n26/pdf/scientificamerican12291906-
493.pdf.

Anonymous:1906:Cbl

[9533] Anonymous. Correspondence. Scientific American, 95(3):47, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-47.pdf.

Anonymous:1906:RPIbt

[9534] Anonymous. Recently patented inventions. Scientific American, 95(3):53, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n3/pdf/scientificamerican07211906-
53.pdf.

Anonymous:1906:NQbn

[9535] Anonymous. Notes and queries. Scientific American, 95(3):53–55, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-53b.pdf.

Anonymous:1906:BNBbc

[9536] Anonymous. Books: New books, etc. Scientific American, 95(3):55–56, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-55.pdf.

Anonymous:1906:IIbt

[9537] Anonymous. Index of inventions. Scientific American, 95(3):56–59, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-56.pdf.

Anonymous:1906:PNPbs

[9538] Anonymous. Patent notes: Patents. Scientific American, 95(3):59, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-59.pdf.

Anonymous:1906:RML

[9539] Anonymous. Reclamation of marsh lands in the South. Scientific American, 95(3):42, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n3/pdf/scientificamerican07211906-
42.pdf.

Anonymous:1906:ART

[9540] Anonymous. Automobile racing and touring in France and America. Scientific American, 95(3):42–43, July 21, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n3/pdf/scientificamerican07211906-42a.pdf.

Anonymous:1906:EEE

[9541] Anonymous. An electric examination of European mineral water. Scientific American, 95(3):43, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n3/pdf/scientificamerican07211906-
43.pdf.

Anonymous:1906:KOL

[9542] Anonymous. A kerosene-oil lamp. Scientific American, 95(3):43, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-43a.pdf.

Anonymous:1906:IAH

[9543] Anonymous. Industrial alcohol: How it is made and how it is used. Scientific American, 95(3):43, July 21, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n3/pdf/scientificamerican07211906-43b.pdf.

Anonymous:1906:ASE

[9544] Anonymous. An advance in steam engine design. Scientific American, 95(3):44, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n3/pdf/scientificamerican07211906-
44.pdf.

Anonymous:1906:ILW

[9545] Anonymous. An improved light-weight ignition storage battery. Scientific American, 95(3):44, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n3/pdf/scientificamerican07211906-
44a.pdf.

Anonymous:1906:NAF

[9546] Anonymous. A new adjustable float-feed carbureter for use with gasoline or alcohol. Scientific American, 95(3):44, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n3/pdf/scientificamerican07211906-44b.pdf.

Willey:1906:EEC

[9547] Day Allen Willey. The enlargement of the Erie Canal. Scientific American, 95(3):45–46, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n3/pdf/scientificamerican07211906-
45.pdf.

VanBrussel:1906:NJB

[9548] J. B. Van Brussel. The new Japanese battleships “Katori” and “Kishima”. Scientific American, 95(3):47–48, July 21, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n3/pdf/scientificamerican07211906-47a.pdf.

Fawcett:1906:NSG

[9549] Waldon Fawcett. New sources of gold. Scientific American, 95(3):49–50, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-49.pdf.

Anonymous:1906:FTS

[9550] Anonymous. The fighting tribes of Southern Africa. Scientific American, 95(3):50–52, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n3/pdf/scientificamerican07211906-
50.pdf.

Anonymous:1906:IV

[9551] Anonymous. An improved valve. Scientific American, 95(3):52, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-52.pdf.

Anonymous:1906:ADR

[9552] Anonymous. An automatic depression recorder for the scientific control of ventilation in mines. Scientific American, 95(3):52, July 21, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n3/pdf/scientificamerican07211906-52a.pdf.

Anonymous:1906:BPWbt 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n3/pdf/scientificamerican07211906-
53a.pdf.

Russell:1906:HAb

[9554] Henry Norris Russell. The heavens in August. Scientific American, 95(4):63, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
63.pdf.

Anonymous:1906:Cbn

[9555] Anonymous. Correspondence. Scientific American, 95(4):67, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n4/pdf/scientificamerican07281906-67.pdf.

Anonymous:1906:RPIbu

[9556] Anonymous. Recently patented inventions. Scientific American, 95(4):72, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
72.pdf.

Anonymous:1906:NQbo

[9557] Anonymous. Notes and queries. Scientific American, 95(4):72, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n4/pdf/scientificamerican07281906-72b.pdf.

Anonymous:1906:IIbu

[9558] Anonymous. Index of inventions. Scientific American, 95(4):72–75, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n4/pdf/scientificamerican07281906-72c.pdf.

Anonymous:1906:PNPbt

[9559] Anonymous. Patent notes: Patents. Scientific American, 95(4):75, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n4/pdf/scientificamerican07281906-75.pdf.

Anonymous:1906:ECR

[9560] Anonymous. English cars in railroad wrecks. Scientific American, 95(4):62, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
62.pdf.

Anonymous:1906:EON

[9561] Anonymous. Early opening of New York Central Electric Service. Scientific American, 95(4):62, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
62a.pdf.

Anonymous:1906:FFS

[9562] Anonymous. The first foreign salute of the American flag. Scientific American, 95(4):62, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
62b.pdf.

Anonymous:1906:PPS

[9563] Anonymous. The production of precious stones in 1905. Scientific American, 95(4):62, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
62c.pdf.

Anonymous:1906:OSG

[9564] Anonymous. The original selden gasoline automobile and its bearing on the patent litigation. Scientific American, 95(4):63, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n4/pdf/scientificamerican07281906-63a.pdf.

Bates:1906:GCB

[9565] Lindon W. Bates, Jr. Galveston — a city built upon sand. Scientific American, 95(4):64, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
64.pdf.

Anonymous:1906:WCG

[9566] Anonymous. The winning car of the Grand Prix, and the new detachable rim. Scientific American, 95(4):65–66, July 28, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n4/pdf/scientificamerican07281906-65.pdf.

Shepstone:1906:OHF

[9567] H. J. Shepstone. A one-hundred-and-fifty-ton electric crane. Scientific American, 95(4):67, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
67a.pdf.

Mayner:1906:MMS

[9568] William Mayner. Model municipal slaughtering establishment at Berlin — a lesson in sanitary meat dressing. Scientific American, 95(4):68–69, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n4/pdf/scientificamerican07281906-68.pdf.

Baltimore:1906:MSC

[9569] J. Mayne Baltimore. A monster shark captured. Scientific American, 95(4):70, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
70.pdf.

Collins:1906:FAL

[9570] A. Frederick Collins. The flaming arc light. Scientific American, 95(4):70–71, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
70a.pdf.

Anonymous:1906:PSGb

[9571] Anonymous. Proposed system of guard rails for the subway. Scientific American, 95(4):71, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
71.pdf.

Anonymous:1906:BPWbu

[9572] Anonymous. Business and personal wants. Scientific American, 95(4):72, July 28, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n4/pdf/scientificamerican07281906-
72a.pdf.

Anonymous:1906:Cbo

[9573] Anonymous. Correspondence. Scientific American, 95(5):83, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n5/pdf/scientificamerican08041906-83.pdf.

Anonymous:1906:RPIbv

[9574] Anonymous. Recently patented inventions. Scientific American, 95(5):88, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
88a.pdf.

Anonymous:1906:IIbv

[9575] Anonymous. Index of inventions. Scientific American, 95(5):88–91, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n5/pdf/scientificamerican08041906-88b.pdf.

Anonymous:1906:PNPbu

[9576] Anonymous. Patent notes: Patents. Scientific American, 95(5):91, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n5/pdf/scientificamerican08041906-91.pdf.

Anonymous:1906:RBS

[9577] Anonymous. Railway block signal legislation. Scientific American, 95(5):78, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
78.pdf.

Anonymous:1906:RLF

[9578] Anonymous. The risk of locks in flight. Scientific American, 95(5):78, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n5/pdf/scientificamerican08041906-78a.pdf.

Anonymous:1906:TEC

[9579] Anonymous. The thickness of the Earth’s crust. Scientific American, 95(5):78, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
78b.pdf.

Anonymous:1906:BD

[9580] Anonymous. Botany up to date. Scientific American, 95(5):79, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n5/pdf/scientificamerican08041906-79.pdf.

Anonymous:1906:NSP

[9581] Anonymous. New sulphur process for the preservation of wood. Scientific American, 95(5):79, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
79a.pdf.

Anonymous:1906:BMW

[9582] Anonymous. Bolometric measurements in wireless telegraphy. Scientific American, 95(5):79, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
79b.pdf.

Willey:1906:SML

[9583] Day Allen Willey. The sulphur mines of Louisiana. Scientific American, 95(5):80–81, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
80.pdf.

Anonymous:1906:RCC

[9584] Anonymous. A railroad coach constructed in forty hours. Scientific American, 95(5):81, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
81.pdf.

Anonymous:1906:RCB

[9585] Anonymous. Recently completed battleship “Nebraska”. Scientific American, 95(5):82, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
82.pdf.

Anonymous:1906:OLB

[9586] Anonymous. Our latest battleships, “South Carolina” and “Michigan”. Scientific American, 95(5):82–83, August 4, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n5/pdf/scientificamerican08041906-82a.pdf.

Anonymous:1906:TER

[9587] Anonymous. Tunneling the East River. Scientific American, 95 (5):83–85, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
83a.pdf.

Fawcett:1906:ARF

[9588] Weldon Fawcett. An automobile for rural free delivery. Scientific American, 95(5):86, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
86.pdf.

Collins:1906:MEP

[9589] A. Frederick Collins. A mechanical electro-plating apparatus. Scientific American, 95(5):86–87, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
86a.pdf.

Anonymous:1906:BPWbv

[9590] Anonymous. Business and personal wants. Scientific American, 95(5):88, August 4, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n5/pdf/scientificamerican08041906-
88.pdf.

Anonymous:1906:RPIbw

[9591] Anonymous. Recently patented inventions. Scientific American, 95(6):104, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
104a.pdf.

Anonymous:1906:IIbw

[9592] Anonymous. Index of inventions. Scientific American, 95(6): 104–107, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
104b.pdf.

Anonymous:1906:Pb

[9593] Anonymous. Patents. Scientific American, 95(6):107, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n6/pdf/scientificamerican08111906-107.pdf.

Anonymous:1906:PGP

[9594] Anonymous. Pushing a good principle too far. Scientific American, 95(6):94, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
94.pdf.

Anonymous:1906:FWC

[9595] Anonymous. Fuel and water consumption on the “Lusitania”. Scientific American, 95(6):94, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
94a.pdf.

Anonymous:1906:CRR

[9596] Anonymous. Canals as rate regulators. Scientific American, 95 (6):94, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
94b.pdf.

Anonymous:1906:BFT

[9597] Anonymous. With the Baltic Fleet at Tsushima. Scientific American, 95(6):94–95, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
94c.pdf.

Anonymous:1906:PE

[9598] Anonymous. Poisonous eggs. Scientific American, 95(6):95, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n6/pdf/scientificamerican08111906-95.pdf.

Anonymous:1906:WGT

[9599] Anonymous. What the Glidden Tour proved. Scientific American, 95(6):95, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
95a.pdf.

Anonymous:1906:CWa

[9600] Anonymous. The color of water. Scientific American, 95(6):95, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n6/pdf/scientificamerican08111906-95b.pdf.

Collins:1906:NSM

[9601] A. Frederic Collins. A new sensitive micro-electroscope. Scientific American, 95(6):96, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
96.pdf.

Gradenwitz:1906:EHF

[9602] Alfred Gradenwitz. Electric heating fabrics. Scientific American, 95(6):96, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
96a.pdf.

Anonymous:1906:NCV

[9603] Anonymous. Novel commercial vehicles in the recent French test. Scientific American, 95(6):97–98, August 11, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n6/pdf/scientificamerican08111906-97.pdf.

Honey:1906:HEO

[9604] Frederic R. Honey. How eclipses occur. Scientific American, 95 (6):98, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
98.pdf.

Anonymous:1906:MS

[9605] Anonymous. The magic sphere. Scientific American, 95(6):99, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n6/pdf/scientificamerican08111906-99.pdf.

Byrn:1906:MPO

[9606] Edward W. Byrn. The Museum of Patent Office Models. Scientific American, 95(6):100–102, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
100.pdf.

Bowdish:1906:MTH

[9607] B. S. Bowdish. The modern taxidermist and his art. Scientific American, 95(6):102–103, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
102.pdf.

Anonymous:1906:BPWbw

[9608] Anonymous. Business and personal wants. Scientific American, 95(6):104, August 11, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n6/pdf/scientificamerican08111906-
104.pdf.

Anonymous:1906:Cbp

[9609] Anonymous. Correspondence. Scientific American, 95(7):115, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-115.pdf.

Anonymous:1906:RPIbx

[9610] Anonymous. Recently patented inventions. Scientific American, 95 (7):120–122, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
120.pdf.

Anonymous:1906:NQbp

[9611] Anonymous. Notes and queries. Scientific American, 95(7):122, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-122a.pdf.

Anonymous:1906:BNBbd

[9612] Anonymous. Books: New books, etc. Scientific American, 95 (7):122, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
122b.pdf.

Anonymous:1906:IIbx

[9613] Anonymous. Index of inventions. Scientific American, 95(7): 122–127, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
122c.pdf.

Anonymous:1906:PNPbv

[9614] Anonymous. Patent notes: Patents. Scientific American, 95(7):127, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-127.pdf.

Anonymous:1906:DE

[9615] Anonymous. Disappearing England. Scientific American, 95(7): 110, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
110.pdf.

Anonymous:1906:ASF

[9616] Anonymous. Aftermath of the San Francisco Fire. Scientific American, 95(7):110, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
110a.pdf.

Anonymous:1906:PSI

[9617] Anonymous. Prof. See’s investigation of the Earth’s rigidity. Scientific American, 95(7):110, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
110b.pdf.

Anonymous:1906:DSS

[9618] Anonymous. Discoveries in the Sargasso Sea. Scientific American, 95(7):110–111, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
110c.pdf.

Larkin:1906:EMG

[9619] Edgar L. Larkin. Electricity and matter in a gaseous state. Scientific American, 95(7):111, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
111.pdf.

Anonymous:1906:DM

[9620] Anonymous. The descent of man. Scientific American, 95(7):111, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-111a.pdf.

Anonymous:1906:DDB

[9621] Anonymous. The death of Daniel B. Wesson. Scientific American, 95(7):111, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
111b.pdf.

Anonymous:1906:RKC

[9622] Anonymous. The race for the King’s Cup. Scientific American, 95(7):112, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
112.pdf.

Smith:1906:ECE

[9623] W. N. Smith. The electric car equipment of the Long Island Railroad. Scientific American, 95(7):113–114, August 18, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n7/pdf/scientificamerican08181906-113.pdf.

Anonymous:1906:RAAb

[9624] Anonymous. Recent aeroplanes and airships in France. Scientific American, 95(7):116–117, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
116.pdf.

Anonymous:1906:PDh

[9625] Anonymous. Patent department. Scientific American, 95(7):118, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-118.pdf.

Anonymous:1906:ADO

[9626] Anonymous. Automatic door opening and closing device. Scientific American, 95(7):118, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
118a.pdf.

Anonymous:1906:CDF

[9627] Anonymous. Car door fastener. Scientific American, 95(7):118, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-118b.pdf.

Anonymous:1906:NRB

[9628] Anonymous. Non-refillable bottle. Scientific American, 95(7):118, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-118c.pdf.

Anonymous:1906:HAC

[9629] Anonymous. Hydraulic apparatus for cane mills. Scientific American, 95(7):118–119, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
118d.pdf.

Anonymous:1906:NTP

[9630] Anonymous. A novel tobacco pipe. Scientific American, 95(7):119, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-119.pdf.

Anonymous:1906:AA

[9631] Anonymous. Anchor for airships. Scientific American, 95(7):119, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-119a.pdf.

Anonymous:1906:ITc

[9632] Anonymous. An improved trolley. Scientific American, 95(7):119, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n7/pdf/scientificamerican08181906-119b.pdf.

Anonymous:1906:BPWbx

[9633] Anonymous. Business and personal wants. Scientific American, 95(7):122, August 18, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n7/pdf/scientificamerican08181906-
122.pdf.

Anonymous:1906:Cbq

[9634] Anonymous. Correspondence. Scientific American, 95(8):135, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n8/pdf/scientificamerican08251906-135.pdf.

Anonymous:1906:RPIby

[9635] Anonymous. Recently patented inventions. Scientific American, 95 (8):141–142, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
141.pdf.

Anonymous:1906:NQbq

[9636] Anonymous. Notes and queries. Scientific American, 95(8):143, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n8/pdf/scientificamerican08251906-143a.pdf.

Anonymous:1906:IIby

[9637] Anonymous. Index of inventions. Scientific American, 95(8): 143–147, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
143b.pdf.

Anonymous:1906:PNPbw

[9638] Anonymous. Patent notes: Patents. Scientific American, 95 (8):147–148, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
147.pdf.

Anonymous:1906:CCRe

[9639] Anonymous. Concrete and crushed rock. Scientific American, 95(8):130, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
130.pdf.

Anonymous:1906:EOR

[9640] Anonymous. Elevation of outer rail on curves. Scientific American, 95(8):130, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
130a.pdf.

Anonymous:1906:BMM

[9641] Anonymous. British Merchant Marine and the Commerce Destroyer. Scientific American, 95(8):130, August 25, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n8/pdf/scientificamerican08251906-130b.pdf.

Anonymous:1906:ESU

[9642] Anonymous. The effect of the sea upon climate. Scientific American, 95(8):130–131, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
130c.pdf.

Stirrup:1906:WBD

[9643] Charles Stirrup. Wanted: Brains to dissect. Scientific American, 95(8):131, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
131.pdf.

Anonymous:1906:SES

[9644] Anonymous. A severe earthquake in South America. Scientific American, 95(8):131, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
131a.pdf.

Anonymous:1906:NOM

[9645] Anonymous. The nickel ore mines of Canada. Scientific American, 95(8):131, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
131b.pdf.

Claudy:1906:NAC

[9646] C. H. Claudy. New asteroid camera at the Naval Observatory. Scientific American, 95(8):132–133, August 25, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n8/pdf/scientificamerican08251906-132.pdf.

Anonymous:1906:FTN

[9647] Anonymous. Finishing touches to the new Croton Reservoir. Scientific American, 95(8):133–134, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
133.pdf.

Takaki:1906:JMS

[9648] Baron K. Takaki. Japanese military sanitation and hygiene. Scientific American, 95(8):136–138, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
136.pdf.

Anonymous:1906:BBD

[9649] Anonymous. The British battleship “Dreadnought”. Scientific American, 95(8):138, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
138.pdf.

Houssay:1906:WWC

[9650] F R. Houssay. A wheel which carries its own track. Scientific American, 95(8):139, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
139.pdf.

Anonymous:1906:ASS

[9651] Anonymous. An automatic street sweeper and sprinkler. Scientific American, 95(8):139, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
139a.pdf.

Anonymous:1906:IVS

[9652] Anonymous. Improved vehicle spring. Scientific American, 95 (8):140, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
140.pdf.

Anonymous:1906:CAW

[9653] Anonymous. Car axle and wheel. Scientific American, 95(8):140, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n8/pdf/scientificamerican08251906-140a.pdf.

Anonymous:1906:CAPb

[9654] Anonymous. Check attachment for pulleys and chocks. Scientific American, 95(8):140, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
140b.pdf.

Anonymous:1906:OSM

[9655] Anonymous. Ore stamp mill. Scientific American, 95(8):140, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n8/pdf/scientificamerican08251906-140c.pdf.

Anonymous:1906:BPWby

[9656] Anonymous. Business and personal wants. Scientific American, 95(8):143, August 25, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n8/pdf/scientificamerican08251906-
143.pdf.

Russell:1906:HS

[9657] Henry Norris Russell. The heavens in September. Scientific American, 95(9):151, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
151.pdf.

Anonymous:1906:Cbr

[9658] Anonymous. Correspondence. Scientific American, 95(9):154, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n9/pdf/scientificamerican09011906-154.pdf.

Anonymous:1906:RPIbz

[9659] Anonymous. Recently patented inventions. Scientific American, 95(9):161, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
161.pdf.

Anonymous:1906:NQbr

[9660] Anonymous. Notes and queries. Scientific American, 95(9):161– 162, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
161b.pdf.

Anonymous:1906:BNBbe

[9661] Anonymous. Books: New books, etc. Scientific American, 95 (9):162, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
162.pdf.

Anonymous:1906:IIbz

[9662] Anonymous. Index of inventions. Scientific American, 95(9): 162–167, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
162a.pdf.

Anonymous:1906:PNPbx

[9663] Anonymous. Patent notes: Patents. Scientific American, 95 (9):167, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
167.pdf.

Anonymous:1906:ARE

[9664] Anonymous. Air resistance of electric cars. Scientific American, 95(9):150, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
150.pdf.

Anonymous:1906:GHR

[9665] Anonymous. Growth of the Hudson River steamboat. Scientific American, 95(9):150, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
150a.pdf.

Anonymous:1906:CBW

[9666] Anonymous. Cement bases for wooden telegraph poles. Scientific American, 95(9):150–151, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
150b.pdf.

Anonymous:1906:UCC

[9667] Anonymous. The use of clay cones in steel heating. Scientific American, 95(9):151, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
151a.pdf.

Anonymous:1906:GZA

[9668] Anonymous. “Galvanizing” with zinc-aluminium alloy. Scientific American, 95(9):151–152, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
151b.pdf.

Winslow:1906:SCWb

[9669] Ben Winslow. Sanitation of the City of Washington — I. Scientific American, 95(9):152–153, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
152.pdf.

Anonymous:1906:RAF

[9670] Anonymous. Review of the Atlantic Fleet by the President. Scientific American, 95(9):155–156, September 1, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n9/pdf/scientificamerican09011906-155.pdf.

Anonymous:1906:WDSb

[9671] Anonymous. Wreck of the Duluth–Superior Drawbridge. Scientific American, 95(9):156, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
156.pdf.

Fuller:1906:BTM

[9672] C. C. Fuller. Brick and tile making in the tropics. Scientific American, 95(9):157–158, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
157.pdf.

Anonymous:1906:SPS

[9673] Anonymous. The screw-propelled swimmer. Scientific American, 95(9):158, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
158.pdf.

Gradenwitz:1906:KAM

[9674] Alfred Gradenwitz. The Kromarograph — an automatic musicrecording apparatus. Scientific American, 95(9):159, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v95/n9/pdf/scientificamerican09011906-159.pdf.

Anonymous:1906:IPB

[9675] Anonymous. An improved puff box. Scientific American, 95 (9):159, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
159a.pdf.

Cleveland:1906:NTS

[9676] W. F. Cleveland. A new type of sixty-foot summit level canal for Panama. Scientific American, 95(9):160, September 1, 1906. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v95/
n9/pdf/scientificamerican09011906-160.pdf.

Anonymous:1906:BPWbz

[9677] Anonymous. Business and personal wants. Scientific American, 95(9):161, September 1, 1906. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v95/n9/pdf/scientificamerican09011906-
161a.pdf.

Russell:1907:HJa

[9678] Henry Norris Russell. The heavens in January. Scientific American, 96(1):9, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n1/pdf/scientificamerican01051907-
9.pdf.

Anonymous:1907:RPIa

[9679] Anonymous. Recently patented inventions. Scientific American, 96(1):14, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n1/pdf/scientificamerican01051907-
14.pdf.

Anonymous:1907:NQa

[9680] Anonymous. Notes and queries. Scientific American, 96(1):14–15, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n1/pdf/scientificamerican01051907-14a.pdf.

Anonymous:1907:IIa

[9681] Anonymous. Index of inventions. Scientific American, 96(1):15–19, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n1/pdf/scientificamerican01051907-15.pdf.

Anonymous:1907:PNPa

[9682] Anonymous. Patent notes: Patents. Scientific American, 96(1):19, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n1/pdf/scientificamerican01051907-19.pdf.

Anonymous:1907:RY

[9683] Anonymous. Review of the year 1906. Scientific American, 96(1):4–5, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n1/pdf/scientificamerican01051907-4.pdf.

Ramakers:1907:IPA

[9684] L. Ramakers. An interesting process for avoiding sponginess in heavy steel castings. Scientific American, 96(1):6, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n1/pdf/scientificamerican01051907-6.pdf.

Anonymous:1907:STB

[9685] Anonymous. Speed trials of the battleship “Vermont”. Scientific American, 96(1):7–8, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n1/pdf/scientificamerican01051907-
7.pdf.

Hawkes:1907:MSP

[9686] L. A. Hawkes. The manufacture of steel pens. Scientific American, 96(1):10–12, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n1/pdf/scientificamerican01051907-
10.pdf.

Anonymous:1907:CBN

[9687] Anonymous. A church blown up by natural gas. Scientific American, 96(1):12–13, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n1/pdf/scientificamerican01051907-
12.pdf.

Anonymous:1907:NDL

[9688] Anonymous. A novel door lock. Scientific American, 96(1):13, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n1/pdf/scientificamerican01051907-13.pdf.

Anonymous:1907:HOS

[9689] Anonymous. Hand-operated shaper. Scientific American, 96(1):13, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n1/pdf/scientificamerican01051907-13a.pdf.

Anonymous:1907:WFO

[9690] Anonymous. Wheel-flange oiler. Scientific American, 96(1):13, January 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n1/pdf/scientificamerican01051907-13b.pdf.

Anonymous:1907:Ca

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-211.pdf.

Anonymous:1907:RPIb

[9692] Anonymous. Recently patented inventions. Scientific American, 96(10):217, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
217.pdf.

Anonymous:1907:NQb

[9693] Anonymous. Notes and queries. Scientific American, 96(10):217–218, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-217a.pdf.

Anonymous:1907:IIb

[9694] Anonymous. Index of inventions. Scientific American, 96(10):218–223, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-218.pdf.

Anonymous:1907:PNPb

[9695] Anonymous. Patent notes: Patents. Scientific American, 96(10):223, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-223.pdf.

Anonymous:1907:AEW

[9696] Anonymous. Army Engineers will build Panama Canal. Scientific American, 96(10):206, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
206.pdf.

Anonymous:1907:SDCa

[9697] Anonymous. Side-door cars and rapid transit. Scientific American, 96(10):206, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
206a.pdf.

Anonymous:1907:GTSa

[9698] Anonymous. Government test of submarines. Scientific American, 96(10):206, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
206b.pdf.

Anonymous:1907:NYC

[9699] Anonymous. New York Central six-track subway. Scientific American, 96(10):206, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
206c.pdf.

Anonymous:1907:BHR

[9700] Anonymous. Bridging the Hudson River. Scientific American, 96(10):206, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
206d.pdf.

Anonymous:1907:ODM

[9701] Anonymous. Orson Desaix Munn. Scientific American, 96(10):207, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-207.pdf.

Anonymous:1907:KLS

[9702] Anonymous. Kites in life-saving operations. Scientific American, 96(10):207, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
207a.pdf.

Mayner:1907:HSB

[9703] William Mayner. A high-speed Bavarian locomotive. Scientific American, 96(10):208–209, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
208.pdf.

MCurdy:1907:PAS

[9704] James J. M’Curdy. The passing of American square-rigged vessels. Scientific American, 96(10):209–210, March 9, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n10/pdf/scientificamerican03091907-209.pdf.

Anonymous:1907:TAE

[9705] Anonymous. The telharmonium — an apparatus for the electrical generation and transmission of music. Scientific American, 96(10):210–211, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-210.pdf.

Anonymous:1907:OIa

[9706] Anonymous. The opium industry. Scientific American, 96(10):212–213, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-212.pdf.

Anonymous:1907:UL

[9707] Anonymous. The Uviol lamp. Scientific American, 96(10):214, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-214.pdf.

Anonymous:1907:MUA

[9708] Anonymous. Modern use of an ancient invention. Scientific American, 96(10):215, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
215.pdf.

Anonymous:1907:ICC

[9709] Anonymous. Improved cable-cutting device. Scientific American, 96(10):215, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
215a.pdf.

Anonymous:1907:ADP

[9710] Anonymous. Anchoring device for posts. Scientific American, 96(10):215, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n10/pdf/scientificamerican03091907-
215b.pdf.

Anonymous:1907:NST

[9711] Anonymous. A novel steam trap. Scientific American, 96(10):215–216, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-215c.pdf.

Anonymous:1907:OIb

[9712] Anonymous. Oddities in invention. Scientific American, 96(10):216, March 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n10/pdf/scientificamerican03091907-216.pdf.

Anonymous:1907:Cb

[9713] Anonymous. Correspondence. Scientific American, 96(11):231, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n11/pdf/scientificamerican03161907-231a.pdf.

Anonymous:1907:RPIc

[9714] Anonymous. Recently patented inventions. Scientific American, 96 (11):237–238, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
237.pdf.

Anonymous:1907:NQc

[9715] Anonymous. Notes and queries. Scientific American, 96(11):238–239, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n11/pdf/scientificamerican03161907-238.pdf.

Anonymous:1907:BNBa

[9716] Anonymous. Books: New books, etc. Scientific American, 96(11):239, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n11/pdf/scientificamerican03161907-239.pdf.

Anonymous:1907:IIc

[9717] Anonymous. Index of inventions. Scientific American, 96(11):239–243, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n11/pdf/scientificamerican03161907-239a.pdf.

Anonymous:1907:PNPc

[9718] Anonymous. Patent notes: Patents. Scientific American, 96(11):243, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n11/pdf/scientificamerican03161907-243.pdf.

Anonymous:1907:GIR

[9719] Anonymous. Government investigation of rail road accidents. Scientific American, 96(11):226, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
226.pdf.

Anonymous:1907:CEP

[9720] Anonymous. Comparative economy of producer gas and steam. Scientific American, 96(11):226, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
226a.pdf.

Anonymous:1907:PEC

[9721] Anonymous. Progress on the Erie Canal. Scientific American, 96(11):226, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
226b.pdf.

Anonymous:1907:CSTa

[9722] Anonymous. Completion of the second tube to Brooklyn. Scientific American, 96(11):226, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
226c.pdf.

Anonymous:1907:PSW

[9723] Anonymous. Preservation of steel work in tall buildings. Scientific American, 96(11):226–227, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
226d.pdf.

Willey:1907:GMO

[9724] Day Allen Willey. Glasgow and municipal ownership. Scientific American, 96(11):227, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
227.pdf.

Anonymous:1907:CSa

[9725] Anonymous. The current supplement. Scientific American, 96(11):227, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n11/pdf/scientificamerican03161907-227a.pdf.

Anonymous:1907:SNb

[9726] Anonymous. Science notes. Scientific American, 96(11):227, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n11/pdf/scientificamerican03161907-227b.pdf.

Boyer:1907:MAF

[9727] Jacques Boyer. Making artificial flowers. Scientific American, 96 (11):228–230, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
228.pdf.

Honey:1907:HSS

[9728] Frederic R. Honey. How to study a star map. Scientific American, 96(11):230–231, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
230.pdf.

Anonymous:1907:HAC

[9729] Anonymous. A hydraulic air-compressing plant. Scientific American, 96(11):231, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
231.pdf.

Anonymous:1907:WCR

[9730] Anonymous. Wooden cars in a railroad wreck. Scientific American, 96(11):232, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
232.pdf.

Jones:1907:TWR

[9731] H. D. Jones. Tablet writings from the Ruins of Nippur. Scientific American, 96(11):232–234, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
232a.pdf.

Collins:1907:CSD

[9732] A. Frederick Collins. Carborundum and silicon detectors for wireless telegraphy. Scientific American, 96(11):234, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n11/pdf/scientificamerican03161907-234.pdf.

Gradenwitz:1907:LAP

[9733] Alfred Gradenwitz. Leduc’s artificial plants and cells. Scientific American, 96(11):234–236, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
234a.pdf.

Michaud:1907:EA

[9734] Gustave Michaud. An experiment in acoustics. Scientific American, 96(11):236, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
236.pdf.

Anonymous:1907:SCCa

[9735] Anonymous. Special camera for copying and enlarging. Scientific American, 96(11):236–237, March 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n11/pdf/scientificamerican03161907-
236a.pdf.

Anonymous:1907:Cc

[9736] Anonymous. Correspondence. Scientific American, 96(12):251, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-251.pdf.

Anonymous:1907:RPId

[9737] Anonymous. Recently patented inventions. Scientific American, 96 (12):257–258, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
257.pdf.

Anonymous:1907:NQd

[9738] Anonymous. Notes and queries. Scientific American, 96(12):258, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-258.pdf.

Anonymous:1907:BNBb

[9739] Anonymous. Books: New books, etc. Scientific American, 96 (12):258–259, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
258a.pdf.

Anonymous:1907:IId

[9740] Anonymous. Index of inventions. Scientific American, 96(12):259–263, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-259.pdf.

Anonymous:1907:PNPd

[9741] Anonymous. Patent notes: Patents. Scientific American, 96(12):263, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-263.pdf.

Anonymous:1907:AIN

[9742] Anonymous. Architectural improvement of New York City. Scientific American, 96(12):246, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
246.pdf.

Anonymous:1907:DTQ

[9743] Anonymous. The demand for technically qualified young men. Scientific American, 96(12):246, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
246a.pdf.

Anonymous:1907:JDS

[9744] Anonymous. The “Jena” disaster and stable gunpowders. Scientific American, 96(12):246, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
246b.pdf.

Anonymous:1907:CFT

[9745] Anonymous. Cohesive fireproof tile construction. Scientific American, 96(12):246–247, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
246c.pdf.

Sidman:1907:IME

[9746] Charles A. Sidman. International Maritime Exposition Commemorating Anniversary of Steam Navigation in Bordeaux, France. Scientific American, 96(12):247, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
247.pdf.

Anonymous:1907:BZ

[9747] Anonymous. Burchell’s zebra. Scientific American, 96(12):247, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-247a.pdf.

Anonymous:1907:CSb

[9748] Anonymous. The current supplement. Scientific American, 96(12):247, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-247b.pdf.

Anonymous:1907:AAM

[9749] Anonymous. An alcohol-acetylene mixture for internal-combustion engines. Scientific American, 96(12):248, March 23, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n12/pdf/scientificamerican03231907-248.pdf.

Fitz-Gerald:1907:LDM

[9750] W. G. Fitz-Gerald. Latest designs of the motor in warfare. Scientific American, 96(12):249–250, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
249.pdf.

Anonymous:1907:HCA

[9751] Anonymous. How compressed air raised a sunken ship. the remarkable salving of the steamship “Bavarian”. Scientific American, 96(12):250, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-250.pdf.

Anonymous:1907:CTW

[9752] Anonymous. Controlling torpedoes by wireless telegraphy. Scientific American, 96(12):250, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
250a.pdf.

Gradenwitz:1907:IGF

[9753] Alfred Gradenwitz. An interesting German flying machine. Scientific American, 96(12):252–253, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
252.pdf.

Boyer:1907:SS

[9754] Jacques Boyer. The story of a silk hat. Scientific American, 96 (12):253–254, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
253.pdf.

Anonymous:1907:AWO

[9755] Anonymous. Autogenous welding with the oxy-acetylene highpressure blowpipe. Scientific American, 96(12):255, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-255.pdf.

Anonymous:1907:ACD

[9756] Anonymous. The aero club dinner. Scientific American, 96(12):256, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-256.pdf.

Anonymous:1907:STW

[9757] Anonymous. Steam trap with water meter attachment. Scientific American, 96(12):256, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n12/pdf/scientificamerican03231907-
256a.pdf.

Anonymous:1907:FCP

[9758] Anonymous. Filter for coffee pots. Scientific American, 96(12):256, March 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n12/pdf/scientificamerican03231907-256b.pdf.

Russell:1907:HAa

[9759] Henry Norris Russell. The heavens in April. Scientific American, 96(13):267, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
267.pdf.

Anonymous:1907:Cd

[9760] Anonymous. Correspondence. Scientific American, 96(13):271, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-271.pdf.

Anonymous:1907:RPIe

[9761] Anonymous. Recently patented inventions. Scientific American, 96 (13):277–278, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
277.pdf.

Anonymous:1907:NQe

[9762] Anonymous. Notes and queries. Scientific American, 96(13):278–279, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-278.pdf.

Anonymous:1907:IIe

[9763] Anonymous. Index of inventions. Scientific American, 96(13):279–283, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-279.pdf.

Anonymous:1907:PNPe

[9764] Anonymous. Patent notes: Patents. Scientific American, 96(13):283, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-283.pdf.

Anonymous:1907:CT

[9765] Anonymous. A city of towers. Scientific American, 96(13):266, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-266.pdf.

Anonymous:1907:NM

[9766] Anonymous. Naval marksmanship. Scientific American, 96(13):266, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-266a.pdf.

Anonymous:1907:ETV

[9767] Anonymous. Electric traction vindicated. Scientific American, 96(13):266, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
266b.pdf.

Anonymous:1907:MSC

[9768] Anonymous. Metallic sodium as a conductor for electric currents. Scientific American, 96(13):266, March 30, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n13/pdf/scientificamerican03301907-266c.pdf.

Anonymous:1907:DPB

[9769] Anonymous. The death of Pierre Berthelot. Scientific American, 96(13):267, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
267a.pdf.

Anonymous:1907:GMS

[9770] Anonymous. The gasoline motor in shallow-draft vessels. Scientific American, 96(13):268, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
268.pdf.

Anonymous:1907:NFAa

[9771] Anonymous. New French aeroplanes. Scientific American, 96 (13):269–270, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
269.pdf.

Michaud:1907:OI

[9772] Gustave Michaud. An optical illusion. Scientific American, 96(13):270, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-270.pdf.

Anonymous:1907:TCC

[9773] Anonymous. A Twentieth Century Campanile. Scientific American, 96(13):270, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
270a.pdf.

Brown:1907:ENE

[9774] Enos Brown. Excavations at Naga-Ed-Der, where prehistoric man first settled in Egypt. Scientific American, 96(13):272–273, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-272.pdf.

Anonymous:1907:NHS

[9775] Anonymous. A new high-speed photographic shutter. Scientific American, 96(13):274, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
274.pdf.

Anonymous:1907:DD

[9776] Anonymous. Dish drainer. Scientific American, 96(13):275, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-275.pdf.

Anonymous:1907:ICW

[9777] Anonymous. An improved concrete wall construction. Scientific American, 96(13):275, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
275a.pdf.

Anonymous:1907:NFH

[9778] Anonymous. A novel flag holder. Scientific American, 96(13):275, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-275b.pdf.

Anonymous:1907:NRS

[9779] Anonymous. A new rifle sight. Scientific American, 96(13):275, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-275c.pdf.

Anonymous:1907:CSc

[9780] Anonymous. Carpet stretcher. Scientific American, 96(13):275–276, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-275d.pdf.

Anonymous:1907:SCa

[9781] Anonymous. Safety cock. Scientific American, 96(13):276, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-276.pdf.

Anonymous:1907:EGL

[9782] Anonymous. Electric gas-lighter for acetylene automobile lamps. Scientific American, 96(13):276, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n13/pdf/scientificamerican03301907-
276a.pdf.

Anonymous:1907:IBBa

[9783] Anonymous. An improved bit brace. Scientific American, 96(13):276, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-276b.pdf.

Anonymous:1907:OIc

[9784] Anonymous. Oddities in inventions. Scientific American, 96(13):276, March 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n13/pdf/scientificamerican03301907-276c.pdf.

Anonymous:1907:RPIf

[9785] Anonymous. Recently patented inventions. Scientific American, 96(14):297–298, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
297.pdf.

Anonymous:1907:NQf

[9786] Anonymous. Notes and queries. Scientific American, 96(14):298, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n14/pdf/scientificamerican04061907-298.pdf.

Anonymous:1907:IIf

[9787] Anonymous. Index of inventions. Scientific American, 96(14):298–303, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n14/pdf/scientificamerican04061907-298a.pdf.

Anonymous:1907:PNPf

[9788] Anonymous. Patent notes: Patents. Scientific American, 96(14):303, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n14/pdf/scientificamerican04061907-303.pdf.

Anonymous:1907:SAG

[9789] Anonymous. The Scientific American gold medal for safety devices. Scientific American, 96(14):286, April 6, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n14/pdf/scientificamerican04061907-286.pdf.

Anonymous:1907:TP

[9790] Anonymous. Theory and practice. Scientific American, 96(14):286, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n14/pdf/scientificamerican04061907-286a.pdf.

Anonymous:1907:FDL

[9791] Anonymous. The four-day liner. Scientific American, 96(14):286, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n14/pdf/scientificamerican04061907-286b.pdf.

Anonymous:1907:TMD

[9792] Anonymous. Three months of denatured alcohol. Scientific American, 96(14):286–287, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
286c.pdf.

Anonymous:1907:ESMa

[9793] Anonymous. The evaporation of solid metals and their compounds. Scientific American, 96(14):287, April 6, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n14/pdf/scientificamerican04061907-287.pdf.

Anonymous:1907:CSd

[9794] Anonymous. The current supplement. Scientific American, 96(14):287, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n14/pdf/scientificamerican04061907-287a.pdf.

Anonymous:1907:GIJ

[9795] Anonymous. A great ice jam on the Susquehanna River. Scientific American, 96(14):288, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
288.pdf.

Anonymous:1907:TCP

[9796] Anonymous. A time-controlled phonograph. Scientific American, 96(14):289, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
289.pdf.

Anonymous:1907:CSCa

[9797] Anonymous. A combined sleeper and chair car. Scientific American, 96(14):289, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
289a.pdf.

Gage:1907:CMD

[9798] Edward G. Gage. The construction of a magnetic detector. Scientific American, 96(14):290–291, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
290.pdf.

Anonymous:1907:CVS

[9799] Anonymous. A continuous variable-speed system of rapid transit. Scientific American, 96(14):291–292, April 6, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n14/pdf/scientificamerican04061907-291.pdf.

Lotz:1907:HCB

[9800] F. P. Lotz. How church bells are cast. Scientific American, 96 (14):292–293, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
292.pdf.

Banks:1907:BVI

[9801] E. J. Banks. A Babylonian vase inlaid with ivory and precious stones dating from 4500 B.C. Scientific American, 96(14):294, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n14/pdf/scientificamerican04061907-294.pdf.

Collins:1907:GC

[9802] A. Frederick Collins. The gyroscope as a compass. Scientific American, 96(14):294–295, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
294a.pdf.

Ramakers:1907:UML

[9803] L. Ramakers. An unsinkable motor lifeboat. Scientific American, 96(14):295, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
295.pdf.

Smith:1907:LRR

[9804] Katherine Louise Smith. A lecture room on a railroad train. Scientific American, 96(14):295–296, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n14/pdf/scientificamerican04061907-
295a.pdf.

Gradenwitz:1907:WRR

[9805] Alfred Gradenwitz. A wire rope railway used in the construction of an Italian fortress. Scientific American, 96(14):296–297, April 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n14/pdf/scientificamerican04061907-296.pdf.

Anonymous:1907:Ce

[9806] Anonymous. Correspondence. Scientific American, 96(15):310–311, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n15/pdf/scientificamerican04131907-310.pdf.

Anonymous:1907:RPIg

[9807] Anonymous. Recently patented inventions. Scientific American, 96(15):317–318, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
317a.pdf.

Anonymous:1907:NQg

[9808] Anonymous. Notes and queries. Scientific American, 96(15):318–319, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n15/pdf/scientificamerican04131907-318.pdf.

Anonymous:1907:BNBc

[9809] Anonymous. Books: New books, etc. Scientific American, 96(15):319, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n15/pdf/scientificamerican04131907-319.pdf.

Anonymous:1907:IIg

[9810] Anonymous. Index of inventions. Scientific American, 96(15):319–323, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n15/pdf/scientificamerican04131907-319a.pdf.

Anonymous:1907:PNPg

[9811] Anonymous. Patent notes: Patents. Scientific American, 96(15):323, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n15/pdf/scientificamerican04131907-323.pdf.

Anonymous:1907:EAC

[9812] Anonymous. Electrolysis in armored concrete. Scientific American, 96(15):306, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
306.pdf.

Anonymous:1907:TCa

[9813] Anonymous. In touch with the conditions. Scientific American, 96(15):306, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
306a.pdf.

Anonymous:1907:NTG

[9814] Anonymous. A new theory of gun erosion. Scientific American, 96(15):306–307, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
306b.pdf.

Anonymous:1907:VNB

[9815] Anonymous. The value of the Nile Barrages to Egypt. Scientific American, 96(15):307, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
307.pdf.

Day:1907:SHI

[9816] Allen Day. Seal hunting on the ice floes of the North. Scientific American, 96(15):308, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
308.pdf.

Wilson:1907:NCV

[9817] H. L. Wilson. The North Carolina–Virginia Inland Waterway. Scientific American, 96(15):309–310, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
309.pdf.

Anonymous:1907:BRT

[9818] Anonymous. Branch railways by the Telpherage system. Scientific American, 96(15):311–312, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
311.pdf.

Willey:1907:CTS

[9819] Day Allen Willey. The Carnegie Technical Schools, Pittsburgh, Pa. Scientific American, 96(15):313–314, April 13, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n15/pdf/scientificamerican04131907-313.pdf.

Anonymous:1907:CFA

[9820] Anonymous. Capt. Ferber’s aeroplane experiments. Scientific American, 96(15):315, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
315.pdf.

Anonymous:1907:ABT

[9821] Anonymous. Adjustable boring tool. Scientific American, 96(15):316, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n15/pdf/scientificamerican04131907-316.pdf.

Anonymous:1907:LSR

[9822] Anonymous. Lubricator for street railway motors. Scientific American, 96(15):316, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
316a.pdf.

Anonymous:1907:SDCb

[9823] Anonymous. Screw driver with a centering device. Scientific American, 96(15):316, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
316b.pdf.

Anonymous:1907:ACR

[9824] Anonymous. Armored-concrete railway tie. Scientific American, 96(15):316, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
316c.pdf.

Anonymous:1907:SCb

[9825] Anonymous. Stanchion for cattle. Scientific American, 96(15):316–317, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n15/pdf/scientificamerican04131907-316d.pdf.

Anonymous:1907:RAS

[9826] Anonymous. Recording auto speed meter. Scientific American, 96(15):317, April 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n15/pdf/scientificamerican04131907-
317.pdf.

Anonymous:1907:Cf

[9827] Anonymous. Correspondence. Scientific American, 96(16):331, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n16/pdf/scientificamerican04201907-331.pdf.

Anonymous:1907:RPIh

[9828] Anonymous. Recently patented inventions. Scientific American, 96(16):337–338, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
337.pdf.

Anonymous:1907:NQh

[9829] Anonymous. Notes and queries. Scientific American, 96(16):338, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n16/pdf/scientificamerican04201907-338.pdf.

Anonymous:1907:BNBd

[9830] Anonymous. Books: New books, etc. Scientific American, 96(16):338, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n16/pdf/scientificamerican04201907-338a.pdf.

Anonymous:1907:IIh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n16/pdf/scientificamerican04201907-338b.pdf.

Anonymous:1907:PNPh

[9832] Anonymous. Patent notes: Patents. Scientific American, 96(16):343, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n16/pdf/scientificamerican04201907-343.pdf.

Anonymous:1907:SAA

[9833] Anonymous. The Scientific American aeronautic trophy. Scientific American, 96(16):326, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
326.pdf.

Anonymous:1907:BRRa

[9834] Anonymous. Broken rails and railroad accidents. Scientific American, 96(16):326, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
326a.pdf.

Anonymous:1907:CFb

[9835] Anonymous. The cruiser of the future. Scientific American, 96(16):326, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n16/pdf/scientificamerican04201907-326b.pdf.

Anonymous:1907:SEO

[9836] Anonymous. Sustained electric oscillations. Scientific American, 96(16):326–327, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
326c.pdf.

Dinger:1907:ATP

[9837] H. C. Dinger. Advantages of turbine propulsion for battleships. Scientific American, 96(16):327, April 20, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n16/pdf/scientificamerican04201907-327.pdf.

Fitz-Gerald:1907:HS

[9838] W. G. Fitz-Gerald. A harvest of the sea. Scientific American, 96(16):328–329, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
328.pdf.

Mclure:1907:MWC

[9839] W. Frank M’clure. The manufacture of worsted cloth. Scientific American, 96(16):329–330, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
329.pdf.

Sterling:1907:UCA

[9840] E. A. Sterling. The use of compressed air in logging. Scientific American, 96(16):330, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
330.pdf.

Anonymous:1907:HMU

[9841] Anonymous. The hippodrome mystery unveiled. Scientific American, 96(16):332–333, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
332.pdf.

Anonymous:1907:TML

[9842] Anonymous. A 250-ton mallet locomotive. Scientific American, 96(16):333, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
333.pdf.

Anonymous:1907:LLS

[9843] Anonymous. Loss of the largest ship ever wrecked. Scientific American, 96(16):334–335, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
334.pdf.

Anonymous:1907:ITCa

[9844] Anonymous. An improved twine cutter. Scientific American, 96 (16):336, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
336.pdf.

Anonymous:1907:ATR

[9845] Anonymous. Adjustable tension rods for pianos. Scientific American, 96(16):336, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
336a.pdf.

Anonymous:1907:NFN

[9846] Anonymous. A novel form of nail. Scientific American, 96(16):336, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n16/pdf/scientificamerican04201907-336b.pdf.

Anonymous:1907:PSTa

[9847] Anonymous. Pipe stopper or test plug. Scientific American, 96 (16):336–337, April 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n16/pdf/scientificamerican04201907-
336c.pdf.

Anonymous:1907:RPIi

[9848] Anonymous. Recently patented inventions. Scientific American, 96(17):357, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
357.pdf.

Anonymous:1907:NQi

[9849] Anonymous. Notes and queries. Scientific American, 96(17):357–358, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-357a.pdf.

Anonymous:1907:BNBe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-358.pdf.

Anonymous:1907:IIi

[9851] Anonymous. Index of inventions. Scientific American, 96(17):358–363, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-358a.pdf.

Anonymous:1907:PNPi

[9852] Anonymous. Patent notes: Patents. Scientific American, 96(17):363, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-363.pdf.

Anonymous:1907:PCa

[9853] Anonymous. The Peace Congress. Scientific American, 96(17):346, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-346.pdf.

Anonymous:1907:SPJ

[9854] Anonymous. Scope and purpose of the Jamestown Exposition. Scientific American, 96(17):346, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
346a.pdf.

Anonymous:1907:RADa

[9855] Anonymous. Raising the Assouan Dam. Scientific American, 96 (17):346, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
346b.pdf.

Anonymous:1907:OPR

[9856] Anonymous. An oppressive postal regulation. Scientific American, 96(17):346–347, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
346c.pdf.

Anonymous:1907:RGCa

[9857] Anonymous. Rules governing the competition for the Scientific American Flying Machine Trophy. Scientific American, 96(17):347, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-347.pdf.

Anonymous:1907:SSAa

[9858] Anonymous. Saving a ship by amputation. Scientific American, 96(17):348, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
348.pdf.

Perry:1907:NHA

[9859] Harry W. Perry. The new home of the automobile club of America. Scientific American, 96(17):349–350, April 27, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n17/pdf/scientificamerican04271907-349.pdf.

Anonymous:1907:IFJ

[9860] Anonymous. The International Fleet at Jamestown. Scientific American, 96(17):350–351, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
350.pdf.

Maginnis:1907:OJE

[9861] John T. Maginnis. Opening of the Jamestown Exposition. Scientific American, 96(17):351–353, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
351.pdf.

Plummer:1907:AS

[9862] John Plummer. Australian sapphires. Scientific American, 96(17):354, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-354.pdf.

Gradenwitz:1907:GC

[9863] Alfred Gradenwitz. A gigantic clock. Scientific American, 96(17):355, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-355.pdf.

Anonymous:1907:IPP

[9864] Anonymous. Improved packing for pistons. Scientific American, 96(17):356, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
356.pdf.

Anonymous:1907:WB

[9865] Anonymous. Wagon brake. Scientific American, 96(17):356, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-356a.pdf.

Anonymous:1907:SBF

[9866] Anonymous. Seal box fastener. Scientific American, 96(17):356, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n17/pdf/scientificamerican04271907-356b.pdf.

Anonymous:1907:AAS

[9867] Anonymous. Alarm attachment for self-binders. Scientific American, 96(17):356, April 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n17/pdf/scientificamerican04271907-
356c.pdf.

Russell:1907:HMa

[9868] Henry Norris Russell. The heavens in May. Scientific American, 96(18):367, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
367.pdf.

Anonymous:1907:Cg

[9869] Anonymous. Correspondence. Scientific American, 96(18):371, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n18/pdf/scientificamerican05041907-371.pdf.

Anonymous:1907:RPIj

[9870] Anonymous. Recently patented inventions. Scientific American, 96(18):377–378, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
377.pdf.

Anonymous:1907:NQN

[9871] Anonymous. Notes and queries: Note and queries. Scientific American, 96(18):378, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
378.pdf.

Anonymous:1907:IIj

[9872] Anonymous. Index of inventions. Scientific American, 96(18):378–383, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n18/pdf/scientificamerican05041907-378a.pdf.

Anonymous:1907:PNPj

[9873] Anonymous. Patent notes: Patents. Scientific American, 96(18):383, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n18/pdf/scientificamerican05041907-383.pdf.

Anonymous:1907:ARG

[9874] Anonymous. Accelerating rate of growth of our commerce. Scientific American, 96(18):366, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
366.pdf.

Anonymous:1907:CSNa

[9875] Anonymous. The commercial supremacy of New York. Scientific American, 96(18):366, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
366a.pdf.

Anonymous:1907:HCP

[9876] Anonymous. Health conditions at Panama. Scientific American, 96(18):366, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
366b.pdf.

Anonymous:1907:FES

[9877] Anonymous. Full electric service on the New York Central. Scientific American, 96(18):366, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
366c.pdf.

Anonymous:1907:PIW

[9878] Anonymous. The President’s Inland Waterways Commission. Scientific American, 96(18):366, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
366d.pdf.

Anonymous:1907:NBN

[9879] Anonymous. No bids for the new subway. Scientific American, 96(18):366, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
366e.pdf.

Anonymous:1907:QGI

[9880] Anonymous. Quarrying glacier ice. Scientific American, 96(18):366, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n18/pdf/scientificamerican05041907-366f.pdf.

Anonymous:1907:DPP

[9881] Anonymous. Developing photographic plates in daylight. Scientific American, 96(18):367, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
367a.pdf.

Beasley:1907:NFW

[9882] Walter L. Beasley. Naosaurus: A fossil wonder. Scientific American, 96(18):368, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
368.pdf.

Collins:1907:MWT

[9883] A. Frederick Collins. A military wireless telegraph equipment. Scientific American, 96(18):369–370, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
369.pdf.

Anonymous:1907:FFS

[9884] Anonymous. The felling of factory stacks. Scientific American, 96(18):372, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
372.pdf.

Tsunetaro:1907:JSJ

[9885] Saito Tsunetaro. The Japanese squadron at Jamestown. Scientific American, 96(18):373, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
373.pdf.

Washington:1907:CRC

[9886] W. D. H. Washington. The Colorado River closure. Scientific American, 96(18):374–377, May 4, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n18/pdf/scientificamerican05041907-
374.pdf.

Anonymous:1907:Ch

[9887] Anonymous. Correspondence. Scientific American, 96(19):391, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-391.pdf.

Anonymous:1907:RPIk

[9888] Anonymous. Recently patented inventions. Scientific American, 96(19):397–398, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
397.pdf.

Anonymous:1907:NQj

[9889] Anonymous. Notes and queries. Scientific American, 96(19):398, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-398.pdf.

Anonymous:1907:BNBf

[9890] Anonymous. Books: New books, etc. Scientific American, 96 (19):398–399, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
398a.pdf.

Anonymous:1907:IIk

[9891] Anonymous. Index of inventions. Scientific American, 96(19):399–403, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-399.pdf.

Anonymous:1907:PNPk

[9892] Anonymous. Patent notes: Patents. Scientific American, 96(19):403, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-403.pdf.

Anonymous:1907:PDa

[9893] Anonymous. The peril of the derelict. Scientific American, 96(19):386, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-386.pdf.

Anonymous:1907:LHE

[9894] Anonymous. The limits of heavy electric traction. Scientific American, 96(19):386, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
386a.pdf.

Anonymous:1907:TEE

[9895] Anonymous. Turbo-electric engines for ships? Scientific American, 96(19):386, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
386b.pdf.

Anonymous:1907:SMA

[9896] Anonymous. The Spring Meeting of the American Society of Mechanical Engineers. Scientific American, 96(19):386, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-386c.pdf.

Dunlap:1907:DGA

[9897] Orrin E. Dunlap. Deflocculated graphite and the “Acheson Effect”. Scientific American, 96(19):387, May 11, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n19/pdf/scientificamerican05111907-387.pdf.

Anonymous:1907:AHa

[9898] Anonymous. An automobile hearse. Scientific American, 96(19):388, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-388.pdf.

Anonymous:1907:BA

[9899] Anonymous. The Blériot aeroplane. Scientific American, 96(19):388, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-388a.pdf.

Gradenwitz:1907:TSD

[9900] Alfred Gradenwitz. The talking siren of Dr. Marage. Scientific American, 96(19):389–390, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
389.pdf.

Anonymous:1907:TAT

[9901] Anonymous. A theater auditorium turned into a ballroom at a moment’s notice. Scientific American, 96(19):391, May 11, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n19/pdf/scientificamerican05111907-391a.pdf.

Boyer:1907:MMF

[9902] Jacques Boyer. The manufacture of matches in France. Scientific American, 96(19):392–394, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
392.pdf.

Anonymous:1907:RGCb

[9903] Anonymous. Reconstructing a great city railroad terminal. Scientific American, 96(19):394, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
394.pdf.

Anonymous:1907:IVJ

[9904] Anonymous. An improved vehicle jack. Scientific American, 96(19):395, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-395.pdf.

Anonymous:1907:CLP

[9905] Anonymous. Coupling for large pipes. Scientific American, 96(19):395, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-395a.pdf.

Anonymous:1907:QAW

[9906] Anonymous. Quick-acting wrench. Scientific American, 96(19):395, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-395b.pdf.

Anonymous:1907:PSO

[9907] Anonymous. Portable saw-operating mechanism. Scientific American, 96(19):395, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
395c.pdf.

Anonymous:1907:NFM

[9908] Anonymous. A new form of mushroom bullet. Scientific American, 96(19):395–396, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
395d.pdf.

Anonymous:1907:PCb

[9909] Anonymous. Portable coal bin. Scientific American, 96(19):396, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n19/pdf/scientificamerican05111907-396.pdf.

Anonymous:1907:INA

[9910] Anonymous. Improved netting for airships. Scientific American, 96(19):396, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
396a.pdf.

Anonymous:1907:ASMa

[9911] Anonymous. Attachment for sewing machines. Scientific American, 96(19):396, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
396b.pdf.

Anonymous:1907:RAR

[9912] Anonymous. Roller appliance for rocking chairs. Scientific American, 96(19):396, May 11, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n19/pdf/scientificamerican05111907-
396c.pdf.

Anonymous:1907:RPIl

[9913] Anonymous. Recently patented inventions. Scientific American, 96(2):37, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
37.pdf.

Anonymous:1907:NQk

[9914] Anonymous. Notes and queries. Scientific American, 96(2):37, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-37a.pdf.

Anonymous:1907:Ci

[9915] Anonymous. Correspondence. Scientific American, 96(2):46, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-46.pdf.

Anonymous:1907:IIl

[9916] Anonymous. Index of inventions. Scientific American, 96(2):56–63, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-56.pdf.

Anonymous:1907:CCC

[9917] Anonymous. Cape Cod Canal. Scientific American, 96(2):22, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-22.pdf.

Anonymous:1907:PNH

[9918] Anonymous. The Pennsylvania and New Haven Connecting Railroad. Scientific American, 96(2):22, January 12, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n2/pdf/scientificamerican01121907-22a.pdf.

Anonymous:1907:SNA

[9919] Anonymous. The Seventh National Automobile Show at Madison Square Garden. Scientific American, 96(2):22–23, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-22b.pdf.

Anonymous:1907:SEA

[9920] Anonymous. Some early American automobiles. Scientific American, 96(2):23, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
23.pdf.

Anonymous:1907:Ea

[9921] Anonymous. Engines. Scientific American, 96(2):24, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-24.pdf.

Anonymous:1907:ECL

[9922] Anonymous. Engine of the car de luxe. Scientific American, 96(2):24, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
24a.pdf.

Anonymous:1907:ELT

[9923] Anonymous. An engine with low-tension magneto ignition. Scientific American, 96(2):25, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
25.pdf.

Anonymous:1907:AMR

[9924] Anonymous. An automobile motor with rotary valve. Scientific American, 96(2):25, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
25a.pdf.

Anonymous:1907:EHP

[9925] Anonymous. Engine of the 60-horse-power Thomas car. Scientific American, 96(2):25, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
25b.pdf.

Anonymous:1907:HPN

[9926] Anonymous. The 50-horse-power Northern engine. Scientific American, 96(2):25, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
25c.pdf.

Anonymous:1907:FDC

[9927] Anonymous. Friction-drive cars. Scientific American, 96(2):26, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-26.pdf.

Anonymous:1907:HRD

[9928] Anonymous. The Holsman rope-drive automobile. Scientific American, 96(2):26, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
26a.pdf.

Anonymous:1907:SFD

[9929] Anonymous. A successful friction-drive automobile. Scientific American, 96(2):26, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
26b.pdf.

Anonymous:1907:Ta

[9930] Anonymous. Transmissions. Scientific American, 96(2):27, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-27.pdf.

Anonymous:1907:NFT

[9931] Anonymous. A new form of two-speed transmission combined with the differential. Scientific American, 96(2):27, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-27a.pdf.

Anonymous:1907:TIJ

[9932] Anonymous. A transmission with individual jaw clutches. Scientific American, 96(2):27, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
27b.pdf.

Anonymous:1907:HTR

[9933] Anonymous. The Haynes transmission and roller bevel drive. Scientific American, 96(2):28, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
28.pdf.

Anonymous:1907:TGD

[9934] Anonymous. Transmission gear with double clutches. Scientific American, 96(2):28, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
28a.pdf.

Anonymous:1907:NCT

[9935] Anonymous. Novel clutch and transmission features of 1907 cars. Scientific American, 96(2):28–29, January 12, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n2/pdf/scientificamerican01121907-28b.pdf.

Anonymous:1907:LPD

[9936] Anonymous. The Lozier protected drive chain. Scientific American, 96(2):29, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
29.pdf.

Anonymous:1907:PTCa

[9937] Anonymous. The Pope–Toledo combined clutch and transmission. Scientific American, 96(2):29, January 12, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n2/pdf/scientificamerican01121907-29a.pdf.

Anonymous:1907:NIC

[9938] Anonymous. A novel individual clutch transmission. Scientific American, 96(2):29, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
29b.pdf.

Anonymous:1907:TBC

[9939] Anonymous. Typical brakes on 1907 cars. Scientific American, 96(2):30, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
30.pdf.

Anonymous:1907:ISG

[9940] Anonymous. An improved sliding-gear transmission. Scientific American, 96(2):30, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
30a.pdf.

Anonymous:1907:DCa

[9941] Anonymous. The dragon chassis. Scientific American, 96(2):30, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-30b.pdf.

Anonymous:1907:WTC

[9942] Anonymous. The Wayne transmission and clutch. Scientific American, 96(2):30–31, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
30c.pdf.

Anonymous:1907:NCS

[9943] Anonymous. A novel combined sliding-gear transmission and multipledisk clutch. Scientific American, 96(2):31, January 12, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n2/pdf/scientificamerican01121907-31.pdf.

Anonymous:1907:SDCc

[9944] Anonymous. The Stoddard–Dayton chassis. Scientific American, 96(2):31, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
31a.pdf.

Anonymous:1907:GHP

[9945] Anonymous. The grout 35-horse-power chassis. Scientific American, 96(2):31–33, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
31b.pdf.

Anonymous:1907:Cj

[9946] Anonymous. Controls. Scientific American, 96(2):34, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-34.pdf.

Anonymous:1907:RTCa

[9947] Anonymous. The Royal touring car and limousine. Scientific American, 96(2):34, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
34a.pdf.

Anonymous:1907:ITLa

[9948] Anonymous. An improved type of live rear axle. Scientific American, 96(2):34, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
34b.pdf.

Anonymous:1907:NFCa

[9949] Anonymous. The Northern four-cylinder 50-horse-power car. Scientific American, 96(2):34, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
34c.pdf.

Anonymous:1907:L

[9950] Anonymous. Lubricators. Scientific American, 96(2):36, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-36.pdf.

Anonymous:1907:Na

[9951] Anonymous. Novelties. Scientific American, 96(2):38, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-38.pdf.

Anonymous:1907:ITSa

[9952] Anonymous. An improved type of steering gear. Scientific American, 96(2):38, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
38a.pdf.

Anonymous:1907:ITSb

[9953] Anonymous. An instrument for testing shock absorbers. Scientific American, 96(2):38, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
38b.pdf.

Anonymous:1907:PSH

[9954] Anonymous. Pneumatic steel hub for automobile wheels. Scientific American, 96(2):39, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
39.pdf.

Anonymous:1907:NR

[9955] Anonymous. A novel reflector. Scientific American, 96(2):39, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-39a.pdf.

Jones:1907:FLD

[9956] H. D. Jones. Food losses during cooking. Scientific American, 96(2):40–42, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
40.pdf.

Anonymous:1907:PBRa

[9957] Anonymous. Photographs with bas-relief effects. Scientific American, 96(2):42, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
42.pdf.

Anonymous:1907:SSE

[9958] Anonymous. The smallest steam engine on Earth. Scientific American, 96(2):42, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
42a.pdf.

Anonymous:1907:IWS

[9959] Anonymous. Improvements in the White steam touring car. Scientific American, 96(2):43, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
43.pdf.

Anonymous:1907:ILD

[9960] Anonymous. An ingenious locking device. Scientific American, 96(2):43, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
43a.pdf.

Anonymous:1907:STA

[9961] Anonymous. A simple type of automobile speedometer. Scientific American, 96(2):43, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
43b.pdf.

Bache:1907:SMI

[9962] Rene Bache. Some modern inventions which are ancient history. Scientific American, 96(2):44, January 12, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n2/pdf/scientificamerican01121907-44.pdf.

Anonymous:1907:SIH

[9963] Anonymous. Some improvements in honeycomb radiators. Scientific American, 96(2):47, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
47.pdf.

Anonymous:1907:MTC

[9964] Anonymous. The Moon touring car. Scientific American, 96(2):48, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-48.pdf.

Anonymous:1907:HRC

[9965] Anonymous. The House recoil check. Scientific American, 96(2):49, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n2/pdf/scientificamerican01121907-49.pdf.

Anonymous:1907:PTCb

[9966] Anonymous. The 1907 Peerless touring car. Scientific American, 96(2):49–50, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
49a.pdf.

Anonymous:1907:ACE

[9967] Anonymous. The autocar combined engine and transmission. Scientific American, 96(2):51–52, January 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n2/pdf/scientificamerican01121907-
51.pdf.

Anonymous:1907:Ck

[9968] Anonymous. Correspondence. Scientific American, 96(20):411, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-411.pdf.

Anonymous:1907:RPIm

[9969] Anonymous. Recently patented inventions. Scientific American, 96(20):417, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
417.pdf.

Anonymous:1907:NQl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-417a.pdf.

Anonymous:1907:BNBg

[9971] Anonymous. Books: New books, etc. Scientific American, 96 (20):418–419, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
418.pdf.

Anonymous:1907:IIm

[9972] Anonymous. Index of inventions. Scientific American, 96(20):419–423, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-419.pdf.

Anonymous:1907:PNPl

[9973] Anonymous. Patent notes: Patents. Scientific American, 96(20):423, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-423.pdf.

Anonymous:1907:SSB

[9974] Anonymous. A speedy solution of the “broken rail” problem. Scientific American, 96(20):406, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
406.pdf.

Anonymous:1907:SDNa

[9975] Anonymous. The subway deadlock in New York. Scientific American, 96(20):406, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
406a.pdf.

Anonymous:1907:NUO

[9976] Anonymous. New uses for an old device — the gyroscopic railroad train. Scientific American, 96(20):406, May 18, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n20/pdf/scientificamerican05181907-406b.pdf.

Anonymous:1907:TAM

[9977] Anonymous. Third annual meeting and run of the automobile engineers. Scientific American, 96(20):407, May 18, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n20/pdf/scientificamerican05181907-407.pdf.

Anonymous:1907:NCP

[9978] Anonymous. New Commissioner of Patents. Scientific American, 96(20):407, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
407a.pdf.

Anonymous:1907:FHP

[9979] Anonymous. Flights by homing pigeons. Scientific American, 96 (20):407, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
407b.pdf.

Anonymous:1907:SNc

[9980] Anonymous. Science notes. Scientific American, 96(20):407, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-407c.pdf.

Bernard:1907:CHS

[9981] Walter Bernard. The construction and handling of submarines. Scientific American, 96(20):408, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
408.pdf.

Anonymous:1907:PBRb

[9982] Anonymous. The peril of the “broken rail” — its cause and cure. Scientific American, 96(20):409–410, May 18, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n20/pdf/scientificamerican05181907-409.pdf.

Hale:1907:CFE

[9983] Allen Hale. Construction of Florida East Coast Railway. Scientific American, 96(20):412–413, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
412.pdf.

Anonymous:1907:EEM

[9984] Anonymous. English Exposition of model flying machines. Scientific American, 96(20):414, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
414.pdf.

Anonymous:1907:PDb

[9985] Anonymous. Patent department. Scientific American, 96(20):415, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-415.pdf.

Anonymous:1907:IB

[9986] Anonymous. An improved bridle. Scientific American, 96(20):415, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-415a.pdf.

Anonymous:1907:MDS

[9987] Anonymous. Motor-driven stump puller. Scientific American, 96 (20):415, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
415b.pdf.

Anonymous:1907:MPPa

[9988] Anonymous. Metallic piston packing. Scientific American, 96(20):415, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-415c.pdf.

Anonymous:1907:IPW

[9989] Anonymous. Improved pipe wrench. Scientific American, 96 (20):415–416, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
415d.pdf.

Anonymous:1907:NSA

[9990] Anonymous. A novel swimming apparatus. Scientific American, 96(20):416, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
416.pdf.

Anonymous:1907:AS

[9991] Anonymous. Attachment for stoves. Scientific American, 96(20):416, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-416a.pdf.

Anonymous:1907:WS

[9992] Anonymous. Weighing scoop. Scientific American, 96(20):416, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n20/pdf/scientificamerican05181907-416b.pdf.

Anonymous:1907:AAL

[9993] Anonymous. Apparatus for agitating liquids. Scientific American, 96(20):416, May 18, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n20/pdf/scientificamerican05181907-
416c.pdf.

Anonymous:1907:Cl

[9994] Anonymous. Correspondence. Scientific American, 96(21):431, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n21/pdf/scientificamerican05251907-431.pdf.

Anonymous:1907:RPIn

[9995] Anonymous. Recently patented inventions. Scientific American, 96(21):437, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
437.pdf.

Anonymous:1907:NQm

[9996] Anonymous. Notes and queries. Scientific American, 96(21):438, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n21/pdf/scientificamerican05251907-438.pdf.

Anonymous:1907:BNBh

[9997] Anonymous. Books: New books, etc. Scientific American, 96(21):438, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n21/pdf/scientificamerican05251907-438a.pdf.

Anonymous:1907:IIn

[9998] Anonymous. Index of inventions. Scientific American, 96(21):438–443, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n21/pdf/scientificamerican05251907-438b.pdf.

Anonymous:1907:PNPm

[9999] Anonymous. Patent notes: Patents. Scientific American, 96(21):443, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n21/pdf/scientificamerican05251907-443.pdf.

Anonymous:1907:NRR

[10000] Anonymous. Needed — a rational railroad track. Scientific American, 96(21):426, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
426.pdf.

Anonymous:1907:EFG

[10001] Anonymous. Excellent foundations for the Gatun Locks. Scientific American, 96(21):426, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
426a.pdf.

Anonymous:1907:SPF

[10002] Anonymous. The smoke peril at fires. Scientific American, 96 (21):426–427, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
426b.pdf.

Tolman:1907:AMS

[10003] William H. Tolman. The American Museum of Safety Devices and Industrial Hygiene. Scientific American, 96(21):427, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n21/pdf/scientificamerican05251907-427.pdf.

Beasley:1907:EIE

[10004] Walter Beasley. Extraordinary insect enlargements. Scientific American, 96(21):428–430, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
428.pdf.

Anonymous:1907:NWS

[10005] Anonymous. The new White Star liner “Adriatic”. Scientific American, 96(21):431–432, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
431a.pdf.

Anonymous:1907:CHH

[10006] Anonymous. Charles Haynes Haswell. Scientific American, 96(21):433, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n21/pdf/scientificamerican05251907-433.pdf.

Honey:1907:ESJ

[10007] Frederic R. Honey. Eclipse of the Sun in July, 1907. Scientific American, 96(21):434, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
434.pdf.

Willey:1907:BTS

[10008] Day Allen Willey. Building a town site by suction dredge. Scientific American, 96(21):435–436, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
435.pdf.

Michaud:1907:CIa

[10009] Gustave Michaud. A curious illusion. Scientific American, 96(21):436, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n21/pdf/scientificamerican05251907-436.pdf.

Gradenwitz:1907:XRS

[10010] Alfred Gradenwitz. An X-ray stereoscope. Scientific American, 96(21):436, May 25, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n21/pdf/scientificamerican05251907-
436a.pdf.

Russell:1907:HJb

[10011] Henry Norris Russell. The heavens in June. Scientific American, 96(22):447, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
447.pdf.

Anonymous:1907:Cm

[10012] Anonymous. Correspondence. Scientific American, 96(22):451, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n22/pdf/scientificamerican06011907-451.pdf.

Anonymous:1907:RPIo

[10013] Anonymous. Recently patented inventions. Scientific American, 96(22):457, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
457.pdf.

Anonymous:1907:NQo

[10014] Anonymous. Notes and queries. Scientific American, 96(22):458, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n22/pdf/scientificamerican06011907-458.pdf.

Anonymous:1907:BNBi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n22/pdf/scientificamerican06011907-458a.pdf.

Anonymous:1907:IIo

[10016] Anonymous. Index of inventions. Scientific American, 96(22):458–463, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n22/pdf/scientificamerican06011907-458b.pdf.

Anonymous:1907:PNPn

[10017] Anonymous. Patent notes: Patents. Scientific American, 96(22):463, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n22/pdf/scientificamerican06011907-463.pdf.

Anonymous:1907:SWP

[10018] Anonymous. Slaughter in war and peace. Scientific American, 96(22):446, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
446.pdf.

Anonymous:1907:SIL

[10019] Anonymous. Speed indicators on locomotives. Scientific American, 96(22):446, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
446a.pdf.

Anonymous:1907:SSP

[10020] Anonymous. Single-service paper milk bottles. Scientific American, 96(22):446, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
446b.pdf.

Anonymous:1907:RMT

[10021] Anonymous. Running municipal trolley cars with garbage and refuse. Scientific American, 96(22):446, June 1, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n22/pdf/scientificamerican06011907-446c.pdf.

Anonymous:1907:DEL

[10022] Anonymous. Domestic electric light plant driven by a windmill. Scientific American, 96(22):448, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
448.pdf.

Anonymous:1907:BGMa

[10023] Anonymous. The Brennan gyroscope monorail. Scientific American, 96(22):449–450, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
449.pdf.

Anonymous:1907:OBW

[10024] Anonymous. Our 735,000,000-bushel wheat crop. Scientific American, 96(22):450, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
450.pdf.

Walsh:1907:JEM

[10025] George Ethelbert Walsh. To the Jamestown Exposition by motor boat. Scientific American, 96(22):450, June 1, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n22/pdf/scientificamerican06011907-450a.pdf.

Parsons:1907:SUA

[10026] T. S. Parsons. Some unknown American natural bridges. Scientific American, 96(22):452, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
452.pdf.

Bastin:1907:HMB

[10027] S. Leonard Bastin. A home-made barometer. Scientific American, 96(22):453–454, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
453.pdf.

Anonymous:1907:WSM

[10028] Anonymous. The wonderful sulphur mines of Louisiana. Scientific American, 96(22):454–456, June 1, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n22/pdf/scientificamerican06011907-
454.pdf.

Anonymous:1907:Cn

[10029] Anonymous. Correspondence. Scientific American, 96(23):471, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n23/pdf/scientificamerican06081907-471.pdf.

Anonymous:1907:RPIp

[10030] Anonymous. Recently patented inventions. Scientific American, 96(23):477, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
477.pdf.

Anonymous:1907:NQp

[10031] Anonymous. Notes and queries. Scientific American, 96(23):477–478, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n23/pdf/scientificamerican06081907-477a.pdf.

Anonymous:1907:BNBj

[10032] Anonymous. Books: New books, etc. Scientific American, 96(23):478, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n23/pdf/scientificamerican06081907-478.pdf.

Anonymous:1907:IIp

[10033] Anonymous. Index of inventions. Scientific American, 96(23):478–483, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n23/pdf/scientificamerican06081907-478a.pdf.

Anonymous:1907:PNPo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n23/pdf/scientificamerican06081907-483.pdf.

Anonymous:1907:TTS

[10035] Anonymous. Tunnel tubes in soft material. Scientific American, 96(23):466, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
466.pdf.

Anonymous:1907:PRE

[10036] Anonymous. The problem of raising an existing dam structure. Scientific American, 96(23):466, June 8, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n23/pdf/scientificamerican06081907-466a.pdf.

Anonymous:1907:SAE

[10037] Anonymous. The Shackleton Antarctic Expedition. Scientific American, 96(23):466, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
466b.pdf.

Anonymous:1907:ILCa

[10038] Anonymous. Influence of light on the conductivity of antimonite. Scientific American, 96(23):466, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
466c.pdf.

Anonymous:1907:GTSb

[10039] Anonymous. Government tests of safety devices for mines. Scientific American, 96(23):467, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
467.pdf.

Anonymous:1907:SNd

[10040] Anonymous. Science notes. Scientific American, 96(23):467, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n23/pdf/scientificamerican06081907-467a.pdf.

Anonymous:1907:LAB

[10041] Anonymous. The largest arch bridge in the world. Scientific American, 96(23):468, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
468.pdf.

Anonymous:1907:TCM

[10042] Anonymous. The two competitors in the motor boat race to Bermuda. Scientific American, 96(23):469–470, June 8, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n23/pdf/scientificamerican06081907-469.pdf.

Willey:1907:MN

[10043] Day Allen Willey. Mining in Newfoundland. Scientific American, 96(23):472–473, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
472.pdf.

Anonymous:1907:TDM

[10044] Anonymous. Teaching deaf-mutes to speak. Scientific American, 96(23):473–474, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
473.pdf.

Collins:1907:NTM

[10045] Percy Collins. Nature’s touch-me-nots. Scientific American, 96 (23):474–475, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
474.pdf.

Anonymous:1907:IGC

[10046] Anonymous. Improved grass cutter. Scientific American, 96(23):476, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n23/pdf/scientificamerican06081907-476.pdf.

Anonymous:1907:NSL

[10047] Anonymous. A novel sash lock. Scientific American, 96(23):476, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n23/pdf/scientificamerican06081907-476a.pdf.

Anonymous:1907:CLAa

[10048] Anonymous. Combined lightning arrester, fuse box, and insulator. Scientific American, 96(23):476, June 8, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n23/pdf/scientificamerican06081907-476b.pdf.

Anonymous:1907:NMH

[10049] Anonymous. A new method of handling acids. Scientific American, 96(23):476, June 8, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n23/pdf/scientificamerican06081907-
476c.pdf.

Anonymous:1907:Co

[10050] Anonymous. Correspondence. Scientific American, 96(24):491, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-491.pdf.

Anonymous:1907:RPIq

[10051] Anonymous. Recently patented inventions. Scientific American, 96(24):497, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
497.pdf.

Anonymous:1907:NQq

[10052] Anonymous. Notes and queries. Scientific American, 96(24):497–498, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-497a.pdf.

Anonymous:1907:BNBk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-498.pdf.

Anonymous:1907:IIq

[10054] Anonymous. Index of inventions. Scientific American, 96(24):498–503, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-498a.pdf.

Anonymous:1907:PNPp

[10055] Anonymous. Patent notes: Patents. Scientific American, 96(24):503, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-503.pdf.

Anonymous:1907:GLD

[10056] Anonymous. Gatun Lake is double the estimated size. Scientific American, 96(24):486, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
486.pdf.

Anonymous:1907:SFE

[10057] Anonymous. San Francisco earthquake and engineering construction. Scientific American, 96(24):486, June 15, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n24/pdf/scientificamerican06151907-486a.pdf.

Anonymous:1907:FCS

[10058] Anonymous. Fluid compression for steel rails. Scientific American, 96(24):486, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
486b.pdf.

Anonymous:1907:SBB

[10059] Anonymous. Sir Benjamin Baker. Scientific American, 96(24):486–487, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-486c.pdf.

Honey:1907:TM

[10060] Frederic R. Honey. The transit of Mercury in 1907. Scientific American, 96(24):487, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
487.pdf.

Anonymous:1907:THM

[10061] Anonymous. The two-hundred-mile automobile endurance test of the New York Motor Club. Scientific American, 96(24):487, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-487a.pdf.

Anonymous:1907:PWA

[10062] Anonymous. The protective works of the Assouan Dam. Scientific American, 96(24):488, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
488.pdf.

Trowbridge:1907:BL

[10063] John Trowbridge. Ball lightning. Scientific American, 96(24):489, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-489.pdf.

Anonymous:1907:NSM

[10064] Anonymous. A new signaling machine system for the automatic control of railroad trains. Scientific American, 96(24):489–490, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n24/pdf/scientificamerican06151907-489a.pdf.

Beasley:1907:DGA

[10065] Walter Beasley. Diplodocus: The greatest of all earthly creatures. Scientific American, 96(24):491–492, June 15, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n24/pdf/scientificamerican06151907-491a.pdf.

Anonymous:1907:ORT

[10066] Anonymous. The ocean race of torpedo-boat destroyers. Scientific American, 96(24):493, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
493.pdf.

Anonymous:1907:PTS

[10067] Anonymous. Practical tests of the Schlick gyrostat for ships. Scientific American, 96(24):494, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
494.pdf.

Stockard:1907:PLS

[10068] C. R. Stockard. A pair of limbless, snake-like lambs. Scientific American, 96(24):495, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
495.pdf.

Anonymous:1907:SHG

[10069] Anonymous. A successful hydroplane gliding boat. Scientific American, 96(24):495–496, June 15, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n24/pdf/scientificamerican06151907-
495a.pdf.

Anonymous:1907:MTW

[10070] Anonymous. A mechanical toy which draws geometrical designs. Scientific American, 96(24):496, June 15, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n24/pdf/scientificamerican06151907-496.pdf.

Anonymous:1907:Cp

[10071] Anonymous. Correspondence. Scientific American, 96(25):511, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-511.pdf.

Anonymous:1907:RPIr

[10072] Anonymous. Recently patented inventions. Scientific American, 96(25):517, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
517.pdf.

Anonymous:1907:NQr

[10073] Anonymous. Notes and queries. Scientific American, 96(25):517–518, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-517a.pdf.

Anonymous:1907:BNBl

[10074] Anonymous. Books: New books, etc. Scientific American, 96(25):518, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-518.pdf.

Anonymous:1907:IIr

[10075] Anonymous. Index of inventions. Scientific American, 96(25):518–523, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-518a.pdf.

Anonymous:1907:PNPq

[10076] Anonymous. Patent notes: Patents. Scientific American, 96(25):523, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-523.pdf.

Anonymous:1907:HLM

[10077] Anonymous. How the law may prevent railroad accidents. Scientific American, 96(25):506, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
506.pdf.

Anonymous:1907:WBL

[10078] Anonymous. We build a locomotive on the metric system. Scientific American, 96(25):506, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
506a.pdf.

Anonymous:1907:BSF

[10079] Anonymous. Brunel and the seven-foot gage. Scientific American, 96(25):506, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
506b.pdf.

Anonymous:1907:ALD

[10080] Anonymous. Another long-distance transmission in California. Scientific American, 96(25):506, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
506c.pdf.

Anonymous:1907:Tb

[10081] Anonymous. Tinol. Scientific American, 96(25):506–507, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-506d.pdf.

Fitz-Gerald:1907:TN

[10082] W. G. Fitz-Gerald. A telephone newspaper. Scientific American, 96(25):507, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
507.pdf.

Anonymous:1907:RMB

[10083] Anonymous. Result of the motor boat race to Bermuda. Scientific American, 96(25):507, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
507a.pdf.

Anonymous:1907:ARR

[10084] Anonymous. Automobile races and record runs. Scientific American, 96(25):507, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
507b.pdf.

Anonymous:1907:GTSc

[10085] Anonymous. The Government tests of submarines. Scientific American, 96(25):508, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
508.pdf.

Anonymous:1907:WPA

[10086] Anonymous. The Wellman Polar Airship Expedition. Scientific American, 96(25):509–510, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
509.pdf.

Anonymous:1907:DDSa

[10087] Anonymous. Double-deck streets as a relief for traffic congestion. Scientific American, 96(25):510, June 22, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n25/pdf/scientificamerican06221907-510.pdf.

Brenner:1907:ACM

[10088] Victor D. Brenner. The art and craft of the medalist. Scientific American, 96(25):511–512, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
511a.pdf.

Gradenwitz:1907:MME

[10089] Alfred Gradenwitz. Mechanical masterpieces of the eighteenth century. Scientific American, 96(25):513–514, June 22, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n25/pdf/scientificamerican06221907-513.pdf.

Camacho:1907:DAR

[10090] L. A. Camacho. Do animals reason? Scientific American, 96(25):515, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-515.pdf.

Hall:1907:AR

[10091] John W. Hall. An artificial respirator. Scientific American, 96(25):515, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-515a.pdf.

Anonymous:1907:ICL

[10092] Anonymous. An improved clothes line hanger. Scientific American, 96(25):516, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
516.pdf.

Anonymous:1907:NPM

[10093] Anonymous. A novel pneumatic massage machine. Scientific American, 96(25):516, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
516a.pdf.

Anonymous:1907:SPM

[10094] Anonymous. Spark-producing mechanism for explosion engines. Scientific American, 96(25):516, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n25/pdf/scientificamerican06221907-
516b.pdf.

Anonymous:1907:NSRa

[10095] Anonymous. Novel safety razor. Scientific American, 96(25):516, June 22, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n25/pdf/scientificamerican06221907-516c.pdf.

Russell:1907:HJc

[10096] Henry Norris Russell. The heavens in July. Scientific American, 96(26):527, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n26/pdf/scientificamerican06291907-
527.pdf.

Anonymous:1907:RPIs 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n26/pdf/scientificamerican06291907-
536.pdf.

Anonymous:1907:NQs

[10098] Anonymous. Notes and queries. Scientific American, 96(26):536–537, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n26/pdf/scientificamerican06291907-536a.pdf.

Anonymous:1907:BNBm

[10099] Anonymous. Books: New books, etc. Scientific American, 96(26):537, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n26/pdf/scientificamerican06291907-537.pdf.

Anonymous:1907:IIs

[10100] Anonymous. Index of inventions. Scientific American, 96(26):537–540, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n26/pdf/scientificamerican06291907-537a.pdf.

Anonymous:1907:Ia

[10101] Anonymous. Index. Scientific American, 96(26):541–542, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n26/pdf/scientificamerican06291907-541.pdf.

Anonymous:1907:PNPr

[10102] Anonymous. Patent notes: Patents. Scientific American, 96(26):543, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n26/pdf/scientificamerican06291907-543.pdf.

Anonymous:1907:CIN

[10103] Anonymous. Costly improvements in New York City. Scientific American, 96(26):526, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n26/pdf/scientificamerican06291907-
526.pdf.

Anonymous:1907:AYS

[10104] Anonymous. American yachting season of 1907. Scientific American, 96(26):526, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n26/pdf/scientificamerican06291907-
526a.pdf.

Anonymous:1907:FFMa

[10105] Anonymous. Forty-four-mile electrification on the West Shore Railroad. Scientific American, 96(26):526, June 29, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n26/pdf/scientificamerican06291907-526b.pdf.

Anonymous:1907:CME

[10106] Anonymous. Converting music into electricity. Scientific American, 96(26):527, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n26/pdf/scientificamerican06291907-
527a.pdf.

Anonymous:1907:NHB

[10107] Anonymous. A new hydroplane boat. Scientific American, 96(26):528, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n26/pdf/scientificamerican06291907-528.pdf.

Anonymous:1907:MPE

[10108] Anonymous. The most powerful express locomotive ever built. Scientific American, 96(26):528, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n26/pdf/scientificamerican06291907-
528a.pdf.

Claudy:1907:SSM

[10109] C. H. Claudy. Sound signals for mariners: An inventive field which is not overworked. Scientific American, 96(26):529–530, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n26/pdf/scientificamerican06291907-529.pdf.

Anonymous:1907:MVF

[10110] Anonymous. The magnitude of Victoria Falls. Scientific American, 96(26):530–531, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n26/pdf/scientificamerican06291907-
530.pdf.

Anonymous:1907:IWC

[10111] Anonymous. Inauguration of work on the Catskill New York Water Supply. Scientific American, 96(26):531, June 29, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n26/pdf/scientificamerican06291907-531.pdf.

Paullin:1907:VGO

[10112] Charles Oscar Paullin. The vessels and guns of the Old Navy, 1775– 1850. Scientific American, 96(26):532–534, June 29, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n26/pdf/scientificamerican06291907-532.pdf.

Boyer:1907:BNS

[10113] Jacques Boyer. Bertillon’s new system of anthropometry. Scientific American, 96(26):534–535, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n26/pdf/scientificamerican06291907-
534.pdf.

Gradenwitz:1907:NAD

[10114] Alfred Gradenwitz. A novel apparatus for demonstrating atmospheric pressure. Scientific American, 96(26):535, June 29, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n26/pdf/scientificamerican06291907-535.pdf.

Anonymous:1907:CCa

[10115] Anonymous. Correspondence: Correspondences. Scientific American, 96(3):70–71, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
70.pdf.

Anonymous:1907:RPIt

[10116] Anonymous. Recently patented inventions. Scientific American, 96(3):77, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
77.pdf.

Anonymous:1907:NQt

[10117] Anonymous. Notes and queries. Scientific American, 96(3):78–79, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n3/pdf/scientificamerican01191907-78.pdf.

Anonymous:1907:BNBn

[10118] Anonymous. Books: New books, etc. Scientific American, 96 (3):79, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
79.pdf.

Anonymous:1907:IIt

[10119] Anonymous. Index of inventions. Scientific American, 96(3):79–83, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n3/pdf/scientificamerican01191907-79a.pdf.

Anonymous:1907:PNPs

[10120] Anonymous. Patent notes: Patents. Scientific American, 96(3):84, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n3/pdf/scientificamerican01191907-84.pdf.

Anonymous:1907:LPC

[10121] Anonymous. Large powder chambers and gun erosion. Scientific American, 96(3):66, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
66.pdf.

Anonymous:1907:VRE

[10122] Anonymous. Value of rare earths for electrical purposes. Scientific American, 96(3):66, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
66a.pdf.

Anonymous:1907:PMN

[10123] Anonymous. Perhydrase milk — a new sterilized milk. Scientific American, 96(3):66–67, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
66b.pdf.

Anonymous:1907:RPFa

[10124] Anonymous. Recent performances of the French airship “Patrie”. Scientific American, 96(3):67, January 19, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n3/pdf/scientificamerican01191907-67.pdf.

Calhoun:1907:FAB

[10125] Katherine B. Calhoun. Facts about black lead pencils. Scientific American, 96(3):67, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
67a.pdf.

Anonymous:1907:YCT

[10126] Anonymous. The yawning cure for throat diseases. Scientific American, 96(3):67, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
67b.pdf.

Anonymous:1907:CPP

[10127] Anonymous. Climate: Past and present. Scientific American, 96(3):67, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
67c.pdf.

Anonymous:1907:CSe

[10128] Anonymous. The current supplement. Scientific American, 96 (3):67, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
67d.pdf.

Fessenden:1907:RPW

[10129] Reginald A. Fessenden. Recent progress in wireless telephony. Scientific American, 96(3):68–69, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
68.pdf.

Anonymous:1907:LS

[10130] Anonymous. The launch of the “Satsuma”. Scientific American, 96(3):69, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
69.pdf.

Willey:1907:MPT

[10131] D. A. Willey. A machine that predicts tides. Scientific American, 96(3):69–70, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
69a.pdf.

Anonymous:1907:OES

[10132] Anonymous. Opening of electric service on the New York Central and the New Haven Railroads. Scientific American, 96(3):72–73, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n3/pdf/scientificamerican01191907-72.pdf.

Lodian:1907:SFA

[10133] L. Lodian. Some facts about tea. Scientific American, 96(3):74, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n3/pdf/scientificamerican01191907-74.pdf.

Anonymous:1907:PDc

[10134] Anonymous. Patent department. Scientific American, 96(3):75, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n3/pdf/scientificamerican01191907-75.pdf.

Anonymous:1907:SSW

[10135] Anonymous. Snow shoes for wagons. Scientific American, 96(3):75, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n3/pdf/scientificamerican01191907-75a.pdf.

Anonymous:1907:ISO

[10136] Anonymous. An improved self-oiling roller bearing. Scientific American, 96(3):75–76, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
75b.pdf.

Anonymous:1907:UFD

[10137] Anonymous. Umbrella frame with detachable ribs. Scientific American, 96(3):76, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
76.pdf.

Anonymous:1907:IRT

[10138] Anonymous. An improved receiver for telephones. Scientific American, 96(3):76, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
76a.pdf.

Collins:1907:HPC

[10139] A. Frederick Collins. A handy portable crane and hoist. Scientific American, 96(3):76, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n3/pdf/scientificamerican01191907-
76b.pdf.

Anonymous:1907:LSH

[10140] Anonymous. Liquid soap holder. Scientific American, 96(3):76, January 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n3/pdf/scientificamerican01191907-76c.pdf.

Anonymous:1907:Cq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n4/pdf/scientificamerican01261907-91.pdf.

Anonymous:1907:RPIu

[10142] Anonymous. Recently patented inventions. Scientific American, 96(4):97, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
97.pdf.

Anonymous:1907:NQu

[10143] Anonymous. Notes and queries. Scientific American, 96(4):97–99, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n4/pdf/scientificamerican01261907-97a.pdf.

Anonymous:1907:BNBo

[10144] Anonymous. Books: New books, etc. Scientific American, 96 (4):99, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
99.pdf.

Anonymous:1907:IIu

[10145] Anonymous. Index of inventions. Scientific American, 96(4):99– 103, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
99a.pdf.

Anonymous:1907:PNPt

[10146] Anonymous. Patent notes: Patents. Scientific American, 96(4): 103, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
103.pdf.

Anonymous:1907:CMBa

[10147] Anonymous. The comedy of the Manhattan Bridge. Scientific American, 96(4):86, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
86.pdf.

Anonymous:1907:MHR

[10148] Anonymous. Mr. Hill on the railroad crisis. Scientific American, 96(4):86, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
86a.pdf.

Anonymous:1907:BBT

[10149] Anonymous. Brooklyn Bridge Terminal Station and Subway Loop. Scientific American, 96(4):86, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
86b.pdf.

Anonymous:1907:NEL

[10150] Anonymous. A new electric lamp filament. Scientific American, 96(4):86–87, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
86c.pdf.

Anonymous:1907:JE

[10151] Anonymous. The Jamaican earthquake. Scientific American, 96 (4):87, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
87.pdf.

Broadwell:1907:MED

[10152] Edward C. Broadwell. Method for electro-deposition and separation. Scientific American, 96(4):87, January 26, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n4/pdf/scientificamerican01261907-87a.pdf.

Anonymous:1907:BGMb

[10153] Anonymous. The bite of a gila monster. Scientific American, 96(4):87, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
87b.pdf.

Anonymous:1907:CSTb

[10154] Anonymous. The Cortlandt Street Tunnels and Terminal Building, New York. Scientific American, 96(4):88–90, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n4/pdf/scientificamerican01261907-88.pdf.

Fawcett:1907:WRS

[10155] Waldon Fawcett. The work of the rotary snow plow. Scientific American, 96(4):92, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
92.pdf.

Anonymous:1907:ESD

[10156] Anonymous. An exposition of safety devices and industrial hygiene. Scientific American, 96(4):93–94, January 26, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n4/pdf/scientificamerican01261907-93.pdf.

Anonymous:1907:SMEa

[10157] Anonymous. Some more electrical devices for the household. Scientific American, 96(4):95, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
95.pdf.

Anonymous:1907:LFB

[10158] Anonymous. Laws of falling bodies. Scientific American, 96(4):96, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n4/pdf/scientificamerican01261907-96.pdf.

Anonymous:1907:STLa

[10159] Anonymous. A safety trip for logging cars. Scientific American, 96(4):96, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
96a.pdf.

Anonymous:1907:SL

[10160] Anonymous. Sheathing lath. Scientific American, 96(4):96, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n4/pdf/scientificamerican01261907-96b.pdf.

Anonymous:1907:IMF

[10161] Anonymous. Improved measure for foaming liquids. Scientific American, 96(4):96, January 26, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n4/pdf/scientificamerican01261907-
96c.pdf.

Russell:1907:HF

[10162] Henry Norris Russell. The heavens in February. Scientific American, 96(5):107, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
107.pdf.

Anonymous:1907:Cr

[10163] Anonymous. Correspondence. Scientific American, 96(5):111, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n5/pdf/scientificamerican02021907-111.pdf.

Anonymous:1907:RPIv

[10164] Anonymous. Recently patented inventions. Scientific American, 96(5):117, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
117.pdf.

Anonymous:1907:NQv

[10165] Anonymous. Notes and queries. Scientific American, 96(5):117– 118, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
117b.pdf.

Anonymous:1907:BNBp (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
118.pdf.

Anonymous:1907:IIv

[10167] Anonymous. Index of inventions. Scientific American, 96(5): 118–123, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
118a.pdf.

Anonymous:1907:PNPu

[10168] Anonymous. Patent notes: Patents. Scientific American, 96(5): 123, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
123.pdf.

Anonymous:1907:ECT

[10169] Anonymous. The English channel tunnel scheme. Scientific American, 96(5):106, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
106.pdf.

Anonymous:1907:QSM

[10170] Anonymous. Quick-steaming marine boilers. Scientific American, 96(5):106, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
106a.pdf.

Anonymous:1907:SIC

[10171] Anonymous. Spontaneous ignition of coal. Scientific American, 96(5):106, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
106b.pdf.

Anonymous:1907:CHa

[10172] Anonymous. The Cornu helicopter. Scientific American, 96(5): 107, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
107a.pdf.

Willey:1907:MWH

[10173] Day Allen Willey. To-morrow’s weather: How it is foretold. Scientific American, 96(5):108–109, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
108.pdf.

Anonymous:1907:ECB

[10174] Anonymous. The “Express” coal-bagging lighter for coaling war vessels in harbor. Scientific American, 96(5):110–111, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n5/pdf/scientificamerican02021907-110.pdf.

Willey:1907:SRN

[10175] Day Allen Willey. The sea-rafts of the Northwest. Scientific American, 96(5):112–113, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
112.pdf.

Fitz-Gerald:1907:BMH

[10176] W. G. Fitz-Gerald. Breathing masks and helmets. Scientific American, 96(5):113–114, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
113.pdf.

Lydekker:1907:CE

[10177] R. Lydekker. Concerning ears. Scientific American, 96(5):115– 116, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
115.pdf.

Anonymous:1907:NTA

[10178] Anonymous. New toy automobiles. Scientific American, 96(5): 116, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
116.pdf.

Anonymous:1907:WBG

[10179] Anonymous. Window-blind guard. Scientific American, 96(5): 116, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
116a.pdf.

Anonymous:1907:CRD

[10180] Anonymous. Chuck for rock drills. Scientific American, 96(5): 116, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
116b.pdf.

Anonymous:1907:BPW

[10181] Anonymous. Business and personal wants. Scientific American, 96(5):117, February 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n5/pdf/scientificamerican02021907-
117a.pdf.

Anonymous:1907:Cu

[10182] Anonymous. Correspondence. Scientific American, 96(6):131, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n6/pdf/scientificamerican02091907-131.pdf.

Anonymous:1907:RPIw

[10183] Anonymous. Recently patented inventions. Scientific American, 96(6):137, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
137.pdf.

Anonymous:1907:NQw

[10184] Anonymous. Notes and queries. Scientific American, 96(6):137– 138, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
137a.pdf.

Anonymous:1907:BNBq

[10185] Anonymous. Books: New books, etc. Scientific American, 96 (6):138, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
138.pdf.

Anonymous:1907:IIw

[10186] Anonymous. Index of inventions. Scientific American, 96(6): 138–143, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
138a.pdf.

Anonymous:1907:PNPv

[10187] Anonymous. Patent notes: Patents. Scientific American, 96(6): 143, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
143.pdf.

Anonymous:1907:CLAb

[10188] Anonymous. The “casualty list” of American industries. Scientific American, 96(6):126, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
126.pdf.

Anonymous:1907:SFP

[10189] Anonymous. Satisfactory foundations for Panama dams and locks. Scientific American, 96(6):126, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
126a.pdf.

Anonymous:1907:PCCa

[10190] Anonymous. The Panama Canal contract. Scientific American, 96(6):126, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
126b.pdf.

Anonymous:1907:SSM

[10191] Anonymous. Should the steel makers subsidize our shipbuilders? Scientific American, 96(6):126, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
126c.pdf.

Anonymous:1907:WBO

[10192] Anonymous. Why the big one-caliber-gun battleships are best. Scientific American, 96(6):126–127, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
126d.pdf.

Anonymous:1907:CTA

[10193] Anonymous. Comparative test of alcohol, kerosene and gasoline as automobile fuels. Scientific American, 96(6):127, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n6/pdf/scientificamerican02091907-127.pdf.

Anonymous:1907:CSf

[10194] Anonymous. The current supplement. Scientific American, 96 (6):127, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
127a.pdf.

Anonymous:1907:TS

[10195] Anonymous. Telephone statistics. Scientific American, 96(6): 127, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
127b.pdf.

Anonymous:1907:ASS

[10196] Anonymous. The accident to the Stanley steam racer. Scientific American, 96(6):128, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
128.pdf.

Anonymous:1907:FMP

[10197] Anonymous. The fastest and most powerful American motor bicycle. Scientific American, 96(6):128, February 9, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n6/pdf/scientificamerican02091907-128a.pdf.

Luehrs:1907:MRV

[10198] Daniel M. Luehrs. Machine for recording the vibration of ships. Scientific American, 96(6):129–130, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
129.pdf.

Honey:1907:MES

[10199] Frederic R. Honey. Morning and evening stars for 1907. Scientific American, 96(6):130–131, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
130.pdf.

Anonymous:1907:SA

[10200] Anonymous. The somersaulting automobile. Scientific American, 96(6):132, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
132.pdf.

Cardullo:1907:BF

[10201] Forrest E. Cardullo. The battleship of the future — I. Scientific American, 96(6):133, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
133.pdf.

Wood:1907:OIP

[10202] R. W. Wood. The optical intensification of paintings. Scientific American, 96(6):134, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
134.pdf.

Anonymous:1907:NLK

[10203] Anonymous. The new lock and key bottle for holding poison. Scientific American, 96(6):135, February 9, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n6/pdf/scientificamerican02091907-135.pdf.

Anonymous:1907:SWN

[10204] Anonymous. The spider’s web as a negative. Scientific American, 96(6):135, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
135a.pdf.

Anonymous:1907:CMS

[10205] Anonymous. A clock made of slate. Scientific American, 96 (6):135, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
135b.pdf.

Anonymous:1907:PWT

[10206] Anonymous. Public weather towers in Vienna. Scientific American, 96(6):135, February 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n6/pdf/scientificamerican02091907-
135c.pdf.

Anonymous:1907:RPIx

[10207] Anonymous. Recently patented inventions. Scientific American, 96(7):156, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
156.pdf.

Anonymous:1907:NQx

[10208] Anonymous. Notes and queries. Scientific American, 96(7):156, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n7/pdf/scientificamerican02161907-156a.pdf.

Anonymous:1907:IIx

[10209] Anonymous. Index of inventions. Scientific American, 96(7): 156–159, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
156b.pdf.

Anonymous:1907:PNPw

[10210] Anonymous. Patent notes: Patents. Scientific American, 96 (7):159, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
159.pdf.

Anonymous:1907:GRN

[10211] Anonymous. Government regulation of Niagara Power. Scientific American, 96(7):146, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
146.pdf.

Anonymous:1907:SWS

[10212] Anonymous. Stupendous water supply scheme. Scientific American, 96(7):146, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
146a.pdf.

Anonymous:1907:SCSa

[10213] Anonymous. A simple-cylinder superheat locomotive test. Scientific American, 96(7):146, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
146b.pdf.

Anonymous:1907:OYW

[10214] Anonymous. One year’s wear of a steam turbine. Scientific American, 96(7):146, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
146c.pdf.

Anonymous:1907:RCC

[10215] Anonymous. Reinforced concrete construction on the Pacific Coast. Scientific American, 96(7):146–147, February 16, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n7/pdf/scientificamerican02161907-146d.pdf.

Anonymous:1907:OM

[10216] Anonymous. The odors of metals. Scientific American, 96(7): 147, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
147.pdf.

Anonymous:1907:ANP

[10217] Anonymous. Aerial navigation prizes. Scientific American, 96 (7):147, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
147a.pdf.

Anonymous:1907:SNe

[10218] Anonymous. Science notes. Scientific American, 96(7):147, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n7/pdf/scientificamerican02161907-147b.pdf.

Grimshaw:1907:KPF

[10219] Robert Grimshaw. Korn’s photographic fac-simile telegraph. Scientific American, 96(7):148, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
148.pdf.

Mclure:1907:CFT

[10220] W. Frank M’clure. The car ferry traffic of lake erie. Scientific American, 96(7):149–150, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
149.pdf.

Cardullo:1907:BFI

[10221] Forrest E. Cardullo. The battleship of the future — II. Scientific American, 96(7):150–151, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
150.pdf.

Wright:1907:TRP

[10222] Hamilton Wright. Tobacco raising in the Philippines. Scientific American, 96(7):152–153, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
152.pdf.

Anonymous:1907:SCBa

[10223] Anonymous. The subsidence of a concrete building. Scientific American, 96(7):154, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
154.pdf.

Anonymous:1907:RA

[10224] Anonymous. A remarkable accident. Scientific American, 96 (7):154, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
154a.pdf.

Anonymous:1907:VSB

[10225] Anonymous. Vehicle spring buffer. Scientific American, 96(7): 155, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
155.pdf.

Anonymous:1907:FAC

[10226] Anonymous. Fresh-air cabinet. Scientific American, 96(7):155, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n7/pdf/scientificamerican02161907-155a.pdf.

Anonymous:1907:MPPb

[10227] Anonymous. Metallic piston packing. Scientific American, 96 (7):155, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
155b.pdf.

Anonymous:1907:NEF

[10228] Anonymous. A new electric furnace. Scientific American, 96 (7):155, February 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n7/pdf/scientificamerican02161907-
155c.pdf.

Anonymous:1907:RPIy

[10229] Anonymous. Recently patented inventions. Scientific American, 96(8):174, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
174.pdf.

Anonymous:1907:NQy

[10230] Anonymous. Notes and queries. Scientific American, 96(8):174, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n8/pdf/scientificamerican02231907-174a.pdf.

Anonymous:1907:IIy

[10231] Anonymous. Index of inventions. Scientific American, 96(8): 174–181, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
174b.pdf.

Anonymous:1907:PNPx

[10232] Anonymous. Patent notes: Patents. Scientific American, 96 (8):183, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
183.pdf.

Anonymous:1907:LD

[10233] Anonymous. The “Larchmont” disaster. Scientific American, 96 (8):162, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
162.pdf.

Anonymous:1907:TMB

[10234] Anonymous. Trend of motor-boat development. Scientific American, 96(8):162, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
162a.pdf.

Anonymous:1907:GAM

[10235] Anonymous. Gyroscopic action of marine turbines. Scientific American, 96(8):162–163, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
162b.pdf.

Anonymous:1907:LDM

[10236] Anonymous. Long-distance motor-boat races for 1907. Scientific American, 96(8):163, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
163.pdf.

Anonymous:1907:PBN

[10237] Anonymous. Power boat notes. Scientific American, 96(8):163, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n8/pdf/scientificamerican02231907-163a.pdf.

Anonymous:1907:RPFb

[10238] Anonymous. Recent performances of French hydroplane boats. Scientific American, 96(8):164–165, February 23, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n8/pdf/scientificamerican02231907-164.pdf.

Anonymous:1907:BCR

[10239] Anonymous. The Branger Cup Race for miniature motor boats. Scientific American, 96(8):166, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
166.pdf.

Anonymous:1907:LTC

[10240] Anonymous. Latest types of cruising power boats. Scientific American, 96(8):167, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
167.pdf.

Stephens:1907:IAL

[10241] W. P. Stephens. Influence of the automobile on launch development. Scientific American, 96(8):167–168, February 23, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n8/pdf/scientificamerican02231907-167a.pdf.

Willits:1907:NTG

[10242] A. B. Willits. New types of gas engines for marine work. Scientific American, 96(8):169–170, February 23, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n8/pdf/scientificamerican02231907-169.pdf.

Anonymous:1907:HSM

[10243] Anonymous. A high-speed motor boat that can be built at home. Scientific American, 96(8):170, February 23, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n8/pdf/scientificamerican02231907-170.pdf.

Anonymous:1907:CAR

[10244] Anonymous. Canovetti’s air-resistance experiments. Scientific American, 96(8):170–171, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
170a.pdf.

Collins:1907:SNU

[10245] A. Frederick Collins. Some new and useful’s electric hand tools. Scientific American, 96(8):172–173, February 23, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n8/pdf/scientificamerican02231907-172.pdf.

Fling:1907:LGM

[10246] M. Glen Fling. Leather goods made of human skin. Scientific American, 96(8):173, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
173.pdf.

Anonymous:1907:HBR

[10247] Anonymous. Huge bowlders raised by bucket dredges. Scientific American, 96(8):173, February 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n8/pdf/scientificamerican02231907-
173a.pdf.

Russell:1907:HMb

[10248] Henry Norris Russell. The heavens in March. Scientific American, 96(9):187, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
187.pdf.

Anonymous:1907:Cv

[10249] Anonymous. Correspondence. Scientific American, 96(9):190–191, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n9/pdf/scientificamerican03021907-190a.pdf.

Anonymous:1907:RPIz

[10250] Anonymous. Recently patented inventions. Scientific American, 96(9):197–198, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
197.pdf.

Anonymous:1907:NQz

[10251] Anonymous. Notes and queries. Scientific American, 96(9):198–199, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n9/pdf/scientificamerican03021907-198.pdf.

Anonymous:1907:IIz

[10252] Anonymous. Index of inventions. Scientific American, 96(9):199–203, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n9/pdf/scientificamerican03021907-199.pdf.

Anonymous:1907:PNPy

[10253] Anonymous. Patent notes: Patents. Scientific American, 96(9):203, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n9/pdf/scientificamerican03021907-203.pdf.

Anonymous:1907:RSC

[10254] Anonymous. Risks of steel-concrete construction. Scientific American, 96(9):186, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
186.pdf.

Anonymous:1907:SSS

[10255] Anonymous. Should surface soil be stripped from reservoirs? Scientific American, 96(9):186, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
186a.pdf.

Anonymous:1907:MWS

[10256] Anonymous. The “Mikasa” was sunk by spontaneous magazine explosion. Scientific American, 96(9):186, March 2, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n9/pdf/scientificamerican03021907-186b.pdf.

Anonymous:1907:PELa

[10257] Anonymous. The peril of the electric locomotive on steam roads. Scientific American, 96(9):186, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
186c.pdf.

Anonymous:1907:PSS

[10258] Anonymous. Pearls from the Sulu Sea. Scientific American, 96(9):187, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n9/pdf/scientificamerican03021907-187a.pdf.

Anonymous:1907:LB

[10259] Anonymous. Light and blood. Scientific American, 96(9):187, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v96/n9/pdf/scientificamerican03021907-187b.pdf.

Collins:1907:UCT

[10260] A. Frederick Collins. A useful combination tool. Scientific American, 96(9):188, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
188.pdf.

Schoop:1907:ASM

[10261] M. U. Schoop. The autogenous soldering of metals. Scientific American, 96(9):189–190, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
189.pdf.

Anonymous:1907:MGI

[10262] Anonymous. The magnitude of the gas industry. Scientific American, 96(9):190, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
190.pdf.

Potter:1907:MBS

[10263] A. E. Potter. The Motor Boat Show at Madison Square Garden. Scientific American, 96(9):191–192, March 2, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n9/pdf/scientificamerican03021907-191.pdf.

Anonymous:1907:WET

[10264] Anonymous. Wreck of an electric train on the New York Central. Scientific American, 96(9):192–193, March 2, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n9/pdf/scientificamerican03021907-192.pdf.

Gradenwitz:1907:EDT

[10265] Alfred Gradenwitz. Electrically-driven target trains in the German Army. Scientific American, 96(9):194, March 2, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v96/
n9/pdf/scientificamerican03021907-194.pdf.

Anonymous:1907:KSA

[10266] Anonymous. The Krieger searchlight automobile. Scientific American, 96(9):194–195, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
194a.pdf.

Anonymous:1907:ANB

[10267] Anonymous. An Algerian natural bridge. Scientific American, 96(9):196, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
196.pdf.

Holder:1907:NFA

[10268] Charles Frederick Holder. A new fish for America. Scientific American, 96(9):196, March 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v96/n9/pdf/scientificamerican03021907-
196a.pdf.

Anonymous:1907:Cw

[10269] Anonymous. Correspondence. Scientific American, 97(1):9, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-9.pdf.

Anonymous:1907:RPIba

[10270] Anonymous. Recently patented inventions. Scientific American, 97(1): 14, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-14.pdf.

Anonymous:1907:NQba

[10271] Anonymous. Notes and queries. Scientific American, 97(1):14, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-14a.pdf.

Anonymous:1907:IIba

[10272] Anonymous. Index of inventions. Scientific American, 97(1):14–19, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-14b.pdf.

Anonymous:1907:PNPz

[10273] Anonymous. Patent notes: Patents. Scientific American, 97(1):18, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-18.pdf.

Anonymous:1907:ONF

[10274] Anonymous. Our Naval forces in the Pacific. Scientific American, 97(1): 4, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-4.pdf.

Anonymous:1907:ESS

[10275] Anonymous. The enemies of structural steel. Scientific American, 97(1): 4, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-4a.pdf.

Anonymous:1907:SMEb

[10276] Anonymous. Six months electric operation of the Grand Central Station. Scientific American, 97(1):4–5, July 6, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n1/pdf/scientificamerican07061907-4b.pdf.

Punnett:1907:OMT

[10277] Milton B. Punnett. Original methods of toning developed photographic prints. Scientific American, 97(1):5, July 6, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n1/pdf/scientificamerican07061907-5.pdf.

Anonymous:1907:SNf

[10278] Anonymous. Science notes. Scientific American, 97(1):5, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-5a.pdf.

Mills:1907:SAP

[10279] W. W. Mills. Simple apparatus for producing the manometric flame. Scientific American, 97(1):6, July 6, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n1/pdf/scientificamerican07061907-6.pdf.

Anonymous:1907:ROS

[10280] Anonymous. A remarkable outcrop of silver. Scientific American, 97(1): 6, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-6a.pdf.

Anonymous:1907:NAR

[10281] Anonymous. A new automobile race track. Scientific American, 97(1): 7, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-7.pdf.

Anonymous:1907:PFT

[10282] Anonymous. Pile foundations for tunnel tubes. Scientific American, 97(1):8, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n1/pdf/scientificamerican07061907-
8.pdf.

Willey:1907:NRC

[10283] Day Allen Willey. New railroad construction in Canada and the Northwest. Scientific American, 97(1):10–11, July 6, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n1/pdf/scientificamerican07061907-10.pdf.

Ward:1907:ESM

[10284] L. B. Ward. The economical surface mining operations of Cuba. Scientific American, 97(1):11–12, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n1/pdf/scientificamerican07061907-
11.pdf.

Anonymous:1907:WWS

[10285] Anonymous. Water-walking shoes. Scientific American, 97(1):12, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n1/pdf/scientificamerican07061907-12.pdf.

Anonymous:1907:SDCd

[10286] Anonymous. Santos Dumont’s combined aeroplane and airship. Scientific American, 97(1):12–13, July 6, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n1/pdf/scientificamerican07061907-
12a.pdf.

Anonymous:1907:RPIbb

[10287] Anonymous. Recently patented inventions. Scientific American, 97(10):173, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
173.pdf.

Anonymous:1907:NQbb

[10288] Anonymous. Notes and queries. Scientific American, 97(10): 174, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
174.pdf.

Anonymous:1907:BNBr

[10289] Anonymous. Books: New books, etc. Scientific American, 97 (10):174, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
174a.pdf.

Anonymous:1907:IIbb

[10290] Anonymous. Index of inventions. Scientific American, 97(10): 174–179, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
174b.pdf.

Anonymous:1907:PNPba

[10291] Anonymous. Patent notes: Patents. Scientific American, 97 (10):179, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
179.pdf.

Anonymous:1907:QBDa

[10292] Anonymous. The Quebec Bridge Disaster. Scientific American, 97(10):162, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
162.pdf.

Anonymous:1907:PCW

[10293] Anonymous. Progress of the Catskill Water Supply. Scientific American, 97(10):162, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
162a.pdf.

Anonymous:1907:COB

[10294] Anonymous. The cruise of our battleships to the Pacific. Scientific American, 97(10):162, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
162b.pdf.

Anonymous:1907:BR

[10295] Anonymous. Better rails for 1908. Scientific American, 97(10): 162, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
162c.pdf.

Anonymous:1907:WEG

[10296] Anonymous. Why erosion is greater in large guns. Scientific American, 97(10):162–163, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
162d.pdf.

Hale:1907:SIZ

[10297] William H. Hale. The Seventh International Zoological Congress — Boston Meeting, August 19 to 23, 1907. Scientific American, 97(10):163, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
163.pdf.

Anonymous:1907:SNg

[10298] Anonymous. Science notes. Scientific American, 97(10):163, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n10/pdf/scientificamerican09071907-163a.pdf.

Anonymous:1907:CSAa

[10299] Anonymous. Completion of the Scientific American Trophy for heavierthan-air flying machines. Scientific American, 97(10):163, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n10/pdf/scientificamerican09071907-163b.pdf.

Anonymous:1907:MEP

[10300] Anonymous. Modern experimental psychology — its methods and apparatus. Scientific American, 97(10):164–167, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n10/pdf/scientificamerican09071907-164.pdf.

Anonymous:1907:EFS

[10301] Anonymous. The erection of the 612-foot Singer Building. Scientific American, 97(10):168–169, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
168.pdf.

Anonymous:1907:ELP

[10302] Anonymous. Electric locomotives of the Pennsylvania Railroad. Scientific American, 97(10):170, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
170.pdf.

Anonymous:1907:MMT

[10303] Anonymous. Moving the Montauk Theater. Scientific American, 97(10):171, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
171.pdf.

Hamner:1907:LCR

[10304] E. C. Hamner. The launch of the caissons recently built at the Norfolk Navy Yard. Scientific American, 97(10):172, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n10/pdf/scientificamerican09071907-172.pdf.

Anonymous:1907:CSFa

[10305] Anonymous. The cold storage of furs. Scientific American, 97 (10):172–173, September 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n10/pdf/scientificamerican09071907-
172a.pdf.

Anonymous:1907:Cx

[10306] Anonymous. Correspondence. Scientific American, 97(11):187, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n11/pdf/scientificamerican09141907-187.pdf.

Anonymous:1907:RPIbc

[10307] Anonymous. Recently patented inventions. Scientific American, 97 (11):193–194, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
193.pdf.

Anonymous:1907:NQbc

[10308] Anonymous. Notes and queries. Scientific American, 97(11): 194–196, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
194.pdf.

Anonymous:1907:BNBs

[10309] Anonymous. Books: New books, etc. Scientific American, 97 (11):196, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
196.pdf.

Anonymous:1907:IIbc

[10310] Anonymous. Index of inventions. Scientific American, 97(11): 196–199, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
196a.pdf.

Anonymous:1907:PNPbb

[10311] Anonymous. Patent notes: Patents. Scientific American, 97 (11):200, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
200.pdf.

Anonymous:1907:PBD

[10312] Anonymous. A portentous bridge disaster. Scientific American, 97 (11):182, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
182.pdf.

Anonymous:1907:SFA

[10313] Anonymous. Successful flight of an aeroplane constructed after Langley’s model. Scientific American, 97(11):182, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n11/pdf/scientificamerican09141907-182a.pdf.

Anonymous:1907:EM

[10314] Anonymous. The evolution of matter. Scientific American, 97(11): 182–183, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
182b.pdf.

Crafts:1907:SPF

[10315] H. A. Crafts. Searching for parasites to fight fruit pests. Scientific American, 97(11):183, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
183.pdf.

Anonymous:1907:SNh

[10316] Anonymous. Science notes. Scientific American, 97(11):183, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n11/pdf/scientificamerican09141907-183a.pdf.

Anonymous:1907:FMR

[10317] Anonymous. A forgotten masterpiece — the Rittenhouse Orrery and its maker. Scientific American, 97(11):184, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n11/pdf/scientificamerican09141907-184.pdf.

Anonymous:1907:QBDb

[10318] Anonymous. The Quebec Bridge Disaster. Scientific American, 97 (11):186–187, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
186.pdf.

Anonymous:1907:NTL

[10319] Anonymous. The new turbine liner “Lusitania”. Scientific American, 97(11):188–190, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
188.pdf.

Anonymous:1907:OCA

[10320] Anonymous. One of the competing aeroplanes for the Scientific American aeronautical trophy. Scientific American, 97(11):190, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n11/pdf/scientificamerican09141907-190.pdf.

Anonymous:1907:SATa

[10321] Anonymous. The Scientific American Trophy for flying machines heavier than air. Scientific American, 97(11):191, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n11/pdf/scientificamerican09141907-191.pdf.

Anonymous:1907:PDd

[10322] Anonymous. Patent department. Scientific American, 97(11): 192, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
192.pdf.

Anonymous:1907:IPC

[10323] Anonymous. Improved pole-climbing shoe. Scientific American, 97 (11):192, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
192a.pdf.

Anonymous:1907:NAD

[10324] Anonymous. A novel advertising device. Scientific American, 97 (11):192, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
192b.pdf.

Anonymous:1907:CCW

[10325] Anonymous. Cushioned car wheel. Scientific American, 97(11): 192, September 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n11/pdf/scientificamerican09141907-
192c.pdf.

Anonymous:1907:Cy

[10326] Anonymous. Correspondence. Scientific American, 97(12):207, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n12/pdf/scientificamerican09211907-207.pdf.

Anonymous:1907:RPIbd

[10327] Anonymous. Recently patented inventions. Scientific American, 97 (12):212, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
212.pdf.

Anonymous:1907:IIbd

[10328] Anonymous. Index of inventions. Scientific American, 97(12): 212–215, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
212a.pdf.

Anonymous:1907:PNPbc

[10329] Anonymous. Patent notes: Patents. Scientific American, 97 (12):215, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
215.pdf.

Anonymous:1907:VTS

[10330] Anonymous. Value of transatlantic speed. Scientific American, 97 (12):202, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
202.pdf.

Anonymous:1907:MPB

[10331] Anonymous. Moving platforms for the Brooklyn Bridge. Scientific American, 97(12):202, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
202a.pdf.

Anonymous:1907:CMBb

[10332] Anonymous. Compression members in bridges. Scientific American, 97(12):202, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
202b.pdf.

Anonymous:1907:NC

[10333] Anonymous. New compounds. Scientific American, 97(12):202– 203, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
202c.pdf.

Anonymous:1907:GCE

[10334] Anonymous. A German chemist’s experiences in America. Scientific American, 97(12):203, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
203.pdf.

Anonymous:1907:BAF

[10335] Anonymous. Balloon, airship, and flying machine competitions at St. Louis. Scientific American, 97(12):203, September 21, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n12/pdf/scientificamerican09211907-203a.pdf.

Anonymous:1907:BRRb

[10336] Anonymous. Butterflies and the Roentgen rays. Scientific American, 97(12):203, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
203b.pdf.

Claudy:1907:RSB

[10337] C. H. Claudy. Regarding spar buoys. Scientific American, 97 (12):204, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
204.pdf.

Collins:1907:BEH

[10338] A. Frederick Collins. A boy’s engines and how he made them. Scientific American, 97(12):205, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
205.pdf.

Anonymous:1907:DDSb

[10339] Anonymous. The double-deck surface and tunnel station of the Hudson Companies in Jersey City. Scientific American, 97(12): 206–207, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
206.pdf.

Sidman:1907:NIG

[10340] Charles A. Sidman. New 14-inch gun for coast defense. Scientific American, 97(12):208, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
208.pdf.

Mclure:1907:CSL

[10341] W. Frank M’clure. The Chicago–St. Louis Waterway. Scientific American, 97(12):209–210, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
209.pdf.

Collins:1907:TEL

[10342] A. Frederick Collins. The “turn-down” electric lamp. Scientific American, 97(12):210, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n12/pdf/scientificamerican09211907-
210.pdf.

Shepstone:1907:BIF

[10343] Harold J. Shepstone. Breaking up the ill-fated British battleship “Montagu”. Scientific American, 97(12):211, September 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n12/pdf/scientificamerican09211907-211.pdf.

Anonymous:1907:Cz

[10344] Anonymous. Correspondence. Scientific American, 97(13):223, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n13/pdf/scientificamerican09281907-223.pdf.

Anonymous:1907:RPIbe

[10345] Anonymous. Recently patented inventions. Scientific American, 97 (13):228, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
228.pdf.

Anonymous:1907:NQbd

[10346] Anonymous. Notes and queries. Scientific American, 97(13): 228, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
228a.pdf.

Anonymous:1907:BNBt

[10347] Anonymous. Books: New books, etc. Scientific American, 97 (13):228, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
228b.pdf.

Anonymous:1907:IIbe

[10348] Anonymous. Index of inventions. Scientific American, 97(13): 228–231, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
228c.pdf.

Anonymous:1907:PNPbd

[10349] Anonymous. Patent notes: Patents. Scientific American, 97 (13):231, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
231.pdf.

Anonymous:1907:OEF

[10350] Anonymous. Our enormous fire losses. Scientific American, 97 (13):218, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
218.pdf.

Anonymous:1907:IAN

[10351] Anonymous. Important addition to New York City’s water supply. Scientific American, 97(13):218, September 28, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n13/pdf/scientificamerican09281907-218a.pdf.

Anonymous:1907:OMD

[10352] Anonymous. Observations of Mars during the recent opposition. Scientific American, 97(13):218, September 28, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n13/pdf/scientificamerican09281907-218b.pdf.

Anonymous:1907:CCAa

[10353] Anonymous. Cracked car axles. Scientific American, 97(13): 218, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
218c.pdf.

Keane:1907:FCS

[10354] Arthur H. J. Keane. Fabled cities submerged. Scientific American, 97(13):219, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
219.pdf.

Anonymous:1907:SNi

[10355] Anonymous. Science notes. Scientific American, 97(13):219, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n13/pdf/scientificamerican09281907-219a.pdf.

Anonymous:1907:SPB

[10356] Anonymous. A simple pipe-bending machine. Scientific American, 97(13):220, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
220.pdf.

Anonymous:1907:SLA

[10357] Anonymous. Simonis liquid-air apparatus. Scientific American, 97 (13):220, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
220a.pdf.

Wade:1907:WTU

[10358] Herbert T. Wade. Wireless telephony for the United States Navy. Scientific American, 97(13):221–223, September 28, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n13/pdf/scientificamerican09281907-221.pdf.

Anonymous:1907:DCb

[10359] Anonymous. Daniel’s Comet. Scientific American, 97(13):224, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n13/pdf/scientificamerican09281907-224.pdf.

Anonymous:1907:IIWa

[10360] Anonymous. If insects were as large as elephants. Scientific American, 97(13):225, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
225.pdf.

Anonymous:1907:NFR

[10361] Anonymous. The new Fastnet Rock lighthouse. Scientific American, 97(13):226–227, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
226.pdf.

Anonymous:1907:DWO

[10362] Anonymous. A device for washing the outside surfaces of windows. Scientific American, 97(13):227, September 28, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n13/pdf/scientificamerican09281907-227.pdf.

Anonymous:1907:NTCa

[10363] Anonymous. A new type of car fender. Scientific American, 97 (13):227, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
227a.pdf.

Anonymous:1907:HWB

[10364] Anonymous. Hand warmer and body heater. Scientific American, 97(13):227, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
227b.pdf.

Anonymous:1907:IFP

[10365] Anonymous. Improved fruit picker. Scientific American, 97(13): 227, September 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n13/pdf/scientificamerican09281907-
227c.pdf.

Russell:1907:HO

[10366] Henry Norris Russell. The heavens in October. Scientific American, 97(14):235, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
235.pdf.

Anonymous:1907:Cba

[10367] Anonymous. Correspondence. Scientific American, 97(14):239, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n14/pdf/scientificamerican10051907-239.pdf.

Anonymous:1907:RPIbf

[10368] Anonymous. Recently patented inventions. Scientific American, 97 (14):245–246, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
245.pdf.

Anonymous:1907:NQbe

[10369] Anonymous. Notes and queries. Scientific American, 97(14):246–247, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n14/pdf/scientificamerican10051907-246.pdf.

Anonymous:1907:BNBu

[10370] Anonymous. Books: New books, etc. Scientific American, 97(14):247, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n14/pdf/scientificamerican10051907-247.pdf.

Anonymous:1907:IIbf

[10371] Anonymous. Index of inventions. Scientific American, 97(14):247–251, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n14/pdf/scientificamerican10051907-247a.pdf.

Anonymous:1907:PNPbe

[10372] Anonymous. Patent notes: Patents. Scientific American, 97(14):251, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n14/pdf/scientificamerican10051907-251.pdf.

Anonymous:1907:PPC

[10373] Anonymous. Progress on the Panama Canal. Scientific American, 97(14):234, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
234.pdf.

Anonymous:1907:ACCa

[10374] Anonymous. The “America” Cup Challenge. Scientific American, 97(14):234, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
234a.pdf.

Anonymous:1907:PHF

[10375] Anonymous. A proposed Hudson–Fulton Celebration. Scientific American, 97(14):234, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
234b.pdf.

Anonymous:1907:TWT

[10376] Anonymous. Transatlantic wireless telegraphy. Scientific American, 97(14):234, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
234c.pdf.

Anonymous:1907:AEU

[10377] Anonymous. Aluminium for electric uses. Scientific American, 97(14):234, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
234d.pdf.

Shepstone:1907:FBM

[10378] Harold J. Shepstone. The first British military airship. Scientific American, 97(14):236–237, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
236.pdf.

Anonymous:1907:SGT

[10379] Anonymous. Sea-going torpedo boats. Scientific American, 97 (14):237–238, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
237.pdf.

Anonymous:1907:MMR

[10380] Anonymous. Manufacture of mechanical rubber goods. Scientific American, 97(14):240–242, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
240.pdf.

Boyer:1907:NPM

[10381] Jacques Boyer. The new photographic Meridian telescope of the Paris Observatory. Scientific American, 97(14):242, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n14/pdf/scientificamerican10051907-242.pdf.

Baldwin:1907:OTB

[10382] T. W. Baldwin. The Outlook Tower of Beinn Bhreagh, the first iron structure built of tetrahedral cells. Scientific American, 97(14):242–243, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n14/pdf/scientificamerican10051907-242a.pdf.

Anonymous:1907:CUC

[10383] Anonymous. Cooking utensil for campers. Scientific American, 97(14):244, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
244.pdf.

Anonymous:1907:ITLb

[10384] Anonymous. Improved track-laying machine. Scientific American, 97(14):244, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
244a.pdf.

Anonymous:1907:LUU

[10385] Anonymous. Ladder for use on uneven ground. Scientific American, 97(14):244, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n14/pdf/scientificamerican10051907-
244b.pdf.

Anonymous:1907:ILR

[10386] Anonymous. Improved leveling rod. Scientific American, 97(14):244, October 5, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n14/pdf/scientificamerican10051907-244c.pdf.

Anonymous:1907:Cbb

[10387] Anonymous. Correspondence. Scientific American, 97(15):259, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n15/pdf/scientificamerican10121907-259.pdf.

Anonymous:1907:RPIbg

[10388] Anonymous. Recently patented inventions. Scientific American, 97(15):264, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
264.pdf.

Anonymous:1907:NQbf

[10389] Anonymous. Notes and queries. Scientific American, 97(15):264, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n15/pdf/scientificamerican10121907-264a.pdf.

Anonymous:1907:IIbg

[10390] Anonymous. Index of inventions. Scientific American, 97(15): 264–267, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
264b.pdf.

Anonymous:1907:PNPbf

[10391] Anonymous. Patent notes: Patents. Scientific American, 97 (15):267, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
267.pdf.

Anonymous:1907:SCCb

[10392] Anonymous. A Swedish challenge for the Cup. Scientific American, 97(15):254, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
254.pdf.

Anonymous:1907:ASE

[10393] Anonymous. Ambitious scheme of electrification. Scientific American, 97(15):254, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
254a.pdf.

Anonymous:1907:OSR

[10394] Anonymous. Our stupendous railroad system. Scientific American, 97(15):254, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
254b.pdf.

Anonymous:1907:FTM

[10395] Anonymous. The formula and the testing machine. Scientific American, 97(15):254–255, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
254c.pdf.

Anonymous:1907:HPF

[10396] Anonymous. How to prevent failure in concrete construction. Scientific American, 97(15):255, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
255.pdf.

Anonymous:1907:ESMb

[10397] Anonymous. The Electrical Show at Madison Square Garden. Scientific American, 97(15):255, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
255a.pdf.

Butten:1907:CHI

[10398] Thomas C. Butten. The so-called hypnotic influence of snakes. Scientific American, 97(15):255, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
255b.pdf.

DiBrazza:1907:NSF

[10399] F. Savorgnan Di Brazza. A new system for the fixation of atmospheric nitrogen. Scientific American, 97(15):256–257, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n15/pdf/scientificamerican10121907-256.pdf.

Anonymous:1907:WQB

[10400] Anonymous. Why the Quebec Bridge failed. Scientific American, 97(15):257–258, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
257.pdf.

Gammeter:1907:GOB

[10401] H. C. Gammeter. The Gammeter orthopter — a beating-wing flying machine. Scientific American, 97(15):258, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n15/pdf/scientificamerican10121907-258.pdf.

MCaughrey:1907:SHM

[10402] L. W. M’Caughrey. A strong home-made turnbuckle. Scientific American, 97(15):259, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
259a.pdf.

Field:1907:WCA

[10403] C. Field. War cars, ancient and modern. Scientific American, 97 (15):260–262, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
260.pdf.

Anonymous:1907:BPM

[10404] Anonymous. Blood pressure and mental conditions. Scientific American, 97(15):262, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
262.pdf.

Gradenwitz:1907:UAC

[10405] Alfred Gradenwitz. A universal astronomical clock. Scientific American, 97(15):262–263, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
262a.pdf.

Knight:1907:GCC

[10406] Willam H. Knight. Gas and coke from crude oil. Scientific American, 97(15):263, October 12, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n15/pdf/scientificamerican10121907-
263.pdf.

Anonymous:1907:Cbc

[10407] Anonymous. Correspondence. Scientific American, 97(16):275, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n16/pdf/scientificamerican10191907-275.pdf.

Anonymous:1907:RPIbh

[10408] Anonymous. Recently patented inventions. Scientific American, 97(16):280, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
280.pdf.

Anonymous:1907:NQbg

[10409] Anonymous. Notes and queries. Scientific American, 97(16):280, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n16/pdf/scientificamerican10191907-280a.pdf.

Anonymous:1907:BNBv

[10410] Anonymous. Books: New books, etc. Scientific American, 97 (16):280, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
280b.pdf.

Anonymous:1907:IIbh

[10411] Anonymous. Index of inventions. Scientific American, 97(16): 280–283, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
280c.pdf.

Anonymous:1907:PNPbg

[10412] Anonymous. Patent notes: Patents. Scientific American, 97 (16):283, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
283.pdf.

Anonymous:1907:OFM

[10413] Anonymous. Old French machinery at Panama does good work. Scientific American, 97(16):270, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
270.pdf.

Anonymous:1907:FDB

[10414] Anonymous. The four-day boat. Scientific American, 97(16):270, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n16/pdf/scientificamerican10191907-270a.pdf.

Anonymous:1907:PRB

[10415] Anonymous. Prospects of relief of Brooklyn Bridge crush. Scientific American, 97(16):270, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
270b.pdf.

Anonymous:1907:BTR

[10416] Anonymous. Battery tunnel ready in two months. Scientific American, 97(16):270–271, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
270c.pdf.

Lydekker:1907:SMA

[10417] R. Lydekker. Some mechanical adaptations in animals. Scientific American, 97(16):271, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
271.pdf.

Anonymous:1907:WSU

[10418] Anonymous. The water supply of the United States. Scientific American, 97(16):271, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
271a.pdf.

Anonymous:1907:CSg

[10419] Anonymous. The current supplement. Scientific American, 97 (16):271, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
271b.pdf.

Lesh:1907:SPE

[10420] L. J. Lesh. Some preliminary experiments in flying. Scientific American, 97(16):272, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
272.pdf.

Dunlap:1907:INO

[10421] Orrin E. Dunlap. Illuminating Niagara with its own power. Scientific American, 97(16):273–274, October 19, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n16/pdf/scientificamerican10191907-273.pdf.

Anonymous:1907:EPC

[10422] Anonymous. Enormous power consumed in our industries. Scientific American, 97(16):274, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
274.pdf.

Anonymous:1907:HSA

[10423] Anonymous. How the Scientific American Trophy was made. Scientific American, 97(16):276, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
276.pdf.

Laking:1907:FVD

[10424] Francis Henry Laking. Food value of dried fruits. Scientific American, 97(16):276–277, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
276a.pdf.

Dunnet:1907:MCS

[10425] J. S. Dunnet. By motor car to the South Pole. Scientific American, 97(16):278, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
278.pdf.

Anonymous:1907:SDNb

[10426] Anonymous. Santos Dumont’s new hydroplane. Scientific American, 97(16):278, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n16/pdf/scientificamerican10191907-
278a.pdf.

Shepstone:1907:PCE

[10427] Harold J. Shepstone. Plant culture by electricity — an interesting experiment. Scientific American, 97(16):279, October 19, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n16/pdf/scientificamerican10191907-279.pdf.

Russell:1907:HN

[10428] Henry Norris Russell. The heavens in November. Scientific American, 97(18):303, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
303.pdf.

Anonymous:1907:Cbd

[10429] Anonymous. Correspondence. Scientific American, 97(18):307, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n18/pdf/scientificamerican11021907-307.pdf.

Anonymous:1907:RPIbi

[10430] Anonymous. Recently patented inventions. Scientific American, 97(18):312, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
312.pdf.

Anonymous:1907:IIbi

[10431] Anonymous. Index of inventions. Scientific American, 97(18): 312–315, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
312a.pdf.

Anonymous:1907:PNPbh

[10432] Anonymous. Patent notes: Patents. Scientific American, 97 (18):315, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
315.pdf.

Anonymous:1907:PCL

[10433] Anonymous. Panama Canal locks to be widened. Scientific American, 97(18):302, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
302.pdf.

Anonymous:1907:HOP

[10434] Anonymous. Hoist with its own petard. Scientific American, 97 (18):302, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
302a.pdf.

Anonymous:1907:HSV

[10435] Anonymous. The humanitarian side of vivisection. Scientific American, 97(18):302, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
302b.pdf.

Anonymous:1907:SUS

[10436] Anonymous. Shipbuilding in the United States. Scientific American, 97(18):302, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
302c.pdf.

Anonymous:1907:AHb

[10437] Anonymous. Artificial honey. Scientific American, 97(18):303, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n18/pdf/scientificamerican11021907-303a.pdf.

Anonymous:1907:SATb

[10438] Anonymous. The Scientific American Trophy and medal. Scientific American, 97(18):303, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
303b.pdf.

Anonymous:1907:CRH

[10439] Anonymous. The Crocco and Ricaldoni hydroplane boat. Scientific American, 97(18):304, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
304.pdf.

Anonymous:1907:CDPa

[10440] Anonymous. Combined dam and power plant. Scientific American, 97(18):305–306, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
305.pdf.

Anonymous:1907:LDB

[10441] Anonymous. The long distance balloon races from St. Louis. Scientific American, 97(18):308–309, November 2, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n18/pdf/scientificamerican11021907-308.pdf.

Collins:1907:CPT

[10442] A. Frederick Collins. A convenient pocket transit. Scientific American, 97(18):310, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
310.pdf.

Dublin:1907:NAM

[10443] Louis I. Dublin. Natural and artificial mixed colonies of ants. Scientific American, 97(18):310–311, November 2, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n18/pdf/scientificamerican11021907-310a.pdf.

Anonymous:1907:ITCb

[10444] Anonymous. An interesting type of cow milker. Scientific American, 97(18):311, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
311.pdf.

Anonymous:1907:FIT

[10445] Anonymous. Folding ironing table. Scientific American, 97(18): 311, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
311a.pdf.

Anonymous:1907:ET

[10446] Anonymous. Extension table. Scientific American, 97(18):311, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n18/pdf/scientificamerican11021907-311b.pdf.

Anonymous:1907:ATO

[10447] Anonymous. Attachment for traction-operated implements. Scientific American, 97(18):311, November 2, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n18/pdf/scientificamerican11021907-
311c.pdf.

Anonymous:1907:IIbj

[10448] Anonymous. Index of inventions. Scientific American, 97(19): 335, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
335.pdf.

Anonymous:1907:PNPbi

[10449] Anonymous. Patent notes: Patents. Scientific American, 97 (19):351, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
351.pdf.

Anonymous:1907:SPE

[10450] Anonymous. Single-phase electric equipment on the erie railroad. Scientific American, 97(19):320, November 9, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n19/pdf/scientificamerican11091907-320.pdf.

Anonymous:1907:ARS

[10451] Anonymous. The automobile at the recent shows. Scientific American, 97(19):320, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
320a.pdf.

Anonymous:1907:NFAb

[10452] Anonymous. A new French aeroplane record. Scientific American, 97(19):320–321, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
320b.pdf.

Walsh:1907:ROR

[10453] George Ethelbert Walsh. Rescuing our roads. Scientific American, 97(19):321, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
321.pdf.

Anonymous:1907:Eb

[10454] Anonymous. Engines. Scientific American, 97(19):322, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n19/pdf/scientificamerican11091907-322.pdf.

Anonymous:1907:IMA

[10455] Anonymous. Improvements in the Marmon air-cooled motor. Scientific American, 97(19):322, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
322a.pdf.

Anonymous:1907:NCE

[10456] Anonymous. The National 6-cylinder engine. Scientific American, 97(19):322, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
322b.pdf.

Anonymous:1907:HCD

[10457] Anonymous. The Hewitt 8-cylinder 90-degree motor. Scientific American, 97(19):323, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
323.pdf.

Anonymous:1907:CTE

[10458] Anonymous. The Carter two-engine car. Scientific American, 97 (19):323, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
323a.pdf.

Anonymous:1907:TCb

[10459] Anonymous. Transmissions and chassis. Scientific American, 97 (19):324, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
324.pdf.

Anonymous:1907:GTSd

[10460] Anonymous. The gearless two-speed transmission. Scientific American, 97(19):324, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
324a.pdf.

Anonymous:1907:PPH

[10461] Anonymous. The power plant of the Holsman car. Scientific American, 97(19):324, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
324b.pdf.

Anonymous:1907:UFT

[10462] Anonymous. A unique form of transmission gear. Scientific American, 97(19):324–325, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
324c.pdf.

Anonymous:1907:IC

[10463] Anonymous. The imperial chassis. Scientific American, 97(19): 325, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
325.pdf.

Anonymous:1907:LFD

[10464] Anonymous. The Lambert friction-drive chassis. Scientific American, 97(19):325, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
325a.pdf.

Greene:1907:TSA

[10465] Howard Greene. Types of shock absorbers for automobiles. Scientific American, 97(19):326, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
326.pdf.

Perry:1907:TMC

[10466] Harry W. Perry. Taximeter motor cabs in America. Scientific American, 97(19):327–329, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
327.pdf.

Haines:1907:HTA

[10467] Harry B. Haines. How to tour in an automobile. Scientific American, 97(19):330–331, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
330.pdf.

Anonymous:1907:Nb

[10468] Anonymous. Novelties. Scientific American, 97(19):332, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n19/pdf/scientificamerican11091907-332.pdf.

Anonymous:1907:AUH

[10469] Anonymous. An automobile used as a hoisting engine. Scientific American, 97(19):332, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
332a.pdf.

Anonymous:1907:ASG

[10470] Anonymous. Automatic speed governor. Scientific American, 97 (19):332, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
332b.pdf.

Anonymous:1907:ASA

[10471] Anonymous. The auto-slip Archimède — a portable apparatus for lifting automobiles for examination and repair. Scientific American, 97(19):332, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
332c.pdf.

Stellmann:1907:NMI

[10472] Louis J. Stellmann. A new musical instrument which produces violin effects. Scientific American, 97(19):334, November 9, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n19/pdf/scientificamerican11091907-334.pdf.

Anonymous:1907:EI

[10473] Anonymous. Electric incubators. Scientific American, 97(19): 336, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
336.pdf.

Anonymous:1907:ACT

[10474] Anonymous. The Aero Club Trophies at the Automobile Show in the Grand Central Palace. Scientific American, 97(19):338, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n19/pdf/scientificamerican11091907-338.pdf.

Anonymous:1907:NRE

[10475] Anonymous. “Newmastic” — a reliable and elastic filling for pneumatic tires. Scientific American, 97(19):339–340, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n19/pdf/scientificamerican11091907-339.pdf.

Anonymous:1907:NGT

[10476] Anonymous. A novel gasoline tester. Scientific American, 97 (19):342, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
342.pdf.

Anonymous:1907:ATG

[10477] Anonymous. Automobile traffic growth in New York. Scientific American, 97(19):343–344, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
343.pdf.

Anonymous:1907:SIM

[10478] Anonymous. Some improvements for the motor bicycle. Scientific American, 97(19):344–345, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
344.pdf.

Anonymous:1907:NFCb

[10479] Anonymous. A novel floatless carbureter. Scientific American, 97(19):345, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
345.pdf.

Anonymous:1907:SVA

[10480] Anonymous. Single vibrator apparatus for synchronizing the ignition sparks of multiple-unit coils. Scientific American, 97(19):346, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n19/pdf/scientificamerican11091907-346.pdf.

Anonymous:1907:NSD

[10481] Anonymous. The new Stoddard–Dayton engine. Scientific American, 97(19):346–347, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
346a.pdf.

Anonymous:1907:PMT

[10482] Anonymous. The 1908 Peerless motors and trans mission. Scientific American, 97(19):347–348, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
347.pdf.

Anonymous:1907:NTCb

[10483] Anonymous. A new two-cycle motor and truck. Scientific American, 97(19):348–349, November 9, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n19/pdf/scientificamerican11091907-
348.pdf.

Anonymous:1907:Cbe

[10484] Anonymous. Correspondence. Scientific American, 97(2):27, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-27.pdf.

Anonymous:1907:RPIbj

[10485] Anonymous. Recently patented inventions. Scientific American, 97(2):33–34, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
33.pdf.

Anonymous:1907:NQbh

[10486] Anonymous. Notes and queries. Scientific American, 97(2):34, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-34.pdf.

Anonymous:1907:BNBw

[10487] Anonymous. Books: New books, etc. Scientific American, 97(2):34–35, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-34a.pdf.

Anonymous:1907:IIbk

[10488] Anonymous. Index of inventions. Scientific American, 97(2):35–39, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-35.pdf.

Anonymous:1907:PNPbj

[10489] Anonymous. Patent notes: Patents. Scientific American, 97(2):39, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-39.pdf.

Anonymous:1907:NBO

[10490] Anonymous. Navy to build “Octopus” type of submarine. Scientific American, 97(2):22, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
22.pdf.

Anonymous:1907:EPP

[10491] Anonymous. Electric power plants are responsible. Scientific American, 97(2):22, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
22a.pdf.

Anonymous:1907:SRP

[10492] Anonymous. The steel rail problem. Scientific American, 97(2):22, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-22b.pdf.

Anonymous:1907:CAT

[10493] Anonymous. Contract awarded for two “Dreadnoughts”. Scientific American, 97(2):22, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
22c.pdf.

Anonymous:1907:BUC

[10494] Anonymous. Board of underwriters on cement construction. Scientific American, 97(2):22–23, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
22d.pdf.

Anonymous:1907:GPI

[10495] Anonymous. The Grand Prix International Automobile Race. Scientific American, 97(2):23, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
23.pdf.

Anonymous:1907:GMC

[10496] Anonymous. Gasoline motor cars adopted for branch line service on the Union Pacific Railroad. Scientific American, 97(2):23, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-23a.pdf.

Anonymous:1907:SNj

[10497] Anonymous. Science notes. Scientific American, 97(2):23, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-23b.pdf.

Anonymous:1907:TWS

[10498] Anonymous. A traction wheel with sinusoidal tread. Scientific American, 97(2):24, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
24.pdf.

Anonymous:1907:SP

[10499] Anonymous. Sand painting. Scientific American, 97(2):24, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-24a.pdf.

Anonymous:1907:RM

[10500] Anonymous. The riddle of Mars. Scientific American, 97(2):25–26, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-25.pdf.

Anonymous:1907:HLO

[10501] Anonymous. How linoleums and oilcloths are made. Scientific American, 97(2):28–30, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
28.pdf.

Anonymous:1907:SIT

[10502] Anonymous. Some interesting tricks. Scientific American, 97(2):30, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-30.pdf.

Gradenwitz:1907:CBG

[10503] Alfred Gradenwitz. Captive balloons in the German Army and Navy. Scientific American, 97(2):30–31, July 13, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n2/pdf/scientificamerican07131907-30a.pdf.

Anonymous:1907:PDe

[10504] Anonymous. Patent department. Scientific American, 97(2):32, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-32.pdf.

Anonymous:1907:ASD

[10505] Anonymous. Apparatus for sustaining and directing balloons. Scientific American, 97(2):32, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n2/pdf/scientificamerican07131907-
32a.pdf.

Anonymous:1907:PP

[10506] Anonymous. The plug puzzle. Scientific American, 97(2):32, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-32b.pdf.

Anonymous:1907:RTCb

[10507] Anonymous. Railway track construction. Scientific American, 97(2):32, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-32c.pdf.

Anonymous:1907:IBL

[10508] Anonymous. An improved button latch. Scientific American, 97(2):32, July 13, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n2/pdf/scientificamerican07131907-32d.pdf.

Anonymous:1907:Cbf

[10509] Anonymous. Correspondence. Scientific American, 97(20):359, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n20/pdf/scientificamerican11161907-359.pdf.

Anonymous:1907:RPIbk 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
365.pdf.

Anonymous:1907:NQbi

[10511] Anonymous. Notes and queries. Scientific American, 97(20): 366–367, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
366.pdf.

Anonymous:1907:BNBx

[10512] Anonymous. Books: New books, etc. Scientific American, 97 (20):367, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
367.pdf.

Anonymous:1907:IIbl

[10513] Anonymous. Index of inventions. Scientific American, 97(20): 367–371, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
367a.pdf.

Anonymous:1907:PNPbk

[10514] Anonymous. Patent notes: Patents. Scientific American, 97 (20):371, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
371.pdf.

Anonymous:1907:MBF

[10515] Anonymous. The man before the furnace. Scientific American, 97(20):354, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
354.pdf.

Anonymous:1907:PSTb

[10516] Anonymous. The present status of transatlantic, wireless telegraphy. Scientific American, 97(20):354, November 16, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n20/pdf/scientificamerican11161907-354a.pdf.

Anonymous:1907:ENA

[10517] Anonymous. The Eighth National Automobile Show in Madison Square Garden. Scientific American, 97(20):354, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n20/pdf/scientificamerican11161907-354b.pdf.

Mitchell:1907:HPP

[10518] S. A. Mitchell. How the planets are photographed. Scientific American, 97(20):355, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
355.pdf.

Anonymous:1907:CSh

[10519] Anonymous. The current supplement. Scientific American, 97 (20):355, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
355a.pdf.

Anonymous:1907:SNk

[10520] Anonymous. Science notes. Scientific American, 97(20):355, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n20/pdf/scientificamerican11161907-355b.pdf.

Anonymous:1907:OMS

[10521] Anonymous. Official Meteorological Summary, New York, N.Y., October, 1907. Scientific American, 97(20):355, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n20/pdf/scientificamerican11161907-355c.pdf.

Anonymous:1907:ESC

[10522] Anonymous. Edison’s system of concrete houses. Scientific American, 97(20):356, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
356.pdf.

Anonymous:1907:BSB

[10523] Anonymous. The Bredsdorff stranding buoy. Scientific American, 97 (20):356–357, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
356a.pdf.

Boneright:1907:NAA

[10524] Jhon Miller Boneright. A new American aeroplane. Scientific American, 97(20):357–358, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
357.pdf.

Anonymous:1907:LFAa

[10525] Anonymous. The latest French aeroplanes and their records. Scientific American, 97(20):358–359, November 16, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n20/pdf/scientificamerican11161907-358.pdf.

Anonymous:1907:VSI

[10526] Anonymous. The Vienna Serum Institute. Scientific American, 97 (20):360, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
360.pdf.

Willey:1907:MQC

[10527] Day Allen Willey. The marble quarries of Carrara. Scientific American, 97(20):361–362, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
361.pdf.

Anonymous:1907:NSB

[10528] Anonymous. The new Susquehanna Bridge, Pennsylvania Railroad. Scientific American, 97(20):362–363, November 16, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n20/pdf/scientificamerican11161907-362.pdf.

Shepstone:1907:LBS

[10529] Harold J. Shepstone. The launch of the bow of the “Suevic”. Scientific American, 97(20):363, November 16, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n20/pdf/scientificamerican11161907-363.pdf.

Smith:1907:CBS

[10530] Katherine Louise Smith. A curious bread school. Scientific American, 97(20):364, November 16, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n20/pdf/scientificamerican11161907-
364.pdf.

Anonymous:1907:Cbg

[10531] Anonymous. Correspondence. Scientific American, 97(21):379, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n21/pdf/scientificamerican11231907-379.pdf.

Anonymous:1907:RPIbl

[10532] Anonymous. Recently patented inventions. Scientific American, 97(21):384, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
384.pdf.

Anonymous:1907:IIbm

[10533] Anonymous. Index of inventions. Scientific American, 97(21): 384–387, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
384a.pdf.

Anonymous:1907:PNPbl

[10534] Anonymous. Patent notes: Patent. Scientific American, 97(21): 387, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
387.pdf.

Anonymous:1907:LSPa

[10535] Anonymous. Labor saving at Panama. Scientific American, 97 (21):374, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
374.pdf.

Anonymous:1907:ONW

[10536] Anonymous. Our national wastefulness. Scientific American, 97 (21):374, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
374a.pdf.

Anonymous:1907:CGL

[10537] Anonymous. Commerce on the Great Lakes. Scientific American, 97 (21):374–375, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
374b.pdf.

MBride:1907:LN

[10538] Raymond M’Bride. Logging in the Northwest. Scientific American, 97(21):375, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
375.pdf.

Anonymous:1907:SNl

[10539] Anonymous. Science notes. Scientific American, 97(21):375, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n21/pdf/scientificamerican11231907-375a.pdf.

Mitchell:1907:KRS

[10540] S. A. Mitchell. “Knots” in the rings of Saturn. Scientific American, 97(21):376, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
376.pdf.

Anonymous:1907:CSM

[10541] Anonymous. The Clifden Station of the Marconi Wireless Telegraph System. Scientific American, 97(21):377, November 23, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n21/pdf/scientificamerican11231907-377.pdf.

Anonymous:1907:LSO

[10542] Anonymous. The largest single office building in the world. Scientific American, 97(21):378, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
378.pdf.

Anonymous:1907:QTS

[10543] Anonymous. The quadruple turbine steamship “Mauretania”. Scientific American, 97(21):380–382, November 23, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n21/pdf/scientificamerican11231907-380.pdf.

Anonymous:1907:PDf

[10544] Anonymous. Patent department. Scientific American, 97(21): 383, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
383.pdf.

Anonymous:1907:ACB

[10545] Anonymous. Automatic coupling for brake pipes. Scientific American, 97(21):383, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
383a.pdf.

Anonymous:1907:WEW

[10546] Anonymous. Water elevator for windmills. Scientific American, 97(21):383, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
383b.pdf.

Anonymous:1907:IDP

[10547] Anonymous. Improved ditching plow. Scientific American, 97 (21):383, November 23, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n21/pdf/scientificamerican11231907-
383c.pdf.

Russell:1907:HD

[10548] Henry Norris Russell. The heavens in December. Scientific American, 97(22):391, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
391.pdf.

Anonymous:1907:Cbh

[10549] Anonymous. Correspondence. Scientific American, 97(22):394– 395, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
394.pdf.

Anonymous:1907:RPIbm

[10550] Anonymous. Recently patented inventions. Scientific American, 97(22):400, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
400.pdf.

Anonymous:1907:IIbn

[10551] Anonymous. Index of inventions. Scientific American, 97(22): 400–403, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
400a.pdf.

Anonymous:1907:PNPbm

[10552] Anonymous. Patent notes: Patents. Scientific American, 97 (22):403, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
403.pdf.

Anonymous:1907:LQB

[10553] Anonymous. Light on the Quebec Bridge Disaster. Scientific American, 97(22):390, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
390.pdf.

Anonymous:1907:TFE

[10554] Anonymous. Train ferries for the English channel. Scientific American, 97(22):390, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
390a.pdf.

Anonymous:1907:NSRb

[10555] Anonymous. National swamp reclamation. Scientific American, 97 (22):390, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
390b.pdf.

Anonymous:1907:UGC

[10556] Anonymous. The use of the gyroscope in connection with airships. Scientific American, 97(22):390, November 30, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n22/pdf/scientificamerican11301907-390c.pdf.

Anonymous:1907:ES

[10557] Anonymous. Electric sunstroke. Scientific American, 97(22): 391, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
391a.pdf.

Stewart:1907:NBW

[10558] J. A. Stewart. The new bridge over the Wissahickon at Philadelphia. Scientific American, 97(22):392–393, November 30, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n22/pdf/scientificamerican11301907-392.pdf.

Mitchell:1907:TM

[10559] S. A. Mitchell. The transit of Mercury. Scientific American, 97 (22):393, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
393.pdf.

Anonymous:1907:ERG

[10560] Anonymous. The enlargement and reconstruction of Governor’s Island. Scientific American, 97(22):395, November 30, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n22/pdf/scientificamerican11301907-395.pdf.

Anonymous:1907:PSR

[10561] Anonymous. The Peters self-registering electric target for rifle ranges. Scientific American, 97(22):395–396, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n22/pdf/scientificamerican11301907-395a.pdf.

Collins:1907:NNI

[10562] Percy Collins. The nests and nurseries of insects. Scientific American, 97(22):396–398, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
396.pdf.

Michaud:1907:SEM

[10563] Gustave Michaud. Some experiments with a mirror. Scientific American, 97(22):398, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
398.pdf.

Anonymous:1907:PDg

[10564] Anonymous. Patent department. Scientific American, 97(22): 399, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
399.pdf.

Anonymous:1907:NTS

[10565] Anonymous. A new type of storage battery. Scientific American, 97(22):399, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
399a.pdf.

Anonymous:1907:CWH

[10566] Anonymous. Combined water heater, filter, and cooler. Scientific American, 97(22):399, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
399b.pdf.

Anonymous:1907:IBBb

[10567] Anonymous. Improved boring bar. Scientific American, 97(22): 399, November 30, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n22/pdf/scientificamerican11301907-
399c.pdf.

Anonymous:1907:RPIbn

[10568] Anonymous. Recently patented inventions. Scientific American, 97(23):419, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
419.pdf.

Anonymous:1907:NQbj

[10569] Anonymous. Notes and queries. Scientific American, 97(23):423, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n23/pdf/scientificamerican12071907-423.pdf.

Anonymous:1907:IIbo

[10570] Anonymous. Index of inventions. Scientific American, 97(23): 423–428, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
423a.pdf.

Anonymous:1907:BNBy

[10571] Anonymous. Books: New books, etc. Scientific American, 97 (23):429, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
429.pdf.

Anonymous:1907:PNPbn

[10572] Anonymous. Patent notes: Patents. Scientific American, 97 (23):435, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
435.pdf.

Anonymous:1907:WCN

[10573] Anonymous. What constitutes Naval strength? Scientific American, 97(23):406, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
406.pdf.

Anonymous:1907:BGF

[10574] Anonymous. The battleship and gun of the future. Scientific American, 97(23):406, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
406a.pdf.

Anonymous:1907:RTP

[10575] Anonymous. Remarkable target practice by the flagship of the Pacific Fleet. Scientific American, 97(23):406, December 7, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n23/pdf/scientificamerican12071907-406b.pdf.

Anonymous:1907:MTB

[10576] Anonymous. The motor torpedo boat — a new type. Scientific American, 97(23):406, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
406c.pdf.

Mahan:1907:TSP

[10577] A. T. Mahan. The true significance of the Pacific cruise. Scientific American, 97(23):407, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
407.pdf.

Anonymous:1907:B

[10578] Anonymous. Battleships. Scientific American, 97(23):408, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n23/pdf/scientificamerican12071907-408.pdf.

Anonymous:1907:BKK

[10579] Anonymous. The battleships “Kearsarge” and “Kentucky”. Scientific American, 97(23):408, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
408a.pdf.

Anonymous:1907:FCBa

[10580] Anonymous. The first-class battleship “Alabama” — class of three ships. Scientific American, 97(23):408–409, December 7, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n23/pdf/scientificamerican12071907-408b.pdf.

Anonymous:1907:FCBb

[10581] Anonymous. First-class battleship “Maine” — class of three ships. Scientific American, 97(23):409, December 7, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n23/pdf/scientificamerican12071907-409.pdf.

Anonymous:1907:FCBc

[10582] Anonymous. First-class battleship “Georgia” — class of five ships. Scientific American, 97(23):409, December 7, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n23/pdf/scientificamerican12071907-409a.pdf.

Anonymous:1907:FCBd

[10583] Anonymous. First-class battleship “Connecticut” — class of six ships. Scientific American, 97(23):409–410, December 7, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n23/pdf/scientificamerican12071907-409b.pdf.

Anonymous:1907:Cbi

[10584] Anonymous. Cruisers. Scientific American, 97(23):411, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n23/pdf/scientificamerican12071907-411.pdf.

Anonymous:1907:PCCb

[10585] Anonymous. Protected cruiser “Charleston” — class of three ships. Scientific American, 97(23):411, December 7, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n23/pdf/scientificamerican12071907-411a.pdf.

Anonymous:1907:UCC

[10586] Anonymous. Unarmored cruiser “Chattanooga” and class. Scientific American, 97(23):411–412, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
411b.pdf.

Anonymous:1907:SCSb

[10587] Anonymous. The Scout cruiser “Salem” and class. Scientific American, 97(23):412, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
412.pdf.

Anonymous:1907:HDM

[10588] Anonymous. Harbor-defense monitors. Scientific American, 97 (23):413, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
413.pdf.

Anonymous:1907:TBD

[10589] Anonymous. Torpedo-boat destroyers. Scientific American, 97 (23):413, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
413a.pdf.

Anonymous:1907:WTP

[10590] Anonymous. Warship tonnage of the Principal Naval Powers. Scientific American, 97(23):414, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
414.pdf.

Beach:1907:USN

[10591] Edward L. Beach. The United States Naval Academy. Scientific American, 97(23):415–417, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
415.pdf.

Taylor:1907:SEM

[10592] D. W. Taylor. Simple explanation of model basin methods. Scientific American, 97(23):418, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
418.pdf.

Anonymous:1907:DOS

[10593] Anonymous. The development of our submarine Fleet. Scientific American, 97(23):420, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
420.pdf.

Anonymous:1907:GA

[10594] Anonymous. Guns and armor. Scientific American, 97(23):422, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n23/pdf/scientificamerican12071907-422.pdf.

Anonymous:1907:FCBe

[10595] Anonymous. First-class battleships “Idaho” and “Mississippi”. Scientific American, 97(23):430, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
430.pdf.

Anonymous:1907:FCBf

[10596] Anonymous. First-class battleships “South Carolina” and “Michigan”. Scientific American, 97(23):430–431, December 7, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n23/pdf/scientificamerican12071907-430a.pdf.

Anonymous:1907:TND

[10597] Anonymous. The 20-000-ton “North Dakota” and “Delaware”. Scientific American, 97(23):431, December 7, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n23/pdf/scientificamerican12071907-
431.pdf.

Anonymous:1907:Cbj

[10598] Anonymous. Correspondence. Scientific American, 97(24):443, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n24/pdf/scientificamerican12141907-443.pdf.

Anonymous:1907:RPIbo

[10599] Anonymous. Recently patented inventions. Scientific American, 97(24):448, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
448.pdf.

Anonymous:1907:BNBz

[10600] Anonymous. Books: New books, etc. Scientific American, 97 (24):448, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
448a.pdf.

Anonymous:1907:IIbp

[10601] Anonymous. Index of inventions. Scientific American, 97(24): 448–451, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
448b.pdf.

Anonymous:1907:PNPbo

[10602] Anonymous. Patent notes: Patents. Scientific American, 97 (24):451, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
451.pdf.

Anonymous:1907:SEE

[10603] Anonymous. Separate entrances and exits on cars. Scientific American, 97(24):438, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
438.pdf.

Anonymous:1907:CSNb

[10604] Anonymous. Comprehensive scheme for national water-ways. Scientific American, 97(24):438, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
438a.pdf.

Anonymous:1907:LFAb

[10605] Anonymous. The loss of the French airship “La Patrie”. Scientific American, 97(24):438, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
438b.pdf.

Anonymous:1907:AWW

[10606] Anonymous. Archæological work in Wyoming. Scientific American, 97(24):438–439, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
438c.pdf.

Anonymous:1907:MRD

[10607] Anonymous. The mysterious railway disasters in England. Scientific American, 97(24):439, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
439.pdf.

Anonymous:1907:SNm

[10608] Anonymous. Science notes. Scientific American, 97(24):439, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n24/pdf/scientificamerican12141907-439a.pdf.

Boyer:1907:NER

[10609] Jacques Boyer. New experiments in radio-activity. Scientific American, 97(24):440, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
440.pdf.

Lowell:1907:ST

[10610] Percival Lowell. Saturn’s Tores. Scientific American, 97(24): 441–442, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
441.pdf.

Baker:1907:EMM

[10611] Thomas R. Baker. An easily made micrometer. Scientific American, 97(24):443, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
443a.pdf.

Gradenwitz:1907:GWT

[10612] Alfred Gradenwitz. A glimpse of Western Tibet — an Austrian Expedition to the Forbidden Land. Scientific American, 97(24):444, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n24/pdf/scientificamerican12141907-444.pdf.

Anonymous:1907:SDL

[10613] Anonymous. Santos Dumont’s latest aeroplane. Scientific American, 97(24):445–446, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n24/pdf/scientificamerican12141907-
445.pdf.

Beasley:1907:CDR

[10614] Walter L. Beasley. A carnivorous dinosaur: A reconstructed skeleton of a huge saurian. Scientific American, 97(24):446–447, December 14, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n24/pdf/scientificamerican12141907-446.pdf.

Anonymous:1907:Cbk

[10615] Anonymous. Correspondence. Scientific American, 97(25):459, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n25/pdf/scientificamerican12211907-459.pdf.

Anonymous:1907:RPIbp

[10616] Anonymous. Recently patented inventions. Scientific American, 97(25):464, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
464.pdf.

Anonymous:1907:NQbk

[10617] Anonymous. Notes and queries. Scientific American, 97(25): 464, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
464a.pdf.

Anonymous:1907:BNBba

[10618] Anonymous. Books: New books, etc. Scientific American, 97(25): 464–465, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
464b.pdf.

Anonymous:1907:IIbq

[10619] Anonymous. Index of inventions. Scientific American, 97(25): 465–467, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
465.pdf.

Anonymous:1907:PNPbp

[10620] Anonymous. Patent notes: Patents. Scientific American, 97 (25):468, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
468.pdf.

Anonymous:1907:PELb

[10621] Anonymous. The President on employers’ liability. Scientific American, 97(25):454, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
454.pdf.

Anonymous:1907:DCG

[10622] Anonymous. Difficulties in constructing the Great Catskill Aqueduct Siphon. Scientific American, 97(25):454, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n25/pdf/scientificamerican12211907-454a.pdf.

Anonymous:1907:MO

[10623] Anonymous. The mastery of the ocean. Scientific American, 97 (25):454, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
454b.pdf.

Anonymous:1907:PNS

[10624] Anonymous. Proposed new ships for the Navy. Scientific American, 97(25):454–455, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
454c.pdf.

Anonymous:1907:RAE

[10625] Anonymous. Recent automobile endurance tests. Scientific American, 97(25):455, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
455.pdf.

Anonymous:1907:MBS

[10626] Anonymous. The Motor Boat Show at the Grand Central Palace. Scientific American, 97(25):455, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
455a.pdf.

Boyer:1907:BPT

[10627] Jacques Boyer. The Belin process of telephotography. Scientific American, 97(25):456–457, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
456.pdf.

Anonymous:1907:AHS

[10628] Anonymous. An artificial horizon for sextants. Scientific American, 97(25):457, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
457.pdf.

Smith:1907:MF

[10629] Katherine Louise Smith. Mexican fiestas. Scientific American, 97 (25):457–458, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
457a.pdf.

Mills:1907:DRT

[10630] James Cooke Mills. The Detroit River Tunnel. Scientific American, 97(25):458, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
458.pdf.

Anonymous:1907:RADb

[10631] Anonymous. Recent archæological discoveries in Egypt. Scientific American, 97(25):460–461, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
460.pdf.

Michaud:1907:CIb

[10632] Gustave Michaud. A chromatic illusion. Scientific American, 97 (25):462, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
462.pdf.

Anonymous:1907:NIF

[10633] Anonymous. New and ingenious French toys. Scientific American, 97 (25):462–463, December 21, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n25/pdf/scientificamerican12211907-
462a.pdf.

Anonymous:1907:RPIbq

[10634] Anonymous. Recently patented inventions. Scientific American, 97(26):481, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
481.pdf.

Anonymous:1907:NQbl

[10635] Anonymous. Notes and queries. Scientific American, 97(26): 482, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
482.pdf.

Anonymous:1907:IIbr

[10636] Anonymous. Index of inventions. Scientific American, 97(26): 482–484, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
482a.pdf.

Anonymous:1907:Id

[10637] Anonymous. Index. Scientific American, 97(26):485–486, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n26/pdf/scientificamerican12281907-485.pdf.

Anonymous:1907:PNPbq

[10638] Anonymous. Patent notes: Patents. Scientific American, 97 (26):487, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
487.pdf.

Anonymous:1907:BGL

[10639] Anonymous. Build Gatun Locks for the future. Scientific American, 97(26):470, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
470.pdf.

Anonymous:1907:PGF

[10640] Anonymous. Proposed Government forest preserves. Scientific American, 97(26):470, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
470a.pdf.

Anonymous:1907:EDR

[10641] Anonymous. Enormous death roll among the miners. Scientific American, 97(26):470, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
470b.pdf.

Anonymous:1907:VPS

[10642] Anonymous. Vibration in passenger ships. Scientific American, 97 (26):470–471, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
470c.pdf.

Anonymous:1907:ACCb

[10643] Anonymous. Anthracite coal centennial. Scientific American, 97 (26):471, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
471.pdf.

Benjamin:1907:ELN

[10644] Marcus Benjamin. Edward Leamington Nichols. Scientific American, 97(26):471, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
471a.pdf.

Anonymous:1907:CSi

[10645] Anonymous. The current supplement. Scientific American, 97 (26):471, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
471b.pdf.

Anonymous:1907:SNn

[10646] Anonymous. Science notes. Scientific American, 97(26):471, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n26/pdf/scientificamerican12281907-471c.pdf.

Anonymous:1907:FAS

[10647] Anonymous. The Foreign Automobile Show in Madison Square Garden. Scientific American, 97(26):471, December 28, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n26/pdf/scientificamerican12281907-471d.pdf.

Anonymous:1907:OS

[10648] Anonymous. To our subscribers. Scientific American, 97(26): 471, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
471e.pdf.

Anonymous:1907:SSAb

[10649] Anonymous. Safety suit of aeronauts. Scientific American, 97 (26):472, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
472.pdf.

Michaud:1907:CP

[10650] Gustave Michaud. A chemical puzzle. Scientific American, 97 (26):472, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
472a.pdf.

Hannum:1907:IDE

[10651] Harry A. Hannum. An instrument for drawing ellipses and hyperbolas. Scientific American, 97(26):472, December 28, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n26/pdf/scientificamerican12281907-472b.pdf.

Boyer:1907:NHD

[10652] Jacques Boyer. The new Henry dynamometer. Scientific American, 97(26):472–473, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
472c.pdf.

Anonymous:1907:OBT

[10653] Anonymous. The oval billiard table. Scientific American, 97 (26):473, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
473.pdf.

Honey:1907:TES

[10654] Frederic R. Honey. The total eclipse of the Sun in January, 1908. Scientific American, 97(26):474, December 28, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n26/pdf/scientificamerican12281907-474.pdf.

Anonymous:1907:LK

[10655] Anonymous. Lord Kelvin. Scientific American, 97(26):475, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n26/pdf/scientificamerican12281907-475.pdf.

Anonymous:1907:MHE

[10656] Anonymous. The manufacture of high explosives. Scientific American, 97(26):475–478, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
475a.pdf.

Gradenwitz:1907:IPS

[10657] Alfred Gradenwitz. The influence of profession on the shape of the hand. Scientific American, 97(26):478, December 28, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n26/pdf/scientificamerican12281907-478.pdf.

Anonymous:1907:DBA

[10658] Anonymous. Dr. Branly’s apparatus for control of distant mechanical effects. Scientific American, 97(26):478–480, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n26/pdf/scientificamerican12281907-478a.pdf.

Gilbert:1907:AVP

[10659] George Gilbert. An automatic violin player. Scientific American, 97(26):480, December 28, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n26/pdf/scientificamerican12281907-
480.pdf.

Anonymous:1907:Cbl

[10660] Anonymous. Correspondence. Scientific American, 97(3):47, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n3/pdf/scientificamerican07201907-47.pdf.

Anonymous:1907:RPIbr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n3/pdf/scientificamerican07201907-
52.pdf.

Anonymous:1907:IIbs

[10662] Anonymous. Index of inventions. Scientific American, 97(3):52–55, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n3/pdf/scientificamerican07201907-52a.pdf.

Anonymous:1907:PNPbr

[10663] Anonymous. Patent notes: Patents. Scientific American, 97(3):55, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n3/pdf/scientificamerican07201907-55.pdf.

Anonymous:1907:AWF

[10664] Anonymous. Around the world in forty days. Scientific American, 97(3):42, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n3/pdf/scientificamerican07201907-
42.pdf.

Anonymous:1907:SIN

[10665] Anonymous. Steady improvement in Naval gunnery. Scientific American, 97(3):42, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n3/pdf/scientificamerican07201907-
42a.pdf.

Anonymous:1907:RRO

[10666] Anonymous. Relative risk of overhead and third-rail conductors. Scientific American, 97(3):42, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n3/pdf/scientificamerican07201907-
42b.pdf.

Anonymous:1907:LSPb

[10667] Anonymous. The Lumière single-plate process of color photography. Scientific American, 97(3):42–43, July 20, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n3/pdf/scientificamerican07201907-42c.pdf.

Smith:1907:RS

[10668] J. Kent Smith. The rail situation. Scientific American, 97(3):43, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n3/pdf/scientificamerican07201907-43.pdf.

Anonymous:1907:EN

[10669] Anonymous. Engineering notes. Scientific American, 97(3):43, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n3/pdf/scientificamerican07201907-43a.pdf.

Boyer:1907:NAB

[10670] Jacques Boyer. A new automobile boat. Scientific American, 97(3):44, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n3/pdf/scientificamerican07201907-44.pdf.

Reese:1907:EFV

[10671] Cara Reese. An exhibition of fruits and vegetables in wax. Scientific American, 97(3):45–46, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n3/pdf/scientificamerican07201907-
45.pdf.

Honey:1907:MPE

[10672] Frederic R. Honey. The Moon’s phases and eclipse in July, 1907. Scientific American, 97(3):46, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n3/pdf/scientificamerican07201907-
46.pdf.

Hale:1907:OCE

[10673] Henry Hale. Osborne College: An English Naval School. Scientific American, 97(3):48–49, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n3/pdf/scientificamerican07201907-
48.pdf.

Morgan:1907:NIS

[10674] H. L. Morgan. A new identification system in the United States Army. Scientific American, 97(3):50, July 20, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n3/pdf/scientificamerican07201907-50.pdf.

Anonymous:1907:ITa

[10675] Anonymous. Ionic therapeutics. Scientific American, 97(3):51, July 20, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n3/pdf/scientificamerican07201907-51.pdf.

Russell:1907:HAb

[10676] Henry Norris Russell. The heavens in August. Scientific American, 97(4):59, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
59.pdf.

Anonymous:1907:Cbm

[10677] Anonymous. Correspondence. Scientific American, 97(4):63, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n4/pdf/scientificamerican07271907-63.pdf.

Anonymous:1907:RPIbs

[10678] Anonymous. Recently patented inventions. Scientific American, 97(4):68, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
68.pdf.

Anonymous:1907:IIbt

[10679] Anonymous. Index of inventions. Scientific American, 97(4):68–71, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n4/pdf/scientificamerican07271907-68a.pdf.

Anonymous:1907:PNPbs

[10680] Anonymous. Patent notes: Patents. Scientific American, 97(4):71, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n4/pdf/scientificamerican07271907-71.pdf.

Anonymous:1907:RGD

[10681] Anonymous. Is a revision of Gatun Dam necessary? Scientific American, 97(4):58, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
58.pdf.

Anonymous:1907:BGD

[10682] Anonymous. Battleship “Georgia” disaster. Scientific American, 97(4):58, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
58a.pdf.

Anonymous:1907:LRM

[10683] Anonymous. Latest results with marine turbines. Scientific American, 97(4):58, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
58b.pdf.

Anonymous:1907:ONH

[10684] Anonymous. Opening of the new haven railroad electric service. Scientific American, 97(4):58, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
58c.pdf.

Anonymous:1907:HCW

[10685] Anonymous. The heating of copper wires by electric currents. Scientific American, 97(4):59, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
59a.pdf.

Anonymous:1907:BRC

[10686] Anonymous. Building railway coaches with side doors. Scientific American, 97(4):59, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
59b.pdf.

Anonymous:1907:FRR

[10687] Anonymous. The first races and records on the New Weybridge Track. Scientific American, 97(4):60–61, July 27, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n4/pdf/scientificamerican07271907-60.pdf.

Anonymous:1907:BSC

[10688] Anonymous. Block signals in the cab of the locomotive. Scientific American, 97(4):61–62, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
61.pdf.

Anonymous:1907:DSW

[10689] Anonymous. Death of Sir William Perkin. Scientific American, 97(4):62, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
62.pdf.

Claudy:1907:PM

[10690] C. H. Claudy. Photomicrographs without a microscope. Scientific American, 97(4):64, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
64.pdf.

Boyer:1907:OA

[10691] Jacques Boyer. Orthopedic appliances. Scientific American, 97 (4):65–66, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
65.pdf.

Williston:1907:SSS

[10692] Arthur L. Williston. Some strange sources of water power on the Coast of Dalmatia. Scientific American, 97(4):66, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n4/pdf/scientificamerican07271907-66.pdf.

Anonymous:1907:PDh

[10693] Anonymous. Patent department. Scientific American, 97(4):67, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n4/pdf/scientificamerican07271907-67.pdf.

Anonymous:1907:ASMb

[10694] Anonymous. Attachment for sewing machines. Scientific American, 97(4):67, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
67a.pdf.

Anonymous:1907:APP

[10695] Anonymous. Answer to the plug puzzle. Scientific American, 97(4):67, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n4/pdf/scientificamerican07271907-67b.pdf.

Anonymous:1907:BBC

[10696] Anonymous. Building block for concrete construction. Scientific American, 97(4):67, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
67c.pdf.

Anonymous:1907:IH

[10697] Anonymous. An improvement in hammers. Scientific American, 97(4):67, July 27, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n4/pdf/scientificamerican07271907-
67d.pdf.

Anonymous:1907:RPIbt

[10698] Anonymous. Recently patented inventions. Scientific American, 97(5):85–86, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
85.pdf.

Anonymous:1907:NQbm

[10699] Anonymous. Notes and queries. Scientific American, 97(5):86–87, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-86.pdf.

Anonymous:1907:BNBbb

[10700] Anonymous. Books: New books, etc. Scientific American, 97(5):87, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-87.pdf.

Anonymous:1907:IIbu

[10701] Anonymous. Index of inventions. Scientific American, 97(5):87–91, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-87a.pdf.

Anonymous:1907:PNPbt

[10702] Anonymous. Patent notes: Patents. Scientific American, 97(5):91, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-91.pdf.

Anonymous:1907:CBo

[10703] Anonymous. A costly blunder. Scientific American, 97(5):74, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-74.pdf.

Anonymous:1907:PLI

[10704] Anonymous. The perils of the Long Island grade crossing. Scientific American, 97(5):74, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
74a.pdf.

Anonymous:1907:PRT

[10705] Anonymous. Perils of record target firing. Scientific American, 97(5):74, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-74b.pdf.

Anonymous:1907:GWO

[10706] Anonymous. Good work by our submarines. Scientific American, 97(5):74, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
74c.pdf.

Anonymous:1907:BOB

[10707] Anonymous. Bright outlook for better rails. Scientific American, 97(5):74, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
74d.pdf.

Anonymous:1907:DFA

[10708] Anonymous. Designer Fife on the “America” Cup Contest. Scientific American, 97(5):74–75, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
74e.pdf.

Anonymous:1907:FFMb

[10709] Anonymous. A Fleet of French military airships. Scientific American, 97(5):75, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
75.pdf.

Crafts:1907:PPD

[10710] H. A. Crafts. Present and prospective docking facilities of the Pacific Coast. Scientific American, 97(5):75, August 3, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n5/pdf/scientificamerican08031907-75a.pdf.

Anonymous:1907:CSj

[10711] Anonymous. The current supplement. Scientific American, 97(5):75, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-75b.pdf.

Anonymous:1907:PGC

[10712] Anonymous. A practical gliding craft with submerged hydroplanes. Scientific American, 97(5):76, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
76.pdf.

Anonymous:1907:RAA

[10713] Anonymous. Results of the American Automobile Association’s Fourth Annual Tour for the Glidden and Hower Trophies. Scientific American, 97(5):77–78, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
77.pdf.

Anonymous:1907:INH

[10714] Anonymous. The inauguration of the new haven railroad electric service. Scientific American, 97(5):79–81, August 3, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n5/pdf/scientificamerican08031907-79.pdf.

Collins:1907:IEW

[10715] A. Frederick Collins. An improved electric welder. Scientific American, 97(5):82, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
82.pdf.

Anonymous:1907:NVM

[10716] Anonymous. The new Vickers–Maxim automatic rifle-caliber gun. Scientific American, 97(5):83, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n5/pdf/scientificamerican08031907-
83.pdf.

Moore:1907:HMA

[10717] H. G. Moore. A home-made airship. Scientific American, 97(5):83–84, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-83a.pdf.

Anonymous:1907:ILCb

[10718] Anonymous. An improved lath cutter. Scientific American, 97(5):84, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-84.pdf.

Anonymous:1907:IVW

[10719] Anonymous. Improved vehicle wheel. Scientific American, 97(5):84, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-84a.pdf.

Anonymous:1907:ITb

[10720] Anonymous. An improvised tandem. Scientific American, 97(5):84, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-84b.pdf.

Anonymous:1907:IBS

[10721] Anonymous. An improved bench stop. Scientific American, 97(5):84, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-84c.pdf.

Anonymous:1907:OId

[10722] Anonymous. Oddities in invention. Scientific American, 97(5):84, August 3, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n5/pdf/scientificamerican08031907-84d.pdf.

Anonymous:1907:Cbp

[10723] Anonymous. Correspondence. Scientific American, 97(6):99, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n6/pdf/scientificamerican08101907-99.pdf.

Anonymous:1907:RPIbu

[10724] Anonymous. Recently patented inventions. Scientific American, 97 (6):105–106, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
105.pdf.

Anonymous:1907:NQbn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n6/pdf/scientificamerican08101907-106.pdf.

Anonymous:1907:BNBbc

[10726] Anonymous. Books: New books, etc. Scientific American, 97 (6):107, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
107.pdf.

Anonymous:1907:IIbv

[10727] Anonymous. Index of inventions. Scientific American, 97(6): 107–111, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
107a.pdf.

Anonymous:1907:PNPbu

[10728] Anonymous. Patent notes: Patents. Scientific American, 97(6):111, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n6/pdf/scientificamerican08101907-111.pdf.

Anonymous:1907:LKL

[10729] Anonymous. At last a 25-knot liner. Scientific American, 97(6):94, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n6/pdf/scientificamerican08101907-94.pdf.

Anonymous:1907:CDPb

[10730] Anonymous. Colorado dam is permanent. Scientific American, 97(6):94, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
94a.pdf.

Anonymous:1907:RE

[10731] Anonymous. Is rusting electrolytic? Scientific American, 97(6):94, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n6/pdf/scientificamerican08101907-94b.pdf.

Anonymous:1907:CCS

[10732] Anonymous. Cause and cure of split rail heads. Scientific American, 97(6):94–95, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
94c.pdf.

Anonymous:1907:RDD

[10733] Anonymous. Ramsay’s discovery of the degradation of copper to lithium. Scientific American, 97(6):95, August 10, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n6/pdf/scientificamerican08101907-95.pdf.

Horwill:1907:SAM

[10734] Herbert W. Horwill and M. A. The story of an ancient mine. Scientific American, 97(6):95, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
95a.pdf.

Anonymous:1907:SNo

[10735] Anonymous. Science notes. Scientific American, 97(6):95, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n6/pdf/scientificamerican08101907-95b.pdf.

Anonymous:1907:KTL

[10736] Anonymous. The 25-knot turbine liner “Lusitania”. Scientific American, 97(6):96, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
96.pdf.

Fergusson:1907:IKA

[10737] S. P. Fergusson. The international kite ascensions. Scientific American, 97(6):97–98, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
97.pdf.

Anonymous:1907:EBI

[10738] Anonymous. The erection of the Blackwell’s Island Bridge. Scientific American, 97(6):100–101, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
100.pdf.

Kane:1907:RPC

[10739] Louis C. Kane. A revolution in postal car construction. Scientific American, 97(6):102–103, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
102.pdf.

Beasley:1907:GM

[10740] Walter L. Beasley. The greatest of mastodons. Scientific American, 97(6):103–104, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
103.pdf.

Anonymous:1907:EDT

[10741] Anonymous. An electric dining table. Scientific American, 97 (6):104–105, August 10, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n6/pdf/scientificamerican08101907-
104.pdf.

Anonymous:1907:Cbq

[10742] Anonymous. Correspondence. Scientific American, 97(7):119, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n7/pdf/scientificamerican08171907-119.pdf.

Anonymous:1907:RPIbv

[10743] Anonymous. Recently patented inventions. Scientific American, 97(7):124, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
124.pdf.

Anonymous:1907:BNBbd

[10744] Anonymous. Books: New books, etc. Scientific American, 97 (7):124, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
124a.pdf.

Anonymous:1907:IIbw

[10745] Anonymous. Index of inventions. Scientific American, 97(7): 124–127, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
124b.pdf.

Anonymous:1907:PNPbv

[10746] Anonymous. Patent notes: Patents. Scientific American, 97(7):127, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n7/pdf/scientificamerican08171907-127.pdf.

Anonymous:1907:CCb

[10747] Anonymous. Concrete and cupidity. Scientific American, 97(7):114, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n7/pdf/scientificamerican08171907-114.pdf.

Anonymous:1907:ACN

[10748] Anonymous. Ambrose Channel now one-half completed. Scientific American, 97(7):114, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
114a.pdf.

Anonymous:1907:SPA

[10749] Anonymous. Smokeless powder and ammunition rooms. Scientific American, 97(7):114, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
114b.pdf.

Anonymous:1907:RCA

[10750] Anonymous. Route of Catskill Aqueduct changed. Scientific American, 97(7):114, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
114c.pdf.

Anonymous:1907:GOE

[10751] Anonymous. Growth of our exports of manufactures. Scientific American, 97(7):114, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
114d.pdf.

Anonymous:1907:PEF

[10752] Anonymous. A prize for the extraction of flax. Scientific American, 97(7):114–115, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
114e.pdf.

Stratton:1907:CI

[10753] George Frederick Stratton. A college of invention. Scientific American, 97(7):115, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
115.pdf.

Anonymous:1907:SNp

[10754] Anonymous. Science notes. Scientific American, 97(7):115, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n7/pdf/scientificamerican08171907-115a.pdf.

Anonymous:1907:CSk

[10755] Anonymous. The current supplement. Scientific American, 97 (7):115, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
115b.pdf.

Anonymous:1907:SSR

[10756] Anonymous. A scientific study of the racehorse, and its use in mounting the Skeleton of Sysonby. Scientific American, 97(7):116–117, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n7/pdf/scientificamerican08171907-116.pdf.

Anonymous:1907:NUS

[10757] Anonymous. The new United States Scout cruiser “Salem”. Scientific American, 97(7):117–118, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
117.pdf.

Brownell:1907:HMA

[10758] Baker Brownell. A home-made air thermometer. Scientific American, 97(7):118, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
118.pdf.

Anonymous:1907:LLE

[10759] Anonymous. The largest locomotive ever constructed. Scientific American, 97(7):119–121, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
119a.pdf.

Anonymous:1907:LDW

[10760] Anonymous. A locomotive driving-wheel recording device. Scientific American, 97(7):122, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
122.pdf.

VanBrussel:1907:NEI

[10761] J. B. Van Brussel. A new Egyptian irrigation canal. Scientific American, 97(7):122–123, August 17, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n7/pdf/scientificamerican08171907-
122a.pdf.

Anonymous:1907:Cbr

[10762] Anonymous. Correspondence. Scientific American, 97(8):135, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n8/pdf/scientificamerican08241907-135a.pdf.

Anonymous:1907:RPIbw

[10763] Anonymous. Recently patented inventions. Scientific American, 97(8):140, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
140.pdf.

Anonymous:1907:BNBbe

[10764] Anonymous. Books: New books, etc. Scientific American, 97 (8):140, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
140a.pdf.

Anonymous:1907:IIbx

[10765] Anonymous. Index of inventions. Scientific American, 97(8): 140–143, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
140b.pdf.

Anonymous:1907:PNPbw

[10766] Anonymous. Patent notes: Patents. Scientific American, 97(8):143, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n8/pdf/scientificamerican08241907-143.pdf.

Anonymous:1907:QTE

[10767] Anonymous. The question of two engineers in the cab. Scientific American, 97(8):130, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
130.pdf.

Anonymous:1907:STLb

[10768] Anonymous. Steam trials of the “Lusitania”. Scientific American, 97(8):130, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
130a.pdf.

Anonymous:1907:CFR

[10769] Anonymous. “Cecilie”: The fastest reciprocating-engine liner afloat. Scientific American, 97(8):130, August 24, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n8/pdf/scientificamerican08241907-130b.pdf.

Anonymous:1907:GFT

[10770] Anonymous. Government fuel tests at Jamestown. Scientific American, 97(8):131, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
131.pdf.

Anonymous:1907:HAR

[10771] Anonymous. The 24-hour automobile race on the Brighton Beach Track. Scientific American, 97(8):131, August 24, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n8/pdf/scientificamerican08241907-131a.pdf.

Anonymous:1907:CSl

[10772] Anonymous. The current supplement. Scientific American, 97 (8):131, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
131b.pdf.

Anonymous:1907:SNq

[10773] Anonymous. Science notes. Scientific American, 97(8):131, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n8/pdf/scientificamerican08241907-131c.pdf.

Anonymous:1907:OLD

[10774] Anonymous. Opening of the large drydock at League Island. Scientific American, 97(8):132, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
132.pdf.

Geare:1907:MC

[10775] Randolph I. Geare. Magical cures. Scientific American, 97(8): 132–133, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
132a.pdf.

Lent:1907:HMA

[10776] Wilmar F. Lent. A home-made alternating-current rectifier. Scientific American, 97(8):134, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
134.pdf.

Anonymous:1907:MWE

[10777] Anonymous. The meats we eat. Scientific American, 97(8):135, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n8/pdf/scientificamerican08241907-135.pdf.

Peltier:1907:TNT

[10778] M. J. Peltier. The “Typhonoid” — a new type of motor boat. Scientific American, 97(8):136, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
136.pdf.

Willey:1907:MMI

[10779] Day Allen Willey. Magnified models of insect pests. Scientific American, 97(8):137–138, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
137.pdf.

Boyer:1907:EHF

[10780] Jacques Boyer. Employment of high-frequency currents in therapeutics. Scientific American, 97(8):138–139, August 24, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n8/pdf/scientificamerican08241907-138.pdf.

Anonymous:1907:ISR

[10781] Anonymous. An improvement in safety razors. Scientific American, 97(8):139, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
139.pdf.

Anonymous:1907:NVS

[10782] Anonymous. A novel vehicle spring. Scientific American, 97(8):139, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n8/pdf/scientificamerican08241907-139a.pdf.

Anonymous:1907:IIBy

[10783] Anonymous. An improved invalid’s bed. Scientific American, 97(8):139, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
139b.pdf.

Anonymous:1907:BST

[10784] Anonymous. Barrel stand and truck. Scientific American, 97 (8):139, August 24, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n8/pdf/scientificamerican08241907-
139c.pdf.

Anonymous:1907:RPIbx

[10785] Anonymous. Recently patented inventions. Scientific American, 97(9):156, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
156.pdf.

Anonymous:1907:BNBbf

[10786] Anonymous. Books: New books, etc. Scientific American, 97 (9):156, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
156a.pdf.

Anonymous:1907:IIbz

[10787] Anonymous. Index of inventions. Scientific American, 97(9): 156–159, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
156b.pdf.

Anonymous:1907:PNPbx

[10788] Anonymous. Patent notes: Patents. Scientific American, 97(9):159, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n9/pdf/scientificamerican08311907-159.pdf.

Anonymous:1907:CRT

[10789] Anonymous. Combined reciprocating and turbine engines. Scientific American, 97(9):146, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
146.pdf.

Anonymous:1907:USB

[10790] Anonymous. United States and British gunnery. Scientific American, 97(9):146, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
146a.pdf.

Anonymous:1907:HCS

[10791] Anonymous. Health conditions in the subway. Scientific American, 97(9):146, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
146b.pdf.

Anonymous:1907:FFC

[10792] Anonymous. Fulton the father of commercial steamboat navigation. Scientific American, 97(9):146, August 31, 1907. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v97/
n9/pdf/scientificamerican08311907-146c.pdf.

Russell:1907:HS

[10793] Henry Norris Russell. The heavens in September. Scientific American, 97(9):147, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
147.pdf.

Anonymous:1907:CSAb

[10794] Anonymous. The competition for the Scientific American Flying Machine Trophy. Scientific American, 97(9):147, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n9/pdf/scientificamerican08311907-147a.pdf.

Anonymous:1907:BAR

[10795] Anonymous. “black art” redivivus. Scientific American, 97(9):148, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n9/pdf/scientificamerican08311907-148.pdf.

Collins:1907:SL

[10796] Percy Collins. Striking a light. Scientific American, 97(9):149–150, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v97/n9/pdf/scientificamerican08311907-149.pdf.

Anonymous:1907:RAK

[10797] Anonymous. The races for the Astor and King’s Cup. Scientific American, 97(9):151, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
151.pdf.

Kuntz:1907:SNG

[10798] F. C. Kuntz. Some notable German bridges. Scientific American, 97(9):151–154, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
151a.pdf.

Marriner:1907:SKK

[10799] George R. Marriner. The sheep-killing kea. Scientific American, 97(9):154, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
154.pdf.

Anonymous:1907:VPM

[10800] Anonymous. Variable pressure mechanism for air brakes. Scientific American, 97(9):155, August 31, 1907. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v97/n9/pdf/scientificamerican08311907-
155.pdf.

Russell:1908:HJa

[10801] Henry Norris Russell. The heavens in January. Scientific American, 98(1):9, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
9.pdf.

Anonymous:1908:RPIa

[10802] Anonymous. Recently patented inventions. Scientific American, 98(1):15–16, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
15.pdf.

Anonymous:1908:NQa

[10803] Anonymous. Notes and queries. Scientific American, 98(1):16, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-16.pdf.

Anonymous:1908:BNBa

[10804] Anonymous. Books: New books, etc. Scientific American, 98(1):16–17, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-16a.pdf.

Anonymous:1908:IIa

[10805] Anonymous. Index of inventions. Scientific American, 98(1):17–19, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-17.pdf.

Anonymous:1908:PNPa

[10806] Anonymous. Patent notes: Patents. Scientific American, 98(1):20, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-20.pdf.

Anonymous:1908:RY

[10807] Anonymous. Retrospect of the year 1907. Scientific American, 98(1):4–5, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
4.pdf.

Willey:1908:NCP

[10808] Day Allen Willey. A North Carolina power dam. Scientific American, 98(1):6, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
6.pdf.

Anonymous:1908:SLB

[10809] Anonymous. Signal lights for breeches buoys. Scientific American, 98(1):7, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
7.pdf.

Anonymous:1908:FWAa

[10810] Anonymous. The fastest warships afloat. Scientific American, 98(1):10, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-10.pdf.

Howell:1908:TPO

[10811] J. F. Howell. Target practice — one element of our national coast defense. Scientific American, 98(1):10–12, January 4, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n1/pdf/scientificamerican01041908-10a.pdf.

Allen:1908:MWT

[10812] William Allen. A motor for winter traction. Scientific American, 98(1):12, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
12.pdf.

Anonymous:1908:DSW

[10813] Anonymous. Device for stretching wire. Scientific American, 98(1):13, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-13.pdf.

Anonymous:1908:NAA

[10814] Anonymous. Novel attachment for agricultural implements. Scientific American, 98(1):13, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
13a.pdf.

Anonymous:1908:SPH

[10815] Anonymous. Safety plate for horseshoes. Scientific American, 98(1):13, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-13b.pdf.

Anonymous:1908:CAT

[10816] Anonymous. Covering for automobile tires. Scientific American, 98(1):13, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
13c.pdf.

Anonymous:1908:MDRa

[10817] Anonymous. Mail deliverer and receiver. Scientific American, 98(1):13–14, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n1/pdf/scientificamerican01041908-
13d.pdf.

Anonymous:1908:LVS

[10818] Anonymous. Lock for vehicle seats. Scientific American, 98(1):14, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-14.pdf.

Anonymous:1908:IMG

[10819] Anonymous. An improved marking gage. Scientific American, 98(1):14, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-14a.pdf.

Anonymous:1908:OIa

[10820] Anonymous. Oddities in invention. Scientific American, 98(1):14, January 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n1/pdf/scientificamerican01041908-14b.pdf.

Russell:1908:HMa

[10821] Henry Norris Russell. The heavens in March. Scientific American, 98(10):163, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
163.pdf.

Anonymous:1908:Ca

[10822] Anonymous. Correspondence. Scientific American, 98(10):167, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n10/pdf/scientificamerican03071908-167.pdf.

Anonymous:1908:IIb

[10823] Anonymous. Index of inventions. Scientific American, 98(10):172–175, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n10/pdf/scientificamerican03071908-172.pdf.

Anonymous:1908:PNPb

[10824] Anonymous. Patent notes: Patents. Scientific American, 98(10):175, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n10/pdf/scientificamerican03071908-175.pdf.

Anonymous:1908:HTT

[10825] Anonymous. Heavy traffic through the newly-opened Hudson Tunnel. Scientific American, 98(10):162, March 7, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n10/pdf/scientificamerican03071908-162.pdf.

Anonymous:1908:DDC

[10826] Anonymous. Double-door cars for subway congestion. Scientific American, 98(10):162, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
162a.pdf.

Anonymous:1908:UNC

[10827] Anonymous. Urgent need for the Catskill Water Supply. Scientific American, 98(10):162, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
162b.pdf.

Anonymous:1908:APO

[10828] Anonymous. The alleged Patent Office frauds. Scientific American, 98(10):162, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
162c.pdf.

Anonymous:1908:SBBa

[10829] Anonymous. Sand building-bricks. Scientific American, 98(10):163, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n10/pdf/scientificamerican03071908-163a.pdf.

Anonymous:1908:ERF

[10830] Anonymous. Experiments with recent French aeroplanes. Scientific American, 98(10):164–165, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
164.pdf.

Anonymous:1908:PCW

[10831] Anonymous. Progress of the Catskill Water Supply. Scientific American, 98(10):165–166, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
165.pdf.

Anonymous:1908:IBN

[10832] Anonymous. Ii — the British Navy of to-day. Scientific American, 98(10):167–170, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
167a.pdf.

Kenney:1908:SWT

[10833] Frank J. Kenney. A simple wireless telegraph detector. Scientific American, 98(10):170, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
170.pdf.

Anonymous:1908:PDa

[10834] Anonymous. Patent department. Scientific American, 98(10):171, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n10/pdf/scientificamerican03071908-171.pdf.

Anonymous:1908:SAM

[10835] Anonymous. Stamp-affixing machine. Scientific American, 98(10):171, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n10/pdf/scientificamerican03071908-171a.pdf.

Anonymous:1908:IPP

[10836] Anonymous. Improved plane provided with adjustable guides. Scientific American, 98(10):171, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
171b.pdf.

Anonymous:1908:IWM

[10837] Anonymous. An improved washing machine. Scientific American, 98(10):171, March 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n10/pdf/scientificamerican03071908-
171c.pdf.

Anonymous:1908:Cb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-183.pdf.

Anonymous:1908:RPIb

[10839] Anonymous. Recently patented inventions. Scientific American, 98 (11):189–191, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
189.pdf.

Anonymous:1908:NQb

[10840] Anonymous. Notes and queries. Scientific American, 98(11):191–192, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-191.pdf.

Anonymous:1908:BNBb

[10841] Anonymous. Books: New books etc. Scientific American, 98(11):192, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-192.pdf.

Anonymous:1908:IIc

[10842] Anonymous. Index of inventions. Scientific American, 98(11):192–195, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-192a.pdf.

Anonymous:1908:PNPc

[10843] Anonymous. Patent notes: Patents. Scientific American, 98(11):195, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-195.pdf.

Anonymous:1908:CQB

[10844] Anonymous. Cause of the Quebec Bridge Failure. Scientific American, 98(11):178, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
178.pdf.

Anonymous:1908:SPE

[10845] Anonymous. Second Pennsylvania East River Tube finished. Scientific American, 98(11):178, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
178a.pdf.

Anonymous:1908:EJP

[10846] Anonymous. Extemporized jet propulsion. Scientific American, 98(11):178, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
178b.pdf.

Anonymous:1908:OMB

[10847] Anonymous. The Ocean Mail Bill. Scientific American, 98(11):178, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-178c.pdf.

Anonymous:1908:CHT

[10848] Anonymous. The crookedness of the Hudson Tunnel. Scientific American, 98(11):178, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
178d.pdf.

Anonymous:1908:SVB

[10849] Anonymous. Submarines versus battleships. Scientific American, 98(11):178–179, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
178e.pdf.

Anonymous:1908:ILS

[10850] Anonymous. An International Life-Saving Congress. Scientific American, 98(11):179, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
179.pdf.

Anonymous:1908:RPFa

[10851] Anonymous. Radiography in pearl fishing. Scientific American, 98(11):179, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
179a.pdf.

Anonymous:1908:OMSa

[10852] Anonymous. Official Meteorological Summary, New York, N.Y., February, 1908. Scientific American, 98(11):179, March 14, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n11/pdf/scientificamerican03141908-179b.pdf.

Anonymous:1908:CSa

[10853] Anonymous. The current supplement. Scientific American, 98(11):179, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-179c.pdf.

Anonymous:1908:NP

[10854] Anonymous. A new planet. Scientific American, 98(11):179, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-179d.pdf.

Anonymous:1908:MIB

[10855] Anonymous. A motor ice boat. Scientific American, 98(11):180, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n11/pdf/scientificamerican03141908-180.pdf.

Anonymous:1908:FSA

[10856] Anonymous. The fastest ship in the American Navy. Scientific American, 98(11):180, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
180a.pdf.

Wade:1908:PSS

[10857] Herbert T. Wade. Poetry and speech subjects of scientific inquiry. Scientific American, 98(11):181–182, March 14, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n11/pdf/scientificamerican03141908-181.pdf.

Boyer:1908:HSA

[10858] Jacques Boyer. A hospital for small animals. Scientific American, 98(11):184, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
184.pdf.

Claudy:1908:OEM

[10859] C. H. Claudy. The Ourdan engraving machine. Scientific American, 98(11):185–188, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
185.pdf.

Macdonald:1908:USM

[10860] Owen Macdonald. The United States Military Academy at West Point — where the American officer is trained. Scientific American, 98(11):188–189, March 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n11/pdf/scientificamerican03141908-
188.pdf.

Anonymous:1908:Cc

[10861] Anonymous. Correspondence. Scientific American, 98(12):203, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-203.pdf.

Anonymous:1908:RPIc

[10862] Anonymous. Recently patented inventions. Scientific American, 98(12):209, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
209.pdf.

Anonymous:1908:NQc

[10863] Anonymous. Notes and queries. Scientific American, 98(12):209–210, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-209a.pdf.

Anonymous:1908:BNBc

[10864] Anonymous. Books: New books, etc. Scientific American, 98 (12):210–211, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
210.pdf.

Anonymous:1908:IId

[10865] Anonymous. Index of inventions. Scientific American, 98(12):211–215, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-211.pdf.

Anonymous:1908:PNPd

[10866] Anonymous. Patent notes: Patents. Scientific American, 98(12):215, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-215.pdf.

Anonymous:1908:JBIa

[10867] Anonymous. Joining the Blackwell’s Island Bridge cantilevers. Scientific American, 98(12):198, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
198.pdf.

Anonymous:1908:NTR

[10868] Anonymous. New transatlantic records. Scientific American, 98 (12):198, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
198a.pdf.

Anonymous:1908:NERa

[10869] Anonymous. A new explanation of raised ocean beaches. Scientific American, 98(12):198, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
198b.pdf.

Anonymous:1908:CSC

[10870] Anonymous. Competition in submarine construction. Scientific American, 98(12):198–199, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
198c.pdf.

Anonymous:1908:RCS

[10871] Anonymous. Reinforced concrete in shipbuilding. Scientific American, 98(12):199, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
199.pdf.

Anonymous:1908:HTA

[10872] Anonymous. How to take advantage of variations in rates of gas and electricity. Scientific American, 98(12):199, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-199a.pdf.

Anonymous:1908:UCM

[10873] Anonymous. The use of the cinematograph in medicine. Scientific American, 98(12):199, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
199b.pdf.

Anonymous:1908:CSb

[10874] Anonymous. The current supplement. Scientific American, 98(12):199, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-199c.pdf.

Anonymous:1908:DJB

[10875] Anonymous. Death of John Burry. Scientific American, 98(12):199, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-199d.pdf.

Anonymous:1908:FST

[10876] Anonymous. The first successful trial of a new American aeroplane. Scientific American, 98(12):200, March 21, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n12/pdf/scientificamerican03211908-200.pdf.

Mclure:1908:LFW

[10877] W. Frank M’clure. The largest fresh-water dock in the world. Scientific American, 98(12):201, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
201.pdf.

Mitchell:1908:WVM

[10878] S. A. Mitchell. Water vapor on Mars. Scientific American, 98 (12):201–202, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
201a.pdf.

Springer:1908:FDS

[10879] J. Springer. The fourth dimension simply explained. Scientific American, 98(12):202–203, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
202.pdf.

Anonymous:1908:USI

[10880] Anonymous. A unique system of illumination for harbor channels. Scientific American, 98(12):203, March 21, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n12/pdf/scientificamerican03211908-203a.pdf.

Gradenwitz:1908:MRG

[10881] Alfred Gradenwitz. The mutual relations of geometry and mechanics and Prof. Reuleaux’s mechanical movements. Scientific American, 98(12):204–205, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
204.pdf.

Michaud:1908:PTE

[10882] Gustave Michaud. Photographs with tapestry effects. Scientific American, 98(12):206, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
206.pdf.

Anonymous:1908:IRSa

[10883] Anonymous. An improved railway spike. Scientific American, 98(12):207, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
207.pdf.

Anonymous:1908:CCCa

[10884] Anonymous. Combined chair and cabinet. Scientific American, 98(12):207, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
207a.pdf.

Anonymous:1908:AAS

[10885] Anonymous. Ax attachment for sawing logs. Scientific American, 98(12):207, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n12/pdf/scientificamerican03211908-
207b.pdf.

Anonymous:1908:CSH

[10886] Anonymous. Combined steam and hot-water heating apparatus. Scientific American, 98(12):207, March 21, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n12/pdf/scientificamerican03211908-207c.pdf.

Anonymous:1908:BP

[10887] Anonymous. Bell piano. Scientific American, 98(12):207–208, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-207d.pdf.

Anonymous:1908:OIb

[10888] Anonymous. Oddities in inventions. Scientific American, 98(12):208, March 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n12/pdf/scientificamerican03211908-208.pdf.

Russell:1908:HAa

[10889] Henry Norris Russell. The heavens in April. Scientific American, 98(13):219, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
219.pdf.

Anonymous:1908:RPId

[10890] Anonymous. Recently patented inventions. Scientific American, 98(13):229, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
229.pdf.

Anonymous:1908:NQd

[10891] Anonymous. Notes and queries. Scientific American, 98(13):229–230, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n13/pdf/scientificamerican03281908-229a.pdf.

Anonymous:1908:BNBd

[10892] Anonymous. Books: New books, etc. Scientific American, 98(13):230, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n13/pdf/scientificamerican03281908-230.pdf.

Anonymous:1908:IIe

[10893] Anonymous. Index of inventions. Scientific American, 98(13):230–235, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n13/pdf/scientificamerican03281908-230a.pdf.

Anonymous:1908:PNPe

[10894] Anonymous. Patent notes: Patents. Scientific American, 98(13):235, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n13/pdf/scientificamerican03281908-235.pdf.

Anonymous:1908:ALS

[10895] Anonymous. Acetylene lamps and smoky automobiles. Scientific American, 98(13):218, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
218.pdf.

Anonymous:1908:CTO

[10896] Anonymous. The comparative trials of our Scout cruisers. Scientific American, 98(13):218, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
218a.pdf.

Anonymous:1908:NYH

[10897] Anonymous. The New York High-Pressure Fire Protection System. Scientific American, 98(13):218, March 28, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n13/pdf/scientificamerican03281908-218b.pdf.

Christadoro:1908:GBC

[10898] Charles Christadoro. “Gluten” bread and cracker frauds. Scientific American, 98(13):219, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
219a.pdf.

Anonymous:1908:RS

[10899] Anonymous. Rubber statistics of 1907. Scientific American, 98(13):219, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n13/pdf/scientificamerican03281908-219b.pdf.

Anonymous:1908:RCE

[10900] Anonymous. A reinforced concrete elevated tank. Scientific American, 98(13):220, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
220.pdf.

Anonymous:1908:NPMa

[10901] Anonymous. New process for making shoes. Scientific American, 98(13):220, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
220a.pdf.

Anonymous:1908:AMSa

[10902] Anonymous. An automatic milk supply. Scientific American, 98(13):220, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
220b.pdf.

Michaud:1908:SEC

[10903] Gustave Michaud. Some experiments with carbon disulphide. Scientific American, 98(13):221, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
221.pdf.

Anonymous:1908:OPU

[10904] Anonymous. An observatory for public use. Scientific American, 98(13):221–222, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
221a.pdf.

Anonymous:1908:CFAa

[10905] Anonymous. Curious facts about numbers. Scientific American, 98 (13):222–223, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
222.pdf.

Anonymous:1908:JLS

[10906] Anonymous. Joining of last span of Blackwell’s Island Bridge. Scientific American, 98(13):223, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n13/pdf/scientificamerican03281908-
223.pdf.

Anonymous:1908:HCB

[10907] Anonymous. How to construct a balloon. — the making, inflating and sailing of gas balloons. Scientific American, 98(13):224–227, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n13/pdf/scientificamerican03281908-224.pdf.

Springer:1908:ESE

[10908] J. F. Springer. The ever-shifting eyeball. — some new experiments and some new conclusions. Scientific American, 98(13):227–228, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n13/pdf/scientificamerican03281908-227.pdf.

Anonymous:1908:NTL

[10909] Anonymous. New types of lifebelts. Scientific American, 98(13):228, March 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n13/pdf/scientificamerican03281908-228.pdf.

Anonymous:1908:Cd

[10910] Anonymous. Correspondence. Scientific American, 98(14):243, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-243a.pdf.

Anonymous:1908:RPIe

[10911] Anonymous. Recently patented inventions. Scientific American, 98(14):248, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n14/pdf/scientificamerican04041908-
248.pdf.

Anonymous:1908:IIf

[10912] Anonymous. Index of inventions. Scientific American, 98(14):248–251, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-248a.pdf.

Anonymous:1908:PNPf

[10913] Anonymous. Patent notes: Patents. Scientific American, 98(14):251, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-251.pdf.

Anonymous:1908:PTD

[10914] Anonymous. The peril of the turnstile door. Scientific American, 98(14):238, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n14/pdf/scientificamerican04041908-
238.pdf.

Anonymous:1908:SHA

[10915] Anonymous. The Senate Hearing on the alleged defects of our Navy. Scientific American, 98(14):238, April 4, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n14/pdf/scientificamerican04041908-238a.pdf.

Anonymous:1908:GADa

[10916] Anonymous. Government aid in the development of aeronautics. Scientific American, 98(14):238–239, April 4, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n14/pdf/scientificamerican04041908-238b.pdf.

Horwill:1908:SP

[10917] Herbert W. Horwill. Scientists at play. Scientific American, 98(14):239, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-239.pdf.

Anonymous:1908:DPA

[10918] Anonymous. Decolorizing power of animal charcoal. Scientific American, 98(14):239, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n14/pdf/scientificamerican04041908-
239a.pdf.

Anonymous:1908:CSd

[10919] Anonymous. The current supplement. Scientific American, 98(14):239, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-239b.pdf.

Wilson:1908:LTT

[10920] L. F. Wilson. Lowering the traction tunnels tinder the Chicago River. Scientific American, 98(14):240, April 4, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n14/pdf/scientificamerican04041908-240.pdf.

Anonymous:1908:JBIb

[10921] Anonymous. Japanese battleship “Iwami” as she was and is — a lesson in “Freeboard”. Scientific American, 98(14):241–242, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-241.pdf.

Anonymous:1908:LMS

[10922] Anonymous. Lippmann’s method of stereophotography without a lens. Scientific American, 98(14):242, April 4, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n14/pdf/scientificamerican04041908-242.pdf.

Anonymous:1908:NAS

[10923] Anonymous. A new apparatus for submarine operations. Scientific American, 98(14):243, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n14/pdf/scientificamerican04041908-
243.pdf.

Boyer:1908:MMO

[10924] Jacques Boyer. The manufacture of mechanical organs. Scientific American, 98(14):244–245, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n14/pdf/scientificamerican04041908-
244.pdf.

Mitchell:1908:SR

[10925] S. A. Mitchell. Saturn’s rings. Scientific American, 98(14):246, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-246.pdf.

Anonymous:1908:PDb

[10926] Anonymous. Patent department. Scientific American, 98(14):247, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-247.pdf.

Anonymous:1908:GEI

[10927] Anonymous. Gas engine igniter. Scientific American, 98(14):247, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-247a.pdf.

Anonymous:1908:CTP

[10928] Anonymous. Casting tool for printing offices. Scientific American, 98(14):247, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n14/pdf/scientificamerican04041908-
247b.pdf.

Anonymous:1908:ISL

[10929] Anonymous. Improved sash lock. Scientific American, 98(14):247, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n14/pdf/scientificamerican04041908-247c.pdf.

Anonymous:1908:IGT

[10930] Anonymous. An improved gasoline trap. Scientific American, 98 (14):247, April 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n14/pdf/scientificamerican04041908-
247d.pdf.

Anonymous:1908:Ce

[10931] Anonymous. Correspondence. Scientific American, 98(15):259, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n15/pdf/scientificamerican04111908-259.pdf.

Anonymous:1908:RPIf

[10932] Anonymous. Recently patented inventions. Scientific American, 98(15):264, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
264.pdf.

Anonymous:1908:IIg

[10933] Anonymous. Index of inventions. Scientific American, 98(15):264–267, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n15/pdf/scientificamerican04111908-264a.pdf.

Anonymous:1908:PNPg

[10934] Anonymous. Patent notes: Patents. Scientific American, 98(15):267, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n15/pdf/scientificamerican04111908-267.pdf.

Anonymous:1908:BAR

[10935] Anonymous. British and American railway accidents. Scientific American, 98(15):254, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
254.pdf.

Anonymous:1908:ERB

[10936] Anonymous. East River Bridge number five. Scientific American, 98(15):254, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
254a.pdf.

Anonymous:1908:GEE

[10937] Anonymous. Great economy of electric traction on steam railroads. Scientific American, 98(15):254, April 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n15/pdf/scientificamerican04111908-254b.pdf.

Anonymous:1908:OFB

[10938] Anonymous. Oil fuel in the British Navy. Scientific American, 98(15):254, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
254c.pdf.

Anonymous:1908:PGDa

[10939] Anonymous. A producer-gas-driven automobile. Scientific American, 98(15):255, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
255.pdf.

Moultrie:1908:LB

[10940] J. W. Moultrie. Lime beton. Scientific American, 98(15):255, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n15/pdf/scientificamerican04111908-255a.pdf.

Anonymous:1908:LIA

[10941] Anonymous. The latest international aeronautic prize. Scientific American, 98(15):255, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
255b.pdf.

Anonymous:1908:EMC

[10942] Anonymous. Extermination of mosquitoes by cactus paste. Scientific American, 98(15):255, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
255c.pdf.

Anonymous:1908:FTP

[10943] Anonymous. The first two-passenger aeroplane. Scientific American, 98(15):256, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
256.pdf.

Anonymous:1908:PNU

[10944] Anonymous. A physician’s novel use of an automobile acetylene light. Scientific American, 98(15):256, April 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n15/pdf/scientificamerican04111908-256a.pdf.

Anonymous:1908:SDSa

[10945] Anonymous. A salvage dock for submarines. Scientific American, 98(15):257, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
257.pdf.

Boyer:1908:NRR

[10946] Jacques Boyer. The new refuse receptacles of Paris. — a suggestion for American municipalities. Scientific American, 98(15):257–258, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n15/pdf/scientificamerican04111908-257a.pdf.

Anonymous:1908:HHM

[10947] Anonymous. The Henry Hudson Memorial Bridge. Scientific American, 98(15):258, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
258.pdf.

Berry:1908:RWL

[10948] Riley M. Fletcher Berry. The recent work of Luther Burbank. Scientific American, 98(15):260–262, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
260.pdf.

Anonymous:1908:UBP

[10949] Anonymous. The use of balloons for the prevention of hail. Scientific American, 98(15):262, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
262.pdf.

Brown:1908:TDB

[10950] Barnum Brown. Trachodont, the duck-billed dinosaur. — skeletons of prehistoric reptiles more than three million years old. Scientific American, 98(15):262–263, April 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n15/pdf/scientificamerican04111908-
262a.pdf.

Anonymous:1908:Cf

[10951] Anonymous. Correspondence. Scientific American, 98(16):275, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-275.pdf.

Anonymous:1908:RPIg

[10952] Anonymous. Recently patented inventions. Scientific American, 98(16):281–282, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
281.pdf.

Anonymous:1908:NQe

[10953] Anonymous. Notes and queries. Scientific American, 98(16):282, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-282.pdf.

Anonymous:1908:BNBe

[10954] Anonymous. Books: New books, etc. Scientific American, 98(16):283, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-283.pdf.

Anonymous:1908:IIh

[10955] Anonymous. Index of inventions. Scientific American, 98(16):283–287, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-283a.pdf.

Anonymous:1908:PNPh

[10956] Anonymous. Patent notes: Patents. Scientific American, 98(16):287, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-287.pdf.

Anonymous:1908:HCP

[10957] Anonymous. Healthy conditions at Panama in 1907. Scientific American, 98(16):270, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
270.pdf.

Anonymous:1908:SBBb

[10958] Anonymous. Safety of the Brooklyn Bridge. Scientific American, 98(16):270, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
270a.pdf.

Anonymous:1908:DDS

[10959] Anonymous. Double-deck subway express stations. Scientific American, 98(16):270, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
270b.pdf.

Anonymous:1908:HMH

[10960] Anonymous. Heroic measures in Hudson Tunnel construction. Scientific American, 98(16):270, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
270c.pdf.

Anonymous:1908:SDR

[10961] Anonymous. Safety devices on railroads. Scientific American, 98 (16):270–271, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
270d.pdf.

Anonymous:1908:ASAa

[10962] Anonymous. The aviation section of the aero club of America. Scientific American, 98(16):271, April 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n16/pdf/scientificamerican04181908-271.pdf.

Anonymous:1908:NYA

[10963] Anonymous. New York’s Automobile Carnival and Hill Climb. Scientific American, 98(16):271, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
271a.pdf.

Wilson:1908:FAU

[10964] L. F. Wilson. Facts about the use of the air brake. Scientific American, 98(16):271, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
271b.pdf.

Anonymous:1908:SES

[10965] Anonymous. Second Exhibition of Safety Devices. Scientific American, 98(16):271, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
271c.pdf.

Anonymous:1908:CSe

[10966] Anonymous. The current supplement. Scientific American, 98(16):271, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-271d.pdf.

Anonymous:1908:SNa

[10967] Anonymous. Science notes. Scientific American, 98(16):271, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-271e.pdf.

Coleman:1908:LDB

[10968] F. C. Coleman. The latest development in British railway locomotive design. Scientific American, 98(16):272, April 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n16/pdf/scientificamerican04181908-272.pdf.

Place:1908:WVA

[10969] William H. Place. Weather vane with attachment for indoor readings. Scientific American, 98(16):272, April 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n16/pdf/scientificamerican04181908-272a.pdf.

Boyer:1908:MSW

[10970] Jacques Boyer. A machine that saws 240,000 wooden paving blocks in a day. Scientific American, 98(16):273–274, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-273.pdf.

Springer:1908:AES

[10971] J. F. Springer. Is the axis of the Earth shifting? Scientific American, 98(16):274–275, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
274.pdf.

Anonymous:1908:STC

[10972] Anonymous. Successful test of the Cornu helicopter. Scientific American, 98(16):276, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
276.pdf.

Anonymous:1908:USBa

[10973] Anonymous. The United States battleships “Delaware” and “North Dakota”. Scientific American, 98(16):276–278, April 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n16/pdf/scientificamerican04181908-276a.pdf.

Anonymous:1908:PCSa

[10974] Anonymous. Puck’s conception of the Scientific American. Scientific American, 98(16):278, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
278.pdf.

Anonymous:1908:PDc

[10975] Anonymous. Patent department. Scientific American, 98(16):279, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-279.pdf.

Anonymous:1908:TH

[10976] Anonymous. Tobacco hanger. Scientific American, 98(16):279, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-279a.pdf.

Anonymous:1908:TC

[10977] Anonymous. Thill coupling. Scientific American, 98(16):279, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-279b.pdf.

Anonymous:1908:RFL

[10978] Anonymous. Revolving flash lamp. Scientific American, 98(16):279, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n16/pdf/scientificamerican04181908-279c.pdf.

Anonymous:1908:PVM

[10979] Anonymous. Pneumatic valve mechanism for piano players. Scientific American, 98(16):279–280, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
279d.pdf.

Anonymous:1908:NTA

[10980] Anonymous. A new type of automobile horn. Scientific American, 98(16):280, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
280.pdf.

Anonymous:1908:MVP

[10981] Anonymous. A machine for vending postage stamps. Scientific American, 98(16):280, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
280a.pdf.

Anonymous:1908:FPF

[10982] Anonymous. Frying pan with folding handle. Scientific American, 98(16):280, April 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n16/pdf/scientificamerican04181908-
280b.pdf.

Anonymous:1908:Cg

[10983] Anonymous. Correspondence. Scientific American, 98(17):295, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n17/pdf/scientificamerican04251908-295.pdf.

Anonymous:1908:IIi

[10984] Anonymous. Index of inventions. Scientific American, 98(17):300–303, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n17/pdf/scientificamerican04251908-300.pdf.

Anonymous:1908:PNPi

[10985] Anonymous. Patent notes: Patents. Scientific American, 98(17):303, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n17/pdf/scientificamerican04251908-303.pdf.

Anonymous:1908:RBSa

[10986] Anonymous. Reinforcement of the Brooklyn Subway Tunnels. Scientific American, 98(17):290, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
290.pdf.

Anonymous:1908:PR

[10987] Anonymous. The panic and the railroads. Scientific American, 98(17):290, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
290a.pdf.

Anonymous:1908:RQBa

[10988] Anonymous. Report on the Quebec Bridge Disaster. Scientific American, 98(17):290, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
290b.pdf.

Anonymous:1908:UNF

[10989] Anonymous. Urgent need of forest preservation. Scientific American, 98(17):290, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
290c.pdf.

Anonymous:1908:BNP

[10990] Anonymous. The British Naval Programme for 1908. Scientific American, 98(17):290–291, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
290d.pdf.

Anonymous:1908:DRA

[10991] Anonymous. Delagrange’s record aeroplane flight. Scientific American, 98(17):291, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
291.pdf.

Munger:1908:IAM

[10992] William P. Munger. Improvement of the American merchant marine. Scientific American, 98(17):291, April 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n17/pdf/scientificamerican04251908-291a.pdf.

Anonymous:1908:NPMb

[10993] Anonymous. A new process in milling. Scientific American, 98(17):291, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n17/pdf/scientificamerican04251908-291b.pdf.

Anonymous:1908:CMTa

[10994] Anonymous. Combined mutoscope and talking machine. Scientific American, 98(17):292, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
292.pdf.

Michaud:1908:SAG

[10995] Gustave Michaud. A simple apparatus for generating hydrogen. Scientific American, 98(17):292, April 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n17/pdf/scientificamerican04251908-292a.pdf.

Anonymous:1908:CMTb

[10996] Anonymous. Convertible military transport motor wagon. Scientific American, 98(17):292, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
292b.pdf.

Newton:1908:CSW

[10997] R. Newton. A clock with a single wheel. Scientific American, 98(17):293, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n17/pdf/scientificamerican04251908-293.pdf.

Gradenwitz:1908:EGO

[10998] Alfred Gradenwitz. The explorations of the German Orient Society. Scientific American, 98(17):293–294, April 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n17/pdf/scientificamerican04251908-293a.pdf.

Willey:1908:PWU

[10999] Day Allen Willey. The practical work of the United States Bureau of Forestry. Scientific American, 98(17):295–296, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n17/pdf/scientificamerican04251908-295a.pdf.

Anonymous:1908:UST

[11000] Anonymous. The United States training ship “Severn”. Scientific American, 98(17):297, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
297.pdf.

Honey:1908:SPP

[11001] Frederic R. Honey. Some phases of the planets in 1908. Scientific American, 98(17):298, April 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n17/pdf/scientificamerican04251908-
298.pdf.

Anonymous:1908:ATJ

[11002] Anonymous. American turbines for the Japanese armored cruiser “Ibuki”. Scientific American, 98(17):299, April 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n17/pdf/scientificamerican04251908-299.pdf.

Russell:1908:HMb

[11003] Henry Norris Russell. The heavens in May. Scientific American, 98(18):307, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
307.pdf.

Anonymous:1908:Ch

[11004] Anonymous. Correspondence. Scientific American, 98(18):311, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n18/pdf/scientificamerican05021908-311.pdf.

Anonymous:1908:RPIh

[11005] Anonymous. Recently patented inventions. Scientific American, 98(18):317–318, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
317.pdf.

Anonymous:1908:NQf

[11006] Anonymous. Notes and queries. Scientific American, 98(18):318, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n18/pdf/scientificamerican05021908-318.pdf.

Anonymous:1908:BNBf

[11007] Anonymous. Books: New books, etc. Scientific American, 98(18):319, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n18/pdf/scientificamerican05021908-319.pdf.

Anonymous:1908:IIj

[11008] Anonymous. Index of inventions. Scientific American, 98(18):319–323, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n18/pdf/scientificamerican05021908-319a.pdf.

Anonymous:1908:PNPj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n18/pdf/scientificamerican05021908-324.pdf.

Anonymous:1908:SWW

[11010] Anonymous. Sir William White on the American Navy. Scientific American, 98(18):306, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
306.pdf.

Anonymous:1908:SSA

[11011] Anonymous. The Senate, the Scientific American, and the New York Sun. Scientific American, 98(18):306–307, May 2, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n18/pdf/scientificamerican05021908-306a.pdf.

Anonymous:1908:EFC

[11012] Anonymous. Erecting the first cored monolith. Scientific American, 98(18):307, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
307a.pdf.

Anonymous:1908:SCR

[11013] Anonymous. The stock car race for the Briarcliff Trophy. Scientific American, 98(18):308–309, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
308.pdf.

Michaud:1908:HSE

[11014] Gustave Michaud. How stereoscopic effects may be produced with a pinhole diaphragm. Scientific American, 98(18):309–310, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n18/pdf/scientificamerican05021908-309.pdf.

Anonymous:1908:CFH

[11015] Anonymous. A Campanile 700 feet high. Scientific American, 98(18):310, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
310.pdf.

Gradenwitz:1908:BOM

[11016] Alfred Gradenwitz. The Berlin Oceanographic Museum. Scientific American, 98(18):312–314, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
312.pdf.

Anonymous:1908:FLG

[11017] Anonymous. Fireless locomotives in Germany. Scientific American, 98(18):314, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
314.pdf.

Anonymous:1908:LDS

[11018] Anonymous. Locking device for screws. Scientific American, 98(18):315, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n18/pdf/scientificamerican05021908-315.pdf.

Anonymous:1908:GAC

[11019] Anonymous. Grinding attachment for check valves. Scientific American, 98(18):315, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
315a.pdf.

Anonymous:1908:CW

[11020] Anonymous. A camp on wheels. Scientific American, 98(18):315, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n18/pdf/scientificamerican05021908-315b.pdf.

Anonymous:1908:ATC

[11021] Anonymous. An attachment for trolley car poles. Scientific American, 98(18):315, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
315c.pdf.

Anonymous:1908:OIc

[11022] Anonymous. Oddities in inventions. Scientific American, 98 (18):315–316, May 2, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n18/pdf/scientificamerican05021908-
315d.pdf.

Anonymous:1908:Ci

[11023] Anonymous. Correspondence. Scientific American, 98(19):331, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n19/pdf/scientificamerican05091908-331.pdf.

Anonymous:1908:RPIi

[11024] Anonymous. Recently patented inventions. Scientific American, 98(19):337, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
337.pdf.

Anonymous:1908:NQg

[11025] Anonymous. Notes and queries. Scientific American, 98(19):337–338, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n19/pdf/scientificamerican05091908-337a.pdf.

Anonymous:1908:BNBg

[11026] Anonymous. Books: New books, etc. Scientific American, 98 (19):338–339, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
338.pdf.

Anonymous:1908:IIk

[11027] Anonymous. Index of inventions. Scientific American, 98(19):339–343, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n19/pdf/scientificamerican05091908-339.pdf.

Anonymous:1908:PNPk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n19/pdf/scientificamerican05091908-343.pdf.

Anonymous:1908:CCW

[11029] Anonymous. Criticism of Catskill Water Supply scheme. Scientific American, 98(19):326, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
326.pdf.

Anonymous:1908:CCL

[11030] Anonymous. Coal consumption, of the “Lusitania”. Scientific American, 98(19):326, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
326a.pdf.

Anonymous:1908:SRTa

[11031] Anonymous. Spraying roads with tar. Scientific American, 98(19):326, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n19/pdf/scientificamerican05091908-326b.pdf.

Anonymous:1908:PSL

[11032] Anonymous. Phenomenal speed of latest torpedoes. Scientific American, 98(19):326, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
326c.pdf.

Anonymous:1908:GGA

[11033] Anonymous. Germany’s great armored cruiser. Scientific American, 98(19):326, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
326d.pdf.

Anonymous:1908:WSa

[11034] Anonymous. The world’s shipbuilding. Scientific American, 98(19):327, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n19/pdf/scientificamerican05091908-327.pdf.

Anonymous:1908:CVL

[11035] Anonymous. Collecting the venom from a lancehead snake. Scientific American, 98(19):327, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
327a.pdf.

Benjamin:1908:NAS

[11036] Marcus Benjamin. National Academy of Sciences. Scientific American, 98(19):327, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
327b.pdf.

Anonymous:1908:PF

[11037] Anonymous. For practical forestry. Scientific American, 98(19):327, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n19/pdf/scientificamerican05091908-327c.pdf.

Boyer:1908:CMO

[11038] Jacques Boyer. Cutting metals with oxygen. Scientific American, 98(19):328, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
328.pdf.

Wade:1908:GPM

[11039] Herbert T. Wade. Gas phosphorescence and meteor trains. Scientific American, 98(19):329–330, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
329.pdf.

Anonymous:1908:SBS

[11040] Anonymous. A submarine boat for sponge fishing. Scientific American, 98(19):330, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
330.pdf.

Anonymous:1908:CHa

[11041] Anonymous. Cheap hydrogen. Scientific American, 98(19):330, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n19/pdf/scientificamerican05091908-330a.pdf.

Anonymous:1908:SNG

[11042] Anonymous. Some notable groups of North American birds. Scientific American, 98(19):332–334, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
332.pdf.

Banks:1908:OSC

[11043] Edgar J. Banks. The origin of the Star and Crescent. Scientific American, 98(19):334, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
334.pdf.

Mclure:1908:NIL

[11044] W. Frank M’clure. A new inspection locomotive. Scientific American, 98(19):335, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
335.pdf.

Anonymous:1908:SSI

[11045] Anonymous. Shipping submarines intact to the Philippines. Scientific American, 98(19):335, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
335a.pdf.

Anonymous:1908:RFD

[11046] Anonymous. The record flight of the Delagrange aeroplane. Scientific American, 98(19):336, May 9, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n19/pdf/scientificamerican05091908-
336.pdf.

Anonymous:1908:Cj

[11047] Anonymous. Correspondence. Scientific American, 98(2):27, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n2/pdf/scientificamerican01111908-27.pdf.

Anonymous:1908:RPIj

[11048] Anonymous. Recently patented inventions. Scientific American, 98(2):32, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
32.pdf.

Anonymous:1908:NQh

[11049] Anonymous. Notes and queries. Scientific American, 98(2):32, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n2/pdf/scientificamerican01111908-32a.pdf.

Anonymous:1908:BNBh

[11050] Anonymous. Books: New books, etc. Scientific American, 98 (2):32, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
32b.pdf.

Anonymous:1908:IIl

[11051] Anonymous. Index of inventions. Scientific American, 98(2):32–35, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n2/pdf/scientificamerican01111908-32c.pdf.

Anonymous:1908:PNPl

[11052] Anonymous. Patent notes: Patents. Scientific American, 98(2):36, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n2/pdf/scientificamerican01111908-36.pdf.

Anonymous:1908:ICP

[11053] Anonymous. Important change of Panama Canal locks. Scientific American, 98(2):22, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
22.pdf.

Anonymous:1908:ODM

[11054] Anonymous. Our decadent merchant marine. Scientific American, 98(2):22, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
22a.pdf.

Anonymous:1908:PFM

[11055] Anonymous. Prevention of “flarebacks” in modern guns. Scientific American, 98(2):22, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
22b.pdf.

Anonymous:1908:WCO

[11056] Anonymous. Warship construction at our Navy yards. Scientific American, 98(2):22–23, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
22c.pdf.

Anonymous:1908:NCC

[11057] Anonymous. A new coin counter. Scientific American, 98(2):23, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n2/pdf/scientificamerican01111908-23.pdf.

Anonymous:1908:EWA

[11058] Anonymous. The economic waste of accidents. Scientific American, 98(2):23, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
23a.pdf.

Hutchinson:1908:HT

[11059] C. E. Hutchinson. Habits of the tarantula. Scientific American, 98(2):23, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
23b.pdf.

Anonymous:1908:OMSb

[11060] Anonymous. Official Meteorological Summary, New York, N.Y., December, 1907. Scientific American, 98(2):23, January 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n2/pdf/scientificamerican01111908-23c.pdf.

Anonymous:1908:CSf

[11061] Anonymous. The current supplement. Scientific American, 98 (2):23, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
23d.pdf.

Willey:1908:WF

[11062] Day Allen Willey. Winter fires. Scientific American, 98(2):24, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n2/pdf/scientificamerican01111908-24.pdf.

Anonymous:1908:AIF

[11063] Anonymous. An apparatus for indicating fire damp in coal mines. Scientific American, 98(2):25–26, January 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n2/pdf/scientificamerican01111908-25.pdf.

Fitz-Gerald:1908:RMT

[11064] W. G. Fitz-Gerald. A railroad in miniature — a toy line which cost. Scientific American, 98(2):28–29, January 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n2/pdf/scientificamerican01111908-28.pdf.

Smith:1908:MWR

[11065] Katherine Louise Smith. The most wonderful railroad of the North. Scientific American, 98(2):29–30, January 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n2/pdf/scientificamerican01111908-29.pdf.

Boyer:1908:TRA

[11066] Jacques Boyer. Tissot’s respiratory apparatus. Scientific American, 98(2):30–31, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
30.pdf.

Mitchell:1908:GSS

[11067] S. A. Mitchell. Great storms on the Sun. Scientific American, 98(2):31, January 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n2/pdf/scientificamerican01111908-
31.pdf.

Anonymous:1908:Ck

[11068] Anonymous. Correspondence. Scientific American, 98(20):351, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-351.pdf.

Anonymous:1908:RPIk

[11069] Anonymous. Recently patented inventions. Scientific American, 98(20):357, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
357.pdf.

Anonymous:1908:NQi

[11070] Anonymous. Notes and queries. Scientific American, 98(20):357, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-357a.pdf.

Anonymous:1908:BNBi

[11071] Anonymous. Books: New books, etc. Scientific American, 98(20):358, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-358.pdf.

Anonymous:1908:IIm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-358a.pdf.

Anonymous:1908:PNPm

[11073] Anonymous. Patent notes: Patents. Scientific American, 98(20):363, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-363.pdf.

Anonymous:1908:IGD

[11074] Anonymous. Investigations for the Gatun Dam. Scientific American, 98(20):346, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
346.pdf.

Anonymous:1908:SBR

[11075] Anonymous. To secure better rails. Scientific American, 98(20):346, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-346a.pdf.

Anonymous:1908:HF

[11076] Anonymous. A heat flywheel. Scientific American, 98(20):346, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-346b.pdf.

Anonymous:1908:FCE

[11077] Anonymous. Former Chief Engineer Stevens on the Canal. Scientific American, 98(20):346, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
346c.pdf.

Anonymous:1908:MTL

[11078] Anonymous. Metchnikoff’s theory of longevity. Scientific American, 98(20):347, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
347.pdf.

Anonymous:1908:CCS

[11079] Anonymous. Copper-clad steel. Scientific American, 98(20):347, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-347a.pdf.

Anonymous:1908:MMW

[11080] Anonymous. Machine for making window glass. Scientific American, 98(20):347, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
347b.pdf.

Anonymous:1908:CSg

[11081] Anonymous. The current supplement. Scientific American, 98(20):347, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-347c.pdf.

Anonymous:1908:MV

[11082] Anonymous. Motors for viticulture. Scientific American, 98(20):347, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-347d.pdf.

Anonymous:1908:WAE

[11083] Anonymous. The Wright aeroplane experiments. Scientific American, 98(20):347, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
347e.pdf.

Anonymous:1908:CTa

[11084] Anonymous. The “Caterpillar” tractor. Scientific American, 98 (20):348, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
348.pdf.

Anonymous:1908:RPFb

[11085] Anonymous. Range and position finders for coast artillery fire control. Scientific American, 98(20):349–350, May 16, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n20/pdf/scientificamerican05161908-349.pdf.

Honey:1908:AES

[11086] Frederic R. Honey. The annular eclipse of the Sun in June, 1908. Scientific American, 98(20):350, May 16, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n20/pdf/scientificamerican05161908-350.pdf.

Anonymous:1908:MBa

[11087] Anonymous. A magazine for the blind. Scientific American, 98 (20):351–352, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
351a.pdf.

Claudy:1908:HCA

[11088] C. H. Claudy. How cave air is used to regulate the temperature of a house. Scientific American, 98(20):353–354, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-353.pdf.

Anonymous:1908:NERb

[11089] Anonymous. A new electric renovator. Scientific American, 98(20):355, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-355.pdf.

Perkins:1908:FUS

[11090] Frank C. Perkins. A French universal system of machine molding. Scientific American, 98(20):355, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
355a.pdf.

Anonymous:1908:PDd

[11091] Anonymous. Patent department. Scientific American, 98(20):356, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-356.pdf.

Anonymous:1908:FWM

[11092] Anonymous. Folding washing machine for stationary tubs. Scientific American, 98(20):356, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
356a.pdf.

Anonymous:1908:HNM

[11093] Anonymous. Hose nozzle with movable spray cap. Scientific American, 98(20):356, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
356b.pdf.

Anonymous:1908:ISP

[11094] Anonymous. Improved switchboard plug. Scientific American, 98 (20):356, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n20/pdf/scientificamerican05161908-
356c.pdf.

Anonymous:1908:FSL

[11095] Anonymous. Folding step ladder. Scientific American, 98(20):356, May 16, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n20/pdf/scientificamerican05161908-356d.pdf.

Anonymous:1908:Cl

[11096] Anonymous. Correspondence. Scientific American, 98(21):371, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-371.pdf.

Anonymous:1908:RPIl

[11097] Anonymous. Recently patented inventions. Scientific American, 98(21):377, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
377.pdf.

Anonymous:1908:NQj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-377a.pdf.

Anonymous:1908:BNBj

[11099] Anonymous. Books: New books, etc. Scientific American, 98 (21):378–379, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
378.pdf.

Anonymous:1908:IIn

[11100] Anonymous. Index of inventions. Scientific American, 98(21):379–383, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-379.pdf.

Anonymous:1908:PNPn

[11101] Anonymous. Patent notes: Patents. Scientific American, 98(21):384, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-384.pdf.

Anonymous:1908:PRC

[11102] Anonymous. President Roosevelt on the conservation of our natural resources. Scientific American, 98(21):366, May 23, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n21/pdf/scientificamerican05231908-366.pdf.

Anonymous:1908:SBI

[11103] Anonymous. Safety of the Blackwell’s Island Bridge. Scientific American, 98(21):366, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
366a.pdf.

Anonymous:1908:CTb

[11104] Anonymous. A city of towers. Scientific American, 98(21):366–367, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-366b.pdf.

Anonymous:1908:PWA

[11105] Anonymous. Progress of the Wright aeroplane experiments. Scientific American, 98(21):367, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
367.pdf.

Anonymous:1908:PRB

[11106] Anonymous. Postal railway in Berlin. Scientific American, 98(21):367, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-367a.pdf.

Anonymous:1908:EPA

[11107] Anonymous. Experiments with a postal automobile in France. Scientific American, 98(21):367, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
367b.pdf.

Anonymous:1908:PPT

[11108] Anonymous. Protection for patents and trade marks of goods to be exhibited. Scientific American, 98(21):367, May 23, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n21/pdf/scientificamerican05231908-367c.pdf.

Anonymous:1908:CSi

[11109] Anonymous. The current supplement. Scientific American, 98(21):367, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-367d.pdf.

Anonymous:1908:SWM

[11110] Anonymous. Sending wireless messages to a balloon. Scientific American, 98(21):367, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
367e.pdf.

Boyer:1908:UAB

[11111] Jacques Boyer. The use of the air brush in the decoration of wall paper and fabrics. Scientific American, 98(21):368, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-368.pdf.

Anonymous:1908:MES

[11112] Anonymous. Maximum economical storage capacity of the Croton watershed. Scientific American, 98(21):369–370, May 23, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n21/pdf/scientificamerican05231908-369.pdf.

Skerrett:1908:CSW

[11113] R. G. Skerrett. Control by sound waves. Scientific American, 98(21):371, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
371a.pdf.

Willey:1908:FDE

[11114] Day Allen Willey. Forest destruction and the erosion of arable lands. Scientific American, 98(21):372–373, May 23, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n21/pdf/scientificamerican05231908-372.pdf.

Anonymous:1908:EVL

[11115] Anonymous. Extracting the venom from lancehead snake for medicinal use. Scientific American, 98(21):374, May 23, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n21/pdf/scientificamerican05231908-374.pdf.

Anonymous:1908:LCW

[11116] Anonymous. The largest clock in the world. Scientific American, 98(21):375–376, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
375.pdf.

Anonymous:1908:NTF

[11117] Anonymous. A new type of filter. Scientific American, 98(21):376, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n21/pdf/scientificamerican05231908-376.pdf.

Anonymous:1908:GLF

[11118] Anonymous. Gathering and loading fork. Scientific American, 98 (21):376, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
376a.pdf.

Anonymous:1908:CCCb

[11119] Anonymous. A computing cheese cutter. Scientific American, 98 (21):376, May 23, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n21/pdf/scientificamerican05231908-
376b.pdf.

Russell:1908:HJb

[11120] Henry Norris Russell. The heavens in June. Scientific American, 98(22):387, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
387.pdf.

Anonymous:1908:Cm

[11121] Anonymous. Correspondence. Scientific American, 98(22):391, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n22/pdf/scientificamerican05301908-391.pdf.

Anonymous:1908:RPIm

[11122] Anonymous. Recently patented inventions. Scientific American, 98(22):397, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
397.pdf.

Anonymous:1908:NQk

[11123] Anonymous. Notes and queries. Scientific American, 98(22):397–398, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n22/pdf/scientificamerican05301908-397a.pdf.

Anonymous:1908:BNBk

[11124] Anonymous. Books: New books, etc. Scientific American, 98 (22):398–399, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
398.pdf.

Anonymous:1908:IIo

[11125] Anonymous. Index of inventions. Scientific American, 98(22):399–403, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n22/pdf/scientificamerican05301908-399.pdf.

Anonymous:1908:PNPo

[11126] Anonymous. Patent notes: Patents. Scientific American, 98(22):404, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n22/pdf/scientificamerican05301908-404.pdf.

Anonymous:1908:GDR

[11127] Anonymous. Gratifying decrease in railway accidents. Scientific American, 98(22):386, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
386.pdf.

Anonymous:1908:MHO

[11128] Anonymous. Mr. Hill on our natural resources. Scientific American, 98(22):386, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
386a.pdf.

Anonymous:1908:LTR

[11129] Anonymous. Lowering the transatlantic record. Scientific American, 98(22):386, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
386b.pdf.

Anonymous:1908:ORS

[11130] Anonymous. Our rank as Second Naval Power. Scientific American, 98(22):386, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
386c.pdf.

Anonymous:1908:SFT

[11131] Anonymous. Severe fire tests for skyscrapers. Scientific American, 98(22):386, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
386d.pdf.

Anonymous:1908:ASD

[11132] Anonymous. Automobile sled for Dr. Charcot’s Expedition. Scientific American, 98(22):388, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
388.pdf.

Hislam:1908:BDC

[11133] Percival A. Hislam. The British “Dreadnought” cruisers. Scientific American, 98(22):389–390, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
389.pdf.

Anonymous:1908:FFA

[11134] Anonymous. First flights of the Aerial Experiment Association’s second aeroplane. Scientific American, 98(22):392, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n22/pdf/scientificamerican05301908-392.pdf.

Anonymous:1908:WATa

[11135] Anonymous. The Wright Aeroplane Test in North Carolina. Scientific American, 98(22):393, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
393.pdf.

Laufer:1908:CPW

[11136] Berthold Laufer. Chinese pigeon whistles. Scientific American, 98(22):394, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
394.pdf.

Anonymous:1908:NAV

[11137] Anonymous. A new automatic violin player. Scientific American, 98(22):395, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
395.pdf.

Michaud:1908:OST

[11138] Gustave Michaud. An oral system of telegraphy. Scientific American, 98(22):395, May 30, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n22/pdf/scientificamerican05301908-
395a.pdf.

Anonymous:1908:IGL

[11139] Anonymous. Incandescent gas lamp with self-generated forced draft. Scientific American, 98(22):396, May 30, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n22/pdf/scientificamerican05301908-396.pdf.

Anonymous:1908:Cn

[11140] Anonymous. Correspondence. Scientific American, 98(23):411, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n23/pdf/scientificamerican06061908-411a.pdf.

Anonymous:1908:RPIn

[11141] Anonymous. Recently patented inventions. Scientific American, 98(23):416, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
416.pdf.

Anonymous:1908:IIp

[11142] Anonymous. Index of inventions. Scientific American, 98(23):416–419, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n23/pdf/scientificamerican06061908-416a.pdf.

Anonymous:1908:PNPp

[11143] Anonymous. Patent notes: Patents. Scientific American, 98(23):419, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n23/pdf/scientificamerican06061908-419.pdf.

Anonymous:1908:BTI

[11144] Anonymous. Battle test of a 12-inch turret. Scientific American, 98(23):406, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
406.pdf.

Anonymous:1908:CSj

[11145] Anonymous. The capacity of the subway. Scientific American, 98(23):406, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
406a.pdf.

Anonymous:1908:RLC

[11146] Anonymous. Rectangular latticed compression members in bridges. Scientific American, 98(23):406, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
406b.pdf.

Manning:1908:PPR

[11147] William J. Manning. Permanent printed records on linen. Scientific American, 98(23):407, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
407.pdf.

Anonymous:1908:AN

[11148] Anonymous. Aeronautical notes. Scientific American, 98(23):407, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n23/pdf/scientificamerican06061908-407a.pdf.

Anonymous:1908:CSk

[11149] Anonymous. The current supplement. Scientific American, 98(23):407, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n23/pdf/scientificamerican06061908-407b.pdf.

Anonymous:1908:ASAb

[11150] Anonymous. The award of the Scientific American medal. Scientific American, 98(23):408, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
408.pdf.

Thompson:1908:LRT

[11151] Grace Agnes Thompson. The largest reflecting telescope in the world. Scientific American, 98(23):409–410, June 6, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n23/pdf/scientificamerican06061908-409.pdf.

Anonymous:1908:PCF

[11152] Anonymous. Panama Canal: Final plan and present condition of the work. Scientific American, 98(23):410–411, June 6, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n23/pdf/scientificamerican06061908-410.pdf.

Anonymous:1908:MAPa

[11153] Anonymous. The mechanics of “autos that pass in the air”. Scientific American, 98(23):411, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
411.pdf.

Anonymous:1908:KPT

[11154] Anonymous. Knudsen’s process of transmitting pictures by wireless telegraphy. Scientific American, 98(23):412, June 6, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n23/pdf/scientificamerican06061908-412.pdf.

Boyer:1908:IEB

[11155] Jacques Boyer. The insect enemies of books. Scientific American, 98(23):413–414, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
413.pdf.

Anonymous:1908:NSD

[11156] Anonymous. A new Sun dial. Scientific American, 98(23):414–415, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n23/pdf/scientificamerican06061908-414.pdf.

Anonymous:1908:ISH

[11157] Anonymous. An improved saw handle. Scientific American, 98(23):415, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n23/pdf/scientificamerican06061908-415.pdf.

Anonymous:1908:RCW

[11158] Anonymous. Rack for coats, wraps, and other garments. Scientific American, 98(23):415, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
415a.pdf.

Anonymous:1908:PTM

[11159] Anonymous. Power transmission mechanism. Scientific American, 98(23):415, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
415b.pdf.

Anonymous:1908:TAJ

[11160] Anonymous. Tongs with adjustable jaws. Scientific American, 98(23):415, June 6, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n23/pdf/scientificamerican06061908-
415c.pdf.

Anonymous:1908:RPIo

[11161] Anonymous. Recently patented inventions. Scientific American, 98(24):432, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
432.pdf.

Anonymous:1908:IIq

[11162] Anonymous. Index of inventions. Scientific American, 98(24):432–435, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n24/pdf/scientificamerican06131908-432a.pdf.

Anonymous:1908:PNPq

[11163] Anonymous. Patent notes: Patents. Scientific American, 98(24):435, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n24/pdf/scientificamerican06131908-435.pdf.

Anonymous:1908:AF

[11164] Anonymous. Airship follies. Scientific American, 98(24):422, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n24/pdf/scientificamerican06131908-422.pdf.

Anonymous:1908:MST

[11165] Anonymous. A marine steam turbine anomaly. Scientific American, 98(24):422, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
422a.pdf.

Anonymous:1908:AER

[11166] Anonymous. Admiral Evans’s report on the needs of our ships. Scientific American, 98(24):422–423, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
422b.pdf.

Wright:1908:OAT

[11167] Orville Wright and Wilbur Wright. Our aeroplane tests at Kitty Hawk. Scientific American, 98(24):423, June 13, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n24/pdf/scientificamerican06131908-423.pdf.

Anonymous:1908:DFB

[11168] Anonymous. Death of Francis B. Stevens. Scientific American, 98(24):423, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
423a.pdf.

Anonymous:1908:DMT

[11169] Anonymous. Death of miss tyler. Scientific American, 98(24):423, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n24/pdf/scientificamerican06131908-423b.pdf.

Anonymous:1908:FFP

[11170] Anonymous. Future foreign postage reduction with Great Britain and Colonies. Scientific American, 98(24):423, June 13, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n24/pdf/scientificamerican06131908-423c.pdf.

Anonymous:1908:CSl

[11171] Anonymous. The current supplement. Scientific American, 98(24):423, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n24/pdf/scientificamerican06131908-423d.pdf.

Anonymous:1908:SNb

[11172] Anonymous. Science notes. Scientific American, 98(24):423, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n24/pdf/scientificamerican06131908-423e.pdf.

Anonymous:1908:NMT

[11173] Anonymous. The new method of trying battleships. Scientific American, 98(24):424–425, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
424.pdf.

Douglas:1908:SDH

[11174] E. M. Douglas. Sun dials: How they are made and used. Scientific American, 98(24):425–427, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
425.pdf.

Willey:1908:TTU

[11175] Day Allen Willey. Towing tank at the University of Michigan. Scientific American, 98(24):428–429, June 13, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n24/pdf/scientificamerican06131908-428.pdf.

Baltimore:1908:MBW

[11176] J. Mayne Baltimore. A machine for burning weeds along railroads. Scientific American, 98(24):429, June 13, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n24/pdf/scientificamerican06131908-429.pdf.

Anonymous:1908:RGC

[11177] Anonymous. Removing the Grand Central Train Shed New York. Scientific American, 98(24):429, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
429a.pdf.

Anonymous:1908:GADb

[11178] Anonymous. A gigantic airship disaster. Scientific American, 98 (24):430–431, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
430.pdf.

Anonymous:1908:BCH

[11179] Anonymous. The British challenger for the Harmsworth Motor-Boat Trophy. — a 400-horse-power racing craft. Scientific American, 98(24):431, June 13, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n24/pdf/scientificamerican06131908-
431.pdf.

Anonymous:1908:RPIp

[11180] Anonymous. Recently patented inventions. Scientific American, 98(25):448, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
448.pdf.

Anonymous:1908:IIr

[11181] Anonymous. Index of inventions. Scientific American, 98(25):448–451, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-448a.pdf.

Anonymous:1908:PNPr

[11182] Anonymous. Patent notes: Patents. Scientific American, 98(25):452, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-452.pdf.

Anonymous:1908:CGa

[11183] Anonymous. The cost of government. Scientific American, 98(25):438, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-438.pdf.

Anonymous:1908:DFT

[11184] Anonymous. The double-flow turbine. Scientific American, 98(25):438, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-438a.pdf.

Anonymous:1908:GTR

[11185] Anonymous. Government tests of railway safety devices. Scientific American, 98(25):438, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
438b.pdf.

Anonymous:1908:CIP

[11186] Anonymous. Congress increases Patent Office salaries. Scientific American, 98(25):438, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
438c.pdf.

Anonymous:1908:INJa

[11187] Anonymous. Improvement of New Jersey–New York suburban service — I. Scientific American, 98(25):438–439, June 20, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n25/pdf/scientificamerican06201908-438d.pdf.

Anonymous:1908:ASAc

[11188] Anonymous. The autogenous soldering of aluminium in aeronautic construction. Scientific American, 98(25):439, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n25/pdf/scientificamerican06201908-439.pdf.

Anonymous:1908:NBM

[11189] Anonymous. A new balloon material. Scientific American, 98(25):439, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-439a.pdf.

Anonymous:1908:CSm

[11190] Anonymous. The current supplement. Scientific American, 98(25):439, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-439b.pdf.

Cox:1908:TN

[11191] John R. Cox. Training in the Navy. Scientific American, 98(25):440, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-440.pdf.

Claudy:1908:NEA

[11192] C. H. Claudy. New electrical apparatus of the United States Signal Corps. Scientific American, 98(25):441–442, June 20, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n25/pdf/scientificamerican06201908-441.pdf.

Anonymous:1908:MBR

[11193] Anonymous. The motor-boat race to Bermuda. Scientific American, 98(25):442, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
442.pdf.

Anonymous:1908:FAFa

[11194] Anonymous. Farman’s aeroplane flights in Belgium. Scientific American, 98(25):443–444, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
443.pdf.

MClure:1908:MCG

[11195] W. Frank M’Clure. Making casts in gelatin molds. Scientific American, 98(25):445, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
445.pdf.

Anonymous:1908:GCT

[11196] Anonymous. A German compass-testing station. Scientific American, 98(25):445, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
445a.pdf.

Gradenwitz:1908:RMR

[11197] Alfred Gradenwitz. A remarkable mountain railway. Scientific American, 98(25):446, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
446.pdf.

Anonymous:1908:ATF

[11198] Anonymous. An aero tube flying machine. Scientific American, 98(25):447, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
447.pdf.

Anonymous:1908:PRL

[11199] Anonymous. Pipe and rod lifting tongs. Scientific American, 98(25):447, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-447a.pdf.

Anonymous:1908:IDB

[11200] Anonymous. An improved dabbing brush. Scientific American, 98(25):447, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n25/pdf/scientificamerican06201908-
447b.pdf.

Anonymous:1908:BHR

[11201] Anonymous. Bracket hose reel. Scientific American, 98(25):447, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-447c.pdf.

Anonymous:1908:SS

[11202] Anonymous. Saw swage. Scientific American, 98(25):447, June 20, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n25/pdf/scientificamerican06201908-447d.pdf.

Russell:1908:HJc

[11203] Henry Norris Russell. The heavens in July. Scientific American, 98(26):455, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
455.pdf.

Anonymous:1908:Co

[11204] Anonymous. Correspondence. Scientific American, 98(26):459, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n26/pdf/scientificamerican06271908-459.pdf.

Anonymous:1908:RPIq

[11205] Anonymous. Recently patented inventions. Scientific American, 98(26):464–465, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
464.pdf.

Anonymous:1908:NQl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n26/pdf/scientificamerican06271908-465.pdf.

Anonymous:1908:BNBl

[11207] Anonymous. Books: New books, etc. Scientific American, 98 (26):465–466, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
465a.pdf.

Anonymous:1908:IIs

[11208] Anonymous. Index of inventions. Scientific American, 98(26):466–471, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n26/pdf/scientificamerican06271908-466.pdf.

Anonymous:1908:PNPs

[11209] Anonymous. Patent notes: Patents. Scientific American, 98(26):472, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n26/pdf/scientificamerican06271908-472.pdf.

Anonymous:1908:INF

[11210] Anonymous. Inspection of Niagara Falls Power Company Tunnel. Scientific American, 98(26):454, June 27, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n26/pdf/scientificamerican06271908-454.pdf.

Anonymous:1908:FCS

[11211] Anonymous. First cable strung on Manhattan Bridge. Scientific American, 98(26):454, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
454a.pdf.

Anonymous:1908:GDW

[11212] Anonymous. The gas driven warship. Scientific American, 98(26):454, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n26/pdf/scientificamerican06271908-454b.pdf.

Anonymous:1908:BMWa

[11213] Anonymous. Brilliant medical work of Col. Gorgas at Panama. Scientific American, 98(26):454, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
454c.pdf.

Anonymous:1908:ILV

[11214] Anonymous. An incandescent lamp without a vacuum. Scientific American, 98(26):454, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
454d.pdf.

Anonymous:1908:RFP

[11215] Anonymous. Reduction of foreign parcel postal rates. Scientific American, 98(26):455, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
455a.pdf.

Anonymous:1908:DCA

[11216] Anonymous. Delagrange to come to America. Scientific American, 98(26):455, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
455b.pdf.

Anonymous:1908:TAM

[11217] Anonymous. Torpedo attack on the monitor “Florida”. Scientific American, 98(26):456, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
456.pdf.

Anonymous:1908:TOR

[11218] Anonymous. The taxicab in old rome. Scientific American, 98(26):456, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n26/pdf/scientificamerican06271908-456a.pdf.

Anonymous:1908:SPR

[11219] Anonymous. Seeing from Paris to Rome. Scientific American, 98(26):457, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
457.pdf.

Perkins:1908:MFT

[11220] Frank C. Perkins. The modern farmer’s tireless horse. Scientific American, 98(26):458, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
458.pdf.

VanBrussel:1908:CEE

[11221] J. B. Van Brussel. A curious electric elevator. Scientific American, 98(26):460, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
460.pdf.

Russell:1908:PLC

[11222] Newton Forest Russell. A Philippine leper colony. Scientific American, 98(26):461–462, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
461.pdf.

Anonymous:1908:NGS

[11223] Anonymous. New grouping of stars in the United States flag. Scientific American, 98(26):463, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
463.pdf.

Masdell:1908:CSD

[11224] H. M. Masdell. A clever substitute for a drydock. Scientific American, 98(26):463, June 27, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n26/pdf/scientificamerican06271908-
463a.pdf.

Anonymous:1908:Cp

[11225] Anonymous. Correspondence. Scientific American, 98(3):43, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n3/pdf/scientificamerican01181908-43.pdf.

Anonymous:1908:RPIr

[11226] Anonymous. Recently patented inventions. Scientific American, 98(3):48, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
48.pdf.

Anonymous:1908:NQm

[11227] Anonymous. Notes and queries. Scientific American, 98(3):48, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n3/pdf/scientificamerican01181908-48a.pdf.

Anonymous:1908:BNBm

[11228] Anonymous. Books: New books, etc. Scientific American, 98 (3):48, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
48b.pdf.

Anonymous:1908:IIt

[11229] Anonymous. Index of inventions. Scientific American, 98(3):48–51, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n3/pdf/scientificamerican01181908-48c.pdf.

Anonymous:1908:PNPt

[11230] Anonymous. Patent notes: Patents. Scientific American, 98(3):52, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n3/pdf/scientificamerican01181908-52.pdf.

Anonymous:1908:RAOa

[11231] Anonymous. The Reuterdahl attack on our Navy. Scientific American, 98(3):38–39, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
38.pdf.

Anonymous:1908:IET

[11232] Anonymous. Inauguration of electric trains under the Hudson and East Rivers. Scientific American, 98(3):39, January 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n3/pdf/scientificamerican01181908-39.pdf.

Anonymous:1908:WAS

[11233] Anonymous. Why not an American “Scientists’ Corner” in the Cathedral of St. John the Divine? Scientific American, 98(3):39, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n3/pdf/scientificamerican01181908-39a.pdf.

Anonymous:1908:UGF

[11234] Anonymous. The use of glass fabrics in the electrical and chemical industries. Scientific American, 98(3):39, January 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n3/pdf/scientificamerican01181908-39b.pdf.

Anonymous:1908:CSn

[11235] Anonymous. The current supplement. Scientific American, 98 (3):39, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
39c.pdf.

Anonymous:1908:NEAa

[11236] Anonymous. New European aeroplanes and airships. Scientific American, 98(3):40–41, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
40.pdf.

Wilson:1908:JCC

[11237] L. F. Wilson. Jet condensers and the Chicago to the Gulf Canal. Scientific American, 98(3):42, January 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n3/pdf/scientificamerican01181908-42.pdf.

Anonymous:1908:NSP

[11238] Anonymous. A novel sand plow for saving ships. Scientific American, 98(3):42, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
42a.pdf.

Beasley:1908:HIM

[11239] Walter L. Beasley. The Huichol Indians of Mexico. Scientific American, 98(3):44–46, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
44.pdf.

Anonymous:1908:HSL

[11240] Anonymous. How the speed of light is measured. Scientific American, 98(3):46, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
46.pdf.

Anonymous:1908:PMC

[11241] Anonymous. Placer mining by a combination of hydraulic and electric methods. Scientific American, 98(3):47, January 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n3/pdf/scientificamerican01181908-47.pdf.

Anonymous:1908:AWM

[11242] Anonymous. Automatic weighing machine. Scientific American, 98(3):47, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
47a.pdf.

Anonymous:1908:TCC

[11243] Anonymous. Traveling cards for car advertising. Scientific American, 98(3):47, January 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n3/pdf/scientificamerican01181908-
47b.pdf.

Anonymous:1908:Cq

[11244] Anonymous. Correspondence. Scientific American, 98(4):59, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n4/pdf/scientificamerican01251908-59.pdf.

Anonymous:1908:RPIs

[11245] Anonymous. Recently patented inventions. Scientific American, 98(4):65–67, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
65.pdf.

Anonymous:1908:IINa

[11246] Anonymous. Index of inventions: Notes and queries. Scientific American, 98(4):67, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
67.pdf.

Anonymous:1908:BNBn

[11247] Anonymous. Books: New books, etc. Scientific American, 98 (4):67–68, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
67a.pdf.

Anonymous:1908:IIu

[11248] Anonymous. Index of inventions. Scientific American, 98(4):68–71, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n4/pdf/scientificamerican01251908-68.pdf.

Anonymous:1908:PNPu

[11249] Anonymous. Patent notes: Patents. Scientific American, 98(4):72, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n4/pdf/scientificamerican01251908-72.pdf.

Anonymous:1908:PCC

[11250] Anonymous. Panama Canal to cost $300,00,00? Scientific American, 98(4):54, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-

54.pdf. Check commas in title??

Anonymous:1908:PTE

[11251] Anonymous. The Pennsylvania tunnel excavation beneath Manhattan completed. Scientific American, 98(4):54, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n4/pdf/scientificamerican01251908-54a.pdf.

Anonymous:1908:GIT

[11252] Anonymous. Great increase in transatlantic travel. Scientific American, 98(4):54, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
54b.pdf.

Anonymous:1908:FAW

[11253] Anonymous. The Farman aeroplane wins the Deutsch Archdeacon Prize. Scientific American, 98(4):54, January 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n4/pdf/scientificamerican01251908-54c.pdf.

Anonymous:1908:PHC

[11254] Anonymous. Pure hydrogen from carbides. Scientific American, 98(4):54, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
54d.pdf.

Anonymous:1908:PO

[11255] Anonymous. The Patent Office in 1907. Scientific American, 98(4):55, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
55.pdf.

Anonymous:1908:PSI

[11256] Anonymous. The Pacific Scientific Institution of Honolulu. Scientific American, 98(4):55, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
55a.pdf.

Anonymous:1908:IIUa

[11257] Anonymous. International investigation of the upper air. Scientific American, 98(4):55, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
55b.pdf.

Anonymous:1908:CSo

[11258] Anonymous. The current supplement. Scientific American, 98 (4):55, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
55c.pdf.

Anonymous:1908:UAT

[11259] Anonymous. A unique aerial tramway used in the construction of Chicago’s Lake Tunnel. Scientific American, 98(4):56, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n4/pdf/scientificamerican01251908-56.pdf.

Gradenwitz:1908:PMP

[11260] Alfred Gradenwitz. A photographic method of producing basreliefs. Scientific American, 98(4):57–58, January 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n4/pdf/scientificamerican01251908-57.pdf.

Anonymous:1908:RAOb

[11261] Anonymous. The Reuterdahl attack on our Navy. Scientific American, 98(4):60–62, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
60.pdf.

Collins:1908:IE

[11262] Percy Collins. Insects that are eaten. Scientific American, 98 (4):63–64, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
63.pdf.

Anonymous:1908:IBT

[11263] Anonymous. An improved bale tie. Scientific American, 98(4):64, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n4/pdf/scientificamerican01251908-64.pdf.

Anonymous:1908:DPS

[11264] Anonymous. Device for pulling spokes. Scientific American, 98 (4):64, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
64a.pdf.

Anonymous:1908:IPC

[11265] Anonymous. An improved portable concrete mixer. Scientific American, 98(4):64, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
64b.pdf.

Anonymous:1908:AET

[11266] Anonymous. An apparatus for electrically transmitting sight. Scientific American, 98(4):64, January 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n4/pdf/scientificamerican01251908-
64c.pdf.

Russell:1908:HF

[11267] Henry Norris Russell. The heavens in February. Scientific American, 98(5):75, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
75.pdf.

Anonymous:1908:Cr

[11268] Anonymous. Correspondence. Scientific American, 98(5):79, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n5/pdf/scientificamerican02011908-79.pdf.

Anonymous:1908:RPIt

[11269] Anonymous. Recently patented inventions. Scientific American, 98(5):84, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
84.pdf.

Anonymous:1908:IIv

[11270] Anonymous. Index of inventions. Scientific American, 98(5): 84–87, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
84a.pdf.

Anonymous:1908:PNPv

[11271] Anonymous. Patent notes: Patents. Scientific American, 98 (5):88, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
88.pdf.

Anonymous:1908:YPP

[11272] Anonymous. A year’s progress on the Panama Canal. Scientific American, 98(5):74, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
74.pdf.

Anonymous:1908:LPB

[11273] Anonymous. The lessons of the Parker Building Fire. Scientific American, 98(5):74, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
74a.pdf.

Anonymous:1908:EAD

[11274] Anonymous. The enlargement of the Assouan Dam. Scientific American, 98(5):74, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
74b.pdf.

Anonymous:1908:FWAb

[11275] Anonymous. The fastest warship afloat. Scientific American, 98 (5):74, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
74c.pdf.

Anonymous:1908:PRD

[11276] Anonymous. Pennsylvania Railroad delays improvements. Scientific American, 98(5):74, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
74d.pdf.

Anonymous:1908:PMS

[11277] Anonymous. A $25,000 prize for a mathematical solution. Scientific American, 98(5):75, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
75a.pdf.

Anonymous:1908:STA

[11278] Anonymous. A strip of territory added to Alaska. Scientific American, 98(5):75, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
75b.pdf.

Anonymous:1908:UBV

[11279] Anonymous. A universal bench vise. Scientific American, 98 (5):76, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
76.pdf.

Anonymous:1908:MCI

[11280] Anonymous. A motor chair invented by a woman. Scientific American, 98(5):76, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
76a.pdf.

Anonymous:1908:CTD

[11281] Anonymous. Chess in three dimensions. Scientific American, 98 (5):76, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
76b.pdf.

Anonymous:1908:EMB

[11282] Anonymous. Erection of the Manhattan Bridge across the East River. Scientific American, 98(5):77–79, February 1, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n5/pdf/scientificamerican02011908-77.pdf.

Anonymous:1908:OAV

[11283] Anonymous. An open-air veneer machine. Scientific American, 98(5):80, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
80.pdf.

Wade:1908:HOM

[11284] Herbert T. Wade. How our measures of length are tested. Scientific American, 98(5):80–82, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
80a.pdf.

Anonymous:1908:HWK

[11285] Anonymous. How we know that the Earth rotates. Scientific American, 98(5):82, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
82.pdf.

Shepstone:1908:FTH

[11286] Harold J. Shepstone. A famous taxidermist and his methods. Scientific American, 98(5):83, February 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n5/pdf/scientificamerican02011908-
83.pdf.

Anonymous:1908:Cu

[11287] Anonymous. Correspondence. Scientific American, 98(6):95, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n6/pdf/scientificamerican02081908-95.pdf.

Anonymous:1908:RPIu

[11288] Anonymous. Recently patented inventions. Scientific American, 98(6):100, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
100.pdf.

Anonymous:1908:NQn

[11289] Anonymous. Notes and queries. Scientific American, 98(6):100, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n6/pdf/scientificamerican02081908-100a.pdf.

Anonymous:1908:BNBo

[11290] Anonymous. Books: New books, etc. Scientific American, 98 (6):100–101, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
100b.pdf.

Anonymous:1908:IIw

[11291] Anonymous. Index of inventions. Scientific American, 98(6): 101–103, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
101.pdf.

Anonymous:1908:PNPw

[11292] Anonymous. Patent notes: Patents. Scientific American, 98(6): 104, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
104.pdf.

Anonymous:1908:HRP

[11293] Anonymous. The Hudson River Pierhead Line should be extended. Scientific American, 98(6):90, February 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n6/pdf/scientificamerican02081908-90.pdf.

Anonymous:1908:PNYa

[11294] Anonymous. Progress on the New York Barge Canal — and a suggestion. Scientific American, 98(6):90, February 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n6/pdf/scientificamerican02081908-90a.pdf.

Anonymous:1908:RCO

[11295] Anonymous. The recent criticisms of our Navy. Scientific American, 98(6):90, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
90b.pdf.

Anonymous:1908:WPC

[11296] Anonymous. The world’s production of cereals. Scientific American, 98(6):90–91, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
90c.pdf.

Stratton:1908:MEE

[11297] George Frederic Stratton. The making of an electrical engineer. Scientific American, 98(6):91, February 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n6/pdf/scientificamerican02081908-91.pdf.

Anonymous:1908:SNc

[11298] Anonymous. Science notes. Scientific American, 98(6):91, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n6/pdf/scientificamerican02081908-91a.pdf.

Anonymous:1908:CSp

[11299] Anonymous. The current supplement. Scientific American, 98 (6):91, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
91b.pdf.

Anonymous:1908:PWC

[11300] Anonymous. The prize-winning circular flight of the Farman aeroplane. Scientific American, 98(6):92, February 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n6/pdf/scientificamerican02081908-92.pdf.

Wade:1908:IBW

[11301] Herbert T. Wade. The International Bureau of Weights and Measures. Scientific American, 98(6):93–94, February 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n6/pdf/scientificamerican02081908-93.pdf.

Anonymous:1908:NTS

[11302] Anonymous. New terminal station and approaches of the Brooklyn Bridge. Scientific American, 98(6):94, February 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n6/pdf/scientificamerican02081908-94.pdf.

Anonymous:1908:GTW

[11303] Anonymous. Gobelin tapestry works. Scientific American, 98 (6):96–97, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
96.pdf.

Perkins:1908:NSP

[11304] Frank C. Perkins. A new solar power plant. Scientific American, 98(6):97–98, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
97.pdf.

Forbes:1908:FPE

[11305] Charles Forbes. Foucault’s Pendulum Experiment repeated at Columbia University. Scientific American, 98(6):98, February 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n6/pdf/scientificamerican02081908-98.pdf.

Anonymous:1908:PDe

[11306] Anonymous. Patent department. Scientific American, 98(6):99, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n6/pdf/scientificamerican02081908-99.pdf.

Anonymous:1908:SLS

[11307] Anonymous. Safety lock for switch stands. Scientific American, 98(6):99, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
99a.pdf.

Anonymous:1908:BHF

[11308] Anonymous. Basket holder for fruit pickers. Scientific American, 98(6):99, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
99b.pdf.

Anonymous:1908:NOS

[11309] Anonymous. A novel observation swing. Scientific American, 98(6):99, February 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n6/pdf/scientificamerican02081908-
99c.pdf.

Anonymous:1908:RPIv

[11310] Anonymous. Recently patented inventions. Scientific American, 98(7):116, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
116.pdf.

Anonymous:1908:IIx

[11311] Anonymous. Index of inventions. Scientific American, 98(7): 116–119, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
116a.pdf.

Anonymous:1908:PNPx

[11312] Anonymous. Patent notes: Patents. Scientific American, 98 (7):119, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
119.pdf.

Anonymous:1908:SPJ

[11313] Anonymous. Southwest Pass jetties of the Mississippi completed. Scientific American, 98(7):106, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
106.pdf.

Anonymous:1908:PKW

[11314] Anonymous. Progress of the Key West Railroad. Scientific American, 98(7):106, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
106a.pdf.

Anonymous:1908:LCR

[11315] Anonymous. The last of the cable roads. Scientific American, 98(7):106, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
106b.pdf.

Anonymous:1908:MBS

[11316] Anonymous. More block signals; fewer accidents. Scientific American, 98(7):106, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
106c.pdf.

Anonymous:1908:PAA

[11317] Anonymous. Precautions against anthracite mining accidents. Scientific American, 98(7):106–107, February 15, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n7/pdf/scientificamerican02151908-106d.pdf.

Anonymous:1908:TUS

[11318] Anonymous. Turbines in the United States Navy. Scientific American, 98(7):107, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
107.pdf.

Anonymous:1908:MBI

[11319] Anonymous. Mine blasting invention. Scientific American, 98 (7):107, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
107a.pdf.

Anonymous:1908:ULS

[11320] Anonymous. The use of the lumière system of color photography in medicine. Scientific American, 98(7):107, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n7/pdf/scientificamerican02151908-107b.pdf.

Anonymous:1908:ESD

[11321] Anonymous. Exposition of safety devices. Scientific American, 98(7):107, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
107c.pdf.

Anonymous:1908:CSq

[11322] Anonymous. The current supplement. Scientific American, 98 (7):107, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
107d.pdf.

Anonymous:1908:GRS

[11323] Anonymous. Growing rubber in Sumatra. Scientific American, 98(7):107, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
107e.pdf.

Anonymous:1908:TMS

[11324] Anonymous. The Tunis Monorail System. Scientific American, 98(7):108, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
108.pdf.

King:1908:HPE

[11325] Charles R. King. A high-power European locomotive. Scientific American, 98(7):108–109, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
108a.pdf.

Willey:1908:RCH

[11326] Day Allen Willey. A reinforced concrete hotel. Scientific American, 98(7):109–110, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
109.pdf.

Honey:1908:MES

[11327] Frederic R. Honey. Morning and evening stars for 1908. Scientific American, 98(7):110–111, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
110.pdf.

Wade:1908:MSP

[11328] Herbert T. Wade. Magnetic survey on the Pacific Ocean. Scientific American, 98(7):112–113, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
112.pdf.

Anonymous:1908:SGE

[11329] Anonymous. The Sampson gas-electric road train. Scientific American, 98(7):114, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
114.pdf.

Anonymous:1908:PDf

[11330] Anonymous. Patent department. Scientific American, 98(7):115, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n7/pdf/scientificamerican02151908-115.pdf.

Anonymous:1908:ITW

[11331] Anonymous. Ice tongs with weighing attachment. Scientific American, 98(7):115, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
115a.pdf.

Anonymous:1908:SCE

[11332] Anonymous. Safety clutch for elevators. Scientific American, 98 (7):115, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
115b.pdf.

Anonymous:1908:IDR

[11333] Anonymous. An improved draft regulator and ventilator. Scientific American, 98(7):115, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
115c.pdf.

Anonymous:1908:CTU

[11334] Anonymous. Cutting tool for use in making paper boxes. Scientific American, 98(7):115, February 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n7/pdf/scientificamerican02151908-
115d.pdf.

Anonymous:1908:Cv

[11335] Anonymous. Correspondence. Scientific American, 98(8):127, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n8/pdf/scientificamerican02221908-127.pdf.

Anonymous:1908:RPIw

[11336] Anonymous. Recently patented inventions. Scientific American, 98(8):132, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
132.pdf.

Anonymous:1908:NQo

[11337] Anonymous. Notes and queries. Scientific American, 98(8):132, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n8/pdf/scientificamerican02221908-132a.pdf.

Anonymous:1908:BNBp

[11338] Anonymous. Books: New books, etc. Scientific American, 98 (8):132, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
132b.pdf.

Anonymous:1908:IIy

[11339] Anonymous. Index of inventions. Scientific American, 98(8): 132–135, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
132c.pdf.

Anonymous:1908:PNPy

[11340] Anonymous. Patent notes: Patents. Scientific American, 98 (8):136, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
136.pdf.

Anonymous:1908:WRE

[11341] Anonymous. Where the reciprocating engine drove out the turbine. Scientific American, 98(8):122, February 22, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n8/pdf/scientificamerican02221908-122.pdf.

Anonymous:1908:PRS

[11342] Anonymous. The Pennsylvania Railroad and the steel rail problem. Scientific American, 98(8):122, February 22, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n8/pdf/scientificamerican02221908-122a.pdf.

Anonymous:1908:NNC

[11343] Anonymous. “The Navy” and Naval criticism. Scientific American, 98(8):122, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
122b.pdf.

Anonymous:1908:DBB

[11344] Anonymous. Designers of the Brooklyn Bridge Terminal. Scientific American, 98(8):122, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
122c.pdf.

Anonymous:1908:AAA

[11345] Anonymous. Award of the Army aeroplane contracts. Scientific American, 98(8):123, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
123.pdf.

Anonymous:1908:WWF

[11346] Anonymous. When was the first pendulum experiment performed in this country? Scientific American, 98(8):123, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n8/pdf/scientificamerican02221908-123a.pdf.

Anonymous:1908:LHA

[11347] Anonymous. Lemoine and his artificial diamonds. Scientific American, 98(8):123, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
123b.pdf.

Anonymous:1908:ASG

[11348] Anonymous. Asbestos slates — a German invention. Scientific American, 98(8):123, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
123c.pdf.

Anonymous:1908:EFP

[11349] Anonymous. Extermination of fruit pests in California. Scientific American, 98(8):123, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
123d.pdf.

Anonymous:1908:CSr

[11350] Anonymous. The current supplement. Scientific American, 98 (8):123, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
123e.pdf.

Anonymous:1908:OHR

[11351] Anonymous. Opening of the Hudson River Tunnel system. Scientific American, 98(8):124–126, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
124.pdf.

Warren:1908:TL

[11352] Frederic Blount Warren. The testing of a locomotive. Scientific American, 98(8):126–127, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
126.pdf.

Brand:1908:CFB

[11353] Paul J. Brand. Closing a 700-foot break in the Mississippi levees. Scientific American, 98(8):128–129, February 22, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n8/pdf/scientificamerican02221908-128.pdf.

Moss:1908:SWS

[11354] L. E. Moss. Sending the whalebacks out to sea. Scientific American, 98(8):129–130, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
129.pdf.

Bache:1908:EHM

[11355] René Bache. The efficiency of the human machine. Scientific American, 98(8):130, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
130.pdf.

Anonymous:1908:SCT

[11356] Anonymous. Shipping can for transporting milk. Scientific American, 98(8):131, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
131.pdf.

Anonymous:1908:IFL

[11357] Anonymous. An improved fish-line float. Scientific American, 98(8):131, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
131a.pdf.

Anonymous:1908:SCG

[11358] Anonymous. Safety connection for gas stoves. Scientific American, 98(8):131, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
131b.pdf.

Anonymous:1908:MAC

[11359] Anonymous. Micrometer attachment for calipers and the like. Scientific American, 98(8):131, February 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n8/pdf/scientificamerican02221908-
131c.pdf.

Anonymous:1908:RPIx

[11360] Anonymous. Recently patented inventions. Scientific American, 98(9):153, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
153.pdf.

Anonymous:1908:IIz

[11361] Anonymous. Index of inventions. Scientific American, 98(9): 153–159, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
153a.pdf.

Anonymous:1908:PNPz

[11362] Anonymous. Patent notes: Patents. Scientific American, 98 (9):159, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
159.pdf.

Anonymous:1908:SAT

[11363] Anonymous. Shall America take the lead in aeronautics? Scientific American, 98(9):138, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
138.pdf.

Anonymous:1908:UWA

[11364] Anonymous. Under-water attack of ships by gunfire. Scientific American, 98(9):138, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
138a.pdf.

Anonymous:1908:SGP

[11365] Anonymous. Speed of Government and private excavation compared. Scientific American, 98(9):138–139, February 29, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n9/pdf/scientificamerican02291908-138b.pdf.

Bowdish:1908:RAM

[11366] B. S. Bowdish. The relation of the Audubon Movement to the sportsman. Scientific American, 98(9):139, February 29, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n9/pdf/scientificamerican02291908-139.pdf.

Wright:1908:FSP

[11367] Wilbur Wright. Flying as a sport — its possibilities. Scientific American, 98(9):139, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
139a.pdf.

Anonymous:1908:SIW

[11368] Anonymous. Some impressive wireless statistics. Scientific American, 98(9):139, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
139b.pdf.

Anonymous:1908:MSL

[11369] Anonymous. Modern seven-league boots. Scientific American, 98 (9):140, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
140.pdf.

Collins:1908:NFM

[11370] Percy Collins. Nature’s flying machines. Scientific American, 98 (9):140–142, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
140a.pdf.

Ashley:1908:IYH

[11371] H. Percy Ashley. An ice yacht with a hollow backbone and 250 square feet sail area. Scientific American, 98(9):142–143, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n9/pdf/scientificamerican02291908-142.pdf.

Anonymous:1908:RBA

[11372] Anonymous. A record-breaking automobile bus. Scientific American, 98(9):144, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
144.pdf.

Willey:1908:MHH

[11373] Day Allen Willey. Mountaineering in the Himalayas — the hardships of climbing at high altitudes. Scientific American, 98(9):144–145, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n9/pdf/scientificamerican02291908-144a.pdf.

Perry:1908:MBM

[11374] Harry Wilkin Perry. Motor boating for the man of small means. Scientific American, 98(9):146–147, February 29, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v98/
n9/pdf/scientificamerican02291908-146.pdf.

Anonymous:1908:NTG

[11375] Anonymous. A new target grip. Scientific American, 98(9):147, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n9/pdf/scientificamerican02291908-147.pdf.

Anonymous:1908:NSU

[11376] Anonymous. The new Syracuse University Stadium — the finest athletic arena in America. Scientific American, 98(9):148–150, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v98/n9/pdf/scientificamerican02291908-148.pdf.

Coleman:1908:FD

[11377] Peter Coleman. Fish doctoring. Scientific American, 98(9):150– 152, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
150.pdf.

Anonymous:1908:IT

[11378] Anonymous. Improvements in timers. Scientific American, 98(9): 152–153, February 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v98/n9/pdf/scientificamerican02291908-
152.pdf.

Anonymous:1908:Cx

[11379] Anonymous. Correspondence. Scientific American, 99(1):9, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-9.pdf.

Anonymous:1908:RPIy

[11380] Anonymous. Recently patented inventions. Scientific American, 99(1):14–15, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
14.pdf.

Anonymous:1908:NQp

[11381] Anonymous. Notes and queries. Scientific American, 99(1):15, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-15.pdf.

Anonymous:1908:BNBq

[11382] Anonymous. Books: New books, etc. Scientific American, 99(1):16, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-16.pdf.

Anonymous:1908:IIba

[11383] Anonymous. Index of inventions. Scientific American, 99(1):16–19, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-16a.pdf.

Anonymous:1908:PNPba

[11384] Anonymous. Patent notes: Patents. Scientific American, 99(1):20, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-20.pdf.

Anonymous:1908:FWAc

[11385] Anonymous. The fastest warship afloat. Scientific American, 99(1):4, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-4.pdf.

Anonymous:1908:PCP

[11386] Anonymous. Present condition of the Pan-American Railway. Scientific American, 99(1):4, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
4a.pdf.

Anonymous:1908:COF

[11387] Anonymous. Conservation of our fuel resources. Scientific American, 99(1):4, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
4b.pdf.

Anonymous:1908:PCR

[11388] Anonymous. Power consumption in the rapid transit subway. Scientific American, 99(1):4, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
4c.pdf.

Anonymous:1908:STR

[11389] Anonymous. Steel ties on roads with heavy traffic. Scientific American, 99(1):4, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
4d.pdf.

Anonymous:1908:IES

[11390] Anonymous. The irrigation of Egypt and the success of the Aswan Dam. Scientific American, 99(1):5, July 4, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n1/pdf/scientificamerican07041908-5.pdf.

Anonymous:1908:SSE

[11391] Anonymous. Special schools for eminently gifted pupils. Scientific American, 99(1):5, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
5a.pdf.

Anonymous:1908:ES

[11392] Anonymous. The emanation of sodium. Scientific American, 99(1):5, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-5b.pdf.

Anonymous:1908:CSs

[11393] Anonymous. The current supplement. Scientific American, 99(1):5, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-5c.pdf.

Anonymous:1908:SNd

[11394] Anonymous. Science notes. Scientific American, 99(1):5, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-5d.pdf.

Anonymous:1908:OLC

[11395] Anonymous. The outcome of the Lemoine Case. Scientific American, 99(1):5, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
5e.pdf.

Boyer:1908:CFH

[11396] Jacques Boyer. The curiosities of feeling, hearing, seeing and smelling. Scientific American, 99(1):6, July 4, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n1/pdf/scientificamerican07041908-6.pdf.

Finley:1908:CC

[11397] William L. Finley. The California condor. Scientific American, 99(1):7– 8, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-7.pdf.

Willey:1908:GRI

[11398] Day Allen Willey. The Great Roosevelt Irrigation Dam. Scientific American, 99(1):10–11, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
10.pdf.

Perkins:1908:GBB

[11399] Frank C. Perkins. The German battleship “Braunschweig”. Scientific American, 99(1):12, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
12.pdf.

Anonymous:1908:JBA

[11400] Anonymous. The “June Bug” aeroplane — a competitor for the Scientific American Trophy. Scientific American, 99(1):13, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n1/pdf/scientificamerican07041908-13.pdf.

Anonymous:1908:ASI

[11401] Anonymous. An Aeronautic Society for Inventors. Scientific American, 99(1):13, July 4, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n1/pdf/scientificamerican07041908-
13a.pdf.

Russell:1908:HS

[11402] Henry Norris Russell. The heavens in September. Scientific American, 99(10):151, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
151.pdf.

Anonymous:1908:Cy

[11403] Anonymous. Correspondence. Scientific American, 99(10):155, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n10/pdf/scientificamerican09051908-155.pdf.

Anonymous:1908:RPIz

[11404] Anonymous. Recently patented inventions. Scientific American, 99(10):160, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
160.pdf.

Anonymous:1908:NQq

[11405] Anonymous. Notes and queries. Scientific American, 99(10): 160, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
160a.pdf.

Anonymous:1908:IIbb

[11406] Anonymous. Index of inventions. Scientific American, 99(10): 160–163, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
160b.pdf.

Anonymous:1908:PNPbb

[11407] Anonymous. Patent notes: Patents. Scientific American, 99 (10):163, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
163.pdf.

Anonymous:1908:OES

[11408] Anonymous. Ornamental elevated structures. Scientific American, 99(10):150, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
150.pdf.

Anonymous:1908:EDP

[11409] Anonymous. Experimental dams at the Panama Canal. Scientific American, 99(10):150, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
150a.pdf.

Anonymous:1908:SNCa

[11410] Anonymous. The street noise crusade and the rail joint. Scientific American, 99(10):150, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
150b.pdf.

Anonymous:1908:DAT

[11411] Anonymous. A delayed-action torpedo. Scientific American, 99 (10):150, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
150c.pdf.

Anonymous:1908:FEY

[11412] Anonymous. Ferment of egg yolk. Scientific American, 99(10): 151, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
151a.pdf.

Anonymous:1908:RCB

[11413] Anonymous. Reinforced concrete as a building material for boats. Scientific American, 99(10):152–154, September 5, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n10/pdf/scientificamerican09051908-152.pdf.

Bayley:1908:RRP

[11414] I. C. Bayley. Repairing railroad picks. Scientific American, 99 (10):155, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
155a.pdf.

Anonymous:1908:BKD

[11415] Anonymous. The British 38-knot destroyer “Swift”. Scientific American, 99(10):156, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
156.pdf.

Anonymous:1908:MKD

[11416] Anonymous. The Mari-Kanave Dam in Southern India. Scientific American, 99(10):156–157, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
156a.pdf.

Willey:1908:ABC

[11417] Day Allen Willey. Antique bronze cannon at Annapolis and Washington. Scientific American, 99(10):158, September 5, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n10/pdf/scientificamerican09051908-158.pdf.

Michaud:1908:TOI

[11418] Gustave Michaud. Three optical illusions. Scientific American, 99(10):158, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
158a.pdf.

Anonymous:1908:NDP

[11419] Anonymous. The new Davis projectile torpedo. Scientific American, 99(10):159, September 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n10/pdf/scientificamerican09051908-
159.pdf.

Anonymous:1908:RPIba

[11420] Anonymous. Recently patented inventions. Scientific American, 99 (11):176, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
176.pdf.

Anonymous:1908:NQr

[11421] Anonymous. Notes and queries. Scientific American, 99(11): 176, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
176a.pdf.

Anonymous:1908:BNBr

[11422] Anonymous. Books: New books, etc. Scientific American, 99 (11):177, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
177.pdf.

Anonymous:1908:IIbc

[11423] Anonymous. Index of inventions. Scientific American, 99(11): 177–178, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
177a.pdf.

Anonymous:1908:PNPbc

[11424] Anonymous. Patent notes: Patents. Scientific American, 99 (11):179, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
179.pdf.

Anonymous:1908:APSa

[11425] Anonymous. Is the aeroplane practicable for scouting? Scientific American, 99(11):166, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
166.pdf.

Anonymous:1908:MUR

[11426] Anonymous. Manufacturing under the requirements of the British Patent Act. Scientific American, 99(11):166, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n11/pdf/scientificamerican09121908-166a.pdf.

Anonymous:1908:PCSb

[11427] Anonymous. Proposals for cooling the subway. Scientific American, 99(11):166, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
166b.pdf.

Anonymous:1908:INJb

[11428] Anonymous. Improvement of New Jersey–New York Suburban Service — III. Scientific American, 99(11):166–167, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n11/pdf/scientificamerican09121908-166c.pdf.

Adams:1908:AGI

[11429] Alton D. Adams. Acetylene gas for isolated street lamps. Scientific American, 99(11):167, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
167.pdf.

Anonymous:1908:OMSc

[11430] Anonymous. Official Meteorological Summary, New York, N.Y., August, 1908. Scientific American, 99(11):167, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n11/pdf/scientificamerican09121908-167a.pdf.

Anonymous:1908:CSt

[11431] Anonymous. The current supplement. Scientific American, 99 (11):167, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
167b.pdf.

Anonymous:1908:EA

[11432] Anonymous. Excavations at Abydos. Scientific American, 99 (11):167, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
167c.pdf.

Wade:1908:STU

[11433] Herbert T. Wade. Sugar testing in the United States Customs Service. Scientific American, 99(11):168, September 12, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n11/pdf/scientificamerican09121908-168.pdf.

Anonymous:1908:FFW

[11434] Anonymous. The first flight of the Wright aeroplane at Fort Myer. Scientific American, 99(11):169–170, September 12, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n11/pdf/scientificamerican09121908-169.pdf.

Anonymous:1908:HBD

[11435] Anonymous. A huge bucket dredger. Scientific American, 99 (11):171, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
171.pdf.

Honey:1908:RE

[11436] Frederic R. Honey. The recurrence of eclipses. Scientific American, 99(11):171, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
171a.pdf.

Gradenwitz:1908:PAR

[11437] Alfred Gradenwitz. A Police Academy at Rome. Scientific American, 99(11):172–173, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
172.pdf.

Anonymous:1908:NGR

[11438] Anonymous. The new German racing schooner “Germania”. Scientific American, 99(11):173, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
173.pdf.

Macdougal:1908:FP

[11439] D. T. Macdougal. The faculties of plants. Scientific American, 99 (11):174–175, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
174.pdf.

Crafts:1908:MID

[11440] H. A. Crafts. Modern improvements in drydock construction. Scientific American, 99(11):175, September 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n11/pdf/scientificamerican09121908-
175.pdf.

Anonymous:1908:Cz

[11441] Anonymous. Correspondence. Scientific American, 99(12):187, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n12/pdf/scientificamerican09191908-187.pdf.

Anonymous:1908:RPIbb

[11442] Anonymous. Recently patented inventions. Scientific American, 99 (12):193, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
193.pdf.

Anonymous:1908:NQs

[11443] Anonymous. Notes and queries. Scientific American, 99(12): 194, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
194.pdf.

Anonymous:1908:BNBs

[11444] Anonymous. Books: New books, etc. Scientific American, 99(12): 194–196, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
194a.pdf.

Anonymous:1908:IIbd

[11445] Anonymous. Index of inventions. Scientific American, 99(12): 196–199, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
196.pdf.

Anonymous:1908:PNPbd

[11446] Anonymous. Patent notes: Patents. Scientific American, 99 (12):200, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
200.pdf.

Anonymous:1908:RQBb

[11447] Anonymous. The rebuilding of the Quebec Bridge. Scientific American, 99(12):182, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
182.pdf.

Anonymous:1908:IRD

[11448] Anonymous. Improved rails by the dry-blast system. Scientific American, 99(12):182, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
182a.pdf.

Anonymous:1908:PGDb

[11449] Anonymous. Progress in the gas-driven ship problem. Scientific American, 99(12):182, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
182b.pdf.

Anonymous:1908:MPE

[11450] Anonymous. More powerful express locomotives. Scientific American, 99(12):182, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
182c.pdf.

Anonymous:1908:FSC

[11451] Anonymous. Future speed of cruisers. Scientific American, 99 (12):182, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
182d.pdf.

Anonymous:1908:MAH

[11452] Anonymous. The making of aeroplane history. Scientific American, 99(12):182, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
182e.pdf.

Anonymous:1908:DCR

[11453] Anonymous. Doubtful cases of radioactivity. Scientific American, 99 (12):182–183, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
182f.pdf.

Springer:1908:FFA

[11454] J. F. Springer. A few facts about fakes. Scientific American, 99 (12):183, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
183.pdf.

Anonymous:1908:CSu

[11455] Anonymous. The current supplement. Scientific American, 99 (12):183, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
183a.pdf.

Anonymous:1908:MCO

[11456] Anonymous. The Morehouse Comet observed. Scientific American, 99(12):183, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
183b.pdf.

Willey:1908:RBO

[11457] Day Allen Willey. Reconstruction of the Baltimore & Ohio Bridge over the Susquehanna River. Scientific American, 99(12):184, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n12/pdf/scientificamerican09191908-184.pdf.

Anonymous:1908:CSLa

[11458] Anonymous. The Collins system of long-distance wireless telephony. Scientific American, 99(12):185–186, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n12/pdf/scientificamerican09191908-185.pdf.

Anonymous:1908:PSC

[11459] Anonymous. Public service car fender and wheel guard tests. Scientific American, 99(12):186, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
186.pdf.

Anonymous:1908:IFN

[11460] Anonymous. IV. — the French Navy of to-day. Scientific American, 99(12):187–189, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
187a.pdf.

Anonymous:1908:FF

[11461] Anonymous. The “flip-flap”. Scientific American, 99(12):190, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n12/pdf/scientificamerican09191908-190.pdf.

MacMurray:1908:PTU

[11462] G. W. MacMurray. Protecting trains under the East River. Scientific American, 99(12):191, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
191.pdf.

Anonymous:1908:SSF

[11463] Anonymous. Shaft sinking by the freezing process. Scientific American, 99(12):191, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
191a.pdf.

Anonymous:1908:PDg

[11464] Anonymous. Patent department. Scientific American, 99(12): 192, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
192.pdf.

Anonymous:1908:OC

[11465] Anonymous. Ore concentrator. Scientific American, 99(12): 192, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
192a.pdf.

Anonymous:1908:IPW

[11466] Anonymous. An improved pipe wrench. Scientific American, 99 (12):192, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
192b.pdf.

Anonymous:1908:ALB

[11467] Anonymous. Attachment for letter boxes. Scientific American, 99 (12):192, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
192c.pdf.

Anonymous:1908:SPT

[11468] Anonymous. A simple package tie. Scientific American, 99(12): 192, September 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n12/pdf/scientificamerican09191908-
192d.pdf.

Anonymous:1908:RPIbc

[11469] Anonymous. Recently patented inventions. Scientific American, 99 (13):212, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
212.pdf.

Anonymous:1908:NQt

[11470] Anonymous. Notes and queries. Scientific American, 99(13): 212, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
212a.pdf.

Anonymous:1908:IIbe

[11471] Anonymous. Index of inventions. Scientific American, 99(13): 213–215, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
213.pdf.

Anonymous:1908:LWA

[11472] Anonymous. Lessons of the Wright aeroplane disaster. Scientific American, 99(13):202, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
202.pdf.

Anonymous:1908:SSS

[11473] Anonymous. Some submarine successes. Scientific American, 99 (13):202, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
202a.pdf.

Anonymous:1908:OPP

[11474] Anonymous. Our ponderous passenger cars. Scientific American, 99(13):202, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
202b.pdf.

Anonymous:1908:RCPa

[11475] Anonymous. Repairing the concrete pier of a bridge by injection of liquid cement mortar. Scientific American, 99(13):203, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n13/pdf/scientificamerican09261908-203.pdf.

Phillips:1908:FSI

[11476] Edwin Phillips. A few suggestions for inventors of safety devices. Scientific American, 99(13):203, September 26, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n13/pdf/scientificamerican09261908-203a.pdf.

Anonymous:1908:SNe

[11477] Anonymous. Science notes. Scientific American, 99(13):203, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n13/pdf/scientificamerican09261908-203b.pdf.

Anonymous:1908:CSv

[11478] Anonymous. The current supplement. Scientific American, 99 (13):203, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
203c.pdf.

Anonymous:1908:NGCa

[11479] Anonymous. The narrowing of the Ganges and construction of the Curzon Bridge. Scientific American, 99(13):204, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n13/pdf/scientificamerican09261908-204.pdf.

Anonymous:1908:SBT

[11480] Anonymous. Sawing a building in two. Scientific American, 99 (13):205–206, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
205.pdf.

Anonymous:1908:CWA

[11481] Anonymous. The construction of the Wright aeroplane. Scientific American, 99(13):209–210, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
209.pdf.

Anonymous:1908:EAP

[11482] Anonymous. An early armor-piercing gun. Scientific American, 99 (13):210, September 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n13/pdf/scientificamerican09261908-
210.pdf.

Anonymous:1908:CFT

[11483] Anonymous. Car fender tests by the Public Service Commission. Scientific American, 99(13):210–211, September 26, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n13/pdf/scientificamerican09261908-210a.pdf.

Russell:1908:HO

[11484] Henry Norris Russell. The heavens in October. Scientific American, 99(14):219, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
219.pdf.

Anonymous:1908:Cba

[11485] Anonymous. Correspondence. Scientific American, 99(14):223, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n14/pdf/scientificamerican10031908-223.pdf.

Anonymous:1908:RPIbd

[11486] Anonymous. Recently patented inventions. Scientific American, 99(14):229, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
229.pdf.

Anonymous:1908:NQNa

[11487] Anonymous. Notes and queries: Note and queries. Scientific American, 99(14):229–231, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
229a.pdf.

Anonymous:1908:BNBt

[11488] Anonymous. Books: New books, etc. Scientific American, 99 (14):231–232, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
231.pdf.

Anonymous:1908:IIbf

[11489] Anonymous. Index of inventions. Scientific American, 99(14):232–235, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n14/pdf/scientificamerican10031908-232.pdf.

Anonymous:1908:NA

[11490] Anonymous. The Naval aeroplane. Scientific American, 99(14):218, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n14/pdf/scientificamerican10031908-218.pdf.

Anonymous:1908:YR

[11491] Anonymous. Yacht racing in 1908. Scientific American, 99(14):218, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n14/pdf/scientificamerican10031908-218a.pdf.

Anonymous:1908:WSD

[11492] Anonymous. Water storage destroys the typhoid bacillus. Scientific American, 99(14):218, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
218b.pdf.

Anonymous:1908:PPE

[11493] Anonymous. Progress on Panama and Erie Canals compared. Scientific American, 99(14):218, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
218c.pdf.

Anonymous:1908:LCL

[11494] Anonymous. The liner “Cleveland” launched. Scientific American, 99(14):219, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
219a.pdf.

Anonymous:1908:MCL

[11495] Anonymous. A motor-car lock is wanted. Scientific American, 99(14):219, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
219b.pdf.

Anonymous:1908:NYS

[11496] Anonymous. The New York State Barge Canal. Scientific American, 99(14):220–222, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
220.pdf.

Anonymous:1908:PW

[11497] Anonymous. The price of water. Scientific American, 99(14):224, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n14/pdf/scientificamerican10031908-224.pdf.

Colins:1908:TTP

[11498] Percy Colins. Table-top photography. Scientific American, 99 (14):224–226, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
224a.pdf.

Carrington:1908:HM

[11499] Hereward Carrington. Hindu magic. Scientific American, 99 (14):226–227, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
226.pdf.

Anonymous:1908:PDh

[11500] Anonymous. Patent department. Scientific American, 99(14):227, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n14/pdf/scientificamerican10031908-227.pdf.

Anonymous:1908:SRTb

[11501] Anonymous. Some recent toys and tricks. Scientific American, 99(14):227, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
227a.pdf.

Anonymous:1908:TAW

[11502] Anonymous. Tire for automobile wheels. Scientific American, 99(14):228, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n14/pdf/scientificamerican10031908-
228.pdf.

Anonymous:1908:AGL

[11503] Anonymous. An automatic gas lighter. Scientific American, 99(14):228, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n14/pdf/scientificamerican10031908-228a.pdf.

Anonymous:1908:OId

[11504] Anonymous. Oddities in invention. Scientific American, 99(14):228, October 3, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n14/pdf/scientificamerican10031908-228b.pdf.

Anonymous:1908:Cbb

[11505] Anonymous. Correspondence. Scientific American, 99(15):243, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n15/pdf/scientificamerican10101908-243.pdf.

Anonymous:1908:RPIbe

[11506] Anonymous. Recently patented inventions. Scientific American, 99(15):248, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
248.pdf.

Anonymous:1908:NQu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n15/pdf/scientificamerican10101908-248a.pdf.

Anonymous:1908:BNBu

[11508] Anonymous. Books: New books, etc. Scientific American, 99 (15):248, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
248b.pdf.

Anonymous:1908:IIbg

[11509] Anonymous. Index of inventions. Scientific American, 99(15): 248–251, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
248c.pdf.

Anonymous:1908:EDI

[11510] Anonymous. Encouraging decrease of idle cars. Scientific American, 99(15):238, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
238.pdf.

Anonymous:1908:PTSa

[11511] Anonymous. Postal tubes in the subways. Scientific American, 99(15):238, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
238a.pdf.

Anonymous:1908:ELF

[11512] Anonymous. The enormous loss by forest fires. Scientific American, 99(15):238, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
238b.pdf.

Anonymous:1908:SCF

[11513] Anonymous. The Schenectady car fender tests. Scientific American, 99(15):238, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
238c.pdf.

Anonymous:1908:FTFa

[11514] Anonymous. A four-track freight tunnel for Manhattan Island. Scientific American, 99(15):238, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
238d.pdf.

Anonymous:1908:NCG

[11515] Anonymous. Nasmyth — the centenary of a great inventor. Scientific American, 99(15):238, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
238e.pdf.

Anonymous:1908:Eb

[11516] Anonymous. Engineering. Scientific American, 99(15):239, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n15/pdf/scientificamerican10101908-239.pdf.

Anonymous:1908:Ec

[11517] Anonymous. Electricity. Scientific American, 99(15):239, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n15/pdf/scientificamerican10101908-239a.pdf.

Anonymous:1908:Sa

[11518] Anonymous. Science. Scientific American, 99(15):239, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n15/pdf/scientificamerican10101908-239b.pdf.

Allen:1908:FBS

[11519] William Allen. Fall of the B. & O. Susquehanna Bridge. Scientific American, 99(15):240, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
240.pdf.

Wade:1908:SRS

[11520] Herbert T. Wade. The Sun’s radiation and its study. Scientific American, 99(15):241–242, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
241.pdf.

Anonymous:1908:PIR

[11521] Anonymous. The paper industry as related to forests and education. Scientific American, 99(15):243, October 10, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n15/pdf/scientificamerican10101908-243a.pdf.

Anonymous:1908:MAPb

[11522] Anonymous. The multiple air propeller. its aeronautic possibilities for dirigible airships, aeroplanes, and helicopters. Scientific American, 99(15):244, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
244.pdf.

Willey:1908:NSI

[11523] Day Allen Willey. The Naval Stores industry. Scientific American, 99(15):245–246, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
245.pdf.

Christadoro:1908:EDB

[11524] Charles Christadoro. The evolution of the dovetailed box. Scientific American, 99(15):247, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
247.pdf.

Anonymous:1908:CWP

[11525] Anonymous. Cutting wood with paper. Scientific American, 99 (15):247, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
247a.pdf.

Anonymous:1908:UJS

[11526] Anonymous. A universal joint of simple design. Scientific American, 99(15):247, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
247b.pdf.

Anonymous:1908:HOF

[11527] Anonymous. How to obtain fresh water from sea water. Scientific American, 99(15):247, October 10, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n15/pdf/scientificamerican10101908-
247c.pdf.

Anonymous:1908:Cbc

[11528] Anonymous. Correspondence. Scientific American, 99(16):259, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n16/pdf/scientificamerican10171908-259.pdf.

Anonymous:1908:RPIbf

[11529] Anonymous. Recently patented inventions. Scientific American, 99(16):265, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
265.pdf.

Anonymous:1908:NQv

[11530] Anonymous. Notes and queries. Scientific American, 99(16): 265–267, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
265a.pdf.

Anonymous:1908:BNBv

[11531] Anonymous. Books: New books, etc. Scientific American, 99 (16):267, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
267.pdf.

Anonymous:1908:IIbh

[11532] Anonymous. Index of inventions. Scientific American, 99(16): 267–271, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
267a.pdf.

Anonymous:1908:STV

[11533] Anonymous. Some theories verified. Scientific American, 99 (16):254, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
254.pdf.

Anonymous:1908:LT

[11534] Anonymous. Langley — a tribute. Scientific American, 99(16):254, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n16/pdf/scientificamerican10171908-254a.pdf.

Anonymous:1908:WSR

[11535] Anonymous. Wanted: A sanitary railway car. Scientific American, 99(16):254, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
254b.pdf.

Anonymous:1908:Ed

[11536] Anonymous. Engineering. Scientific American, 99(16):255, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n16/pdf/scientificamerican10171908-255.pdf.

Anonymous:1908:Ee

[11537] Anonymous. Electricity. Scientific American, 99(16):255, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n16/pdf/scientificamerican10171908-255a.pdf.

Anonymous:1908:Sb

[11538] Anonymous. Science. Scientific American, 99(16):255, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n16/pdf/scientificamerican10171908-255b.pdf.

Anonymous:1908:EP

[11539] Anonymous. The eyes of plants. Scientific American, 99(16):256, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n16/pdf/scientificamerican10171908-256.pdf.

Anonymous:1908:NIM

[11540] Anonymous. A new ice-making machine for domestic use. Scientific American, 99(16):257, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
257.pdf.

Anonymous:1908:RTS

[11541] Anonymous. A rescue training school and experimental gallery for miners. Scientific American, 99(16):260–261, October 17, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n16/pdf/scientificamerican10171908-260.pdf.

Beasley:1908:NFW

[11542] Walter L. Beasley. New fossil wonders from the Fayum of Egypt. Scientific American, 99(16):261–262, October 17, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n16/pdf/scientificamerican10171908-261.pdf.

Anonymous:1908:PGC

[11543] Anonymous. Polishing the great Cullinan Diamond. Scientific American, 99(16):262, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
262.pdf.

Anonymous:1908:HMWa

[11544] Anonymous. Handy man’s workshop. Scientific American, 99 (16):263–264, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
263.pdf.

Fagan:1908:SMM

[11545] A. Fagan. A simple method for making a coiled spring. Scientific American, 99(16):264, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
264.pdf.

Allen:1908:HMW

[11546] John R. Allen. A home-made Wimshurst machine. Scientific American, 99(16):264, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
264a.pdf.

Anonymous:1908:BNR

[11547] Anonymous. A blacksmith’s napkin ring. Scientific American, 99(16):264, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
264b.pdf.

Anonymous:1908:ULJ

[11548] Anonymous. An unbreakable Leyden jar. Scientific American, 99(16):264, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
264c.pdf.

Osborn:1908:SFP

[11549] A. E. Osborn. A simple foot-power compound grinder. Scientific American, 99(16):264, October 17, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n16/pdf/scientificamerican10171908-
264d.pdf.

Anonymous:1908:Cbd

[11550] Anonymous. Correspondence. Scientific American, 99(17):279, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n17/pdf/scientificamerican10241908-279.pdf.

Anonymous:1908:RPIbg

[11551] Anonymous. Recently patented inventions. Scientific American, 99(17):285, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
285.pdf.

Anonymous:1908:NQw

[11552] Anonymous. Notes and queries. Scientific American, 99(17): 285–288, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
285a.pdf.

Anonymous:1908:BNBw

[11553] Anonymous. Books: New books, etc. Scientific American, 99 (17):288, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
288.pdf.

Anonymous:1908:IIbi

[11554] Anonymous. Index of inventions. Scientific American, 99(17): 288–291, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
288a.pdf.

Anonymous:1908:BFW

[11555] Anonymous. The balloon in foggy weather. Scientific American, 99(17):274, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
274.pdf.

Anonymous:1908:OMR

[11556] Anonymous. Our macadamized road fiasco. Scientific American, 99(17):274, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
274a.pdf.

Anonymous:1908:FFN

[11557] Anonymous. Flarebacks in the French Navy. Scientific American, 99(17):274, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
274b.pdf.

Anonymous:1908:MPC

[11558] Anonymous. Manufacturing paper from cornstalks. Scientific American, 99(17):274, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
274c.pdf.

Anonymous:1908:Ef

[11559] Anonymous. Engineering. Scientific American, 99(17):275, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n17/pdf/scientificamerican10241908-275.pdf.

Anonymous:1908:Eg

[11560] Anonymous. Electricity. Scientific American, 99(17):275, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n17/pdf/scientificamerican10241908-275a.pdf.

Anonymous:1908:Sc

[11561] Anonymous. Science. Scientific American, 99(17):275, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n17/pdf/scientificamerican10241908-275b.pdf.

Anonymous:1908:AH

[11562] Anonymous. An automobile for horses. Scientific American, 99 (17):276, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
276.pdf.

Anonymous:1908:OMM

[11563] Anonymous. Original method of making a railroad fill. Scientific American, 99(17):276, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
276a.pdf.

Shepstone:1908:MTS

[11564] Harold J. Shepstone. The mountain telescopes of Switzerland. Scientific American, 99(17):277–278, October 24, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n17/pdf/scientificamerican10241908-277.pdf.

Anonymous:1908:VJN

[11565] Anonymous. V — the Japanese Navy of to-day. Scientific American, 99(17):279–282, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
279a.pdf.

Anonymous:1908:PDi

[11566] Anonymous. Patent department. Scientific American, 99(17):283, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n17/pdf/scientificamerican10241908-283.pdf.

Anonymous:1908:DCC

[11567] Anonymous. Door check and closes. Scientific American, 99 (17):283, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
283a.pdf.

Anonymous:1908:WCM

[11568] Anonymous. Water current motor. Scientific American, 99(17):283, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n17/pdf/scientificamerican10241908-283b.pdf.

Anonymous:1908:IRSb

[11569] Anonymous. Improved railway spike. Scientific American, 99 (17):283, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
283c.pdf.

Anonymous:1908:ITS

[11570] Anonymous. An improved telegraph sounder. Scientific American, 99(17):283, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
283d.pdf.

Anonymous:1908:RCPb

[11571] Anonymous. Report of the Commissioner of Patents. Scientific American, 99(17):284, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
284.pdf.

Anonymous:1908:NSA

[11572] Anonymous. A new shutter attachment. Scientific American, 99 (17):284, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
284a.pdf.

Anonymous:1908:IAP

[11573] Anonymous. An improved air pump. Scientific American, 99 (17):284, October 24, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n17/pdf/scientificamerican10241908-
284b.pdf.

Anonymous:1908:Cbe

[11574] Anonymous. Correspondence. Scientific American, 99(18):299, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n18/pdf/scientificamerican10311908-299.pdf.

Russell:1908:HN

[11575] Henry Norris Russell. The heavens in November. Scientific American, 99(18):302, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
302.pdf.

Anonymous:1908:RPIbh

[11576] Anonymous. Recently patented inventions. Scientific American, 99(18):304, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
304.pdf.

Anonymous:1908:IIbj

[11577] Anonymous. Index of inventions. Scientific American, 99(18): 304–307, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
304a.pdf.

Anonymous:1908:NC

[11578] Anonymous. That Newport Conference. Scientific American, 99 (18):294, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
294.pdf.

Anonymous:1908:FIA

[11579] Anonymous. The force involved in automobile collisions. Scientific American, 99(18):294, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
294a.pdf.

Anonymous:1908:VCR

[11580] Anonymous. The 1908 Vanderbilt Cup Race. Scientific American, 99(18):294, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
294b.pdf.

Anonymous:1908:Eh

[11581] Anonymous. Engineering. Scientific American, 99(18):295, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n18/pdf/scientificamerican10311908-295.pdf.

Anonymous:1908:Ei

[11582] Anonymous. Electricity. Scientific American, 99(18):295, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n18/pdf/scientificamerican10311908-295a.pdf.

Anonymous:1908:Ab

[11583] Anonymous. Aeronautics. Scientific American, 99(18):295, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n18/pdf/scientificamerican10311908-295b.pdf.

Anonymous:1908:KWT

[11584] Anonymous. The Knudsen wireless typewriter. Scientific American, 99(18):296–297, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
296.pdf.

Hislam:1908:SWB

[11585] Percival Hislam. Salving the wrecked British cruiser “Gladiator”. Scientific American, 99(18):297–298, October 31, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n18/pdf/scientificamerican10311908-297.pdf.

Anonymous:1908:FCM

[11586] Anonymous. Fire-control masts. Scientific American, 99(18):299, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n18/pdf/scientificamerican10311908-299a.pdf.

Anonymous:1908:CFAb

[11587] Anonymous. Curious facts about numbers. Scientific American, 99(18):299, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
299b.pdf.

Boyer:1908:MIG

[11588] Jacques Boyer. The manufacture of incandescent gas mantles. Scientific American, 99(18):300–301, October 31, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n18/pdf/scientificamerican10311908-300.pdf.

Ward:1908:LPQ

[11589] L. B. Ward. The “lifting process” in the quarrying of granite. Scientific American, 99(18):301–302, October 31, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n18/pdf/scientificamerican10311908-301.pdf.

Willey:1908:YID

[11590] Day Allen Willey. The Yuma Irrigation Dam. Scientific American, 99(18):302–303, October 31, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n18/pdf/scientificamerican10311908-
302a.pdf.

Anonymous:1908:Cbf

[11591] Anonymous. Correspondence. Scientific American, 99(19):315, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n19/pdf/scientificamerican11071908-315.pdf.

Anonymous:1908:RPIbi

[11592] Anonymous. Recently patented inventions. Scientific American, 99(19):321, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
321.pdf.

Anonymous:1908:NQx

[11593] Anonymous. Notes and queries. Scientific American, 99(19): 321–323, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
321a.pdf.

Anonymous:1908:BNBx

[11594] Anonymous. Books: New books, etc. Scientific American, 99 (19):323, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
323.pdf.

Anonymous:1908:IIbk

[11595] Anonymous. Index of inventions. Scientific American, 99(19): 323–327, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
323a.pdf.

Anonymous:1908:PAW

[11596] Anonymous. A precaution against water famine. Scientific American, 99(19):310, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
310.pdf.

Anonymous:1908:CMA

[11597] Anonymous. Carnegie Meridian Astrometry Expedition. Scientific American, 99(19):310, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
310a.pdf.

Anonymous:1908:DCM

[11598] Anonymous. Detection of carbon monoxide in the air. Scientific American, 99(19):310, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
310b.pdf.

Anonymous:1908:FAFb

[11599] Anonymous. The fallacy of artificial food. Scientific American, 99(19):310, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
310c.pdf.

Anonymous:1908:Ej

[11600] Anonymous. Engineering. Scientific American, 99(19):311, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n19/pdf/scientificamerican11071908-311.pdf.

Anonymous:1908:Ek

[11601] Anonymous. Electricity. Scientific American, 99(19):311, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n19/pdf/scientificamerican11071908-311a.pdf.

Anonymous:1908:Sd

[11602] Anonymous. Science. Scientific American, 99(19):311, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n19/pdf/scientificamerican11071908-311b.pdf.

Anonymous:1908:FVC

[11603] Anonymous. The Fourth Vanderbilt Cup Automobile Race. Scientific American, 99(19):312–313, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
312.pdf.

Anonymous:1908:MGS

[11604] Anonymous. A mammoth gyratory stone crusher. Scientific American, 99(19):314, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
314.pdf.

Anonymous:1908:WA

[11605] Anonymous. Was this an aerolite? Scientific American, 99 (19):315, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
315a.pdf.

Anonymous:1908:PWG

[11606] Anonymous. A paddle wheel for gliding boats. Scientific American, 99(19):315, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
315b.pdf.

Anonymous:1908:ESC

[11607] Anonymous. Electric scrubber for cleaning ships’ bottoms without drydocking. Scientific American, 99(19):316–317, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n19/pdf/scientificamerican11071908-316.pdf.

Springer:1908:MPS

[11608] J. F. Springer. Metallographs, or photographs of the structure of metal specimens. Scientific American, 99(19):317–318, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n19/pdf/scientificamerican11071908-317.pdf.

Lodian:1908:SCU

[11609] L. Lodian. Some commercial uses of the glass-sponges. Scientific American, 99(19):318, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
318.pdf.

Anonymous:1908:HMWb

[11610] Anonymous. Handy man’s workshop. Scientific American, 99 (19):319, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
319.pdf.

Anonymous:1908:PPB

[11611] Anonymous. A painter’s platform bracket. Scientific American, 99(19):319, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
319a.pdf.

Chapman:1908:SSA

[11612] H. D. Chapman. Screw-slotting attachment for lathes. Scientific American, 99(19):319, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
319b.pdf.

Anonymous:1908:GDH

[11613] Anonymous. A guide for drilling horizontal holes. Scientific American, 99(19):319, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
319c.pdf.

Bergstrom:1908:SPW

[11614] J. A. Bergstrom. Substitutes for a pipe wrench. Scientific American, 99(19):319, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
319d.pdf.

Searing:1908:OGP

[11615] A. V. Searing. Old glue pot as a metal pot and ladle. Scientific American, 99(19):319, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
319e.pdf.

Anonymous:1908:SDSb

[11616] Anonymous. Starting device for siphons. Scientific American, 99 (19):319, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
319f.pdf.

Anonymous:1908:RCN

[11617] Anonymous. Rag carpet needle. Scientific American, 99(19): 319–320, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
319g.pdf.

C:1908:HMV

[11618] W. J. C. Home-made vacuum cleaner. Scientific American, 99 (19):320, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
320.pdf.

Michaud:1908:SSP

[11619] Gustave Michaud. A stereoscope for single photographs. Scientific American, 99(19):320, November 7, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n19/pdf/scientificamerican11071908-
320a.pdf.

Anonymous:1908:RPIbj

[11620] Anonymous. Recently patented inventions. Scientific American, 99(2):32, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
32.pdf.

Anonymous:1908:NQy

[11621] Anonymous. Notes and queries. Scientific American, 99(2):32, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-32a.pdf.

Anonymous:1908:BNBy

[11622] Anonymous. Books: New books, etc. Scientific American, 99(2):32, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-32b.pdf.

Anonymous:1908:IIbl

[11623] Anonymous. Index of inventions. Scientific American, 99(2):32–35, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-32c.pdf.

Anonymous:1908:PNPbe

[11624] Anonymous. Patent notes: Patents. Scientific American, 99(2):36, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-36.pdf.

Anonymous:1908:SDSc

[11625] Anonymous. Side doors for subway cars. Scientific American, 99(2):22, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-22.pdf.

Anonymous:1908:GRN

[11626] Anonymous. Government restrictions of Niagara water power. Scientific American, 99(2):22, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
22a.pdf.

Anonymous:1908:MEE

[11627] Anonymous. Marine engine endurance. Scientific American, 99(2):22, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-22b.pdf.

Anonymous:1908:EEG

[11628] Anonymous. Extensive electrification of German railroads. Scientific American, 99(2):22, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
22c.pdf.

Anonymous:1908:QS

[11629] Anonymous. A question of speed. Scientific American, 99(2):22, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-22d.pdf.

Anonymous:1908:EPN

[11630] Anonymous. Electro-plating non-metallic articles. Scientific American, 99(2):22–23, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
22e.pdf.

Field:1908:FNS

[11631] C. Field. Further notes on the Star and Crescent. Scientific American, 99(2):23, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
23.pdf.

Anonymous:1908:KSF

[11632] Anonymous. The Kaiser’s Silver Flotilla at the Berlin Shipbuilding Exhibition. Scientific American, 99(2):23, July 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n2/pdf/scientificamerican07111908-23a.pdf.

Anonymous:1908:OMSd

[11633] Anonymous. Official Meteorological Summary, New York, N.Y., June, 1908. Scientific American, 99(2):23, July 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n2/pdf/scientificamerican07111908-23b.pdf.

Anonymous:1908:CSw

[11634] Anonymous. The current supplement. Scientific American, 99(2):23, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-23c.pdf.

Anonymous:1908:NAP

[11635] Anonymous. No award of the prize for humane slaughtering. Scientific American, 99(2):23, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
23d.pdf.

Anonymous:1908:STN

[11636] Anonymous. Successful test of New York’s New High-Pressure Fire Service. Scientific American, 99(2):24, July 11, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n2/pdf/scientificamerican07111908-24.pdf.

Anonymous:1908:TSC

[11637] Anonymous. The trials of the Scout cruiser “Salem”. Scientific American, 99(2):25–26, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
25.pdf.

Luyties:1908:EH

[11638] Otto G. Luyties. Experiments with a helicopter. Scientific American, 99(2):26–27, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
26.pdf.

Anonymous:1908:GTM

[11639] Anonymous. Government tests of mine explosives. Scientific American, 99(2):27–28, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
27.pdf.

Shepstone:1908:CMO

[11640] Harold J. Shepstone. A collection of models of Old London. Scientific American, 99(2):29–30, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
29.pdf.

Anonymous:1908:FTFb

[11641] Anonymous. The first trial flight of an aeroplane for the Scientific American Trophy. Scientific American, 99(2):30, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-30.pdf.

Anonymous:1908:BFH

[11642] Anonymous. Book for fish hooks. Scientific American, 99(2):31, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-31.pdf.

Anonymous:1908:APSb

[11643] Anonymous. Attachment for printing stamps. Scientific American, 99(2):31, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n2/pdf/scientificamerican07111908-
31a.pdf.

Anonymous:1908:HG

[11644] Anonymous. Hook for garments. Scientific American, 99(2):31, July 11, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n2/pdf/scientificamerican07111908-31b.pdf.

Anonymous:1908:Cbg

[11645] Anonymous. Correspondence. Scientific American, 99(20):335, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n20/pdf/scientificamerican11141908-335.pdf.

Anonymous:1908:RPIbk

[11646] Anonymous. Recently patented inventions. Scientific American, 99(20):341, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
341.pdf.

Anonymous:1908:NQz

[11647] Anonymous. Notes and queries. Scientific American, 99(20): 341–343, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
341a.pdf.

Anonymous:1908:BNBz

[11648] Anonymous. Books: New books, etc. Scientific American, 99 (20):343, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
343.pdf.

Anonymous:1908:IIbm

[11649] Anonymous. Index of inventions. Scientific American, 99(20): 343–347, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
343a.pdf.

Anonymous:1908:AUE

[11650] Anonymous. Accuracy and uniformity in electrical standards. Scientific American, 99(20):330, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
330.pdf.

Anonymous:1908:TVT

[11651] Anonymous. The tactical value of torpedo craft. Scientific American, 99(20):330, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
330a.pdf.

Anonymous:1908:El

[11652] Anonymous. Engineering. Scientific American, 99(20):331, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n20/pdf/scientificamerican11141908-331.pdf.

Anonymous:1908:Em

[11653] Anonymous. Electricity. Scientific American, 99(20):331, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n20/pdf/scientificamerican11141908-331a.pdf.

Anonymous:1908:Se

[11654] Anonymous. Science. Scientific American, 99(20):331, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n20/pdf/scientificamerican11141908-331b.pdf.

Wade:1908:STN

[11655] Herbert T. Wade. Successful test of New York’s High-Pressure Fire Service. Scientific American, 99(20):332, November 14, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n20/pdf/scientificamerican11141908-332.pdf.

Winley:1908:MFF

[11656] Day Allen Winley. The menace of the forest fire. Scientific American, 99(20):333–334, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
333.pdf.

Anonymous:1908:LOF

[11657] Anonymous. The launch of our first “Dreadnought”. Scientific American, 99(20):334–335, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
334.pdf.

Anonymous:1908:MCB

[11658] Anonymous. A multiple cycle for the blind. Scientific American, 99(20):336, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
336.pdf.

Anonymous:1908:AAF

[11659] Anonymous. Action aloft; or fighting tops, past and passing. Scientific American, 99(20):336–338, November 14, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n20/pdf/scientificamerican11141908-336a.pdf.

Anonymous:1908:ASF

[11660] Anonymous. The Aeronautic Society’s First Exhibition. Scientific American, 99(20):338, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
338.pdf.

Anonymous:1908:PDj

[11661] Anonymous. Patent department. Scientific American, 99(20): 339, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
339.pdf.

Anonymous:1908:CPT

[11662] Anonymous. Combined poncho and tent. Scientific American, 99 (20):339, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
339a.pdf.

Anonymous:1908:PTSb

[11663] Anonymous. Plow for turning the surface soil under the subsoil. Scientific American, 99(20):339, November 14, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n20/pdf/scientificamerican11141908-339b.pdf.

Anonymous:1908:IW

[11664] Anonymous. An improved winch. Scientific American, 99(20): 339, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
339c.pdf.

Anonymous:1908:CSLb

[11665] Anonymous. Collapsible step-ladder. Scientific American, 99 (20):339, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
339d.pdf.

Anonymous:1908:BBS

[11666] Anonymous. Brake for bob sleds. Scientific American, 99(20): 340, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
340.pdf.

Anonymous:1908:WFC

[11667] Anonymous. Waterproof fuse cap. Scientific American, 99(20): 340, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
340a.pdf.

Anonymous:1908:NTP

[11668] Anonymous. Novel type of plow. Scientific American, 99(20): 340, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
340b.pdf.

Anonymous:1908:OIe

[11669] Anonymous. Oddities in invention. Scientific American, 99(20): 340, November 14, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n20/pdf/scientificamerican11141908-
340c.pdf.

Anonymous:1908:RPIbl

[11670] Anonymous. Recently patented inventions. Scientific American, 99(21):361, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
361.pdf.

Anonymous:1908:NQba

[11671] Anonymous. Notes and queries. Scientific American, 99(21): 361–362, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
361a.pdf.

Anonymous:1908:BNBba

[11672] Anonymous. Books: New books, etc. Scientific American, 99 (21):362, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
362.pdf.

Anonymous:1908:IIbn

[11673] Anonymous. Index of inventions. Scientific American, 99(21): 362–367, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
362a.pdf.

Anonymous:1908:MBb

[11674] Anonymous. A monumental blunder. Scientific American, 99 (21):350, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
350.pdf.

Anonymous:1908:NEAb

[11675] Anonymous. A new era in aeroplane transport. Scientific American, 99(21):350, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
350a.pdf.

Anonymous:1908:WWR

[11676] Anonymous. Wilbur Wright’s record for height. Scientific American, 99(21):350, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
350b.pdf.

Anonymous:1908:HMWc

[11677] Anonymous. The handy man’s workshop. Scientific American, 99 (21):350, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
350c.pdf.

Anonymous:1908:En

[11678] Anonymous. Engineering. Scientific American, 99(21):351, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n21/pdf/scientificamerican11211908-351.pdf.

Anonymous:1908:Eo

[11679] Anonymous. Electricity. Scientific American, 99(21):351, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n21/pdf/scientificamerican11211908-351a.pdf.

Anonymous:1908:Sf

[11680] Anonymous. Science. Scientific American, 99(21):351, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n21/pdf/scientificamerican11211908-351b.pdf.

Grimshaw:1908:NFT

[11681] Robert Grimshaw. A new form of telautograph. Scientific American, 99(21):352, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
352.pdf.

Mitchell:1908:CS

[11682] S. A. Mitchell. Cyclones on the Sun. Scientific American, 99(21): 353–355, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
353.pdf.

Anonymous:1908:Cbh

[11683] Anonymous. Correspondence. Scientific American, 99(21):355, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n21/pdf/scientificamerican11211908-355.pdf.

Boyer:1908:AF

[11684] Jacques Boyer. An aeroplane factory. Scientific American, 99 (21):356, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
356.pdf.

Anonymous:1908:LEA

[11685] Anonymous. The latest European airships and aeroplanes and their performances. Scientific American, 99(21):357–359, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n21/pdf/scientificamerican11211908-357.pdf.

Anonymous:1908:HMWd

[11686] Anonymous. Handy man’s workshop. Scientific American, 99 (21):359, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
359.pdf.

Johns:1908:EMP

[11687] B. A. Johns. Ejector made out of pipe fittings. Scientific American, 99(21):359, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
359a.pdf.

Harrison:1908:SBC

[11688] Walton Harrison. Storage battery without chemicals. Scientific American, 99(21):359, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
359b.pdf.

Tiede:1908:CHS

[11689] Edward J. Tiede. Convenient holder for sandpaper. Scientific American, 99(21):359, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
359c.pdf.

Bayley:1908:CW

[11690] I. C. Bayley. The construction of a workshop. Scientific American, 99(21):359–360, November 21, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n21/pdf/scientificamerican11211908-
359d.pdf.

Anonymous:1908:Cbi

[11691] Anonymous. Correspondence. Scientific American, 99(22):374, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n22/pdf/scientificamerican11281908-374.pdf.

Russell:1908:HD

[11692] Henry Norris Russell. The heavens in December. Scientific American, 99(22):375, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
375.pdf.

Anonymous:1908:RPIbm

[11693] Anonymous. Recently patented inventions. Scientific American, 99(22):382, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
382.pdf.

Anonymous:1908:NQbb

[11694] Anonymous. Notes and queries. Scientific American, 99(22): 382–383, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
382a.pdf.

Anonymous:1908:BNBbb

[11695] Anonymous. Books: New books, etc. Scientific American, 99(22): 383–384, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
383.pdf.

Anonymous:1908:IIbo

[11696] Anonymous. Index of inventions. Scientific American, 99(22): 384–387, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
384.pdf.

Anonymous:1908:THT

[11697] Anonymous. Technical heads for technical city departments. Scientific American, 99(22):370, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
370.pdf.

Anonymous:1908:MRM

[11698] Anonymous. Making rocks and minerals artificially. Scientific American, 99(22):370, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
370a.pdf.

Anonymous:1908:Eq

[11699] Anonymous. Engineering. Scientific American, 99(22):371, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n22/pdf/scientificamerican11281908-371.pdf.

Anonymous:1908:Er

[11700] Anonymous. Electricity. Scientific American, 99(22):371, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n22/pdf/scientificamerican11281908-371a.pdf.

Anonymous:1908:Sg

[11701] Anonymous. Science. Scientific American, 99(22):371, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n22/pdf/scientificamerican11281908-371b.pdf.

Honey:1908:ESL

[11702] Frederic R. Honey. The eclipse of the Sun and lunar appulse in December, 1908. Scientific American, 99(22):372, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n22/pdf/scientificamerican11281908-372.pdf.

Anonymous:1908:LND

[11703] Anonymous. Launch of the “North Dakota”. Scientific American, 99(22):372, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
372a.pdf.

Fitzpatrick:1908:CSP

[11704] J. A. Fitzpatrick. Is concrete steel a permanent construction? Scientific American, 99(22):373, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
373.pdf.

Anonymous:1908:HIF

[11705] Anonymous. How India is fighting the plague — the manufacture and use of anti-plague vaccine. Scientific American, 99(22):373–374, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n22/pdf/scientificamerican11281908-373a.pdf.

Boyer:1908:AH

[11706] Jacques Boyer. An automobile hoe. Scientific American, 99(22): 376–377, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
376.pdf.

Claudy:1908:GLO

[11707] C. H. Claudy. The guiding lights of our coasts. Scientific American, 99(22):377, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
377.pdf.

Anonymous:1908:HW

[11708] Anonymous. Handyman’s workshop. Scientific American, 99(22): 378, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
378.pdf.

Naylor:1908:RMC

[11709] George W. Naylor. Roller mounting for the Christmas tree. Scientific American, 99(22):378, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
378a.pdf.

Searing:1908:FTS

[11710] A. V. Searing. Folding tree stand. Scientific American, 99 (22):378, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
378b.pdf.

L:1908:PCT

[11711] H. G. L. To prevent the Christmas tree from upsetting. Scientific American, 99(22):378, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
378c.pdf.

Handy:1908:CTG

[11712] L. Gessford Handy. A Christmas tree garden. Scientific American, 99(22):379, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
379.pdf.

Rutherford:1908:PCT

[11713] G. H. Rutherford. Protection for Christmas tree candles. Scientific American, 99(22):379, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
379a.pdf.

Anonymous:1908:SHM

[11714] Anonymous. Some home-made Christmas presents. Scientific American, 99(22):379, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
379b.pdf.

C:1908:CPF

[11715] W. J. C. Copper-plating flowers and other perishable articles. Scientific American, 99(22):379–380, November 28, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n22/pdf/scientificamerican11281908-379c.pdf.

Davison:1908:OCF

[11716] Ralph C. Davison. Ornamental concrete flower pots and how to make them. Scientific American, 99(22):380, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n22/pdf/scientificamerican11281908-380.pdf.

Johns:1908:HMM

[11717] B. A. Johns. Home-made metal lamp shade. Scientific American, 99 (22):380–381, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
380a.pdf.

Anonymous:1908:CEa

[11718] Anonymous. The Christmas entertainment. Scientific American, 99(22):381, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
381.pdf.

Michaud:1908:TT

[11719] Gustave Michaud. A trick with tourmaline. Scientific American, 99(22):381, November 28, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n22/pdf/scientificamerican11281908-
381a.pdf.

Anonymous:1908:IIbp

[11720] Anonymous. Index of inventions. Scientific American, 99(23): 418–423, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
418.pdf.

Anonymous:1908:ECa

[11721] Anonymous. The ethics of criticism. Scientific American, 99 (23):392, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
392.pdf.

Anonymous:1908:SEL

[11722] Anonymous. Some early light-weight motors. Scientific American, 99(23):392, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
392a.pdf.

Anonymous:1908:SIY

[11723] Anonymous. Schooners for international yacht racing. Scientific American, 99(23):392, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
392b.pdf.

Anonymous:1908:GPA

[11724] Anonymous. The Grand Prize Automobile Road Race and its probable successor. Scientific American, 99(23):392, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n23/pdf/scientificamerican12051908-392c.pdf.

Anonymous:1908:Et

[11725] Anonymous. Engineering. Scientific American, 99(23):393, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n23/pdf/scientificamerican12051908-393.pdf.

Anonymous:1908:Eu

[11726] Anonymous. Electricity. Scientific American, 99(23):393, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n23/pdf/scientificamerican12051908-393a.pdf.

Anonymous:1908:AA

[11727] Anonymous. Aeronautical and automobile. Scientific American, 99(23):393, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
393b.pdf.

Anonymous:1908:WSb

[11728] Anonymous. Water supply. Scientific American, 99(23):394–396, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n23/pdf/scientificamerican12051908-394.pdf.

Anonymous:1908:CWS

[11729] Anonymous. The Catskill Water Supply. Scientific American, 99 (23):396, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
396.pdf.

Anonymous:1908:CGB

[11730] Anonymous. The city’s giant bridges. Scientific American, 99 (23):397, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
397.pdf.

Anonymous:1908:WSBa

[11731] Anonymous. The Williamsburg Suspension Bridge. Scientific American, 99(23):397, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
397a.pdf.

Anonymous:1908:MSBa

[11732] Anonymous. The Manhattan Suspension Bridge. Scientific American, 99(23):397–400, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
397b.pdf.

Anonymous:1908:QCB

[11733] Anonymous. The Queensborough Cantilever Bridge. Scientific American, 99(23):400, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
400.pdf.

Anonymous:1908:TBN

[11734] Anonymous. Tall buildings of New York. Scientific American, 99 (23):400–403, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
400a.pdf.

Anonymous:1908:TS

[11735] Anonymous. Tunnels and subways. Scientific American, 99(23): 404–407, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
404.pdf.

Anonymous:1908:HDI

[11736] Anonymous. Harbor and dock improvements. Scientific American, 99(23):408, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
408.pdf.

Anonymous:1908:RT

[11737] Anonymous. Railway terminals. Scientific American, 99(23):409, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n23/pdf/scientificamerican12051908-409.pdf.

Anonymous:1908:PRT

[11738] Anonymous. Pennsylvania Railroad Terminal Station. Scientific American, 99(23):409–410, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
409a.pdf.

Anonymous:1908:NGCb

[11739] Anonymous. New Grand Central Terminal Station. Scientific American, 99(23):410–412, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
410.pdf.

Anonymous:1908:FDS

[11740] Anonymous. Freight distribution by subway. Scientific American, 99(23):413, December 5, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n23/pdf/scientificamerican12051908-
413.pdf.

Anonymous:1908:Cbj

[11741] Anonymous. Correspondence. Scientific American, 99(24):431, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n24/pdf/scientificamerican12121908-431.pdf.

Anonymous:1908:RPIbn

[11742] Anonymous. Recently patented inventions. Scientific American, 99 (24):437–438, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
437.pdf.

Anonymous:1908:NQbc

[11743] Anonymous. Notes and queries. Scientific American, 99(24): 438–439, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
438.pdf.

Anonymous:1908:BNBbc

[11744] Anonymous. Books: New books, etc. Scientific American, 99 (24):439, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
439.pdf.

Anonymous:1908:IIbq

[11745] Anonymous. Index of inventions. Scientific American, 99(24): 439–443, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
439a.pdf.

Anonymous:1908:LPF

[11746] Anonymous. Legislation to prevent forest fires. Scientific American, 99(24):426, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
426.pdf.

Anonymous:1908:NCA

[11747] Anonymous. Newport Conference approves designs of new battleships. Scientific American, 99(24):426, December 12, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n24/pdf/scientificamerican12121908-426a.pdf.

Anonymous:1908:DSA

[11748] Anonymous. Detection of sulphurous acid in foodstuffs. Scientific American, 99(24):426, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
426b.pdf.

Anonymous:1908:Ev

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n24/pdf/scientificamerican12121908-427.pdf.

Anonymous:1908:Ew

[11750] Anonymous. Electricity. Scientific American, 99(24):427, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n24/pdf/scientificamerican12121908-427a.pdf.

Anonymous:1908:Sh

[11751] Anonymous. Science. Scientific American, 99(24):427, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n24/pdf/scientificamerican12121908-427b.pdf.

Anonymous:1908:BBM

[11752] Anonymous. Brazilian battleship “Minas Geraes” — most powerful fighting ship afloat. Scientific American, 99(24):428–429, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n24/pdf/scientificamerican12121908-428.pdf.

Anonymous:1908:HDS

[11753] Anonymous. A hydraulic drive system. Scientific American, 99 (24):429–430, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
429.pdf.

Mills:1908:NDA

[11754] James Cooke Mills. New discoveries about lightning. Scientific American, 99(24):430, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
430.pdf.

Anonymous:1908:PPC

[11755] Anonymous. Progress on the Panama Canal. Scientific American, 99 (24):431–433, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
431a.pdf.

Anonymous:1908:SDL

[11756] Anonymous. Santos Dumont’s latest aeroplane. Scientific American, 99(24):433, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
433.pdf.

Anonymous:1908:FAAa

[11757] Anonymous. The fourth aeroplane of the Aerial Experiment Association. Scientific American, 99(24):433–434, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n24/pdf/scientificamerican12121908-433a.pdf.

Aaron:1908:MOD

[11758] S. Frank Aaron. Mechanism of the ovipositor drill. Scientific American, 99(24):435, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
435.pdf.

Anonymous:1908:PDk

[11759] Anonymous. Patent department. Scientific American, 99(24): 436, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
436.pdf.

Anonymous:1908:GWM

[11760] Anonymous. Grain-washing machine. Scientific American, 99 (24):436, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
436a.pdf.

Anonymous:1908:NBP

[11761] Anonymous. Novel boat-propelling device. Scientific American, 99 (24):436, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
436b.pdf.

Anonymous:1908:BMWb

[11762] Anonymous. Brake mechanism for wagons. Scientific American, 99(24):436, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
436c.pdf.

Anonymous:1908:IRT

[11763] Anonymous. Improved railroad tie. Scientific American, 99(24): 436, December 12, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n24/pdf/scientificamerican12121908-
436d.pdf.

Anonymous:1908:RPIbo

[11764] Anonymous. Recently patented inventions. Scientific American, 99(25):457, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
457.pdf.

Anonymous:1908:NQbd

[11765] Anonymous. Notes and queries. Scientific American, 99(25): 457–458, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
457a.pdf.

Anonymous:1908:BNBbd

[11766] Anonymous. Books: New books, etc. Scientific American, 99 (25):458, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
458.pdf.

Anonymous:1908:PNPbf

[11767] Anonymous. Patent notes: Patent causes. Scientific American, 99(25):458, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
458a.pdf.

Anonymous:1908:IIbr

[11768] Anonymous. Index of inventions. Scientific American, 99(25): 458–463, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
458b.pdf.

Anonymous:1908:PRW

[11769] Anonymous. President Roosevelt would reorganize the Navy Department. Scientific American, 99(25):446, December 19, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n25/pdf/scientificamerican12191908-446.pdf.

Anonymous:1908:TCI

[11770] Anonymous. Tribute to a civilian inventor. Scientific American, 99(25):446, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
446a.pdf.

Anonymous:1908:Ac

[11771] Anonymous. Aeronautics. Scientific American, 99(25):447, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n25/pdf/scientificamerican12191908-447.pdf.

Anonymous:1908:Ex

[11772] Anonymous. Electricity. Scientific American, 99(25):447, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n25/pdf/scientificamerican12191908-447a.pdf.

Anonymous:1908:SNf

[11773] Anonymous. Science notes. Scientific American, 99(25):447, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n25/pdf/scientificamerican12191908-447b.pdf.

Hislam:1908:BPB

[11774] Percival Hislam. Battle practice in the British Navy. Scientific American, 99(25):448–449, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
448.pdf.

Coleman:1908:HER

[11775] F. C. Coleman. How English royalty travels. Scientific American, 99(25):449, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
449.pdf.

Anonymous:1908:AMSb

[11776] Anonymous. The aeroplane military Scout. Scientific American, 99(25):450, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
450.pdf.

Anonymous:1908:CFAc

[11777] Anonymous. Curious facts about numbers. Scientific American, 99(25):451, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
451.pdf.

Anonymous:1908:GE

[11778] Anonymous. Gun erosion. Scientific American, 99(25):451, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n25/pdf/scientificamerican12191908-451a.pdf.

Anonymous:1908:AS

[11779] Anonymous. Artificial sapphires. Scientific American, 99(25): 452, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
452.pdf.

Boyer:1908:NMT

[11780] Jacques Boyer. New mechanical toys. Scientific American, 99(25): 453–454, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
453.pdf.

Anonymous:1908:DFW

[11781] Anonymous. Destructive fungus of white cedar. Scientific American, 99(25):454–455, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
454.pdf.

Anonymous:1908:Ad

[11782] Anonymous. Announcement. Scientific American, 99(25):455, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n25/pdf/scientificamerican12191908-455.pdf.

Whitemore:1908:LCO

[11783] H. L. Whitemore. Let the clock open the furnace drafts. Scientific American, 99(25):455, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
455a.pdf.

Honey:1908:HUS

[11784] Frederic R. Honey. How to use the slide rule. Scientific American, 99(25):455–456, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
455b.pdf.

Handy:1908:SDS

[11785] L. G. Handy. Simple driver for small drills. Scientific American, 99(25):456, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
456.pdf.

Brouillet:1908:DHM

[11786] J. O. Brouillet. Drilling holes in marbles. Scientific American, 99(25):456, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
456a.pdf.

Bayley:1908:FW

[11787] I. G. Bayley. Furnishing the workshop. Scientific American, 99 (25):456, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
456b.pdf.

Sweet:1908:DFC

[11788] F. D. Sweet. Device for finding centers of round work. Scientific American, 99(25):456, December 19, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n25/pdf/scientificamerican12191908-
456c.pdf.

Anonymous:1908:Cbk

[11789] Anonymous. Correspondence. Scientific American, 99(26):471, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n26/pdf/scientificamerican12261908-471.pdf.

Anonymous:1908:RPIbp

[11790] Anonymous. Recently patented inventions. Scientific American, 99(26):476, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
476.pdf.

Anonymous:1908:NQbe

[11791] Anonymous. Notes and queries. Scientific American, 99(26): 476–478, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
476a.pdf.

Anonymous:1908:BNBbe

[11792] Anonymous. Books: New books, etc. Scientific American, 99 (26):478, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
478.pdf.

Anonymous:1908:IIbs

[11793] Anonymous. Index of inventions. Scientific American, 99(26): 478–480, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
478a.pdf.

Anonymous:1908:I

[11794] Anonymous. Index. Scientific American, 99(26):481–482, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n26/pdf/scientificamerican12261908-481.pdf.

Anonymous:1908:GDF

[11795] Anonymous. Gatun Dam facts perverted. Scientific American, 99 (26):466, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
466.pdf.

Anonymous:1908:FES

[11796] Anonymous. Fashions in electric systems. Scientific American, 99 (26):466, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
466a.pdf.

Anonymous:1908:MSBb

[11797] Anonymous. Manhattan suspension bridge to be investigated. Scientific American, 99(26):466–467, December 26, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n26/pdf/scientificamerican12261908-466b.pdf.

Anonymous:1908:Ey

[11798] Anonymous. Engineering. Scientific American, 99(26):467, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n26/pdf/scientificamerican12261908-467.pdf.

Anonymous:1908:Ez

[11799] Anonymous. Electricity. Scientific American, 99(26):467, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n26/pdf/scientificamerican12261908-467a.pdf.

Anonymous:1908:Si

[11800] Anonymous. Science. Scientific American, 99(26):467, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n26/pdf/scientificamerican12261908-467b.pdf.

Hislam:1908:TS

[11801] P. A. Hislam. The transportation of submarines. Scientific American, 99(26):468, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
468.pdf.

Anonymous:1908:FEH

[11802] Anonymous. Farman’s experiments with his triple-surface aeroplane. Scientific American, 99(26):468–469, December 26, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n26/pdf/scientificamerican12261908-468a.pdf.

Gradenwitz:1908:MIS

[11803] Alfred Gradenwitz. Models for illustrating the strain on structures. Scientific American, 99(26):469–470, December 26, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n26/pdf/scientificamerican12261908-469.pdf.

Anonymous:1908:MSI

[11804] Anonymous. Magnitude of the salt industry. Scientific American, 99 (26):470–471, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
470.pdf.

Anonymous:1908:DWS

[11805] Anonymous. A defense of the Wright system of propellers. Scientific American, 99(26):471–472, December 26, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n26/pdf/scientificamerican12261908-471a.pdf.

Willey:1908:NSC

[11806] Day Allen Willey. A novel system of concrete construction. Scientific American, 99(26):472–473, December 26, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n26/pdf/scientificamerican12261908-472.pdf.

Kretschmer:1908:FSB

[11807] Otto Kretschmer. Fast steamers built on “Tetrahedral” lines. Scientific American, 99(26):473–474, December 26, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n26/pdf/scientificamerican12261908-473.pdf.

Keane:1908:NIM

[11808] Arthur H. J. Keane. A new and interesting motor cycle. Scientific American, 99(26):474–475, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
474.pdf.

Anonymous:1908:NIG

[11809] Anonymous. A new incandescent gas mantle. Scientific American, 99(26):475, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
475.pdf.

Marriott:1908:MC

[11810] Crittenden Marriott. The measles cannibal. Scientific American, 99(26):475, December 26, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n26/pdf/scientificamerican12261908-
475a.pdf.

Anonymous:1908:Cbl

[11811] Anonymous. Correspondence. Scientific American, 99(3):43, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-43.pdf.

Anonymous:1908:RPIbq

[11812] Anonymous. Recently patented inventions. Scientific American, 99(3):48, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
48.pdf.

Anonymous:1908:IIbt

[11813] Anonymous. Index of inventions. Scientific American, 99(3):48–51, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-48a.pdf.

Anonymous:1908:PNPbg

[11814] Anonymous. Patent notes: Patents. Scientific American, 99(3):52, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-52.pdf.

Anonymous:1908:SAA

[11815] Anonymous. The Scientific American aeronautical trophy. Scientific American, 99(3):38, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
38.pdf.

Anonymous:1908:FKL

[11816] Anonymous. The first 25-knot liner. Scientific American, 99(3):38, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-38a.pdf.

Anonymous:1908:AWR

[11817] Anonymous. Aeronautics in war — a Reductio ad Absurdum. Scientific American, 99(3):38, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
38b.pdf.

Anonymous:1908:PQN

[11818] Anonymous. Peary’s quest of the North Pole. Scientific American, 99(3):38, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
38c.pdf.

Anonymous:1908:TSD

[11819] Anonymous. Test of safety devices by the Public Service Commission. Scientific American, 99(3):38–39, July 18, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n3/pdf/scientificamerican07181908-38d.pdf.

Anonymous:1908:PPD

[11820] Anonymous. The present position of the Darwinian theory. Scientific American, 99(3):39, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
39.pdf.

Anonymous:1908:SNh

[11821] Anonymous. Science notes. Scientific American, 99(3):39, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-39a.pdf.

Anonymous:1908:CSx

[11822] Anonymous. The current supplement. Scientific American, 99(3):39, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-39b.pdf.

Bayley:1908:NMS

[11823] I. G. Bayley. A novel music stand or book rest. Scientific American, 99(3):40, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
40.pdf.

Anonymous:1908:MHC

[11824] Anonymous. A machine for handling coke in the yard. Scientific American, 99(3):40, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
40a.pdf.

Mills:1908:MMM

[11825] James Cooke Mills. A multiple molding machine. Scientific American, 99(3):41, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
41.pdf.

Coleman:1908:NBL

[11826] F. C. Coleman. A new British locomotive. Scientific American, 99(3):41–42, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
41a.pdf.

Anonymous:1908:SP

[11827] Anonymous. A selenium photometer. Scientific American, 99(3):42, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-42.pdf.

Anonymous:1908:NFG

[11828] Anonymous. The new French and German airships “République” and “Zeppelin IV”. Scientific American, 99(3):43–44, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-43a.pdf.

Anonymous:1908:FSF

[11829] Anonymous. First successful flights of Blériot’s no. 8 monoplane. Scientific American, 99(3):44, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
44.pdf.

Anonymous:1908:WFJ

[11830] Anonymous. The winning flight of the “June Bug” aeroplane for the Scientific American Trophy. Scientific American, 99(3):45, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-45.pdf.

Anonymous:1908:HS

[11831] Anonymous. Hatching snakes. Scientific American, 99(3):46, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-46.pdf.

Anonymous:1908:PDl

[11832] Anonymous. Patent department. Scientific American, 99(3):47, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n3/pdf/scientificamerican07181908-47.pdf.

Foster:1908:HMR

[11833] C. M. Foster. Home-made rotary printing press built of junk. Scientific American, 99(3):47, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
47a.pdf.

Anonymous:1908:RBSb

[11834] Anonymous. Rope bracket for swing seats. Scientific American, 99(3):47, July 18, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n3/pdf/scientificamerican07181908-
47b.pdf.

Anonymous:1908:Cbm

[11835] Anonymous. Correspondence. Scientific American, 99(4):59, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n4/pdf/scientificamerican07251908-59a.pdf.

Anonymous:1908:RPIbr

[11836] Anonymous. Recently patented inventions. Scientific American, 99(4):64, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n4/pdf/scientificamerican07251908-
64.pdf.

Anonymous:1908:IIbu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n4/pdf/scientificamerican07251908-64a.pdf.

Anonymous:1908:PNPbh

[11838] Anonymous. Patent notes: Patents. Scientific American, 99(4):68, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n4/pdf/scientificamerican07251908-68.pdf.

Anonymous:1908:WWM

[11839] Anonymous. The wreck of the White Mountain Express. Scientific American, 99(4):54, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n4/pdf/scientificamerican07251908-
54.pdf.

Anonymous:1908:PCM

[11840] Anonymous. The pros and cons of the marine gas engine. Scientific American, 99(4):54, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n4/pdf/scientificamerican07251908-
54a.pdf.

Anonymous:1908:FAAb

[11841] Anonymous. The Fifth Annual Automobile Tour for the Glidden Trophy. Scientific American, 99(4):54–55, July 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n4/pdf/scientificamerican07251908-54b.pdf.

Mitchell:1908:SRT

[11842] S. A. Mitchell. Some results from the 1908 total solar eclipse. Scientific American, 99(4):55, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n4/pdf/scientificamerican07251908-
55.pdf.

Stockard:1908:HDE

[11843] Charles R. Stockard. How does an egg develop into an animal? Scientific American, 99(4):56–58, July 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n4/pdf/scientificamerican07251908-56.pdf.

Anonymous:1908:CSDa

[11844] Anonymous. Comparison of a steam-driven and a gas-driven battleship of the same size and speed. Scientific American, 99(4):58–59, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n4/pdf/scientificamerican07251908-58.pdf.

Anonymous:1908:LHT

[11845] Anonymous. Limit of height for tall buildings. Scientific American, 99(4):59, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n4/pdf/scientificamerican07251908-
59.pdf.

Anonymous:1908:WSC

[11846] Anonymous. The widening of the Suez Canal. Scientific American, 99(4):60–62, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n4/pdf/scientificamerican07251908-
60.pdf.

Anonymous:1908:PGO

[11847] Anonymous. Preserving grand opera records for future generations. Scientific American, 99(4):62, July 25, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n4/pdf/scientificamerican07251908-62.pdf.

Boyer:1908:SFB

[11848] Jacques Boyer. The soaring flight of birds attained mechanically. Scientific American, 99(4):63, July 25, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n4/pdf/scientificamerican07251908-
63.pdf.

Russell:1908:HAb

[11849] Henry Norris Russell. The heavens in August. Scientific American, 99(5):71, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
71.pdf.

Anonymous:1908:Cbn

[11850] Anonymous. Correspondence. Scientific American, 99(5):75, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n5/pdf/scientificamerican08011908-75.pdf.

Anonymous:1908:RPIbs

[11851] Anonymous. Recently patented inventions. Scientific American, 99(5):80, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
80.pdf.

Anonymous:1908:NQbf

[11852] Anonymous. Notes and queries. Scientific American, 99(5):80, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n5/pdf/scientificamerican08011908-80a.pdf.

Anonymous:1908:BNBbf

[11853] Anonymous. Books: New books, etc. Scientific American, 99(5):80, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n5/pdf/scientificamerican08011908-80b.pdf.

Anonymous:1908:IIbv

[11854] Anonymous. Index of inventions. Scientific American, 99(5):80–83, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n5/pdf/scientificamerican08011908-80c.pdf.

Anonymous:1908:PNPbi

[11855] Anonymous. Patent notes: Patents. Scientific American, 99(5):84, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n5/pdf/scientificamerican08011908-84.pdf.

Anonymous:1908:PNYb

[11856] Anonymous. The pollution of New York Harbor. Scientific American, 99(5):70, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
70.pdf.

Anonymous:1908:GWR

[11857] Anonymous. Good work in the rainy season at Panama. Scientific American, 99(5):70, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
70a.pdf.

Anonymous:1908:MDRb

[11858] Anonymous. More daylight for recreation. Scientific American, 99(5):70, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
70b.pdf.

Anonymous:1908:RGP

[11859] Anonymous. Relation of the gas producer to low-grade fuels. Scientific American, 99(5):70, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
70c.pdf.

Anonymous:1908:NPMc

[11860] Anonymous. A new process for the manufacture of copper wire. Scientific American, 99(5):70, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
70d.pdf.

Anonymous:1908:IMO

[11861] Anonymous. Immortality of micro-organisms. Scientific American, 99(5):71, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
71a.pdf.

Bastin:1908:SHM

[11862] S. Leonard Bastin. Some home-made weather instruments. Scientific American, 99(5):72, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
72.pdf.

Willey:1908:IYV

[11863] Day Allen Willey. Irrigation of the Yuma Valley. Scientific American, 99(5):73–74, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
73.pdf.

Anonymous:1908:ECF

[11864] Anonymous. Erecting the construction footways of the Manhattan Bridge. Scientific American, 99(5):76–77, August 1, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n5/pdf/scientificamerican08011908-76.pdf.

Anonymous:1908:WNC

[11865] Anonymous. The work of a Nebraska cyclone. Scientific American, 99(5):78, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
78.pdf.

Anonymous:1908:PDm

[11866] Anonymous. Patent department. Scientific American, 99(5):79, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n5/pdf/scientificamerican08011908-79.pdf.

Anonymous:1908:ILC

[11867] Anonymous. Improved lock for cars. Scientific American, 99(5):79, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n5/pdf/scientificamerican08011908-79a.pdf.

Anonymous:1908:NRB

[11868] Anonymous. Non-refillable bottle. Scientific American, 99(5):79, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n5/pdf/scientificamerican08011908-79b.pdf.

Anonymous:1908:ERO

[11869] Anonymous. Elevator for raising ore concentrates. Scientific American, 99(5):79, August 1, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n5/pdf/scientificamerican08011908-
79c.pdf.

Anonymous:1908:RPIbt

[11870] Anonymous. Recently patented inventions. Scientific American, 99(6):96, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n6/pdf/scientificamerican08081908-
96.pdf.

Anonymous:1908:NQbg

[11871] Anonymous. Notes and queries. Scientific American, 99(6):96, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-96a.pdf.

Anonymous:1908:BNBbg

[11872] Anonymous. Books: New books, etc. Scientific American, 99(6):97, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-97.pdf.

Anonymous:1908:IIbw

[11873] Anonymous. Index of inventions. Scientific American, 99(6):97–99, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-97a.pdf.

Anonymous:1908:PNPbj

[11874] Anonymous. Patent notes: Patents. Scientific American, 99(6):100, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-100.pdf.

Anonymous:1908:RSA

[11875] Anonymous. The railroad spike — an anachronism of the Twentieth Century. Scientific American, 99(6):86, August 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n6/pdf/scientificamerican08081908-86.pdf.

Anonymous:1908:THF

[11876] Anonymous. A talk with Henry Farman. Scientific American, 99(6):86–87, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n6/pdf/scientificamerican08081908-
86a.pdf.

Anonymous:1908:CTc

[11877] Anonymous. Clothing for the tropics. Scientific American, 99(6):87, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-87.pdf.

Anonymous:1908:JE

[11878] Anonymous. The Japanese Exposition. Scientific American, 99(6):87, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-87a.pdf.

Anonymous:1908:CIV

[11879] Anonymous. The calendar of the inhabitants of Venus. Scientific American, 99(6):87, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n6/pdf/scientificamerican08081908-
87b.pdf.

Anonymous:1908:SNi

[11880] Anonymous. Science notes. Scientific American, 99(6):87, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-87c.pdf.

Anonymous:1908:CSy

[11881] Anonymous. The current supplement. Scientific American, 99(6):87, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-87d.pdf.

Anonymous:1908:ONQ

[11882] Anonymous. Our “Notes and Queries” column. Scientific American, 99(6):87, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n6/pdf/scientificamerican08081908-
87e.pdf.

Anonymous:1908:MPT

[11883] Anonymous. A machine for piling timber. Scientific American, 99(6):88, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-88.pdf.

Anonymous:1908:ASW

[11884] Anonymous. Artificial sea water. Scientific American, 99(6):88, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-88a.pdf.

Springer:1908:MC

[11885] J. F. Springer. The mystery of comets. Scientific American, 99(6):89–90, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n6/pdf/scientificamerican08081908-89.pdf.

Anonymous:1908:IGN

[11886] Anonymous. III. — the German Navy of to-day. Scientific American, 99(6):91–94, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n6/pdf/scientificamerican08081908-
91.pdf.

Anonymous:1908:PEC

[11887] Anonymous. Progress of excavation of Culebra Cut. Scientific American, 99(6):94, August 8, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n6/pdf/scientificamerican08081908-
94.pdf.

Boyer:1908:OBR

[11888] Jacques Boyer. The opacity of bones to Roentgen rays as an indication of age. Scientific American, 99(6):95, August 8, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n6/pdf/scientificamerican08081908-95.pdf.

Anonymous:1908:Cbo

[11889] Anonymous. Correspondence. Scientific American, 99(7):107, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n7/pdf/scientificamerican08151908-107.pdf.

Anonymous:1908:RPIbu

[11890] Anonymous. Recently patented inventions. Scientific American, 99(7):112, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
112.pdf.

Anonymous:1908:NQbh

[11891] Anonymous. Notes and queries. Scientific American, 99(7):112– 113, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
112a.pdf.

Anonymous:1908:BNBbh

[11892] Anonymous. Books: New books, etc. Scientific American, 99 (7):113, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
113.pdf.

Anonymous:1908:IIbx

[11893] Anonymous. Index of inventions. Scientific American, 99(7): 113–115, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
113a.pdf.

Anonymous:1908:PNPbk

[11894] Anonymous. Patent notes: Patents. Scientific American, 99(7):116, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n7/pdf/scientificamerican08151908-116.pdf.

Anonymous:1908:SDI

[11895] Anonymous. The successful defender of the international motor boat cup. Scientific American, 99(7):102, August 15, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n7/pdf/scientificamerican08151908-102.pdf.

Anonymous:1908:CBHa

[11896] Anonymous. A cruiser-battleship holds the transatlantic record. Scientific American, 99(7):102, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
102a.pdf.

Anonymous:1908:ZAD

[11897] Anonymous. The Zeppelin airship disaster. Scientific American, 99(7):102, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
102b.pdf.

Anonymous:1908:MMB

[11898] Anonymous. May a manufacturer buy a patent, never use it, and sue for infringement? Scientific American, 99(7):102–103, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n7/pdf/scientificamerican08151908-102c.pdf.

Springer:1908:CB

[11899] J. F. Springer. Color-blindness. Scientific American, 99(7):103, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n7/pdf/scientificamerican08151908-103.pdf.

Anonymous:1908:MBJ

[11900] Anonymous. Military ballooning in Japan. Scientific American, 99(7):103, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
103a.pdf.

Anonymous:1908:CSz

[11901] Anonymous. The current supplement. Scientific American, 99 (7):103, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
103b.pdf.

Shepstone:1908:MMF

[11902] Harold J. Shepstone. A municipal “microbe farm”. Scientific American, 99(7):104, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
104.pdf.

Collins:1908:EB

[11903] Percy Collins. The evolution of the blow. Scientific American, 99 (7):105–106, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
105.pdf.

Anonymous:1908:MJC

[11904] Anonymous. The mechanical joys of Coney Island. Scientific American, 99(7):108–110, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
108.pdf.

Anonymous:1908:NSG

[11905] Anonymous. A no-segment graphite-resistance controller. Scientific American, 99(7):110, August 15, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n7/pdf/scientificamerican08151908-
110.pdf.

Anonymous:1908:IMB

[11906] Anonymous. The International Motor Boat Race for the Harmsworth Trophy. Scientific American, 99(7):111, August 15, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n7/pdf/scientificamerican08151908-111.pdf.

Anonymous:1908:Cbp

[11907] Anonymous. Correspondence. Scientific American, 99(8):123, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n8/pdf/scientificamerican08221908-123.pdf.

Anonymous:1908:RPIbv

[11908] Anonymous. Recently patented inventions. Scientific American, 99(8):128, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
128.pdf.

Anonymous:1908:NQbi

[11909] Anonymous. Notes and queries. Scientific American, 99(8):128, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n8/pdf/scientificamerican08221908-128a.pdf.

Anonymous:1908:BNBbi

[11910] Anonymous. Books: New books, etc. Scientific American, 99 (8):129, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
129.pdf.

Anonymous:1908:IIby

[11911] Anonymous. Index of inventions. Scientific American, 99(8): 129–131, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
129a.pdf.

Anonymous:1908:PNPbl

[11912] Anonymous. Patent notes: Patents. Scientific American, 99(8):132, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n8/pdf/scientificamerican08221908-132.pdf.

Anonymous:1908:PME

[11913] Anonymous. Paucity of monuments to engineers. Scientific American, 99(8):118, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
118.pdf.

Anonymous:1908:IEU

[11914] Anonymous. International electrical units. Scientific American, 99(8):118, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
118a.pdf.

Anonymous:1908:INY

[11915] Anonymous. Improvement of New York–New Jersey Suburban Service — II. Scientific American, 99(8):118–119, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n8/pdf/scientificamerican08221908-118b.pdf.

Horwill:1908:SSM

[11916] H. W. Horwill. Science and the schoolboy mind. Scientific American, 99(8):119, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
119.pdf.

Anonymous:1908:SNj

[11917] Anonymous. Science notes. Scientific American, 99(8):119, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n8/pdf/scientificamerican08221908-119a.pdf.

Anonymous:1908:CSba

[11918] Anonymous. The current supplement. Scientific American, 99 (8):119, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
119b.pdf.

Mclure:1908:DFS

[11919] W. Frank M’clure. Dry farming in semi-arid districts. Scientific American, 99(8):120–121, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
120.pdf.

Anonymous:1908:EPM

[11920] Anonymous. The Ekenberg process for manufacturing dried milk. Scientific American, 99(8):121–122, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
121.pdf.

Anonymous:1908:SLU

[11921] Anonymous. Side launching of the U.S.S. “Patuxent”. Scientific American, 99(8):122–123, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
122.pdf.

Anonymous:1908:RFA

[11922] Anonymous. Recent foreign aeroplanes. Scientific American, 99 (8):124–125, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
124.pdf.

Anonymous:1908:USBb

[11923] Anonymous. The United States and British Government airships. Scientific American, 99(8):125, August 22, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n8/pdf/scientificamerican08221908-125.pdf.

Anonymous:1908:BMN

[11924] Anonymous. A boat made of newspapers. Scientific American, 99(8):126, August 22, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n8/pdf/scientificamerican08221908-
126.pdf.

Bastin:1908:WDR

[11925] B. Leonard Bastin. Why do roots grow down and shoots grow up? Scientific American, 99(8):126–127, August 22, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n8/pdf/scientificamerican08221908-126a.pdf.

Anonymous:1908:Cbq

[11926] Anonymous. Correspondence. Scientific American, 99(9):139, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n9/pdf/scientificamerican08291908-139.pdf.

Anonymous:1908:RPIbw

[11927] Anonymous. Recently patented inventions. Scientific American, 99(9):144, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
144.pdf.

Anonymous:1908:NQbj

[11928] Anonymous. Notes and queries. Scientific American, 99(9):144– 145, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
144a.pdf.

Anonymous:1908:BNBbj

[11929] Anonymous. Books: New books, etc. Scientific American, 99 (9):145, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
145.pdf.

Anonymous:1908:IIbz

[11930] Anonymous. Index of inventions. Scientific American, 99(9): 145–148, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
145a.pdf.

Anonymous:1908:PNPbm

[11931] Anonymous. Patent notes: Patents. Scientific American, 99(9):148, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n9/pdf/scientificamerican08291908-148.pdf.

Anonymous:1908:MBP

[11932] Anonymous. Is the motor bus profitable? Scientific American, 99(9):134, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
134.pdf.

Anonymous:1908:PTF

[11933] Anonymous. Proposed test of full-sized Blackwell’s Island Bridge Chord. Scientific American, 99(9):134, August 29, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n9/pdf/scientificamerican08291908-134a.pdf.

Anonymous:1908:PTH

[11934] Anonymous. The pneumatic tire and the heavy commercial automobile. Scientific American, 99(9):134, August 29, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n9/pdf/scientificamerican08291908-134b.pdf.

Anonymous:1908:Cbr

[11935] Anonymous. On courtesy. Scientific American, 99(9):134–135, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n9/pdf/scientificamerican08291908-134c.pdf.

Plummer:1908:SAE

[11936] John Plummer. The Shackleton Antarctic Expedition. Scientific American, 99(9):135, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
135.pdf.

Anonymous:1908:WATb

[11937] Anonymous. The Wright aeroplane tests. Scientific American, 99(9):135, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
135a.pdf.

Anonymous:1908:RMD

[11938] Anonymous. Recent military dirigible balloons. Scientific American, 99(9):135, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
135b.pdf.

Anonymous:1908:SNk

[11939] Anonymous. Science notes. Scientific American, 99(9):135, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n9/pdf/scientificamerican08291908-135c.pdf.

Salvage:1908:WSF

[11940] W. L. Salvage. Wire suspension foot bridge at Platt National Park. Scientific American, 99(9):136, August 29, 1908. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v99/
n9/pdf/scientificamerican08291908-136.pdf.

Springer:1908:TH

[11941] J. F. Springer. Testing for hardness. Scientific American, 99(9):136, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n9/pdf/scientificamerican08291908-136a.pdf.

Anonymous:1908:NTC

[11942] Anonymous. A new type of clam-shell bucket. Scientific American, 99(9):137, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
137.pdf.

Anonymous:1908:SIF

[11943] Anonymous. Some ingenious foreign inventions. Scientific American, 99(9):137–139, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
137a.pdf.

Anonymous:1908:GPG

[11944] Anonymous. Gasoline-propelled gunboats for river service. Scientific American, 99(9):140, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
140.pdf.

Anonymous:1908:WBA

[11945] Anonymous. The Wright Brothers’ aeroplane in France and the United States. Scientific American, 99(9):140–141, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n9/pdf/scientificamerican08291908-140a.pdf.

Anonymous:1908:PIH

[11946] Anonymous. The propagating influence of high-tension electric currents upon plants. Scientific American, 99(9):142, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n9/pdf/scientificamerican08291908-142.pdf.

Anonymous:1908:PDn

[11947] Anonymous. Patent department. Scientific American, 99(9):143, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v99/n9/pdf/scientificamerican08291908-143.pdf.

Anonymous:1908:SNT

[11948] Anonymous. Some novel types of shaving brushes. Scientific American, 99(9):143, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
143a.pdf.

Anonymous:1908:HBS

[11949] Anonymous. Holder for brooms and similar articles. Scientific American, 99(9):143, August 29, 1908. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v99/n9/pdf/scientificamerican08291908-
143b.pdf.